A Complete Bibliography of Publications in Marine Ecology Progress Series: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

21 August 2024 Version 1.05

Title word cross-reference

− [CATAV18, FSE^{+}17]. ^{13} [CCJR14, CBR19, CSC^{+}11, FSE^{+}17, GFC^{+}13, GCS^{+}18, GCS^{+}19, GMT^{+}11, HHS^{+}12a, KBA13, KSK07, KBP^{+}10, MPB13, MHS14, MFS14, NMPM15, OPN10, OFT^{+}11, PPD^{+}14, PCW^{+}12, PHWC13, RP12, RLAGL12, RCVG10, SML^{+}14, UDF^{+}18, VRSS18, VSM11, VB14, VLM14]. ^{1}5 [AGNH^{+}19, CPEE18, GCS^{+}18, GCS^{+}19, HHS^{+}12a, UDF^{+}18, VRSS18, CCJR14, CBR19, CSC^{+}11, FSE^{+}17, GFC^{+}13, GM13b, HU13, KV13, MPB13, MHS14, MFS14, NMPM15, OFT^{+}11, PCW^{+}12, RP12, RET^{+}19, ROE17, RLAGL12, RCVG10, SE11a, SML^{+}14, VSM11, VB14, VWEP18]. ^{18} [AAAE17, DH18a, KSK07, KIU^{+}13, PHWC13]. ^{2} [SHG^{+}15]. ^{2+} [PZL^{+}17]. ^{3} [CWR^{+}18]. ^{34} [MHS14, MFS14, VRSS18]. ^{∘} [WAP10]. _{12} [BLHRP^{+}19]. _{2} [AM19, BRF11, BPM^{+}18b, CPEE18, CG16, CH15, DMW^{+}16, DB11a, DDP13b, DWL19, ENW^{+}14, EES15, FRC13a, FCR14, FRRÁSL11, GSA^{+}18, GMG^{+}16, GCP^{+}17, HJN^{+}10, HP13, HSB12, HYC^{+}14, KJW^{+}15, LGW^{+}15, LERS13, LMH12, MKM^{+}16, MKS18, MDM18, MSM^{+}12, PC12, PMS15, PKB16, RWM^{+}14, RMC^{+}14, RT17, SGS^{+}17, SHP^{+}14, SLIGG18, SCSC16, SSX^{+}17, TG12a, TZF11, TFÁS^{+}12, TDM16, TBL16, TPKC18, VJM^{+}16, YKG^{+}15, dlSGS17]. _{3}^{−} [DB11a]. _{4}^{+} [DB11a]. β [AAGGG19, BPSG11, WHB^{+}14]. : [KJW^{+}15, LMH12, MC10a, MFI11, VTZ^{+}13]. Δ [MSM^{+}12, CCJR14, CBR19, CSC^{+}11, DH18a, FSE^{+}17, GFC^{+}13, GCS^{+}18, GCS^{+}19, HHS^{+}12a, HU13, KSK07, KV13, KIU^{+}13, KBP^{+}10, MPB13, MHS14, MFS14, NMPM15, OPN10, OFT^{+}11, PPD^{+}14, PCW^{+}12, PHWC13, RP12, RET^{+}19, RLAGL12, RCVG10, SE11a, SML^{+}14, UDF^{+}18, VRSS18, VB14, VWEP18, VLM14]. F_{v}/F_{m} [SOB16].

-ATPase [PZL^{+}17]. -dependent [MO13]. -diversity [AAGGG19, BPSG11, WHB^{+}14]. -encoding [EH13]. -induced [HP13]. -recapture [PBM^{+}11].

/Mg [PZL^{+}17].

0-group [FWPP12, vdVFK^{+}10].

18S [HPMZ11]. 1970s [SSB^{+}16]. 1981 [GE14]. 1990 [AMLR^{+}17]. 1990s [AF14]. 1993 [RQN^{+}14]. 1997 [GFGGDSM12]. 1997/1998 [RCP^{+}13]. 1998 [GFGGDSM12].

2 [CRG^{+}16, GBH^{+}10, JPL^{+}15, RAB^{+}15, WNNH10]. 2000s [LOES11]. 2001 [DPO15]. 2002 [PPEH^{+}14]. 2004 [DFGH^{+}15, PPEH^{+}14]. 2005 [DWD^{+}18]. 2007 [DFGH^{+}15]. 2008 [RQN^{+}14]. 2009 [SSB^{+}16, SSdW^{+}12, WRDF12]. 2010 [ACA^{+}10b, Ano10f, Ano12c, DPO15, PPEH^{+}14, SSdW^{+}12]. 2011 [RUP11b]. 2012 [AMLR^{+}17, PPEH^{+}14]. 2013 [GE14]. 2014 [Ano15b]. 2014a [SB15a]. 2015 [Ano15a, Ano16, DWD^{+}18]. 2016 [BW16b]. 2100 [CATAV18]. 23^{o} [RHRB^{+}17]. 25 [BBG^{+}16]. 25-year-old [BMV^{+}17]. 25^{o} [GDB^{+}16].

3-dimensional [FCFD15, LSJ^{+}15]. 300 [CATAV18]. 349 [Ano10e]. 393 [Ano10d]. 3D [AAM^{+}12].

401 [Ano10f]. 404 [Ano12c]. 410 [ACA^{+}10b]. 432 [MHO^{+}11]. 435 [RUP11b]. 438 [Ano14b]. 441 [Ano12d]. 442 [Ano12d]. 454-pyrosequencing [VBŽ^{+}17].

500 [Ano14c]. 507 [Ano15b]. 534 [Ano15a]. 538 [Ano16]. 54 [WAP10]. 559 [Ano19d]. 561 [BW16b]. 589 [BS18b]. 595 [Ano19a]. 598 [Ano19c].

614 [Ano19b]. 627 [Ano21].

79 [WAP10].

A. [HCM^{+}11, RGWA11]. aalge [ERB^{+}13b, ERB^{+}13a]. abalone [AM13, BRB11b, CHB^{+}13, MKK^{+}11]. abandon [Chi11]. abatement [MZC^{+}10]. Abatus [CBKM16]. Abbott [HW14]. abbotti [HW14]. abbreviated [OT13]. ABEC [TLB15]. Ability [SZ13, MBBBC12, MCUO12, NMS^{+}18, PT13, She12, SKW^{+}14, TSCSB15, Ued14, WHL11]. Abiotic [KK17, MSK^{+}19, AOW12, BAL^{+}11, EDS14, HMKF11, HCA11, LMG12, MYS^{+}16, PML^{+}10]. ablation [HJ13]. abnormalities [OCDD16]. above [PMYH13]. aboveground [GP12]. Abrupt [CPR^{+}16, OV10, BHE14]. Absence [BBC12b, GDSPRA13]. absorbing [EMG^{+}15]. Absorption [BTS^{+}14, MSS^{+}14a, TZF11]. Absorption-based [BTS^{+}14]. Abundance [BS11, BVAB^{+}17, CJS^{+}12, HRAM13, MSC13, PWCvO^{+}14, SDB12, ASR^{+}19, AMSdL^{+}18, Alb15, AAGGFG^{+}12, AJ16, BGB^{+}17, BMM17, BBBF15, BCdV13, Ber12a, Ber12b, BJMS12, BS12, BD14, BDDvdM17, BBS^{+}17, BSC17, BDH^{+}17, BAW11, BDVP19, CBC^{+}10, DB14, DA10, DHH10, DMJ11, ET11, EBDB18, EFOM18, FKZTT18, FNA^{+}16, FJD13, GHES18, GRCP^{+}10, GVW^{+}12, GJSR18, Goy14, GKR^{+}15, GGB^{+}18, HLS17, HGL15, HS15a, HHH^{+}14, HFR^{+}13, HSMB17, HMAC^{+}11, HFN14, HORRLS^{+}14, IMS^{+}11, JUMS10, KSA^{+}13, KB12, KWDD10, LPC^{+}11a, LRHP18, LSH^{+}19b, LMG12, LLBP15, LANER15, LCL^{+}13, LRS17, MDLM17, MRC10a, MAP^{+}12, MJ10, MHFN10, Met11, MRB^{+}11, MGB^{+}19, MPG^{+}12, MJHS13, MSK^{+}19, MS10, NTN^{+}13, NMG^{+}13, NUO^{+}19, OCDD16, PSCS16, PS19, PT13, RPP^{+}13, RLO18, RL11, RZPH14, RPCL19, RDF^{+}19, SGC^{+}13, SPH^{+}17, SB18]. abundance [SHB^{+}10a, SHB^{+}13b, SNNS^{+}12, SBMÅ12, TBH^{+}18, TM10, TSS^{+}12, TYG12, TBS^{+}17, TKIO10, TPCN12, TBMdV12, TNN11, VDG^{+}15, VM14, WHvdK11, WCP^{+}14, WHB^{+}19, Wil14, WPB^{+}14, YAL12, ZAC11, ZBC10]. abundances [KHBL11, RCU^{+}11, TAAN15, WK10]. Abundant [HAF^{+}17, SCP16, BYAA13, BH15, BHF^{+}16, HPL10, MMB^{+}12b, SL15b]. abyssal [ALSO^{+}18, DBH^{+}16, ENO^{+}11, FKRJ13, KKS^{+}14]. Acanthaster [BBW^{+}19, KHH18, RPB17, WGLDB15]. acanthias [SGB^{+}10, TNB^{+}15]. Acanthina [SSBC11, SHMB15]. acanthistius [BJCB^{+}14]. Acanthopagrus [RA10, RGWA11, WJHS13]. Acartia [ACDJ11, BBRBG10, CLC12a, ET11, LME^{+}19, PFPP18]. Acartiella [SIK16]. Accelerated [DdR10]. accelerates [KCB18, RG14]. acceleration [VGR^{+}16]. accelerometer [BGDC14, WMHW14]. accelerometers [WPS^{+}19]. acclimated [ENW^{+}14]. Acclimation [LME^{+}19, dlSBB^{+}10, ATO^{+}14, FMC^{+}11, LH14]. acclimatised [CA14]. acclimatization [HAF^{+}17, SMOH19, ZRBV15]. accompanied [PSC^{+}13a]. according [BPB18, DGC12, VBT^{+}17]. account [HDP^{+}17]. accounting [FBPCC^{+}19, PMKL11]. accretion [CLL15, SGTD14, WSK^{+}16]. accumulation [ATS14, HSM13, HDP^{+}17, MHN^{+}12, NCS^{+}10, PLF^{+}12b, PLZL14, SM10b, TB10]. accumulations [LBH18, QNG^{+}15]. Accurate [MWB13, DOH^{+}17]. accurately [YIKG10]. achromatic [LCCT19]. acid [AR14, ATP^{+}12, BSG^{+}10, BCGWS19, BPH^{+}19, BADQ^{+}15, BNE^{+}19, BDS^{+}18b, CWD11, CBR19, CMW^{+}15, CBD14, CL10c, DWPH11, DE17, DPH14, FRGI15, GED^{+}14, GHNF10, HRPBLT18, JBS11, KKMA^{+}10, KJW^{+}15, KKS^{+}14, KBP^{+}10, LLv13, LRR^{+}10, LFM^{+}16, LBL^{+}12, LBE^{+}10, MVP^{+}16, MHHF13, NW15, OBI^{+}16, PLG^{+}19, PBM^{+}10, PBAP^{+}14, PIPD10, QSB16, RLG14, RVF14, RCR^{+}13, RT12, TTS17, TFM^{+}18, THC12, VBT^{+}17, VLM14, VMC^{+}19, WBI^{+}15, WNNN12, WVD^{+}17, YSC^{+}15, ZdVA^{+}16, vdBUPP^{+}11]. acid-base [ATP^{+}12]. acid-carbon [WBI^{+}15]. acid-related [BSG^{+}10]. Acidification [BKHW13, LZ12, MBBK16, ACB12, ATP^{+}12, ATO^{+}14, BGJ^{+}19, BDPD14, BRIC13, BGBD^{+}13, CCM^{+}14, CSJP17, CLMB13b, CG16, CAG12, CESC14, CS13, DBG15, EKE16, EVB^{+}16, EVHA^{+}18, FRRÁS^{+}12, GHHH12, GLHM^{+}18, GLW^{+}17, GB10b, HMP15, HM17, HBBH12, JDX^{+}17, KCB18, LME^{+}19, LBP^{+}19, LTG17, LHB19, MP15, MFG^{+}13, MJT^{+}14, MJM^{+}14, MTM^{+}17, MWBMH13, MCR^{+}13, MW16a, MGD^{+}11, MMGB14, OTS^{+}10, OSL18, PL15, PSPR18, PZL^{+}17, PRP^{+}16, PHTS11, RMB^{+}14, RFM^{+}15, RSW13, SGL^{+}17, SSM^{+}16a, SSM^{+}17, SG18a, SG18b, TRL^{+}15, TBH^{+}17, VJM^{+}16, VHM16, VMWU15, WAP10, WDL^{+}18, Whi11, WNP^{+}14, YLPG19, ZGH^{+}17]. acidification-mediated [PRP^{+}16]. acidifying [CBD14]. acidosis [HP13]. acids [BMJH12, BLPP16, BGS^{+}16, BFEMLÁJ19, CMIMGS15, CCG^{+}10, CLC12a, DBC^{+}12a, DSG^{+}13, EMG^{+}15, GKHM13, GNHLN11, ILDL10, KS12a, LAL^{+}18, NLK^{+}19, OCMT13, PFG10, RKM^{+}10, VTZ^{+}13, VWEP18]. Acipenser [SNS^{+}13, ZBK^{+}14]. acorn [BSK^{+}11, FSR10, SM17b, TCWC13]. acoroides [PBC^{+}15]. acoupa [VWEP18]. Acoustic [BNAK11, BRLM17, BFG^{+}11a, DFD^{+}19, DGP^{+}19, EFOM13, GSÁCM^{+}16, GJ13, TB14b, VVB^{+}10, WGS16, AMP^{+}11, BSB^{+}18, BBKH^{+}11, BMJ16, BHL^{+}19, CCS^{+}18, EFOM18, GOO14, GSJS^{+}19, LCG^{+}17, LOBW^{+}11, LBJI14, LEB14a, MGM^{+}18, MOMF14, MHS15, MBKM17, PGWS10, PIGK16, PCLB18, RIM^{+}13, RCD^{+}13, RNSA15, SAWK18, SBG^{+}16, SKK12, SWH^{+}11, SD17b, TBB^{+}19, TS11b, USR^{+}16, WSLM13, WWGD13]. acoustical [SRK11]. acoustically [AvBBL^{+}16, LBJI14]. acoustics [SMSB19, TRJ^{+}11]. acquire [LRC^{+}13]. Acquisition [HGPM10, CFPC^{+}13, ORJH19, SH10]. Acropora [CESC14, DSVM10, LSBB17, LPC11b, MMRDS18, PWCvO^{+}14, RBL^{+}17, SWB^{+}10, SAH11, SYK^{+}13, SKFY18, TPB^{+}16, TPB^{+}18, YSS^{+}18]. acroporid [Fog12, NKS^{+}15]. Acroporidae [WMS^{+}11]. across [ARMA15, ASCHM19, AY10, BG19, BWV^{+}12, BCOMKJ18, BO15, BGSM13, BPN11, BPT15, BTLF17, BWH^{+}19, BOKJ18, BMPC15, BPB11, BWCZ11, BAE^{+}10, BBH^{+}12, BDB^{+}12, CKJOM14, CBJB19, CWAS17, DR16, DBE^{+}12, DZDB10, DHR^{+}13, EFOM13, FGOS16, FPM14, Fog12, FL17, FBET15, FKRJ13, GOC^{+}18, GE17, GBH^{+}10, GGCB16, GKR^{+}15, GDPR18, GHT^{+}15a, HBK^{+}17, HS17, HCL^{+}11b, HEM^{+}12, HM13, JMB^{+}17, KHW^{+}10, LMG12, LYH^{+}17, LKB^{+}14, LCB^{+}12, LGT^{+}14, MC10a, MJR^{+}14, OPN10, PDC^{+}17, PMOAC^{+}17, PLZL14, RRL^{+}19, RM19, SAWK18, SLL^{+}18, SMM^{+}15, SCB^{+}14, SST19, SPP15, SH14, SFL^{+}18, SMK^{+}17, SNT^{+}18, TL18, TdRH17, ULCAZG17, VW10, Vig15, VJJ^{+}19, WW15, WAA^{+}19, YSH^{+}16, ZO10, ZKS17, vDBH^{+}15]. act [GBD13, HFM12]. acting [SMR14]. Actinia [DMK11]. action [GPB^{+}15, Kat16]. actions [HT10]. Active [AAM^{+}12, BBKH^{+}11, DVV^{+}18, AAK^{+}15, BST19b, BMVAS19, LKL^{+}19, SWH^{+}18]. activities [BVS^{+}11, CRL13, CAB^{+}16, DGMB10, OGA13, TCJ^{+}19, VPW^{+}12]. Activity [MOK^{+}11, AFC^{+}16, AN11a, ASS^{+}13, AGGAO^{+}18, ARG^{+}10, BWV^{+}12, Ben16, BPT15, BNS^{+}14, BGDC14, Car12, CLC^{+}12b, CASG^{+}19, FSY^{+}10a, HHS12c, KRC17, LFPA^{+}11, LTG17, MH11b, MIA^{+}11, MG14, MCS^{+}15, MAS^{+}10b, MR16, PGWS10, PK13, PZL^{+}17, PVYS16, RB15, SHUS17, SHT^{+}18, SS10b, SK10, SRM^{+}15, SPP15, SGTP16, SKR^{+}10, ST14, SBE^{+}19, SJO16, TIN^{+}18, VJM^{+}16, VRAP^{+}14, WEAG^{+}15, WJWC16, WNNH10, WMHW14, ZHTA13, ZVR18, ZFC^{+}13, ZAZS12, MPS^{+}10]. activity-specific [WMHW14]. actual [KP18]. aculeatus [SML17]. acuminata [NKH12]. acute [BPLT14, BPM^{+}18b]. acutidens [CBL^{+}10b]. acutorostrata [dB10]. acutus [WPNK12]. adaptation [BSP^{+}10, BSCH14, CBF^{+}19, HDWF10, LME^{+}19, PHHL^{+}11]. Adaptations [CIM^{+}13, BCOMKJ18, OT14, SUPCB11]. Adaptive [EBOM11, FDP^{+}16, LAK^{+}19, MPW12, VBV^{+}10, WTD^{+}19]. add [HKL12]. Adding [ARW12, BC14]. Addition [GSF^{+}13, WBLS^{+}11, BVGH10, CLJF12, KCR^{+}10, OMM^{+}12]. additional [Miu12]. additions [HCK12, LMR^{+}12, MGFÁS^{+}10, ROE17, TMGCM11, WWO^{+}15]. Additive [FAT^{+}15, HGKB14, CFA16, HWTH12]. address [Nob11]. Addressing [DSG^{+}13, PCS^{+}15]. adeliae [CHP^{+}11]. Adélie [ADL^{+}18, BDNA10, BST^{+}19a, CHP^{+}11, CFPF^{+}14, EPS11, SDM^{+}13, SE16, SEFS15, TIN^{+}18, WLB^{+}15]. adequacy [PSdA^{+}17]. adequate [DCW^{+}10]. adhaerens [GRG^{+}12]. Adjacent [RSJ14, ASK^{+}15, GMB^{+}16, LDVP^{+}17, MKII19, NUO^{+}19, PSÓ^{+}15, RPL^{+}18, TL18, TT16]. adjust [MMB^{+}19]. Adjusted [MSV^{+}11]. adjusting [BDNA10]. adjustment [BCGWS19, KCCG12]. adjustments [CdTK^{+}16, KPW^{+}19]. administration [CMK^{+}14]. adriani [HGFA^{+}11]. Adriatic [CSPA10, CFC^{+}12, KLMR14, MGB^{+}14, SCD^{+}17, TFFPM18, VBŽ^{+}17]. Adult [Rit17, BWM^{+}15, BJRH10, CWS13, DAC^{+}12, FHMG19, GS11, GDJdBSL19, HCW^{+}15, HAHM16, JWHR11, JJF14, KRdVG^{+}18, LA14, LHHG11, MH19, MAS^{+}10b, RBSL11, REAN^{+}14, RBGT10, SES12, SL11b, SB18, SFKQ15, TGP^{+}16, WTC^{+}15, WRW^{+}16b, WNA^{+}15, WHD^{+}14, vdWTWS15]. adults [APC^{+}14b, GGMGR12, PCF^{+}19]. advancement [LFN^{+}12]. advances [BPSK11, FPC^{+}18]. advantage [PMKL11]. Advection [SRG14, DWD^{+}18, FPM14]. Advection-diffusion [SRG14]. adverse [VSB^{+}15]. aeglefinus [AGIW16, BCSB13, RP12, SKN^{+}15, WTGG10]. aequinoctialis [PDP^{+}10]. aequituberculata [EES15]. Aequorea [KBR^{+}11, Pur18]. Aerial [BBBF15, LPC^{+}11a, SPHK17, BMVBC10, KMGH16, SHP^{+}18]. aerobic [BSN^{+}18, CTS^{+}15]. aerosol [GLY^{+}14, LWD13]. Aethia [DKB^{+}12]. affect [AMAM^{+}18, BVS^{+}11, BFKR17, BJS^{+}16, CPTL11, CRRL17, CDMB^{+}18, CvOB18, DRT13, GSU12, HCK12, HBG^{+}17, JH11, JWB^{+}17, KJCN^{+}11, LZ12, LMd12c, MBR^{+}13, MD16a, MGD^{+}11, NM18, PG15, RB18, RKT16, SRB^{+}17, SWB^{+}10, WR12, WK16, ZdVA^{+}16]. affected [BPWF14, BMMS10, BAVSP11, KSA^{+}13, PSPR18, RSE18, SMK^{+}17, WJHS13]. affecting [BG19, BW13, CHP^{+}11, CGEGC^{+}19, DAC^{+}12, HBS17, JMB10, KAAI15, KLNJ15, MUC12, MOMF14, OFFQGG10, PML^{+}10, RLBC10, WS12]. affects [AvBBL^{+}16, AFMR^{+}13, BFS10, BLD^{+}16b, BSN^{+}18, BIL^{+}10, BNW14, CMR12, CF12, CMO^{+}10, CLMB13b, CMAJ12, Col10a, CME^{+}12, DKB^{+}12, EM13, FGP17, HLL^{+}17, HRK^{+}18, KTP^{+}18, LAM13, MJM^{+}14, MYK^{+}16, MH11b, MP12, PMS15, PCH^{+}14, PSC13b, RLG14, REAN^{+}14, RMB18, SES12, SHP^{+}14, SML^{+}15, SBW16, TBC^{+}17, WLM^{+}13]. affinis [PL10, PFG10]. affinities [WGD10]. affinity [JRW^{+}14, TRMC11]. Affordable [DOH^{+}17]. Africa [CP10, ARvdW^{+}16, AFH11, ASF13, BP12b, CGM^{+}19, DFS13, DPC13, ECM^{+}12, HVJ^{+}18, INK^{+}17, McC19, MKG10, MMB^{+}19, OCS14, PFB^{+}17, PSHG16, RFBH^{+}11, RMNP17, TMS^{+}16, ZBP18]. African [CMF^{+}19, DKV^{+}16, DAC^{+}12, DCW^{+}10, DBS^{+}19, KCH^{+}11, MRL^{+}17, OCSW16, PBG^{+}10, PW19a, PBCA10, PJMP10, SUB^{+}13, TP19, WPRS18, vERRP16]. africana [CP10]. africanum [CLMM^{+}14]. Afropinnotheres [DMHS^{+}14]. after [BK14, CSM12, DB14, GFGGDSM12, HSC^{+}15, HJ16, KKL13, KHW^{+}19, LS12, LPSL14a, MMBA13, NMS^{+}18, OSPWB15, RM13, RCP^{+}13, TKV^{+}19, WEA^{+}17, vWSGL11, vdPLdV^{+}11]. against [BC14, Goc10, HCM^{+}11, KPW^{+}19, MDPG13, QYS^{+}12, SKD^{+}11a, SJO16, TBG15]. Age [AGIW16, BCB11, DH17, JWR^{+}12, JWB^{+}17, LABM17, PSAK12, PAC^{+}15, SK11, AGH^{+}12, BD11b, BBL11, BMT17, BMF^{+}15, CPB^{+}11, CCM^{+}15, CS11a, CGBGS18, DC15, DHR^{+}13, FCR12, HLS17, HLC^{+}18, HBP^{+}16, LRS^{+}12, MRE14, MH12a, MGAV15, MWHM18, PPML17, PRNR^{+}18, PRBR11, SH12, SBL^{+}10, TMP^{+}11]. Age- [SK11]. age-0 [LRS^{+}12, MGAV15]. age-based [HBP^{+}16]. age-biased [CGBGS18]. Age-related [AGIW16, TMP^{+}11]. Age-structure [DH17, CS11a]. Age-structured [BCB11]. ageing [RGCT^{+}10]. Agent [TMT^{+}12, VBR^{+}14]. Agent-based [TMT^{+}12, VBR^{+}14]. Aggregate [FOL12, BBH^{+}12, FRJ10, GGF^{+}12a, GL12, WS10]. aggregate-species [GGF^{+}12a]. aggregates [MLP^{+}14, PSF^{+}19, WA14]. aggregating [FCFD15, RDF^{+}13, RTCF^{+}17]. Aggregation [Cra12, NCK^{+}14, RPB17, Bea11a, BN16, CCS^{+}18, CNJ^{+}18, DMd14, DFD^{+}19, FM14, GGF^{+}12b, KH10, LGS10, LBWB^{+}13, MSH^{+}14, RTM11, RPM^{+}13b, RSA^{+}12, SBFP10, SDM16, SMF^{+}11, TMM^{+}19, TGH15, WOWG10, YChH11, ZBB^{+}19]. Aggregations [GBE^{+}17, MBR^{+}13, SK10, ALK17, BPP^{+}18, BJCB^{+}14, BDB^{+}13, DRT13, DP14a, DSGR^{+}15, EZZ^{+}12, GSÁCM^{+}16, HCL10, JGM^{+}16, KNY^{+}15, KCLP16, RNSA15, SHP^{+}18, SHM^{+}10a, WLBB^{+}13, YYW14]. Aggregative [FSLG12]. aggression [EMJ18]. agonistic [BMSS14]. agricultural [TKV^{+}19]. Agulhas [BG12]. ahead [BFG^{+}19]. Aialik [HMAC^{+}11]. aids [AAE14]. air [CNSW16, DHSL11, MSM^{+}12, PJDC^{+}13, TBS^{+}17, dJdBdJB12]. air-breathing [CNSW16]. Akashiwo [JPP^{+}17, QYS^{+}12]. Akodon [PCA^{+}15]. al [AvdHH^{+}17, Ano10e, Ano12d, Ano14c, Ano14b, Ano15a, Ano15b, Ano16, Ano19b, Ano19a, Ano19c, Ano19d, Ano21, MAD18, MAB^{+}12a, TSTP13b]. al. [AF14, BBRM14, BD11a, BSBR^{+}12, BM10b, Car12, CSS^{+}13, DL13, FLC^{+}14, Fry13b, GBH12, GMR13a, HS11a, HG13, HBCK13, INK^{+}17, LGL11, MPM^{+}18, MBS^{+}14, POL^{+}11, SB15a, SHM^{+}12, SL15a, TSS11a, dQSC14]. alalunga [MIM^{+}12]. Alaska [AMM^{+}12, BJS^{+}16, Ran18, SNVH19, BZC^{+}12, BLMD^{+}11, BWH^{+}19, BWP^{+}11, CRS^{+}18, DIB15, DPH13, FGTS16, HAHM16, MHR19a, MIA^{+}11, MBB^{+}12b, NRK^{+}19, PK13, PCH11, RZPH14, SGC17, SQS^{+}13, SE11b, SOC13, SMF10, SB14, TSS^{+}12, Urb12]. Alaskan [CMC^{+}15, HA15, HMAC^{+}11, JS12b, SAM18]. albacares [KRR^{+}18, ODK^{+}14, RTCF^{+}17]. albacore [Gla10, MIM^{+}12]. Albatross [NSY^{+}18, AGHC10, CXdG^{+}17, CRJ^{+}18, MTB^{+}17, PWM^{+}19, SCW^{+}18, VMC^{+}19]. Albatross-borne [NSY^{+}18]. Albatrosses [CPW15, CAB^{+}19b, CMB^{+}14, DMT^{+}14, MPS^{+}10, TTB^{+}11, XWE^{+}14]. albicostatus [FC19]. Alboran [MSS^{+}14a]. alcohol [CWD11]. Aldabra [HSH12]. Alderdice [Kin10b]. Aleutian [RZPH14]. Alexandrium [BFL^{+}11, RZH^{+}16b, STT14, SE11a, SBKA14, TLM^{+}11, VTZ^{+}13]. alfredi [CNJ^{+}18, DBB11, PSD^{+}19]. alga [BMMS10, CSJP17, DDP^{+}13a, DRDV15, GS18, GTF10, KS12b, LKDP17, MAS10a, TG12a, TBKG10]. Algae [KKMM16, AS14, AS15, BWV^{+}12, BMVBC10, BRF11, CCM^{+}14, CLJF12, CTMU10, CLL13, DZDB10, DDP13b, EMG^{+}15, GGL^{+}17, HWTH12, JKM11, KOLDP17, KEH10, LWS^{+}10, LCC12, MWA17, PNPF14, PMSGBL16, PHTS11, TW16, TTS17, VDS11, WP10]. algal [ASAG17, ATS14, AS15, AFMR^{+}13, ASM10, ASK^{+}15, BHF^{+}10, BMSB13, BDPD14, BSD^{+}14, BBG^{+}11, BC13, BOS18, CMAJ12, CAB^{+}14, CFA14, CRF13, DHAV13, ESBG14, FCR14, FPC^{+}12, FTH12, GCC^{+}14, HLS15a, HVT12, HPP^{+}10, HMM14, JT10b, KKS12a, KF13b, LBM^{+}18, LRR^{+}10, LWP^{+}13, LMP12, MJR^{+}14, MTM^{+}11, MTMM^{+}15, MHF^{+}14, NOT^{+}11, OSE13, PLG^{+}16, PQAH^{+}17, PC19, RLG14, SLS18, SLL^{+}18, SLT^{+}19, SCLB11, SS12b, SBMÅ12, TS12, TCA^{+}11, TMP^{+}13, VHKT15, ZAMB10]. algal-dominated [DHAV13]. algal-dwelling [HVT12]. algorithm [BTS^{+}14]. Algorithms [MTMM^{+}15]. algosus [ZBP18]. alien [BFG^{+}19, HTHH11, LFMC11, ZBP18]. alive [HH15]. Alkaline [PK13]. alkalinity [GUC^{+}15]. Alle [KHW^{+}10]. Allee [RPB17]. allele [McD13]. allelochemicals [QYS^{+}12]. Allelopathic [PSPK10, SHSK11, TG10]. Allelopathy [QYS^{+}12]. allocation [BJM19, CLSS17, GDA^{+}18, HPMB11, HH15, JS12a, KBA13, MRC10b, PO13, PDC^{+}17, TMB^{+}12]. allochthonous [BW16b, BW16a, FK10]. allometry [MFI11]. allopatric [GDSPRA13]. allow [PNPF14]. alone [BMVAS19, CVM^{+}17]. along [ANRT10, AFH11, BHT15, BSD^{+}14, BPSG11, BLM^{+}18, BDF^{+}11, BAvW12, CAR^{+}19, CRS^{+}11, CDCZ16, DAL^{+}10, DRT13, Det10, DHGW15, Ebe10, FFS^{+}15, FTS^{+}13, FSS^{+}18, GHGF12, GMSC14, HHS10, HMS10, HVT12, HU13, HTB^{+}15, HBK^{+}12, JTTP19, KO12, KAD^{+}15, KWW^{+}14, KFW^{+}18, KCH^{+}11, KRDV18, LAK^{+}19, LAB10, LAM11, LLC^{+}14, LLF^{+}18, MZ10, MBB^{+}15, MF15, MWWS17, MZF^{+}19, MU13, MGOL19, MF10, NKM^{+}15, NKS^{+}15, NBGGST^{+}15, NCS^{+}10, NYK^{+}13, NDD^{+}13, NCK^{+}14, OBA^{+}18, PGR^{+}10, PHD^{+}13, PML^{+}10, PB10, PÁIIS^{+}19, PYBB11, RMKR13, RSC14, SAL^{+}12, SGTD14, SWW11, SBM^{+}18, SA17, TRLS^{+}11, TKS^{+}14, TT16, TMP^{+}13, VDM^{+}10, WHB^{+}14, WPRS18, WMB^{+}14, WNA^{+}15, XS19, YHR16, ZRR^{+}13, dB10]. alongshore [DEM15]. Alopias [CWK^{+}10, CWW^{+}16a]. Alosa [SSS19]. alpina [KMSW10]. alter [BAT14, CGC15, DS10, FMG^{+}18, GGNL^{+}17, LHB19, LFN^{+}12, RWM^{+}14, RLO18, SGS^{+}17, SK13, SSM^{+}17, Sma13, TTB^{+}11, TBH^{+}17, WWO^{+}15]. Alteration [Col10a, CMK^{+}14, DdRH14, GHGF12, KMV18, PGR^{+}17]. alterations [MVCD14, PUP16, SA17]. Altered [Nor13, CLS^{+}16, EM13, KS12b, MD11, OE11, RBD^{+}16, TPC^{+}15, TT16]. alternate [FMH^{+}19, HGL15, HT10]. alternate-stage [HGL15]. Alternating [SDAH12]. Alternative [Fry13a, AdCM^{+}19, FDS14, FJL12, OV10, PDT^{+}11, SAM14, WHK^{+}11]. alternatives [DMD^{+}12]. alterniflora [BPB18, FCR^{+}15, HL14, KM10b, MA13a, MKMZ11, SWWK13, WWO^{+}15, ZH13]. alters [BGM^{+}18, BKHW13, BSD17, BGM^{+}17, CHDH^{+}19, CS13, JC17, LBM^{+}18, MWD10, MF13, OTS^{+}10, PZL^{+}17, PSA17, PT13, RBHR19, RG14, TDKS10, TVH^{+}10, VBR19, WF18]. Álvarez [BBRM14]. Alvinocarididae [YWiIK15]. always [QSvN^{+}11, TB14a]. ambient [KMPS15, KIU^{+}13]. amboinensis [KHS^{+}11]. ameliorates [EF16]. America [CBC^{+}16, GHL11]. American [BTO^{+}18, BPOS^{+}16, BPSCD17, BSRDR10, CCHG10, CS11b, DCC^{+}10, DFTZ^{+}15, GW15, LS10b, MLCC19, MCC^{+}16, MR16, OCE^{+}12, RS13, SSS19, STS^{+}18, SPS16, SH17, VCG^{+}15, WPNK12]. americana [TLAT13]. americanus [CCHG10, CS11b, LS10b, MKS18, MCC^{+}16, MLF^{+}18, MR16, RS13, SPS16]. Amiantis [MPD11]. Amino [CBR19, HRPBLT18, BLPP16, DWPH11, DPH14, EMG^{+}15, GKHM13, NW15, OCMT13, VTZ^{+}13, VWEP18, VLM14, VMC^{+}19, YSC^{+}15]. Amirante [PSD^{+}19]. Ammodytes [DHG16, EvDB^{+}14, WHN^{+}18]. ammonia [BKHW13, TKW^{+}14, HSRP11]. ammonia-oxidising [TKW^{+}14]. ammonia-oxidizing [BKHW13]. Ammonium [STT14, FUOG^{+}16, FF12, VPBV13, VBJR^{+}13, VSB^{+}15, VBGO^{+}16]. Amoebophrya [SS10a]. Among [SKR^{+}13, ADF^{+}18, AS15, AMM^{+}12, AAvS^{+}13, Ano17c, BJCB^{+}14, BBC^{+}17, CP11, CMB^{+}14, CCKB^{+}15, CLSS17, DGC18, DHJ12, DMC^{+}16, DGFH10, EN11, EPF^{+}17, FSS^{+}18, GMG^{+}16, GL16, GLR^{+}14, GSS13, GBJ^{+}14, GB12, GPA^{+}16, Goy14, GRN14, HKI^{+}17, HTH10, HDP^{+}17, HDT^{+}12, HSWS16, JW10, JWHR11, JBL^{+}16, JWBS10, KRR^{+}18, KHN15, KNB^{+}13, LH16, LPI^{+}14, LSH^{+}15, LP10, LS19, LBC^{+}19, LFN^{+}12, Lyo14, MGP12, MH19, MGPC13, MGL^{+}16, NRF11, NBM12, NJJ^{+}11, NvOH13, ORM^{+}13, OKLM14, PBCA10, PKBS14, PCW^{+}12, RR18, RUPT^{+}11, RVF14, RBD^{+}16, RMT^{+}13, SAZHAAHS14, SBB^{+}12a, SGGL15, SH10, TKEV^{+}19, TPB^{+}16, TRSV^{+}12, TB12, TUK13, TSF^{+}10, WSHY16, WOHJ^{+}15, WJ12, YSS^{+}18, YMD^{+}10, dG14]. Among-year [SKR^{+}13]. amongst [GHNF10]. amount [PMKL11]. Amphibolis [GLM^{+}10]. amphidromous [FZI^{+}16, HS16, LLD^{+}12, MKII19]. Amphipholis [BEFC15]. Amphipod [MH13, ADO^{+}12, AAvS^{+}13, BFS10, BS14a, DB11b, DBC^{+}12b, FKRJ13, GFGGDSM12, LHPP16, LBC^{+}19, MHHF13, NPRAGG^{+}12, NWC^{+}10, PMS10a, RCB12, SSM^{+}17, SLPP^{+}14]. amphipods [CVP11, FLAL10, FHG^{+}14, SSM^{+}16a, ZAMB10]. Amphiprion [SBB18]. Amphistegina [MWBMH13, PUP16]. Amphiura [CPSC^{+}15, CPSC^{+}18]. Amplification [KSS13]. amplifies [VSB^{+}15]. amplify [VMWU15]. amplitude [KF13b, NKA16, RFJ^{+}10]. Amur [KN12]. amurensis [GSTK11, MS13]. Anableps [KCE14]. anadromous [ABT^{+}15, AHØ^{+}15, JJJ16]. analogue [FKT15]. Analyses [GBMW12, AAJJ^{+}19, AR14, AGP^{+}13, BBB^{+}15, BADQ^{+}15, CLD^{+}13, CGG10a, CMB^{+}14, DEWK15, EAB^{+}19, FC10, GCS^{+}19, GSSM^{+}11, GPC^{+}15, HB12, JBS11, dDTC^{+}17, KKS^{+}14, LKC^{+}15, LLBT18, LAL^{+}18, LPK^{+}16, LCB^{+}12, MGAV15, MBS^{+}13, MLDG12, NUO^{+}19, OBI^{+}16, PLG^{+}19, PMKL11, RKT16, RCR^{+}13, SCB^{+}16, SWH^{+}18, SH17, SJL13, TAEB^{+}11, WBI^{+}15, WLF^{+}17, YSKH11, ZO10]. Analysis [BFEMLÁJ19, CLP18, ALV12, ADF^{+}18, ALE^{+}13b, ASWB10, Ano17c, AFH11, AGRPSVL13, BCM^{+}11a, BLPP16, BBC^{+}17, BLM^{+}18, BKC^{+}19, BDC10, BBH^{+}12, BLM^{+}12, BCB11, CKK^{+}10, CHE^{+}17, CPS^{+}14, DFS13, DBC^{+}12a, DM11, DFDC18, DCIG18, DBHMG12a, EMH^{+}19, FPC^{+}13, FLC^{+}14, FH18, FMJA^{+}18, GCSF18, GMR^{+}19, GTM14, HGL15, HBF^{+}10, HTSA16, HPMZ11, HCGP13, HLS^{+}15b, HT15, HHNI17, HFR^{+}13, HEM^{+}12, HBP^{+}16, HTC17, HT11b, JTH^{+}12, JCF^{+}12, KKMA^{+}10, KBR^{+}11, KHÁB^{+}17, KHF^{+}11, KTR^{+}16, KHT^{+}17, KCLP16, KBP^{+}10, LCGS^{+}19, LHT^{+}16, LMV^{+}19, LCS^{+}10, LSD^{+}13, MMH^{+}17, MVP^{+}16, MIIB16, MWZ^{+}16, MTP^{+}10, MJ10, MBRO10, MBS^{+}14, MHS14, MBG^{+}15, NGA18, NHTS14, OFDG14, PKS11a, PSA12, PSL^{+}12, PBCA10, PMF^{+}17, PLZL14, RMP^{+}19, RSS^{+}16, RTM11, RLAGL12, RFRD10, RMT^{+}13, SKB15, SKMS^{+}18, SAMW14, SRA^{+}17, SD17b, SGD^{+}12, SBS10]. analysis [TSTP13c, TSTP13b, UTS^{+}18, VSG^{+}11, VMC^{+}19, WWAS18, WDM^{+}19, WHI^{+}18, WF18, WWAL13, YMD^{+}10, YBB^{+}12, ZAC11, ZBB^{+}11, ZSSK11, ZKS17, dlSOV^{+}16, vdVFK^{+}10]. analytic [BFD^{+}16]. analytical [BFM^{+}13]. anammox [ETAE13]. ananas [PPDW16]. anatomical [FRGI15]. anatomy [TPM19]. anchoita [FKZTT18]. anchor [JROS17]. anchored [RTCF^{+}17]. anchoveta [CCG^{+}10]. anchovies [SGT^{+}14]. Anchovy [PAP^{+}12, ARB^{+}12, ASEB11, HPH^{+}18, NIS14, OTK^{+}19, QLCUG15, RNB^{+}12, RLN^{+}13, STK^{+}13, TGSM12, WB12, WM10, YHT16, YKT^{+}14, vdSRT^{+}18]. Ancient [ZBK^{+}14, TKL^{+}11, VMC^{+}19]. Andaman [SPJ^{+}12]. andesite [PML^{+}10]. Andreotti [INK^{+}17]. anemone [HDT^{+}12, LMP12, NC11, RCE^{+}13, RM19]. anemonefish [HTR19a, HDT^{+}12, LF14, SBB18]. anemones [BFPONS^{+}12, DMK11, HSS^{+}19, OC17, VJM^{+}16]. angasi [PSP^{+}19]. angiosperm [MORM15, SD17a]. anglerfish [CALB10]. anglers [APC^{+}14b]. anglica [CSN12]. Anguilla [CKK^{+}10, HHLT10, HLW^{+}16, HVD^{+}16, JT10a, TWY^{+}19, ZSSK11]. anguillicolosis [LC12]. anguillid [SMD^{+}16]. anguilliform [LFM^{+}16]. angular [NdBKR13]. angulosus [MvdHBM12]. angustirostris [AGNH^{+}19]. anhydrase [VJM^{+}16]. animal [CH15, NWB^{+}19, ST10b, SRK13, SSM^{+}19, VMN^{+}12, WPS^{+}19, ZJ17]. animal-borne [SSM^{+}19, WPS^{+}19]. animals [AvBBL^{+}16, CGG10a, CS13, KKMM16, NTS^{+}18]. Anisakis [KBK^{+}10]. Annelida [Mus12]. annotation [DBS^{+}16]. Annual [BGL^{+}11, CPF^{+}14, Chi11, GAB^{+}14, KKPL14, NWC^{+}10, TTW^{+}12, AJ16, ACG^{+}11, BDDvdM17, CRRL17, DGDJ14, DRH15, FTB14, KTP^{+}18, LMCD19, MPB14a, OKF^{+}18, PNPF14, RKM^{+}10, TDB^{+}14, WCG^{+}14, dJdBdJB12]. annularis [DZH^{+}11, EPL^{+}14, HIKK12, HI14]. annulata [HSS^{+}19, OC17]. anomalies [McC17, PBM^{+}10]. Anomalous [DBA17, FPBL16, RBL^{+}17]. anomaly [CBJB19, GUC^{+}15]. Anomura [BHP11]. anonymous [VBV^{+}10]. anophagefferens [TG12a]. anoxia [HSZR10, RZS12]. anoxic [AWJ12, KdEPW18, VPW^{+}12]. antagonistic [AS13]. Antarctic [CDCZ16, CDC18, NHSG11, RSC14, ZSEV17, ASR^{+}19, ARGB13, AAvS^{+}13, BS14a, BBH^{+}13, BLJP14, BCR15, CAR^{+}19, CLGCA^{+}19, CHP^{+}11, CDCZ16, CDC18, CBF^{+}19, DMW^{+}16, DHK^{+}17, EBOM11, EPRGO13, EZZ^{+}12, FTT^{+}17, FKT15, FAK^{+}12, FTS^{+}13, GSJS13, GSD12, GPW^{+}10, GMES18, GBD13, HPL^{+}16, HGFA^{+}11, HPR^{+}12, HBK^{+}12, JR15, KAD^{+}15, KTF10, KSK^{+}12, KPH^{+}15, Lap19, LRQ^{+}12, LKB^{+}14, MEDS11, MAT^{+}11, MBB^{+}12a, MWWS17, MTMM^{+}15, MBvO^{+}18, MSM^{+}12, NPRAGG^{+}12, OMBP10, OCMT13, PR11, PSL^{+}12, PHD^{+}13, QGM^{+}14, RWG15, RCS^{+}17, RWS^{+}15, RQN^{+}14, SDM^{+}13, SRLV10, SV13, SKD^{+}17b, SSM^{+}16a, SSM^{+}17, SGCF^{+}12, SLL^{+}18, SKR^{+}13, SLW15, SMR^{+}11b, SOT15, SA17, THP^{+}16, TDM16, TTM19, TMP^{+}13, UFP^{+}10, VVB^{+}10, VKY^{+}12, WVD^{+}17, YRM^{+}12, YKG^{+}15, ZAMB10, vdPLdV^{+}11]. Antarctica [GMSL^{+}14, PMIG19, REP18, ABJ^{+}15, BLM^{+}18, BKM^{+}18, BSA^{+}16, CBKM16, DHK^{+}17, DRH15, GMK18, HS14a, KHD19, KBK^{+}10, LOS^{+}19, LS15, SEFS15, TIN^{+}18, WRR10, SLL^{+}18, SLW15, ZXF^{+}16, vdPLdV^{+}11]. antarctica-dominated [SLL^{+}18]. antarcticum [BLJP14]. antenna [RBD^{+}14]. antennal [JBCGC19]. antennatus [FHM^{+}11]. Anthoathecatae [PCD^{+}18]. Anthometrina [HGFA^{+}11]. Anthopleura [LMP12, QSB16].

Anthozoa [BBG^{+}11, TH11]. anthropic [GP11a]. Anthropogenic [DHG11, TMB^{+}12, BFS10, BHP12, BWN^{+}14, COK10, GKR^{+}15, LRS^{+}12, MLB19, NGR^{+}17, OSPWB15, PBG^{+}10, RGdLB^{+}11, TCJ^{+}19, YTM^{+}12]. anthropogenically [BH12b, CSS^{+}17, CLS^{+}16, RCBE15, VSG^{+}11]. Anti [SPP15, BHASP18, BDT^{+}16, FWW^{+}16, MPLLE17]. Anti-bacterial [SPP15]. anti-barnacle [FWW^{+}16]. anti-herbivory [BDT^{+}16]. anti-parasite [BHASP18]. anti-predatory [MPLLE17]. antibacterial [SJO16]. antifouling [WSD^{+}10]. antigrazing [WRO^{+}11]. antillarum [BWH^{+}17, IHP10, KB14]. antimicrobial [WSD^{+}10]. antioxidant [PHD^{+}14]. Antipatharia [BBG^{+}11]. Antipathes [HYL^{+}13]. antipredator [CD14, MDM18, OF10, TOW10]. anxiety [CGH^{+}16]. aomoriensis [HSRP11]. apart [CMC^{+}16, OWH^{+}19]. apetala [FWW^{+}16]. apex [AB11, BAS^{+}10, CWW^{+}16b, FHRC16, IDT^{+}18, JSCM15, JWT^{+}14, RSSvV19, SKFH18, SPBS14]. aplanochytrids [DR10]. Aplysina [GEF^{+}12]. apodus [MDA^{+}10]. Aposematism [LCCT19]. Apostichopus [SGD^{+}12]. apparatus [GRCH^{+}18]. Apparent [HVK^{+}13, AC11, FPBL16, KP18, LGJH12, MM11]. appear [MSLR13]. appearance [KEBB10]. appendicularians [LLP^{+}10]. applicability [HPP^{+}10, Løk11]. applicable [MDS14]. Application [AMS^{+}13, CBYH13, PKSN11, BZC^{+}12, GBL^{+}15, HP11b, IMU17, KKMA^{+}10, ME11, RFRD10, SMBC18, SBKA14, WRP^{+}17]. applications [LS14, SFW^{+}16, SSM^{+}19]. applied [CKV12, GM10, OGGP14, SDC^{+}16]. applying [VBR^{+}14]. appraising [KNS^{+}15]. Approach [HGW^{+}11, TTC11, ASM^{+}12, ABN14, ARvdW^{+}16, BT16, CSG11, CLSS17, EM11, FNF^{+}18, FBC^{+}15, FC12, GMT^{+}17, GOM^{+}19, GBC^{+}16, GKRG^{+}19, HCGP13, HMM^{+}15, KdWV^{+}18, KM10a, KTR^{+}16, KW12, KDB^{+}10, LMRRB17, LRK^{+}15, MMH^{+}17, MKI^{+}15, MBC^{+}11, MPG^{+}12, Mus12, NHAS15, OLD^{+}15, PBH^{+}13, PCP^{+}15, PMB^{+}15, PDCH15, PCD^{+}18, PMM^{+}11, SDM^{+}13, SLS11, ST10b, STRP18, SJ10, SLPP^{+}14, SSS12, TLS^{+}19, TGRH17, TPSS12, Vig15]. approaches [CCS^{+}18, EAB^{+}19, Fry13a, MBCB^{+}19, Nob11, PMM^{+}15]. approaching [HT10, LAM10]. Appropriate [CSS^{+}13]. Aqaba [BAE^{+}10, GKBEF18, ZAZS11]. aquaculture [BDLW10, FCDC12, HU13, JCZ^{+}12, KKY15, MLF^{+}18, PBC11, SCM13, SFW^{+}16]. aquaria [OFPR^{+}11]. aquatic [AHSG16, BFRM13, BVC^{+}16, CWGG15, CS13, KEH10, KPLD^{+}12, LMTD16, MV12, PW15, Ree14, SLS11, VBR18]. Arab [BBB^{+}14]. Arabian [PQAH^{+}17]. Aragonite [LRH14]. araneus [SHP^{+}14, WAP10]. Arbacia [ZdVA^{+}16]. arborescent [EL19, GMM10]. arbuscula [LFPC^{+}14]. Arc [SJL^{+}15]. archaea [TKW^{+}14]. archetypes [MKT^{+}18]. Archipelago [FRC^{+}13b, PFVFF^{+}15, GLS^{+}15, HNK^{+}18, HS17, HIKK12]. architecture [HCK12, WDM^{+}19]. Archival [FLK12, ADHP12, MAW^{+}11, MDNL^{+}11, WV12]. Arctic [CLSS17, GLS^{+}15, SLRS13, ÅCAC16, AFMR^{+}13, AFRAG16, AGM^{+}11, AHSG16, BFS10, BVS^{+}16, BLV^{+}19, BRS10, BCSR10, BSH13, BDS^{+}18b, BRF11, CRS^{+}11, CPTL11, Chr17, CAG12, DKL^{+}18, DPY^{+}16, DLRF15, DGMB10, DMC^{+}17, EMG^{+}15, FLK12, GMT^{+}15, GMT^{+}17, GGL^{+}17, GBH^{+}10, GCS^{+}18, GCS^{+}19, GAB^{+}14, GWK^{+}12, Hal15, HLIK16, HKI^{+}17, HKL19, HHS^{+}12a, HPN^{+}17, HLPS^{+}15, HG10, HCLB18, HTR^{+}19b, HB15a, JPI^{+}19, JKM11, JPAM^{+}15, KHW^{+}10, KBW^{+}11, KFBM19, KNB^{+}13, LSY^{+}18, LGB^{+}13, MBR^{+}13, MWA17, MTG^{+}10, MFD12, MBS^{+}18, MADI13, MKT^{+}18, OVF11, PLG^{+}19, PHH^{+}12, PSS^{+}12, PMF^{+}17, PLF^{+}12a, PGOG12, PJMA18, RZH^{+}16b, RNH15, SBH^{+}19, SHP^{+}14, SdBAH^{+}14, SRER16, STP13, STBP13, SNK^{+}10, SKN^{+}15, SKR^{+}10, SDMR19, SLW15, SBM^{+}18, SLNG18, TAAN15, TRL^{+}15, TJS17, TTS17, TBS^{+}17, VCMS^{+}18, WKX^{+}19, WCG^{+}14, WKRW^{+}12]. Arctic [WEK14]. Arctic-boreal [PSS^{+}12]. Arctica [PPML17, BNAN12, FRMS17]. Arcto [DH17]. Arcto-boreal [DH17]. Arctocephalus [FHMG19, MMB^{+}19, VCG^{+}15]. Arctodiaptomus [JMGG13]. Area [NHTS14, PDM12, ABH18, AFSL^{+}15, ABB^{+}14, BCG10, BMV^{+}17, BFD^{+}16, BCM^{+}11a, BP12a, BD19, BAFB12, BDH^{+}17, CGK13b, Dao13, DKF^{+}15, DMPG11, FCRD18, HVBG^{+}19, KCS^{+}17, LBM^{+}13, LNF^{+}11, MUHC11, MRC10a, MWB13, MUOS13, MJA^{+}18, NMS^{+}18, NCW^{+}19, PSdA^{+}17, PJMP10, QHG^{+}19, RLN17, SPHK17, SFM^{+}19, SHB^{+}19, SDB12, SLPP^{+}14, TBB^{+}19, VGGF12, WEAG^{+}15, YFC^{+}13]. Area-restricted [PDM12, FCRD18]. Areas [LBT^{+}10, MC16, AN11b, ALS18, BDM^{+}13, BBC^{+}17, CEH18, CMC10b, CHRB10, CHB11, CBGDS12, CME^{+}12, CVM^{+}17, DvV^{+}16, DMH15, DMT^{+}14, DBHMG12a, FNC13, FTWA17, FGW^{+}15, GSF18, GAB^{+}14, GCRP15, GMES18, HS11b, HIH^{+}17, HFH10, JW10, JFK12, JBS12, KBG12, KLBT^{+}16, KRR^{+}18, LKC^{+}17, LRHP18, LMLB12, LCBK^{+}13, LBS^{+}15, MM13a, MADI13, MHFN10, MHSG18, MSS^{+}14a, MDKM13, MS11, MLSK^{+}15, NUO^{+}19, PCP^{+}15, PSHG16, RKD18, REAN^{+}14, RKM^{+}10, SKD17a, SMD^{+}16, TRSV^{+}12, TM10, TCTB11, TMH^{+}18, TUK13, VHV13, VMN^{+}12, WYI^{+}16, WPF^{+}16, WTdB16, WTGG10, YTO^{+}11, dJH11, DHB18, GPA^{+}16, LCR^{+}16]. arena [CWGG15]. arenaria [CLG19a, CBC^{+}16, GSO^{+}18]. arenarius [WLBB^{+}13]. Arenicola [AOW11, EBK^{+}12]. Ares [PLFR^{+}11]. Ares-Betanzos [PLFR^{+}11]. Argentina [GZYS11, RUS^{+}13, SWFQ12]. Argentine [BGIA19, SLD^{+}13]. argentiventris [TEAO^{+}14]. Argopecten [HW11b, WMMC14]. Argos [DBHMG12a]. argument [GNJJ^{+}17]. argus [AB13, BFEMLÁJ19, GFC^{+}14, KPS14, MD11, PPDW16]. Argyrosomus [NCD^{+}12]. Ariake [YMMO14]. arid [WAMS^{+}14]. Aristeus [FHM^{+}11]. aristotelis [LANE18]. arm [BBW^{+}19, Bus12]. armiger [BH16b]. armoring [MCT15]. armouring [BGO^{+}18]. arrays [MOMF14, RBD^{+}14, TBP11]. arrival [RB18, ZBY^{+}14]. arriving [VK11]. arrowtooth [ZAC11]. artefacts [Bet13, War14]. Arthropod [BT14]. Artificial [CRHV14, JOFP18, MHQ17, WTP^{+}18, BJ14, BC14, CSS15, DPS16, DHJ10, LL18, LÁBFÁF^{+}10, MBC18, ME11, MRC10a, PTP^{+}19, PK15, RB17, RPL^{+}18, SRR15, SJC10, WGS16]. artificially [SKFY18]. artisanal [AMR^{+}10]. asbestinum [PMH^{+}13].

Ascension [OWW^{+}17]. ascent [DHSL11]. ascidian [BSRDR10, LS15, MDLM17, MC10b, PPTB12, RMKR13, SF14, YAFV12]. ascidians [DS16, OHP18, RBGT10, RCF17]. Ascophyllum [BAVSP11, SCRL11]. Asexual [SMJ^{+}14, CG17]. Asia [TCWC13, WYI^{+}16]. Asian [DHGW15, GVGH15, JMB^{+}18, MKK13]. aspect [GHCC15]. aspects [MG18, SNT^{+}16]. Aspidochirotida [EBdlTC12]. Aspiring [TPC^{+}15]. assays [BCW19]. assemblage [AN11a, BRC^{+}15, CGR^{+}15, CGEGC^{+}19, DMS14, DBS14, GSMK15, GSTK11, HPLS13, KSB11, LYH^{+}17, LDT^{+}13, NSI^{+}15, OMT^{+}17, PTS^{+}18, PCH11, PBC^{+}12a, PLF^{+}12a, PCNM17, RH12, SNW^{+}18, Sma13, SDF^{+}19, SJHH^{+}17, SLR^{+}14, TBH^{+}18, THW^{+}16, VDM^{+}10, ZBK^{+}14]. assemblages [ATS14, AAIB15, ABB^{+}14, AAGGG19, AHB^{+}14, AOW11, AFH11, BHF^{+}10, BWS^{+}12, BH16a, BCW12, BDPD14, BHT15, BAM^{+}10, BMVBC10, BAVSP11, BRA^{+}10, BHDC15, BSG^{+}17, BKHW13, BBC14, BTM^{+}18, BAvW12, CLD^{+}13, CW12, CBR^{+}18, CGK13a, CMO^{+}10, CMAJ12, CC12, CH12, CATAV18, CSN12, DFGH^{+}15, DMJ11, DRDV15, DDC^{+}15, EZZ^{+}12, Fis10, FHL^{+}15, FJW14, FSDS10, FSS^{+}18, FLSG^{+}15, Fuj16, GWT18, GKC11, GKBEF18, GOC^{+}18, GGCB16, GC13a, GSU12, GCCCM16, HLS15a, HFR^{+}18, HOL^{+}17, HBM^{+}12, HAF^{+}14, HWB^{+}18, JDB^{+}10, JGL12, KdWV^{+}18, KMPS15, KPB16, KHG^{+}19, LLBT18, LN10, MK13, MSFM12, MLPC18, MJR^{+}14, MHGW13, MHQ17, MF13, MPB15b, MVOPB^{+}16, MLR12, MCTHGV12, MKT^{+}18, OCPM12, OSPWB14, Par15, POPH12, PS15a, PMB^{+}10, PA13, RKD18, RWMP12, RWO13, RCGE12, ROLA15, RBD^{+}16, RCP^{+}13]. assemblages [SCD^{+}17, SJC10, She12, She13, SBCA15, SLT^{+}19, SLO^{+}17, SW12, SLPP^{+}14, SKS18, TS10, TS11a, TAB^{+}14, TT16, TMP^{+}13, TPCN12, VBT14, VJG^{+}13, VHKT15, WFBM14, WH11, WLM^{+}13]. Assembly [GGF^{+}12a, AAGGG19, GA11, PTL16, RMBM14, VVB19]. assess [BLIM11, COK10, DE17, DP14b, KPP^{+}16, LR18, LPK^{+}16, MMH^{+}17, NSC^{+}19, PL10, SMVM^{+}12, VGGF12, dJSB19]. assessed [AGRPSVL13, CMIMGS15, MASW^{+}11, MDA^{+}10, MMFP^{+}15, PFB^{+}17, Pro17, TBP11]. Assessing [BWN^{+}14, FMJA^{+}18, HBK^{+}17, KTR^{+}16, LBCS^{+}15, MS14, MGCH^{+}16, MWMD17, PLM^{+}16, PSdA^{+}17, PPC^{+}19, SDSA14, WTGG10, ZVR18, BBC14, FBC^{+}15, FPL^{+}12, GFC^{+}14, LMRRB17, PJMP10, PLF^{+}12a]. Assessment [GKRG^{+}19, HSH18, NWD13, SMOH19, AAE14, BT16, BSB^{+}11b, CWR^{+}18, CSPA10, DPS16, DL13, ET11, FRJ10, FRJ11, INK^{+}17, JS12a, JLPG17, LALCM18, MF15, MBW^{+}19, MWB13, OGGP14, RNFG10, RCF17, SLKN14, SBL^{+}10, TSTP13a, TMH^{+}13, TNS^{+}12, WFH11]. assessments [AHSG16, BMVAS19, MPRPM^{+}18, PMB^{+}15, WMCC15]. assign [SWK19]. assignment [Coo11]. assimilate [CKK^{+}10]. assimilates [RCB^{+}10]. Assimilation [FF12, FSE^{+}16, HGPM10, KCOP13, MWP10, PWO^{+}16, PBWM18]. associate [SMVC^{+}15, SPH^{+}13]. associated [AFC^{+}16, BCW12, BCvH^{+}15, BHT15, BAVSP11, BRA^{+}10, BPLT14, Col10a, CM11, CSRS15, DRG^{+}12, DBCS16, DEWK15, EMG^{+}15, EMH^{+}19, FCFD15, FFS^{+}15, HKB^{+}19a, HFM^{+}15, HSdBS10, HS14b, HIKK12, KMPS15, KHG^{+}19, KW12, LEB14b, MMG^{+}16, MA14, MGGSP15, MBP^{+}12, MHNL14, MF13, MPB15b, MLP^{+}13, MRC^{+}19, OHST19, OGA13, PMK^{+}15, PBV^{+}11, PWCvO^{+}14, PHHL^{+}11, RMB^{+}14, RU11, RSP17, RdGv^{+}18, ROLA15, RNSA15, SFLG10, SSM^{+}16a, SL11a, SLIGG18, SCWV12, SKMS^{+}18, SCM13, SWBMJ17, SGT12, SBH^{+}18, SO15, SJO16, SLNG18, SBW16, TBC^{+}17, TTB^{+}11, TKL^{+}11, TVRG^{+}13, VB19b, WK10, WH14, WvLF18, WJW^{+}18, ZBB^{+}11]. associates [HD10, MHCRR18]. Association [DR10, SGC17, FPC^{+}12, KBS16, LHPP16, LP12, Lyo14, RSO^{+}13, RA10, SRLV10, TNB^{+}15, ZRBV15, dG14]. associational [Goc10]. Associations [BLE^{+}13, CD19, KPKA13, LJB14, PHM11, AN11a, AEGH11, BMEL11, CLB^{+}14, CWS13, CDCS12, CMLC12, DR16, DH18b, DAM^{+}17, ESBG14, FMP13, Fis10, GSF18, GKR^{+}15, HLPS^{+}15, KBK^{+}15, LHJH^{+}10, LET^{+}12, NSK^{+}14, PMSGBL16, SHK14, SCO^{+}12b, SMJBØ11, TBH^{+}18, TDKS10, WRFM16]. Associative [RTCF^{+}17]. Assortative [GGD^{+}11]. assumptions [DSG^{+}13]. Astarte [DMC^{+}17, GMT^{+}17]. Asterias [ATP^{+}12]. Asterionellopsis [GSA^{+}18]. Asteroidea [BLD^{+}16b]. asteroids [WHB^{+}14]. Asteroplanus [YMMO14]. astreoides [CSJP17, Edm10, PKW^{+}11]. Asymmetric [DDDR10, PP18, MPF^{+}15]. asymmetrical [AN11b, BT16, JSMO14, NDD^{+}13]. asymmetry [PLMN10]. asynchrony [OBM15]. At-sea [BBH^{+}13, HWR^{+}14, LMSC^{+}19, ZHM^{+}10, ARW12, LBVH13, PHHL19, RHW12]. Atherestes [ZAC11]. Atlantic [BS18b, CC19, DCHI12, GHT^{+}15a, HGB^{+}13, KLD18, LABM17, MAB^{+}10, RNFG10, SBCT17, SDF^{+}19, SGB^{+}10, TMH^{+}13, VWW^{+}16, tHHR10, AED^{+}10, AFOVRB14, ALNB17, ACG12, AHH^{+}11, AJ16, BVM^{+}19, BDM^{+}13, BNAN12, BWLP17, Bea11a, Bea11b, BSN^{+}18, BPWF14, BCM^{+}11a, BGM^{+}18, BDH^{+}17, BAW11, BS18a, BDC10, BKAM12, CMP^{+}11, CFF^{+}18, CWS13, CAG13, CDR12, CLMM^{+}14, CBF^{+}13, CBK^{+}12, CMH14, CHLR16, DSS10, DSG15, DMMW12, DKAC^{+}12, DBG15, DM17, DZB^{+}18, DDP19, DAD^{+}11, DFAH11, EOMS15, ECC^{+}15, FMP13, FSD^{+}14, FWPP12, FJ12, FRMS17, FSP^{+}17, FTM^{+}13, GDB^{+}16, GN11, GM13a, GLR^{+}14, GBL^{+}15, GME^{+}19, GLM16, GPM^{+}13, GSM^{+}11a, GTM14, GCR16, GLM^{+}19, HGG^{+}18, HWET13, HS11a, HBB^{+}18, HMO^{+}12, HRS^{+}17, HW11b, HWE^{+}12, HDB^{+}10, HGA^{+}16b, HGA^{+}16c, HRC12, HBP^{+}16, JMK12, JP15, JBS12, JJF14, KAAI15]. Atlantic [KPKA13, KR16, KRR^{+}18, KHM^{+}12, KBB^{+}10a, KSK^{+}12, KBC^{+}12, LBN11, LCB10, LGBSC11, LRR^{+}10, LC12, LMB^{+}19, LNP12, LCBK^{+}13, LMSC^{+}19, LCL^{+}17, LD17b, LBCS^{+}15, LFRF13, LA12b, LBCL10, LPB^{+}16, MFG^{+}13, MMDM12, MVTS^{+}12, MAD17, MBR^{+}18, MSH^{+}14, MGGSP15, MW16b, MMFP^{+}15, MVOPB^{+}16, MMNR^{+}12, MRC10b, MO13, MLSK^{+}15, NBGS^{+}12, NJJ^{+}11, NSK^{+}14, OBA^{+}18, OSD^{+}13, PPML17, PNI14, PXG^{+}10, PBAS^{+}10, PSÓ^{+}15, PPH11, PFVFF^{+}15, PLA^{+}12, PSS19, PHHL^{+}11, PB19, PRK^{+}12, QHN^{+}10, RMKR13, RBA^{+}18, RWMP12, RSDN15, RZH16a, RVPG17, RAN^{+}10, RTJ^{+}13, RJH14, RSP^{+}16, RAF^{+}14, RGDNGS11, RFRD10, SBV^{+}15, SSBA12, SNS^{+}13, SBB^{+}12a, SHF12a, SML^{+}14, STS^{+}18, SJ10, SG18b, SCHO11a, SCHO11b, SHM10b, SST^{+}18, SB15b, TNB^{+}15, TFB^{+}17, TPT^{+}12, TPM19, TVM16, VCL^{+}15, VHP^{+}11, WHB^{+}14, WBW^{+}14, WWAS18, WHH^{+}13, WFB16]. Atlantic [WNA^{+}15, WL19, WHJ^{+}12, WFP15, ZMGS17, dQSC13]. atlantica [BLHL15]. atlantid [WPML^{+}18]. atlas [MMWG13a]. atmospheric [GCP^{+}17, GLY^{+}14, LERS13, TZF11, THRBL^{+}16]. Atol [GDJdBSL19]. Atoll [HSPS14, RL11, TBG^{+}17, EC14, FSDS10, TLAT13]. atolls [LCM^{+}16]. ATPase [PZL^{+}17]. Atractoscion [SCSC16]. atresia [SNK^{+}10]. attachment [BA13a, EKOÅ13, GRD11, MWR^{+}12, ZGH^{+}17]. attack [BKR^{+}15, BFG^{+}11b, vBBW^{+}14]. attacks [RLH^{+}19]. attendance [CRJ^{+}18, GZYS11, LAH^{+}15]. attenuate [DTC^{+}17]. attenuation [CSS^{+}17, KSS13, PBA12]. attitude [MM10]. attracted [FG16a]. attraction [CRJ^{+}18, RKFD15]. attractive [DP14a]. attractors [SAM14]. attracts [MMD^{+}13]. attributes [BDLW10, TPVCV19]. Audouin [CBC^{+}13]. auklet [SHD^{+}19]. auklets [BCH^{+}11, CEC^{+}18, DKB^{+}12, TIS^{+}15, WJW^{+}18]. auks [FCH^{+}10, GWH^{+}18, GWK^{+}12, HWR^{+}14, JGWJ^{+}11, JTWJ^{+}13, KHW^{+}10, MJA^{+}18, PRM17]. aurata [MVCD14, MMB^{+}12a]. auratus [CMLC12, LMC14, MJHS11, MJHS13]. Aurelia [CL17, CGC15, CG17, FWS^{+}18, FSM^{+}14, GJSR18, HSS18, HMBP18, HFN14, IMU17, SNT^{+}18, TPKC18]. Aureococcus [TG12a]. auretenra [MWP10]. aurifrons [HPL12]. aurita [FSM^{+}14, GJSR18, HMBP18, HFN14, IMU17, SNT^{+}18, TPKC18]. auritus [PRLR^{+}18]. austral [SOT15, VKY^{+}12, EZZ^{+}12]. Australasian [ABA16, CVW^{+}11, CB17]. australasicus [ST10a]. Australia [RUP11b, BPZ^{+}14, BDF^{+}11, BHL^{+}19, BSH^{+}14, CPA^{+}11, CAB^{+}19b, CBL^{+}10c, EPF^{+}17, GHGF12, GMP^{+}11, GGCB16, HCF^{+}16, HSMB17, HvEBC12, JRW^{+}14, JMB^{+}17, JBD10, KHN15, LAM11, LRHP18, LPK^{+}16, LDP^{+}17, MJW19, MVP^{+}16, MMC^{+}18, MMCL19, PS15a, PBC^{+}12a, RUP11a, SRSH^{+}18, SSB^{+}18, SDCC12, SCLB11, TKS^{+}14, THB^{+}12, TPA^{+}13, WFBM14]. australiae [GRN^{+}16]. Australian [LLS^{+}18, ACG^{+}11, BPB^{+}12, BCM11b, BBMM14, CSN12, DMR10, ETD^{+}18, FGOS16, FSR17, GFKC13, HMW^{+}12, HA14, HCW^{+}15, KBA17, LHHG11, ME11, MHAP18, OWN^{+}15, RABS19, RLH^{+}19, SCMSB^{+}11, SMSE19, SHK14, SYM^{+}16, STSS^{+}16, SBW16, ZPPM17, dlSBO^{+}12]. australis [CHG^{+}18, HCK12, JSOK18, LBM^{+}18, MRB^{+}11, MBP^{+}12, RA10, RGWA11, RUS^{+}13, SKA^{+}14, VRSS18, VCG^{+}15]. Austrohelice [NPLT10]. Austrovenus [NDBL16]. Author [Ano10a, Ano10b, Ano10c, Ano12a, Ano12b, Ano13a, Ano13b, Ano14a]. autocorrelation [KWDD10]. automated [BOS16, YChH11]. autonomous [ETD^{+}18, HBM^{+}15, RBD^{+}14, WLM12]. autotomization [HG17]. autotomy [Bus12]. Autotrophic [SKR^{+}10, WS10, CL15b, SVIH^{+}10, VAHL11]. autumn [FTS^{+}13, HBF^{+}10, PPH11]. Auxis [LCQT^{+}13]. availability [ACDJ11, AvDL^{+}19, AAIB15, BRS10, BKAM12, BKAS14, CDWC17, CPG^{+}14, CEK^{+}16, DMM^{+}10, DKB^{+}12, FJ11, FA10a, FOJF15, GE14, GGPDS19, Goy14, GCRP15, HJN^{+}10, HP11a, HJMN^{+}12, KLN^{+}13, LÁBFÁF^{+}10, MH10, MC12b, MGGSP15, MJHS13, NKH12, ODP^{+}12, OFT^{+}11, PRA10, RLN^{+}13, RDM18, SNW^{+}18, SUB^{+}13, SQRDLC18, SSAB^{+}15, TMS^{+}16, TGM^{+}17, TR13, VSB^{+}15, VMWU15, WSF^{+}12, WAW^{+}12, WNNH10, WKK18, WL11, YKT^{+}14, vdMPRM12]. availability-mediated [YKT^{+}14]. Available [TLB15, BB14a, ØLH^{+}13, RBSL11, ZKS17]. avian [AL12b, DGMB10, KCH^{+}11, LCL^{+}13, LCL^{+}17, VHBG15]. Avicennia [DKV^{+}16, GZB16]. avicularia [CGG10a]. Avoidance [Mar17, BWH^{+}17, CMGS19, Cra12, CSV^{+}15, KSA12, LPJ^{+}13, MJM^{+}14, MCT16, NWH^{+}12, Rit17, ZAMB10]. avoided [MD16b]. Avoiding [LWM15, UHG12]. away [DZL16]. axial [BGDC14]. Axinella [HCM^{+}11]. azarae [PCA^{+}15]. Azores [CSABH^{+}13, NFN^{+}15, NBGS^{+}12, VHP^{+}11].

B [LRTLF12, GGLT15, HGB^{+}13, SB15a, BLHRP^{+}19]. Ba [BAH^{+}15, MC10a]. Ba/Ca [BAH^{+}15]. back [HU13, HBM^{+}12, LEB^{+}19]. back-reef [HU13, LEB^{+}19]. back-tracking [HBM^{+}12]. backdrop [FEA^{+}15]. backed [TRSB^{+}18]. background [MAP^{+}12]. backs [PO13]. backscatter [EFOM18]. backscattering [MVTS^{+}12]. Bacteria [BAFP^{+}13, MSLR13, SJO16, AOW11, BHMCM16, CESAM12, CLMM^{+}14, GBU15, GLWI10, GGF^{+}14, IVD^{+}10, KLD18, LMR^{+}12, LRSB19, MGFÁS^{+}10, MVTS^{+}12, MDE13, OAGV11, PDBG14, PDGL17, TKW^{+}14, TH11, VARO12, VCQM^{+}16]. Bacterial [CR19, GRD11, MLP^{+}13, SWR11, SO15, VCMS^{+}18, YYD^{+}15, AGTS^{+}12, AOW11, BTM^{+}18, CMM^{+}18, DZH^{+}11, GEF^{+}12, HCM^{+}11, LKL^{+}19, LLRC^{+}12, NWB^{+}19, PSL13, PRP^{+}16, RP13, RRG^{+}11, RKPB^{+}15, SPP15, SO16, TFÁS^{+}12]. Bacterioplankton [THRBL^{+}16, GBW^{+}12, RNK^{+}11, TMGCM11, THMF^{+}15]. Bacterivorous [HSR^{+}10, SP13a]. Báez [dQSC14]. Bahama [DZDB10, MBB^{+}15]. Bahamas [GCSF18, KM14, LA12a, SLF^{+}19]. Bahamian [Alb15, CM10, Dav19, GAC11, SBB17]. Bairdiella [SL15b]. Bait [BG12, WLM^{+}13, WD14]. Baited [KWC^{+}17, YD11, BSB^{+}18, CS10a, ETHB17, WFBM14, WLM^{+}13]. Baited-camera [YD11]. Baiting [BM10b]. Baja [LLCM^{+}12, TKEV^{+}19]. Balaena [FDLL10, PLF^{+}12a]. Balaenoptera [SRP^{+}19, TCSA17, dB10]. balance [BW16b, BW16a, dGCF^{+}17, TVH^{+}10]. Balanced [PMB^{+}17]. Balancing [PGVF14]. balanoides [BJRH10, FSR10, GPS13, HWVM19, SM17b]. Balanus [BSK^{+}11, NC14, ORG19]. Balearic [LCQT^{+}13]. Baleen [SMJBØ11, BMV^{+}11, CSV^{+}15, GLR^{+}14, LHJH^{+}10, MF15, RMT^{+}13, SRLV10, TBB^{+}19, VHP^{+}11]. Balgzand [vdVKA^{+}11]. Balistes [BSB^{+}18, BFN^{+}19]. ballast [DMR10, SWR11, VK11]. balthica [BDD17].

Baltic [BPM^{+}18a, AOW12, AWJ12, BBD^{+}14, BMV19, BSST13, BPSG11, BBG^{+}13, CBMK10, DAL^{+}10, DMvB10, DvBH12, EBH16, EKOÅ13, FQ11, HGL15, HB13, HMH^{+}12, HHB^{+}14, JNN12, KKL13, KOPW15, LGB^{+}13, MBC^{+}11, MFG16, OALK14, OTK^{+}13, PSPR18, PDCH15, RLTL14, RJTS12, SHKR15, SBH^{+}18, SHSK11, TSS11a, TSS11b, UDF^{+}18, VLK^{+}15, WL10, WNG^{+}19]. Baluarte [MVVGAV10]. Banc [CCA^{+}14, vdGvdLvG^{+}19]. band [JJNP12, NHPW18]. banded [GKW10]. bands [BdSC^{+}14]. Bank [DHG16, RJH14, YIKG10, KKL13, dB10, ASF13, BCSB13, DGDJ14, DZDB10, LFG^{+}12, LA14, MBB^{+}15, PSA12, STT14, SCL13, TPM19]. banks [GA11, PÁIIS^{+}19, SFL^{+}18]. banksii [BMMS10]. Barau [LBJ^{+}16, PJP^{+}11]. baraui [LBJ^{+}16, PJP^{+}11]. Barbados [KCL^{+}11]. Barbara [CSRD11, MPB13]. barbatus [BCH^{+}18, GWK^{+}19, HLF^{+}19, KAP^{+}12]. barcoding [BVD^{+}14, CGM^{+}13]. Bardiella [SGP^{+}17]. bare [AED^{+}10, BCW12, BDM^{+}14, HPQ15, MRC^{+}12]. Barents [HKL15, DVSM10, DJL^{+}16, EBDB18, JLM^{+}12, ØLH^{+}13, SMJBØ11, WAC^{+}14]. barium [WWG12]. Barkley [HLS17]. Barnacle [RCMB18, ARMA15, BBF^{+}19, BGP^{+}16, BSK^{+}11, DFPH11, DZL16, FWW^{+}16, FSR10, FC19, FTF15, GHCC15, GJ10, GPS13, HWVM19, KF13a, LAM10, LAM13, NC14, ORG19, SM17b, TCWC13, VGL16]. barnacles [BMTF10, LCD11, PBW^{+}11, SM15]. baroli [NBGS^{+}12, NBGS^{+}12]. Barolo [NBGS^{+}12]. barracuda [DSC13, DSC13]. barred [TWL15]. barren [ESW14]. barrens [FDS14, FJL12, HCG^{+}10]. barrens-forming [HCG^{+}10]. Barrier [BP12a, BB11, BKB11, BSB^{+}11b, BDF^{+}11, BK14, CED^{+}19, ETB13, ETAE13, GGCB16, HSMB17, HPLS13, KHH18, LKDP17, LPC11b, RUP11b, RUP11a, SGG^{+}12, SWB^{+}10, Sti13a, TPB^{+}16, TPB^{+}18, WSC13, ARMA15, MSV^{+}11]. Barriers [YBH^{+}11, DBS^{+}13, FHM^{+}11, TPBM13]. Barrow [SOC13]. Bartholomea [HSS^{+}19, OC17]. bartramii [IMS^{+}11, VSW^{+}15]. basalt [PML^{+}10]. base [ATP^{+}12, SDKS19, VWH18, WBJ12]. based [ASM^{+}12, ARB^{+}12, AFSL^{+}15, AAJJ^{+}19, AHD12, Ano18a, AWH19, BAL14b, BTS^{+}14, BLIM11, BMVAS19, BBG^{+}16, CMV^{+}11, CBGDS12, CMW^{+}15, DSG15, EM11, FL10, FC10, FUOG^{+}16, FHMG19, GSR10, GKBEF18, GWH^{+}18, HHNI17, HR19, HBP^{+}16, HS15b, HOE^{+}12, IS16, Ji11, KMP14, KCLP16, LNS10, LMRRB17, LCB^{+}16, MRLD19, MSG^{+}13, MW16a, MCL^{+}17, NMPM15, NWS^{+}14, PDO13, PB19, PCLB18, RWM18, RWM^{+}10, SOB16, SKSH19, SCB^{+}14, SKR^{+}13, SLKN14, SFM^{+}19, ST10b, SS13, STRP18, SSM^{+}16b, SJ10, SSFB11, SAD^{+}16, SPP18, TAEB^{+}11, TMT^{+}12, TP16, TRJ^{+}11, VBR^{+}14, VMC^{+}19, WWB14, WY15, Wir14, WEK14, ZFRG12, ZSSK11, ZB16, vDJF15]. baseline [DRWQ12, GKBEF18, MBKM17, SWK19]. baselines [MSF^{+}19, MDKM13, VWEP18]. basic [SDC^{+}16]. Basin [MSH^{+}14, RWP^{+}14, DRT13, KdEPW18, LKL^{+}17, SSB^{+}16, ZKS17]. basins [CBYH13]. basis [BD13, DBC^{+}12a, MCR^{+}13, SWP^{+}13]. Basking [WHJ^{+}12, CQR^{+}17]. bass [CKM^{+}16, CPS15, FMP13, GFS15, MH10, TWL15]. bassanus [SBB^{+}12b]. Batch [NRF11, JMO^{+}19, YKT^{+}14]. batch-tagging [JMO^{+}19]. Batesian [GWGC16]. Bathurst [STP13, STBP13]. bathyal [GMT^{+}11]. Bathyarca [GMT^{+}15]. Bathymetric [GN11, CLB^{+}14, TRLS^{+}11, TFB^{+}17, WHB^{+}14]. bathymetry [UW18, YIKG10]. bathymetry-derived [YIKG10]. Bathymodiolinae [FC13]. bathymodioline [RT17]. Bathymodiolus [RCF16]. bathypelagic [BHT10, UMF^{+}10]. Bathytoshia [PFKB18]. Batoid [MHCS18, FABH^{+}16]. Bay [BSG^{+}10, BWP^{+}11, BHN^{+}12, CRWG16, DWD^{+}18, FWS^{+}18, HFMD16, HFR^{+}18, MSH^{+}14, MMB^{+}19, NOT^{+}11, RLO18, RCD^{+}13, RLH^{+}13, SPH^{+}17, SMP10, SS11a, SMSB19, SMR10, SKD^{+}11b, TRL^{+}15, THB^{+}12, TT16, WYS^{+}10, WLBB^{+}13, WMB^{+}14, dlGMD^{+}17, ARB^{+}12, ASI15, BFH13b, GKE^{+}19, GLB^{+}10, HW11b, HNMK15, JCZ^{+}12, LMV^{+}19, LLH^{+}11, LDT^{+}13, MBJ16, OMH12, OMM^{+}12, OLM^{+}12, PA15, PS15b, RWM^{+}12, TPC^{+}13, TPC^{+}15, WMMC14, ÁMLUN16, AHH^{+}11, BI15, BBGL13, CR16, CB11, CSF13, DMd14, DSK^{+}10, GE14, GSO^{+}18, HJMN^{+}12, HMAC^{+}11, IAV^{+}10, JHL^{+}19, JL12, KVR^{+}18, KDS^{+}12, LGB17, LKCH12, LOBW^{+}11, LS16, LBCL10, MSC13, MH10, NKH^{+}10, NGA18, ORR^{+}14, PDGR^{+}13, PLG^{+}16, PW15, PS19, PFPP18, PRR^{+}16, RWM18, RSO^{+}13, RCB12, RHRB^{+}17, SL15b, SBE^{+}19, TR13, TNM^{+}11b, WMH17, WLB^{+}11, ZOM^{+}10]. Bay-scale [DMd14]. Bayesian [ASM^{+}12, Bre14, GED^{+}14, PCP^{+}15, SAMW14, USR^{+}16, YChH11]. Bayesian-based [ASM^{+}12]. Bays [RHH^{+}15, TCKY19]. be [HCL10, KBS16, McD13, RTMJ11, RUPT^{+}11, SMSL17, SCHO11b, dJSB19]. Beach [RPE^{+}13, BSG^{+}17, CGD10, CD16a, DM11, FRP^{+}17, LNH^{+}17, MAD18, MHPvdH19, OCDD16, PHM11, RHMR18, RTDUM16, ROLA15, SMM^{+}15, TBH10, TT16]. beaches [APH^{+}10, BGL^{+}11, BSG^{+}17, MPB14a, MAD17, MFDW17, MBT14, QNG^{+}15, RHMR18, ST13]. bead [TBP11]. bead-arrays [TBP11]. beaked [BPZ^{+}14, PBB^{+}15, RVPG17, WWB^{+}17]. beaks [CBR19]. bear [BDR19, DAW^{+}11, JHL^{+}19, ORE^{+}17, dlGMD^{+}17, SDF^{+}19]. bearded [BCH^{+}18, GWK^{+}19, HLF^{+}19, KAP^{+}12, SUPCB11]. bearing [MWBMH13, RUP11b, RUP11a]. bears [FKA^{+}12]. Beaufort [BBI16, BDR19, BGS^{+}16, CRRL17, CMC^{+}15, DIB15, HIH^{+}17, JPMG10, LKL^{+}19, ORE^{+}17, SSM^{+}16b]. Beaugrand [AF14]. Beauty [MBC^{+}11]. Becker [HGS15]. becomes [HG10]. becoming [MFSRR14]. bed [AY10, ASK^{+}15, BRA^{+}10, BH12a, FSLG12, FMC^{+}10, HNMK15, JW11, JPI^{+}19, LPSK17, OSK15, PFVFF^{+}15, PMYH13, RK15, VBT14, WO11, vdGSE^{+}14]. bedload [VNN10]. bedrock [FFLS16]. beds [ATT^{+}16, AL12b, BMV^{+}17, BBR12, CEW15, DEWK15, FS14, FDS12, GBK^{+}14, HR12, KW12, KS11, KS12b, LGS10, LHC^{+}15, MMPH^{+}12, MHCRR18, MLB18, MG18, MAL15, OMD11, PAG19, PRA10, SCK^{+}17, SRK13, SBLM14, SH14, TNIH12]. before [GFGGDSM12]. begins [HJ16]. Behavior [BSB^{+}18, BPG^{+}11, ADF^{+}18, BLS^{+}11, BHASP18, BBKH^{+}11, BBG12, BBBNT13, Ber12b, BKR^{+}15, BFL^{+}11, BSA^{+}16, BJS^{+}16, BGM^{+}17, CLB19, CTZ^{+}16, CGC15, CL16, CZH^{+}13, CBK^{+}12, DFPH11, DT15, DBCP15, DWM15, EC14, FTT^{+}17, FSLG12, FMG^{+}18, FCRD18, FTS^{+}13, GLF^{+}17, GP11b, GMK18, Goy14, GTM14, HLPS^{+}15, HTR19a, HGC15, HJ16, HR15, HNG^{+}11, KBW^{+}11, KTP^{+}18, KPS14, LCB10, LD17a, LBWB^{+}13, MH12a, MRLD19, MKI^{+}15, MBG^{+}15, MF10, MCT16, NC14, NCK^{+}14, OAB^{+}19, PFL13, PAK15, PVYS16, RCD^{+}13, Rit17, RTDUM16, RTCF^{+}17, RBH12, SL11a, SBA12, SKK12, SS19, SPBS14, TFB^{+}17, TTB^{+}11, VSW^{+}15, WLH^{+}17, WHB^{+}19, YOWK17, ZHM^{+}10, vDCR13]. Behavioral [ACF^{+}10, HBC15, LH10, Lyo14, RKT16, AAK^{+}15, BMPC15, DAM^{+}18, HW12a, HTK^{+}17, INvdB^{+}13, KRA17, KML13, RDF^{+}13, SPP12, TGH15, ZH13]. behaviors [CG10, GDPR18, RAF^{+}14, RNSA15, TB14b, WHK^{+}11]. Behaviour [HSZR10, KBGA19, PPSR15, WSL^{+}14, ABK^{+}14, AJF^{+}12, ACRLB12, ATP^{+}12, ACMD11, ADRA19, BBC^{+}13, BTO^{+}18, BCW12, BBH^{+}13, BWH^{+}17, BGDC14, BK14, BFG^{+}11a, CPM^{+}17, CH15, CBL^{+}10a, CPG^{+}14, CMK^{+}14, CD14, DIG^{+}11, DGMB10, DVV^{+}18, EHG16, EKOÅ13, FG10, FBA11, FJ11, FCFD15, FOJF15, FCS15, FC19, dlÁGRB^{+}19, GCC^{+}16, GAC11, HBF^{+}10, HKL15, HRS^{+}17, HCDR^{+}15, HW14, HSR^{+}13, HCW^{+}15, KBA17, LSH19a, LAK^{+}19, MCPC19, MJS17, MJO^{+}13, MJM^{+}14, MF15, MM10, MTP^{+}10, MPG^{+}10, MDM18, MBC^{+}17, MHVV13, MLB^{+}14, MHR^{+}19b, MMB^{+}19, NGJHB12, NVA^{+}19, NCC^{+}14, OHCG14, ORM^{+}13, OWN^{+}15, PK15, PZL^{+}17, PDP^{+}10, PHHL19, PWM^{+}19, QHN^{+}10, QSvN^{+}11, RAKB^{+}12, RP15, RBGT10, RMNP17, SKSH19, SRK^{+}18, SØA^{+}13, SBG^{+}16, SWK18, SGL^{+}17, SEF^{+}15, SFTG14, SMSL17, SMR^{+}11b, SPS16, SBB^{+}12b, TPT^{+}12, TRGF18, TFW^{+}12, VCG^{+}10, VSG^{+}15]. behaviour [WBG^{+}14, WPS^{+}19, WOHJ^{+}15, WLL^{+}14, WHD^{+}14, vdMPM^{+}10]. Behavioural [BPSCD17, CLG19a, CBL^{+}10b, HLF^{+}19, KHW15, LBP^{+}19, MC11, NSB^{+}19, RAB^{+}10, SUPCB11, BPC17, CdTK^{+}16, KCCG12, LSQ^{+}13, MSH^{+}14, MMCL19, PGR^{+}13, PLM^{+}16, SMF^{+}12, YRM^{+}12]. behind [CFF^{+}18]. beholder [MBC^{+}11]. Beibu [BW13]. being [DZL16, WF18]. believing [AWH19]. Belize [ASWB10, CVM^{+}17, KH10, TBG^{+}17]. bellied [BL13]. below [GP12]. below- [GP12]. Belt [KRDV18, PLG^{+}19]. Beluga [HIH^{+}17, CRRL17, HLPS^{+}15, LMLB12, MMFF12, SWM^{+}19, SOC13]. bemisi [DGFH10]. benchmarks [MSF^{+}19]. beneath [LOS^{+}19]. beneficial [DEGF^{+}10, FTWA17]. benefit [LP12]. Benefits [KMO19, TBG^{+}17, BWP^{+}11, FM14, HKJ17, HGW^{+}11, JVSD10, LKC^{+}15, OMN17, PMS10a]. Bengal [KVR^{+}18]. Benguela [HWE^{+}12, KW17, PBW^{+}11, SUPCB11, SOJ14, vdSRT^{+}18]. Benthic [APC14a, ASK^{+}15, AHSG16, BMW^{+}14, BD11a, CW12, CF12, CFC^{+}12, CCA^{+}14, CRS^{+}18, DDB^{+}16, DZDB10, GBH^{+}10, KBB14, ME11, MHGW13, MTM^{+}11, NCT^{+}14, PDGR^{+}13, PMML12, PP13, PMIG19, RAB^{+}15, SHB^{+}13a, SBM^{+}18, TLB15, WR12, ALS18, AFH11, AGC12, BH16a, BMV^{+}18, BMV19, BST^{+}10, Bet13, BW12, BOU13, CLD^{+}13, CLG19a, CGR^{+}15, CWEE19, CSRF16, CPS^{+}14, CHG^{+}18, CSN12, DMd14, DFS13, DIB15, DDP15, DRDV15, DBS^{+}16, EL16, ES12b, EDS14, EFR12, ENO^{+}11, EBK^{+}12, ESH^{+}10, FC12, FCR^{+}15, FRJ10, FDDD13, FPN17, GHES18, GSJS13, GLM16, GS16, GMBL11, GWM^{+}15, GCCCM16, GMES18, HSH12, HCAV10, HSZR10, HSRP11, HPN11, HWB^{+}18, HSS17, JMB^{+}17, JU16, JM17, JNN12, JC15, JMPZ10, KS12a, KPB16, KKS^{+}14, KTH^{+}18, KFBM19, KHD^{+}11, KB10, KFB15, LBP^{+}16, LS12, LMW12, LDS^{+}15, LHL11]. benthic [LPKT^{+}19, LGD10, LJB14, LMd12c, LMd12b, LDVP^{+}17, LNK^{+}16, LM18, LTCH10, LRS17, MBG^{+}16, MTM^{+}17, MWBMH13, MKS18, MWD10, MRL^{+}10, MHR^{+}19b, MOH^{+}10, NOT^{+}11, NLK^{+}19, OMBP10, OSPWB14, OFT^{+}11, ORM^{+}13, PFB^{+}13, PDBG14, POPH12, PLS^{+}10, PWW12, PFR11, PPM11, PMS10b, PT13, PBH^{+}17, QOGS^{+}16, RMHP14, RP13, RIM^{+}13, RWMP12, RNB^{+}16, RUD^{+}16, RZS12, RTD14, RBS16, RMRA15, RCP^{+}13, SDMB15, SVG16, SRR15, SSB10, SHDR13, SCL13, SGP13, SMJP^{+}15, TPS10, WS13, WKRW^{+}12, WEK14, WWZ^{+}19, XLXS17, XS19, YMCV10, ZBK^{+}14, ZSEV17, dG14, vDBH^{+}15, vdMPRM12]. Benthic-pelagic [SBM^{+}18, PMS10b, SGP13]. benthivorous [BGDC14]. benthopelagic [CBR^{+}18, DIG^{+}11]. benthos [BPN11, CTMU10, HBE^{+}13, HCL^{+}11b, RCK^{+}11, SSB^{+}16, THC12]. Benthosema [STYT10]. Berau [BCdV13]. Bering [Ano18a, BLE^{+}13, BBMH13, DLC^{+}13, DRD^{+}18, DKB^{+}12, FSY^{+}10b, HTCH^{+}18, Hof10, HEAI10, HRK^{+}18, HMF^{+}18, JUK^{+}10, KSA^{+}13, KH18, KTP^{+}18, KLN^{+}13, KLNJ15, LRC^{+}13, LNK^{+}16, MMDM12, MBW10, MHF^{+}14, OBI^{+}16, RDK14, RSG^{+}16, RHS^{+}19, SDAM11, SRH^{+}15, SDAH12, SKPS^{+}16, TB12, WBI^{+}15, WKK18, WK18, ZBK^{+}15]. Beringraja [FGTS16]. Bermuda [EPM^{+}16, FMC^{+}10]. Bernard [MLDG12]. berndti [MTG^{+}11]. Beroe [HTHH11]. berried [MLF^{+}18]. Berryteuthis [HEAI10]. Bertalanffy [BDC10]. berteroana [OAD^{+}14]. Best [Løk11]. BestAgg [BT16]. Bestiolina [JSLG17]. Beta [CLB^{+}18, LRB^{+}12a]. betaine [HAF^{+}17]. betaines [HAF^{+}17]. Betanzos [PLFR^{+}11]. Bett [War14]. better [CFA14, GBU15, HCL10, TM15, TTPM12]. between [ALNB17, AP14, AHD12, Ano18b, AWH19, BHA^{+}11, BH16a, BLE^{+}13, Bea11b, BHE14, BB15, BRBH19, BHN^{+}14, BMJ16, BJH13, BBW^{+}17, BH10, BBR^{+}14, BCWSR14, BCB11, BVD^{+}14, CEH18, CPF^{+}13, CGOC11, CPP10, CMV^{+}11, CPTL11, CL10a, CAG13, CDC18, CS10a, CGM^{+}19, CSC13b, DGCH18, DKAC^{+}12, DM11, DDP15, DEM15, DBC^{+}12b, ES12b, ERG^{+}11, EES15, FCR12, FGRM12, FK10, FCS15, FWL15, FB14, GFL^{+}17, GHHH12, GP11a, GLWI10, GMBL11, GLB^{+}10, GDH^{+}19, GCL^{+}16, HTT^{+}12, HW11a, HW11b, HN13, HW13, HSB12, HGW^{+}11, HTHH11, HA14, HTR^{+}19b, HPQ15, ITT14, JSK19, KB12, KWM12, KLD18, KA16, KBS16, KPH^{+}15, KDO^{+}14, LSBB17, LHPP16, LLBT18, LHJH^{+}10, LHS^{+}14, LN11, Lec18, LRB^{+}12b, LLRC^{+}12, LP10, LOBW^{+}11, LGC10, LL18, LET^{+}12, LCBK^{+}13, LHB19, LDVP^{+}17, LIT13, LÁBFÁF^{+}10, LTCH10]. between [MMDM12, MPM^{+}14, MBB^{+}12b, MRC^{+}13, MKII19, MLSK^{+}15, NMPM15, NMCP17, NDBL16, OSK15, OS18, ODGL^{+}12, ØLH^{+}13, OLM^{+}12, PCA^{+}15, PKW^{+}11, PBIO^{+}18, PH10, PFPP18, PWCvO^{+}14, PHM11, PRB^{+}17, PP18, RNB^{+}12, RAH^{+}10, RSO^{+}13, RHW12, RSTV^{+}13, RGWA11, RSE18, RPM12, RGB^{+}10, RPR16, RPL^{+}18, SGMG^{+}14, SRLV10, SLRM^{+}15, SCAM^{+}14, SMC^{+}14, SS12a, SLKN14, SLO^{+}17, SMK^{+}10, SPH^{+}13, SDVL16, SNNS^{+}12, TAB^{+}14, TSR17, TPCN12, TPVCV19, TVH^{+}10, VBJR^{+}13, VNK^{+}12, WPF^{+}16, WPRS18, Wir12, YAFV12, YSS^{+}18, YOWK17, YRL^{+}10, ZHTA13, ZIT^{+}17, ZFC^{+}13, dMA18, dDTA^{+}18, vdHTT^{+}15]. between-habitat [CAG13]. between-mudflat [DBC^{+}12b]. Beyond [DZL16, TM10, BCW19, CCC^{+}18, GBD^{+}12, MFDW17, MS11, SESS^{+}18]. bi [RNB^{+}16]. bi-phasic [RNB^{+}16]. bias [Bre14, RWR^{+}15]. biased [CGBGS18, GMR13a, GMR13b, HBCK13, SY10, SWNSF14]. biennial [RSSvV19]. Biennium [TPB^{+}18]. biformis [HCBV^{+}11]. big [FGTS16]. bigeye [RTCF^{+}17]. bigger [TTPM12]. Bight [MVS^{+}10, SKR^{+}10, DFS13, FKZTT18, ALE^{+}13b, BAA12, CWK^{+}10, DEJN10, FMP13, GSJS^{+}19, HS11a, HYL^{+}13, MASMH19, PPC^{+}19, SHF12a, SRB^{+}17, SCHO11a, TBP11, TFB^{+}17, TPM19, WMS^{+}10]. billed [CAC^{+}12, KTP^{+}18, SG12, YOWK17]. billfishes [RLNLAF17]. billion [SCHO11a, SCHO11b]. binational [NCW^{+}19]. binoculata [FGTS16]. Bio [LHW^{+}19, CCA^{+}17, GH11, HCL^{+}11b, SMVM^{+}12, VSG^{+}15]. bio-geographical [HCL^{+}11b]. bio-logging [CCA^{+}17]. bio-oceanography [SMVM^{+}12]. Bio-optical [LHW^{+}19, GH11]. bio-telemetric [VSG^{+}15]. Bioaccumulation [CLG^{+}19b, OHA10, SLF^{+}19]. Bioacoustic [SOA^{+}17]. bioacoustics [JBCGC19]. bioassay [KDB^{+}10]. bioassays [HU13, NPRAGG^{+}12]. bioavailable [SLE^{+}13, WJS^{+}14]. biochemical [BCGWS19, KTKM14, SSM^{+}17, SKD^{+}11b, XGLH14]. biochronologies [DAB17]. biochronology [CZB10]. bioconvection [EOA15]. Bioconversion [DBC^{+}12a]. biodeposit [PFL18]. biodiverse [DBE^{+}12]. Biodiversity [LLBT18, LCF^{+}14, PHB^{+}11a, RMSPA16, AHD12, Bar13, BB14a, Bar14, BD12, BLM^{+}18, CPMR11, CB12, CSRS15, FFL10, GC13b, HKB^{+}19a, HC13, HPP^{+}10, KBH11b, KMW17, MJ17, McC19, MS11, RAA^{+}16, RJH14, RDF^{+}19, SHS^{+}17, SDSA14, SGEO13, SOA^{+}17, TBS^{+}14]. biodiversity-ecosystem [FFL10]. bioeconomic [CS15]. Bioenergetic [IMU17, FNF^{+}18]. Bioenergetics [NUJ12, RGB^{+}10, AHH^{+}11, SEFS15]. bioengineer [CM11, FB14]. bioengineers [Col10a]. bioeroding [GRFSM12]. bioerosion [Hug11, LMB^{+}19, WSK^{+}16, YPWP18]. biofilm [CAC11, GC13a, RHF19, SLRM^{+}15, SCWV12, TH11, WSZS16, vdMPM^{+}10]. biofilm-associated [SCWV12]. biofilms [GVR^{+}14, PLS^{+}10, SVIH^{+}10]. biofouling [DGQ10]. Biogenic [MLP^{+}14, BWH11, BPSK11, DT11, LR15, POPH12, PFdP^{+}12, SGTP16]. biogeochemical [BJG^{+}11, BPG^{+}10, LBVH13, MMFP^{+}15, NTN^{+}13, STK^{+}13, TNSN11]. biogeochemistry [HBBH12, PT13, VNK^{+}12, WWO^{+}15, ZFC^{+}13].

Biogeographic [Ber12b, CIF11, HLW^{+}16, MWW^{+}14, SPS16, ARMA15, GFKC13, HRAM13, HAF^{+}14, SRCE^{+}10]. biogeographical [GMSC14]. biogeography [GHT^{+}15a, HP11b, LAM10, PML^{+}19, SFL^{+}18, VCL^{+}15]. bioinvader [GB12]. bioinvasion [LCFR10]. bioirrigation [KRDV18, RF14]. biologgers [BGDC14]. Biologging [BA12, dDTC^{+}17, SHB^{+}12, YNK18]. Biological [BHT15, BH12b, KAAI15, PRCG^{+}17, SHUS17, SGTP16, TNSN11, YPK15, ZGK^{+}19, BGP^{+}16, Bea12, BWH11, CSRF16, CKV12, FHJ15, GC13b, HS14b, HRÁB^{+}14, HWB^{+}18, KVR^{+}18, LMd12c, LKN^{+}14, MN10, PC12, SGS^{+}12, TYG12, VJJ^{+}19]. Biologically [GB10a, LC18]. biologist [FPL^{+}12]. biology [BFEMLÁJ19, CPSC^{+}15, DKL^{+}18, Edm10, GKE^{+}19, MFSRR14, SDST16, SMSL17, TKA10, TGLC10]. Biomagnification [MSS^{+}14b, PBC^{+}12a]. biomarker [BBG^{+}16, CMC10a, GMT^{+}17, KR16]. biomarkers [GCC^{+}14, VLK^{+}15]. Biomass [Fox16, FSY^{+}10b, PLZL14, BMM17, BPN11, BD13, CSRD11, CWW^{+}16b, CCP^{+}14, CKJOM14, CVP11, CRHV14, CWAS17, DABP13, DRD^{+}18, DBH^{+}16, GLY^{+}14, HS17, HFR^{+}13, HCF^{+}16, HBS10, KSA12, KAD^{+}15, LTN19, LMW12, LRB^{+}12b, LCM^{+}16, LBR^{+}19, LRSB19, LRK^{+}15, MJ17, MSF^{+}19, MTMM^{+}15, MH13, NMCP17, OHB^{+}15, OKJMC15, PH16, PMS15, PNF^{+}12, PPB10b, PPC^{+}19, PLF^{+}12b, Pro17, QGM^{+}14, RS16, RLH^{+}13, RPCL19, STYT10, SCRGP11, SVIH^{+}10, SGGL15, STRP18, TFP14, TM10, TL18, TKBN10, TBGW^{+}11, VUL^{+}11, WY15]. biomasses [EM11, GDA^{+}18]. biomechanic [FHS^{+}13]. biomechanical [Wir14, dlSGS17]. biomes [EPF^{+}17]. Biometry [HSRP11, PYBB11]. biominerals [DCIG18]. biophonies [LDVP^{+}17]. Biophysical [HBS^{+}13, LMC14, ASCHM19, BHUCMO19, KCLP16, THM^{+}13, TR13]. BIOPOLE [SMBC18]. biopsy [SMBC18]. bioregion [BG12]. bioregions [PKBS14]. bioremediation [DC15, YYD^{+}15]. biosynthesis [DWV^{+}12]. biota [DCW^{+}14, TVRG^{+}13]. biotelemetry [Ued14]. Biotic [BAvW12, GAO14, HMKF11, HPH^{+}15, HCA11, PML^{+}10, SDM16, SO12, AOW12, GMB^{+}16, HYL^{+}13, KK17, MYS^{+}16, ORM^{+}13, PRA^{+}19, RPHB15, dJH11]. biotopes [GMBL11]. biotoxin [LRR^{+}10]. bioturbated [AOW11]. bioturbating [EBK^{+}12, NPLT10, VCQM^{+}16]. bioturbation [AED^{+}10, BST^{+}10, BBG^{+}13, BPG^{+}10, DLdM19, GME^{+}19, KPLD^{+}12, MBG^{+}16, MFG16, PWW12, RJ12, SK10, Sti13a]. bioturbator [GGNL^{+}17, JSOK18]. bioturbators [Pil10]. biphenyl [MHN^{+}12]. bipolar [KB10]. Bird [HGS14a, HGS14b, HGS15, MC16, SB15a, BWN^{+}14, DPH13, FHH^{+}17, MPS^{+}10, TP16, DBE^{+}12, WHPT12]. birds [BDP10, HS15a, JPP^{+}17]. Birgus [SY10]. birth [RSSvV19]. Biscay [ÁMLUN16, BI15, IAV^{+}10, LOBW^{+}11]. Biscayne [LRH14]. bite [ADF^{+}18, BFM^{+}13]. Bivalve [KN12, BDP10, BD14, BCR15, CLDC^{+}13, DMC^{+}17, DB14, FBC^{+}15, GMT^{+}17, NLKZ12, NDBL16, OBI^{+}16, PZL^{+}17, SAL^{+}12, SPM^{+}13, SHH10, SWPK13, TCN^{+}10, TYG12, TKIO10, TG12b, WPK10, vdGSE^{+}14]. bivalves [ADvdM13, BB17, BDDvdM17, BD19, CG16, CTvdM^{+}08, CKR^{+}11, FS18, GS17, GLB^{+}10, JS12a, LPKT^{+}19, LCFR10, MGP11, RL18, SG18b, TWM^{+}15, YSC^{+}15, vdGvdLvG^{+}19]. Bivalvia [BRS10, SCO^{+}12a]. Black [MB11, SHB^{+}10b, BER17, BBG^{+}11, CSABH^{+}13, CLC^{+}12b, CDMB^{+}18, FMP13, GAB^{+}14, GS11, HLS17, HHS12c, HCDR^{+}15, JJNP12, KCH^{+}11, LZF^{+}14, MATM^{+}19, MIA^{+}11, NHPW18, PRBR11, REAN^{+}14, TRSB^{+}18, WJHS13, YTM^{+}12, DBS^{+}13, OV10, SLKN14, SHI^{+}10]. black-backed [TRSB^{+}18]. black-capped [HHS12c]. black-legged [BER17, CLC^{+}12b, CDMB^{+}18, GAB^{+}14, HCDR^{+}15, MATM^{+}19, MIA^{+}11, REAN^{+}14]. black-tailed [YTM^{+}12]. blackfin [CCS^{+}18]. blacktip [GMHS19]. bladder [WSD^{+}10, WRO^{+}11]. blades [TRSB^{+}18]. Blanco [MJTC^{+}12]. blanket [ZF19]. bleached [HGPM10]. Bleaching [MAS10a, AZSKWL11, BMÁN^{+}18, Car12, CMGS19, CRWG16, DSVM10, KLMR14, LPC11b, LF14, LHW^{+}19, MBB^{+}18, MSLR13, SBO^{+}18, WKvW10, ZAZS11, ZAZS12, vWSGL11]. blennies [HLN^{+}19]. Blenniidae [HLN^{+}19]. blenny [Car12, ZAZS12]. Blood [CSRS15, CCJR14, Lap19]. bloom [ALANGT10, CLG^{+}19b, DSK^{+}10, EVB^{+}16, FW11, FPC^{+}12, GTF10, HGL15, HEB^{+}11, HCJ^{+}11, KAP^{+}14, LBM^{+}18, LHS^{+}14, LS10a, MMEM14, MGG^{+}14, NMCP17, NYK^{+}13, PMF^{+}17, RSS^{+}16, RNH15, SPH^{+}17, STT14, SLL12, SHSK11, VHP^{+}11, YSK^{+}10, GGMTGC18]. bloom-forming [EVB^{+}16, GTF10, GGMTGC18]. blooming [SMJ^{+}14]. blooms [ADD18, AMAM^{+}18, BSG^{+}10, Chi11, CLPU14, CLPiU14, DLL^{+}18a, FTH12, FPC^{+}18, LLRC^{+}12, LWD13, LD17a, OS12, PLG^{+}16, PA15, PQAH^{+}17, Pro17, RG14, SJS^{+}11, SS12b, SBMÅ12, TG10, TSF^{+}10, WRGP12, WWKS15, YMMO14]. blotchfin [GB12]. blow [AL12a]. blowout [BMC^{+}15]. Blubber [MRB^{+}11, BDS^{+}18b, GNHLN11, LBL^{+}12]. Blue [CCG^{+}19, GMP^{+}11, KLSJ10, BSB14, BH10, BL17, BQ16, DEWK15, FHM^{+}11, GFC^{+}13, GLM^{+}19, HW13, JRvdM15, JE10, JL12, JGG^{+}15, KW12, KOPW15, LMGJ13, MGP12, MRH11, NLLS15, NCW^{+}19, ORR^{+}14, ORJH19, QHN^{+}10, RWM18, RSO^{+}13, RLTL14, RLBC10, SvdMK^{+}12, SDT15, SN17, VNK^{+}12, WMH17, WBJ12]. bluefin [BBBF15, DFAH11, FCDC12, GRL^{+}15, IS16, KIU^{+}13, KF17, KHT^{+}17, LCB10, MCW^{+}16, MGAV15, MRC^{+}13, RBA^{+}18, RAF^{+}14, Sat10, SWY^{+}10, WHI^{+}18]. bluefish [JCM13]. Boat [HSR^{+}13]. boats [CRJ^{+}18]. bocaccio [HRT12]. Bocas [AHB^{+}14]. Bodega [RBS16]. bodied [HVK^{+}13, SGS^{+}19, WRW^{+}16a]. bodies [MKII19]. Bodkin [HG13]. Body [FSM^{+}14, GHCC15, MBP^{+}12, AFOVRB14, BDNA10, BC16, BSC17, BMF^{+}15, CRRL17, CVC^{+}18, DGC18, ESD16, FFL10, HTR^{+}19b, JGWJ^{+}11, JWR^{+}12, KLN^{+}13, MHB14, MAS^{+}10b, MD11, NOB^{+}19, NSC^{+}19, NWC^{+}10, ØLH^{+}13, OCDD16, POR^{+}15, QGM^{+}14, RB18, RH12, RSTM14, STSS^{+}16, SF14, VGR^{+}16]. Boersch [MAB^{+}12a]. Boersch-Supan [MAB^{+}12a]. Bohadschia [PPDW16]. Bohai [YYD^{+}15]. Bonamia [EKR^{+}10, HSMS^{+}15]. bonasus [OBA^{+}18]. bone [RMP^{+}19]. bonito [RBA^{+}18]. bony [GRBH15]. boobies [KWPL11, YSM^{+}10, ZHM^{+}10]. booby [HW14, MBC^{+}17, WBS^{+}12]. boom [WGLDB15]. boom-and-bust [WGLDB15]. boosted [CMLC12]. boosts [SVIH^{+}10]. Bopyrid [MH11b]. borders [KLKM12]. boreal [DH17, PSS^{+}12, WCG^{+}14, CLSS17]. borealis [HP13, JBS12, Ric12, WS12]. Boreogadus [SGL^{+}17]. borer [DdRH14]. borers [NHAS15]. boring [CAD^{+}10, CMP^{+}11, YSC^{+}15]. borne [LBJI14, NSY^{+}18, RCBE15, SSM^{+}19, TCKY19, WPS^{+}19]. Both [GP12, Ano18a, BFH13b, CLM^{+}13, CBT13, DFDC18, DHAV13, FLE^{+}11, GSMK15, KCR^{+}10, LZ12, MM11, MP17, MAK^{+}15, MSLR13, SES12, VGL16, Wil14]. Botryllus [BSRDR10, RMKR13]. bottle [RGB^{+}11, RGP11]. bottleneck [JSOK18]. Bottlenecks [PPTB12]. bottlenose [BT10, EC14, FFS^{+}15, PBC11, SKGS11, SHT^{+}13, TLK^{+}12, TNN11, ZPPM17]. Bottom [DMS14, Gar13, PLJ^{+}12, BHE^{+}17, CHRB10, DvV^{+}16, EHR12, GBH^{+}10, HKL12, HB13, HWB^{+}18, KCW^{+}17, KTR^{+}16, KH10, KPB11, KJCN^{+}11, LS10b, LEB14a, MBvR12, MTV^{+}11, MCW11, Mus12, NGP^{+}15, POH16, RMC^{+}14, RZPH14, SGJ^{+}18, SCL13, WPK10, vDBH^{+}15]. Bottom-up [DMS14, Gar13, PLJ^{+}12, EHR12, MTV^{+}11]. bottomfish [MDKM13]. boulder [GC13b, Lec18, LC18]. boulder-field [Lec18]. boundaries [PBH^{+}17, SKD17a]. boundary [GCFW14, HCL^{+}11b, HBM^{+}12, HLD12, JLL^{+}12, LSH^{+}19b, MM10, NMG^{+}13, RLN^{+}13, SJB16a, SBBB^{+}18]. Bouvetøya [BBH^{+}13]. Bowhead [FDLL10, MF15, PLF^{+}12a, RKT16]. box [BEEH15]. Boyer [Ano12d]. Brachyura [APH^{+}10]. brachyuran [OFFQGG10]. brackish [GBW^{+}12, OT14, SDMB15]. brackish-water [OT14]. brain [BVM^{+}19]. Bramanti [Bru10]. branching [AP14, CSM12, LH14, MMRDS18, SPH^{+}13, TBL16]. brasiliensis [CD16a]. Brazil [ABB^{+}14, GDJdBSL19, MLML^{+}19, MSS^{+}14b, PSAK12, dGCF^{+}17, TCSA17, TCJ^{+}19, dSdSdCO^{+}11]. Brazilian [CPS^{+}14, DVB^{+}16, FKZTT18, GKW10, KCE14, LFRF13, PAC^{+}15, RPCL19, VJG^{+}13]. braziliensis [MAD17]. breaching [BP12b]. break [WJW^{+}18]. break-associated [WJW^{+}18]. breakage [BL14]. breakdown [WMS^{+}11]. breaks [SRCE^{+}10]. breakwaters [MLSK^{+}15]. bream [GCL^{+}16, MMB^{+}12a, WJHS13]. breath [MW16b]. breathing [CNSW16]. bred [ODGL^{+}12, TSR17]. breeder [PWB^{+}19]. Breeding [BPT15, GGB^{+}18, OWH^{+}19, ALWD15, AGHC10, BEMD11, BBH^{+}13, BBW^{+}17, BPG^{+}13, BCH^{+}11, BDH^{+}17, BANE^{+}13, CD19, CTZ^{+}16, DMT^{+}14, DZB^{+}18, DJJ^{+}16, DFTZ^{+}15, DCW^{+}10, EPS11, FSF^{+}17, FBA11, GMK18, GAB^{+}14, HTR^{+}19b, JCF^{+}12, KTP^{+}18, KPM^{+}11, KPr^{+}18, LSN^{+}19, LBVH13, LBM^{+}13, LCR^{+}16, LANER15, LFN^{+}12, MC12a, MC16, MIA^{+}11, MJA^{+}18, NWHD15, OMN17, PBC^{+}18, PLNE11, RGEA^{+}15, RCP^{+}16, RB18, REAN^{+}14, RKM^{+}10, RGDNGS11, SPHK17, SPP12, SMC^{+}14, SHT^{+}18, SDKS19, SUB^{+}13, SGG^{+}12, SNS12, TBB^{+}19, TTT^{+}15, TP19, VAB^{+}13, WI12, WK18, XWE^{+}14, ZHTA13, ZJ17, vBMH^{+}17]. Brevetoxin [FFS^{+}15]. Brevetoxin-associated [FFS^{+}15]. brevicaudata [PFKB18]. brevirostris [GGD^{+}11, GGF^{+}12b, KCF^{+}14, NHG10]. brevis [Car12, LWD13, SBS^{+}13, WHT12, WLBB^{+}13, ZAZS12, vdPLdV^{+}11]. Brevoortia [AHH^{+}11, LBCL10, OFDG14]. Brey [BD13]. Briareum [PMH^{+}13]. bridge [PBH^{+}13]. Bridging [RG11]. bridled [CBT13]. brief [ARW12]. brightness [dG14]. brine [GBMW12]. brine-preserved [GBMW12]. brink [KFW^{+}18]. Britain [RGdLB^{+}11]. British [BCH^{+}11, HLS17, HPB^{+}18, FTB14, HHP^{+}17, WOWG10]. brittle [BS11, GBE^{+}17]. brittlestar [CPSC^{+}15, CPSC^{+}18]. Broad [AGC12, BHL^{+}19, CM11, RSO^{+}13, BRLM17, ODK^{+}14, PMSGBL16]. Broad-scale [AGC12, BHL^{+}19, CM11, RSO^{+}13, BRLM17, ODK^{+}14, PMSGBL16]. broadcast [BCW19, MBvO^{+}18, NvOH13, SHH10, TPBM13, TSCvS^{+}15]. broadcast-spawning [NvOH13]. Broader [VMC^{+}19, BGO^{+}18]. broadly [BSH^{+}18]. broadnose [AB11]. Bronx [KKY15]. bronze [IHD^{+}16]. brood [Col10b, LANER15, ORR^{+}14]. brooded [SPM^{+}16]. brooder [KB10]. brooders [HVT12]. brooding [BEFC15, FHL^{+}17, HRC12, JWB^{+}17, SPM^{+}16, WJW^{+}18]. broods [PHB11b]. Brown [TBMdV12, ABT^{+}15, ASM^{+}12, AHØ^{+}15, BFL^{+}11, CSJP17, CBC^{+}10, CMC^{+}16, Cra12, DWPH11, HT11a, HT11b, KHK^{+}17, LERS13, MHS17, OHA10, SHT^{+}18, TGRH17]. Brownian [PBH^{+}13]. Brünnich [FSP^{+}17]. Brunnschweiler [CBL^{+}10a]. Bryde [TCSA17]. Bryozoan [OOK11, CGG10a, DHG11, FKT15, GMM10, LKB^{+}14, PHB11b, SWMK15, SM10a, SG18a, ZOM^{+}10]. bryozoan-nudibranch [SG18a]. bryozoans [KP11, LKN^{+}14]. Bubble [DTC^{+}17]. buccinid [AJF^{+}12]. Buccinum [HLC^{+}18]. budegassa [CALB10]. Budget [SvdMK^{+}12, JR15, NWC^{+}10, RCP17, SPM^{+}13]. budgets [BPT15, CLC^{+}12b, GLW^{+}17, LHW^{+}19, áNGGS11, SdBAH^{+}14]. buffer [GWK^{+}12, HPR^{+}12, RPL^{+}18]. Bugio [RPM^{+}13a]. Bugula [DHG11]. Builder [AMAJ^{+}17]. Building [FUOG^{+}16, AHPKS17, Ber12b, ELM^{+}19, LKDP17, PGG12, RPGP^{+}15]. bulk [DWPH11, SWR11, VMC^{+}19, YSC^{+}15]. bull [BMJH12, BTLF17, HS11b, SHB^{+}19]. Buller [PWM^{+}19]. bullet [LCQT^{+}13]. buoyancy [EBOM11]. buoyant [WC18]. burdens [LRC^{+}13]. Burial [JNN12, BL14, BO13, CL15c, HBB^{+}12, MATR11, MO12, SMSE19, SDH13]. Buried [HH15, BKC14]. Burmeister [CASG^{+}19]. burrow [MC12a]. burrow-nesting [MC12a]. Burrowing [BST^{+}10, CAD^{+}10, DdR10, DBC^{+}12b, GME^{+}19, PZL^{+}17, PWW12]. burrows [POX^{+}17]. bursts [AAM^{+}12]. bust [WGLDB15]. butcheri [RGWA11, WJHS13]. butterflyfish [FBGH18]. by-catch [BBRM14, dQSC14]. by-caught [dQSC13]. Bycatch [RLBC10, ACP^{+}16, BRJ^{+}12, BPZ^{+}14, CAGS17, HW12a, Joh10, KLBT^{+}16, LLMA^{+}10, Løk11, LMSC^{+}19, MTB^{+}17, MO13, OMW^{+}16, SBLM14, TRA^{+}14, WFS10]. byssal [ZGH^{+}17]. Byssus [BC11, BA13a].

C [MFH^{+}11, ACG12, CCJR14, CBR19, CSC^{+}11, DTVH10, FSE^{+}17, GFC^{+}13, GCS^{+}18, GCS^{+}19, GMT^{+}11, HHS^{+}12a, KBA13, KSK07, KJW^{+}15, KBP^{+}10, LMH12, MWP10, MFI11, MPB13, MHS14, MFS14, NMPM15, OPN10, OFT^{+}11, PPD^{+}14, PCW^{+}12, PHWC13, RP12, RLAGL12, RCVG10, SML^{+}14, TMP^{+}13, UDF^{+}18, VRSS18, VTZ^{+}13, VSM11, VB14, VLM14, YSC^{+}15]. C-cycles [MFH^{+}11]. C-labelled [GMT^{+}11]. C. [KDS^{+}12, PPJT15, SKM^{+}15, SKD^{+}11b, YBH^{+}11]. CA [GHES18, BAH^{+}15, KSK07, MC10a, PZL^{+}17]. cable [BHMCM16]. cabled [ASS^{+}13, RJ12]. Cabo [CPS^{+}14, MJTC^{+}12, SML^{+}14]. cabrilla [AMP^{+}11]. CaCO [SMSE19, SDMR19]. Cádiz [PDGR^{+}13, PPS^{+}13, PBV^{+}11]. Cadmium [NCS^{+}10, SPSP10]. Caesio [APC^{+}13]. caespitosa [KCLLL13, KLMR14, TPFP11]. cage [HHH^{+}11c, MLF^{+}18, PBC11]. Caged [GSL17]. calanoid [CLC12a, KFGSMPG14, LME^{+}19, PFL13]. Calanoida [ACDJ11, OHST19]. Calanus [BLS^{+}11, DSM11a, DKL^{+}18, DRT13, HTHM18, Hal15, HJMN^{+}12, Ji11, KDS^{+}12, KDO^{+}14, MPZ10, OHST19, PPJT15, PPH11, PPD^{+}14, PSS^{+}12, RSS^{+}16, SKM^{+}15, SLNG18, SRRS12, SHI^{+}10, SKD^{+}11b, WWB14, YBH^{+}11]. Calcareous [CGR^{+}15, GBU15, PHTS11, RJB^{+}14]. Calcification [BRF11, WDL^{+}18, CTRBLC^{+}13, GUC^{+}15, HYC^{+}14, LKDP17, MP15, MFG^{+}13, MEKC15, PKB16, RC12, WBG10]. calcifier [DPY^{+}16]. calcifying [BPM^{+}18b, BGBD^{+}13, CCM^{+}14, HSB12, JDX^{+}17, LZ12, LLF^{+}18, MM16, VMWU15]. calcite [DPY^{+}16, Fie14, FKT15, LKB^{+}14, SCL^{+}16]. calcium [GUC^{+}15, McC12]. calculations [GKHM13]. calderas [SNW^{+}18]. Caledonian [OGGP14]. calf [CVC^{+}18, FBA11, TFW^{+}12]. Calibration [AHH^{+}11, GHT15b, KKMA^{+}10, RT12]. Calidris [BEMD11, KMSW10]. California [ALE^{+}13b, CWK^{+}10, GMSC14, GSJS^{+}19, HYL^{+}13, LLCM^{+}12, PPC^{+}19, RBS16, SRB^{+}17, TBP11, TKEV^{+}19, WMS^{+}10, YIKG10, AGGAO^{+}18, AGRPSVL13, BSG^{+}10, BFF^{+}10, BO15, BS16, BBR^{+}14, BBC12b, BHN^{+}12, CTRBLC^{+}13, CLH^{+}15, CSRD11, DABP13, DFGH^{+}15, DSGR^{+}15, DEM15, DPO15, ED12, ESW14, FHRC16, FK10, FWL15, FSS^{+}18, Gla10, GB16, GGMGR12, HCL^{+}11a, HFR^{+}18, HKB^{+}19b, HRT12, HM09, HHWW11, JGM^{+}16, KBG12, KLKM12, KCW^{+}17, KFW^{+}18, KHW^{+}19, KGLLW11, KMM15, KMW17, LPKT^{+}19, LFL11, LBE^{+}10, LH10, LMH12, LBS^{+}15, MSF18, MVVGAV10, MMSC17, MBRO10, MBS^{+}13, NCK^{+}14, PSG^{+}13, PTS^{+}18, SPH^{+}17, SS11b, SSX^{+}17, SHF^{+}12b, TL18, TFS14, TAB^{+}14, THW^{+}16, Tho15, TEAO^{+}14, TS16, VMC^{+}19, WM10, WY15, WSF^{+}12, WSS^{+}16, WJW^{+}18, WZP^{+}14, YPK15, YD11]. Californian [SWW11]. californiana [McD13]. californianus [GLW^{+}17, LKS12]. californicus [ABM12, TB14a]. californiensis [VPW^{+}12]. callianassid [UTS^{+}18]. Callichirus [PWW12]. Callinectes [JC17, MRH11, ORR^{+}14, ORJH19, SN17, WMH17]. calling [FMG^{+}18, MBKM17]. Callionymus [GB12]. Callorhinus [LJM^{+}10, SBA12, ZO10]. calls [AHR^{+}18]. calves [NSB^{+}19, RUS^{+}13]. calving [SGG^{+}12]. Calyptraeidae [MC12b]. camera [KWC^{+}17, RST15, YD11]. cameras [DIG^{+}11, ETHB17, FBPCC^{+}19, RJ12, WPS^{+}19]. campana [MPLLE17].

Campbell [SCW^{+}18]. campechanus [LBCS^{+}15, MH19, TS11b, WGS16]. camtschaticus [FPN17, PES12]. Can [FKT15, AS13, BMÁN^{+}18, BSC^{+}15, DSG^{+}13, FDVL18, FFS^{+}13, GBE^{+}17, GRL^{+}15, GKPT16, HPR^{+}12, dDTC^{+}17, KBS16, LSH^{+}15, LGR^{+}10, MV12, RSW^{+}16, RLM^{+}14, SGS^{+}17, SWK18, STM^{+}14, TMSG19, TVA^{+}19, VDS11, VMWU15, ZJ17]. Canada [PPD^{+}14, AB12, AJ16, BWS^{+}12, BOCB10, FTB14, FHH^{+}17, GLM^{+}19, HLS17, KWW^{+}14, MMDM12, MF15, MSH^{+}14, MGG^{+}14, MSK^{+}19, PGOG12, RBP^{+}10, RJ12, SWM^{+}19, TBH^{+}18, TR13, dlGMD^{+}17]. Canadian [KFBM19, AGM^{+}11, BDR19, DMMW12, EMG^{+}15, GMT^{+}15, GLS^{+}15, HIH^{+}17, LKL^{+}19, MTG^{+}10, MFD12, MKT^{+}18, PMF^{+}17, PLF^{+}12a, SHH10, STP13, STBP13, VK11]. Canary [KFW^{+}18, BSD^{+}14, SMVC^{+}15]. Cancer [GPM^{+}13, MHS17, SE11b]. Candidate [GCL^{+}16]. canine [MBSG11]. Cannibalism [LH16, RUPT^{+}11, BL17, GSP^{+}15]. cannot [HFM12, STS^{+}18, dJSB19]. canopies [BSD^{+}14, MBBBC12, WHP^{+}17]. Canopy [BAM^{+}10, CPHH15, PSC13b, CSS^{+}17, FNA^{+}16, FJW14, GL16, PNPF14, SLT^{+}19, TS11a, WS13]. Canopy-forming [BAM^{+}10, SLT^{+}19]. Cantabrian [VDMLU13]. canyon [GLM16]. canyons [BWS^{+}12, KAD^{+}15, KHN15, SRH^{+}15]. capable [SBS^{+}13]. capacities [DDD^{+}11, NM18]. Capacity [LH14, BD19, DC15, FDP^{+}16, HP13, PHD^{+}14, TBH10]. Cape [BNAK11, BRT^{+}17, CPG^{+}14, DAC^{+}12, GCRP15, KW17, MMB^{+}19, MvdHBM12, STP13, STBP13, TYG12]. capelin [BKAM12, CD14, CD16b, LBR^{+}19, MHR19a, MP17, OA10, RDHM10]. capensis [BNAK11, PW19a]. capitata [CRWG16, PGG12]. capped [HHS12c, TTB^{+}11]. capping [NSG^{+}12]. caprellids [Hos14]. capriscus [BSB^{+}18]. capsule [HA18]. capsules [LMZ16, TKL^{+}11]. captive [CLP18, JCF^{+}12, ODGL^{+}12]. captive-bred [ODGL^{+}12]. captive-breeding [JCF^{+}12]. capture [BCPC18, CUCS17, DCHI12, FH18, GSCH14, MASW^{+}11, RL10, TOR^{+}10, VGGD^{+}15]. captured [CWR^{+}18, DHB18, HIMH15]. Capturing [YMCV10, RSH^{+}19]. carangid [BBKP10, LHT^{+}16]. Caranx [BBKP10, DFD^{+}19]. carbohydrates [SPSP10]. Carbon [BKM^{+}18, CE11, DVSM10, ESH^{+}10, GGK^{+}16, KRDV18, MDE13, OCSW16, Rei12, SDM^{+}13, VARO12, BW16b, BW16a, BST19b, BIL^{+}10, BS11, BBRBG10, BMF^{+}12, BBG^{+}16, BVC^{+}16, DAL^{+}10, DvV^{+}16, EBT^{+}15, FBC^{+}15, FSE^{+}16, FLSG^{+}15, GWME16, HHS^{+}12a, HCAV10, HN13, HTCH^{+}18, HLL^{+}17, HGPM10, JBS11, JS12b, KBA13, KMP15, LBN11, LMW^{+}13, LR15, LGW^{+}15, LLCM^{+}12, LLP^{+}10, LSRRCM13, ME11, MBSG11, MWD10, MG12, MGB^{+}14, MDJN19, NW15, áNGGS11, OFT^{+}11, OFDG14, OMD11, PSA17, RCP17, RMHP14, SMSE19, SBE^{+}16, SDMR19, SMJP^{+}15, SSX^{+}17, SO16, STRP18, SJL^{+}15, SCW^{+}15b, SOBNL13, SGD^{+}12, TGI14, TNS^{+}12, TNM^{+}11b, VKH10, WKE^{+}17, WBI^{+}15, WBLS^{+}11, WP10, YYU^{+}17, YKG^{+}15, ZIT^{+}17]. carbon-concentrating [YKG^{+}15]. carbonate [BvHC^{+}15, BWH11, DLL^{+}18a, Mat10b, MTH15, TGM^{+}17, TTF^{+}11, TS16, WDL^{+}18]. carbonates [PFdP^{+}12]. carbonation [ZHJ^{+}17]. carbonic [VJM^{+}16]. carcass [ALSO^{+}18, ET11, HGS14b]. carcasses [BLMD^{+}11, ITT14]. carcharhinid [DCHI12]. Carcharhinus [BMJH12, DOH^{+}17, FCFD15, GMHS19, HS11b, KHS^{+}11, LSH^{+}19b, MBB^{+}15, MH12a, SHB^{+}19]. Carcharias [HBB^{+}18, KCS12, RABS19, BHL^{+}19, MMB^{+}19, SBCT17, SKFH18, TKEV^{+}19]. Carcharodon [BHL^{+}19, MMB^{+}19, SBCT17, SKFH18, TKEV^{+}19]. carcini [WGD^{+}16]. Carcinus [ATP^{+}12, BSH^{+}14, DT15, DCT^{+}10, HBHF^{+}11, MMG^{+}16, MEK11, WGD^{+}16, WNA^{+}15, YPK15]. cardinalfish [GJ16]. care [BDS^{+}18b]. Caretta [CCC^{+}18, ECM^{+}12, LPC^{+}11a, LKL^{+}17, MO13, PPM^{+}15, PDT^{+}11, RAB^{+}10, SMVC^{+}15, WHK^{+}11, dQSC13, dQSC14]. Caribbean [ABRP15, ADF^{+}18, ASI15, ACL^{+}16, BHUCMO19, BFPONS^{+}12, BFEMLÁJ19, CSB14, CL16, CSRS15, DAM^{+}17, DAM^{+}18, Edm10, EL19, Fog12, GL16, GEF^{+}12, GHL11, GCCCM16, GCCM18, HJN^{+}10, HPN11, HPL12, HSS^{+}19, HB15b, KHBL11, LP10, LSFQ19, LÁBFÁF^{+}10, ŁCM^{+}18, MLPC18, MWP10, MHC11, MSE^{+}18, MLP^{+}13, MSE^{+}12, ORWPR14, PMEZ15a, PMEZ15b, RMME10, RWAP16, RNSA15, SSB10, SD17b, TLAT13, TBL16, TMS14, WMCC15]. Caridea [KHS10b]. Carijoa [CKC^{+}10]. Carlson [ZAZS12]. carmabi [MWP10]. carnatus [LBS^{+}15]. carneipes [WPF^{+}16]. carniepes [RHW12]. carnivore [SKSH19, WHK^{+}11]. carnivorous [NLK^{+}19, WvLF18]. carnivory [WTGC^{+}11]. Carolina [ARA^{+}10, CR14, WRB^{+}15, WMB^{+}14]. carriers [SWR11]. Carry [VHM16, Ano17c, BBW^{+}17, SMC^{+}14]. Carry-over [VHM16, Ano17c, BBW^{+}17, SMC^{+}14]. carrying [BD19, TBH10]. Carryover [CPS15, MLB^{+}14]. cascade [BBTW14, Ber12a, NT10]. cascades [ABJ^{+}15, BS12, JGS14, MSE^{+}12, OE11]. Cascading [DTM^{+}12, GP11b]. case [ACDJ11, APH^{+}10, BDLW10, BRB11a, CLB19, DOH^{+}17, DDB^{+}14, DMR10, FSP^{+}17, GHGF12, GSJS13, HYC^{+}14, JSBP15, KHW^{+}12, LAB10, LA12a, LEB14b, LD10, MATM^{+}19, MPRPM^{+}18, NGA18, PBM^{+}11, RR18, RCK^{+}11, RWM^{+}12, RUD^{+}16, RZH16a, RvOP12, SMSE19, SKGS11, SABMG16, SP12, VGGF12]. Cassin [SHD^{+}19]. Cassiopea [FSE^{+}16, JWR^{+}10, FSE^{+}17]. caste [Miu12]. Castellanos [She13]. Castellanos-Galindo [She13]. catadromy [DMD^{+}12]. Catalina [GHES18, TS16]. Catch [MA14, BBRM14, GOM^{+}19, KW17, KB12, KFW^{+}18, MVVGAV10, MS10, RLBC10, TS14, ZAC11, dQSC14]. catch-per-unit-effort [KB12]. catchability [VRAP^{+}14]. catches [GHD^{+}10, MLF^{+}18]. catching [BBC14]. catchments [WRW^{+}16a]. category [YChH11]. catfish [AdCM^{+}19]. Catomerus [ARMA15]. catshark [TPVCV19]. caught [TBMG17, TWL15, WRW^{+}16b, dQSC13]. Caulerpa [BAC^{+}11, BOU13, GP11a, TBKG10]. cauliformis [GEF^{+}12]. causality [KLD18]. cause [GB10b, HBHS^{+}18, Ing16, KS13, KJC^{+}15, SVG16, SFLG10]. caused [BWLP17, BRB17, BRIC13, CAD^{+}10, HHH^{+}11c, JPP^{+}17, KCCG12, McD13, RTDUM16]. Causes [PLGSD17, RK15, ADD18, Ano17c, BCM11b, Car12, KSA12, KKA12, KRA17, LTP13, NDAP19, SMF10, TMM15, TBMdV12, ZAZS12, DDM^{+}16]. causing [GMR13b, LRBS15, MSLR13]. caution [AvdHH^{+}17, SH17]. cave [JCA13, NBGGST^{+}15, RHVRC11, SGEO13]. caveats [AHD12]. cavirostris [WWB^{+}17]. cavolini [BAFP^{+}13]. Cd [CSC^{+}11, LRC^{+}13]. CDOM [CAKD14]. Celestial [BNW14, SPD^{+}14]. Cell [AFRAG16, BOS16, Dao13, DTVH10, LWD13, LAALH10, LSRRCM13, NDD^{+}13, PO13, PLS^{+}10, PMF^{+}17, QYS^{+}12, SH10]. cell-quota [PO13]. Cellana [DHGW15]. cells [MDAPH11]. Cellular [KR16, OSL18, VTZ^{+}13, VMN^{+}12]. census [BGB^{+}17, CS10a, KB12]. censuses [AWH19]. Center [PML^{+}10]. Central [HVJ^{+}18, SCAM^{+}14, AAGGFG^{+}12, ALS18, BBF^{+}19, BW13, BOÁNF19, CGR^{+}15, CBR^{+}18, CBMK10, CPID13, CD13, CWW^{+}16b, DLRF15, DSGR^{+}15, DEM15, FK10, FSDS10, GBMW12, HSPS14, HHWW11, HSS17, IMS^{+}11, KSA^{+}13, LBS^{+}15, MHGW13, MZF^{+}19, PPEH^{+}14, RBEGG13, RLNLAF17, RPR16, SBH^{+}18, TBDM11, VDMLU13, VBŽ^{+}17, WKX^{+}19, WY15, WBLRN18, WSF^{+}12, YSM^{+}10, dB10, BPM^{+}18a, CSPA10, DKRS12]. central-northern [BBF^{+}19, DSGR^{+}15]. central-place [RPR16]. central-southern [PPEH^{+}14, RBEGG13]. centre [FHJ15]. Centropages [SMR14]. century [CCP^{+}14, CHFS12, LPKT^{+}19, WWAL13]. cepedianus [AB11, IDT^{+}18]. Cephalopholis [GFC^{+}14, MD11]. Cephalopod [DFGH^{+}15, XWE^{+}14, GCS^{+}19, KHÁB^{+}17, PPX17, SDF^{+}19]. Cephalopoda [BB14b, GCS^{+}18, ICT^{+}11]. cephalopods [CBR19, CSC^{+}11, GPC^{+}15, PSG11, PHG^{+}14, RGDS12]. cephalus [DBS^{+}13]. Cerastoderma [BD15, TCN^{+}10, dML11, dMBC12]. Cercopagis [KKL13]. cervicornis [LSBB17, MMRDS18]. cessation [WD14]. cetacean [BFF^{+}10, FYG^{+}18, JTH^{+}12, PSG^{+}13, RNFG10, SCB^{+}16, TCJ^{+}19, TRHM14]. cetacean-habitat [BFF^{+}10]. Cetaceans [ECC^{+}15, CRJ^{+}18, DER^{+}19, LMR^{+}10, NKM^{+}15]. Cetorhinus [CQR^{+}17]. Chaenogobius [HIKK12, HI14]. Chaetoceros [vdPLdV^{+}11]. chaetognaths [KBB10b]. Chagos [GS15]. chain [DMH15, NT10, PLG^{+}16, YKBB12]. chalcogramma [BBMH13, ZBC10]. chalcogrammus [WDM^{+}18]. challenge [CAC^{+}12, PCS^{+}15, SSM^{+}16a]. challenges [BPSK11, FPC^{+}18, GS15, KNS^{+}15, LS14, Nob11, Pur18]. chamber [BWJB16]. Chandeleur [KCMH^{+}17]. Change [LFMC11, Ano17c, AB12, BRJ^{+}12, BFDM18, BDPL19, BDR19, BW15, BVQ12, BJL^{+}17, BWN^{+}14, CDMW12, CA14, CLB^{+}18, CPR^{+}14, CHFS12, DMW^{+}16, DKAC^{+}12, DED^{+}15, DS16, DFB^{+}17, FAK^{+}12, FDP^{+}16, FT10, FTH12, FGH^{+}13, GTL^{+}16, GKR^{+}15, GWK^{+}12, GCR16, HHS12c, HWVM19, HCL^{+}11b, HLB^{+}16, HA15, IMS^{+}11, KTNO19, KHW^{+}10, KAP^{+}12, KHD19, LS10a, LEKT12, LBG17, ŁCM^{+}18, MBB11, MHF^{+}19, MUC12, MRC^{+}12, MC12b, MBCB^{+}19, MOK^{+}18, NdBKR13, OCDD16, PMB^{+}17, PBG^{+}10, Pil10, RIM^{+}13, RPE^{+}13, RFMT12, RGdLB^{+}11, RK15, SKB15, SH12, SS11b, SH17, SA17, STK12, SK12, TSS^{+}12, VCL^{+}15, VMN^{+}12, VM14, WNG^{+}19, WNP^{+}14, WL19, YKG^{+}15]. change-induced [SA17]. change-points [SH12]. Changes [BO15, BDLW10, BD12, BPF^{+}18, BANE^{+}13, CY10, CPG^{+}14, Coo11, GBTS11, HBF^{+}10, HFR^{+}18, IMS^{+}11, JMB^{+}17, KHW^{+}19, KV13, McC17, RWM^{+}14, SSS19, TMGCM11, TFS14, TPM19, VJJ^{+}19, AGGAO^{+}18, AFOVRB14, AFLM12, AGIW16, AKF14, ASEB11, AMLR^{+}17, BNAN12, BJG^{+}11, BCSH18, Bea12, BMVBC10, BDDvdM17, BDD17, BADH^{+}12, BGIA19, BLPP16, BLC^{+}16, BCM11b, CBR^{+}18, CLGCA^{+}19, CWEE19, CPR^{+}16, CLB^{+}18, CLGC16, CAG13, CMC^{+}15, CNÁCP12, CHLR16, DHG16, DLC^{+}13, Edm13, EPL^{+}14, FKT15, dSKFV12, FA10b, GLM^{+}10, GE14, GP11b, GBS^{+}10, GCS^{+}18, HBHS^{+}18, HBI^{+}18, HH15, HRM^{+}11, HIKK12, HFR^{+}13, HCF^{+}16, JCRC10, KFBM19, KDB^{+}10, LKC^{+}15, LBJ^{+}16, LALCM18, LNK^{+}16, LD17b, LRSB19, MVCD14, MTS^{+}17, MGGSP15, MBP^{+}12, MPG^{+}12, MvdHBM12, MCTHGV12, MN10, MMGB14, OSC19, ODK^{+}14, OSL18, OMH12, PDGR^{+}13, PBC^{+}18]. changes [PvdMS^{+}15, PBAP^{+}14, Pör12, RGEA^{+}15, RP12, RLAGL12, RSTM14, RCP^{+}13, RM19, SBV^{+}15, Sat17, SLSC16, STK^{+}13, SKB15, SSB^{+}18, SOJ14, SK10, SSHB16, SSB^{+}16, SGP^{+}17, SESS^{+}18, SLR^{+}14, TDM16, UCK^{+}13, VCG^{+}15, VTZ^{+}13, WL10, WJWC16, WKK18, XGLH14, YTO^{+}11, ZBK^{+}14, dG14, tHHR10]. Changing [BSY13, KRA17, PUP16, RSDN15, SAM18, BMV^{+}11, DS13, EW16, HLIK16, HBK^{+}12, JCM10, KPW^{+}13, KOLDP17, KH18, LM12, LJB14, MGFW13, MCC^{+}16, MTH15, OBÁSV18, PSP^{+}19, PGOG12, RLN17, WAM^{+}15, vdMPM^{+}10]. Changjiang [MYK^{+}16]. Channel [CSRD11, KBG12, MBM^{+}14, BLHL15, FDDD13, JGL12, KHG^{+}19, LPC19, MVE^{+}10, MPB13, NFR^{+}14, dGCF^{+}17, SMND16, TKW^{+}14, XTW^{+}15]. channels [McC12]. Chaotic [CFMN16, HTH10, MHS17]. characterisation [AOW11, FSB15]. characterised [LHT^{+}16]. Characterising [LCR^{+}16]. Characteristic [Bet13]. Characteristics [DBB11, HHB^{+}14, ABF^{+}14, ABM12, AGH^{+}12, ADB10, BBM^{+}14, BQ16, CMAJ12, DMT^{+}14, HPA17a, HPA17b, HD10, KTF10, KSK^{+}12, KJCN^{+}11, LEB14a, LEB^{+}19, MRC^{+}13, OSL18, TLK^{+}12, VB19a, WH14]. Characterization [JMK12, PIPD10, RRRF^{+}17, SPSP10, THM^{+}13, VTSH^{+}19, CAKD14, EFOM13, HOA14, IMU17, NMPM15]. characterize [SBT12]. characterizes [CTvdM^{+}08, HCL^{+}11a]. Characterizing [LA12a, NHTS14, BST^{+}19a]. Charlie [RWMP12]. Charlie-Gibbs [RWMP12]. charruana [GFL^{+}17]. Chasing [BBC14]. Chatham [GOO14, BPN11, DMT^{+}14]. cheater [AAvS^{+}13]. Cheirimedon [NPRAGG^{+}12]. Chelonia [BHTF11, CLS^{+}16, GDB^{+}16, HNK^{+}18, HSMB17, JBL^{+}16, LABM17, LKL^{+}17, MLM^{+}12, PRK^{+}12, SMBH15, TGH15, TSA^{+}18, WBB^{+}10, WHFM17, vdMHW^{+}10]. Chemical [DP14a, GLWI10, HCM^{+}11, WRW^{+}16b, AAvS^{+}13, AEM^{+}18, BTTK10, BAT14, CLM^{+}13, IMD11, LP10, MK13, MPLLE17, MB11, SWWK13, VLK^{+}15, WWAS18, WRFM16]. Chemically [DL14, PKW^{+}11]. chemistry [ACW18, AdCM^{+}19, IHD^{+}16, KRR^{+}18, MTH15, NSS10, SW18, TTF^{+}11, WDL^{+}18, WDM^{+}18, WWG^{+}13b, ZAMB10]. Chemoautotrophic [SNW^{+}18, VCQM^{+}16, VCMS^{+}18, WBJ12]. chemosymbiotic [vdGSE^{+}14]. chemosynthesis [ALSO^{+}18]. Chemosynthetic [LKL^{+}19, GSD10, SJL^{+}15]. Chesapeake [AHH^{+}11, BBGL13, CB11, GSO^{+}18, JL12, LGB17, LKCH12, LBCL10, MSC13, MH10, ORR^{+}14, PW15, PS19, PFPP18, RWM18, RSO^{+}13, RHH^{+}15, RLO18, SL15b, TNM^{+}11b, WMH17, WLB^{+}11]. chesapeakei [SB18]. chick [BM13, BRT^{+}17, CHP^{+}11, CDMB^{+}18, CFPF^{+}14, Hat13, MGPC13, RP15, WHN^{+}18, WLB^{+}15, WJW^{+}18]. chick-brooding [WJW^{+}18]. chicks [ADL^{+}18, FRMS17, PWM^{+}19]. childressi [RCF16]. Chile [BBF^{+}19, BOÁNF19, BQ16, CGOC11, FKRJ13, GCC^{+}10, HVT12, JDQS11, LAB10, LCC^{+}11, MC10b, OAD^{+}14, PPEH^{+}14, RBEGG13, VHHGF11, WBLRN18]. Chilean [CMC10a, GH11, GGK^{+}16, MFH^{+}11, SFNP11, VHHGF11]. chilensis [MS18a]. Chiloé [GCC^{+}10]. chimaeroid [DGFH10]. chimney [LZF^{+}14]. China [BW13, DLL^{+}18b, GFW^{+}14, LLC^{+}14, ST10a, STYT10, SLS16, WLF^{+}17, XLXS17, YYO14, YHTC19, ZCC^{+}12, DWD^{+}18, FWS^{+}18, NLKZ12, XS19, YYD^{+}15]. Chinese [JCZ^{+}12, QCX^{+}10]. chinned [PDP^{+}10]. Chinook [BWRP14, DBA17, HTB^{+}15, JGM^{+}16, MGM10, RWP^{+}14, RRG^{+}11, SCAM^{+}14, SAM18, SFKQ15, TFS14, TTPM12, TTW^{+}12, WSF^{+}12, WSS^{+}16, WWW^{+}13, YBB^{+}12]. chinstrap [BBH^{+}13]. Chionoecetes [DMMW12, TB12]. Chironex [SWK18]. chitobiase [SAD^{+}16]. chitobiase-based [SAD^{+}16]. Chlamys [BRS10]. chloride [RJB^{+}14]. chlorophyll [BW13, FMH^{+}15, dSKFV12, IAV^{+}10, MTG^{+}10, SSR^{+}10, SYX^{+}10, SSS12, WKX^{+}19, WP10, dJdBdJB12, vdPLdV^{+}11]. chlorophyll-a [BW13, dSKFV12, IAV^{+}10, SYX^{+}10, SSS12, WKX^{+}19, dJdBdJB12]. Chlorophyta [PCNB16]. chloroplast [ZFRG12]. chloroplasts [BH16b]. Chlorostoma [CS11a, GB16]. chloroticus [GFP19, RP17, WW15]. Choerodon [SKK^{+}18]. choice [GVGH15, LRRM^{+}11, SML17, ZAMB10]. chondrichthyans [APNC^{+}15, PNI14, PBC^{+}12a]. chop [KCD13]. Christmas [CW12, HYL^{+}13]. chromatic [HDWF10, LCCT19]. Chromis [BTTK10, BS14b]. chromophoric [CAKD14]. chromosomes [HLN^{+}19]. Chronic [CMR12, BVGH10, SGJ^{+}18, SCL13, TTF^{+}11, VHG^{+}18]. chronologies [MBW10]. chronology [DDP^{+}13a, DCHI12]. Chrysaora [BB14c, BBR^{+}14, SB18]. chrysolophus [WHPT12]. Chrysophrys [MJHS11, MJHS13]. chrysoura [SGP^{+}17, SL15b]. Chthamalus [SSBA12]. chub [KTNO19, YYO14, YNK18]. Chukchi [CMM^{+}18, JFK12]. chum [FSY^{+}10b, KSA^{+}13]. ciguatera [LRBS15, RL11]. ciguatera-causing [LRBS15]. ciliary [RL18]. ciliata [PAF^{+}10]. ciliate [KLSJ10, SP13a]. ciliates [GSMK15, MAB^{+}10, SCWV12, WSZS16, XS19]. Ciliophora [SP13a]. Ciona [MSK^{+}19]. circular [EH13, PBM^{+}11]. circulation [DEM15, HCJ^{+}18, MvdHBM12, RAA^{+}16]. circum [MBvO^{+}18]. circum-Antarctic [MBvO^{+}18]. Circumpolar [NFSW10, TTM19]. citations [TGLC10]. citizen [HSMB17, PCD^{+}18]. citizen-science [HSMB17]. clade [DBS^{+}13]. cladoceran [KKL13]. Cladocora [KCLLL13, KLMR14, TPFP11]. clam [BD13, CLG19a, DBE^{+}12, GBMW12, GSTK11, MS13, MW16a, MPD11, MAL15, OCE^{+}12, OCDD16, TKDM16, WBG10]. clams [AAIB15, BFL^{+}11, FRRÁSL11, GLF^{+}17, GSO^{+}18, GP12, HAF^{+}17, HP14, OPN10, PMS13, WO11]. clans [EPC^{+}19]. Clarification [MHO^{+}11]. clarify [MKG10]. clarity [HTT^{+}12]. class [DC15, HNK^{+}18, YPWP18]. classes [HLS17, JBL^{+}16, RUPT^{+}11]. classification [GMBL11, MRLD19, TAEB^{+}11, YChH11]. clathratus [GS18]. clavata [LD10, TMB^{+}12]. cleaner [NFN^{+}15]. cleaning [NVA^{+}19, SPHK17, Tut17]. cleaning-station [SPHK17]. clear [PRM17]. clearance [KBPD10, SDT15]. client [NFN^{+}15]. client-fish [NFN^{+}15]. Climate [DKB^{+}12, GBS^{+}10, HHS12c, HWVM19, KTNO19, LS10a, MBW10, MGGSP15, NdBKR13, OCDD16, RFMT12, SES12, WL19, WPB^{+}14, BHP12, BRJ^{+}12, BSG^{+}10, BNAN12, BER17, BFDM18, BJS^{+}11, BDVP19, BWN^{+}14, BANE^{+}13, CDMW12, CBMK10, CD16a, CFPF^{+}14, CMHH17, CHFS12, DMP12, DED^{+}15, DFB^{+}17, DH17, FAK^{+}12, FWB^{+}15, FGH^{+}13, GL12, GTL^{+}16, GKR^{+}15, GWK^{+}12, GCR16, HRM^{+}11, HCL^{+}11b, HA15, HSA^{+}15, IMS^{+}11, JCM10, KHW^{+}10, KAP^{+}12, KCB18, KPW^{+}13, KOLDP17, KTP^{+}18, KMW17, KPr^{+}18, LB15a, LM12, LPC11b, LE13, LBG17, LNK^{+}16, LD17b, LFN^{+}12, MBB11, MPM^{+}14, MHF^{+}19, MMA18, MTB^{+}17, MBCB^{+}19, NGP^{+}15, NTB^{+}17, OSPWB14, OVF11, OCE^{+}12, OKLM14, PNI14, PMB^{+}17, PSP^{+}19, PSS^{+}12, PLJ^{+}12, PPM11, Pör12, RU11, RPE^{+}13, REG^{+}11, RGdLB^{+}11, RBHR19, RWH14, SHS^{+}17, SKM12, SKB15, SSB^{+}18, SKFH18, SA17, SNS12]. climate [STK12, SK12, TSS^{+}12, TvdMM^{+}17, VCL^{+}15, VMN^{+}12, VM14, WNG^{+}19, WNP^{+}14, YFC^{+}13, ZHJ^{+}17]. Climate-associated [MGGSP15, RU11]. Climate-driven [GBS^{+}10, MBW10, LNK^{+}16]. climate-induced [LPC11b, MMA18, TvdMM^{+}17]. climate-mediated [SSB^{+}18]. climate-related [OSPWB14, Pör12]. Climatic [AOPMOS10, SHAS15, WI12, HT10, PHG^{+}14, SSB^{+}16]. cline [Det10]. clines [BCC^{+}11, MKII19]. clingfishes [FG10]. clinically [MBB^{+}12b]. Clinocardium [LD17a]. Clione [BNE^{+}19]. clonal [RM19, TVRG^{+}13]. clone [BL14, JRK13]. cloning [ARD^{+}19]. close [GSL17, KMO19, RA10]. closed [ALS18, CPP10, MHFN10, WTGG10]. closed-mouth [CPP10]. closely [BBBNT13, BJH13, GMG^{+}16, GTL^{+}16, HRAM13, JCA13, MVL^{+}16, YSS^{+}18]. closer [JTWJ^{+}13]. closest [LNF^{+}11]. closure [MHC11, SHB^{+}13a]. closures [BS19a, KWW^{+}14, MH12b, McC14, MSF^{+}19, RvOP12]. clothing [WTGC^{+}11]. clove [RSV10]. clown [SBB18]. clownfish [RSH^{+}19]. Clua [BM10b]. clues [PEB^{+}14]. Clupea [BSST13, BPM^{+}18a, CBMK10, EOMS15, FESPR14, LFJ^{+}14, PSÓ^{+}15]. cluster [PBH^{+}17]. clustering [LPC19]. clusters [CGH^{+}16]. clutch [CCE^{+}14]. Cnemidocarpa [LS15]. Cnidaria [BLHL15, IMU17, PCD^{+}18, PMH^{+}13, WFB16]. cnidarian [PHD^{+}14, TDKS10]. cnidarian-dinoflagellate [TDKS10]. cnidarians [PC19]. Co [BA13a, lLChTC11, RLH^{+}13, VB19b, AVB13, BADH^{+}12, BADQ^{+}15, CTvdM^{+}08, FBNJ13, GGMTGC18, HLS18, HFH10, KBW^{+}15, LFG^{+}12, RLO18, RBH12, SGL^{+}17, SHSK11, SMR14, TG10, WF18, dlSBB^{+}10, BRF11, BPM^{+}18b, CG16, CH15, DMW^{+}16, DDP13b, ENW^{+}14, FRC13a, FCR14, FRRÁSL11, GSA^{+}18, GMG^{+}16, GCP^{+}17, HP13, HSB12, KJW^{+}15, LERS13, LMH12, MKS18, MDM18, PC12, PMS15, RMC^{+}14, RT17, SGS^{+}17, SLIGG18, SSX^{+}17, TG12a, TZF11, TFÁS^{+}12, TDM16, TBL16, YKG^{+}15, dlSGS17, PWO^{+}16]. co-acting [SMR14]. Co-existence [BA13a]. co-management [WF18]. Co-occurrence [lLChTC11, AVB13, CTvdM^{+}08, FBNJ13]. Co-occurring [RLH^{+}13, VB19b, BADH^{+}12, BADQ^{+}15, GGMTGC18, HLS18, RLO18, RBH12, SGL^{+}17, SHSK11, TG10, dlSBB^{+}10]. co-occurs [HFH10]. co-predictability [LFG^{+}12]. co-varying [KBW^{+}15]. coarse [LMLB12]. coarse-scale [LMLB12]. Coast [WRR10, ZBP18, AFH11, BHE^{+}17, BS16, BD11b, BMF^{+}12, CWW^{+}16a, CD19, CH12, CD16b, DHGW15, ETO^{+}10, FNC^{+}14, FFS^{+}15, FHH^{+}17, GHGF12, GSP^{+}15, GBH12, GBL^{+}15, IEM16, KO12, KWW^{+}14, KFW^{+}18, KBB^{+}10a, LAB10, LAM11, MKK^{+}11, MF10, NKM^{+}15, NKS^{+}15, NCS^{+}10, NGP^{+}15, NCK^{+}14, OBA^{+}18, PBCA10, RDF^{+}17, RTCF^{+}17, SAL^{+}12, SHM^{+}12, SHH10, SWW11, TBDM11, WY15, WPRS18, WNA^{+}15, XS19, YYW14, ZTNS13, TSTP13a]. Coastal [BSK^{+}11, HLIK16, LC18, SBBB^{+}18, SOC13, WHD^{+}14, AMSdL^{+}18, ASS^{+}13, AHØ^{+}15, AFOVRB14, APC14a, APC^{+}14b, AFMR^{+}13, AAGGFG^{+}12, AFF^{+}15, BW16b, BW16a, BVS^{+}16, BPP^{+}18, BTS^{+}14, BdSC^{+}14, BJG^{+}11, BCdV13, BMW^{+}14, BADQ^{+}15, BPSK11, BCB^{+}17, BGS^{+}16, BSS18, BHTF11, BGM^{+}17, CDD^{+}16, CRG^{+}16, CSS15, CAB^{+}19a, CL10a, CL15b, CHST16, CLGC16, CMAJ12, CCA^{+}14, CGM^{+}19, DMS14, DBS14, DDB^{+}16, DLL^{+}18a, DQC^{+}13, DEM15, DHM^{+}14, DDC^{+}15, EH13, EW16, EOMS15, EPF^{+}17, FJC10, FGOS16, FRRÁSL11, FF12, FTB14, FBPCC^{+}19, FSW10, GELJ19, GMB^{+}16, GSJS13, GBS^{+}10, GA11, GW15, GMD16, HEMJ18, HPL10, HGS14b, HBM^{+}15, HBB^{+}18, HW11a, HHP^{+}17, HSdBS10, HWE^{+}12, HRPBLT18, HTB^{+}15, HS16, HGA^{+}16b, HGA^{+}16c, HGW^{+}11, HP14, HORRLS^{+}14, JGM^{+}16, KNIN17, KdEPW18, KCP^{+}11, KEH10, KMM15, KRD^{+}12]. coastal [KHD^{+}11, LBAK13, LSJ^{+}15, LME^{+}19, LGD10, LLCM^{+}12, LCS^{+}10, LSD^{+}13, LNP12, LANER15, LCL^{+}13, LRM^{+}14, MdSW^{+}16, MCPC19, MTV^{+}11, MGFÁS^{+}10, MRGG^{+}13, MOK^{+}11, MDS^{+}13, MVOPB^{+}16, MMGB14, NRF11, NLKZ12, NMG^{+}13, Nob11, áNGGS11, NLGE18, ORW15, OAGV11, OMH12, OMM^{+}12, OMZ10b, OBÁSV18, PDGL17, PHB^{+}11a, PSHG16, PDB15, PLA^{+}12, PPB^{+}17, PHAZ18, PPS^{+}13, PKS11b, PW16, PB10, PGVF14, PBH^{+}17, PJMA18, RST^{+}10, RM13, dGCF^{+}17, RSTV^{+}13, REG^{+}11, RWM^{+}12, RZS12, RBEGG13, RG14, RB17, ROKS14, RPM^{+}13b, RPM12, RGB^{+}10, RSW13, RBD^{+}16, RFEJK11, RG11, SWFQ12, SCB15, SMND16, SDKD13, SHPS^{+}16, SM10a, SS10b, SJL^{+}16, SCRGP11, SLL12, SSX^{+}17, SFMB10, SMF^{+}11, SD17b, SMF10, SOBNL13, TLB15, TLS^{+}19, TKW^{+}14, TFFPM18, TZF11, TLFS14, TMGCM11, THMF^{+}15, THRBL^{+}16, TSCvS^{+}15, TP16, TDM16, TSA^{+}18, TTC11]. coastal [TTW^{+}12, TBMdV12, TSF^{+}10, TNN11, UCK^{+}13, UTS^{+}18, VARO12, VAHL11, VBD^{+}11, VB19a, VHV13, VFMdW12, VJJ^{+}19, VMN^{+}12, VC14, VLM14, VMC^{+}19, WRW^{+}16a, WAM^{+}15, WSF^{+}12, WRFM16, WOWG10, WS11, WWG^{+}13b, YMMO14, vdPLdV^{+}11, vdSRT^{+}18]. coastal-offshore [CAB^{+}19a]. coastally [FWF^{+}10]. coastline [CS11a, KCH^{+}11, vdMPM12, vdMPRM12]. coasts [BSRDR10, MLSK^{+}15, RAA^{+}16, RMKR13, SKN^{+}17, TNIH12]. cobble [JH11]. coccinea [RTJ^{+}13]. coccolithophore [CPTL11, EM13, Fie14, FSY^{+}10a, LBM^{+}18, RC12, SLIGG18, WDL^{+}18]. coccolithophores [BP10]. coccolithophorid [TGI14]. coccoliths [PYBB11]. Coccolithus [DPY^{+}16]. Cocconeis [JMPZ10]. Cochlodinium [GG16, RTC^{+}14, TG10]. cockle [BD15, dML11, dMBC12]. cockles [LD17a]. coconut [SY10]. Cocos [BMEL11]. Cod [EBDB18, ALNB17, ACG12, BSN^{+}18, BPWF14, BBS^{+}17, BLM10, BKAS14, CL10c, DSG15, ETO^{+}10, FWPP12, FOJF15, FQ11, GTM14, HMH^{+}12, HDB^{+}10, HFR^{+}13, HEM^{+}12, HBP^{+}16, HPH^{+}18, HMF^{+}18, HHB^{+}14, JLM^{+}12, JBS12, KBB^{+}10a, LDM10, LE13, LNP12, LOES11, MFG^{+}13, MRE14, MFD12, MBC^{+}11, MMNR^{+}12, MRC10b, NOB^{+}19, NJJ^{+}11, OVF11, OMSP10, PLA^{+}12, PHHL^{+}11, RAN^{+}10, RJTS12, RFRD10, SGM^{+}14, SNK^{+}10, SMK^{+}10, SBH^{+}18, SJB^{+}15, SNT^{+}16, TSGW10, TPT^{+}12, TDPS13, Urb12, VCFF13, WBW^{+}14, WWAL13, WRDF12, WBJ12, WFP15, TYG12]. Codium [DMJ11, KS12b, LSV10b]. Coefficient [TLB15, MVTS^{+}12]. coefficients [RT12]. coelestis [LDAE12]. coelurea [FWS^{+}18]. coeruleoalba [AMLR^{+}17]. coexist [LGJH12]. Coexistence [OSE13, AVB13, ES12a, MdNW12, VNK^{+}12]. coexisting [ABN14, KDS^{+}12, KH16, RLNLAF17]. cogradient [BBC12b]. coherence [SKPS^{+}16]. coherent [KBR^{+}11]. cohesiveness [CDD^{+}16]. Coho [SFKQ15]. cohort [GHGG11, HT11a, RRB13, ZSSK11]. cohorts [WDM^{+}18]. coincide [PBM^{+}10]. coincident [BCM11b, PH16]. coincides [HCJ^{+}18]. cold [ÅCAC16, AY10, CLGCA^{+}19, DIG^{+}11, DRG^{+}12, DB14, DJL^{+}12, GSC14, GDE17, KN12, LKC^{+}15, LLv13, Mat10b, OKJMC15, OFPR^{+}11, RAB^{+}15, TRMC11, TKL^{+}11, TOR^{+}10, VHM16]. cold-seep [AY10]. cold-water [CLGCA^{+}19, DIG^{+}11, DJL^{+}12, GSC14, GDE17, KN12, LKC^{+}15, LLv13, Mat10b, OFPR^{+}11, RAB^{+}15, TRMC11, VHM16]. coldwater [MBW^{+}19]. colias [WBJ12]. collaboration [JMK^{+}18]. collapse [BMA^{+}19, CDMW12, FPBL16, GLR^{+}14, HLIK16]. collapsed [FDS14, SB15b]. collapses [MATM^{+}19]. Collar [ORE^{+}17]. collected [CMV^{+}11]. collecting [RSV10]. collection [TKBN10]. Collective [JROS17]. collectors [RWR^{+}15]. Colombia [CGKW10]. Colomesus [GKW10]. colonial [DS16, JMS^{+}15, KLSJ10, RMKR13, WBG^{+}14]. colonies [ADL^{+}18, AGHC10, BPG^{+}13, CE14, DAC^{+}12, KEH10, LBVH13, MAP^{+}12, OGA13, SDM^{+}13, SMR^{+}11b, SBO^{+}18, WLB^{+}15]. colonisation [MBB^{+}12a].

Colonization [GLM16, GCCM18, NBGGST^{+}15, HFLI11, KAAI15, Par15, VHG^{+}18]. colonized [SZ13]. colony [ABA16, Ano18a, BBW^{+}17, BPG^{+}13, BGM^{+}17, DWLH13, DKF^{+}15, DAC^{+}12, JTWJ^{+}13, KBW^{+}11, MDLM17, MC12a, MWC18, MMRDS18, NWHD15, PM16, RSTM14, SRK^{+}18, TMH^{+}13, VM14, WBG^{+}14, WHN^{+}18, YTO^{+}11]. colony-based [Ano18a]. Colorado [PPM11]. coloration [HPL12]. colossal [CBR19]. colour [Bur12, TDPS13, dG14]. colours [BMSS14]. Columbia [FTB14, HLS17, BCH^{+}11, BWRP14, HPB^{+}18, HHP^{+}17, HPMZ11, KEBB10, RWP^{+}14, TTPM12, WOWG10]. columella [GBU15]. column [ALNB17, EOA15, LCB10, OFM^{+}18, THM^{+}13, TKB^{+}13, VB19a, WJ12, ZFC^{+}13]. comber [MPA^{+}10]. Combination [SKS18]. combine [WK16, YLPG19]. Combined [AAIB15, Ano10a, Ano10b, Ano10c, Ano12a, Ano12b, Ano13a, Ano13b, Ano14a, CMB^{+}14, CSC13b, FC12, LRTLF12, ORWPR14, OD15, ROLA15, ZHTA13, AMR^{+}10, CCA^{+}17, DB11a, FMJA^{+}18, FAWBF10, HWSK19, LD10, PMM^{+}15]. combines [Ano18a]. Combining [AAE14, GSSM^{+}11, dDTC^{+}17, LYH^{+}17, PBA12, PMM^{+}11, SMHB14]. combtooth [HLN^{+}19]. comes [TP19]. comfort [LAK^{+}19]. commander [HEAI10]. Comment [AF14, BBRM14, BD11a, BSBR^{+}12, BM10b, Car12, CGK13a, CSS^{+}13, DL13, FLC^{+}14, Fry13b, GBH12, HS11a, HG13, HBCK13, INK^{+}17, LGL11, MPM^{+}18, RL18, SB15a, SWP^{+}13, SL15a, TW10, War14, POL^{+}11]. Commercial [SBLM14, AHA^{+}15, ETB13, MAL15, RJR18, SGGL15, TKBN10, WPF^{+}16]. commercially [BFDM18, BRLM17, MDKM13, TvdMM^{+}17]. Commerson [GDP^{+}11]. Common [BBH^{+}12, DMK11, BHA^{+}11, BMEL11, BPZ^{+}14, BMPH14, BKAM12, CT10, CWW^{+}16a, CP11, CD14, DDP13b, ERB^{+}13b, ERB^{+}13a, EKOÅ13, FCR12, GCR^{+}17, GRFSM12, Goy14, GNBM11, GDH^{+}19, HG10, HCA11, HMW^{+}12, KDB^{+}10, LBM^{+}18, MCUO12, MHHF13, ORWPR14, PBB^{+}15, RDHM10, SGC^{+}13, SHUS17, SKGS11, SMSL17, SRB^{+}17, TSTP13a, TBC^{+}17, UW18, WMPM13, YYW14]. commonly [TBMG17]. Communal [KHF^{+}11]. Communicating [SOJ14]. communities [AS15, ANJ^{+}11, AGTS^{+}12, ALS18, BMSB13, BCM^{+}14, BvHC^{+}15, BA13b, BFC14, BSB^{+}11b, BPB11, BDF^{+}11, BMVAS19, BPF15, BVD^{+}14, Bur12, CRRD^{+}13, CLB^{+}18, CPS^{+}14, COK10, CVM^{+}17, DZH^{+}11, DMW^{+}16, DP11, DFLC17, DBCS16, DGP^{+}19, DNO17, DGQ10, DDC^{+}15, DHJ10, EPL^{+}14, EM13, EMG^{+}15, EFR12, EBK^{+}12, FJC10, FFLS16, FT10, FAWBF10, FFS^{+}13, GHES18, GYRJ15, GBTS11, GSD10, GMB^{+}16, GKBEF18, GSJS13, GA11, GGG18, GSM^{+}11b, GGvW^{+}15, GMR^{+}19, GWM^{+}16, Gre10, GBW^{+}12, HHI^{+}14, Hv14, HHH^{+}11c, HORRLS^{+}14, JOFP18, JCA13, JREVB18, JGS14, JT10b, JC15, JDG^{+}17, KO12, KSF18, KHÁB^{+}17, KKMM16, KCD13, KW12, KMM15, LFMC11, LJ10, LJK^{+}11, LCL^{+}13, MBvR12, MHB14, MKPS13, MAT^{+}11, MMG^{+}16, McC19, MWD10, MLC14, MYS^{+}16, MHNL14, MD16a, MOH^{+}10, MLSK^{+}15, NSH^{+}15, NWB^{+}19, NCR^{+}14, NGP^{+}15, NNT^{+}10]. communities [OCS14, OMZ10a, OMW^{+}10, PK13, PBS10, PMIG19, PMH^{+}13, POH16, PSL13, PKBS14, PCLB18, PGOG12, PPTK19, QGM^{+}14, RMHP14, RP13, RUD^{+}16, RHH^{+}15, RNK^{+}11, RSP17, RDC^{+}12, RBS16, RAB^{+}15, RMB18, SOB16, SLS11, SCST17, SLS18, SSB10, SLL^{+}18, SdBAH^{+}14, SGEO13, SPP17, SSM^{+}16b, SSvD^{+}15, SO15, TRLS^{+}11, TS12, TCA^{+}11, TNSN11, TB14b, TPSS12, VVB19, VCL^{+}15, VDEO18, WKX^{+}19, WGD^{+}19, WADKT13, WGD10, YYD^{+}15, ZFZ16, ZKS17, vDBH^{+}15]. Community [BMC^{+}15, BP12b, BPM^{+}18b, DLRF15, DRG^{+}12, HVBG^{+}19, JDQS11, LKCH12, OMD11, PBV^{+}11, PPB^{+}17, RSW13, SHR^{+}11, SDAM11, TYVR^{+}15, TNN11, WHN^{+}18, ÅCAC16, AAAE17, AFLM12, ÁMLUN16, AAGGFG^{+}12, AGM^{+}11, ANRT10, BVS^{+}16, BBTW14, BSB13, BDPL19, BD12, BWH^{+}19, BC15, BGM^{+}18, BLM^{+}18, BAL^{+}11, BOS16, BBGL13, BOCB10, BWN^{+}14, BCM19, CRGT19, CLD^{+}13, CLB^{+}14, CLS19, CR16, CR19, CL15a, CPTL11, CPR^{+}16, CL10b, CBGDS12, CFBG14, CHLR16, DPS16, DM11, DBCV^{+}12, DS10, DSJNH18, DPO15, Edm13, Edm14, EL16, ES12a, EM13, ES12b, ENO^{+}11, ETD^{+}18, EBdlTC12, FH10, FGW^{+}15, FMJA^{+}18, FSR17, FHE^{+}15, FAZ^{+}14, FF10, FKRJ13, GEAOH17, GHGF12, GL12, Gar13, GFGGDSM12, GMT^{+}11, GMBL11, GJ10, GM10, GLY^{+}14, GFW^{+}14, HCAV10, HHH^{+}11b, HCGP13, HTCH^{+}18, HT10, HC13, HDWF10, HRÁB^{+}14, HT15, HTR^{+}19b]. community [HOA14, HGKB14, HD10, JMB^{+}17, JBS11, JCZ^{+}12, KAD^{+}15, KHD19, KSBM^{+}11, KKS^{+}14, KTR^{+}16, KMW17, KFGSMPG14, KNB^{+}13, LMD12a, LHL11, LLRC^{+}12, LCG^{+}17, LLH^{+}11, LLCM^{+}12, LRTLF12, LFG^{+}12, ŁCM^{+}18, LTCH10, LGT^{+}14, LEB^{+}19, MJ17, MMA18, MRGG^{+}13, MLB19, MSS^{+}14a, MHS14, MRL^{+}10, MSM^{+}12, MH11c, MCHC^{+}16, MER14, MCT16, Mus12, NBGGST^{+}15, NC11, NLGE18, OE11, OTK^{+}13, OMW^{+}10, PDGR^{+}13, PSG^{+}13, PTL16, PMOAC^{+}17, PPEH^{+}14, PFR11, QHG^{+}19, QBW18, RWM^{+}14, RFM^{+}15, RNFG10, RRH^{+}16, RMBM14, RSK^{+}14, RCG^{+}15, RLH^{+}13, RFRD10, SK13, SCW^{+}15a, SPJ^{+}12, SRR15, SML^{+}15, SMK^{+}18, SHB^{+}13a, SLIGG18, SS10b, SRER16, SS11b, SSX^{+}17, SRRS12, TFFPM18, TNM^{+}11a, TBKG10, TMGCM11, TRL^{+}15, TDM16, TBDM11, TCKY19, TDSS15, TPS^{+}13, VH11, VWH18, WS17, WMB^{+}14, WAC^{+}14, WHH12, WK16, XTW^{+}15, YSKH11, ZTNS13, ZIT^{+}17]. community [dJTH^{+}14, dML11]. community-based [CBGDS12]. community-habitat [CLB^{+}14]. Community-level [BPM^{+}18b, OMD11, SDAM11, GHGF12, Mus12]. community-managed [LCG^{+}17]. Community-wide [HVBG^{+}19, WHN^{+}18, BDPL19]. company [HSWS16]. comparable [AKF14, RBSL11, SJHH^{+}17]. Comparative [ADF^{+}18, BPN11, BDM^{+}14, BLM^{+}12, CBK^{+}12, DBHMG12a, DTL^{+}15, FZI^{+}16, HEM^{+}12, HA15, LCB^{+}12, MLDG12, PRNR^{+}18, RBSL11, WHI^{+}18, WRP^{+}17, WFH11, BCPC18, BBH^{+}12, EAB^{+}19, HLS^{+}15b, JPF12, KPB16, TAB^{+}14, ZXF^{+}16]. compared [LEB14a, LOES11, MCUO12]. Comparing [BGB^{+}17, BFF^{+}10, CBYH13, LAM11, LGM^{+}12, MSF^{+}19, ROE17].

Comparison [ABK^{+}14, ABF^{+}14, BBB^{+}15, BLT14, CR11, DBS^{+}16, DWL19, FMH^{+}15, GUC^{+}15, GRBH15, JRK13, KOLDP17, KHG^{+}19, MRC^{+}13, PPML17, RBD^{+}16, WCP^{+}14, YTM^{+}15, BDS18a, BCOMKJ18, BSC^{+}16, BPG^{+}13, CPP10, CS10a, DKAC^{+}12, DHH10, FPC^{+}13, JOFP18, JBCGC19, JMGG13, KBW^{+}11, LAH^{+}15, LLF^{+}18, ML11, MEKC15, MPM^{+}14, NIS14, PL10, SNS^{+}13, TS10, TP16, TPFP11, WWB^{+}17]. Comparisons [HSDB13, KB12, ZBP18, CFBG14, MMDM12, SCB^{+}16, SMF^{+}17, TT16]. compensation [MBC18, MPD11, SPP12]. compensatory [HM17]. competent [CMF^{+}19, MC11, MWHM18, MD16b, OFFQGG10, WHA^{+}13]. Competition [CGOC11, HSB12, JSK19, ODGL^{+}12, AHPKS17, BWV^{+}12, BSST13, CSB14, ESBG14, ES12a, EES15, FGRM12, FB14, GFL^{+}17, GRFSM12, ICUM10, JD18, KWS^{+}13, Kin16, LGMB^{+}16, LR15, LGJH12, MdNW12, MLP^{+}13, NRF11, NGSF13, NGJHB12, OS18, PP18, RSE18, SWMK15, SLRM^{+}15, SMK^{+}10, TMM15, WPB^{+}14, dMA18]. Competitive [CLW^{+}11, GSS13, GCP^{+}17, MBBBC12, MH19, BMMS10, DP11, FGP17, KHF^{+}11, MCUO12, PBH^{+}17]. competitors [KP18, RNB^{+}12]. complement [SOA^{+}17]. complementarity [SLRM^{+}15]. Complementary [AGP^{+}13, AL12b]. Complete [Ano17a]. completely [Chi11]. Complex [CTMU10, ESBG14, HBP^{+}16, MBvO^{+}18, TCN^{+}10, ADF^{+}18, AHØ^{+}15, BSP^{+}10, BEFC15, BBKP10, DVV^{+}18, DJL^{+}12, GCL^{+}16, HGC15, HOA14, JWT^{+}14, JCF^{+}12, KWM^{+}17, KDJP10, KGJ^{+}16, KBK^{+}10, KFB15, MAD18, PPB^{+}17, SJO16, SP12, TTC11, YTM^{+}15, vdWHG11]. complexation [SPSP10]. Complexities [SFMB10]. Complexity [BSG^{+}10, AM13, ADO^{+}12, CMK^{+}14, DAM^{+}17, FSR17, GSU12, GNBM11, HW13, KSF18, LEB^{+}19, MSG11, MD16a, NJP12, PCS^{+}15, PGVF14, TSGW10, TDSS15, TPVCV19, WKRW^{+}12]. component [DP16, RUPT^{+}11, YMPR13]. components [FBS14, HRÁB^{+}14, SJB^{+}15]. Composition [GWT18, MVVGAV10, MMFF12, RHH^{+}15, RCG^{+}15, AFMR^{+}13, BCGWS19, BCdV13, BHT15, BHF^{+}16, BD11a, Bol17, BRA^{+}10, BKHW13, BP12b, BLM^{+}18, CR16, CWD11, CPTL11, CWW^{+}16b, CDMB^{+}18, CLGC16, CMAJ12, CPR^{+}14, CMW^{+}15, CMC^{+}15, DEJN10, DE17, FSB15, FRGI15, FMJA^{+}18, FSR17, GGL^{+}17, GOO14, GME^{+}19, GNC^{+}10, HNK^{+}18, HWET13, HPN^{+}17, HC13, HWB^{+}18, HBG^{+}17, HPLS13, IMD11, JUMS10, JMB^{+}17, KO12, KAD^{+}15, KSK07, KJW^{+}15, KHT^{+}17, KTKM14, KLSJ10, LBN11, LSJ17, LLv13, LMD12a, LFM^{+}16, LBL^{+}12, LBE^{+}10, MHGW13, MWD10, MHHF13, MSS^{+}14a, MAS^{+}10b, MFS14, MSM^{+}12, NWC^{+}10, PSCS16, PB18, PWNP^{+}11, QSB16, RLG14, RDC^{+}12, RLH^{+}13, RFRD10, SK13, SPJ^{+}12, SSM^{+}17, SNW^{+}18, SS10b, SWY^{+}10, SST19, TMGCM11, TDM16, TMP^{+}11, TFM^{+}18, TPCN12, UDF^{+}18, UFP^{+}10, WKvW10, WS13, WGD^{+}19, XS19, YSC^{+}15]. composition [YSKH11, ZIT^{+}17, vDBH^{+}15]. compositional [PBAP^{+}14, SS11b]. compositions [dDTC^{+}17]. Compound [BLPP16, BKC^{+}19, DBC^{+}12a, GKHM13, OBI^{+}16]. Compound-specific [BLPP16, BKC^{+}19, DBC^{+}12a, GKHM13, OBI^{+}16]. compounded [OSPWB15]. compounds [EMG^{+}15, JMPZ10, PSPK10]. comprehensive [dlSOV^{+}16]. compression [CR14]. compromised [PRB^{+}17]. Compton [BD11a]. con [MB11]. con- [MB11]. concentrating [YKG^{+}15]. concentration [ACA^{+}10b, ACA^{+}10a, BW13, BHAC17, CSR^{+}14, CPHH15, KKO^{+}14, PMS15, SYX^{+}10, TFÁS^{+}12, WYI^{+}16]. concentrations [CAB^{+}19a, CSC^{+}11, CG16, DMW^{+}16, GSA^{+}18, GGF^{+}14, KM10b, SGS^{+}17, VPBV13]. concept [HBS17, PBWM18, TSS11b]. conceptual [FBNJ13]. concerns [GJ13, MLASTS13]. conch [BMEL11, BBdG^{+}19, DH18b, KCS^{+}17, SD17b, SDB12]. conchilega [DvV^{+}16]. Concholepas [MC11, MGO^{+}12, MJO^{+}13, MJT^{+}14]. conclusions [DSG^{+}13]. concomitant [MH13]. concurrent [HFR^{+}13]. Condition [RFH16, BDNA10, BSC17, BMPH14, BMF^{+}15, CRRL17, CVC^{+}18, CRS^{+}18, CTS^{+}15, CLN^{+}13, DEP19, FA10a, HLL^{+}17, HTR^{+}19b, JGWJ^{+}11, LGB17, LWG14, LRQ^{+}12, LRSB19, MD11, MBP^{+}12, MRC10b, NGJHB12, NSC^{+}19, ØLH^{+}13, OLF11, PPH11, PDCH15, PL10, QHG^{+}19, RB18, RJR18, RSTM14, RGB^{+}10, RT17, RVH^{+}19, STK^{+}13, SF16, SFTG14, SDT15, SST^{+}18, Tho15, TKIO10, WLB^{+}15, WA19, YAL12, dMBC12]. Conditional [NNT^{+}10]. conditioning [LCQT^{+}13]. conditions [ASR^{+}19, AFRAG16, ASEB11, BBM^{+}14, BNAN12, BCK15, BBC^{+}10, BJRH10, CED^{+}19, CPP10, CSC13a, CLW^{+}11, CDMB^{+}18, DBA17, DWLH13, DE17, DCW^{+}10, DWL19, EHR12, EMMSL13, ECZD14, FRMS17, dlÁGRB^{+}19, GVGP^{+}14, GHT15b, HER^{+}19, HBG^{+}17, KEBB10, KSF18, KMV18, KMSW10, LST11, LLv13, LMD12a, LPC19, LE10, LBE^{+}10, LMG12, LMZ16, MMSC17, MRL^{+}17, MMB^{+}19, MH11c, MSK^{+}19, NMCP17, NMHB16, OBÁSV18, PDL^{+}10, PNPF14, PBAP^{+}14, PPEH^{+}14, RAA^{+}16, RB18, RCS^{+}17, RMME10, RNB^{+}16, RRB13, ST10a, SCL^{+}19, TYG12, TvdMM^{+}17, TPCN12, VJM^{+}16, WL11, YBD^{+}16, ZHTA13, ZVL^{+}16]. coney [BJCB^{+}14]. confamilial [CR19]. confidence [YChH11]. configuration [Col10a, RDPR15]. confirms [KKMA^{+}10]. conflict [OALK14]. conflicts [OGH18]. Confluences [TMSG19]. confound [DSG^{+}13]. confounds [DL13]. congener [MH19]. Congeneric [MH11c, CPPM12, KA16, Kin16, MM13b, WFH11]. congeners [BB15, PPS^{+}13]. conger [PSdA^{+}17]. connecting [TIS^{+}15]. connections [EPF^{+}17, LCBK^{+}13]. Connectivity [FWL15, JWBS10, RSTV^{+}13, RWH14, SAZHAAHS14, SLO^{+}17, BHUCMO19, BJH13, BOÁNF19, BAFB12, BWCZ11, BBBHC17, CAR^{+}19, CCNB15, CCKB^{+}15, CCG^{+}19, CNSW16, CGG^{+}10b, DSH14, DGDJ14, DMS14, DKV^{+}16, DMC^{+}16, EHD^{+}17, FTO^{+}15, FCV^{+}16, GPA^{+}16, GSU12, HTH10, HS15b, KPW^{+}13, KHN15, KCLP16, KPP^{+}16, LFU^{+}19, LMRRB17, LMW^{+}13, LGBSC11, LL18, LBC^{+}19, MOP^{+}15, MMWG13a, MMWG13b, MKK^{+}11, MS14, NSS10, NDD^{+}13, NJJ^{+}11, OLD^{+}15, OBA^{+}18, OCPM12, ORW15, PHD^{+}13, RBSL11, ROKS14, RSHM10, RMV13, SCO^{+}12a, ST11a, SDC^{+}16, SWH^{+}18, SKW^{+}14, SMMFH16, SMVM^{+}12, TSCvS^{+}15, WWAS18, WMS^{+}10, WWZ^{+}19]. consequence [BDLW10, CWB^{+}15]. Consequences [KPW^{+}13, PGR^{+}17, TvdMM^{+}17, AH13, Ano17c, BCV^{+}18, BL13, BBWG12, CPSC^{+}18, CFSM^{+}13, CCDP15, DDDR10, EPS11, HG17, JDG^{+}17, KdEPW18, KHS10a, LBP^{+}16, MFSRR14, MMCW14, MWMD17, OBM15, PLGSD17, PHTS11, RGEA^{+}15, RK15, SS13, TKBN10, TMM15, WBG^{+}14]. Conservation [MHSG18, RAKB^{+}12, BBD^{+}14, BJ14, BWH^{+}17, CM14, DFDC18, FEA^{+}15, GLW^{+}12, GLJ^{+}18, GJ13, HFG^{+}19, Kat16, LKC^{+}15, LCF^{+}14, LKL^{+}17, LBT^{+}10, MLASTS13, MLCC19, MJ17, MCW^{+}17, NLKZ12, PvdMLC17, PW19a, QBW18, SCB15, SDST16, SJL^{+}16, SHT^{+}16, SS15b, SS15a, TF12, TDB^{+}14, WTdB16, WMPM13]. conserve [SDSA14]. conserved [BWH^{+}19]. conserves [RLBC10]. Consideration [YSKH11]. consistencies [She12]. Consistency [CMC^{+}16, HPM14, TPT^{+}12, BPSG11, CPF^{+}14, HBS17, MBvR12, MJR^{+}14, RDGC17, TP19, WS13]. Consistent [KHK^{+}17, PTP^{+}19, CLMB13a, GLHM^{+}18, LPI^{+}14, RSTV^{+}13, VW10, vBMH^{+}17]. consistently [Tho15]. conspecific [GMM10, WRW^{+}16b, vdMPM^{+}10]. conspecifics [AB13]. conspicuous [HCG15]. constitutive [GB16]. Constrained [HRP18, GCFW14]. Constraining [YPWP18]. constrains [BPG^{+}11, LRRM^{+}11, SWBMJ17]. constraint [CG11]. constraints [BC16, CALB10, PGVF14]. constricta [PZL^{+}17]. constructed [DPM15]. Constructing [Cha13]. construction [BHD^{+}13, BDD^{+}18, CAB^{+}16, DTC^{+}17, NGA18]. construction-related [CAB^{+}16]. Consumer [FSH^{+}16, GNC^{+}10, DMH15, VW10]. consumer-diet [VW10]. Consumers [BPF15, ES12b, DMC^{+}16, LA12a, MAK^{+}15, ODCE11, OHR^{+}13, PH11, WRW^{+}16a]. consumes [RCB12]. Consumption [DD15, RB17, Sti13b, BB11, BS11, BFL^{+}19, CLJF12, GAC11, HFM^{+}15, LS15, MP12, MPBSB18, PBB^{+}15, SGGL15, SEFS15, TG12b, VGR^{+}16]. consumptive [DGA^{+}15, GG12]. Contact [BSD17, CSJP17, QYS^{+}12, TMM15]. contagiously [CKV12]. contain [DDK^{+}18]. contaminant [MJ10]. contaminated [CFC^{+}12, Mar17, NSG^{+}12]. contamination [BFS10, HSC^{+}15, MCL^{+}17, NKM^{+}15, VLK^{+}15]. Contemporary [BCC^{+}11, GVG^{+}18, AKF14, JSK19, KGJ^{+}16]. content [CMB^{+}14, DHG16, DTVH10, dSKFV12, HTSA16, KSA^{+}13, LBE^{+}10, LSV10b, MA13a, MGAV15, PBAP^{+}14, PMF^{+}17, RSS^{+}16, RPLT10, RCR^{+}13, WHB^{+}19]. Contents [Ano17b, BCM^{+}14, DWPH11, FC10, GCS^{+}19, IS16, MCW11, NLK^{+}19, VBT^{+}17]. contests [WFP15]. Context [CHB11, HC13, SP13b, SST^{+}18, AFF^{+}15, BB17, BGO^{+}18, CRJ13, CFH11, GAO14, HOL^{+}17, LDS^{+}15, PSPK10, SAMW14, TGP^{+}16, WNP^{+}14, YAL12]. Context-dependent [CHB11, HC13, SP13b, SST^{+}18, AFF^{+}15, PSPK10, TGP^{+}16]. Continental [BVP11, HTB^{+}15, LN10, PFF^{+}17, BRMT11, BSD^{+}14, CAR^{+}19, CL15a, DKRS12, FSS^{+}18, GSD12, GFW^{+}14, HKL12, HGFA^{+}11, KAD^{+}15, KPKA13, LRB^{+}12a, LPKT^{+}19, MSS^{+}14b, NHSG11, PPML17, SS11a, TMBP13, TMP^{+}13, VVB19]. Continental-scale [HTB^{+}15]. continental-shelf [GFW^{+}14]. contingent [BGO^{+}18, GFS15]. continued [BMA^{+}19]. continues [Kin10a]. Continuous [HTK^{+}17, PLJ^{+}12, SKK12, ENW^{+}14, NLV^{+}15]. continuum [BJH13, War14]. Contracaecum [SBH^{+}18]. contrary [SJL13]. contrast [DKV^{+}16]. Contrasting [BEFC15, BVS^{+}11, DAM^{+}18, EMMSL13, GTL^{+}16, HTC17, Jon11, KÖHSO12, KHD^{+}11, PTP^{+}15, PÁIIS^{+}19, RNB^{+}16, SKM^{+}15, vdHTT^{+}15, AFRAG16, AGM^{+}11, CRG^{+}16, CCG^{+}10, CBYH13, CLN^{+}13, DZB^{+}18, GMT^{+}15, GGPDS19, IAV^{+}10, Lec18, MSS^{+}14a, MSPO12, MCHC^{+}16, NMHB16, RVH^{+}19, SATS13, SRCE^{+}10, TBMdV12, WBLRN18, YRM^{+}12]. contrasts [CDCS12]. contribute [Ano17c, Col10b, NMHB16, TLK^{+}12]. contributes [DM17, HPMB11]. contributing [HPMZ11, MN10]. Contribution [GEAOH17, GC13a, MBS^{+}13, NRK^{+}19, BDP10, GHD^{+}10, KNIN17, LR18, LS16, MGM10, RTD14, SK10, WTGG10, YYU^{+}17, dJTH^{+}14]. Contributions [MVTS^{+}12, MLCC19, STSS^{+}16, SOBNL13, ALSO^{+}18, BFRM13, GLJ^{+}18, KRD^{+}12, LKC^{+}17, LMP12, PHWC13, RSTV^{+}13, SDC^{+}16]. Control [MPZ10, BWH11, BCH^{+}11, BW10, BAW11, BB14c, BCM19, CGR^{+}15, CGS^{+}19, CPHH15, DMS14, FS14, FHJ15, FSH^{+}16, Gar13, HKS^{+}12, HBHG10, HOE^{+}12, JT10a, LH14, LTP^{+}18, LC12, LRSB19, MSGGE10, MH13, RMHP14, RDK14, RSK^{+}14, RDC^{+}12, SDMB15, SBB^{+}12a, SLKN14, SBMÅ12, WHT12]. controlled [DHF^{+}13, KAP^{+}14, NKA16, SSAB^{+}15, dSM11b]. controlling [AY10, BW16b, BW16a, BC16, BCHG18, CEW15, DFS13, GMSL^{+}14, WWZ^{+}19]. Controls [JM17, BWM^{+}15, CSRD11, CWGG15, GJSR18, GS16, GG16, HMKF11, HWE^{+}12, HvEBC12, KK17, LCTD^{+}15, MMEM14, MCHC^{+}16, PGWS10, RC12, SWR11, SCRGP11, SLNG18]. conurbation [CPR^{+}14]. convergence [ANJ^{+}11, GOO14, RNK^{+}11]. convergent [SRCE^{+}10]. conversion [MWD10, vDCR13]. conversions [DBH^{+}16]. Coordinated [BBG12, JP15]. Coos [PRR^{+}16, SS11a]. cope [BDNA10]. Copepod [PPEH^{+}14, AMLS15, ABM12, BTS10, BRB11a, BLER15, BAT14, BBRBG10, CWD11, CLC12a, DTS14, GSM^{+}11a, GvSGCK14, GK10, Hal15, JCA13, JSLG17, KK17, KFGSMPG14, LME^{+}19, MBR^{+}18, OSC19, OKLM14, PSS^{+}12, PFL13, PSER17, PFPP18, PMBH11, RSH^{+}19, RHRB^{+}17, SNTN15, SM10b, SNW^{+}18, SK11, SIK16, SLNG18, SRRS12, SMR14, SHB^{+}10b, TB14a, VLK^{+}15, vDJF15]. Copepoda [ACA^{+}10b, ACDJ11, ACA^{+}10a, CWD11, DKL^{+}18, OHST19, PPJT15]. copepods [ANJ^{+}11, AGC12, BC16, CMW^{+}15, FHS^{+}13, GMR13a, GMR13b, HPL^{+}16, HT15, HBCK13, Kiø16, KDS^{+}12, KBPD10, KDB^{+}10, LSY^{+}18, PTF^{+}14, PDBB14, RGCT^{+}10, TWW^{+}15, YOST11]. Coping [HW14]. copper [OHP18, RP17, SPSP10]. copulatory [PVYS16]. coquimbensis [BVQ12]. Coral [Ber12a, BMVAS19, CPR^{+}14, DAM^{+}17, FG16a, GSM^{+}11b, GWM^{+}15, KMPS15, MHC11, McC19, NJP12, PM16, SJB16a, SPJ^{+}12, WSK^{+}16, APC^{+}13, ABRP15, Alb15, ARD^{+}19, ACL^{+}16, ASM10, AHB^{+}14, BMÁN^{+}18, BWS^{+}12, BHA^{+}11, BVM^{+}19, BMM17, BWV^{+}12, BBB^{+}14, BAFP^{+}13, BCSH18, BA13b, BTTK10, BJMS12, BJH13, BPH^{+}19, BS14b, BBG^{+}11, BSB^{+}11b, BLPP16, BDF^{+}11, BAE^{+}10, BSD17, BJL^{+}17, BFL^{+}19, BTM^{+}18, BK14, BAvW12, CTRBLC^{+}13, CED^{+}19, CSJP17, CSM12, CSABH^{+}13, CCM^{+}15, CSB14, CMC10b, CESC14, CHG^{+}18, CM10, CMGS19, CVM^{+}17, CBK^{+}12, DIG^{+}11, DPO^{+}17, Dav19, DWK13, DSVM10, DBCS16, DDD^{+}11, DQM^{+}19, DMM^{+}10, DDP13b, DHAV13, DBTV13, DWV^{+}12, DABP10, DJL^{+}12, ELA10, Edm13, Edm14, EAB^{+}19, EGH^{+}15, ESD16, ETB^{+}11, EMJ18, FTWA17, Fog12, FJD13, FF16, GJ16, Gea10, GSS13, GSC14, GEA^{+}14, GB10a, GGvW^{+}15, GAC11]. coral [GNBM11, GRN14, HSPS14, HBI^{+}18, HMKF11, HHI^{+}14, HLS15a, HSC^{+}15, HAF^{+}17, HPN11, HRC12, Hoe10, HSR^{+}13, HPM14, HM13, HHH^{+}11c, HYC^{+}14, HDP^{+}17, HYL^{+}13, HNG^{+}11, HMM14, IHP10, INvdB^{+}13, JGJ15, JWR^{+}10, JHNL12, JDB^{+}10, JBD10, Joh10, JRK13, JWBS10, JJNP12, JMK^{+}18, KCLLL13, Kin16, KLT^{+}12, KMGH16, KBK^{+}15, KMO19, KRC^{+}11, KBB14, KBB16, KCL^{+}11, KLMR14, LFMC11, LJ10, LBH18, LLv13, LFPC^{+}14, LWP^{+}13, LB11, LP10, LCM^{+}16, LBJK14, LS19, LKN^{+}13, LBT^{+}10, LPC11b, LRH14, MBJ16, MBB^{+}18, MOP^{+}15, MTS^{+}17, MHGW13, MH12b, MM16, McC17, MJ17, MMM^{+}15, MMRDS18, MHVV13, MBW^{+}19, MMCW14, MZF^{+}19, MSLR13, MTK^{+}18, MLP^{+}13, NHS^{+}17, NTN^{+}13, NAY^{+}11, NMS^{+}18, NAG^{+}16, NSH^{+}15, NvOH13, OPSM13, OCPM12, OLF11, ORWPR14, OPW^{+}11, OFPR^{+}11, PGG12, PMK^{+}15, Par15, PKW^{+}11, PWM18, PMB^{+}10, PA13]. coral [PCH^{+}14, PWCvO^{+}14, PBM^{+}18, PML^{+}19, PPDW16, PPTK19, RMB^{+}14, RSJ14, RET^{+}19, RMME10, RRH^{+}16, RTJ^{+}13, RBC^{+}15, RdGv^{+}18, RSV10, RZPH14, RKvW10, RLM^{+}14, RAB^{+}15, RDF^{+}19, RMRA15, SPSG14, SAZHAAHS14, SL11a, SSB10, SM17a, SGEO13, SWBMJ17, SWB^{+}10, SAM14, Sti13b, SKS18, SKFY18, SBO^{+}18, TPB^{+}18, TBC^{+}17, TBL16, TPFP11, TNS^{+}12, TB14b, TYVR^{+}15, Tut17, VRMKH^{+}15, VDS11, VDG^{+}15, VHG^{+}18, VMWU15, WKvW10, WH14, WGD10, WNNH10, WLP^{+}10, WLHW10, YPWP18, YAL12, YSKH11, vWSGL11]. coral- [DHAV13]. coral-associated [DBCS16, SWBMJ17, WH14]. coral-derived [BTTK10, RdGv^{+}18]. coral-dwelling [PM16]. coral-feeding [BSD17]. coral-macroalgal [MLP^{+}13]. coral-population [RKvW10]. coral-reef [Alb15, BLPP16, HMKF11, Kin16, Tut17]. coralgrouper [MHS15]. coralline [AS15, ASK^{+}15, BMMS10, BT14, BHT15, BRF11, CPHH15, DDP^{+}13a, DDP13b, KOLDP17, LKDP17, MAS10a, VDS11]. corallines [WNP^{+}14]. Corallium [Bru10, SB10]. corallivorous [FBGH18, NHPW18]. corallivory [KMO19, LE10]. Corals [NAY^{+}11, AHS^{+}13, AP14, AZSKWL11, BÁNC^{+}18, BMSB13, BVM^{+}19, BHE^{+}17, BCvH^{+}15, BKB11, BPLT14, BFKR17, Car12, CLGCA^{+}19, CLMB13a, CLMB13b, CP11, CBJB19, CvOB18, DFLC17, EES15, ELM^{+}19, FHL^{+}17, FGRM12, GKE^{+}19, Goc10, GSM^{+}11b, GDE17, HAF^{+}17, HCA11, HGPM10, HB15b, ILDL10, KHA^{+}12, LKC^{+}15, LH14, LGW^{+}15, LCC12, LP12, LHW^{+}19, MWP10, McC12, MTK^{+}18, NKS^{+}15, PRBR11, QBW18, RRL^{+}19, RWAP16, RFJ^{+}10, RPGP^{+}15, SMOH19, SCLB11, SPH^{+}13, SAH11, SYK^{+}13, TRR^{+}17, TGM^{+}17, TTF^{+}11, TRMC11, TPA^{+}13, TOR^{+}10, VPDPMRC11, WSK^{+}16, WBB13, WA14, ZRBV15, ZAZS12]. Corbula [GSTK11, MS13]. Cordell [YIKG10]. Cordgrass [FNA^{+}16, CAD^{+}10, HBHG10, MKMZ11, PCA^{+}15, QCX^{+}10]. cordiformis [BB14b]. Core [JWT^{+}14, DKV^{+}16, NvOH13]. coregonids [VHV13]. corer [BEEH15]. cores [LPKT^{+}19]. coriiceps [EBOM11]. Coris [BMSS14]. Cork [TLM^{+}11]. corkscrew [OC17]. Cormorant [GYRJ15, MCW^{+}17, WBS^{+}12]. Cormorant-induced [GYRJ15]. cormorants [VSG^{+}15]. corrected [BAL14a]. Correction [BS18b, UDF^{+}18]. correctly [AAAE17]. correlate [BANE^{+}13, CAB^{+}19a, CS11a]. correlated [CRRD^{+}13, KS13, MMC^{+}16, SBO^{+}18, dlSBO^{+}12]. correlates [CTS^{+}15, DDD^{+}11, DCB^{+}16, HWR^{+}13, HCA11, LRHP18, MPLLE17, RBH12, TMP^{+}11, VGL16, dJdBdJB12]. Correlation [FK10, SOC13, GBH^{+}10, KTH^{+}18, RAB^{+}15]. correlations [DHK^{+}17, FWPP12, PMKL11]. correlative [BD14]. correspond [WOHJ^{+}15]. corresponds [CUCS17]. corridors [DPO^{+}17]. Corrigendum [Ano12c, Ano19b, Ano19a, Ano19c, Ano19d, Ano21]. corrugata [BRB11b, CHB^{+}13]. Corticosterone [CMK^{+}14, SMC^{+}14, EMTB^{+}12, FRMS17, GAB^{+}14]. cortisol [MLB^{+}14]. Cory [DGC12, PXG^{+}10]. Coryphaenoides [DDK^{+}18, KJBS12, LTN^{+}11]. coseismic [CMC10a]. cospecialization [TOW10]. cost [MPM^{+}18, ORG19, OMW^{+}16, SPM^{+}16, SBT15]. Costa [HCBV^{+}11]. costly [MD16b]. costs [BWP^{+}11, BPLT14, FDVL18, KMO19]. Coulter [RPS12]. count [BG12, ZB16]. count-based [ZB16]. Counter [MKII19, RPS12]. Counter-directional [MKII19]. countergradient [BBC12b]. counting [HHG10]. counts [RPS12]. coupled [ARB^{+}12, BST^{+}10, FHRC16, GWH^{+}18]. Coupling [DM11, GSF18, SGMG^{+}14, CMV^{+}11, FSE^{+}17, GTP^{+}17, HBS^{+}13, KFBM19, MDE13, PMML12, PMS10b, SGP13, SBM^{+}18, THMF^{+}15]. covariance [ASK^{+}15, AHSG16, LAB10, LBMZ15, VBR18]. Cove [WVD^{+}17]. cover [AFMR^{+}13, BCdV13, BMVAS19, BNW14, CPR^{+}14, HSH12, JFK12, MBB^{+}18, RSO^{+}13, TAEB^{+}11]. covered [AFRAG16, Ano18b, SKD^{+}17b, SDMR19]. Covering [BSA^{+}16]. cownose [OBA^{+}18]. COZIGAM [YBB^{+}12]. Crab [AED^{+}10, SCTA10, BHUCMO19, BAW11, BL17, BBMM14, CSG11, CMF^{+}19, CFA16, CGD10, CD16a, CMO^{+}10, CFMN16, CCG^{+}19, DMMW12, DDDR10, DDM^{+}16, DT15, DAD^{+}11, DMHS^{+}14, FRC^{+}13b, FPN17, GFC^{+}13, GBL^{+}15, GPM^{+}13, HG17, HW12a, HJ16, HW13, HBHG10, JL12, JGG^{+}15, JDD^{+}13, KCR^{+}10, LAB10, MMDM12, MMG^{+}16, MEK11, MHS17, MP12, MRH11, MBT14, MS10, NdBKR13, ORJH19, PDB15, PES12, QCX^{+}10, RWM18, RSO^{+}13, RS14, RWR^{+}15, RLBC10, SY10, SHP^{+}14, SHC^{+}10, SE11b, SFNP11, SN17, TB10, TG12b, WAP10, WRB^{+}15, WMH17, WNA^{+}15, WPK10, YPK15, dGR11]. crab-bivalve [WPK10]. Crab-tiling [SCTA10]. crabeater [BKC^{+}19, HBK^{+}12]. crabs [AAIB15, BSB14, BSA^{+}16, CAD^{+}10, CY10, CRS^{+}18, DD15, FFS14, GME^{+}19, GVGH15, HN13, HTC17, JE10, JMB^{+}18, KCCG12, KLN^{+}13, KLNJ15, LHB19, MGP12, MP17, ORR^{+}14, OFFQGG10, POX^{+}17, RLM^{+}14, SZ13, SHF12a, SPH^{+}13, TB12, WGD^{+}16, WJWC16, WRW^{+}16b]. Crangon [CBC^{+}10, HT11a, HT11b, MASMH19, TGRH17]. crash [TFS14]. crassa [NPLT10]. crassirostris [JSLG17]. Crassostrea [ALV12, ALK17, ATP14, BG19, BIL^{+}10, BPHW16, BSCH14, CB11, CL15c, DWL19, FAM^{+}17, GLF^{+}17, GFL^{+}17, HPA17b, HR15, KBW^{+}15, KW10, KPP^{+}16, LdU^{+}17, MB16, MWHM18, NCS^{+}10, PFL18, RHF19, SCM13, SFW^{+}16, VTSH^{+}19, WDH^{+}12, WHA^{+}13, WLH^{+}17, WR12]. create [MSG11, MHQ17]. created [CSS^{+}17]. creates [ZOM^{+}10]. creating [KMV18]. Creatures [LHHG11]. credits [BMV^{+}18]. creek [SO16]. creeks [CR11, GKW10, KCE14, RBD^{+}16, RMB18, SSPKF16]. Crepidula [BVQ12, DP13, DBCP15, LC13, MC12b, PDC^{+}17, VBD^{+}11, dMA18]. Crepipatella [CMP^{+}13, PDC^{+}17, PCF^{+}19]. Crepuscular [RDHM10]. crested [MPG^{+}10]. crispata [LKN^{+}13]. Critical [LFFL15, BBC^{+}10, FWF^{+}10, KSFL14, KWM^{+}17, RPB17, SFTG14, WCRCR17]. Critically [DCB^{+}16, FKP^{+}15, TNB^{+}15, CME^{+}12, GLW^{+}12, LNF^{+}11, RABS19]. Crittercams [SSM^{+}19]. croaker [MW16b]. crocodile [WPNK12]. Crocodylus [WPNK12]. Crocosphaera [DB11a, RCP17]. Croix [BFN^{+}19, DH18b, OLF11]. crop [JM17, SWWK13]. Cross [WTdB16, DKF^{+}15, GTP^{+}17, HBM^{+}12, HLD12, JGCRS12, PS15a]. cross-boundary [HLD12]. cross-colony [DKF^{+}15]. cross-ecosystem [GTP^{+}17]. Cross-seasonal [WTdB16]. cross-sectional [JGCRS12]. cross-shelf [HBM^{+}12, PS15a]. Crossing [RAF^{+}14]. crowd [HSWS16]. crowding [BS19a, GPB^{+}15]. crown [ARD^{+}19, BBW^{+}19, KHH18, MPB^{+}14b, RPB17, WGLDB15]. crown-of-thorns [ARD^{+}19, BBW^{+}19, KHH18, MPB^{+}14b, RPB17, WGLDB15]. Croxall [TSTP13b]. Crozet [CXdG^{+}17]. crucial [SBH^{+}19]. Crude [HSC^{+}15]. Crustacea [KBB14, AGC12, BHP11, GGMGR12, HS14a]. Crustacean [HSS^{+}19, CLDC^{+}13, DdR10, DMP12, DVS13, KBB16, LRBSJ15, OT13, SHMOS17, SCK^{+}17, SDVL16, SAD^{+}16]. crustacean-feeding [KBB16]. crustaceans [AR14, BFPONS^{+}12, DKAC^{+}12, HSZR10, JPF12, KHS10b, NFSW10, Whi11]. Crustose [VDS11, DDP13b, LKDP17]. crusts [OSE13]. Cruz [GMSC14]. Cryptic [Ale13a, EL16, BCB11, DVV^{+}18, GB12, HDP^{+}17, LDAE12, MKK13, MPM^{+}18, ORM^{+}13, PFPP18, RB11]. cryptobenthic [GEAOH17, GGCB16]. cryptofauna [ETB^{+}11]. crysos [BBKP10]. CS [HHG10]. Ctenochaetus [GB10a]. Ctenophora [EBDB18]. ctenophore [BFH^{+}13a, DTM^{+}12, LGB^{+}13, MBR^{+}13, MLC14, RG14, SM15, TLG14]. ctenophores [HTHH11, PDBB14, SLKN14]. Cuban [KCLP16]. Cubozoa [SWK18]. cubs [FKA^{+}12]. cucumber [CLP18, EBdlTC12, ETB13, RP13, SK10, SGD^{+}12]. cucumbers [LCCT19]. cue [LEB14b, VE18, ZFZ16]. cues [BTTK10, BAT14, CGH^{+}16, DF15, DL14, DAM^{+}18, INvdB^{+}13, LWP^{+}13, LCCT19, Mat10a, PW19b, SA14, UCK^{+}13, WRW^{+}16b, vERRP16]. cuescape [SAWK18]. Culebra [PVZD^{+}11]. cultivated [FBC^{+}15]. culturally [EPC^{+}19]. Culture [LFE^{+}14, AFSL^{+}15, FAZ^{+}14, HHH^{+}11c, JSS^{+}11, KTR^{+}16, STT14]. cultured [EGH^{+}15, HP14, NAY^{+}11, PRNR^{+}18, WFP15]. cultures [ENW^{+}14]. culturing [VBŽ^{+}17]. Cumulative [MBB^{+}18, VDEO18, dJTH^{+}14, AvBBL^{+}16, FRJ10, FRJ11, GKRG^{+}19, SLSR10]. cuning [APC^{+}13]. cup [TFW^{+}12]. cuprea [Ber12b]. Current [ASEB11, HM09, ICT^{+}11, LBJ^{+}16, LFL11, PBC^{+}18, RHRB^{+}17, SMVC^{+}15, SOC13, TKA10, WBS^{+}12, BRG^{+}14, CPP10, CRJ13, CCGP14, EHG16, GWK^{+}12, HBM^{+}12, LSH^{+}19b, LHB^{+}15, NCR^{+}14, OFM^{+}18, PP10, PBA12, PG15, TBG15, BFF^{+}10, BO15, BBR^{+}14, CLH^{+}15, lLChTC11, DABP13, DSGR^{+}15, DPO15, FHRC16, FSS^{+}18, Gla10, HKB^{+}19b, KCW^{+}17, KGLLW11, KMW17, LLS^{+}18, LBE^{+}10, LMH12, MBRO10, MBS^{+}13, OMT^{+}17, SHF^{+}12b, TL18, TAB^{+}14, Tho15, WM10, WCG^{+}14, WJW^{+}18, WZP^{+}14, YPK15]. current-induced [EHG16]. currently [MKM^{+}16]. currents [CC19, DRT13, DPH13, FPL^{+}12, HLW^{+}16, HG10, LFRF13, PBH^{+}17, UW18]. curryfish [ETB13]. curtains [DTC^{+}17]. curves [VHBG15]. Cuvier [WWB^{+}17, NWP^{+}11]. Cyanea [HFN14]. cyanobacteria [SCRGP11, WL10, DB11a]. cyanobacterial [BDPD14, WRGP12]. cyanobacterium [HvEBC12, PSPR18, RCP17]. cyb [DSH14]. cycle [AOW11, BGL^{+}11, CPP10, DDDR10, DMvB10, HLS^{+}15b, OKLM14, OMD11, RNB^{+}16, RJTS12, SP12, TGRH17, TDB^{+}14, ZdVA^{+}16, dlGMD^{+}17, vdPLdV^{+}11]. cycles [Chi11, CE13, LKC^{+}17, MF15, MFH^{+}11, NTN^{+}13, OAB^{+}19, RB11, RQN^{+}14, SBB18, SRB^{+}17, vdWHG11]. Cyclical [RBHR19]. cycling [BRK^{+}12, BST^{+}10, BCK15, BT13, DTG16, DWM15, DBDK14, HCAV10, KNIN17, KBW^{+}15, LNH^{+}17, LTG17, RHF19, SGTP16, SDMR19, SMJP^{+}15, SLNG18, TLG14, VKH10]. Cyclone [WC18, FRV^{+}13, FA10b, GDH^{+}19]. Cyclone-driven [WC18]. cyclones [MBB^{+}18]. cyclopoid [GK10]. Cyclopoida [ACA^{+}10b, ACA^{+}10a]. cygnus [KBG^{+}13]. cylindracea [GP11a]. Cymodocea [BVL10, OD15, SSL^{+}14, dlSGS17]. Cynoscion [BT13, CLB19, LTG17, LWG14, VWEP18, WLBB^{+}13]. cyphonautes [SM10a]. cyprids [DFPH11]. cypris [SM17b]. cysts [CIF11, RZH^{+}16b, SBKA14]. cytometry [BOS16, MRSD15].

D [LKN^{+}13]. Daily [ASS^{+}13, MBV^{+}11, MMEM14, NPMS11, QLCUG15, SGT^{+}14, WBLRN18]. Damage [DdRH14, BBW^{+}19, RGCT^{+}10, VPDPMRC11]. damicornis [TKS^{+}14, TH11]. dampen [RSW13]. dampens [JGS14]. damselfish [EBB10, LDAE12, NJP12, VDG^{+}15]. damselfishes [CSB14, EMJ18, GFKC13, Goc10]. danger [GGF^{+}12b]. dangerous [KBGA19, LRHP18, VPBV13]. Danish [BDBN11, DBCV^{+}12, LRSB19, PAF^{+}10]. d’Arguin [CCA^{+}14, vdGvdLvG^{+}19]. dark [DP16, ROE17]. Dartfish [HED^{+}13]. Darwin [PBV^{+}11]. Dasyatis [DWPH11, TLAT13]. data [AMAJ^{+}17, BLIM11, BFF^{+}10, BG12, BD13, BD14, BS18b, BS18a, BBC14, BRB11b, CGK13a, CMV^{+}11, CSS^{+}13, CFBG14, DBHMG12a, DCHI12, DBS^{+}16, FTKK14, FH18, FGH^{+}13, GTM14, GKRG^{+}19, GMES18, HSH12, HW12a, HMB^{+}12, IME12, JGCRS12, KW17, KLBT^{+}16, KM17, LAM11, LCF^{+}14, LFL11, LMSC^{+}19, LAH^{+}15, LCB^{+}16, MPKO11, MSG^{+}13, MLS^{+}14, MO13, OALK14, ORW15, OTK^{+}13, ORE^{+}17, PBM^{+}11, PMM^{+}11, PLJ^{+}12, PMKL11, RB11, SKD^{+}11a, SKR^{+}13, SJLH12, TSTP13b, TW10, TRA^{+}14, WLP^{+}10, WWAL13, ZBC10, ZKS17, vDJF15]. data-driven [SJLH12]. data-limited [GKRG^{+}19]. data-poor [GMES18]. data-sparse [CGK13a]. datasets [BMTC^{+}15, BLT14]. date [MS18b, RB18]. Davis [Ano14c]. davisae [ACA^{+}10b, ACA^{+}10a]. Dawson [NWD13]. day [DGC12, LPI^{+}14]. daylength [VKY^{+}12]. days [HBG^{+}17]. daytime [ADHP12]. dead [CSM12, HKB^{+}19b, HCL10, SBCA15]. dealing [IKS^{+}12]. death [BLD^{+}16a, GKBEF18, HBHS^{+}18, LWD13, LAALH10, SCHO11b]. DEB [SPM^{+}13, SvdMK^{+}12]. debris [CD13]. Debunking [WEF19]. Decadal [DMPG11, Edm13, KCLP16, MLR12, ODK^{+}14, PH16, PTS^{+}18, RDF^{+}19, SLSC16, BADH^{+}12, BJS^{+}11, ORW15, SMD^{+}13, Sti13a, SLR^{+}14]. Decadal-scale [Edm13, MLR12, Sti13a]. decade [BVAB^{+}17, CMC^{+}16, CPG^{+}14, CQR^{+}17, KHW^{+}19, MEKC15, MGB^{+}19, SDC^{+}16, vWSGL11]. decades [ALS18, BEP^{+}12, CW12, FFS^{+}13, HMAC^{+}11, HBG^{+}17, JS12b, KCS^{+}17, RSP^{+}16, SSS19]. Decapod [BdSC^{+}14, AR14, HEMJ18, JPF12, LA14, SCK^{+}17]. Decapoda [KHS10b]. decapods [BW12]. decipiens [SD17a]. decision [SSFB11]. decision-making [SSFB11]. decisions [PW19b]. decline [BMÁN^{+}18, BNAN12, Bru10, CBKM16, CCP^{+}14, DRDV15, FEE^{+}11, HMAC^{+}11, JBD10, KPW^{+}19, KHW^{+}12, MMG^{+}16, MHFN10, MDF10, MTB^{+}17, MS18a, MGB^{+}19, PH16, PSC^{+}13a, PML^{+}19, SB10, WHN^{+}18]. declines [BCSH18, BCM11b, ESD16, GJSR18, HBI^{+}18, KMM15, MGP12, STS^{+}18, SDB12]. Declining [CB11, Har16, CMHH17, KGLLW11, SPM^{+}16, dDTA^{+}18]. Decoding [TMP^{+}11]. decomposition [JNN12, KCB18, NOB12]. deconstructive [DM11]. décor [KBS16]. Decorating [HJ16]. decorator [HJ16, KBS16]. decouple [KF13b]. Decoupled [LSY11]. decoupling [BPM^{+}18a, LHS^{+}14]. decrease [ATO^{+}14, GGMTGC18, MUOS13]. Decreased [MKS18, VMWU15, LMZ16, MWHM18, VJM^{+}16]. decreases [BAE^{+}10, CPF^{+}13, GHES18, NRF11, OSK15, RB17, SLR^{+}14, SJL13, ZGH^{+}17]. decreasing [VKY^{+}12]. decussatus [FRRÁSL11]. deduce [GHT15b]. Deep [CATAV18, FKRJ13, GGF^{+}12b, HP13, LRM^{+}14, WWKS15, AJF^{+}12, AAM^{+}12, BMC^{+}15, BCG10, BWS^{+}12, BPB11, BSA^{+}16, DIG^{+}11, DRT13, DRG^{+}12, DDK^{+}18, DBCS16, DGFH10, DBH^{+}16, FC10, FPC^{+}13, GWT18, GSD10, GFF16, GGFB15, GSMK15, HBE^{+}13, IVD^{+}10, JRW^{+}14, JMB^{+}17, JCA13, JGL12, JLPG17, KHA^{+}12, LSY11, LRB^{+}12b, LTN^{+}11, MBJ16, MWA17, MHPvdH19, MNKB11, MYS^{+}16, NHSG11, NTS^{+}18, NSY^{+}18, NOT^{+}11, Par15, PRBR11, PPTK19, QSvN^{+}11, RWMP12, RCB^{+}10, RDC^{+}12, REP18, SBB17, SNW^{+}18, SBB^{+}14, SDF^{+}19, SHM10b, TBMG17, TRLS^{+}11, TRHM14, TTF^{+}11, TKL^{+}11, WHB^{+}14, WFH11, YYU^{+}17, YSC^{+}15, YD11, YOST11]. deep-diving [AAM^{+}12, BCG10, TRHM14]. deep-layer [YOST11]. Deep-sea [CATAV18, FKRJ13, AJF^{+}12, BMC^{+}15, BWS^{+}12, DIG^{+}11, DRG^{+}12, DBCS16, DGFH10, DBH^{+}16, FC10, FPC^{+}13, GSD10, GFF16, GGFB15, HBE^{+}13, IVD^{+}10, JCA13, LRB^{+}12b, LTN^{+}11, MYS^{+}16, NSY^{+}18, NOT^{+}11, PRBR11, RWMP12, SBB^{+}14, SDF^{+}19, SHM10b, TBMG17, TTF^{+}11, WHB^{+}14, YSC^{+}15, YD11]. Deep-water [HP13, WWKS15, BPB11, BSA^{+}16, DDK^{+}18, JGL12, KHA^{+}12, PPTK19, TKL^{+}11]. deeper [Ano18b, BST^{+}10]. Deepwater [MDKM13, KJBS12, KGLLW11, PHH^{+}12, PBC^{+}12a, SKD17a, BMC^{+}15, BSF17, DBCS16, FCR^{+}15, GFF16, GPRR17, HGS14a, HGS14b, HGS15, JGG^{+}15, KCMH^{+}17, PGR^{+}17, PPCH17, RFH16, RBS^{+}17, SB15a, ZWP^{+}17]. defence [DH16]. defense [BDT^{+}16, FWW^{+}16, Goc10, HCM^{+}11, JL17, MPLLE17, MDPG13, SH10, WRO^{+}11]. defenses [AAvS^{+}13, Fis10, LP10, LTP13, LIT13, PW19b, SG18a, SWWK13, TOW10, TdRH17]. Defensive [BFG^{+}11b, HKJ17]. deficiency [GCBB13]. deficits [SHG^{+}15]. define [FSR17, HRT12, HBM^{+}12]. defined [BGS^{+}16]. defines [DTS14]. Defining [ECZD14, HM09]. definition [KPLD^{+}12, MHCS18]. definitive [TP16]. defoliation [FDS12]. deforestation [VGBJ^{+}13]. Degradation [RLG14, SRRS12, FL10, FDFS16, LWP^{+}13, SMR14, VQKF14, dJTH^{+}14]. dehydrogenase [FUOG^{+}16]. del [AHB^{+}14]. Delaware [RHH^{+}15, TCKY19, TT16]. Delayed [MATM^{+}19, MPG^{+}12, PZA^{+}13, SCP^{+}14, HTB^{+}17, LBD^{+}15, RWP^{+}14]. deletron [Bus12]. delicatula [POS14]. delineate [TTT^{+}15]. delineation [LJM^{+}12]. deliver [OHB^{+}15]. delivery [FDS12, SBE^{+}19]. Delphinapterus [HLPS^{+}15, HIH^{+}17, SWM^{+}19, SOC13]. delphinid [SMSB19]. Delphinus [PBB^{+}15, SRB^{+}17]. delphis [PBB^{+}15, SRB^{+}17]. delta [RMV13]. demand [BS11, BCGG19]. Demersal [KHÁB^{+}17, APNC^{+}15, ACW18, AFH11, BS16, BSB13, BBGL13, BL10, CBR^{+}18, CMLC12, CAGS17, Cra12, FC10, FOL12, KWW^{+}14, KCW^{+}17, KPKA13, LHDH17, LYH^{+}17, LLBP15, MVE^{+}10, MBF16, NM18, PNI14, PH16, SL11a, SCL13, TBDM11]. demersus [SUB^{+}13]. demissa [ZZG^{+}19]. democratica [HEB^{+}11, PNA^{+}16]. Demographic [Ebe10, KH16, PDT^{+}11, BBF^{+}19, CCDP15, DAM^{+}17, HRM^{+}11, ICT^{+}11, KLNJ15, MMRDS18, PHG^{+}14, RTM11, RSP^{+}16, SRLV10, WFH11]. demographically [RSE18]. Demographics [SCMSB^{+}11, MLASTS13, MEK11]. Demography [KNS^{+}15, BÁNC^{+}18, CGD10, DHS^{+}19, ESD16, FSP^{+}17, GBL^{+}15, HCBV^{+}11, IDT^{+}18, JAPK16, MM12, MLML^{+}19, PNF^{+}12, SKM12, SG12, SABMG16, TM10, vdHTT^{+}15]. demonstrate [KGJ^{+}16]. demonstrates [IHD^{+}16, RSS^{+}16]. demosponge [KRC17, LRSB19, PBDMRB13, SIHH10]. dendrochristos [HYL^{+}13]. Dendronephthya [CHG^{+}18]. Denitrification [ETAE13, HTB^{+}17, KCOP13, SLRP^{+}12, AM18, BMW^{+}14, BKC14, FBNJ13, GS16, POKGRP10, XLXS17]. Denmark [BDS^{+}18b, OGH18, WKE^{+}17]. dense [OKJMC15]. densiflora [PCA^{+}15]. densities [Ben16, CBT13, CSRF16, KWC^{+}17, KMGH16, LHJ11, LDT^{+}13, MG14, TSGW10, WMB^{+}14]. Density [BD15, CPSC^{+}18, CRGT19, DEP19, DWLH13, DBTV13, GJSR18, KP18, LK15, LGJH12, PMS10a, VBR^{+}14, WDH^{+}12, vdGvdLvG^{+}19, ANWG18, BVL10, Bea11b, BCK15, BPG^{+}10, BVS^{+}11, BDC10, BJRH10, BJS^{+}16, CPSC^{+}15, CGG10a, CEK^{+}16, CD14, DQM^{+}19, DH18b, DCC^{+}10, FJW14, FBPCC^{+}19, FHH^{+}17, FTB^{+}17, GMM10, GSL17, GLM^{+}19, IW15, JTH^{+}12, JE10, KW10, LH11, MCUO12, MHR19a, MMB^{+}19, OSK15, OKLM14, PHSBO16, PPB^{+}17, PPC^{+}19, PZA^{+}13, PUP16, RSO^{+}13, RSS^{+}16, REAN^{+}14, RHE14, RFRE19, RHvD19, RSA^{+}12, SSB^{+}18, SLT^{+}19, Tho15, UW18, VHBG15, WY15, WFH^{+}18, WHH12, dB10]. Density- [PMS10a]. Density-dependent [CPSC^{+}18, DEP19, DWLH13, KP18, LGJH12, WDH^{+}12, vdGvdLvG^{+}19, DCC^{+}10, IW15, KW10, OSK15]. Density-driven [GJSR18]. dentatus [CFA16, DTG16, SFNP11]. denticles [DNO17]. dentine [MBSG11]. deoxygenation [PSER17]. depend [AB17, BS12, GDPR18, LK15, TBS^{+}14, TVRG^{+}13, VDG^{+}15]. dependence [BD15, BDGB19, CCDP15, CEK^{+}16, GRCP^{+}10, HLW^{+}16, KRW18, LSRRCM13, NPLT10, PPB^{+}17, RHvD19, TSTP13b]. dependencies [Wir14]. dependency [BMJH12, KMP11, RH12]. Dependent [MHSG18, ÁMLUN16, ADvdM13, AFF^{+}15, BMSB13, BA12, BCS^{+}12, CPSC^{+}18, CRGT19, CCNB15, CHB11, DEP19, DWLH13, DLL^{+}18b, DCC^{+}10, DRH15, FSD^{+}14, GTP^{+}17, GGNL^{+}17, HC13, HSS^{+}10, IW15, KP18, KNY^{+}15, KLD18, KW10, KPB11, KdVvdP^{+}12, LFJ^{+}14, LOBW^{+}11, LKLC11, LE13, LGJH12, MFI11, MDF10, MO13, NSS10, OSK15, PPH11, PSS^{+}12, PSPK10, PCW^{+}12, RFBH^{+}11, RSE18, SRLV10, SP13b, SK11, SST^{+}18, TFP14, THP^{+}16, TGP^{+}16, WDH^{+}12, WBG10, WWW^{+}13, ZRBV15, vdGvdLvG^{+}19]. depends [GPW^{+}10, MOP^{+}15, MWB13]. depicts [AAAE17]. deplete [YMMO14]. depleted [BSY13, FG16b, HBI^{+}18, JCM13, SFL^{+}18, WYS^{+}10]. Depletion [CV11, GGB^{+}18, LDT^{+}13]. deployments [EMTB^{+}12]. deposit [PFR11, RP13, VDM^{+}10]. deposition [HA18, LS12, TPVCV19]. depredating [CRJ^{+}18]. depredation [GJ13, MMC^{+}18, MMCL19]. depressed [LCJ^{+}12]. depressions [HKL12]. depressus [WTP^{+}18]. deprivation [MEDS11, VCFF13]. Depth [BCGWS19, CME^{+}12, DEWK15, FKT15, JGJ15, MBJ16, MHS15, PHH^{+}12, AOW12, AMM^{+}12, BW16b, BW16a, BMÁN^{+}18, BPN11, BDS^{+}18b, BAE^{+}10, BJS^{+}16, CAB^{+}14, CBJB19, EKOÅ13, FWPP12, FMC^{+}11, FT10, GBR11, GBW^{+}12, HGL15, HJ13, HSWS16, JTTB^{+}18, KJCN^{+}11, KRDV18, KRdVG^{+}18, LH14, LHDH17, LLC^{+}14, LRBS15, LAH^{+}15, MM13c, NYK^{+}13, PCLB18, RB19, RL11, SBT12, SWBMJ17, SFKQ15, SSM^{+}16b, TWW^{+}15, VTB12, VGR^{+}16, WMB^{+}14, YYU^{+}17, ZRR^{+}13]. Depth-induced [BCGWS19]. depth-keeping [TWW^{+}15]. Depth-related [DEWK15]. depth-temperature [WMB^{+}14]. depths [FKRJ13, SL15a, WFH11]. derivation [Wir14]. derived [BTTK10, BBD^{+}14, BADQ^{+}15, HLD12, KMP15, LAM11, LNH^{+}17, MMFF12, NDD^{+}13, RdGv^{+}18, SOJ14, SBS^{+}13, TKV^{+}19, VSG^{+}11, YIKG10]. Dermal [DNO17]. describe [JDB^{+}10]. described [HTR^{+}12]. describes [WDM^{+}19]. describing [JGCRS12, WCP^{+}14]. description [AHR^{+}18, BBG^{+}11, CFA14]. Descriptive [ANWG18]. descriptors [NMCP17, TGSM12]. desertion [DDM^{+}16]. Desiccation [DBCP15, JG13, MAS10a]. design [ABH18, BWJB16, GRL^{+}14, GSF18, GS11, GHNF10, MOK^{+}11, NMS^{+}18, ST10b]. designation [OE11]. designing [BAFB12]. designs [BT16]. Desperate [MWHM18]. despite [ALK17, ARMA15, MAP^{+}12, OSE13, PCNM17, RBSL11, RA10]. Destructive [BS15, FNC^{+}14, FSLG12]. detached [PYBB11]. detail [GMBL11]. Detailed [RTM11]. details [ST11b]. detect [GSR10, JDB^{+}10, SGP^{+}17]. detectability [FBPCC^{+}19, IKS^{+}12]. detectable [ENW^{+}14]. detected [PK13, VK11, WHA^{+}13]. Detecting [BHE14, NWB^{+}19, RIM^{+}13, CBLM^{+}17, JTH^{+}12]. Detection [BSS18, MKK13, vBBW^{+}14, ATF12, ARGB13, BBM^{+}14, BWM^{+}15, BCS^{+}12, BFD^{+}16, DL13, EH13, GSÁCM^{+}16, MJM^{+}14, MOMF14, NWH^{+}12, NWD13, PCH^{+}14, WBS^{+}17]. detention [SCRGP11]. deter [BHD^{+}13, SWWK13]. determinant [FTO^{+}15, VM10]. Determinants [BAC^{+}11, SPH^{+}13, AED^{+}12, ETB^{+}11, EMJ18, MBJ16, MN18, PM16, PLF^{+}12b, RHW12, RGDS12]. Determinate [OFG^{+}18]. determination [HN13, HLC^{+}18, VUL^{+}11]. determine [CCE^{+}14, CVC^{+}18, GFC^{+}13, GSF18, HNG^{+}11, KCD13, LSJ17, LWG14, LPJ^{+}13, OKLM14, PBW^{+}11, PRP^{+}16, SGM^{+}14, SF14, TMS^{+}16, Wir12]. determined [AZVB13, BSB^{+}18, BPH^{+}19, DWPH11, FF10, JCM10, LRTLF12, MRSD15, MHW^{+}12, PWNP^{+}11, WFBM14]. determines [CRR^{+}19, CWW^{+}16b, FJ11, HMJ^{+}13, MMRDS18, RWG15, SKS18]. Determining [MS10, LCL^{+}13, MORM15, MLP19, MWZ^{+}16, MMWG13a, PH11, RGB^{+}11, RGP11, SMBH15]. determinism [KGJ^{+}16]. deterrence [SV15]. deterrency [NPRAGG^{+}12]. deterrent [GJ13]. deterrents [WFS10]. Detrital [KS11, KS12b]. detritivores [FESPR14, PMBH11]. detritus [CMP^{+}11, FDS12, FRP^{+}17, GGNL^{+}17, HS15b, KS12c, MHGW13, SMM^{+}15, SS13, TBKG10, YMPR13]. detritus-based [HS15b, SS13]. developed [BFF^{+}10, FLS^{+}18]. developer [SSBC11]. developers [FSB15]. Developing [BCW19, WFS10, FEA^{+}15, LMZ16, MO12, NC11, NCT^{+}14, PC19, SBT15]. Development [BBL11, GK10, MKM^{+}16, ACA^{+}10b, ACA^{+}10a, BAT14, BFL^{+}11, BPHW16, BCR15, CR19, CvOB18, DSM11a, DHJ12, DFB^{+}17, DWL19, EHR12, FW11, GCC^{+}16, GSF^{+}13, GJ10, HSH18, HWVM19, JKM11, LCR^{+}16, MGFW13, MJT^{+}14, MM13b, MMM^{+}15, MW16a, MGD^{+}11, NMC^{+}16, NSB^{+}19, OT13, OLF11, PSB17, PPTK19, RNB^{+}16, SKD^{+}11a, SHP^{+}14, SPJ^{+}12, SRR15, SW12, Sma13, SZG11, TRGF18, VSW^{+}15, WAP10, WBB13, YHT16, vdMHW^{+}10]. developmental [DZL16, KLT^{+}12, KDO^{+}14, MGPC13, RCE^{+}13]. device [BKTD10, RLBC10, SMBC18]. devices [EKOÅ13, FCFD15, GJ13, RDF^{+}13, RTCF^{+}17, VSG^{+}15]. devil [CNW^{+}12]. diacanthus [SBFP10]. Diadema [BWH^{+}17, BGK17, CLMM^{+}14, IHP10, KB14]. diadromous [SS12a]. Diamante [SJL^{+}15]. diamondback [HW12a, RLBC10]. diapause [Ji11]. diapausing [DTS14]. Diatom [ALANGT10, LLH^{+}11, CRG^{+}16, CAB^{+}19a, CRS^{+}18, DSK^{+}10, FHS^{+}13, GSA^{+}18, GHT^{+}15a, JMPZ10, MAB12b, MPBSB18, MDS^{+}13, MLP^{+}14, PL15, PSF^{+}19, POS14, PMF^{+}17, YMMO14]. diatom-dominated [CRG^{+}16]. diatom-sourced [CRS^{+}18]. diatoms [lLChTC11, GFR^{+}18, GMT^{+}11, HLL^{+}17, LGD10, LG18, PR11, VK11, XGLH14, YSK^{+}10, ZXF^{+}16]. Diazotroph [SBE^{+}19, DB11a]. Diazotrophic [LMR^{+}12, BMD^{+}11, PSPR18, RCP17, RWM^{+}14, SS10b, WL10]. diazotrophy [PDO13]. Dicentrarchus [VMN^{+}12]. dichotoma [ZAZS11]. dichotomy [ECM^{+}12, ZBB^{+}11]. Dictyota [PKW^{+}11, RMME10, YMCV10]. die-off [HFMD16]. dieffenbachii [JT10a]. Diego [DHB18]. Diel [ACRLB12, ALNB17, BLS^{+}11, CESC14, ETO^{+}10, HPN11, HSS17, MJHS11, NMS^{+}18, NPMS11, SBG^{+}16, SSHB16, YNK18, BQRB12, BT13, CGK13b, DTG16, DHM^{+}14, EOA15, FTS^{+}13, HTHM18, HCLB18, KBW^{+}15, LSY^{+}18, LTG17, MPA^{+}10, MG12, MGB^{+}14, PGWS10, RWR^{+}19, RDHM10, RCD^{+}13, SDKD13, SFKQ15]. diel-cycling [BT13, DTG16, KBW^{+}15, LTG17]. Diet [BBdG^{+}19, CJM^{+}13, CGM^{+}13, CWAS17, EPM^{+}16, HBK^{+}12, KHT^{+}17, KM14, LSJ17, MBB^{+}15, MMC^{+}16, MCW11, NFC^{+}11, NHG10, NIPS12, PCW^{+}12, RCR^{+}13, WHPT12, WWB^{+}17, AB11, ASR^{+}19, AGV^{+}11, ACAD12, ASCHM19, ACG^{+}11, AMLR^{+}17, BI15, BPJ^{+}15, BKAM12, CLD^{+}13, CP10, CSF13, CG10, CDMB^{+}18, CBL^{+}10c, DHG16, DMH15, EBDB18, GSP^{+}15, GE14, GCSF18, GWH^{+}18, GNHLN11, GVGH15, HHNI17, HRS^{+}16, HBG^{+}17, JW11, dDTC^{+}17, KF17, LCJ^{+}12, LMR^{+}10, LANE18, LMCD19, LRO^{+}17, MPH^{+}11, MMH^{+}17, MMFF12, MVP^{+}16, MHHF13, MBRO10, MBS^{+}14, MPRPM^{+}18, NMHB16, ODK^{+}14, PKS11a, PRNR^{+}18, PLF^{+}12a, PHWC13, PGOG12, RR18, RGL^{+}16, RKM^{+}10, SGC^{+}13, SKD^{+}17b, SHB^{+}10a, SGP^{+}17, SG12, TFS14, THB^{+}12, TGR^{+}19, UTS^{+}18, VW10, WAW^{+}12, WWAL13, ZO10, ZBK^{+}14, ZLCW15]. Diet- [CJM^{+}13]. Diet-dependent [PCW^{+}12]. Dietary [DGCH18, DMC^{+}17, GMT^{+}15, OTK^{+}19, PWNP^{+}11, RNB^{+}12, VH11, vdBUPP^{+}11, AMS^{+}13, Ano17c, BCV^{+}18, BLPP16, BLC^{+}16, CLC12a, CEC^{+}18, DPC13, HVJ^{+}18, JMB^{+}18, KS12a, KBP^{+}10, LS19, RSTM14, RT12, SNTN15, VCG^{+}15, WOHJ^{+}15]. Diets [CPID13, ACG12, ACS12, BBB^{+}15, BM16, HBK^{+}17, Hat13, KHI12, KLN^{+}13, MCT15, PLNE11, QCX^{+}10, RLG14, RVF14, RSTM14, SDR^{+}15, SCL13, STSS^{+}16, SCW^{+}15b, TRR^{+}17, WHN^{+}18, YFC^{+}13, ZLGH16]. differ [GHGG11, LIT13, MLSK^{+}15, NBM12, PWCvO^{+}14, WFP15]. Differences [ANRT10, PPM^{+}15, PDC^{+}17, WK10, WOHJ^{+}15, WS11, ABA16, BHN^{+}14, CPB^{+}11, CSC^{+}11, CDC18, CBF^{+}19, CS10a, DHS^{+}19, DFDC18, FRGI15, FPM14, HVJ^{+}18, HTR19a, HCL^{+}11b, HHB^{+}14, JNYT13, JD18, KCCG12, LLBT18, LKS12, Lyo14, MCBR18, MM13c, NMG^{+}13, OFT^{+}11, PBAS^{+}10, RBSL11, RSC14, RCVG10, SS19, SPBS14, SKD^{+}11b, SCW^{+}18, THMF^{+}15, VTSH^{+}19, VFMdW12, YTO^{+}11]. Different [CTRBLC^{+}13, CL17, HPL^{+}16, AFS19, ACG12, CIM^{+}13, CLJF12, CUCS17, DHJ12, DBCB11, DAD^{+}11, DAM^{+}18, DBS^{+}19, EOM^{+}14, FMH^{+}15, GRG^{+}12, GC13a, GGFB15, GMR^{+}19, HKL19, Hos14, dDTC^{+}17, KKPL14, KCD13, KPM^{+}11, LSV^{+}10a, LSQ^{+}13, LZF^{+}14, LDAE12, LJB14, MPKO11, MH11a, MHSG18, MFG16, OFFQGG10, OHR^{+}13, PXG^{+}10, PP13, RSJ14, SMK^{+}17, VBT14, WAP10, WHI^{+}18, WVD^{+}17, WS17, WFH11, WLB^{+}15]. Differential [ANB13, GPV17, GAF^{+}16, GB12, HPMB11, HMM14, ICUM10, LFN^{+}12, MGFÁS^{+}10, MH12b, OHP18, SDG^{+}12, ZLGH16, BBC19, BBC^{+}17, FRGI15, JKM11, LCFR10, PPH11, SGS^{+}19, VHG^{+}18]. differentially [BVS^{+}11, GSU12]. differentiated [LPB^{+}16]. differentiating [LMGJ13, WMCC15]. differentiation [ADD18, ALK17, ARMA15, AVB13, BJH13, CTvdM^{+}08, GDSPRA13, GFR^{+}18, NM18, NvOH13, RVF14]. Differing [JSCM15, MTH15, NOT^{+}11, SHC^{+}10, SPP15, SS16, SMR^{+}11b]. differs [BPB18, HNK^{+}18, RAH^{+}10, WSHY16, YSS^{+}18]. difficult [PGR^{+}13]. diffuse [PML^{+}10]. diffusion [PZA^{+}13, SRG14]. diffusion-limited [PZA^{+}13]. diffusivity [SPWK13]. Digestion [PTF^{+}14]. digestive [RHF19]. Digital [FH18, THM^{+}13]. dilution [LFL11]. dimension [DC17, DBH^{+}16, WSHY16]. dimensional [FCFD15, JLPG17, LSH19a, LSJ^{+}15, LHDH17, MRLD19, PLM^{+}16, PRLR^{+}18, TRSB^{+}18, WDM^{+}19]. dimensions [BBSM19, DOH^{+}17]. dimethyl [ARGB13]. dimethylsulfoniopropionate [GLS^{+}15, KM10b]. Dimethylsulfoxide [ST14]. dimethylsulphide [FSY^{+}10a]. Dimethylsulphoniopropionate [FSY^{+}10a]. diminishing [MS18a]. dimorphic [BOC^{+}16, CLB19, MTP^{+}10, QSvN^{+}11]. dimorphism [GRCH^{+}18]. Dinitrogen [BvHC^{+}15, CPEE18, BCvH^{+}15, RMB^{+}14, RBC^{+}15, TBC^{+}17]. dinner [CD13]. Dinoflagellata [SHMOS17]. dinoflagellate [AEM^{+}18, BH16b, BBRBG10, BQRB12, BCM19, CIF11, EVB^{+}16, EOA15, GG16, LLH^{+}11, LRBS15, PSPK10, QYS^{+}12, RTC^{+}14, SS10a, SE11a, TDKS10, YTM^{+}15]. dinoflagellates [CIM^{+}13, LLS^{+}18, LGR^{+}10, MDAPH11, PMBH11, RL11, SHSK11, SMR14, ZFRG12]. Dinophyceae [RZH^{+}16b, TLM^{+}11]. Dinophysis [NKH12]. dioecious [VMW^{+}10]. Diomedea [CXdG^{+}17, SPP15]. Diopatra [Ber12b]. dioxide [BIL^{+}10, HLL^{+}17, Rei12]. diphenyl [MHN^{+}12]. Diplodus [AHA^{+}15, BDG^{+}16, NVA^{+}19]. diploproa [GCM10]. Diploria [Edm10]. Direct [EGH^{+}15, MHF^{+}19, NOB12, SFW^{+}16, Tut17, AM19, BHP12, CPEE18, FRC13a, HSC^{+}15, LBG17, MER14, NIS14, SSBC11, SPWK13]. directed [AL12a]. directional [CPF^{+}16, MKII19, MCW^{+}17]. directions [LDS^{+}15]. disaggregation [MLS^{+}14]. Discard [SCHO11a]. discarded [TBMG17]. Discards [GSC17]. discern [HS15b]. discerning [CGS^{+}19]. discharge [AMAM^{+}18, IAV^{+}10, LTN19, MYK^{+}16, SKN^{+}17, WK10]. Discharged [EFR12]. discharges [TNSN11]. disciplinary [HDB^{+}10]. discloses [DMD^{+}12]. discoidea [PKB16]. disconnect [RG11]. discovery [LdU^{+}17]. Discrete [SHM^{+}10a, FWL15, War14]. discriminate [AMS^{+}13]. discriminates [LDVP^{+}17, TRSV^{+}12]. Discriminating [KRR^{+}18, MMRC16, NLK^{+}19]. discrimination [AdCM^{+}19, BLV^{+}19, CE11, SWK19]. Disease [FS14, GEF^{+}12, SFLG10, BFL^{+}11, GSO^{+}18, JTTB^{+}18, JTTP19, JJNP12, KBW^{+}15, LC12, MKS18, MPF^{+}15, NHPW18, OPW^{+}11, RSE18, SBW16, WLB^{+}11, YSKH11]. diseased [AB13]. diseases [SBL^{+}10]. Disentangling [PLMN10, PRA^{+}19, SGT^{+}14, UMF^{+}10, VANB15, Vig15, DAWS19]. disjunct [NvOH13]. Disko [DSK^{+}10, HJMN^{+}12, KDS^{+}12, SKD^{+}11b, TRL^{+}15]. disparate [LD10]. Disparity [BVD^{+}14]. Dispersal [DKV^{+}16, PFB^{+}17, PBIO^{+}18, TPBM13, AN11b, AGHC10, BOÁNF19, BPB^{+}12, BFH^{+}13a, BSCH14, BPG^{+}11, CDD^{+}16, CM14, CMK12, CL16, CME^{+}12, CCG^{+}19, DXF^{+}18, DVV^{+}18, DCT^{+}10, FMP13, FZI^{+}16, GCFW14, GBD^{+}12, GBL^{+}15, HVT12, HSMS^{+}15, HBP^{+}16, HNMK15, HR15, JBL^{+}16, JMO^{+}19, LLS^{+}18, LMC14, LTCH10, MORM15, MPW12, MIA^{+}11, MPF^{+}15, MTG^{+}11, NFSW10, PPSR15, PPS^{+}13, PÁIIS^{+}19, RMKR13, ST11a, SWH^{+}18, SLO19, SO12, SS10c, TSCSB15, TBG15, VNN10, VDV^{+}15, WC18, WBHL10, WTP^{+}18]. dispersal-diversity [SLO19]. dispersing [SGM^{+}14]. Displacement [WN16, KJC^{+}15, NHTS14, vdWTWS15]. displacements [VDM^{+}10]. displayed [MSH^{+}14, RDF^{+}13]. displays [HP13]. disposal [MPB14a]. disrupt [WHT12]. Disruption [KF13a, BRH13]. disruptive [SS12b, WRO^{+}11]. disrupts [CRL13, HSR^{+}13, LF14, SWK^{+}16, VGR^{+}16, WTP^{+}18]. Dissimilarity [CPF^{+}13]. dissipative [CD16a]. dissolution [BO15, HSRP11, MLP^{+}14]. Dissolved [CHLR16, IAV^{+}10, VBM^{+}13, BW16b, BW16a, CAB^{+}19a, CAKD14, CG16, FJC10, FLSG^{+}15, GBW^{+}12, LMW^{+}13, ROE17, RCB^{+}10, SLIGG18, SBS^{+}13, SSX^{+}17, TNM^{+}11b, UFP^{+}10, VAHL11, VBM^{+}11, WADKT13]. distance [AFF^{+}15, BHMCM16, DL13, FDVL18, JHNL12, KLKM12, LGBSC11, MCW^{+}17, NWD13, PCH^{+}14, RAA^{+}16, SCO^{+}12a, vBMH^{+}17]. distances [RS14, TPF^{+}13]. distant [ADvdM13, TPC^{+}16]. Distinct [GSMK15, HORRLS^{+}14, PS15a, EPC^{+}19, GNBM11, GHT^{+}15a, JWHR11, LBP^{+}19, LEB14a, MHQ17, MDS14, OWH^{+}19, PK13, RST^{+}10, SHM^{+}10a, VB19b]. Distinction [AHD12]. distinctness [COK10]. distinguish [ITT14]. distributed [BB14a, BSH^{+}18, CRGM18, CKV12, CKC^{+}10]. Distribution [AN11a, ANJ^{+}11, BMD^{+}11, CCE^{+}14, DIG^{+}11, DMT^{+}14, FBA11, FDDD13, GMG^{+}16, HEB^{+}11, ILDL10, KWW^{+}14, LGB^{+}13, LMP12, LDP^{+}17, MAD17, OPW^{+}11, RZH^{+}16b, RNK^{+}11, ST10a, SLL^{+}18, UFP^{+}10, dML11, AAE14, AGK13, AMSdL^{+}18, AVB13, AOW12, AGIW16, ASEB11, BS14a, BHA^{+}11, BAS^{+}10, BLD^{+}16b, BFG^{+}19, Bea11a, BJMS12, BBS^{+}17, BSBR^{+}12, BEG^{+}13, BDH^{+}17, BGK17, BCHG18, BAC^{+}11, BSY13, CBC^{+}13, CWW^{+}16a, CGD10, CLGCA^{+}19, CHDH^{+}19, CCHG10, CIF11, CAKD14, CDCZ16, CME^{+}12, CUCS17, CBYH13, Cra12, CLS^{+}16, CD16b, DXF^{+}18, DMd14, DTS14, Dav19, DBE^{+}12, DCB^{+}16, DH18b, DWD^{+}18, DB11b, DT11, DC17, EC14, EBdlTC12, EFOM13, EFOM18, FKZTT18, FRGI15, FMJA^{+}18, FHH^{+}17, FJD13, GFKC13, GMP^{+}11, Goy14, GCRP15, GMES18, GVGP^{+}14, GDE17, HMH^{+}18, HEMJ18, HLW^{+}16, HS14a, HHS^{+}12b, HCG15, HCBV^{+}11, HSMS^{+}15]. distribution [HMJ^{+}13, HWR^{+}14, HYL^{+}13, HRK^{+}18, ICUM10, JGJ15, JCZ^{+}12, KN12, KO12, KFW^{+}18, KF13b, KKO^{+}14, KPB11, KAJP^{+}15, KPr^{+}18, KFB15, LSN^{+}19, LRS^{+}12, LZF^{+}14, LABG^{+}12, LMd12c, LMd12b, LMSC^{+}19, LLD^{+}12, LMTD16, LGT^{+}14, MSGGE10, MBG^{+}16, MSC13, MHF^{+}19, Mat10b, MYK^{+}16, MLP19, MHR19a, MAB^{+}12a, MHW^{+}12, MG14, MG12, MGB^{+}14, MOK^{+}18, ORE^{+}17, OMT^{+}17, OWW^{+}17, OAD^{+}14, PPX17, PR11, PSL^{+}12, PSER17, PJP^{+}11, PSC13b, PLNE11, POX^{+}17, PLF^{+}12a, POH16, PAK15, PRBR11, PHG^{+}14, RPM^{+}13a, RGEA^{+}15, RPE^{+}13, RBA^{+}18, RHW12, RCK^{+}11, RPP^{+}13, RDK14, RLO18, RB19, RL11, RBGT10, RJH14, RKT16, RZPH14, RSG^{+}16, RNGUR^{+}11, RHRB^{+}17, SAL^{+}12, STYT10, Sat10, SHK14, SHKR15, SHT^{+}18, SSB^{+}18, SSB^{+}16, SHAS15, SCM13, SHT^{+}16, SWBMJ17, SFKQ15, SBKA14, SHB^{+}19, SS19, SA11, SPP17, SNT^{+}18, TM15, TMS^{+}16, TMSG19]. distribution [TTF^{+}11, TLAT13, TPM19, TS16, TW10, TRA^{+}14, TLK^{+}12, ULCAZG17, UHG12, VCFF13, WPML^{+}18, WLBB^{+}13, WHvdK11, WLL^{+}14, WGD^{+}19, WJHS13, WFH^{+}18, WWZ^{+}19, WZP^{+}14, XLXS17, YYU^{+}17, YIKG10, YBB^{+}12, ZPPM17, dB10]. distribution-shifts [DWD^{+}18]. Distributional [BWS^{+}12, CTvdM^{+}08, LD17b, BdSC^{+}14, BBC19, RBH12]. Distributions [DHK^{+}17, KBC^{+}12, TAAN15, AP14, BO15, CGKW10, HRP18, HHS10, HFK^{+}13, KHW^{+}10, LH14, MBJ16, NSY^{+}18, PPML17, PML^{+}10, PDBB14, QCX^{+}10, RCAB^{+}12, RTÁB^{+}14, SHPS^{+}16, SBBB^{+}18, SM10a, SMSB19, SMJBØ11, SNNS^{+}12, TBP11, TTT^{+}15, WHH^{+}13, WTC^{+}15]. Disturbance [BDD^{+}18, LB15b, WK16, BHE^{+}17, BAM^{+}10, BKF15, CF12, CWMP13, FNF^{+}18, FT10, HH15, HCL10, JGL12, KPB16, LRS^{+}12, LCD11, LCL^{+}13, LTCH10, MBBBC12, MKPS13, MAT^{+}11, MV12, MJR^{+}14, MA14, MLB18, MMBA13, MF13, NBM^{+}14, NNT^{+}10, OSK15, OSPWB14, PSA17, PBG^{+}10, SCB^{+}16, SYM^{+}16, SWB^{+}10, TS16, VHG^{+}18, vDBH^{+}15]. disturbances [GEA^{+}14, HBB^{+}12, HHI^{+}14, JC15, LD10, OSPWB15, SMR11a]. disturbed [DDP15, HCL10]. ditches [CR11]. Diurnal [CG16, SHDR13, WBS^{+}17, KCGC14, MG14, SA11, TGSM12, WPML^{+}18]. diurnally [WDL^{+}18]. Dive [CVW^{+}11, AAM^{+}12, BDM^{+}13, GMK18, HBF^{+}10, JHNL12, QSvN^{+}11, SPBS14, VGGD^{+}15, WOHJ^{+}15]. Diver [HCL10, Dav19, AWH19]. diverged [vdHTT^{+}15]. Divergence [MGPC13, EBOM11, GDSPRA13, GE17, MPW12, MHPvdH19, PFPP18, VBV^{+}10]. divergens [NUJ12]. Divergent [JUK^{+}10, MKK13]. Diverse [SGS^{+}19, FS18, LC18, OCS^{+}13, TGI^{+}11, VRSS18]. diversicolor [CML^{+}12, HKS^{+}12, KHD^{+}11]. diversification [FTKK14]. diversions [ASM^{+}12]. Diversity [BHTF11, MSH^{+}14, MDS^{+}13, RRH^{+}16, VDM^{+}10, YHTS12, ADF^{+}18, AGGAO^{+}18, AAGGG19, AGC12, BMM17, Ber12a, BD11a, BPSG11, BPG^{+}13, BSB^{+}11b, BVS^{+}11, BB19, CPF^{+}13, CAB^{+}19a, CIF11, CWEE19, CLB^{+}18, CHRB10, DEGF^{+}10, DP11, DGP^{+}19, Det10, DEWK15, DDP15, DVB^{+}16, DMJ11, DDC^{+}15, EDS14, EHD^{+}17, FLAL10, FMC^{+}11, FDDD13, GHES18, GWT18, GSD10, GKHS16, GRN^{+}16, GS17, GGFB15, GBD13, HMH^{+}18, HHH^{+}14, HMM^{+}15, HDP^{+}17, HBS10, HL14, ICT^{+}11, JGJ15, JD18, KWS^{+}13, KTH^{+}18, KPB11, LTN19, LMRRB17, LRB^{+}12a, LRB^{+}12b, LS19, LKN^{+}13, LDVP^{+}17, LCL^{+}13, MPM^{+}18, MHPvdH19, McC19, MTM^{+}11, MVS^{+}10, MMFP^{+}15, MGOL19, Mus12, NFR^{+}14, NHSG11, NOB^{+}19, NGP^{+}15, NvOH13, PHSBO16, PPTB12, POL^{+}11, RR18, RWM^{+}12, RK15, RZPH14, RNGUR^{+}11, SSBC11, SLRM^{+}15, SCTA10, SCLB11, SKA^{+}14, SLW15, SLO19, TSS11a, TSS11b]. diversity [TCA^{+}11, TG12b, VPOA12, VRMKH^{+}15, WHB^{+}14, WS13, WS17, WAC^{+}14, WKRW^{+}12, WA14, XS19, vdHTT^{+}15]. dives [BSBR^{+}12, MAB^{+}12a]. Diving [HRS^{+}17, LJM^{+}10, MRLD19, AAM^{+}12, BCG10, BBC^{+}13, BBKH^{+}11, BPC17, CMK^{+}14, DTL^{+}15, GPW^{+}10, HLPS^{+}15, HG10, HA14, JRW^{+}14, JLPG17, KBW^{+}11, KBA17, LCB10, LDT^{+}13, MTP^{+}10, MGB^{+}19, NTS^{+}18, OAB^{+}19, POR^{+}15, QHN^{+}10, RDHM10, RMNP17, TIN^{+}18, TRHM14, VSG^{+}15, VGR^{+}16, WPMP14, ZHM^{+}10]. DLA [FSY^{+}10a]. DMS [FSY^{+}10a]. DMSO [MA13a]. DMSP [FSY^{+}10a, KM10b, LSV10b, MA13a]. DMSP-lyase [FSY^{+}10a]. DNA [BS18b, AGGAO^{+}18, BBB^{+}15, BCSR10, BS18a, BVD^{+}14, CR16, CML^{+}12, CGM^{+}13, DHJ^{+}17, GBMW12, HBK^{+}17, HTSA16, dDTC^{+}17, MIM^{+}12, PCNB16, PFG10, PBB^{+}15, PHHL^{+}11, TWY^{+}19, vBBW^{+}14]. DNRA [AM18]. Do [CRGM18, GZB16, RPR16, ADL^{+}18, AS12, CMC10b, CKK^{+}10, CG16, Col10b, COK10, CRHV14, GNHLN11, MAB^{+}12a, MDS14, PAFC13, QSvN^{+}11, SHM^{+}10a, Sma13, TPBM13]. document [CCJR14, JJP^{+}13]. Dodging [TRSB^{+}18]. Does [FPM14, MLB^{+}14, PHB11b, POL^{+}11, CVM^{+}17, Det10, EM13, HLS17, JGWJ^{+}11, KS13, LH14, MBG^{+}16, MDM18, MGD^{+}11, PG15, PBM^{+}18, RDK14, TSS11a, TPKC18, WFP15, ZH13]. dogfish [BM16, SGB^{+}10]. dogtooth [BHN^{+}14]. dogwhelk [BVO16, MKK13]. dollar [CG10]. dolphin [DCB^{+}16, GDP^{+}11, Nor13, PBB^{+}15, SGC^{+}13, SKGS11, SRB^{+}17, SWH^{+}11, ZPPM17]. dolphins [ARG^{+}10, ACP^{+}16, AMLR^{+}17, BT10, BBSM19, BNAK11, BPZ^{+}14, BAH^{+}15, EC14, FFS^{+}15, GCR^{+}17, KSBM^{+}11, MHAP18, MMM^{+}19, NWB^{+}19, PBC11, SCO^{+}12b, SHT^{+}13, TLK^{+}12, TNN11, VHHGF11]. domesticated [DWL19]. dominance [BDPD14, EM13, JLL^{+}12, MRC^{+}12, MBRO10, MOH^{+}10, WCG^{+}14]. dominans [CIM^{+}13]. Dominant [MPB14a, BCvH^{+}15, HJN^{+}10, HFLI11, KM10a, PAF^{+}10, RWM^{+}14, RPGP^{+}15, SK13, SBW16, WTGC^{+}11]. dominate [BAFP^{+}13]. dominated [BSB12, CRG^{+}16, DVSM10, DHAV13, HSPS14, MMEM14, ROE17, SMK^{+}18, SLIGG18, SLL^{+}18, SDM16, WHH12]. dominates [GSMK15, KAMA13]. dominating [vdGvdLvG^{+}19]. domoic [BSG^{+}10, LRR^{+}10]. Don [Kin10b]. donate [GZB16]. don’t [Chi11]. dormancy [MPZ10]. dorsalis [MSF18]. Doryteuthis [GBH12, SHM^{+}12]. dose [PLS^{+}10]. Dosidicus [BBG12, GRCH^{+}18, ICT^{+}11, MBS^{+}13, RRRF^{+}17, RCVG10, SRCE^{+}10, SZG11, SHF^{+}12b, TKA10]. Dotilla [APH^{+}10]. Doto [YFC^{+}13]. double [JWB^{+}17, RBL^{+}17]. doubly [BRH13]. dougallii [Goy14]. Douro [ADB10]. down [BW10, BAW11, CWGG15, HOE^{+}12, JWR^{+}10, LGMB^{+}16, MTV^{+}11, MSE^{+}12, PS15b, PJMP10, RDK14, RSK^{+}14, STP13, ST11b, TSTP13c, WHT12, YKG^{+}15]. down-regulate [YKG^{+}15]. Downs [DVA^{+}18]. downsizing [APC^{+}14b]. downstream [ABT^{+}15]. Downward [BEMD11, LBN11]. downwelling [AAGGFG^{+}12, DFPH11]. Drag [DHSL11, VSG^{+}15]. dragonet [GB12]. Dramatic [KMM15, TBH^{+}18, vdVKA^{+}11]. drastic [MLCC19]. draw [STP13]. draw-down [STP13]. dredging [DBCB11, HTR^{+}11, SHB^{+}13a]. Dreissena [SSS19]. Drift [HWTH12, KKS12a, GBD^{+}12, LKC^{+}17, WHH12]. drifter [PRK^{+}12]. drifters [RWS^{+}15]. drifting [HTK^{+}17, SS11a]. drilling [EFR12, JGL12, SES17, TNSN11]. drive [BCH^{+}18, BOU13, CBMK10, CFPF^{+}14, CBF^{+}19, Det10, DB11b, DZL16, DAM^{+}17, FCS15, GLM^{+}10, GRN14, HTCH^{+}18, Ing16, MHGW13, MGGSP15, MAK^{+}15, RKD18, RBM14, TR13, VBGO^{+}16, WC18]. driven [BMSB13, EPS11, GDB^{+}16, GBS^{+}10, GJSR18, HR12, LNK^{+}16, MBW10, MCTHGV12, PvdMS^{+}15, PFL13, PZA^{+}13, Rai12, RCAB^{+}12, RPHB15, SJLH12, TTM19, WC18, WADKT13, WJ12]. driver [FSP^{+}17, HFK^{+}13, MTB^{+}17, SNNS^{+}12]. Drivers [AFOVRB14, HMR^{+}16, LMD12a, NAG^{+}16, NDG15, RPCL19, TRLS^{+}11, AAAE17, AFLM12, BGP^{+}16, BFDM18, Bol17, BBGL13, BBH^{+}12, CSB14, DJL^{+}16, FGL^{+}12, GWM^{+}15, GCCM18, HPN^{+}17, HS14b, JREVB18, KHW15, LBAK13, LHU^{+}18, LGB17, LRB^{+}12a, LSH^{+}19b, LGT^{+}14, MWC18, MPB13, OPW^{+}11, PNA^{+}16, PTL16, PP13, POX^{+}17, PPC^{+}19, Pro17, SWM^{+}19, TMSG19, TP16, VJG^{+}13, WAM^{+}15, WRP^{+}17]. drives [BMF^{+}12, CMP^{+}11, CLL17, DGC18, FWB^{+}15, GBM^{+}12, GMSC14, LFU^{+}19, MS18b, MWW^{+}14, SS10c, ZFZ16]. Driving [CFF^{+}18, BDBN11, BKAS14, CAGS17, DGA^{+}15, GFKC13, LdU^{+}17, ME11, RKvW10, SS19, WLB^{+}15, dJH11]. droebachiensis [CD16b, FNC^{+}14, FG16b, JH14, LH16, RMR12]. Droop [PO13]. dropstone [MYS^{+}16]. dropstones [ZSEV17]. Drought [CHDH^{+}19, CPP10, HLB^{+}16, MPM^{+}14, PH16, PPB10b, PP13, WJHS13]. drum [ARA^{+}10, DR16, NRF11, OF10, RSHM10]. dry [DCW^{+}14, FA11, HSF^{+}12]. Dual [HLD12, AAJJ^{+}19, RCB^{+}10]. Duality [LFU^{+}19]. Dubai [BBB^{+}14]. dubius [DHG16]. duck [LDT^{+}13]. ducks [BSY13]. due [CLGCA^{+}19, CY10, CLB^{+}18, HGS15, LFMC11, LBJK14, OWW^{+}17, SB15a, TKBN10, TNM^{+}11b, WHP^{+}17]. dufresnii [ZdVA^{+}16]. dumpling [SWNSF14]. dune [RTDUM16]. Dungeness [RS14, SE11b]. dunlin [KMSW10]. Duration [KNY^{+}15, AGP^{+}13, BM11, CBF^{+}13, HLW^{+}16, SWMK15, ST11a, ZSSK11]. during [ABT^{+}15, AHØ^{+}15, ACRLB12, AFMR^{+}13, AFRAG16, AAM^{+}12, BTS10, BEMD11, BMD^{+}11, BPW^{+}18, BRT^{+}17, BGIA19, BDD^{+}18, BMSS14, BPHW16, BK14, BFG^{+}11b, CRAP^{+}12, CPP10, CR19, CE11, CCG^{+}10, CKK^{+}10, CMM^{+}18, CMGS19, DSC13, DKAC^{+}12, DVSM10, DFGH^{+}15, DRD^{+}18, DCC^{+}10, DFTZ^{+}15, DPO15, DAWS19, FW11, FBA11, FKA^{+}12, FTB^{+}17, FTS^{+}13, GLS^{+}15, GGL^{+}17, GFGGDSM12, HHI^{+}14, HFR^{+}18, HHP^{+}17, HMO^{+}12, HEB^{+}11, HLB^{+}16, HCLB18, JFK12, JGG^{+}15, KPW^{+}13, KAP^{+}14, KTF10, KFBM19, KCS^{+}17, KBP^{+}10, KRdVG^{+}18, LSY^{+}18, LSN^{+}19, LHS^{+}14, LHL11, LBL^{+}12, LWG14, LJB14, LCL^{+}13, LRQ^{+}12, LHW^{+}19, MSGGE10, MPS^{+}10, MGM^{+}18, MFD12, MRC^{+}12, MWW^{+}14, MCHC^{+}16, MOK^{+}18, NMC^{+}16, POR^{+}15, PPB10b, PP13, PPEH^{+}14, PMF^{+}17, PGR^{+}17, PSB17, RCP^{+}16, RPLT10, RNB^{+}16, RCG^{+}15, SMOH19, SBH^{+}19, SMK^{+}18, SWR11, SP15, SGP^{+}17, SHF^{+}12b, SSdW^{+}12, SOT15]. during [SLR^{+}14, TNS^{+}12, VKY^{+}12, VBT^{+}17, WRGP12, WSL^{+}14, WMMC14, WCRCR17, WWW^{+}13, WFP15, YHT16, YSK^{+}10, YOST11, YOWK17, ZTNS13, ZCC^{+}12, ZGK^{+}19, vBAP^{+}19]. durum [DDP^{+}13a]. dusky [BNAK11, GDP^{+}11]. dust [LMR^{+}12]. Dutch [BDD17, CBC^{+}10, vdVKA^{+}11]. Dwarfism [RLTL14]. dwelling [HVT12, NFSW10, PM16]. Dynamic [BBSM19, RPM12, SvdMK^{+}12, WP10, FHTY19, GGG18, GMSC14, GLB^{+}10, JR15, LSH^{+}19b, LFG^{+}12, MSG11, PFG10, RWS^{+}15, RMNP17, SPM^{+}13, SHB^{+}13a, TTW^{+}12, VTZ^{+}13, VGR^{+}16]. Dynamics [KCP^{+}11, LKC^{+}17, NFR^{+}14, TS11a, TAB^{+}14, AB17, ABM12, AL12b, AB13, ASCHM19, ARA^{+}10, AGH^{+}12, AGP^{+}13, BZBR^{+}11, BS16, BDB^{+}13, BMR17, BBC19, BFH^{+}13a, BCSB13, BCH^{+}18, BMVAS19, BC13, BKAS14, CLG19a, CR19, CLL15, CBMK10, CSC13a, CGD10, CLW^{+}11, CF12, CZB10, CLL13, CGH12, CCS^{+}18, CL16, CLL17, CMM^{+}18, CFBG14, DFD^{+}19, DDM^{+}16, DBC^{+}12b, DvBH12, ET11, EBDB18, FFLS16, FPBL16, FL17, FG16b, GEAOH17, GCC^{+}10, HB13, HHH^{+}11b, HSS18, Hv14, HTCH^{+}18, HT10, HRM^{+}11, HMBP18, HEM^{+}12, HPM14, Hos14, JBL^{+}16, JLM^{+}12, JD18, JLQ^{+}14, JPL^{+}15, KBH11b, KM10a, KGJ^{+}16, KFBM19, KRC17, LCQT^{+}13, LFJ^{+}14, LSJ^{+}15, LTP^{+}18, LGS10, LLRC^{+}12, LFL11, LLCM^{+}12, LCB^{+}12, LHT^{+}17, MDLM17, MATR11, MASMH19, MWZ^{+}16, MTM^{+}11, MGGSP15, MPD11, MMEM14, MN10, NW15, NNT^{+}10, OC17, OE11, ORJH19]. dynamics [OFM^{+}18, OCSW16, OMZ10b, OSD^{+}13, OKLM14, PGG12, PNI14, PDGR^{+}13, PNA^{+}16, PDvG16, PvdMS^{+}15, PBF10, PFVFF^{+}15, PSA17, PBC^{+}15, PMS13, PFL18, PSB17, RU11, RWM^{+}10, RSDN15, RMME10, RZH16a, RRB13, RJTS12, RGB^{+}10, RGB^{+}11, RGP11, RBBG12, RKvW10, SDM^{+}13, SS10a, SES12, SCAM^{+}14, SF16, SBE^{+}14, SPH^{+}17, SWR11, SJC10, SHMOS17, STP13, STBP13, ST08, SDMR19, SS14, SP12, SIHH10, TFP14, TMT^{+}12, TGH15, TLM^{+}11, VSM11, VCDP11, VM14, WKE^{+}17, WB12, WWB14, WSF^{+}12, WK18, YMCV10, YYW14, YOWK17].

E. [MBP^{+}12]. each [BB14a]. Early [AHA^{+}15, MJR^{+}14, MD16b, ORM^{+}13, PMB^{+}10, SK13, SHMB15, SYK^{+}13, SNT^{+}16, TB14a, ABK^{+}14, AS14, ABW14, BVL10, BMJH12, BPHW16, CSJP17, CPPM12, CLMB13a, CLMB13b, CTB13, CG16, CPS15, DSG15, DABP13, DBG15, DQC^{+}13, FDVL18, FLS^{+}18, FTB^{+}17, GM13a, GJ10, GNBM11, GHT15b, HHP^{+}17, HSH18, HHNI17, HB15b, HHB^{+}14, JH11, JH14, KKMM16, KLT^{+}12, LCQT^{+}13, LAM10, LAM13, MJT^{+}14, MJM^{+}14, MTM^{+}17, MKS18, MWC^{+}12, MGD^{+}11, NRF11, ORWPR14, PSG11, PES12, PRP^{+}16, RPLT10, RWP^{+}14, RBC^{+}14, RTC^{+}14, SVG16, SKD^{+}11a, SPMN19, SW12, SDAH12, SLL12, TFH^{+}19, VSG^{+}11, VCFF13, WHI^{+}18, WWW^{+}13, vdMPM12, TTPM12]. early-developed [FLS^{+}18]. early-life [BMJH12]. early-life-stage [CG16]. early-stage [SPMN19]. earthquake [CMC10a]. EASI [GKRG^{+}19]. EASI-Fish [GKRG^{+}19]. East [TNB^{+}15, WRR10, FFS^{+}15, KO12, MVOPB^{+}16, OWN^{+}15, RABS19, YYW14, CMF^{+}19, CBKM16, DKV^{+}16, GFW^{+}14, HGFA^{+}11, KAJP^{+}15, LOS^{+}19, LLS^{+}18, MLM^{+}12, MKK13, McC19, MMBA13, RSW^{+}16, ST10a, STYT10, SEFS15, SJL^{+}15, TIN^{+}18, TBH10, XLXS17, YYO14, ZCC^{+}12]. Eastern [CGK13b, GGG18, PBC^{+}12b, ALK17, ACAD12, BS15, BG19, BZBR^{+}11, BHL^{+}19, BSCH14, BHT10, CGKW10, CRRL17, CLMM^{+}14, CL15c, CSRS15, CATAV18, CHLR16, DLC^{+}13, DRD^{+}18, DPM15, DAD^{+}11, DHJ10, ETB^{+}11, FHTY19, FDDD13, FQ11, GLF^{+}17, GLJ^{+}18, GBH12, GMK18, GGMTGC18, HTCH^{+}18, Hof10, HBM^{+}12, HR15, HEAI10, JRW^{+}14, JPMG10, KBW^{+}15, KR16, KW10, KPP^{+}16, KLMR14, LBN11, LSY11, LPC19, MB16, MMRC16, MVE^{+}10, MAP^{+}12, MWHM18, MKT^{+}18, ODK^{+}14, PPM^{+}15, PS15a, PCH11, PBC11, PLF^{+}12a, PGOG12, RHF19, RDK14, RDC^{+}12, RSG^{+}16, RHS^{+}19, RLNLAF17, RSSvV19, SSBA12, SRSH^{+}18, SCO^{+}12b, SKGS11, SHM^{+}12, SDCC12, SRH^{+}15, SRCE^{+}10, SNS12, TLFS14, TB12, TSA^{+}18, TFH^{+}19, WLM12, WLH^{+}17, ZO10, tHHR10, BB19, CNW^{+}12, GFGGDSM12, GEA^{+}14, JMB^{+}17, KFGSMPG14, LCL^{+}17, MF15, PML^{+}10, PFL18, PW19b]. easternmost [BMV19]. eaten [DZL16]. eating [ARGB13, BDP10, FM14, PTF^{+}14, SVAV17, TMT^{+}12, Wir12]. echinoderm [MTM^{+}17, PSB17]. Echinodermata [BLD^{+}16b, HGFA^{+}11, JMB10, SK10, ZdVA^{+}16]. echinoderms [FSB15, REP18]. Echinogammarus [ADO^{+}12]. echinoid [BGK17, DDDR10, JDD^{+}13, SCL^{+}16]. Echinoidea [JMB10, LHU^{+}18, ZdVA^{+}16]. echinoids [BSA^{+}16, BS19b, CR19, DDM^{+}16, PSL^{+}12]. echinoplutei [MPB10]. Echoes [KRW18]. Ecklonia [BVG13, DBS^{+}19, MGFW13, RWR^{+}19, TW16]. Eclipse [KFBM19]. eclosion [SLC11]. eco [LBP^{+}19, LEKT12, SPM^{+}13, SBH^{+}19, SHG^{+}15]. eco-evolutionary [LEKT12]. eco-physiological [LBP^{+}19, SPM^{+}13]. eco-physiology [SBH^{+}19]. Ecol [ACA^{+}10b, BW16b, RUP11b]. Ecological [ASI15, BGSM13, BJMS12, BJH13, BH15, BSC17, BW12, BGS^{+}16, BFPONS^{+}12, CLGCA^{+}19, DFDC18, DHJ10, GKRG^{+}19, GHL11, KdEPW18, KSBM^{+}11, LGD10, MMA18, OMZ10a, RvOP12, SLS18, TLK^{+}12, WD14, BCN15, BW10, BW15, BMPC15, BDC10, CFMFFJ18, CGK13a, CPR^{+}16, CLMB13a, CHFS12, CLPU14, CLPiU14, CPMR11, CE13, ESBG14, FL10, FYG^{+}18, FB14, GL12, GSC14, GBE^{+}17, GNJJ^{+}17, GCS^{+}18, GMR^{+}19, GLB^{+}10, GCCCM16, HKSD14, HWTH12, HM13, HTC17, JMK12, KHW15, LSN^{+}19, LGS10, LPK^{+}16, LC18, LLP^{+}10, MUC12, MLS^{+}14, NVA^{+}19, PMK^{+}15, PIPD10, PSL^{+}12, PMM^{+}15, PEB^{+}14, Pro17, RSP17, SKB15, She13, SBE^{+}16, SBL^{+}10, SHG^{+}15, SPP18, TMSG19, VJG^{+}13, WBG^{+}14, Whi11, WMCC15, ZPPM17, ZB16, vDCR13]. ecologically [EBOM11, GS18, HKL12, MTM^{+}17, POS14]. ecologies [OBI^{+}16, SATS13]. Ecology [FBS14, Kin10a, Mus12, WSD^{+}10, APNC^{+}15, APBC^{+}19, AGV^{+}11, ABA16, BAS^{+}10, BYAA13, BWLP17, BFD^{+}16, BMV^{+}11, BLHL15, BPSK11, BB14b, BKC^{+}19, BL10, CCB^{+}17, CPF^{+}14, CCJR14, CXdG^{+}17, CBR19, DSS10, DWPH11, DAW^{+}11, DZDB10, DZB^{+}18, DBS^{+}16, EPRGO13, FHRC16, Fuj16, GDB^{+}16, GLW^{+}12, GSP^{+}15, GCSF18, GSV14, GPC^{+}15, HCL^{+}11a, HBK^{+}17, HBB^{+}18, HFG^{+}19, HTB^{+}15, HB12, HA15, HOA14, IS16, JJP^{+}13, JU16, JHL^{+}19, KBH11b, KMP11, KCH^{+}11, KW12, KRD^{+}12, LS14, LAL^{+}18, LA12a, LFPC^{+}14, LPSL14a, LPSL14b, LBS^{+}15, MCW^{+}16, MTS^{+}17, MWZ^{+}16, MD11, MCL^{+}17, MJV^{+}11, NBGS^{+}12, NJP12, NCW^{+}19, OMBP10, PXG^{+}10, POR^{+}15, PKS11b, PKSN11, PW16, PSS19, PWB^{+}19, Pur18, RWG15, RLNLAF17, REP18, RVH^{+}19, SLS11, SL15a, SDF^{+}19, SN17, TWY^{+}19, TF12, VCRUQ15, VVB^{+}10, VAB^{+}13, WvLF18, WCRCR17, WWZ^{+}19]. ecology [YD11, YSH^{+}16]. Economic [CS15, LCC^{+}11, RvOP12, SHG^{+}15, WF18]. economically [GS18, MTM^{+}17]. Economics [SS15b, SS15a, ON15]. Economy [Bus12]. Ecophysiological [LG18, FAT^{+}15, LLP^{+}10]. Ecosystem [AH13, AKF14, BL10, CAD^{+}10, CSPA10, DKRS12, EBT^{+}15, FCDC12, HGW^{+}11, LN10, MLDG12, MKG10, NSG^{+}12, PS19, PBG^{+}10, PEB^{+}14, PRB^{+}17, RFRE19, SCRL11, She12, TTC11, WBB^{+}10, AMR^{+}10, AB12, ASF13, BAA12, BdSC^{+}14, BJG^{+}11, BB17, BHE14, Ber12b, BH15, BHF^{+}16, BLT14, BKF15, BRB17, BLC^{+}16, BSS18, BPM^{+}18b, BDGB19, BNW14, CRRD^{+}13, CMV^{+}11, CF12, CPR^{+}16, CFC^{+}12, CSRF16, CAKD14, CV11, CB12, CFH11, DLdM19, DGMB10, DKAC^{+}12, DDB^{+}16, DVB^{+}16, EOM^{+}14, FLAL10, FBS14, FHRC16, FRV^{+}13, FDFS16, FBC^{+}15, FFL10, FPBL16, FF16, GHGF12, GKHS16, GTP^{+}17, GGNL^{+}17, GC13a, GS15, GKPT16, GNC^{+}10, GDA^{+}18, HVK^{+}13, HPN11, HLD12, KN12, KNIN17, KVY^{+}13, KCW^{+}17, KSB11, KLD18, KTR^{+}16, KRC^{+}11, LFJ^{+}14, LSJ^{+}15, LFFL15, LB11, LFD13, LBMZ15, LRM^{+}14, MTV^{+}11, MBC^{+}11]. ecosystem [MLR12, NMPM15, NGA18, OSK15, OPN10, OV10, PAG19, POPH12, PTL16, PHH^{+}12, PMB^{+}15, PR11, PCS^{+}15, PMS10b, Pör12, PRA^{+}19, dGCF^{+}17, RWM^{+}10, RBM14, RRG^{+}11, RSTM14, RGB^{+}10, SDM^{+}13, SDMB15, SHS^{+}17, SBB17, SGT^{+}14, SP13b, SOJ14, SLT^{+}19, SLRP^{+}12, SFL^{+}18, SS15a, SH10, SS12b, SPP18, SB15b, TPS10, TRJ^{+}11, TGI^{+}11, VBM^{+}13, VF13, VBR19, WRP^{+}17, WSS^{+}16, Wil14, YHR16, vdBUPP^{+}11, LFL11, SMVC^{+}15]. ecosystem-based [CMV^{+}11, RWM^{+}10, TRJ^{+}11]. Ecosystem-level [AH13, Pör12]. ecosystems [BSF17, BW12, BMF^{+}12, BBH^{+}12, CCA^{+}14, CS10b, DMMW12, EAB^{+}19, FTB14, FDS14, FGL^{+}12, GBJ^{+}14, HW11a, HOL^{+}17, HGA^{+}16b, HGA^{+}16c, HEM^{+}12, HPM14, HFK^{+}13, HPQ15, JDB^{+}10, JSK19, KNIN17, KCP^{+}11, LNH^{+}17, LGM^{+}12, LCB^{+}12, NTN^{+}13, PHB^{+}11a, PPCH17, PLF^{+}12b, RDPR15, RFRD10, SCB^{+}14, STRP18, SBMÅ12, TZF11, TL18, TTC11, TVA^{+}19, TYVR^{+}15, VB19a, dJSB19]. ecotone [GDPR18, HHS10, HMS10, WW15]. ecotourism [CBL^{+}10b, HCL10]. ecotypes [MTH15, RDF^{+}17, vDCFM10]. ectoenzyme [OGA13]. ectoparasites [NFN^{+}15]. Ecuador [FBA11]. eddies [BCG10, Sat10]. eddy [ASK^{+}15, AHSG16, GBH^{+}10, HSM13, JRW^{+}14, LBMZ15, PCH11, RAB^{+}15, VBR18]. edeni [TCSA17]. edge [AVB13, BH16a, Cra12, DRDV15, FHJ15, dlÁGRB^{+}19, GDPR18, JLL^{+}12, PSC13b, RRH^{+}16, SJB16a]. edges [BCW12, MGP12]. edible [PAFC13]. eDNA [BCW19]. education [FEA^{+}15]. edule [BD15, dML11, dMBC12]. edulis [BSH13, EKR^{+}10, GCBS19, JRvdM15, JSS^{+}11, LIT13, PRP^{+}16, RPLT10, RJL^{+}11, RCBE15, SvdMK^{+}12, SDT15, SSAB^{+}15, vBJV^{+}15]. edwardsii [CB17, JW10, JWHR11]. eel [BPOS^{+}16, CKK^{+}10, HHLT10, HLW^{+}16, HVD^{+}16, MAW^{+}11, MSG^{+}13, MSV^{+}11, RBSL11, RBC^{+}14, TWY^{+}19]. Eelgrass [IEM16, MSPO12, OKJMC15, OCS^{+}13, OMH12, RWM^{+}12, BCOMKJ18, BO13, BD12, BOKJ18, BBR12, BPF15, CDMW12, CKJOM14, CM12, DMJ11, DAWS19, FLAL10, HBHS^{+}18, HR12, JMK12, JS10, JBS11, KBA13, KHGS12, KJCN^{+}11, MO12, MMG^{+}16, MRC^{+}12, MHS14, NOB12, OFT^{+}11, OS12, OMM^{+}12, RCB12, RBM14, SRB15, SCRL11, SHKR15, SCM13, SO12, TSGW10, VPBV13, VSB^{+}15, VNK^{+}12, WO11, WL19, ZHJ^{+}17]. eels [BPSCD17, FBS14, LC12, LÁBFÁF^{+}10, SØA^{+}13, SØK^{+}15, SMD^{+}16, WSL^{+}14]. Effect [ACAD12, BT10, BAA12, BVL10, BDAH10, BBRBG10, BBG^{+}13, BBW^{+}19, CLL13, CGH12, CMHH17, CZH^{+}13, CAB^{+}16, DMH15, FJ12, FLSG^{+}15, GCC^{+}16, GMK^{+}19, HCG^{+}10, HJ13, IOS^{+}12, KAD^{+}15, KCR^{+}10, LD17a, MRC10a, MLS^{+}14, OAD^{+}14, OWN^{+}15, PA15, PMS10b, QSB16, RHB13, SBL^{+}10, TRL^{+}15, VCFF13, WLL^{+}14, WHB^{+}19, XTW^{+}15, YOST11, ZS17, dMBC12, vdWHG11, ARB^{+}12, BC11, BGL^{+}11, CBC^{+}13, Cha13, CBF^{+}13, CHLR16, ENW^{+}14, HR12, HMKF11, HWVM19, KM10b, LTSDHL12, LHT^{+}17, MDM18, MER14, MDS14, OD15, PBA12, PRNR^{+}18, RCP17, RPLT10, RGB^{+}11, RGP11, SFTG14, SM17b, UMF^{+}10, VSB^{+}15, WBG10, ZVR18]. Effective [MOP^{+}15, RABS19, AHD12, BPB^{+}12, DH18a, EGH^{+}15, MDPG13, NMS^{+}18]. effectiveness [AAE14, ABB^{+}14, HVBG^{+}19, LYH^{+}17]. Effects [AvDL^{+}19, ATS14, AS14, ACA^{+}10a, ACB12, AB12, AFH11, BÁNC^{+}18, BRJ^{+}12, BSB13, BSD^{+}14, BCPC18, BMR17, BBS^{+}17, BHDC15, BCSB13, BFL^{+}11, BKTD10, BDC10, BGBD^{+}13, CSR^{+}14, CE14, CR14, CSJP17, CMC10a, CSC13a, CESAM12, CGD10, CD16a, CLC12a, CAB^{+}14, CBGDS12, CBL^{+}10a, CAG12, CRF13, DSM11a, DC15, DMMW12, DMP12, DLL^{+}18a, DGQ10, DMM^{+}10, DPH13, DRDV15, DWV^{+}12, DCW^{+}14, EHG16, EBK^{+}12, EVHA^{+}18, EES15, FRV^{+}13, FHL^{+}15, FMC^{+}10, FOJF15, FTB^{+}17, FA10b, GRG^{+}12, Gea10, GJ10, GCBB13, Gre10, GBW^{+}12, HJMN^{+}12, HFLI11, HTR^{+}11, HWB^{+}18, HCJ^{+}11, HR15, JT10b, KLKM12, KBW^{+}15, KJW^{+}15, KB14, KLT^{+}12, KKS^{+}12b, KDS^{+}12, KEH10, KBPD10, KLN^{+}13, LSH19a, LRS^{+}12, LTG17, LCC12, LBE^{+}10, LRBS15, LMTD16, MJS17, MFG^{+}13, MJO^{+}13, MJT^{+}14, MTM^{+}17, MMDM12, MH12a, MM13b, MHR19a, MCR^{+}13, MWC^{+}12]. Effects [MLB18, MSS^{+}14a, MC12b, MB18, MLML^{+}19, MCK^{+}13, MCT15, NGSF13, NWHD15, NKH12, NGP^{+}15, NLLS15, OCE^{+}12, PNI14, PSCS16, PKB16, PNF^{+}12, PPB^{+}10a, PFVFF^{+}15, PFKB18, PMS13, QCX^{+}10, QNG^{+}15, RWO13, RMME10, RF14, RDPR15, RL11, RB15, RT12, RKvW10, RMV13, SLRM^{+}15, SRR15, SMP10, SGJ^{+}18, SCL13, SAMW14, SDMR19, SML17, SS16, SZG11, SSD13, SNS12, SDH13, TG12a, TGM^{+}17, TTF^{+}11, TLAT13, VUL^{+}11, VSW^{+}15, VNK^{+}12, WLBB^{+}13, WAP10, WBW^{+}14, WRO^{+}11, WFB16, WVM^{+}14, WNF^{+}11, YAL12, ZMS^{+}14, ZLCW15, vDJF15, AHA^{+}15, ACDJ11, ASM^{+}12, ANB13, AMR^{+}10, ADO^{+}12, AOW12, AS13, AAIB15, Ano17c, ATP^{+}12, ASM10, AKF14, ASF13, BHF^{+}10, BSF17, BGJ^{+}19, BFDM18, BCS^{+}12, BCSH18, BAM^{+}10, BPJ^{+}15, BBC19, BSC17, BBC^{+}12a, BWH^{+}17, BBW^{+}17, BCK15, BH12a, BPC17, BPF15, BGO^{+}18, BDVP19]. effects [BJRH10, BBMM14, CPSC^{+}15, CLM^{+}13, CDMW12, CFA16, CGK13b, CLHW14, CG17, CLMB13a, CHB11, CL15c, CGQB10, CPS15, CV11, CVP11, Cra12, CFH11, DGA^{+}15, DTG16, DP11, DFLC17, DBCB11, DKRS12, DTM^{+}12, DWM15, DMHS^{+}14, DHR^{+}13, DH17, EKE16, EVB^{+}16, EMMSL13, EBH16, EFR12, EF16, FRC13a, FSD^{+}14, FAT^{+}15, FNA^{+}16, FC12, FSR10, FA10a, FCDC12, FBGT12, FAWBF10, FGH^{+}13, GHGF12, GL12, GP11b, GKC11, GGNL^{+}17, GFP19, GG12, GLHM^{+}18, GC13b, GCP^{+}17, GVGH15, GNC^{+}10, GDPR18, GHL11, HPA17a, HWSK19, HMP15, HRS^{+}17, HVBG^{+}19, HDWF10, HRM^{+}11, HLS^{+}15b, HFR^{+}13, HWTH12, HDT^{+}12, HGKB14, IAV^{+}10, JBT^{+}16, JS10, JG13, JDB^{+}10, JMPZ10, KP18, KÖHSO12, KW17, KSF18, KAP^{+}12, KPB16, Kin16, KP11, KHD^{+}11, KPP^{+}16, KDO^{+}14, LGMB^{+}16, LH11, LBP^{+}16, LC12, LK15]. effects [LC13, LS10b, LCCT19, LD10, LE13, LRTLF12, LPK^{+}16, LJB14, LMd12b, LCD11, LGJH12, LNK^{+}16, LMT15, LSV10b, MP15, MBB^{+}18, MBBBC12, ML11, MB16, MO12, MHF^{+}19, MAS10a, MWBMH13, MTP^{+}10, MLB^{+}14, MBW^{+}19, MMBA13, MDR16, MDAPH11, MVOPB^{+}16, MOH^{+}10, MWMD17, MH13, NBM12, NAG^{+}16, NSG^{+}12, NOB12, NJP12, NSI^{+}15, ORR^{+}14, OCPM12, ORWPR14, OGH18, OS12, OSD^{+}13, OKLM14, PMS10a, PSS^{+}12, PDCH15, PSER17, PBS10, PWW12, PW19a, PLJ^{+}12, PFR11, PJMP10, Pör12, PIGK16, PRA^{+}19, RHVRC11, RWP^{+}14, RHE14, RRB13, RL18, RTDUM16, ROLA15, RMSPA16, RLH^{+}13, ROKS14, RPB17, RSW13, RP17, RMRA15, SGT^{+}14, SHUS17, SMC^{+}14, SA14, SP13b, SCP^{+}14, SG18a, SWH^{+}18, SMF^{+}17, SAM18, SSX^{+}17, SF18, SO16, SSvD^{+}15, SG18b, STP10, SCF^{+}17, SHSK11, SA17, TM15, TG10, TM10, TWM^{+}15, TKB^{+}13]. effects [TGP^{+}16, TNSN11, TBGW^{+}11, TMB^{+}12, Tut17, TVH^{+}10, TVRG^{+}13, VANB15, VGBJ^{+}13, VGGF12, VCQM^{+}16, VPOA12, VTB12, VHM16, VDG^{+}15, VB19b, VLK^{+}15, VDEO18, WDH^{+}12, WI12, WS13, WAW^{+}12, WNNH10, WL11, WFH^{+}18, WPK10, WBB13, YHTS12, YMCV10, YHTC19, ZWP^{+}17, ZXF^{+}16, dGR11, vDBH^{+}15, vDGK^{+}12, vdMHW^{+}10, vdWTWS15, ACA^{+}10b, HT11a, SMS15, VSG^{+}15]. Efficacy [KCS^{+}17]. efficiency [BBBF15, BCB^{+}17, CME^{+}12, DAL^{+}10, GD15, GAF^{+}16, GJ13, KAP^{+}14, Løk11, MC10b, PHD^{+}14, PBWM18, SOB16, TS11a, dDTA^{+}18, vDCR13]. Efficient [FLE^{+}11, KSA12, KHA^{+}12, SMSL17]. effluent [BMMS10, LLH^{+}11]. effluent-affected [BMMS10]. effluents [HU13]. effort [BOS18, CVW^{+}11, FSF^{+}17, HA14, JGWJ^{+}11, KB12, MS10, SWFQ12, SBLM14, TMB^{+}12]. efforts [TPC^{+}13]. egestion [SPP17]. egg [ARB^{+}12, BNAN12, BER17, BVD^{+}14, CCM^{+}15, Col10b, DED^{+}15, DHH10, FSB15, FTvdK^{+}12, GGMGR12, HA18, HJMN^{+}12, KKL13, LK15, LMZ16, MM13b, MP12, NGSF13, RHS^{+}19, SM10b, SS12a, SFW^{+}14, SPP15, SKD^{+}11b, TKL^{+}11, TPVCV19]. egg-ratio [DHH10]. eggs [AGGAO^{+}18, CCG^{+}10, DHB18, FKZTT18, FM14, LJM^{+}12, MP17, NAY^{+}11, NRK^{+}19, RHRB^{+}17, SHB^{+}10b, TAAN15, TVM16, WB12, WJHS13, YOST11]. eider [GDH^{+}19]. eiders [HG10, LRC^{+}13]. EIFEX [LHS^{+}14]. Eilat [ZAZS11]. Elasmobranch [TKL^{+}11, VH11]. elasmobranchs [BCN15, EPF^{+}17, MVE^{+}10, SMF^{+}17]. electra [MHHB^{+}17]. electron [BHMCM16, RWR^{+}19, ZIT^{+}17]. electronic [KLBT^{+}16]. elegans [GSV14]. elegantissima [QSB16]. element [BWH11, CGS^{+}19, KRC^{+}11, MGO^{+}12, MMB^{+}12a, MMWG13a, MMWG13b, PBWM18, RMP^{+}19]. Elemental [PSG11, SMMFH16, BWS11, ENW^{+}14, JWHR11, KPP^{+}16, MDA^{+}10, MS14, NDBL16, TMP^{+}11]. elements [LMR^{+}10, LCC^{+}11, LCBK^{+}13, MCL^{+}17, PWO^{+}16]. Elephant [BSBR^{+}12, MAB^{+}12a, AGNH^{+}19, BCG10, BPG^{+}13, CMHH17, GVGP^{+}14, JLPG17, LCB^{+}16, LRO^{+}17, MBSG11, MRLD19, MTP^{+}10, NCC^{+}14, NFC^{+}11, OHCG14, OGLH17, RLAGL12, RDGC17, VGGD^{+}15, DRH15]. elephas [BFG^{+}11a, BFG^{+}11b, JBCGC19]. Elevated [BIL^{+}10, MD11, VPBV13, WMMC14, AHR^{+}18, BSB^{+}11a, BRF11, BCR15, DMM^{+}10, GCP^{+}17, HWTH12, HSB12, HLL^{+}17, HYC^{+}14, LGW^{+}15, MKS18, MDM18, MCR^{+}13, ORWPR14, PKB16, RNH15, SLIGG18, SSX^{+}17, SDH13, TPKC18, UHG12]. elevation [BG19, CY13, NOB12, PK15, SHB^{+}10a, WGD^{+}16, WS13]. elevations [MH11a]. elicits [FNA^{+}16]. eliminating [JPF12]. elizabethae [MH11b]. elkhorn [DSVM10, PWM18]. ellipse [RSW^{+}16]. elliptica [BCR15, GMT^{+}17]. Ellisella [MLASTS13]. elongatus [DHH10, KDB^{+}10]. elucidate [BPJ^{+}15, KBA13]. elucidated [BMV^{+}11, OBI^{+}16]. Elucidating [RWMP12]. elusive [LRBSJ15, PPX17]. Elysia [CRR^{+}19]. embayment [EZZ^{+}12, HLB^{+}16, LMd12c, LMd12b, MHAP18, MCT15, OPN10, SM10a]. embayments [BFNB16]. Embiotoca [JVSD10]. embryo [CSG11]. embryonic [CR19, JJF14, SZG11, VSW^{+}15, WBB13]. embryos [BGBD^{+}13, PCF^{+}19, SCP^{+}14, SPM^{+}16]. emergence [SCL^{+}19]. emergent [LGM^{+}12, MSE^{+}18, TFP14]. emerges [LAK^{+}19]. emerging [BPP^{+}18]. Emerita [CD16a]. emersion [BNW14, FAT^{+}15, LHL11, ORG19]. Emigration [BAL14a, BBC19]. Emigration-corrected [BAL14a]. Emiliania [ENW^{+}14, GAF^{+}16, HWE^{+}12, JDX^{+}17, KAP^{+}14, LBM^{+}18, MTH15, PYBB11, SLIGG18, TGI14]. Emirates [BBB^{+}14]. emperor [DHSL11, GMK18, LOS^{+}19, WRR10, WPMP14]. emphasis [OCSW16, RZH^{+}16b]. Empirical [AFSL^{+}15, HT10, CCDP15, DP14b, WSS^{+}16]. enables [BPW^{+}18, CPF^{+}16]. encased [GCDH19]. enclosed [KdEPW18, LFJ^{+}14, LSJ^{+}15]. encoding [EH13]. encounter [ML11]. encounter-model [ML11]. encounters [LBJI14, RHW12]. encrasicolus [QLCUG15, RLN^{+}13, TGSM12, vdSRT^{+}18]. end [NIPS12]. end-to-end [NIPS12]. Endangered [DCB^{+}16, FKP^{+}15, KHW^{+}12, TNB^{+}15, ACP^{+}16, AdCM^{+}19, FNF^{+}18, FPC^{+}12, GLW^{+}12, HHS12c, KKMA^{+}10, LNF^{+}11, LDP^{+}17, MRL^{+}17, MMCW14, PJP^{+}11, RABS19, RSSvV19, SDST16, WRP^{+}17, ZBK^{+}15]. Endemic [SDMB15, BH10, FCV^{+}16, KBK^{+}10, LBJ^{+}16, SHK14]. endemism [LLD^{+}12]. endocrinology [Ued14]. endogenous [HTHM18, KF13a]. endophytic [BBR12]. endosymbiosis [BBG^{+}11]. endosymbiotic [PC19, RCB^{+}10]. endotrophy [OT13]. Endozoicomonas [BAFP^{+}13]. ends [HSC^{+}15, HKB^{+}19b]. endures [MGB^{+}19]. enemies [Gre10]. enemy [FFS14]. Energetic [FTM^{+}13, GVGH15, HG17, OT14, BPHW16, JVSD10, MCR^{+}13, PBM^{+}10, TGSM12]. Energetically [SMSL17]. energetics [FRRÁSL11, GLM^{+}19, HBF^{+}10, HG10, MS13]. Energy [MMSC17, SvdMK^{+}12, SPP18, TLB15, GLW^{+}17, HLL^{+}17, JR15, JCM13, KNY^{+}15, KBB10b, LKCH12, LRS17, LHW^{+}19, MM13b, MRC10b, NWC^{+}10, QNG^{+}15, RSDN15, SPM^{+}13, SNVH19, SNK^{+}10, TMB^{+}12, WMHW14, vDCFM10, vDCR13]. Energy-based [SPP18]. energy/predation [vDCFM10]. enforcement [BS15]. engineer [BPB18, BNW14, GHGF12, GKHS16, RTD14, SP13b, YHR16]. engineering [AAIB15, BB17, Ber12b, CAD^{+}10, EBT^{+}15, LC18, PRA^{+}19, SLT^{+}19, Sti13a]. engineers [CFH11, PRB^{+}17, Wil14]. England [BG19, Dan11, FMH^{+}15, FGW^{+}15, GAB11, GSMK15, OHP18, POS14, RWO13]. English [BLHL15, FDDD13, KHG^{+}19, LPC19, MVE^{+}10, NFR^{+}14, SMND16, TKW^{+}14, XTW^{+}15]. Engraulis [CCG^{+}10, FKZTT18, QLCUG15, RLN^{+}13, TGSM12, WB12, WM10, vdSRT^{+}18]. Enhalus [PBC^{+}15]. enhance [ACS12, Ber12a, GSF^{+}13, LH19, WAA^{+}19, ZZG^{+}19]. Enhanced [DMW^{+}16, JWR^{+}10, LCM^{+}16, MAL15, WHP^{+}17, AHS^{+}13, BMTC^{+}15, BQRB12, DBDK14, OHB^{+}15, OHP18, PMS15, SSAB^{+}15]. Enhancement [LMW12, LÁBFÁF^{+}10, ODGL^{+}12, SYX^{+}10]. enhances [AM18, EBT^{+}15, GME^{+}19, LCD11, PHI^{+}12, SYK^{+}13, WBLS^{+}11, WS17]. Enhancing [JPF12, ZSEV17]. Enhydra [MBB^{+}12b]. enriched [AGNH^{+}19, FRC13a, SBMÅ12, TRR^{+}17, VQKF14]. Enrichment [NTN^{+}13, SUPJ^{+}10, AS14, AS15, AHS^{+}13, ACG12, BHF^{+}10, BLHRP^{+}19, BPF15, BPM^{+}18b, CRS^{+}18, FCR14, FC12, GP11a, GEF^{+}12, HHH^{+}11c, KdWV^{+}18, KVY^{+}13, KPB16, LG18, LFD13, PFB^{+}13, SBB^{+}14, SO16, SCW^{+}15b, TPF^{+}13, TVRG^{+}13, VKH10, VBGO^{+}16, VDEO18, YKBB12, dlSGS17]. Ensemble [ZPPM17]. ensembles [FHH^{+}17]. ENSO [SBE^{+}14]. ensure [PAFC13]. entanglement [KHM^{+}12]. enticing [RKFD15]. entire [PMOAC^{+}17, dJSB19]. entities [War14]. entoproct [SIHH10]. entrained [Sat10, SLC11]. entrainment [PZA^{+}13]. entry [CWMP13, SCAM^{+}14]. envelope [Pil10]. Environment [CMAJ12, TCJ^{+}19, Ano18b, BA13a, BGL^{+}11, BBB^{+}14, BMV^{+}11, CSF13, DS13, EPS11, FRV^{+}13, FNA^{+}16, FHG^{+}14, GSM^{+}11b, GRD11, HS11b, HBK^{+}12, JLM^{+}12, JWB^{+}17, JAPK16, KW12, MCC^{+}16, MCRB11, PFG10, PRM17, PBM^{+}18, RMHP14, SUPCB11, SGMG^{+}14, SMBH15, SKA^{+}14, SBA12, SE16, SLSR10, TGP^{+}16, VHV13, WH11, WRDF12, YTO^{+}11, vERRP16]. environment-driven [EPS11]. Environmental [AN11b, AGM^{+}11, BBM^{+}14, BBF^{+}19, BFDM18, BJF17, BCH^{+}11, CSRD11, CAC11, DCB^{+}16, ETB^{+}11, FH10, FSS^{+}18, FSW10, HWSK19, HWE^{+}12, HvEBC12, HYL^{+}13, KPP^{+}16, LST11, LRHP18, LSH^{+}19b, LGT^{+}14, MOMF14, áNGGS11, PNA^{+}16, PBM^{+}10, RC12, RHW12, RGDS12, RFRD10, RLH^{+}19, SWM^{+}19, SG12, TBH^{+}18, VM10, WS12, YHTC19, ALWD15, AFSL^{+}15, AFLM12, AED^{+}12, Ano17c, AB12, BDNA10, BB17, BPWF14, BHT15, BDPL19, BT16, BD12, BCSB13, BDR19, BWCZ11, BWRP14, BMF^{+}15, CED^{+}19, CPF^{+}13, CGS^{+}19, CLB^{+}18, CRRL17, CO16, CKR^{+}11, CFH11, DLRF15, DWLH13, DHK^{+}17, DLC^{+}13, DDD^{+}11, DMT^{+}14, DPD^{+}12, ECZD14, FKT15, FTF15, FGL^{+}12, GMB^{+}16, GAO14, GW15, GS16, GHT^{+}15a, HER^{+}19, HS17, HS14b, HA14, HBG^{+}17, JUMS10, JMB^{+}17, JWR^{+}12, JCZ^{+}12, KW17, KSF18, KSK07, KHD19, KKS^{+}12b, KMSW10, LdU^{+}17, LFFL15, LMD12a]. environmental [LRB^{+}12a, LE10, LEKT12, LPK^{+}16, LCL^{+}13, LRQ^{+}12, LKN^{+}14, MLR17, MAT^{+}11, MH10, MN18, MA14, MM16, MMSC17, MWGH15, MPG^{+}12, MRC10b, MRC^{+}13, MGG^{+}14, MCTHGV12, MWMD17, NBGGST^{+}15, OOK11, OBÁSV18, PGR^{+}13, PDGR^{+}13, PSG^{+}13, PHHL19, RPRN^{+}16, RAA^{+}16, RTÁB^{+}14, RMME10, RRB13, RMBM14, RDC^{+}12, RPM^{+}13b, SHUS17, SPJ^{+}12, SPH18, SWR11, SDAM11, SGTD14, SSB^{+}16, SGP^{+}17, SHAS15, SBE^{+}16, SPS16, SHB^{+}13b, SMF10, STM^{+}14, TKA10, TWY^{+}19, TKIO10, TPCN12, UCK^{+}13, VANB15, VTSH^{+}19, WHH^{+}13, WKK18, WK18, WJHS13, WL11, YMCV10, YBD^{+}16, ZHTA13, ZVR18, ZBC10, dJH11]. Environmentally [GOG14, RCAB^{+}12, TTM19]. environments [ÅCAC16, CPF^{+}13, CPPM12, CMW^{+}14, CCKB^{+}15, DZB^{+}18, DHR^{+}13, GBC^{+}16, HGW^{+}11, KHS10a, LRB^{+}18, LWM15, LMTD16, OTK^{+}19, PXG^{+}10, RB17, SHT^{+}16, TPKC18, WSHY16, WS17, WR12]. ephemeral [AS15, PHM11]. ephyrae [CL17, FSM^{+}14]. epibenthic [DB14, DHJ10, LLS^{+}18, MKPS13, SS11b, vdVFK^{+}10]. epibiont [MUC12]. epibionts [GCCCM16, GCCM18]. epibiosis [Mou17]. epibiotic [CESAM12]. Epidemiology [EKR^{+}10, HGG^{+}10]. epifauna [NT10, SCL13]. Epifaunal [DSJNH18, KHD19, CLS19, JOFP18, LJK^{+}11, NBGGST^{+}15, RWO13, SJHH^{+}17]. epilithic [BTM^{+}18, JUMS10]. Epinephelus [NMS^{+}18, RTM11, SDST16]. epipelagic [HT15]. epiphyte [CSC13a, CVP11, MH13]. Epiphytes [MMPH^{+}12, CESAM12, JS10, JREVB18, MHS14]. epiphytic [BHF^{+}10, BT14, BCM11b, JT10b, PNPF14, PW19a]. episode [MSGGE10]. episodes [SLL12]. epizooic [DRG^{+}12]. epizootic [WVM^{+}14]. EPS [UFP^{+}10]. equally [OPSM13]. equation [HMM^{+}15, Pro17]. Equatorial [OMT^{+}17, DR10, HMO^{+}12, LSY11, MC16, RCP^{+}16]. equilibrium [CLW^{+}11, RP13]. equina [DMK11]. equivalence [CGK13a, DPM15, She12, She13]. era [DS16]. eradication [MK13]. ercodes [HMJ^{+}13]. Erect [BTM^{+}18]. eremita [DMT^{+}14]. Eretmochelys [BP12a, GLW^{+}12]. Erignathus [BCH^{+}18, GWK^{+}19, HLF^{+}19]. erodibility [BPB18, HPH^{+}15]. eroding [CFSM^{+}13]. erosion [DdR10, HBB^{+}12, MCK^{+}13, RTDUM16, SGTD14]. Errata [Ano12d, Ano14b, Ano15a]. Erratum [ACA^{+}10b, Ano10d, Ano10e, Ano10f, Ano14c, Ano15b, Ano16, BW16b, ERB^{+}13b, RUP11b]. Erring [AvdHH^{+}17]. error [KCCG12, SN17, SJL13]. errors [HRT^{+}14, LMR^{+}10]. eruption [MMBA13, ODL^{+}13]. erythraea [CLC12a]. erythrinus [FCR12]. erythrogramma [FDP^{+}16]. escalated [Fis10]. escape [EN11, Joh10]. Escaping [CPM^{+}17, ETS^{+}11]. esculentum [FWS^{+}18]. essential [AHD12, CLC12a, DE17, Har16, LCR^{+}16, PWO^{+}16]. EST [VBV^{+}10]. EST-microsatellites [VBV^{+}10]. established [GPM^{+}13, KCS^{+}17, MMD^{+}13, SK13]. Establishing [GSJS13, MDKM13]. Establishment [CVM^{+}17, HFH10, BBH^{+}11, CvOB18, FAM^{+}17, FTWA17, FSH^{+}16, MO12, MKK^{+}11, ZBY^{+}14]. estimate [ADD18, ARvdW^{+}16, AGNH^{+}19, BWM^{+}15, BC15, BD11b, GUC^{+}15, MTMM^{+}15, NCC^{+}14, RS14, RST15, SKW^{+}14, TLB15, TSTP13a]. estimated [BGDC14, BRB11b, HBM^{+}15, RAB^{+}15, TS11b, UTS^{+}18, VBR^{+}14]. Estimates [SGD^{+}12, AWH19, BBG^{+}16, BM11, CSC13a, ET11, GED^{+}14, GS15, HS17, HSMB17, KB12, LPC^{+}11a, MBKM17, POX^{+}17, SPWK13, SJLH12, TNM^{+}11a, WEK14, YPWP18, vdMPM12, STBP13]. Estimating [ALS18, BBRM14, BRB11a, CCHG10, FBPCC^{+}19, HLS^{+}15b, MC12a, MO13, RBD^{+}14, SH12, TBH10, VK11, WPNK12, WFH^{+}18, WMHW14, ZBB^{+}19, ZB16, BFM^{+}13, GEAOH17, JHNL12, KHI12, LBH18, NUJ12, SL15b, SEFS15, STRP18]. Estimation [ARG^{+}10, HRT^{+}14, Mat10b, ZSSK11, CPB^{+}11, DBH^{+}16, IKS^{+}12, JMO^{+}19, NKM^{+}15, PBM^{+}11, PRNR^{+}18]. estimations [TMM^{+}19]. estimators [LHDH17]. estuaries [BSB12, BD11b, Bri12, BBMM14, CWEE19, CHG^{+}18, DHJ12, DVB^{+}16, FWL15, GLF^{+}17, GS16, IAV^{+}10, LD17a, LMT15, ME11, MV12, MGCH^{+}16, NDBL16, OCS14, OCSW16, RSTV^{+}13, RA10, SMSE19, SJC10, SDM16, TG10, VQKF14, VGBJ^{+}13, VJG^{+}13, dJTH^{+}14]. Estuarine [DMC^{+}16, HS11b, LEB14a, VC14, AAAE17, ACB12, AMAM^{+}18, BRK^{+}12, BA12, BAL14b, BCC^{+}11, BDT^{+}11, BT13, BGO^{+}18, CLL15, CGD10, CSM^{+}19, CdTK^{+}16, CLG^{+}19b, CPR^{+}16, CLB^{+}18, CMW^{+}15, CCKB^{+}15, Det10, DTM^{+}12, EH13, EBH17, FWF^{+}10, FSW10, FAWBF10, GOC^{+}18, GS17, GLB^{+}10, GB10b, HBS17, HLS^{+}15b, HS15b, HOA14, JCM10, JSCM15, JE10, JGS14, JAPK16, JCM13, KdWV^{+}18, KSK07, LB15a, LKC^{+}17, LKCH12, LEB14b, LWG14, LDT^{+}13, LHT^{+}17, MC10a, MSC13, MPM^{+}14, Mar17, MH12a, MHNL14, MB18, MMM^{+}19, MS14, MGOL19, MDE13, MSV^{+}11, MCT15, NCD^{+}12, NDG15, NBM^{+}14, NDBL16, NSI^{+}15, OFFQGG10, OHR^{+}13, OFT^{+}11, OLM^{+}12, PDL^{+}10, PS19, PL10, PP13, PPS^{+}13, RSTV^{+}13, REG^{+}11, REBL16, RGWA11, RCU^{+}11, RTC^{+}14, RMV13, SRSH^{+}18, SF16, SCTA10, SQS^{+}13, SS16, SLE^{+}13, SMR10, TLS^{+}19, TRSV^{+}12, VSM11]. estuarine [VJJ^{+}19, VMN^{+}12, WEF19, WWG12, WDL^{+}18, WWGD13, WD14, WPP11, WS11, WWZ^{+}19, ZGK^{+}19, vdHTT^{+}15]. estuarine-associated [MHNL14]. Estuarine-coastal [VC14]. estuarine-dependent [BA12]. Estuary [BP12b, CR14, CP10, DPC13, LB15a, PPM11, RWR^{+}15, SWM^{+}19, SIK16, ASI15, AR14, ASCHM19, ADB10, BGM^{+}18, CCNB15, CAG13, CHFS12, Dan11, DBCV^{+}12, DPM15, DDP15, DAD^{+}11, DCW^{+}14, FHE^{+}15, Gal14a, Gal14b, GVW^{+}12, GP11b, HP11a, HTB^{+}17, HPMZ11, HHWW11, KN12, KMV18, KCS12, LBAK13, LGD10, MPB15a, MTM^{+}11, MBKM17, MAHK12, MMEM14, NMCP17, PAF^{+}10, PS15b, PRR^{+}16, Pro17, SPMN19, SPM^{+}16, SS16, SAD^{+}16, SLR^{+}14, TYG12, TNM^{+}11b, VGBJ^{+}13, WH11, WJHS13, CPP10, GK10, GSTK11, GLM^{+}19, KK17, KKY15, LRTLF12, MS13, PPJT15, PPB10b]. estuary-dependent [CCNB15]. ether [MHN^{+}12]. Etrumeus [HHNI17]. Eubalaena [KHM^{+}12, LRR^{+}10, MRB^{+}11, MBP^{+}12, RSP^{+}16, RUS^{+}13, VRSS18, ZBK^{+}15]. Eucalanidae [CWD11]. Eudendrium [RBBG12]. Eudyptes [WHPT12]. Eudyptula [ALWD15, PHHL19, SDCC12]. eukaryotic [HRPBLT18, KdVvdP^{+}12]. Eulerian [BDS18a]. Eumetopias [SDR^{+}15]. Eunicella [BAFP^{+}13, FMC^{+}11]. Euphausia [CDCZ16, CDC18, GGK^{+}16, JR15, KTF10, MKM^{+}16, MAS^{+}10b, PKS11a, RWG15, RBEGG13, SRLV10, SKD^{+}17b, SGCF^{+}12, SKR^{+}13, THP^{+}16]. Euphausiacea [GGMGR12, HS14a, KHS10b]. euphausiid [AAGGFG^{+}12, PKS11a, RDM18, SB14]. euphausiids [RDK14, SDKD13]. euphotic [JPMG10, LSY11]. Euprymna [SWNSF14]. Eurasian [DLRF15, KMV18, MJS17, MV12, SSS^{+}10]. Euro [Chr17]. Europe [BFH^{+}13a, CBC^{+}16, GBS^{+}10, PPTB12]. European [ADB10, BBR12, BFG^{+}11a, BFG^{+}11b, BANE^{+}13, CRAP^{+}12, CSR^{+}14, CLM^{+}13, CKM^{+}16, CTB13, DMD^{+}12, DHJ^{+}17, DT15, EHD^{+}17, EKR^{+}10, HVD^{+}16, JBCGC19, JMB^{+}18, LANE18, LMCD19, MMG^{+}16, MSG^{+}13, MBV^{+}11, MOS10, MOK^{+}11, MUOS13, PSdA^{+}17, PMM^{+}15, RLN^{+}13, RVH^{+}19, SAL^{+}12, SRM^{+}15, SABMG16, TMH^{+}18, TGSM12, TBMdV12, WSL^{+}14, YBH^{+}11]. euryhaline [BSC17, EPF^{+}17]. Eurytemora [PFG10]. eusocial [MH11b]. eutrophic [AGM^{+}11, CSC13a, Gal14a, Gal14b, JCZ^{+}12, MMEM14, OV10, PAF^{+}10, SDMB15, SAD^{+}16, VQKF14, ZGK^{+}19, ZZG^{+}19]. Eutrophication [BRIC13, PLG^{+}16, GM10, HTT^{+}12, HHWW11, Joh11, LSD^{+}13, MLC14, NGP^{+}15, SKB15]. euxinus [SHI^{+}10, YBH^{+}11]. evaluate [HW12a, MB11, SGGL15, WRP^{+}17]. evaluated [SKD^{+}11a]. Evaluating [AJ16, BSB^{+}11a, BWRP14, CPB^{+}11, DTG16, PBH^{+}13, RFM^{+}15, RVPG17, SRSH^{+}18, WAW^{+}12, WRFM16, WPP11, HA15, LET^{+}12, MM11, SESS^{+}18]. Evaluation [BD13, HCGP13, RWM18, SWK19, SMBC18, CC12, DFLC17, EAMT10, GGF^{+}12a, JBD10, WYS^{+}10, YSC^{+}15]. evasive [RSH^{+}19]. Evechinus [GFP19, RP17, WW15]. even [VHBG15, YSS^{+}18]. evenness [RMSPA16]. event [BGM^{+}18, BP12b, CMGS19, DAWS19, FPC^{+}12, FFS^{+}15, MH12a, NWHD15, NTB^{+}17, PBM^{+}10, TNS^{+}12]. events [AZSKWL11, BSG^{+}10, BCW19, BAS16, BOKJ18, BPB^{+}12, BPLT14, FPR18, GD15, MPM^{+}14, SMOH19, SMM^{+}14]. Everglades [RJB^{+}14]. Every [MHPvdH19]. Evidence [BSH^{+}18, CMW^{+}14, CMGS19, FS18, FBNJ13, GOM^{+}19, LP10, MKPS13, NKA16, RMHP14, VBD^{+}11, WTD^{+}19, AAK^{+}15, BWS11, BD14, BSCH14, CWEE19, CKK^{+}10, CFSM^{+}13, CL10c, CBC^{+}16, DB14, FCV^{+}16, GDB^{+}16, Gar13, GGPDS19, GMR13b, HBI^{+}18, HT10, JDD^{+}13, KHF^{+}11, KNB^{+}13, LSY^{+}18, LJB14, LP12, MJ10, Met11, MBRO10, MER14, ODL^{+}13, PMEZ15a, PMEZ15b, PPH11, PBB^{+}15, RGW17, RWR^{+}15, RTD14, RKPB^{+}15, RKM^{+}10, SKD17a, SMVC^{+}15, SØK^{+}15, SGS^{+}19, SE11b, SBT15, SS13, SWNSF14, SOA^{+}17, SV15, SWPK13, SKD^{+}11b, TSS11b, TFM^{+}18, TKDM16, TCWC13, VCG^{+}10, VMW^{+}10, WRB^{+}15, WRW^{+}16b, WS10, WDM^{+}18, YNK18, vdBUPP^{+}11, RPR16]. evidenced [CRHV14]. Evolution [BS19b, CB17, FSB15, KBH11b, MCTHGV12, PCH11, FTKK14, MDR16, MPF^{+}15, TF12]. Evolutionary [AMC^{+}11, OBM15, SH10, APC^{+}14b, LGBSC11, LEKT12, OT13, Pil10]. evolve [ZJ17]. ex [ROE17]. exacerbate [TPKC18]. exacerbates [KMV18, TBL16]. Exallias [Car12, ZAZS12]. examination [BBKH^{+}11]. examine [BA12, BFEMLÁJ19, FGTS16, OALK14]. examined [OFDG14]. Examining [TdRH17, HKB^{+}19b]. example [DSG15, GMES18, LBCS^{+}15, ZBB^{+}19]. examples [Pur18]. exceeds [SGTP16]. exceptional [LRC^{+}13, PBM^{+}10]. Exceptionally [LSFQ19]. excess [vdPLdV^{+}11]. Excessive [SJL13]. exchange [ASK^{+}15, DDB^{+}16, GBH^{+}10, GJSR18, LGC10, RWR^{+}19, RDPR15]. exchanges [FT10]. Excirolana [MAD17]. excluding [FMC^{+}10]. exclusion [BMMS10]. Exclusive [LCC^{+}11]. excretion [FUOG^{+}16, LTSDHL12]. excursions [JRW^{+}14]. exhalant [RJL^{+}11]. exhibit [BPT15, GWM^{+}15, SJB16a, SJO16]. exhibits [DAD^{+}11, WHK^{+}11]. existence [BA13a, PRP^{+}18]. existing [BC14, KHBL11, RPL^{+}18]. exitiosa [HSMS^{+}15]. Exmouth [MMC^{+}18]. exogenous [MM13b]. exopolymer [BZBR^{+}11, HSR^{+}10, KCP^{+}11, KKO^{+}14, MSC13, PPB^{+}10a, YYU^{+}17].

Exotic [JL12, BB19, Gre10, LCFR10, MK13, PH16]. exotics [DS10]. Expanded [SBB^{+}12a]. Expanding [Pil10, TSA^{+}18, APBC^{+}19, BL14, EvDB^{+}14, MH19, TNIH12]. Expansion [PMM^{+}15, BLJP14, BB14c, BGK17, BOU13, DHJ^{+}17, DMHS^{+}14, FCR14, GALS16, GCR16, HI14, JC17, KMP14, MZF^{+}19, NdBKR13, PAP^{+}12, SPSG14, SHF^{+}12b, TJS17]. expectations [SJL13]. expected [BOC^{+}16, TSCSB15]. expenditure [WMHW14]. experience [BNS^{+}14, DHM^{+}14, MPB14a, OAGV11]. experienced [KSK07, SCP^{+}14]. experiencing [DHJ12]. Experiment [SAM14, ALANGT10, AHPKS17, BWH11, BSCH14, CSRF16, Gar13, JMO^{+}19, LFL11, MJS17, MAT^{+}11, PFL18, RZB18, SRR15, SMK^{+}18, SF14, VMW^{+}10, dMBC12]. Experimental [AS13, CL10c, DPS16, DFLC17, EAMT10, GFC^{+}14, GA11, HN13, MBW^{+}19, SLL12, TNS^{+}12, ZHJ^{+}17, BCOMKJ18, BD12, BVGH10, CC12, CLSS17, FRRÁS^{+}12, FC12, FBET15, GSCH14, HHH11a, ICUM10, KdWV^{+}18, KVY^{+}13, ML11, MD16a, NHAS15, OSPWB14, PMS10b, RMC^{+}14, SBCA15, SBT15, SGTP16, SLO19, SCF^{+}17, TPSS12, VCFF13]. Experimentally [PSL13, BFKR17, HSZR10]. experiments [BHF^{+}10, BH12a, GBR11, JU16, KBA13, LGL11, MWA17, OMZ10a, STT14, SV15]. explain [BWM^{+}15, BVG13, BSB^{+}11a, DBE^{+}18, FNC13, FPM14, LSN^{+}19, MMA18, MJA^{+}18, STS^{+}18, TPBM13]. explained [STK^{+}13, Tho15, VB14]. explains [BMT17, DXF^{+}18, Hos14, WWG13a]. explanation [RL18]. explants [KRC17]. explicit [GDA^{+}18, WMPM13]. Explicitly [DC17]. exploit [SJL^{+}15, WVD^{+}17]. Exploitation [PDL^{+}10, APC^{+}14b, BGSM13, BHDC15, DDDR10, WEA^{+}17, vdWHG11]. exploited [BBB^{+}15, BOÁNF19, CBGDS12, LdU^{+}17, MdSW^{+}16, MA14, NBM12, RFRD10, SCB15, TM10, WLP^{+}10]. Exploiting [LNF^{+}11, JDD^{+}13]. exploration [FCRD18]. explore [GL12]. explored [FEE^{+}11]. explores [Ano18a]. Exploring [BMV^{+}18, CBR^{+}18, DVB^{+}16, USR^{+}16, WLP^{+}10, Ano18b, RMSPA16]. explosions [AvBBL^{+}16]. export [BVGH10, ETS^{+}11, JPMG10, LBN11, LLP^{+}10, OT14, SOBNL13, WS10]. exported [Sti13b]. exposed [AvBBL^{+}16, ACW18, ATO^{+}14, BVGH10, DWL19, HPH^{+}15, MBC18, ODCE11, PK15, ST13, WADKT13]. Exposure [HPQ15, LBM^{+}18, LRR^{+}10, NDAP19, TPKC18, ASAG17, AS13, ARG^{+}10, BJM19, BLD^{+}16b, BPWF14, BRG^{+}14, BMVBC10, BMR17, BBC^{+}12a, BSG^{+}17, BLJP14, BCM11b, CMR12, DF15, FPC^{+}12, HGS14b, HSC^{+}15, HGC15, HPP^{+}10, KCMH^{+}17, KS13, KOPW15, LPJ^{+}13, MWC^{+}12, MW16b, NCR^{+}14, NMHB16, OVF11, OCMT13, PNF^{+}12, PJDC^{+}13, PBW^{+}11, RP17, RBS^{+}17, SHB^{+}10a, WMMC14, WWG^{+}13b, HGS14a]. exposures [BPT15]. expressed [WJWC16]. expression [BMD^{+}11, BBC^{+}10, GAF^{+}16, GB16, HHWW11, KP18, OTS^{+}10, PZL^{+}17, SGCF^{+}12, WNF^{+}11]. extended [BRK^{+}12, RTMJ11]. Extending [EM11, HB12, BT16, WLP^{+}11]. extends [KTNO19, MFDW17]. Extension [DHF^{+}13, WNG^{+}19, RRRF^{+}17]. Extensive [CSQRA11, HCL^{+}11a]. extent [PSHG16]. external [LBAK13, VSG^{+}15]. extinction [HHS12c, PBC^{+}12b]. extirpation [Ing16, WLB^{+}11]. extracellular [KCP^{+}11, MPBSB18, UFP^{+}10]. extractable [BD11b]. extraction [CHFS12, FRJ10, KKY15, LMR^{+}10, SMBC18]. extratropical [GDH^{+}19]. Extreme [FTS^{+}13, LMZ16, SGT12, SB15b, vBAP^{+}19, BBB^{+}14, CED^{+}19, DAWS19, FC19, HTHM18, LHJ11, MPM^{+}14, MH12a, NWHD15, NTB^{+}17, RBL^{+}17, SUPCB11]. extremely [BAS16, BS19a]. extremes [DCW^{+}14]. Extrinsic [JMB10, GDB^{+}16]. exudates [SPSP10]. Exudation [LSRRCM13, GRG^{+}12]. exulans [CXdG^{+}17]. Exxon [BBM13, BBC^{+}12a, HG13]. eye [MBC^{+}11, RST15, SST19]. eyed [KHW^{+}12, KCE14]. Ezo [MKK^{+}11].

fabricii [GCS^{+}18]. face [PSP^{+}19, SBH^{+}19, WHB^{+}19]. facilitate [DS10, Goc10, ORJH19, PTP^{+}15, RdGv^{+}18, WO11, ZF19]. facilitated [NdBKR13]. facilitates [BRBH19, CMH14, GTF10, OMM^{+}12]. Facilitation [BBTW14, NSI^{+}15, SGEO13, TWM^{+}15, ATP14, Ber12a, ESBG14, GAO14, PRB^{+}17, TS12, dMA18]. facilitators [MM11]. facing [PGR^{+}13]. factor [BLV^{+}19, GALS16, LRQ^{+}12]. Factoring [GVW^{+}12]. Factors [BG19, BW13, BCHG18, CAB^{+}19b, CEW15, DAC^{+}12, GFKC13, GMSL^{+}14, SHK14, WLB^{+}15, AFSL^{+}15, AOW12, APC14a, AWJ12, Ano17c, ASCHM19, AY10, BW16b, BW16a, BBF^{+}19, BWRP14, CZB10, CHP^{+}11, CRRL17, CM11, CGEGC^{+}19, CAGS17, DLC^{+}13, FH10, HBS17, HWR^{+}14, HYL^{+}13, HHWW11, JMB10, KAAI15, LSN^{+}19, LdU^{+}17, LMd12c, LCL^{+}13, LRC^{+}13, LKN^{+}14, MOMF14, MLP19, MYS^{+}16, MGG^{+}14, MN10, PML^{+}10, PH11, SMBH15, SHAS15, SCW^{+}15b, TKIO10, VLK^{+}15, WS12, ZVR18, dJH11]. Facultative [MBvR12, NVA^{+}19, SW18, LFPC^{+}14, NFN^{+}15]. facultatively [BFKR17]. FADs [FCFD15, RDF^{+}13, RTCF^{+}17]. faecal [SRRS12, SMR14]. fails [GSF^{+}13, OHB^{+}15, RPHB15]. failure [GGB^{+}18, KHW^{+}12]. falciformis [FCFD15]. Fall [JCM13, ERB^{+}13b, ERB^{+}13a, EZZ^{+}12, HFR^{+}18, HCJ^{+}11, LRQ^{+}12, SCAM^{+}14, SBB^{+}14, SOT15, VKY^{+}12]. fallen [HED^{+}13]. falls [AJF^{+}12, HBE^{+}13]. false [LMBM18]. Family [BST19b, DSS10, FTKK14, KSFL14]. famine [CIM^{+}13]. fan [Kat16]. farm [BSB13, BDBN11, DTC^{+}17, SSvD^{+}15, WN16, vDGK^{+}12]. farmed [SMK^{+}10, TWM^{+}15]. farming [CHDH^{+}19, LPK^{+}16, MRC10a, áNGGS11, SWWK13]. farms [BDD^{+}18, CLD^{+}13, HMH^{+}18, TRSB^{+}18, TPF^{+}13]. Faroe [BDS^{+}18b, ERG^{+}11, JGL12, áNGGS11]. Fast [ATT^{+}16, DHSL11, FKA^{+}12, PBF10, TFH^{+}19, WEA^{+}17]. fast- [PBF10]. Fast-spreading [ATT^{+}16]. fast-warming [TFH^{+}19]. faster [JTWJ^{+}13]. Fate [HCAV10, DSK^{+}10, KAMA13, KS12c, MMRDS18]. Fatty [BMJH12, CCG^{+}10, FRGI15, GNHLN11, KKMA^{+}10, KS12a, RVF14, THC12, AR14, BCGWS19, BPH^{+}19, BADQ^{+}15, BNE^{+}19, BDS^{+}18b, BGS^{+}16, BFEMLÁJ19, CMIMGS15, CWD11, CLC12a, CMW^{+}15, CL10c, DBC^{+}12a, DSG^{+}13, DE17, GED^{+}14, GHNF10, ILDL10, JBS11, KJW^{+}15, KKS^{+}14, KBP^{+}10, LLv13, LAL^{+}18, LFM^{+}16, LBL^{+}12, LBE^{+}10, MVP^{+}16, MHHF13, NLK^{+}19, OBI^{+}16, PLG^{+}19, PBM^{+}10, PBAP^{+}14, PIPD10, PFG10, QSB16, RLG14, RCR^{+}13, RKM^{+}10, RT12, TTS17, TFM^{+}18, VBT^{+}17, WBI^{+}15, WNNN12, WVD^{+}17, ZdVA^{+}16, vdBUPP^{+}11]. fauna [ACF^{+}10, ALSO^{+}18, CM11, CKV12, DIG^{+}11, DLRF15, HTR^{+}11, JCA13, JPI^{+}19, JNN12, KN12, KPLD^{+}12, LJB14, MVVGAV10, NCR^{+}14, RdGv^{+}18, SCP16, SJL^{+}15, TVH^{+}10, XWE^{+}14, vMVSW12]. Faunal [KSF18, SRH^{+}15, SLRS13, BH16a, DMJ11, HPA17a, KLSJ10, MNKB11, RDC^{+}12, ST13]. faunated [VKH10]. faveolata [MEKC15]. Favia [HRC12]. favorable [BMTF10, LdU^{+}17]. favouring [ASEB11]. feasibility [JMK^{+}18, PGVF14]. feast [CIM^{+}13]. Feather [WYI^{+}16, CCJR14]. feathers [WKK18]. features [AMM^{+}12, BT10, BST^{+}19a, FHB^{+}12, GCRP15, HS15a, HSRP11, KLKM12, LGM^{+}12, LFG^{+}12, PHHL19, RT17, SMD^{+}16, SRH^{+}15, TFB^{+}17, vdMPM12]. fecal [GSD12, PMBH11]. feces [ELM^{+}19, vBJV^{+}15]. fecunda [CMP^{+}13]. fecundity [BFS10, CG17, CS11b, DvBH12, HRAM13, JW10, KWM12, SNK^{+}10, TPB^{+}16]. fed [ACG12, LHW^{+}19, SD17b]. feed [DMH15, MP17]. Feedback [ZFC^{+}13, KWS^{+}13, OSK15, SS12b]. feeder [PKBS14, WS17]. feeders [RVF14, VDM^{+}10]. Feeding [APNC^{+}15, ASWB10, BDS^{+}18b, BB14b, Car12, CXdG^{+}17, CDC18, DCHI12, DGFH10, GLF^{+}17, GFL^{+}17, JLM^{+}12, MHW^{+}12, MGAV15, NBGS^{+}12, NPRAGG^{+}12, OBI^{+}16, PW16, PMBH11, RHVRC11, RLNLAF17, SLNG18, TUK13, VCRUQ15, ZAZS12, AJF^{+}12, AEGH11, ATP^{+}12, ATO^{+}14, BBC^{+}13, BLMD^{+}11, BCM^{+}11a, BMJH12, BS12, BD19, BCPC18, BSD17, BQ16, BS19b, BL10, CPF^{+}14, CLL13, CCJR14, CRL13, CDR12, CBL^{+}10b, CNÁCP12, CRF13, DVA^{+}18, DDP19, DFAH11, DPD^{+}12, DCW^{+}10, FC10, FSLG12, FTS^{+}13, FF10, FS18, Fuj16, GSR10, GDB^{+}16, GVR^{+}14, GKW10, GvSGCK14, GRCH^{+}18, GG12, GLW^{+}17, GHT15b, HJMN^{+}12, HTB^{+}15, Hoe10, JJP^{+}13, dDTC^{+}17, Kiø16, KHT^{+}17, KBB16, LBP^{+}19, LAL^{+}18, LCS^{+}10, LLBP15, LCB^{+}12, MBvR12, MGP12, MZC^{+}10, MTS^{+}17, MSH^{+}14, MAS^{+}10b, MD11, MPB10, MB18, MBV^{+}11, MCL^{+}17, NMC^{+}16, NHPW18, NPMS11]. feeding [NC14, NSY^{+}18, OSC19, OMSP10, OWN^{+}15, OWH^{+}19, PP10, PBV^{+}11, PDC^{+}17, PPH11, PDCH15, PBAP^{+}14, PFR11, RP13, RWG15, RSDN15, RL10, RL18, RLAGL12, SPWK13, SSM^{+}17, SSS19, SNVH19, SHC^{+}10, SF18, SV15, SSM^{+}19, SF14, TNM^{+}11a, TGP^{+}16, WCRCR17, WS10, WJ12, Wir12, WLM^{+}13, YAFV12, ZLGH16, ZRR^{+}13, dDTA^{+}18, vdGvdLvG^{+}19]. Feeding-related [SLNG18]. Female [GMR13a, HBCK13, ACG^{+}11, Col10b, FSF^{+}17, FHMG19, FKA^{+}12, GMR13b, HBHF^{+}11, HA14, HCW^{+}15, JT10a, LHHG11, MRLD19, NUJ12, NCK^{+}14, ORR^{+}14, OHST19, PVYS16, RB18, SL11b, SWNSF14, TKB^{+}13]. Female-biased [GMR13a, HBCK13, GMR13b]. females [CTRBLC^{+}13, NGSF13, SPM^{+}16, WMH17]. feminised [TMM^{+}19]. femoratus [NPRAGG^{+}12]. Feral [PSA17]. Fernando [PFVFF^{+}15]. ferruginea [PSA12]. ferrugineus [AZVB13]. Fertilisation [BCR15]. fertility [SK11]. Fertilization [MC10b, TKB^{+}13, BWJB16, BCW19, CSC13a, EKE16, KV13, ODL^{+}13, OBM15, SMK^{+}18, WMMC14]. Fertilizer [TKV^{+}19]. Fertilizer-derived [TKV^{+}19]. fetal [HAHM16]. fetch [Bur12, HPP^{+}10]. few [ECC^{+}15]. fibropapillomatosis [PDvG16]. Fiddler [HBHG10, KCCG12, RWR^{+}15, TB10, WRB^{+}15, WRW^{+}16b]. fidelity [ACP^{+}16, BSC^{+}13, BDB^{+}13, BMPH14, Dan11, GJ16, GGFB15, HVJ^{+}18, KCS12, LBWB^{+}13, LHHG11, MBB^{+}15, MHAP18, NTS^{+}18, OALK14, RHB13, SGM^{+}14, SCW^{+}18, TWL15, THB^{+}12, TS11b, TMS14, WTdB16]. Field [BM11, GBR11, GGPDS19, KAP^{+}12, KML13, AHPKS17, BHF^{+}10, BWJB16, BMTC^{+}15, CSKR10, DBA17, DL14, FWS^{+}18, GFR^{+}18, GC13b, JU16, KWM12, LH16, Lec18, LRK^{+}15, MJS17, MJ10, MSG^{+}13, MLB^{+}14, NIPS12, PDCH15, VMW^{+}10, WRW^{+}16b, MBS^{+}14]. field-caught [WRW^{+}16b]. fields [DABP13, HPH^{+}18]. Fifteen [KKL13]. fight [CSV^{+}15]. Fijian [CBGDS12]. Filamentation [HSM13]. filamentosus [GB12]. Filchner [PMIG19]. filefish [HMJ^{+}13]. filiformis [CPSC^{+}15, CPSC^{+}18]. filter [PKBS14, VJJ^{+}19]. filter-feeder [PKBS14]. filtering [PSL13, VANB15]. filtration [KRC17, PMYH13, WS17]. Fin [AdCM^{+}19, BMV^{+}11, DPD^{+}12, GVG^{+}18, GRN^{+}16, RMT^{+}13, SRP^{+}19]. find [BBWG12]. Finding [CWW^{+}16b, PK15]. Fine [AGK13, BST^{+}19a, BSC^{+}15, BRBH19, CDWC17, FCFD15, GMHS19, KW17, KHGS12, KBB^{+}10a, MCT16, PWM^{+}19, PCLB18, PDBB14, SRM^{+}15, TR13, VHHGF11, ZAC11, ZVL^{+}16, vERRP16, vdMPM^{+}10, BSB^{+}18, BT10, BPB^{+}12, BJS^{+}16, CLB^{+}14, DPH13, EMJ18, HSWS16, KLKM12, LMLB12, LR18, LCB^{+}16, LGT^{+}14, MPB14a, MRL^{+}17, MHS17, OWN^{+}15, SSR^{+}10, ST08, TPC^{+}16, TTB^{+}11, ZBC10]. fine-resolution [CLB^{+}14]. Fine-scale [AGK13, BST^{+}19a, BSC^{+}15, BRBH19, CDWC17, FCFD15, GMHS19, KW17, KHGS12, KBB^{+}10a, MCT16, PWM^{+}19, PCLB18, PDBB14, SRM^{+}15, TR13, VHHGF11, ZAC11, ZVL^{+}16, vERRP16, vdMPM^{+}10, BSB^{+}18, BT10, BPB^{+}12, BJS^{+}16, DPH13, EMJ18, HSWS16, KLKM12, LMLB12, LCB^{+}16, LGT^{+}14, MRL^{+}17, OWN^{+}15, SSR^{+}10, ST08, TPC^{+}16, TTB^{+}11]. fine-tuned [ZBC10]. finescale [PHHL19]. finfish [SGGL15]. fingerprints [HDB^{+}10, MDA^{+}10, MMB^{+}12a, MJV^{+}11, NDBL16, TMP^{+}11]. finmarchicus [BLS^{+}11, HTHM18, Ji11, KDS^{+}12, KDO^{+}14, MPZ10, PPJT15, PPH11, SKM^{+}15, SRRS12, SKD^{+}11b]. finned [HER^{+}19, MMFP^{+}15, OAB^{+}19, TFB^{+}17, WVM^{+}14]. Fiordland [DW12, PWB^{+}19, JW10, WJ12]. fireworm [MMCW14]. First [AHR^{+}18, BBG^{+}11, BQ16, FGTS16, HSMB17, KNB^{+}13, KB10, LOS^{+}19, MGM^{+}18, MAW^{+}11, TWY^{+}19, AFC^{+}16, BSB^{+}11b, BDD^{+}18, CM14, DHH10, DTL^{+}15, HLF^{+}19, HFH10, JKM11, KBG12, OMSP10, SH12, SDCC12, SKR^{+}10]. first-year [JKM11, SDCC12, SKR^{+}10]. Fish [AS12, BDVP19, CDCS12, LWP^{+}13, LCG^{+}17, LEB^{+}19, OPSM13, PRFA11, RM13, RCGE12, RPL^{+}18, RNSA15, SSPKF16, TNIH12, AHA^{+}15, AOPMOS10, APC^{+}13, AGGAO^{+}18, ADL^{+}18, AL12a, APC^{+}14b, ACW18, AAGGG19, ASCHM19, AHR^{+}18, BBM^{+}14, BSB^{+}18, BI15, BA12, BAL14a, BAL14b, BMM17, BNAN12, BFDM18, BCS^{+}12, BFD^{+}16, BWS11, BGSM13, BA13b, BFC14, BSB^{+}11a, BSH^{+}18, Ber12a, BSB13, BJH13, BG12, BDPL19, BDB^{+}13, BSC17, BS14b, BAFB12, BSG^{+}17, BCC^{+}11, BDT^{+}11, BLPP16, BGO^{+}18, BSD17, BMF^{+}15, BBKP10, BJL^{+}17, BGDC14, BBGL13, BBC14, BMA^{+}19, BVD^{+}14, BANE^{+}13, BL10, CDD^{+}16, CE14, CR14, CLDC^{+}13, CCM^{+}15, CGKW10, CGK13a, CGK13b, CMV^{+}11, CSM^{+}19, CSS15, CZB10, CL17, CPR^{+}16, CLB^{+}18, CCP^{+}14, CMW^{+}14, CAB^{+}14, CGEGC^{+}19, CMLC12, CV11, CNÁCP12, CBJB19, CATAV18, CBL^{+}10c, CEC^{+}18, DXF^{+}18]. fish [DGCH18, DPS16, DH18a, DMS14, DBS14, DPO^{+}17, DP14a, DBG15, DRD^{+}18, DDK^{+}18, DMH15, DGP^{+}19, DQC^{+}13, DVB^{+}16, DMM^{+}10, DBTV13, DHB18, DHJ10, DS13, DH17, DAB17, EBOM11, ETS^{+}11, EM11, EHG16, EAMT10, ERB^{+}13b, ERB^{+}13a, EFOM13, EVHA^{+}18, FDVL18, FTWA17, FJ11, FCFD15, FPC^{+}12, FFL10, FHL^{+}15, FBPCC^{+}19, FWF^{+}10, FSR17, FTvdK^{+}12, FHB^{+}12, FM14, FSDS10, Fuj16, GL12, GSÁCM^{+}16, GVW^{+}12, GRL^{+}14, GSP^{+}15, Gea10, GKW10, GOC^{+}18, GS18, GRBH15, GEA^{+}14, GGCB16, GSF18, GSU12, GKR^{+}15, GCDH19, GNBM11, GRN14, GGB^{+}18, HMH^{+}18, HCL^{+}11a, HSPS14, HPL10, HHS10, HTT^{+}12, HBS17, HP11b, HFM^{+}15, HS17, HOL^{+}17, HCJ^{+}18, HCL10, HVBG^{+}19, HRM^{+}11, HTH10, HSR^{+}13, HBM^{+}12, HM10, HAF^{+}14, HPH^{+}18, HDT^{+}12, HNG^{+}11, INvdB^{+}13, JHNL12, JCM10, JSCM15, JVSD10, JWBS10, KROA19, KEBB10]. fish [KHBL11, KBG12, KNY^{+}15, KPKA13, KMP11, KSK07, KNS^{+}15, KWDD10, KMO19, KJBS12, KMM15, KCE14, KFB15, LJ10, LCJ^{+}12, LAK^{+}19, LGB17, LTP^{+}18, LHDH17, LK15, LPI^{+}14, LSH^{+}15, LJM^{+}12, LMV^{+}19, LOBW^{+}11, LBJI14, LABG^{+}12, LBE^{+}10, LFG^{+}12, LYH^{+}17, LTN^{+}11, LANER15, LMT15, MBJ16, MK13, MC10a, MSFM12, MLPC18, MOP^{+}15, MMG^{+}16, MHGW13, MJ17, McC19, MSF^{+}19, MRC10a, MRL^{+}17, MJ10, MMA18, MDM18, MHVV13, MLB^{+}14, MMC^{+}16, MHNL14, MHQ17, MS18b, MBKM17, MMM^{+}19, MGB^{+}19, MPG^{+}12, MVOPB^{+}16, MRB^{+}15, MLR12, MGD^{+}11, MCT16, MKII19, MMGB14, MTG^{+}11, NCD^{+}12, NMPM15, NHS^{+}17, NFN^{+}15, NT10, NGA18, áNGGS11, NLLS15, OCPM12, OSD^{+}13, PBH^{+}13, PBM^{+}11, PHSBO16, PCP^{+}15, PH16, PSCS16, PDL^{+}10, PvdMS^{+}15, PS15a, PTL16, PMS10a, PMSGBL16, PFF^{+}17, PDCH15, PBAP^{+}14, PIPD10, PWW12, PLJ^{+}12]. fish [PPC^{+}19, PBWM18, PRA10, PAK15, PTF^{+}14, QHG^{+}19, RTMJ11, Rai12, RKFD15, RKD18, RSTV^{+}13, RHE14, RFRE19, REG^{+}11, RH12, RHH^{+}15, RZH16a, RST15, Rit17, RDF^{+}13, RTCF^{+}17, RCG^{+}15, RGB^{+}10, RCU^{+}11, RTC^{+}14, RSC14, RBD^{+}14, RBD^{+}16, RVH^{+}19, RFRD10, SUPCB11, SCMSB^{+}11, SHPS^{+}16, SBBB^{+}18, SKK^{+}18, SCW^{+}15a, SCST17, SL11a, SML^{+}15, SH12, SW18, SS12a, SJC10, She12, She13, SJB16b, SDAM11, SBM^{+}16, SWBMJ17, SSS^{+}10, SML^{+}14, SL15b, STSS^{+}16, SPW^{+}12, SA11, SRK11, SSM^{+}16b, SGS^{+}12, SSvD^{+}15, SJO16, SHB^{+}13b, SUPJ^{+}10, SCW^{+}15b, SKS18, SB15b, SS10c, TLB15, TLS^{+}19, TM10, TMSG19, TP16, TAB^{+}14, Tho15, TvdMM^{+}17, TPCN12, TDSS15, TB14b, TPF^{+}13, TVM16, VVB19, VM10, VJG^{+}13, VRAP^{+}14, VHKT15, WSLM13, WEF19, WWG12, WGD^{+}19, WvLF18, WFH11, WMB^{+}14, WAC^{+}14, WLP^{+}10, WWGD13, WS11, WWZ^{+}19]. fish [WLM^{+}13, WHL11, WMHW14, WLF^{+}17, YAL12, YRM^{+}12, ZBK^{+}14, ZJ17, ZKS17, tHHR10, vDJF15, vdWTWS15, vdBUPP^{+}11, vBAP^{+}19, vdHTT^{+}15, HHG10, GKRG^{+}19]. FISH-CS [HHG10]. Fish-habitat [CDCS12]. fish/invertebrate [SCW^{+}15b]. fished [OGCHS18, TMH^{+}18]. Fisheries [CS10b, GKRG^{+}19, AB17, AOPMOS10, ABH18, AMR^{+}10, AB12, BS15, BRJ^{+}12, BYAA13, BPM^{+}13, BPZ^{+}14, BBBW17, BLM^{+}12, BKAS14, CBMK10, CGBGS18, ECC^{+}15, FGL^{+}12, GBM^{+}12, GOM^{+}19, GSC17, GZYS11, Joh10, JMK^{+}18, JSK19, KGJ^{+}16, LGM^{+}12, Løk11, LHB^{+}15, LANE18, MHC11, MA14, McC14, McC19, MSF^{+}19, Nor13, OMW^{+}16, RWM^{+}10, RPP^{+}13, RvOP12, RGB^{+}10, SRG14, SWK19, SDSA14, TS15, TPC^{+}15, TGR^{+}19, WFS10, WPF^{+}16, WF18]. fishers [MHK11]. Fishery [CBC^{+}13, ARG^{+}10, BLD^{+}16a, BHE^{+}17, BMV^{+}18, BAS16, CLB19, CXdG^{+}17, CCNB15, CCG^{+}19, CSS^{+}13, FOL12, GHD^{+}10, HIMH15, JROS17, LKC^{+}15, LLMA^{+}10, LCG^{+}17, MLCC19, MA14, MMC^{+}18, MMCL19, MRH11, MPB15b, MHK11, MO13, OALK14, OGH18, RWM18, RCS^{+}17, RTM11, SBLM14, TSTP13c, TSTP13b, TRA^{+}14, ZS17]. fishery-independent [MO13]. fishery-induced [RWM18]. fishery-related [CXdG^{+}17]. Fishes [RSP17, AGK13, ASS^{+}13, Alb15, BGB^{+}17, BFNB16, BCW12, BST19b, BS16, BB11, BBB^{+}15, BDG^{+}16, BGS^{+}16, BKTD10, BBR^{+}14, BB19, CLB^{+}14, CPID13, CD13, Chr17, CFSM^{+}13, CCDP15, CBGDS12, CB12, CSRS15, CGM^{+}19, Cra12, DA11, DKAC^{+}12, DWK13, DA10, DGFH10, EW16, FC10, FWL15, GEAOH17, GSR10, GSS13, GWGC16, HMS10, HMKF11, HPN11, HDP^{+}17, IMD11, JGJ15, JD18, KSA12, KKA12, KWW^{+}14, KCW^{+}17, KA16, Kin16, KGLLW11, KMW17, KS13, LB11, LB15b, MCPC19, Mar17, MM13a, MH12b, MMM^{+}15, MWWS17, MBBK16, MZF^{+}19, MIV^{+}10, MRC^{+}19, NRF11, NM18, NHPW18, PSHG16, PM16, PS19, PSS19, RGWA11, RSV10, RGB^{+}10, RMRA15, SAWK18, SJB16a, SHK14, SF16, SATS13, SSB^{+}18, SLF^{+}19, SKMS^{+}18, SCL13, SMBC18, SS16, SS19, Sti13b, TUK13, TBDM11, WLM12, WH14, WAW^{+}12, YRM^{+}12]. fishes [ZMS^{+}14]. Fishing [DP16, HS17, MSE^{+}12, ST11b, YYO14, Ale13a, ALS18, AKF14, BS15, BDVP19, CRJ^{+}18, CWGG15, CV11, DKRS12, FHL^{+}15, FRJ10, FFS^{+}13, GL12, GKRG^{+}19, HRM^{+}11, KWW^{+}14, LJK^{+}11, LFFL15, MM16, MHFN10, MTB^{+}17, MMCL19, MDR16, OCE^{+}12, PNI14, PBM^{+}11, POH16, RGEA^{+}15, RHW12, SKD17a, SKK^{+}18, SCB^{+}14, SCL13, SGGL15, SESS^{+}18, SAM14, SCHO11b, TTB^{+}11, TBGW^{+}11, WB12, WBW^{+}14, ZS17, ZJ17]. fishing-induced [HRM^{+}11, SESS^{+}18]. Fistulobalanus [FC19]. Fit [Bea11a]. fitness [CPG^{+}14, CCGP14, FSR10, FC19, NCC^{+}14, PPJT15, PM16, WG11]. Fitting [DDB^{+}14]. Five [ZWP^{+}17, CPID13, GSCH14, Hug11, KBG12, LERS13, RBSL11]. five-fold [RBSL11]. fixation [AFC^{+}16, AFS19, AM19, BvHC^{+}15, BCvH^{+}15, BST^{+}10, CPEE18, CM12, CEW15, DB11a, FBNJ13, RMB^{+}14, RWM^{+}14, RBC^{+}15, TBC^{+}17, WBLS^{+}11, ZCC^{+}12, ZIT^{+}17]. fixed [lLChTC11, MOMF14]. fixing [HvEBC12]. fjord [AHSG16, BWH11, BSH13, BPB11, CRS^{+}11, GMT^{+}17, GWT18, GBH^{+}10, GJSR18, GMD16, HMH^{+}18, HGG^{+}18, HKL19, HMAC^{+}11, JPAM^{+}15, KWM^{+}17, KAJP^{+}15, KBGA19, LOES11, MFH^{+}11, MBS^{+}18, MAHK12, QOGS^{+}16, SHPS^{+}16, SGM^{+}14, SHAS15, SMJP^{+}15, SSAB^{+}15, SDVL16, SB14, TNM^{+}11a, WL11, GCC^{+}10]. fjord-like [MAHK12]. fjordic [WBJ12, WKRW^{+}12]. fjords [CDC18, DMC^{+}17, DDK^{+}18, OCS^{+}13, VHHGF11, ZSEV17]. FL [LRH14, SQRDLC18]. flamingo [GPV17]. flat [BD14, BD19, EKR^{+}10, PBTB15, PSP^{+}19, Sti13b, VDM^{+}10, VBR19, WHvdK11, WD14]. flatback [PK15, TRGF18, WTP^{+}18]. Flatfish [ZF19, ATF12, FKvdV12, KRD^{+}12, LRS^{+}12, MBW10, PHH^{+}12, RvOP12, RLS10, SBL^{+}10, vdVKA^{+}11]. flatfishes [RBH12]. flats [BBH^{+}11, CKR^{+}11, DPO^{+}17, RPL^{+}18, SDM16]. flatworm [Miu12]. flavescens [ETS^{+}11]. flavolineatum [MDA^{+}10]. fleckeri [SWK18]. fledging [ADL^{+}18, AGHC10, CFPF^{+}14]. fledgling [WRR10]. fleets [SGGL15]. flesh [RHW12, WPF^{+}16]. flesh-footed [RHW12, WPF^{+}16]. flesus [CLM^{+}13, Dan11]. Flexibility [vBMH^{+}17, HW14, INvdB^{+}13, MU13, OWW^{+}17, ZRBV15]. Flexible [BB15, HPR^{+}12, QSvN^{+}11, ADA^{+}19, GKRG^{+}19, SKK12]. flight [CPW15, JHNL12]. flights [UAK^{+}14]. floating [BAL^{+}11, KF13b]. flocking [BSY13]. flood [DPC13]. flooding [BSC17, TKV^{+}19]. flora [BS18b, BS18a, YTM^{+}15]. floral [MCK^{+}13]. Florida [HSF^{+}12, JOFP18, SRK13, ALV12, ASWB10, BRBH19, BAvW12, DSS10, FA11, FFS^{+}15, HFMD16, KBK^{+}15, MEKC15, OC17, RJB^{+}14, RPE^{+}13, RCP^{+}13, SQRDLC18, SPW^{+}12, SSdW^{+}12, TYVR^{+}15, VRMKH^{+}15, WKvW10, WSLM13, WLBB^{+}13, WPNK12, YYW14]. flotation [ALANGT10]. flotsam [HRC12]. flounder [CLM^{+}13, Dan11, DMD^{+}12, DTG16, MB18, PSA12, TUK13, WWAS18, ZAC11]. Flow [VBGO^{+}16, ABF^{+}14, BRG^{+}14, BC13, BOS16, Bur12, CPS15, DCW^{+}14, EHG16, GGK^{+}16, HCJ^{+}18, JCM10, JSCM15, KBB10b, LLCM^{+}12, LRS17, MBM^{+}14, MRSD15, MHS17, MBW^{+}19, NC14, NDD^{+}13, OHR^{+}13, OFT^{+}11, OCSW16, PG15, PAFC13, PML^{+}10, REG^{+}11, SF14, SH17, TKB^{+}13, TVRG^{+}13, WADKT13, WS17, WHA^{+}13, WLH^{+}17, WR12]. flow-on [TVRG^{+}13]. flowering [ZH13]. Flows [LR15, CWGG15, DvV^{+}16, ESH^{+}10, HR12, KBR^{+}11, PPTK19, TWW^{+}15, WJHS13]. fluctuating [PLMN10, WDL^{+}18]. Fluctuations [CBC^{+}10, AN11b, BDP10, CG16, FKZTT18, HMBP18, HA14, MOK^{+}18]. fluid [ML11, NMC^{+}16, PZA^{+}13]. flume [ABF^{+}14, DFPH11, SF14]. Fluorescence [MAB12b, DVA^{+}18, VUL^{+}11, vdPLdV^{+}11]. Fluorescent [RPGP^{+}15]. fluviatilis [SSS^{+}10]. flux [BST19b, BMV19, BBRBG10, Bol17, GSD12, HHH11a, LBN11, LMW12, MG12, MGB^{+}14, OMD11]. fluxes [APC14a, BCvH^{+}15, BMW^{+}14, BPG^{+}10, CPSC^{+}18, EBH16, GB10a, JWR^{+}10, LGW^{+}15, LRH14, MDE13, RF14, SDM^{+}13, TNS^{+}12, WAA^{+}19, WLHW10]. flying [IMS^{+}11, LCC^{+}11, PVYS16, VSW^{+}15, WLF^{+}17, YHTC19]. focal [TBG^{+}17]. focus [HSH18, SLF^{+}19]. focusing [WBB^{+}10]. foe [LP12]. fold [RBSL11]. Foliar [MA13a]. Folliculinopsis [KLSJ10]. Follow [SA11, PMM^{+}11]. following [ASF13, FTWA17, FCR^{+}15, FPBL16, GSCH14, GLR^{+}14, GEA^{+}14, HLIK16, HPMB11, IW15, KCMH^{+}17, LPC11b, MPM^{+}14, NTB^{+}17, PBS10, RP13, SV12, SAH11, TBH^{+}18, TPC^{+}13, TPC^{+}15, WD14]. Food [ACDJ11, AZVB13, ALSV16, GMT^{+}17, GCS^{+}19, KHS10b, MDM18, NMPM15, PLG^{+}19, RTD14, TP16, Urb12, VM14, vdVFK^{+}10, AJF^{+}12, ACA^{+}10b, ACA^{+}10a, AAIB15, BEP^{+}12, BFRM13, BLV^{+}19, BBI16, BDP10, BDDvdM17, BRS10, BGIA19, BMF^{+}12, BKAS14, BHT10, BDB^{+}12, BSY13, CF12, CL10a, CLG^{+}19b, CWGG15, CEK^{+}16, CHG^{+}18, DSM11a, DVSM10, DFS13, DBC^{+}12a, DMH15, DSG^{+}13, DMC^{+}16, DTM^{+}12, DIB15, DVB^{+}16, DMM^{+}10, DRWQ12, DKB^{+}12, DJL^{+}12, DCW^{+}10, EvDB^{+}14, ESH^{+}10, FJ11, FRMS17, FOJF15, FL17, FG16b, FHS^{+}13, FPN17, GSJS13, GGK^{+}16, GGPDS19, GGB^{+}18, HHS10, HHS^{+}12a, HJMN^{+}12, HBE^{+}13, HWR^{+}14, HS15b, HT11a, HEAI10, JRvdM15, JSMO14, JDG^{+}17, KÖHSO12, KF13a, KSS13, KS12a, KCR^{+}10, KFBM19, KDB^{+}10, KLSJ10, LLv13, LR15, LMV^{+}19, LWS^{+}10, LLC^{+}14, LCL^{+}17, LDT^{+}13, LNK^{+}16, MEDS11, MFH^{+}11, MM13b, MBS^{+}18, MWWS17]. food [MC12b, MBRO10, MHS14, MFS14, MBV^{+}11, MRB^{+}15, MSE^{+}12, MSS^{+}14b, NDG15, NT10, NKH12, NRK^{+}19, NBM^{+}14, NCT^{+}14, NUJ12, OFT^{+}11, ODK^{+}14, OKLM14, PHI^{+}12, PFB^{+}13, PDGL17, PMEZ15a, PMEZ15b, PFdP^{+}12, PAF^{+}10, PBDMRB13, PMOAC^{+}17, PFL13, PFKB18, PWO^{+}16, PCW^{+}12, PGVF14, RLN^{+}13, RBC^{+}14, SBV^{+}15, SNW^{+}18, SMS15, SESS^{+}18, SS13, SBM^{+}18, TLB15, TLS^{+}19, TMS^{+}16, TFP14, TTS17, VRSS18, VC14, VCFF13, WKK18, WL11, WPP11, YTM^{+}12, YKBB12, YLPG19, dJSB19]. food-depleted [FG16b]. food-driven [PFL13]. Food-limited [ALSV16, PAF^{+}10]. food-supplemented [FRMS17]. food-web [MBS^{+}18, NCT^{+}14, PGVF14]. Food-web-based [TP16]. footed [MBC^{+}17, RHW12, WPF^{+}16]. footprint [LKC^{+}15]. footprints [RUD^{+}16]. Forage [LTP^{+}18, LANER15, BNAN12, BBR^{+}14, BMA^{+}19, BANE^{+}13, CEC^{+}18, DGCH18, DBG15, DRD^{+}18, FPBL16, FSS^{+}18, GKR^{+}15, KEBB10, LBE^{+}10, MBBK16, MGB^{+}19, OWH^{+}19, PBAP^{+}14, PIPD10, REG^{+}11, RTC^{+}14, SHS^{+}17, vBAP^{+}19]. forager [OWW^{+}17, TSA^{+}18, WEAG^{+}15]. foragers [RPR16]. Foraging [ACMD11, BDNA10, BBBNT13, BMV^{+}11, BGDC14, CLC^{+}12b, CMO^{+}10, DF15, DZB^{+}18, ECM^{+}12, EC14, EKOÅ13, FCRD18, GCSF18, GAC11, GCRP15, GLM^{+}19, HGC15, HSCA17, JGWJ^{+}11, JTWJ^{+}13, KHW^{+}10, Kiø16, KBB16, KMSW10, MM13a, MPG^{+}10, OHCG14, PXG^{+}10, PGR^{+}10, POR^{+}15, PLA^{+}12, PSS19, PWB^{+}19, SWFQ12, SEF^{+}15, SMR^{+}11b, TDB^{+}14, TMS14, UW18, WBS^{+}12, YTO^{+}11, ADF^{+}18, AED^{+}12, ABA16, Ano17c, ADA^{+}19, BT10, BCG10, BST^{+}19a, BSG^{+}10, BAS^{+}10, BWLP17, BTO^{+}18, BSN^{+}18, BP12a, BSC^{+}15, Ben16, BBKH^{+}11, BBG12, BBSM19, BBC^{+}12a, BSBR^{+}12, BRT^{+}17, BBC^{+}17, BKC^{+}19, BBL11, BHTF11, BJS^{+}16, CCB^{+}17, CCA^{+}17, CVW^{+}11, CCE^{+}14, CBLM^{+}17, CSB14, CY10, CCJR14, CBR19, CPM^{+}17, CPG^{+}14, CWGG15, CMK^{+}14, CD14, CTS^{+}15, CEC^{+}18, DWPH11, DWLH13, DGMB10, DP14a, DKF^{+}15, DMT^{+}14, DJJ^{+}16, DL14, DFTZ^{+}15, DBHMG12a, DBHMG12b, DAM^{+}18]. foraging [FNF^{+}18, FTT^{+}17, FSF^{+}17, FHRC16, FRMS17, FHMG19, FMH^{+}19, GDB^{+}16, GLJ^{+}18, GCC^{+}16, GP11b, GPV17, GPW^{+}10, GAB^{+}14, Goy14, GWK^{+}19, GVGP^{+}14, HAM^{+}15, HBF^{+}10, HNK^{+}18, HHS10, HVJ^{+}18, HHS^{+}12b, HG10, HW14, HSMB17, HA15, HIH^{+}17, HCW^{+}15, HWR^{+}14, HSS17, IS16, JWR^{+}12, JLM^{+}12, JS12b, JHL^{+}19, KKMA^{+}10, KHW^{+}12, KWH^{+}18, KTP^{+}18, LSH19a, LCBK^{+}13, LMCD19, LNF^{+}11, LHHG11, MLM^{+}12, MCW^{+}16, MJS17, MVP^{+}16, MYK^{+}16, MF15, MWC18, MWZ^{+}16, MAB^{+}12a, MBC^{+}17, MCP^{+}18, MHR^{+}19b, MJHS11, NWP^{+}11, NGJHB12, OALK14, OWW^{+}17, PMK^{+}15, PHI^{+}12, PPM^{+}15, PVZD^{+}11, PRLR^{+}18, PDT^{+}11, PS19, PLGSD17, PHHL19, PFL13, PWW12, PTP^{+}15, PLF^{+}12a, PWM^{+}19, PWA^{+}17, PRM17, QSvN^{+}11, RLN17, RDHM10, RPR16, SKSH19, SRK^{+}18, SMBH15, SBG^{+}16, SBT12, SSR^{+}10, SJL^{+}16, SGC17, SRB^{+}17, SWK^{+}16, SBA12, SRK13, SHT^{+}13, SQRDLC18]. foraging [SBB^{+}12b, SCW^{+}18, TLJ^{+}11, THB^{+}12, TGH15, TFB^{+}17, TTB^{+}11, TKDM16, TP19, TFW^{+}12, VF13, VMC^{+}19, WBG^{+}14, WEAG^{+}15, WHK^{+}11, WPF^{+}16, WSC13, WTdB16, WHB^{+}19, WHFM17, ZMRCP^{+}17, dSdSdCO^{+}11, dDTA^{+}18, vERRP16, vMVSW12, DGC12]. foraging-site [OALK14]. foraminifer [HSRP11, MWBMH13, RUP11b, RUP11a]. Foraminifera [CGR^{+}15, JM17, MBG^{+}16, PUP16, XLXS17]. Foraminiferal [CW12, ENO^{+}11, GA11, PDGR^{+}13]. Foraminiferan [POKGRP10, SHDR13]. Foray [PFL13]. forbesi [KK17]. forces [CFF^{+}18, KPB11, ME11, SLSC16]. forcing [AGM^{+}11, BGP^{+}16, HCJ^{+}11, PMML12, PRA10, RPM12, SM10a, WK18]. forecast [ZBC10]. Forecasting [LBR^{+}19, LD10, SABMG16, BZC^{+}12, MBCB^{+}19]. foreshore [BDLW10]. forest [BKR^{+}15, DA10, FNC13, dSKFV12, KS13, MHNL14, PMGQM17, RS16, SKD^{+}11a, SJB16b, TPVCV19, WFD19]. forests [CMF^{+}19, DSJNH18, ESW14, GBK^{+}14, LRL^{+}13, MA13b, NCR^{+}14, SBM^{+}16, SBE^{+}16, VDV^{+}15, YMPR13]. form [AEM^{+}18, MDS14, VPBV13]. format [Ano17a]. formation [BIL^{+}10, BPHW16, GGF^{+}12b, LGS10, MDF10, Miu12, MMEM14]. Forming [FJL12, AS14, BB17, BMMS10, BAM^{+}10, BTLF17, DDP^{+}13a, EVB^{+}16, GGMTGC18, GTF10, HCG^{+}10, KNY^{+}15, MGFW13, MHF^{+}19, OS18, PBC^{+}12b, RSG^{+}16, SLT^{+}19, TW16, TSCSB15, TRMC11, VHKT15]. Formosa [RCGE12, MWP10]. forms [DB11a]. fornicata [DP13, DBCP15, PDC^{+}17, VBD^{+}11, dMA18]. forsteri [FHMG19]. Forsterygion [NSS10]. Forty [GKR^{+}15]. fouling [FWW^{+}16, Gre10, HCM^{+}11, KO12, KKMM16, KCD13, LBC^{+}19, NC11, OMW^{+}10, PW19a, SK13]. Foundation [MD16a, ATP14, BRB17, FSH^{+}16, HGKB14, HL14, MHCRR18, OS18, PRA^{+}19, SRB15, VB19b]. foundational [MORM15, RLO18]. Four [JS12b, MS18a, PRM17, AHS^{+}13, ASCHM19, BBSM19, BDG^{+}16, BL10, CFPC^{+}13, CO16, CC12, GS16, KP11, KCE14, LCGS^{+}19, MRE14, OFPR^{+}11, RBSL11, Rei12, SWK19, SWB^{+}10, SOA^{+}17, SG18b, TGLC10, TOR^{+}10, WFB16, YWiIK15]. four-eyed [KCE14]. Four-year [MS18a]. Foveaux [MS18a]. Fractal [TSGW10]. fractionated [BTS^{+}14, SGMG^{+}14]. fractionation [DPH14, EAMT10, GKHM13, HN13, PCW^{+}12, RMHP14, TGI14, TMP^{+}13]. fractioned [WKX^{+}19]. fractions [JKM11]. Fracture [RWMP12, dlSBO^{+}12]. fragile [DMJ11, KS12b, LSV10b, DMJ11, KS12b, LSV10b]. fragment [PWM18]. Fragmentation [RMB18, DA10, MRH11, NHAS15, SLS11]. fragmented [PA13, TCN^{+}10]. fragum [HRC12]. Fram [LBN11]. framework [CSPA10, EW16, ON15, RBC^{+}15, SLSR10]. frameworks [HGW^{+}11]. Francisco [GK10, GSTK11, KK17, MS13, RCB12, SIK16, ZOM^{+}10]. Fratercula [BNAN12, FRMS17]. Free [MPS^{+}10, ASAG17, AGNH^{+}19, CWR^{+}18, GGD^{+}11, KAAI15, KdWV^{+}18, LLRC^{+}12, MRLD19, SBG^{+}16, VBT^{+}17]. free-living [KdWV^{+}18, LLRC^{+}12]. free-ranging [ASAG17, AGNH^{+}19, GGD^{+}11, MRLD19, SBG^{+}16]. free-swimming [CWR^{+}18]. freely [VGR^{+}16]. Freeze [GRCP^{+}10]. Freeze-lysable [GRCP^{+}10]. French [ELA10, TPA^{+}13, CdTK^{+}16, MDA^{+}10, TBH10, TRA^{+}14, VVB19]. Frenulata [LKL^{+}19]. frequencies [McD13]. Frequency [BCV^{+}18, AP14, DMMW12, FA10b, LST11, PPML17, SKR^{+}13]. frequent [CD13, KKA12]. freshly [WMS^{+}11]. freshly-isolated [WMS^{+}11]. Freshwater [BSB14, ABT^{+}15, ASM^{+}12, CHFS12, DCW^{+}14, EPF^{+}17, JCM10, JSCM15, JGM^{+}16, KRD^{+}12, LGD10, OHR^{+}13, RPRN^{+}16, REG^{+}11, RLH^{+}13, SLS11]. freshwater-induced [RPRN^{+}16, SLS11]. freshwater-mixohaline [LGD10]. Friend [LP12]. Frigate [TBH10]. frigatebird [ADA^{+}19, OWW^{+}17]. fringing [CAB^{+}14, HHH^{+}11b, WLHW10]. Frio [CPS^{+}14, SML^{+}14]. front [GMSC14, LA14, LGT^{+}14, NYK^{+}13, TGBM^{+}16]. Frontal [FLE^{+}11, KPH^{+}15, SBT12, BO15, HRÁB^{+}14, RBC^{+}14, TWW^{+}15]. frontier [HBK^{+}17]. fronts [AIA18, FKA^{+}12, KVR^{+}18, SBBB^{+}18, SMVC^{+}15]. frustules [FHS^{+}13]. fry [MWC^{+}12, SWP^{+}13]. fry-to-smolt [MWC^{+}12]. FTIR [PMF^{+}17]. Fuchs [RL18]. fucoid [ATS14, AS14, BMMS10, FHJ15]. fucoids [AS15]. Fucus [AOW12, BSP^{+}10, DRDV15, MCRB11, TPC^{+}16, WSD^{+}10, WRO^{+}11]. full [BDS^{+}18b, KWC^{+}17, YYU^{+}17]. full-depth [BDS^{+}18b, YYU^{+}17]. full-spherical [KWC^{+}17]. fully [FHRC16]. fumatus [RJR18]. function [AGM^{+}11, BAA12, BMSS14, EHG16, GGNL^{+}17, HTT^{+}12, HTK^{+}17, LR18, RFRE19, SKB15, SSD13, TMSG19, TPS10, ZPPM17, vDBH^{+}15, vdVKA^{+}11]. Functional [ADO^{+}12, Bea11b, BCB^{+}17, CLGC16, CHG^{+}18, DPM15, FLAL10, LS19, MHK11, ØLH^{+}13, TVM16, VVB19, VHBG15, WAC^{+}14, ADF^{+}18, AGV^{+}11, AAGGG19, BH16a, BPG^{+}10, BMPC15, CFMFFJ18, CHRB10, DDP15, EOM^{+}14, GMG^{+}16, GWT18, GS17, GH11, GKPT16, JBD10, JD18, LMB^{+}19, LM18, LCB^{+}12, McC14, MTM^{+}11, MGOL19, MRB^{+}15, MDS14, MWMD17, NPLT10, OSC19, OBÁSV18, SPM^{+}13, TBS^{+}14, VPOA12, WDM^{+}19, WKRW^{+}12, XS19]. functionally [OHB^{+}15]. functioning [CF12, CFC^{+}12, CSRF16, CB12, DLdM19, DFDC18, DVB^{+}16, EM13, FFL10, GC13a, GCP^{+}17, GKPT16, GNC^{+}10, LRL^{+}13, LRM^{+}14, dJSB19]. functions [AL12b, KNIN17, LS12, MCT16, PAG19, POPH12, Pro17]. fundamental [Ano17c, LGM^{+}12, ST11b]. Fundulus [JMO^{+}19, LFD13]. Fundy [MSH^{+}14, TR13, LS16]. fundyense [RZH^{+}16b, STT14]. funebralis [CS11a, GB16]. Fungal [VBŽ^{+}17, CHDH^{+}19, SWWK13]. fungal-farming [CHDH^{+}19]. furrow [SCO^{+}12a]. Further [LSY^{+}18]. fuscescens [BBR^{+}14]. fuscus [TRSB^{+}18]. Future [FRC13a, BM13, CDMW12, CLGCA^{+}19, Chr17, CLMB13a, CLMB13b, CCGP14, FL10, FKT15, FABH^{+}16, FTH12, GAB11, KHW^{+}10, KCB18, KDS^{+}12, KPr^{+}18, LS14, LDS^{+}15, MOK^{+}11, RU11, SGS^{+}17, STK12, VCL^{+}15]. Fyllas [DHG16].

Gabès [HS15a, FCR12]. gadfly [KPr^{+}18]. gadid [CLSS17, MMDM12, MRE14, SGL^{+}17]. gadoid [ETHB17]. Gadus [ALNB17, ACG12, BSN^{+}18, BPWF14, DSG15, FWPP12, HDB^{+}10, JBS12, LNP12, MFG^{+}13, MMNR^{+}12, MRC10b, NJJ^{+}11, OMSP10, PLA^{+}12, PHHL^{+}11, RAN^{+}10, RJTS12, SGL^{+}17, SGM^{+}14, SMK^{+}10, SBH^{+}18, SJB^{+}15, SNT^{+}16, WBW^{+}14, WDM^{+}18, WFP15]. gait [Nor13]. Galapagos [AMSdL^{+}18, JWR^{+}12, BBBW17, EPC^{+}19]. Galatheidae [BHP11]. Galaxias [HS16, SS12a]. galbana [BBRBG10]. Galeocerdo [NWP^{+}11]. Galicia [BVP11, THMF^{+}15]. Galician [GSÁCM^{+}16]. Galindo [She13]. Gall [BTLF17]. Gall-forming [BTLF17]. galloprovincialis [BC11, CGS^{+}19, EKE16, FRRÁS^{+}12, FRGI15, FAZ^{+}14, LBP^{+}19, PLFR^{+}11, ZFC^{+}13, vdMPM^{+}10]. Galveston [RLH^{+}13, SMR10]. Gambierdiscus [LRBS15]. Gambusia [PL10]. gamefishes [LL18]. gamete [FDP^{+}16, MC10b]. Gametogenesis [TJS17, AZSKWL11, NWS^{+}14]. gametophytes [MWW^{+}14]. gammarus [MOS10, MOK^{+}11, SRM^{+}15]. gannet [BCV^{+}18, CVW^{+}11, DHS^{+}19, DAC^{+}12, LSH19a]. gannets [ABA16, BRT^{+}17, CPG^{+}14, GCRP15, SBB^{+}12b]. gap [LGS10]. gaps [ABRP15, BPSK11, CSS^{+}17, HMB^{+}12, Joh10]. gardening [VDG^{+}15]. gas [EFR12, Fuj16, MRC^{+}19, RCMB18]. Gasterosteus [SML17]. gastropod [BVQ12, CBD14, Fis10, HCG15, LC13, MP15, MCR^{+}13, PCF^{+}19, SCP^{+}14, SPM^{+}16, TMBP13, VBD^{+}11, WG11, YLPG19]. Gastropoda [BBdG^{+}19, DP13, DBCP15, MEDS11, MC12b, SSBC11, SHMB15]. gastropods [AJF^{+}12, BGJ^{+}19, Col10b, CO16, FTKK14, LCL^{+}11, LMd12c, LHB19, MCUO12, PDC^{+}17, PP18]. gaudichaudii [KPH^{+}15]. Gauging [LPKT^{+}19]. gauntlet [ASM10]. gear [CEH18, JHNL12, MMCL19, ZJ17]. gelatinous [BGM^{+}18, GCDH19, HKB^{+}19b, LGT^{+}14, PJMA18]. Gender [FM14, RGCT^{+}10, CCC^{+}18, HT11a, KLN^{+}13]. Gender-specific [FM14, RGCT^{+}10]. gene [BMD^{+}11, GAF^{+}16, GB16, GCL^{+}16, HPMZ11, MHS17, OTS^{+}10, PAFC13, PZL^{+}17, PHHL^{+}11, RB11, SGCF^{+}12, SH17, WNF^{+}11]. gene-associated [PHHL^{+}11]. gene-flow [SH17]. genera [YMCV10]. General [ST10b, TW10, BC16, EW16, JOFP18]. generalise [She12]. generalist [BPJ^{+}15, CCB^{+}17, Cha13, HMR^{+}16, LA12a, OMN17, PTP^{+}15, TGR^{+}19]. generalists [THB^{+}12]. generalized [AN11b, YIKG10]. generated [DBS^{+}16, KBR^{+}11, RCBE15, TNOY17]. generation [MSGGE10]. generic [JR15, TLB15]. genes [JGM^{+}16]. Genetic [ARMA15, BHP11, BCM^{+}11a, BJCB^{+}14, BHASP18, BPG^{+}13, BSH^{+}14, CLM^{+}13, CRJ13, CML^{+}12, CBC^{+}16, DCT^{+}10, FHM^{+}11, GM13a, GRN^{+}16, GPM^{+}13, GE17, HVT12, ICT^{+}11, MZ10, NvOH13, SØK^{+}15, VCG^{+}10, YWiIK15, ADD18, APC^{+}13, ALK17, ARvdW^{+}16, BSC^{+}13, BBF^{+}19, BSK^{+}11, BHN^{+}14, BEFC15, BCC^{+}11, BCB11, CT10, CBF^{+}19, CFMN16, CNSW16, DEGF^{+}10, DSH14, DBS^{+}13, DBE^{+}18, EHD^{+}17, FRC^{+}13b, GDSPRA13, GSM^{+}11a, HHLT10, HPL12, HTH10, JSBP15, JBL^{+}16, JRK13, KHGS12, KBG^{+}13, KTH^{+}18, KJBS12, KBB^{+}10a, LMRRB17, LGBSC11, LBC^{+}19, MHPvdH19, MBB^{+}12a, MHS17, MBvO^{+}18, MMFP^{+}15, MIM^{+}12, NWS^{+}14, NM18, PHSBO16, PFB^{+}17, PPTB12, PMM^{+}15, RBSL11, RMKR13, RWM^{+}12, SSBC11, ST11a, SHH10, SKA^{+}14, SWB^{+}10, SLW15, SH17, TCN^{+}10, TPC^{+}16, TPBM13, TKS^{+}14, TSR17, VM10, WWG13a, vdHTT^{+}15]. genetically [OCS^{+}13, SWK19, SHM^{+}10a]. Genetics [MAD18, BHUCMO19, BSRDR10, BPB^{+}12, BFH^{+}13a, HBP^{+}16, LKN^{+}13, MdSW^{+}16, SDC^{+}16, SMVM^{+}12, SN17, SH17, TF12, WWG13a, WFH11]. geniculate [WNP^{+}14]. genome [GCBB13]. genomes [EH13]. Genotype [SRB15, FWPP12, LSBB17, PWM18, PHD^{+}14]. Genotype-specific [SRB15, LSBB17]. genotypes [DAWS19, FSR10]. Genotypic [GKHS16, HBS10, HL14]. genotyping [CJS^{+}12, MHS17]. Gentile [BBM13]. gentoo [CCB^{+}17]. Genus [NNJ^{+}13, AMC^{+}11, BAFP^{+}13, CWD11, CKC^{+}10, LP12, SWB^{+}10]. Genus-specific [NNJ^{+}13]. geo [TPT^{+}12]. geo-location [TPT^{+}12]. Geochemical [ODL^{+}13, BDLW10, CLDC^{+}13, KPP^{+}16, SKW^{+}14]. geochemistry [BDT^{+}11, LTN^{+}11, TRSV^{+}12]. Geographic [BTO^{+}18, CC12, FCH^{+}10, KCH^{+}11, LFM^{+}16, MGO^{+}12, MM12, McD13, RAA^{+}16, ADD18, BCB11, FSS^{+}18, HCL^{+}11a, KO12, KMP14, LAM10, LDS^{+}15, MLR17, RGDNGS11, RCVG10, SDR^{+}15, SMMFH16]. Geographical [MBC^{+}17, BC11, CDCS12, CSQRA11, DMHS^{+}14, HCL^{+}11b, LLBP15, LNF^{+}11]. Geographically [RCAB^{+}12, LAB10]. Geography [HDP^{+}17, SST19]. geolocation [DH18a, GWH^{+}18, GSSM^{+}11, LNS10]. geolocators [EMTB^{+}12]. geology [LCL^{+}13]. geometry [Bre14, PAG19]. geomorphic [TMSG19]. geomorphology [KH10]. George [WVD^{+}17]. Georges [BCSB13, DGDJ14, LFG^{+}12, LA14, PSA12, STT14, SCL13, TPM19]. Georgia [RWM^{+}10, SBT12, WHPT12]. Geospatial [PSA12, SKW^{+}14, SLSR10]. geostatistical [DEWK15, LJM^{+}12, WFH^{+}18]. geothermal [PDGL17]. Gerlach [SHM^{+}12]. German [DEJN10, MASMH19]. Germany [BDD^{+}18]. germination [AED^{+}10, MDS^{+}13]. gestating [SDR^{+}15]. get [GSL17]. Getting [BHE^{+}17]. Geukensia [ZZG^{+}19]. ghost [BD11b, MBT14, PPS^{+}13, POX^{+}17, VPW^{+}12]. giant [BBMM14, CSKR10, CSRD11, CBR19, DFD^{+}19, DBE^{+}12, FK10, Fox16, GBMW12, GVR^{+}14, HAF^{+}17, KMP15, MJW19, MZ10, MCW^{+}16, MLASTS13, MU13, MA13b, PMS10a, TDB^{+}14, YMPR13]. giant-kelp [CSRD11]. gibbosa [MWBMH13]. Gibbs [RWMP12]. Gibraltar [GVG^{+}18, MSGGE10]. gigas [BBG12, BBdG^{+}19, BPHW16, DH18b, DWL19, FAM^{+}17, GRCH^{+}18, ICT^{+}11, KCS^{+}17, LdU^{+}17, MBS^{+}13, NCS^{+}10, RRRF^{+}17, RCVG10, SRCE^{+}10, SZG11, SHF^{+}12b, SD17b, TKA10, VTSH^{+}19, WDH^{+}12]. gilchristi [SRG14]. gill [MDAPH11, WFS10]. gillnet [CASG^{+}19, Løk11]. gillnets [DL13, NWH^{+}12, NWD13, OMW^{+}16, RCU^{+}11]. gilthead [GCL^{+}16, MMB^{+}12a]. Gironde [CPR^{+}16]. give [BCM^{+}14]. gives [OVF11]. Glacial [ZSEV17, BWP^{+}11, HLIK16, HKL19, HI14, HMAC^{+}11, QOGS^{+}16]. glaciale [BRF11]. glacialis [DSM11a, GMT^{+}15, GSA^{+}18, KDS^{+}12, KHM^{+}12, KHS10b, LRR^{+}10, MRB^{+}11, MBP^{+}12, PPJT15, PSS^{+}12, RSP^{+}16, SKM^{+}15, SKD^{+}11b]. glaciations [SLW15]. glacier [FKA^{+}12, JPAM^{+}15, BWP^{+}11]. gladius [RCVG10]. glandula [BSK^{+}11, NC14, ORG19]. Glass [CMYL11, HKB^{+}19a, AHR^{+}18, CL10b, GMK^{+}19, MSV^{+}11]. glauca [NCW^{+}19, QHN^{+}10]. glaucum [TCN^{+}10]. glider [WLM12]. gliders [SMSB19]. Global [FFL10, FTH12, HFG^{+}19, JTH^{+}12, MSF^{+}19, MER14, PML^{+}19, PIGK16, ADD18, BMÁN^{+}18, Bea12, DS16, DFDC18, DP14b, GSC17, HTC17, LBJ^{+}16, LEKT12, MHHB^{+}17, MS11, PCLB18, SBH^{+}19, She12, STP10, TS15]. globally [CKC^{+}10, KKMA^{+}10, PBCA10]. Globicephala [MCL^{+}17, NSC^{+}19, WVM^{+}14]. Globobulimina [POKGRP10]. glomerata [GB10b, TSR17]. Glover [TBG^{+}17]. glow [PK15]. glucose [NOT^{+}11]. glutamate [FUOG^{+}16]. Glutamine [HKT10]. glycine [AGNH^{+}19]. glycoprotein [ZFZ16]. Glyphis [LDP^{+}17]. go [DDM^{+}16, GRL^{+}15, MCPC19]. gobies [PB10]. Gobiesocidae [FG10]. Gobioidei [LLD^{+}12]. goby [CBT13, FA10a, HIKK12, HI14, Lyo14, SUPCB11]. Godthåbsfjord [ANRT10, CRS^{+}11, TAAN15]. going [GBD^{+}12]. gold [Par15]. Golfo [GZYS11]. gonad [EHR12]. gonads [RLG14, ZdVA^{+}16]. Gonatus [GCS^{+}18]. Goneplax [NdBKR13]. good [HB15b]. goods [TTC11]. gopher [LBS^{+}15]. gorbuscha [KSA^{+}13]. Gordon [Ras10]. gorgonian [BAFP^{+}13, CFPC^{+}13, CCM^{+}12, EPRGO13, KLT^{+}12, LCC12, MLASTS13, SGT12, WMCC15]. gorgonian-associated [SGT12]. gorgonians [LD10]. governing [BL17]. GPS [DBHMG12a, EKOÅ13, HCDR^{+}15, LBJI14, SRK13]. GPS-loggers [HCDR^{+}15]. Gracilaria [HWR^{+}13, JL12]. gracilis [HJ16, HTHH11]. gradient [BG19, BWV^{+}12, BHT15, BPSG11, BAE^{+}10, CBJB19, DAL^{+}10, DR16, Det10, Ebe10, FL17, FBET15, JTTP19, JMB^{+}17, LAK^{+}19, LAB10, MC10a, MWWS17, MCP^{+}18, MZF^{+}19, MU13, MGOL19, NBGGST^{+}15, OPN10, PGR^{+}10, PS15a, PDC^{+}17, PÁIIS^{+}19, RRL^{+}19, SAWK18, SHK14, SBT12, SGTD14, SHAS15, SWPK13, VARO12, VDM^{+}10, VC14, WHB^{+}14, WMB^{+}14, YHR16, ZMGS17, ZRR^{+}13]. gradients [BSD^{+}14, BOKJ18, BWCZ11, CGD10, CWEE19, CKJOM14, CS11b, FGOS16, FOJF15, HWSK19, JGJ15, LLC^{+}14, LLF^{+}18, MWGH15, PMYH13, RS13, SBM^{+}18, SH14, SHB^{+}10b, TL18, TRLS^{+}11, TNM^{+}11b, VGBJ^{+}13, VCDP11, VJJ^{+}19, VCFF13]. grafting [JRK13]. grandis [JMO^{+}19]. grani [OSC19]. granosa [NLKZ12]. granulatus [MHPvdH19]. Grapsidae [CY10]. grass [BMR17, CSN12, DD15]. grasses [HKT10]. gratilla [BJM19, JMB10]. gravis [PWA^{+}17]. gray [BSB^{+}18, BLMD^{+}11, BM16, GSJS^{+}19]. Grazer [PW19a, AN11a, ANB13, BHF^{+}10, BHDC15, CB14, DGA^{+}15, FCS15, HBS10, LGS10, SLRM^{+}15, SSM^{+}17, WHT12]. grazer-plant [DGA^{+}15]. grazers [CL10a, CLL13, CFH11, EOM^{+}14, EBT^{+}15, EN11, LGMB^{+}16, LM18, PMS10a, SDMB15]. Grazing [BAE^{+}10, DQM^{+}19, GSTK11, SKD^{+}11b, AS14, AWJ12, ADRA19, BLER15, BKB11, BRIC13, BCM19, CRS^{+}11, CRGT19, CSC13a, CHDH^{+}19, CAC11, CP11, DP11, DH16, DLL^{+}18b, FSLG12, GTF10, GLY^{+}14, Hv14, JSLG17, LGMB^{+}16, LSY11, LHS^{+}14, LMD12a, LPSK17, LS19, LFL11, LRBS15, LSV10b, ML11, Mat10a, MDF10, OSK15, OPSM13, OSL18, OAD^{+}14, PMML12, PBS10, RSS^{+}16, SVIH^{+}10, SPP17, SV12, SH10, TLFS14, ZLCW15, vDCR13]. Great [BKB11, ETB13, NGA18, RUP11b, DSC13, PWA^{+}17, VSG^{+}15, BP12a, BB11, BSB^{+}11b, BDF^{+}11, BK14, CED^{+}19, DZDB10, ETAE13, GGCB16, HSMB17, HPLS13, KHH18, LKDP17, LPC11b, MBB^{+}15, RUP11a, SGG^{+}12, SWB^{+}10, Sti13a, TPB^{+}16, TPB^{+}18, WSC13]. greater [Ben16, GPV17, GLHM^{+}18, GKR^{+}15]. Greece [KAP^{+}12]. Green [BBC19, GDB^{+}16, KAMA13, PRK^{+}12, ATT^{+}16, AGH^{+}12, BBC^{+}13, BBC^{+}17, BHTF11, CC19, CdTK^{+}16, CLS^{+}16, CD16b, DT15, FG16b, GBL^{+}15, GCSF18, GDJdBSL19, HNK^{+}18, HVJ^{+}18, Hv14, HSMB17, HRS^{+}16, JH14, JBL^{+}16, KS12b, LH16, LABM17, LKL^{+}17, LD17a, MLM^{+}12, MMG^{+}16, MEK11, OMW^{+}16, PVZD^{+}11, PDvG16, PFB^{+}17, PHTS11, SMBH15, SZ13, SWH^{+}18, TGH15, TBH10, TSA^{+}18, VAB^{+}13, WBB^{+}10, WHB^{+}19, WNA^{+}15, WHFM17, vdMHW^{+}10]. Greenland [ANRT10, BRS10, BS11, BSH13, BNS^{+}14, BDS^{+}18b, CRS^{+}11, DHG16, DDP19, DSK^{+}10, FLK12, HJMN^{+}12, JPAM^{+}15, KHW^{+}10, KBW^{+}11, KDS^{+}12, KAJP^{+}15, KHS10b, LHJH^{+}10, MF15, NTS^{+}18, OKJMC15, PHH^{+}12, RSDN15, RZH^{+}16b, SGM^{+}14, SKR^{+}10, SKD^{+}11b, TAAN15, TRL^{+}15, TJS17, VCMS^{+}18, WS12, ZTNS13]. Greenlandic [SMJP^{+}15, SNT^{+}16, TNM^{+}11a]. greenshell [WWG13a]. gregarious [RBGT10]. grenadier [DDK^{+}18]. Grevelingen [EKR^{+}10]. grey [BSB^{+}11a, BPT15, DHJ^{+}17, HHS^{+}12b, KBP^{+}10, OALK14, RABS19, vBBW^{+}14]. gridded [BLT14]. griffithii [GLM^{+}10]. griseus [HMO^{+}12, TTT^{+}15]. ground [ATF12, BQ16, BHTF11, CCE^{+}14, CRL13, GDJdBSL19, KTNO19, KRD^{+}12, MLM^{+}12, PAC^{+}15, SKD^{+}11a, WHFM17, YYO14]. ground-truth [SKD^{+}11a]. groundfish [BWH^{+}19, CH12, GLR^{+}14, HA15]. groundfishes [OGCHS18, TSTP13a, Tho15]. grounds [ASF13, BDG^{+}16, GLJ^{+}18, HKB^{+}19a, KPKA13, LJM^{+}12, LNF^{+}11, MCW^{+}16, MMRC16, PVZD^{+}11, RB18, WHI^{+}18, WHD^{+}14]. groundwater [BMW^{+}14, GBTS11, LTN19, SD17b, SKN^{+}17, TCKY19, WK10]. groundwater-borne [TCKY19]. groundwater-fed [SD17b]. groundwater-influenced [BMW^{+}14]. group [Bea11a, Bea11b, CWMP13, FWPP12, SHK14, SKS18, vdVFK^{+}10]. Grouper [FA11, ABB^{+}14, EF16, KA16, MD11, NMS^{+}18, SDST16, WDNM11]. groupers [RNSA15]. Grouping [BCGG19]. groups [CWW^{+}16b, GKPT16, LSV^{+}10a, LPC19, LCB^{+}12, McC14, MDS14, PIPD10, SESS^{+}18]. growing [PBF10, WNG^{+}19]. grows [MGP12]. Growth [ADvdM13, AHPKS17, BM10a, BCM19, CY13, DDP^{+}13a, EOA15, GVR^{+}14, GCM10, GRN14, HT11a, HT11b, HAHM16, MCC^{+}16, MPD11, OA10, PRBR11, RBH12, SNVH19, THP^{+}16, ASM^{+}12, AHS^{+}13, ACA^{+}10b, ACA^{+}10a, AHH^{+}11, ATP^{+}12, ATO^{+}14, ACS12, AGNH^{+}19, AGH^{+}12, AGP^{+}13, BA13a, BG19, BVL10, BGJ^{+}19, BFDM18, BJF17, BVG13, BP12a, BSC^{+}15, BDAH10, BS19a, BDP10, BD15, BDDvdM17, BAT14, BBC19, BRS10, BCSR10, BBdG^{+}19, BDVP19, BDC10, BLM10, CRS^{+}11, CR14, CC19, CSABH^{+}13, CLB19, CBMK10, CSC13a, CESAM12, CHP^{+}11, CLHW14, CGC15, CDMB^{+}18, CVC^{+}18, CTB13, CBF^{+}13, CPS15, CL10c, CRS^{+}18, CEK^{+}16, CHLR16, DEP19, DTG16, DWLH13, DTVH10, DBCB11, DVS13, DLL^{+}18b, DCC^{+}10, DRH15, DWV^{+}12, DHR^{+}13, DH17, FDVL18, FCR12, FLS^{+}18, FTB14, FJ11, FJ12, FRMS17, Fox16, FEE^{+}11, FTB^{+}17]. growth [GM13a, GCC^{+}16, GGMGR12, GRFSM12, GG12, GSL17, GK10, GLHM^{+}18, GGMTGC18, GTF10, GHT15b, HBHS^{+}18, HSMB17, HSR^{+}10, HPH^{+}18, HMF^{+}18, IHP10, JW10, JTWJ^{+}13, JH11, JNYT13, KJW^{+}15, KKS^{+}12b, KCL^{+}11, KdVvdP^{+}12, LTN19, LSBB17, LKC^{+}17, LSY11, LLv13, LHS^{+}14, LAM13, LN11, LABM17, LP10, LD17a, LKDP17, LTG17, LFE^{+}14, LH19, LRQ^{+}12, LCB^{+}16, MEKC15, MBSG11, MBW10, MH11a, MWZ^{+}16, MKS18, MFI11, MWC^{+}12, MAS^{+}10b, MLB^{+}14, MBP^{+}12, MAK^{+}15, MPB10, MS18b, MRC10b, MJHS13, NKH^{+}10, NBM12, NKH12, NWC^{+}10, ODP^{+}12, OOK11, OFG^{+}18, OFPR^{+}11, PDO13, PL15, PSC^{+}13a, PKB16, PFdP^{+}12, PAF^{+}10, PSAK12, PAC^{+}15, PL10, PC19, PB19, PPDW16, QLCUG15, Rai12, RS13, RPLT10, RRL^{+}19, RCMB18, RJR18, RUP11b, RUP11a, RZH^{+}16b, RLTL14, RLAGL12, RP15, RG14, SDMB15, SWMK15, ST10a, SP13a, STK^{+}13]. growth [SGT^{+}14, SH12, SKR^{+}13, SAM18, SP15, SBS^{+}13, SWBMJ17, SH14, SG18b, SLS16, SHM10b, SGD^{+}12, SNT^{+}16, TG12a, TLFS14, TTF^{+}11, TS16, TPA^{+}13, VPDPMRC11, VCDP11, VDS11, WHI^{+}18, WWO^{+}15, WEA^{+}17, WSK^{+}16, WP10, WWW^{+}13, YKG^{+}15, ZLCW15, ZLGH16, ZGK^{+}19, ZZG^{+}19, dMBC12, dlSGS17, vDJF15, vdGvdLvG^{+}19, RGB^{+}11]. growth-increment [MBW10]. growth-selective [JNYT13]. grubei [CRJ13]. grunion [BBC12b]. grunt [MDA^{+}10]. grypus [HHS^{+}12b, vBBW^{+}14]. Guam [PBC^{+}15, RBL^{+}17]. Guanay [WBS^{+}12]. guano [WBS^{+}12]. guano-producing [WBS^{+}12]. guarantee [CMC10b]. guarding [MS18b]. Guay [MHO^{+}11]. Guiana [CdTK^{+}16, VVB19]. guide [FPL^{+}12, JGCRS12]. guides [LANE18]. guild [AGV^{+}11, DW12, DWK13, EMJ18, HBS17, LD17b, NGA18, SGS^{+}19]. guilds [PJMP10, WLM^{+}13]. guillemot [ERB^{+}13b, ERB^{+}13a]. guillemots [FSP^{+}17]. Guinardia [POS14]. Gulf [BW13, BBB^{+}14, Cra12, FCR12, GLM^{+}19, MMC^{+}18, TEAO^{+}14, WDNM11, BJCB^{+}14, JMO^{+}19, OFDG14, AGGAO^{+}18, ANWG18, ALS18, AGRPSVL13, AGH^{+}12, BRMT11, BCM^{+}14, BPOS^{+}16, BSB^{+}11a, BRB11a, BWH^{+}19, BW10, BAE^{+}10, BMF^{+}15, BLC^{+}16, BDVP19, BOCB10, BHT10, BM16, CTRBLC^{+}13, CL15a, CDR12, DGCH18, DPS16, DEP19, DD15, DFGH^{+}15, DPM15, DRWQ12, DHJ10, DAB17, ECZD14, FNF^{+}18, FGTS16, FOL12, GSC14, GKBEF18, GW15, GGMGR12, GPRR17, HS15a, HGS14a, HGS14b, HWSK19, HFM^{+}15, HW11b, HSS18, HRS^{+}16, HCJ^{+}11, HAHM16, JGG^{+}15, KAP^{+}12, KMV18, LCGS^{+}19, LABG^{+}12, MVVGAV10, MH19, MLCC19, MHR19a, MHW^{+}12, MPB15b, MBF16, MLR12, MRC^{+}13, NRK^{+}19, PBV^{+}11, PSG^{+}13, PK13, PCH11, PPS^{+}13, PW16, PWA^{+}17, PRBR11, RBP^{+}10, RCMB18, RFMT12, RCF16, RSHM10, SPSG14, SAZHAAHS14, SRP^{+}19, SDKS19, SL11b, STT14, SM17a]. Gulf [SBKA14, SJS^{+}11, SMF10, TBH^{+}18, TPS^{+}13, TMS14, TNN11, Urb12, WLM12, WWKS15, WGS16, WWAL13, ZMS^{+}14, ZAZS11]. Gull [RSTM14, CAC^{+}12, CBC^{+}13, CPF^{+}14, OMN17, RGEA^{+}15]. Gullmarsfjorden [HTHH11]. gulls [MPRPM^{+}18, TRSB^{+}18, WPB^{+}14, YTM^{+}12]. guns [JHNL12]. Gusmão [HBCK13]. Gut [DVA^{+}18, GRBH15, NLK^{+}19, RSS^{+}16]. guts [LRBSJ15, WS10]. guttatus [KB14]. gymnodinioid [MDAPH11]. Gymnosarda [BHN^{+}14]. gymnosomatous [SOT15]. Gyre [DA11, DBDK14, LKLC11, RKC16, CPM^{+}17, FSP^{+}17, TSS^{+}12]. gyres [PPSR15]. Gyrodinium [CIM^{+}13].

habit [LHHG11]. Habitat [AM13, ABM12, AGRPSVL13, BTS10, BEG^{+}13, BSG^{+}17, BJL^{+}17, CSM^{+}19, CCNB15, CTS^{+}15, CMH14, DA10, EPC^{+}19, EMJ18, FMP13, GHGF12, GJ16, GMK18, GSU12, GDPR18, GPC^{+}15, HS15a, HBM^{+}15, HW13, HMJ^{+}13, HNG^{+}11, IDT^{+}18, KBA17, LWG14, LMBM18, LF14, MPH^{+}11, MUHC11, MWD10, MIV^{+}10, MRC^{+}19, NPLT10, PCH^{+}14, PPC^{+}12, PES12, RKD18, ROKS14, SL11a, SFW^{+}14, SM17a, SFM^{+}19, SWBMJ17, SSS^{+}10, Sti13a, TRHM14, TCA^{+}11, TRMC11, TGI^{+}11, WY15, ZAMB10, ADHP12, ALV12, AB11, AHØ^{+}15, AS14, ADO^{+}12, AMM^{+}12, AMP^{+}11, ACRLB12, AED^{+}12, ABA16, Ano17c, AHPKS17, ASK^{+}15, ACMD11, BBM^{+}14, BGB^{+}17, BAL14a, BAL14b, BB17, BMV^{+}18, BOC^{+}16, BFF^{+}10, BMEL11, BMMS10, BDvGW16, BRBH19, BWH^{+}17, BDH^{+}17, BAH^{+}15, BMPH14, BL17, BK14, CLB^{+}14, CR14, CLH^{+}15, CMC^{+}16, CWK^{+}10, CLP18, CGD10, CdTK^{+}16]. habitat [CCHG10, Cha13, CWEE19, CGH11, CHST16, CAG13, CPM^{+}17, CWB^{+}15, CAB^{+}19b, CWMP13, CM11, CDCS12, CTvdM^{+}08, CMLC12, CH12, CR11, CNJ^{+}18, CNW^{+}12, DTC^{+}17, DWPH11, DR16, DGA^{+}15, DTS14, DER^{+}19, DCB^{+}16, DT15, DQM^{+}19, DGC12, DH18b, DPD^{+}12, DAM^{+}17, DBE^{+}18, EOMS15, ETHB17, EBdlTC12, FCR14, FA11, FGTS16, FHG^{+}14, FWF^{+}10, FSR17, FCS15, FKvdV12, GDP^{+}11, GLM^{+}10, GMP^{+}11, GS17, GSO^{+}18, GGFB15, GSF18, Goy14, GGF^{+}12b, HFM12, HLIK16, HPL10, HPA17a, HPA17b, HMKF11, HSF^{+}12, HMO^{+}12, HBC15, HGFA^{+}11, HRT12, Hof10, HIH^{+}17, HRS^{+}16, HDT^{+}12, HSS^{+}19, HSWS16, IME12, JW11, JRW^{+}14, JGJ15, JH11, JWT^{+}14, JL12, JC17, JT10b, KSHS19, KSF18, KLKM12, KWM^{+}17, KNS^{+}15, KHW^{+}12, KBK^{+}15, KCS12, KHS^{+}11, KBPD10, KCL^{+}11, KJC^{+}15, LWM15, LI18, LH11, LDS^{+}15, LBJ^{+}16, LMR^{+}10, LKL^{+}17, LR18, LEB14b, LCR^{+}16]. habitat [LC18, LNP12, LHHG11, LA12b, LDP^{+}17, LEB^{+}19, MGFW13, MBJ16, MGP12, MSF18, ME11, MPKO11, MHF^{+}19, MC16, MHW^{+}12, MPG^{+}10, MSG11, MHAP18, MRH11, MRC^{+}13, MCT16, NGJHB12, NAG^{+}16, NGR^{+}17, NHTS14, OAB^{+}19, PAG19, PSCS16, PLM^{+}16, PSHG16, PWM18, PRLR^{+}18, PSA12, PM16, PMSGBL16, PDP^{+}10, PL10, PBCA10, PJP^{+}11, PMM^{+}11, PBC^{+}12b, PSL13, QHG^{+}19, RPM^{+}13a, RMP^{+}19, RRRF^{+}17, RBSL11, RHMR18, RCS^{+}17, RCGE12, RNFG10, RL11, Rit17, RBHR19, RMNP17, RHS^{+}19, RT17, RMB18, RMRA15, SAWK18, SJB16a, SHK14, SWM^{+}19, SBT12, SRP^{+}19, SNS^{+}13, SATS13, SHB^{+}13a, SSR^{+}10, SW18, SBCA15, SSB^{+}11, SBM^{+}16, SGT12, SGG^{+}12, SWH^{+}11, STSS^{+}16, SPBS14, SD17b, SCF^{+}17, SMR10, SKS18, TCSA17, TCJ^{+}19, TCN^{+}10, TW16, TSCSB15, TSGW10, TMP^{+}11, TLAT13, TSA^{+}18, TMH^{+}13, TDSS15, USR^{+}16, WKvW10, WJWC16]. habitat [WM10, WTD^{+}19, WLB^{+}11, WK18, WHH12, WWZ^{+}19, YYO14, ZO10, ZMS^{+}14, ZMRCP^{+}17, vDGK^{+}12, vdWHG11, vdWTWS15, vdGvdLvG^{+}19, vdMPRM12]. Habitat-associated [SL11a, LEB14b]. Habitat-based [WY15]. habitat-building [AHPKS17]. Habitat-diversity [TCA^{+}11]. Habitat-forming [TRMC11, AS14, BB17, BMMS10, MGFW13, MHF^{+}19, PBC^{+}12b, TW16, TSCSB15]. habitat-modifying [JT10b]. Habitat-modulated [LMBM18]. Habitats [MBF16, AIA18, AL12b, BHE^{+}17, BSC^{+}16, BWP^{+}11, BBdG^{+}19, BPSK11, CEH18, CTSH12, CCT12, DLRF15, DPO^{+}17, DGP^{+}19, DPM15, DHAV13, DPH13, DFAH11, ECZD14, FHTY19, FDS12, FRJ10, FRJ11, FHMG19, FG16b, GRN14, GCL^{+}16, HKL12, HKL19, Har16, HMJ^{+}13, INvdB^{+}13, JWBS10, KKS12a, KFB15, LHT^{+}16, LEB14a, LYH^{+}17, MdNW12, MUC12, MUHC11, MIIB16, MD11, MGCH^{+}16, NHS^{+}17, NMG^{+}13, PDM12, PMK^{+}15, PHI^{+}12, PTP^{+}19, PLA^{+}12, PMM^{+}15, PWA^{+}17, RST^{+}10, RSJ14, RIM^{+}13, RTÁB^{+}14, RCGE12, RTD14, RB17, RMV13, SF16, SCRL11, SML^{+}15, SJL^{+}16, SLO^{+}17, ST10b, SPP15, SOA^{+}17, TS16, WJW^{+}18, WBJ12, WPP11, WS11, YHTS12, ZPPM17, vDBH^{+}15]. habits [BDS^{+}18b, CNÁCP12, DGFH10, FC10, HEAI10, KHT^{+}17, LLBP15, RLAGL12, Urb12, ZRR^{+}13]. hadal [FKRJ13]. haddock [AGIW16, BBS^{+}17, BCSB13, BLM10, MRE14, RP12, SKN^{+}15, TDPS13, WTGG10, WGGM11]. haematological [DTL^{+}15]. Haematopus [GCBS19]. haemorrhagic [JBM^{+}13]. Haemulon [MDA^{+}10, YAL12]. Hagerty [Ano19a]. Hagfish [ZRR^{+}13]. Hagihara [Ano12c]. Hainan [SLS16]. hair [RWR^{+}15]. hake [BGIA19, KW17, MBV^{+}11, RVH^{+}19]. half [CBT13, CHFS12]. halibut [BNS^{+}14, NHTS14, PHH^{+}12, WYS^{+}10]. Halichoerus [HHS^{+}12b, vBBW^{+}14]. Halichondria [KRC17, LRSB19]. Halimeda [ATT^{+}16, CCM^{+}14, PKB16, PHTS11, YMCV10]. Haliotis [BRB11b, CHB^{+}13]. haloclines [BLD^{+}16b]. Halodule [LMRRB17, MH13]. Halophila [BCGWS19, GALS16, SD17a, SMK^{+}17]. halophyte [PÁIIS^{+}19]. haloplanktis [SP15]. halos [DQM^{+}19]. hamatus [SMR14]. handle [SZ13]. Haney [SB15a]. haphazard [DAD^{+}11]. Haplosporidium [CB11]. haplotype [MPLLE17]. haplotypes [PRK^{+}12]. Haptophyta [HGL15]. harak [LCG^{+}17]. Harbor [LCJ^{+}12, RBS16, TLJ^{+}11, BMJ16, BWP^{+}11, BM16, GKHM13, HMAC^{+}11, HAHM16]. harbors [Bri12]. Harbour [AvBBL^{+}16, IME12, TLM^{+}11, BDvGW16, BBD^{+}14, BEG^{+}13, BDBN11, BHD^{+}13, BDD^{+}18, DTC^{+}17, KLBT^{+}16, KJC^{+}15, MMH^{+}17, MMD^{+}13, NWH^{+}12, NTS^{+}18, RLN17, SBG^{+}16, ST08, SNNS^{+}12, WBS^{+}17, vBBW^{+}14]. Harbouring [FFS14]. hard [BBWG12, GLF^{+}17, GKE^{+}19, GBH^{+}10, GP12, HP14, HB15b, dDTC^{+}17, MW16a, MCW11, MAL15, Mus12, NGP^{+}15, PMS13, TP19, WBG10]. hard-bottom [GBH^{+}10, Mus12]. hard-part [dDTC^{+}17]. Hardison [Ano10f]. hardwicke [Gea10]. harengus [BSST13, BPM^{+}18a, CBMK10, EOMS15, LFJ^{+}14, PSÓ^{+}15]. hares [TOW10]. Harmful [LMT15, FPC^{+}12, FTH12, LBM^{+}18, LLS^{+}18, RTC^{+}14, TG12a, WRGP12]. harming [GCBS19]. Harmonizing [OTK^{+}13]. harp [BKAS14, GNHLN11, ØLH^{+}13]. harpacticoid [AGC12, CMW^{+}15]. harpacticoids [WVD^{+}17]. Harriotta [DGFH10]. harsh [FNA^{+}16]. harshness [CPF^{+}13]. Hart [SCHO11a]. harvest [AB17, BFM^{+}13, GCBS19, HDB^{+}10, MKK^{+}11]. Harvested [DS13, BSH^{+}18, CFSM^{+}13, JSK19, ORJH19, OAD^{+}14]. harvesting [BCS^{+}12, CHE^{+}17, DHR^{+}13, DH17, JDB^{+}10, KHA^{+}12, MCTHGV12, OSPWB14, SY10, WD14]. Harwell [BBM13]. hasitata [HHS12c]. Hatch [MS18b, MGO^{+}12, PSHG16, Ric12]. hatchery [ABK^{+}14, EHD^{+}17, HHP^{+}17]. hatchery-reared [ABK^{+}14]. hatching [LANER15, MJT^{+}14, RGW17, SHP^{+}14]. hatchling [NSS10, PK15]. haul [HKL15]. haul-out [HKL15]. haulout [ED12]. haumeaae [Par15]. Hawai’i [CRWG16, FHB^{+}12, GFC^{+}14, SMOH19, WGD^{+}19, DBB11, ETS^{+}11, SKD17a, SMP10]. Hawaiian [LFPA^{+}11, PIPD10, BHP12, DWK13, FJD13, HS17, MDKM13, ODP^{+}12, OAB^{+}19, PGG12, Par15, PPTK19, TB14b, VDS11, WRP^{+}17, WRI12]. Hawkes [Ano19b]. hawksbill [BP12a, GLW^{+}12, GLJ^{+}18, HSF^{+}12, HTR^{+}12, KCL^{+}11, LSFQ19]. Head [DBE^{+}12]. headed [MHHB^{+}17]. Heading [HVD^{+}16]. headlands [CGH11, QHG^{+}19]. Health [FEA^{+}15, RSP^{+}16, GLHM^{+}18, GCP^{+}17, NWB^{+}19, NJP12, SGEO13, SBL^{+}10, VDG^{+}15]. healthy [BLD^{+}16a, FG16a, PWCvO^{+}14]. healthy-looking [FG16a]. heard [RTMJ11]. hearing [CMR12, WHL11]. Heart [WPMP14, CBKM16]. Heat [SMF^{+}12, BPF15, DSVM10, DHGW15, FC19, HGC15, JG13, KF13a, MMB^{+}12b, TB14a, WNF^{+}11]. Heat-seeking [SMF^{+}12]. heatwave [NDAP19, vBAP^{+}19]. heavily [LG18, MA14, NBM12, SCB15]. heavy [CFC^{+}12, FAWBF10]. heavy-metal [CFC^{+}12]. Hebrides [DJL^{+}12]. Hediste [CML^{+}12, HKS^{+}12]. height [CAB^{+}14, CLL17, GHCC15, Hoe10, SS19]. heightened [HHS12c]. Helgoland [SdBAH^{+}14]. helgolandicus [YBH^{+}11]. helianthus [EN11]. Heliaster [EN11]. helicina [BNE^{+}19, CAG12, SLW15]. Heliocidaris [FDP^{+}16]. helminth [TP16]. help [HPH^{+}18, TMSG19, TDPS13]. helping [HFG^{+}19]. hematocrit [CTS^{+}15]. Hematodinium [SHMOS17]. Heme [GGLT15, HGB^{+}13]. Heme-b [GGLT15, HGB^{+}13]. hemi [BB15]. hemi-sessile [BB15]. Hemigrapsus [CFMN16]. Hemimycale [GBU15]. Hemisphere [HER^{+}19, WNNN12, HEM^{+}12, LGM^{+}12, LCB^{+}12]. hemocyte [BFL^{+}11]. hemorrhagic [HGG^{+}10]. henlei [SCB15]. herbivore [CFF^{+}18, CSB14, DVSM10, DQM^{+}19, DAM^{+}17, JL17, SBMÅ12]. herbivore- [DVSM10]. herbivore-induced [JL17]. Herbivores [DHAV13, ACS12, BDT^{+}16, CCM^{+}14, DH16, FMC^{+}10, Hoe10, HMM14, JMPZ10, WBB^{+}10, YRL^{+}10]. herbivorous [BGSM13, BDB^{+}13, CVP11, DBTV13, HS17, KMO19, LS19, MHVV13, PBH^{+}13, PH11, PP18, Sti13b, VHKT15]. Herbivory [ABRP15, EOM^{+}14, FCR14, FWB^{+}15, LHC^{+}15, AED^{+}10, BDT^{+}16, CDWC17, CAB^{+}14, DBTV13, FRC13a, HKJ17, HSPS14, JREVB18, KAMA13, LB11, MTV^{+}11, MHVV13, NAG^{+}16, PRFA11, PRA10, SSL^{+}14, SV15, VTB12, YLPG19, dlSBO^{+}12]. Herding [KROA19]. herds [LMLB12]. Herfort [Ano14b]. Heritability [LALÁC11]. heritable [BPB18]. hermatypic [RMB^{+}14]. Hermes [SGC14]. hermit [PDB15]. herring [BCM^{+}11a, BSST13, BPM^{+}18a, BDC10, CBMK10, DVA^{+}18, EOMS15, FESPR14, FPR18, HGG^{+}10, HHNI17, HFR^{+}13, HEM^{+}12, HFH10, JBM^{+}13, KRW18, LFJ^{+}14, LSJ^{+}15, LET^{+}12, PSÓ^{+}15, PSC^{+}13a, RNB^{+}12, Ran18, SVAV17, SNVH19, SFW^{+}14, SPP18, SNNS^{+}12, TLJ^{+}11]. herring-eating [SVAV17]. herrmanni [ETB13]. hetero [VAHL11]. hetero- [VAHL11]. heteroclitus [LFD13]. heterococcolithophore [CE13]. heterogeneities [WBHL10]. heterogeneity [BS14a, BVL10, BBMH13, BAVSP11, DZH^{+}11, GLB^{+}10, HKL12, PJDC^{+}13, PLNE11, RBC^{+}14, dML11]. heterogeneous [BA13a, DDB^{+}16, DHR^{+}13, JLM^{+}12, MUHC11, RJH14, SNT^{+}18]. heteroplasmy [BRH13]. heteropods [WPML^{+}18]. Heterosigma [QYS^{+}12]. heterospecific [MB11]. heterospecifics [Gea10]. heterotrophic [ALSO^{+}18, BBRBG10, CIM^{+}13, CL15b, FBNJ13, JDQS11, LLRC^{+}12, MGFÁS^{+}10, OAGV11, OMSP10, SKR^{+}10, VARO12]. Heterotrophy [TPFP11]. heterozygosity [FCR12]. Hexactinellida [CL10b]. Hexechamaesipho [TCWC13]. Heye [EN13]. heyseana [MKK13]. Hg [CSC^{+}11]. Hichirippu [KTK^{+}12]. Hidden [BFL^{+}19]. ’hiding [FWB^{+}15]. Hierarchical [KWH^{+}18, TAEB^{+}11, BBBNT13, PMKL11, USR^{+}16, Vig15]. hierarchies [GSS13]. High [AM19, AGM^{+}11, AHSG16, BDM^{+}13, BWCZ11, CL10b, DDP13b, FLK12, FRMS17, GSC14, HBHS^{+}18, HTR^{+}19b, JDX^{+}17, JBM^{+}13, Joh11, KF13b, LPC19, TTS17, VSB^{+}15, AGGAO^{+}18, AZSKWL11, BVS^{+}16, BCGWS19, BAS16, BS19a, BMF^{+}12, BH10, BSCH14, BS19b, CR16, CGK13b, CBT13, CH15, DGC18, DRD^{+}18, DWL19, FRRÁSL11, GRN^{+}16, GSV14, HGG^{+}18, HP13, HS14a, HSRP11, Hv14, KWC^{+}17, KLBT^{+}16, LBN11, LMLB12, LSFQ19, LE13, LERS13, LMH12, LD17b, MTV^{+}11, MHS17, MMB^{+}19, NvOH13, PC12, PCNM17, PAK15, QNG^{+}15, QGM^{+}14, RT17, RUS^{+}13, SCSC16, SCP16, SLL12, SH17, TMS^{+}16, TPB^{+}18, TFÁS^{+}12, TSS11b, TBL16, VHBG15, VJM^{+}16, WB12, WGD10, Wil14, WNF^{+}11, YKG^{+}15, vDCR13, HB15a]. high- [WNF^{+}11]. high-Arctic [BVS^{+}16, HB15a]. high-energy [QNG^{+}15]. high-latitude [GSV14, NvOH13, WGD10]. high-light [SLL12]. High-marsh [Joh11]. high-quality [BMF^{+}12, BS19b]. high-rainfall [CGK13b]. High-resolution [GSC14, LD17b]. high-risk [KLBT^{+}16]. high-throughput [CR16]. High-use [BDM^{+}13]. Higher [DMJ11, CLG^{+}19b, CRGM18, DMH15, ERG^{+}11, RGdLB^{+}11, ZJ17]. highlight [BBC^{+}17]. highlights [CCG^{+}19]. Highly [JCM10, NHS^{+}17, TMM^{+}19, APC^{+}13, BOC^{+}16, CEH18, CCM^{+}12, JDQS11, KCF^{+}14, LJM^{+}10, MLM^{+}12, MKK13, MAP^{+}12, PFG10, SCL^{+}19, SMMFH16, SAD^{+}16, TCN^{+}10, TM10, TTW^{+}12, TGI^{+}11, WHA^{+}13, vERRP16]. Hills [CBKM16, GSJS13]. hind [GFC^{+}14, RSA^{+}12]. hindcasts [ORW15]. hinders [SES17]. hippoglossoides [BNS^{+}14, PHH^{+}12]. Hirst [GMR13a]. hirundo [Goy14]. hispida [BVM^{+}19, CSF13, HLIK16]. histocompatibility [JCF^{+}12]. Histopathological [MVCD14]. historic [GKBEF18, WWAL13]. Historical [ŁCM^{+}18, BW15, CCJR14, DHJ^{+}17, KGJ^{+}16, RK15, vdHTT^{+}15]. histories [LPB^{+}16, OFFQGG10, PFdP^{+}12, VM14]. history [ARW12, AMC^{+}11, AS14, AHB^{+}14, AKF14, BWV^{+}12, BDS18a, BFD^{+}16, BOKJ18, BEFC15, CAR^{+}19, CRJ13, CLMB13a, CLMB13b, DMD^{+}12, DP13, Fog12, GFC^{+}13, GSV14, HWET13, HS14a, HEB^{+}11, HJ13, HDP^{+}17, ICT^{+}11, JMK12, Ji11, JS12b, KFW^{+}18, KPW^{+}13, KCGC14, KKPL14, LAB10, LGBSC11, LALÁC11, LLF^{+}18, LOES11, MM12, MH12b, MMA18, MRB^{+}11, MDR16, MGD^{+}11, NWS^{+}14, NWC^{+}10, OHA10, ORWPR14, OGH18, OMW^{+}10, PKW^{+}11, PSG11, RWP^{+}14, RBC^{+}14, SRSH^{+}18, SKMS^{+}18, SAM18, SMK^{+}17, VRAP^{+}14, WSHY16, WBW^{+}14, WHI^{+}18, WFH11, ZBP18]. history-based [Ji11]. hitchhikers [MBB11]. Hitchhiking [LLS^{+}18]. hog [PSA17, RWR^{+}15]. hoki [AIA18]. Hokkaido [KTK^{+}12, NKM^{+}15]. holdfasts [FFS14]. holes [WDNM11]. Holo [CE13]. Holo-heterococcolithophore [CE13]. holobiont [RET^{+}19]. holography [THM^{+}13]. Holothuria [CLP18, SK10]. holothurians [WHB^{+}14]. holothuroid [VHM16]. Holothuroidea [SK10]. holothuroids [PPDW16]. Holt [Ano10d]. Homarus [CCHG10, CS11b, LS10b, MKS18, MCC^{+}16, MLF^{+}18, MOS10, MOK^{+}11, MR16, RS13, SRM^{+}15, SPS16]. Home [BFN^{+}19, HSF^{+}12, MLM^{+}12, SJB^{+}15, AAE14, AHA^{+}15, BBKP10, CHB^{+}13, MPA^{+}10, NMS^{+}18, PBM^{+}11, SGM^{+}14, SJLH12, SJL^{+}16, SRM^{+}15, TWL15]. home-ranges [SRM^{+}15]. Homing [Ued14, Dan11, DHM^{+}14, GJ16, LSFQ19, RAF^{+}14]. homogenization [BAvW12]. Hong [GSM^{+}11b]. Honokōhau [WBB^{+}10]. Hopen [DAW^{+}11]. Hoque [Ano12c]. horizon [PK15, BMC^{+}15, BSF17, DBCS16, FCR^{+}15, GFF16, GPRR17, HGS14a, HGS14b, HGS15, JGG^{+}15, KCMH^{+}17, PGR^{+}17, PPCH17, RFH16, RBS^{+}17, SB15a, ZWP^{+}17]. Horizontal [DBHMG12b, Sat10, TRGF18, BL14, BPSCD17, BOU13, MAHK12, SSB^{+}11]. Hormosira [BMMS10]. Horns [BDBN11]. horohalinicum [TSS11a, TSS11b]. horribile [TPB^{+}18]. horse [BNAK11, KAP^{+}12, YAFV12]. horseshoe [WJWC16]. Host [DMHS^{+}14, HDT^{+}12, LKN^{+}13, PC19, RR18, TDKS10, DLdM19, DDDR10, DDM^{+}16, FTO^{+}15, FSE^{+}17, HD10, JDD^{+}13, LRRM^{+}11, MB16, MPLLE17, PHD^{+}14, RLM^{+}14, STP10, SWPK13, ZRBV15, dML11]. host-parasite [STP10]. host-symbiont [ZRBV15]. hosts [HWSK19, JDD^{+}13, KBK^{+}10, SS10a, SJHH^{+}17, TMBP13, TP16]. hot [AMLS15, DRG^{+}12, FC19, SGC17, SSM15, SMS15]. hotspot [DPH13, FDFS16]. hotspots [BFNB16, HBS^{+}13, LKL^{+}17, SRLV10, SV13, SDSA14, SSR^{+}10, SDM16, SSS12, TMSG19, UW18, YYO14]. HP15 [GRG^{+}12]. HPLC [LCS^{+}10]. hubbsi [BGIA19]. Hudson [CSF13, GE14, JHL^{+}19, dlGMD^{+}17]. Human [FRJ11, RGEA^{+}15, RUD^{+}16, BWV^{+}12, BWLP17, BAVSP11, BPT15, BHDC15, CW12, DHJ12, EOMS15, HS15a, HTR19a, PMOAC^{+}17, SLSR10]. human-caused [BWLP17]. human-mediated [CW12]. human-modified [EOMS15]. human-oriented [HTR19a]. humans [LWD13, MF13]. Humboldt [ASEB11, PBC^{+}18, FLC^{+}14, ICT^{+}11, MBS^{+}13, MBS^{+}14, RHRB^{+}17, SZG11, SHF^{+}12b, TKA10, WBS^{+}12]. humidity [RB15]. Humpback [FMG^{+}18, LFPA^{+}11, ACP^{+}16, BSC^{+}13, BDH^{+}17, BJS^{+}16, CTZ^{+}16, CGH11, CJS^{+}12, FBA11, FTS^{+}13, KWH^{+}18, MHAP18, OGGP14, OWN^{+}15, RBP^{+}10, SWK^{+}16, SGG^{+}12, TFW^{+}12, VCG^{+}10, WNNN12, WEA^{+}17]. humpbacks [KWM^{+}17]. hundreds [VW10]. hungry [GRL^{+}15, MW16b]. Hunted [SCF^{+}17]. hunters [SCF^{+}17]. Hunting [DGC18, WPS^{+}19, AAM^{+}12, AAK^{+}15, BNAK11, JLPG17]. Hurricane [FDS12, SHT^{+}13, HHS12c, BGM^{+}18]. Hurricane-mediated [FDS12]. hurricanes [SFLG10]. huxleyi [ENW^{+}14, GAF^{+}16, HWE^{+}12, JDX^{+}17, KAP^{+}14, LBM^{+}18, MTH15, PYBB11, SLIGG18, TGI14]. hyacinthus [CESC14, LPC11b, PWCvO^{+}14]. Hyalinoecia [CLN^{+}13]. Hyas [SHP^{+}14, WAP10]. hybrid [LHU^{+}18, RGWA11, SHH10]. Hybridization [BH10, FCV^{+}16, LHU^{+}18, TBP11, TDKS10]. hybrids [Fog12, PPJT15]. hydrate [ÅCAC16]. Hydraulic [VPW^{+}12]. hydro [PHG^{+}14]. hydro-climatic [PHG^{+}14]. hydrobiological [ZVL^{+}16]. hydrocarbon [ACF^{+}10, CFC^{+}12, JGL12, YYD^{+}15]. Hydrocharitaceae [SD17a]. Hydrodynamic [BFRM13, LBP^{+}16, KCD13, LFU^{+}19, LS10b, STK^{+}13, VDM^{+}10, VHG^{+}18, dlSBB^{+}10]. hydrodynamics [DBE^{+}18, LTN19, PBIO^{+}18, VBR19]. Hydrographic [HRÁB^{+}14, LCQT^{+}13, SMD^{+}16, VLK^{+}15]. hydrographical [FHM^{+}11]. hydrographically [FMH^{+}15, HOA14]. hydrography [TNM^{+}11a, WDM^{+}18]. Hydrolagus [DGFH10]. Hydrological [KMV18, LPC19, PP13, RMV13]. hydromedusa [KBR^{+}11]. hydromedusae [CUCS17]. hydrostatic [SBT15, YOST11]. hydrothermal [CKV12, FC13, GGFB15, KM10a, KLSJ10, LZF^{+}14, Met11, NUO^{+}19, RSW^{+}16, RCB^{+}10, SNW^{+}18, SJL^{+}15, SLRS13, YWiIK15]. hydrothermal-vent [CKV12]. Hydroxamate [SP15]. hydrozoan [RBBG12]. Hydrurga [MCS^{+}15]. Hymenodora [KHS10b]. Hyperbenthic [MBS^{+}18]. hyperborea [BRG^{+}14, NCR^{+}14, PNF^{+}12, PNPF14]. hyperboreus [Hal15, HJMN^{+}12, SLNG18, SKD^{+}11b]. Hyperiid [GFGGDSM12, FHG^{+}14]. hypersaline [DE17]. Hyporthodus [BJCB^{+}14]. Hyposalinity [FLS^{+}18]. hypotheses [RTÁB^{+}14]. hypothesis [DHSL11, DHF^{+}13, GFS15, JGJ15, LET^{+}12, MSE^{+}18, RLS10, TLS^{+}19]. hypoxia [ARB^{+}12, ACW18, BRK^{+}12, BSN^{+}18, BT13, CPSC^{+}15, CPSC^{+}18, CR14, CG16, Cra12, DTG16, HSZR10, KBW^{+}15, KR16, KBPD10, LBAK13, LSQ^{+}13, LC13, LTG17, LCL^{+}11, LM18, MBBK16, MW16b, PDBB14, RHB13, RZS12, SHPS^{+}16, SNS^{+}13, SCP^{+}14, SBE^{+}19, SG18b, SSD13, ZMS^{+}14]. hypoxia-induced [CR14]. hypoxic [BRMT11, EMMSL13, FHE^{+}15, KMV18]. hystrix [RMB^{+}14].

ianthina [YTM^{+}15]. Iberia [MCTHGV12]. Iberian [APC14a, MGFÁS^{+}10, TMGCM11, AFSL^{+}15, FAZ^{+}14, GSP^{+}15, MATM^{+}19, NDD^{+}13, RCG^{+}15]. IBMs [DSG15]. Ice [ANRT10, MWA17, AFMR^{+}13, AFRAG16, Ano18a, BCH^{+}18, BBG^{+}16, DLRF15, DAW^{+}11, DJJ^{+}16, EMG^{+}15, FDLL10, FKA^{+}12, GLS^{+}15, GGL^{+}17, GCC^{+}14, HLIK16, HKI^{+}17, HPN^{+}17, HRK^{+}18, JUK^{+}10, JFK12, JKM11, KFBM19, LSY^{+}18, LOS^{+}19, LHJ11, LWS^{+}10, MAT^{+}11, MWWS17, MCS^{+}15, MTMM^{+}15, MGG^{+}14, MLSK^{+}15, ORE^{+}17, PMF^{+}17, RCS^{+}17, SKD^{+}17b, SLL^{+}18, SKR^{+}10, SDMR19, TTS17, TMP^{+}13, UFP^{+}10, YSH^{+}16, dlGMD^{+}17]. ice-associated [EMG^{+}15]. ice-benthic [KFBM19]. ice-breeding [DJJ^{+}16]. ice-covered [AFRAG16, SKD^{+}17b]. ice-scoured [MLSK^{+}15]. iceberg [KHD19, RJH14]. iceberg-mediated [KHD19]. iceberg-scoured [RJH14]. icefish [Lap19]. Iceland [GPM^{+}13, JBS12, RZH^{+}16b, SVAV17]. Icelandic [OA10]. Ichii [Ano12d]. Ichthyoplankton [BOCB10, DABP13, HCGP13, KBPD10, PTS^{+}18]. ichthyotoxic [GG16]. Ichthyotoxicity [MDAPH11]. iconic [BFH13b]. ideal [Wir12]. Identification [FTvdK^{+}12, HBHF^{+}11, Hof10, KLBT^{+}16, OHST19, SGG^{+}12, WLF^{+}17, CJS^{+}12, DHB18, LJM^{+}12, MKK13, PMM^{+}15, WRB^{+}15]. identified [ALV12, CR16, CMLC12, FGW^{+}15, PW11, THW^{+}16]. identify [CWR^{+}18, CSC13b, GPA^{+}16, GM10, SSS12, TMSG19, WWG^{+}13b]. Identifying [CHRB10, FHMG19, FWF^{+}10, HHWW11, LNH^{+}17, LMLB12, ADD18, JRK13, PCP^{+}15, ZBY^{+}14]. Identity [FGP17, MHVV13, BHF^{+}10, DAM^{+}17, FCS15, HBS10, KKS^{+}12b, MUHC11, MD16a, RHE14, ROLA15, SS19]. Idiosepius [Sat17]. Idiosyncratic [FBGT12, SV12]. Idioteuthis [BB14b]. if [BB14a, Bar14]. ignobilis [DFD^{+}19]. ignoring [MPM^{+}18]. II [ANJ^{+}11, BDBN11, Gal14b, HGS14b, HT11b, KAP^{+}14, PCNB16, RGP11, SMS15, STBP13, WMS^{+}11]. Illegal [MTB^{+}17]. illuminated [OMW^{+}16]. illuminates [SPP18]. image [DBS^{+}16, RST15]. imagery [FMH^{+}15]. images [TAEB^{+}11]. imaging [BOS16, FH18, PMF^{+}17]. imbalances [RPM12]. imbricata [BP12a, GLW^{+}12]. immature [PVZD^{+}11, SL11b]. immediately [HJ16]. immense [Ano18a]. immersion [PSG11]. immigrant [MMNR^{+}12]. immigration [SCST17]. immune [GB10b]. immune-suppression [GB10b]. Impact [AEM^{+}18, CSS^{+}17, DFB^{+}17, FAK^{+}12, GGvW^{+}15, JRvdM15, JPI^{+}19, KGLLW11, KLMR14, LBG17, MRSD15, MTK^{+}18, MRC10b, PMGQM17, POH16, PFL18, RNH15, SKK^{+}18, SGL^{+}17, SPH18, SBM^{+}16, SLW15, VDV^{+}15, VMN^{+}12, AH13, BHP12, BMV^{+}18, BST^{+}10, BT16, BKB11, CB11, CPG^{+}14, FRJ10, FRJ11, FAZ^{+}14, GFF16, GSTK11, GWK^{+}12, HMH^{+}18, KTR^{+}16, LJK^{+}11, MBR^{+}13, MHCRR18, NMG^{+}13, OFM^{+}18, PMML12, RBBG12, SCW^{+}15a, SCB^{+}16, SIK16, TVA^{+}19, WWO^{+}15, ZH13]. impacted [BADQ^{+}15, LLH^{+}11, VKH10]. Impacts [BKF15, CSN12, DBCS16, GSO^{+}18, GRFSM12, GKRG^{+}19, HHI^{+}14, HCG15, MK13, MJ10, NKH^{+}10, POPH12, PWM18, RJR18, SHS^{+}17, SHPS^{+}16, TBKG10, TBS^{+}14, ÅCAC16, BGM^{+}18, BWN^{+}14, CHE^{+}17, CWGG15, CHFS12, CLSS17, COK10, DBA17, DHJ10, GP11a, GKHS16, GPV17, GPRR17, GLW^{+}17, HFM^{+}15, HA15, HTC17, HSA^{+}15, HMM14, JDX^{+}17, JSLG17, KMV18, KFBM19, LPSK17, MPB14a, MMG^{+}16, MM16, MFDW17, MBCB^{+}19, MLF^{+}18, MWMD17, áNGGS11, ODL^{+}13, PMOAC^{+}17, PSPK10, RFM^{+}15, RGdLB^{+}11, SCB^{+}14, SHT^{+}13, SABMG16, SLSR10, TRR^{+}17, VMWU15, WSC13, WEK14, ZHJ^{+}17]. impaired [MKM^{+}16]. impairment [GSCH14]. imperfect [IKS^{+}12]. implementation [HGW^{+}11]. implication [HJN^{+}10, JMK12, MDF10, TDB^{+}14]. Implications [BLV^{+}19, BAS16, BS19a, GMR^{+}19, MBB11, ADD18, ABRP15, AGIW16, BL14, BWH^{+}17, BRC^{+}15, BH12a, BW15, BWCZ11, CLG19a, CCM^{+}14, CMK12, CSC13a, CGD10, CF12, CTSH12, CLGC16, CHB^{+}13, CL16, CZH^{+}13, CCG^{+}19, CE13, CEC^{+}18, DC15, DBB11, EHD^{+}17, ESW14, Fis10, FRTB14, FABH^{+}16, FTM^{+}13, GLW^{+}12, GRL^{+}14, GKHM13, GTP^{+}17, GFR^{+}18, GW15, GMBL11, GS11, GVGH15, GHNF10, HWVM19, HYC^{+}14, JGJ15, JCM10, KVR^{+}18, KHW^{+}10, KBW^{+}11, KHH18, KBC^{+}12, KM17, LCJ^{+}12, LMW^{+}13, LWP^{+}13, LRB^{+}12b, LWD13, LÁBFÁF^{+}10, MM13a, MFH^{+}11, McC12, MC16, MLP19, MWZ^{+}16, MRL^{+}17, MMM^{+}15, MPB13, MOK^{+}11, MBF16, MHR^{+}19b, MCRB11, MCW^{+}17, NTN^{+}13, NMS^{+}18, NWB^{+}19, NSS10, NLKZ12, NYK^{+}13, NSY^{+}18, Nor13, NNT^{+}10, NTB^{+}17, ODGL^{+}12, OALK14, OT13, PDT^{+}11, PW19a, POX^{+}17, PCW^{+}12, PB19, RRRF^{+}17, RHB13]. implications [RU11, RKD18, RWM^{+}10, RCS^{+}17, RSDN15, RBC^{+}14, RPL^{+}18, RT12, SPM^{+}16, SB18, SJL^{+}16, SFMB10, SBB^{+}12b, SBL^{+}10, TFB^{+}17, TGR^{+}19, TMH^{+}13, VQKF14, VJJ^{+}19, VM14, WTdB16, WBS^{+}17, WMPM13, WPMP14, YPWP18, dQSC13, dlSBO^{+}12, vDCR13]. implied [MVP^{+}16]. imply [RDK14]. Importance [BVO16, BHF^{+}16, CWEE19, DAB17, FKA^{+}12, FGH^{+}13, GCC^{+}14, HS14b, MCP^{+}18, NCR^{+}14, ODCE11, RGW17, SCK^{+}17, TB10, TTS17, TGR^{+}19, WBI^{+}15, WMCC15, YMPR13, BWM^{+}15, BRG^{+}14, BJMS12, BGIA19, CSM^{+}19, CHRB10, DP16, DCW^{+}10, FJW14, FGL^{+}12, GH11, HLIK16, JSBP15, JSCM15, KHW15, KHS10a, KTKM14, KBP^{+}10, LCL^{+}17, LRM^{+}14, ME11, MHCS18, MM12, MM10, MADI13, MLB18, MHS14, MJA^{+}18, MN10, PA13, PHWC13, TS11a, TFB^{+}17, dJH11, vdGvdLvG^{+}19]. Important [MC16, BFDM18, BKR^{+}15, BRLM17, BVC^{+}16, CPID13, CHE^{+}17, DAM^{+}17, GS18, GCS^{+}18, MTM^{+}17, MBBK16, MSLR13, OHB^{+}15, POS14, RZH16a, SMH^{+}12, SS19, TvdMM^{+}17, ZPPM17]. improve [LJM^{+}12, MSV^{+}11, PMB^{+}15, RG11]. improved [PLG^{+}16]. improvement [BEEH15, RSG^{+}16]. improves [CWGG15, LYH^{+}17]. Improving [DBH^{+}16, GDA^{+}18, SJLH12, GMES18]. In-depth [HGL15]. in-forest [SJB16b]. in-river [WHD^{+}14]. in-shore [CDCZ16]. inadequacy [WHT12]. inappropriate [TGLC10]. incentives [CS15]. incidence [MHR^{+}19b, SBB^{+}12b]. Incidental [SCHO11b, GZYS11]. Including [CWGG15, Ano18a, CBR19, HAF^{+}17, WBB^{+}10]. income [MA14, YKT^{+}14]. Incongruence [BHA^{+}11]. inconsistencies [AWH19]. inconsistency [HPM14]. inconsistent [CTMU10]. Incorporating [JMK^{+}18, LNS10, Rai12, LC18, WBHL10]. Incorporation [PSF^{+}19, BWH11, CGS^{+}19, CJM^{+}13, HLD12]. incrassata [ATT^{+}16]. increase [Ben16, BAT14, CRHV14, EMTB^{+}12, ELM^{+}19, GGF^{+}14, LMH12, MCBR18, PML^{+}19, RLM^{+}14, SBB^{+}12a, TZF11, TBS^{+}17, VBD^{+}11]. Increased [FDVL18, LWS^{+}10, OMN17, SWMK15, SMR14, BRIC13, Ing16, JREVB18, KSS13, LRTLF12, MKS18, MAK^{+}15, RMC^{+}14]. increases [BG12, CAD^{+}10, CPF^{+}13, CLJF12, DFTZ^{+}15, IW15, JCM13, MGP12, MLCC19, MBBK16, OHB^{+}15, RTMJ11, SB15b, TBH^{+}18, TS14, WDL^{+}18, XTW^{+}15]. Increasing [GCBS19, MCUO12, RDM18, TDM16, BTLF17, GSA^{+}18, JROS17, NNT^{+}10, PLS^{+}10, TVM16, WSC13]. increasingly [DAM^{+}17]. increment [MBW10]. incremental [FLS^{+}18]. incubation [DHH10, FJ12, dlÁGRB^{+}19]. incubations [MHF^{+}14, RGB^{+}11, RGP11]. indeed [MAB^{+}12a]. Independence [MPW12]. independent [BVM^{+}19, LCCT19, MO13, RSG^{+}16]. independently [BB14a]. indeterminate [OFG^{+}18]. Index [Ano10a, Ano10b, Ano10c, Ano12a, Ano12b, Ano13a, Ano13b, Ano14a, KMP14, CS15, CTS^{+}15, LRSB19, SHG^{+}15, dMBC12]. Indian [EBdlTC12, KVR^{+}18, KAMA13, MTG^{+}11, OPW^{+}11, PRP^{+}18, ASWB10, CTZ^{+}16, DR10, ECC^{+}15, FLSG^{+}15, HSH12, HBM^{+}12, JOFP18, MM16, SMD^{+}13, SNS12, TDB^{+}14]. indica [FHTY19]. indicate [CW12, CM14, HW11b, SWH^{+}18, WH11, ZO10, vdVKA^{+}11]. indicated [ASWB10, BVD^{+}14, HTH10, RBBG12, SSR^{+}10]. indicates [ABB^{+}14, AAK^{+}15, ACMD11, BFH13b, KNB^{+}13, LCGS^{+}19, LDAE12, LTN^{+}11, PGOG12, YAFV12]. indications [MJT^{+}12]. indicator [CPR^{+}16, EBDB18, HB12, KIU^{+}13, LANE18, RSA^{+}12, SOB16]. indicators [BMJH12, CDMW12, FF16, LMR^{+}10, LHB^{+}15, MA13a, MLS^{+}14, SOJ14, SCB^{+}14, TRJ^{+}11, VSG^{+}11, VMN^{+}12, YPK15]. indices [HM09, PL10, SHB^{+}13b, ZHTA13]. indigenous [CGOC11, GAO14, LD17b, MDLM17, MUC12, PP18, VM10]. Indirect [HKS^{+}12, PWW12, CV11, FRC13a, LGJH12, LBG17, MHF^{+}19, MGL^{+}16, NOB12, PMS10a, Tut17, YRL^{+}10]. indirectly [EF16, OSL18]. Individual [BSN^{+}18, CC19, CEC^{+}18, DJJ^{+}16, EOMS15, FAWBF10, JW11, RRB13, SG18a, WWB14, ARB^{+}12, Ano17c, ACG^{+}11, BPWF14, BCV^{+}18, CCB^{+}17, CBLM^{+}17, CPA^{+}11, CMAJ12, DSG15, DED^{+}15, FBPCC^{+}19, FCS15, HAM^{+}15, JCRC10, KHK^{+}17, LSJ^{+}15, LBWB^{+}13, MSG^{+}13, NGR^{+}17, OWW^{+}17, OCDD16, OBÁSV18, PFdP^{+}12, PLGSD17, RWM18, RDGC17, RDF^{+}13, RSP^{+}16, SBB18, SCW^{+}18, dJTH^{+}14]. Individual-based [WWB14, ARB^{+}12, DSG15, MSG^{+}13, RWM18]. individual-level [SBB18]. Individuality [MHR^{+}19b]. individually [OWH^{+}19]. individuals [Ano17c, LMLB12, SWK19, WRW^{+}16b]. Indo [BHN^{+}14, CM10, CGM^{+}13, ETS^{+}11, GAC11, HDP^{+}17, HLN^{+}19, KBC^{+}12, RRL^{+}19, RBSL11, RDF^{+}19, WMB^{+}14, WA14]. Indo-Pacific [ETS^{+}11]. Indonesia [BS15, BCdV13, CPR^{+}14, DBE^{+}12]. induce [BDT^{+}16, MDPG13, PUP16, VPW^{+}12]. Induced [DH16, BCGWS19, BPB18, BP12b, CR14, EHG16, FHL^{+}17, GYRJ15, HP13, HA18, HSZR10, HRM^{+}11, JL17, LHS^{+}14, LBJ^{+}16, LLRC^{+}12, LPC11b, LAALH10, LMT15, LIT13, MMA18, PSL13, RWM18, RPRN^{+}16, RPS12, SLS11, SES17, SESS^{+}18, SA17, TvdMM^{+}17, VE18, VUL^{+}11, WMS^{+}11, WRO^{+}11]. induces [OMH12, SSM15, VTZ^{+}13, ZGH^{+}17]. inducible [PW19b, SG18a, TdRH17]. induction [SLL12]. industrial [TRGF18]. inevitable [BB14a]. infauna [BO13, KPB11, SCTA10]. infaunal [AAIB15, FAWBF10, MH11a, SBCA15, TCA^{+}11]. infect [BTLF17, WA14]. infected [GAF^{+}16, PWCvO^{+}14]. infecting [EKR^{+}10]. infection [ADRA19, BMR17, MP15, OD15, POS14, PRP^{+}16, RRG^{+}11, SBH^{+}18, SWPK13, WGD^{+}16]. infections [LK15, SDT15]. infectivity [LGR^{+}10]. inference [JCF^{+}12, MPB13]. inferences [IME12]. inferred [DCHI12, DAB17, FTKK14, GWH^{+}18, GCS^{+}19, KKS^{+}14, MSF18, MGAV15, RCR^{+}13, SW18, SWY^{+}10, STP13, SST19, ZRR^{+}13]. Inferring [CFMFFJ18, GWK^{+}19, HAM^{+}15, DBHMG12a]. Infestation [GKE^{+}19, HGG^{+}18, MH11b]. infested [DdR10, SST^{+}18]. inflate [RPS12]. inflated [SAMW14]. inflorescences [RCB12]. inflow [RLH^{+}13]. inflows [BSB14]. Influence [AAGGFG^{+}12, BRK^{+}12, BCW12, BFD^{+}16, BBI16, BMV19, BOÁNF19, BKC14, BMF^{+}15, CLS19, CRJ^{+}18, CO16, CD16b, DLdM19, DGMB10, DP11, DCW^{+}10, ET11, FMH^{+}19, HHS10, HBBH12, HCW^{+}15, JHNL12, KMW17, KFB15, LBM^{+}13, LÁBFÁF^{+}10, MSM^{+}12, MJV^{+}11, NGJHB12, PK15, PLFR^{+}11, PWO^{+}16, QHG^{+}19, QLCUG15, RLN^{+}13, dGCF^{+}17, SY10, SUB^{+}13, SBA12, SLE^{+}13, VRMKH^{+}15, YSH^{+}16, AOPMOS10, ABF^{+}14, AN11a, AS14, ABA16, BBM^{+}14, BHP12, BBF^{+}19, BB17, BLS^{+}11, BSB14, BD12, BAL^{+}11, BPB11, BWRP14, BTM^{+}18, CGS^{+}19, CAC11, CM11, Coo11, CPHH15, CFBG14, DQM^{+}19, DHAV13, FPR18, FHS^{+}13, Fuj16, GP12, GSL17, GDH^{+}19, HPA17b, HPH^{+}15, HCDR^{+}15, HHH^{+}11c, HYL^{+}13, HBS10, HHWW11, JGWJ^{+}11, KHBL11, KMP11, LRB^{+}18, LBN11, LST11, LS10a, LABG^{+}12, LMG12, LKN^{+}14, MBG^{+}16, MM10, MDM18, MHVV13, MLB^{+}14]. influence [MRH11, MH11c, NC14, OCSW16, OAB^{+}19, PMSGBL16, PMYH13, RGEA^{+}15, RPE^{+}13, RFRE19, REG^{+}11, RH12, RNB^{+}16, RCGE12, RMBM14, RBEGG13, SWMK15, SNS^{+}13, SCL^{+}19, SHB^{+}10a, SNK^{+}10, SDAH12, SH14, TCJ^{+}19, ULCAZG17, UW18, WKvW10, WL10, WWG12, WH14, WPK10, ZVL^{+}16]. influenced [ADA^{+}19, BMW^{+}14, CL15a, DBTV13, LTN19, SSS19, TNM^{+}11a, VWW^{+}16, VHKT15]. Influences [Bur12, FHB^{+}12, PDGL17, SCWV12, SQRDLC18, SSM^{+}16b, BJF17, BDDvdM17, BLJP14, CCM^{+}15, CGKW10, CGH11, CD14, DF15, EBH17, FJC10, FYG^{+}18, FSW10, GGF^{+}12b, HS17, HHH^{+}14, Hoe10, LLMA^{+}10, LGC10, MPH^{+}11, MGP11, MUHC11, MAB^{+}12a, MPB10, MOS10, OCDD16, POR^{+}15, PB18, PPSR15, PES12, PKBS14, RR18, RPM^{+}13b, RSTM14, SGCF^{+}12, SNW^{+}18, SLT^{+}19, SLO^{+}17, SW12, STM^{+}14, YRM^{+}12]. influencing [CAB^{+}19b, SHK14, SHAS15, SGTP16]. inform [BHUCMO19, BRB11b, HFG^{+}19]. information [HB12, JBL^{+}16, XWE^{+}14]. Informing [ABH18]. informs [LCR^{+}16, SCB15]. Infrared [LS14]. ingestion [CSR^{+}14, CXdG^{+}17, CD13, DA11, LBCL10, Wir14]. ingredients [AS12]. ingress [RWR^{+}15]. inhabited [DdRH14, JLQ^{+}14]. inhabiting [ALSO^{+}18, Ano18a, GCR^{+}17, SIHH10]. Inherent [AC11, SGT^{+}14]. inheritance [BRH13, RGCT^{+}10]. inhibit [GGMTGC18, LH19]. Inhibited [RUP11a, SP13a, RUP11b]. Inhibition [WSZS16, BPW^{+}18, OSE13]. inhibitory [ASM10]. Initial [DWD^{+}18, EM13, KWM12, CLJF12, HFMD16]. initiation [JHNL12, Ji11, MGG^{+}14, TPC^{+}13]. initiator [EH13]. injured [MB11]. injuries [MFDW17]. Inland [TCKY19]. inlet [SDKD13, SD17b]. innovation [HW12c, HW12b, WV12]. Innovative [HLH18, Ano18b]. Inorganic [TMP^{+}13, BCK15, FF12, FC12, GRCP^{+}10, MGFÁS^{+}10, ROE17, SDMR19, SSX^{+}17, WBLS^{+}11, WADKT13]. input [FK10, GBTS11, GLY^{+}14, HBE^{+}13, RWH14]. inputs [BMF^{+}12, THRBL^{+}16]. insects [CAD^{+}10]. insemination [HPMB11]. inshore [AAGGFG^{+}12, CRAP^{+}12, CAC^{+}12, CTSH12, DDP19, IMD11, JBS12, LHT^{+}16, LSH^{+}19b, MEKC15, MOP^{+}15, PKBS14, SHUS17, SJ10, SNT^{+}16]. inshore-offshore [MEKC15]. inshore/offshore [DDP19]. insight [AGP^{+}13, BCM^{+}14, BDS^{+}18b, GKPT16, GVGP^{+}14, MVL^{+}16, SHMOS17]. insights [ASS^{+}13, Ano17c, BDG^{+}16, BW10, BMPC15, BQ16, CTZ^{+}16, DKL^{+}18, DBC^{+}12a, GCSF18, GCC^{+}14, HGG^{+}10, HVBG^{+}19, HVD^{+}16, HEAI10, KWC^{+}17, KBP^{+}10, LLMA^{+}10, LAL^{+}18, LMSC^{+}19, MLP^{+}14, MCW11, OT13, PDvG16, PHHL^{+}11, RAKB^{+}12, She12, SJB16b, SAM14, SRA^{+}17, TFFPM18, TRSB^{+}18, TFW^{+}12, Ued14, WH11]. instantaneous [WP10]. instrument [KW17, SBA12]. insular [FPC^{+}13, LSFQ19, PA13]. intact [JBD10]. intake [AZVB13, VHBG15]. intakes [PWNP^{+}11]. Integrated [BSF17, SSFB11, ARvdW^{+}16, BMTC^{+}15, FBC^{+}15, KHT^{+}17, MBCB^{+}19, RBD^{+}14, SMJP^{+}15]. Integrating [ASCHM19, Dav19, LJM^{+}12, MSG^{+}13, Pör12, SDST16, ACP^{+}16, DP16, DVB^{+}16, DC17, OMW^{+}10]. integration [SBO^{+}18, TVRG^{+}13]. Intense [OGA13]. intensification [vdMPRM12]. intensify [DGA^{+}15]. intensity [BMVBC10, CMR12, DBTV13, KCR^{+}10, MLB18, MRL^{+}10, NPMS11, OLF11, PCA^{+}15, PBW^{+}11, SSHB16, WK16, ZS17]. intensive [TPC^{+}13, TPC^{+}15, ZGK^{+}19]. Inter [CdTK^{+}16, CSC^{+}11, CRRL17, CP11, KBW^{+}11, KTP^{+}18, RKM^{+}10, TCTB11, ABA16, AJ16, ACG^{+}11, CCB^{+}17, CML^{+}12, DAC^{+}12, DRH15, FTB14, JJP^{+}13, JSMO14, LL18, PBAP^{+}14, RMNP17, SMD^{+}13, SMF^{+}17, WCG^{+}14, YTO^{+}11]. inter- [JJP^{+}13]. Inter-annual [CRRL17, KTP^{+}18, RKM^{+}10, AJ16, ACG^{+}11, DRH15, FTB14, WCG^{+}14]. Inter-colony [KBW^{+}11, ABA16, DAC^{+}12, YTO^{+}11]. inter-decadal [SMD^{+}13]. inter-individual [ACG^{+}11, CCB^{+}17]. Inter-nesting [CdTK^{+}16, RMNP17]. Inter-population [TCTB11]. inter-reef [LL18]. inter-simple [CML^{+}12]. Inter-specific [CSC^{+}11, CP11, JSMO14, PBAP^{+}14]. inter-tissue [SMF^{+}17]. interact [PBH^{+}17]. Interacting [BWH^{+}17, FTF15, PFR11]. Interaction [GMBL11, MLB19, VBJR^{+}13, Ano18b, BLD^{+}16a, CB14, CLPiU14, ESBG14, LZ12, MMG^{+}15, OLM^{+}12, PCA^{+}15, PBIO^{+}18, SKA^{+}14, YRL^{+}10, dSM11b]. Interactions [BBBW17, CLB19, GL16, HTHH11, LTCH10, OKLM14, SLKN14, AS15, ACB12, ALNB17, BHE^{+}17, BMJ16, BAW11, BVQ12, BMSS14, BC13, BB19, BHTF11, CFF^{+}18, CBMK10, CSRF16, CAB^{+}14, CHB11, CLPU14, CGM^{+}19, CAGS17, CSS^{+}13, CFBG14, DGA^{+}15, DKAC^{+}12, ERB^{+}13b, ERB^{+}13a, ECC^{+}15, FRGI15, FCS15, GL12, GHHH12, GMB^{+}16, GBJ^{+}14, GGNL^{+}17, Gla10, GLWI10, GLB^{+}10, GCP^{+}17, GGD^{+}11, HPH^{+}15, HSM13, HW13, HM13, HOE^{+}12, KW10, LH11, LCFR10, LBG17, LHB19, LÁBFÁF^{+}10, MB16, MH19, MMCL19, MGL^{+}16, Nor13, ORM^{+}13, PKW^{+}11, PPB^{+}17, PES12, RZH16a, RSW13, SGS^{+}17, SP13b, SG18a, SPH^{+}13, SF14, SNNS^{+}12, THM^{+}13, TR13, TFM^{+}18, TGR^{+}19, TSTP13c, UDF^{+}18, VGGF12, YHR16]. Interactive [CFH11, EKE16, EVB^{+}16, FA10a, MP15, SSX^{+}17, SG18b, VTB12, VPOA12, ZXF^{+}16]. interactively [HBS10, HNG^{+}11]. Interannual [DPO15, HEMJ18, JNYT13, MMWG13a, MJHS13, OGCHS18, PW15, SJS^{+}11, SDVL16, YFC^{+}13, BCSB13, BHN^{+}12, CFPF^{+}14, DHB18, HCJ^{+}11, JPAM^{+}15, KFGSMPG14, MSK^{+}19, NLGE18, RQN^{+}14, TBDM11, Gal14b]. intercept [ZB16]. interception [BVGH10]. Interdecadal [Gla10]. interface [BH16a, BMV19, CLSS17, WAA^{+}19]. interfaces [AWJ12]. interference [MOMF14, PP18, RSE18, YAFV12]. Intergenerational [BCSH18]. Interior [GCC^{+}10]. interisland [KPW^{+}13]. intermedia [APH^{+}10]. intermediate [CLL13, KGLLW11, POL^{+}11, TSGW10, TPF^{+}13, dML11]. intermediate-water [KGLLW11]. intermedius [ZVL^{+}16, ZVR18]. intermittent [PLFR^{+}11, SLL12]. intermittent-upwelling [PLFR^{+}11]. intermonsoon [BMD^{+}11]. Internal [KKA12, LBAK13, ACF^{+}10, FHL^{+}17, KdEPW18, MSGGE10, RS14, RFJ^{+}10, SMM^{+}14, VGL16]. international [JP15]. Interplay [OSPWB14]. interpopulation [GGPDS19]. interpretation [BH12a, FH18, GHNF10]. Interpreting [PGWS10, KM17]. interrupts [HSC^{+}15, PVYS16]. interruptus [LH10, Mat10a]. Interspecific [MM13c, TOW10, AN11a, CLDC^{+}13, FB14, GCR^{+}17, KCCG12, Kin16, LBP^{+}16, VFMdW12, WHL11]. Intertidal [CHE^{+}17, KCE14, MLSK^{+}15, ULCAZG17, AN11a, ADO^{+}12, AAIB15, AAJJ^{+}19, ACS12, BG19, BBF^{+}19, BGJ^{+}19, BDLW10, BSH13, BBC^{+}12a, BH12a, BMF^{+}12, BC14, BNW14, CGOC11, CGK13b, CMC10a, CF12, CY13, CAC11, CMW^{+}15, CGEGC^{+}19, CTvdM^{+}08, CC12, COK10, CNSW16, CSQRA11, DMK11, DPO^{+}17, DvV^{+}16, DP13, DPM15, DHGW15, DB11b, DBC^{+}12b, DCW^{+}14, EMMSL13, EN11, FAT^{+}15, Fis10, FC12, FPR18, FC19, FSH^{+}16, FTF15, FB14, GRCP^{+}10, Gar13, GPV17, GHCC15, GME^{+}19, GGF^{+}14, GG12, GJ10, GLB^{+}10, GC13b, GKPT16, GPB^{+}15, GPS13, HPA17a, HPA17b, HMP15, HKS^{+}12, HIKK12, HI14, HR19, ICUM10, JAPK16, JC15, JDG^{+}17, KF13a, KCR^{+}10, KRD^{+}12, LLBT18, LAM11, LAM13, LHL11, LCFR10, LCL^{+}17, MMG^{+}15, MJR^{+}14, MAS10a, MCUO12, MH11a, MN18, MSH^{+}14, MSG11, MCR^{+}13, MLB18, MBCB^{+}19]. intertidal [MAK^{+}15, MF13, MPBSB18, MCRB11, NLV^{+}15, NSI^{+}15, OMN17, OSPWB15, OLM^{+}12, PDGR^{+}13, PBTB15, PDC^{+}17, PCF^{+}19, PLS^{+}10, PJDC^{+}13, PBW^{+}11, PJMP10, PB10, PÁIIS^{+}19, RPRN^{+}16, RMC^{+}14, RFBH^{+}11, RCGE12, Rit17, RK15, SLSC16, SKB15, SA14, SSPKF16, SDM16, SHC^{+}10, SHB^{+}10a, SLC11, STP10, SP12, TMS^{+}16, TS10, TJS17, TBS^{+}17, VBT14, VUL^{+}11, VBR18, WDH^{+}12, WSHY16, WKE^{+}17, WK10, WS13, WHvdK11, WR12, WK16, ZPLW11, ZFZ16, dJH11, vdGSE^{+}14, vdGvdLvG^{+}19, vdVFK^{+}10, vdVKA^{+}11, vdMPM12]. intestinalis [MSK^{+}19, PCNB16]. intimidators [GG12]. Intra [CCB^{+}17, GCR^{+}17, RDF^{+}13, SVAV17, CCNB15, GGF^{+}12b, JJP^{+}13, JHL^{+}19, LFJ^{+}14, RZH16a, VFMdW12, MCPC19]. Intra- [CCB^{+}17, GCR^{+}17, VFMdW12]. Intra-individual [RDF^{+}13]. Intra-population [SVAV17, CCNB15, JHL^{+}19, MCPC19]. intra-school [LFJ^{+}14]. intra-specific [GGF^{+}12b, JJP^{+}13, RZH16a]. intracellular [GBU15, HKS^{+}12]. intraguild [GSP^{+}15, HW13]. Intrapopulation [AB11, HPL10, HMR^{+}16]. Intraspecific [BPW^{+}18, DS16, KP11, MVP^{+}16, MCBR18, SPBS14, TGH15, YAFV12, dMA18, Col10b, SL11a, SGS^{+}19, YHR16]. Intrinsic [MCW^{+}16, GDB^{+}16, HT10, HCW^{+}15, JMB10]. introduced [ATT^{+}16, CB11, GC13b, HKT10, ORM^{+}13, PPTB12, PP18, SIK16, TdRH17, WDH^{+}12, dGR11]. Introduction [KBH11a, STK12, KBK^{+}10, PH16, KH18]. introgression [BH10, SHH10, TSR17]. intrusion [HLS^{+}15b]. Inundation [BVC^{+}16]. invade [ATT^{+}16]. invaded [CBK^{+}12, SO15]. invader [CGOC11, KNB^{+}13, OHP18, PRA^{+}19, RPHB15]. invaders [BFG^{+}19, HCG15, TBS^{+}14, ZBK^{+}14]. invading [GB12, SSS19]. Invasion [HWR^{+}13, HGA^{+}16b, HGA^{+}16c, KSB11, TSCSB15, BVO16, CMH14, DBCV^{+}12, DM17, DMJ11, GP11a, GBL^{+}15, HW11a, HGA^{+}16a, IW15, JP15, KKL13, KBC^{+}12, LFMC11, LFRF13, MV12, MMG^{+}16, RCF17, RBS16, TS12, VPOA12, WHB^{+}19]. Invasion-mediated [KSB11]. invasions [Bri12, CKC^{+}10, PRCG^{+}17]. Invasive [AL12a, Alb15, BVM^{+}19, Ben16, DS10, KBS16, MP12, MP17, PHSBO16, RTJ^{+}13, ACL^{+}16, BWM^{+}15, BAA12, BMA^{+}10, BSRDR10, BFH^{+}13a, BBdG^{+}19, BH10, BGK17, BOU13, BGBD^{+}13, CCT12, CGM^{+}13, CWAS17, CSN12, DEP19, Dav19, DFLC17, DHG11, EPM^{+}16, EF16, FAM^{+}17, FPN17, GFL^{+}17, GHGF12, GFC^{+}14, GS18, GSTK11, GCR16, GVGH15, HTSA16, HH15, HWTH12, JT10b, JLQ^{+}14, JMB^{+}18, KMV18, KNIN17, KHG^{+}19, KA16, KS12b, LA12a, LMTD16, LSV10b, MEK11, MS13, MTK^{+}18, MCW11, NLLS15, NSI^{+}15, OHP18, QCX^{+}10, RAB^{+}15, SPSG14, SZ13, SLS18, SLKN14, SJHH^{+}17, TBKG10, TWM^{+}15, Tut17, VBD^{+}11, VBR19, VDEO18, WO11, WNA^{+}15, WPB^{+}14, YPK15, ZBP18, ZBK^{+}14, dMA18]. invasive-species-mediated [ZBK^{+}14]. inventory [HFG^{+}19]. inverse [SDM^{+}13]. Invertebrate [KW12, ALSV16, BHT15, BSCH14, BM11, CHE^{+}17, CMAJ12, ES12a, EVHA^{+}18, GTP^{+}17, GOG14, GMSC14, GGPDS19, HGKB14, JDB^{+}10, LWM15, MRC10a, MHNL14, MA13b, Pil10, RNB^{+}16, RBS16, RBD^{+}16, SDM16, SLT^{+}19, SCWV12, SCW^{+}15b, TSCSB15, TMM15, WSZS16, WRDF12]. invertebrates [Ale13a, BMV19, BMVBC10, BOÁNF19, DMd14, DBH^{+}16, EMMSL13, GKC11, GLM16, HSdBS10, JH11, Joh11, KO12, KTH^{+}18, KEH10, LM12, LBT^{+}10, LMd12b, MMB^{+}12b, PIPD10, PT13, RL10, RSG^{+}16, SVG16, WJ12, dG14]. investigate [CBR19, KMGH16, ZMRCP^{+}17]. investigated [BOS16, DH16]. Investigating [CMV^{+}11, CPSC^{+}15]. investigation [CCA^{+}17, HSH18, ICUM10, LME^{+}19, LRO^{+}17]. investigations [CF12]. investigators [DBS^{+}16]. investment [BER17, CRR^{+}19, DED^{+}15, JJF14, MN18, OT13, WGGM11, YKT^{+}14, YBD^{+}16]. investments [PRFA11]. Invisible [LCL^{+}17]. involved [FMC^{+}11]. involving [MGB^{+}19]. inwelling [WAMS^{+}14]. Ionian [CBR^{+}18]. IP [BBG^{+}16]. Iran [SSPKF16]. Ircinia [MPLLE17]. Ireland [OLD^{+}15, TLM^{+}11]. Irion [AvdHH^{+}17]. Irish [CMAJ12, MMRC16, PPSR15]. Irminger [MAB^{+}10]. Iron [LHT^{+}17, TBH^{+}17, AvDL^{+}19, LHS^{+}14, ODL^{+}13, SMK^{+}18, SLIGG18, SP15, SPSP10, WJS^{+}14, ZXF^{+}16]. iron-induced [LHS^{+}14]. Irradiance [CPTL11, FBET15, GMC^{+}12, LN11, LWS^{+}10, MTMM^{+}15, PKB16, RCP17]. irradians [HW11b, WMMC14]. irradiation [BSD^{+}14, PCNB16]. irrigation [DWM15, WAA^{+}19]. irrorata [KM10b, ZWP^{+}17]. irroratus [GPM^{+}13]. Isefjord [PAF^{+}10]. Ishigaki [SAH11, HHI^{+}14]. Island [BLMD^{+}11, BKB11, BCH^{+}11, CAB^{+}19b, CRS^{+}18, DAW^{+}11, DRH15, GHES18, GMSC14, HPB^{+}18, HVJ^{+}18, HHI^{+}14, HTR19a, KKY15, LBJ^{+}16, MOH^{+}10, RGL^{+}16, SSPKF16, SLR^{+}14, TG10, TBH10, TLG14, WHPT12, WFBM14, ATT^{+}16, HP11b, KSBM^{+}11, MSFM12, MHPvdH19, RRH^{+}16, SFL^{+}18, BBC^{+}13, CW12, DDD^{+}11, LRO^{+}17, MBB^{+}18, PDP^{+}10, SHB^{+}19, SLS16, TS16, WVD^{+}17]. islandica [BRS10, PPML17]. Islands [GMSL^{+}14, KCMH^{+}17, PSD^{+}19, RZPH14, SPJ^{+}12, SBA12, SRK13, WGD10, KAMA13, MYS^{+}16, QHG^{+}19, WLF^{+}17, ZSEV17, BWV^{+}12, BMEL11, BSD^{+}14, BDS^{+}18b, BFN^{+}19, CXdG^{+}17, DH18b, Edm13, Edm14, EPL^{+}14, EL16, FJD13, HM13, KBG12, KBK^{+}10, LFPA^{+}11, LRBS15, MDKM13, áNGGS11, ODP^{+}12, OLF11, OAB^{+}19, PIPD10, WSK^{+}16, WRI12]. Isle [MHK11]. Islotes [TEAO^{+}14]. Isochrysis [BBRBG10]. isolated [CBL^{+}10c, MSFM12, OCS^{+}13, PBM^{+}18, RRH^{+}16, WMS^{+}11]. isolates [TG10, TFÁS^{+}12]. Isolation [SCO^{+}12a, AFF^{+}15, FRC^{+}13b, LGBSC11, PFPP18]. isopod [GED^{+}14, MAD17, MHPvdH19]. Isopoda [HVT12]. isoscapes [BKM^{+}18, KM17, VLM14]. Isotope [DRWQ12, ALV12, ABN14, AMS^{+}13, ALE^{+}13b, ASWB10, AR14, AEM^{+}18, AGRPSVL13, AGP^{+}13, BLV^{+}19, BCM^{+}14, BSC^{+}15, BPJ^{+}15, BADQ^{+}15, BBC^{+}17, BKC^{+}19, Bre14, BPHW16, CLD^{+}13, CKM^{+}16, CCA^{+}17, CE11, CKK^{+}10, CMW^{+}15, CMB^{+}14, CS13, CPS^{+}14, DEJN10, DFS13, DBC^{+}12a, DCIG18, DPH14, DBS^{+}19, EAMT10, EMH^{+}19, FC10, FPC^{+}13, FTKK14, FLC^{+}14, Fry13a, Fry13b, GSR10, GKHM13, GCSF18, GWH^{+}18, GCS^{+}18, GCS^{+}19, GSSM^{+}11, GPC^{+}15, GHNF10, HHS^{+}12a, HWET13, HN13, HKB^{+}19b, HB12, JBS11, JS12b, KSK07, KHF^{+}11, KKS^{+}14, KRC^{+}11, KM17, LLBT18, LCGS^{+}19, LMV^{+}19, LCB^{+}16, LRO^{+}17, MWA17, MBSG11, MFH^{+}11, MWZ^{+}16, MGAV15, MBS^{+}13, MFS14, NW15, NUO^{+}19, OFDG14, OBI^{+}16, PLG^{+}19, PBAS^{+}10, PKS11a, PvdMLC17, PBCA10, RMHP14, RP12, RGDNGS11, RMT^{+}13, SDR^{+}15, SWP^{+}13, SWY^{+}10, SMF^{+}17, SST19, SRA^{+}17, SCW^{+}15b, SGD^{+}12]. isotope [TTT^{+}15, TFM^{+}18, TMP^{+}13, TMS14, UTS^{+}18, VSG^{+}11, VLM14, VMC^{+}19, WBI^{+}15, WRW^{+}16a, WLF^{+}17, YAFV12, YMD^{+}10, ZBB^{+}11, ZO10, vdBUPP^{+}11]. isotope-based [GWH^{+}18]. Isotopes [JGM^{+}16, ACG12, ACG^{+}11, BMJH12, BMV^{+}11, BDR19, BGS^{+}16, CMIMGS15, CJM^{+}13, CAG13, CCKB^{+}15, CRHV14, DWPH11, DSG^{+}13, FLAL10, FHMG19, HAM^{+}15, HVJ^{+}18, HFG^{+}19, HRT^{+}14, HS15b, HEAI10, JCRC10, JJP^{+}13, dDTC^{+}17, LMR^{+}10, MWP10, MMFF12, MBRO10, MCL^{+}17, MCW11, NFC^{+}11, NLK^{+}19, OHR^{+}13, PLNE11, RKM^{+}10, VRSS18, VW10, ZRR^{+}13, ZMRCP^{+}17]. Isotopic [RSW^{+}16, RNFG10, SS13, TKEV^{+}19, TGI14, BLPP16, CXdG^{+}17, FK10, FCH^{+}10, HTR^{+}19b, LFM^{+}16, MF15, MJV^{+}11, MPRPM^{+}18, NMPM15, PLF^{+}12a, RVPG17, RSTM14, SVAV17, SGS^{+}19, SKMS^{+}18, UDF^{+}18, WK18, YSC^{+}15, ZMGS17, THB^{+}12]. Israel [BPP^{+}18]. ISSR [CML^{+}12]. Istiblennius [HLN^{+}19]. Istiompax [FHTY19]. Isurus [NCW^{+}19, PBB^{+}15, TKEV^{+}19, VWW^{+}16]. items [BAL^{+}11, CXdG^{+}17, PWNP^{+}11]. Ixodes [MBB^{+}12a].

Jakarta [CPR^{+}14]. Jamaican [IHP10]. Japan [HHI^{+}14, HTR19a, KKS^{+}14, KHT^{+}17, KTK^{+}12, MKK^{+}11, MOK^{+}18, NKM^{+}15, NKS^{+}15, NOT^{+}11, SAH11, TNIH12, WYS^{+}10, WI12]. Japanese [CKK^{+}10, HHLT10, HHNI17, HIKK12, HI14, MAW^{+}11, Mat10b, MKI^{+}15, MOK^{+}18, MKII19, NYK^{+}13, PVYS16, Sat17, TWY^{+}19, TUK13, YHT16, YKT^{+}14]. japonica [CKK^{+}10, HHLT10, HH15, TWY^{+}19, ZBK^{+}15]. japonicus [HPL^{+}16, NCD^{+}12, ST10a, SGD^{+}12, WB12, YYO14]. Jasus [CB17, JW10, JWHR11]. jawfish [HPL12]. Jellyfish [CLPU14, CLPiU14, DRD^{+}18, FPC^{+}18, HMH^{+}18, KTKM14, ADD18, ANWG18, AMAM^{+}18, CL17, CGQB10, CFBG14, DGCH18, DLC^{+}13, FHG^{+}14, FSM^{+}14, GJSR18, GKR^{+}15, HTT^{+}12, HSH18, HSS18, JWR^{+}10, LFE^{+}14, MAHK12, NMPM15, Pur18, SHP^{+}18, SWK18, SNT^{+}18, TFM^{+}18, TPKC18, UAK^{+}14]. jet [RJL^{+}11]. jets [AL12a]. Jiaozhou [FWS^{+}18]. John [Edm13, EPL^{+}14, EL16]. Johnston [RL11]. Joint [CFA16]. Jorge [GZYS11]. journey [WRR10]. jubatus [SDR^{+}15]. julis [BMSS14]. jumbo [BBG12, GRCH^{+}18, LCC^{+}11, RCVG10, TKA10]. jumping [DHSL11]. Jusufovski [Ano21]. Juvenile [ATO^{+}14, ETHB17, RLM^{+}14, ASM^{+}12, AM13, ACG12, ARA^{+}10, AGH^{+}12, AGP^{+}13, BM10a, BDM^{+}13, BBMH13, BBdG^{+}19, BT13, BMPH14, BCH^{+}18, BHL^{+}19, BCWSR14, CMR12, CR14, CC19, CWS13, CLH^{+}15, CPPM12, CWK^{+}10, CWW^{+}16a, CMK12, CSM^{+}19, CPF^{+}16, CTB13, CAG12, CLSS17, CRS^{+}18, CGG^{+}10b, DSC13, DABP13, DR16, DTG16, ETO^{+}10, FFS14, FRRÁSL11, FRRÁS^{+}12, FTB14, FDP^{+}16, FWL15, FCRD18, FTB^{+}17, GBD^{+}12, GSS13, GMHS19, GSF^{+}13, GBR11, GCSF18, Gla10, GFP19, GRCH^{+}18, GHT15b, GGD^{+}11, GGF^{+}12b, HAM^{+}15, HVJ^{+}18, HBC15, HTB^{+}15, HM10, HRS^{+}16, HMF^{+}18, HIMH15, HHB^{+}14, JG13, KSFL14, KHH18, KNS^{+}15, KF17, KCS12, KHS^{+}11, KP11, KRD^{+}12, LOS^{+}19, LABM17, LTG17, LCC12, LNP12, LMSC^{+}19, LMT15, LRO^{+}17, MGP12, MVCD14, MMRC16, MH19, MCC^{+}16, MWC^{+}12, MGM10, MBV^{+}11, MGCH^{+}16, MR16]. juvenile [MCT15, NVA^{+}19, NFC^{+}11, NHG10, OTK^{+}19, PDT^{+}11, PLA^{+}12, PWNP^{+}11, PES12, PMH^{+}13, PRP^{+}18, PW19b, RLN^{+}13, RSO^{+}13, RJR18, RRG^{+}11, RP17, RBH12, SVG16, SKD^{+}17b, SNVH19, SS11a, SGM^{+}14, SWW11, SHB^{+}13b, SCHO11a, SCHO11b, SCF^{+}17, SDH13, TSGW10, TPA^{+}13, TTW^{+}12, VPDPMRC11, VMN^{+}12, VCFF13, WAW^{+}12, WHFM17, WDM^{+}18, WS11, WWW^{+}13, WWG^{+}13b, YHT16, ZBC10, vdWTWS15, vdVKA^{+}11]. juvenile-adult [CWS13]. juveniles [BS19b, GGMGR12, GNBM11, HTK^{+}17, JNYT13, LI18, MJO^{+}13, NAY^{+}11].

Kāne’ohe [CRWG16]. Kaldbaksfjørdur [áNGGS11]. Kaloko [WBB^{+}10]. Kandelia [CY10, HSS^{+}10]. Kangaroo [WFBM14]. Karenia [LWD13, SBS^{+}13, WHT12, WLBB^{+}13]. karianus [YMMO14]. Karlodinium [BH16b]. Kasatochi [ODL^{+}13]. Katsuwonus [RTCF^{+}17]. Kattegat [KRDV18, LE13, POH16, SGJ^{+}18]. Kawabata [Ano15a]. Keeling [BMEL11]. keeping [TWW^{+}15]. Kelp [SSM15, SMS15, BVG13, BRG^{+}14, BBTW14, BKR^{+}15, BTLF17, BOS18, BNW14, CSKR10, CSRD11, DSJNH18, DA10, DBE^{+}18, ESW14, FNC13, FSLG12, FS14, FFS14, FDS12, FDS14, FDFS16, FJW14, FK10, Fox16, FWB^{+}15, HLD12, JW11, KKY15, KMP15, KS13, KS11, KS12b, KS12c, LGS10, LRL^{+}13, MGFW13, MJW19, MZ10, MWW^{+}14, MA13b, NFSW10, NCR^{+}14, OSK15, OAD^{+}14, PLG^{+}19, PNF^{+}12, PNPF14, PMS10a, PMSGBL16, RWR^{+}19, RRS15, RS16, SBH^{+}19, SLS18, SLT^{+}19, SBM^{+}16, SBE^{+}16, SWW11, SH14, SFNP11, SJHH^{+}17, TW16, TPVCV19, VB14, WC18, WHP^{+}17, YMPR13]. kelp-dwelling [NFSW10]. kelps [OSE13]. kelts [ABT^{+}15]. Kemp [SL11b]. Kenya [HMM14, KKS^{+}12b]. Kenyan [KMO19]. Keppel [Ano16]. Kerguelen [CCA^{+}17, CXdG^{+}17, GMES18, MBSG11, OHCG14, OGLH17, PDP^{+}10]. Kermadec [FKRJ13, WGD10]. Kerr [Ano10e]. keta [KSA^{+}13]. Key [FGW^{+}15, Ano17c, BHDC15, CSS15, CFH11, DPY^{+}16, DKAC^{+}12, FBS14, FSP^{+}17, GGK^{+}16, HC13, LLBT18, LCL^{+}17, LM18, NPLT10, OKLM14, OBI^{+}16, PBH^{+}13, RKvW10, SUPCB11, SKM^{+}15, SPBS14, ZBY^{+}14, vdBUPP^{+}11, vBAP^{+}19]. Keys [SQRDLC18, MEKC15, OC17, RCP^{+}13, SPW^{+}12, TYVR^{+}15, VRMKH^{+}15]. keystone [ARD^{+}19, BSS18, BMA^{+}19, CHRB10, FPBL16, FCS15, GG12, MJM^{+}14, TTS17, GG12]. KEYSTONEin [ZFZ16]. kill [SGS^{+}12]. killed [BLD^{+}16a]. killer [BLMD^{+}11, BDS^{+}18b, MHN^{+}12, NSK^{+}14, RGL^{+}16, RDF^{+}17, RSSvV19, SVAV17, TMT^{+}12, TGR^{+}19]. killifish [JMO^{+}19]. kilometres [VW10]. Kin [ALK17]. Kinematics [JVSD10]. Kinetics [HGG^{+}10, FUOG^{+}16, Ree14, SMHB14]. King [WVD^{+}17, BHUCMO19, BSA^{+}16, FPN17, HBF^{+}10, MP12, MP17, PES12, SBT12, SSHB16, SDH13]. kingsejongensis [HPL^{+}16]. Kiritimati [CW12]. Kittiwake [Hat13, BER17, CLC^{+}12b, CDMB^{+}18, GSSM^{+}11, HCDR^{+}15, SMC^{+}14, VM14, WKK18]. kittiwakes [GAB^{+}14, MATM^{+}19, MIA^{+}11, REAN^{+}14]. Kittlitz [AMM^{+}12]. km [BLM^{+}18]. knoll [SNW^{+}18]. knowledge [ABRP15, BW10, CCGP14, PMKL11, QBW18]. known [SJO16].

Kobbefjord [VCMS^{+}18]. Kodiak [CRS^{+}18]. Kong [GSM^{+}11b]. Kongsfjord [HB15a]. Kongsfjorden [SdBAH^{+}14]. Korea [KHG^{+}19]. kraussi [PWW12]. Krill [CDR12, LTSDHL12, RWM^{+}10, ASR^{+}19, ABN14, ARGB13, BBH^{+}13, CDCZ16, CDC18, DHK^{+}17, DSGR^{+}15, EZZ^{+}12, FAK^{+}12, HB15a, JR15, KTF10, KSK^{+}12, LRQ^{+}12, MKM^{+}16, MAS^{+}10b, PHD^{+}13, RCS^{+}17, RWS^{+}15, RSC14, SRLV10, SKD^{+}17b, SGCF^{+}12, SKR^{+}13, THP^{+}16, TAAN15, TTM19, Wil14]. krill-eating [ARGB13]. Krumme [She13]. Kulamanamana [Par15]. Kuroshio [lLChTC11, LGC10, NYK^{+}13]. KwaZulu [DFS13]. KwaZulu-Natal [DFS13].

L [BOKJ18]. L. [BNE^{+}19, CALB10, FG10, PP18]. labelled [GMT^{+}11]. labelling [MWA17]. Laboratory [GHT15b, CSR^{+}14, CLL13, CLSS17, LME^{+}19, LH16, LFE^{+}14, RMR12, STT14, TS10, UTS^{+}18]. Labrador [WRDF12, GDE17]. labrax [VMN^{+}12]. labrid [CFMFFJ18, PBM^{+}11]. labriform [JVSD10]. Labyrinthula [BMR17, BBR12, BRB17, OD15]. Lack [DM17, FNC13, TPC^{+}16, TSR17, WNA^{+}15, PHM11]. lactating [SBA12, dDTA^{+}18]. lactuca [BKC14]. Lacuna [YLPG19]. laevigatus [BFDM18]. laevis [LP12, MTS^{+}17]. lag [MPG^{+}12]. Lagenorhynchus [BNAK11]. lagged [ZHTA13]. Lagoon [JOFP18, KTK^{+}12, BRBH19, BMW^{+}14, CFC^{+}12, ETAE13, GCL^{+}16, HJN^{+}10, HSH12, HvEBC12, HP14, KMGH16, LdU^{+}17, MMB^{+}12a, PDM12, RCGE12, TCN^{+}10, WRGP12, SKFY18]. lagoonal [DP11, GNBM11]. lagoons [CED^{+}19, CLGC16, CMC^{+}15, MRGG^{+}13]. Lagrangian [BDS18a, KBR^{+}11, ZSSK11]. lake [CLL15, CGC15, PP13, EKR^{+}10]. lakes [BCdV13, FGOS16]. Lakshadweep [KAMA13]. Laminaria [BRG^{+}14, DBS^{+}19, NCR^{+}14, PNF^{+}12, PNPF14]. Laminariaceae [LS16]. Laminariales [DBS^{+}19]. laminarin [SBH^{+}19]. lance [DHG16]. Land [AHB^{+}14, CWEE19, CHFS12, DJJ^{+}16, KTF10, MS11, ORE^{+}17, RPE^{+}13, TKV^{+}19]. land-locked [DJJ^{+}16]. land-use [CWEE19, CHFS12]. Landfast [KFBM19, MGG^{+}14]. landings [GOM^{+}19, MLS^{+}14, PLZL14, WYS^{+}10]. Landscape [ELA10, Edm14, HHH^{+}14, AMM^{+}12, BB14c, GBJ^{+}14, HGC15, JWT^{+}14, KW12, MSG11, OLF11, PKSN11, RDPR15, RJH14, RBD^{+}16, SLS11, SHH10, VDV^{+}15]. Landscape-scale [ELA10, Edm14, GBJ^{+}14]. landscapes [NGR^{+}17, TBCG^{+}12]. lane [WEA^{+}17]. Lanice [DvV^{+}16]. Lanka [SKD^{+}11a]. lapillum [NSS10]. lapillus [BVO16, DF15]. Large [BL10, CFPF^{+}14, DHB18, FSD^{+}14, FDFS16, LN10, MHFN10, PSL^{+}12, SMSE19, SMVC^{+}15, SWPK13, WSLM13, AGK13, AFF^{+}15, BSC17, BW12, BBR^{+}14, CS10b, DER^{+}19, EBT^{+}15, EMH^{+}19, Gre10, HVK^{+}13, HRT12, JJP^{+}13, JD18, JMK^{+}18, JCM13, KSA12, KHÁB^{+}17, KCW^{+}17, KCF^{+}14, LHJH^{+}10, LJ10, LSH^{+}19b, LCR^{+}16, LL18, MFSRR14, MHVV13, OSK15, OTK^{+}13, PDBG14, PTL16, PSS19, PGR^{+}17, QSvN^{+}11, RFJ^{+}10, SKK^{+}18, SHR^{+}11, SRR15, SGS^{+}19, SFL^{+}18, TNOY17, WRW^{+}16a, WHK^{+}11, WTD^{+}19]. large-bodied [HVK^{+}13, SGS^{+}19, WRW^{+}16a]. Large-scale [CFPF^{+}14, FSD^{+}14, FDFS16, PSL^{+}12, WSLM13, DER^{+}19, JD18, JMK^{+}18, KHÁB^{+}17, LCR^{+}16, MHVV13, OSK15, PGR^{+}17, SRR15, WTD^{+}19]. larger [Gea10, TMH^{+}18]. largest [BB14b, KHW^{+}19, KPW^{+}19, PFB^{+}17, RPE^{+}13, RAB^{+}10, TMM^{+}19]. largha [GWK^{+}19]. Larsen [FS18]. larseni [DSH14]. Larus [CPF^{+}14, TRSB^{+}18]. Larvae [HRT12, TH11, AGGAO^{+}18, ALSV16, BLD^{+}16b, BCSB13, BLM10, BPG^{+}11, BGBD^{+}13, BS19b, CSR^{+}14, CMF^{+}19, CLDC^{+}13, CGS^{+}19, CZB10, CL17, CL10c, CME^{+}12, DMd14, DVA^{+}18, DAD^{+}11, DZL16, EGH^{+}15, EMMSL13, EVHA^{+}18, FKZTT18, FJ11, FG16a, FRP^{+}17, GCC^{+}16, GFP19, GOG14, GCDH19, HEMJ18, HSC^{+}15, HHNI17, JNYT13, JGG^{+}15, KPW^{+}13, KCD13, LCQT^{+}13, LCGS^{+}19, LBD^{+}15, LWP^{+}13, LPI^{+}14, LSH^{+}15, LEB14b, LMd12c, ML11, MC11, MGO^{+}12, MKM^{+}16, MSG^{+}13, MWHM18, MKI^{+}15, MDAPH11, MD16b, MJHS11, OF10, OA10, OFFQGG10, OVF11, OMSP10, PAFC13, PS15b, PSB17, PTF^{+}14, QLCUG15, RTMJ11, RUPT^{+}11, RTÁB^{+}14, RBA^{+}18, ST10a, Sat10, SM10a, SPMN19, STK^{+}13, SMM^{+}15, SCSC16, SCWV12, SDAM11, SFTG14, SE11b, SM15, SFW^{+}16, SKFY18, TG12a, TAAN15, TYG12, UTS^{+}18, VGL16, VE18, WHA^{+}13, WR12, WJHS13]. larvae [WHL11, YSS^{+}18]. Larval [ARD^{+}19, CL16, DSS10, GWGC16, HBM^{+}12, KS13, LHU^{+}18, MF10, NM18, PAF^{+}10, RB11, RBGT10, WGLDB15, ARB^{+}12, BZC^{+}12, BdSC^{+}14, BCSH18, BOÁNF19, BS14b, BFEMLÁJ19, BPHW16, BM11, BVD^{+}14, BCR15, CDD^{+}16, CR19, CLDC^{+}13, CMK12, CG10, CB17, CG11, CBF^{+}13, CCG^{+}19, DSC13, DGDJ14, DT15, DCT^{+}10, DEM15, DWL19, ETS^{+}11, ES12a, FHL^{+}17, FPM14, FOJF15, FTF15, FZI^{+}16, GSÁCM^{+}16, GVR^{+}14, GTP^{+}17, GW15, GLW^{+}17, GGMTGC18, GPS13, HRP18, HLW^{+}16, HSM13, HS16, HSDB13, HBP^{+}16, HR15, HPH^{+}18, KP18, KSFL14, KF13a, KHH18, KHT^{+}17, KHN15, KP11, KCLP16, KML13, LAM13, LMC14, LEB14b, LABG^{+}12, LMd12b, LRQ^{+}12, MDLM17, MSFM12, MFG^{+}13, MJT^{+}14, MM13b, MLP19, MMM^{+}15, Met11, MM13c, MMWG13b, MW16a, MMBA13, MKI^{+}15, MWM^{+}11, MLR12, MRC^{+}13, MPF^{+}15, MJHS13, NMG^{+}13, NDD^{+}13, OTS^{+}10]. larval [OLD^{+}15, OTK^{+}19, OT13, OT14, ORW15, ORM^{+}13, PSC^{+}13a, PDCH15, PPSR15, PPS^{+}13, PRR^{+}16, PW19b, RFH16, RM13, RWR^{+}15, RNB^{+}16, RHH^{+}15, RWAP16, RSH^{+}19, RCG^{+}15, RWH14, RP17, RMR12, SAWK18, SWMK15, SKD^{+}17b, ST11a, SS12a, SFTG14, SMVM^{+}12, SS16, SPW^{+}12, SPS16, SFNP11, SS10c, TSCvS^{+}15, TPC^{+}13, TAB^{+}14, THW^{+}16, TTM19, TGP^{+}16, VNN10, VBT^{+}17, VCFF13, WBG10, WAP10, WMS^{+}10, WSZS16, WLL^{+}14, WLH^{+}17, WBHL10, WMMC14, WBB13, YHT16, YRM^{+}12, ZSSK11]. laser [HJ13, VUL^{+}11]. laser-induced [VUL^{+}11]. lasting [RP17]. lata [DWPH11]. Late [PPH11, SB14, AF14, HIH^{+}17, NSK^{+}14, STK^{+}13, SE11b]. late-stage [SE11b]. Latent [LC13, GFP19, RP17]. Lateral [LGC10]. lateralis [JVSD10]. Laternula [BCR15]. latest [Ano17c]. Latin [GHL11]. latissima [NDAP19]. Latitude [CS11a, BOKJ18, lLChTC11, CKJOM14, GSV14, HS14a, HSCA17, NvOH13, NMHB16, RWM^{+}14, RB17, RM19, SS16, TPCN12, WGD10, WHFM17, WNF^{+}11]. latitudes [Hos14, PAK15, SCP16, WAP10]. Latitudinal [BB11, DHGW15, FKvdV12, JUMS10, LKS12, MMG^{+}15, MPLLE17, MMM^{+}15, SBV^{+}15, VCDP11, CFF^{+}18, CS11b, DR16, Ebe10, FPM14, HBM^{+}12, JMB^{+}17, LAK^{+}19, LAB10, MKII19, PRP^{+}18, RS13, RRL^{+}19, RGDS12, SHK14, SWPK13, YSH^{+}16]. Latitudinal- [MMG^{+}15]. latro [SY10]. Lau [PML^{+}10]. Lauzon [MHO^{+}11]. Lauzon-Guay [MHO^{+}11]. lava [PPTK19]. Lavaca [PPM11]. lavage [KHI12]. law [CGH^{+}16]. Lawrence [GLM^{+}19, LRTLF12, PPJT15, PPD^{+}14, RBP^{+}10, BPOS^{+}16, BSB^{+}11a, BDVP19, BOCB10, LMLB12, SRP^{+}19, TBH^{+}18]. lay [SBCA15]. layer [NSG^{+}12, NMG^{+}13, NYK^{+}13, NKA16, SGMG^{+}14, SRRS12, TPS10, YOST11]. layers [BBRBG10, GMD16, MBSG11, MJT^{+}12, PIGK16, PCLB18, RLAGL12, SBB17, SMP10]. laying [DED^{+}15]. Lazarev [HS14a, MAS^{+}10b]. lead [APC^{+}14b, LHJ11]. leadership [GGD^{+}11]. leading [CDMW12]. leads [CV11, CMK^{+}14, FSLG12, JTWJ^{+}13, KTH^{+}18, PRFA11, Rit17, SRK^{+}18, YHR16]. Leaf [GBK^{+}14, dlSBO^{+}12, CESAM12, dSKFV12, GVR^{+}14, GZB16]. Leaf-fracture [dlSBO^{+}12]. leaping [KBP^{+}10]. learn [RSW^{+}16]. learned [LGM^{+}12]. learning [FHH^{+}17, SUPJ^{+}10]. least [DKB^{+}12]. leatherback [AJ16, DP14b, GBD^{+}12, HKB^{+}19b, LMSC^{+}19, PMGQM17, RTDUM16, RMNP17, SSB^{+}11]. leaves [CY10, FWW^{+}16, GZB16, HN13]. lecithotrophic [FSB15, VHM16]. led [FMC^{+}10]. legged [BER17, CPF^{+}14, CLC^{+}12b, CDMB^{+}18, GAB^{+}14, HCDR^{+}15, MATM^{+}19, MIA^{+}11, REAN^{+}14, WKK18]. legislative [HGW^{+}11]. leidyi [BFH^{+}13a, BB14c, DTM^{+}12, HTHH11, KBPD10, MLC14, PDBB14, RG14, TLG14]. Leiopathes [CSABH^{+}13]. Leiostomus [BT13]. lemon [CBL^{+}10b, GGD^{+}11, GGF^{+}12b, NHG10, PWNP^{+}11, SCF^{+}17]. Length [BRG^{+}14, RST15, SKR^{+}13, YBD^{+}16]. length-frequency [SKR^{+}13]. lengths [BGB^{+}17]. lenses [SST19]. leonina [CMHH17, JLPG17, LCB^{+}16, LRO^{+}17, RDGC17]. leopard [MHS15, MCS^{+}15, NCK^{+}14]. leopardus [BK14, CCM^{+}15, MHS15, MTS^{+}17]. Lepadogaster [FG10]. Lepeophtheirus [GSM^{+}11a]. Lepetodrilus [NWS^{+}14]. Lepidochelys [PAC^{+}15, RAKB^{+}12]. Lepidonotothen [DSH14]. Lepidopsetta [LBD^{+}15]. leptocephali [LFM^{+}16, OMT^{+}17]. leptonyx [MCS^{+}15]. Lesion [DDD^{+}11, BFKR17]. less [HCG15]. Lessepsian [GB12]. Lesser [PMEZ15a, FEE^{+}11, JROS17, RGW17, TRSB^{+}18, WHN^{+}18]. Lessonia [OAD^{+}14]. lessonii [LMBM18]. lessons [LGM^{+}12]. Let [PG15]. Lethal [CL15c, LM18, AdCM^{+}19, BPHW16, LAALH10, MWBMH13, SMBC18, TM15]. lethrinid [MHFN10]. Lethrinus [LCG^{+}17]. Leuca [DIG^{+}11]. leucas [BMJH12, HLPS^{+}15, HS11b, HIH^{+}17, LSH^{+}19b, MH12a, SWM^{+}19, SHB^{+}19, SOC13]. Leuresthes [BBC12b]. Levant [LKL^{+}17]. Levantine [SKGS11]. level [AH13, AGRPSVL13, BIL^{+}10, BSP^{+}10, BPM^{+}18b, EBOM11, EFOM18, GHGF12, JGWJ^{+}11, JBT^{+}16, LRO^{+}17, MWGH15, Mus12, OMD11, Pör12, PLF^{+}12b, RPE^{+}13, RP17, SG18a, SBB18, SCB^{+}14, SDAM11, VBGO^{+}16, WK18]. level-based [SCB^{+}14]. levels [AHR^{+}18, BLV^{+}19, BH10, BOU13, BDB^{+}12, CB14, CLG^{+}19b, CV11, CNÁCP12, DHJ12, ERG^{+}11, EMTB^{+}12, FC10, GAB^{+}14, GM10, Hal15, HSB12, JCRC10, JDX^{+}17, JCM13, KCW^{+}17, MH11b, MKM^{+}16, MR16, PTP^{+}19, PLZL14, PUP16, RHE14, SWPK13, SDH13, TTF^{+}11, YBD^{+}16]. Leveraging [QBW18]. Lévy [FCRD18, UAK^{+}14]. Liaodong [DWD^{+}18]. Libya [SMH^{+}12]. lice [HGG^{+}18, SST^{+}18]. LiDAR [HR19]. LiDAR-based [HR19]. Life [BOKJ18, DMvB10, dlÁGRB^{+}19, LHJ11, LOES11, NWS^{+}14, SJB16a, VRAP^{+}14, AS14, AAJJ^{+}19, ABW14, Ano18b, AKF14, BA12, BDS18a, BMJH12, BEFC15, BL13, CAR^{+}19, CRJ13, CB17, CLMB13a, CLMB13b, CG16, CE13, DSG15, DMD^{+}12, DDDR10, DBG15, DQC^{+}13, DP13, Fog12, GSV14, HLF^{+}19, HWET13, HS14a, HEB^{+}11, HJ13, HDP^{+}17, HDT^{+}12, HHB^{+}14, JMK12, Ji11, KFW^{+}18, KPW^{+}13, KCGC14, KKPL14, KDJP10, LCQT^{+}13, LAB10, LSQ^{+}13, LALÁC11, LTN^{+}11, LLF^{+}18, LPB^{+}16, MH19, MM12, MH12b, MMA18, MRB^{+}11, MDR16, MGD^{+}11, NRF11, NWC^{+}10, OHA10, OFFQGG10, ORWPR14, OKLM14, PKW^{+}11, PSG11, POKGRP10, PRP^{+}16, RB11, RWP^{+}14, RNB^{+}16, RBC^{+}14, RTC^{+}14, SRSH^{+}18, SDAH12, SMK^{+}17, SP12, SKD^{+}11b, SNT^{+}16, TB14a, TGRH17, VM14, WSHY16, WBW^{+}14, ZBP18, vdWHG11]. Life-history [VRAP^{+}14, AKF14, CLMB13a, GSV14, HWET13, HEB^{+}11, LAB10, LLF^{+}18, MM12, MH12b, MDR16, ORWPR14, PKW^{+}11, PSG11, WSHY16, ZBP18]. life-long [LTN^{+}11]. life-style [BL13]. lifespan [SPH18]. lifetime [MMB^{+}12a, VWEP18]. lifting [DBCP15]. Light [CRG^{+}16, GLM^{+}10, LAK^{+}19, MPB10, SMF10, WLH^{+}17, AvDL^{+}19, BVG13, BBC^{+}10, BOU13, EOA15, FGOS16, GWH^{+}18, GFR^{+}18, GCBB13, GVGP^{+}14, HP11a, HRS^{+}17, KROA19, KHA^{+}12, KKPL14, LNS10, LH14, LS10a, LSV10b, MJW19, MBC18, MAS10a, MRC10a, MSS^{+}14a, NKH12, NKA16, OFM^{+}18, PL15, PNPF14, PK15, PLS^{+}10, PUP16, RMHP14, ROE17, RS16, SRB15, SMR11a, SSHB16, SLL12, SLC11, SA11, SMK^{+}17, TS11a, TW16, VB14, VPBV13, VBGO^{+}16, VMWU15, VCFF13, WGD^{+}19, WTP^{+}18, ZIT^{+}17, dlSBB^{+}10]. light-based [GWH^{+}18, LNS10]. light-controlled [NKA16]. light-environment [RMHP14]. light-use [TS11a]. Ligurian [TRHM14]. like [AMLS15, EMG^{+}15, Fis10, MAHK12, OCMT13, PWCvO^{+}14]. likely [PCH^{+}14]. Limacina [BNE^{+}19, CAG12, SLW15]. Limanda [PSA12]. Limecola [BDD17]. Limfjorden [PMYH13]. limit [BPF15, CLM^{+}13, CWMP13, CPW15, DKV^{+}16, JMK12, LAM10, LE13, PHB11b, RLTL14, RBS16]. limitation [BRB17, CMF^{+}19, Dao13, DKV^{+}16, EVB^{+}16, GHGG11, LNIH16, Met11, ORJH19, OAGV11, PMEZ15a, PMEZ15b, PBF10, RWM18, SFW^{+}14, SP15, SMF10, TLS^{+}19, TJS17, TPSS12, VMW^{+}10, WRO^{+}11, YLPG19, vdVFK^{+}10]. limitations [AGV^{+}11, CL17, MWP10, RWH14]. Limited [DXF^{+}18, DED^{+}15, GS18, PRP^{+}18, SO16, ALK17, ALSV16, BMW^{+}14, BFH13b, CL10c, DSH14, DB14, DTM^{+}12, ENW^{+}14, GFF16, GKRG^{+}19, HCG15, LGBSC11, LMH12, NTS^{+}18, PAF^{+}10, PZA^{+}13, SGMG^{+}14, SSR^{+}10, TSCSB15, WTD^{+}19, WLP^{+}10]. limiting [FGP17, GALS16, VHG^{+}18]. limits [GPB^{+}15, HB15a, KJCN^{+}11, LDT^{+}13, MVCD14, MMB^{+}12b, RAN^{+}10, SPWK13, TSCvS^{+}15]. Limnocalanus [VLK^{+}15]. Limnoithona [GK10]. Limnoria [HVT12]. limpet [CMP^{+}13, DHGW15, LMBM18, MCTHGV12, SHC^{+}10]. limpets [AVB13, ICUM10, NBM12]. Lindeman [LPSL14a, NBM^{+}14]. line [JTH^{+}12, RAF^{+}14, BWV^{+}12]. line-transect [JTH^{+}12]. lineages [DHJ^{+}17, PFPP18, WNA^{+}15]. linear [CBYH13, GWM^{+}15, HMP15, LRK^{+}15, RRH^{+}16, VBGO^{+}16, YIKG10]. lingering [BBC^{+}12a]. Lingulodinium [BCM19]. link [AMLS15, BPM^{+}13, ERG^{+}11, HKB^{+}19b, KEBB10, LRB^{+}12b, MJT^{+}12, MCW^{+}17, PBH^{+}13, WKK18]. linkage [LG18]. linkages [ASI15, BCWSR14, CL10a, DRDV15, JWBS10]. linked [AHA^{+}15, BEP^{+}12, BSK^{+}11, BRS10, FCR^{+}15, Har16, HRAM13, HPN11, JW11, KF13a, LPJ^{+}13, LMCD19, Mou17, MvdHBM12, PPM^{+}15, RHMR18, ST14, TSTP13a, WBJ12]. Linking [BLC^{+}16, BKAM12, CBMK10, DLC^{+}13, HLB^{+}16, JJF14, LBVH13, MMG^{+}16, RJTS12, SKM12, SKB15, SBE^{+}16, SMVM^{+}12, WJHS13, AHH^{+}11, LRL^{+}13, PMK^{+}15, PHB^{+}11a, Pil10, Pör12, QGM^{+}14]. links [ASR^{+}19, BD14, CLH^{+}15, CSC13b, GOG14, LCL^{+}17, MBT14, RB11]. Lion [ECZD14, BSG^{+}10, BOC^{+}16, ED12, SCW^{+}15b]. Lionfish [HGA^{+}16a, Ing16, AL12a, Alb15, ACL^{+}16, BWM^{+}15, BMA^{+}10, Ben16, CCT12, CM10, CGM^{+}13, CBK^{+}12, CMH14, CWAS17, DPS16, DEP19, Dav19, EPM^{+}16, EF16, GAC11, GCR16, HTSA16, HGA^{+}16b, HGA^{+}16c, IW15, JP15, KA16, LA12a, LFRF13, MCW11, PHSBO16, SO15, SJO16, Tut17, WMB^{+}14]. lionfishes [KBC^{+}12, MG14]. lions [ASAG17, ACMD11, DCC^{+}10, DFTZ^{+}15, JWR^{+}12, LHHG11, MMSC17, SDR^{+}15, SGC17, VGR^{+}16, BMF^{+}15, BLC^{+}16, MBF16]. Lipid [BNE^{+}19, CWD11, KSA^{+}13, TRR^{+}17, WNNN12, BS19b, CMB^{+}14, CLSS17, CRS^{+}18, DHG16, Hal15, JCM13, JMB^{+}18, LMR^{+}10, LBE^{+}10, MPZ10, MASMH19, PBAP^{+}14, PSB17, RJTS12, SMF^{+}17, VBT^{+}17, ZdVA^{+}16]. Lipid-enriched [TRR^{+}17]. lipid-rich [JCM13]. Lipids [PDCH15, ILDL10, KHI12]. lipofuscin [BD11b]. Lithothamnion [BRF11]. Litopenaeus [BM10a, DPH14]. litoralis [NWC^{+}10]. litter [GVR^{+}14, TKDM16]. Little [GWK^{+}12, ALWD15, FCH^{+}10, HWR^{+}14, JGWJ^{+}11, JTWJ^{+}13, KHW^{+}10, MJA^{+}18, PHHL19, SDCC12]. littoral [GYRJ15, HCL^{+}11b, MHHF13, VLM14]. Littoraria [KM10b, ZWP^{+}17]. littorea [PP18]. Littorina [PP18, RB15, VE18, WG11]. littorinid [MVL^{+}16]. Live [XLXS17, ASK^{+}15, BOS16, NLLS15]. live-cell [BOS16]. lived [BBL11, BPG^{+}11, CSABH^{+}13, GSM^{+}11b, JWB^{+}17, KHW^{+}19, LD10, LBT^{+}10, MFSRR14, PCH11, RBBG12, SKM12, SLC11, YBD^{+}16]. Liver [SNK^{+}10, SBH^{+}18, SBL^{+}10]. lividus [CT10, PAFC13, PPC^{+}12]. Living [JLL^{+}12, TS16, BHE^{+}17, DDM^{+}16, JMPZ10, KdWV^{+}18, LLRC^{+}12, LLF^{+}18, MBG^{+}16, NGR^{+}17, NSH^{+}15, RSP17, WBG^{+}14, WSK^{+}16, WFH11, vdGSE^{+}14]. Liza [KBP^{+}10, PWW12]. Lizard [BKB11, MBB^{+}18]. Loa [PPTK19]. load [JREVB18, LB15a, MH13, SPP15]. loading [VARO12]. loadings [SLR^{+}14]. Lobatus [BBdG^{+}19, KCS^{+}17, SD17b]. lobifera [PUP16]. Lobophora [RMME10]. lobster [AB13, BW10, BPF^{+}18, BFEMLÁJ19, BFG^{+}11a, BFG^{+}11b, BPG^{+}11, CCHG10, CB17, CS11b, DMPG11, ESW14, JW10, KLKM12, KBG^{+}13, KB14, KPS14, LS10b, LH10, LÁBFÁF^{+}10, MPH^{+}11, Mat10a, MLCC19, MCC^{+}16, MLF^{+}18, MKI^{+}15, MOS10, MOK^{+}11, MUOS13, OLD^{+}15, RS13, SRG14, SRM^{+}15, SPS16, TFH^{+}19]. lobsters [BHP11, BCWSR14, EHD^{+}17, FRTB14, dlÁGRB^{+}19, GELJ19, GW15, HBC15, JWHR11, JBCGC19, LÁBFÁF^{+}10, MR16, ODP^{+}12, TMH^{+}18]. Local [AFF^{+}15, BFS10, BGP^{+}16, BHN^{+}14, BSCH14, MNKB11, NDBL16, PBTB15, THMF^{+}15, VJG^{+}13, AGC12, Ben16, BSP^{+}10, BW10, BVQ12, BJRH10, CFPF^{+}14, CBF^{+}19, DCW^{+}10, Ing16, LTN19, LB15b, MMG^{+}15, MHF^{+}19, MvdHBM12, PHI^{+}12, PHD^{+}13, RBHR19, SMBH15, SUB^{+}13, dJH11]. Local-scale [MNKB11, NDBL16, PBTB15, MMG^{+}15]. Localised [RST^{+}10, RBC^{+}14, SWK18]. localized [NHS^{+}17]. locate [LRBSJ15, SRK13]. Location [SWB^{+}10, CUCS17, CEC^{+}18, GWH^{+}18, HDB^{+}10, HFN14, TPT^{+}12]. locations [BOU13, GWK^{+}19, LPI^{+}14, OWH^{+}19, PSHG16, SSR^{+}10, THW^{+}16]. loch [Hug11]. loci [HW11b]. locked [DJJ^{+}16]. loco [MC11, MGO^{+}12, MJO^{+}13, MJT^{+}14]. locomotion [WJWC16]. locus [FWPP12]. Loggerhead [BLD^{+}16a, SMH^{+}12, ALE^{+}13b, AGP^{+}13, BBRM14, BDM^{+}13, CCE^{+}14, CCC^{+}18, ECM^{+}12, FCRD18, HAM^{+}15, LPC^{+}11a, LKL^{+}17, MASW^{+}11, MLML^{+}19, MO13, PBAS^{+}10, PPM^{+}15, PDT^{+}11, PSAK12, RPE^{+}13, RAB^{+}10, SMVC^{+}15, SBB^{+}12a, TMM^{+}19, THB^{+}12, TMS14, WRGP12, WHK^{+}11, WFH^{+}18, dQSC13, dQSC14]. loggers [HCDR^{+}15, LAM11, NSY^{+}18, WMHW14]. logging [CCA^{+}17, PMGQM17]. LOHAFEX [SMK^{+}18]. loliginid [SHM^{+}10a]. Loligo [SHM^{+}12, CRAP^{+}12, SJ10]. Long [ASEB11, AMLR^{+}17, BMSB13, BNAN12, BDP10, BDDvdM17, BEEH15, BBC^{+}12a, BHMCM16, BBMM14, EPL^{+}14, Gal14a, Gal14b, HI14, JDG^{+}17, LBJK14, MHAP18, MUOS13, MBKM17, MMFP^{+}15, MVOPB^{+}16, MOK^{+}18, OFPR^{+}11, PBC^{+}18, PPM11, RU11, RRL^{+}19, RMRA15, SGC^{+}13, SKD^{+}11a, SJL^{+}16, SRK11, SSvD^{+}15, SS14, WAM^{+}15, dJdBdJB12, vdVKA^{+}11, BBM13, BD13, BD14, BMSS14, BBL11, BGK17, BPG^{+}11, CR16, CBKM16, CSABH^{+}13, CTZ^{+}16, CB17, DL13, EMTB^{+}12, FDVL18, FKZTT18, FHMG19, FMH^{+}19, FEE^{+}11, GNJJ^{+}17, GSM^{+}11b, GTL^{+}16, HGL15, HAM^{+}15, HER^{+}19, HG13, HA14, HBG^{+}17, JWB^{+}17, KHW^{+}19, KV13, KB14, LME^{+}19, LD10, LBT^{+}10, LBL^{+}12, LPK^{+}16, LFD13, LTN^{+}11, ŁCM^{+}18, MFSRR14, NDAP19, NCC^{+}14, NWH^{+}12, NWD13, ORE^{+}17, OKLM14, PFB^{+}13, PCH11, RBBG12, RP17, SKM12, SKB15, SLKN14, SKFH18, SBLM14, TSA^{+}18, TS11b, WVM^{+}14]. long [YYO14, YBD^{+}16, dGR11, vBMH^{+}17, vDGK^{+}12, KKY15, MOH^{+}10, SLR^{+}14, TG10, TLG14]. Long-distance [BHMCM16, DL13, FDVL18, NWD13, vBMH^{+}17]. Long-finned [MMFP^{+}15, HER^{+}19, WVM^{+}14]. long-lasting [RP17]. long-lived [BBL11, BPG^{+}11, CSABH^{+}13, GSM^{+}11b, JWB^{+}17, KHW^{+}19, LD10, LBT^{+}10, MFSRR14, PCH11, RBBG12, SKM12, YBD^{+}16]. long-nosed [FHMG19, FMH^{+}19]. long-range [CTZ^{+}16]. long-spined [BGK17, KB14]. long-tailed [vBMH^{+}17]. Long-term [ASEB11, AMLR^{+}17, BMSB13, BNAN12, BDP10, BDDvdM17, BBC^{+}12a, BBMM14, EPL^{+}14, Gal14a, Gal14b, HI14, JDG^{+}17, LBJK14, MHAP18, MUOS13, MBKM17, MVOPB^{+}16, MOK^{+}18, OFPR^{+}11, PBC^{+}18, PPM11, RU11, RRL^{+}19, RMRA15, SGC^{+}13, SKD^{+}11a, SJL^{+}16, SRK11, SSvD^{+}15, SS14, WAM^{+}15, dJdBdJB12, vdVKA^{+}11, BBM13, BD13, BD14, CR16, CBKM16, FKZTT18, FEE^{+}11, GNJJ^{+}17, GTL^{+}16, HGL15, HAM^{+}15, HG13, HA14, HBG^{+}17, KV13, LME^{+}19, LBL^{+}12, LPK^{+}16, LFD13, ŁCM^{+}18, NDAP19, NCC^{+}14, ORE^{+}17, OKLM14, PFB^{+}13, SKB15, SLKN14, SKFH18, SBLM14, TSA^{+}18, TS11b, YYO14, dGR11, vDGK^{+}12]. longevity [KWM12]. longfin [SJ10]. longibranchiata [CLN^{+}13]. longicarpus [PDB15]. longicornis [CA14, DMvB10, DvBH12, GvSGCK14, KDB^{+}10, SK11]. longimanus [DOH^{+}17, MBB^{+}15]. longitudinal [BPHW16, CAB^{+}19a, TRLS^{+}11]. longline [CEH18, CGBGS18, DCHI12, LLMA^{+}10, Løk11]. longliners [CAGS17]. longlines [TBMG17]. Longstanding [VWH18]. look [ST11a]. looking [FG16a, RR18]. loons [WMPM13]. loop [DVSM10, MSE^{+}18, RdGv^{+}18]. loop-dominated [DVSM10]. Lophelia [BJ14, GSC14, LLv13]. Lophius [CALB10]. lophophore [KP11]. losers [GMG^{+}16, vWSGL11]. losing [CFH11]. loss [ADvdM13, ACP^{+}16, ASF13, BSD^{+}14, BC14, DTC^{+}17, FMC^{+}10, Fox16, GHES18, GKBEF18, Har16, HBC15, KGJ^{+}16, LSBB17, ŁCM^{+}18, MHPvdH19, MWD10, MMRDS18, MRH11, MS11, NCS^{+}10, PPTB12, PBG^{+}10, SSM15, SCHO11a, TKBN10, WLB^{+}11]. losses [MO12, SKK^{+}18]. lost [CM14, PRB^{+}17]. Louisiana [KCMH^{+}17, GFC^{+}13, MRC^{+}19]. Lovushki [SBA12]. Low [DGMB10, FWW^{+}16, GFP19, LRO^{+}17, LBC^{+}19, SPMN19, SMF^{+}17, SM15, TMS^{+}16, BLV^{+}19, BLD^{+}16b, BPB^{+}12, BMA^{+}19, CMR12, lLChTC11, DGC18, DBG15, DRD^{+}18, EBOM11, FRMS17, GGLT15, HWR^{+}13, HED^{+}13, HS11b, HP14, KKS12a, KPB11, LWM15, LMW12, LMZ16, MPW12, MAP^{+}12, NOB^{+}19, NSI^{+}15, PNPF14, RLTL14, RP17, SCO^{+}12a, TW16, TTF^{+}11, TvdMM^{+}17, TPKC18, VHBG15, VPBV13, VSB^{+}15, WNF^{+}11, DEGF^{+}10]. low-abundance [MAP^{+}12]. low-diversity [NOB^{+}19]. low-intensity [CMR12]. low-latitude [lLChTC11, WNF^{+}11]. low-mortality [HS11b]. low-oxygen [LWM15]. low-productivity [KKS12a]. Low-salinity [SPMN19]. low-tide [HED^{+}13]. lower [BCM^{+}14, CV11, KRA17, LH14, PMF^{+}17, RHE14, RLTL14, SDM16, SL15b, TSCSB15, TYVR^{+}15]. lower-than-expected [TSCSB15]. lowered [BDAH10]. Loxosomella [SIHH10]. lucerna [NMC^{+}16]. Lucia [BP12b, CP10, CPP10, DPC13, PPB10b]. Lucie [KBK^{+}15]. lugworm [DWM15, EBK^{+}12]. lugworms [PMM^{+}15]. Lumps [WDNM11]. lumpsucker [MP12]. Lunar [OAB^{+}19, SBB18, SRB^{+}17, FHL^{+}17, LJM^{+}10, MMB^{+}19]. lurida [PRR^{+}16]. Lusseau [NWD13]. lutea [DDD^{+}11, EES15]. Lutjanidae [DSS10, DSC13, KCLP16]. Lutjanus [LBCS^{+}15, MH19, MDA^{+}10, TEAO^{+}14, TS11b, WGS16]. lutris [MBB^{+}12b]. luzonensis [HLW^{+}16]. lyase [FSY^{+}10a]. Lychnorhiza [NMC^{+}16]. Lyons [MHO^{+}11]. lysable [GRCP^{+}10]. lysis [AWJ12]. Lytechinus [PCW^{+}12].

M [Ano12c, CBJB19, CATAV18, SL15a]. M. [BHP11, LIT13, MWP10, RPLT10]. Māui [DCB^{+}16]. macaroni [BBH^{+}13, CTS^{+}15, HMR^{+}16, TCTB11, WHPT12]. maccoyii [IS16]. machine [FHH^{+}17]. Mackenzie [SWM^{+}19, STBP13]. mackerel [BNAK11, KTNO19, NSK^{+}14, RFH16, SNNS^{+}12, YYO14, YNK18]. Macoma [BDD17]. Macquarie [CAB^{+}19b]. macro [BC11, BDC10, Jon11, OLM^{+}12, PWNP^{+}11, SDM16]. macro- [BC11]. macro-ecological [BDC10]. macro-intertidal [OLM^{+}12]. macro-invertebrate [SDM16]. macro-nutrient [Jon11, PWNP^{+}11]. macroalga [GGF^{+}14, GM13b, JL12, LH19, VBR19]. Macroalgae [NMHB16, VPDPMRC11, BCSH18, BTM^{+}18, CPHH15, FGRM12, GLWI10, GLHM^{+}18, GGMTGC18, HCAV10, HSB12, JBD10, JT10b, LKC^{+}17, LB11, LERS13, MAL15, OFT^{+}11, PKW^{+}11, SHAS15, SS13, Sti13b, TKV^{+}19, TBC^{+}17, ULCAZG17, VBM^{+}11, WDM^{+}19, WvLF18]. macroalgae-associated [WvLF18]. Macroalgal [HP14, HPLS13, LTN19, SS19, TPS^{+}13, AAvS^{+}13, BB11, BAM^{+}10, BS18b, BS18a, GC13a, GRFSM12, HKJ17, Hoe10, LD17a, MBBBC12, MHC11, MHVV13, MLP^{+}13, ORWPR14, PCNM17, QNG^{+}15, RMSPA16, SSM^{+}16a, SSM^{+}17, TS10, TS11a, VPOA12, VDS11, YMCV10]. macroalgal-associated [SSM^{+}16a]. Macrobenthic [DBCV^{+}12, MGOL19, ÅCAC16, Bar13, BVS^{+}11, DDC^{+}15, GBD13, JH11, KSB11, MBvR12, MPB14a, QGM^{+}14, SKB15, SSD13, dJH11]. Macrobenthos [WHvdK11, EBH16, War14]. Macrobioerosion [CSM12]. macrocephalus [EPC^{+}19, FNF^{+}18]. macroconsumers [RSW^{+}16]. Macrocystis [FK10, Fox16, HHH11a, KMP15, MJW19, MZ10, PMS10a, SH14, YMPR13]. macroecological [MPM^{+}18, SBT15]. Macroecology [Bet14, RNGUR^{+}11]. macrofauna [BPG^{+}10, DM11, FC12, GGFB15, MAD18, NUO^{+}19, NLLS15, POPH12, PP13, PPM11, RMC^{+}14, RZS12, SGJ^{+}18, VFMdW12]. Macrofaunal [WAA^{+}19, BPB11, CSN12, DS10, DEWK15, FHE^{+}15, GWT18, MBG^{+}16, MRGG^{+}13, ROLA15, SBB^{+}14, VJJ^{+}19, WK10]. macrogeographic [WWG13a]. macrograzers [BHF^{+}16]. macroinfauna [MWA17]. macroinfaunal [CLS19]. Macroinvertebrate [WPRS18, BRA^{+}10, DHJ10, GLM^{+}10, MTM^{+}11, MH11c, ULCAZG17]. macroinvertebrates [NLK^{+}19]. macromedusae [ADD18]. macroparasites [FTO^{+}15]. Macrophyte [RHMR18, LBP^{+}16, MCRB11]. macrophytes [DSG^{+}13, FESPR14, LMW12]. macrophytic [MBB11]. macrorhynchus [NSC^{+}19]. macroscale [CGEGC^{+}19]. macrotidal [QNG^{+}15]. macrozoobenthic [CBYH13, NSI^{+}15, WH11]. macrozoobenthos [PLG^{+}19, UDF^{+}18]. macrozooplankton [LOBW^{+}11]. macrura [DRH15, HS14a]. Macruronus [AIA18]. macrurus [VLK^{+}15]. maculata [TWM^{+}15]. maculatus [HS16, RFH16, SS12a]. Madagascar [CTZ^{+}16, MJ17]. made [HRC12, LBC^{+}19]. Madracis [MWP10]. maenas [ATP^{+}12, BSH^{+}14, DT15, DCT^{+}10, HBHF^{+}11, MMG^{+}16, MEK11, WGD^{+}16, WNA^{+}15, YPK15]. maeotica [CGQB10]. Maerl [BMV^{+}17, SBCA15]. Magellanic [RB18, SWFQ12]. magellanicus [AIA18, BCW19, CWS13, HS11a]. magister [HEAI10, SE11b]. magnesium [SCL^{+}16]. magnificent [ADA^{+}19]. magnitude [BB17, DBCB11]. Maguimithrax [BHUCMO19]. Maiden [WRR10]. main [FTO^{+}15, MDKM13, OAB^{+}19, PAC^{+}15]. Maine [ALS18, BW10, BM16, CDR12, FOL12, GW15, HCJ^{+}11, MLCC19, PWA^{+}17, RFMT12, SDKS19, STT14, SBKA14, SJS^{+}11, WWAL13]. mainland [FPC^{+}13]. maintain [BDT^{+}16, PBM^{+}18, SWK18]. maintained [OFPR^{+}11, VHBG15, Wil14]. maintaining [GS17]. maintains [BBTW14, MBM^{+}14, RWM^{+}12]. Maintenance [JW10, MDF10, PA13, STT14]. major [BCSH18, CPR^{+}14, CNJ^{+}18, ED12, HBE^{+}13, HGA^{+}16a, JCF^{+}12, JJNP12, PMB^{+}17, SBLM14, SCHO11a, TKDM16, WBS^{+}12]. Makaira [FHTY19]. make [JSBP15, ZJ17]. make-up [JSBP15]. making [SSFB11]. mako [PBB^{+}15, VWW^{+}16]. makos [NCW^{+}19]. Malacanthus [PFVFF^{+}15]. Malaclemys [HW12a]. Malayan [PB10]. malayensis [HSH18]. Maldives [CMGS19]. Maldivian [YPWP18]. Male [PVYS16, BPT15, BMSS14, BBWG12, HKL15, KBA17, NCK^{+}14, VCG^{+}15, WFP15]. male-male [BMSS14]. Malene [SKR^{+}10]. males [CTRBLC^{+}13, DKL^{+}18, GDJdBSL19, KNY^{+}15, MS18b, QSvN^{+}11, SKK^{+}18]. Mallorca [ATT^{+}16]. Mallotus [CD16b, LBR^{+}19, OA10]. Mamala [SMP10]. mammal [BHN^{+}12, GDA^{+}18, LJM^{+}10, MTP^{+}10, RNGUR^{+}11, SGGL15, TMT^{+}12]. mammal-eating [TMT^{+}12]. mammals [AGRPSVL13, CAB^{+}16, JSK19, KHW15, MGM^{+}18, OGGP14, PBCA10, SHR^{+}11, WJS^{+}14]. Man [MHK11, HRC12, LBC^{+}19]. man-made [HRC12, LBC^{+}19]. manage [BMV^{+}18, CSS^{+}13, MPB^{+}14b]. managed [BDS^{+}18b, LCG^{+}17]. managed-care [BDS^{+}18b]. management [ABRP15, Ano18b, AGIW16, BPM^{+}13, BBBF15, BPZ^{+}14, CEH18, CBC^{+}13, CM14, CCG^{+}19, DFLC17, DBB11, EW16, FABH^{+}16, GGF^{+}12a, GLW^{+}12, HT10, HLH18, LHB^{+}15, LANE18, MM13a, MA14, MRH11, MBC^{+}11, MBF16, MSV^{+}11, NAG^{+}16, NLKZ12, Nob11, OMZ10a, OMW^{+}10, PLG^{+}16, PHB^{+}11a, PRP^{+}18, RWM^{+}10, RCS^{+}17, RvOP12, RPB17, RFEJK11, RG11, SMH^{+}12, SCB15, SRG14, SDSA14, SDST16, SHT^{+}16, SGGL15, SFMB10, SSFB11, TFB^{+}17, TSTP13b, TTC11, TRJ^{+}11, VQSH14, WF18, WBS^{+}17, WPMP14]. managers [AHD12]. Managing [MJ17, GFC^{+}14, MBCB^{+}19]. Manatee [LPSK17]. manatees [ASWB10, FFS^{+}15, SRK13, SSdW^{+}12]. maneuvers [AAM^{+}12]. mangle [ATP14]. Mangrove [BMA^{+}10, CED^{+}19, CGK13a, GZB16, JC17, She13, ATP14, BBH^{+}11, BJH13, BHMCM16, CMF^{+}19, CGK13b, CY10, CY13, CCT12, DKV^{+}16, FWW^{+}16, dSKFV12, GVR^{+}14, GKW10, GBK^{+}14, GSU12, GCCCM16, GCCM18, HHS10, HMS10, HHH^{+}11b, HED^{+}13, HN13, HFK^{+}13, HPQ15, JOFP18, JWBS10, KCE14, MOP^{+}15, PDBG14, PB18, SKD^{+}11a, SSPKF16, She12, SDM16, SKMS^{+}18, VDV^{+}15, WAMS^{+}14]. mangrove-dominated [SDM16]. mangrove-seagrass [HHS10, HMS10]. mangrove-seagrass-coral [BJH13]. mangroves [BCdV13, DdRH14, DS10, HSS^{+}10, KSHS19, KKS^{+}12b, MK13, MDA^{+}10, SJB16b, SLO^{+}17]. Manila [BFL^{+}11]. Manipulating [LM12]. manipulation [MLB^{+}14]. manipulative [BH12a]. manta [DBB11, JRW^{+}14, PDM12, PSD^{+}19, RPM^{+}13b, SSM^{+}19, DBB11]. mantis [HCBV^{+}11, NKH^{+}10]. Mantua [Ano10d]. many [DBS^{+}13]. Maoricolpus [RJR18]. map [KMP11, LMRRB17, PPX17]. map-based [LMRRB17]. Mapping [GKPT16, LDVP^{+}17, MVS^{+}10, RFEJK11, SHKR15, SRK13, CL10b, GCSF18, HR19, LDS^{+}15, LJM^{+}12, WLM12, WY15]. Mar [ACA^{+}10b, BW16b, RUP11b]. marbled [JNYT13]. Marenzelleria [BBG^{+}13, DBCV^{+}12, JLQ^{+}14, KHD^{+}11, RF14]. margin [CGR^{+}15, CL15a, DRT13, LGC10]. Marginal [BAVSP11, WEK14, ACMD11, GSM^{+}11b, HKI^{+}17, WGD10, WBJ12]. Marginopora [MU13, RUP11b, RUP11a, SHDR13]. margins [HRAM13]. Maria [DIG^{+}11]. Mariana [SJL^{+}15]. mariculture [TZF11]. marina [AM18, AOW11, BO13, BD12, BOKJ18, BBRBG10, BBR12, BRB17, BPF15, CIM^{+}13, DKV^{+}16, EBK^{+}12, HBHS^{+}18, HNMK15, IEM16, JTTB^{+}18, JAPK16, JPL^{+}15, KBA13, KHGS12, KKPL14, KJCN^{+}11, MHS14, MFS14, OKJMC15, OCS^{+}13, OS12, OMD11, RFRE19, RBM14, SRB15, SCRL11, SHKR15, SCM13, SO12, VCVK^{+}10, VPBV13, VSB^{+}15, WDH^{+}12, WL19, WNF^{+}11, ZHJ^{+}17]. marinas [BDG^{+}16]. Marine [AMSdL^{+}18, AIA18, Bri12, BHN^{+}12, BL10, CHP^{+}11, CH15, DHB18, HOL^{+}17, Kin10a, KMP14, LBT^{+}10, LN10, MMC^{+}18, MLDG12, MJT^{+}12, NHTS14, OE11, OGLH17, OAGV11, PPD^{+}14, PBC11, SMVC^{+}15, ST11a, SHH10, SHF^{+}12b, WDNM11, WSS^{+}16, WGD10, ABK^{+}14, ÅCAC16, AB11, ABH18, AN11b, AGV^{+}11, ABB^{+}14, Ano18b, AEM^{+}18, ACMD11, AMAJ^{+}17, AGRPSVL13, AGC12, BAS^{+}10, BCS^{+}12, BTS10, BOC^{+}16, BFD^{+}16, BCdV13, BFC14, BBI16, BSH^{+}18, BDAH10, BKR^{+}15, BBB^{+}15, BDPL19, BAT14, BPJ^{+}15, BOÁNF19, BVP11, BAFB12, BCV^{+}18, BLT14, BBBW17, BVAB^{+}17, BW12, BKHW13, BW15, BVS^{+}11, BH10, BVQ12, BBL11, BWCZ11, BS18b, BS18a, BGDC14, BM10b, BBBHC17, BSS18, BLM^{+}12, BSCH14, BM11, BWN^{+}14, BBMM14, CIM^{+}13, CBLM^{+}17, CLW^{+}11, CRRD^{+}13, CMV^{+}11, CZB10, CLHW14, CGC15, CHST16, CFSM^{+}13, CMW^{+}14, CWB^{+}15, CHB11]. marine [CBGDS12, Col10b, CO16, CKC^{+}10, CS10b, CCKB^{+}15, CME^{+}12, CVM^{+}17, CCGP14, CAB^{+}16, CSQRA11, DGC18, DdRH14, DMW^{+}16, DGDJ14, DGMB10, DVV^{+}18, DBC^{+}12a, DB11a, DMH15, DSG^{+}13, DGC12, DMC^{+}16, DMPG11, DZB^{+}18, DL14, DBHMG12b, DHM^{+}14, DPH13, DC17, DZL16, DAB17, ES12b, EVHA^{+}18, EPF^{+}17, FTKK14, FTO^{+}15, FUOG^{+}16, FTB14, FTWA17, FFL10, FEA^{+}15, FSR17, FTvdK^{+}12, FHH^{+}17, FBGT12, FTB^{+}17, FGL^{+}12, FZI^{+}16, FGH^{+}13, FBNJ13, FFS^{+}13, GBM^{+}12, GSR10, GHHH12, GRL^{+}14, GFC^{+}14, GBJ^{+}14, GTP^{+}17, GB16, GBS^{+}10, GLWI10, GHD^{+}10, GOG14, GSF18, GGPDS19, GRD11, GCRP15, Gre10, GBC^{+}16, GDA^{+}18, GHL11, GCL^{+}16, GMR13a, GMR13b, HHS^{+}12a, HWET13, Har16, HPN^{+}17, HMP15, HSM13, HFG^{+}19, HMB^{+}12, HBS^{+}13, HHP^{+}17, HSdBS10, HVK^{+}13, HHG10, HVBG^{+}19, HSMS^{+}15, HBCK13, HLB^{+}16, HRT12, HGA^{+}16a, HSB12, HTH10, HEM^{+}12]. marine [HGB^{+}13, HTC17, HBG^{+}17, HHH^{+}11c, HSR^{+}10, HSWS16, HW12c, HW12b, HLH18, IKS^{+}12, JW10, JW11, JCA13, JWT^{+}14, JMS^{+}15, JPP^{+}17, JSK19, KÖHSO12, KBH11b, KHBL11, KBG12, KLKM12, KSS13, KCW^{+}17, KHW^{+}19, KCF^{+}14, KJW^{+}15, KHW15, KTH^{+}18, KCS^{+}17, KRdVG^{+}18, KM17, LWM15, LJ10, LSH19a, LSJ^{+}15, LRRM^{+}11, LM12, LS14, LCTD^{+}15, LN11, LRR^{+}10, LR15, LPSL14a, LPSL14b, LRTLF12, LNIH16, LMG12, LHB^{+}15, LSRRCM13, LBS^{+}15, LMTD16, LAH^{+}15, LCB^{+}12, MP15, MBvR12, MBC18, MGM^{+}18, MORM15, MPA^{+}10, MPW12, MJR^{+}14, MM13a, MRC10a, MWC18, MJ10, MMA18, MMC^{+}16, MCP^{+}18, MAB12b, MWB13, MOK^{+}11, MUOS13, MIV^{+}10, MDKM13, MRL^{+}10, MHR^{+}19b, MD16b, MGD^{+}11, MMGB14, NBM12, NMS^{+}18, NM18, NBGGST^{+}15, NVA^{+}19, NOB^{+}19, NHAS15, NPRAGG^{+}12, OMBP10, OPN10, OSC19, OGGP14, OMZ10a, OMZ10b, OMW^{+}10, PGR^{+}10, PGR^{+}13, PRCG^{+}17]. marine [PAG19, PDC^{+}17, PCF^{+}19, PTL16, PHH^{+}12, PMB^{+}15, PSdA^{+}17, PR11, PBG^{+}10, PBCA10, PCS^{+}15, PKSN11, PJMP10, PBC^{+}12b, PSL13, Pör12, PZA^{+}13, PLF^{+}12b, PGOG12, QHG^{+}19, RHVRC11, RWP^{+}14, RKD18, Rei12, RCK^{+}11, RH12, RNB^{+}16, RUD^{+}16, RPS12, RGdLB^{+}11, RA10, RGWA11, RB17, RGCT^{+}10, RNGUR^{+}11, RFRD10, RMRA15, RG11, SRB15, SKSH19, SLRM^{+}15, SBV^{+}15, SHR^{+}11, SD17a, SML^{+}15, SP13b, SCP^{+}14, SDC^{+}16, SCB^{+}14, SCWV12, SGS^{+}19, SFM^{+}19, SKW^{+}14, ST10b, SHT^{+}16, SKFH18, SAMW14, SGGL15, SHB^{+}19, SPD^{+}14, SOA^{+}17, SBM^{+}18, SPBS14, SLSR10, SSFB11, SBL^{+}10, SLNG18, SDB12, SFL^{+}18, SLPP^{+}14, SS15b, SS15a, SH10, SNNS^{+}12, TBP11, TM10, TFÁS^{+}12, TPBM13, TSCSB15, TMBP13, TSS^{+}12, TMM15, TBS^{+}14, TMH^{+}18, TGLC10, TTPM12, TTC11, TRA^{+}14, TMH^{+}13, TVA^{+}19, TKL^{+}11, TPSS12, VKH10, VVB19, VGGF12]. marine [VTB12, VF13, VWH18, WHK^{+}11, WC18, WLP^{+}11, WTdB16, WAW^{+}12, WFD19, WPNK12, WBHL10, WHB^{+}19, WHD^{+}14, WJS^{+}14, WS11, WHL11, XS19, YTO^{+}11, ZBB^{+}11, vERRP16, vMVSW12, vBAP^{+}19, vdHTT^{+}15, GPA^{+}16, LCR^{+}16]. marine-terrestrial [RG11]. Marinobacter [GRG^{+}12]. marinoi [RSS^{+}16]. marinum [SP13a]. marinus [ADO^{+}12, EvDB^{+}14, WHN^{+}18]. Marion [LRO^{+}17, RGL^{+}16]. maritima [JPL^{+}15]. maritimus [BDR19]. mark [ARvdW^{+}16, BRB11b, CWR^{+}18, PBM^{+}11]. mark-recapture [ARvdW^{+}16, BRB11b]. mark-release- [PBM^{+}11]. marked [OPN10]. marker [PPD^{+}14]. markers [CML^{+}12, HLN^{+}19, JWHR11, LBVH13, MMFP^{+}15, MIM^{+}12, OHST19, PFB^{+}17, PHHL^{+}11, SSBA12, TTS17, WVD^{+}17]. market [MLF^{+}18, OCE^{+}12]. marking [PSG11]. Marmara [SHI^{+}10, SHB^{+}10b]. marmoratus [FRC^{+}13b]. marsh [BM10a, BFRM13, BPB18, BVGH10, CPF^{+}13, CHDH^{+}19, CAKD14, CR11, EBT^{+}15, FH10, FCR^{+}15, FSH^{+}16, GM10, GSU12, HKT10, HL14, JE10, Joh11, KAAI15, KR16, KV13, LHT^{+}17, MCK^{+}13, NCT^{+}14, PCA^{+}15, QCX^{+}10, RMB18, SO16, TB10, TNM^{+}11b, ZWP^{+}17]. Marshall [HM13]. marshes [AED^{+}10, CMP^{+}11, CFF^{+}18, CGQB10, GAB11, JBT^{+}16, JC17, MDJN19, PSA17, SLRP^{+}12, ZBY^{+}14, ZZG^{+}19]. marshlands [MVOPB^{+}16]. Martínez [MPM^{+}18]. Martinique [GRL^{+}14]. mask [BMVAS19]. Mass [JPP^{+}17, SH14, CSM12, CFPF^{+}14, CMHH17, CLMM^{+}14, CMGS19, DTS14, EGH^{+}15, FFS^{+}15, HHH11a, HRT^{+}14, JWR^{+}12, MGB^{+}19, NAY^{+}11, PSG11, RH12, SCST17, SFLG10, SBM^{+}18, WRFM16, WLB^{+}15]. mass-bleaching [CMGS19]. mass-cultured [EGH^{+}15, NAY^{+}11]. mass-marking [PSG11]. Mass-transfer [SH14]. Massachusetts [KCS12, PDB15, RCD^{+}13, SMSB19]. masses [GWK^{+}19, JWHR11, SPP15]. massive [BKB11, BRC^{+}15, BC13, CE14, CKV12, DDD^{+}11, JROS17, RRL^{+}19, SGS^{+}12, SLS16]. mastigoteuthid [BB14b]. Mastigoteuthidae [BB14b]. mat [MPBSB18]. Match [SCAM^{+}14, DHF^{+}13, PvdMLC17, PDCH15, RPM12]. Match-mismatch [SCAM^{+}14, DHF^{+}13, RPM12]. matches [BBBNT13]. Mate [CCA^{+}17, JCF^{+}12]. material [HGB^{+}13, JPMG10]. materials [KOLDP17, LBH18, NSG^{+}12, RDPR15]. Maternal [CCM^{+}15, CVC^{+}18, MOS10, BSC^{+}13, BMJH12, BS19b, CPPM12, DHJ^{+}17, FSB15, GM13a, HPMB11, SA14, vdMHW^{+}10]. mates [JRK13]. mathematical [SS10a]. Mating [SMK^{+}10, DEGF^{+}10, FC19, SK11, TPC^{+}16, WMH17]. matrix [GCDH19, KM10a, LAB10]. mats [GBD13, KF13b, KLSJ10, RJB^{+}14, SS10b]. matter [BBI16, BLER15, BVP11, CGR^{+}15, CAKD14, CMC^{+}15, DMC^{+}16, FJC10, FYG^{+}18, GBW^{+}12, HJN^{+}10, HCG15, HU13, HPMZ11, KRC^{+}11, KRD^{+}12, LRL^{+}13, LFM^{+}16, LNK^{+}16, MPB13, NOT^{+}11, QOGS^{+}16, RCB^{+}10, RdGv^{+}18, SBS^{+}13, SGP^{+}17, SE16, SBE^{+}19, SBM^{+}18, SDH13, TB10, TNSN11, VAHL11, VBM^{+}11, VBM^{+}13, WNNH10, YMPR13]. matters [FGRM12, GGFB15, Hoe10, JS12a, VBT14]. Mattos [MAD18]. maturation [JT10a, MRE14, NKH^{+}10, SAM18]. mature [ORR^{+}14]. maturity [SH12]. Maui [DBB11]. Mauna [PPTK19]. mauri [BEMD11]. Mauritania [CCA^{+}14, vdGvdLvG^{+}19]. Mauritius [RTCF^{+}17]. Maurolicus [PAK15, SA11, SRK11]. Mawson [WRR10]. maxima [HPMZ11, MTG^{+}10, DBS^{+}19, GBMW12]. maximized [TPF^{+}13]. maximum [LKCH12, MSC13, TSS11b, TSGW10, VQSH14]. maximus [CQR^{+}17, LAL^{+}18, PWO^{+}16, SDH13]. may [APC^{+}14b, AMAM^{+}18, AZSKWL11, CWMP13, FNC13, FRC13a, FMC^{+}10, GGF^{+}14, JT10a, KPr^{+}18, LHB19, LFN^{+}12, McD13, MJA^{+}18, OPN10, OAGV11, RTMJ11, RUPT^{+}11, RPS12, RDM18, SMSL17, SCHO11b, TDPS13, vMVSW12, MBKM17]. Mayotte [BBC^{+}13, EBdlTC12]. mccullochi [VHKT15]. McIntyre [BSBR^{+}12]. McKenzie [CBL^{+}10a]. McMurdo [KHD19, LS15]. meadow [ABF^{+}14, BDM^{+}14, CPEE18, CSS^{+}17, CVP11, DDB^{+}16, GMC^{+}12, IOS^{+}12, PG15, RU11]. meadow-scale [ABF^{+}14]. meadows [ABF^{+}14, AM18, BHF^{+}10, BS12, BKF15, CM12, DMJ11, FMC^{+}10, GOC^{+}18, GWME16, HHH^{+}14, OKJMC15, RFRE19, RBM14, SMSE19, SKA^{+}14, WFBM14, WWGD13, ZKS17]. Mean [TS14, GOM^{+}19, KDJP10, dJdBdJB12]. meandrina [CE14]. Measure [GSD10, BPSG11, MLCC19, SHP^{+}18]. measured [AM19, CPEE18, JAPK16, LBMZ15, VBR18]. Measurement [TNOY17, CGK13a]. measurements [AFS19, BSC^{+}15, CSKR10, RS16, SOA^{+}17, WEK14]. Measures [COK10, BBBF15, CLB^{+}14, DTL^{+}15, GNBM11, NCC^{+}14, TSGW10, TFH^{+}19, ZB16]. Measuring [BWJB16, JGCRS12, MVL^{+}16, MH11a, DOH^{+}17, LAH^{+}15, RRS15, SMMFH16]. meat [GBMW12]. mechanical [dlSOV^{+}16]. mechanics [ADF^{+}18, BNW14]. mechanism [LWM15, LLS^{+}18, MPD11, VHG^{+}18]. Mechanisms [BL17, FEE^{+}11, GvSGCK14, MG18, SKSH19, SNT^{+}18, CRHV14, DP11, HS15b, KBH11b, KLNJ15, NMC^{+}16, OFFQGG10, PRCG^{+}17, RL10, SLKN14, SS12b, WHT12, WBLRN18, WLHW10, YKG^{+}15]. Mechanistic [GLW^{+}17, DP14b, SPM^{+}13, Ued14]. mediate [AL12b, BAM^{+}10, CPPM12, ES12b, HW13, MMPH^{+}12, PRA10]. mediated [Ber12a, BB14c, BC13, CW12, CAB^{+}14, DL14, FDS12, GB10a, GOG14, GM13b, KKA12, KSB11, KHD19, LGW^{+}15, MGL^{+}16, PKW^{+}11, PMS10a, PRP^{+}16, QYS^{+}12, SSB^{+}18, SGP13, TPSS12, YKT^{+}14, YRL^{+}10, ZBK^{+}14]. mediates [AM13, BJL^{+}17, CB14, FCR14, HPQ15, KWM12, SMC^{+}14]. mediation [RFRD10]. mediator [WS13]. Mediterranean [ATT^{+}16, DKRS12, ECZD14, FCDC12, FRC^{+}13b, PTF^{+}14, TBCG^{+}12, VGGF12, AFC^{+}16, APNC^{+}15, APBC^{+}19, AMR^{+}10, AMP^{+}11, AMLR^{+}17, BBRM14, BZBR^{+}11, BCN15, BAFP^{+}13, BMV^{+}11, BSC^{+}16, BFH^{+}13a, BDG^{+}16, BGK17, BB19, BAC^{+}11, CT10, CDD^{+}16, CCE^{+}14, CMIMGS15, CBR^{+}18, CM14, CFPC^{+}13, CGQB10, CAGS17, CGBGS18, CML^{+}12, DQC^{+}13, DDB^{+}14, DFAH11, DPD^{+}12, DBS^{+}13, FC10, FPC^{+}13, FMC^{+}11, FCV^{+}16, GALS16, GCR^{+}17, HCM^{+}11, HSA^{+}15, JCA13, KKMA^{+}10, Kat16, KHÁB^{+}17, KCLLL13, LdU^{+}17, LLMA^{+}10, LSV^{+}10a, LPC^{+}11a, LCF^{+}14, LBT^{+}10, LAALH10, MJTC^{+}12, MGAV15, MDS^{+}13, MPRPM^{+}18, MMNR^{+}12, MCHC^{+}16, MRC^{+}13, Mus12, NLGE18, OD15, OFPR^{+}11, PPM^{+}15, PSdA^{+}17, PBAP^{+}14, PBC11, PCS^{+}15, PLZL14, PGVF14, PHG^{+}14, QLCUG15, RAA^{+}16, RCAB^{+}12, RDC^{+}12, RVH^{+}19, SMH^{+}12, SSBA12, SKGS11, SLPP^{+}14, TRLS^{+}11, TFM^{+}18, TPFP11, TS14, TOR^{+}10, TMB^{+}12]. Mediterranean [TGI^{+}11, UMF^{+}10, VCRUQ15, WTGC^{+}11, WVM^{+}14, ZKS17, dQSC13, dQSC14]. medusa [BBR^{+}14, NMC^{+}16]. medusae [CG17, CGC15]. meet [MdSW^{+}16]. mega [CMC10a, MKPS13]. mega-earthquake [CMC10a]. mega-epibenthic [MKPS13]. Megabenthic [BRC^{+}15, DJL^{+}16, LLC^{+}14]. megabenthos [DBH^{+}16, TRLS^{+}11, TTF^{+}11]. Megadisturbances [FB14]. megafauna [FPC^{+}13, PML^{+}10, ZSEV17]. Megafaunal [RJH14, JGL12, YD11]. megagrazers [BHF^{+}16]. megalopae [RS14, RWR^{+}15, SN17, WRB^{+}15, WRW^{+}16b]. megalopal [BSB12]. Megalorchestia [McD13]. Meganyctiphanes [KHS10b]. Megaptera [BJS^{+}16, FMG^{+}18, LFPA^{+}11, RBP^{+}10, TFW^{+}12, WNNN12]. megazoobenthic [SCD^{+}17]. Megrey [MLDG12]. meio [GGFB15, NUO^{+}19]. meio- [GGFB15, NUO^{+}19]. Meiobenthic [BLM^{+}18, JCA13]. Meiobenthos [War14, Bet14, DRG^{+}12]. meiofauna [BMC^{+}15, BP12b, FCR^{+}15, POPH12]. Mejillones [RHRB^{+}17]. Mekong [BMD^{+}11]. Melanochlamys [SPP15]. Melanogrammus [AGIW16, BCSB13, RP12, SKN^{+}15, WTGG10]. melanopterus [GMHS19]. melanostictus [HHNI17]. Melarhaphe [GDSPRA13]. melas [WVM^{+}14]. Melbourne [RPE^{+}13]. melon [MHHB^{+}17]. melon-headed [MHHB^{+}17]. melt [GLS^{+}15]. meltwater [TNM^{+}11a]. meltwater-influenced [TNM^{+}11a]. membranacea [DM17]. Membranipora [DM17]. Memoriam [EN13, Kin10b, Ras10, SGC14]. Mendelian [RGCT^{+}10]. menhaden [AHH^{+}11, LBCL10, OFDG14]. Menidia [GM13a, PB19]. Menippe [HG17]. MEPS [BS18b, Kin10a, MHO^{+}11]. Mercenaria [CMK12, GLF^{+}17, MW16a, MAL15, OPN10, SF18, WBG10]. mercury [BLJP14, MSS^{+}14b, PBC^{+}12a, SLF^{+}19, WYI^{+}16]. Merlangius [LTP^{+}18]. merlangus [LTP^{+}18]. Merluccius [BGIA19, LLBP15]. meroplankton [HRÁB^{+}14, SRER16, WLL^{+}14]. Mertensia [LGB^{+}13, MBR^{+}13]. mesh [RCU^{+}11]. Meso [vMVSW12, YRL^{+}10, vdMPM12]. Meso-fauna [vMVSW12]. meso-herbivores [YRL^{+}10]. meso-scale [vdMPM12]. mesocosm [AOW11, CPSC^{+}15, LS10a, PPB^{+}10a, PFL18, SV15, VSM11]. mesocosms [BKHW13, PMS13]. mesograzer [BS12, CRF13, LRRM^{+}11, SV15]. mesograzers [BT14, JS10, JBS11]. mesohaline [SPMN19]. mesopelagic [BST19b, Bus12, CPID13, DA11, GOO14, KSA12, KRA17, KROA19, LAK^{+}19, PAK15, SNTN15, SBB17, SCST17, SCP16, SA11, SRK11, WS10]. mesophotic [BSB^{+}11b, BLPP16, BDF^{+}11, PMK^{+}15, PS15a, RPGP^{+}15, SM17a, SL15a, WGD^{+}19, WSK^{+}16]. Mesopodopsis [CP10]. mesopredator [CJM^{+}13, CWAS17, HSWS16, VH11]. mesopredators [BPH^{+}19]. Mesoscale [BCG10, EDS14, LABG^{+}12, LGT^{+}14, PCH11, Sat10]. mesotrophic [DFS13]. Mesozooplankton [LCS^{+}10, MG12, MGB^{+}14, AHS^{+}13, BPM^{+}18a, CLL13, DVS13, EZZ^{+}12, LS10a, SOBNL13, WS10]. Meta [SCB^{+}16, WF18, CHE^{+}17, DM11, HTC17, MIIB16, MJ10, PMKL11, ZKS17, HT11b]. Meta-analyses [SCB^{+}16, PMKL11]. Meta-analysis [WF18, CHE^{+}17, DM11, HTC17, MIIB16, MJ10, ZKS17, HT11b]. metabarcoding [AAAE17, BBB^{+}15, BLM^{+}18, HBK^{+}17, HTSA16]. Metabolic [MEDS11, BPF^{+}18, BCGG19, CPHH15, HBHS^{+}18, MAS^{+}10b, PHWC13, dGCF^{+}17, SBT15]. Metabolism [IMD11, SOT15, BIL^{+}10, BBG^{+}13, BDGB19, CWEE19, CCA^{+}14, FSE^{+}17, GMC^{+}12, GLS^{+}15, HLL^{+}17, HORRLS^{+}14, JDQS11, KHD^{+}11, LKCH12, LS19, LBMZ15, ME11, MPZ10, MSM^{+}12, OPN10, PZL^{+}17, RBM14, SGD^{+}12, TYVR^{+}15, VBR18, VBR19]. metabolite [DWV^{+}12]. metabolites [GBU15]. metacommunity [DBS14, SLO19, VANB15]. Metagenetic [HT15, KHT^{+}17]. metal [CAB^{+}19a, CFC^{+}12, HBBH12, MA13a, PWNP^{+}11]. metals [FAWBF10]. metamorphic [JMB10, TGP^{+}16]. metamorphose [TH11]. metamorphosis [BGIA19, EVHA^{+}18, HJ16, LRB^{+}18, MD16b]. Metapopulation [KDJP10, OLD^{+}15, SCB15, AN11b, ABM12, BHUCMO19, DBC^{+}12b, MPF^{+}15]. metazoan [BMC^{+}15, OAGV11]. Metazooplankton [TNM^{+}11a]. meteorological [FHB^{+}12, ORW15, SGS^{+}12]. methane [ÅCAC16, PBV^{+}11]. methanol [SMND16]. methanotrophic [RCF16]. method [AM19, CPEE18, CKV12, HHG10, NMCP17, NSC^{+}19, ZB16]. methodological [WBG^{+}14]. methodology [SHMOS17]. Methods [CBLM^{+}17, GMES18, Ano17c, CS10a, DBHMG12a, DHH10, GFC^{+}14, GDPR18, HA15, JMGG13, LS14, LPI^{+}14, LYH^{+}17, MWR^{+}12, NHTS14, SWK19, SNS^{+}13]. Methyl [RJB^{+}14]. methylation [GCBB13]. metric [MWB13]. metrics [BMPH14, ETB13, EPF^{+}17, GMR^{+}19, LMCD19, PLM^{+}16, ST13, ST11a, WLP^{+}11]. Metridium [NC11]. Meuse [WH11]. Mexican [AAGGFG^{+}12, dSKFV12, RNGUR^{+}11]. Mexico [BRB11a, FNF^{+}18, ANWG18, AGH^{+}12, BRMT11, BCM^{+}14, BHT10, CWW^{+}16a, CL15a, Cra12, CATAV18, DGCH18, DPS16, DEP19, DD15, DPM15, DRWQ12, DHJ10, DAB17, GSC14, GPRR17, HGS14a, HGS14b, HWSK19, HFM^{+}15, HW11b, HSS18, HRS^{+}16, JGG^{+}15, KMV18, KWW^{+}14, LCGS^{+}19, LLCM^{+}12, LABG^{+}12, MH19, MHW^{+}12, MPB15b, MLR12, MRC^{+}13, PW16, PRBR11, RCMB18, RCF16, RSHM10, SPSG14, SAZHAAHS14, SL11b, SM17a, TKEV^{+}19, TMS14, TNN11, WDNM11, WLM12, WWKS15, WGS16, ZMS^{+}14]. Mg [PZL^{+}17, FKT15, LKB^{+}14]. Mg-calcite [FKT15, LKB^{+}14]. MHC [JCF^{+}12]. Mianzan [SGC14]. michahellis [CPF^{+}14]. micranthus [SPSG14]. Micro [CGEGC^{+}19, BC11, CTMU10, CSQRA11, FLS^{+}18, GDSPRA13, Jon11, Pil10]. Micro- [CGEGC^{+}19, Jon11]. micro-algae [CTMU10]. micro-allopatric [GDSPRA13]. micro-evolutionary [Pil10]. micro-geographical [BC11, CSQRA11]. micro-zooplankton [FLS^{+}18]. microalgae [FCR^{+}15, ROE17, WEK14]. microalgae-dominated [ROE17]. Microalgal [PPB10b, DP11, PFR11, PT13, RMHP14, SDMB15]. microbe [FBET15]. microbes [MZC^{+}10, POPH12, SS13]. Microbial [DBDK14, FT10, FL17, FAZ^{+}14, LZF^{+}14, MRB^{+}15, WA14, AC11, BLHRP^{+}19, BKHW13, CL10a, DMW^{+}16, DVSM10, ELM^{+}19, FH10, GLS^{+}15, GBTS11, HCAV10, HSR^{+}10, JRvdM15, LR15, LLCM^{+}12, MZC^{+}10, MAT^{+}11, MSM^{+}12, NTN^{+}13, SVG16, SMND16, SPH^{+}17, SS10b, SDMR19, SSX^{+}17, SLNG18, SRRS12, VQKF14, WNNH10]. microbiome [BAFP^{+}13, MPLLE17]. microbiomes [BCSH18, GRBH15]. microbubbles [RPS12]. microcephalus [FLK12]. Microchemical [CGG^{+}10b]. Microchemistry [CMK12, Coo11, HMH^{+}12, HS16, LR18, MMRC16, STM^{+}14]. microcosms [HSR^{+}10]. Microeconomic [ON15]. microenvironments [CRG^{+}16]. microfouling [WSD^{+}10]. Microgeographical [PHHL^{+}11]. Microhabitat [WH14, WvLF18, Ale13a, BB15, BTTK10, EC14, GHCC15, KBB16, RB15, dG14]. microhabitats [GNBM11, KCD13, MAK^{+}15, WPRS18]. microlayer [NTN^{+}13]. micronekton [BHT10, CWW^{+}16b]. Micronesia [BGSM13]. microorganisms [BMD^{+}11]. microphytobenthic [FNA^{+}16, GME^{+}19, SLRM^{+}15, SdBAH^{+}14]. microphytobenthos [BRMT11, FSFS18, GRCP^{+}10, GFR^{+}18, JUMS10, KTK^{+}12, LS12, OLM^{+}12, PMS13, RTD14, VUL^{+}11, WK10, dJdBdJB12]. microphytoplankton [CGH12, LGC10]. microplankton [GSTK11]. Microsatellite [HW11b, MHS17, CJS^{+}12, MIM^{+}12]. microsatellites [DSH14, VBV^{+}10]. microscopy [VBŽ^{+}17]. microstructure [ZSSK11]. microtidal [BJG^{+}11, HP11a, MPB15a]. Microtopographic [LH19]. microtopography [LBP^{+}16]. Microzooplankton [BOS16, LFL11, OSL18, CRS^{+}11, CRGT19, GLY^{+}14, LSY11, LHS^{+}14, MAB^{+}10]. mid [CWW^{+}16b, EFOM18, FMP13, LOES11, MO13, RKFD15, Sti13a, SCHO11b, BGM^{+}18, HS11a, HBB^{+}18, MVTS^{+}12, RWMP12, SHF12a, SCHO11a, TFB^{+}17, TPM19]. mid-2000s [LOES11]. mid-Atlantic [FMP13, MO13, SCHO11b, BGM^{+}18, HS11a, HBB^{+}18, MVTS^{+}12, RWMP12, SHF12a, SCHO11a, TFB^{+}17, TPM19]. mid-shelf [Sti13a]. mid-trophic [CWW^{+}16b, EFOM18]. mid-water [RKFD15]. middle [TTPM12]. midge [SLC11]. midshipman [HPB^{+}18]. midwater [KBB10b]. migrants [MSF18]. migrating [BMJ16, DH18a, JT10a, LA14, OWN^{+}15, SMD^{+}16, WNNN12]. Migration [AHØ^{+}15, HGG^{+}18, PJP^{+}11, ABK^{+}14, ABT^{+}15, ARMA15, Ano17c, BLS^{+}11, BPOS^{+}16, BSN^{+}18, BEMD11, BNS^{+}14, BQRB12, BCH^{+}18, BFN^{+}19, CMC^{+}16, CKK^{+}10, CMW^{+}14, CPS15, DHS^{+}19, DT15, EOA15, ETO^{+}10, FDVL18, FSF^{+}17, FTB^{+}17, GFS15, GVG^{+}18, GKW10, GSSM^{+}11, GSJS^{+}19, HP11a, HHP^{+}17, HMO^{+}12, HCLB18, HVD^{+}16, HSS17, JJJ16, KNY^{+}15, KCF^{+}14, KPM^{+}11, KHK^{+}17, KRdVG^{+}18, KCE14, LSY^{+}18, LET^{+}12, MBM^{+}14, MAW^{+}11, MC16, MADI13, MSG^{+}13, MKI^{+}15, MBG^{+}15, MF10, MSV^{+}11, MKII19, MJHS11, MCW^{+}17, NMS^{+}18, NKA16, OVF11, OKF^{+}18, PLGSD17, PSER17, RDHM10, SDKD13, SØA^{+}13, SMC^{+}14, SRK11, SHF^{+}12b, TWL15, TPT^{+}12, TGSM12, TTW^{+}12, Ued14, VUL^{+}11, WPML^{+}18, WSL^{+}14, WHD^{+}14, YNK18]. migrations [BPSCD17, Dan11, HSF^{+}12, JT10a, KKA12, MMB^{+}12a, SRG14, SFKQ15]. migrators [MLP19]. Migratory [HTR^{+}12, OBA^{+}18, TIS^{+}15, TIN^{+}18, ZBB^{+}11, ALE^{+}13b, CEH18, CSM^{+}19, CGM^{+}19, GFC^{+}13, GTL^{+}16, KBK^{+}10, LJM^{+}10, LBM^{+}13, LSH^{+}19b, MGM10, NJJ^{+}11, PFL13, RA10, RGWA11, RKM^{+}10, RAF^{+}14, SØK^{+}15, SMMFH16, SHD^{+}19, TR13, VCG^{+}10, VHP^{+}11, vBMH^{+}17]. miles [EPM^{+}16, SJO16]. Milewski [Ano19c]. milfoil [KMV18, MV12]. milk [SCW^{+}15b]. Millepora [ZAZS11, TPB^{+}16, TPB^{+}18]. Miller [FLC^{+}14]. mimic [GBK^{+}14, GZB16]. mimicry [GWGC16]. mimics [SAM14]. Minas [MSH^{+}14]. mineral [OAGV11]. mineralization [AHSG16]. mineralogy [LKN^{+}14, WNP^{+}14]. Mingulay [DJL^{+}12]. minimizes [HGC15]. minimum [dlÁGRB^{+}19, GE17, JLL^{+}12, RPL^{+}18, RHRB^{+}17]. mining [BRC^{+}15]. minirhizotron [JAPK16]. Minke [RCD^{+}13, AEGH11, AED^{+}12, CRL13, dB10]. Minmax [Fry13b]. minnow [MDAPH11]. minor [KBP^{+}10, ALWD15, PHHL19, SDCC12]. minutus [PSS^{+}12]. MinxEnt [DER^{+}19]. Mirounga [AGNH^{+}19, CMHH17, JLPG17, LCB^{+}16, LRO^{+}17, RDGC17]. mis [GSD10]. mis-measure [GSD10]. Misidentification [SN17]. mismatch [BDB^{+}12, DHF^{+}13, PDCH15, RGW17, RPM12, SCAM^{+}14]. Mismatches [HGW^{+}11, DB14, SBMÅ12]. Missing [WFD19, CKR^{+}11]. Mississippi [SHT^{+}13]. mitigate [BPF15, DPS16, Løk11, RHVRC11]. Mitigating [BC14]. mitigation [LL18, MMCW14, OALK14]. Mitochondrial [BS18a, DHJ^{+}17, BCC^{+}11, DBS^{+}13, SBB^{+}12a, BS18b]. mitogenome [FKP^{+}15]. mitsukurii [DEGF^{+}10]. Mixed [ACS12, BPM^{+}13, SWH^{+}18, BMV^{+}18, BCM^{+}11a, BBdG^{+}19, GDJdBSL19, HPMB11, LPSK17, MTP^{+}10, NYK^{+}13, OTK^{+}19, PRK^{+}12, dGCF^{+}17, RvOP12, SWK19, SAMW14, TPS10, WFH^{+}18, ZLGH16]. mixed-effects [MTP^{+}10]. mixed-stock [BCM^{+}11a, SWK19]. Mixing [BW16a, KLD18, AMS^{+}13, BDS18a, BDM^{+}14, BPJ^{+}15, Bre14, Fry13a, Fry13b, GED^{+}14, LPB^{+}16, MBS^{+}14, MAB12b, PMYH13, RBC^{+}14, RCG^{+}15, RGB^{+}11, RGP11, SPWK13, SSR^{+}10, SWP^{+}13, BW16b]. mixohaline [LGD10]. mixotrophic [CRR^{+}19]. mixotrophs [FGP17]. Mixotrophy [RCB^{+}10]. Miyako [WYS^{+}10]. Mlalazi [OCS14]. MMSY [VQSH14]. Mn [PWO^{+}16]. Mnemiopsis [BFH^{+}13a, BB14c, DTM^{+}12, HTHH11, KBPD10, MLC14, PDBB14, RG14, TLG14]. mobile [APC^{+}13, CHST16, FHTY19, GKC11, KCF^{+}14, NBGGST^{+}15, PBH^{+}13, SW18, SHB^{+}10a]. mobility [VBT14]. Mobula [CNJ^{+}18, PSD^{+}19]. mobulid [SRA^{+}17]. mode [Kiø16, KMSW10, OBM15, WJ12, vdGvdLvG^{+}19]. Model [AMAJ^{+}17, CFBG14, FJ11, MPB^{+}14b, NDD^{+}13, PP10, STRP18, AMR^{+}10, ADO^{+}12, AHH^{+}11, AZSKWL11, BDS18a, BLIM11, BD13, BSP^{+}10, BSB12, Bre14, DGDJ14, DEM15, DPD^{+}12, FUOG^{+}16, FHRC16, FGW^{+}15, FOJF15, FEE^{+}11, FBNJ13, GED^{+}14, GDB^{+}16, GL12, GSJS13, GCCCM16, HHH11a, HCJ^{+}11, JR15, Ji11, LNS10, LLP^{+}10, MdNW12, ML11, MPZ10, MFI11, MW16a, MHO^{+}11, NLV^{+}15, NUJ12, PO13, PDO13, RWM18, RWS^{+}15, RHS^{+}19, SDM^{+}13, SS10a, SvdMK^{+}12, STK^{+}13, SGT^{+}14, SOJ14, ST10b, SEFS15, TGRH17, TTM19, TSTP13a, VBR^{+}14, WMPM13, YChH11, YIKG10, vDJF15]. Model-based [STRP18]. Model-derived [NDD^{+}13, SOJ14]. model-simulated [SGT^{+}14]. Model-to-data [CFBG14]. Modeled [HPH^{+}18, RPP^{+}13, BSN^{+}18, SPW^{+}12]. Modeling [ADHP12, AOW12, BPOS^{+}16, CDMW12, DSGR^{+}15, DWM15, FHTY19, FKZTT18, GDP^{+}11, HSS18, KWDD10, LB15a, MPB15a, MCRB11, MOH^{+}10, PHD^{+}13, SNS^{+}13, SS12b, TCSA17, TRA^{+}14, VGGF12, BHUCMO19, BFG^{+}19, CMK12, CWGG15, CCKB^{+}15, FJD13, Gal14b, GBD^{+}12, HMM^{+}15, HSS^{+}19, KM10a, KCLP16, LMV^{+}19, MPKO11, MKI^{+}15, MPF^{+}15, Pro17, SHKR15, SM17a, SWH^{+}11, TMT^{+}12, WSS^{+}16]. modelled [FL10, OLD^{+}15, WEK14]. Modelling [ASF13, BMEL11, CHFS12, GLB^{+}10, HDWF10, PPS^{+}13, PCS^{+}15, PMM^{+}11, RTMJ11, RS16, RGB^{+}11, RGP11, TBGW^{+}11, VCL^{+}15, vdWTWS15, AFSL^{+}15, AMAJ^{+}17, CLGCA^{+}19, CBYH13, DC17, FYG^{+}18, GSC14, GMBL11, GDE17, HSH12, HMBP18, HT11b, KTR^{+}16, LGS10, LMC14, LRK^{+}15, MC16, MTP^{+}10, MHW^{+}12, MIV^{+}10, MLP^{+}14, PLG^{+}16, PSHG16, PSL^{+}12, PEB^{+}14, RWR^{+}19, RCK^{+}11, SATS13, SSB^{+}16, SSS^{+}10, SLSR10, SSFB11, SPP18, TGI^{+}11, WWB14, WHH^{+}13, ZPPM17]. models [AAE14, ARB^{+}12, AMS^{+}13, ANWG18, BLV^{+}19, Bea11a, BFF^{+}10, BPJ^{+}15, BDH^{+}17, DSG15, DLC^{+}13, DP14b, FOL12, GGF^{+}12a, GDA^{+}18, GMES18, HPP^{+}10, KHS10a, LFL11, LHB^{+}15, ME11, MSG^{+}13, MHNL14, MBS^{+}14, MKG10, PLM^{+}16, PCW^{+}12, PGVF14, Rai12, RFM^{+}15, RZPH14, RSG^{+}16, RLH^{+}19, SRG14, SPM^{+}13, SWP^{+}13, SFM^{+}19, ST10b, SHT^{+}16, SAMW14, SJL13, TW10, USR^{+}16, VF13, WM10, WRP^{+}17, WCP^{+}14, WFH^{+}18, ZBC10]. Moderate [BRB17, BSN^{+}18, CPSC^{+}18, GCP^{+}17, Hv14, SHPS^{+}16]. moderately [CSC13a]. modern [TKL^{+}11, WWAL13]. modes [CUCS17, RMKR13]. Modest [LMH12]. modifications [EHR12, PBM^{+}10]. Modified [MUC12, BH12b, CG10, EOMS15, KLKM12]. modifies [DVS13, FJW14, GME^{+}19, LH11, SGP13, SLO19]. modify [DPO^{+}17, PT13, SLT^{+}19]. modifying [JT10b]. Modiolus [KAP^{+}12]. modular [CGG10a]. modulate [CESC14, FTF15, KPr^{+}18]. modulated [LMBM18, NLV^{+}15]. modulates [PCA^{+}15]. Modulation [GFR^{+}18, WBLRN18]. modulations [CL15b]. moerchi [DMC^{+}17]. Mohn [SLRS13]. mole [CGD10, CD16a]. Molecular [ATF12, FJC10, HHNI17, HOA14, RSS^{+}16, DHB18, FTvdK^{+}12, HLN^{+}19, JRK13, LJM^{+}12, MVCD14, OHST19, SSBA12, TF12, ZGH^{+}17, TCWC13]. Molecular-based [HHNI17]. molecule [Pör12]. mollissima [GDH^{+}19]. mollusc [COK10, MPW12, dMA18]. Mollusca [MEDS11, SSBC11, SHMB15]. molluscan [GKBEF18]. molluscivores [VHBG15]. Molluscs [CBD14, BMT17, Cha13, LMG12]. molluskan [EN11]. mollusks [SLW15]. molt [DMMW12, GGMGR12, RBEGG13]. molting [DMMW12, WRW^{+}16b]. Monachus [MJTC^{+}12]. monitor [BBBF15].

Monitoring [KBG12, KHM^{+}12, ASS^{+}13, BMSB13, BT16, BRLM17, CPMR11, GNJJ^{+}17, GSJS^{+}19, KSHS19, KOLDP17, KLBT^{+}16, KCS^{+}17, LRK^{+}15, MGM^{+}18, MWB13, MBKM17, SWH^{+}11, TS11b, USR^{+}16, WSLM13]. monk [BHP12, MJTC^{+}12, WRP^{+}17]. monoceros [LHJ11, WOHJ^{+}15]. monodi [DMHS^{+}14]. Monodon [LHJ11, WOHJ^{+}15, GVR^{+}14, SSBC11, SHMB15]. monogamous [FSF^{+}17]. monospecific [ZLGH16]. monsoonal [PA15]. montagui [FCV^{+}16, MASMH19]. Montastraea [DZH^{+}11, LP12]. Monterey [BSG^{+}10, BHN^{+}12, HFR^{+}18, SPH^{+}17]. month [HBF^{+}10]. Montipora [CRWG16, EES15, PGG12, WMS^{+}11]. moon [FSM^{+}14, PIGK16, SNT^{+}18]. Moorea [ELA10, TPA^{+}13]. Morakot [CGH12]. moray [LÁBFÁF^{+}10, PSdA^{+}17, RBSL11]. morbillivirus [WVM^{+}14]. mordax [BDT^{+}11, WM10]. morhua [ALNB17, ACG12, BSN^{+}18, BPWF14, DSG15, FWPP12, HDB^{+}10, JBS12, LNP12, MFG^{+}13, MMNR^{+}12, MRC10b, NJJ^{+}11, OMSP10, PLA^{+}12, PHHL^{+}11, RAN^{+}10, RJTS12, SGL^{+}17, SMK^{+}10, SBH^{+}18, SJB^{+}15, SNT^{+}16, WBW^{+}14, WFP15]. Mori [Ano12c]. Morone [CPS15]. Morpho [GH11]. Morpho-functional [GH11]. morphodynamic [CGD10]. Morphological [CG17, FGOS16, BBB^{+}15, CFMFFJ18, GDSPRA13, HBI^{+}18, KHT^{+}17, LALÁC11, MJO^{+}13, PMM^{+}15, PFPP18, PW19b, SSBC11, dlSOV^{+}16]. morphologically [TMP^{+}11]. morphologies [NMC^{+}16, SJHH^{+}17, VE18]. morphology [BJF17, BRG^{+}14, Ber12a, BCPC18, CHDH^{+}19, CG17, CG11, dSKFV12, KP18, LCTD^{+}15, LMB^{+}19, MGFW13, MH11b, PWCvO^{+}14, RHMR18, SNTN15, SBA12, SWBMJ17, SHMB15, YMCV10, ZAMB10, Wir12]. morphology-mediated [Ber12a]. morphometric [CGG10a, TGSM12]. morphometrics [HPL12]. morphospecies [TKS^{+}14]. morphotypes [GDSPRA13, HWE^{+}12]. mortalities [ECC^{+}15]. Mortality [HS11a, KHH18, dQSC13, ARB^{+}12, AHØ^{+}15, AWJ12, BA13a, BM10a, BWLP17, BSB^{+}11a, BDP10, BRB11a, BS14b, BLM10, CAD^{+}10, CSM12, CMW^{+}14, CLMM^{+}14, CGBGS18, DSC13, DZL16, ET11, EBH17, FPC^{+}12, FFS^{+}15, FSY^{+}10b, GZYS11, GJ10, HGS14a, HGS14b, HGS15, HSZR10, HS11b, IW15, Ing16, JG13, JH11, JGCRS12, JPP^{+}17, LBH18, LAM10, LBVH13, LE13, MUOS13, NBM12, PMB^{+}10, PBM^{+}10, PRP^{+}16, RWP^{+}14, RUPT^{+}11, RUS^{+}13, RSSvV19, SB15a, SVG16, SFLG10, SNVH19, SCHO11a, SCHO11b, SB15b, TBMG17, TPB^{+}18, TPA^{+}13, WB12, WWW^{+}13, vdMPM12]. mortem [NLLS15]. Morus [CVW^{+}11, SBB^{+}12b]. mosaic [BAL14b, PB18]. mosaics [CSM^{+}19]. mossambicus [DPC13]. mossbacks [MSF18]. Mossel [MMB^{+}19]. most [BMÁN^{+}18, BG12, FTO^{+}15, RUD^{+}16, TB14a]. mostly [HNMK15]. moth [CMP^{+}11]. mother [FBA11, TFW^{+}12]. mother-calf [FBA11]. mothers [Nor13]. motile [BRA^{+}10, ETB^{+}11, SHDR13]. motility [EKE16, FSFS18, SM17b]. motility-physiology [FSFS18]. motion [RL11]. motorboat [MJS17]. Moult [GWH^{+}18, WCRCR17]. moulting [BPG^{+}13, HT11a]. mounted [KW17]. mouse [PCA^{+}15]. mouth [BP12b, CPP10, MVVGAV10, OCSW16]. mouth-breaching [BP12b]. mouthpart [SNTN15]. move [MS11]. Movement [BWM^{+}15, BT13, BBKP10, BK14, CRAP^{+}12, CWK^{+}10, CHB^{+}13, CNW^{+}12, DB11b, GRL^{+}14, LHT^{+}16, NCW^{+}19, PSD^{+}19, PPDW16, SL11b, TFB^{+}17, AvBBL^{+}16, AH13, Ano17c, AMAJ^{+}17, BN16, BCH^{+}18, CBLM^{+}17, CPA^{+}11, CLS^{+}16, DCL^{+}17, DDB^{+}14, DAC^{+}12, DBC^{+}12b, FGTS16, FCFD15, IHD^{+}16, JE10, KCF^{+}14, KWH^{+}18, KHS^{+}11, LB15b, LA12b, MCPC19, MAHK12, NHTS14, OALK14, PLS^{+}10, RWS^{+}15, RBD^{+}14, RMB18, SRM^{+}15, SMF^{+}11, SPBS14, TMH^{+}18, TRGF18, TLAT13, VHHGF11, WGS16, WLP^{+}10, YOWK17, YSH^{+}16, ZHM^{+}10]. Movements [CWW^{+}16a, GS11, JRW^{+}14, LCB10, MR16, PMK^{+}15, PWA^{+}17, RMNP17, SBCT17, BT10, BAL14b, BDM^{+}13, Ben16, BBW^{+}17, BRLM17, BHL^{+}19, CTZ^{+}16, CPM^{+}17, DJJ^{+}16, DBHMG12b, FA11, FWL15, GLJ^{+}18, GMHS19, GW15, HHP^{+}17, HPN11, JUK^{+}10, LMLB12, LCG^{+}17, LDP^{+}17, MASW^{+}11, MRLD19, NCD^{+}12, NWP^{+}11, NBGS^{+}12, NTS^{+}18, QHN^{+}10, RHB13, SKMS^{+}18, SFMB10, SGB^{+}10, TIS^{+}15, TIN^{+}18, TEAO^{+}14, TS11b, VWW^{+}16, ZMGS17, vBMH^{+}17]. Moving [CAC^{+}12, SESS^{+}18, BCW19, BDvGW16, TCWC13]. Mozambique [DPC13]. MPA [AAE14, AMR^{+}10, DMH15, GRL^{+}14, LMC14, PSdA^{+}17, WD14]. MPAs [DMH15, GS11, TMH^{+}18]. Mpenjati [OCS14]. Mpi [FSR10]. mtDNA [PRK^{+}12]. MTI [KMP14]. much [WNA^{+}15]. mucronata [RBEGG13]. mucus [DZH^{+}11]. mud [BBMM14, HA18, HW13, LKL^{+}19, MYS^{+}16, PBV^{+}11, POH16, RDC^{+}12]. mud-bottom [POH16]. muddy [PDGR^{+}13, RTD14]. mudflat [BDM^{+}14, DB11b, DBC^{+}12b, SP12, ZBY^{+}14]. mudflats [BVC^{+}16, DFDC18, FSH^{+}16, GME^{+}19, GGF^{+}14, PFB^{+}13, ZOM^{+}10]. muelleri [FGOS16, PAK15, SA11, SRK11]. Muggiaea [BLHL15]. Mugil [DBS^{+}13]. Mugilidae [KBP^{+}10]. mullet [KBP^{+}10]. Mullica [NGA18]. Multi [BADH^{+}12, BBW^{+}17, EMH^{+}19, GHNF10, HDB^{+}10, MMH^{+}17, MMB^{+}12a, MBG^{+}15, MPRPM^{+}18, OS12, Sma13, ADRA19, BN16, BMR17, DFD^{+}19, DP14a, DKF^{+}15, GMT^{+}17, GSJS^{+}19, HMKF11, KH10, LHC^{+}15, ORW15, RCD^{+}13, RK15, RCU^{+}11, SLS11, SMF^{+}11, TFW^{+}12, VQSH14]. Multi-approach [MMH^{+}17]. multi-biomarker [GMT^{+}17]. Multi-colony [BBW^{+}17, DKF^{+}15]. Multi-decadal [BADH^{+}12, ORW15]. Multi-disciplinary [HDB^{+}10]. Multi-element [MMB^{+}12a]. Multi-isotopic [MPRPM^{+}18]. multi-mesh [RCU^{+}11]. Multi-population [MBG^{+}15]. Multi-scale [GHNF10, Sma13, HMKF11, SLS11]. multi-seasonal [DFD^{+}19]. multi-sensor [TFW^{+}12]. multi-site [ADRA19]. multi-species [BN16, DP14a, KH10, LHC^{+}15, SMF^{+}11, VQSH14]. multi-stress [BMR17]. Multi-tissue [EMH^{+}19]. multi-trophic [RK15]. Multi-year [OS12, GSJS^{+}19, RCD^{+}13]. multibeam [MIV^{+}10]. multidirectional [BAL14b]. Multifocal [ZAZS11, Car12, ZAZS12]. multimarker [DMC^{+}17]. Multinational [MBCB^{+}19]. Multiple [AAJJ^{+}19, BPZ^{+}14, GM10, PH10, RHE14, RBM14, SWWK13, SWNSF14, TMH^{+}13, TFH^{+}19, WMH17, dGR11, ADL^{+}18, BA12, BH12a, BMPC15, BPB11, BDB^{+}12, CB14, CTMU10, DEGF^{+}10, DAWS19, DBS^{+}16, GBC^{+}16, HWSK19, HHI^{+}14, HPMZ11, HB12, HL14, INvdB^{+}13, JPI^{+}19, JC15, KPB16, LMG12, MM12, MW16a, MGOL19, MLR12, OMW^{+}10, PHI^{+}12, PL15, PKW^{+}11, PTP^{+}19, SE11b, SHD^{+}19, Vig15, WPK10, YKT^{+}14, dSdSdCO^{+}11]. multiple-batch-spawning [YKT^{+}14]. Multiple-colony [TMH^{+}13]. multiple-tissue [HB12]. multiplex [WRB^{+}15]. Multiscale [REG^{+}11, SLPP^{+}14, VSG^{+}11]. multispecies [AAE14, HPM14, LLP^{+}10, TSTP13a]. Multivariate [MHNL14, YIKG10, AFH11, CSPA10, JGCRS12, MPG^{+}12, RFRD10]. Mummichog [LFD13]. Munguia [LGL11]. Munida [BHP11]. Muraenidae [RBSL11]. muricid [BFL^{+}19]. Murphy [CPM^{+}17]. murre [YOWK17]. murrelet [AMM^{+}12]. murres [BBKH^{+}11, BKAM12, CD14, EMTB^{+}12, EKOÅ13, KTP^{+}18, RDHM10, SG12, TMH^{+}13]. muscle [EAMT10]. Mussel [BMF^{+}12, SCL^{+}19, AFSL^{+}15, AL12b, AY10, BH12a, CGOC11, DEWK15, EKE16, FB14, GFL^{+}17, GBE^{+}17, GAO14, GLW^{+}17, GPB^{+}15, HFLI11, JRvdM15, JSS^{+}11, Kat16, KAP^{+}12, KW12, LBP^{+}19, MLB18, PAG19, PMYH13, PJMP10, RFBH^{+}11, RCF16, RCBE15, RT17, SvdMK^{+}12, SPWK13, TM15, TPBM13, TBS^{+}17, VBT14, VNK^{+}12, WWG13a, WBLRN18, YAFV12, ZGH^{+}17, vBJV^{+}15]. mussels [BH10, CGH^{+}16, FC13, GPA^{+}16, GCBS19, KR16, KOPW15, LMGJ13, MN18, NLV^{+}15, NLLS15, PBW^{+}11, RPRN^{+}16, RJL^{+}11, RLTL14, SSS19, SES17, SDT15, TMS^{+}16, VLM14, WL11, ZBP18, ZZG^{+}19, ZFZ16, vdMPM^{+}10, vdMPM12]. Mussismilia [BVM^{+}19]. Mustelus [MdSW^{+}16, SCB15]. muticum [KHG^{+}19]. mutualism [FSE^{+}16, FSE^{+}17, LGJH12]. mutualisms [FBET15]. mutualist [Lyo14]. Mutualistic [GFF16, BBG^{+}11, KBS16, SPH^{+}13]. mutualists [AAvS^{+}13, Ber12a, Tut17]. Mya [CLG19a, CBC^{+}16, GSO^{+}18]. Mycale [LP12]. mycosporine [EMG^{+}15, OCMT13]. mycosporine-like [EMG^{+}15, OCMT13]. Mycteroperca [TEAO^{+}14]. myctophid [SCW^{+}15a]. Myctophidae [BST19b, STYT10]. mydas [BHTF11, CLS^{+}16, GDB^{+}16, HNK^{+}18, HSMB17, JBL^{+}16, LABM17, LKL^{+}17, MLM^{+}12, PRK^{+}12, SMBH15, TGH15, TSA^{+}18, WBB^{+}10, WHFM17, vdMHW^{+}10]. mykiss [HHP^{+}17]. Myoxocephalus [TB12]. Myriophyllum [KMV18]. Myripristis [MTG^{+}11]. Mysid [GHGG11, CP10, LA14]. mysids [RHVRC11]. mysticetus [FDLL10, PLF^{+}12a]. Mytella [GFL^{+}17]. mytilid [BA13a, RCB^{+}10, ZBP18]. Mytilus [BC11, BSH13, BRH13, CGS^{+}19, EKE16, FRRÁS^{+}12, FRGI15, FAZ^{+}14, GCBS19, GLW^{+}17, JRvdM15, JSS^{+}11, JG13, LBP^{+}19, LMGJ13, LKS12, LIT13, PLFR^{+}11, RPLT10, RJL^{+}11, RCBE15, SvdMK^{+}12, SS11a, SHH10, SDT15, SSAB^{+}15, TJS17, TTS17, ZFC^{+}13, vBJV^{+}15, vdMPM^{+}10].

N [BFH13b, KJW^{+}15, MFH^{+}11, VTZ^{+}13, ACG12, AM19, AGNH^{+}19, CPEE18, CLGCA^{+}19, CCJR14, CBR19, CSC^{+}11, CHLR16, DC15, DTVH10, DB11a, FGP17, FSE^{+}17, GRG^{+}12, GFC^{+}13, GCS^{+}18, GCS^{+}19, GM13b, HHS^{+}12a, HU13, KV13, LMH12, MWP10, MFI11, MPB13, MHS14, MFS14, NMPM15, OAGV11, OFT^{+}11, PCW^{+}12, RWM^{+}14, RP12, RET^{+}19, ROE17, RLAGL12, RCVG10, SE11a, SML^{+}14, TKV^{+}19, UDF^{+}18, VRSS18, VSM11, VB14, VWEP18, YSC^{+}15, WAP10]. N- [MFH^{+}11]. N-enriched [AGNH^{+}19]. N-limited [BFH13b, LMH12]. N. [EBOM11]. naiveté [ACL^{+}16]. Namibia [BNAK11, DER^{+}19, TBB^{+}19]. nanism [SKGS11]. nano [TDM16]. nano-sized [TDM16]. nanoflagellate [SMK^{+}18]. nanoflagellate-dominated [SMK^{+}18]. nanoflagellates [UMF^{+}10]. Narragansett [CR16, SBE^{+}19]. narrow [BSB12]. narwhal [WOHJ^{+}15]. narwhals [LHJ11]. nassariid [LCL^{+}11]. Nassau [SDST16]. natal [BSC^{+}13, BDT^{+}11, Coo11, LSFQ19, SWW11, DFS13, SWY^{+}10]. Natator [WTP^{+}18]. National [HSF^{+}12, LRH14]. Native [BB19, EF16, IW15, MG14, RBS16, WO11, WMB^{+}14, Alb15, AHPKS17, BBdG^{+}19, BOU13, CGOC11, CBK^{+}12, CMH14, CWAS17, DdR10, DFLC17, DT15, DCT^{+}10, GFL^{+}17, GKHS16, GGF^{+}14, GM13b, HKT10, HFM^{+}15, HCG15, HTHH11, HTC17, Ing16, JL12, JMB^{+}18, KHG^{+}19, KA16, KBS16, KBC^{+}12, LHU^{+}18, LCFR10, MEK11, MD11, MTK^{+}18, NAY^{+}11, NdBKR13, NLLS15, ORM^{+}13, OHP18, PHSBO16, PRA^{+}19, RCB12, SO15, SJHH^{+}17, TWM^{+}15, TS12, TdRH17, WDH^{+}12, WPB^{+}14, ZOM^{+}10]. Native-exotic [BB19]. Natural [BWV^{+}12, JWHR11, MHSG18, POX^{+}17, AJF^{+}12, BSB^{+}11a, BCM19, CRGT19, CKC^{+}10, DH18a, DPM15, ES12b, FMG^{+}18, FL17, GTP^{+}17, GPA^{+}16, GWM^{+}15, GWM^{+}16, HKJ17, IMU17, JTTB^{+}18, LBH18, LRS^{+}12, LC12, LL18, LLF^{+}18, MLB19, MW16b, MUOS13, MU13, MLSK^{+}15, NUO^{+}19, OMW^{+}10, PG15, PCNB16, RPS12, RB17, SOB16, SMOH19, SJC10, SGTD14, SKW^{+}14, SHT^{+}16, SPP17, SB15b, TPC^{+}16, TPSS12, VTSH^{+}19, WB12, WJWC16, YTM^{+}12, ZIT^{+}17, vDBH^{+}15]. natural-abundance [NUO^{+}19]. naturalized [DWL19]. naturally [JTTP19, LMZ16, TS11a, WSL^{+}14]. nature [MR16]. naupliar [ACA^{+}10b, ACA^{+}10a, DSM11a]. nauplii [BAT14, JSLG17]. nauplius [SM17b]. Naval [SWK^{+}16]. NC [RWR^{+}15]. NCMA [ST14]. Near [BZC^{+}12, CLMB13a, LS10b, PMYH13, BPP^{+}18, BSC^{+}16, BPLT14, CBKM16, CGH11, CLMB13b, CPHH15, FHG^{+}14, JMK12, KCB18, KCW^{+}17, MGO^{+}12, MWC18, MHAP18, NHS^{+}17, SDM^{+}13, SV13, SOC13, SDF^{+}19, TKB^{+}13]. Near-bed [PMYH13]. Near-bottom [LS10b, KCW^{+}17]. near-colony [MWC18]. Near-future [CLMB13a, KCB18]. near-hatch [MGO^{+}12]. near-shore [BSC^{+}16, BPLT14, CBKM16, FHG^{+}14]. near-surface [CPHH15]. near-urban [MHAP18]. Nearshore [RRG^{+}11, RBS^{+}17, AAGGG19, BSF17, BLV^{+}19, DCB^{+}16, FPR18, FHE^{+}15, GMHS19, HRP18, KHS10a, KHS^{+}11, LRHP18, LMTD16, MLR17, MF10, MCRB11, MRC^{+}19, MCT16, OGCHS18, PPCH17, RUP11b, RUP11a, SMF^{+}17, SLO^{+}17, SKN^{+}17, VH11, WLL^{+}14, WTP^{+}18]. nebouxii [ACAD12]. nebulosus [CLB19]. necessarily [GNHLN11, MDS14]. Necora [SHC^{+}10]. need [KPLD^{+}12, MS11, PGWS10, RL18]. needed [JDB^{+}10, PvdMLC17]. needs [BBD^{+}14, GLJ^{+}18, SGGL15]. Negaprion [CBL^{+}10b, GGD^{+}11, GGF^{+}12b, KCF^{+}14, NHG10]. Negative [VDG^{+}15, EF16, GLHM^{+}18, HYC^{+}14, KMV18, LCD11, RZH16a, TVH^{+}10, VMWU15]. negatively [MRL^{+}10, PSPR18, WWO^{+}15]. neighboring [CPF^{+}13]. neighbouring [TMH^{+}18]. neither [NHPW18]. Nekton [TT16, CR11, HHH^{+}14, RMB18, SLO^{+}17, SMR10]. Nektonic [WFBM14, FRTB14]. nelsoni [CB11]. Nematocyst [CUCS17]. Nematode [LRB^{+}12a, BVS^{+}11, DVV^{+}18, KdWV^{+}18, KHG^{+}19, MVS^{+}10, dSM11b]. nematodes [HSR^{+}10, IVD^{+}10, LRB^{+}12b, PBV^{+}11, dSM11b]. Nematostella [RCE^{+}13]. Nemopilema [DWD^{+}18, FLS^{+}18, FWS^{+}18]. Neocalanus [YOST11]. neon [IMS^{+}11, LDAE12, VSW^{+}15]. neonate [MWR^{+}12, WTP^{+}18]. Neotrypaea [VPW^{+}12]. Nephrops [MLP19, OLD^{+}15, PPSR15]. NEPTUNE [RJ12]. nereid [BCK15]. Nereididae [CML^{+}12]. Nereis [CRJ13, KHD^{+}11]. neritic [AGH^{+}12, LCBK^{+}13]. neritina [DHG11]. neritoides [GDSPRA13]. nest [MS18b]. nest-guarding [MS18b]. Nested [BBMH13, BSP^{+}10]. nesting [BGM^{+}17, CdTK^{+}16, KEH10, MC12a, POR^{+}15, PPM^{+}15, PAC^{+}15, PRM17, RAKB^{+}12, RTDUM16, RMNP17, SMH^{+}12, SSB^{+}11, TMM^{+}19, TBH10, WCP^{+}14, vdMHW^{+}10]. nestlings [LSJ17]. nests [RPE^{+}13]. Net [ACP^{+}16, GHD^{+}10, LBCL10, DZDB10, EH13, GUC^{+}15, NHTS14, PNPF14, PR11, SGTD14, STP10, WFS10]. Netherlands [EKR^{+}10, JRvdM15]. nets [ACP^{+}16]. network [HLS^{+}15b, HPM14, LHT^{+}16, ME11, RFM^{+}15, SKB15, VC14]. networks [BAFB12, KDJP10, LCG^{+}17, WBHL10]. neumayeri [MBvO^{+}18]. neural [ME11]. neurophysiology [Ued14]. Neuse [CR14]. Neutral [ITT14, MPW12]. Newfoundland [BWS^{+}12, CD19, CD16b, GDE17, LBR^{+}19, MDLM17, SGM^{+}14, WRDF12]. newly [ADvdM13, GPM^{+}13, MEK11, WKX^{+}19]. Newport [RWR^{+}15]. next [CPR^{+}14]. NH [DB11a]. Niche [DW12, EPF^{+}17, LDT^{+}13, MZF^{+}19, RCP^{+}16, WSHY16, YMD^{+}10, ZRBV15, AVB13, AAJJ^{+}19, BFG^{+}19, BJH13, BH15, BADQ^{+}15, BGIA19, BLJP14, BGS^{+}16, CLGCA^{+}19, CBT13, CXdG^{+}17, DDP19, DPD^{+}12, DGFH10, FYG^{+}18, FCH^{+}10, FPN17, GLR^{+}14, GSC14, GCR^{+}17, GFR^{+}18, GRCH^{+}18, GMR^{+}19, HSCA17, JCRC10, KSBM^{+}11, LCGS^{+}19, MTS^{+}17, NMPM15, PRLR^{+}18, PSL^{+}12, PSS19, PTP^{+}15, RP12, RSW^{+}16, RWO13, RVPG17, RAN^{+}10, RPR16, SVAV17, SNTN15, TKEV^{+}19, VH11, WK18, WS11]. niche-based [AAJJ^{+}19]. niches [BPH^{+}19, LEKT12, LLP^{+}10, MBR^{+}18, NLK^{+}19, PBCA10, WVD^{+}17]. Nielsen [DL13]. nifH [BMD^{+}11]. niger [UHG12]. night [DGC12, HCLB18, MBC18, SBH^{+}19, SLC11, UAK^{+}14]. nigra [GKE^{+}19]. nigricans [CNSW16, FHTY19]. nigripes [LS16]. nine [JJP^{+}13, dlSBO^{+}12]. Ningaloo [JBD10, MVP^{+}16, MMC^{+}18, OPW^{+}11, TKS^{+}14]. Niño [ACAD12, GFGGDSM12, GEA^{+}14, RCP^{+}13]. nitrate [DEJN10, ETAE13, FF12, HKS^{+}12, HTB^{+}17, HMS^{+}12, POKGRP10, SSX^{+}17, TGM^{+}17, XLXS17]. nitrates [MJW19]. Nitrogen [BVGH10, CM12, DPH14, GKHM13, GM13b, HLL^{+}17, MKMZ11, PBC^{+}15, RHF19, SOB16, WRW^{+}16a, ZCC^{+}12, AFC^{+}16, AFS19, AEM^{+}18, BST^{+}10, BVP11, BCK15, BKM^{+}18, BCB^{+}17, CLG19a, lLChTC11, CEW15, DTVH10, DB11a, DWM15, EVB^{+}16, ESH^{+}10, FW11, FF12, FSE^{+}16, FBNJ13, GM10, HHS^{+}12a, HCAV10, HN13, HTCH^{+}18, HLS^{+}15b, HLB^{+}16, HvEBC12, HP14, JS12b, LG18, LBCL10, MBSG11, áNGGS11, OFDG14, OCSW16, PMS13, RET^{+}19, Rei12, ROE17, SBS^{+}13, SE11a, SGP13, SGTP16, SLNG18, SCW^{+}15b, SLR^{+}14, VARO12, VKH10, VWEP18, WBLS^{+}11, WADKT13, WP10, YSC^{+}15, ZZG^{+}19]. nitrogen-fixing [HvEBC12]. nitrogen-removal [ZZG^{+}19]. nitrogenase [SS10b]. Nitrous [HSdBS10]. nitzschia [GHT^{+}15a, HOA14, LBM^{+}18]. No [BPC17, Chr17, ENW^{+}14, GMR13b, JDD^{+}13, PMEZ15a, ATO^{+}14, BMV^{+}17, BHN^{+}14, BKTD10, CLMB13a, DDB^{+}14, FSM^{+}14, PSdA^{+}17, RL18, RMRA15, SWNSF14, SWPK13, YAFV12, YKG^{+}15, vdMPM12, DB11a, LP12]. no-take [BMV^{+}17, DDB^{+}14, PSdA^{+}17, RMRA15]. nobilis [Kat16, SCSC16]. Noctiluca [ZLCW15, ZLGH16, LFE^{+}14, PTF^{+}14, TFM^{+}18]. nocturnal [BBG12, DP16, HHS10, MC12a, PAK15, PIGK16]. nodosa [BVL10, OD15, SSL^{+}14, dlSGS17]. nodosum [BAVSP11, SCRL11]. Nodularia [HvEBC12]. noise [BGM^{+}17, CMR12, DTC^{+}17, FMG^{+}18, HSR^{+}13, MJS17, RTMJ11, RCBE15]. noltei [BDM^{+}14, DDB^{+}16, JSBP15, VBGO^{+}16]. noltii [HBB^{+}12, OMD11, VBJR^{+}13, dlSBB^{+}10]. nomurai [DWD^{+}18, FLS^{+}18, FWS^{+}18]. Non [AS15, DGA^{+}15, GGF^{+}14, HMP15, MWBMH13, RP13, SBE^{+}14, SHT^{+}18, TMM15, vdPLdV^{+}11, vdSRT^{+}18, AHPKS17, AdCM^{+}19, BSY13, BGBD^{+}13, BS19b, CGOC11, CD19, CLW^{+}11, CBYH13, DdR10, DP11, DT15, GKHS16, GAO14, GMK18, GG12, GWM^{+}15, GM13b, HCG15, HTC17, HGPM10, JDX^{+}17, KBS16, KPr^{+}18, LCL^{+}17, LD17b, MDLM17, MUC12, MC16, MHS17, MD11, NMCP17, NdBKR13, NIS14, RGEA^{+}15, RCP^{+}16, REAN^{+}14, RCB12, RRH^{+}16, RGCT^{+}10, RKM^{+}10, RAB^{+}15, SGMG^{+}14, SMBC18, TTT^{+}15, VNN10, VM10, VBGO^{+}16, ZOM^{+}10, vBMH^{+}17]. non-bleached [HGPM10]. non-bloom [NMCP17]. Non-breeding [SHT^{+}18, CD19, GMK18, KPr^{+}18, MC16, RGEA^{+}15, RCP^{+}16, REAN^{+}14, RKM^{+}10, TTT^{+}15, vBMH^{+}17]. non-calcifying [BGBD^{+}13, JDX^{+}17]. non-chaotic [MHS17]. non-competitive [DP11]. Non-consumptive [DGA^{+}15, GG12]. Non-contact [TMM15]. Non-equilibrium [RP13, CLW^{+}11]. non-feeding [BS19b]. non-indigenous [CGOC11, GAO14, LD17b, MDLM17, MUC12, VM10]. non-invasive [RAB^{+}15]. Non-lethal [MWBMH13, AdCM^{+}19, SMBC18]. Non-linear [HMP15, CBYH13, GWM^{+}15, RRH^{+}16, VBGO^{+}16]. non-Mendelian [RGCT^{+}10]. Non-native [GGF^{+}14, AHPKS17, DdR10, DT15, GKHS16, GM13b, HCG15, HTC17, KBS16, MD11, NdBKR13, RCB12, ZOM^{+}10]. non-nutrient-limited [SGMG^{+}14]. Non-photochemical [vdPLdV^{+}11]. non-regenerating [BSY13]. non-standard [LCL^{+}17]. Non-stationary [SBE^{+}14, vdSRT^{+}18]. non-tidal [VNN10]. Non-trophic [AS15]. non-upwelling [NIS14]. nonbreeding [HMO^{+}12, MPS^{+}10]. noncalcifying [HSB12]. nonlethal [LHB19]. Nonlinear [LFG^{+}12, DLC^{+}13, KGJ^{+}16]. nonphotochemical [MAB12b]. nor [NHPW18]. Nordic [HMH^{+}18]. normal [MBB^{+}12b]. normalization [LMR^{+}10]. Normalized [QGM^{+}14]. norms [SAM18]. Noronha [PFVFF^{+}15]. North [BDBN11, CR14, SdBAH^{+}14, WMB^{+}14, BCC^{+}11, BAC^{+}11, CBR^{+}18, DAL^{+}10, GKW10, HVD^{+}16, RWMP12, ADHP12, AFLM12, AF14, BSC^{+}13, BLE^{+}13, BWLP17, BHE14, BJ14, BJS^{+}11, BDB^{+}12, BWN^{+}14, CJM^{+}13, CPTL11, CPID13, CD13, CWW^{+}16b, CDR12, CSPA10, CEK^{+}16, CBC^{+}16, CBL^{+}10c, CGG^{+}10b, CHLR16, DSG15, DA11, DHH10, DCHI12, DBDK14, ENO^{+}11, FSD^{+}14, FSP^{+}17, FTM^{+}13, FEE^{+}11, Fuj16, Gla10, GWH^{+}18, GLM^{+}19, HB13, HCJ^{+}18, HFR^{+}13, HGB^{+}13, HSA^{+}15, HPH^{+}18, HVD^{+}16, IMS^{+}11, JMK12, KTNO19, KLD18, KHM^{+}12, KKO^{+}14, KDB^{+}10, KCE14, LTP^{+}18, LRR^{+}10, LJM^{+}12, LKLC11, LBC^{+}19, LPB^{+}16, MMH^{+}17, MASMH19, MRE14, MBR^{+}18, MWZ^{+}16, MVS^{+}10, MGGSP15, MJA^{+}18, MDR16, MCTHGV12, NdBKR13, NWHD15, NTS^{+}18, OTK^{+}19, OMT^{+}17, OSD^{+}13, PXG^{+}10, PSC^{+}13a, PDT^{+}11, PAP^{+}12, RNB^{+}12, RCK^{+}11, RKC16, RHvD19, RSP^{+}16, RSSvV19]. North [SBB^{+}12a, SSB^{+}16, SBCT17, STS^{+}18, SSvD^{+}15, SG18b, SHD^{+}19, SGB^{+}10, SH17, TGRH17, TTT^{+}15, TNB^{+}15, TVM16, UDF^{+}18, VWW^{+}16, VCL^{+}15, VHP^{+}11, WHN^{+}18, WBLS^{+}11, WHH^{+}13, WRB^{+}15, WN16, WWB^{+}17, WGGM11, ZBK^{+}15, dB10, tHHR10, vDJF15]. North-Central [CSPA10]. north-south [DAL^{+}10]. North-West [CBL^{+}10c]. north-western [BAC^{+}11, CBR^{+}18]. Northeast [BL10, CAB^{+}19a, CD16b, GWT18, GCM10, HDB^{+}10, JBT^{+}16, RGDNGS11, SBE^{+}14, TNN11, WHJ^{+}12, CBF^{+}13, Hat13, HTB^{+}15, JPP^{+}17, KLSJ10, LN10, MKM^{+}16, NJJ^{+}11, NSK^{+}14, OVF11, PFF^{+}17, SNK^{+}10, SKN^{+}15, SBB^{+}14, SPP18]. northeastern [AGH^{+}12, CD19, LLC^{+}14, QHN^{+}10, SM17a, WFB16, VM14]. Northern [BM13, GCC^{+}10, ANWG18, AGNH^{+}19, BAA12, BBF^{+}19, BRMT11, BCOMKJ18, BCvH^{+}15, BP12a, BBBNT13, BRB11a, BBS^{+}17, BBR12, BVP11, BCV^{+}18, BBR^{+}14, BQ16, BKAS14, CL15a, CATAV18, DGCH18, DPS16, DEP19, DABP13, DHG16, DHS^{+}19, DPM15, DLL^{+}18b, DSGR^{+}15, DHH10, DPO15, DAB17, EPF^{+}17, GE14, GB16, GSTK11, HWSK19, HFR^{+}18, HFM^{+}15, HSMB17, HEM^{+}12, IDT^{+}18, JDQS11, JUK^{+}10, JGG^{+}15, JBS12, KW17, KMV18, KLN^{+}13, KLNJ15, LCJ^{+}12, LSH19a, LBD^{+}15, LJM^{+}10, LLCM^{+}12, LABG^{+}12, LGM^{+}12, LBE^{+}10, LNK^{+}16, LCB^{+}12, LDP^{+}17, MH19, MHW^{+}12, MBRO10, MBS^{+}13, MKK^{+}11, MLR12, OCS^{+}13, OAD^{+}14, PK13, RLN^{+}13, RCMB18, RFMT12, Ric12, RLAGL12, RBC^{+}15, RSHM10, SUPCB11, SPSG14, SRP^{+}19, SBA12, SCL13, SBB^{+}12b, SHF^{+}12b, SLS16, TKA10, TRHM14, TBC^{+}17, Urb12, VHHGF11, VLK^{+}15, WI12, WM10, WS12, WGS16, WNA^{+}15, WWAL13]. northern [YPK15, ZHM^{+}10, ZO10, ZMS^{+}14, ZMRCP^{+}17, dDTA^{+}18, BvHC^{+}15, Cra12, HKL15, PBC^{+}18]. northernmost [RRRF^{+}17]. northward [KTNO19]. northwards [RRRF^{+}17]. Northwest [BS18b, PMF^{+}17, RHMR18, SWM^{+}19, YWiIK15, BAW11, BOCB10, CWS13, DM17, GBL^{+}15, KPKA13, PW16, SJ10, SHM10b, TMS14, TBMdV12, Bea11a, Bea11b, BS18a, BKAM12, CGH12, DKAC^{+}12, DBG15, GLR^{+}14, GCR16, MKG10, MLSK^{+}15, RSDN15, SDF^{+}19, SB15b, TMH^{+}13, WL19]. northwestern [APNC^{+}15, CAGS17, CML^{+}12, ECM^{+}12, HBB^{+}18, HRS^{+}16, LSV^{+}10a, LPC^{+}11a, LCB10, MPB15b, MCHC^{+}16, ODP^{+}12, OV10, PPH11, PQAH^{+}17, SL11b, SNW^{+}18, TIS^{+}15, LFPA^{+}11, PIPD10, YTO^{+}11]. norvegica [KHS10b]. norvegicus [MLP19, OLD^{+}15, PPSR15]. Norway [FNC13, FLK12, FKA^{+}12, MRE14, OLD^{+}15, PBF10]. Norwegian [BANE^{+}13, DDK^{+}18, FNC^{+}14, JBM^{+}13, NKA16, OCS^{+}13, PSÓ^{+}15]. nosed [FHMG19, FMH^{+}19]. notable [AZSKWL11]. NOTE [KA16, LF14, PJMA18, SHP^{+}18, YNK18]. Notorynchus [AB11, IDT^{+}18]. Notothenia [EBOM11]. nourishment [Jon11]. Nova [FSLG12, FS14, KS11, MSK^{+}19, SFLG10, SJS^{+}11]. novaeangliae [BJS^{+}16, FMG^{+}18, LFPA^{+}11, RBP^{+}10, TFW^{+}12, WNNN12]. novaezealandiae [DGFH10]. novaezelandiae [WA19]. Novel [PDvG16, SCLB11, SSM^{+}19, TLS^{+}19, YMMO14, CB14, CWD11, DHSL11, HLC^{+}18, HVD^{+}16, MPG^{+}12, NMCP17, OTK^{+}13, PMM^{+}11, RRS15, SHMOS17, SRK11, TdRH17, WDM^{+}19, WNA^{+}15, ZOM^{+}10]. NSW [GHGF12]. Nucella [BVO16, DF15, MKK13]. nuclear [BCC^{+}11, DBS^{+}13, LLH^{+}11]. nudibranch [SG18a]. nuisance [MV12]. number [AvBBL^{+}16, PHB11b]. Numerical [FRP^{+}17, DEM15, KRW18]. numerically [KM10a]. nurse [RABS19]. nurseries [FKvdV12, JE10, KMO19, MHCS18, MDA^{+}10, PCP^{+}15, RLS10, TLB15, TLS^{+}19]. Nursery [DWPH11, WRI12, ATF12, ASF13, BAL14a, BDG^{+}16, CWK^{+}10, CCDP15, GMHS19, GRN14, HS11b, Hof10, JL12, KHF^{+}11, KRR^{+}18, KRD^{+}12, LR18, MMRC16, MIIB16, MCT16, NHS^{+}17, NCW^{+}19, RHS^{+}19, RMV13, SRSH^{+}18, SF16, TRSV^{+}12, TKL^{+}11, VMN^{+}12, WS11, WTGG10, vdVKA^{+}11]. nursery-dependence [CCDP15]. nutricline [RB19]. Nutrient [BHF^{+}10, BCM11b, CLJF12, CFPC^{+}13, EBH16, HHH^{+}11c, JREVB18, MHCRR18, MCHC^{+}16, PRA10, SCRGP11, STP13, STBP13, VGBJ^{+}13, VTZ^{+}13, VFMdW12, WA19, vBJV^{+}15, AC11, AS14, AS15, APC14a, BVL10, BRK^{+}12, BLHRP^{+}19, BHAC17, BMW^{+}14, BPB18, BRB17, BVGH10, BPF15, CPSC^{+}18, CL17, Dao13, DTVH10, DVS13, DTM^{+}12, ENW^{+}14, FCR14, FGOS16, FGP17, FT10, GP11a, GEF^{+}12, HP11a, HCK12, JSS^{+}11, JWR^{+}10, Jon11, KNIN17, KVY^{+}13, KCOP13, KWS^{+}13, KF13b, KPB16, KKY15, KTK^{+}12, LMW12, LRH14, LNIH16, LFD13, MTV^{+}11, MGFÁS^{+}10, MTM^{+}11, MSS^{+}14a, MMEM14, MAL15, PFB^{+}13, PWNP^{+}11, PFL18, RF14, RPM12, SGMG^{+}14, SMHB14, ST14, SO16, SH10, SBMÅ12, TFP14, TMGCM11, TLG14, TPSS12, TPF^{+}13, TVRG^{+}13, VARO12, VJJ^{+}19, VDEO18, WHB^{+}19, WAA^{+}19, WLHW10, YSC^{+}15, YKBB12]. nutrient-enriched [SBMÅ12]. nutrient-high [MTV^{+}11]. nutrient-induced [BPB18]. nutrient-limited [DTM^{+}12, ENW^{+}14]. Nutrients [BDT^{+}16, BP10, BVC^{+}16, FC12, FF10, GRCP^{+}10, HLD12, LNH^{+}17, LRBS15, MGFW13, PPB^{+}10a, PMYH13, RHE14, RPM12, TCKY19, VSG^{+}11, XTW^{+}15, YMMO14]. nutrification [CW12]. nutrition [CKK^{+}10, MPH^{+}11]. Nutritional [CRR^{+}19, KBP^{+}10, NUO^{+}19, vdGSE^{+}14, BCM^{+}14, DVS13, FA10a, GTP^{+}17, HLL^{+}17, KTP^{+}18, LMP12, PWNP^{+}11, SFTG14, WHT12, ZLCW15, dlSBO^{+}12, vBAP^{+}19]. nutritive [MBP^{+}12]. Nuttallia [TKIO10]. nuttallii [LD17a]. nux [NWS^{+}14]. NW [APC14a, ECZD14, THRBL^{+}16, AFSL^{+}15, APBC^{+}19, AMP^{+}11, ATT^{+}16, BC11, lLChTC11, FPC^{+}13, FCDC12, FAZ^{+}14, GSÁCM^{+}16, GLM16, JCA13, LD17a, LD17b, MGFÁS^{+}10, Mat10b, MAB^{+}10, PBAP^{+}14, PBCA10, PTF^{+}14, RCG^{+}15, TBCG^{+}12, TMGCM11, THMF^{+}15, TJS17, TFM^{+}18, VGGF12]. Nyctiphanes [GGMGR12].

O [CRG^{+}16, DH18a, GBH^{+}10, HJN^{+}10, JPL^{+}15, KSK07, KIU^{+}13, PHWC13, RAB^{+}15, WNNH10]. O. [KSA^{+}13]. obesus [RTCF^{+}17]. Obligate [SJL^{+}15, FBGH18]. obovata [CY10, HSS^{+}10]. obscure [SS12a]. obscurus [BNAK11]. observation [HTK^{+}17]. observations [BNAK11, CSKR10, Dav19, ETHB17, GCDH19, HFMD16, MAW^{+}11, MRLD19, PNA^{+}16, RS14, SLKN14, SRK11, YD11, ZVR18, CD13]. observatory [ASS^{+}13, ETD^{+}18, MPKO11, RJ12]. Observed [SPW^{+}12, DIG^{+}11, MKM^{+}16, WLM^{+}13]. observers [JHNL12]. obsoleta [HA18]. obsolete [BJ14]. obtained [FF16, GBMW12]. obtaining [LCB^{+}16]. obtusata [PP18]. Occupancy [IKS^{+}12, CNJ^{+}18, FGW^{+}15, HSS^{+}19, PHAZ18, TMBP13]. occur [ALK17, LMZ16, PTP^{+}19, TSGW10]. occurred [CW12]. Occurrence [CQR^{+}17, SM10b, AVB13, BLS^{+}11, BLIM11, BBR12, lLChTC11, CTSH12, CTvdM^{+}08, DWD^{+}18, FSW10, FBNJ13, PA15, RDF^{+}17, SHAS15, WHN^{+}18]. occurring [BADH^{+}12, BADQ^{+}15, GGMTGC18, HLS18, JTTP19, RLO18, RLH^{+}13, RBH12, SGL^{+}17, SHSK11, TG10, VB19b, dlSBB^{+}10]. occurs [HFH10]. Ocean [BER17, EHR12, ESD16, GLHM^{+}18, HM17, KCB18, KAMA13, LHB19, MPKO11, MJM^{+}14, MGD^{+}11, OTS^{+}10, PL15, PZL^{+}17, PBH^{+}17, RMB^{+}14, RSW^{+}16, SSM^{+}17, STS^{+}18, YLPG19, ZGH^{+}17, AMLS15, ARW12, AHPKS17, BGJ^{+}19, BRB11a, BJS^{+}11, BSG^{+}17, BH12b, BRIC13, Bur12, BGBD^{+}13, CCM^{+}14, CSJP17, CMC10b, CCP^{+}14, Chr17, CLMB13b, CLMM^{+}14, CH15, CAG12, CESC14, CB12, DBA17, DBG15, EKE16, EH13, EVB^{+}16, EVHA^{+}18, FRRÁSL11, FDFS16, FRTB14, FDP^{+}16, FPL^{+}12, FWB^{+}15, FCRD18, FTH12, GHHH12, GOM^{+}19, GLW^{+}17, HMP15, HBB^{+}18, HDWF10, HBBH12, HORRLS^{+}14, IS16, JDX^{+}17, JGM^{+}16, Jon11, KEBB10, KDS^{+}12, LME^{+}19, LHDH17, LBJI14, MP15, MLPC18, MFG^{+}13, MJT^{+}14, MTM^{+}17, MWBMH13, MMM^{+}15, MCR^{+}13, MWC^{+}12, MKI^{+}15, MH11c, MMGB14, ORW15, OSL18, PPML17, PHTS11, PCLB18, RBSL11, RHMR18, RFM^{+}15, RB19, SGMG^{+}14, SHS^{+}17]. ocean [SCAM^{+}14, ST13, SGL^{+}17, SG18a, SLF^{+}19, SDCC12, SMSB19, SABMG16, TL18, TBG15, TBH^{+}17, TDPS13, TTW^{+}12, TFH^{+}19, VJM^{+}16, VHM16, VMWU15, WAP10, WDL^{+}18, Whi11, WNP^{+}14, WWW^{+}13, XGLH14, YPK15, YYU^{+}17, ZHJ^{+}17, ADHP12, AFMR^{+}13, ALSV16, BRJ^{+}12, BDM^{+}13, BMTC^{+}15, BST19b, BPN11, BTLF17, BAW11, CC19, CTZ^{+}16, CLGCA^{+}19, CPTL11, CMB^{+}14, CNW^{+}12, DSS10, DR10, DSH14, DPY^{+}16, DLRF15, DCHI12, ETD^{+}18, EBdlTC12, EFOM13, FHTY19, FLE^{+}11, FLSG^{+}15, GLJ^{+}18, GBMW12, GCM10, GGLT15, GHT^{+}15a, GPC^{+}15, HSH12, HTB^{+}15, HBM^{+}12, HGB^{+}13, IMS^{+}11, KVR^{+}18, KLD18, KRR^{+}18, KSK^{+}12, KBB10b, LHS^{+}14, LABM17, LDAE12, LMSC^{+}19, MASW^{+}11, MM16, MBR^{+}18, MAB^{+}10, MMB^{+}12b, MTH15, MTG^{+}11, ODK^{+}14, OPW^{+}11, PDT^{+}11, PRP^{+}18, PJMA18, QHN^{+}10, RCP^{+}16, RGL^{+}16, RNFG10, RDM18, RL11, RNGUR^{+}11, RLNLAF17]. Ocean [RSSvV19, SBV^{+}15, SCW^{+}15a, SCST17, SMK^{+}18, SCO^{+}12b, SMD^{+}13, SSB^{+}11, SBCT17, SDF^{+}19, SGB^{+}10, SNS12, SLRS13, TDB^{+}14, TTT^{+}15, TSA^{+}18, TMH^{+}13, TBH^{+}17, VWW^{+}16, WKX^{+}19, Wil14, WJS^{+}14, YSM^{+}10, tHHR10]. ocean-entry [SCAM^{+}14]. ocean-exposed [ST13]. ocean-wide [RBSL11]. oceana [SLC11]. Oceania [CJS^{+}12]. Oceanic [GBD^{+}12, KTK^{+}12, MvdHBM12, NTS^{+}18, SMVC^{+}15, SØA^{+}13, WHFM17, AGP^{+}13, BPSCD17, BBC12b, CKK^{+}10, CWW^{+}16b, CLN^{+}13, DOH^{+}17, FBS14, HLW^{+}16, JBL^{+}16, JM17, KMM15, KdVvdP^{+}12, LCBK^{+}13, MSFM12, MBB^{+}15, MAW^{+}11, ODL^{+}13, PXG^{+}10, RRH^{+}16, SWR11, SCP16, TWY^{+}19, TLFS14, ZMGS17]. oceanica [AFC^{+}16, AFS19, BRA^{+}10, BBC^{+}10, CESAM12, CSS^{+}17, DBCB11, GMC^{+}12, GCBB13, IOS^{+}12, JMPZ10, MATR11, PRFA11, SMR11a, VBŽ^{+}17, ZKS17]. Oceanographic [CCS^{+}18, HCJ^{+}18, KLNJ15, MWC18, RB18, ASR^{+}19, Ano18a, BT10, BST^{+}19a, BKM^{+}18, BQ16, BJRH10, CCKB^{+}15, DCT^{+}10, EDS14, EFOM18, FMH^{+}19, FHB^{+}12, GCFW14, GMSC14, GCRP15, LBE^{+}10, LRC^{+}13, MRL^{+}17, PBAS^{+}10, PDP^{+}10, PCD^{+}18, RNB^{+}16, ST10a, SGS^{+}12]. oceanographical [FHM^{+}11]. oceanographically [RMNP17]. Oceanography [HTCH^{+}18, WSC13, KS13, SMVM^{+}12]. oceans [CBD14, KPM^{+}11, MdSW^{+}16, TBGW^{+}11, WTC^{+}15, ECC^{+}15, MAD17, RVPG17]. ocellatus [ARA^{+}10, NRF11]. ochraceus [BLD^{+}16b, MWGH15]. octocoral [CL16, CKC^{+}10, GGG18, TMB^{+}12]. Octocorallia [PMH^{+}13]. octocorals [EL19, GL16, GFF16, KBL11, Mat10b, SHM10b, VBT^{+}17]. octolineatus [BH15]. octopods [SPH18]. Octopoteuthis [Bus12]. octopus [BRBH19, HSDB13, CPB^{+}11, KM14, NGSF13, PRNR^{+}18]. ocular [SCSC16]. Oculina [AZSKWL11, LFPC^{+}14, MSLR13]. Ocypode [POX^{+}17]. Ocypodidae [APH^{+}10]. Odessia [CGQB10]. Odobenus [HKL15, NUJ12]. odontocete [NSC^{+}19]. odorabile [PW11]. Odour [JMPZ10]. odyssey [LOS^{+}19]. off [BWS^{+}12, BS16, BNAK11, BPZ^{+}14, BJL^{+}17, CTZ^{+}16, CHP^{+}11, CH12, CATAV18, DER^{+}19, DBB11, DDP19, ECM^{+}12, FBA11, FMH^{+}15, FLK12, FSFS18, GSP^{+}15, GKC11, GMP^{+}11, GB10a, HFMD16, JDQS11, JRW^{+}14, KWM12, KBK^{+}10, KMM15, KBB^{+}10a, LP10, LLC^{+}14, LLCM^{+}12, LCC^{+}11, MVVGAV10, MGFÁS^{+}10, MSM^{+}12, MKG10, PBCA10, PPEH^{+}14, RDF^{+}17, RBEGG13, RTCF^{+}17, RHRB^{+}17, SVAV17, SS11a, SMHB14, TCSA17, TMGCM11, THMF^{+}15, TBB^{+}19, TBDM11, WY15, WZP^{+}14, YBB^{+}12, vDCFM10]. off-reef [GB10a]. offer [AGP^{+}13]. offs [BJM19, BL17, HKJ17, RPRN^{+}16, RvOP12, RT17, SPP12, SS12a, SGGL15, VQSH14, dMA18, dDTA^{+}18]. offset [FDVL18, JDX^{+}17]. offsets [BRIC13]. offshore [ÅCAC16, ANRT10, BJ14, BSB13, BDBN11, BHD^{+}13, BDD^{+}18, CAB^{+}19a, DTC^{+}17, DDP19, FDS12, Fuj16, GFC^{+}13, HGS14a, JRW^{+}14, JBS12, KAJP^{+}15, LHJ11, MEKC15, MF10, NSK^{+}14, SSvD^{+}15, TRSB^{+}18, WN16, dB10, vDGK^{+}12, vdWTWS15]. Offspring [CPPM12, MMGB14, FDVL18, HBG^{+}17, KWM12, MOS10, OT13]. ogac [SGM^{+}14]. Oil [MFDW17, BMC^{+}15, BSF17, BBM13, BBC^{+}12a, DBCS16, EFR12, FCR^{+}15, Fuj16, GFF16, GPRR17, HGS14a, HGS14b, HGS15, HSC^{+}15, HG13, JGG^{+}15, KCMH^{+}17, LBM^{+}18, Mar17, MRC^{+}19, PSF^{+}19, PGR^{+}17, PRB^{+}17, PPCH17, RFH16, RCMB18, RSV10, RBS^{+}17, SB15a, ZWP^{+}17]. Oithona [ACA^{+}10b, ACA^{+}10a, DHH10, ZTNS13]. Okinawa [HHI^{+}14, NKS^{+}15, HTR19a, NWS^{+}14, YWiIK15]. Oksanen [Ano15b]. old [BMV^{+}17, JW10, MFSRR14]. old-growth [JW10]. older [BCWSR14, HLS17]. Olfactory [ARGB13]. Oligobrachia [LKL^{+}19]. oligohaline [GM10, SW18]. oligotrophic [AGTS^{+}12, AGM^{+}11, CPM^{+}17, HTB^{+}17, HW14, HDWF10, JSS^{+}11, KKO^{+}14, SSAB^{+}15, WBLS^{+}11]. oligotrophication [TPS^{+}13]. oligotrophy [RHVRC11]. olivacea [PAC^{+}15, RAKB^{+}12, TKIO10]. olivaceus [CL17]. olive [PAC^{+}15, RAKB^{+}12, dSdSdCO^{+}11]. Olympia [PS15b]. Oman [RAKB^{+}12]. Ommastrephes [IMS^{+}11, VSW^{+}15]. Ommastrephid [SRCE^{+}10]. Ommastrephidae [ICT^{+}11]. omnivore [JW11, NPRAGG^{+}12]. omnivorous [CMO^{+}10, PH11]. Omnivory [JGS14]. On-shelf [TSCvS^{+}15]. Onchidella [CNSW16]. Oncorhynchus [HHP^{+}17, JGM^{+}16, KSA^{+}13, MGM10, WSF^{+}12, WWW^{+}13]. One [HFM12, TGLC10, ATP14, GGB^{+}18, PL15, RAB^{+}10, ETB13]. Ongoing [MS11]. ongus [NMS^{+}18]. Onisimus [NWC^{+}10]. online [AS12]. only [BLIM11, BSN^{+}18, GSL17, MIV^{+}10, PNPF14, PWM^{+}19, TKB^{+}13]. Onna [NKS^{+}15]. onset [ALANGT10, SPH^{+}13]. Onshore [SMM^{+}14, FRP^{+}17, VGL16]. Onslow [WMB^{+}14]. Ontogenetic [BLJP14, CMLC12, CNÁCP12, CBL^{+}10c, GCS^{+}18, HRS^{+}16, JG13, KHS^{+}11, OSC19, PKS11a, RP12, RLAGL12, RCVG10, SST19, VCG^{+}15, WW15, WHL11, BL17, CSG11, DFB^{+}17, DZL16, ETHB17, GRN14, HCL^{+}11a, KF17, MM10, RMP^{+}19, Rit17, SQRDLC18, SPS16, dG14, vdMPM^{+}10]. ontogenic [CSC^{+}11, DWD^{+}18]. Ontogeny [FG10, HMB^{+}12, KSFL14, SFTG14, HLF^{+}19, HNG^{+}11, JWR^{+}12, JWBS10, LJM^{+}10, MJM^{+}14, MTM^{+}17, RPLT10, SHB^{+}12]. onyx [LC13]. oocyte [DE17, GGPDS19, HPMB11, SNK^{+}10]. Open [IS16, AFRAG16, ALS18, BRB11a, BMF^{+}12, CPP10, CDC18, EH13, FCRD18, HORRLS^{+}14, INvdB^{+}13, LBJI14, McC12, MHFN10, MKI^{+}15, MF10, MN10, NCK^{+}14, PBC11, RB19, SLF^{+}19, WKRW^{+}12, WWG^{+}13b]. open- [CPP10]. open-coast [BMF^{+}12]. Open-ocean [IS16, HORRLS^{+}14, MKI^{+}15, SLF^{+}19]. open-water [AFRAG16]. opened [WKX^{+}19]. opens [OPSM13]. operated [KBG12]. operating [LMG12]. operation [MGM^{+}18]. operations [RKT16]. ophiuroid [BADH^{+}12, BADQ^{+}15, FSB15]. ophiuroids [GFF16]. opilio [DMMW12]. Opinion [Chr17, DC17, Lap19, MBCB^{+}19, QBW18, TVA^{+}19, WFD19]. Opistognathidae [HPL12]. Opistognathus [HPL12]. opportunism [ELM^{+}19]. opportunistic [AS14, CPF^{+}14, DS10, IME12, TKV^{+}19]. opportunities [AGV^{+}11, CCGP14]. opportunity [BBH^{+}11]. opposing [RLH^{+}13]. Optical [CAKD14, GFR^{+}18, GH11, GMD16, LHW^{+}19, MVTS^{+}12, MRSD15]. optics [DZDB10]. optima [LGD10, RAN^{+}10]. Optimal [JJJ16, PO13, VF13, BLM10, CLP18, CRGM18, FJ11, GHT15b, OTK^{+}19, PP10]. Optimality [PDO13, Wir14]. Optimality-based [PDO13, Wir14]. optimizes [WP10]. option [EGH^{+}15]. options [FQ11, VQSH14]. Orbicella [EPL^{+}14, LBJK14, MEKC15, WMCC15]. orbicellid [WSK^{+}16]. orca [BDS^{+}18b, NSK^{+}14, RGL^{+}16]. Orcinus [BDS^{+}18b, NSK^{+}14, RGL^{+}16]. order [Col10b, SDKD13]. Oregon [SS11a, DHJ12, OGCHS18, PRR^{+}16, TAB^{+}14, TBDM11, TSF^{+}10, WZP^{+}14, YBB^{+}12]. oregonensis [CFMN16]. Oregonia [HJ16]. Oreochromis [DPC13]. Organic [BLER15, HJN^{+}10, LGW^{+}15, WNNH10, BW16b, BW16a, BBI16, BP10, BVP11, BBG^{+}16, CGR^{+}15, CIF11, CAKD14, CMC^{+}15, DMC^{+}16, FJC10, FC12, FLSG^{+}15, GBW^{+}12, HPMZ11, JKM11, JPMG10, KdWV^{+}18, KRC^{+}11, KRD^{+}12, LBN11, LMW^{+}13, LBM^{+}13, LRL^{+}13, LFM^{+}16, LSRRCM13, LNK^{+}16, MGFÁS^{+}10, MPB13, MDJN19, NOT^{+}11, QOGS^{+}16, ROE17, RCB^{+}10, RdGv^{+}18, SMSE19, SBS^{+}13, SGP^{+}17, SBE^{+}19, STRP18, SBM^{+}18, TB10, TNSN11, TNS^{+}12, VAHL11, VKH10, VQKF14, VBM^{+}11, VBM^{+}13, WBLS^{+}11, YYU^{+}17, YMPR13, vdMHW^{+}10]. organic-enriched [VQKF14]. organic-walled [CIF11]. organisation [CHST16]. organism [CLD^{+}13, PBA12]. organisms [EFOM18, GOO14, MM16, NOT^{+}11, NPRAGG^{+}12, OSPWB15, Pör12, VMWU15, VB19b, WNNH10]. organization [AN11a, Miu12]. organize [CGH^{+}16]. organotin [OHA10]. organs [FRGI15]. orientalis [KIU^{+}13, KHT^{+}17, Sat10, SWY^{+}10, WHI^{+}18]. Orientation [FCS15, INvdB^{+}13, LPI^{+}14, HSR^{+}13, KPS14, LSH^{+}15]. oriented [HTR19a, HCJ^{+}11]. Origin [FAM^{+}17, MDJN19, ALE^{+}13b, BPWF14, Coo11, HPN^{+}17, LBVH13, SWY^{+}10, WRI12, WWW^{+}13]. origin-dependent [WWW^{+}13]. origins [BS18b, BS18a, JGM^{+}16, LDAE12, PPD^{+}14, SMMFH16]. ornata [BDT^{+}16]. ornithogenic [WLF^{+}17]. ornithology [HW12c, HW12b]. oscillation [RBHR19]. oscillations [CESC14, VPW^{+}12]. osculatum [SBH^{+}18]. Osculum [KRC17]. Oslofjord [PBF10]. Osmerus [BDT^{+}11]. Osprey [CBJB19]. ostenfeldii [TLM^{+}11]. ostralegus [GCBS19]. Ostrea [EKR^{+}10, MS18a, PSP^{+}19, PRP^{+}16, PRR^{+}16]. ostreae [EKR^{+}10]. other [BB14a, BWN^{+}14, CSM^{+}19, DGMB10, EMG^{+}15, HFM12, JKM11, Kiø16, LCL^{+}13, MAB^{+}10, SBL^{+}10, TTF^{+}11, TPFP11, ZBP18]. otherwise [vBMH^{+}17]. Otolith [BWS11, CZB10, HS16, KIU^{+}13, LTN^{+}11, PB19, SCSC16, TRSV^{+}12, WDM^{+}18, ACW18, BDT^{+}11, Coo11, DH18a, FH18, HMH^{+}12, HJ13, KRR^{+}18, LR18, MMRC16, MDA^{+}10, MBW10, MMB^{+}12a, MMC^{+}16, MS14, NSS10, PLMN10, SGT^{+}14, SWY^{+}10, SKMS^{+}18, SGP^{+}17, STM^{+}14, VM10, Vig15, WWAS18, WWG12, WWG^{+}13b, ZSSK11]. Otolith-based [PB19]. otoliths [AdCM^{+}19, CLDC^{+}13, CGG^{+}10b, EAMT10, GHT15b, HWET13, JGCRS12, KSK07, MC10a, MFG^{+}13, SWW11, VWEP18]. otter [BBC^{+}12a]. otters [BBM13, HG13, MBB^{+}12b, TKDM16]. oualaniensis [SRCE^{+}10]. Oulastrea [LKN^{+}13]. our [Ano17c, GBC^{+}16]. outbreak [PBS10, SBW16]. outbreaks [BPW^{+}18, KHH18, MPB^{+}14b, SFLG10]. Outcomes [MPF^{+}15, BKTD10, SKD^{+}11a, WF18]. Outer [DJL^{+}12, RDF^{+}17]. outfall [JJNP12]. outlet [JPAM^{+}15]. outmigration [MLB^{+}14]. outplanted [PWM18]. output [BBW^{+}19, CGC15, CCM^{+}12, DMPG11, FTF15, Hal15, PPM^{+}15, SKK^{+}18, TP19, VBD^{+}11]. outputs [SWP^{+}13]. outward [AHØ^{+}15]. outweigh [CM11, KMO19, STM^{+}14]. Outwelling [WAMS^{+}14, TNM^{+}11b]. ovalis [SMK^{+}17]. overall [VGR^{+}16]. Overcoming [PGR^{+}13]. overdue [BEEH15]. overestimation [AGTS^{+}12]. overfished [vdBUPP^{+}11]. Overfishing [WLB^{+}11, BEP^{+}12, CS15]. overgrazed [FNC13]. Overgrazing [FMC^{+}10, FJL12]. overgrowth [BT14]. Overlap [CEH18, WPF^{+}16, AGV^{+}11, BI15, BBR^{+}14, CTvdM^{+}08, DGCH18, DRD^{+}18, GSP^{+}15, HKI^{+}17, KBPD10, LCGS^{+}19, LDT^{+}13, OTK^{+}19, RNB^{+}12, RWO13, SRA^{+}17, TRGF18, VH11]. overlying [MHGW13]. overshadows [MTB^{+}17]. Overview [PBWM18, BBW^{+}19, LPC19]. overwhelms [MBBBC12]. Overwinter [RCS^{+}17, KPM^{+}11]. Overwintering [PJMA18, vDCFM10, CAG12, SKK12, WHFM17]. ovum [LGB^{+}13, MBR^{+}13]. oxic [AWJ12, VPW^{+}12]. oxidation [KRDV18, TKW^{+}14]. Oxidative [VLK^{+}15, PUP16, RGCT^{+}10]. oxide [HSdBS10]. oxidising [TKW^{+}14]. oxidizing [BKHW13]. oxudercine [PB10]. Oxygen [JLQ^{+}14, SdBAH^{+}14, VBR18, APC14a, ASK^{+}15, BCvH^{+}15, BS11, BCK15, BC13, CLG19a, CG16, CPHH15, DBG15, DDB^{+}16, DTL^{+}15, dlÁGRB^{+}19, GE17, HHH11a, HB13, IAV^{+}10, JLL^{+}12, JNN12, KCW^{+}17, KGLLW11, LWM15, LMZ16, OFM^{+}18, RWR^{+}19, RHRB^{+}17, SWY^{+}10, SKMS^{+}18, TvdMM^{+}17, TPKC18, VGR^{+}16, WKE^{+}17, WMS^{+}11, WPMP14]. oxygen-minimum [dlÁGRB^{+}19]. oxygenation [EBH16]. oxyrinchus [NCW^{+}19, PBB^{+}15, SNS^{+}13, TKEV^{+}19, VWW^{+}16]. Oxyrrhis [BBRBG10, CIM^{+}13]. oxytaxis [LWM15]. Oyster [SGP13, ALV12, ALK17, BG19, BPW^{+}18, BPHW16, BSCH14, CB11, CLL15, CL15c, CLL17, DPM15, DWL19, FAM^{+}17, GFL^{+}17, GSF^{+}13, GGMTGC18, GB10b, HPA17a, HPA17b, HWSK19, HSMS^{+}15, HW13, HR19, HR15, KSF18, KCOP13, KBW^{+}15, KW10, KPP^{+}16, LdU^{+}17, LFU^{+}19, LGBSC11, LEB14a, LRS17, MGP11, MGP12, MB16, MCBR18, MWHM18, MRB^{+}15, NCS^{+}10, PSP^{+}19, PS15b, PS19, PFL18, PGR^{+}17, RHF19, SPMN19, SCM13, SGTP16, SMR10, SH17, TG12a, TG12b, VTSH^{+}19, VBR18, WDH^{+}12, WHA^{+}13, WR12]. Oyster-mediated [SGP13]. oystercatcher [KCH^{+}11]. oystercatchers [GCBS19]. Oysters [ATP14, BHASP18, BIL^{+}10, DC15, EKR^{+}10, GLF^{+}17, GSF^{+}13, GSL17, GPRR17, GC13b, GCP^{+}17, KR16, LRB^{+}18, MPF^{+}15, NLLS15, PBM^{+}10, PW19b, TSR17, WSHY16, WLH^{+}17, WLB^{+}11].

P [BFH13b, BBC^{+}10, CHLR16, FGP17, GRG^{+}12, KJW^{+}15, OAGV11, VTZ^{+}13]. P-limited [BFH13b]. P. [EPM^{+}16, HCK12, MTS^{+}17, PRLR^{+}18]. Pachygrapsus [CMO^{+}10, FRC^{+}13b]. Pacific [ADHP12, BPN11, CNW^{+}12, EFOM13, HTB^{+}15, IMS^{+}11, LDAE12, PDT^{+}11, RL11, RSSvV19, SSB^{+}11, TTT^{+}15, Alb15, AAGGFG^{+}12, ACAD12, BHP12, BSC^{+}13, BLE^{+}13, BHN^{+}14, BJS^{+}11, BPHW16, CWW^{+}16a, CGKW10, CGK13b, CAB^{+}19a, lLChTC11, CPID13, CD13, CWW^{+}16b, CGH12, CL10c, CLSS17, CM10, CGM^{+}13, CATAV18, CBK^{+}12, CMH14, DA11, DPH14, DBDK14, DWL19, ETS^{+}11, ENO^{+}11, ETB^{+}11, FHTY19, FAM^{+}17, dSKFV12, FESPR14, FHH^{+}17, FPR18, FSDS10, GLJ^{+}18, GBMW12, GWT18, GFGGDSM12, GCM10, Gla10, GEA^{+}14, GGG18, GAC11, HSPS14, Hat13, HLPS^{+}15, HVT12, HCBV^{+}11, HGG^{+}10, HAF^{+}17, HT15, HHNI17, HGA^{+}16b, HGA^{+}16c, HDP^{+}17, HLN^{+}19, HMF^{+}18, JBT^{+}16, JPP^{+}17, KN12, KTNO19, KRW18, KIU^{+}13, KF17, KKO^{+}14, KHT^{+}17, KLSJ10, KFGSMPG14, KBC^{+}12, LdU^{+}17, LRB^{+}18, LFU^{+}19, LSY11, LD17a, LKLC11, LKN^{+}13, LFM^{+}16, LCR^{+}16, LET^{+}12, MLM^{+}12, MZ10]. Pacific [MCW^{+}16, MASW^{+}11, MAD17, Mat10b, MHGW13, MNKB11, MAP^{+}12, MWZ^{+}16, MADI13, MKM^{+}16, MZF^{+}19, MMBA13, MKK^{+}11, MGCH^{+}16, MCT15, NCS^{+}10, NHTS14, NUJ12, NLLS15, OTK^{+}19, ODL^{+}13, ODK^{+}14, OMT^{+}17, PBAS^{+}10, PBC^{+}12b, Ran18, RCP^{+}16, RRL^{+}19, RBSL11, RHMR18, RKC16, RNGUR^{+}11, RLNLAF17, RDF^{+}19, Sat10, SØA^{+}13, SØK^{+}15, SBE^{+}14, SCO^{+}12b, SNW^{+}18, SNVH19, SFW^{+}14, SWY^{+}10, SBB^{+}14, SWH^{+}11, SRCE^{+}10, SHD^{+}19, SH17, SPP18, TIS^{+}15, TLFS14, TLJ^{+}11, TSA^{+}18, Ued14, Urb12, VCG^{+}10, VTSH^{+}19, VM14, WSHY16, WHI^{+}18, WBLS^{+}11, WM10, WWB^{+}17, WMB^{+}14, WA14, XWE^{+}14, YWiIK15, YTO^{+}11, YSK^{+}10, YSM^{+}10, ZO10, ZBK^{+}15, vBAP^{+}19]. pacifica [MKM^{+}16, PKS11a]. pacificus [PVYS16]. pack [LHJ11, ORE^{+}17]. package [KW17]. page [Ano17b]. Pagellus [FCR12]. Pagrus [CMLC12, LMC14]. Pagurus [PDB15, MHS17]. painted [MPA^{+}10]. pair [FSF^{+}17]. paired [WFP15]. pairs [FBA11]. Pakistan [JLL^{+}12]. Palaemonetes [OT14]. palaemonine [OT14]. palaeoclimatic [MvdHBM12]. Palaeoecological [LSD^{+}13]. palatabilities [CLJF12]. palatability [TOW10]. palatibility [HWR^{+}13]. Palau [CCS^{+}18, GGvW^{+}15]. Paleo [LPK^{+}16].

Paleo-ecological [LPK^{+}16]. paleoenvironmental [HIKK12]. Palinurus [BFG^{+}11a, BFG^{+}11b, JBCGC19, SRG14]. pallasii [FESPR14]. pallida [DBS^{+}19]. palmaeformis [TKBN10]. palmata [DSVM10]. Palmyra [HSPS14]. Palos [TWL15]. palpebrosa [GCS^{+}19]. Pamlico [GSF^{+}13]. Pan [FWPP12, HCLB18]. Pan-Arctic [HCLB18]. Panamá [AHB^{+}14]. Panamanian [VGBJ^{+}13]. panamensis [CTRBLC^{+}13]. Pandalus [HP13, JBS12, MASMH19, Ric12, WS12]. pandora [FCR12]. panicea [KRC17, LRSB19]. Panmictic [RA10]. panmixia [HHLT10]. pantropical [MBC^{+}17]. Panulirus [AB13, BFEMLÁJ19, KBG^{+}13, KB14, KPS14, LH10, Mat10a]. Papasula [HW14]. papers [TGLC10]. Paracartia [OSC19]. Paracentrotus [CT10, PAFC13, PPC^{+}12]. paradigm [ALSV16, CCC^{+}18, TSA^{+}18]. paradigms [WEF19]. paradise [ETS^{+}11]. paradox [GRL^{+}15, KWS^{+}13]. paradoxus [Sat17]. Paralabrax [GS18]. paralarvae [PRNR^{+}18]. paralarval [DFGH^{+}15, VSW^{+}15]. Paralichthys [CL17, DTG16]. Paralithodes [FPN17, PES12]. Parallel [BDD17]. paralytic [VTZ^{+}13]. parameter [PMKL11, ST10b]. Parameterisation [HB13, SPM^{+}13]. Parameterising [JR15]. parameterizing [LFL11]. parameters [BDC10, DTL^{+}15, EFOM18, FCR12, GDJdBSL19, HU13, JCZ^{+}12, LCB^{+}16, MHO^{+}11, SGT^{+}14, VGGD^{+}15, ZBB^{+}19]. Paramuricea [LD10, TMB^{+}12]. parapatric [FCV^{+}16]. Parapercis [WBJ12]. paraplexauroides [MLASTS13]. Parasagitta [GSV14]. Parasesarmea [CY10]. Parasite [Fis10, BHASP18, CSRS15, HMP15, HSMS^{+}15, LGR^{+}10, MB16, MH11b, RB11, RR18, STP10, TP16, WGD^{+}16]. Parasite-like [Fis10]. Parasites [CGC15, STP10, BTLF17, HWSK19, dML11, dMBC12]. Parasitic [POS14, DDDR10, DDM^{+}16, GSM^{+}11a, JDD^{+}13, MP15, Miu12]. parasitism [DLdM19, LHPP16, Mou17, SWPK13, WGD^{+}16]. parasitoids [SS10a]. Parechinus [MvdHBM12]. parent [HNMK15]. parentage [BHP11, JCF^{+}12]. Parental [OVF11, RP15, YBD^{+}16]. Park [HSF^{+}12, LRH14, PPD^{+}14, BBMM14, SML^{+}15, MMC^{+}18]. Parma [VHKT15]. Parrotfish [DCL^{+}17, LMB^{+}19, AZVB13, BKB11, CDWC17, GRFSM12, SQRDLC18, SAM14, VPDPMRC11, YPWP18]. parrotfishes [ADF^{+}18, NGJHB12, OHB^{+}15, RPCL19]. parsimony [TW10]. part [FBC^{+}15, dDTC^{+}17]. Partial [GFS15, CPS15]. Particle [CMP^{+}13, RL10, SF14, BZBR^{+}11, BDM^{+}14, FMJA^{+}18, HBM^{+}12, KCP^{+}11, KKO^{+}14, MG12, MGB^{+}14, RPS12, SPP17]. particles [MSC13, PPB^{+}10a, PMYH13, PWCvO^{+}14, YYU^{+}17]. particular [OCSW16]. Particulate [WLHW10, BBG^{+}16, CMC^{+}15, FLSG^{+}15, HGB^{+}13, JPMG10, KRC^{+}11, KRD^{+}12, LBN11, LFM^{+}16, MRSD15, MPB13, NOT^{+}11, RCB^{+}10, SDH13, YMPR13]. Partitioning [BHT10, BB15, BRBH19, BLER15, CMB^{+}14, DW12, EMH^{+}19, EMJ18, GLR^{+}14, GCR^{+}17, GGCB16, GRCH^{+}18, HKT10, HVK^{+}13, HSS^{+}19, HSWS16, JMS^{+}15, KWPL11, KHF^{+}11, LCGS^{+}19, MCT16, NMPM15, PBTB15, PRLR^{+}18, PBCA10, PTP^{+}15, RHVRC11, RCP^{+}16, RSTM14, RMT^{+}13, SNTN15, SST19, TB12, Vig15, YMD^{+}10, YSM^{+}10, ZO10]. partner [PC19]. partnerships [MU13]. Parvocalanus [JSLG17]. Passage [MGG^{+}14, PMF^{+}17]. Passive [SSdW^{+}12, TBG15, GBD^{+}12, MGM^{+}18, MHS15, RWS^{+}15, RBD^{+}14, SMSB19, SWH^{+}11, USR^{+}16, WSLM13]. Past [SVIH^{+}10, BOCB10, KGJ^{+}16, TFS14]. Patagonia [GCC^{+}10, BQ16, GDP^{+}11, GGK^{+}16, IDT^{+}18, MPD11, SWFQ12, VCG^{+}15]. Patagonian [AIA18, LMBM18, PYBB11, QOGS^{+}16, VHHGF11]. patagonica [AZSKWL11, MSLR13]. Patch [Hv14, SLT^{+}19, ABF^{+}14, BVL10, BBBNT13, CDD^{+}16, Cha13, DA10, KB12, KMO19, LEB^{+}19, MPH^{+}11, MHC11, RMME10, SCK^{+}17, VRMKH^{+}15]. patch-scale [ABF^{+}14]. patches [AED^{+}10, AMS^{+}13, KAAI15, MDF10, MBW^{+}19, PB18, Sat10]. Patchiness [TWW^{+}15, BRK^{+}12, BHN^{+}14, CFMN16, DBE^{+}18, GMD16, HTH10, KBG^{+}13, MGP11, MGP12, MHS17, TGBM^{+}16]. Patchy [vDCR13, BBKH^{+}11, CWMP13, KW12]. Patella [BHDC15, NMHB16, SHC^{+}10]. paternity [DEGF^{+}10, Sat17, SWNSF14]. pathogen [BSS18, CB11, HSMS^{+}15, SHMOS17]. pathogenic [MZC^{+}10]. pathogens [SVG16, SJO16]. pathology [SBL^{+}10]. pathosystem [BMR17]. pathway [BBC^{+}12a]. pathways [CAG13, LCQT^{+}13, LNH^{+}17, QLCUG15, RWMP12, RVH^{+}19, SE11b, SYM^{+}16]. Patiriella [BGBD^{+}13]. patronus [OFDG14]. Pattern [CWR^{+}18, FRC^{+}13b, AHD12, BBC^{+}10, BB19, BFH13b, DMHS^{+}14, FJL12, GCM10, HHH^{+}14, MADI13, RSSvV19, SSBC11, SBT15, SWW11, WLP^{+}11]. pattern-based [AHD12]. patterned [WHvdK11]. Patterns [BBB^{+}14, BW15, BBBHC17, BBGL13, CAB^{+}19a, CRWG16, DBE^{+}12, DJL^{+}16, DQC^{+}13, JMS^{+}15, KMP11, Lec18, LMd12b, OSPWB15, PTL16, PBC^{+}12b, RWAP16, SKFH18, SS11b, TWL15, WH11, AMSdL^{+}18, AN11a, AVB13, AMP^{+}11, AAGGG19, AGC12, BWS^{+}12, BDM^{+}13, BSG^{+}10, BCM^{+}14, BSC^{+}15, BSB13, Ber12b, BN16, BOÁNF19, Bol17, BAH^{+}15, BB14c, BBKP10, BFN^{+}19, BBH^{+}12, BNW14, CRAP^{+}12, CLG19a, CCE^{+}14, CDWC17, CWK^{+}10, CM14, CHDH^{+}19, CRJ13, CTMU10, CGC15, CTSH12, CRJ^{+}18, DMd14, DHJ12, DMD^{+}12, DMS14, DCL^{+}17, DWK13, DQM^{+}19, DDB^{+}14, DEWK15, DHM^{+}14, DB11b, Ebe10, EOA15, ETO^{+}10, FKT15, FDDD13, FCS15, FSDS10, Gal14a, GRL^{+}14, GKW10, GB16, GW15, GGFB15, GPA^{+}16, GH11, GDPR18, GSJS^{+}19, GCL^{+}16, HP11a, HSPS14, HTR^{+}12, HW11a, HLB^{+}16, HMJ^{+}13, HORRLS^{+}14, ICUM10, JUMS10, JH14, JHL^{+}19]. patterns [JJNP12, KBA13, KHBL11, KF13b, KCLLL13, KJBS12, KPM^{+}11, KS11, LFU^{+}19, LCQT^{+}13, LAM10, LM12, LHT^{+}16, LRB^{+}12a, LMP12, LOBW^{+}11, LBJI14, LWG14, LMG12, LAH^{+}15, MPS^{+}10, MORM15, MPA^{+}10, MHHB^{+}17, MVE^{+}10, MJR^{+}14, MGPC13, MIA^{+}11, MSH^{+}14, MCS^{+}15, MSG11, MGAV15, MRGG^{+}13, Met11, MHVV13, MMBA13, MOK^{+}11, MMM^{+}19, MPB15b, MDS^{+}13, MAHK12, MA13b, MER14, NBGGST^{+}15, NWH^{+}12, NDD^{+}13, OTK^{+}13, ORE^{+}17, OKF^{+}18, Par15, PTP^{+}19, PGWS10, PSD^{+}19, PPTB12, PLFR^{+}11, PFR11, PTP^{+}15, PKBS14, PAK15, QLCUG15, QOGS^{+}16, RIM^{+}13, RFBH^{+}11, RKD18, RHH^{+}15, REBL16, RRH^{+}16, RCD^{+}13, RDC^{+}12, RNSA15, RMR12, SRLV10, SAZHAAHS14, SBG^{+}16, SRSH^{+}18, SRP^{+}19, SHT^{+}18, SL11b, SNVH19, SBB18, SJL^{+}16, SRM^{+}15, Sma13, SHT^{+}13, SBH^{+}18, SMF^{+}11, SPD^{+}14, SRK11, SWPK13, SS10c, SB14, TMSG19, TBB^{+}19, TTM19, TTB^{+}11, TPSS12]. patterns [TGSM12, TCWC13, TTW^{+}12, VVB19, VCDP11, VTB12, VHHGF11, VHG^{+}18, WSD^{+}10, WGD^{+}16, WGD^{+}19, WGD10, WGS16, WLP^{+}10, XS19, YYU^{+}17, YRM^{+}12, ZHM^{+}10, ZBK^{+}15]. paucispinis [HRT12]. Pawlik [SL15a]. PBDE [MHN^{+}12]. PCB [MHN^{+}12]. PCBs [CLG^{+}19b]. pCO [MSM^{+}12, CESC14, DWL19, EES15, HSRP11, HYC^{+}14, LGW^{+}15, MKM^{+}16, PKB16, SHP^{+}14, SCSC16, TPKC18, VJM^{+}16, WMMC14]. PCR [NNJ^{+}13, PW11, WRB^{+}15]. pdf [Ano17a]. pea [DDDR10, DMHS^{+}14, JDD^{+}13]. pea-crab [JDD^{+}13]. peacock [GFC^{+}14]. peak [POL^{+}11, TSS11a]. peaks [HFH10]. Peale [VHHGF11]. pealeii [GBH12, SHM^{+}12, SJ10]. pearl [LPK^{+}16]. pearly [ACRLB12]. Pecten [LAL^{+}18, RJR18, SDH13, WA19]. pejus [JPF12]. Pelagia [LFE^{+}14, PTF^{+}14, TFM^{+}18]. Pelagic [SCO^{+}12b, AGK13, Ano18a, BI15, BPM^{+}18a, BBW^{+}17, BGIA19, BL13, BCHG18, BMF^{+}15, BLC^{+}16, BBKP10, BDB^{+}12, CMIMGS15, CMV^{+}11, CD13, CNÁCP12, CSC13b, DWLH13, DVSM10, DKF^{+}15, DRWQ12, EFOM13, FHTY19, FBS14, FTB14, FRV^{+}13, FSY^{+}10b, GSP^{+}15, GGK^{+}16, GRD11, GCDH19, GMR13a, GMR13b, HSDB13, HBCK13, KSFL14, KPW^{+}13, KWDD10, KHS10b, KBB10b, LBVH13, LR15, LJM^{+}12, LCM^{+}16, LALÁC11, LRM^{+}14, MM13a, MC16, MRL^{+}17, MBRO10, OWW^{+}17, PGR^{+}10, PPX17, PDCH15, PSS19, PMS10b, PCLB18, RKFD15, RPP^{+}13, RP15, RKM^{+}10, RGB^{+}10, RLNLAF17, RGDNGS11, RQN^{+}14, SDM^{+}13, SHS^{+}17, SHR^{+}11, ST11a, SWH^{+}18, SK11, SGP13, SMJP^{+}15, SLW15, SSM^{+}16b, SBM^{+}18, TSS^{+}12, TTS17, TSA^{+}18, TMP^{+}13, WEAG^{+}15, WHH12, YBH^{+}11, ZMS^{+}14, vdMPRM12]. pelagics [CGM^{+}19, GRL^{+}15]. pelagicus [DPY^{+}16]. Pelagos [LPC^{+}11a]. pelamis [RTCF^{+}17]. Pelates [BH15]. pellets [GSD12, PMBH11, SRRS12, SMR14]. Penaeus [GVR^{+}14, ODGL^{+}12]. penetration [LAALH10]. pengoi [KKL13]. penguin [ARGB13, BST^{+}19a, CCB^{+}17, CHP^{+}11, CFPF^{+}14, EPS11, HMR^{+}16, KHW^{+}12, MRL^{+}17, PWB^{+}19, SDM^{+}13, SSHB16, SE16, SEFS15, WLB^{+}15]. penguins [ALWD15, ADL^{+}18, BDNA10, BBH^{+}13, BM13, CTS^{+}15, DHSL11, DED^{+}15, DCW^{+}10, GMK18, HBF^{+}10, HPR^{+}12, LOS^{+}19, LFN^{+}12, PHHL19, PLNE11, PTP^{+}15, RB18, RPR16, SWFQ12, SBT12, SUB^{+}13, SDCC12, TIN^{+}18, TCTB11, TP19, WHPT12, WCRCR17, WRR10, WPMP14, vERRP16]. penicillatus [PRLR^{+}18]. peninsula [MJTC^{+}12, APC14a, CHP^{+}11, CDCZ16, CDC18, EZZ^{+}12, FTS^{+}13, GSD12, HPR^{+}12, HBK^{+}12, KAD^{+}15, MGFÁS^{+}10, MBB^{+}12a, MSM^{+}12, NHSG11, NDD^{+}13, PHD^{+}13, RWG15, RQN^{+}14, RSC14, SDM^{+}13, SV13, SLL^{+}18, SOT15, SA17, TMGCM11, VKY^{+}12, ZSEV17, RUS^{+}13]. penis [HWVM19]. Peponocephala [MHHB^{+}17]. peppery [SCO^{+}12a]. Perca [SSS^{+}10]. perch [MJS17, SSS^{+}10]. Perciformes [HPL12]. percula [SBB18]. perennial [AS15, GM13b, KKPL14]. perennis [AED^{+}10]. perezi [SHMOS17]. perfect [SGS^{+}12]. perform [GBE^{+}17]. Performance [MDNL^{+}11, AFRAG16, CPPM12, CG11, DMM^{+}10, DAWS19, FSF^{+}17, FAT^{+}15, FJ12, FDP^{+}16, FTF15, dlÁGRB^{+}19, GFP19, GOG14, HWTH12, LRB^{+}18, MWW^{+}14, NDAP19, Nor13, OF10, PLA^{+}12, RNB^{+}16, RP17, SUB^{+}13, SK11, SNS12, SJL13, TGP^{+}16, TFH^{+}19, VNK^{+}12, WHPT12]. period [Ano18a, BK14, DKAC^{+}12, DFTZ^{+}15, GLS^{+}15, KTNO19, MPS^{+}10, PPB10b, RCG^{+}15, SGP^{+}17, WCRCR17]. periodic [LBAK13]. periodicity [DT15, RWR^{+}15, SHM10b]. periods [BAA12, DRD^{+}18]. Periphylla [UAK^{+}14]. periphyton [ANB13, RJB^{+}14]. peritidal [RSP17, WPRS18]. Periwinkle [Mou17, ADRA19]. periwinkles [BMTF10, ZWP^{+}17]. permeability [ZPLW11]. permeable [BKC14]. Perna [TMS^{+}16, vdMPM^{+}10]. Persian [BBB^{+}14]. persistence [AN11b, BPW^{+}18, BOS18, DM17, ESW14, GVGH15, HSS^{+}19, KW10, LD10, LF14, LTCH10, MWWS17, PCP^{+}15, SCD^{+}17, SV13, SCK^{+}17, WBHL10, ZBY^{+}14]. Persistent [DWK13, HMAC^{+}11, MMBA13, OKF^{+}18, TGP^{+}16, vdMHW^{+}10, BPC17, HEMJ18, LBM^{+}13, PHAZ18, SGC17]. persisting [CED^{+}19]. persists [WHA^{+}13]. ‘Personality’ [EBB10]. perspective [HSMS^{+}15, HMBP18, KO12, LEB14b, NBM^{+}14, SAWK18, dlGMD^{+}17, vDCFM10]. Perspectives [LFRF13, WSS^{+}16]. perturbations [PA15]. pertusa [BJ14, GSC14, LLv13]. Peru [FKRJ13, LCC^{+}11, ZHM^{+}10, ZMRCP^{+}17]. Peruvian [ASEB11, CGR^{+}15, PBC^{+}18, WBS^{+}12, ZHM^{+}10]. peruviana [PDC^{+}17, PCF^{+}19]. Pessimistic [INK^{+}17]. Peter [Ras10]. petrel [HHS12c, KPr^{+}18, LBJ^{+}16, PJP^{+}11, RPM^{+}13a, RCP^{+}16, SABMG16]. petrels [CCJR14, CPM^{+}17, FTT^{+}17, PDP^{+}10, TDB^{+}14]. petroleum [BBKP10]. pH [BDAH10, BCR15, CPTL11, CLMB13a, CPHH15, DTG16, DWV^{+}12, HP13, HSA^{+}15, KBW^{+}15, LMZ16, LLF^{+}18, RNH15, WBG10, ZGK^{+}19]. Phaeocystis [RSS^{+}16, SLL^{+}18, ZXF^{+}16, vdPLdV^{+}11]. Phaeophyceae [LS16]. phagotrophic [BC15]. Phalacrocorax [LANE18, PRLR^{+}18]. Phalaropus [TR13]. Phallusia [GKE^{+}19]. pharmaceuticals [KOPW15]. Phase [DABP10, KCGC14, KRdVG^{+}18, LGS10, MKS18, PIGK16, RSS^{+}16, SHB^{+}10a, SQRDLC18, ZBC10]. phasic [RNB^{+}16]. phenolic [SV12]. Phenological [BDB^{+}12, Ric12]. Phenology [DB14, PPJT15, TSS^{+}12, DMvB10, EPS11, FTT^{+}17, FEE^{+}11, HFN14, LFN^{+}12, MJA^{+}18, MLML^{+}19, OBM15, RWO13, SNS12, YSH^{+}16]. phenomenon [BPP^{+}18, POS14]. phenotype [BWH^{+}17, CRF13, PHD^{+}14]. phenotypes [MGM10]. Phenotypic [LAB10, SD17a, BVO16, BWCZ11, DED^{+}15, GRN^{+}16, GOG14, LHPP16, WGLDB15, ZBB^{+}11]. pheromone [HBHF^{+}11]. Philippine [FTWA17]. Philopatry [CCC^{+}18, BSC^{+}13, FABH^{+}16, NCK^{+}14, OBA^{+}18]. Phlebobranchia [GKE^{+}19]. phlorotannins [HKJ17]. Phoca [CSF13, GWK^{+}19, HMAC^{+}11]. Phocoena [BEG^{+}13, KLBT^{+}16, KJC^{+}15, MMH^{+}17, MMD^{+}13, NWH^{+}12, ST08, vBBW^{+}14]. Phoenicopterus [GPV17]. phosphatase [PK13]. phosphate [TGM^{+}17, VBJR^{+}13]. Phosphorus [PBF10, BMV19, DBDK14, LGC10, RHF19, Rei12, VARO12, WBLS^{+}11]. photic [ESH^{+}10, GSMK15]. Photo [BPLT14, CJS^{+}12, PUP16, SVIH^{+}10, TNS^{+}12]. photo-autotrophic [SVIH^{+}10]. photo-identification [CJS^{+}12]. photo-oxidative [PUP16]. Photo-physiological [BPLT14]. Photoacclimation [GWM^{+}16, VDMLU13, ÁMLUN16]. Photochemical [GD15, PHD^{+}14, VAHL11, GAF^{+}16, vdPLdV^{+}11]. photogrammetric [NSC^{+}19]. Photoinactivation [FSFS18]. photoinhibition [RAH^{+}10]. photoperiods [HTHM18]. photophysiological [PJDC^{+}13]. photophysiology [KdVvdP^{+}12, WNF^{+}11]. photoprotection [GGL^{+}17]. photorepair [RAH^{+}10]. photorespirometry [TS10]. photosymbiont [RUP11b, RUP11a]. photosymbiont-bearing [RUP11b, RUP11a]. Photosynthesis [PR11, BHA^{+}11, NKH12, PKB16, RRS15, RS16, SCM13, XTW^{+}15]. Photosynthetic [HHH11a, AvDL^{+}19, AFRAG16, ADB10, GH11, LKLC11, SOB16, SHDR13, SJL^{+}15, YSK^{+}10]. Photosystem [WMS^{+}11, KAP^{+}14, PCNB16]. phycotoxins [TSF^{+}10]. Phyllopteryx [SCMSB^{+}11]. Phyllospadix [BBWG12]. phyllosphere [AFC^{+}16]. Phylogenetic [AOW11, FC13]. Phylogeny [LGBSC11, RT12]. Phylogeographic [HSMS^{+}15]. Phylogeography [HIKK12, LDAE12, NLKZ12, FCV^{+}16, FZI^{+}16, HMW^{+}12, LGBSC11, LLD^{+}12, MvdHBM12, RBSL11]. phylotypes [RAH^{+}10]. physalus [SRP^{+}19]. Physeter [EPC^{+}19, FNF^{+}18]. Physical [BWH11, FHJ15, GMD16, GC13b, LSN^{+}19, LMd12c, MSGGE10, MN10, SM10a, YRM^{+}12, BGL^{+}11, BKF15, CLB^{+}14, DKB^{+}12, FRV^{+}13, GWM^{+}15, HWR^{+}14, KAAI15, KPB16, LPB^{+}16, MV12, PMML12, PMYH13, RMME10, SRR15, SMP10, TYG12, TPVCV19, VCVK^{+}10, YPK15]. physical-complexity [TPVCV19]. Physiological [FRRÁSL11, FSR10, GSCH14, LSQ^{+}13, LCL^{+}11, MJW19, RCE^{+}13, RCF17, SLIGG18, STM^{+}14, SBO^{+}18, Whi11, XGLH14, BMPC15, BPLT14, CGS^{+}19, CY13, DHGW15, dlÁGRB^{+}19, GMSL^{+}14, LH14, LBP^{+}19, LWG14, LRC^{+}13, MVCD14, MVL^{+}16, RPRN^{+}16, SUPCB11, SPM^{+}13, SMHB14, SMF10, TFP14, ZFC^{+}13]. physiologically [MD16b]. physiology [CvOB18, DTL^{+}15, FSFS18, GLW^{+}17, HCDR^{+}15, LBM^{+}18, LKS12, MRSD15, OMBP10, SBH^{+}19, YSK^{+}10]. phytodetrital [CLS19]. phytodetritus [ENO^{+}11]. phytoflagellate [BBRBG10]. phytoflagellates [HP11a]. Phytoplankton [BHAC17, CRS^{+}11, CL15a, CMM^{+}18, FW11, JCZ^{+}12, LCTD^{+}15, LEKT12, MFI11, MMEM14, Pro17, RNK^{+}11, SGS^{+}17, SP13a, SPH^{+}17, SLD^{+}13, SKN^{+}17, SLR^{+}14, TL18, TCKY19, UTS^{+}18, WKX^{+}19, ALANGT10, AvDL^{+}19, AFRAG16, ÁMLUN16, AGM^{+}11, ADB10, BTS^{+}14, BDS18a, BDAH10, BW15, BCB^{+}17, CLW^{+}11, CRRD^{+}13, Chi11, CHLR16, Dao13, DVS13, DLL^{+}18b, DPO15, ETD^{+}18, FJC10, FSD^{+}14, FMJA^{+}18, FLSG^{+}15, GMSL^{+}14, GGLT15, GRD11, GWM^{+}16, GMD16, GLB^{+}10, GLY^{+}14, GFW^{+}14, HU13, HDWF10, HMM^{+}15, HGB^{+}13, HCJ^{+}11, HORRLS^{+}14, IVD^{+}10, JDQS11, JPAM^{+}15, KAD^{+}15, KWS^{+}13, KLD18, KJW^{+}15, LSY11, LSV^{+}10a, LHS^{+}14, LN11, LPC19, LLRC^{+}12, LS10a, LLH^{+}11, LRTLF12, LSD^{+}13, LNIH16, LSRRCM13, LMH12, LRSB19, LBCL10, MBM^{+}14, MWA17, MGFÁS^{+}10, MVTS^{+}12, MPB15a, MRSD15, MSS^{+}14a, MDE13, MER14, MGG^{+}14, NFR^{+}14, NMCP17, NLGE18, OGLH17]. phytoplankton [OTK^{+}13, OBÁSV18, OCMT13, PDO13, PA15, PMS15, PK13, PCH11, PFL18, PSPK10, Rei12, RB19, RLH^{+}13, RGB^{+}11, RGP11, SOB16, SGMG^{+}14, SLIGG18, SLL^{+}18, SMHB14, SJS^{+}11, SPSP10, SSAB^{+}15, SHSK11, SH10, TFFPM18, TG10, TMGCM11, THMF^{+}15, TRL^{+}15, TTS17, TBH^{+}17, TPSS12, TSF^{+}10, VLM14, WNG^{+}19, WS10, WLHW10, XTW^{+}15, YKG^{+}15, ZGK^{+}19, ZIT^{+}17, vdPLdV^{+}11]. phytoplankton-rich [GMD16]. pico [TDM16]. pico- [TDM16]. picophytoplankton [HRPBLT18, KdVvdP^{+}12, RKC16, TLFS14, VDMLU13]. Picoplankton [LS15, CLHW14, MCHC^{+}16, WS10]. Piece [Chr17, DC17, Lap19, MBCB^{+}19, QBW18, TVA^{+}19, WFD19]. pier [AGK13]. pigeye [KHS^{+}11]. Pigment [GGL^{+}17, ZFRG12, AFMR^{+}13, BCGWS19, DVA^{+}18, FMJA^{+}18, LN11, LRTLF12, LCS^{+}10]. Pigment-based [ZFRG12]. pigment-determined [LRTLF12]. pigments [AFRAG16, JNN12, LAL^{+}18]. pilchardus [CSR^{+}14, GN11, GCC^{+}16, NPMS11, NIPS12, SFTG14, TGI^{+}11]. pile [BDBN11]. pileus [SM15]. pilot [HER^{+}19, MMFP^{+}15, NSC^{+}19, OAB^{+}19, TFB^{+}17, WVM^{+}14]. pilsbryi [TCWC13]. Pingers [CASG^{+}19, KJC^{+}15]. pink [BRB11b, CHB^{+}13, CZH^{+}13, KSA^{+}13, LCCT19]. Pinna [Kat16]. pinnatifida [SJHH^{+}17, TS12]. pinniger [KFW^{+}18]. pinniped [DJJ^{+}16, GJ13, RT12]. pinnipeds [MWZ^{+}16, VVB^{+}10]. pioneer [MLDG12, ZBY^{+}14]. Pisaster [BLD^{+}16b, MWGH15]. piscatorius [CALB10]. Pisces [STYT10]. piscivory [PSS19]. pistillata [TNS^{+}12]. pits [TKDM16]. PIV [WHA^{+}13]. place [GPW^{+}10, HSS17, RPR16]. placed [KMO19]. Placopecten [BCW19, CWS13, HS11a]. plaice [ATF12, CTB13, CBF^{+}13, GBR11, MMRC16, MDR16, vdVFK^{+}10]. plain [LBAK13]. plainfin [HPB^{+}18]. plains [WMCC15]. plakorticola [SIHH10]. Plakortis [SIHH10]. plana [SCO^{+}12a, SAL^{+}12, VCDP11]. planci [KHH18, RPB17, WGLDB15]. planifrons [VDG^{+}15]. planktivores [BPM^{+}18a]. planktivorous [BJS^{+}11, DKB^{+}12, SBBB^{+}18, ZMS^{+}14]. Plankton [FF10, MJT^{+}12, PLJ^{+}12, TGBM^{+}16, TPF^{+}13, VSM11, ZTNS13, AAAE17, AFLM12, AF14, ANJ^{+}11, AGTS^{+}12, ANRT10, BLHRP^{+}19, BHE14, BCM19, CL15b, CFBG14, EH13, FAZ^{+}14, GCC^{+}10, GMD16, GWGC16, KWS^{+}13, LMD12a, LLCM^{+}12, MDF10, MMBA13, POL^{+}11, RBC^{+}14, RNK^{+}11, RPM12, RSW13, SMK^{+}18, SMM^{+}14, SJS^{+}11, TSS11b].

Planktonic [DAL^{+}10, KCGC14, PAFC13, RWR^{+}15, TFP14, BQRB12, BM11, BVD^{+}14, DTM^{+}12, EMMSL13, GMB^{+}16, GSMK15, HT15, KAJP^{+}15, MLC14, SWMK15, SCRGP11, SAD^{+}16, VARO12, Wir12, Wir14]. planktotrophic [CSQRA11, FSB15, PSB17]. planktotrophs [MWHM18]. planning [JMK^{+}18, LCF^{+}14]. plant [AFS19, BWCZ11, CPF^{+}13, CFF^{+}18, DGA^{+}15, FWW^{+}16, GM10, HBHS^{+}18, JBT^{+}16, JNN12, LLH^{+}11, MRL^{+}10, PRA10, SRB15, VNK^{+}12, WHB^{+}19, ZZG^{+}19]. plant-herbivore [CFF^{+}18]. planted [KKS^{+}12b]. plants [BPP^{+}18, CS13, FSH^{+}16, RZB18]. Planulation [SHM10b]. Plastic [CD13, DA11, BVO16, SCL^{+}19, SHMB15]. Plasticity [DMD^{+}12, FHL^{+}17, RPRN^{+}16, Ano17c, BCGWS19, BPWF14, BPB18, BVQ12, BWCZ11, CCJR14, CG17, CMH14, DMC^{+}17, DED^{+}15, FBGH18, FGOS16, GPV17, HA18, KP18, MCR^{+}13, PGR^{+}10, RCF16, SD17a, SES17, WGLDB15]. Plata [DFTZ^{+}15, LGD10, PNI14]. plate [BMV^{+}11]. Plateau [GMES18, OHCG14]. platessa [CTB13, CBF^{+}13, GBR11, MMRC16, MDR16, vdVFK^{+}10]. platform [BBKP10, MRC^{+}19]. platform-associated [MRC^{+}19]. platforms [EFR12, Fuj16, RCMB18]. Platichthys [CLM^{+}13, Dan11]. Platycephalus [BFDM18]. play [MSLR13]. playbacks [CSV^{+}15]. played [FHM^{+}11, SCHO11a]. plebeius [GSO^{+}18]. plebejus [ODGL^{+}12]. Plectropomus [BK14, CCM^{+}15, MHS15, MTS^{+}17]. Pleistocene [BCC^{+}11]. plenty [GRL^{+}15]. Pleuragramma [BLJP14]. Pleurobrachia [SM15]. Pleurobranchaea [TWM^{+}15]. Pleuronectes [CTB13, CBF^{+}13, GBR11, MMRC16, MDR16, vdVFK^{+}10]. plicata [SF14]. plicatum [CY10]. plots [BVGH10]. plume [BMD^{+}11, BWRP14, DFTZ^{+}15, KEBB10, KVR^{+}18, OGLH17, PHAZ18, VSM11]. plumieri [PFVFF^{+}15]. plumierii [YAL12]. plutei [CG11]. PoCHL [BBC^{+}10]. Pocillopora [CE14, LE10, LHW^{+}19, MAP^{+}12, TKS^{+}14, TH11]. pocilloporid [FHL^{+}17]. pocket [FRP^{+}17, SMM^{+}15]. pockmarks [DDC^{+}15]. podocyst [FLS^{+}18]. point [FSM^{+}14, SFM^{+}19, ZB16]. point-intercept [ZB16]. points [LFFL15, PEB^{+}14, SLS18, SH12]. poisoning [VTZ^{+}13]. Polar [FLE^{+}11, KPH^{+}15, BDR19, DAW^{+}11, FL17, FKA^{+}12, HTHM18, HCLB18, IVD^{+}10, JHL^{+}19, KPM^{+}11, ORE^{+}17, SBH^{+}19, WTC^{+}15, dlGMD^{+}17, TGBM^{+}16]. policy [CPMR11, KPW^{+}19, PHB^{+}11a]. Pollachius [AFOVRB14, CEK^{+}16, LANE18]. Pollen [VMW^{+}10, vMVSW12]. pollination [vMVSW12]. pollock [BZC^{+}12, BBMH13, NRK^{+}19, RDK14, SDAH12, Urb12, WDM^{+}18, YFC^{+}13, ZBC10]. pollutants [LBM^{+}13, vdMHW^{+}10]. pollution [LPKT^{+}19, MZC^{+}10]. polyamine [LNIH16]. polybrominated [MHN^{+}12]. Polychaeta [CML^{+}12, DBCV^{+}12]. Polychaete [NHSG11, Ber12b, BCK15, CRJ13, HKS^{+}12, JLQ^{+}14, LRS^{+}12, LSQ^{+}13, MH11a, OBI^{+}16, PAF^{+}10]. polychaetes [GMG^{+}16, KHD^{+}11, LLF^{+}18, VCQM^{+}16]. polychlorinated [MHN^{+}12]. polycystina [Bol17]. Polydora [PAF^{+}10]. polyedrum [BCM19]. polyembryonous [PHB11b]. polygon [Bre14]. polygynous [SKSH19]. polykrikoides [GG16, RTC^{+}14, TG10]. polylepis [HGL15]. polymer [GRG^{+}12]. polymeric [KCP^{+}11, MPBSB18, UFP^{+}10]. polymerus [ARMA15]. polymorpha [SSS19]. Polymorphism [CGG10a, CSQRA11, RAB^{+}10]. polynemids [HMW^{+}12]. Polynesia [ELA10, TPA^{+}13]. polynya [MJA^{+}18, STP13, STBP13]. polyp [FLS^{+}18, IMU17]. polyphekadion [RTM11]. polypoides [HCM^{+}11]. polyps [CG17, HMBP18, IMU17, SB18, TPKC18]. polystigma [FCV^{+}16]. polyxystra [LBD^{+}15]. pomacentrid [SSB^{+}18]. Pomacentridae [VHKT15]. Pomacentrus [LDAE12]. Pomatomus [JCM13, SRSH^{+}18]. Pond [HU13, BM10a, JMGG13]. ponds [SCRGP11]. Pontomyia [SLC11]. pool [HKS^{+}12, MK13, RWH14, VDEO18, YYU^{+}17]. pools [BC14, FSR17, GC13a, GNC^{+}10, HS16, KCP^{+}11, LNIH16, RET^{+}19, RSP17, SGP13]. poor [FRMS17, GMES18, SMSL17]. pop [ADHP12, FLK12, MAW^{+}11, MDNL^{+}11]. pop-off [FLK12]. pop-up [ADHP12, MAW^{+}11, MDNL^{+}11]. Population [AGH^{+}12, BHUCMO19, BSRDR10, BLHL15, BSH13, BPB^{+}12, BFH^{+}13a, CBKM16, CPA^{+}11, CHST16, DSH14, DBS^{+}13, DHR^{+}13, Edm10, EHD^{+}17, ETB13, FCV^{+}16, GBH12, GSM^{+}11a, GDJdBSL19, HHLT10, HCBV^{+}11, HPL12, JE10, KJBS12, MJTC^{+}12, MFSRR14, MBB^{+}12a, MEK11, MKK^{+}11, NJJ^{+}11, OC17, RMKR13, RSHM10, SPSG14, SHM^{+}12, TKS^{+}14, TKBN10, WWAS18, WHK^{+}11, WTC^{+}15, WSF^{+}12, WBHL10, ARvdW^{+}16, AFF^{+}15, BGP^{+}16, BAL14a, BCS^{+}12, Bea11b, BWS11, BPWF14, BD13, BD15, BEFC15, BVAB^{+}17, BPW^{+}18, BVQ12, BDVP19, BDC10, BBBHC17, BANE^{+}13, CPSC^{+}15, CLL15, CBT13, CCNB15, CGH12, CFSM^{+}13, CMHH17, CHB^{+}13, CWMP13, CCG^{+}19, CCM^{+}12, DXF^{+}18, DBB11, DCC^{+}10, DH17, EL19, ET11, ERB^{+}13a, FTO^{+}15, FKP^{+}15, FTWA17, FABH^{+}16, GPM^{+}13, GAB^{+}14, GHGG11, GGD^{+}11, HW11b, HSS18, HRM^{+}11, HSDB13, HMBP18, HSMB17, HPM14, INK^{+}17, IHD^{+}16, JCRC10, dDTC^{+}17]. population [JBL^{+}16, JHL^{+}19, JWBS10, KHH18, KM10a, KHW^{+}12, KW10, KLNJ15, KPM^{+}11, KCS^{+}17, KPP^{+}16, KCL^{+}11, LTP^{+}18, LTN^{+}11, LCBK^{+}13, LBWB^{+}13, MDLM17, MdSW^{+}16, MKK13, MCPC19, MC12a, MPZ10, MHHB^{+}17, MHPvdH19, MGGSP15, MTB^{+}17, MGM10, MMWG13a, MMWG13b, MMFP^{+}15, MBG^{+}15, MPD11, MOK^{+}18, MN10, NKH^{+}10, NSS10, OGGP14, PNA^{+}16, PMB^{+}17, PNF^{+}12, PAP^{+}12, PHHL^{+}11, PBM^{+}18, Rai12, RABS19, RWM^{+}10, RA10, ROKS14, RWH14, RSP^{+}16, RBBG12, RKvW10, RMT^{+}13, SS10a, SVAV17, SES12, SSBA12, SHT^{+}18, SG18a, SBB18, SWH^{+}18, SDST16, SCK^{+}17, SKW^{+}14, SGT12, SMMFH16, SMVM^{+}12, SWB^{+}10, SEFS15, SRCE^{+}10, SMR^{+}11b, SS14, SDB12, SIHH10, SN17, SBS10, SNT^{+}16, TCN^{+}10, TSCvS^{+}15, TFS14, TCTB11, TJS17, VAB^{+}13, VCDP11, VBV^{+}10, WWB14, WEA^{+}17, WMH17, WFH11, WHH12, YWiIK15, YBH^{+}11, dDTA^{+}18, vdMHW^{+}10, vdWTWS15, ERB^{+}13b]. population-level [SBB18]. population-scale [SWH^{+}18]. Population-specific [WTC^{+}15]. Populations [dSM11b, APC^{+}14b, ARA^{+}10, ACMD11, BCOMKJ18, BCM^{+}11a, BBD^{+}14, BHN^{+}14, BAVSP11, BHF^{+}16, BDD17, BADH^{+}12, BOKJ18, BD11b, BJL^{+}17, Bru10, BBMM14, CKM^{+}16, CED^{+}19, CWS13, CDCZ16, CS11a, CBL^{+}10c, DMP12, DLC^{+}13, DDK^{+}18, DDD^{+}11, DBE^{+}18, ETB13, FS14, FHJ15, GDH^{+}19, HPA17b, Har16, HW11b, HTH10, IMU17, JTTB^{+}18, KKPL14, KNS^{+}15, KBK^{+}10, LPB^{+}16, MPB14a, MCBR18, MGPC13, MEK11, MAP^{+}12, MHSG18, MDKM13, MN10, NHS^{+}17, NvOH13, OCS^{+}13, PHD^{+}13, PLJ^{+}12, PJMP10, PTP^{+}15, PGR^{+}17, RAB^{+}10, RSTV^{+}13, REG^{+}11, SVG16, SB10, SWK18, STT14, SHM^{+}10a, SBM^{+}16, STS^{+}18, SH17, SKPS^{+}16, SB15b, TPB^{+}16, TWM^{+}15, TPC^{+}16, TPC^{+}15, TdRH17, VMC^{+}19, WWKS15, WOHJ^{+}15, WRP^{+}17, WNF^{+}11, YTM^{+}15, dQSC13, vdWHG11]. porbeagle [RB11]. pore [CGR^{+}15, JWR^{+}10]. pore-water [JWR^{+}10]. porewater [CAKD14, JLQ^{+}14, RF14]. Porifera [CL10b]. Porites [BKB11, BC13, CTRBLC^{+}13, CE14, CSJP17, CSM12, DDD^{+}11, Edm10, EES15, JRK13, PKW^{+}11, RRL^{+}19, SLS16, TBL16]. Porites-algal [BC13]. porpoise [AvBBL^{+}16, BBD^{+}14, BEG^{+}13, CASG^{+}19, IME12, KLBT^{+}16, KJC^{+}15, ST08, SNNS^{+}12, WBS^{+}17]. porpoises [BDvGW16, BDBN11, BHD^{+}13, BDD^{+}18, DTC^{+}17, DL13, MMH^{+}17, MMD^{+}13, NWH^{+}12, NWD13, NTS^{+}18, SBG^{+}16, TW10, vBBW^{+}14]. portfolio [YHTS12]. ports [VK11]. Portugal [ADB10, BAVSP11, BSD^{+}14, RCGE12]. Portuguese [BAM^{+}10, CQR^{+}17]. pose [BVM^{+}19]. Posidonia [AFC^{+}16, AFS19, AMC^{+}11, BRA^{+}10, BBC^{+}10, CESAM12, CSS^{+}17, DBCB11, GMC^{+}12, GCBB13, HCK12, IOS^{+}12, JSOK18, JMPZ10, MATR11, OFM^{+}18, PRFA11, SMR11a, SKA^{+}14, VBŽ^{+}17, ZKS17]. Posidoniaceae [SKA^{+}14]. position [ABN14, APNC^{+}15, BKC^{+}19, CAB^{+}19a, DGC18, GKHM13, GCS^{+}19, ILDL10, JW11, MBJ16, MBM^{+}14, MWWS17, QGM^{+}14, TBS^{+}14, VWEP18]. positional [TGSM12]. positioning [WGS16]. positions [LZF^{+}14]. Positive [HYC^{+}14, JS10, KVY^{+}13, OSK15, GBJ^{+}14, KTH^{+}18, SS12b, TVH^{+}10]. possibility [CKR^{+}11]. possible [BSA^{+}16, FWW^{+}16, GJ13, HPN^{+}17, HLPS^{+}15, KNB^{+}13, KPB11, MHPvdH19, PR11, SKGS11]. Post [AB17, ADL^{+}18, AGHC10, DPC13, HIMH15, MASW^{+}11, SHF12a, SKFY18, TPA^{+}13, AED^{+}10, AY10, BGIA19, BJRH10, GVR^{+}14, GJ10, HGG^{+}18, HHNI17, HI14, HBP^{+}16, JH11, JH14, LSJ^{+}15, MJT^{+}14, MIA^{+}11, MW16a, NLLS15, PSHG16, PMB^{+}10, PA13, RWAP16, SWMK15, SSB^{+}11, SYK^{+}13, TBMG17, TGP^{+}16, VNN10, WYS^{+}10, WBG10, vdMPM12]. post-breeding [MIA^{+}11]. Post-capture [MASW^{+}11]. Post-fledging [ADL^{+}18, AGHC10]. Post-flood [DPC13]. post-germination [AED^{+}10]. post-glacial [HI14]. Post-harvest [AB17]. post-hatching [MJT^{+}14]. post-larvae [HHNI17]. post-larval [GVR^{+}14, HBP^{+}16, MW16a, VNN10, WBG10]. post-metamorphic [TGP^{+}16]. post-mortem [NLLS15]. post-nesting [SSB^{+}11]. post-recruitment [SWMK15]. Post-release [HIMH15, TBMG17, WYS^{+}10]. Post-settlement [SHF12a, SKFY18, TPA^{+}13, AY10, BJRH10, GJ10, JH11, JH14, PMB^{+}10, PA13, RWAP16, SYK^{+}13, vdMPM12]. post-smolts [HGG^{+}18]. post-spawning [LSJ^{+}15]. Postelsia [TKBN10]. postglacial [BS18b, BS18a, DHJ^{+}17]. postlarvae [CZH^{+}13, KPS14]. postlarval [LS10b]. postsettlement [HCG^{+}10]. Potential [AGTS^{+}12, BDG^{+}16, BM16, DBC^{+}12b, DFAH11, DPD^{+}12, Fuj16, GKC11, GBJ^{+}14, GPRR17, HOE^{+}12, JBT^{+}16, RTDUM16, SSB^{+}18, VBM^{+}11, VSG^{+}11, ASR^{+}19, ATF12, Ano18b, ATO^{+}14, BFG^{+}19, BMJH12, BSB^{+}11a, Bet13, BSCH14, DSG15, DE17, DCIG18, DMHS^{+}14, FBC^{+}15, GEF^{+}12, GB12, GKPT16, GRN14, HFM^{+}15, HMF^{+}18, KFBM19, LGMB^{+}16, LHU^{+}18, LNH^{+}17, LMd12b, MBJ16, MZC^{+}10, MBF16, NNT^{+}10, ORR^{+}14, ODL^{+}13, PMK^{+}15, PMS10a, PKBS14, RNB^{+}12, RGEA^{+}15, RTJ^{+}13, SMD^{+}16, SMJ^{+}14, SNS^{+}13, SLL^{+}18, SN17, TJS17, TBMdV12, WMS^{+}10, WLB^{+}11, WNNH10, WBS^{+}17, XLXS17, YYD^{+}15, YHTS12, ZPPM17]. potentially [LRHP18]. potentials [PTF^{+}14]. pots [HW12a]. Potter [WVD^{+}17]. pouchetii [RSS^{+}16]. pout [MRE14]. Power [GSR10, BPP^{+}18, CGH^{+}16, LLH^{+}11]. power-law [CGH^{+}16]. powered [DFB^{+}17]. powerful [CGG^{+}10b]. practical [JGCRS12, Løk11, MS11]. practices [Løk11, RGEA^{+}15]. Practicing [PKS11b]. praeputialis [MC10b]. prawn [ASF13, HP13]. Pre [AY10, KHBL11, SHP^{+}14, AED^{+}10, DFTZ^{+}15, GAB^{+}14, LFJ^{+}14, LSJ^{+}15, PLNE11, WCRCR17]. Pre- [AY10, AED^{+}10, LSJ^{+}15]. pre-breeding [DFTZ^{+}15, GAB^{+}14, PLNE11]. Pre-existing [KHBL11]. Pre-hatching [SHP^{+}14]. pre-moult [WCRCR17]. pre-spawning [LFJ^{+}14]. precise [KPLD^{+}12, NMS^{+}18]. precision [BG12, LSFQ19]. Preconditioning [TBL16, RWG15]. Predation [BLMD^{+}11, CSB14, CM10, JBS12, LRRM^{+}11, MBV^{+}11, MRL^{+}10, RPHB15, RSH^{+}19, SIK16, ARB^{+}12, AM13, BSB^{+}11a, BS12, BC15, BPF^{+}18, BL17, BBW^{+}19, CE14, CGH^{+}16, CUCS17, DMP12, DQM^{+}19, DHG11, GBU15, GSP^{+}15, GRFSM12, GP12, GRN14, GMR13b, GGF^{+}12b, HHS10, HMS10, HMJ^{+}13, KCGC14, KB14, LST11, LH10, LCFR10, MGP11, MGP12, MBR^{+}13, MORM15, MJO^{+}13, MMDM12, MFD12, MMCW14, MB11, MCT16, NGSF13, ORM^{+}13, PMSGBL16, PT13, PW19b, PTF^{+}14, PDBB14, RM13, RFRE19, RDK14, RLM^{+}14, SRLV10, SHF12a, SES17, SKFH18, SHMB15, SM15, SJ10, SCF^{+}17, TKIO10, VTB12, WL11, ZFZ16, vDCFM10]. Predation-driven [RPHB15]. Predator [BAS^{+}10, BMJ16, BAT14, BQRB12, BB14c, CSV^{+}15, GSL17, HBI^{+}18, HA18, LBJI14, LCFR10, LIT13, MB16, NOB^{+}19, SES17, VE18, AB17, AB11, ARD^{+}19, ACB12, AL12b, AAJJ^{+}19, AH13, AWH19, BCG10, BHP12, BYAA13, BOC^{+}16, BMTC^{+}15, BKR^{+}15, BD14, BPJ^{+}15, BCV^{+}18, BWH^{+}17, BPW^{+}18, BDS^{+}18b, BBL11, BDVP19, BKAM12, BFG^{+}11b, CWW^{+}16b, CWB^{+}15, CGQB10, DGA^{+}15, DGMB10, DF15, DBHMG12b, DHF^{+}13, EvDB^{+}14, Gla10, HSPS14, HKI^{+}17, HHS^{+}12b, HM17, HRT^{+}14, HW13, HMR^{+}16, HBG^{+}17, HOE^{+}12, IDT^{+}18, JWT^{+}14, JLPG17, JDG^{+}17, KÖHSO12, KBW^{+}11, KCF^{+}14, KCH^{+}11, LZ12, LSH19a, LH11, LPJ^{+}13, LLBP15, LHB19, MM11, MJM^{+}14, MATM^{+}19, MH12a, MHR^{+}19b, MB11, NGR^{+}17, PGR^{+}13, PES12, PJMP10, RH12, RGdLB^{+}11, RSSvV19, SV13, SA14, SSR^{+}10, SCO^{+}12b, SP13b, SG18a, SGS^{+}19, SKFH18, SPBS14, SHF^{+}12b, SFL^{+}18]. predator [SBMÅ12, TB12, VDG^{+}15, WPK10, ZAMB10, ZBC10]. Predator-borne [LBJI14]. predator-controlled [DHF^{+}13]. predator-depleted [SFL^{+}18]. predator-dominated [HSPS14]. Predator-enhanced [BQRB12]. predator-herbivore [SBMÅ12]. Predator-induced [HA18, LIT13, SES17, VE18]. Predator-mediated [BB14c]. Predators [FNC^{+}14, MM11, ZKS17, ATF12, BHF^{+}16, BBH^{+}13, BH12a, BFL^{+}19, BL10, CMIMGS15, CBLM^{+}17, CMV^{+}11, CHST16, CHRB10, DGC18, DGMB10, DB14, DL14, EMH^{+}19, FHTY19, FHRC16, FRV^{+}13, FTvdK^{+}12, GHCC15, GS17, GSO^{+}18, Goc10, GCC^{+}14, HCM^{+}11, HBK^{+}17, HVK^{+}13, HKSD14, HW13, HSS17, JBS11, JMS^{+}15, KMP15, LCFR10, LCL^{+}17, LAH^{+}15, ŁCM^{+}18, MBR^{+}13, MRB^{+}15, OGLH17, OMBP10, PMK^{+}15, PMS10a, Ran18, RHE14, RCS^{+}17, RFRE19, RSP17, RBS16, SLF^{+}19, SHB^{+}10a, SNNS^{+}12, TKDM16, TKL^{+}11, TdRH17, Wir12]. Predatory [HEAI10, SCW^{+}15a, TG12b, BVO16, BFD^{+}16, BH16b, BSA^{+}16, CPID13, CD13, CHB11, CV11, CGM^{+}19, DKAC^{+}12, DAM^{+}18, FTKK14, FFS14, FWL15, GSV14, HFLI11, HM10, KKL13, MPLLE17, MWGH15, NKS^{+}15, SIK16, UHG12, WAW^{+}12, dGR11, dSM11b]. predict [BGP^{+}16, BFG^{+}19, Dav19, FSF^{+}17, FHRC16, GJ16, HLS17, HPH^{+}18, LM12, MSK^{+}19, TDPS13, YPK15, YBD^{+}16]. predictability [DMT^{+}14, LFG^{+}12, WAA^{+}19]. Predictable [KCF^{+}14, DAD^{+}11]. predictably [Sma13]. Predicted [MHN^{+}12, BFDM18, GHGG11, LBJ^{+}16, MKT^{+}18, RTMJ11]. Predicting [ASM^{+}12, AGV^{+}11, DER^{+}19, EFOM18, FL10, Fie14, HW11a, HGFA^{+}11, NTB^{+}17, PSHG16, SAWK18, SS10a, SHB^{+}13b, VGGD^{+}15, FGH^{+}13, HPP^{+}10, LWD13, LMR^{+}10, LMTD16, WHH^{+}13]. Prediction [LLP^{+}10, GMBL11, GMD16, LJK^{+}11]. Predictions [FHH^{+}17, FZI^{+}16, ML11, SKD^{+}11a, SMD^{+}13, STSS^{+}16]. Predictive [FJD13, GDE17, RZPH14, SATS13, WM10, DSG15, GMBL11, RHS^{+}19, RLH^{+}19, WY15, YChH11]. predictor [DABP13, LANER15]. predictors [DBHMG12b, HS15a, MLR17, RRG^{+}11, TM15]. predicts [Bre14, FJ11, YIKG10]. predominantly [DMK11]. prefer [LWP^{+}13]. Preference [GNBM11, MGP12, DF15, ESW14, GMK18, HMJ^{+}13, HDT^{+}12, LBD^{+}15, LS19, MM13c, NHG10, PBV^{+}11, PMM^{+}11, RBA^{+}18, SV15, WWAL13, ZH13]. Preferences [KA16, LSV^{+}10a, ASWB10, BEG^{+}13, CMC^{+}16, CPM^{+}17, HGFA^{+}11, IME12, JMB^{+}18, KRdVG^{+}18, MPKO11, PHH^{+}12, PPC^{+}12, RPM^{+}13a, RNFG10, RWAP16, SSM^{+}17, SSB^{+}11, SQRDLC18, SPBS14, SHM10b, SKS18, TRHM14]. Preferential [SHC^{+}10]. Preferred [IVD^{+}10, AIA18, GWK^{+}19]. preliminary [SOA^{+}17, YNK18]. presence [BLIM11, CBT13, CLL13, GSL17, HS15a, HS14b, ILDL10, MLC14, MIV^{+}10, ORWPR14, SBG^{+}16, SOC13, TBB^{+}19, vdMPM^{+}10]. presence-only [BLIM11, MIV^{+}10]. present [BOCB10, CLGCA^{+}19, KHW^{+}10, SGS^{+}17, SVIH^{+}10]. presents [SSM^{+}16a]. preservation [LMR^{+}10]. preserved [GBMW12, UDF^{+}18]. preserving [CHRB10]. pressure [AS14, AHR^{+}18, BAVSP11, BAE^{+}10, HGG^{+}18, Hv14, LMD12a, MPF^{+}15, RB17, SBT15, VPOA12, VK11, WB12, YOST11]. pressures [FRJ11, GMB^{+}16, LFFL15, SLSR10]. presumed [TBB^{+}19]. Prevailing [CDMB^{+}18]. Prevalence [MTG^{+}10, MMCW14, dG14, BTLF17, JBM^{+}13, MAP^{+}12, TMBP13]. prevent [GJ13, HM17, PLG^{+}16, SB15b]. preventing [MSLR13]. Previous [LCD11, LS14]. Prey [ACL^{+}16, BMTC^{+}15, BJS^{+}16, CBF^{+}19, EvDB^{+}14, LNP12, MBRO10, NSK^{+}14, PLF^{+}12a, REAN^{+}14, TOR^{+}10, WPK10, AL12a, ADO^{+}12, ACB12, AL12b, ASCHM19, BAS^{+}10, BLE^{+}13, BBKH^{+}11, BBBNT13, BMJ16, BSST13, BSBR^{+}12, BPW^{+}18, BB14c, BSD17, BKAM12, CSR^{+}14, CMIMGS15, CRRL17, CHE^{+}17, CPG^{+}14, CV11, CUCS17, CSC13b, CD14, DABP13, DBA17, DRT13, DGMB10, EC14, ES12b, FLC^{+}14, GE14, GP11b, GCSF18, Gla10, GvSGCK14, Goy14, GAC11, GCRP15, GVGH15, GLM^{+}19, GVGP^{+}14, HKI^{+}17, HBI^{+}18, HHS^{+}12b, HM17, HLPS^{+}15, HRT^{+}14, HW13, HM10, HPH^{+}18, HOE^{+}12, HRK^{+}18, IW15, Ing16, JCM13, JDG^{+}17, KCGC14, KA16, Kiø16, KWH^{+}18, KLN^{+}13, LHJH^{+}10, LZ12, LH11, LST11, LRBSJ15, LGJH12, LLBP15, LCFR10, LHB19, LDT^{+}13, MM11, MBR^{+}13, MJM^{+}14, MATM^{+}19, MH10, MMSC17, MAB^{+}12a, MGGSP15, MRB^{+}11, MBS^{+}13, MB18]. prey [MB11, MJHS13, NHG10, NOB^{+}19, ODP^{+}12, ØLH^{+}13, OTK^{+}19, OSL18, OWN^{+}15, PWNP^{+}11, PES12, RLN17, RET^{+}19, RSS^{+}16, RH12, RNFG10, RDM18, RSH^{+}19, RT12, RSC14, SGC^{+}13, SHPS^{+}16, SZ13, SPP12, SDKS19, SP13b, SG18a, SUB^{+}13, SGC17, SMJBØ11, SML17, SEFS15, SKPS^{+}16, SNNS^{+}12, TB12, TMT^{+}12, TR13, UW18, VGGD^{+}15, VHBG15, WSF^{+}12, WAW^{+}12, WJHS13, WL11, Wir12, Wir14, ZBK^{+}14, dSM11b]. Prey-capture [TOR^{+}10]. Prey-field [BMTC^{+}15]. prey-size [Wir14]. price [OCE^{+}12]. Primacy [OCPM12]. Primary [GCC^{+}10, HMS^{+}12, TS10, TFFPM18, VKY^{+}12, ATS14, AHH^{+}11, AHSG16, BTS^{+}14, BvHC^{+}15, CGOC11, CS10b, DZDB10, ERG^{+}11, FMH^{+}15, Gal14a, Gal14b, GHHH12, GME^{+}19, HJN^{+}10, HBHG10, HWB^{+}18, KSS13, KAMA13, KRD^{+}12, LMGJ13, LTSDHL12, LRM^{+}14, MCW^{+}16, MBC18, MZF^{+}19, MCW^{+}17, NFR^{+}14, PMB^{+}17, PLS^{+}10, PR11, Ree14, RB19, RKC16, RBC^{+}15, RS16, SLD^{+}13, SMF10, SDVL16, SKN^{+}17, SA17, RGP11]. Priming [TPC^{+}13]. Prince [AMM^{+}12, MIA^{+}11, MBB^{+}12b, Ran18, SNVH19]. principal [UTS^{+}18]. Principles [TTC11, Pör12]. Prionace [NCW^{+}19, QHN^{+}10]. Prior [BLD^{+}16b, BLD^{+}16a]. priorities [PvdMLC17]. prioritise [BM13]. priority [MJ17, MOH^{+}10]. pristine [BHTF11, FSDS10, HHH^{+}11b, HVK^{+}13, NMPM15, OV10]. Pristis [FKP^{+}15]. probabilities [BBM^{+}14, BWM^{+}15]. probability [BLIM11, HGS14a, HGS14b, PVZD^{+}11]. probes [FTvdK^{+}12]. problems [Fry13a, Fry13b]. Procellaria [PDP^{+}10]. Procellariiformes [DTL^{+}15, JJP^{+}13]. process [CSS15, HCJ^{+}11, NHAS15, SFM^{+}19, ST10b, ST11b, TPS10, TKB^{+}13, VPOA12, YSS^{+}18]. process-based [ST10b]. process-oriented [HCJ^{+}11]. Processes [DFS13, ASM10, BPB11, CSC13b, DBS14, DKB^{+}12, DAM^{+}17, GLB^{+}10, HBP^{+}16, HLD12, JH14, KW10, KJBS12, KS11, LB11, MUC12, MHGW13, MLR12, PA13, PH10, PHG^{+}14, RP13, RFBH^{+}11, RBM14, REBL16, RKvW10, SMP10, SGP13, SDMR19, SS11b, SLE^{+}13, WBB^{+}10, ZBY^{+}14]. Processing [GMT^{+}11, KTKM14]. Prochlorococcus [LAALH10]. produce [RSJ14]. produced [GBU15, HT10, Mat10b, PHB11b, SWWK13]. producers [GHHH12, HJN^{+}10, MBC18, Ree14]. producing [SE11a, WBS^{+}12]. Production [KS12c, MPBSB18, OCMT13, ASM^{+}12, ANJ^{+}11, AHH^{+}11, AHSG16, BTS^{+}14, BPM^{+}18a, BD13, BD15, BBH^{+}12, BLM^{+}12, CMP^{+}11, CSS15, CMW^{+}14, CRHV14, DAL^{+}10, DPY^{+}16, DJL^{+}16, DZDB10, DDP15, DHH10, EM11, ERG^{+}11, EMG^{+}15, FMH^{+}15, FOL12, FLSG^{+}15, GGF^{+}12a, GBM^{+}12, Gal14a, Gal14b, GL12, GME^{+}19, GGMGR12, GCC^{+}10, GCC^{+}14, Hat13, HSdBS10, HJMN^{+}12, HMS^{+}12, HBHG10, HEM^{+}12, HWB^{+}18, HSR^{+}10, JDQS11, JMGG13, JPAM^{+}15, KVR^{+}18, KSS13, KAMA13, KRD^{+}12, KS11, LHL11, LTSDHL12, LGM^{+}12, LCB^{+}12, LRM^{+}14, MHCRR18, MZF^{+}19, NKH12, ORR^{+}14, PPB^{+}10a, PCH^{+}14, RJB^{+}14, RWR^{+}19, RB19, RNK^{+}11, RBEGG13, RS16, RGP11, RSA^{+}12, RNSA15, SLD^{+}13, STP13, STBP13, SL15b, SV12, SMF10, SSD13, SAD^{+}16, SA17, SKD^{+}11b, TFFPM18, VAHL11, VCDP11, VKY^{+}12, VLM14, WLM12, WSLM13, WBI^{+}15, WBJ12, WEK14]. production [WPP11, WLHW10, YAL12, ZS17]. production-based [EM11]. production/biomass [BD13]. productive [BTS10, CPF^{+}16, FL17, JDQS11, LNF^{+}11, RUD^{+}16]. productivities [TIS^{+}15]. productivity [ATS14, BVG13, BvHC^{+}15, BMA^{+}19, CCE^{+}14, CRRD^{+}13, CS10b, CLN^{+}13, GKHS16, HMM^{+}15, JBT^{+}16, KKS12a, LRB^{+}12b, MEDS11, MPB15a, MCP^{+}18, MCW^{+}17, NFR^{+}14, PGR^{+}10, PNF^{+}12, PNPF14, PR11, RKC16, RBC^{+}15, RMSPA16, RLH^{+}13, SKM12, SCM13, SJS^{+}11, SDVL16, SKN^{+}17, TS10, TS11a, TFFPM18, VB14, WAMS^{+}14, ZJ17, vdWHG11]. profiles [AMM^{+}12, AAM^{+}12, BPH^{+}19, CMW^{+}15, GHNF10, MBB^{+}12b, MS14, RF14, RVF14, WNNN12, ZdVA^{+}16]. profitability [PLA^{+}12, TLB15, VF13]. Prog [ACA^{+}10b, BW16b, RUP11b]. progeny [LHU^{+}18]. Progress [Kin10a]. Progression [CCT12, ABT^{+}15, NHPW18, WAM^{+}15]. Progressive [LHS^{+}14]. projected [FA10b, XGLH14]. Projecting [LNK^{+}16]. projections [CLGCA^{+}19, KAP^{+}12, LD17b, WL19]. prokaryote [JDQS11]. prokaryotes [TDM16, UMF^{+}10]. prokaryotic [AWJ12, KdVvdP^{+}12]. prolifera [DLL^{+}18a]. proliferation [MV12]. prolonged [BLMD^{+}11, PH16]. promote [AS12, BDPD14, SH10]. promotes [CRS^{+}18, DHSL11, IHP10, NHPW18, SUPJ^{+}10]. Pronounced [APC^{+}13, ABA16]. prop [ATP14, JOFP18]. propagation [IOS^{+}12, PG15]. Propagule [CDD^{+}16, VPOA12, BMTF10, GA11, PÁIIS^{+}19, VANB15, VDV^{+}15, VK11]. propagules [LH19, PBIO^{+}18]. properties [AvDL^{+}19, BAL^{+}11, CY10, DLRF15, FNA^{+}16, FHS^{+}13, FF10, GFR^{+}18, GGG18, GMD16, LHW^{+}19, MTG^{+}10, MRSD15, SPP15, VC14, WDM^{+}19, dlSBO^{+}12]. property [TFP14]. Proportions [ACW18]. proposals [TRJ^{+}11]. proposed [NSG^{+}12]. prospects [SMMFH16]. protandric [BVQ12]. protect [FFS14, GBU15, Kat16, VPDPMRC11]. Protected [DHB18, GPA^{+}16, LBT^{+}10, LCR^{+}16, NHTS14, ABH18, AN11b, ABB^{+}14, BFD^{+}16, BAFB12, CEH18, CHB11, CBGDS12, CME^{+}12, CVM^{+}17, DMH15, DMPG11, ETB13, FTWA17, GSF18, GCRP15, HVBG^{+}19, HNMK15, JW10, KBG12, KCS^{+}17, LBS^{+}15, LFRF13, MM13a, MWB13, MUOS13, MDKM13, MS11, NMS^{+}18, OGCHS18, PSdA^{+}17, PJMP10, QHG^{+}19, RKD18, SFM^{+}19, SHB^{+}19, SDB12, SLPP^{+}14, TM10, TMH^{+}18, VGGF12, WFBM14]. protecting [BPM^{+}13]. Protection [OHB^{+}15, RP15, Ale13a, FBGT12, HVBG^{+}19, KHW^{+}19, MOP^{+}15, RLM^{+}14, VTB12, WSD^{+}10]. protects [MLCC19]. Protein [MHF^{+}14, EH13, OAGV11]. proteins [RPGP^{+}15]. Proteomic [GDSPRA13, MVL^{+}16, VTSH^{+}19]. Protist [AWJ12, HPMZ11, BRB17, HPN^{+}17, MU13]. Protista [Bol17]. Protistan [TSS11a, BTLF17, EBK^{+}12, GMB^{+}16, LMD12a, MDF10, TSS11b]. protists [BC15, DR10, KAJP^{+}15, NNJ^{+}13, OMSP10, TDM16]. Protoceratium [EOA15]. protogynous [CCM^{+}15, EW16, SKK^{+}18]. Protonibea [SBFP10]. Protozoa [SP13a]. protozoan [OAGV11]. Protozooplankton [ANJ^{+}11, AMLS15, RSK^{+}14]. protracted [Coo11]. provide [Ano17c, Ano18b, BMÁN^{+}18, CG16, DMC^{+}16, EPF^{+}17, FKT15, FHG^{+}14, GKPT16, HGG^{+}10, HS11b, WWGD13]. provided [SLRP^{+}12, SSdW^{+}12]. provides [BVC^{+}16, JBL^{+}16, JL12, KHF^{+}11, LP12, MBKM17, OHP18, PEB^{+}14, WRB^{+}15, YHTS12]. providing [Goc10, HFM12, WNP^{+}14]. province [DIG^{+}11, MJR^{+}14]. provisioning [BM13, BMPC15, BS19b, CBL^{+}10a, FSB15, LSJ17, PFKB18, RP15]. provisions [PWM^{+}19]. proxies [KSK07]. Proximate [LTP13]. Proximity [PHI^{+}12, TVH^{+}10, DBTV13, KTF10, RHMR18]. proxy [MS10, PDCH15, TPS10, WPP11]. Prymnesiales [HGL15]. Prymnesium [HGL15]. Pseudanthias [CBL^{+}10c]. Pseudo [DL13, GHT^{+}15a, HOA14, LBM^{+}18]. Pseudo-nitzschia [GHT^{+}15a, HOA14, LBM^{+}18]. Pseudo-replication [DL13]. Pseudoalteromonas [SP15]. Pseudoboletia [LHU^{+}18]. Pseudocalanus [DHH10, KDB^{+}10, PSS^{+}12]. Pseudodiaptomus [KK17]. pseudofeces [vBJV^{+}15]. pseudonana [ST14]. Pseudopleuronectes [JNYT13]. PSII [RAH^{+}10]. psittacus [GKW10]. Ptacnik [TSS11a]. Pteraeolidia [YTM^{+}15]. Pterodroma [HHS12c, LBJ^{+}16, PJP^{+}11, RCP^{+}16]. Pterois [AL12a, Alb15, ACL^{+}16, BWM^{+}15, BMA^{+}10, CBK^{+}12, CMH14, EPM^{+}16, HTSA16, KBC^{+}12, PHSBO16, SJO16, WMB^{+}14]. Pteropod [HSA^{+}15, BO15]. pteropods [BNE^{+}19, CAG12, MEDS11, SOT15]. pterotum [STYT10]. Puakō [GFC^{+}14]. puber [SHC^{+}10]. Puerto [PVZD^{+}11, ASWB10, BA13b, HTSA16]. puerulus [FRTB14]. PUFA [MDAPH11]. puffer [GKW10, LMV^{+}19]. puffin [FRMS17]. puffins [BNAN12, SEF^{+}15]. Puffinus [HMO^{+}12, NBGS^{+}12, PWA^{+}17, RHW12, RKM^{+}10, TTT^{+}15, WPF^{+}16]. Puget [GKR^{+}15, HOA14, LCJ^{+}12, MBG^{+}15, SFKQ15]. pugnax [RWR^{+}15, TB10]. pulchella [RMME10]. pull [AM19]. pulse [ENO^{+}11, TLJ^{+}11, VTZ^{+}13]. pulsed [FDS12]. pump [PC12, TPC^{+}13]. pumping [GMK^{+}19]. pup [SDR^{+}15, dDTA^{+}18]. pupping [MHSG18]. Purple [Mat10a, Ebe10, MPD11]. purpleback [YHTC19]. purposes [CBL^{+}10b]. purpurata [MPD11]. purpuratus [Ebe10, EHR12, Mat10a]. purpurea [FG10]. purse [ECC^{+}15, HIMH15, Nor13]. purse-seine [Nor13]. Pursuit [MGB^{+}19, POR^{+}15, RDHM10]. Pursuit-diving [MGB^{+}19, POR^{+}15, RDHM10]. Pusa [HLIK16]. push [AM19]. push-pull [AM19]. pusilla [DKB^{+}12]. pusillus [MMB^{+}19, MMB^{+}19]. putative [NCW^{+}19]. pycnocline [SNT^{+}18]. pygmy [Sat17, SGT12]. pygoscelid [LFN^{+}12]. Pygoscelis [CHP^{+}11, HPR^{+}12, PLNE11, PTP^{+}15]. pyrifera [FK10, Fox16, HHH11a, KMP15, MJW19, MZ10, PMS10a, SH14, YMPR13]. pyrosequencing [VBŽ^{+}17]. Pyrrhophyta [RAH^{+}10]. Pyura [MC10b].

Qeshm [SSPKF16]. QFASA [RT12]. Qingdao [DLL^{+}18a]. quadrata [POX^{+}17]. quahogs [PPML17]. qualitative [RFM^{+}15]. qualitatively [PMS15]. quality [ARB^{+}12, AFSL^{+}15, AHB^{+}14, BMF^{+}12, BS19b, CLS19, CCM^{+}15, DSM11a, DVS13, FKvdV12, HMKF11, HM17, HU13, HCF^{+}16, HCA11, JC17, KÖHSO12, KHW^{+}12, KDB^{+}10, LWS^{+}10, PW15, PDB15, PL10, PCH^{+}14, RSC14, SAWK18, SPP12, SMS15, SMK^{+}17, WKvW10, WTD^{+}19, ZMS^{+}14, ZLCW15]. Quantification [KBR^{+}11, LJK^{+}11, Ano17c, BSS18, RT12]. quantified [GBH^{+}10, MR16, PKS11a, RJ12]. quantify [BCW19, DVA^{+}18, GKRG^{+}19, JCRC10]. Quantifying [BBM13, Bru10, HSPS14, HLS15a, HG13, HPP^{+}10, HR19, LGMB^{+}16, MGM10, MMC^{+}18, RET^{+}19, SB10, SLS11, SLSR10, WLP^{+}11, BFD^{+}16, CBLM^{+}17, CCM^{+}12, KNIN17, SJ10, Vig15, LCL^{+}17]. quantile [CBYH13, NMCP17]. Quantitative [BBG^{+}16, GED^{+}14, GNJJ^{+}17, KBB14, NNJ^{+}13, PW11, SJL13]. quantitatively [PMS15]. quantity [CMC^{+}15, KÖHSO12, KDB^{+}10, SPP12]. Québec [PPD^{+}14]. queen [BBdG^{+}19, BFN^{+}19, DH18b, KCS^{+}17, SD17b, SDB12]. quenching [MAB12b, vdPLdV^{+}11]. Quevedo [BBRM14]. quickly [TS14]. quinquecirrha [BB14c]. quoianum [DdR10]. quota [MFI11, PO13]. quotas [GGLT15]. quotient [RKPB^{+}15].

R [Ano12c]. R. [FCV^{+}16]. racemosa [GP11a, BAC^{+}11]. racemosum [RBBG12]. radiata [BVG13, MGFW13, RWR^{+}19, TW16]. radiation [CFA16, DGQ10, EBOM11, GHHH12, GJ10, HPL^{+}16, JDX^{+}17, LAALH10, MWC^{+}12, OCMT13, ROLA15]. radiative [LHW^{+}19]. radiocarbon [NUO^{+}19]. Radiocesium [NKM^{+}15]. Radiolaria [Bol17]. radula [CHDH^{+}19]. Rafting [BAL^{+}11, GCFW14, CMAJ12, HVT12, HRC12, LLS^{+}18, NFSW10, WC18]. rafts [CMAJ12, MBB11, WC18]. Rain [PvdMS^{+}15, Fie14]. Rain-driven [PvdMS^{+}15]. rainbow [BDT^{+}11, MDAPH11]. Rainfall [PCA^{+}15, CGK13b, VGBJ^{+}13]. raised [ADL^{+}18]. Raja [FCV^{+}16]. Rajidae [SST19]. raleighana [DGFH10]. ramifications [GFKC13]. Ramsey [MGM^{+}18]. random [AAGGG19, MHNL14]. Range [GCR16, AAE14, APBC^{+}19, ABW14, Ano17c, BCOMKJ18, BPF15, BFN^{+}19, CLM^{+}13, CTZ^{+}16, CHB^{+}13, CPW15, CWAS17, DKV^{+}16, DCT^{+}10, DRDV15, FHJ15, GFKC13, HSF^{+}12, HRAM13, JPF12, KHG^{+}19, LAM10, MPA^{+}10, MH19, MOMF14, NMS^{+}18, NM18, NdBKR13, NvOH13, PBM^{+}11, PBH^{+}17, RTMJ11, RRRF^{+}17, SGM^{+}14, SJLH12, SKW^{+}14, SHF^{+}12b, TJS17, TPM19, TCWC13, VMC^{+}19, WL19, YSH^{+}16]. range-expanding [APBC^{+}19, MH19]. range-wide [SKW^{+}14]. ranges [AHA^{+}15, BBKP10, CKC^{+}10, GSL17, KBC^{+}12, MLM^{+}12, MVP^{+}16, NWH^{+}12, SJL^{+}16, SRM^{+}15, SJB^{+}15, SO15]. ranging [ASAG17, AGNH^{+}19, BCV^{+}18, CWB^{+}15, DHS^{+}19, GGD^{+}11, LSH19a, MCPC19, MRLD19, MPG^{+}10, MHR^{+}19b, PRP^{+}18, SBG^{+}16, TWL15, WEAG^{+}15]. raphid [GFR^{+}18]. raphidophyte [QYS^{+}12]. Rapid [KF17, LRK^{+}15, SAH11, HHG10, WC18, WEA^{+}17, YChH11]. Rapidly [FF16, RMB^{+}14]. Rappahannock [SSD13]. rare [CKR^{+}11, MLASTS13, MCUO12]. rarely [NBM12]. rarity [BBWG12, WKRW^{+}12]. rasps [JBCGC19]. rate [BGP^{+}16, BPF^{+}18, CC19, CA14, CSC13a, CLHW14, CTB13, HSMB17, JMB10, KJW^{+}15, LLv13, LS19, MS18b, PRBR11, RWR^{+}19, SPWK13, SGT^{+}14, SH12, SWBMJ17, TNM^{+}11a, VGR^{+}16, WP10, dDTA^{+}18]. Rates [VNN10, ABT^{+}15, AHØ^{+}15, ACA^{+}10b, ACA^{+}10a, ACG12, AM19, AGTS^{+}12, BST19b, BP12a, BDAH10, BS12, BRB11a, BC15, BRB11b, CTRBLC^{+}13, CSR^{+}14, CRGT19, CPEE18, CSABH^{+}13, CLJF12, CLC12a, CL15b, CVC^{+}18, CAGS17, CM10, DTVH10, DPH14, DHH10, GUC^{+}15, GGMGR12, GS16, GMK^{+}19, HPH^{+}18, HB15b, Ing16, KW17, KBA13, KHH18, KHM^{+}12, KBPD10, KCL^{+}11, LTN19, LSJ17, LA12b, Mat10a, MH11a, MA14, MAK^{+}15, NHPW18, OFPR^{+}11, PSC^{+}13a, RS13, RFRE19, RRS15, RGB^{+}11, RBD^{+}14, SKFH18, SM15, SDT15, SLS16, SAD^{+}16, SF14, SGD^{+}12, TFP14, TLFS14, TPA^{+}13, TOR^{+}10, VGGD^{+}15, VHBG15, VBJR^{+}13, WS17, Wir14, WPMP14, YPWP18, YKG^{+}15, ZAC11, dQSC13]. rather [CTvdM^{+}08, HMJ^{+}13, McD13, MMRDS18]. ratio [BCSR10, CHLR16, DHH10, Fie14, HFH10, LMH12, MA13a, PFG10, TMM^{+}19]. ration [MDM18, MBV^{+}11, NPMS11]. ratios [BAH^{+}15, CS13, FGH^{+}13, GRG^{+}12, GMR13a, GMR13b, HHS^{+}12a, HRT^{+}14, HBCK13, HTR^{+}19b, LN11, MC10a, MHS14, PMB^{+}17, RH12]. ray [BMPC15, CNW^{+}12, DBB11, GRN^{+}16, KMGH16, LPB^{+}16, SSM^{+}19]. rays [JRW^{+}14, OBA^{+}18, PDM12, PSD^{+}19, RPM^{+}13b, SRA^{+}17]. razor [GSO^{+}18]. razorbills [SEF^{+}15]. razorfish [ACRLB12]. Re [HKB^{+}19b, MMD^{+}13, SBS10, YSC^{+}15, SBCA15, TPS^{+}13, MM11]. Re-analysis [SBS10]. Re-established [MMD^{+}13]. Re-evaluating [MM11]. Re-evaluation [YSC^{+}15]. Re-examining [HKB^{+}19b]. re-lay [SBCA15]. re-oligotrophication [TPS^{+}13]. reach [CPF^{+}16]. Reaction [NBM^{+}14, SAM18]. reactive [MLP^{+}14, WMS^{+}11]. Real [SHT^{+}16, BZC^{+}12]. Real-time [SHT^{+}16, BZC^{+}12]. Realignment [PvdMLC17]. realistic [RZS12]. Realized [WMS^{+}10, HMF^{+}18]. realizes [CBT13]. realm [WLP^{+}11]. reanalysis [ORW15]. reappraised [NFC^{+}11]. reared [ABK^{+}14]. rearing [BRT^{+}17, FTB^{+}17, SQS^{+}13, UTS^{+}18]. Reassessing [MIIB16]. Reassessment [Dao13, MZC^{+}10]. rebuilding [FQ11, TPC^{+}15]. recapture [ARvdW^{+}16, BRB11b, CWR^{+}18, PBM^{+}11]. receiver [MOMF14, RBD^{+}14]. recently [ATT^{+}16, PPTB12, SZ13, vdHTT^{+}15]. recipes [AS12]. reciprocal [BSCH14]. recognition [EN11]. Recolonisation [LPC11b, CBC^{+}16, HBK^{+}17, LB15b]. recolonising [ACMD11]. Reconciling [CPMR11]. reconstruct [DNO17, WRFM16]. Reconstructing [CLH^{+}15, HWET13, MDR16, PFdP^{+}12, RMP^{+}19, VWEP18, VLM14]. Reconstruction [SKMS^{+}18, OOK11]. reconstructions [PB19]. recorded [EKOÅ13, MRLD19, RCVG10, WPS^{+}19]. Recorder [MJT^{+}12, PLJ^{+}12, EKOÅ13, LAH^{+}15]. recording [ZVR18]. recover [FFS^{+}13]. recoveries [SCAM^{+}14, SMR^{+}11b]. recovering [CCM^{+}12]. Recovery [ABB^{+}14, FCR^{+}15, FRJ10, IHP10, JGL12, LS12, McC14, MRC^{+}12, SYM^{+}16, VQKF14, AB17, BFKR17, BRB11b, CHE^{+}17, CHB^{+}13, CMGS19, CRWG16, DTVH10, DCC^{+}10, FNC13, FTM^{+}13, FB14, GEA^{+}14, JMK^{+}18, KOPW15, LCJ^{+}12, LPKT^{+}19, LDM10, LE13, MJR^{+}14, NTB^{+}17, OSPWB15, OAD^{+}14, OMH12, OMM^{+}12, SAM14, SB15b, TBMG17, VHG^{+}18, dJTH^{+}14]. recreational [AMR^{+}10, APC^{+}14b, BBC14, DP16, MMC^{+}18, MMCL19, PBM^{+}11]. Recruit [KW10]. Recruit-recruit [KW10]. recruited [WSL^{+}14]. recruiting [JC17]. Recruitment [CMF^{+}19, LdU^{+}17, MH10, PSC^{+}13a, ZBC10, AOPMOS10, AY10, ASM10, ASF13, BZC^{+}12, BG19, BVL10, BFNB16, BAS16, BS19a, BDP10, BD14, BBC^{+}17, BTM^{+}18, BOS18, CHP^{+}11, DGDJ14, DWK13, Det10, DA10, DHAV13, DvBH12, ELA10, FHJ15, FPM14, FHB^{+}12, GTP^{+}17, GS18, GSO^{+}18, HLS17, HCJ^{+}18, HS16, HRAM13, HFR^{+}13, HPH^{+}18, ICUM10, JCM10, JSCM15, JSOK18, JCM13, KOLDP17, KF13b, KS13, LRB^{+}18, LFU^{+}19, LM12, LMGJ13, LB15b, LCC12, LE13, LH19, LANE18, MDLM17, MORM15, MMDM12, MLR17, Met11, MS18a, MP12, MTK^{+}18, MN10, NKS^{+}15, OHB^{+}15, OSE13, OSD^{+}13, PA13, PFF^{+}17, PB18, PBW^{+}11, PML^{+}19, PRR^{+}16, RFBH^{+}11, RWO13, RZH16a, RFMT12, SK13, SWMK15, SHB^{+}13b, SS10c, TW16, TPC^{+}13, TTM19, TSTP13a, TDPS13, VDS11, VB19b, WSS^{+}16, WR12, WS12, WTGG10, vdSRT^{+}18]. recruitment-linked [TSTP13a]. recruits [FWB^{+}15, MGO^{+}12, MS18a, OPSM13, RSP17]. Recurrence [HFMD16]. recurrent [OGLH17, POS14]. recurring [RUS^{+}13]. recycling [FLE^{+}11, MHF^{+}14, VFMdW12]. Red [BvHC^{+}15, BAE^{+}10, GKBEF18, ITT14, WKK18, ZAZS11, AL12a, ATP14, ARA^{+}10, BRF11, CBK^{+}12, CMH14, DR16, DDP^{+}13a, DHJ10, FHM^{+}11, FG16a, FPN17, JW10, JWHR11, LBT^{+}10, LBCS^{+}15, MH19, MBC^{+}17, MP12, MP17, NRF11, OF10, PES12, PSPK10, RLG14, RSHM10, RSA^{+}12, SBS^{+}13, TS11b, TPVCV19, WDNM11, WLBB^{+}13, WWKS15, WGS16, BCvH^{+}15, BDGB19, HDT^{+}12, JWR^{+}10, MG14, RBC^{+}15, TBC^{+}17, WNNH10, ZRBV15]. red-footed [MBC^{+}17]. Red-legged [WKK18]. red-spotted [TPVCV19]. redirect [CPW15]. redistribute [WJS^{+}14]. redox [CGR^{+}15]. reduce [AED^{+}10, Alb15, BT14, BTM^{+}18, CASG^{+}19, DTC^{+}17, HWTH12, LGR^{+}10, LTCH10, LRS17, MV12, Mat10a, MPB^{+}14b, OPSM13, PHSBO16, SDT15, TRR^{+}17, TW16, TVA^{+}19, VPDPMRC11, WFS10]. Reduced [GB10b, ASF13, BS19a, BCR15, EBH16, HBHS^{+}18, HM17, LBJK14, PSC^{+}13a, RMV13]. reduces [BCGG19, DDP13b, Fox16, LWS^{+}10, NHPW18, RMB^{+}14, SCTA10, WMMC14]. Reducing [Joh10, KCCG12, OMW^{+}16]. Reduction [DCC^{+}10, KAP^{+}14, ABW14, BST^{+}10, DHSL11, ETAE13, FSM^{+}14, GAF^{+}16, MHC11, RLBC10, ST14, SST^{+}18, vdVKA^{+}11, vBAP^{+}19]. Reductions [SCM13, CLMB13a, GLM^{+}10, LB15a, NDAP19]. Reef [CTSH12, CLL17, DJL^{+}12, ETB13, FTWA17, GGCB16, HSMB17, KBK^{+}15, MVP^{+}16, RUP11b, RUP11a, RdGv^{+}18, TKS^{+}14, TPA^{+}13, AOPMOS10, APC^{+}13, Alb15, AP14, AAJJ^{+}19, AHR^{+}18, BBM^{+}14, BGB^{+}17, BÁNC^{+}18, BHA^{+}11, BCSH18, BvHC^{+}15, BWS11, BGSM13, BJMS12, BJH13, BG12, BPH^{+}19, BDB^{+}13, BS14b, BAFB12, BVAB^{+}17, BLPP16, BPC17, BDF^{+}11, BMVAS19, BKTD10, BSD17, BJL^{+}17, BAvW12, CLB^{+}14, CCM^{+}15, CSS15, CHST16, CLMB13a, CAB^{+}14, CCT12, CCDP15, CBGDS12, CP11, CV11, CSRS15, CVM^{+}17, CBJB19, CBL^{+}10c, DXF^{+}18, DPS16, DPO^{+}17, DWK13, DPM15, DMM^{+}10, DHAV13, DWV^{+}12, DHJ10, ETS^{+}11, Edm10, Edm14, EAB^{+}19, EGH^{+}15, EHG16, ETB^{+}11, EMH^{+}19, EMJ18, ELM^{+}19, FMP13, FHL^{+}17, FFLS16, FA10a, FHB^{+}12, FM14, FF16, FSDS10, GL16, GJ16, Gea10, GSS13, GMHS19, GSF^{+}13, GEA^{+}14, GB10a, GSF18, GSM^{+}11b, GGvW^{+}15, GWM^{+}15]. reef [GNBM11, GRN14, GHNF10, HJN^{+}10, HCL^{+}11a, HSPS14, HPA17a, HMKF11, HU13, HS14b, HLS18, HCL10, HVBG^{+}19, HPP^{+}10, HAF^{+}17, HTR^{+}11, HPN11, HRC12, Hoe10, HSR^{+}13, HM10, HPM14, HDP^{+}17, Hug11, HNG^{+}11, HB15b, HMM14, HLD12, INvdB^{+}13, JRW^{+}14, JMB^{+}17, JGJ15, JWR^{+}10, JDB^{+}10, JBD10, Joh10, JD18, JRK13, JWBS10, JMK^{+}18, KMPS15, KB12, KCOP13, KMP11, Kin16, KMO19, KRC^{+}11, KCL^{+}11, LJ10, LWP^{+}13, LHDH17, LB11, LB15b, LP10, LKDP17, LS19, LRH14, LL18, ŁCM^{+}18, LEB^{+}19, MBJ16, MGP11, MGP12, MLPC18, MWP10, MOP^{+}15, MM12, MHGW13, MH12b, MJ17, McC19, McC12, MMM^{+}15, MDM18, MSE^{+}18, MHVV13, MMD^{+}13, MZF^{+}19, MS18b, MTG^{+}11, NHS^{+}17, NTN^{+}13, NAY^{+}11, NMS^{+}18, NJP12, OCPM12, OLF11, OPW^{+}11, PGG12, PHSBO16, PMEZ15a, PS15a, PSD^{+}19, PA13, PMSGBL16, PMOAC^{+}17, PPC^{+}19]. reef [PBM^{+}18, PPDW16, QHG^{+}19, QBW18, RTMJ11, RSJ14, RBC^{+}15, RSV10, RPL^{+}18, RPGP^{+}15, RDF^{+}19, RMRA15, SAWK18, SJB16a, SAZHAAHS14, SHK14, SLS18, SHUS17, SPJ^{+}12, SATS13, SL11a, SRR15, SSB10, SSB^{+}18, SLF^{+}19, SCLB11, SGEO13, SWBMJ17, SESS^{+}18, SMF^{+}11, SMF^{+}12, SPW^{+}12, SAM14, Sti13b, SMR10, SKS18, SS10c, TPB^{+}18, TM10, TBC^{+}17, TDSS15, TB14b, TYVR^{+}15, Tut17, VBR^{+}14, VHG^{+}18, VMWU15, WBB^{+}10, WR12, WNNH10, WLP^{+}10, WLHW10, YAL12, BP12a, BB11, BKB11, BSB^{+}11b, BK14, CED^{+}19, CBJB19, ETB13, ETAE13, HPLS13, KHH18, LKDP17, LPC11b, SGG^{+}12, SWB^{+}10, Sti13a, TPB^{+}16, TPB^{+}18, WSC13]. reef-associated [CSRS15, EMH^{+}19, HS14b]. reef-building [ELM^{+}19, LKDP17, PGG12, RPGP^{+}15]. reef-derived [HLD12]. Reef-scale [GGCB16]. Reef-wide [FTWA17]. Reefs [SGTD14, ALV12, ABRP15, ALK17, Ale13a, ACL^{+}16, BMM17, BFNB16, BA13b, BJ14, BAE^{+}10, BFL^{+}19, BBC14, CLL15, CSB14, CMC10b, CL10b, CMYL11, CPR^{+}14, CLL17, CM10, CRHV14, CBK^{+}12, DPS16, Dav19, DQM^{+}19, DBTV13, DABP10, DJL^{+}12, Edm13, EL16, EL19, EBdlTC12, FTWA17, FG16a, GEAOH17, GBE^{+}17, GB10a, GAC11, HCL^{+}11a, HBI^{+}18, HLS15a, HW13, HR19, HM13, IHP10, JHNL12, KSF18, KBB14, KBB16, LTN19, LBH18, LMB^{+}19, LEB14a, LL18, LFRF13, LRS17, LEB^{+}19, MPH^{+}11, MBJ16, MOP^{+}15, MHC11, MM16, MSF^{+}19, MHQ17, MCW11, NAG^{+}16, NSH^{+}15, OS18, PMK^{+}15, PSCS16, PMEZ15b, PS19, PCH^{+}14, RMME10, RPL^{+}18, RCP^{+}13, SGTP16, SBW16, TG12b, TVH^{+}10, VRMKH^{+}15, VDS11, VHKT15, VBR18, WSK^{+}16, WGS16, WMCC15, YPWP18]. reefscape [FHL^{+}15]. reference [CSPA10, MBB^{+}12b]. refines [FKP^{+}15]. reflect [ABW14, BCC^{+}11, BKM^{+}18, DAM^{+}18, GE14, GNHLN11, MDE13, OPN10, VHV13, WRW^{+}16a]. reflected [RP12]. reflective [SMM^{+}15]. reflects [DGP^{+}19, MWWS17]. reflex [GSCH14, SCSC16]. refuge [BMÁN^{+}18, BPW^{+}18, CMC10b, CG16, DHG11, FA10a, Gea10, HED^{+}13, HNG^{+}11, JGJ15, MBJ16, MWGH15, RLS10, WDM^{+}19, WWGD13, WL11]. refuge-seeking [HNG^{+}11]. refuges [LH19, Sti13b]. refugia [HLIK16, MKK^{+}11, SSdW^{+}12]. Refugio [LBM^{+}18]. regalis [BT13, LTG17, LWG14]. regardless [CGH^{+}16]. regenerated [STBP13]. regenerating [BSY13]. regeneration [DDD^{+}11, FF12, JSS^{+}11, MAL15, vBJV^{+}15]. Regenerative [GME^{+}19]. Regime [LDM10, PFF^{+}17, AF14, BMSB13, EL16, Hat13, HCJ^{+}18, KLD18, LLMA^{+}10, LSD^{+}13, MGB^{+}19, OS18, PBC^{+}18, PTS^{+}18, SDKS19, VM14]. regime-dependent [KLD18]. regimes [FPM14, GWM^{+}15, GHT^{+}15a, HRS^{+}17, KKPL14, LBP^{+}19, MBBBC12, MWM^{+}11, NMC^{+}16, OHR^{+}13, TFP14]. Region [KMP14, PMIG19, Ano18a, BBR^{+}14, CBKM16, DBE^{+}12, DKV^{+}16, GDE17, HEMJ18, HPR^{+}12, HPL12, KN12, KHS^{+}11, MSH^{+}14, OTK^{+}19, OMT^{+}17, PBW^{+}11, RBC^{+}14, STYT10, SZ13, SBB^{+}12a, TRMC11, TGBM^{+}16, TFH^{+}19, WSF^{+}12, WRFM16].

Region-based [KMP14]. Regional [DIB15, FTB14, GB16, HLPS^{+}15, MLPC18, RSC14, SMBH15, SSB10, SS10c, VQSH14, WS13, tHHR10, ALK17, BHN^{+}14, CSKR10, DFDC18, GMP^{+}11, HLIK16, HM09, JD18, JMK^{+}18, KO12, MEDS11, MdSW^{+}16, MPKO11, MTMM^{+}15, OGCHS18, PLNE11, PRP^{+}18, SUB^{+}13, SBE^{+}16, TMSG19, TMBP13, VJG^{+}13, WGD10, WDM^{+}18, YPWP18, dJH11]. regional-scale [SBE^{+}16, YPWP18]. regions [ADvdM13, AGM^{+}11, AWH19, BFH13b, CEH18, CGK13a, CBF^{+}13, CLN^{+}13, GMT^{+}15, IVD^{+}10, LSH^{+}15, MJR^{+}14, MIIB16, OVF11, TLFS14]. regression [BBC19, CMLC12, CBYH13, RFRD10]. regrowths [LRO^{+}17]. regular [MPM^{+}14]. regularis [BGBD^{+}13]. regulate [DBS14, GG12, YKG^{+}15]. regulated [CLB19, KCF^{+}14]. regulates [DMS14, MTM^{+}11, MF10]. regulating [APC14a, GLY^{+}14]. Regulation [EL19, BFRM13, BRH13, CMP^{+}13, KML13, MTV^{+}11, MM13c, Mou17, QBW18, Rei12, SNK^{+}10, WBLRN18, WP10, ZGK^{+}19, ZJ17]. regulatory [HP13, MPD11]. rehabilitation [EGH^{+}15]. Reinhardtius [BNS^{+}14, PHH^{+}12]. related [AGIW16, ADRA19, BSG^{+}10, BNAN12, BJH13, BJS^{+}11, BBC^{+}10, CXdG^{+}17, CAB^{+}16, DEWK15, GMG^{+}16, GFR^{+}18, GTL^{+}16, GNBM11, LMR^{+}10, MVL^{+}16, MFDW17, MRB^{+}11, MRL^{+}10, MJHS11, MJHS13, OSPWB14, POPH12, Pör12, RAKB^{+}12, RBA^{+}18, SLNG18, TMP^{+}11, TGSM12, YTO^{+}11, YSS^{+}18]. relatedness [HBP^{+}16, KHGS12]. relates [TP19]. Relating [OLF11, PFG10]. relation [ACF^{+}10, BMM17, BCSR10, BDR19, BT13, CSF13, CEK^{+}16, DLRF15, DMT^{+}14, GHCC15, GCRP15, GGB^{+}18, GVGP^{+}14, HTHM18, HP11a, HFM12, HHS^{+}12b, IMS^{+}11, JCZ^{+}12, JJNP12, KWW^{+}14, LCB10, MCS^{+}15, MG12, MGB^{+}14, MKK^{+}11, OHA10, PMML12, PHD^{+}14, RFH16, RPP^{+}13, RB19, RBGT10, RCBE15, SGC^{+}13, SSHB16, SPSP10, TKA10, TYG12, TKIO10, TUK13, VHP^{+}11, WDM^{+}18, Wir12, WA19, YTO^{+}11, YFC^{+}13]. Relationship [DDP15, DEM15, Goy14, HTT^{+}12, HA14, LLRC^{+}12, YOWK17, ZIT^{+}17, BS14a, BYAA13, Bea11b, HW11a, HMM^{+}15, KPH^{+}15, LRB^{+}12b, MM11, MS18a, ØLH^{+}13, PDGR^{+}13, PW15, PH10, RGB^{+}10, SCAM^{+}14, SS12a, SDVL16, VW10, vdMPM12]. Relationships [AMM^{+}12, FCR12, HTR^{+}19b, LN11, TPCN12, TSF^{+}10, BAS^{+}10, BCN15, BB19, EMH^{+}19, FC13, FSS^{+}18, HYL^{+}13, JSMO14, LOBW^{+}11, LHB19, MUHC11, McC14, MHS14, NOB^{+}19, PLG^{+}19, PHM11, RBHR19, RLNLAF17, SMBH15, SLO19, TB12, TMBP13, TCA^{+}11, TUK13, TPVCV19, WRDF12, ZHTA13]. Relative [BHP12, BKB11, BPB11, DAWS19, FGL^{+}12, HMS10, PA13, PCNB16, RMC^{+}14, RMBM14, SMR10, dJH11, AGK13, AMSdL^{+}18, ARG^{+}10, AJ16, BGB^{+}17, BJM19, BRG^{+}14, EFOM18, GMB^{+}16, HWR^{+}14, JL17, KWH^{+}18, KTF10, LRHP18, LS16, LKN^{+}14, MN10, MS10, OFT^{+}11, PDBB14, RST15, RLO18, SCAM^{+}14, SCW^{+}15b, TSS^{+}12, WHH^{+}13, WSF^{+}12, WHB^{+}19, WFH^{+}18]. relatively [BHTF11, HP13, HHH^{+}11b, HVK^{+}13, NMPM15]. relatives [CRGM18]. relaxation [SMM^{+}14]. relaxes [MSE^{+}12]. Release [LMW^{+}13, BMTF10, CLP18, DHSL11, FHL^{+}17, GSCH14, GW15, GPS13, HJN^{+}10, HIMH15, KF13a, MWM^{+}11, PBM^{+}11, RM13, TBMG17, VBT^{+}17, WYS^{+}10, WNNH10]. relevance [CC19, GGK^{+}16, GCCCM16, PMB^{+}15, TFS14, WTGG10]. relevant [WLHW10]. Reliability [LAH^{+}15]. reliable [INvdB^{+}13]. reliably [COK10]. relocation [NLV^{+}15]. Reloncaví [GCC^{+}10]. rely [KMP15]. remain [CPF^{+}16, FT10]. remained [BVAB^{+}17]. remains [ATF12, IW15]. Remane [TSS11b]. Remarkable [SMK^{+}18]. remnant [ACMD11]. Remote [KWC^{+}17, BGP^{+}16, BG12, BSC^{+}16, CCS^{+}18, CS10a, Dav19, FBGT12, GSF18, HSH12, HR19, KLBT^{+}16, MSF^{+}19, MZF^{+}19, NAY^{+}11, PHD^{+}13, RIM^{+}13, RKFD15, RSTM14, SKD^{+}11a, WLM^{+}13]. remote-sensing [BSC^{+}16]. remotely [BFF^{+}10, KBG12, SSS12]. removal [DC15, GFC^{+}14, GB10a, LBCL10, SBCA15, ZZG^{+}19]. removals [DPS16]. remove [NFN^{+}15]. reoxidation [VCMS^{+}18]. reoxygenation [BBG^{+}13]. rep [EH13]. repair [CBD14, FSFS18]. repatriation [RWH14]. repeat [CML^{+}12, HFH10, SDB12]. Repeatability [Col10b, TPT^{+}12, TP19]. Repeated [AZSKWL11, GD15, SMOH19, VKH10]. replacement [Cha13]. replenishment [ATS14, NHS^{+}17, TBG^{+}17]. replicate [JWHR11]. Replicated [TMH^{+}18]. replication [DL13, EH13]. Reply [AvdHH^{+}17, BBM13, Bet14, Bru10, CBL^{+}10a, Fry13b, FS18, HGS15, MAD18, MAB^{+}12a, MBS^{+}14, MHO^{+}11, NWD13, PMEZ15a, SHM^{+}12, She13, SCHO11a, TSS11a, TSTP13b, ZAZS12, dQSC14]. report [AFC^{+}16]. represent [SHM^{+}10a]. representative [TW10]. represents [Ano18b, UHG12, WEAG^{+}15]. Reproducibility [LGL11]. reproducing [BEFC15]. Reproduction [SIHH10, VHV13, BM13, CLC12a, CO16, CFMN16, FLS^{+}18, GK10, KBL11, KNY^{+}15, KCLLL13, LGB^{+}13, MGFW13, MRB^{+}11, MRC10b, MCRB11, PDC^{+}17, PRFA11, Rai12, SES12, SY10, SMJ^{+}14, SBB18, SB14, VCDP11, ZH13]. Reproductive [ABW14, BJS^{+}11, DT15, EPS11, LMCD19, MN18, PFPP18, RM19, TKA10, WGGM11, ACDJ11, BFS10, BSH^{+}18, BBW^{+}19, BK14, CPSC^{+}15, CLB19, CRR^{+}19, CGC15, CLC^{+}12b, CCM^{+}12, DMPG11, DMM^{+}10, DE17, DMHS^{+}14, FHG^{+}14, FTF15, Hal15, HH15, HWVM19, HPR^{+}12, Hos14, HAHM16, JJF14, KCCG12, KCR^{+}10, KHW^{+}12, KNB^{+}13, LSJ17, LBWB^{+}13, LBCS^{+}15, MGGSP15, MBP^{+}12, MS18b, MBKM17, OBM15, PPM^{+}15, PMGQM17, RJTS12, RT17, Sat17, SKK^{+}18, SE16, TPBM13, TPKC18, TMB^{+}12, VBD^{+}11, WHPT12, WA19, YKT^{+}14, ZdVA^{+}16, vdGSE^{+}14]. repulsion [BKTD10]. require [FFS^{+}13, ST10b, TW10]. required [EAB^{+}19, JP15]. requirement [MG12, MGB^{+}14, ZIT^{+}17]. requirements [CPMR11, FTM^{+}13, GLM^{+}19, JDG^{+}17, LCF^{+}14, MMSC17, NUJ12, RTÁB^{+}14, SSS^{+}10]. requires [MBC^{+}11]. RES [WHH^{+}13]. research [ARW12, Ano17c, BW10, BBBHC17, CCGP14, FABH^{+}16, RSV10, STM^{+}14, WNP^{+}14, WV12]. resembles [JCA13]. Reserve [AMSdL^{+}18, TEAO^{+}14, WDNM11, WGD10, AHA^{+}15, BBBW17, BVAB^{+}17, FBGT12, GRL^{+}14, GHD^{+}10, HRT12, KHBL11, KNY^{+}15, KLKM12, KHW^{+}19, LJ10, LCG^{+}17, MPA^{+}10, MOK^{+}11, OE11, PBG^{+}10, REBL16, RMRA15, WBHL10]. reserves [BKR^{+}15, GTP^{+}17, GHL11, HOL^{+}17, KLKM12, MOK^{+}11, NBM12, PBM^{+}10, RUD^{+}16, SAMW14, TMB^{+}12, WAW^{+}12]. residence [AGRPSVL13, MPA^{+}10, TTPM12, TS11b, WWW^{+}13]. Residency [LL18, SHB^{+}19, TEAO^{+}14, ACP^{+}16, BDB^{+}13, KCF^{+}14, KCS12, LSH^{+}19b, LMLB12, LWG14, PSD^{+}19, TNB^{+}15]. resident [APBC^{+}19, CSM^{+}19, FBPCC^{+}19, FM14, MCPC19, MHN^{+}12, RMBM14, RSSvV19, SGM^{+}14]. residents [MGL^{+}16]. residual [MBM^{+}14]. Resilience [FNF^{+}18, HBB^{+}12, VJM^{+}16, ABRP15, BRK^{+}12, CDMW12, CWMP13, CC12, GKPT16, LDM10, LBCS^{+}15, MJTC^{+}12, MW16b, OSK15, PGR^{+}13, PHI^{+}12, PMB^{+}17, PRP^{+}16, SBCA15, SCL^{+}16, WBB^{+}10, WK16]. Resistance [BMVBC10, GELJ19, JC15, TS12, CMGS19, GAO14, MLB19, MPF^{+}15, RPHB15, SMK^{+}18, SBO^{+}18, WSD^{+}10]. Resolute [MGG^{+}14]. resolution [CLB^{+}14, CL10b, GSC14, JDB^{+}10, KMP11, LBN11, LPC19, LMTD16, LD17b, SBB^{+}12a, SJL13]. resolved [BSST13]. Resolving [BDT^{+}11, CKC^{+}10, ST08, KMP14]. Resource [BJM19, Bre14, LMV^{+}19, TB12, WBJ12, WJ12, YSM^{+}10, Ano18b, BMJH12, CDWC17, CHE^{+}17, ES12b, EMH^{+}19, Fox16, GED^{+}14, GHGG11, HVK^{+}13, HMR^{+}16, HLH18, JS12a, KWPL11, KHF^{+}11, LCGS^{+}19, LP10, MWA17, NW15, PBTB15, PDC^{+}17, PBC11, PRM17, RHVRC11, RMT^{+}13, SGS^{+}19, SST19, SQRDLC18, TKEV^{+}19, TLJ^{+}11, VARO12, WPNK12, WF18, WW15, YKT^{+}14]. resource-use [SGS^{+}19]. resources [BHT10, BSY13, FOL12, GMT^{+}17, KKMA^{+}10, OMN17, PDL^{+}10, RLN17, SDG^{+}12, SHT^{+}16, SRK13, YTM^{+}12]. Respiration [BST19b, RKPB^{+}15, AGTS^{+}12, CPSC^{+}15, CE11, CA14, HSS^{+}10, LLv13, LHL11, MAK^{+}15, RRS15, SG18b]. respiratory [NWB^{+}19]. respond [Ano18a, LMR^{+}12]. responds [OSL18]. Response [BCOMKJ18, CBJB19, DTVH10, DBCB11, ENO^{+}11, GMR13a, KdWV^{+}18, KPB16, KOPW15, LE10, LNIH16, MATR11, MAT^{+}11, PFB^{+}13, PPCH17, SSAB^{+}15, TFÁS^{+}12, ALWD15, AB13, Ano17c, BVO16, BMC^{+}15, BL14, BLHRP^{+}19, BLER15, BSA^{+}16, BJL^{+}17, BRF11, BHN^{+}12, CLS19, CBL^{+}10b, CESC14, CLS^{+}16, DBCV^{+}12, DS16, DT11, DBDK14, FMG^{+}18, GSCH14, GBTS11, GD15, GM10, GLY^{+}14, HP13, HM17, HSZR10, HBC15, HH15, HCL^{+}11b, Hoe10, HLL^{+}17, KCMH^{+}17, KHD19, LB15a, LSD^{+}13, LERS13, LMH12, LRQ^{+}12, LFN^{+}12, MEDS11, MJW19, MVCD14, Mat10a, MCBR18, MKS18, MHHF13, MFDW17, MW16a, MPG^{+}12, MHK11, OSC19, OFT^{+}11, OBÁSV18, PDM12, PL15, PLS^{+}10, PJDC^{+}13, PHHL19, PPEH^{+}14, PPM11, RLN17, Ran18, RDHM10, RFJ^{+}10, SOB16, SGS^{+}17, ST10a, SPJ^{+}12, SA14, SHB^{+}13a, SLIGG18, SMR11a, SDAM11, SSB^{+}16, SHDR13, SKM^{+}15, TMGCM11]. response [TNIH12, TLJ^{+}11, TCKY19, TBL16, TH11, TBH^{+}17, TWW^{+}15, TPF^{+}13, TPS^{+}13, UCK^{+}13, VHBG15, VTSH^{+}19, VCL^{+}15, VBGO^{+}16, WNG^{+}19, WK18, WD14, ZBK^{+}14]. response-related [MFDW17]. Responses [BDBN11, CCM^{+}14, CL15b, GSA^{+}18, GHHH12, GMM10, WRGP12, AS15, ADO^{+}12, BH16a, BDPL19, BH12b, BMPC15, BFL^{+}11, CPSC^{+}18, CY13, CSV^{+}15, DSVM10, DHGW15, DAM^{+}18, EN11, Fie14, FDP^{+}16, FS18, GEA^{+}14, GWM^{+}15, GHGG11, GTL^{+}16, HER^{+}19, HYC^{+}14, JBT^{+}16, KRA17, KHBL11, KVY^{+}13, KCW^{+}17, KRW18, KOLDP17, KR16, KH18, KH16, LST11, LFFL15, LBP^{+}19, LJM^{+}10, LSQ^{+}13, LEKT12, LCL^{+}11, LFD13, LH10, LF14, MGFÁS^{+}10, MHC11, MH12b, MRL^{+}17, MAB12b, MRB^{+}15, MLR12, MTH15, MWMD17, NNT^{+}10, RCE^{+}13, RKT16, SDG^{+}12, SRB15, SJB16a, SSL^{+}14, SBE^{+}14, SL11a, SDKS19, SCB^{+}16, SHMB15, SV12, TB14a, THRBL^{+}16, TVM16, TVRG^{+}13, Whi11, XGLH14, YTM^{+}12, YKBB12, ZGH^{+}17, dlSGS17, vdPLdV^{+}11, vdSRT^{+}18, vdMPM^{+}10]. responsible [SCHO11b]. responsiveness [WTD^{+}19]. resting [FCS15, MDS^{+}13, SM10b, SBKA14]. Restoration [AM18, BMMS10, CLL15, CHB^{+}13, HC13, IEM16, JSBP15, LC18, QBW18, RWM^{+}12, SFW^{+}16, TPC^{+}13, TPC^{+}15]. restorative [BWH^{+}17]. restore [CVM^{+}17]. restored [CLL17, CM12, FWL15, GWME16, KSF18, KCOP13, LSBB17, PS19, ZZG^{+}19]. Restoring [PSP^{+}19]. restrict [HGW^{+}11, RPHB15]. Restricted [MTG^{+}11, BH10, FCRD18, PDM12, SKD17a]. restriction [MVOPB^{+}16]. restrictions [MA14]. restructuring [LJ10]. result [AS13, AZSKWL11, BLM10, KF13b]. resulting [BRH13]. results [BBD^{+}14, BH10, ML11, STT14]. resurgence [TPC^{+}13]. resuspension [PFR11, PMS10b, PMS13, PFL18]. retain [GZB16]. Retention [BMTF10, BdSC^{+}14, BCSB13, EDS14, JKM11, LA14, MLP19, MF10, PZA^{+}13, SJLH12, SS10c, TSCvS^{+}15, YRM^{+}12]. Retention-favorable [BMTF10]. retentive [PS15b]. Rethinking [GBC^{+}16]. reticulatum [EOA15]. retreat [Ano18a, HPN^{+}17, HRK^{+}18]. retreating [DRDV15]. retrospective [KHM^{+}12]. retroversa [BNE^{+}19]. return [FSM^{+}14, KFW^{+}18]. returning [NMS^{+}18]. Reunion [SHB^{+}19, DDD^{+}11, LBJ^{+}16]. Rev [BDBN11]. reveal [AAJJ^{+}19, AAGGG19, ACG^{+}11, AWH19, BSC^{+}15, BFH^{+}13a, BAH^{+}15, BGIA19, BS18b, BS18a, BPHW16, BB19, CKM^{+}16, CCS^{+}18, CAG13, CMB^{+}14, CCKB^{+}15, CFBG14, CSV^{+}15, GCS^{+}18, GSSM^{+}11, HVJ^{+}18, HVBG^{+}19, HDB^{+}10, HCL^{+}11b, HHB^{+}14, dDTC^{+}17, JGM^{+}16, JMS^{+}15, LKC^{+}15, LLBT18, LBJI14, MCW^{+}16, MWP10, MHSG18, MMWG13b, MW16b, OFT^{+}11, ORE^{+}17, ORM^{+}13, PLNE11, RB11, RPP^{+}13, RVF14, RSTM14, RNSA15, SDR^{+}15, SBE^{+}14, SPHK17, SNVH19, SMMFH16, SKK12, SRK11, SSM^{+}19, SGB^{+}10, VRSS18, VVB19, VW10, WL19, YYO14]. revealed [AMP^{+}11, ACW18, BBSM19, BMJ16, BKC^{+}19, CVW^{+}11, CML^{+}12, CGM^{+}13, DMC^{+}17, DFS13, DEWK15, FLAL10, HBP^{+}16, HCJ^{+}11, JRW^{+}14, LMV^{+}19, LCS^{+}10, LCBK^{+}13, MWA17, MFI11, MBS^{+}13, MBvO^{+}18, MIM^{+}12, NW15, PLG^{+}19, PBC^{+}15, PRK^{+}12, RLAGL12, SPP12, TBG^{+}17, TTS17, VM10, VBŽ^{+}17, WBI^{+}15, WWAS18, YMD^{+}10, ZBB^{+}11]. revealing [SSBA12]. reveals [ADF^{+}18, AGGAO^{+}18, ALE^{+}13b, AHR^{+}18, BMSB13, BDPL19, BBC19, BPB^{+}12, BBW^{+}17, BRLM17, BLPP16, BCB11, CEH18, CTZ^{+}16, CZB10, DFD^{+}19, DPY^{+}16, DKF^{+}15, DHJ^{+}17, DFDC18, DCC^{+}10, EMH^{+}19, GCL^{+}16, HTK^{+}17, HPMZ11, HLS18, KSHS19, KNS^{+}15, LSBB17, LFPA^{+}11, MHS17, MADI13, NMPM15, NMCP17, OTK^{+}13, PBAS^{+}10, PMOAC^{+}17, PBCA10, RBSL11, RMT^{+}13, SFKQ15, SHD^{+}19, TCWC13, WF18, ZAC11, ZPPM17, dSdSdCO^{+}11]. Reversal [YHR16]. reversed [HCJ^{+}18]. Review [MDNL^{+}11, Ued14, BRJ^{+}12, BW12, CFSM^{+}13, EVHA^{+}18, FABH^{+}16, FPL^{+}12, GS15, HFG^{+}19, LDS^{+}15, MJ10, MS11, NJJ^{+}11, PMEZ15b, WCP^{+}14]. revisited [ABJ^{+}15, DSH14, GCM10, SHAS15]. Revisiting [JBCGC19, TSS11b, vWSGL11, GFS15, RTÁB^{+}14, OFG^{+}18]. Rhincalanus [CWD11]. Rhincodon [MHW^{+}12, RCR^{+}13]. Rhine [WH11]. Rhine-Meuse [WH11]. rhinoceros [BCH^{+}11, CEC^{+}18, TIS^{+}15, WJW^{+}18]. Rhinoptera [OBA^{+}18]. rhizocephalan [WGD^{+}16]. rhizomes [BL14, OFM^{+}18]. Rhizophora [ATP14, GZB16]. Rhizosphere [JPL^{+}15]. Rhizostomeae [DWD^{+}18]. Rhodolith [GHES18, DDP^{+}13a, MFSRR14, MHCRR18, MG18, PFVFF^{+}15]. rhodolith-forming [DDP^{+}13a]. rhodoliths [TS16]. rhomboides [NdBKR13]. Rhopaloeides [PW11]. Rhopilema [FWS^{+}18]. Rhynchobatus [GRN^{+}16]. rhythm [SLC11]. rhythmicity [HTHM18]. Rhythms [WJWC16, ACF^{+}10, ASS^{+}13, BBC^{+}13, KF13a]. Ria [APC14a, RCGE12, BC11, FAZ^{+}14, PLFR^{+}11, THRBL^{+}16]. rías [BVP11]. Ribbed [ZZG^{+}19, KR16]. Rica [HCBV^{+}11]. rich [GMD16, JCM13]. richardsi [HMAC^{+}11]. richardsonii [PWW12]. richness [Alb15, CRRD^{+}13, CGM^{+}13, ES12b, GNC^{+}10, HPLS13, KKS^{+}12b, LRB^{+}12b, MBB^{+}18, MJ10, MRL^{+}10, RMSPA16, TP16, TPCN12, VVB19, WS13, WEF19, ZSEV17, dJH11, tHHR10]. Rico [ASWB10, BA13b, HTSA16, PVZD^{+}11]. Ridge [RSW^{+}16, SLRS13, MVTS^{+}12, RWMP12]. Riding [BDvGW16, NCD^{+}12]. ridley [PAC^{+}15, RAKB^{+}12, SL11b, dSdSdCO^{+}11]. right [BLS^{+}11, BLE^{+}13, BWLP17, CPA^{+}11, CVC^{+}18, DRT13, DTS14, FTM^{+}13, KHM^{+}12, LRR^{+}10, MGGSP15, MRB^{+}11, MBP^{+}12, NSB^{+}19, RSP^{+}16, RUS^{+}13, VRSS18, ZBK^{+}15]. rights [ABH18]. rigida [CLG^{+}19b]. Riisgård [FS18]. rim [FBS14]. ring [BFL^{+}11]. ringed [CSF13, HLIK16, YSH^{+}16]. ringens [CCG^{+}10]. Rio [MVVGAV10, DFTZ^{+}15, LGD10, PNI14]. Rippled [HKL12]. riprap [TT16]. riprap-altered [TT16]. Rise [BPN11, MMBA13, ERB^{+}13b, ERB^{+}13a, JBT^{+}16, RPE^{+}13, GOO14]. Rising [TBS^{+}17, BLER15, MCBR18, VBD^{+}11, WEK14]. Risk [EN11, SCL^{+}16, BGSM13, BPF^{+}18, CSB14, DQM^{+}19, DAM^{+}18, GRN14, GGF^{+}12b, HHS10, HMS10, HW12a, HHS12c, HMJ^{+}13, JS12a, KCGC14, KLBT^{+}16, LST11, LH10, MGP12, MJO^{+}13, Mat10a, MB11, MO13, PMSGBL16, PBC^{+}12b, PW19b, RCF17, SHMB15, SSFB11, SCF^{+}17, TMH^{+}13, YSKH11]. risk-based [SSFB11]. risks [BBM13, HG13, SNVH19]. Rissa [BER17, HCDR^{+}15, MIA^{+}11, REAN^{+}14, SMC^{+}14]. Risso [BBSM19, SWH^{+}11]. rival [FRC13a]. River [CR14, JOFP18, KVR^{+}18, KKY15, LBAK13, LB15a, NGA18, RWR^{+}15, SSD13, AMAM^{+}18, BSC17, BSB12, CL15a, CPS15, IAV^{+}10, OCSW16, PHAZ18, RMV13, WHD^{+}14, BMD^{+}11, BWRP14, HPMZ11, KEBB10, KN12, MYK^{+}16, MBKM17, RWP^{+}14, TTPM12]. river-dominated [BSB12]. river-influenced [CL15a]. Riverine [PKBS14, ABK^{+}14, FJC10, HTB^{+}17, THRBL^{+}16, WWZ^{+}19]. rivers [HS16]. RNA [BCSR10, PFG10]. roach [MJS17]. roadmap [STK12]. Robust [OGGP14, DF15, IME12, KLKM12]. Rocas [GDJdBSL19]. rochei [LCQT^{+}13]. rock [CB17, GPM^{+}13, GC13a, GB10b, GNC^{+}10, JW10, JWHR11, KBG^{+}13, LBD^{+}15, LJB14, MPH^{+}11, MCBR18, TSR17, TFH^{+}19, VDEO18, WJ12]. Rockall [RJH14]. Rockfish [THW^{+}16, BCB11, DHJ12, DT11, GCM10, GS11, HLS17, HRT12, LBS^{+}15, RHB13, SWW11, WY15, WWG^{+}13b, YIKG10]. rockfishes [JCF^{+}12, KSFL14, KFW^{+}18, KHW^{+}19, MLR17, WRFM16]. rockhopper [BM13, DED^{+}15]. rockweed [SCRL11]. Rocky [LAM11, MYS^{+}16, ANB13, Ale13a, ACS12, BAM^{+}10, BMVBC10, BHDC15, BDG^{+}16, BKTD10, BBC14, Bur12, CGOC11, CGKW10, CMC10a, CMO^{+}10, CAC11, CHRB10, CGEGC^{+}19, CC12, DGP^{+}19, EMMSL13, EHG16, Fis10, Gar13, HCL^{+}11a, HI14, KSB11, KCH^{+}11, LGMB^{+}16, LFMC11, LLBT18, LM12, LCFR10, MMG^{+}15, MN18, MSG11, MLB19, MF13, MGL^{+}16, MLSK^{+}15, ODCE11, OSPWB14, PSCS16, PMOAC^{+}17, PPC^{+}19, QHG^{+}19, RAA^{+}16, RVF14, RPHB15, SLS18, TCA^{+}11, TVH^{+}10, VANB15, WBLRN18, ZFZ16]. rocky-bottom [CHRB10]. roi [GFC^{+}14]. Role [AEGH11, BLHRP^{+}19, BTTK10, BH16b, BPG^{+}10, BVQ12, CSG11, DDC^{+}15, GS17, GLY^{+}14, KBB10b, LHB^{+}15, LLBP15, LCL^{+}13, MHB14, MDA^{+}10, MLC14, MAB^{+}10, MGG^{+}14, NC11, OBÁSV18, OMSP10, PSG^{+}13, PHG^{+}14, SQS^{+}13, TLG14, WWZ^{+}19, AOW11, BCN15, BWLP17, BO13, BW12, BL17, BKAS14, BM16, CMIMGS15, CMW^{+}14, CLMM^{+}14, CHG^{+}18, EOM^{+}14, EvDB^{+}14, EOA15, FHM^{+}11, FBC^{+}15, FFL10, FB14, GCS^{+}18, HKJ17, HKSD14, HEAI10, JNN12, LB15b, MMPH^{+}12, MIIB16, MWGH15, MSPO12, MKG10, MAL15, NVA^{+}19, OS18, OFDG14, PDBG14, PBG^{+}10, PMBH11, QOGS^{+}16, REG^{+}11, RSK^{+}14, RP15, SBH^{+}19, SS13, SCHO11a, SGD^{+}12, SPP18, SRRS12, TDSS15, WBB^{+}10, WLL^{+}14, ZTNS13, vdMPRM12]. Roles [MORM15, MKI^{+}15, CFMFFJ18, DAWS19, GBE^{+}17, GNC^{+}10, ICUM10, KHW15, LMGJ13, LR15, MSLR13, RHB13, RMSPA16, ZAMB10]. Rømø [BAA12]. rookeries [LSFQ19, SBB^{+}12a]. Rookery [GLJ^{+}18, PFB^{+}17, RPE^{+}13]. root [HCK12, JAPK16, VBŽ^{+}17]. roots [ATP14, GZB16, GCCCM16, GCCM18, JOFP18]. rope [JSS^{+}11]. rorquals [RMT^{+}13]. rosacea [TEAO^{+}14]. rosea [HRAM13]. roseate [Goy14, UW18]. roseus [GPV17, RJR18]. rosmarus [NUJ12, HKL15]. Ross [BST^{+}19a, WLB^{+}15, ABJ^{+}15, AvDL^{+}19, BKC^{+}19, DHK^{+}17, GMK18, LSN^{+}19, ZXF^{+}16]. Rossia [GCS^{+}19]. rossii [EBOM11]. round [ASAG17, APBC^{+}19, CCJR14, CPM^{+}17, HKL15, HHNI17, HB12, RPM^{+}13a, SHD^{+}19, ZMGS17]. roundnose [DDK^{+}18]. Route [HHP^{+}17, FAM^{+}17, FSF^{+}17, WHD^{+}14]. Route-specific [HHP^{+}17]. routes [AHØ^{+}15, NKM^{+}15]. routines [NWC^{+}10]. roving [AWH19]. rRNA [HPMZ11]. rubens [ATP^{+}12]. rubrizonatus [CBL^{+}10c]. rubrum [Bru10, SB10]. Rudarius [HMJ^{+}13]. Ruditapes [FRRÁSL11]. rugosa [BHP11]. Rules [RKFD15, AAGGG19, CWMP13, GGF^{+}12a, VVB19]. Rumohr [EN13]. run [BJL^{+}17, GKC11, SCAM^{+}14]. run-off [BJL^{+}17, GKC11]. Running [ASM10, WEA^{+}17]. rupestris [DDK^{+}18, KJBS12, LTN^{+}11]. Ruppia [JPL^{+}15]. rus [JRK13]. Russia [SBA12]. RUVS [RKFD15]. Ryukyu [HNK^{+}18].

S [Ano12c, AAAE17, CWR^{+}18, GDB^{+}16, MHS14, MFS14, RHRB^{+}17, VRSS18, YSC^{+}15]. S. [Goy14, ST10a, TRSV^{+}12]. s.l [SNT^{+}18]. Saccharina [BNW14, LS16, NDAP19]. Saccostrea [GB10b, TSR17]. Sacculina [WGD^{+}16]. Sackmann [HGS15]. Sagami [NOT^{+}11]. sagax [WM10]. Saguenay [PPD^{+}14]. Saguenay-St. [PPD^{+}14]. Saharan [LMR^{+}12]. sahulensis [MHAP18]. saida [SGL^{+}17]. saithe [CEK^{+}16, LANE18]. salar [DDP19, JJF14]. saliens [KBP^{+}10]. saline [BVC^{+}16, JMGG13, RLTL14]. salinities [POL^{+}11]. Salinity [LC12, SHI^{+}10, BLD^{+}16b, BSB14, BPSG11, CGS^{+}19, DS16, EOA15, GFP19, GB10b, JTTB^{+}18, JTTP19, JCM10, JSCM15, KSK07, LK15, LNIH16, MC10a, PGR^{+}17, RLTL14, SPMN19, SPM^{+}16, SKMS^{+}18, SHB^{+}10b, VSB^{+}15, WWG12, WFB16, WL11]. salinus [JMGG13]. Salish [ED12, SG18a]. Salmo [ABK^{+}14, ABT^{+}15, AHØ^{+}15, DDP19, JJF14, KRdVG^{+}18, KBGA19, OHA10]. salmon [BWRP14, CLD^{+}13, CLH^{+}15, CSM^{+}19, DABP13, DBA17, DDP19, DHM^{+}14, FTB14, FJ12, FTB^{+}17, FSS^{+}18, FSY^{+}10b, HGG^{+}18, HWET13, HRS^{+}17, HTB^{+}15, JS12b, JGM^{+}16, JJF14, KSA^{+}13, MRC10a, MWC^{+}12, MGM10, MGCH^{+}16, MCT15, ODL^{+}13, RWP^{+}14, RSDN15, RRG^{+}11, SCAM^{+}14, SAM18, SQS^{+}13, SFKQ15, STS^{+}18, SST^{+}18, TFS14, TTPM12, TTW^{+}12, Ued14, WSF^{+}12, WSS^{+}16, WHD^{+}14, WWW^{+}13, YBB^{+}12]. salmonid [JJJ16]. salmonis [GSM^{+}11a]. Salp [HBE^{+}13, HEB^{+}11, PNA^{+}16, SS14, SBS10]. Salp-falls [HBE^{+}13]. Salpa [KPH^{+}15]. salt [AED^{+}10, BFRM13, BPB18, BVGH10, CPF^{+}13, CFF^{+}18, CHDH^{+}19, CAKD14, CGQB10, CR11, EBT^{+}15, FH10, FCR^{+}15, GAB11, GSU12, HKT10, JE10, KAAI15, KV13, MCK^{+}13, MDJN19, NCT^{+}14, PSA17, QCX^{+}10, RMB18, SLRP^{+}12, SO16, TB10, ZBY^{+}14, ZZG^{+}19]. salt-marsh [MCK^{+}13]. saltatrix [JCM13, SRSH^{+}18]. saltmarsh [AMS^{+}13, CSN12, PFB^{+}13, PB18, RZB18]. saltmarshes [DdR10]. saltwater [RBD^{+}14]. Salvador [KM14]. same [PMKL11]. Sample [CS13, KTKM14]. sampled [RCU^{+}11]. samplers [CEC^{+}18]. samples [RPS12, SWK19]. Sampling [CKR^{+}11, BGB^{+}17, Bet13, BEEH15, CB12, ETD^{+}18, GMD16, GDPR18, HGS14b, MJT^{+}12, RWR^{+}15, THW^{+}16, War14]. San [GZYS11, DHB18, GK10, GSTK11, KK17, KM14, MS13, RCB12, SIK16, ZOM^{+}10]. Sanctuary [LPC^{+}11a]. sand [BH16a, BCW12, BvHC^{+}15, CG10, CBT13, DHG16, KCS12, MFDW17, PFVFF^{+}15, RPL^{+}18, TWL15, WLBB^{+}13, vDGK^{+}12]. Sandeel [ERG^{+}11, EvDB^{+}14, HCJ^{+}18, JROS17, RGW17, RHvD19, vDCFM10, vDGK^{+}12, vDCR13, vDJF15]. sandeels [FEE^{+}11, WHN^{+}18]. Sanderia [HSH18]. sandflat [PBS10, WRB^{+}15, ZPLW11]. sandflats [GKPT16, HPH^{+}15, LHL11]. sandhoppers [PHM11]. sandpiper [BEMD11]. sandprawn [PWW12]. sandy [APH^{+}10, ASK^{+}15, BGL^{+}11, CGD10, ESH^{+}10, FT10, HBHG10, KHD^{+}11, MPB14a, MAD17, MAD18, MHPvdH19, MBT14, OCDD16, QNG^{+}15, RHMR18, ST13, TUK13, BGM^{+}18]. sanguinea [JPP^{+}17, QYS^{+}12]. Santa [CSRD11, DIG^{+}11, GHES18, GMSC14, MPB13, TS16]. Santangelo [Bru10]. Santos [MSS^{+}14b]. São [HVJ^{+}18, dGCF^{+}17]. sapidissima [SSS19]. sapidus [MRH11, ORR^{+}14, ORJH19, SN17, WMH17]. saplings [CY13]. Sarcocornia [AED^{+}10]. Sarda [RBA^{+}18]. Sardina [CSR^{+}14, GN11, GCC^{+}16, NPMS11, NIPS12, SFTG14, TGI^{+}11]. sardine [ASEB11, CSR^{+}14, CMIMGS15, GN11, HHNI17, MOK^{+}18, NPMS11, NIPS12, NIS14, NYK^{+}13, OTK^{+}19, SFTG14, TNOY17, TGI^{+}11, WM10]. sardine-generated [TNOY17]. sardinella [APBC^{+}19]. sardines [MATM^{+}19]. Sardinops [HHNI17, WM10]. Sargasso [ANJ^{+}11, OGA13, RBC^{+}14, RNK^{+}11, SS14]. Sargassum [BCHG18, HCF^{+}16, KHG^{+}19, LH19, TNIH12, WHH12]. Sargassum-dominated [WHH12]. sargus [AHA^{+}15, NVA^{+}19]. Saronikos [TPS^{+}13]. sarsi [BHP11]. Satellite [BBC^{+}17, CTZ^{+}16, MWR^{+}12, dSdSdCO^{+}11, BMV^{+}11, BHL^{+}19, CSKR10, FGTS16, FMH^{+}15, HTR^{+}12, JRW^{+}14, KLBT^{+}16, LAM11, MASW^{+}11, MDNL^{+}11, QHN^{+}10, SGB^{+}10, WFH^{+}18, ZBB^{+}11]. satellite-derived [LAM11]. satellite-tagged [WFH^{+}18]. satellite-tracked [QHN^{+}10]. Satellite-tracking [dSdSdCO^{+}11]. saturation [Fie14, LRH14, TTF^{+}11]. saving [OMW^{+}16]. sawfish [FKP^{+}15]. saxatilis [CPS15, WG11]. Saxitoxin [FPC^{+}12, SE11a]. scabra [CLP18]. Scale [BCS^{+}12, GTP^{+}17, KPB11, NSS10, PSS^{+}12, RFBH^{+}11, RHvD19, TSTP13b, AGK13, ABF^{+}14, AMS^{+}13, AED^{+}12, AGC12, BSB^{+}18, BT10, BST^{+}19a, BRK^{+}12, BSC^{+}15, BRBH19, BPB^{+}12, BRLM17, BHL^{+}19, BJS^{+}16, CKM^{+}16, CW12, CDWC17, CFPF^{+}14, CASG^{+}19, CM11, DMd14, DH18a, DCL^{+}17, DER^{+}19, DPH13, ELA10, Edm13, Edm14, EMJ18, FSD^{+}14, FYG^{+}18, FDFS16, FCFD15, FHL^{+}15, GMB^{+}16, GMHS19, GBJ^{+}14, GGCB16, GCCM18, GHNF10, HMKF11, HTB^{+}15, HSWS16, JU16, JD18, JM17, JMK^{+}18, KW17, KHGS12, KLKM12, KHÁB^{+}17, KMP11, KRC^{+}11, KBB^{+}10a, LLMA^{+}10, LAM13, LDS^{+}15, LMLB12, LR18, LOBW^{+}11, LCR^{+}16, LGR^{+}10, LBWB^{+}13, LCB^{+}16, LGT^{+}14, MSFM12, MGP11, MMG^{+}15, MOP^{+}15, MNKB11, MRL^{+}17, MHS17, MHVV13, MWB13, MJT^{+}12, MLR12, MCT16, MTG^{+}11, NDBL16, OSK15, ODK^{+}14, OMW^{+}16, OWN^{+}15, PDM12, PBTB15, PMSGBL16, PSL^{+}12]. scale [PWM^{+}19, PGR^{+}17, PCLB18, PDBB14, RSO^{+}13, SLS11, SRR15, SSR^{+}10, SW18, SWH^{+}18, SBM^{+}16, SSB^{+}16, SRM^{+}15, ST08, Sma13, SBE^{+}16, SYM^{+}16, Sti13a, SWPK13, TPC^{+}16, TPBM13, TR13, TLAT13, TTB^{+}11, VHHGF11, WLM12, WSLM13, WWG13a, WTD^{+}19, YAFV12, YPWP18, ZAC11, ZVL^{+}16, dML11, vERRP16, vdMPM^{+}10, vdMPM12]. Scale-dependence [TSTP13b]. scale-dependency [KMP11]. Scale-dependent [BCS^{+}12, GTP^{+}17, KPB11, NSS10, PSS^{+}12, RFBH^{+}11, LOBW^{+}11]. Scale-specific [RHvD19]. Scales [GCCCM16, BC11, Bar14, BBBNT13, BBMH13, BJS^{+}11, BMPC15, BPB11, CSKR10, CTMU10, CSS^{+}13, DAD^{+}11, DBS^{+}19, GVW^{+}12, HHLT10, HLS15a, HL14, JMS^{+}15, KBG^{+}13, LMG12, LKB^{+}14, MTMM^{+}15, MGOL19, MFG16, NNT^{+}10, PMH^{+}13, SHR^{+}11, SH14, TMBP13, VGL16, Vig15, WLF^{+}17, WLHW10]. Scaling [CSKR10, Ree14, TSTP13c, BHAC17, CWS13, CS11b, HORRLS^{+}14, RS13, SMHB14]. scallop [BWJB16, BCW19, BAS16, CWS13, HW11b, HTR^{+}11, KTR^{+}16, MHK11, SHB^{+}13a, SHF12a, TG12a, TPC^{+}13, TPC^{+}15, WMMC14, WA19]. scallops [BCSR10, DGDJ14, HS11a, RJR18, SCHO11a, SCHO11b, SDH13]. Scandinavia [FAM^{+}17]. Scandinavian [RMKR13]. scanning [MRSD15]. scans [HJ13]. scarce [GVGH15]. scared [GSL17]. scarers [BHD^{+}13]. scarps [RTDUM16]. scars [BKB11]. Scarus [AZVB13]. scat [ZO10]. scats [dDTC^{+}17]. scattering [NKA16, PIGK16, PCLB18, SBB17]. scavenger [HKI^{+}17, YD11]. scavengers [DMK11, OMBP10]. scenario [LCFR10, RZS12]. scenarios [CDMW12, KCB18, NDAP19, ULCAZG17]. schedules [DGC12]. schlosseri [BSRDR10, RMKR13]. schoenleinii [SKK^{+}18]. school [LFJ^{+}14, MHSG18]. schooling [GE14, HNG^{+}11, JVSD10]. schoolmaster [MDA^{+}10]. schools [JROS17, TGSM12]. sciaenid [NCD^{+}12, SBFP10]. Sciaenops [ARA^{+}10, NRF11]. science [HMB^{+}12, HSMB17, PHB^{+}11a, PCD^{+}18, RG11, TRJ^{+}11]. sciences [KPLD^{+}12]. Scientific [Nob11, TVA^{+}19]. scintillans [ZLCW15, ZLGH16]. Scleractinia [WMS^{+}11]. scleractinian [AHS^{+}13, HCA11, IHP10, KCLLL13, LKN^{+}13, LHW^{+}19, TGM^{+}17, TPA^{+}13, TPFP11, TNS^{+}12, WBB13]. sclerochronology [Ano18b, HLH18]. sclerotized [CBR19]. Scomber [ST10a, YYO14]. Scomberomorus [RFH16]. scope [BSN^{+}18]. Scophthalmus [PWO^{+}16]. Scorpaenidae [DT11, KSFL14, MG14]. scoter [LET^{+}12]. Scotia [MSK^{+}19, RSW^{+}16, BST19b, KTF10, SCW^{+}15a, SCST17, SFLG10]. Scotian [AB12, DTS14, FSLG12, FS14, HKB^{+}19a, KS11, SJS^{+}11]. Scotland [DJL^{+}12, BEG^{+}13]. Scottish [Hug11]. scour [HKL12]. scoured [MLSK^{+}15, RJH14]. screening [SJLH12]. screwshell [RJR18]. scriba [MPA^{+}10]. Scrobicularia [SCO^{+}12a, SAL^{+}12, VCDP11]. SCUBA [BPC17]. sculpins [TB12]. Scurria [AVB13]. scutellum [JMPZ10]. scutulata [VE18]. Scyliorhinus [CJM^{+}13]. Scylla [BBMM14]. scyphistomae [WFB16]. scyphomedusae [BCPC18]. Scyphozoa [DWD^{+}18, IMU17, SMJ^{+}14, WFB16]. Scyphozoan [FHG^{+}14, ANWG18, HSH18, HSS18, MAHK12]. scyphozoans [FWS^{+}18]. SE [HVT12, HWE^{+}12, HvEBC12, LRC^{+}13]. Sea [ABJ^{+}15, BST19b, CLMM^{+}14, CMM^{+}18, DHK^{+}17, DAW^{+}11, EBdlTC12, ECZD14, FDS14, FRC^{+}13b, GMK18, GCC^{+}10, HS14a, HPN^{+}17, JBD10, KH10, KJCN^{+}11, LCBK^{+}13, MLF^{+}18, PMIG19, RPE^{+}13, REP18, TKDM16, UHG12, YYD^{+}15, ZXF^{+}16, dlGMD^{+}17, ABK^{+}14, ABT^{+}15, AJF^{+}12, ARW12, ASAG17, ARD^{+}19, Ano18a, ATO^{+}14, AJ16, ACMD11, AGP^{+}13, BMC^{+}15, BWS^{+}12, BBM13, BJM19, BDM^{+}13, BSG^{+}10, BCW19, BOC^{+}16, BFF^{+}10, BAS16, BBH^{+}13, BBC^{+}12a, BLT14, BP12b, BCH^{+}18, BFPONS^{+}12, BL13, BBG^{+}16, BHTF11, BRIC13, BSY13, BGBD^{+}13, CT10, CWR^{+}18, CAC^{+}12, CKM^{+}16, CWS13, CRR^{+}19, CM14, CLP18, CPF^{+}16, CRGM18, CLS^{+}16, CD16b, CATAV18, DIG^{+}11, DMK11, DLRF15, DRG^{+}12, DBCS16, DFB^{+}17, DCC^{+}10, DFTZ^{+}15, DP14b, DGFH10, DAM^{+}18, DBH^{+}16, Ebe10, EHR12, ECM^{+}12, ED12, EMG^{+}15, ETB13, FMP13, FNC13, FNC^{+}14, FC10]. sea [FPC^{+}13, FSLG12, FS14, FFS14, FDLL10, FJL12, FDP^{+}16, FMC^{+}11, FMC^{+}10, FCRD18, FG16b, FA10b, FKRJ13, GGL^{+}17, GSA^{+}18, GSD10, GFF16, GFP19, GRL^{+}15, GGFB15, GG12, GCC^{+}14, GDA^{+}18, GCL^{+}16, GFW^{+}14, HAM^{+}15, HLIK16, HPMB11, HS11a, HG13, HRS^{+}17, HBE^{+}13, HCG^{+}10, Hv14, HMS^{+}12, HSMB17, HWR^{+}14, HRS^{+}16, HDT^{+}12, HSS^{+}19, Hug11, HRK^{+}18, IHP10, IVD^{+}10, JBT^{+}16, JCA13, JUK^{+}10, JFK12, JWR^{+}12, JH11, JH14, JPI^{+}19, JKM11, KKS12a, KB14, KFBM19, KRdVG^{+}18, KBGA19, LSY^{+}18, LOS^{+}19, LNS10, LI18, LPC^{+}11a, LBVH13, LH16, LRB^{+}12b, LWS^{+}10, LSFQ19, LMP12, LKL^{+}17, LS19, LCCT19, LTN^{+}11, LMSC^{+}19, LHHG11, MWA17, ML11, MWR^{+}12, MMPH^{+}12, MAT^{+}11, Mat10a, MM13b, MMSC17, MWWS17, MCS^{+}15, MTMM^{+}15, MMB^{+}12a, MYS^{+}16, MLB^{+}14, MBB^{+}12b, MBvO^{+}18, MDPG13, MWGH15, MLML^{+}19, MS11, MSM^{+}12, MB11]. sea [MvdHBM12, MGG^{+}14, NTN^{+}13, NC11, NSY^{+}18, NOT^{+}11, OTS^{+}10, OPSM13, OC17, OFG^{+}18, OMW^{+}16, PMB^{+}17, PvdMLC17, PAFC13, PK15, PSS^{+}12, PSAK12, PAC^{+}15, PPSR15, PHHL19, PPC^{+}12, PSC13b, PMF^{+}17, PCW^{+}12, PHWC13, PRBR11, RP13, RMP^{+}19, RAB^{+}10, RAKB^{+}12, RHW12, RWMP12, RCS^{+}17, RCE^{+}13, RCB^{+}10, RSE18, RP17, RM19, SFLG10, SDR^{+}15, SLSC16, SPHK17, SLL^{+}18, SL11b, SNW^{+}18, SHF12a, SK10, SJL^{+}16, SGC17, SBB^{+}14, SKR^{+}10, SDMR19, SDF^{+}19, SV12, SCHO11b, SCW^{+}15b, SHM10b, SGD^{+}12, SST^{+}18, SBW16, TOW10, TBMG17, TMM^{+}19, TWM^{+}15, TKB^{+}13, TTF^{+}11, TMP^{+}13, USR^{+}16, UFP^{+}10, VHV13, VJM^{+}16, VGR^{+}16, WBB^{+}10, WHB^{+}14, WFS10, WTGC^{+}11, WSC13, WWG13a, WFH11, WCP^{+}14, WW15, WFH^{+}18, WHH12, WPMP14, WHP^{+}17, YSC^{+}15, YD11, YTM^{+}15, YSH^{+}16, ZHM^{+}10, ZVL^{+}16, ZVR18, vdMHW^{+}10, AvDL^{+}19, AOW12, ATT^{+}16, AFLM12]. Sea [AF14, ANJ^{+}11, AWJ12, ADvdM13, Ano18a, BAA12, BST^{+}19a, BZBR^{+}11, BCN15, BW13, BLE^{+}13, BHE14, BvHC^{+}15, BCvH^{+}15, BBI16, BBD^{+}14, BBMH13, BMV19, BJ14, BSC^{+}16, BPM^{+}18a, BD15, BDDvdM17, BDD17, BPSG11, BFH^{+}13a, BBG^{+}13, BDR19, BDBN11, BGS^{+}16, BAE^{+}10, BGK17, BDGB19, BDB^{+}12, BWN^{+}14, CAR^{+}19, CBC^{+}10, CMIMGS15, CBR^{+}18, CBMK10, CJM^{+}13, CPTL11, CRRL17, CSPA10, CMC^{+}15, CEK^{+}16, CGG^{+}10b, DSG15, DAL^{+}10, DVSM10, DLC^{+}13, DRD^{+}18, DJL^{+}16, DKRS12, DIB15, DLL^{+}18b, DKB^{+}12, DHH10, DFAH11, DPD^{+}12, DMvB10, DvBH12, ED12, EBH16, EBDB18, EKOÅ13, FC10, FCDC12, FEE^{+}11, FQ11, Fuj16, FSY^{+}10b, GALS16, GKBEF18, GCR^{+}17, GWH^{+}18, GFW^{+}14, HGL15, HKL15, HB13, HCJ^{+}18, HBE^{+}13, HTCH^{+}18, HFR^{+}13, HMM^{+}15, Hof10, HIH^{+}17, HPH^{+}18, HDT^{+}12, HVD^{+}16, HEAI10, HRK^{+}18, HMF^{+}18, JRvdM15, JWR^{+}10, JUK^{+}10]. Sea [JFK12, JLM^{+}12, JNN12, JDG^{+}17, JPMG10, KSA^{+}13, KHW^{+}10, KBW^{+}11, KKL13, KKMM16, KH18, KKS^{+}14, KTF10, KHT^{+}17, KTP^{+}18, KLN^{+}13, KLNJ15, KDB^{+}10, KHS10b, KRDV18, KLMR14, KOPW15, LSN^{+}19, LCQT^{+}13, LSV^{+}10a, LTP^{+}18, LPC^{+}11a, LKL^{+}19, LGB^{+}13, LJM^{+}12, LCF^{+}14, LLC^{+}14, LAALH10, LRC^{+}13, LNK^{+}16, LBC^{+}19, MMH^{+}17, MKPS13, MMDM12, MMRC16, MASMH19, MRE14, MBW10, MG14, MSS^{+}14a, MVS^{+}10, MAS^{+}10b, MG12, MGB^{+}14, MDR16, MAB^{+}10, MHF^{+}14, MLP^{+}13, MFG16, MCHC^{+}16, MDJN19, MRC^{+}13, MOK^{+}18, NdBKR13, NWHD15, NTS^{+}18, NKA16, OV10, ØLH^{+}13, OTK^{+}13, ORE^{+}17, OGA13, OBI^{+}16, PSÓ^{+}15, PPM^{+}15, PSPR18, PSC^{+}13a, PAP^{+}12, PPSR15, PQAH^{+}17, PBC11, PLZL14, PTF^{+}14, RNB^{+}12, RCAB^{+}12, RCK^{+}11, RDK14, RBC^{+}14, RNK^{+}11, RLTL14, RHvD19, RDC^{+}12, RBC^{+}15, RSG^{+}16, RHS^{+}19, SCD^{+}17, ST10a, STYT10, SCW^{+}15a, SCST17, SPJ^{+}12]. Sea [SHKR15, SLD^{+}13, SG18a, SdBAH^{+}14, SLKN14, SDAM11, SRH^{+}15, SSB^{+}16, SMJBØ11, SDAH12, SBH^{+}18, SSM^{+}16b, SSvD^{+}15, SS14, SLS16, SHSK11, SKPS^{+}16, TFFPM18, TBCG^{+}12, TSS11a, TSS11b, TGRH17, TB12, TRHM14, TBC^{+}17, TS14, TGI^{+}11, UMF^{+}10, UDF^{+}18, VGGF12, VDMLU13, VBŽ^{+}17, VLK^{+}15, WL10, WWB14, WBI^{+}15, WHN^{+}18, WNG^{+}19, WI12, WN16, WLB^{+}15, WAC^{+}14, WNNH10, WKK18, WK18, WGGM11, WLF^{+}17, XLXS17, YMMO14, YYO14, YHTC19, ZBK^{+}15, ZCC^{+}12, ZRBV15, ZAZS11, dB10, dQSC13, dQSC14, vDJF15, vdVKA^{+}11]. Sea-cage [MLF^{+}18]. sea-finding [PK15]. sea-ice [Ano18a, DLRF15, GCC^{+}14, HLIK16, HRK^{+}18, MAT^{+}11, PMF^{+}17, RCS^{+}17]. sea-induced [BP12b]. sea-level [JBT^{+}16]. Sea [?]NorthHansen:2013:POD. sea-spawning [VHV13]. sea-surface [LNS10, NTN^{+}13, WSC13]. seabass [SCSC16, VMN^{+}12]. seabed [FRJ10, FRJ11, HPLS13, JROS17, SHB^{+}13a, TAEB^{+}11]. Seabird [CCJR14, ERB^{+}13a, GE14, GZYS11, MFD12, PGOG12, ARW12, ACAD12, AEGH11, ADA^{+}19, BBW^{+}17, BGM^{+}17, CCB^{+}17, CMK^{+}14, CSS^{+}13, DWLH13, DHS^{+}19, DKF^{+}15, DTL^{+}15, FSF^{+}17, FGW^{+}15, FSP^{+}17, GDH^{+}19, HHS12c, HB12, HTR^{+}19b, JWB^{+}17, KKMA^{+}10, KEH10, LSJ17, LLMA^{+}10, LBM^{+}13, LBJ^{+}16, Løk11, LANER15, LNF^{+}11, MHB14, MBB^{+}12a, MWC18, MPG^{+}10, MCP^{+}18, MGB^{+}19, NWHD15, OKF^{+}18, PGR^{+}10, PHI^{+}12, POR^{+}15, PRLR^{+}18, PHAZ18, PJP^{+}11, QSvN^{+}11, RDHM10, RPP^{+}13, RP15, SRK^{+}18, SES12, SKM12, SBE^{+}14, SABMG16, SHD^{+}19, SSS12, SKPS^{+}16, TF12, TSS^{+}12, TSTP13c, TSTP13b, WBG^{+}14, WHN^{+}18, WSC13, WV12, YBD^{+}16, ZHTA13, ZMGS17, vBMH^{+}17, ERB^{+}13b, KH18]. seabird-fishery [TSTP13b]. seabird-minke [AEGH11]. Seabirds [CAGS17, LANE18, STK12, SK12, WJS^{+}14, ARW12, Ano17c, Ano18a, BRJ^{+}12, Bea11a, Bea11b, BJS^{+}11, CGM^{+}19, CSC13b, GTL^{+}16, GGB^{+}18, HRK^{+}18, JCRC10, KHI12, MC12a, MYK^{+}16, MM13a, MJV^{+}11, OWH^{+}19, PBC^{+}18, PPX17, PLGSD17, RKM^{+}10, RGDNGS11, SDG^{+}12, SPP12, SDKS19, SNS12, WI12, WHH^{+}13, WBS^{+}12, WN16, WK18, YMD^{+}10]. seabream [DDB^{+}14, KPW^{+}19]. seadragon [SCMSB^{+}11]. seafloor [BRC^{+}15, CKV12, MNKB11, SGJ^{+}18, ZSEV17]. Seagrass [BL14, DPO^{+}17, GOC^{+}18, HFM^{+}15, JTTB^{+}18, JTTP19, KCMH^{+}17, MRH11, PH11, SV15, SMK^{+}17, SLPP^{+}14, TVRG^{+}13, WWGD13, ABF^{+}14, AFC^{+}16, AFS19, AMC^{+}11, AL12b, AM18, AM19, BEP^{+}12, BVL10, Bar13, BH16a, BCGWS19, BCW12, BJH13, BH15, BHF^{+}16, BS12, BBdG^{+}19, BRA^{+}10, BKF15, BFH13b, BDGB19, BRIC13, CPEE18, CB14, CESAM12, CLGC16, CCT12, CSS^{+}17, CVP11, CEW15, DZDB10, EOM^{+}14, FWF^{+}10, FMC^{+}10, GMC^{+}12, GLM^{+}10, GKBEF18, GBK^{+}14, GWME16, GDPR18, HFMD16, HFM12, HHS10, HMS10, HVK^{+}13, HH15, HOL^{+}17, HHH^{+}14, Hv14, HLB^{+}16, HWTH12, HMJ^{+}13, HNMK15, HBS10, HLD12, IOS^{+}12, JSBP15, JREVB18, JSOK18, KVY^{+}13, KCB18, KAMA13, LH11, LMRRB17, LMW^{+}13, LHC^{+}15, LPSK17, LBMZ15, LRK^{+}15, MMPH^{+}12, MIIB16, MD16a, NTB^{+}17, OD15, OFM^{+}18, PG15, PBIO^{+}18, PJDC^{+}13, PBG^{+}10, PW19a, PBC^{+}15]. seagrass [PSC13b, PRFA11, PRA10, RSO^{+}13, RU11, RDPR15, RK15, SMSE19, SSL^{+}14, SM10b, SCK^{+}17, SKA^{+}14, SLO^{+}17, SRK13, SYM^{+}16, SL15b, SBLM14, STRP18, SV12, SLO19, SBMÅ12, TYVR^{+}15, TVH^{+}10, VBM^{+}13, VMW^{+}10, VBJR^{+}13, VBGO^{+}16, VBŽ^{+}17, WDH^{+}12, WAMS^{+}14, WADKT13, WHB^{+}19, WFBM14, YKBB12, dlSBB^{+}10, dlSBO^{+}12, dlSOV^{+}16, dlSGS17, vMVSW12, vdGSE^{+}14]. seagrass-unvegetated [GDPR18]. seagrasses [BCM11b, HCK12, RLO18, VBM^{+}11]. seahorse [Har16]. seahorses [SGT12]. Seal [BHD^{+}13, BCG10, BHP12, BSB^{+}11a, BSBR^{+}12, BM16, CSF13, DHJ^{+}17, FHMG19, GNHLN11, GVGP^{+}14, HHS^{+}12b, HSCA17, HBK^{+}12, dDTC^{+}17, JLPG17, LCJ^{+}12, LJM^{+}10, LBL^{+}12, MJTC^{+}12, MAB^{+}12a, ØLH^{+}13, OALK14, OGH18, SBA12, SBH^{+}18, SMR^{+}11b, TLJ^{+}11, WRP^{+}17, vBBW^{+}14]. seal-associated [SBH^{+}18]. seal-fishery [OALK14, OGH18]. seals [ASR^{+}19, ACG^{+}11, AGNH^{+}19, BTO^{+}18, BBBNT13, BMJ16, BPT15, BBH^{+}13, BWP^{+}11, BPG^{+}13, BKC^{+}19, BCH^{+}18, BKAS14, CMHH17, CBF^{+}19, FMH^{+}19, GKHM13, GPW^{+}10, GWK^{+}19, HLIK16, HLF^{+}19, HKL19, HMAC^{+}11, HA14, HCW^{+}15, HAHM16, KBA17, LSN^{+}19, LCB^{+}16, LRO^{+}17, MBSG11, MRLD19, MTP^{+}10, MCS^{+}15, MMB^{+}19, NCC^{+}14, NFC^{+}11, OHCG14, OGLH17, OALK14, RLN17, RLAGL12, RDGC17, VGGD^{+}15, VCG^{+}15, WTdB16, YSH^{+}16, ZO10, dDTA^{+}18]. seamount [Mat10b, TTF^{+}11, SDF^{+}19]. seamounts [BRC^{+}15, HP11b, Met11, TRMC11]. search [FCRD18]. Searching [HMW^{+}12, PDM12]. seas [RZS12, SSR^{+}10, DBS^{+}13, SHI^{+}10, SHB^{+}10b]. Seascape [BPSK11, DBS14, DBE^{+}18, PKSN11, WWG13a, ASCHM19, HOL^{+}17, HRÁB^{+}14, HPN11, KMP11, OCPM12, PKS11b, RKD18, ROKS14, SDC^{+}16, SBM^{+}16, SH17, TMSG19, WLP^{+}11, WvLF18, YAL12, TAEB^{+}11]. seascapes [GOC^{+}18]. seashore [MF13]. seashores [OSPWB14]. Season [HCK12, RH12, SGM^{+}14, ACRLB12, BTS10, CRAP^{+}12, Coo11, FBA11, FSD^{+}14, GMP^{+}11, HMO^{+}12, LSN^{+}19, LHT^{+}17, RCGE12, Tho15, WNG^{+}19, WSC13, vdGvdLvG^{+}19]. season-dependent [FSD^{+}14]. Seasonal [AL12b, BBC^{+}13, BVS^{+}16, BPP^{+}18, BRMT11, BCvH^{+}15, BRS10, BCSR10, BNS^{+}14, BH12a, BCH^{+}18, BLM10, CA14, CMC^{+}15, CLS^{+}16, DHG16, DHH10, DvBH12, ETD^{+}18, FFLS16, dSKFV12, GW15, GGMGR12, GSV14, GFW^{+}14, HS14a, HCF^{+}16, HMBP18, HSS^{+}10, JSS^{+}11, JBS11, JKM11, JMB^{+}18, JPAM^{+}15, KNIN17, KBK^{+}15, KCS12, KTH^{+}18, KDB^{+}10, KFGSMPG14, LKDP17, LERS13, LRSB19, MASMH19, MF15, MAS^{+}10b, MS13, NW15, NLGE18, OMBP10, OF10, OHR^{+}13, OWW^{+}17, PDP^{+}10, PJDC^{+}13, PAK15, QOGS^{+}16, SMND16, Sat17, STK^{+}13, SGCF^{+}12, SMR11a, SDCC12, SML^{+}14, SMJP^{+}15, SP12, TBDM11, WL10, YHT16, AGGAO^{+}18, APC14a, AH13, Ano17c, AGIW16, BAL14a, BAL14b, BDM^{+}13, BB14c, BCHG18, CPB^{+}11, CGK13b, CPF^{+}14, CO16, CNÁCP12, DFD^{+}19, DJJ^{+}16, FDLL10, FCH^{+}10, GJSR18, HTHM18, HW12a, HMS^{+}12, JCRC10, KEBB10, LAL^{+}18]. seasonal [LMD12a, LALCM18, LDP^{+}17, MGB^{+}14, NCK^{+}14, NWC^{+}10, PSER17, PRA10, Rai12, RCD^{+}13, SDKD13, SBG^{+}16, SJL^{+}16, SFKQ15, SL15b, SKR^{+}10, SA11, SSD13, TIS^{+}15, TWL15, TAAN15, TLJ^{+}11, TEAO^{+}14, VCMS^{+}18, WJWC16, WTdB16, YTO^{+}11, Gal14a]. Seasonality [Bol17, CKJOM14, RBC^{+}15, TBC^{+}17, ABW14, IDT^{+}18, NKH^{+}10, PSG^{+}13, PBC^{+}18]. seasonally [AZVB13, FHE^{+}15, MMGB14, PFL13, SE16]. seasons [CCG^{+}10, DPO15, MCHC^{+}16, OCSW16]. seastar [WGLDB15]. seatrout [CLB19, EBH17, LBWB^{+}13, WLBB^{+}13, ZBB^{+}19]. seawall [MCT15]. seawalls [BC14, ICUM10, LLBT18]. seaward [JJJ16]. Seawater [BA13b, SSM^{+}16a, ATP^{+}12, BPWF14, BSC^{+}16, FRRÁS^{+}12, GP11a, HLS^{+}15b, MTH15, SHP^{+}14, SSM15, SMS15, VBD^{+}11]. Seaweed [LPJ^{+}13, YRL^{+}10, ABW14, BSD17, CE11, CLG^{+}19b, HWR^{+}13, JL17, KBS16, LTP13, LSV10b, MHF^{+}19, NT10, TZF11, TKBN10, WS13]. seaweed-epifauna-fish [NT10]. Seaweed-mediated [YRL^{+}10]. seaweeds [AHPKS17, FAT^{+}15, GH11, HTK^{+}17, PBF10, SJHH^{+}17]. Sebastes [GCM10, HRT12, JCF^{+}12, KSFL14, KFW^{+}18, LBS^{+}15]. Sebastião [dGCF^{+}17]. Second [SDKD13, dML11]. Second-order [SDKD13]. Secondary [JMGG13, BFH^{+}13a, CRHV14, DJL^{+}16, DDP15, GBU15, LMGJ13, VCDP11, VB19b, WSK^{+}16, WPP11, YAL12]. secretion [BC11, SP15]. Section [Ano17a, Ano17c, Ano18a, Ano18b, BH12b, BLM^{+}12, CLPiU14, DMP12, DS13, HBS^{+}13, HGA^{+}16c, KBH11b, LPSL14b, PHB^{+}11a, SS15a, SK12]. sectional [JGCRS12]. sector [KSK^{+}12]. securis [BVO16]. sedentary [VRAP^{+}14]. Sediment [MF13, SML^{+}15, TPS10, ATS14, AS13, BMC^{+}15, BS14a, BVL10, Bar14, BMV19, BPN11, BDM^{+}14, BD11a, BBG^{+}13, BCK15, BVS^{+}11, BPLT14, CTMU10, CY10, CAB^{+}14, CDR12, CL15c, CTvdM^{+}08, DBCS16, DWM15, EBH16, ESH^{+}10, FGOS16, FFLS16, FNA^{+}16, GGNL^{+}17, GFP19, GB10a, GWME16, HR12, HPH^{+}15, HKS^{+}12, HH15, HCK12, HSR^{+}10, ITT14, JSOK18, KHD^{+}11, LS12, LMW12, LPKT^{+}19, LMT15, MRC^{+}12, MFG16, NNT^{+}10, OLM^{+}12, PDBG14, PT13, QOGS^{+}16, ROE17, RJ12, SPP15, SGTP16, TBKG10, TNSN11, VNN10, VCQM^{+}16, VNK^{+}12, WKE^{+}17, WAA^{+}19, ZPLW11, ZF19]. sediment-induced [LMT15]. sedimentary [CGR^{+}15, FDS12]. sedimentological [MG18]. sediments [AM18, AM19, BRK^{+}12, BKC14, BHMCM16, CRG^{+}16, CFC^{+}12, ETAE13, FH10, FT10, FBNJ13, GRCP^{+}10, GPV17, GMK^{+}19, HBHG10, JNN12, JLQ^{+}14, KV13, KRDV18, LKC^{+}17, LRH14, MPB14a, Mar17, NSG^{+}12, PDGR^{+}13, RZB18, SM10b, SBB^{+}14, SBE^{+}19, SO16, TKW^{+}14, VKH10, VQKF14, VCMS^{+}18, VK11, VPW^{+}12, WLF^{+}17]. SEE [FQ11, AS12, Bri12, CFA14, FBC^{+}15, GBC^{+}16, LFRF13, MBB11, MM11, ST11b]. Seed [BO13, HNMK15, OMM^{+}12, ZBY^{+}14, BVL10, MO12, PÁIIS^{+}19, RWM^{+}12, SCL^{+}19, SLL^{+}18, WWKS15]. seeding [EGH^{+}15]. Seedling [MO12, RZB18, BBH^{+}11, JSOK18, MORM15, YHR16, ZBY^{+}14]. seedlings [DD15, MO12, VCVK^{+}10]. seeds [DD15, IEM16]. Seeing [FG16a, AWH19]. seeking [HNG^{+}11, SMF^{+}12]. Seep [RDC^{+}12, ACF^{+}10, AY10, BCM^{+}14]. seepage [PBV^{+}11]. seeps [ÅCAC16, DRG^{+}12, TKL^{+}11]. segmentation [TAEB^{+}11]. segmentation-based [TAEB^{+}11]. Segregation [MTP^{+}10, BOC^{+}16, CWB^{+}15, CALB10, DFTZ^{+}15, GN11, HDT^{+}12, KSBM^{+}11, LMBM18, LNF^{+}11, PvdMS^{+}15, PRM17, RPR16, SKSH19, SRK^{+}18, SJB^{+}15, TCTB11, WCRCR17]. sei [BLS^{+}11]. Seiches [BJG^{+}11, KdEPW18]. seine [ECC^{+}15, HIMH15, Nor13, PLG^{+}16]. seismic [RKT16]. Sekisei [SKFY18]. selecting [RPL^{+}18]. Selection [GPW^{+}10, RLN17, ADO^{+}12, AHD12, Ano18b, BLV^{+}19, BTS10, BTTK10, BS19b, CMP^{+}13, CGH11, CWB^{+}15, CTS^{+}15, EMJ18, FDLL10, FYG^{+}18, GBR11, GMP^{+}11, GCSF18, GvSGCK14, GCL^{+}16, HBM^{+}15, HW11b, JCF^{+}12, JGCRS12, LWM15, MHS17, MPF^{+}15, NSK^{+}14, PH11, RCS^{+}17, RB15, ST10b, USR^{+}16, WAW^{+}12, Wir14, WPK10]. selection-based [AHD12]. Selective [APC^{+}14b, DSC13, MZC^{+}10, NHPW18, PHAZ18, AB17, BSB14, BS14b, CMW^{+}14, CZH^{+}13, FF10, JNYT13, JGCRS12, LGMB^{+}16, LCS^{+}10, PB19, SMSL17, WRB^{+}15]. selectively [NFN^{+}15, ORJH19, TSR17]. selectivity [BSST13, BMF^{+}12, BSD17, CLL13, HM10, JCM13, LNP12, MWHM18, SSS^{+}10, VF13, WvLF18, ZS17]. selenium [SLE^{+}13]. Selenka [SGD^{+}12]. self [BEFC15, CGH^{+}16]. self-organize [CGH^{+}16]. self-reproducing [BEFC15]. Semi [DGDJ14, FESPR14, KdEPW18, LFJ^{+}14, LSJ^{+}15, YChH11]. Semi-annual [DGDJ14]. semi-automated [YChH11]. semi-enclosed [KdEPW18, LFJ^{+}14, LSJ^{+}15]. semi-terrestrial [FESPR14]. Semibalanus [BJRH10, FSR10, GPS13, HWVM19, SM17b]. semifasciata [NCK^{+}14]. semilunar [RWR^{+}15, SLC11]. Semimytilus [ZBP18]. Semmens [Fry13b]. senegalensis [TRSV^{+}12]. senescence [BSH^{+}18]. senile [NC11]. sensed [BFF^{+}10, SSS12]. sensing [BSC^{+}16, Dav19, GSF18, HSH12, HR19, RIM^{+}13, SKD^{+}11a]. sensitive [RAH^{+}10, RSS^{+}16, TB14a]. Sensitivity [DB11a, GBL^{+}15, NGA18, RCBE15, ABW14, BPM^{+}18b, MLPC18, McC17, MBBK16, MMGB14, PCNB16, RMC^{+}14, WBW^{+}14]. sensor [ORE^{+}17, TFW^{+}12]. Sensory [CALB10]. sentinel [SHMOS17, SHD^{+}19]. separate [WYI^{+}16]. separating [BC16, EOM^{+}14]. Separation [CGH^{+}16, CKM^{+}16, DGFH10, HMH^{+}12, HLS18, HSCA17]. Sepiolida [GCS^{+}19]. septicaemia [JBM^{+}13]. septicemia [HGG^{+}10]. sequence [BS18b, BS18a, CML^{+}12, GM10, HPMZ11, RB11]. sequences [SBB^{+}12a]. sequencing [AGGAO^{+}18, CR16, FKP^{+}15]. sequential [BMPH14, GGvW^{+}15, HG17]. sequestered [GWME16]. sequestration [AAvS^{+}13]. Ser [ACA^{+}10b, BW16b, RUP11b]. Seriatopora [RMB^{+}14]. Series [ETD^{+}18, Kin10a, CR16, DLC^{+}13, EAB^{+}19, EM11, FGL^{+}12, GSD12, HSA^{+}15, NGA18, SBKA14, SS14, SBS10]. Seriola [HTK^{+}17, MSF18]. serpulid [Hug11, LSQ^{+}13]. Serranus [AMP^{+}11, MPA^{+}10]. serrata [BBMM14]. serrator [CVW^{+}11]. serratus [DRDV15]. serve [vMVSW12]. service [SLRP^{+}12]. services [GS15, PMB^{+}15, PS19, PRB^{+}17, SCRL11, TTC11, ZZG^{+}19]. sesarmid [CMF^{+}19]. Sesarminae [CY10]. sessile [BB15, DMK11, ES12a, RPHB15, SLT^{+}19, SW12, Sma13, TMM15]. sessilis [BNW14]. seston [AFSL^{+}15, PFG10, PMS13, PFL18, THC12, WL10]. set [BD13, WBW^{+}14]. setiferus [BM10a]. setosum [BGK17]. sets [EDS14, OTK^{+}13]. setting [SFW^{+}16]. settings [GKRG^{+}19]. settle [TH11]. settled [ADvdM13]. Settlement [DHJ12, Par15, SB18, WRB^{+}15, vdMPRM12, AY10, BGP^{+}16, BBB^{+}14, BGIA19, BFEMLÁJ19, BJRH10, CSJP17, CC19, CMF^{+}19, CL16, DDP13b, DEM15, ES12a, EVHA^{+}18, FNC13, FWF^{+}10, GMSC14, GJ10, HSC^{+}15, HCG^{+}10, HR15, HB15b, INvdB^{+}13, JH11, JH14, KHH18, KBP^{+}10, LAM10, LBD^{+}15, LMGJ13, LS10b, LEB14b, MWHM18, MA13b, NLV^{+}15, OPSM13, OGCHS18, PMB^{+}10, PA13, PLFR^{+}11, PW19b, RCMB18, RWAP16, RBGT10, SHF12a, SCWV12, SMS15, SPW^{+}12, SHM10b, SAH11, SYK^{+}13, SKFY18, TPA^{+}13, VGL16, WSZS16, WBLRN18, WRFM16, vdMPM^{+}10, vdMPM12]. settlement-stage [WRFM16]. settler [vdMPM^{+}10]. Settling [FWPP12, BTTK10, BCWSR14, GELJ19, KCD13, PZA^{+}13, SCL^{+}19]. Settling-depth [FWPP12]. Seven [SBB^{+}14, BDD^{+}18, EPRGO13, KJW^{+}15, RWAP16]. Seven-year [SBB^{+}14]. sevengill [AB11, IDT^{+}18]. severe [SCP^{+}14, SSdW^{+}12, UCK^{+}13]. severity [EBH17, MW16a]. sewage [JJNP12]. Sex [BRT^{+}17, CGBGS18, DHS^{+}19, HLN^{+}19, RBP^{+}10, SBB^{+}12b, SCW^{+}18, ZdVA^{+}16, ADA^{+}19, BPJ^{+}15, BVQ12, EW16, FC19, FGH^{+}13, GMR13a, GMR13b, HBHF^{+}11, HBCK13, MC12b, MUOS13, PMB^{+}17, PB19, SKSH19, SY10, TMM^{+}19, THP^{+}16, TP19, VCG^{+}10, YTO^{+}11, YSM^{+}10]. Sex- [CGBGS18]. sex-based [SKSH19]. sex-biased [SY10]. sex-change [BVQ12]. sex-changing [EW16]. sex-dependent [THP^{+}16]. sex-influenced [ADA^{+}19]. sex-ratio [TMM^{+}19]. sex-related [YTO^{+}11]. sex-selective [PB19]. Sex-specific [BRT^{+}17, RBP^{+}10, SBB^{+}12b, HLN^{+}19, MUOS13, VCG^{+}10]. sexes [KCR^{+}10]. Sexual [BOC^{+}16, CWB^{+}15, CCM^{+}12, DFTZ^{+}15, GRCH^{+}18, KBL11, FRGI15, HTR19a, KCLLL13, PRFA11]. sexually [MTP^{+}10, QSvN^{+}11]. Seychelles [PSD^{+}19]. SGDC [KTH^{+}18]. Shad [SSS19, TRA^{+}14]. shading [GD15, KKA12]. shadow [MF10, SPH^{+}17]. shag [BANE^{+}13, LANE18, LMCD19]. shags [CCA^{+}17]. shallow [AS13, BCW12, BRBH19, BVS^{+}11, BBC14, CLGC16, CCA^{+}14, DJL^{+}12, Edm10, EL16, EL19, FT10, FKvdV12, GLF^{+}17, GMHS19, GP11b, GJSR18, HP11a, HPP^{+}10, IHP10, JMGG13, KMGH16, LKC^{+}17, LMd12c, LMd12b, MPB15a, MTM^{+}11, NHAS15, PSHG16, PS15b, QHG^{+}19, RWO13, RCGE12, RZS12, RMV13, RLS10, SSR^{+}10, SL15a, SOA^{+}17, TUK13, TYVR^{+}15, VB19a, VHV13, WADKT13, WFBM14]. shallow-water [AS13, FKvdV12]. Shank [SCHO11a]. shape [BSC^{+}13, BSG^{+}17, CS11a, FFLS16, GOC^{+}18, GBD13, HWSK19, HOL^{+}17, HD10, JWR^{+}12, LMBM18, MMC^{+}16, MBP^{+}12, PLMN10, RAA^{+}16, VM10, Vig15, VRAP^{+}14, WG11, WH14]. shaped [CDWC17]. shapes [FTT^{+}17, LDVP^{+}17, MBRO10, PSS19, SBB17, SHH10, WvLF18]. shaping [FH10, MHB14, TDSS15]. share [MDS14, RPR16]. Shared [LI18, DKF^{+}15, KS13, SDG^{+}12, TS15]. sharing [BEFC15, TKEV^{+}19]. Shark [BMPC15, MMCL19, WOWG10, AB11, AMSdL^{+}18, AFF^{+}15, ACP^{+}16, BMM17, BMJH12, BVAB^{+}17, BFM^{+}13, BKTD10, CJM^{+}13, CBL^{+}10a, CQR^{+}17, DNO17, GSCH14, GRN^{+}16, GRBH15, GGD^{+}11, HBM^{+}15, HS14b, IDT^{+}18, INK^{+}17, KHF^{+}11, KMGH16, LSH^{+}19b, LDP^{+}17, MdSW^{+}16, MHW^{+}12, MHSG18, MMC^{+}18, MMB^{+}19, PPB^{+}17, PW16, PBB^{+}15, RB11, RABS19, RLH^{+}19, SCB15, SHUS17, SBCT17, SKFH18, SMMFH16, SFMB10, SMF^{+}11, SCF^{+}17, YHTS12, THB^{+}12]. Sharks [KHS10a, TPVCV19, ARvdW^{+}16, BPP^{+}18, BPB^{+}12, BRLM17, BPC17, BHL^{+}19, BM10b, CEH18, CLH^{+}15, CWK^{+}10, CWW^{+}16a, CTSH12, CHST16, CBL^{+}10b, DOH^{+}17, DCHI12, FCFD15, FLK12, FSW10, GMHS19, GGF^{+}12b, HBI^{+}18, HS11b, HKSD14, HLS18, HIMH15, KWC^{+}17, KCS12, KHS^{+}11, LRHP18, MBB^{+}15, MVP^{+}16, NWP^{+}11, NHG10, NCK^{+}14, NCW^{+}19, PWNP^{+}11, PRP^{+}18, QHN^{+}10, RR18, RCR^{+}13, RPM^{+}13b, RLNLAF17, SMD^{+}13, SHB^{+}19, SMF^{+}12, TBMG17, TKEV^{+}19, TMP^{+}11, UCK^{+}13, VBR^{+}14, VWW^{+}16, WPS^{+}19, WHJ^{+}12]. Sharma [Ano12c]. sharp [PZA^{+}13, SNT^{+}18]. Shaw [GBH12]. shear [ML11, PMS13]. shearwater [CD19, CGBGS18, LBVH13, MGPC13, NBGS^{+}12, WPF^{+}16]. shearwaters [DGC12, HMO^{+}12, MC16, PXG^{+}10, PWA^{+}17, RHW12, TTT^{+}15, WYI^{+}16, YTO^{+}11]. Sheaves [CGK13a]. shedding [HGG^{+}10]. sheep [WTGC^{+}11]. sheepshead [MDAPH11]. Sheet [ANRT10]. Shelf [AB12, BL10, CBL^{+}10c, LN10, OFFQGG10, POS14, PFF^{+}17, PYBB11, SJS^{+}11, TWL15, WLM12, WSLM13, WRDF12, ÅCAC16, AIA18, ÁMLUN16, BRMT11, BBI16, BS11, CAR^{+}19, CPS^{+}14, DH18a, DRT13, DRD^{+}18, DJL^{+}16, DIB15, ERG^{+}11, FMH^{+}19, FL17, FSS^{+}18, GSD12, GSMK15, GFW^{+}14, HKL12, HGFA^{+}11, HEB^{+}11, HMS^{+}12, HBM^{+}12, KAD^{+}15, KFGSMPG14, LPKT^{+}19, LBR^{+}19, LGC10, MSS^{+}14b, NHSG11, OV10, PPML17, PS15a, PPSR15, STYT10, SS11a, SRER16, Sti13a, SOT15, TSCvS^{+}15, TMP^{+}13, VVB19, WJW^{+}18, WKRW^{+}12, CMM^{+}18, DTS14, HKB^{+}19a, STBP13, TGBM^{+}16]. shelf-basin [DRT13]. shelf-margin [LGC10]. Shelf-scale [WLM12, DH18a]. shell [BA13a, BGJ^{+}19, BO15, BIL^{+}10, BPHW16, CLG19a, CLL15, CMK12, CBD14, CSQRA11, GSO^{+}18, LMBM18, MKS18, MCR^{+}13, NWS^{+}14, NLLS15, PDB15, PUP16, RHF19, SCO^{+}12a, SES17, SHMB15, TKDM16, WMMC14, WG11]. Shellfish [TZF11, BDLW10, BDP10, LdU^{+}17, SFW^{+}16, TSF^{+}10, VTZ^{+}13]. shellfish-eating [BDP10]. shellfish-exploited [LdU^{+}17]. shells [CLDC^{+}13, CGS^{+}19, FBC^{+}15, GBMW12, KPP^{+}16]. shelter [LÁBFÁF^{+}10, NGSF13]. sheltered [BFNB16, ODCE11]. Shetland [GMSL^{+}14, JGL12, KBK^{+}10]. Shift [BEP^{+}12, CLD^{+}13, WCG^{+}14, AF14, BHE14, BEMD11, DGC12, EL16, ETHB17, GKBEF18, Hat13, JG13, KF17, KMSW10, LLH^{+}11, MBB^{+}15, MMA18, MGB^{+}19, ODK^{+}14, PKS11a, RDM18, SDKS19, SSHB16, SGTD14, TPM19, TCWC13]. Shifting [AJF^{+}12, LN10, NGR^{+}17, KWM^{+}17, SBE^{+}14]. Shifts [BKR^{+}15, FBET15, MPM^{+}14, AFLM12, BFG^{+}19, BHE14, BL17, BLC^{+}16, BPHW16, CPR^{+}16, CLS^{+}16, CBL^{+}10c, CWAS17, DPS16, DGA^{+}15, DWD^{+}18, DRWQ12, DAWS19, DABP10, FTWA17, GYRJ15, GFKC13, GRN14, Hv14, HRS^{+}16, JDG^{+}17, KTNO19, KSB11, KHS^{+}11, KNB^{+}13, LGS10, LDM10, LSD^{+}13, MSF18, MHHF13, MLP^{+}13, NKH^{+}10, NMCP17, OS18, OWW^{+}17, ORM^{+}13, OSD^{+}13, PBC^{+}18, PTS^{+}18, PDB15, PFF^{+}17, PRFA11, RMP^{+}19, Ric12, Rit17, SGP^{+}17, SPP17, SKPS^{+}16, VDM^{+}10, VM14, WL19, WW15, ZBC10, dG14, vdMPM^{+}10]. Shining [WGD^{+}19]. ship [DMR10, KCD13]. ships [VK11]. Shipworms [MdNW12]. shipwrecks [Sti13a]. shoalgrass [GP12, MH13]. Shoals [TBH10]. shocking [BKTD10]. shoot [RFRE19]. shoots [MKMZ11]. Shore [MWGH15, ANB13, BSC^{+}16, BMVBC10, BSP^{+}10, BHDC15, BPLT14, CBKM16, CMC10a, CDCZ16, CFMN16, FHG^{+}14, GHCC15, GVGH15, JMB^{+}18, KCH^{+}11, LGMB^{+}16, LFMC11, MHPvdH19, MSG11, MGL^{+}16, RVF14, RPHB15, TCA^{+}11, VANB15, WGD^{+}16]. Shore-level [MWGH15]. shorebird [BEMD11]. shorebirds [HS15a]. shoreline [BGO^{+}18, MFDW17, MCK^{+}13, NMG^{+}13]. shorelines [TT16]. shores [BAM^{+}10, CGOC11, CGKW10, CMO^{+}10, CAC11, LLBT18, LM12, MLB19, ODCE11, VBT14, WBLRN18]. Short [BCK15, DDP19, MPA^{+}10, MCW^{+}17, QHN^{+}10, RP17, SW12, SLC11, SAD^{+}16, dlSGS17, vDGK^{+}12, BPZ^{+}14, CL15b, ENO^{+}11, FHG^{+}14, HBG^{+}17, ITT14, LH14, MPG^{+}10, MHR^{+}19b, NCC^{+}14, OAB^{+}19, PFB^{+}13, PFKB18, PBB^{+}15, PRA^{+}19, RBBG12, TFB^{+}17, VMC^{+}19]. Short- [vDGK^{+}12, HBG^{+}17, PFB^{+}13]. short-beaked [BPZ^{+}14, PBB^{+}15]. short-finned [OAB^{+}19, TFB^{+}17]. Short-lived [SLC11, RBBG12]. short-ranging [MPG^{+}10, MHR^{+}19b]. short-tail [PFKB18]. short-tailed [VMC^{+}19]. Short-term [BCK15, DDP19, MPA^{+}10, QHN^{+}10, SW12, SAD^{+}16, dlSGS17, CL15b, ENO^{+}11, FHG^{+}14, ITT14, LH14, NCC^{+}14, PRA^{+}19]. shortages [PHI^{+}12]. shortcomings [MS11]. shortfin [NCW^{+}19, PBB^{+}15, VWW^{+}16]. shortfinned [NSC^{+}19]. shortnose [FPC^{+}12]. Shortspine [DT11, DEGF^{+}10]. Should [DDM^{+}16, MCPC19]. show [KPM^{+}11, NWH^{+}12, NTS^{+}18, NSY^{+}18, PRM17, SRCE^{+}10, SDB12, TRMC11, ZHTA13, vdMPM12]. showcase [Ano18b]. showcases [Ano17c]. shows [HS16]. shrimp [ASM^{+}12, BM10a, BBS^{+}17, BD11b, CBC^{+}10, Cra12, CZH^{+}13, DPH14, FHM^{+}11, GVR^{+}14, GZYS11, HCBV^{+}11, HT11a, HT11b, HD10, JBS12, LA14, Lyo14, MH11b, MMM^{+}19, MPB15b, NKH^{+}10, OT14, PPS^{+}13, RFMT12, Ric12, SBLM14, TBH^{+}18, TGRH17, TBMdV12, UTS^{+}18, VPW^{+}12, WS12, YWiIK15]. shrimps [FZI^{+}16, MASMH19]. shrinkage [THP^{+}16]. shy [AGHC10]. Si [DTVH10]. sibling [FCV^{+}16, RF14, TDB^{+}14]. siblings [DXF^{+}18]. sicklefin [CBL^{+}10b]. Sicydiinae [LLD^{+}12]. Sicyopterus [LLD^{+}12]. side [AvdHH^{+}17, LAM10]. sided [SWH^{+}11]. siderophore [SP15]. Siganus [BDB^{+}13]. sightings [RPM^{+}13b, WHJ^{+}12]. signal [Hat13]. Signalling [BMSS14]. signals [LMD12a, MB11, STM^{+}14, VWH18, vdHTT^{+}15]. signature [AEM^{+}18, CWD11, KKMA^{+}10, MVP^{+}16, RCR^{+}13]. signatures [BSC^{+}15, BDS^{+}18b, CKM^{+}16, CLDC^{+}13, DP14a, GOO14, HVT12, JS12b, KPP^{+}16, LKL^{+}19, MGO^{+}12, MMWG13a, MMWG13b, PIPD10, PHWC13, RST^{+}10, RSJ14, RP12, RET^{+}19, RSTM14, RGDNGS11, SMF^{+}17, SMMFH16, SWW11, TFM^{+}18, WRFM16, YAFV12, vdBUPP^{+}11]. Significance [Hal15, PR11, SM10b]. Significant [AFC^{+}16, HKL12, JSLG17, SDB12, WCRCR17]. significantly [CS13]. signposts [SMD^{+}16]. silica [MLP^{+}14]. silicate [BvHC^{+}15]. silicified [LG18]. Silicon [LALCM18, CMYL11, DTVH10, FLE^{+}11, GBD13]. silky [FCFD15, HIMH15]. silver [BPSCD17, HVD^{+}16, LET^{+}12]. silverfish [BLJP14, CAR^{+}19, DHK^{+}17]. silverside [BBC12b, GM13a]. silversides [PB19]. Similan [SPJ^{+}12]. Similar [MM16, vDBH^{+}15, BPT15, CGH^{+}16, EBOM11, FT10, GBE^{+}17, JGCRS12, KPM^{+}11, SJHH^{+}17, TMP^{+}11]. similarities [NTS^{+}18]. similarity [ALS18, CCA^{+}17, FSF^{+}17, MVL^{+}16, TBS^{+}14]. similis [DHH10, JSLG17, ZTNS13]. Simple [BSB12, CML^{+}12, EM11, LRK^{+}15, SBKA14, SHG^{+}15]. simplex [GGMGR12, KBK^{+}10]. Simplifying [TTC11]. Simulated [KHN15, AJF^{+}12, BHDC15, CE14, ENO^{+}11, MAB12b, NDAP19, OSL18, SSL^{+}14, SGT^{+}14, TWW^{+}15, WNF^{+}11]. Simulating [ARB^{+}12, LWD13, LRQ^{+}12]. Simulation [AMR^{+}10, CLL13, CCDP15, GL12, LGL11, NRF11, STK^{+}13, TGRH17]. Simulations [SGGL15, FRP^{+}17, GGF^{+}12a, JSMO14, ZSSK11]. Simultaneous [AFS19, DKF^{+}15, JMO^{+}19, MHS14, BGB^{+}17, EKOÅ13]. sinensis [SIK16]. single [DBS^{+}13, JLL^{+}12, KPM^{+}11, KRC17, PBM^{+}11, RST15]. single-osculum [KRC17]. single-species [JLL^{+}12]. singularis [FMC^{+}11]. sinicus [OHST19, WWB14]. sink [BFH^{+}13a, CGD10, CMYL11, HTB^{+}17, NDD^{+}13]. Sinking [JPMG10, BBG^{+}16, LCTD^{+}15]. Sinonovacula [PZL^{+}17]. sinuosa [HCK12, OFM^{+}18]. Siphonaria [LMBM18]. Siphonophora [BLHL15]. sister [BEFC15]. Site [BMPH14, Dan11, GGNL^{+}17, KHW^{+}12, LBWB^{+}13, SAL^{+}12, THB^{+}12, TS11b, ACP^{+}16, ADRA19, BC11, BMSB13, BDB^{+}13, CL15b, CCS^{+}18, CNJ^{+}18, GJ16, HVJ^{+}18, HGKB14, KCS12, MBB^{+}15, MWC18, MHAP18, MDS^{+}13, NTS^{+}18, NLGE18, OALK14, PPM^{+}15, PFKB18, RHB13, SGM^{+}14, SCW^{+}18, TWL15, TDM16, TNB^{+}15, TR13, TMS14, WTdB16, WGD10, ZBB^{+}19, TWL15]. Site-dependent [GGNL^{+}17]. site-fidelity [MBB^{+}15]. Site-specific [KHW^{+}12, SAL^{+}12, BMSB13, HGKB14, ZBB^{+}19]. sites [BMMS10, BHD^{+}13, FM14, HPB^{+}18, KH10, SBFP10, SNW^{+}18, TPVCV19]. Sitka [ADRA19, BJS^{+}16]. situ [ALANGT10, AHS^{+}13, AHSG16, BFF^{+}10, BWH11, BSC^{+}16, BPM^{+}18b, CSRF16, GSF18, GMK^{+}19, GCDH19, GHT15b, HJN^{+}10, HSZR10, HMBP18, HB15b, JBCGC19, KBA13, KH16, KPS14, LAM11, LHL11, LPI^{+}14, LSH^{+}15, MGM^{+}18, MAT^{+}11, Mat10b, MH11a, OFM^{+}18, PNA^{+}16, RWR^{+}19, RS14, ROE17, RRS15, RS16, RGB^{+}11, RGP11, SW12, SMBC18, SDVL16, SAH11, TS10, THM^{+}13, WKE^{+}17, WNNH10, WEK14]. six [DHJ12, DCHI12, PB10, RBGT10]. Size [ÁMLUN16, BAL14b, DLL^{+}18b, FGRM12, FHS^{+}13, GKW10, GGFB15, Hoe10, HM10, JS12a, LKLC11, PBM^{+}11, PB10, RSE18, SA14, TLFS14, WBG10, WB12, WWW^{+}13, AB17, AZVB13, AFOVRB14, AL12b, AP14, ARvdW^{+}16, ADvdM13, ASEB11, BC16, BAL14a, BMM17, BTS^{+}14, BNAN12, Bea11b, BHAC17, Bet13, BPJ^{+}15, BVAB^{+}17, BPW^{+}18, BCK15, BMT17, BJRH10, CCE^{+}14, CPPM12, CLB19, CCM^{+}15, Cha13, CIF11, CVC^{+}18, CMW^{+}14, CDCZ16, Col10b, CS11b, DGC18, DMMW12, DA10, DQM^{+}19, DDB^{+}14, DCC^{+}10, EL19, ESD16, ECZD14, FJC10, FWPP12, FFL10, FMJA^{+}18, FEE^{+}11, FTB^{+}17, FSM^{+}14, FF10, FFS^{+}13, GSR10, GvSGCK14, GGPDS19, HNK^{+}18, HPMB11, HW13, HSS^{+}10, HORRLS^{+}14, HYL^{+}13, HT11a, HOE^{+}12, JW10, JDQS11, JBL^{+}16, JJJ16, KHH18, KWM12, Kiø16, KLN^{+}13, KP11, LRB^{+}18, LSRRCM13, LANER15, LRS17, MC12a, MHB14]. size [MPA^{+}10, MRE14, MRSD15, MFI11, MMRDS18, MSS^{+}14a, MD11, MOS10, MWGH15, MER14, MKII19, NKH^{+}10, NWS^{+}14, NMS^{+}18, NM18, NOB^{+}19, NJP12, OOK11, OCDD16, POR^{+}15, PPM^{+}15, PWM18, PPH11, PM16, PSdA^{+}17, QGM^{+}14, RLG14, RUPT^{+}11, RABS19, RH12, RST15, RGDS12, SGMG^{+}14, SRLV10, SY10, SS12a, SKGS11, SLT^{+}19, SCSC16, SCK^{+}17, SK11, SBA12, SMHB14, STSS^{+}16, SPP17, SSM^{+}16b, SJ10, SF14, SH10, TFP14, TL18, TG12b, TDSS15, WKX^{+}19, War14, WH14, WMMC14, WLB^{+}15, Wir12, Wir14, WPK10, YPWP18, vDJF15]. size- [SY10]. size-at-age [BMT17]. size-at-terminal [DMMW12]. Size-based [BAL14b, GSR10, SSM^{+}16b, SJ10]. Size-dependent [ÁMLUN16, DLL^{+}18b, LKLC11, RSE18, WBG10, ADvdM13, HSS^{+}10, PPH11, SRLV10, SK11, TFP14]. size-fractionated [BTS^{+}14, SGMG^{+}14]. size-fractioned [WKX^{+}19]. size-frequency [AP14]. size-regulated [CLB19]. size-scaling [HORRLS^{+}14, SMHB14]. size-selective [AB17, FF10]. Size-selectivity [HM10]. Size-specific [SA14, TLFS14, CS11b]. size-structure [CIF11]. Size-structured [GKW10, BAL14a]. size-trends [RGDS12]. sized [BS16, TDM16]. sizes [Bea11a, BPB^{+}12, SHC^{+}10, WLB^{+}15]. Sizing [HKSD14, SMHB14]. Sizing-up [SMHB14]. Skagerrak [ETO^{+}10, HFN14, KRDV18, LOES11, NGP^{+}15, SBH^{+}18]. Skallingen [WKE^{+}17]. Skate [RHS^{+}19, FCV^{+}16, Hof10]. skates [FGTS16, SST19]. Skeletal [AHS^{+}13, LLv13, WNP^{+}14, FKT15, LBH18, LKN^{+}14, SCL^{+}16, TTF^{+}11]. skeletochronology [AGP^{+}13]. skeletogenesis [OTS^{+}10]. skeleton [PHWC13, TGM^{+}17]. Skeletonema [CR16, RSS^{+}16]. skeletons [LKB^{+}14]. skin [VRSS18]. skipjack [RTCF^{+}17]. Skipped [SKN^{+}15]. Skov [TW10]. skua [SHT^{+}18, vBMH^{+}17]. skuas [CMC^{+}16, KPM^{+}11, KHK^{+}17, WTC^{+}15]. skull [DCC^{+}10]. Slattery [PMEZ15a]. slender [CAC^{+}12]. slender-billed [CAC^{+}12]. slicks [WLL^{+}14]. slipper [CMP^{+}13]. slope [BCM^{+}14, BBI16, DKRS12, LRB^{+}12a, SRH^{+}15, Sti13b, YD11]. slopes [FPC^{+}13, KPKA13]. slow [HCL^{+}11b, MH12b, OKJMC15, PBF10, WEA^{+}17]. slow-growing [PBF10]. slug [CRR^{+}19, CNSW16, TWM^{+}15, YTM^{+}15]. slugs [DFB^{+}17]. Small [GMB^{+}16, JH11, LGR^{+}10, MSFM12, MRL^{+}17, SKGS11, SSB^{+}16, AHA^{+}15, AMS^{+}13, BI15, BRK^{+}12, BMF^{+}15, BLC^{+}16, CKM^{+}16, CE14, CMV^{+}11, CCJR14, CASG^{+}19, CBGDS12, CGM^{+}19, CNÁCP12, EMTB^{+}12, GS11, GCCM18, IMD11, JBS11, KHH18, KRC17, LKC^{+}15, LLMA^{+}10, LB15b, MGP11, MJO^{+}13, MJT^{+}12, OCSW16, OMW^{+}16, PBM^{+}11, PSdA^{+}17, PDCH15, Rai12, RPS12, RGB^{+}10, RBD^{+}14, SYM^{+}16, SH14, TPBM13, TBDM11, VGL16, WS10, YAFV12, ZMGS17]. small-particle [RPS12]. Small-scale [GMB^{+}16, LGR^{+}10, MSFM12, SSB^{+}16, AMS^{+}13, BRK^{+}12, CKM^{+}16, CASG^{+}19, GCCM18, LLMA^{+}10, MGP11, MJT^{+}12, OMW^{+}16, SYM^{+}16, TPBM13, YAFV12]. smaller [BOC^{+}16]. smelt [BDT^{+}11]. smolt [MWC^{+}12]. smolts [ABK^{+}14, BMJ16, HGG^{+}18, HHP^{+}17]. smooth [MKMZ11]. smoothing [CKR^{+}11]. SMS [CKV12]. Snail [GTF10, ZH13, CHDH^{+}19, CSQRA11, GB16, HA18, HWTH12, MM12, SA14]. snails [SES17, SWWK13]. snakes [BL13, USR^{+}16]. snapper [BN16, CMR12, CCS^{+}18, FA11, KCLP16, LMC14, LBCS^{+}15, MH19, MJHS11, MJHS13, TS11b, WGS16]. snappers [DSS10, DSC13]. snapping [MH11b, MMM^{+}19]. snapshot [WVD^{+}17]. snook [BMPH14, YYW14].

Snow [AFMR^{+}13, BAW11, DMMW12, KLN^{+}13, KLNJ15, LCTD^{+}15, MMDM12, MJT^{+}12, PZA^{+}13, SDMR19]. snow-covered [SDMR19]. Social [CWMP13, Miu12, AB13, BVQ12, BCGG19]. sociality [GJ16]. socio [WF18]. socio-economic [WF18]. socioeconomic [FEA^{+}15]. sockeye [DHM^{+}14, FTB^{+}17, JS12b, ODL^{+}13, SQS^{+}13, WHD^{+}14]. Socotra [MCW^{+}17]. soft [BMC^{+}15, Bar14, BCvH^{+}15, CLG19a, CTvdM^{+}08, CHG^{+}18, DvV^{+}16, GPV17, GGNL^{+}17, GSO^{+}18, HKL12, KPB11, LEB14a, MBvR12, MMWG13b, NNT^{+}10, RMC^{+}14, SGJ^{+}18, TBKG10, TPCN12, VNN10]. soft-bottom [DvV^{+}16, HKL12, KPB11, LEB14a, MBvR12, RMC^{+}14]. soft-seafloor [SGJ^{+}18]. soft-sediment [BMC^{+}15, Bar14, GGNL^{+}17, NNT^{+}10, VNN10]. soft-shell [CLG19a, GSO^{+}18]. software [TAEB^{+}11]. Soil [BPB18, CAKD14, TB10, WWO^{+}15]. solar [DFB^{+}17, DGQ10, FLSG^{+}15, GHHH12, GHCC15, GJ10, JDX^{+}17, MWC^{+}12, PCNB16, ROLA15]. solar-powered [DFB^{+}17]. sole [JNYT13, LBD^{+}15]. Solea [CGG^{+}10b, TRSV^{+}12]. solidissima [TPM19]. solitary [SIHH10, SF14]. soluble [KCP^{+}11]. solute [KHD^{+}11, RF14]. solutions [Fry13b, GJ13, NSS10]. solving [Fry13a]. Somateria [GDH^{+}19]. Somatic [BP12a, KCL^{+}11, RBEGG13, BBC19, CC19, DVS13, SDVL16, SST^{+}18]. Some [AMLS15]. Somniosus [FLK12]. sonar [MIV^{+}10, SWK^{+}16]. song [LFPA^{+}11]. Sonneratia [FWW^{+}16]. sooty [HMO^{+}12, TTT^{+}15]. Soritinae [MU13]. sorting [HHG10]. SOTS [ETD^{+}18]. Sound [AMM^{+}12, BJS^{+}16, GKR^{+}15, HLS17, HOA14, KHD19, KFBM19, LS15, MGM^{+}18, MIA^{+}11, MBB^{+}12b, MMM^{+}19, Ran18, RSA^{+}12, SNVH19, TLG14, AHR^{+}18, CSV^{+}15, DGP^{+}19, LEB14b, NKA16, PCH^{+}14, PCLB18, RST^{+}10, RTMJ11, RSJ14, RNSA15, WLM12, WSLM13, BS11, GSF^{+}13, KKY15, LCJ^{+}12, MBG^{+}15, MOH^{+}10, SHT^{+}13, SFKQ15, SLR^{+}14]. sounds [FMH^{+}15, FMG^{+}18]. Soundscape [LEB14b, AHR^{+}18, HPB^{+}18, LEB^{+}19, MMM^{+}19, REBL16]. Soundscapes [NSH^{+}15, FF16, KMPS15, LEB14a, SPD^{+}14]. Sour [ATP^{+}12]. source [BFH^{+}13a, BVC^{+}16, CGD10, NDD^{+}13, RTMJ11, RHMR18, SMS15, SBS^{+}13, SE11a, SN17]. source-sink [BFH^{+}13a, CGD10, NDD^{+}13]. sourced [CRS^{+}18]. Sources [GWME16, KRC^{+}11, WWG12, AMS^{+}13, BCM^{+}14, BVP11, HPMZ11, HRT^{+}14, HLB^{+}16, JBS11, KF13b, KBP^{+}10, LCL^{+}17, MJ10, MWD10, NUO^{+}19, PLG^{+}19, SO16, SJL^{+}15, TTS17, TBH^{+}17, VRSS18, Vig15, YSC^{+}15]. Sousa [MHAP18]. South [BW13, BP12b, CP10, Dan11, DFS13, DPC13, KPM^{+}11, MMB^{+}19, OCS14, WHPT12, WFBM14, ZBP18, BTLF17, DAL^{+}10, HVD^{+}16, MMRC16, OCMT13, RWMP12, SSB^{+}18, SDCC12, WTC^{+}15, ARvdW^{+}16, ARA^{+}10, ASF13, BTO^{+}18, BSRDR10, BDH^{+}17, CC19, DKV^{+}16, DAC^{+}12, DLL^{+}18b, DCC^{+}10, DFTZ^{+}15, DBS^{+}19, EFOM13, GMSL^{+}14, GH11, GHT^{+}15a, HMO^{+}12, INK^{+}17, KHG^{+}19, KBK^{+}10, LABM17, LFM^{+}16, LLC^{+}14, LCR^{+}16, LBCS^{+}15, LFRF13, OCSW16, OCE^{+}12, PSHG16, PBG^{+}10, PJMP10, PSS19, PRK^{+}12, RWM^{+}10, RCGE12, RNFG10, RVPG17, RMNP17, SBT12, SSB^{+}11, SLS16, TMS^{+}16, VCG^{+}15, VCG^{+}10, WWAS18, WPRS18, WLF^{+}17, XWE^{+}14, YHTC19]. south-eastern [MMRC16, SDCC12]. south-western [SSB^{+}18]. southeast [BSH^{+}14, EBH17, LAM11, ME11, PPB^{+}17, SGC17, WYI^{+}16, DPH13, KSK^{+}12, MCW11, SB14]. southeastern [AGGAO^{+}18, Ano18a, BLE^{+}13, BBMH13, BPZ^{+}14, CPS^{+}14, DKB^{+}12, FKZTT18, HRK^{+}18, HMF^{+}18, KTP^{+}18, LRHP18, MZ10, MVVGAV10, MASW^{+}11, MBKM17, MMM^{+}19, PSA17, PBC^{+}12a, SDAM11, SDAH12, SE11b, SKPS^{+}16, TCSA17, TCJ^{+}19, WK18]. Southern [BST19b, ETD^{+}18, GVGP^{+}14, HER^{+}19, MKPS13, MVS^{+}10, RSW^{+}16, RGL^{+}16, WNNN12, vdSRT^{+}18, AIA18, ÁMLUN16, AFH11, ADB10, BCG10, BG19, BBB^{+}14, BOC^{+}16, BSB^{+}11a, BPZ^{+}14, BTLF17, BPG^{+}13, BPF15, BK14, CAR^{+}19, CLM^{+}13, CPA^{+}11, CVC^{+}18, CMHH17, CQR^{+}17, DVSM10, DED^{+}15, DSJNH18, DRDV15, ETAE13, FMH^{+}15, FGW^{+}15, FWL15, GMP^{+}11, GB16, GW15, GGK^{+}16, HCL^{+}11a, HRT12, IS16, JJP^{+}13, JMK12, JLPG17, KCH^{+}11, KMW17, KNB^{+}13, LPKT^{+}19, LCB^{+}16, LRO^{+}17, LBC^{+}19, MMH^{+}17, MBSG11, MRLD19, MTP^{+}10, MB18, MGB^{+}14, MHN^{+}12, NCC^{+}14, NFC^{+}11, NSB^{+}19, NCK^{+}14, OHCG14, ORE^{+}17, OVF11, PTS^{+}18, PSAK12, PBW^{+}11, PPEH^{+}14, PMM^{+}15, PWM^{+}19, RFBH^{+}11, dGCF^{+}17, RWO13, RBEGG13, RDGC17, RUS^{+}13, RSSvV19, SAZHAAHS14, ST10a, SOJ14, SAH11, SLRS13, TL18, TAB^{+}14, THW^{+}16, TLAT13, VGGD^{+}15, VRSS18, WKvW10]. southern [WWAS18, WPNK12, ZPPM17, ZCC^{+}12, ALE^{+}13b, ALSV16, ABB^{+}14, BRJ^{+}12, BMTC^{+}15, BTLF17, CWK^{+}10, CLGCA^{+}19, CMB^{+}14, DSH14, FLE^{+}11, GGLT15, GSJS^{+}19, GCCM18, GPC^{+}15, HYL^{+}13, KSK^{+}12, KBB10b, LHS^{+}14, MBR^{+}18, MMB^{+}12b, MTH15, PPC^{+}19, RVPG17, RDM18, SCW^{+}15a, SCST17, SMK^{+}18, SRB^{+}17, TBP11, TBH^{+}17, WMS^{+}10, WLB^{+}15, Wil14, WJS^{+}14]. Southwest [BPN11, ALV12, CTZ^{+}16, HCF^{+}16, KHN15, RTJ^{+}13]. Southwestern [BVM^{+}19, BDM^{+}13, SBV^{+}15, CAG13, GCR^{+}17, MFG16, PSG^{+}13, RNFG10, RVPG17, SML^{+}14]. sp [BMR17, BBW^{+}19, CSABH^{+}13, CL17, JWR^{+}10, KLSJ10, LKL^{+}19, LHW^{+}19, OD15, OS12, PPD^{+}14, SIHH10, TLK^{+}12, ZPPM17]. Space [ZBK^{+}15, BRT^{+}17, CGOC11, DQM^{+}19, DZDB10, EHR12, HS14b, HMW^{+}12, JMS^{+}15, KDO^{+}14, LHDH17, LRK^{+}15, MHS15, OPSM13, Pro17, SHUS17, SD17b, TRSB^{+}18, UCK^{+}13]. Spain [BC11, BVP11, GSÁCM^{+}16, PDGR^{+}13, THMF^{+}15, THRBL^{+}16]. Spanish [RFH16, WVM^{+}14]. spanning [Ano18a]. sparid [RA10]. Sparidae [AHA^{+}15, FCR12, MJHS13, NVA^{+}19]. sparse [CGK13a, JFK12]. Spartina [BPB18, CSN12, FCR^{+}15, HL14, KM10b, MA13a, MKMZ11, PCA^{+}15, SWWK13, WWO^{+}15, ZH13]. Sparus [MVCD14, MMB^{+}12a]. spat [DB14, vBJV^{+}15]. Spatial [AMSdL^{+}18, AVB13, AMP^{+}11, AB13, AED^{+}12, ARA^{+}10, BS14a, Bar13, BB14a, BCM^{+}14, BN16, BWH^{+}19, BS14b, BRA^{+}10, BBR^{+}14, BCWSR14, BFH13b, BJRH10, CAR^{+}19, CWS13, CP10, CPF^{+}16, CTB13, CCDP15, CS11b, DZH^{+}11, DH18b, DJL^{+}12, EW16, FG16b, FSDS10, FHE^{+}15, GLW^{+}12, GS16, HKI^{+}17, HER^{+}19, HSH12, HHS^{+}12a, HBB^{+}18, HFR^{+}13, HMF^{+}18, JH14, JBL^{+}16, JJNP12, KBG^{+}13, KAJP^{+}15, KM17, LKC^{+}15, LHJH^{+}10, LGS10, LDS^{+}15, LOBW^{+}11, LSV10b, MCS^{+}15, MRGG^{+}13, Met11, MPB15b, MMB^{+}12b, MA13b, NKS^{+}15, NMG^{+}13, ODP^{+}12, OLM^{+}12, RS13, RDF^{+}17, SSBC11, SCD^{+}17, SRLV10, SBB17, ST13, SBFP10, SMD^{+}13, SS10b, SRER16, SE11b, SBH^{+}18, SMF^{+}11, SWW11, SNNS^{+}12, TPB^{+}16, TKIO10, TNM^{+}11b, WDNM11, WGD10, XS19, YBB^{+}12, AN11a, AY10, AGIW16, ADB10, BBF^{+}19, Bar14, BVG13, BFD^{+}16, BWS11, BS16, BBMH13]. spatial [BSB13, BAVSP11, BS19a, BCSR10, BPB^{+}12, BDH^{+}17, BB14c, BPB11, CKM^{+}16, CDWC17, CLDC^{+}13, CHDH^{+}19, CCHG10, CTMU10, CAC11, CFSM^{+}13, CM11, CMB^{+}14, CCKB^{+}15, CH12, CSC13b, Cra12, DXF^{+}18, DMd14, DWK13, DRD^{+}18, DEWK15, DRH15, DHG11, DBS^{+}19, EES15, FTT^{+}17, Gal14a, GVW^{+}12, GLB^{+}10, GSU12, GDA^{+}18, GCCCM16, GCCM18, GHNF10, HEMJ18, HHLT10, HW12a, HLS15a, HCBV^{+}11, HLS18, HL14, JLM^{+}12, JMS^{+}15, KHBL11, KWW^{+}14, KWDD10, KBB^{+}10a, LFU^{+}19, LFJ^{+}14, LBP^{+}16, Lec18, LRB^{+}12a, LR18, LZF^{+}14, LMBM18, LMG12, LTN^{+}11, LS16, LMTD16, LKB^{+}14, MSFM12, MLR17, MH10, MHVV13, MS13, MJV^{+}11, MBF16, MFG16, MCRB11, MSK^{+}19, NSH^{+}15, NDG15, NSK^{+}14, ORR^{+}14, PDM12, PSHG16, PTP^{+}19, PvdMS^{+}15, PFR11, PMH^{+}13, PLF^{+}12a, RKD18, RDK14, RVF14, RDC^{+}12, RCU^{+}11, RNSA15, SRK^{+}18, SRG14, SHR^{+}11, SGCF^{+}12]. spatial [SFW^{+}14, SMSB19, SRK13, Sma13, SAMW14, SJB^{+}15, SBKA14, SHB^{+}19, SSFB11, SH14, SKPS^{+}16, SJL13, SNT^{+}16, TBP11, TMBP13, TRGF18, TPM19, TW10, VANB15, VGL16, WLBB^{+}13, WLP^{+}11, WTD^{+}19, WCG^{+}14, WBHL10, WWZ^{+}19, WGGM11, YSK^{+}10, ZAC11, dML11]. spatial-recruitment [MLR17]. Spatially [WMPM13, APC^{+}14b, GDA^{+}18, KS13, WHvdK11, WDM^{+}18]. Spatio [BDPL19, MDLM17, MVE^{+}10, SV13, STYT10, SKD^{+}17b, SRP^{+}19, TKW^{+}14, Tho15, WRDF12, YTM^{+}12, BBW^{+}17, CBR^{+}18, CCS^{+}18, FGH^{+}13, HHB^{+}14, IME12, MTS^{+}17, MPRPM^{+}18, PNI14, PCP^{+}15, PPTB12, PP13, PHHL^{+}11, PHG^{+}14, SDG^{+}12, SSB10, ST08, WHJ^{+}12]. Spatio-temporal [BDPL19, MDLM17, MVE^{+}10, SV13, STYT10, SKD^{+}17b, SRP^{+}19, TKW^{+}14, Tho15, WRDF12, YTM^{+}12, BBW^{+}17, CBR^{+}18, CCS^{+}18, FGH^{+}13, HHB^{+}14, IME12, MTS^{+}17, MPRPM^{+}18, PNI14, PCP^{+}15, PPTB12, PP13, PHHL^{+}11, PHG^{+}14, SDG^{+}12, SSB10, ST08, WHJ^{+}12]. Spatiotemporal [AMAJ^{+}17, LGB17, LKL^{+}17, LET^{+}12, YYW14, BHE^{+}17, DFD^{+}19, DCL^{+}17, GCL^{+}16, LBJI14, MN18, MWB13, MCT16, NMCP17, SLE^{+}13, YHTC19]. spawing [TKB^{+}13]. spawn [FPR18, LET^{+}12]. spawned [FWF^{+}10, MP17]. spawned-out [MP17]. spawner [DE17, HFH10, MS18a, TSCvS^{+}15].

Spawning [GTM14, PGG12, WFP15, AGGAO^{+}18, ACRLB12, BCW19, BJCB^{+}14, BN16, BDB^{+}13, BFN^{+}19, CRAP^{+}12, CCG^{+}10, CKK^{+}10, Coo11, CD16b, Dan11, DGDJ14, DDP13b, DFAH11, DHB18, FDVL18, FM14, GN11, GFC^{+}13, HCL10, HRT12, HFR^{+}13, JBM^{+}13, KTNO19, KNY^{+}15, KH10, KCLP16, LFJ^{+}14, LSJ^{+}15, LJM^{+}12, LBJK14, LBWB^{+}13, MCW^{+}16, MLCC19, MBvO^{+}18, NKH^{+}10, NRF11, NMS^{+}18, NvOH13, OVF11, OSD^{+}13, PSHG16, PVYS16, RRRF^{+}17, RCAB^{+}12, RTÁB^{+}14, RTM11, RvOP12, RSA^{+}12, RNSA15, SØK^{+}15, SMD^{+}16, SBFP10, SHM^{+}10a, SHH10, SKN^{+}15, SSS^{+}10, SJB^{+}15, SAH11, TWY^{+}19, TWL15, TPBM13, TLJ^{+}11, THW^{+}16, VHV13, WLBB^{+}13, WHI^{+}18, WM10, WHD^{+}14, YKT^{+}14, YOST11, YYW14, ZBB^{+}19, ZVL^{+}16, ZVR18]. spawns [KRW18]. spear [JHNL12]. spearfishing [BBC14]. Spechiospongia [WRGP12]. Specht [RL18]. specialisation [Ano17c, BCV^{+}18, CBLM^{+}17, GJ16, JCRC10, JGJ15, PW19a, SGT12, WJ12]. specialist [ACAD12, Cha13, CRGM18, HBK^{+}12, OKF^{+}18, PTP^{+}15]. specialists [THB^{+}12]. specialization [AB17, CEC^{+}18, JDD^{+}13, KM14, NGR^{+}17, PLGSD17, RDGC17]. specialized [MdNW12]. speciation [HP11b]. specie [BFG^{+}19, CLM^{+}13, JMK12]. Species [BMM17, BT16, CRRD^{+}13, CSRF16, DPY^{+}16, FSR17, GOO14, HTR19a, JSLG17, KO12, KCW^{+}17, ML11, MDS14, NBM12, PB18, PHTS11, RCK^{+}11, TYG12, VCQM^{+}16, WKvW10, WKRW^{+}12, AMSdL^{+}18, ABN14, ANB13, AGGAO^{+}18, Alb15, APBC^{+}19, ACS12, ATP14, ASCHM19, BA13a, BBM^{+}14, BI15, BAA12, BC16, BAL14b, BB14a, BFG^{+}19, BLIM11, BCdV13, BRBH19, BSB^{+}11a, BJH13, BAM^{+}10, BD11a, BN16, BSP^{+}10, BADQ^{+}15, BPSG11, BEFC15, Bol17, BDG^{+}16, BB14b, BRB17, Bri12, BC14, BMA^{+}19, BCB11, CAC^{+}12, CPF^{+}13, CBKM16, CE11, CD19, CWD11, CMC10a, CPF^{+}14, CLW^{+}11, CSM^{+}19, CY13, CCNB15, CLGC16, CFPC^{+}13, CALB10, CKR^{+}11, CMLC12, CLSS17, CGBGS18, CNÁCP12, CSC13b, CBYH13, DWLH13, DBE^{+}12, DP14a, DDDR10, DVV^{+}18, DGP^{+}19, DDB^{+}14, DVB^{+}16, DDP13b, DCHI12, DDC^{+}15, DC17, EVB^{+}16, EPRGO13]. species [EBB10, FUOG^{+}16, FTM^{+}13, FPR18, FPBL16, FSY^{+}10a, FB14, GGF^{+}12a, GSCH14, GMG^{+}16, GSP^{+}15, GSD10, GLR^{+}14, GSS13, GFC^{+}14, GRBH15, GB12, GGK^{+}16, GMK^{+}19, Gre10, GNC^{+}10, GDPR18, GHNF10, GMES18, HGL15, HFM12, HPL10, HBM^{+}15, HFM^{+}15, HHG10, Hv14, HC13, HSDB13, HSCA17, HYC^{+}14, HOA14, HB15a, HGKB14, HL14, HSWS16, HPLS13, IKS^{+}12, JBT^{+}16, JLL^{+}12, JSCM15, JL17, KKMA^{+}10, KMPS15, KHBL11, KNIN17, KM10a, KWS^{+}13, KHG^{+}19, KJW^{+}15, KKS^{+}12b, KTH^{+}18, KH10, KH16, LBM^{+}18, LCGS^{+}19, LI18, LN11, LGBSC11, LHT^{+}16, LRB^{+}12b, LHC^{+}15, LPSK17, LBJK14, LDAE12, LERS13, LLD^{+}12, LBG17, LMTD16, LN10, LRS17, Lyo14, MBB^{+}18, MKK13, MVL^{+}16, MPKO11, MJM^{+}14, MV12, MAD18, MRE14, MUHC11, MCUO12, MTS^{+}17, MBW10, MHCRR18, MJ10, MHW^{+}12, MWD10, MC12b, MBCB^{+}19, MF13, Miu12, MPB15b]. species [MCL^{+}17, MD16a, MRL^{+}10, MBF16, MOH^{+}10, MWMD17, NAG^{+}16, NHSG11, NPLT10, NSI^{+}15, OS18, OMN17, ORM^{+}13, OFPR^{+}11, OCE^{+}12, OBÁSV18, OBI^{+}16, PGR^{+}10, PSCS16, PRLR^{+}18, PZL^{+}17, PBDMRB13, PPX17, PSS^{+}12, PJDC^{+}13, PHAZ18, PIPD10, PFL13, PBCA10, PW16, PB10, PBC^{+}12b, PSPK10, PCNM17, PBH^{+}17, PRA^{+}19, RCP^{+}16, RBSL11, RCAB^{+}12, RGW17, Rei12, RWR^{+}15, RF14, RMBM14, RTD14, RBHR19, RMSPA16, RSE18, RSG^{+}16, RNGUR^{+}11, RT12, RQN^{+}14, RBBG12, SUPCB11, SPSG14, SGS^{+}17, SNTN15, SHS^{+}17, SGL^{+}17, SMSB19, SKW^{+}14, SHT^{+}16, SE11a, SKM^{+}15, SM15, SWB^{+}10, SMF^{+}11, SS19, SSM^{+}19, SH17, SBW16, TBH^{+}18, TSS11b, TRHM14, TBG^{+}17, TP16, TDB^{+}14, TBB^{+}19, TTS17, TPCN12, TPFP11, VBT14, VCRUQ15, VQSH14, VB19b, VDEO18, WYS^{+}10, WMS^{+}11, WS13, WEF19, WRB^{+}15, WCG^{+}14, WFH11, WO11, WFB16, WS11, YWiIK15, YPWP18]. species [YSKH11, YSM^{+}10, YRL^{+}10, ZBK^{+}14, ZSEV17, dJH11, dJTH^{+}14, dlSBO^{+}12, dlSOV^{+}16, tHHR10, vDGK^{+}12, vdVFK^{+}10, vdHTT^{+}15]. Species-specific [DPY^{+}16, JSLG17, KCW^{+}17, ML11, NBM12, PHTS11, TYG12, VCQM^{+}16, BC16, DVV^{+}18, GDPR18, JBT^{+}16, KHBL11, LERS13, PSPK10, RWR^{+}15, SE11a]. specific [ACG12, BMSB13, BC16, BRB11a, BRT^{+}17, BLPP16, BKC^{+}19, CJM^{+}13, CSC^{+}11, CP11, CS11b, DR16, DPY^{+}16, DVV^{+}18, DBC^{+}12a, FM14, GKHM13, GDPR18, GGF^{+}12b, HHP^{+}17, HGKB14, HLN^{+}19, JJP^{+}13, JBT^{+}16, JS12b, JSMO14, JSLG17, KHBL11, KCW^{+}17, KHW^{+}12, LSBB17, LERS13, ML11, MUOS13, NNJ^{+}13, NBM12, OBI^{+}16, PBAP^{+}14, PSPK10, PHTS11, RBP^{+}10, RWR^{+}15, RZH16a, RHvD19, RGCT^{+}10, SRB15, SAL^{+}12, SA14, SE11a, SBB^{+}12b, SYK^{+}13, TM15, TFFPM18, TLFS14, TYG12, TTW^{+}12, VCG^{+}10, VCQM^{+}16, WTC^{+}15, WMHW14, ZBB^{+}19]. specificity [DM17, TDKS10, VE18]. spectra [Bet13, Kiø16, MTMM^{+}15, QGM^{+}14, TFP14]. spectral [LPC19, LDVP^{+}17]. spectrum [GCS^{+}19]. speed [AHØ^{+}15, AAM^{+}12, GBL^{+}15, OFM^{+}18, SFTG14, SF14]. speeds [KSFL14, SSHB16]. Sperm [BPWF14, ORJH19, AAM^{+}12, AAK^{+}15, EKE16, EPC^{+}19, FNF^{+}18, ORR^{+}14, PHB11b, PMM^{+}11, RWM18, SWNSF14, ZMRCP^{+}17]. Sphaeroma [DdR10]. Spheniscus [SUB^{+}13]. spherical [KWC^{+}17]. spherules [GBU15]. Sphyraena [DSC13]. spicatum [KMV18]. spicule [GBD13]. spider [BMEL11, SHP^{+}14, WAP10]. spill [BMC^{+}15, BSF17, BBM13, BBC^{+}12a, DBCS16, FCR^{+}15, GFF16, GPRR17, HGS14a, HGS14b, HGS15, HG13, JGG^{+}15, LBM^{+}18, MFDW17, PGR^{+}17, PPCH17, RFH16, SB15a, ZWP^{+}17]. Spilling [SKD17a]. spillover [GHD^{+}10, SKD17a]. spine [AdCM^{+}19, CNW^{+}12]. spine-tail [CNW^{+}12]. spined [BGK17, KB14, SML17]. spinosissimus [BHUCMO19]. spiny [BPF^{+}18, BFEMLÁJ19, BFG^{+}11a, BFG^{+}11b, BPG^{+}11, BM16, ESW14, FRTB14, GELJ19, HBC15, JBCGC19, KLKM12, KB14, KPS14, LH10, LÁBFÁF^{+}10, Mat10a, MKI^{+}15, SRG14, SGB^{+}10]. spiralis [BSP^{+}10]. spirolide [TLM^{+}11]. Spirorbinae [LLF^{+}18]. Spisula [TPM19].

Spitsbergen [HB15a, TGBM^{+}16, WCG^{+}14]. split [WDM^{+}18]. splitnose [GCM10]. Sponge [BCdV13, HD10, MKT^{+}18, AHR^{+}18, BHA^{+}11, CL10b, CMYL11, ESBG14, FBET15, GBU15, GEF^{+}12, GRFSM12, GMK^{+}19, GBD13, HKB^{+}19a, HFK^{+}13, KPKA13, LP12, ŁCM^{+}18, MPLLE17, MSE^{+}18, PW11, RdGv^{+}18, RZPH14, RDF^{+}19, SGEO13, TLK^{+}12]. sponge-algal [ESBG14]. sponge-loop [MSE^{+}18]. sponge-microbe [FBET15]. Sponges [VB19a, BHE^{+}17, DWV^{+}12, HCM^{+}11, LKC^{+}15, LP10, LALCM18, MZC^{+}10, MSE^{+}18, PMEZ15a, PMEZ15b, RdGv^{+}18, SL15a, WRGP12]. Sporolithon [DDP^{+}13a]. spot [BT13]. spots [SGC17]. spotted [CLB19, EBH17, GWK^{+}19, KB14, LBWB^{+}13, TPVCV19, WYS^{+}10, ZBB^{+}19]. spp [BBR12, BBG^{+}13, BRH13, CE14, CCM^{+}14, CBKM16, lLChTC11, DKL^{+}18, DRT13, GGMTGC18, HG17, HTK^{+}17, HVT12, HSS18, HPR^{+}12, HFN14, JCF^{+}12, JC17, KSFL14, KBC^{+}12, LMGJ13, MAP^{+}12, MTK^{+}18, PKW^{+}11, PLNE11, Pur18, RF14, RKM^{+}10, RLM^{+}14, SS11a, SCLB11, TR13, TMH^{+}13, WBG10, ZSSK11]. sprat [ALNB17, BSST13, BPM^{+}18a, GHT15b, PDCH15, RNB^{+}12, SKK12]. Sprattus [ALNB17, BSST13, BPM^{+}18a, GHT15b, SKK12]. spread [BL14, RPHB15]. Spreading [PML^{+}10, ATT^{+}16, GBL^{+}15]. Spring [EMG^{+}15, FTT^{+}17, dB10, ALANGT10, AFMR^{+}13, AFRAG16, Chi11, CMM^{+}18, DBA17, DVSM10, DSK^{+}10, DMvB10, DAB17, FKA^{+}12, GGL^{+}17, HEB^{+}11, HM09, JBM^{+}13, KFBM19, LS10a, LET^{+}12, NRK^{+}19, NYK^{+}13, PMF^{+}17, RNH15, RG14, SLL^{+}18, STP13, SLL12, SHSK11, SB14, TTPM12, VHP^{+}11, YSK^{+}10, YBB^{+}12, ZTNS13]. spring-spawning [JBM^{+}13]. spring-summer [CMM^{+}18]. spumigena [HvEBC12]. spurdog [DEGF^{+}10, TNB^{+}15]. Squalus [DEGF^{+}10, SGB^{+}10, TNB^{+}15]. squamata [BEFC15]. squared [NHTS14]. squat [BHP11, dlÁGRB^{+}19]. squeeze [MHF^{+}19]. Squid [ASR^{+}19, BBG12, BB14b, BCGG19, Bus12, CRAP^{+}12, FLC^{+}14, GBH12, GCS^{+}18, GRCH^{+}18, HEAI10, IMS^{+}11, LCC^{+}11, LMZ16, MBS^{+}13, MBS^{+}14, NSY^{+}18, PVYS16, RCVG10, Sat17, SHM^{+}10a, SHM^{+}12, SML^{+}14, SWNSF14, SZG11, SJ10, SHF^{+}12b, TKA10, VCRUQ15, VSW^{+}15, YHTC19]. squids [CBR19, NSY^{+}18, SRCE^{+}10]. Squilla [HCBV^{+}11]. Squillidae [HCBV^{+}11]. squirt [HPMB11]. Sr [KSK07, MC10a]. Sr/Ca [KSK07]. Sri [SKD^{+}11a]. ssDNA [EH13]. ssp. [DMJ11, KS12b, LSV10b]. SST [BSC^{+}16, LAM11]. St [BPOS^{+}16, BSB^{+}11a, BDVP19, BOCB10, GLM^{+}19, RBP^{+}10, SRP^{+}19, TBH^{+}18, BP12b, DPC13, PPB10b]. St. [BFN^{+}19, CP10, CPP10, DH18b, Edm13, EPL^{+}14, EL16, KBK^{+}15, LMLB12, LRTLF12, OLF11, PPJT15, PPD^{+}14]. Stability [BFC14, BL10, GGG18, LBL^{+}12, AN11b, BMSB13, Bar13, CDMW12, GHES18, HM13, KBK^{+}15, MG18, OCS14, PCNM17, RZB18, RGWA11, SH10]. stabilizes [MBC18]. Stable [ALE^{+}13b, ACG^{+}11, BPJ^{+}15, CKM^{+}16, CAG13, CCKB^{+}15, DEJN10, FLC^{+}14, HVJ^{+}18, JJP^{+}13, KSK07, LCGS^{+}19, LMR^{+}10, MWP10, MBSG11, MFH^{+}11, PBAS^{+}10, PBCA10, PLNE11, RMT^{+}13, SDR^{+}15, SCW^{+}15b, TTT^{+}15, VRSS18, VW10, ZO10, ALV12, ABN14, ASWB10, ACG12, AR14, AGRPSVL13, AGP^{+}13, BLV^{+}19, BCM^{+}14, BMJH12, BSC^{+}15, BMV^{+}11, BADQ^{+}15, BVAB^{+}17, BDR19, BBC^{+}17, Bre14, BGS^{+}16, CLD^{+}13, CCA^{+}17, CMIMGS15, CE11, CKK^{+}10, CMW^{+}15, CMB^{+}14, CS13, CPS^{+}14, CRHV14, DWPH11, DFS13, DBC^{+}12a, DSG^{+}13, DABP10, DBS^{+}19, EAMT10, EMH^{+}19, FC10, FPC^{+}13, FLAL10, FTKK14, FDS14, FHMG19, GSR10, GCSF18, GCS^{+}18, GCS^{+}19, GSSM^{+}11, GPC^{+}15, GHNF10, HAM^{+}15, HHS^{+}12a, HWET13, HFG^{+}19, HN13, HRT^{+}14, HKB^{+}19b, HB12, HS15b, HEAI10, JCRC10, JBS11, dDTC^{+}17, JS12b, KKS^{+}14, KRC^{+}11, KM17, LLBT18, LMV^{+}19, LFM^{+}16]. stable [LCB^{+}16, LRO^{+}17, MMFF12, MAP^{+}12, MWZ^{+}16, MBRO10, MBS^{+}13, MFS14, MCL^{+}17, MCW11, NFC^{+}11, NW15, NUO^{+}19, NLK^{+}19, OHR^{+}13, OFDG14, OBI^{+}16, PLG^{+}19, PKS11a, PH10, RP12, RKM^{+}10, RGDNGS11, SWP^{+}13, SMF^{+}17, SST19, SRA^{+}17, SGD^{+}12, TPC^{+}15, TFM^{+}18, TMS14, UTS^{+}18, VSG^{+}11, WBI^{+}15, WRW^{+}16a, WLF^{+}17, YAFV12, YMD^{+}10, ZBB^{+}11, ZRR^{+}13, ZMRCP^{+}17, vdBUPP^{+}11, KHF^{+}11]. stable-isotope [HB12]. Stage [DR16, AGP^{+}13, BRB11a, CG16, DSG15, DBG15, EDS14, FRTB14, HGL15, HSDB13, HHB^{+}14, KCGC14, MW16a, PSB17, SPMN19, SNK^{+}10, SE11b, TGP^{+}16, TFH^{+}19, WRFM16]. stage-based [MW16a]. Stage-specific [DR16, BRB11a]. stages [ACA^{+}10b, ACA^{+}10a, BA12, CLMB13a, CLMB13b, DSC13, DP13, Fog12, GN11, HEB^{+}11, HDT^{+}12, KLT^{+}12, KDO^{+}14, LSQ^{+}13, MH19, MDS^{+}13, NRF11, ORWPR14, PKW^{+}11, PSG11, PRP^{+}16, RTC^{+}14, SHP^{+}14, SHB^{+}12, SDAH12, SM17b, SFNP11, SMK^{+}17, TB14a, WSHY16, YHT16]. staggered [LFN^{+}12]. staghorn [MMCW14, RBL^{+}17]. staging [RKM^{+}10]. staining [BOS16]. stakeholder [RFEJK11]. standard [HU13, LCL^{+}17]. standardised [KTKM14]. standing [JM17, PCNM17, SBE^{+}16]. stands [OAD^{+}14]. star [ARD^{+}19, BGBD^{+}13, EN11, GG12, MWGH15, RSE18]. starfish [AM13, BBW^{+}19, DP11, HFLI11, KHH18, MPB^{+}14b, NKS^{+}15, PH16, PBS10, RPB17]. starry [LPB^{+}16]. stars [ATO^{+}14, BS11, GBE^{+}17, JH11, MDPG13, SHF12a, TKDM16, UHG12]. start [OVF11]. starvation [DTVH10, FSM^{+}14, GELJ19, OHST19, PDCH15]. State [LFJ^{+}14, RDF^{+}17, CRR^{+}19, DPO^{+}17, Fie14, GMSL^{+}14, HT10, KTP^{+}18, MRC^{+}12, OV10, OCSW16, OMH12, QSB16, RLG14, SOJ14, SP15, TPC^{+}15, TRJ^{+}11, WH11]. State-dependent [LFJ^{+}14]. States [Edm14, DABP10, FDS14, HPQ15, LRH14, OMW^{+}10, PH10, PP13, PWCvO^{+}14, SDAH12, PPB^{+}17]. static [TTW^{+}12]. station [SPHK17]. Stationary [KSHS19, NJJ^{+}11, SBE^{+}14, vdSRT^{+}18]. stations [GMB^{+}16, LAM11, WLM^{+}13]. Statistical [SWP^{+}13, RBHR19, YSKH11]. statistics [AGV^{+}11, SKW^{+}14]. Statoliths [HLC^{+}18, HSH18, LCC^{+}11, MGO^{+}12]. status [ATP^{+}12, BBD^{+}14, CLN^{+}13, INK^{+}17, LSV^{+}10a, McC19, MRB^{+}11, MBP^{+}12, MCL^{+}17, PRA10, SBL^{+}10]. stay [DDM^{+}16, MCPC19, PS15b]. Staying [BFG^{+}19]. steady [SP15]. steady-state [SP15]. Steamboat [WDNM11]. steelhead [BMJ16, HHP^{+}17, MBG^{+}15]. steep [FRP^{+}17, SMM^{+}15]. Stegastes [VDG^{+}15]. stellaris [CJM^{+}13]. stellatus [SSBA12]. Steller [ED12, SDR^{+}15, SGC17, SCW^{+}15b, VGR^{+}16]. stem [CAD^{+}10, CMP^{+}11]. stem-boring [CAD^{+}10, CMP^{+}11]. stems [ZH13]. Stenella [AMLR^{+}17]. step [GMD16, Joh10]. Stepping [MADI13]. Sterechinus [MBvO^{+}18]. stereo [DOH^{+}17, ETHB17]. stereo-video [DOH^{+}17, ETHB17]. sterility [BH10]. Sterna [Goy14]. sterols [LAL^{+}18]. Sthenoteuthis [SRCE^{+}10]. Stichopus [ETB13]. stickleback [SML17]. still [GSL17]. stimulate [BJG^{+}11, HSR^{+}10, WRW^{+}16b]. stimulated [SBE^{+}19]. stimulates [WLH^{+}17]. stingray [DWPH11, PFKB18, TLAT13]. stingrays [KSHS19, PBTB15]. stipulacea [BCGWS19, GALS16]. Stochastic [BGM^{+}18, RWS^{+}15, TPBM13]. stochasticity [PGR^{+}13]. Stock [HNK^{+}18, PSÓ^{+}15, AFOVRB14, AdCM^{+}19, BCM^{+}11a, FQ11, HMH^{+}12, HFR^{+}13, JS12b, LDM10, MLCC19, MS18a, MBC^{+}11, NYK^{+}13, ODGL^{+}12, OSD^{+}13, PCNM17, SMH^{+}12, SWK19, SBB^{+}12a, SWH^{+}18, SBE^{+}16, TSTP13a, TTW^{+}12]. stock-recruitment [MS18a]. stock-specific [JS12b, TTW^{+}12]. stocked [WSL^{+}14]. stocking [EHD^{+}17, TUK13, WYS^{+}10]. stocks [BNAN12, BDP10, DS13, DH17, DWL19, EBT^{+}15, GLR^{+}14, GFC^{+}13, LCJ^{+}12, OSD^{+}13, PRK^{+}12, RZH16a, ZJ17]. stoichiometry [BLER15, BFH13b, DVS13, ENW^{+}14, MFI11, WL10]. Stokesbury [HS11a]. stomach [CMB^{+}14, DWPH11, FC10, GCS^{+}19, HTSA16, IS16, MGAV15, MCW11, RCR^{+}13]. stomach-content [MGAV15]. stomachs [ATF12]. Stomatopoda [HCBV^{+}11]. stomias [ZAC11]. stone [HG17, MADI13]. stony [AZSKWL11, MMD^{+}13, VRMKH^{+}15]. stopover [MGCH^{+}16, TR13]. storage [GTM14, Hal15, JMB^{+}18, LBL^{+}12, ORJH19, POKGRP10, SNVH19, WA19, XLXS17]. store [WPMP14]. stores [DTL^{+}15, ORR^{+}14]. storm [GKC11, SABMG16, SGS^{+}12, VWH18]. storm-water [GKC11]. storminess [BJM19]. storms [BÁNC^{+}18, UCK^{+}13, WC18]. stormwater [SCRGP11]. story [Kin10a]. stout [GSO^{+}18]. strains [CIM^{+}13, DB11a, JDX^{+}17]. Strait [MS18a, GVG^{+}18, LBN11, MSGGE10]. Straits [DSS10]. stramenopilan [DR10]. stranded [NKM^{+}15, vBBW^{+}14]. strandings [BSG^{+}10, CLH^{+}15]. strategic [GMD16]. strategies [AEGH11, BDNA10, BC16, BVS^{+}16, BBL11, BFG^{+}11b, BS19b, CVW^{+}11, CMC^{+}16, CBLM^{+}17, DJJ^{+}16, DAM^{+}18, FWW^{+}16, FMH^{+}19, HWET13, JMK12, dDTC^{+}17, JLPG17, KWM^{+}17, KKPL14, KHK^{+}17, MRH11, OMW^{+}10, PLG^{+}16, PDT^{+}11, PS15b, PLGSD17, RHVRC11, RSH^{+}19, SMJ^{+}14, SBB18, SGS^{+}19, SKD^{+}11b, dSdSdCO^{+}11, vdGSE^{+}14]. strategist [CSV^{+}15]. strategy [AvBBL^{+}16, ACDJ11, ADA^{+}19, GGF^{+}12a, GPW^{+}10, GWGC16, GSV14, JP15, MBV^{+}11, NWP^{+}11, OFG^{+}18, POKGRP10, RM19, UHG12, VBT^{+}17, vDCFM10, vDCR13]. stratification [EOA15, GBW^{+}12, HRP18, HRS^{+}17, JCM10, JSCM15, PDBB14]. stratified [ALNB17, HMS^{+}12]. streaked [WYI^{+}16, YTO^{+}11]. stream [BDvGW16, BSB14, CZH^{+}13, SMSL17]. strength [BA13a, BGJ^{+}19, BRG^{+}14, CBD14, HW11a, JL17, PBH^{+}17, SPH^{+}13, ZGH^{+}17]. strengthen [SKW^{+}14]. strengthens [DGDJ14]. Streptophyta [SD17a]. Stress [TBMG17, BMR17, BPF15, BPLT14, BPHW16, CLM^{+}13, DSVM10, DS16, DHGW15, DEM15, FSR10, FC19, GSCH14, GB16, GFP19, JGWJ^{+}11, JG13, KF13a, KLT^{+}12, LSBB17, LM12, LBJK14, LKN^{+}13, LGC10, MVCD14, MCBR18, MM16, MA13a, PL15, PWM18, PMS13, PUP16, RMC^{+}14, RKvW10, SMOH19, SRB15, SKM12, SL11a, SLIGG18, SKM^{+}15, ST14, SG18b, TRR^{+}17, TB14a, TBL16, TNS^{+}12, ULCAZG17, VCVK^{+}10, VDM^{+}10, VLK^{+}15, WRO^{+}11, WL11, YHR16, YBD^{+}16]. stress-tolerant [LKN^{+}13]. stresses [JC15, KCD13, dlSBB^{+}10]. stressful [PDL^{+}10]. stressor [GBC^{+}16, Pör12]. stressors [BD12, FAT^{+}15, FDP^{+}16, FTF15, GEF^{+}12, JPI^{+}19, KPB16, MLB19, MW16a, MWGH15, PL15]. striatus [GB10a, SDST16]. Strict [PW11]. strigosa [Edm10]. striped [AMLR^{+}17, CPS15, GFS15, GCR^{+}17, JVSD10, MH10]. Strobilation [FWS^{+}18]. stromatolite [RSP17]. stromatolites [WPRS18]. Strombidae [BBdG^{+}19]. Strombus [DH18b]. Strong [BSC^{+}13, CVP11, JSMO14, LHPP16, SYX^{+}10, CBC^{+}13, CSV^{+}15, DKV^{+}16, HCJ^{+}18, PRA^{+}19, SJB16a, SPJ^{+}12, SOC13]. strongly [DHAV13, VWW^{+}16, VHKT15]. Strongylocentrotus [CD16b, Ebe10, EHR12, FNC^{+}14, FG16b, JH14, LH16, Mat10a, RMR12, ZVL^{+}16, ZVR18]. strontium [WWG12]. Structural [HMM^{+}15, BDT^{+}16, DAM^{+}17, GHES18, Gea10, GNBM11, MRB^{+}15, NJP12, Pro17, SMK^{+}18]. structurally [KFB15]. Structure [MDF10, PPTK19, ÅCAC16, APC^{+}13, Ale13a, ALSO^{+}18, AAGGFG^{+}12, AGM^{+}11, ANRT10, AFF^{+}15, AGH^{+}12, BEP^{+}12, BAA12, BSC^{+}13, BBF^{+}19, BMM17, BJCB^{+}14, BD12, BWH^{+}19, BSH13, BPB^{+}12, BGM^{+}18, BVAB^{+}17, BRC^{+}15, BD11b, BKF15, BP12b, BB14c, BOS16, BOCB10, BBC14, BSH^{+}14, CLD^{+}13, CLM^{+}13, CLB^{+}14, CLS19, CBKM16, CPA^{+}11, CL15a, CRJ13, CIF11, CCNB15, CMO^{+}10, CL10b, CFSM^{+}13, CDCZ16, CM11, CGEGC^{+}19, CS11a, CPS^{+}14, CML^{+}12, CCM^{+}12, CHLR16, DSH14, DLRF15, DMW^{+}16, DTS14, DMS14, DBS14, DM11, DS10, DSJNH18, DIB15, DCT^{+}10, DPO15, DBS^{+}13, DHR^{+}13, DH17, Edm13, Edm14, EL16, EM13, EHD^{+}17, EBK^{+}12, ETD^{+}18, ECZD14, FHM^{+}11, FKP^{+}15, FTWA17, FH10, FL17, FSDS10, FHE^{+}15, FKRJ13, FFS^{+}13, Gar13, GBH12, GSM^{+}11a, GP12, GMSC14, GLY^{+}14, HHLT10, HVK^{+}13, HW11b, HTCH^{+}18, HDWF10, HRÁB^{+}14]. structure [HPN11, HPL12, HOA14, HGKB14, HD10, HOE^{+}12, IHD^{+}16, JW10, JDQS11, KHGS12, KBR^{+}11, KWM^{+}17, KWH^{+}18, KKS^{+}14, KJBS12, KLNJ15, KMW17, KLSJ10, KBB^{+}10a, KFGSMPG14, LH11, LMRRB17, LCB10, LRL^{+}13, LLRC^{+}12, LLC^{+}14, LCBK^{+}13, LEB^{+}19, MBJ16, MZ10, MBvR12, MKK13, MHB14, MHHB^{+}17, MTG^{+}10, MH12a, MBB^{+}12a, MBR^{+}18, MBS^{+}18, MWWS17, MSG11, MRGG^{+}13, MSS^{+}14a, MBRO10, MBvO^{+}18, MMFP^{+}15, MIM^{+}12, MH11c, MCHC^{+}16, MER14, NFR^{+}14, NSI^{+}15, NLGE18, OTK^{+}13, OMT^{+}17, PSÓ^{+}15, PBV^{+}11, PNF^{+}12, PMB^{+}10, PFVFF^{+}15, PM16, PPTB12, PBC^{+}12a, PSC13b, PES12, PFR11, PHHL^{+}11, PCLB18, QHG^{+}19, RMKR13, RSDN15, RRH^{+}16, RA10, RCG^{+}15, RSG^{+}16, RGDNGS11, RMT^{+}13, SCB15, SBV^{+}15, SSBA12, SHR^{+}11, SLSC16, ST13, SCRL11, SRR15, SKB15, SML^{+}15, SMP10, SBB^{+}12a, SHM^{+}12, SHH10, SBM^{+}16, SRER16, Sma13, SBE^{+}16, SGT12]. structure [SGTP16, SBKA14, SWB^{+}10, SSM^{+}16b, SDB12, SLRS13, TS11a, TFFPM18, TBCG^{+}12, TBH^{+}18, TNM^{+}11a, TCN^{+}10, TL18, TPC^{+}16, TPBM13, TSGW10, TKS^{+}14, THW^{+}16, TBDM11, TCKY19, TDSS15, TNN11, VANB15, VB19a, VDM^{+}10, VDV^{+}15, VC14, WBB^{+}10, WHK^{+}11, WLP^{+}11, WMB^{+}14, WJ12, WWZ^{+}19, XTW^{+}15, YWiIK15, ZTNS13, ZS17, ZKS17]. Structure-dependent [MDF10]. structure-forming [RSG^{+}16]. structured [APC^{+}14b, BAL14a, BCB11, GKW10, PSS19, TS11a]. structures [BJ14, BEFC15, DvV^{+}16, DT11, DHG11, JOFP18, JROS17, JGCRS12, LABG^{+}12, LBC^{+}19, NLLS15, POPH12, RTD14, RPL^{+}18, SRCE^{+}10, YMCV10, ZFZ16]. structuring [BWS11, BPB11, CAR^{+}19, DBE^{+}18, HSDB13, KPB11, LTN^{+}11, MLC14, MYS^{+}16, NC11, OCPM12, RP13, VVB19, VWH18, YBH^{+}11]. studies [Ano17c, Ano18a, Ano18b, BLT14, DSG^{+}13, GHGG11, GBC^{+}16, GHNF10, HYC^{+}14, KAP^{+}12, MOMF14, MPM^{+}18, NSS10, PvdMLC17, SCB^{+}16, SKK12, SN17, WSS^{+}16]. study [ABF^{+}14, ARW12, APH^{+}10, ASK^{+}15, ADRA19, BDLW10, BDM^{+}14, BCPC18, BRB11a, BPHW16, CPSC^{+}15, CLB19, CCDP15, CPR^{+}14, CPS^{+}14, DGDJ14, DMC^{+}17, DRG^{+}12, DOH^{+}17, DDB^{+}14, DMR10, FSP^{+}17, GHGF12, GSJS13, GM13b, HAF^{+}17, Hug11, JSBP15, JD18, JPF12, KV13, KRC^{+}11, LAB10, LA12a, Lec18, LS10a, LOBW^{+}11, LD10, MPF^{+}15, NRF11, NGA18, NIPS12, NSY^{+}18, NJJ^{+}11, OS12, PBM^{+}11, PPB^{+}10a, PRNR^{+}18, PMS10b, RWR^{+}19, RR18, RCK^{+}11, RWM^{+}12, RUD^{+}16, RZH16a, RvOP12, SMSE19, SKR^{+}10, SABMG16, SMJP^{+}15, SS14, TWY^{+}19, TAB^{+}14, VGGF12, VSM11, VCMS^{+}18, YNK18, ZVL^{+}16, ZXF^{+}16]. studying [AN11b, CGG^{+}10b]. sturgeon [FPC^{+}12, MSH^{+}14, SNS^{+}13, ZBK^{+}14]. stutchburyi [NDBL16]. Styela [SF14]. style [BL13]. Stylophora [TNS^{+}12]. Stypopodium [CSJP17]. Sub [LBMZ15, PNPF14, SSR^{+}10, TM15, AP14, BHP12, BMW^{+}14, BRS10, BCSR10, BSH13, BPHW16, CRS^{+}11, CMO^{+}10, DDK^{+}18, dSKFV12, GMT^{+}17, GEAOH17, Gal14a, Gal14b, GBH^{+}10, GMT^{+}11, GMES18, HCL^{+}11b, JPAM^{+}15, SRER16, SFKQ15, TAAN15, VCMS^{+}18, WBG^{+}14, WRGP12, OCMT13]. sub-adult [SFKQ15]. sub-Antarctic [GMES18, OCMT13]. sub-Arctic [BRS10, BCSR10, BSH13, CRS^{+}11, GMT^{+}17, GBH^{+}10, JPAM^{+}15, SRER16, TAAN15, VCMS^{+}18]. Sub-canopy [PNPF14]. sub-colony [WBG^{+}14]. sub-estuary [Gal14a, Gal14b]. Sub-lethal [TM15, BPHW16]. sub-littoral [HCL^{+}11b]. sub-populations [DDK^{+}18]. sub-specific [TM15]. Sub-surface [SSR^{+}10]. Sub-tropical [LBMZ15, AP14, BHP12, BMW^{+}14, CMO^{+}10, dSKFV12, GEAOH17, WRGP12]. sub-zero [GMT^{+}11]. Subantarctic [FLE^{+}11, CCB^{+}17, NFSW10, TGR^{+}19, WTdB16]. Subarctic [BTS10, DMP12, OKJMC15, RFRE19, RSK^{+}14, YSK^{+}10]. subestuaries [PW15]. subject [BPZ^{+}14]. subjected [RPRN^{+}16, WDL^{+}18]. sublethal [ACW18, BBW^{+}19, CL15c, TKIO10]. sublittoral [CHRB10, DEWK15, LCL^{+}11]. submarine [BWS^{+}12, GBTS11, GLM16, KHN15, LTN19, SRH^{+}15, SJL^{+}15, SFL^{+}18, SKN^{+}17]. submerged [BLMD^{+}11, PW15, PBA12, SLS11]. subphotic [WGD^{+}19]. subpolar [FSP^{+}17]. subpopulations [DP13, WBJ12]. subregional [CM14]. subrotundata [RB15]. subsea [RT17]. subsequent [CG17, SMOH19, WMMC14]. Subsets [dJSB19]. subsidies [HLD12, KKS12a, MCRB11, PKBS14, SGS^{+}19, SSM^{+}16b]. subsidize [FESPR14]. subsidy [FK10, KS12b]. substances [KCP^{+}11, MPBSB18, UFP^{+}10]. substantial [RUPT^{+}11]. substitute [dJSB19]. substrata [BAL^{+}11, Lec18, TPCN12]. Substrate [HFK^{+}13, LBD^{+}15, DM17, HFLI11, RCBE15, SMS15, SSS^{+}10, ZOM^{+}10]. substrate-borne [RCBE15]. substrate-spawning [SSS^{+}10]. substrates [GBH^{+}10, RBC^{+}15]. substratum [DT11, PBIO^{+}18]. Subsurface [PQAH^{+}17, AM18, MTG^{+}10]. subtidal [Ale13a, BG12, BT14, BVS^{+}11, Bur12, DRT13, DP13, DL14, EDS14, ESH^{+}10, FT10, FPR18, GPRR17, JH11, KSB11, KS12b, NC11, OSE13, PDC^{+}17, PCF^{+}19, PMOAC^{+}17, PGR^{+}17, RTD14, SdBAH^{+}14, TBCG^{+}12, WL11]. Subtle [GP11b, WCRCR17, BCSH18]. subtorquata [SWMK15]. subtraction [WHA^{+}13]. Subtropical [DA11, DBDK14, LKLC11, RKC16, ANJ^{+}11, BHF^{+}10, BGP^{+}16, BH15, BHF^{+}16, BKF15, BFH13b, CPEE18, CWD11, CL10a, CLC12a, CL15b, CWEE19, CAC11, CGH12, CLMM^{+}14, CHFS12, EOM^{+}14, EHG16, GVW^{+}12, GOO14, GKE^{+}19, GLHM^{+}18, HMS10, HHH^{+}11b, HT15, HHH^{+}11c, JCZ^{+}12, KLD18, KBK^{+}15, LSJ17, LHL11, LMV^{+}19, LCS^{+}10, LHT^{+}17, MPB15a, MGOL19, MSS^{+}14b, NMCP17, PS15a, PML^{+}19, Pro17, dGCF^{+}17, RNK^{+}11, RG14, SML^{+}15]. subzero [DSM11a]. Success [TP19, ALWD15, AED^{+}10, BVO16, BFS10, BWJB16, BCPC18, BBW^{+}17, BJS^{+}11, BCH^{+}11, CLC^{+}12b, DM17, DBHMG12b, DMHS^{+}14, FGP17, GJ16, Gre10, GVGP^{+}14, HWR^{+}13, HWVM19, HC13, HPR^{+}12, HTR^{+}19b, HPH^{+}18, Kin10a, KNB^{+}13, LSJ17, LMCD19, MWB13, MS18b, MJHS11, NWHD15, NC14, OMN17, PMGQM17, PWM18, SSS19, SE16, TPBM13, TSCSB15, TS12, Ued14, WFP15, YPK15, YHR16, vdBUPP^{+}11]. Successes [Pur18]. Successful [WYS^{+}10, BST^{+}19a, CMH14, DBHMG12a, NGR^{+}17]. Succession [DSK^{+}10, VBT14, BAL^{+}11, ETD^{+}18, FW11, GCCM18, HMM14, LJB14, NCT^{+}14, Par15, RNH15, SBB^{+}14]. successional [NBGGST^{+}15]. succumb [RBL^{+}17]. such [Ano17c]. suction [TFW^{+}12]. sufficiency [HCGP13]. sufficient [MSK^{+}19]. sugar [KKY15]. suggest [BMPH14, DSH14, EPF^{+}17, NWP^{+}11]. suggests [BMÁN^{+}18, BCGWS19, FSE^{+}17, MS18a, ODK^{+}14, PRP^{+}18, SRLV10, SH17, TOW10, VBV^{+}10]. Suitability [HMH^{+}12, DER^{+}19, MIV^{+}10, MRC^{+}19, SM17a, TGI^{+}11, WHH^{+}13]. suitable [CEH18, MS10]. suite [DSS10, OFFQGG10]. Sula [ACAD12, ZHM^{+}10]. sulfate [BST^{+}10]. sulfide [BRC^{+}15, LZF^{+}14]. Sulfur [VKH10, CS13, DCIG18, PDBG14, PDGL17]. sulphide [ARGB13]. sulphides [CKV12]. sulphur [AOW11]. sulphur-cycle [AOW11]. Summer [TLM^{+}11, AFRAG16, AMAM^{+}18, CMM^{+}18, CH12, DABP13, DTG16, DVSM10, FLE^{+}11, HMH^{+}18, HBF^{+}10, HIH^{+}17, JLM^{+}12, KTF10, LSY^{+}18, LLH^{+}11, MFD12, McD13, MSH^{+}14, MSPO12, NSK^{+}14, SBE^{+}19, SMF10, SB14, WKX^{+}19, WVD^{+}17, ZCC^{+}12, vdVFK^{+}10]. sun [BVM^{+}19, EN11]. Sunda [Ano10f]. Supan [MAB^{+}12a]. super [GGvW^{+}15]. superba [CDCZ16, CDC18, JR15, KTF10, MAS^{+}10b, RWG15, SRLV10, SKD^{+}17b, SGCF^{+}12, SKR^{+}13, THP^{+}16]. Superfund [TWL15]. superior [JL12]. superoxide [MDAPH11]. supplemented [FRMS17]. supplied [SKFY18]. supplies [DCW^{+}10]. Supply [DAD^{+}11, LAM10, ADO^{+}12, BGL^{+}11, DJL^{+}12, FLS^{+}18, HHS10, HWR^{+}14, JDG^{+}17, KF13a, KTK^{+}12, LFU^{+}19, LRL^{+}13, LNK^{+}16, MCHC^{+}16, NMG^{+}13, OFFQGG10, PHB11b, PRR^{+}16, VARO12, VANB15, WLHW10]. Supply-side [LAM10]. Support [SFL^{+}18, CED^{+}19, GGF^{+}12a, GDB^{+}16, HP14, HIH^{+}17, OHP18, RCF17, SDSA14, SMF^{+}12, SSFB11, TMH^{+}18, WPP11]. supported [RTD14]. supporting [SBS^{+}13]. supports [CBC^{+}16, LS15, SBT15, VAHL11]. suppress [VDS11, YLPG19]. suppression [GB10b, TPKC18]. supralittoral [LHPP16]. supratidal [BVC^{+}16]. surf [BSG^{+}17, LET^{+}12, NLV^{+}15, SMM^{+}15, WBLRN18]. Surface [RJ12, BFF^{+}10, BCPC18, BLT14, CPHH15, Dao13, DEM15, FMH^{+}15, HCJ^{+}11, JNN12, LNS10, NTN^{+}13, Ran18, SLSC16, SSR^{+}10, SWH^{+}18, SBE^{+}19, TH11, VGL16, WSC13, WWG13a, WLL^{+}14, WHP^{+}17]. surface-biofilm [TH11]. surface-pelagic [SWH^{+}18]. Surface-sediment [RJ12]. surfaces [UHG12]. surfclam [TPM19]. surfgrass [BBWG12]. surfgrasses [MH11c]. surfperch [JVSD10]. Surgeonfish [ELM^{+}19, GB10a]. surplus [BBH^{+}12, EM11, FOL12, LCB^{+}12]. surprising [MMNR^{+}12]. surrogacy [BT16]. surrogate [HFM12]. Surrogates [BAFB12, AHD12]. surrounding [MPH^{+}11]. survey [BBD^{+}14, CKV12, CS10a, CQR^{+}17, HHI^{+}14, KW17, KBB14, LPC^{+}11a, LCG^{+}17, RSG^{+}16, ZBC10]. surveys [BZC^{+}12, BBBF15, BW10, CMV^{+}11, DP16, GSF18, JTH^{+}12, KWC^{+}17, LLBT18, LYH^{+}17, MVE^{+}10, RZPH14, SPHK17, SOA^{+}17, SDB12, TRA^{+}14, VBR^{+}14, WYS^{+}10]. Survival [ABT^{+}15, PVZD^{+}11, POKGRP10, ABK^{+}14, ALWD15, ADL^{+}18, AGHC10, ACA^{+}10b, ACA^{+}10a, BG19, BAL14a, BCSH18, BWRP14, BLM10, CSJP17, CDMB^{+}18, CL10c, DABP13, DTG16, DAC^{+}12, FLS^{+}18, FJ11, FA10a, FOJF15, GSCH14, GCC^{+}16, GVR^{+}14, Gea10, Goc10, GGMTGC18, GWGC16, GDH^{+}19, HHP^{+}17, HCG^{+}10, HIMH15, HHB^{+}14, ICUM10, JMO^{+}19, JNYT13, JE10, LAM13, LK15, LD17a, LTG17, LCC12, LH19, MJO^{+}13, MM10, MWC^{+}12, MLB^{+}14, MSPO12, MBG^{+}15, OPSM13, OA10, PES12, PB19, Rai12, RBP^{+}10, RPLT10, REAN^{+}14, RP15, RWAP16, RLM^{+}14, RBD^{+}14, SMOH19, SWMK15, SES12, SHP^{+}14, SB18, SFW^{+}14, SDCC12, SG18b, SYK^{+}13, SNT^{+}16, TG12a, VB19b, WMMC14, WBB13]. survivors [HHB^{+}14]. survivorship [EMMSL13, HB15b, IHP10, PC19, SKFY18]. susceptibilities [HPL^{+}16]. susceptibility [BS12, CP11, KBW^{+}15, dlSBO^{+}12]. susceptible [Joh11]. suspended [CMC^{+}15, GMK^{+}19, HU13, QOGS^{+}16, SDH13, YMPR13]. Suspension [WS17, BD19, FS18, GvSGCK14, HR12, Kiø16, LMW12, PDC^{+}17, RVF14, RL10, RL18, SF18, WJ12]. suspension-feeders [RVF14]. suspension-feeding [BD19, FS18, GvSGCK14, Kiø16, PDC^{+}17, RL10, RL18, WJ12]. sustain [KWS^{+}13]. sustainability [Joh10]. Sustainable [GKRG^{+}19, AS12, VQSH14]. sustained [JP15, MPB14a, SNTN15, SCST17, WAMS^{+}14]. sustaining [KCS^{+}17]. sutor [BDB^{+}13]. Suwa [Ano12c]. Svalbard [ÅCAC16, BNE^{+}19, DAW^{+}11, FLK12, FSP^{+}17, FKA^{+}12, SGL^{+}17, SdBAH^{+}14]. Sverdrup [Chi11]. SW [PNI14, PDGR^{+}13, SKR^{+}10, TAAN15, AED^{+}10, BST^{+}19a, BSH13, CMP^{+}11, CFF^{+}18, GME^{+}19, JPAM^{+}15, KAAI15, LMB^{+}19, MTG^{+}11]. swamp [PB10]. Swarm [KTF10, KSK^{+}12, RGWA11]. swartzii [LH19]. Sweden [IEM16]. sweepstakes [CFMN16]. swept [DZL16]. swim [SSHB16]. swimmer [SMSL17]. swimmers [ITT14]. Swimming [CG11, DFPH11, NWH^{+}12, RMR12, SWK18, CWR^{+}18, CG10, CPF^{+}16, FG10, HR15, KSFL14, KPS14, LEB14b, NM18, Nor13, SFTG14, WHA^{+}13, WLH^{+}17]. switch [PvdMS^{+}15]. switches [LR15]. switching [VF13]. swordfish [ADHP12]. Sydney [GB10b, MCBR18, TSR17]. Syllidae [Mus12]. Sylt [BAA12]. Symbiodinium [BSB^{+}11b, CvOB18, DBE^{+}12, EPL^{+}14, FSE^{+}16, FSE^{+}17, KBK^{+}15, LKN^{+}13, MAP^{+}12, PMH^{+}13, PHD^{+}14, QBW18, RAH^{+}10, SCLB11, WMS^{+}11, YSS^{+}18, YTM^{+}15]. Symbiont [FMC^{+}11, FSE^{+}16, BMSB13, FSE^{+}17, MWBMH13, SCLB11, ZRBV15]. symbiont-bearing [MWBMH13]. symbionts [FC13, GEF^{+}12, LMP12, SYK^{+}13, VBŽ^{+}17]. Symbiosis [MU13, YSS^{+}18, BFKR17, CvOB18, GFF16, HSS^{+}19]. symbiotic [BFPONS^{+}12, BFKR17, GBU15, LFPC^{+}14, MU13, PHD^{+}14, QSB16, SHDR13, TPFP11]. Symmetrical [Kin16]. sympagic [WBI^{+}15]. Sympatric [HKL19, MCL^{+}17, APNC^{+}15, APBC^{+}19, BPH^{+}19, CCJR14, CAB^{+}19b, CMB^{+}14, GSCH14, GLR^{+}14, HSDB13, HSCA17, HSWS16, JMS^{+}15, KWPL11, KSBM^{+}11, LIT13, LFN^{+}12, MTS^{+}17, MPRPM^{+}18, PRLR^{+}18, PTP^{+}15, RPR16, RMT^{+}13, SDG^{+}12, SØK^{+}15, SEF^{+}15, SST19, VHHGF11, WCRCR17, YMD^{+}10, YSM^{+}10]. sympatrically [OWH^{+}19, PLNE11, PRM17]. sympatry [BRBH19]. Synalpheus [MH11b]. synchronized [RM13]. Synchronous [TFW^{+}12, VGL16, STS^{+}18]. synchrony [ARA^{+}10, MBW10, MWM^{+}11]. syndrome [PWCvO^{+}14]. syndrome-infected [PWCvO^{+}14]. Synechococcus [LAALH10, TBP11].

Synergism [GP11a]. Synergistic [HRM^{+}11, PSER17, WHT12]. synergy [VPBV13]. synthesis [WSS^{+}16]. Synthesizing [LGM^{+}12]. synthetase [HKT10]. syntopic [RVF14]. System [BPSG11, Gla10, SHF^{+}12b, TAB^{+}14, WJW^{+}18, AFSL^{+}15, APC14a, AS13, BO15, BD12, BVC^{+}16, CPS^{+}14, DOH^{+}17, DLL^{+}18a, DTM^{+}12, DAD^{+}11, EOMS15, EBH17, FF12, FBC^{+}15, FL17, FAZ^{+}14, GMG^{+}16, GMP^{+}11, GJSR18, GCCCM16, HWE^{+}12, HRPBLT18, HRÁB^{+}14, HvEBC12, HOA14, JS10, JBD10, KB12, KFBM19, KBGA19, LLCM^{+}12, LGJH12, LEB^{+}19, MHS17, MM13c, NOB^{+}19, NIS14, OCSW16, OMM^{+}12, PA13, RM13, RHF19, RWM^{+}12, RUD^{+}16, RRS15, RCG^{+}15, RSW13, SKB15, SG18a, SRER16, SML^{+}14, STP10, SOBNL13, THRBL^{+}16, VNN10, VLM14, WAM^{+}15, WSS^{+}16, WL11, WPK10, WWG^{+}13b, ASEB11, HM09, ICT^{+}11, PBC^{+}18, RHRB^{+}17, TKA10, WBS^{+}12]. systematic [LCF^{+}14]. systems [AAIB15, BADQ^{+}15, CTvdM^{+}08, DKAC^{+}12, DHF^{+}13, EW16, GBS^{+}10, HCAV10, HPP^{+}10, JSS^{+}11, JCZ^{+}12, KPB11, KM17, LRM^{+}14, MWC18, MSPO12, OMZ10b, PHM11, RKFD15, RGB^{+}10, SAZHAAHS14, SL15b, WAMS^{+}14, WKRW^{+}12, vdVFK^{+}10].

T [Ano12c]. T. [HPL^{+}16, RTCF^{+}17]. tactics [RDHM10, SHD^{+}19]. taeniolatus [SCMSB^{+}11]. tag [FLK12, MSF18, MWR^{+}12]. tag-inferred [MSF18]. Tagelus [GSO^{+}18]. tagged [LBJI14, WFH^{+}18]. Tagging [SHB^{+}12, HMB^{+}12, JMO^{+}19, KPP^{+}16, WLP^{+}10]. tags [ADHP12, BDT^{+}11, DH18a, FGTS16, GPA^{+}16, GTM14, MAW^{+}11, MHSG18, MW16b, MDNL^{+}11, RBD^{+}14, SKW^{+}14, SGB^{+}10, TFW^{+}12, VSG^{+}15, WV12]. tail [CNW^{+}12, PFKB18]. tailed [MGPC13, VMC^{+}19, YTM^{+}12, vBMH^{+}17]. Taiwan [LLC^{+}14]. take [BMV^{+}17, DDB^{+}14, PSdA^{+}17, RMRA15]. takes [GPW^{+}10, LRBSJ15]. Taking [ZJ17]. Taliepus [CFA16, SFNP11]. Tamar [Dan11]. tamarense [BFL^{+}11, SE11a, VTZ^{+}13]. Tampa [WLBB^{+}13]. tang [ETS^{+}11]. tank [TNOY17, VK11]. tanks [JBCGC19]. Tanner [CRS^{+}18]. tapeworm [RR18]. tapeworms [RB11]. targeted [MHFN10, PBM^{+}11]. targets [BFM^{+}13]. tarpon [LA12b]. Tasman [CNSW16, HBE^{+}13]. Tasmania [AGHC10, MJW19, OCMT13]. Tasmanian [WPB^{+}14]. tasmanica [SWNSF14]. taurus [HBB^{+}18, KCS12, RABS19]. taxa [AS12, BMÁN^{+}18, Col10a, HFM12, PMF^{+}17]. taxifolia [BOU13, TBKG10]. taxon [BMSB13, LSRRCM13, TFFPM18]. taxon-specific [TFFPM18]. Taxonomic [AAGGG19, JDB^{+}10, BHT15, COK10, EBOM11, HCGP13, ILDL10]. taxonomy [MAD18, STSS^{+}16]. technical [MS11, ST11b]. technique [DVA^{+}18, JRK13, PSG11, RT12, RAB^{+}15]. techniques [FH18, GSR10, GUC^{+}15, HSH18]. Technological [HW12c, HW12b, WV12]. technologies [HMB^{+}12, KWC^{+}17]. technology [JAPK16, WLM12]. teeth [BAH^{+}15, LBH18, RLAGL12]. Tegillarca [NLKZ12]. Tegula [CS11a]. tehuelchus [NGSF13]. telemetric [VSG^{+}15]. Telemetry [HLS18, AMP^{+}11, BWM^{+}15, BMJ16, BHL^{+}19, CTZ^{+}16, DFD^{+}19, JRW^{+}14, KLBT^{+}16, LAH^{+}15, MOMF14, MR16, PGWS10, RNSA15, SRK13, SFKQ15, SD17b]. teleost [BH15, BHF^{+}16, IMD11, MS14, SLF^{+}19, WMHW14]. Teleostei [LLD^{+}12]. teleplanic [MKI^{+}15]. Telesh [POL^{+}11]. Telomere [YBD^{+}16, YOWK17]. Temora [CA14, DMvB10, DvBH12, GvSGCK14, KDB^{+}10, SK11]. Temperate [NFN^{+}15, PMOAC^{+}17, ANB13, ASS^{+}13, ÁMLUN16, AR14, BGB^{+}17, BAS^{+}10, BFNB16, BWS11, BG12, BCC^{+}11, BCB^{+}17, BCM11b, BFKR17, CLB^{+}14, CV11, CVP11, CEW15, CHG^{+}18, CSN12, DDB^{+}16, DMPG11, EBB10, FG10, FGOS16, FFLS16, FHG^{+}14, FA10a, GFKC13, Gar13, GHNF10, HCL^{+}11a, HPL^{+}16, HPP^{+}10, HTR^{+}11, HLB^{+}16, HWTH12, HSB12, HvEBC12, HAF^{+}14, HSCA17, HCK12, HPQ15, JMB^{+}17, JD18, JC17, LMC14, LKN^{+}14, LRSB19, MLM^{+}12, MPA^{+}10, MM12, MIIB16, MMA18, MFS14, MS18b, NVA^{+}19, NHAS15, OS18, OLM^{+}12, PSCS16, PBF10, PBDMRB13, PMSGBL16, PW19a, PRA10, QHG^{+}19, RKD18, SK13, SHK14, SPHK17, SHT^{+}18, SdBAH^{+}14, SYM^{+}16, STSS^{+}16, SESS^{+}18, SDVL16, SS10c, SBW16, TNIH12, TS16, TDSS15, VBM^{+}13, VBT^{+}17, VHKT15, WHFM17, YKBB12, dlSOV^{+}16, dlSGS17, vdVFK^{+}10]. temperate-latitude [HSCA17]. Temperature [BVG13, BFKR17, CGS^{+}19, CG10, CLHW14, CLMB13b, CBF^{+}13, CLSS17, CvOB18, DRH15, GALS16, GG16, Hos14, KRdVG^{+}18, KdVvdP^{+}12, KDO^{+}14, LAM13, NC14, RPLT10, RPS12, YKT^{+}14, AZVB13, AS14, AMM^{+}12, ACA^{+}10b, ACA^{+}10a, ABW14, BCOMKJ18, BFF^{+}10, BSC^{+}16, BSD^{+}14, BDDvdM17, BLER15, BRS10, BOKJ18, BWH^{+}17, BLT14, BSBR^{+}12, BNS^{+}14, BAW11, BDC10, BOU13, CFA16, CLGCA^{+}19, CL15b, CKJOM14, CEK^{+}16, DMMW12, DMM^{+}10, DEM15, DWV^{+}12, FSD^{+}14, FJ12, FOJF15, FL17, GCC^{+}16, GOM^{+}19, GGPDS19, GCP^{+}17, GVGP^{+}14, HG17, HRS^{+}17, HJMN^{+}12, HCG^{+}10, HWVM19, HFR^{+}13, HWTH12, HT11a, IAV^{+}10, JD18, JJF14, KCCG12, KCF^{+}14, KIU^{+}13, LNS10, LAM11, LGW^{+}15, LBJK14, LS10a, LRTLF12, LNIH16, LSV10b, MGFW13, MJW19, MMDM12, MAS10a, MH12a, MHR19a, MAB^{+}12a, MSG11, MCR^{+}13, MW16a, MWW^{+}14, MU13, MLML^{+}19, MRC10b, MER14]. temperature [MJHS13, ORWPR14, ORE^{+}17, OSD^{+}13, PL15, PHH^{+}12, PRNR^{+}18, PSER17, PW19a, RLN^{+}13, Rai12, RHB13, RBA^{+}18, RWO13, RCE^{+}13, REG^{+}11, RDM18, RB15, SBT12, SLSC16, SPH18, SDCC12, SSM15, SMS15, SDAH12, SM17b, SSX^{+}17, SF18, SS16, SZG11, STP10, SHB^{+}10b, TM15, TMS^{+}16, Tho15, TBL16, TS14, TVM16, VWW^{+}16, VRMKH^{+}15, VSW^{+}15, VCFF13, WBW^{+}14, WWG12, WSC13, WWG13a, WMB^{+}14, WFB16, WEK14, WBB13, WHP^{+}17, XTW^{+}15, ZXF^{+}16, dJdBdJB12]. Temperature- [YKT^{+}14]. Temperature-dependent [DRH15, KdVvdP^{+}12]. temperature-depth [AMM^{+}12]. temperature-driven [Rai12]. Temperature-induced [RPS12]. temperature-regulated [KCF^{+}14]. temperatures [BPWF14, BD14, BCR15, CRGM18, CLS^{+}16, DSM11a, GHCC15, GMT^{+}11, HBHS^{+}18, MCBR18, McD13, MSPO12, RBL^{+}17, RLO18, TBS^{+}17, VBD^{+}11].

Template [AMAJ^{+}17]. tempo [HW11a]. Temporal [AFLM12, ADB10, BS16, BDB^{+}13, CT10, CR16, CLDC^{+}13, CSF13, CH12, FC10, HCL^{+}11b, JU16, JHL^{+}19, LRL^{+}13, LB11, LLCM^{+}12, MLR17, MRE14, MFS14, ORR^{+}14, OSD^{+}13, PBDMRB13, PBAP^{+}14, PCNM17, REBL16, RKC16, RDGC17, RGWA11, RCP^{+}13, SF16, SJC10, SMSB19, SLO19, SH17, SKPS^{+}16, SBMÅ12, TBP11, TBB^{+}19, YSK^{+}10, ZPLW11, dQSC14, AN11a, AMP^{+}11, ARA^{+}10, BS14a, BSK^{+}11, BSC^{+}15, BSB13, BMVBC10, BAVSP11, BDPL19, BN16, BBW^{+}17, BCB11, CBR^{+}18, CP10, CCS^{+}18, CAG13, CPR^{+}14, CALB10, CPS^{+}14, DDP19, DAD^{+}11, DBS^{+}19, FYG^{+}18, FHE^{+}15, FGH^{+}13, GVW^{+}12, HHLT10, HHB^{+}14, IME12, JBL^{+}16, JJNP12, KBB^{+}10a, KAJP^{+}15, KM17, LBN11, LS16, LSV10b, MDLM17, MSFM12, MMFF12, MVE^{+}10, MH10, MTS^{+}17, MPRPM^{+}18, MMB^{+}12b, NSH^{+}15, NDG15, OCS14, OLM^{+}12, PNI14, PCP^{+}15, PvdMS^{+}15, PPTB12, PP13]. temporal [PMH^{+}13, PHHL^{+}11, PHG^{+}14, RHH^{+}15, RDF^{+}17, RVF14, RCU^{+}11, SDG^{+}12, SV13, STYT10, SKD^{+}17b, SRP^{+}19, SPHK17, SSB10, SBFP10, SMD^{+}13, SS10b, SK10, SRER16, ST08, SAMW14, SMF^{+}11, SNT^{+}16, TKW^{+}14, TPB^{+}16, Tho15, WDNM11, WRDF12, WHJ^{+}12, WGGM11, YTM^{+}12, ZVL^{+}16]. Temporally [ES12a, SNTN15]. temporary [DTC^{+}17, KJC^{+}15]. ten [MBF16, SCHO11b, SRK13]. tenuis [BBC12b]. TEP [BZBR^{+}11, KKO^{+}14, YYU^{+}17]. Terebridae [FTKK14]. teredinid [HED^{+}13]. teres [HHNI17]. term [ASEB11, AMLR^{+}17, BMSB13, BBM13, BNAN12, BDP10, BD13, BD14, BDDvdM17, BBC^{+}12a, BCK15, BBMM14, CR16, CBKM16, CL15b, DDP19, EPL^{+}14, ENO^{+}11, FKZTT18, FHG^{+}14, FEE^{+}11, Gal14a, Gal14b, GNJJ^{+}17, GTL^{+}16, HGL15, HAM^{+}15, HG13, HI14, HA14, HBG^{+}17, ITT14, JDG^{+}17, KV13, LME^{+}19, LH14, LBJK14, LBL^{+}12, LPK^{+}16, LFD13, ŁCM^{+}18, MPA^{+}10, MHAP18, MUOS13, MBKM17, MVOPB^{+}16, MOK^{+}18, NDAP19, NCC^{+}14, ORE^{+}17, OFPR^{+}11, OKLM14, PFB^{+}13, PBC^{+}18, PPM11, PRA^{+}19, QHN^{+}10, RU11, RRL^{+}19, RMRA15, SGC^{+}13, SKD^{+}11a, SKB15, SLKN14, SJL^{+}16, SKFH18, SW12, SRK11, SBLM14, SSvD^{+}15, SS14, SAD^{+}16, TSA^{+}18, TS11b, WAM^{+}15, YYO14, dJdBdJB12, dGR11, dlSGS17, vDGK^{+}12, vdVKA^{+}11]. terminal [DMMW12]. termination [LWD13]. terms [RP15]. tern [MADI13, MPG^{+}10]. terns [Goy14, UW18]. terrapin [RLBC10, HW12a]. terrapins [HW12a]. terrestrial [BBI16, BWP^{+}11, BJL^{+}17, CHP^{+}11, CCKB^{+}15, DMC^{+}16, FESPR14, LAM11, QOGS^{+}16, RG11, VAHL11, WLP^{+}11]. terrestrialisation [CSG11]. Territorial [Goc10, ABH18, CSB14, EMJ18, VHKT15]. territoriality [AZVB13]. Territories [SWM^{+}19]. Tesseropora [HRAM13]. test [DL14, FZI^{+}16, KKMA^{+}10, KML13, MPZ10, MSE^{+}18, MLB^{+}14, SMSL17, SCF^{+}17]. Testing [RWP^{+}14, RLS10, SVG16, SBCA15, SBKA14, WBG^{+}14, ALSV16, TLS^{+}19]. tests [BWJB16, MCUO12]. testudinum [BHF^{+}10, BJF17, GD15, HFMD16, MMPH^{+}12]. Tetraodontidae [GKW10]. Tetrapygus [UHG12]. tetraspina [GK10]. Texas [PPM11, SMR10]. th [LPKT^{+}19]. Thalassarche [DMT^{+}14, PWM^{+}19]. Thalassia [BHF^{+}10, BJF17, GD15, HFMD16, MMPH^{+}12]. Thalassiosira [DTVH10, GRG^{+}12, MAB12b, PL15, ST14]. Thalassoma [Gea10]. Thalia [HEB^{+}11, PNA^{+}16]. thallus [BDT^{+}16]. their [BHF^{+}10, BB17, BLE^{+}13, BDP10, BAL^{+}11, CPW15, CRGM18, CME^{+}12, CAGS17, DBA17, DB14, DEWK15, FC13, FKA^{+}12, GNHLN11, GCCCM16, HR12, HMKF11, HKB^{+}19a, HYC^{+}14, HD10, HRK^{+}18, ILDL10, KSA12, KVR^{+}18, KWH^{+}18, KBC^{+}12, LHJH^{+}10, LZ12, LMd12b, MBB11, MCW^{+}16, McC19, MP17, NAY^{+}11, NKM^{+}15, NFN^{+}15, NTB^{+}17, OHB^{+}15, PDGR^{+}13, PDM12, PMK^{+}15, PPJT15, PR11, PMM^{+}15, PRM17, PB19, RdGv^{+}18, SS10a, SNTN15, ST13, SF16, SPM^{+}16, SLL^{+}18, SWWK13, SFMB10, SO15, SPSP10, SNNS^{+}12, TB12, TMT^{+}12, TDB^{+}14, UDF^{+}18, VBT^{+}17, WJWC16, WWB^{+}17, Wir12, YYD^{+}15, YKG^{+}15, YSH^{+}16, ZPPM17, dG14, dML11, dSM11b, vDGK^{+}12]. Thelenota [PPDW16]. them [VPDPMRC11]. Theme [Ano17a, Ano17c, Ano18a, Ano18b, BH12b, BLM^{+}12, CLPiU14, DMP12, DS13, HBS^{+}13, HGA^{+}16c, KBH11b, LPSL14b, PHB^{+}11a, SS15a, SK12]. Themisto [KPH^{+}15]. Theory [CFSM^{+}13, ABH18, AN11b, GBL^{+}15, HP11b, OMZ10a, SFL^{+}18]. Theragra [BBMH13, ZBC10]. there [GGB^{+}18]. Thermaikos [KAP^{+}12]. Thermal [BMT17, BDGB19, DP13, FRTB14, HB15a, LSBB17, PCF^{+}19, RAN^{+}10, RL18, SF18, SFNP11, BMSB13, CRGM18, CBJB19, DHM^{+}14, GB16, HRP18, HGC15, JPF12, KLT^{+}12, LHU^{+}18, LM12, LLH^{+}11, LKS12, LHW^{+}19, MVCD14, McC17, MWGH15, OVF11, PW11, PWM18, RKvW10, SMOH19, SMVC^{+}15, SKM^{+}15, SG18b, SSdW^{+}12, TRR^{+}17, TM15, TNS^{+}12, TFH^{+}19]. thermally [RAH^{+}10]. thermocline [HMS^{+}12, SSM^{+}19]. thermography [LS14]. thermoregulation [SMF^{+}12, ZF19]. These [Ano17c, DMC^{+}16]. thick [KTP^{+}18, SG12, YOWK17]. thick-billed [KTP^{+}18, SG12, YOWK17]. thickness [MRB^{+}11]. thin [BBRBG10, MJT^{+}12, NSG^{+}12, SMP10]. thin-layer [NSG^{+}12]. Thinning [FJW14]. third [DC17, TMM^{+}19]. third-largest [TMM^{+}19]. thompsoni [KPH^{+}15]. Thomsen [TW10]. thorns [ARD^{+}19, BBW^{+}19, KHH18, MPB^{+}14b, RPB17, WGLDB15]. thornyhead [DT11]. those [AKF14, GHGG11]. though [KPM^{+}11]. Thousand [SRK13]. thraustochytrid [NNJ^{+}13]. Thread [SS11a]. Thread-drifting [SS11a]. threadfins [HMW^{+}12]. threads [HMW^{+}12]. Threat [PT13, CGH^{+}16, LF14, PBC^{+}12b, RBL^{+}17]. threatened [BCN15, GRN^{+}16, LMLB12, MCW^{+}17, PBCA10]. threats [BVM^{+}19, DP14b, WHH12, dSdSdCO^{+}11]. Three [FQ11, JLPG17, PRLR^{+}18, PSB17, APNC^{+}15, AGHC10, ACG12, AR14, BAA12, BWS^{+}12, BSB^{+}11a, BBH^{+}13, BRC^{+}15, BPLT14, BBMM14, CBKM16, CR19, CE11, CWEE19, CL10b, CBF^{+}13, CMB^{+}14, CGBGS18, DBE^{+}12, DAC^{+}12, DDP13b, DTL^{+}15, FC10, FWS^{+}18, FZI^{+}16, GSS13, GMK^{+}19, HMAC^{+}11, HSWS16, JCZ^{+}12, JPF12, KCS^{+}17, LI18, LSH19a, LHDH17, LNP12, LL18, LLD^{+}12, MBW10, MRLD19, Miu12, MGB^{+}19, MCHC^{+}16, MTH15, OCSW16, PLM^{+}16, PBDMRB13, PR11, PW16, RCP^{+}16, RF14, RSP^{+}16, RT12, RQN^{+}14, REP18, RTC^{+}14, RBH12, SDG^{+}12, SMSE19, SSS19, SMF^{+}17, SML17, SM15, TRSB^{+}18, WDM^{+}19, WOHJ^{+}15, WTC^{+}15, WLF^{+}17, ZBP18, vDGK^{+}12]. three-decade [MGB^{+}19]. Three-dimensional [JLPG17, PRLR^{+}18, LSH19a, LHDH17, MRLD19, PLM^{+}16, TRSB^{+}18, WDM^{+}19]. three-spined [SML17]. Three-stage [PSB17]. thresher [CWK^{+}10, CWW^{+}16a]. Threshold [LE13, OMZ10b, CBMK10, CLL17, GWM^{+}15, GGB^{+}18, LDT^{+}13, NNT^{+}10, PW11]. Threshold-dependent [LE13]. Thresholds [DMR10, GMC^{+}12, OMW^{+}10, BBH^{+}11, CMR12, FL10, JG13, KHH18, LGS10, LM18, LTCH10, OMZ10a, PH10, RPB17]. throughout [GALS16, GSM^{+}11a, HNG^{+}11, PCLB18, TDB^{+}14, VKY^{+}12, WA14]. throughput [CR16]. Thukela [ASF13]. Thunnus [FCDC12, IS16, KIU^{+}13, KRR^{+}18, KHT^{+}17, LCB10, MIM^{+}12, MRC^{+}13, ODK^{+}14, RBA^{+}18, RTCF^{+}17, Sat10, SWY^{+}10, WHI^{+}18]. thus [Ano17c]. thynnus [FCDC12, FCDC12, LCB10, MRC^{+}13, RBA^{+}18]. Thysanoessa [DRH15, HS14a]. tick [MBB^{+}12a]. Tidal [CGKW10, CGK13b, CGH11, DRT13, WGD^{+}16, BBH^{+}11, BDvGW16, BD14, BD19, BSB14, BT13, Bur12, CKR^{+}11, CSC13b, CZH^{+}13, DPO^{+}17, DJL^{+}12, EHG16, EBT^{+}15, HG10, JBT^{+}16, KF13a, LA14, LHT^{+}17, MBM^{+}14, MGM^{+}18, MVOPB^{+}16, NOB12, OMD11, PBA12, PFR11, PMS10b, PMS13, RBD^{+}16, RMB18, SCL^{+}19, SSR^{+}10, SHB^{+}10a, SMSL17, SSB^{+}16, SO16, SFW^{+}16, TNM^{+}11b, UW18, VNN10, VDM^{+}10, VBR19, WD14, ZBY^{+}14]. tidal-stream [BDvGW16]. tidally [HR12, MMEM14]. tide [BDvGW16, DHJ10, FHL^{+}17, HED^{+}13, JP15, MK13, NCD^{+}12, PSPK10, SBS^{+}13, WLBB^{+}13, WWKS15]. tide-induced [FHL^{+}17]. tidepools [MGL^{+}16]. tides [ACF^{+}10, BCW12, DPH13, KF13b, RBL^{+}17, SMM^{+}14]. tidewater [HMAC^{+}11, JPAM^{+}15]. tied [GAF^{+}16]. tiger [GVR^{+}14, KCS12, NWP^{+}11]. Tigriopus [ABM12, HPL^{+}16, TB14a]. tilapia [DPC13]. tile [KOLDP17]. tilefish [PFVFF^{+}15]. tiling [SCTA10]. Time [ETD^{+}18, GSD12, WEAG^{+}15, BZC^{+}12, BAA12, BRT^{+}17, BMT17, BNW14, CR16, CLC^{+}12b, DLC^{+}13, EHR12, EAB^{+}19, EM11, EKOÅ13, HMW^{+}12, HSA^{+}15, HFH10, KDJP10, KDO^{+}14, LPI^{+}14, LAH^{+}15, LCB^{+}16, MPG^{+}12, NGA18, Pro17, RSTV^{+}13, SHT^{+}16, Sma13, SBKA14, SS14, SBS10, TRSV^{+}12, TS11b, VKY^{+}12]. time-activity [CLC^{+}12b]. time-based [LCB^{+}16]. time-depth [EKOÅ13]. Time-in-area [WEAG^{+}15]. time-lag [MPG^{+}12]. time-series [EAB^{+}19, NGA18, SS14, SBS10]. time/repeat [HFH10]. timelines [MBKM17]. times [ATP^{+}12, NTB^{+}17, PGR^{+}13, PTF^{+}14, TP19]. timescale [RBM14]. Timing [HRK^{+}18, KEBB10, VHP^{+}11, Ano18a, BMTF10, CMW^{+}14, DED^{+}15, FHL^{+}17, FYG^{+}18, GAB^{+}14, HTHM18, HPR^{+}12, JWB^{+}17, KCCG12, KCR^{+}10, LANER15, LBWB^{+}13, MWM^{+}11, Ric12, RG14, SDKD13, SCAM^{+}14, SE16, TPT^{+}12]. tipping [PEB^{+}14, SLS18]. Tissue [ACG12, MMRDS18, BCM^{+}14, CJM^{+}13, EMH^{+}19, HB12, HCA11, LSBB17, RMP^{+}19, SMF^{+}17, SSM15, SMBC18, VMC^{+}19]. Tissue-specific [ACG12, CJM^{+}13]. tissues [AFS19, MMWG13b, PCW^{+}12, TGM^{+}17, WA14]. Title [Ano10a, Ano10b, Ano10c, Ano12a, Ano12b, Ano13a, Ano13b, Ano14a]. TMB [AMAJ^{+}17]. Todarodes [PVYS16]. together [OWH^{+}19]. Tokyo [NKH^{+}10]. Tolerance [AAvS^{+}13, FRRÁS^{+}12, RZS12, SSL^{+}14, AZSKWL11, DP13, DBCP15, FRTB14, JG13, JPF12, MMB^{+}12b, OHP18, PCF^{+}19, RAN^{+}10, RCF17, SPMN19, SFNP11, SHI^{+}10, TM15]. tolerances [CL17, CRGM18, LGD10]. tolerant [LKN^{+}13, RAH^{+}10]. Tomé [HVJ^{+}18]. Tonga [PML^{+}10]. tonne [SBB^{+}14]. tonsa [ACDJ11, BBRBG10, ET11, LME^{+}19, PFPP18]. Too [FC19, HCG15]. tool [BHD^{+}13, BBC14, CWR^{+}18, CBR19, CGG^{+}10b, DNO17, HLC^{+}18, LC12, PLF^{+}12a, RCF17, TLB15, TLK^{+}12]. toolkit [BCW19]. tools [BFD^{+}16, CB12, LKL^{+}17, SSFB11]. toothed [SFM^{+}19]. toothfish [TGR^{+}19]. Top [BW10, BAW11, CGQB10, KMP15, MTV^{+}11, PJMP10, AH13, BHP12, BMTC^{+}15, CMV^{+}11, CWW^{+}16b, CWGG15, FRV^{+}13, GCC^{+}14, HBK^{+}17, HHS^{+}12b, HBG^{+}17, HOE^{+}12, KÖHSO12, KBW^{+}11, LGMB^{+}16, LLBP15, MH12a, MRB^{+}15, RCS^{+}17, RDK14, RSK^{+}14, SV13, SSR^{+}10, WHT12]. Top-down [BW10, BAW11, MTV^{+}11, PJMP10, CWGG15, HOE^{+}12, LGMB^{+}16, RDK14, RSK^{+}14, WHT12]. Top-predator [CGQB10]. Topics [Ano18b]. Topographic [MSG11, WBLRN18, vdMPM12]. topography [vdMPRM12]. topsoil [MDJN19]. toreuma [DHGW15]. Toro [AHB^{+}14]. Torres [CSS^{+}13]. torreyi [BBWG12]. Tortugas [FA11, HSF^{+}12]. total [BTS^{+}14]. toughness [BDT^{+}16]. tourism [BM10b]. towed [DIG^{+}11]. toxic [BSG^{+}10, BFL^{+}11, EOA15, SP13a, SBS^{+}13]. toxicant [AS13]. Toxicity [RTC^{+}14, WHT12, AEM^{+}18, GG16, LWD13, OHP18]. Toxigenicity [GHT^{+}15a]. toxin [ASAG17, NKH12, SE11a]. toxin-producing [SE11a]. toxins [RHB13, TLM^{+}11, VTZ^{+}13]. Trace [LCC^{+}11, MMWG13b, CAB^{+}19a, HBBH12, JWHR11, KRC^{+}11, LMR^{+}10, LCBK^{+}13, MGO^{+}12, MMWG13a, MCL^{+}17, NDBL16, PBWM18, RMP^{+}19]. trace-element [MGO^{+}12]. traced [HU13]. tracer [BC15, CPEE18, GM13b, KBA13, MSF18, SE11a, VSM11]. tracer- [MSF18]. tracers [GMT^{+}15, KS12a, MCW^{+}16]. Trachurus [BNAK11]. Tracing [PPD^{+}14, DMR10]. track [BBBF15, BSBR^{+}12, MAB^{+}12a, SCB^{+}14, VGGD^{+}15]. tracked [HSF^{+}12, QHN^{+}10, TMS14]. Tracking [CCGP14, KPW^{+}19, LSJ^{+}15, WWAL13, BSB^{+}18, BMV^{+}11, BBW^{+}17, BBC^{+}17, DKF^{+}15, DBHMG12a, FLK12, HTR^{+}12, HMB^{+}12, HBM^{+}12, LBJI14, MASW^{+}11, MWR^{+}12, MHS15, SHD^{+}19, ZBB^{+}11, dSdSdCO^{+}11]. TrackIt [LNS10]. tract [BAvW12]. tracts [HAHM16]. Trade [RT17, SPP12, SS12a, dDTA^{+}18, BJM19, BL17, FSFS18, GRN^{+}16, HKJ17, KWM12, LP10, RPRN^{+}16, RvOP12, SMHB14, SGGL15, VQSH14, dMA18, vDCFM10]. trade-off [FSFS18, KWM12, LP10, SMHB14, vDCFM10]. Trade-offs [RT17, SPP12, SS12a, dDTA^{+}18, BJM19, BL17, HKJ17, RPRN^{+}16, RvOP12, SGGL15, VQSH14, dMA18]. tradeoff [LSBB17]. tradeoffs [SH10]. trading [FBC^{+}15]. traditional [Ji11]. traffic [CAB^{+}16]. Trait [LMG12, MGL^{+}16, BPB18, BCB^{+}17, MGOL19, OBÁSV18, PMS10a, VDM^{+}10]. Trait-mediated [MGL^{+}16, PMS10a]. Traits [BC16, MBR^{+}18, AC11, AGV^{+}11, BVO16, BB17, BHT15, BDPL19, BOÁNF19, BS14b, BEFC15, BPG^{+}10, BVQ12, BFPONS^{+}12, CFMFFJ18, DQC^{+}13, DVB^{+}16, DAWS19, GMG^{+}16, GH11, HWB^{+}18, HDP^{+}17, HDT^{+}12, KPW^{+}13, LAB10, LEKT12, LALÁC11, LPJ^{+}13, LLF^{+}18, MC11, MJO^{+}13, MJT^{+}14, MM12, MMA18, MDS14, NWS^{+}14, NPLT10, PBA12, RPRN^{+}16, RMSPA16, SPM^{+}13, Sat17, SS12a, SMHB14, SS16, VJJ^{+}19, WAA^{+}19, ZBP18, dlSBO^{+}12, dlSOV^{+}16]. trajectories [MRC^{+}12, PRK^{+}12]. trampling [BAVSP11, LCD11, MLB18, MLB19, MF13]. Trans [CNSW16, HMO^{+}12, HRC12, MC16, SWR11]. Trans-Atlantic [HRC12]. Trans-equatorial [HMO^{+}12, MC16]. trans-oceanic [SWR11]. Trans-Tasman [CNSW16]. transatlantic [LTP13]. transceivers [LBJI14]. transcript [MBB^{+}12b]. transcriptional [MCBR18]. transcriptome [SKM^{+}15]. Transcriptomic [DSVM10]. transect [BPN11, BLM^{+}18, CPTL11, JTH^{+}12, KRDV18, PMM^{+}11]. transfer [ALV12, lLChTC11, CLG^{+}19b, DAL^{+}10, DVSM10, HHH11a, LKCH12, PBWM18, RET^{+}19, RdGv^{+}18, SL15b, SH14, vdMHW^{+}10]. transfers [GM13b]. transformation [THC12, VAHL11]. transformations [VGBJ^{+}13]. transient [BLMD^{+}11, BJG^{+}11, CMK^{+}14, SP15]. Transiting [VGR^{+}16]. transition [AFRAG16, CAB^{+}19a, CMM^{+}18, DAB17, EZZ^{+}12, HB13, HFR^{+}18, HM09, HAF^{+}14, OV10, YHT16, ZTNS13, vdMPRM12]. Transitions [NMC^{+}16, BKM^{+}18, DZL16, EOMS15, HPQ15, LBP^{+}16, PZA^{+}13]. translate [EM13]. translocation [Fox16, KBA13]. transmission [HMP15, SHMOS17]. transmitting [MAW^{+}11]. transmontanus [ZBK^{+}14]. transoceanic [BPB^{+}12]. Transparent [BZBR^{+}11, YYU^{+}17, KCP^{+}11, KKO^{+}14, MSC13, PPB^{+}10a]. transplant [BSCH14, Gar13, MAT^{+}11, dMBC12]. Transplantation [Kat16]. transplanted [NAY^{+}11]. transponder [RBD^{+}14]. Transport [JGG^{+}15, LA14, SMM^{+}15, BSB14, BSB12, BCSB13, BHMCM16, CZH^{+}13, FRP^{+}17, GBK^{+}14, HSM13, JLQ^{+}14, KHD^{+}11, KML13, LMd12b, MM13c, MMWG13b, MKI^{+}15, MLP^{+}14, OLD^{+}15, OFFQGG10, PCD^{+}18, RWR^{+}19, RS14, Ree14, SMM^{+}14, SMSL17, SE11b, WBLRN18, WZP^{+}14]. transport-retention [BCSB13]. transportation [GZB16, GB10a]. transversus [CMO^{+}10]. trap [BBM^{+}14, BGB^{+}17, GBD13, Joh10]. Trapezia [RLM^{+}14]. traps [BSB^{+}18, ITT14, RSP17]. travel [NKM^{+}15]. traversing [MMB^{+}19]. trawl [KSA12, KW17, LKC^{+}15, Løk11, MVVGAV10, MKPS13, MVE^{+}10, MPB15b, POH16, RZPH14]. trawl-mounted [KW17]. trawlable [CH12]. Trawling [LLMA^{+}10, AFH11, ASF13, DT11, HWB^{+}18, SGJ^{+}18, SBLM14, TVA^{+}19, vDBH^{+}15]. treatment [MFDW17]. Tree [ETB13, HYL^{+}13, RFRD10]. trees [CMLC12, WFD19]. Trematode [SDT15, TMBP13, ADRA19, SP12, SWPK13, dMBC12]. Tremblay [Ano12d]. Tremblay-Boyer [Ano12d]. trenches [FKRJ13]. trend [DMPG11]. Trends [HFN14, RPM^{+}13b, RQN^{+}14, TS15, APC14a, BHE^{+}17, BBBHC17, BLM10, BDB^{+}12, FTB14, FKvdV12, HGL15, HAM^{+}15, JTH^{+}12, LGB17, MRE14, MHAP18, OGGP14, PPM11, RRL^{+}19, RPP^{+}13, RSP^{+}16, RGDS12, SSB10, SMD^{+}13, SST19, SBLM14, TBG^{+}17, TSTP13a, TBMdV12, WHJ^{+}12, WPB^{+}14, ZAC11, dQSC14, vdVKA^{+}11]. trevally [DFD^{+}19]. tri [BGDC14]. tri-axial [BGDC14]. triacylglycerol [RPLT10]. triage [PMB^{+}15]. Triakis [NCK^{+}14]. trials [SV15]. Triangle [APC^{+}13, BCH^{+}11]. tributary [BT13, LBAK13, PS19, SFW^{+}16]. Trichodesmium [lLChTC11, OGA13, SBS^{+}13]. Trichodesmium-derived [SBS^{+}13]. Tridacna [GBMW12]. Tridacnidae [HAF^{+}17]. tridactyla [BER17, HCDR^{+}15, MIA^{+}11, REAN^{+}14, SMC^{+}14]. triggerfish [BSB^{+}18, BFN^{+}19]. triggers [GW15, GPS13]. trip [CLC^{+}12b, LMCD19]. Tripneustes [BJM19, JMB10]. Tritia [HA18]. tritrophic [DL14]. Trophic [ALV12, ABN14, ABJ^{+}15, APBC^{+}19, ALSO^{+}18, BYAA13, BPM^{+}18a, BS12, BPH^{+}19, BADQ^{+}15, BGIA19, BKC^{+}19, BL13, CL10a, CPS^{+}14, CGM^{+}19, CLN^{+}13, DKAC^{+}12, EPRGO13, FPC^{+}13, FBGH18, FTKK14, FPN17, GSP^{+}15, GLR^{+}14, HBS17, HHH^{+}11b, HM13, KKS^{+}14, KMP14, KRD^{+}12, LRB^{+}18, LJ10, LFPC^{+}14, LLC^{+}14, LBS^{+}15, MBC18, MTS^{+}17, MWWS17, NT10, OFDG14, PDBG14, PBC^{+}12a, PLF^{+}12b, RCF16, RSK^{+}14, RFJ^{+}10, RGDNGS11, RBBG12, REP18, SCB^{+}14, SL15a, SL15b, SDF^{+}19, SRA^{+}17, TFM^{+}18, VAB^{+}13, YKBB12, APNC^{+}15, AMR^{+}10, AGV^{+}11, AS15, AH13, ACG12, AGRPSVL13, BW16b, BW16a, BMM17, BCN15, BLV^{+}19, BEMD11, BHTF11, BDB^{+}12, CB14, CBMK10, CLG^{+}19b, CWW^{+}16b, CAG13, CWGG15, CMB^{+}14, CCKB^{+}15, CV11, CNÁCP12, DGC18, DDP19, DRDV15, ERG^{+}11, EFOM18, EMH^{+}19, EPF^{+}17, FC10, FRGI15, GMT^{+}15]. trophic [GKHM13, GCS^{+}19, GPC^{+}15, GHNF10, HCL^{+}11a, HBK^{+}17, HVK^{+}13, HKB^{+}19b, HB12, JW11, JCRC10, JGS14, JSLG17, KTR^{+}16, KCH^{+}11, KPH^{+}15, LFU^{+}19, LCQT^{+}13, LCGS^{+}19, LSV^{+}10a, LMW^{+}13, LA12a, LRL^{+}13, LHB^{+}15, LCL^{+}17, LRO^{+}17, MBvR12, MHB14, MMPH^{+}12, McC14, MBRO10, MPB13, MHS14, MW16b, MJT^{+}12, MD16a, MJV^{+}11, MKG10, MBT14, MSE^{+}12, NLK^{+}19, OE11, PLG^{+}19, PDL^{+}10, PTP^{+}19, PBCA10, PBC11, PSS19, PHM11, PKBS14, PBWM18, PMBH11, PLZL14, QLCUG15, QGM^{+}14, RB11, RGW17, RWMP12, RHE14, RSDN15, RDGC17, RK15, RLNLAF17, RVH^{+}19, SATS13, SKB15, SOJ14, SST19, SESS^{+}18, SSM^{+}16b, SBM^{+}18, SFL^{+}18, SCW^{+}15b, SKS18, SLRS13, TL18, TBS^{+}14, TTS17, UDF^{+}18, VGGF12, VWEP18, VBT^{+}17, WRP^{+}17, WVD^{+}17, WK18, WS11, ZS17]. Trophic-level [PLF^{+}12b]. Trophodynamic [SA17, FGL^{+}12]. Trophodynamics [AR14, BI15, CMW^{+}15, LPSL14a, LPSL14b, NIS14, SLF^{+}19, AB12]. trophy [BS16]. trophy’-sized [BS16]. Tropical [GFGGDSM12, KFGSMPG14, APH^{+}10, AAGGFG^{+}12, ACAD12, AP14, ADA^{+}19, BÁNC^{+}18, BHP12, BCGWS19, BFNB16, BMW^{+}14, BDB^{+}13, CWD11, CGKW10, CLJF12, CMO^{+}10, CO16, CRGM18, CHLR16, DZB^{+}18, DGQ10, DCW^{+}14, EC14, FRV^{+}13, dSKFV12, FTF15, FA10b, GEAOH17, GKBEF18, GBJ^{+}14, GCCCM16, HBS17, HHS12c, HW14, HT15, HOE^{+}12, HIMH15, IMD11, KWPL11, KSBM^{+}11, KHS^{+}11, KKO^{+}14, KRC^{+}11, LTN19, LFMC11, LLBT18, LHC^{+}15, LH19, LBMZ15, LEB^{+}19, MHB14, MH19, MA14, McC14, MIIB16, MC16, MWC18, MHFN10, MCP^{+}18, MS14, MGD^{+}11, OS18, ODK^{+}14, OWW^{+}17, PDGL17, PA15, PMGQM17, PFVFF^{+}15, PJDC^{+}13, PJP^{+}11, PHTS11, PML^{+}19, RU11, RCF17, SDMB15, SØA^{+}13, SØK^{+}15, SCO^{+}12b, SBFP10, SJC10, STRP18, SAD^{+}16, SNS12, TNIH12, TPCN12, UCK^{+}13, VB19a, VM10, WRGP12, WvLF18, YMD^{+}10, ZMGS17, dlSOV^{+}16, vdGSE^{+}14]. Tropical [CGK13b, GGG18, PBC^{+}12b]. Tropicalization [HAF^{+}14, HFM^{+}15]. tropically [HFM^{+}15]. tropicbirds [DZB^{+}18]. trossulus [JG13, LIT13, RPLT10, TJS17, TTS17]. Trough [YWiIK15, NWS^{+}14]. trout [ABK^{+}14, ABT^{+}15, AHØ^{+}15, BMJ16, BK14, KRdVG^{+}18, KBGA19, MTS^{+}17, MLB^{+}14, MDAPH11, áNGGS11, OHA10]. trumps [TMS^{+}16]. truncate [PSER17]. truncatus [EC14, SKGS11, SHT^{+}13, TNN11]. Trunk [TPA^{+}13]. truth [SKD^{+}11a]. trutta [ABK^{+}14, ABT^{+}15, AHØ^{+}15, KRdVG^{+}18, KBGA19, OHA10]. tshawytscha [JGM^{+}16, MGM10, WSF^{+}12, WWW^{+}13]. Tubastraea [MTK^{+}18, RTJ^{+}13, SPSG14]. tube [Ber12b, Hug11]. tube-building [Ber12b]. tubes [LRS^{+}12]. tubeworm [DRG^{+}12, LKL^{+}19]. tubeworm-associated [DRG^{+}12]. tug [SLRM^{+}15]. Tumbling [NLV^{+}15]. tuna [BBBF15, BHN^{+}14, BBBW17, DFAH11, ECC^{+}15, FCDC12, GSC17, GRL^{+}15, HIMH15, IS16, KIU^{+}13, KF17, KRR^{+}18, KHT^{+}17, LCQT^{+}13, LCGS^{+}19, LCB10, MCW^{+}16, MGAV15, MCP^{+}18, MRC^{+}13, Nor13, ODK^{+}14, RUPT^{+}11, RCAB^{+}12, RTÁB^{+}14, RBA^{+}18, RDF^{+}13, RTCF^{+}17, RAF^{+}14, Sat10, SCO^{+}12b, SWY^{+}10, WHI^{+}18, WRI12]. tunas [HOE^{+}12]. tuned [ZBC10]. Tunicata [GKE^{+}19]. tunicate [LALÁC11]. tunicates [LD17b]. Tunisia [FCR12, HS15a]. tunnels [HED^{+}13]. turbid [BPLT14, HHH^{+}11c]. Turbidity [GPS13, BA13b, HPMZ11, LKCH12, LMT15, MSC13, MB18, MSPO12, SML17]. Turbinaria [BDT^{+}16]. turbine [MGM^{+}18]. turbot [PWO^{+}16, VBV^{+}10]. turbulence [CZH^{+}13, GSA^{+}18, LGR^{+}10, PPB^{+}10a, RPM12, RMR12, SHMB15, TNOY17]. Turbulent [BDS18a, SPWK13, WHA^{+}13, WLH^{+}17]. Turf [OS18, VHKT15, BDPD14, BT14, BHT15, BC13, BOS18, CAB^{+}14, HLS15a, LCC12, TCA^{+}11]. Turf-forming [VHKT15, OS18]. turfs [ATS14, AS15, ASM10, BMMS10, CFA14]. turgida [POKGRP10]. turn [JP15]. turnover [ACG12, BPSG11, BNE^{+}19, DEJN10, DPH14, MNKB11, OKJMC15, PSG^{+}13, PCNM17, SGD^{+}12, VVB19, WKE^{+}17]. Tursiops [EC14, SKGS11, SHT^{+}13, TLK^{+}12, TNN11, ZPPM17]. turtle [AGP^{+}13, BMR17, BBC19, DD15, FMC^{+}10, FGH^{+}13, GDB^{+}16, GDJdBSL19, GDA^{+}18, KAMA13, KCL^{+}11, LI18, LPC^{+}11a, LNH^{+}17, LCBK^{+}13, MO13, OMW^{+}16, PDvG16, PMB^{+}17, PvdMLC17, PRK^{+}12, RMP^{+}19, RPE^{+}13, RAKB^{+}12, SMBH15, SBB^{+}12a, SJL^{+}16, TGH15, TBH10, VAB^{+}13, WBB^{+}10, WFS10, WHK^{+}11, WCP^{+}14, dQSC13, vdMHW^{+}10]. turtle-derived [LNH^{+}17]. turtlegrass [BHF^{+}10]. turtles [ALE^{+}13b, AJ16, AGH^{+}12, BBRM14, BBC^{+}13, BDM^{+}13, BLD^{+}16a, BP12a, BBC^{+}17, BHTF11, CWR^{+}18, CC19, CCE^{+}14, CM14, CdTK^{+}16, CPF^{+}16, CCC^{+}18, CLS^{+}16, DP14b, ECM^{+}12, FCRD18, FA10b, GLW^{+}12, GLJ^{+}18, GBD^{+}12, GCSF18, HAM^{+}15, HNK^{+}18, HVJ^{+}18, HSF^{+}12, HTR^{+}12, HFG^{+}19, Hv14, HKB^{+}19b, HSMB17, HRS^{+}16, JBL^{+}16, LABM17, LSFQ19, LKL^{+}17, LMSC^{+}19, MLM^{+}12, MASW^{+}11, MWR^{+}12, MLML^{+}19, OFG^{+}18, PBAS^{+}10, PPM^{+}15, PMGQM17, PVZD^{+}11, PFB^{+}17, PDT^{+}11, PK15, PSAK12, PAC^{+}15, RAB^{+}10, RTDUM16, RMNP17, SMH^{+}12, SMVC^{+}15, SPHK17, SL11b, SWH^{+}18, SSB^{+}11, TMM^{+}19, THB^{+}12, TRGF18, TSA^{+}18, TMS14, WHB^{+}19, WHFM17, WTP^{+}18, WFH^{+}18, WHH12, ZBB^{+}11, dSdSdCO^{+}11, dQSC14]. Tuscan [FRC^{+}13b]. Twilight [MLP19]. two [ADvdM13, ACS12, AHPKS17, BA13a, BGJ^{+}19, BCvH^{+}15, BRBH19, BJH13, BN16, BADH^{+}12, BADQ^{+}15, BOÁNF19, BRA^{+}10, BCM11b, BCB11, CMC^{+}16, CPPM12, CRG^{+}16, CCG^{+}10, CLW^{+}11, CY13, CLGC16, Col10b, CALB10, CS10a, CNÁCP12, CQR^{+}17, CBYH13, CLN^{+}13, DMMW12, DKAC^{+}12, DDDR10, DMC^{+}17, DB11a, DZB^{+}18, DVB^{+}16, DAM^{+}18, EVB^{+}16, EBB10, FHL^{+}17, FG10, FUOG^{+}16, FAT^{+}15, GFKC13, GMB^{+}16, GP11a, GHCC15, GGvW^{+}15, GTL^{+}16, HPB^{+}18, HPL10, HYC^{+}14, HGKB14, JD18, JDD^{+}13, JSLG17, KHS10b, LHT^{+}16, LHDH17, LCL^{+}11, LMG12, LGJH12, LCFR10, LM18, MdSW^{+}16, MTS^{+}17, MCBR18, MWZ^{+}16, MC12b, MSPO12, MPRPM^{+}18, MMB^{+}12b, MKII19, NFSW10, OCE^{+}12, PLM^{+}16, PVZD^{+}11, PDC^{+}17, PRLR^{+}18, PSS^{+}12, PJDC^{+}13, PHAZ18, PFL13, PBCA10, PJMP10, PRB^{+}17, RLO18, RDC^{+}12, RSE18, RPR16, RAB^{+}15, RNSA15, SKD17a, SDM^{+}13, SNTN15]. two [SNS^{+}13, SGL^{+}17, SSM^{+}16a, SLF^{+}19, SGT12, SKM^{+}15, SS16, SMR^{+}11b, SSM^{+}19, SHM10b, SBW16, TFÁS^{+}12, TRHM14, TDB^{+}14, TMP^{+}11, TBC^{+}17, TNSN11, TPFP11, VCRUQ15, VCQM^{+}16, VBT^{+}17, WFH11, WS11, WBB13, YRM^{+}12, YRL^{+}10, ZGK^{+}19, dSM11b, vdHTT^{+}15, HSWS16]. two- [LHDH17, PLM^{+}16]. two-decade [CQR^{+}17]. two-prey [LGJH12]. Tylos [MHPvdH19]. type [BSG^{+}17, CMW^{+}14, DM11, HNG^{+}11, OWN^{+}15, PBIO^{+}18, PHWC13, RH12, RT12, Wir12, WPK10, WK16, WLM^{+}13]. types [BRA^{+}10, GMBL11, LJB14, LNP12, MFG16, SLD^{+}13, SS16, TPT^{+}12, YSS^{+}18, ZFRG12]. typhoon [SYX^{+}10, CGH12]. typhoons [GGvW^{+}15]. typus [MHW^{+}12, RCR^{+}13]. tyrannus [AHH^{+}11, LBCL10]. Tyrrhenian [MKPS13, MG12].

UAVs [KMGH16, SHP^{+}18]. Ubiquitous [BRH13, CL16, LBM^{+}18]. Uca [RWR^{+}15, TB10]. UK [FRJ10, FRJ11, NMHB16, SBE^{+}16, WHJ^{+}12]. ultimate [LTP13]. ultraplankton [CL10a]. Ultrasonic [SFKQ15, MR16]. ultraviolet [HPL^{+}16, MWC^{+}12, OCMT13]. Ulva [BKC14, CLG^{+}19b, DLL^{+}18a, GGMTGC18, OS12, PCNB16, WWO^{+}15]. Unanticipated [Bea12]. unavoidable [DEGF^{+}10]. Uncertain [GAB11]. uncertainty [PGVF14]. uncommon [BCN15, DH16]. unconsolidated [SML^{+}15]. uncoupling [PMML12]. Uncovering [KPH^{+}15]. uncovers [MHS14]. Undaria [SJHH^{+}17, TS12]. undatum [HLC^{+}18]. Under-ice [GLS^{+}15, DLRF15, MTMM^{+}15]. under-investigated [DH16]. undercount [KWC^{+}17]. underdetermined [Fry13a, Fry13b]. underestimation [KSA12]. undergoing [ABW14]. underlying [CZB10, PG15, SNT^{+}18]. underscores [CSM^{+}19]. understand [HFG^{+}19]. Understanding [CLB^{+}14, KM10a, TGRH17, VHG^{+}18, Ano17c, Ano18b, BMTC^{+}15, CB12, GLW^{+}17, NHAS15, PEB^{+}14, RHB13, TF12, WBG^{+}14, vdMPRM12]. understorey [CPHH15, SLT^{+}19]. Understory [PMSGBL16, TW16, FJW14]. Underwater [CD14, KWC^{+}17, TRJ^{+}11, AvBBL^{+}16, BGB^{+}17, BG12, CVW^{+}11, CBL^{+}10b, CS10a, DOH^{+}17, FBPCC^{+}19, HBM^{+}15, KB12, LYH^{+}17, RST^{+}10, RSJ14, RKFD15, SPHK17, WFBM14, WLM^{+}13]. Unexpected [KCLLL13, MHHB^{+}17, MHHF13, PBS10]. Unexplained [RUS^{+}13]. unfed [LHW^{+}19]. unfished [HS17]. unicolor [BHN^{+}14]. uniformity [BB14a, Bar14]. unifying [Pör12]. Unimak [BLMD^{+}11]. Unimodal [LRB^{+}12b, DP11]. uninhabited [HSPS14]. uniparental [BRH13]. unique [Kin10a, PBCA10, SCMSB^{+}11]. unit [KB12, MS10, SMH^{+}12]. United [BBB^{+}14, Edm14, PPB^{+}17]. units [BPZ^{+}14, CM14, PRP^{+}18]. universal [GGB^{+}18, SMSL17]. universally [MDS14]. unmanned [KMGH16, SHP^{+}18]. Unprecedented [RSSvV19]. unpredictable [GD15, MWMD17]. unprofitable [HG10]. unprotected [WFBM14]. unravel [MMB^{+}12a]. Unravelling [BCN15]. Unrecognized [GKBEF18]. unsteady [WADKT13]. unstructured [HS16]. unsupervised [LPC19]. Untangling [PCD^{+}18, RCU^{+}11, MAS10a]. Untapping [DCIG18]. unusually [HCJ^{+}18]. unvegetated [BH16a, GDPR18, SDM16]. uplift [CMC10a]. Upogebia [PPS^{+}13]. upon [BGO^{+}18, HLW^{+}16, KB14, NWS^{+}14]. upper [CGR^{+}15, GK10, KK17, MS13, SGMG^{+}14, SRRS12, TTPM12, WLB^{+}11]. upside [JWR^{+}10]. upside-down [JWR^{+}10]. upstream [MKII19]. Uptake [ROE17, WADKT13, AC11, CMP^{+}13, Dao13, DTVH10, FHS^{+}13, HMS^{+}12, KTK^{+}12, MSS^{+}14a, MKMZ11, PSG11, RAB^{+}15, SMHB14, SSX^{+}17, SYK^{+}13, VBM^{+}11, VBM^{+}13, VBJR^{+}13]. Upward [WHA^{+}13]. upwelled [dGCF^{+}17, WWG^{+}13b]. Upwelling [CMC10b, PBW^{+}11, AFSL^{+}15, APC14a, AAGGFG^{+}12, BSK^{+}11, BdSC^{+}14, CLH^{+}15, CPS^{+}14, DAD^{+}11, DPO15, FHL^{+}17, FW11, FF12, FPM14, FAZ^{+}14, GMP^{+}11, HEMJ18, HWE^{+}12, HRPBLT18, JDQS11, LLCM^{+}12, LGC10, MADI13, MM13c, MF10, MWM^{+}11, MvdHBM12, NIS14, OBÁSV18, PLFR^{+}11, PPEH^{+}14, RM13, RHMR18, REG^{+}11, RUD^{+}16, RBEGG13, RCG^{+}15, RGB^{+}10, SUPCB11, SHS^{+}17, SBBB^{+}18, SPH^{+}17, SMM^{+}14, SML^{+}14, SSAB^{+}15, SOBNL13, THMF^{+}15, THRBL^{+}16, TSF^{+}10, VGBJ^{+}13, VLM14, WSS^{+}16, WRFM16, vdSRT^{+}18]. upwelling-relaxation [SMM^{+}14]. urban [AGK13, MHAP18, MCT15]. urbanisation [BGO^{+}18]. urbanization [CHFS12, WRW^{+}16a]. urbanized [MLM^{+}12, SLR^{+}14]. urchin [BJM19, BRIC13, CT10, CLMM^{+}14, Ebe10, EHR12, ESW14, FNC13, FNC^{+}14, FS14, FDS14, FJL12, FDP^{+}16, FG16b, GFP19, HCG^{+}10, IHP10, JH14, KB14, ML11, MBvO^{+}18, MvdHBM12, OPSM13, PAFC13, PPC^{+}12, PSC13b, PCW^{+}12, PHWC13, RLG14, RP17, SV12, SBW16, TKB^{+}13, UHG12, WW15, ZVL^{+}16, ZVR18]. urchins [CBKM16, CRGM18, CD16b, DAM^{+}18, ESW14, FSLG12, FFS14, JH11, JBD10, KKS12a, LH16, LS19, MMPH^{+}12, Mat10a, MM13b, MB11, OTS^{+}10, OPSM13, SFLG10, WTGC^{+}11]. urea [SMF^{+}17, SSX^{+}17]. Uria [ERB^{+}13b, ERB^{+}13a, TMH^{+}13]. uriae [MBB^{+}12a]. Uronema [SP13a]. ursinus [LJM^{+}10, SBA12, ZO10]. Ursus [BDR19]. Uruguay [CD16a, GOM^{+}19, MRGG^{+}13]. USA [BS16, GKR^{+}15, HCL^{+}11a, HSPS14, KCMH^{+}17, SMSB19, SQS^{+}13, SQRDLC18, TS16, WRB^{+}15, AMM^{+}12, ARA^{+}10, ALS18, BLD^{+}16a, BRB11a, BD11b, BWRP14, BJS^{+}16, CR14, CWW^{+}16a, CH12, CRWG16, FNF^{+}18, FFS^{+}15, FK10, GBH12, GMSC14, Gre10, HFMD16, HSF^{+}12, HOA14, JOFP18, KO12, KBB^{+}10a, LB15a, LRH14, MBKM17, MMM^{+}19, MCW11, NCS^{+}10, NCK^{+}14, OBA^{+}18, OGCHS18, PRR^{+}16, Ran18, RPE^{+}13, RHH^{+}15, RDF^{+}17, RLO18, RCD^{+}13, RLH^{+}13, RBS16, SNVH19, SS11a, SHM^{+}12, SIK16, SRK13, SSX^{+}17, SSD13, SLR^{+}14, TT16, TCKY19, TLG14, WPNK12, WMB^{+}14, YIKG10]. usage [PLM^{+}16]. Use [AGNH^{+}19, BWP^{+}11, BD11b, CB12, HTSA16, ITT14, JCRC10, KKY15, MBS^{+}14, RSV10, SJB16b, SHB^{+}10a, SGB^{+}10, TF12, ALV12, AB11, ABH18, AMM^{+}12, AMP^{+}11, AED^{+}12, Ano18b, AS12, AHB^{+}14, BMA^{+}10, BDM^{+}13, BOC^{+}16, BMTC^{+}15, BDvGW16, BDH^{+}17, BAH^{+}15, BDG^{+}16, BMPH14, BCB^{+}17, CLH^{+}15, CSM^{+}19, CWEE19, CHST16, CAB^{+}19b, CHFS12, CNJ^{+}18, CNW^{+}12, DWPH11, DPO^{+}17, DMHS^{+}14, DPH13, ED12, EPC^{+}19, EC14, ES12b, FA11, FGTS16, GDP^{+}11, GTM14, GDPR18, GGF^{+}12b, HLIK16, HKL19, HPL10, HSF^{+}12, HMO^{+}12, HED^{+}13, HS14b, HIH^{+}17, HMR^{+}16, HLH18, INvdB^{+}13, IVD^{+}10, IDT^{+}18, JRW^{+}14, JFK12, JWT^{+}14, JMS^{+}15, KSHS19, KWM^{+}17, KCS12, KHS^{+}11, KBA17, KBB16, LI18, LHDH17, LMR^{+}10, LKL^{+}17, LA12b, MHCS18, MHS15, MWC18, MPG^{+}10, MHSG18, MHAP18, MWGH15, MCT16, NGSF13, NW15, OMN17, OGLH17, OAB^{+}19, PRLR^{+}18]. use [PSA12, PM16, PMSGBL16, PJP^{+}11, PFKB18, PW19b, RPE^{+}13, Rit17, RPL^{+}18, RMV13, SRG14, SWM^{+}19, SRSH^{+}18, SRP^{+}19, SPHK17, SW18, SBFP10, SSPKF16, SGS^{+}19, SFM^{+}19, SWNSF14, SD17b, SCF^{+}17, TS11a, TWY^{+}19, TCSA17, TCJ^{+}19, TRSB^{+}18, TMP^{+}11, TSA^{+}18, TMH^{+}13, TLK^{+}12, UCK^{+}13, WPNK12, WK18, WW15, WHH12, WS11, ZBK^{+}15, ZMRCP^{+}17]. used [RS14, SE11a, VRSS18, WV12, dJSB19]. useful [MPKO11]. Using [DE17, DP14b, GL12, GFC^{+}13, HW12a, HS15b, KMGH16, LR18, NCC^{+}14, OALK14, ORW15, PPX17, PL10, RST15, SHP^{+}18, SGP^{+}17, TSTP13a, ZMRCP^{+}17, ARB^{+}12, AMR^{+}10, APH^{+}10, Ano17c, AR14, AMAJ^{+}17, BWM^{+}15, BGB^{+}17, BFG^{+}19, BFF^{+}10, BMTC^{+}15, BJ14, BDT^{+}11, BGDC14, CLB^{+}14, CR16, CMIMGS15, CLB19, CSKR10, CMV^{+}11, CCHG10, CMLC12, CJS^{+}12, CPS^{+}14, DIG^{+}11, DER^{+}19, DOH^{+}17, EM11, ETHB17, EFOM18, ECZD14, FGW^{+}15, FBPCC^{+}19, FMJA^{+}18, FTvdK^{+}12, FEE^{+}11, GED^{+}14, GPA^{+}16, GMR^{+}19, GCC^{+}14, HAM^{+}15, HSH12, HBK^{+}17, HRT^{+}14, HCJ^{+}11, IEM16, JAPK16, KW17, KLBT^{+}16, KRR^{+}18, KPP^{+}16, LAL^{+}18, LBVH13, LHT^{+}16, LPC19, LPKT^{+}19, LLP^{+}10, LRO^{+}17, LRK^{+}15, MMRC16, MOMF14, MHS15, MHW^{+}12, MIV^{+}10, MB11, MR16, MO13, NRF11, NHTS14, OMW^{+}16, PKS11a, PPD^{+}14, PIPD10, PSL^{+}12, PMM^{+}15, Pro17, RWM18, RFM^{+}15]. using [RS16, RBD^{+}14, SS10a, SLS11, SKD^{+}11a, SWK19, SKB15, SOJ14, SHMOS17, SKMS^{+}18, SMSB19, SLO^{+}17, SAMW14, SGD^{+}12, SSS12, THM^{+}13, TRA^{+}14, USR^{+}16, Vig15, WLM12, WRB^{+}15, WVD^{+}17, WFH^{+}18, WPP11, WMHW14, YSC^{+}15, YBB^{+}12]. utilisation [ACRLB12, LDP^{+}17, PBDMRB13, SMND16, VARO12]. utilised [HLB^{+}16]. utility [PLM^{+}16, WHH^{+}13]. Utilization [BP10, BK14, DKF^{+}15, EBdlTC12, FW11, FF12, GED^{+}14, HPA17a, HRPBLT18, KHS10b, LMV^{+}19, LWG14, LALCM18, MWA17, NOT^{+}11, TMP^{+}13, WJW^{+}18, WA19]. utilize [DMC^{+}16, GWGC16, ZF19]. utilizing [HR19]. UV [DGQ10, EMG^{+}15, HPL^{+}16, LRTLF12, LAALH10]. UV-absorbing [EMG^{+}15]. UV-B [LRTLF12]. UVB [CFA16]. UVR [FLSG^{+}15, LAALH10, vdPLdV^{+}11].

V9 [AAAE17]. vagile [FTO^{+}15]. Valdez [BBM13, BBC^{+}12a, HG13]. Valdís [RUS^{+}13]. validated [SWK19]. validating [PCP^{+}15]. Validation [DH18a, NSC^{+}19, RSG^{+}16, SvdMK^{+}12, CPB^{+}11, HHH11a]. vallentini [GGK^{+}16]. Valley [SCAM^{+}14]. valuable [MDKM13]. value [GRN^{+}16, SMR10, vBAP^{+}19]. values [CCJR14, CSC^{+}11, DH18a, DBS^{+}19, FSE^{+}17, GCS^{+}18, HKB^{+}19b, KNS^{+}15, NW15, OPN10, RFEJK11, SDR^{+}15, TTT^{+}15, TMS14, VWEP18, VLM14, WRW^{+}16a]. Valuing [GS15]. valve [GFR^{+}18]. Vancouver [HPB^{+}18]. vannamei [DPH14]. Vanuatu [LCG^{+}17, SØA^{+}13, SØK^{+}15]. var. [GP11a]. Variability [BD14, BDR19, BDF^{+}11, CSABH^{+}13, DBS^{+}19, HTH10, KWPL11, KM10b, LKB^{+}14, LKN^{+}14, MC10a, MFG16, OCS14, PHHL19, TGSM12, WK18, AAGGFG^{+}12, Ano17c, Ano18a, ADB10, BHP12, BDNA10, BSK^{+}11, BPSCD17, Ber12b, BMVBC10, BDP10, BCSB13, BMF^{+}15, BHN^{+}12, CT10, CC19, CAC11, CFPF^{+}14, CMHH17, DMd14, DR16, DQC^{+}13, DCT^{+}10, DHGW15, DRH15, EDS14, EN11, FK10, FMH^{+}19, FCH^{+}10, GDB^{+}16, GMB^{+}16, GCM10, Gla10, GCCM18, HER^{+}19, HHS^{+}12a, HPN^{+}17, HLS15a, HTB^{+}15, HT10, HBM^{+}12, HSA^{+}15, HCJ^{+}11, JW11, JSS^{+}11, KNIN17, KTP^{+}18, KFGSMPG14, KM17, LAM11, LAM13, Lec18, LKLC11, LS16, LRQ^{+}12, MSFM12, MCPC19, MLPC18, MVP^{+}16, MLR17, MPM^{+}14, MH10, MFH^{+}11, MN18, MMWG13a, MG12, MGB^{+}14, MPRPM^{+}18, MMEM14, NWS^{+}14, NKS^{+}15, NDG15, NSS10, NLGE18, ODP^{+}12, OF10, OHR^{+}13, OBÁSV18, OGCHS18]. variability [OMD11, PNI14, PLGSD17, PPM11, PLF^{+}12a, QNG^{+}15, RKD18, RSW^{+}16, RFMT12, RRB13, RKC16, RBEGG13, RDF^{+}13, RKPB^{+}15, RKM^{+}10, RQN^{+}14, RDF^{+}19, SDG^{+}12, SMSE19, SGMG^{+}14, SHS^{+}17, SMND16, SDKD13, SBB17, SKM12, SKD^{+}17b, ST13, SML^{+}15, SS10b, SDAM11, SKFH18, Sma13, SBE^{+}16, SML^{+}14, SJS^{+}11, SPS16, SDVL16, SAD^{+}16, SNS12, TKA10, TKW^{+}14, TBDM11, VRMKH^{+}15, WDNM11, WHPT12, WB12, WPRS18, WCG^{+}14, YPWP18, YHT16, YHTC19]. Variable [DHM^{+}14, LH11, UCK^{+}13, WJW^{+}18, BWM^{+}15, DS16, GKHS16, HW14, JCM10, JWB^{+}17, LCQT^{+}13, LBE^{+}10, LH10, MBBBC12, MMSC17, MFI11, PHAZ18, PFL13, PPEH^{+}14, YHR16, vERRP16]. variables [CCHG10, DHK^{+}17, KKS^{+}12b, LLBP15, SBE^{+}16, WG11]. variably [RBD^{+}16]. varians [OT14]. Variation [ALWD15, AP14, BGJ^{+}19, CB14, CNJ^{+}18, Det10, FSE^{+}17, LMGJ13, MGFW13, MBT14, PMH^{+}13, PRR^{+}16, RGL^{+}16, TBCG^{+}12, TDPS13, VB14, VBT^{+}17, VBV^{+}10, WL11, AH13, ABA16, AY10, ACG^{+}11, BTO^{+}18, BVG13, BCvH^{+}15, BSN^{+}18, BHASP18, BB11, BPB18, BS12, BDPL19, BRS10, BS14b, BBW^{+}17, BRA^{+}10, BH12a, BCB^{+}17, BMT17, BBC12b, BCB11, BJRH10, CCB^{+}17, CR16, CLDC^{+}13, CGG10a, CSF13, CTMU10, CRRL17, CPF^{+}16, CTB13, CP11, Col10b, CO16, CKR^{+}11, CC12, CGG^{+}10b, DSG^{+}13, DDP19, DJJ^{+}16, DAD^{+}11, DJL^{+}12, DHB18, DPC13, ELA10, Edm14, ES12b, EPS11, FFL10, FTS^{+}13, FHE^{+}15, GM13a, GVW^{+}12, GTP^{+}17, GPM^{+}13, GS16, GGPDS19, GAB^{+}14, GCCCM16, GDH^{+}19, GHNF10, GCL^{+}16, GFW^{+}14, HEMJ18, HCL^{+}11a, HPL10, HPMB11, HRM^{+}11, HSS^{+}10, HMR^{+}16, HCW^{+}15, HBG^{+}17, HMF^{+}18, JJP^{+}13, JBS11]. variation [JWR^{+}12, JBL^{+}16, JS12b, KCCG12, KTH^{+}18, KCH^{+}11, KP11, KHK^{+}17, LHPP16, LRL^{+}13, LB11, LFM^{+}16, LEB14b, LNIH16, LTP13, LMCD19, LSV10b, MGP11, MGO^{+}12, MMFF12, MPLLE17, MM12, MM13b, McD13, MMM^{+}15, MBC^{+}17, MAS^{+}10b, MS13, MFS14, MWW^{+}14, MOS10, MSM^{+}12, MJV^{+}11, MMB^{+}12b, MH11c, MJHS13, MSK^{+}19, MWMD17, NMS^{+}18, NSH^{+}15, NDBL16, ORR^{+}14, OOK11, PBAS^{+}10, PW15, PBDMRB13, PDP^{+}10, PPS^{+}13, PSS19, PB10, Rai12, RPRN^{+}16, RMBM14, RCU^{+}11, RCVG10, SVAV17, SK13, SRK^{+}18, SGC^{+}13, SDR^{+}15, SPHK17, SL11a, SKR^{+}13, SFW^{+}14, SCK^{+}17, SGS^{+}19, SE11b, SG12, SAMW14, SFKQ15, SLE^{+}13, SLO19, SS10c, TOW10, TPB^{+}16, THB^{+}12, Tho15, TKIO10, Vig15, WBG^{+}14, WG11, WWG13a, WTD^{+}19, WBS^{+}17, WHL11, WGGM11, YAFV12, YFC^{+}13, YYO14, YKT^{+}14, ZBB^{+}11, ZPLW11, ZMRCP^{+}17, dJdBdJB12, vdMHW^{+}10]. Variations [MBB^{+}12b, SLS16, AB11, BCSR10, CFF^{+}18, CP10, CHST16, DPO15, FC10, Gal14a, Gal14b, MMG^{+}15, MBSG11, OPN10, PBAP^{+}14, SPJ^{+}12, SCL^{+}16, SA11, WRDF12, YSK^{+}10, YBB^{+}12]. Varied [SDKS19, SRK11]. variegata [RMME10, ZHM^{+}10]. variegatus [PCW^{+}12, WYS^{+}10]. varies [AZVB13, GD15, JTTB^{+}18, JL17, KCGC14, LPI^{+}14, LE10, PC19]. various [MFG16]. vary [Ale13a, CFH11, DCL^{+}17, DGC12, GAO14, JWHR11, KDO^{+}14, LSH^{+}15]. Varying [LABG^{+}12, BPT15, BCK15, DWLH13, ES12a, KBW^{+}15, KMSW10, LLv13, RCS^{+}17, SE16]. vectensis [RCE^{+}13]. vector [HW11a]. vectors [LWD13]. vegetated [HPQ15]. vegetation [KAAI15, LHT^{+}17, PW15, PBA12, SLS11]. vegetation-free [KAAI15]. vehicle [HBM^{+}15, KBG12]. vehicles [KMGH16, SHP^{+}18]. Velella [PCD^{+}18]. veliger [VE18]. velocities [DFPH11]. velocity [LCTD^{+}15, OFM^{+}18, VBGO^{+}16]. Venice [BPSG11]. vent [CKV12, FC13, GMG^{+}16, SNW^{+}18, SJL^{+}15, YWiIK15]. vents [DRG^{+}12, GGFB15, KM10a, KLSJ10, Met11, NUO^{+}19, RSW^{+}16, SLRS13]. Verasper [WYS^{+}10]. Verdes [TWL15]. Verity [Ras10]. vermetids [BFL^{+}19]. vermiculophylla [HWR^{+}13, JL12]. verrucosa [HCM^{+}11, LE10, LS15]. versatility [BJMS12]. versus [AP14, ACMD11, BDS18a, BWH11, CKR^{+}11, CBK^{+}12, JOFP18, JRK13, JMB^{+}18, KKMM16, KBK^{+}10, LBAK13, LHDH17, LB15b, LBWB^{+}13, MWA17, RGDNGS11, SCL^{+}19, ST11b, TKDM16, TPA^{+}13, VFMdW12, VMN^{+}12, WFBM14, WKRW^{+}12, dJH11]. vertebrae [TMP^{+}11]. Vertebral [IHD^{+}16]. vertebralis [MU13, RUP11b, RUP11a, SHDR13]. vertebrate [DH16]. vertebrates [CCGP14, PBC^{+}12b]. Vertical [HP11a, JT10a, KKO^{+}14, LA12b, MBM^{+}14, MAHK12, OMT^{+}17, PLS^{+}10, RBA^{+}18, RB19, RHRB^{+}17, SSB^{+}11, VWW^{+}16, WPML^{+}18, ADHP12, ALNB17, BVS^{+}16, BLD^{+}16b, BLS^{+}11, BPSCD17, BSN^{+}18, BBRBG10, BNS^{+}14, BQRB12, Bol17, DGA^{+}15, DT15, DBHMG12b, EOA15, ETO^{+}10, FOJF15, GWT18, GSD12, HCLB18, HSS17, KKA12, KBPD10, LSY^{+}18, LMd12c, LMd12b, MBG^{+}16, MPKO11, MLP19, MAB12b, MG12, MGB^{+}14, MKI^{+}15, MJHS11, NWP^{+}11, PSER17, PDBB14, RDHM10, RKC16, RGB^{+}11, RGP11, SDKD13, Sat10, SPWK13, SFKQ15, SA11, SRK11, SBM^{+}18, SPBS14, SOBNL13, TRGF18, TGSM12, VCFF13, XLXS17, YNK18]. Vertically [BSST13, LA14]. very [BH16a, GSL17, NWH^{+}12, TPB^{+}18]. vesiculosus [AOW12, BSP^{+}10, MCRB11, TPC^{+}16, WSD^{+}10, WRO^{+}11]. vesparium [WRGP12]. vessel [BLD^{+}16a, CAB^{+}16, FMG^{+}18]. vessels [CASG^{+}19, CPW15, RHW12, TTB^{+}11]. Vestfold [CBKM16, GSJS13]. vestibulo [SCSC16]. vestibulo-ocular [SCSC16]. veteran [KBGA19]. vetula [BFN^{+}19]. VHF [LAH^{+}15]. VHR [SKD^{+}11a]. VHSV [JBM^{+}13]. via [AM19, DLC^{+}13, OLD^{+}15, PSG11, QYS^{+}12, RdGv^{+}18, SES12, VUL^{+}11]. viability [AFRAG16, MC10b, MOS10, RZH^{+}16b, VK11]. viable [Fog12]. vibration [RCBE15]. Vibrio [GGF^{+}14]. vibrissa [AGNH^{+}19, FHMG19]. vibrissae [SDR^{+}15, SCW^{+}15b]. Vibrissal [LCB^{+}16, LRO^{+}17]. vicariance [BCC^{+}11, HI14]. vicinity [RKT16]. victoria [KBR^{+}11]. Video [KWC^{+}17, MJT^{+}12, ASS^{+}13, AWH19, BBM^{+}14, BGB^{+}17, BG12, CS10a, DOH^{+}17, ETHB17, KSHS19, LYH^{+}17, MRLD19, RKFD15, VBR^{+}14, WFBM14, WLM^{+}13, ZVR18]. video-based [AWH19]. video-recorded [MRLD19]. videography [CVW^{+}11]. videos [BBC14, SJB16b]. view [Ji11]. Vigo [APC14a, FAZ^{+}14, THRBL^{+}16, BC11]. villosus [CD16b, LBR^{+}19, OA10]. vincta [YLPG19]. viral [AWJ12, EH13, HGG^{+}10, JBM^{+}13]. virally [GAF^{+}16]. virens [CEK^{+}16, LANE18]. Virgin [BFN^{+}19, DH18b, Edm13, Edm14, EPL^{+}14, EL16, LRBS15, OLF11, WSK^{+}16]. Virginia [LB15a, SSD13, BLD^{+}16a, LBAK13]. virginica [ALV12, ALK17, ATP14, BG19, BIL^{+}10, BSCH14, CB11, CL15c, GLF^{+}17, GFL^{+}17, HPA17b, HR15, KBW^{+}15, KW10, KPP^{+}16, MB16, MWHM18, PFL18, RHF19, SCM13, SFW^{+}16, WHA^{+}13, WLH^{+}17, WR12]. viridis [BTTK10, BS14b, CRR^{+}19, DBCV^{+}12, JLQ^{+}14, KHD^{+}11]. virulence [BRB17]. virus [JBM^{+}13, KAP^{+}14, PWCvO^{+}14]. virus-controlled [KAP^{+}14]. virus-like [PWCvO^{+}14]. viruses [UMF^{+}10]. viscosity [FS18]. viscous [RL18, SF18]. visible [DQM^{+}19]. Visitor [BGM^{+}17]. Visual [AAK^{+}15, BGB^{+}17, CPW15, CS10a, GSJS^{+}19, KB12, SOA^{+}17, WFS10]. Vital [BRB11b]. vitamin [BLHRP^{+}19]. vitulina [HMAC^{+}11]. vocalization [RCD^{+}13]. volcanic [ODL^{+}13]. volcano [LKL^{+}19, PBV^{+}11, RT17, SJL^{+}15]. volcanoes [RDC^{+}12]. volitans [AL12a, Alb15, ACL^{+}16, BWM^{+}15, BMA^{+}10, CBK^{+}12, CMH14, EPM^{+}16, HTSA16, PHSBO16, SJO16, WMB^{+}14]. volitans/miles [SJO16]. Volumes [Ano10a, Ano10b, Ano10c, Ano12a, Ano12b, Ano13a, Ano13b, Ano14a, Kin10a]. voyages [SWR11]. vs [FWPP12, FGP17, RB17, WAP10]. vulgaris [CRAP^{+}12, CPB^{+}11, KM14, PRNR^{+}18]. vulgata [BHDC15, NMHB16, SHC^{+}10]. Vulnerability [DBG15, SHB^{+}10b, VCVK^{+}10, ABW14, BWN^{+}14, CS15, CTSH12, GKRG^{+}19, JROS17, SCMSB^{+}11]. Vulnerable [KPr^{+}18, AS12, CEH18]. vulpinus [CWK^{+}10, CWW^{+}16a].

Wadden [BDD17, ADvdM13, BAA12, BD15, BDDvdM17, CBC^{+}10, HMM^{+}15, JRvdM15, JDG^{+}17, MDJN19, vdVKA^{+}11]. wading [HS15a]. wakes [KCD13]. Wales [MGM^{+}18]. walled [CIF11]. walleye [BBMH13, NRK^{+}19, RDK14, SDAH12, Urb12, WDM^{+}18, YFC^{+}13, ZBC10]. walls [GWT18, LZF^{+}14, WJ12]. Walrus [JFK12, JUK^{+}10]. walruses [HKL15, NUJ12]. Wandering [GPA^{+}16, CXdG^{+}17, XWE^{+}14]. war [SLRM^{+}15]. Warm [MAK^{+}15, ACAD12, BG12, FWB^{+}15, HWTH12, KN12, OVF11, SJL^{+}16, SSdW^{+}12]. warm-temperate [BG12, HWTH12]. warm-water [KN12]. warmer [AHPKS17, KDS^{+}12, LB15a]. Warming [BDPD14, PMS15, PRP^{+}16, RLO18, RG14, AMLS15, AS13, BMÁN^{+}18, BVM^{+}19, Bea12, BCOMKJ18, BRIC13, BGBD^{+}13, CMC10b, CLMM^{+}14, DAWS19, DP14b, EKE16, ESD16, FDFS16, FRTB14, GHHH12, GELJ19, GOM^{+}19, GLHM^{+}18, HM17, HBBH12, JPI^{+}19, KBW^{+}11, KCB18, KH16, LZ12, LBJ^{+}16, MLPC18, MTM^{+}17, MMM^{+}15, MWW^{+}14, OD15, PSPR18, SBH^{+}19, SGL^{+}17, SSM^{+}16a, SSM^{+}17, SMS15, SW12, SABMG16, STS^{+}18, STP10, TFH^{+}19, WAP10, ZHJ^{+}17, tHHR10, SSM15]. wars [OS18]. warty [LCCT19]. Warwick [Bet14]. Washington [GKR^{+}15, HOA14, Gre10, KFW^{+}18, RDF^{+}17, YBB^{+}12]. waste [EFR12]. Wasted [BOS18]. wasting [JTTB^{+}18, JTTP19, RSE18]. watching [CRL13, SCB^{+}16]. Water [BC13, BOU13, DTS14, DVS13, LBN11, MJA^{+}18, WWG^{+}13b, ARB^{+}12, AL12a, AFRAG16, AS13, ALNB17, AHB^{+}14, BRG^{+}14, BMV19, BLER15, BPB11, BSA^{+}16, BPLT14, CGR^{+}15, CLGCA^{+}19, CLS^{+}16, DIG^{+}11, DDK^{+}18, DMR10, DMM^{+}10, DWV^{+}12, DJL^{+}12, EOA15, FKvdV12, GSA^{+}18, GSC14, GKC11, GJSR18, GWK^{+}19, GDE17, HP13, HB13, HTT^{+}12, HU13, HCF^{+}16, INvdB^{+}13, JWHR11, JWR^{+}10, JGL12, KHA^{+}12, KN12, KIU^{+}13, KGLLW11, LKC^{+}15, LLv13, LCB10, LWP^{+}13, LGC10, MC10a, Mat10b, MBW^{+}19, MSPO12, MKII19, NC14, OTK^{+}19, OT14, OFM^{+}18, OFPR^{+}11, PW15, PPTK19, RAA^{+}16, RKFD15, RL11, RAB^{+}15, RLS10, SWR11, SSM15, SMS15, SHMB15, SBM^{+}18, SSdW^{+}12, THM^{+}13, TKB^{+}13, TRMC11, TKL^{+}11, TOR^{+}10, VB19a, VRMKH^{+}15, VHM16, WKvW10, WWKS15, WRFM16, WJ12, WAA^{+}19, ZFC^{+}13]. water-column [VB19a]. Water-flow [BC13]. water-temperature [VRMKH^{+}15]. waterbird [CHE^{+}17]. waterbirds [KRW18]. waterborne [CLMM^{+}14]. waterlogged [RZB18]. waters [AFOVRB14, ÁMLUN16, ANRT10, BTS^{+}14, BRB11a, BNE^{+}19, BNS^{+}14, CWW^{+}16a, CL10a, CPF^{+}16, CMAJ12, CDC18, CQR^{+}17, DHJ^{+}17, DLL^{+}18a, DHM^{+}14, FLK12, FSW10, GELJ19, GSÁCM^{+}16, GW15, GSMK15, HS14a, HW14, HEB^{+}11, HU13, HHH^{+}11c, IMD11, KEH10, KFGSMPG14, KAJP^{+}15, LSV^{+}10a, LGC10, LCS^{+}10, LRSB19, MTG^{+}10, MGFÁS^{+}10, MMFP^{+}15, MMB^{+}19, MRC^{+}19, NVA^{+}19, NHAS15, OLD^{+}15, OKJMC15, OAGV11, OGCHS18, PSÓ^{+}15, dGCF^{+}17, RG14, REP18, RLH^{+}19, SMND16, SKD^{+}17b, SS11a, SOT15, SB14, TMGCM11, TS16, TBMdV12, TNN11, UMF^{+}10, UTS^{+}18, VAHL11, VHV13, VMC^{+}19, WKX^{+}19, WPF^{+}16, WS12, WVM^{+}14, WOWG10, WHJ^{+}12, WWG^{+}13b, XWE^{+}14, YMMO14, YBH^{+}11, YHTC19, ZGK^{+}19]. watershed [ARB^{+}12, VGBJ^{+}13]. watersheds [OLF11]. Watersipora [SWMK15]. watsonii [DB11a]. Wave [BFNB16, GPB^{+}15, HR12, LRS17, PBA12, WBLRN18, BJM19, BRG^{+}14, Bur12, CSS^{+}17, DF15, FT10, HPH^{+}15, HPP^{+}10, IOS^{+}12, KKA12, KWM^{+}17, LET^{+}12, LPJ^{+}13, MSGGE10, NLV^{+}15, NCR^{+}14, NMHB16, OFM^{+}18, PG15, PNF^{+}12, PBW^{+}11, SLSC16, SHB^{+}10a, VWH18, WADKT13, WNF^{+}11]. wave-base [VWH18]. wave-driven [WADKT13]. wave-exposed [HPH^{+}15]. wave-mediated [KKA12]. wave-modulated [NLV^{+}15]. Wave-modulation [WBLRN18]. Wave-sheltered [BFNB16]. wavelength [BMSS14]. waves [RS14, RFJ^{+}10, SCL^{+}19, VGL16]. way [CWW^{+}16b]. Weak [MWM^{+}11, BB19, FNA^{+}16, SYX^{+}10]. weaken [LHB19]. Weakening [FSP^{+}17, SSM15]. weakest [BPM^{+}13]. weakfish [BT13, LTG17]. weakly [LEB14b, RM13]. wealth [FEA^{+}15]. weaning [CMHH17]. weather [CDMB^{+}18, CFPF^{+}14, GD15, KRA17, LAM11, NWHD15]. web [BEP^{+}12, BLV^{+}19, BDB^{+}12, CF12, CL10a, CLG^{+}19b, CHG^{+}18, DFS13, DSG^{+}13, DTM^{+}12, DIB15, EvDB^{+}14, ESH^{+}10, FL17, FPN17, GSJS13, GGK^{+}16, HHS^{+}12a, JRvdM15, JSMO14, KSS13, KFBM19, KLSJ10, LMV^{+}19, MMG^{+}15, MFH^{+}11, MBS^{+}18, MWWS17, MBRO10, MHS14, MFS14, MRB^{+}15, MSE^{+}12, MSS^{+}14b, NMPM15, NRK^{+}19, NBM^{+}14, NCT^{+}14, ODK^{+}14, OKLM14, PFB^{+}13, PDGL17, PMOAC^{+}17, PCW^{+}12, PGVF14, RBC^{+}14, RTD14, SBM^{+}18, TFP14, TP16, VC14, WPP11]. webs [BFRM13, BBI16, CWGG15, DVSM10, DBC^{+}12a, DMC^{+}16, DVB^{+}16, DRWQ12, FPC^{+}13, HS15b, KS12a, LR15, LLC^{+}14, LNK^{+}16, NDG15, OFT^{+}11, SBV^{+}15, SESS^{+}18, SS13, dJSB19]. Weddell [PMIG19, BKC^{+}19, CAR^{+}19, LSN^{+}19, REP18]. wedge [MGPC13]. wedge-tailed [MGPC13]. wedgefish [GRN^{+}16]. weedy [SCMSB^{+}11]. weight [DBH^{+}16]. Weinstein [Ano19d]. weissflogii [DTVH10, GRG^{+}12, MAB12b, PL15]. well [ADL^{+}18, WF18]. well-being [WF18]. West [BHN^{+}14, Dan11, DHG16, HVJ^{+}18, HDP^{+}17, PFVFF^{+}15, TRL^{+}15, AFH11, BHE^{+}17, BEG^{+}13, BD11b, CRS^{+}11, CH12, EZZ^{+}12, IEM16, KWW^{+}14, KFW^{+}18, NCS^{+}10, RQN^{+}14, SHH10, ASWB10, BNS^{+}14, BKM^{+}18, CDCZ16, CDC18, CGM^{+}19, CBL^{+}10c, DDP19, LHJH^{+}10, LKN^{+}13, LDAE12, MF15, NHSG11, RSDN15, SNT^{+}16, TSTP13a, TGBM^{+}16, WSLM13, WCG^{+}14, WS12, ZBP18, ZSEV17]. Western [EBdlTC12, HSA^{+}15, JBD10, MVP^{+}16, MMC^{+}18, TKS^{+}14, THB^{+}12, ABJ^{+}15, AB12, AMLR^{+}17, BCN15, BEMD11, BSC^{+}16, BDH^{+}17, BAC^{+}11, BWN^{+}14, CBC^{+}10, CC19, CMIMGS15, CBR^{+}18, CSF13, CHP^{+}11, CBC^{+}16, DSS10, DD15, DHK^{+}17, DCHI12, DPD^{+}12, DSK^{+}10, FC10, FTS^{+}13, GDB^{+}16, GSD12, GBS^{+}10, GLM^{+}19, HSH12, HJMN^{+}12, HAF^{+}17, HBK^{+}12, JRvdM15, JMK12, JHL^{+}19, KTNO19, KAD^{+}15, KBG^{+}13, KDS^{+}12, KKO^{+}14, LSH^{+}19b, LABM17, MPH^{+}11, MBB^{+}12a, MGAV15, MSM^{+}12, MPRPM^{+}18, MRC^{+}13, NDD^{+}13, OTK^{+}19, OMT^{+}17, PPML17, PHD^{+}13, PRK^{+}12, PHG^{+}14, QLCUG15, RCAB^{+}12, RGdLB^{+}11, RVH^{+}19, SDM^{+}13, SMND16, SHKR15, SLL^{+}18, SSB^{+}18, STP13, STBP13, SGB^{+}10, SOT15, SA17, SKD^{+}11b, TKW^{+}14, VCRUQ15, VCG^{+}10, VWW^{+}16, VKY^{+}12, WNG^{+}19, XTW^{+}15, YSK^{+}10, ZTNS13, dlGMD^{+}17, BLHL15, BLM^{+}18, FSR17, LPK^{+}16, MMCL19, PRP^{+}18, RSC14, SHK14]. Western [SCLB11]. wet [DCW^{+}14, DBH^{+}16]. wet-dry [DCW^{+}14]. wetlands [DMS14, DBS14]. Whale [BDH^{+}17, CRL13, KWM^{+}17, ALSO^{+}18, AEGH11, AED^{+}12, BLS^{+}11, BWLP17, BQ16, CGH11, CSV^{+}15, DRT13, DTS14, EPC^{+}19, FNF^{+}18, FBA11, FDLL10, GLR^{+}14, GMP^{+}11, GSJS^{+}19, HIH^{+}17, KHM^{+}12, LFPA^{+}11, LBH18, LRR^{+}10, LMLB12, MVP^{+}16, MHW^{+}12, MGGSP15, MMFP^{+}15, NSB^{+}19, NSC^{+}19, PMM^{+}11, PLF^{+}12a, PRP^{+}18, RBP^{+}10, RDF^{+}17, RCD^{+}13, RCR^{+}13, RPM^{+}13b, RUS^{+}13, RSSvV19, SWM^{+}19, SRP^{+}19, SCB^{+}16, SMD^{+}13, SMJBØ11, SGG^{+}12, SBB^{+}14, TCSA17, TBB^{+}19, ZBK^{+}15, dB10]. whale-watching [SCB^{+}16]. whaler [IHD^{+}16]. Whales [Wil14, AAM^{+}12, AAK^{+}15, BSC^{+}13, BLMD^{+}11, BLE^{+}13, BMV^{+}11, BDH^{+}17, BDS^{+}18b, BJS^{+}16, CPA^{+}11, CTZ^{+}16, CRRL17, CRL13, CVC^{+}18, CJS^{+}12, DPD^{+}12, FTM^{+}13, FMG^{+}18, FTS^{+}13, GVG^{+}18, GLM^{+}19, HER^{+}19, HLPS^{+}15, KWH^{+}18, LHJH^{+}10, LCR^{+}16, MMFF12, MHHB^{+}17, MF15, MRB^{+}11, MBP^{+}12, MHN^{+}12, MKG10, NSK^{+}14, ON15, OGGP14, OWN^{+}15, OAB^{+}19, RGL^{+}16, RVPG17, RKT16, RSP^{+}16, RMT^{+}13, SVAV17, SRLV10, SFM^{+}19, SWK^{+}16, SOC13, TMT^{+}12, TFB^{+}17, TGR^{+}19, TFW^{+}12, VRSS18, VCG^{+}10, VHP^{+}11, WNNN12, WEA^{+}17, WWB^{+}17, WVM^{+}14, ZMRCP^{+}17]. whelk [GHCC15, GP12, HLC^{+}18]. Whelks [MDPG13, CPPM12, DF15, GG12]. Where [BBWG12, GPW^{+}10, TCWC13, Hug11]. whether [MS10]. which [RTMJ11]. whip [FMC^{+}11]. Whisker [MWZ^{+}16, BSC^{+}15]. whiskers [NFC^{+}11]. White [Lap19, TTB^{+}11, ARvdW^{+}16, BM10a, BPB^{+}12, BHL^{+}19, DPH14, INK^{+}17, MMB^{+}19, PDP^{+}10, PWCvO^{+}14, SCSC16, SBCT17, SKFH18, SWH^{+}11, WPS^{+}19, ZBK^{+}14, KKMM16]. White-capped [TTB^{+}11]. white-chinned [PDP^{+}10]. white-sided [SWH^{+}11]. whitetip [DOH^{+}17, MBB^{+}15]. whiting [LTP^{+}18, MRE14]. whitmaei [SK10]. Who [TCWC13, Wir12]. Whole [FKP^{+}15, LFD13, PCS^{+}15, VMN^{+}12]. whole-animal [VMN^{+}12]. whole-ecosystem [LFD13]. whom [Wir12]. wide [BAS^{+}10, BDPL19, BCV^{+}18, BAE^{+}10, CWB^{+}15, DHS^{+}19, FTWA17, HVBG^{+}19, LSH19a, MCPC19, RBSL11, SKW^{+}14, WHN^{+}18, WEAG^{+}15, WWB^{+}17]. wide-ranging [BCV^{+}18, CWB^{+}15, DHS^{+}19, LSH19a, MCPC19, WEAG^{+}15]. widely [CRGM18]. Widespread [BBR12, EH13, LPB^{+}16, Bar14, DFDC18, LAB10, LLD^{+}12, PBM^{+}18]. width [MZF^{+}19]. wild [ABK^{+}14, CCM^{+}15, HWET13, KKMA^{+}10, MLB^{+}14, MOS10, MOK^{+}11, NWB^{+}19, NWH^{+}12, ODGL^{+}12, PRNR^{+}18, SVG16, SMK^{+}10, SST^{+}18, TKBN10, TSR17, WFP15, ZJ17]. will [CLMB13a, TBS^{+}17]. William [AMM^{+}12, MIA^{+}11, MBB^{+}12b, Ran18, SNVH19]. Wind [KCD13, BSB13, BDBN11, BDD^{+}18, DTC^{+}17, DEM15, SSvD^{+}15, TRSB^{+}18, WN16, vDGK^{+}12]. window [BLM10, LdU^{+}17]. windows [LHU^{+}18, BBH^{+}11]. winds [VGL16]. windscape [LSH19a]. winners [GMG^{+}16, vWSGL11]. Winter [AMAM^{+}18, CDCZ16, DABP13, EBH17, GDH^{+}19, HHS^{+}12b, MIA^{+}11, NYK^{+}13, RWG15, TNB^{+}15, BD14, BBW^{+}17, BNS^{+}14, CCG^{+}10, ED12, EZZ^{+}12, FSF^{+}17, JCM13, LG18, LCL^{+}13, LRQ^{+}12, MO12, McD13, PWB^{+}19, PWM^{+}19, RGW17, RCG^{+}15, SKD^{+}17b, SNVH19, SDMR19, SSdW^{+}12, TIN^{+}18, TMH^{+}13, VWH18, YOWK17, ZTNS13]. winter-hatching [RGW17]. winter-spring [ZTNS13]. Wintering [LRC^{+}13, HS15a, HG10, HFH10, KRW18, LFPA^{+}11, LBM^{+}13, LBJ^{+}16, PJP^{+}11, RB18, TCTB11, WYI^{+}16, WTC^{+}15, WMPM13]. winters [CMC^{+}16, DB14]. Within [FSF^{+}17, SRK^{+}18, ABF^{+}14, BC11, BdSC^{+}14, BKR^{+}15, BS19a, BVAB^{+}17, BC13, BCB11, CMF^{+}19, CLDC^{+}13, CTZ^{+}16, CL10a, DGC18, DSJNH18, DCT^{+}10, DB11b, FRC^{+}13b, GMP^{+}11, GCDH19, GDH^{+}19, HCAV10, HBS17, HMS^{+}12, HRT12, HTH10, HNMK15, JOFP18, dDTC^{+}17, KCW^{+}17, KHW^{+}19, KDJP10, KHF^{+}11, KSBM^{+}11, KM17, LSJ^{+}15, LR15, LCG^{+}17, LZF^{+}14, LMZ16, MJR^{+}14, MHS17, MUOS13, NBM12, NMG^{+}13, NDBL16, NCW^{+}19, OT14, RPS12, RVF14, RPR16, RMV13, SDSA14, SHB^{+}13a, SNW^{+}18, SKA^{+}14, SBE^{+}16, TMSG19, TWL15, VGBJ^{+}13, WDNM11, WSC13, WA14, YMD^{+}10]. Within-colony [SRK^{+}18]. within-estuary [VGBJ^{+}13]. within-guild [HBS17]. within-mudflat [DB11b]. Within-pair [FSF^{+}17]. within-season [GMP^{+}11, WSC13]. without [CFMN16, GCBS19, HSDB13, RLBC10]. wolf [WTGC^{+}11]. wood [DdRH14, HED^{+}13, NHAS15, YSC^{+}15]. wood-borer [DdRH14]. wood-boring [YSC^{+}15]. Working [DGC12]. world [FEA^{+}15, KPW^{+}19, PC12, RAB^{+}10, TMM^{+}19, TBGW^{+}11, WOHJ^{+}15, WWB^{+}17]. world-wide [WWB^{+}17]. Worldwide [MIM^{+}12, RTÁB^{+}14, RUD^{+}16]. worm [KBS16, LRS^{+}12, SBH^{+}18]. wounded [BFKR17]. wounds [vBBW^{+}14]. wrack [BGL^{+}11, FESPR14, LMW^{+}13, PHM11, RHMR18, ROLA15, WSD^{+}10, WRO^{+}11]. wrasse [BMSS14, DW12]. wrasses [BJMS12, KBB16, NFN^{+}15]. wrightii [LMRRB17, MH13].

xanthogrammica [LMP12]. xanthophyll [vdPLdV^{+}11]. xanthurus [BT13]. Xenostrobus [BVO16]. Xisha [YHTC19]. Xisha-Zhongsha [YHTC19].

Year [ASAG17, EMTB^{+}12, HKL15, RPM^{+}13a, SHD^{+}19, ZMGS17, BMV^{+}17, BCWSR14, CM14, CCJR14, CPM^{+}17, DHJ12, GSJS^{+}19, HLS17, HLF^{+}19, HHI^{+}14, HB12, Hug11, HB15b, JKM11, MS18a, NCW^{+}19, OS12, PB19, RCD^{+}13, SKR^{+}13, SDCC12, SBB^{+}14, SKR^{+}10, SLL12]. Year-long [EMTB^{+}12]. Year-round [ASAG17, HKL15, RPM^{+}13a, SHD^{+}19, ZMGS17, CCJR14, CPM^{+}17, HB12]. yearling [BWRP14, RWP^{+}14, YBB^{+}12]. Yearly [PDB15]. years [Ano18a, CSM12, ED12, GKR^{+}15, KBG12, KKL13, LPI^{+}14, LPSL14a, MPM^{+}14, NJJ^{+}11, RCP^{+}13, TFS14, ZWP^{+}17]. yellow [BL13, CPF^{+}14, ETS^{+}11, KHW^{+}12, WWB14, XLXS17, ZCC^{+}12]. yellow-bellied [BL13]. yellow-eyed [KHW^{+}12]. yellow-legged [CPF^{+}14]. yellowfin [KRR^{+}18, ODK^{+}14, RTCF^{+}17, WRI12]. yellowhead [HPL12]. yellowtail [MSF18, PSA12]. yield [CS10b, VQSH14]. yields [SGS^{+}12]. yo [NWP^{+}11, NWP^{+}11]. Yo-yo [NWP^{+}11]. yokohamae [JNYT13]. York [TLG14, LBAK13, LB15a, TG10]. Young [HLS17, WHH12, AIA18, BCWSR14, CG17, DHJ12, HS11b, JPL^{+}15, MD16b, NCW^{+}19, PB19, SGT^{+}14, SSS19, TKEV^{+}19, YHTS12, BS11]. Young-of-the-year [HLS17, DHJ12, NCW^{+}19]. young-of-year [BCWSR14, PB19].

Z. [OMD11]. Zealand [GOO14, ACMD11, BWH11, BRC^{+}15, CPA^{+}11, CDCS12, DSJNH18, DTL^{+}15, GS16, JW10, LRB^{+}12a, MS18a, RJR18, SLSC16, TRMC11, WPF^{+}16, WWG13a, WGD10, WJ12, WA19, XWE^{+}14]. zebra [SSS19]. Zebrasoma [ETS^{+}11]. zero [GMT^{+}11, SAMW14]. zero-inflated [SAMW14]. Zhongsha [YHTC19]. Ziphius [WWB^{+}17]. Zn [PWO^{+}16]. zoea [SHP^{+}14]. zoeae [CFA16]. zonale [CSJP17]. Zonation [APH^{+}10, GH11, MCK^{+}13, ST13]. Zone [KPH^{+}15, ANJ^{+}11, BSH13, CAB^{+}19a, CGEGC^{+}19, FB14, GOO14, GE17, GJ10, HKI^{+}17, HB13, JDQS11, JLL^{+}12, JPMG10, LHU^{+}18, LSY11, LAK^{+}19, LGD10, OBÁSV18, PHB^{+}11a, PB10, RNK^{+}11, RFEJK11, RG11, RHRB^{+}17, SMM^{+}15, SHH10, SHB^{+}10a, TBG^{+}17, WK10, WBLRN18, WS10, vdGvdLvG^{+}19, OCMT13, RWMP12]. Zones [FLE^{+}11, BSG^{+}17, DDB^{+}14, EMJ18, HAF^{+}14, RPL^{+}18, SBT12, TDB^{+}14, LCC^{+}11]. zoning [BBMM14, WD14]. zoobenthos [QNG^{+}15, WAM^{+}15]. zooid [OOK11]. zoophilous [vMVSW12]. zooplankter [GSV14]. Zooplanktivory [CSS15]. Zooplankton [CPP10, MJA^{+}18, SPP17, TPSS12, WZP^{+}14, ASEB11, BVS^{+}16, BGM^{+}18, BKM^{+}18, DR10, ECZD14, FSD^{+}14, FLS^{+}18, FUOG^{+}16, GSD12, HFR^{+}18, HHG10, HTCH^{+}18, HKB^{+}19b, ITT14, Kiø16, KNB^{+}13, LGT^{+}14, MSGGE10, MPM^{+}14, MH10, MJT^{+}12, NW15, OAGV11, PP10, PFL18, PJMA18, RRB13, SHPS^{+}16, SCW^{+}15a, SMP10, SKM^{+}15, THC12, VCL^{+}15, WBI^{+}15, WCG^{+}14, YChH11, vDCR13]. Zooplankton-mediated [TPSS12]. zooplanktonic [BLE^{+}13]. Zooxanthellae [McC12, ILDL10]. zooxanthellate [BCvH^{+}15, KHA^{+}12]. Zostera [AM18, BDM^{+}14, BO13, BD12, BOKJ18, BBR12, BRB17, BPF15, DDB^{+}16, FGOS16, HBHS^{+}18, HBB^{+}12, HH15, HNMK15, IEM16, JSBP15, JTTB^{+}18, JAPK16, JPL^{+}15, KBA13, KHGS12, KKPL14, KJCN^{+}11, MHS14, MFS14, OKJMC15, OCS^{+}13, OS12, OMD11, PW19a, RFRE19, RBM14, SRB15, SCRL11, SHKR15, SCM13, SO12, VCVK^{+}10, VPBV13, VBJR^{+}13, VSB^{+}15, VBGO^{+}16, WDH^{+}12, WL19, WNF^{+}11, ZHJ^{+}17, dlSBB^{+}10]. zosterae [BRB17]. Zvuloni [Car12].

References

Abad:2017:VMC

[AAAE17] D. Abad, A. Albaina, M. Aguirre, and A. Estonba. ^{18}S V9 metabarcoding correctly depicts plankton estuarine community drivers. Marine Ecology Progress Series, 584:31– 43, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

              articles/meps2017/584/m584p031.pdf.

Abecasis:2014:CMH

[AAE14] D. Abecasis, P. Afonso, and K. Erzini. Combining multispecies home range and distribution models aids assessment of MPA effectiveness. Marine Ecology Progress Series, 513:155–169, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

              com/articles/meps2014/513/m513p155.pdf.

Ambriz-Arreola:2012:ICU

[AAGGFG^{+}12] I. Ambriz-Arreola, J. Gómez-Gutiérrez, C. Franco-Gordo, B. E. Lavaniegos, and E. Godínez-Domínguez. Influence of coastal upwelling–downwelling variability on tropical euphausiid abundance and community structure in the inshore Mexican central Pacific. Marine Ecology Progress Series, 451:119–136, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/451/m451p119.pdf.

Araujo:2019:TFD

[AAGGG19] F. G. Araújo, M. C. C. Azevedo, R. S. Gomes-Gonçalves, and A. P. P. Guedes. Taxonomic and functional β-diversity patterns reveal random assembly rules in nearshore fish assemblages. Marine Ecology Progress Series, 627:109–123, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2019/627/m627p109.pdf.

Alvarez:2015:CEE

[AAIB15] M. F. Alvarez, M. Addino, O. Iribarne, and F. Botto. Combined engineering effects of clams and crabs on infaunal assemblages and food availability in intertidal systems. Marine Ecology Progress Series, 540:57–71, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/540/
m540p057.pdf.

Andrades:2019:MNB

[AAJJ^{+}19] R. Andrades, J. M. Andrade, P. S. Jesus-Junior, R. M. Macieira, A. F. Bernardino, T. Giarrizzo, and J. C. Joyeux. Multiple niche-based analyses reveal the dual life of an intertidal reef predator. Marine Ecology Progress Series, 624: 131–141, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/624/m624p131.pdf.

Aoki:2015:VBE

[AAK^{+}15] K. Aoki, M. Amano, T. Kubodera, K. Mori, R. Okamoto, and K. Sato. Visual and behavioral evidence indicates active hunting by sperm whales. Marine Ecology Progress Series, 523:233–241, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m523p233.pdf.

Aoki:2012:AHD

[AAM^{+}12] K. Aoki, M. Amano, K. Mori, A. Kourogi, T. Kubodera, and N. Miyazaki. Active hunting by deep-diving sperm whales: 3D dive profiles and maneuvers during bursts of speed. Marine Ecology Progress Series, 444:289–301, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/444/
m444p289.pdf.

Amsler:2013:TSM

[AAvS^{+}13] M. O. Amsler, C. D. Amsler, J. L. von Salm, C. F. Aumack, J. B. McClintock, R. M. Young, and B. J. Baker. Tolerance and sequestration of macroalgal chemical defenses by an Antarctic amphipod: a ‘cheater’ among mutualists. Marine Ecology Progress Series, 490:79–90, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m490p079.pdf.

Abrantes:2011:IVD

[AB11] K. G. Abrantes and A. Barnett. Intrapopulation variations in diet and habitat use in a marine apex predator, the broadnose sevengill shark Notorynchus cepedianus. Marine Ecology Progress Series, 442:133–148, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/442/m442p133.pdf.

Araujo:2012:EEC

[AB12] J. N. Araújo and A. Bundy. Effects of environmental change, fisheries and trophodynamics on the ecosystem of the western Scotian Shelf, Canada. Marine Ecology Progress Series, 464:51–67, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m464p051.pdf.

Anderson:2013:SDS

[AB13] J. R. Anderson and D. C. Behringer. Spatial dynamics in the social lobster Panulirus argus in response to diseased conspecifics. Marine Ecology Progress Series, 474:191– 200, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/474/m474p191.pdf.

Aalto:2017:PHR

[AB17] E. A. Aalto and M. L. Baskett. Post-harvest recovery dynamics depend on predator specialization in size-selective fisheries. Marine Ecology Progress Series, 564:127–143, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/564/
m564p127.pdf.

Angel:2016:PIC

[ABA16] L. P. Angel, M. Berlincourt, and J. P. Y. Arnould. Pronounced inter-colony variation in the foraging ecology of Australasian gannets: influence of habitat differences. Marine Ecology Progress Series, 556:261–272, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/556/m556p261.pdf.

Anderson:2014:RGA

[ABB^{+}14] A. B. Anderson, R. M. Bonaldo, D. R. Barneche, C. W. Hackradt, F. C. Félix-Hackradt, J. A. García-Charton, and S. R. Floeter. Recovery of grouper assemblages indicates effectiveness of a marine protected area in Southern Brazil. Marine Ecology Progress Series, 514:207–215, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/514/
m514p207.pdf.

Adhitya:2014:CIP

[ABF^{+}14] A. Adhitya, T. J. Bouma, A. M. Folkard, M. M. van Katwijk, D. Callaghan, H. H. de Iongh, and P. M. J. Herman. Comparison of the influence of patch-scale and meadow-scale characteristics on flow within seagrass meadows: a flume study. Marine Ecology Progress Series, 516:49–59, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/516/
m516p049.pdf.

Aceves-Bueno:2018:IDT

[ABH18] E. Aceves-Bueno and B. S. Halpern. Informing the design of territorial use rights in fisheries from marine protected area theory. Marine Ecology Progress Series, 596: 247–262, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/596/m596p247.pdf.

Ainley:2015:TCW

[ABJ^{+}15] D. G. Ainley, G. Ballard, R. M. Jones, D. Jongsomjit, S. D. Pierce, W. O. Smith, Jr., and S. Veloz. Trophic cascades in the western Ross Sea, Antarctica: revisited. Marine Ecology Progress Series, 534:1–16, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Aarestrup:2014:CRE

[ABK^{+}14] K. Aarestrup, H. Baktoft, A. Koed, D. del Villar-Guerra, and E. B. Thorstad. Comparison of the riverine and early marine migration behaviour and survival of wild and hatchery-reared sea trout Salmo trutta smolts. Marine Ecology Progress Series, 496:197–206, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/496/m496p197.pdf.

Altermatt:2012:HCM

[ABM12] F. Altermatt, A. Bieger, and S. G. Morgan. Habitat characteristics and metapopulation dynamics of the copepod Tigriopus californicus. Marine Ecology Progress Series, 468:85– 93, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/468/m468p085.pdf.

Agersted:2014:TPC

[ABN14] M. D. Agersted, A. Bode, and T. G. Nielsen. Trophic position of coexisting krill species: a stable isotope approach. Marine Ecology Progress Series, 516:139–151, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m516p139.pdf.

Adam:2015:HRC

[ABRP15] T. C. Adam, D. E. Burkepile, B. I. Ruttenberg, and M. J. Paddack. Herbivory and the resilience of Caribbean coral reefs: knowledge gaps and implications for management. Marine Ecology Progress Series, 520:1–20, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Aarestrup:2015:SPR

[ABT^{+}15] K. Aarestrup, H. Baktoft, E. B. Thorstad, J. C. Svendsen, J. Höjesjö, and A. Koed. Survival and progression rates of anadromous brown trout kelts Salmo trutta during downstream migration in freshwater and at sea. Marine Ecology Progress Series, 535:185–195, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/535/m535p185.pdf.

Andrews:2014:RSE

[ABW14] S. Andrews, S. Bennett, and T. Wernberg. Reproductive seasonality and early life temperature sensitivity reflect vulnerability of a seaweed undergoing range reduction. Marine Ecology Progress Series, 495:119–129, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/495/
m495p119.pdf.

Aksnes:2011:IAT

[AC11] D. L. Aksnes and F. J. Cao. Inherent and apparent traits in microbial nutrient uptake. Marine Ecology Progress Series, 440:41–51, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m440p041.pdf.

Almeda:2010:ETF

[ACA^{+}10a] R. Almeda, A. Calbet, M. Alcaraz, L. Yebra, and E. Saiz. Effects of temperature and food concentration on the survival, development and growth rates of naupliar stages of Oithona davisae (Copepoda, Cyclopoida). Marine Ecology Progress Series, 410:97–109, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/410/m410p097.pdf. See erra-

tum [ACA^{+}10b].

Almeda:2010:EET

[ACA^{+}10b] Rodrigo Almeda, Albert Calbet, Miquel Alcaraz, Lidia Yebra, and Enric Saiz. Erratum: Effects of temperature and food concentration on the survival, development and growth rates of naupliar stages of Oithona davisae (Copepoda, Cyclopoida), Mar Ecol Prog Ser 410:97–109, 2010. Marine Ecology Progress Series, 419:302–302a, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/419/m419p302.pdf. See

[ACA^{+}10a].

Aastrom:2016:ACS

[ÅCAC16] E. K. L. Åström, M. L. Carroll, W. G. Ambrose, Jr., and J. Carroll. Arctic cold seeps in marine methane hydrate environments: impacts on shelf macrobenthic community structure offshore Svalbard. Marine Ecology Progress Series, 552: 1–18, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ancona:2012:END

[ACAD12] S. Ancona, I. Calixto-Albarrán, and H. Drummond. Effect of El Niño on the diet of a specialist seabird, Sula nebouxii, in the warm eastern tropical Pacific. Marine Ecology Progress Series, 462:261–271, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/462/m462p261.pdf.

Amaral:2012:EEA

[ACB12] V. Amaral, H. N. Cabral, and M. J. Bishop. Effects of estuarine acidification on predator–prey interactions. Marine Ecology Progress Series, 445:117–127, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/445/
m445p117.pdf.

Acheampong:2011:FAE

[ACDJ11] E. Acheampong, R. W. Campbell, A. B. S. Diekmann, and M. A. St John. Food availability effects on reproductive strategy: the case of Acartia tonsa (Copepoda: Calanoida). Marine Ecology Progress Series, 428:151–159, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/428/
m428p151.pdf.

Aguzzi:2010:BRH

[ACF^{+}10] J. Aguzzi, C. Costa, Y. Furushima, J. J. Chiesa, J. B. Company, P. Menesatti, R. Iwase, and Y. Fujiwara. Behavioral rhythms of hydrocarbon seep fauna in relation to internal tides. Marine Ecology Progress Series, 418:47–56, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2010/418/m418p047.pdf.

Arnould:2011:SIR

[ACG^{+}11] J. P. Y. Arnould, Y. Cherel, J. Gibbens, J. G. White, and C. L. Littnan. Stable isotopes reveal inter-annual and interindividual variation in the diet of female Australian fur seals. Marine Ecology Progress Series, 422:291–302, 2011. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m422p291.pdf.

Ankjaero:2012:TST

[ACG12] T. Ankjærø, J. T. Christensen, and P. Grønkjær. Tissuespecific turnover rates and trophic enrichment of stable N and C isotopes in juvenile Atlantic cod Gadus morhua fed three different diets. Marine Ecology Progress Series, 461: 197–209, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/461/m461p197.pdf.

Anton:2016:PNI

[ACL^{+}16] A. Anton, K. Cure, C. A. Layman, R. Puntila, M. S. Simpson, and J. F. Bruno. Prey naiveté to invasive lionfish Pterois volitans on Caribbean coral reefs. Marine Ecology Progress Series, 544:257–269, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/544/m544p257.pdf.

Auge:2011:FBI

[ACMD11] A. A. Augé, B. L. Chilvers, A. B. Moore, and L. S. Davis. Foraging behaviour indicates marginal marine habitat for New Zealand sea lions: remnant versus recolonising populations. Marine Ecology Progress Series, 432:247–256, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/432/
m432p247.pdf.

Atkins:2016:NLE

[ACP^{+}16] S. Atkins, M. Cantor, N. Pillay, G. Cliff, M. Keith, and G. J. Parra. Net loss of endangered humpback dolphins: integrating residency, site fidelity, and bycatch in shark nets. Marine Ecology Progress Series, 555:249–260, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/555/
m555p249.pdf.

Alos:2012:DBH

[ACRLB12] J. Alós, M. Cabanellas-Reboredo, and S. Lowerre-Barbieri. Diel behaviour and habitat utilisation by the pearly razorfish during the spawning season. Marine Ecology Progress Series, 460:207–220, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/460/m460p207.pdf.

Aquilino:2012:MSD

[ACS12] K. M. Aquilino, M. E. Coulbourne, and J. J. Stachowicz. Mixed species diets enhance the growth of two rocky intertidal herbivores. Marine Ecology Progress Series, 468:179– 189, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/468/m468p179.pdf.

Altenritter:2018:PDF

[ACW18] M. E. Altenritter, A. Cohuo, and B. D. Walther. Proportions of demersal fish exposed to sublethal hypoxia revealed by otolith chemistry. Marine Ecology Progress Series, 589:193–208, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m589p193.pdf.

Austin:2019:SIF

[ADA^{+}19] R. E. Austin, F. De Pascalis, J. P. Y. Arnould, J. Haakonsson, S. C. Votier, G. Ebanks-Petrie, T. Austin, G. Morgan, G. Bennett, and J. A. Green. A sex-influenced flexible foraging strategy in a tropical seabird, the magnificent frigatebird. Marine Ecology Progress Series, 611:203–214, 2019. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m611p203.pdf.

Azevedo:2010:TSV

[ADB10] I. C. Azevedo, P. Duarte, and A. A. Bordalo. Temporal and spatial variability of phytoplankton photosynthetic characteristics in a southern European estuary (Douro, Portugal). Marine Ecology Progress Series, 412:29–44, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/412/
m412p029.pdf.

Avigliano:2019:FSC

[AdCM^{+}19] E. Avigliano, B. Maichak de Carvalho, N. Miller, S. Córdoba Gironde, A. Tombari, K. Limburg, and A. V. Volpedo. Fin spine chemistry as a non-lethal alternative to otoliths for stock discrimination in an endangered catfish. Marine Ecology Progress Series, 614:147–157, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/614/m614p147.pdf.

Abboud:2018:GEG

[ADD18] S. S. Abboud, L. Gómez Daglio, and M. N. Dawson. A global estimate of genetic and geographic differentiation in macromedusae — implications for identifying the causes of jellyfish blooms. Marine Ecology Progress Series, 591:199– 216, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/591/m591p199.pdf.

Adam:2018:CAF

[ADF^{+}18] T. C. Adam, A. Duran, C. E. Fuchs, M. V. Roycroft, M. C. Rojas, B. I. Ruttenberg, and D. E. Burkepile. Comparative analysis of foraging behavior and bite mechanics reveals complex functional diversity among Caribbean parrotfishes. Marine Ecology Progress Series, 597:207–220, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/597/
m597p207.pdf.

Abecassis:2012:MSD

[ADHP12] M. Abecassis, H. Dewar, D. Hawn, and J. Polovina. Modeling swordfish daytime vertical habitat in the North Pacific Ocean from pop-up archival tags. Marine Ecology Progress Series, 452:219–236, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/452/m452p219.pdf.

Ainley:2018:PFS

[ADL^{+}18] D. G. Ainley, K. M. Dugger, M. La Mesa, G. Ballard, K. J. Barton, S. Jennings, B. J. Karl, A. Lescroël, P. O. Lyver, A. Schmidt, and P. Wilson. Post-fledging survival of Adélie penguins at multiple colonies: chicks raised on fish do well.

Marine Ecology Progress Series, 601:239–251, 2018. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m601p239.pdf.

Alexander:2012:FRI

[ADO^{+}12] M. E. Alexander, J. T. A. Dick, N. E. O’Connor, N. R. Haddaway, and K. D. Farnsworth. Functional responses of the intertidal amphipod Echinogammarus marinus: effects of prey supply, model selection and habitat complexity. Marine Ecology Progress Series, 468:191–202, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/468/
m468p191.pdf.

Ayala-Diaz:2019:SPG

[ADRA19] M. Ayala-Díaz, J. M. L. Richardson, and B. R. Anholt. Sitka periwinkle grazing behaviour related to trematode infection: a multi-site study. Marine Ecology Progress Series, 632:123–130, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/632/m632p123.pdf.

Andresen:2013:GSD

[ADvdM13] H. Andresen, I. Dorresteijn, and J. van der Meer. Growth and size-dependent loss of newly settled bivalves in two distant regions of the Wadden Sea. Marine Ecology Progress Series, 472:141–154, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/472/m472p141.pdf.

Alberti:2010:CBH

[AED^{+}10] J. Alberti, M. Escapa, P. Daleo, A. Méndez Casariego, and O. Iribarne. Crab bioturbation and herbivory reduce preand post-germination success of Sarcocornia perennis in bare patches of SW Atlantic salt marshes. Marine Ecology Progress Series, 400:55–61, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/400/m400p055.pdf.

Anderwald:2012:SSE

[AED^{+}12] P. Anderwald, P. G. H. Evans, R. Dyer, A. Dale, P. J. Wright, and A. R. Hoelzel. Spatial scale and environmental determinants in minke whale habitat use and foraging. Marine Ecology Progress Series, 450:259–274, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/450/m450p259.pdf.

Anderwald:2011:RFS

[AEGH11] P. Anderwald, P. G. H. Evans, L. Gygax, and A. R. Hoelzel. Role of feeding strategies in seabird-minke whale associations. Marine Ecology Progress Series, 424:219–227, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/424/
m424p219.pdf.

Armstrong:2018:INC

[AEM^{+}18] C. T. Armstrong, D. L. Erdner, J. W. McClelland, M. P. Sanderson, D. M. Anderson, C. J. Gobler, and J. L. Smith. Impact of nitrogen chemical form on the isotope signature and toxicity of a marine dinoflagellate. Marine Ecology Progress Series, 602:63–76, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/602/m602p063.pdf.

Alvarez-Fernandez:2014:NSP

[AF14] Santiago Alvarez-Fernandez. The North Sea plankton regime shift in the late 1990s: Comment on Beaugrand et al. (2014). Marine Ecology Progress Series, 514:279–280, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/514/
m514p279.pdf. See [BHE14].

Agawin:2016:SNF

[AFC^{+}16] N. S. R. Agawin, P. Ferriol, C. Cryer, E. Alcon, A. Busquets, E. Sintes, C. Vidal, and G. Moyà. Significant nitrogen fixation activity associated with the phyllosphere of Mediterranean seagrass Posidonia oceanica: first report. Marine Ecology Progress Series, 551:53–62, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/551/m551p053.pdf.

Ashe:2015:LPS

[AFF^{+}15] J. L. Ashe, K. A. Feldheim, A. T. Fields, E. A. Reyier, E. J. Brooks, M. T. O’Connell, G. Skomal, S. H. Gruber, and D. D. Chapman. Local population structure and context-dependent isolation by distance in a large coastal shark. Marine Ecology Progress Series, 520:203–216, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/520/m520p203.pdf.

Atkinson:2011:EDT

[AFH11] L. J. Atkinson, J. G. Field, and L. Hutchings. Effects of demersal trawling along the west coast of southern Africa: multivariate analysis of benthic assemblages. Marine Ecology Progress Series, 430:241–255, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m430p241.pdf.

Alvarez-Fernandez:2012:TCP

[AFLM12] S. Alvarez-Fernandez, H. Lindeboom, and E. Meesters. Temporal changes in plankton of the North Sea: community shifts and environmental drivers. Marine Ecology Progress Series, 462:21–38, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/462/m462p021.pdf.

Alou-Font:2013:SCA

[AFMR^{+}13] E. Alou-Font, C. J. Mundy, S. Roy, M. Gosselin, and S. Agustí. Snow cover affects ice algal pigment composition in the coastal Arctic Ocean during spring. Marine Ecology Progress Series, 474:89–104, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/474/m474p089.pdf.

Alonso-Fernandez:2014:DBS

[AFOVRB14] A. Alonso-Fernández, J. Otero, D. Villegas-Ríos, and R. Bañón. Drivers of body size changes in a Pollachius pollachius stock in NE Atlantic coastal waters. Marine Ecology Progress Series, 511:223–235, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/511/m511p223.pdf.

Alou-Font:2016:CVP

[AFRAG16] E. Alou-Font, S. Roy, S. Agustí, and M. Gosselin. Cell viability, pigments and photosynthetic performance of Arctic phytoplankton in contrasting ice-covered and open-water conditions during the spring–summer transition. Marine Ecology Progress Series, 543:89–106, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/543/m543p089.pdf.

Agawin:2019:SMN

[AFS19] N. S. R. Agawin, P. Ferriol, and E. Sintes. Simultaneous measurements of nitrogen fixation in different plant tissues of the seagrass Posidonia oceanica. Marine Ecology Progress Series, 611:111–127, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/611/m611p111.pdf.

Aguiar:2015:EMS

[AFSL^{+}15] E. Aguiar, I. Fuentes-Santos, U. Labarta, X. A. Álvarez-Salgado, and M. J. Fernández-Reiriz. Empirical modelling of seston quality based on environmental factors in a mussel culture area (NW Iberian upwelling system). Marine Ecology Progress Series, 536:89–105, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/536/m536p089.pdf.

Azovsky:2012:BSP

[AGC12] A. I. Azovsky, L. A. Garlitska, and E. S. Chertoprud. Broadscale patterns in local diversity of marine benthic harpacticoid copepods (Crustacea). Marine Ecology Progress Series, 460:63–77, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/460/m460p063.pdf.

Ahern:2018:DSF

[AGGAO^{+}18] A. L. M. Ahern, J. Gómez-Gutiérrez, O. Aburto-Oropeza, R. J. Saldierna-Martínez, A. F. Johnson, A. E. Harada, A. R. Sánchez-Uvera, B. Erisman, D. I. Castro Arvizú, and R. S. Burton. DNA sequencing of fish eggs and larvae reveals high species diversity and seasonal changes in spawning activity in the southeastern Gulf of California. Marine Ecology Progress Series, 592:159–179, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/592/m592p159.pdf.

Avens:2012:PCA

[AGH^{+}12] L. Avens, L. R. Goshe, C. A. Harms, E. T. Anderson, A. Goodman Hall, W. M. Cluse, M. H. Godfrey, J. Braun-McNeill, B. Stacy, R. Bailey, and M. M. Lamont. Population characteristics, age structure, and growth dynamics of neritic juvenile green turtles in the northeastern Gulf of Mexico. Marine Ecology Progress Series, 458:213–229, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m458p213.pdf.

Alderman:2010:PFS

[AGHC10] R. Alderman, R. Gales, A. J. Hobday, and S. G. Candy. Postfledging survival and dispersal of shy albatross from three breeding colonies in Tasmania. Marine Ecology Progress Series, 405:271–285, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/405/m405p271.pdf.

Asjes:2016:ARS

[AGIW16] A. Asjes, J. M. González-Irusta, and P. J. Wright. Agerelated and seasonal changes in haddock Melanogrammus aeglefinus distribution: implications for spatial management. Marine Ecology Progress Series, 553:203–217, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/553/
m553p203.pdf.

Able:2013:FSD

[AGK13] K. W. Able, T. M. Grothues, and I. M. Kemp. Fine-scale distribution of pelagic fishes relative to a large urban pier. Marine Ecology Progress Series, 476:185–198, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/476/
m476p185.pdf.

Ardyna:2011:EFP

[AGM^{+}11] M. Ardyna, M. Gosselin, C. Michel, M. Poulin, and J. É. Tremblay. Environmental forcing of phytoplankton community structure and function in the Canadian high Arctic: contrasting oligotrophic and eutrophic regions. Marine Ecology Progress Series, 442:37–57, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m442p037.pdf.

Aurioles-Gamboa:2019:UEG

[AGNH^{+}19] D. Aurioles-Gamboa, S. D. Newsome, J. L. Hassrick, T. Acosta-Pachón, F. Aurioles-Rodríguez, and D. P. Costa. Use of ^{1}5N-enriched glycine to estimate vibrissa growth in free-ranging northern elephant seals Mirounga angustirostris. Marine Ecology Progress Series, 614:199–207, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/614/
m614p199.pdf.

Avens:2013:CSS

[AGP^{+}13] L. Avens, L. R. Goshe, M. Pajuelo, K. A. Bjorndal, B. D. MacDonald, G. E. Lemons, A. B. Bolten, and J. A. Seminoff. Complementary skeletochronology and stable isotope analyses offer new insight into juvenile loggerhead sea turtle oceanic stage duration and growth dynamics. Marine Ecology Progress Series, 491:235–251, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/491/m491p235.pdf.

Aurioles-Gamboa:2013:HTL

[AGRPSVL13] D. Aurioles-Gamboa, M. Y. Rodríguez-Pérez, L. Sánchez-Velasco, and M. F. Lavín. Habitat, trophic level, and residence of marine mammals in the Gulf of California assessed by stable isotope analysis. Marine Ecology Progress Series, 488:275–290, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/488/m488p275.pdf.

Aranguren-Gassis:2012:POB

[AGTS^{+}12] M. Aranguren-Gassis, E. Teira, P. Serret, S. Martínez-García, and E. Fernández. Potential overestimation of bacterial respiration rates in oligotrophic plankton communities. Marine Ecology Progress Series, 453:1–10, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Albouy:2011:PTG

[AGV^{+}11] C. Albouy, F. Guilhaumon, S. Villéger, M. Mouchet, L. Mercier, J. M. Culioli, J. A. Tomasini, F. Le Loc’h, and D. Mouillot. Predicting trophic guild and diet overlap from functional traits: statistics, opportunities and limitations for marine ecology. Marine Ecology Progress Series, 436:17–28, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2011/436/m436p017.pdf.

Andrews:2013:ELC

[AH13] K. S. Andrews and C. J. Harvey. Ecosystem-level consequences of movement: seasonal variation in the trophic impact of a top predator. Marine Ecology Progress Series, 473:247–260, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/473/m473p247.pdf.

Abecasis:2015:ERE

[AHA^{+}15] D. Abecasis, B. Horta e Costa, P. Afonso, E. J. Gonçalves, and K. Erzini. Early reserve effects linked to small home ranges of a commercial fish, Diplodus sargus, Sparidae. Marine Ecology Progress Series, 518:255–266, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/518/m518p255.pdf.

Aronson:2014:LUW

[AHB^{+}14] R. B. Aronson, N. L. Hilbun, T. S. Bianchi, T. R. Filley, and B. A. McKee. Land use, water quality, and the history of coral assemblages at Bocas del Toro, Panamá. Marine Ecology Progress Series, 504:159–170, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/504/m504p159.pdf.

Andrefouet:2012:DBE

[AHD12] S. Andréfouët, M. A. Hamel, and M. Dalleau. Distinction between effective pattern-based and selection-based biodiversity surrogates is essential: caveats for managers. Marine Ecology Progress Series, 452:287–295, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/452/m452p287.pdf.

Annis:2011:CBM

[AHH^{+}11] E. R. Annis, E. D. Houde, L. W. Harding, Jr., M. E. Mallonee, and M. J. Wilberg. Calibration of a bioenergetics model linking primary production to Atlantic menhaden Brevoortia tyrannus growth in Chesapeake Bay. Marine Ecology Progress Series, 437:253–267, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2011/437/m437p253.pdf.

Aldven:2015:MSR

[AHØ^{+}15] D. Aldvén, R. Hedger, F. Økland, P. Rivinoja, and J. Höjesjö. Migration speed, routes, and mortality rates of anadromous brown trout Salmo trutta during outward migration through a complex coastal habitat. Marine Ecology Progress Series, 541:151–163, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/541/m541p151.pdf.

Armitage:2017:GCW

[AHPKS17] C. S. Armitage, V. Husa, E. A. Petelenz-Kurdziel, and K. Sjøtun. Growth and competition in a warmer ocean: a field experiment with a non-native and two native habitatbuilding seaweeds. Marine Ecology Progress Series, 573: 85–99, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/573/m573p085.pdf.

Archer:2018:FDG

[AHR^{+}18] S. K. Archer, W. D. Halliday, A. Riera, X. Mouy, M. K. Pine, J. W. F. Chu, A. Dunham, and F. Juanes. First description of a glass sponge reef soundscape reveals fish calls and elevated sound pressure levels. Marine Ecology Progress Series, 595:245–252, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m595p245.pdf.

Alldredge:2013:SGF

[AHS^{+}13] A. L. Alldredge, S. J. Holbrook, R. J. Schmitt, A. J. Brooks, and H. Stewart. Skeletal growth of four scleractinian corals is not enhanced by in situ mesozooplankton enrichment. Marine Ecology Progress Series, 489:143–153, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/489/
m489p143.pdf.

Attard:2016:BPP

[AHSG16] K. M. Attard, K. Hancke, M. K. Sejr, and R. N. Glud. Benthic primary production and mineralization in a high Arctic fjord: in situ assessments by aquatic eddy covariance. Marine Ecology Progress Series, 554:35–50, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m554p035.pdf.

Alemany:2018:MFP

[AIA18] D. Alemany, O. O. Iribarne, and E. M. Acha. Marine fronts as preferred habitats for young Patagonian hoki Macruronus magellanicus on the southern Patagonian shelf. Marine Ecology Progress Series, 588:191–200, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/588/
m588p191.pdf.

Archibald:2016:EIA

[AJ16] D. W. Archibald and M. C. James. Evaluating inter-annual relative abundance of leatherback sea turtles in Atlantic Canada. Marine Ecology Progress Series, 547:233–246, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/547/
m547p233.pdf.

Aguzzi:2012:SFB

[AJF^{+}12] J. Aguzzi, A. J. Jamieson, T. Fujii, V. Sbragaglia, C. Costa, P. Menesatti, and Y. Fujiwara. Shifting feeding behaviour of deep-sea buccinid gastropods at natural and simulated food falls. Marine Ecology Progress Series, 458:247–253, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2012/458/m458p247.pdf.

Audzijonyte:2014:EEC

[AKF14] A. Audzijonyte, A. Kuparinen, and E. A. Fulton. Ecosystem effects of contemporary life-history changes are comparable to those of fishing. Marine Ecology Progress Series, 495:219–231, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/495/m495p219.pdf.

Albins:2012:IRL

[AL12a] M. A. Albins and P. J. Lyons. Invasive red lionfish Pterois volitans blow directed jets of water at prey fish. Marine Ecology Progress Series, 448:1–5, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anderson:2012:SDP

[AL12b] E. M. Anderson and J. R. Lovvorn. Seasonal dynamics of prey size mediate complementary functions of mussel beds and seagrass habitats for an avian predator. Marine Ecology Progress Series, 467:219–232, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/467/m467p219.pdf.

Acuna:2010:DFO

[ALANGT10] J. L. Acuña, M. López-Alvarez, E. Nogueira, and F. González-Taboada. Diatom flotation at the onset of the spring phytoplankton bloom: an in situ experiment. Marine Ecology Progress Series, 400:115–125, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/400/m400p115.pdf.

Albins:2015:IPL

[Alb15] M. A. Albins. Invasive Pacific lionfish Pterois volitans reduce abundance and species richness of native Bahamian coral-reef fishes. Marine Ecology Progress Series, 522:231– 243, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/522/m522p231.pdf.

Alexander:2013:CIS

[Ale13a] T. J. Alexander. Cryptic invertebrates on subtidal rocky reefs vary with microhabitat structure and protection from fishing. Marine Ecology Progress Series, 481:93–104, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/481/
m481p093.pdf.

Allen:2013:SIA

[ALE^{+}13b] C. D. Allen, G. E. Lemons, T. Eguchi, R. A. LeRoux, C. C. Fahy, P. H. Dutton, S. H. Peckham, and J. A. Seminoff. Stable isotope analysis reveals migratory origin of loggerhead turtles in the Southern California Bight. Marine Ecology Progress Series, 472:275–285, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/472/m472p275.pdf.

Adrian:2017:KAO

[ALK17] A. J. Adrian, C. E. Lack, and S. J. Kamel. Kin aggregations occur in eastern oyster Crassostrea virginica reefs despite limited regional genetic differentiation. Marine Ecology Progress Series, 584:79–90, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/584/m584p079.pdf.

Andersen:2017:DVI

[ALNB17] N. G. Andersen, B. Lundgren, S. Neuenfeldt, and J. E. Beyer. Diel vertical interactions between Atlantic cod Gadus morhua and sprat Sprattus sprattus in a stratified water column. Marine Ecology Progress Series, 583:195–209, 2017. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m583p195.pdf.

Asci:2018:ESB

[ALS18] S. C. Asci, R. W. Langton, and K. D. E. Stokesbury. Estimating similarity in benthic communities over decades and in areas open and closed to fishing in the central Gulf of Maine, USA. Marine Ecology Progress Series, 595:15–26, 2018. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m595p015.pdf.

Alfaro-Lucas:2018:TSC

[ALSO^{+}18] J. M. Alfaro-Lucas, M. Shimabukuro, I. V. Ogata, Y. Fujiwara, and P. Y. G. Sumida. Trophic structure and chemosynthesis contributions to heterotrophic fauna inhabiting an abyssal whale carcass. Marine Ecology Progress Series, 596: 1–12, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ameneiro:2016:FLI

[ALSV16] J. Ameneiro, L. M. Lubián, P. Sangrà, and E. Vázquez. Food-limited invertebrate larvae in the Southern Ocean: testing a paradigm. Marine Ecology Progress Series, 554:71– 80, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/554/m554p071.pdf.

Abeels:2012:TTH

[ALV12] H. A. Abeels, A. N. Loh, and A. K. Volety. Trophic transfer and habitat use of oyster Crassostrea virginica reefs in southwest Florida, identified by stable isotope analysis. Marine Ecology Progress Series, 462:125–142, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/462/
m462p125.pdf.

Agnew:2015:VBS

[ALWD15] P. Agnew, C. Lalas, J. Wright, and S. Dawson. Variation in breeding success and survival of little penguins Eudyptula minor in response to environmental variation. Marine Ecology Progress Series, 541:219–229, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/541/m541p219.pdf.

Aguirre:2013:HCM

[AM13] J. D. Aguirre and D. C. McNaught. Habitat complexity mediates predation of juvenile abalone by starfish. Marine Ecology Progress Series, 487:101–111, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/487/m487p101.pdf.

Aoki:2018:RED

[AM18] L. R. Aoki and K. J. McGlathery. Restoration enhances denitrification and DNRA in subsurface sediments of Zostera marina seagrass meadows. Marine Ecology Progress Series, 602: 87–102, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/602/m602p087.pdf.

Aoki:2019:HRF

[AM19] L. R. Aoki and K. J. McGlathery. High rates of N fixation in seagrass sediments measured via a direct 30 N_{2} push-pull method. Marine Ecology Progress Series, 616:1–11, 2019. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Auger-Methe:2017:SMM

[AMAJ^{+}17] M. Auger-Méthé, C. M. Albertsen, I. D. Jonsen, A. E. Derocher, D. C. Lidgard, K. R. Studholme, W. D. Bowen, G. T.

Crossin, and J. Mills Flemming. Spatiotemporal modelling of marine movement data using template model builder (TMB). Marine Ecology Progress Series, 565:237–249, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/565/
m565p237.pdf.

Amorim:2018:WRD

[AMAM^{+}18] K. Amorim, R. M. Mattmüller, M. Algueró-Muñiz, C. L. Meunier, S. Alvarez-Fernandez, M. Boersma, P. Morais, and M. A. Teodósio. Winter river discharge may affect summer estuarine jellyfish blooms. Marine Ecology Progress Series, 591:253–265, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/591/m591p253.pdf.

Aires:2011:EHS

[AMC^{+}11] T. Aires, N. Marbà, R. L. Cunha, G. A. Kendrick, D. I. Walker, E. A. Serrão, C. M. Duarte, and S. Arnaud-Haond. Evolutionary history of the seagrass genus Posidonia. Marine Ecology Progress Series, 421:117–130, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/421/m421p117.pdf.

Aznar:2017:LTC

[AMLR^{+}17] F. J. Aznar, R. Míguez-Lozano, B. Ruiz, A. B. de Castro, J. A. Raga, and C. Blanco. Long-term changes (1990– 2012) in the diet of striped dolphins Stenella coeruleoalba from the western Mediterranean. Marine Ecology Progress Series, 568:231–247, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/568/m568p231.pdf.

Aberle:2015:SLI

[AMLS15] N. Aberle, A. M. Malzahn, A. M. Lewandowska, and U. Sommer. Some like it hot: the protozooplankton–copepod link in a warming ocean. Marine Ecology Progress Series, 519: 103–113, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/519/m519p103.pdf.

Alvarez:2016:SDP

[ÁMLUN16] E. Álvarez, X. A. G. Morán, Á. López-Urrutia, and E. Nogueira. Size-dependent photoacclimation of the phytoplankton community in temperate shelf waters (southern Bay of Biscay). Marine Ecology Progress Series, 543:73–87, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/543/m543p073.pdf.

Allyn:2012:RAK

[AMM^{+}12] A. J. Allyn, A. McKnight, K. McGarigal, C. R. Griffin, K. J. Kuletz, and D. B. Irons. Relationships among Kittlitz’s murrelet habitat use, temperature-depth profiles, and landscape features in Prince William Sound, Alaska, USA. Marine Ecology Progress Series, 466:233–247, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/466/
m466p233.pdf.

Alos:2011:STP

[AMP^{+}11] J. Alós, D. March, M. Palmer, A. Grau, and B. Morales-Nin. Spatial and temporal patterns in Serranus cabrilla habitat use in the NW Mediterranean revealed by acoustic telemetry. Marine Ecology Progress Series, 427:173–186, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/427/
m427p173.pdf.

Albouy:2010:SCE

[AMR^{+}10] C. Albouy, D. Mouillot, D. Rocklin, J. M. Culioli, and F. Le Loc’h. Simulation of the combined effects of artisanal and recreational fisheries on a Mediterranean MPA ecosystem using a trophic model. Marine Ecology Progress Series, 412: 207–221, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/412/m412p207.pdf.

Alderson:2013:AIM

[AMS^{+}13] B. Alderson, D. Mazumder, N. Saintilan, K. Zimmerman, and P. Mulry. Application of isotope mixing models to discriminate dietary sources over small-scale patches in saltmarsh. Marine Ecology Progress Series, 487:113–122, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/487/
m487p113.pdf.

Acuna-Marrero:2018:SPD

[AMSdL^{+}18] D. Acuña-Marrero, A. N. H. Smith, P. Salinas de León, E. S. Harvey, M. D. M. Pawley, and M. J. Anderson. Spatial patterns of distribution and relative abundance of coastal shark species in the Galapagos marine reserve. Marine Ecology Progress Series, 593:73–95, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/593/m593p073.pdf.

Aguilera:2011:DAP

[AN11a] M. A. Aguilera and S. A. Navarrete. Distribution and activity patterns in an intertidal grazer assemblage: influence of temporal and spatial organization on interspecific associations. Marine Ecology Progress Series, 431:119–136, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/431/
m431p119.pdf.

Aiken:2011:EFA

[AN11b] C. M. Aiken and S. A. Navarrete. Environmental fluctuations and asymmetrical dispersal: generalized stability theory for studying metapopulation persistence and marine protected areas. Marine Ecology Progress Series, 428:77–88, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/428/
m428p077.pdf.

Aguilera:2013:DEG

[ANB13] M. A. Aguilera, S. A. Navarrete, and B. R. Broitman. Differential effects of grazer species on periphyton of a temperate rocky shore. Marine Ecology Progress Series, 484: 63–78, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/484/m484p063.pdf.

Noridi:2011:EIC

[áNGGS11] Gunnvor á Nordi, Ronnie N. Glud, Eilif Gaard, and Knud Simonsen. Environmental impacts of coastal fish farming: carbon and nitrogen budgets for trout farming in Kaldbaksfjørdur (Faroe Islands). Marine Ecology Progress Series, 431:223– 241, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/431/m431p223.pdf.

Andersen:2011:DPP

[ANJ^{+}11] N. G. Andersen, T. G. Nielsen, H. H. Jakobsen, P. Munk, and L. Riemann. Distribution and production of plankton communities in the subtropical convergence zone of the Sargasso Sea. II. protozooplankton and copepods. Marine Ecology Progress Series, 426:71–86, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/426/m426p071.pdf.

Anonymous:2010:CATa

[Ano10a] Anonymous. Combined author and title index: Volumes 391 to 400 (2009–2010). Marine Ecology Progress Series, 400:289–302, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/400/m400p289.pdf.

Anonymous:2010:CATb

[Ano10b] Anonymous. Combined author and title index: Volumes 401 to 410 (2010). Marine Ecology Progress Series, 410: 289–300, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/410/m410p289.pdf.

Anonymous:2010:CATc

[Ano10c] Anonymous. Combined author and title index: Volumes 411 to 420 (2010). Marine Ecology Progress Series, 420: 289–301, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/420/m420p289.pdf.

Anonymous:2010:EHM

[Ano10d] Anonymous. Erratum: Holt & Mantua, 393:285–299 (2009). Marine Ecology Progress Series, 398:305–306, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/398/
m398p305.pdf. See [HM09].

Anonymous:2010:EKA

[Ano10e] Anonymous. Erratum: Kerr et al., 349:245–253, (2007). Marine Ecology Progress Series, 417:301, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/417/
m417p301.pdf. See [KSK07].

Anonymous:2010:ESH

[Ano10f] Anonymous. Erratum: Sunda & Hardison, 401:63–76, 2010. Marine Ecology Progress Series, 410:288, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/410/
m410p288.pdf. See [SH10].

Anonymous:2012:CATa

[Ano12a] Anonymous. Combined author and title index: Volumes 451 to 460 (2012). Marine Ecology Progress Series, ??: ??, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/460/m460p289.pdf.

Anonymous:2012:CATb

[Ano12b] Anonymous. Combined author and title index: Volumes 461 to 470 (2012). Marine Ecology Progress Series, 470: 291–301, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m470p291.pdf.

Anonymous:2012:CHM

[Ano12c] Anonymous. Corrigendum: Hoque, A. T. M. R., Sharma, S., Suwa, R., Mori, S., Hagihara, A., 404:31–37, 2010. Marine Ecology Progress Series, 467:303, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/467/

m467p303.pdf. See [HSS^{+}10].

Anonymous:2012:EIA

[Ano12d] Anonymous. Errata: Ichii et al., 441:151–164 (2011) Tremblay-Boyer et al., 442:169–185 (2011). Marine Ecology Progress Series, 457:302–304, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m457p302.pdf. See

[IMS^{+}11, TBGW^{+}11].

Anonymous:2013:CATa

[Ano13a] Anonymous. Combined author and title index: Volumes 471 to 480 (2012–2013). Marine Ecology Progress Series, 480:289–301, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/480/m480p289.pdf.

Anonymous:2013:CATb

[Ano13b] Anonymous. Combined author and title index: Volumes 481 to 490 (2013). Marine Ecology Progress Series, 490: 295–308, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/490/m490p295.pdf.

Anonymous:2014:CAT

[Ano14a] Anonymous. Combined author and title index: Volumes 491 to 500 (2013–2014). Marine Ecology Progress Series, 500:291–304, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/500/m500p291.pdf.

Anonymous:2014:EHA

[Ano14b] Anonymous. Errata: Herfort et al., 438:19–31 (2011). Marine Ecology Progress Series, 502:303, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/502/

m502p303.pdf. See [HPMZ11].

Anonymous:2014:EDA

[Ano14c] Anonymous. Erratum: Davis et al., 500:175–186 (2014). Marine Ecology Progress Series, 502:304, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/502/
m502p304.pdf. See [DMS14].

Anonymous:2015:EKA

[Ano15a] Anonymous. Errata: Kawabata et al., 534:149–161, 2015. Marine Ecology Progress Series, 536:281, 2015. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m536p281.pdf. See [KNY^{+}15].

Anonymous:2015:EOA

[Ano15b] Anonymous. Erratum: Oksanen et al., 507:297–308, 2014. Marine Ecology Progress Series, 536:282, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/536/

m536p282.pdf. See [OALK14].

Anonymous:2016:EKA

[Ano16] Anonymous. Erratum: Keppel et al., 538:169–183, 2015. Marine Ecology Progress Series, 547:287, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/547/

m547p287.pdf. See [KBW^{+}15].

Anonymous:2017:CTS

[Ano17a] Anonymous. Complete theme section in pdf format. Marine Ecology Progress Series, 576:105–234, 2017. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m576_ThemeSection.pdf.

Anonymous:2017:CP

[Ano17b] Anonymous. Contents page. Marine Ecology Progress Series, 576:105–??, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m576p105.pdf.

Anonymous:2017:TSS

[Ano17c] Anonymous. This theme section showcases the latest research on individual variability in foraging and migration in seabirds. the studies thus contribute to our understanding of the causes and consequences of movement, dietary and habitat variability among individuals, using a range of methods for quantification and analysis. these studies provide fundamental insights into key factors such as individual specialisation, plasticity in response to environmental change and seasonal carry-over effects. Marine Ecology Progress Series, ??:??, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2018:TSE

[Ano18a] Anonymous. This theme section explores how seabirds inhabiting the southeastern Bering Sea region respond to oceanographic variability, including the timing of sea-ice retreat. it combines both colony-based and pelagic studies of seabirds, spanning a period of 100 years and an immense oceanographic region. Marine Ecology Progress Series, ??:??, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2018:TSR

[Ano18b] Anonymous. This theme section represents a selection of studies exploring the innovative use of sclerochronology in resource management. topics covered showcase the potential for sclerochronology to provide a deeper understanding of the interaction between marine life and its environment. Marine Ecology Progress Series, ??:??, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2019:CHAb

[Ano19a] Anonymous. Corrigendum: Hagerty et al., 595:105–122 (2018). Marine Ecology Progress Series, 267:233, 2019. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m627p233.pdf. See [HRP18].

Anonymous:2019:CHAa

[Ano19b] Anonymous. Corrigendum: Hawkes et al., 614:91–109 (2019). Marine Ecology Progress Series, 622:231, 2019. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m622p231.pdf. See [HKB^{+}19a].

Anonymous:2019:CMA

[Ano19c] Anonymous. Corrigendum: Milewski et al., 598:85–97 (2018). Marine Ecology Progress Series, 625:233, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/625/

m625p233.pdf. See [MLF^{+}18].

Anonymous:2019:CWA

[Ano19d] Anonymous. Corrigendum: Weinstein et al., 559:45–63 (2016). Marine Ecology Progress Series, 619:215–217, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m619p215.pdf. See [WSK^{+}16].

Anonymous:2021:CJA

[Ano21] Anonymous. Corrigendum: Jusufovski et al., 627:207– 232 (2019). Marine Ecology Progress Series, 658:253, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/658/m658p253.pdf. See [JSK19].

Arendt:2010:DPC

[ANRT10] K. E. Arendt, T. G. Nielsen, S. Rysgaard, and K. Tönnesson. Differences in plankton community structure along the Godthåbsfjord, from the Greenland Ice Sheet to offshore waters. Marine Ecology Progress Series, 401:49–62, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2009/401/m401p049.pdf.

Aleksa:2018:DDM

[ANWG18] K. T. Aleksa, R. W. Nero, J. D. Wiggert, and W. M. Graham. Descriptive density models of scyphozoan jellyfish in the northern Gulf of Mexico. Marine Ecology Progress Series, 591:71–85, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/591/m591p071.pdf.

Aburto-Oropeza:2010:CIR

[AOPMOS10] O. Aburto-Oropeza, G. Paredes, I. Mascareñas-Osorio, and E. Sala. Climatic influence on reef fish recruitment and fisheries. Marine Ecology Progress Series, 410:283–287, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2010/410/m410p283.pdf.

Ashforth:2011:PCB

[AOW11] E. J. Ashforth, P. J. W. Olive, and A. C. Ward. Phylogenetic characterisation of bacterial assemblages and the role of sulphur-cycle bacteria in an Arenicola marina bioturbated mesocosm. Marine Ecology Progress Series, 439: 19–30, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/439/m439p019.pdf.

Alexandridis:2012:MEA

[AOW12] N. Alexandridis, A. Oschlies, and M. Wahl. Modeling the effects of abiotic and biotic factors on the depth distribution of Fucus vesiculosus in the Baltic Sea. Marine Ecology Progress Series, 463:59–72, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/463/m463p059.pdf.

Anderson:2014:VSF

[AP14] K. D. Anderson and M. S. Pratchett. Variation in sizefrequency distributions of branching corals between a tropical versus sub-tropical reef. Marine Ecology Progress Series, 502:117–128, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/502/m502p117.pdf.

Albo-Puigserver:2019:TER

[APBC^{+}19] M. Albo-Puigserver, D. Borme, M. Coll, V. Tirelli, I. Palomera, and J. Navarro. Trophic ecology of range-expanding round sardinella and resident sympatric species in the NW Mediterranean. Marine Ecology Progress Series, 620:139– 154, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/620/m620p139.pdf.

Ackiss:2013:PGS

[APC^{+}13] A. S. Ackiss, S. Pardede, E. D. Crandall, M. C. A. Ablan-Lagman, Ambariyanto, N. Romena, P. H. Barber, and K. E. Carpenter. Pronounced genetic structure in a highly mobile coral reef fish, Caesio cuning, in the coral triangle. Marine Ecology Progress Series, 480:185–197, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/480/
m480p185.pdf.

Alonso-Perez:2014:BON

[APC14a] F. Alonso-Pérez and C. G. Castro. Benthic oxygen and nutrient fluxes in a coastal upwelling system (Ria de Vigo, NW Iberian Peninsula): seasonal trends and regulating factors. Marine Ecology Progress Series, 511:17–32, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/511/
m511p017.pdf.

Alos:2014:SES

[APC^{+}14b] J. Alós, M. Palmer, I. A. Catalan, A. Alonso-Fernández, G. Basterretxea, A. Jordi, L. Buttay, B. Morales-Nin, and R. Arlinghaus. Selective exploitation of spatially structured coastal fish populations by recreational anglers may lead to evolutionary downsizing of adults. Marine Ecology Progress Series, 503:219–233, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m503p219.pdf.

Allen:2010:ZST

[APH^{+}10] C. J. Allen, G. L. J. Paterson, L. E. Hawkins, C. Hauton, P. F. Clark, and C. Aryuthaka. Zonation on sandy tropical beaches: a case study using Dotilla intermedia (Brachyura: Ocypodidae). Marine Ecology Progress Series, 408:97–107, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2010/408/m408p097.pdf.

Albo-Puigserver:2015:FET

[APNC^{+}15] M. Albo-Puigserver, J. Navarro, M. Coll, J. Aguzzi, L. Cardona, and R. Sáez-Liante. Feeding ecology and trophic position of three sympatric demersal chondrichthyans in the northwestern Mediterranean. Marine Ecology Progress Series, 524:255–268, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/524/m524p255.pdf.

Antonio:2014:TTD

[AR14] E. S. Antonio and N. B. Richoux. Trophodynamics of three decapod crustaceans in a temperate estuary using stable isotope and fatty acid analyses. Marine Ecology Progress Series, 504:193–205, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/504/m504p193.pdf.

Arnott:2010:SST

[ARA^{+}10] S. A. Arnott, W. A. Roumillat, J. A. Archambault, C. A. Wenner, J. I. Gerhard, T. L. Darden, and M. R. Denson. Spatial synchrony and temporal dynamics of juvenile red drum Sciaenops ocellatus populations in South Carolina, USA. Marine Ecology Progress Series, 415:221–236, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/415/
m415p221.pdf.

Adamack:2012:SEH

[ARB^{+}12] A. T. Adamack, K. A. Rose, D. L. Breitburg, A. J. Nice, and W. S. Lung. Simulating the effect of hypoxia on bay anchovy egg and larval mortality using coupled watershed, water quality, and individual-based predation models. Marine Ecology Progress Series, 445:141–160, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/445/m445p141.pdf.

Allen:2019:LCC

[ARD^{+}19] J. D. Allen, E. L. Richardson, D. Deaker, A. Agüera, and M. Byrne. Larval cloning in the crown-of-thorns sea star, a keystone coral predator. Marine Ecology Progress Series, 609:271–276, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m609p271.pdf.

Archer:2010:ERE

[ARG^{+}10] F. I. Archer, J. V. Redfern, T. Gerrodette, S. J. Chivers, and W. F. Perrin. Estimation of relative exposure of dolphins to fishery activity. Marine Ecology Progress Series, 410:245–255, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/410/m410p245.pdf.

Amo:2013:ODD

[ARGB13] L. Amo, M. Á. Rodríguez-Gironés, and A. Barbosa. Olfactory detection of dimethyl sulphide in a krill-eating Antarctic penguin. Marine Ecology Progress Series, 474:277–285, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/474/m474p277.pdf.

Aguilar:2015:GDB

[ARMA15] L. A. Aguilar, D. G. Roberts, T. E. Minchinton, and D. J. Ayre. Genetic differentiation in the barnacle Catomerus polymerus despite migration across a biogeographic barrier. Marine Ecology Progress Series, 524:213–224, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/524/
m524p213.pdf.

Andreotti:2016:IMR

[ARvdW^{+}16] S. Andreotti, M. Rutzen, S. van der Walt, S. Von der Heyden, R. Henriques, M. Meÿer, H. Oosthuizen, and C. A. Matthee. An integrated mark-recapture and genetic approach to estimate the population size of white sharks in South Africa. Marine Ecology Progress Series, 552:241–253, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/552/m552p241.pdf. See comment [INK^{+}17] and

reply [AvdHH^{+}17].

Ainley:2012:AOS

[ARW12] D. G. Ainley, C. A. Ribic, and E. J. Woehler. Adding the ocean to the study of seabirds: a brief history of at-sea seabird research. Marine Ecology Progress Series, 451:231–243, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m451p231.pdf.

Apostolidis:2012:WSI

[AS12] C. Apostolidis and Konstantinos I. Stergiou. As we see it: Fish ingredients in online recipes do not promote the sustainable use of vulnerable taxa. Marine Ecology Progress Series, 465:299–304, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m465p299.pdf.

Alsterberg:2013:EWT

[AS13] C. Alsterberg and K. Sundbäck. Experimental warming and toxicant exposure can result in antagonistic effects in a shallow-water sediment system. Marine Ecology Progress Series, 488:89–101, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/488/m488p089.pdf.

Alestra:2014:EOA

[AS14] T. Alestra and D. R. Schiel. Effects of opportunistic algae on the early life history of a habitat-forming fucoid: influence of temperature, nutrient enrichment and grazing pressure. Marine Ecology Progress Series, 508:105–115, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/508/
m508p105.pdf.

Alestra:2015:NTR

[AS15] T. Alestra and D. R. Schiel. Non-trophic responses of algal communities to nutrient enrichment: interactions among coralline turfs, ephemeral algae and perennial fucoids. Marine Ecology Progress Series, 538:145–156, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/538/
m538p145.pdf.

Akmajian:2017:YRA

[ASAG17] A. M. Akmajian, J. J. Scordino, and A. Acevedo-Gutiérrez. Year-round algal toxin exposure in free-ranging sea lions. Marine Ecology Progress Series, 583:243–258, 2017. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m583p243.pdf.

Arbeider:2019:IPD

[ASCHM19] M. Arbeider, C. Sharpe, C. Carr-Harris, and J. W. Moore. Integrating prey dynamics, diet, and biophysical factors across an estuary seascape for four fish species. Marine Ecology Progress Series, 613:151–169, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m613p151.pdf.

Ayon:2011:LTC

[ASEB11] P. Ayón, G. Swartzman, P. Espinoza, and A. Bertrand. Longterm changes in zooplankton size distribution in the Peruvian Humboldt Current System: conditions favouring sardine or anchovy. Marine Ecology Progress Series, 422:211– 222, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/422/m422p211.pdf.

Ayers:2013:MEE

[ASF13] M. J. Ayers, U. M. Scharler, and S. T. Fennessy. Modelling ecosystem effects of reduced prawn recruitment on the Thukela Bank trawling grounds, South Africa, following nursery loss. Marine Ecology Progress Series, 479:143–161, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/479/
m479p143.pdf.

Andrade:2015:ELC

[ASI15] H. Andrade, J. Santos, and M. J. Ixquiac. Ecological linkages in a Caribbean estuary bay. Marine Ecology Progress Series, 533:29–46, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/533/m533p029.pdf.

Attard:2015:BOE

[ASK^{+}15] K. M. Attard, H. Stahl, N. A. Kamenos, G. Turner, H. L. Burdett, and R. N. Glud. Benthic oxygen exchange in a live coralline algal bed and an adjacent sandy habitat: an eddy covariance study. Marine Ecology Progress Series, 535: 99–115, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m535p099.pdf.

Arnold:2010:RGI

[ASM10] S. N. Arnold, R. S. Steneck, and P. J. Mumby. Running the gauntlet: inhibitory effects of algal turfs on the processes of coral recruitment. Marine Ecology Progress Series, 414:91–105, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/414/m414p091.pdf.

Adamack:2012:PEF

[ASM^{+}12] A. T. Adamack, C. A. Stow, D. M. Mason, L. P. Rozas, and T. J. Minello. Predicting the effects of freshwater diversions on juvenile brown shrimp growth and production: a Bayesian-based approach. Marine Ecology Progress Series, 444:155–173, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m444p155.pdf.

Abreu:2019:SDA

[ASR^{+}19] J. Abreu, I. Staniland, C. F. Rodrigues, J. P. Queirós, J. M. Pereira, and J. C. Xavier. Squid in the diet of Antarctic fur seals: potential links to oceanographic conditions and Antarctic krill abundance. Marine Ecology Progress Series, 628:211–221, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/628/m628p211.pdf.

Aguzzi:2013:DAR

[ASS^{+}13] J. Aguzzi, V. Sbragaglia, G. Santamaría, J. Del Río, F. Sardà, M. Nogueras, and A. Manuel. Daily activity rhythms in temperate coastal fishes: insights from cabled observatory video monitoring. Marine Ecology Progress Series, 486:223– 236, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/486/m486p223.pdf.

Alves-Stanley:2010:FPW

[ASWB10] C. D. Alves-Stanley, G. A. J. Worthy, and R. K. Bonde. Feeding preferences of West Indian manatees in Florida, Belize, and Puerto Rico as indicated by stable isotope analysis. Marine Ecology Progress Series, 402:255–267, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/402/
m402p255.pdf.

Albaina:2012:MDP

[ATF12] A. Albaina, M. I. Taylor, and C. J. Fox. Molecular detection of plaice remains in the stomachs of potential predators on a flatfish nursery ground. Marine Ecology Progress Series, 444:223–238, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/444/m444p223.pdf.

Appelhans:2014:JSS

[ATO^{+}14] Y. S. Appelhans, J. Thomsen, S. Opitz, C. Pansch, F. Melzner, and M. Wahl. Juvenile sea stars exposed to acidification decrease feeding and growth with no acclimation potential. Marine Ecology Progress Series, 509:227–239, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/509/
m509p227.pdf.

Appelhans:2012:STS

[ATP^{+}12] Y. S. Appelhans, J. Thomsen, C. Pansch, F. Melzner, and M. Wahl. Sour times: seawater acidification effects on growth, feeding behaviour and acid-base status of Asterias rubens and Carcinus maenas. Marine Ecology Progress Series, 459:

85–98, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m459p085.pdf.

Aquino-Thomas:2014:OCV

[ATP14] J. Aquino-Thomas and C. E. Proffitt. Oysters Crassostrea virginica on red mangrove Rhizophora mangle prop roots: facilitation of one foundation species by another. Marine Ecology Progress Series, 503:177–194, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/503/m503p177.pdf.

Alestra:2014:EAT

[ATS14] T. Alestra, L. W. Tait, and D. R. Schiel. Effects of algal turfs and sediment accumulation on replenishment and primary productivity of fucoid assemblages. Marine Ecology Progress Series, 511:59–70, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/511/m511p059.pdf.

Alos:2016:FSG

[ATT^{+}16] J. Alós, F. Tomas, J. Terrados, H. Verbruggen, and E. Ballesteros. Fast-spreading green beds of recently introduced Halimeda incrassata invade Mallorca island (NW Mediterranean Sea). Marine Ecology Progress Series, 558:153–158, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/558/m558p153.pdf.

Aguilera:2013:SND

[AVB13] M. A. Aguilera, N. Valdivia, and B. R. Broitman. Spatial niche differentiation and coexistence at the edge: co-occurrence distribution patterns in Scurria limpets. Marine Ecology Progress Series, 483:185–198, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/483/m483p185.pdf.

Aarts:2016:HPM

[AvBBL^{+}16] G. Aarts, A. M. von Benda-Beckmann, K. Lucke, H. Ö. Sertlek, R. van Bemmelen, S. C. V. Geelhoed, S. Brasseur, M. Scheidat, F. P. A. Lam, H. Slabbekoorn, and R. Kirkwood. Harbour porpoise movement strategy affects cumulative number of animals acoustically exposed to underwater explosions. Marine Ecology Progress Series, 557:261–275, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m557p261.pdf.

Andreotti:2017:ESC

[AvdHH^{+}17] S. Andreotti, S. von der Heyden, R. Henriques, M. Rutzen, M. Meÿer, and C. A. Matthee. Erring on the side of caution: Reply to Irion et al. (2017). Marine Ecology Progress Series, 577:257–262, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/577/m577p257.pdf. See

[ARvdW^{+}16, INK^{+}17].

Alderkamp:2019:EIL

[AvDL^{+}19] A. C. Alderkamp, G. L. van Dijken, K. E. Lowry, K. M. Lewis, H. L. Joy-Warren, W. van de Poll, P. Laan, L. Gerringa, T. O. Delmont, B. D. Jenkins, and K. R. Arrigo. Effects of iron and light availability on phytoplankton photosynthetic properties in the Ross Sea. Marine Ecology Progress Series, 621:33–50, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m621p033.pdf.

Asher:2019:SBD

[AWH19] J. Asher, I. D. Williams, and E. S. Harvey. Is seeing believing? Diver and video-based censuses reveal inconsistencies in roving predator estimates between regions. Marine Ecology Progress Series, 630:115–136, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m630p115.pdf.

Anderson:2012:PGV

[AWJ12] R. Anderson, C. Winter, and K. Jürgens. Protist grazing and viral lysis as prokaryotic mortality factors at Baltic Sea oxic– anoxic interfaces. Marine Ecology Progress Series, 467:1–14, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Arellano:2010:PPS

[AY10] S. M. Arellano and C. M. Young. Pre- and post-settlement factors controlling spatial variation in recruitment across a cold-seep mussel bed. Marine Ecology Progress Series, 414:

131–144, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/414/m414p131.pdf.

Armoza-Zvuloni:2011:RBE

[AZSKWL11] R. Armoza-Zvuloni, R. Segal, E. Kramarsky-Winter, and Y. Loya. Repeated bleaching events may result in high tolerance and notable gametogenesis in stony corals: Oculina patagonica as a model. Marine Ecology Progress Series, 426:149–159, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/426/m426p149.pdf.

Afeworki:2013:FIP

[AZVB13] Y. Afeworki, Z. A. Zekeria, J. J. Videler, and J. H. Bruggemann. Food intake by the parrotfish Scarus ferrugineus varies seasonally and is determined by temperature, size and territoriality. Marine Ecology Progress Series, 489:213– 224, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/489/m489p213.pdf.

Barbour:2012:BEM

[BA12] A. B. Barbour and A. J. Adams. Biologging to examine multiple life stages of an estuarine-dependent fish. Marine Ecology Progress Series, 457:241–250, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m457p241.pdf.

Babarro:2013:CET

[BA13a] J. M. F. Babarro and M. J. Abad. Co-existence of two mytilid species in a heterogeneous environment: mortality, growth and strength of shell and byssus attachment. Marine Ecology Progress Series, 476:115–128, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/476/m476p115.pdf.

Bejarano:2013:STF

[BA13b] I. Bejarano and R. S. Appeldoorn. Seawater turbidity and fish communities on coral reefs of Puerto Rico. Marine Ecology Progress Series, 474:217–226, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/474/m474p217.pdf.

Baird:2012:EIS

[BAA12] D. Baird, H. Asmus, and R. Asmus. Effect of invasive species on the structure and function of the Sylt–Rømø Bight ecosystem, northern Wadden Sea, over three time periods. Marine Ecology Progress Series, 462:143–161, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/462/
m462p143.pdf.

Bulleri:2011:DCR

[BAC^{+}11] F. Bulleri, T. Alestra, G. Ceccherelli, L. Tamburello, S. Pinna, N. Sechi, and L. Benedetti-Cecchi. Determinants of Caulerpa racemosa distribution in the north-western Mediterranean. Marine Ecology Progress Series, 431:55–67, 2011. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m431p055.pdf.

Blanchet-Aurigny:2012:MDC

[BADH^{+}12] A. Blanchet-Aurigny, S. F. Dubois, C. Hily, S. Rochette, E. Le Goaster, and M. Guillou. Multi-decadal changes in two cooccurring ophiuroid populations. Marine Ecology Progress Series, 460:79–90, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m460p079.pdf.

Blanchet-Aurigny:2015:TNT

[BADQ^{+}15] A. Blanchet-Aurigny, S. F. Dubois, C. Quéré, M. Guillou, and F. Pernet. Trophic niche of two co-occurring ophiuroid species in impacted coastal systems, derived from fatty acid and stable isotope analyses. Marine Ecology Progress Series, 525:127–141, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/525/m525p127.pdf.

Brokovich:2010:GPC

[BAE^{+}10] E. Brokovich, I. Ayalon, S. Einbinder, N. Segev, Y. Shaked, A. Genin, S. Kark, and M. Kiflawi. Grazing pressure on coral reefs decreases across a wide depth gradient in the Gulf of Aqaba, Red Sea. Marine Ecology Progress Series, 399: 69–80, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/399/m399p069.pdf.

Bode:2012:SRF

[BAFB12] M. Bode, P. R. Armsworth, H. E. Fox, and L. Bode. Surrogates for reef fish connectivity when designing marine protected area networks. Marine Ecology Progress Series, 466: 155–166, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/466/m466p155.pdf.

Bayer:2013:BGE

[BAFP^{+}13] T. Bayer, C. Arif, C. Ferrier-Pagès, D. Zoccola, M. Aranda, and C. R. Voolstra. Bacteria of the genus Endozoicomonas dominate the microbiome of the Mediterranean gorgonian coral Eunicella cavolini. Marine Ecology Progress Series, 479: 75–84, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m479p075.pdf.

Botta:2015:BCR

[BAH^{+}15] S. Botta, C. Albuquerque, A. A. Hohn, V. M. F. da Silva, M. C. O. Santos, C. Meirelles, L. Barbosa, A. P. M. Di Beneditto, R. M. A. Ramos, C. Bertozzi, M. J. Cremer, V. Franco-Trecu, N. Miekeley, and E. R. Secchi. Ba/Ca ratios in teeth reveal habitat use patterns of dolphins. Marine Ecology Progress Series, 521:249–263, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/521/
m521p249.pdf.

Bravo:2011:RAS

[BAL^{+}11] M. Bravo, J. C. Astudillo, D. Lancellotti, G. Luna-Jorquera, N. Valdivia, and M. Thiel. Rafting on abiotic substrata: properties of floating items and their influence on community succession. Marine Ecology Progress Series, 439:1–17, 2011. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Barbour:2014:ECS

[BAL14a] A. B. Barbour, A. J. Adams, and K. Lorenzen. Emigrationcorrected seasonal survival of a size-structured fish population in a nursery habitat. Marine Ecology Progress Series, 514:191–205, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/514/m514p191.pdf.

Barbour:2014:SBS

[BAL14b] A. B. Barbour, A. J. Adams, and K. Lorenzen. Size-based, seasonal, and multidirectional movements of an estuarine fish species in a habitat mosaic. Marine Ecology Progress Series, 507:263–276, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/507/m507p263.pdf.

Bertocci:2010:CFS

[BAM^{+}10] I. Bertocci, F. Arenas, M. Matias, S. Vaselli, R. Araújo, H. Abreu, R. Pereira, R. Vieira, and I. Sousa-Pinto. Canopyforming species mediate the effects of disturbance on macroalgal assemblages on Portuguese rocky shores. Marine Ecology Progress Series, 414:107–116, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/414/m414p107.pdf.

Baird:2018:ETS

[BÁNC^{+}18] A. H. Baird, M. Álvarez-Noriega, V. R. Cumbo, S. R. Connolly, M. Dornelas, and J. S. Madin. Effects of tropical storms on the demography of reef corals. Marine Ecology Progress Series, 606:29–38, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/606/m606p029.pdf.

Bustnes:2013:CNB

[BANE^{+}13] J. O. Bustnes, T. Anker-Nilssen, K. E. Erikstad, S. H. Lorentsen, and G. H. Systad. Changes in the Norwegian breeding population of European shag correlate with forage fish and climate. Marine Ecology Progress Series, 489:235– 244, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/489/m489p235.pdf.

Barnes:2013:SSM

[Bar13] R. S. K. Barnes. Spatial stability of macrobenthic seagrass biodiversity. Marine Ecology Progress Series, 493:127– 139, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/493/m493p127.pdf.

Barnes:2014:SUS

[Bar14] R. S. K. Barnes. Is spatial uniformity of soft-sediment biodiversity widespread and, if so, over what scales? Marine Ecology Progress Series, 504:147–158, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/504/
m504p147.pdf.

Barnett:2010:PPR

[BAS^{+}10] A. Barnett, K. Abrantes, J. D. Stevens, J. L. Yick, S. D. Frusher, and J. M. Semmens. Predator–prey relationships and foraging ecology of a marine apex predator with a wide temperate distribution. Marine Ecology Progress Series, 416: 189–200, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/416/m416p189.pdf.

Bethoney:2016:IEH

[BAS16] N. D. Bethoney, S. Asci, and K. D. E. Stokesbury. Implications of extremely high recruitment events into the US sea scallop fishery. Marine Ecology Progress Series, 547:137– 147, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/547/m547p137.pdf.

Bjaerke:2014:PCC

[BAT14] O. Bjærke, T. Andersen, and J. Titelman. Predator chemical cues increase growth and alter development in nauplii of a marine copepod. Marine Ecology Progress Series, 510: 15–24, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m510p015.pdf.

Bertocci:2011:MPU

[BAVSP11] I. Bertocci, R. Araújo, S. Vaselli, and I. Sousa-Pinto. Marginal populations under pressure: spatial and temporal heterogeneity of Ascophyllum nodosum and associated assemblages affected by human trampling in Portugal. Marine Ecology Progress Series, 439:73–82, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/439/m439p073.pdf.

Burman:2012:BHC

[BAvW12] S. G. Burman, R. B. Aronson, and R. van Woesik. Biotic homogenization of coral assemblages along the Florida reef tract. Marine Ecology Progress Series, 467:89–96, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/467/
m467p089.pdf.

Boudreau:2011:TIT

[BAW11] S. A. Boudreau, S. C. Anderson, and B. Worm. Top-down interactions and temperature control of snow crab abundance in the northwest Atlantic Ocean. Marine Ecology Progress Series, 429:169–183, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m429p169.pdf.

Bennett:2011:LVM

[BB11] S. Bennett and D. R. Bellwood. Latitudinal variation in macroalgal consumption by fishes on the Great Barrier Reef. Marine Ecology Progress Series, 426:241–252, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/426/
m426p241.pdf.

Barnes:2014:SUB

[BB14a] R. S. K. Barnes and M. K. S. Barnes. Spatial uniformity of biodiversity is inevitable if the available species are distributed independently of each other. Marine Ecology Progress Series, 516:263–266, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/516/m516p263.pdf.

Braid:2014:FEL

[BB14b] H. E. Braid and K. S. R. Bolstad. Feeding ecology of the largest mastigoteuthid squid species, Idioteuthis cordiformis (Cephalopoda, Mastigoteuthidae). Marine Ecology Progress Series, 515:275–279, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2014/515/m515p275.pdf.

Breitburg:2014:PML

[BB14c] D. Breitburg and R. Burrell. Predator-mediated landscape structure: seasonal patterns of spatial expansion and prey control by Chrysaora quinquecirrha and Mnemiopsis leidyi. Marine Ecology Progress Series, 510:183–200, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m510p183.pdf.

Beermann:2015:FMP

[BB15] J. Beermann and K. Boos. Flexible microhabitat partitioning between hemi-sessile congeners. Marine Ecology Progress Series, 520:143–151, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/520/m520p143.pdf.

Bateman:2017:ECT

[BB17] D. C. Bateman and M. J. Bishop. The environmental context and traits of habitat-forming bivalves influence the magnitude of their ecosystem engineering. Marine Ecology Progress Series, 563:95–110, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/563/m563p095.pdf.

Buba:2019:NED

[BB19] Y. Buba and J. Belmaker. Native-exotic diversity relationships for Eastern Mediterranean fishes reveal a weak pattern of interactions. Marine Ecology Progress Series, 611: 215–220, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/611/m611p215.pdf.

Bauman:2014:PCS

[BBB^{+}14] A. G. Bauman, A. H. Baird, J. A. Burt, M. S. Pratchett, and D. A. Feary. Patterns of coral settlement in an extreme environment: the southern Persian Gulf (Dubai, United Arab Emirates). Marine Ecology Progress Series, 499:115– 126, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/499/m499p115.pdf.

Berry:2015:CMD

[BBB^{+}15] O. Berry, C. Bulman, M. Bunce, M. Coghlan, D. C. Murray, and R. D. Ward. Comparison of morphological and DNA metabarcoding analyses of diets in exploited marine fishes. Marine Ecology Progress Series, 540:167–181, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/540/
m540p167.pdf.

Bauer:2015:ASM

[BBBF15] R. K. Bauer, S. Bonhommeau, B. Brisset, and J. M. Fromentin. Aerial surveys to monitor bluefin tuna abundance and track efficiency of management measures. Marine Ecology Progress Series, 534:221–234, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/534/m534p221.pdf.

Bryan-Brown:2017:PTM

[BBBHC17] D. N. Bryan-Brown, C. J. Brown, J. M. Hughes, and R. M. Connolly. Patterns and trends in marine population connectivity research. Marine Ecology Progress Series, 585:243– 256, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/585/m585p243.pdf.

Benoit-Bird:2013:FBN

[BBBNT13] K. J. Benoit-Bird, B. C. Battaile, C. A. Nordstrom, and A. W. Trites. Foraging behavior of northern fur seals closely matches the hierarchical patch scales of prey. Marine Ecology Progress Series, 479:283–302, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m479p283.pdf.

Boerder:2017:ITF

[BBBW17] K. Boerder, A. Bryndum-Buchholz, and B. Worm. Interactions of tuna fisheries with the Galápagos marine reserve. Marine Ecology Progress Series, 585:1–15, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Bruno:2010:PPE

[BBC^{+}10] A. Bruno, L. Bruno, A. Chiappetta, D. Giannino, and M. B. Bitonti. PoCHL P expression pattern in Posidonia oceanica is related to critical light conditions. Marine Ecology Progress Series, 415:61–71, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/415/m415p061.pdf.

Bodkin:2012:LTE

[BBC^{+}12a] James L. Bodkin, Brenda E. Ballachey, Heather A. Coletti, George G. Esslinger, Kimberly A. Kloecker, Stanley D. Rice, John A. Reed, and Daniel H. Monson. Long-term effects of the ‘Exxon Valdez’ oil spill: sea otter foraging in the intertidal as a pathway of exposure to lingering oil. Marine Ecology Progress Series, 447:273–287, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m447p273.pdf. See

comment [HG13].

Brown:2012:ACC

[BBC12b] E. E. Brown, H. Baumann, and D. O. Conover. Absence of countergradient and cogradient variation in an oceanic silverside, the California grunion Leuresthes tenuis. Marine Ecology Progress Series, 461:175–186, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/461/m461p175.pdf.

Ballorain:2013:SDB

[BBC^{+}13] K. Ballorain, J. Bourjea, S. Ciccione, A. Kato, N. Hanuise, M. Enstipp, S. Fossette, and J. Y. Georges. Seasonal diving behaviour and feeding rhythms of green turtles at Mayotte Island. Marine Ecology Progress Series, 483:289–302, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/483/m483p289.pdf.

Bulleri:2014:CFC

[BBC14] F. Bulleri and L. Benedetti-Cecchi. Chasing fish and catching data: recreational spearfishing videos as a tool for assessing the structure of fish assemblages on shallow rocky reefs. Marine Ecology Progress Series, 506:255–265, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m506p255.pdf.

Bradshaw:2017:STS

[BBC^{+}17] P. J. Bradshaw, A. C. Broderick, C. Carreras, R. Inger, W. Fuller, R. Snape, K. L. Stokes, and B. J. Godley. Satellite tracking and stable isotope analysis highlight differential recruitment among foraging areas in green turtles. Marine Ecology Progress Series, 582:201–214, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m582p201.pdf.

Bjorndal:2019:GTS

[BBC19] K. A. Bjorndal, A. B. Bolten, and M. Chaloupka. Green turtle somatic growth dynamics: distributional regression reveals effects of differential emigration. Marine Ecology Progress Series, 616:185–195, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m616p185.pdf.

Benke:2014:BSH

[BBD^{+}14] H. Benke, S. Bräger, M. Dähne, A. Gallus, S. Hansen, C. G. Honnef, M. Jabbusch, J. C. Koblitz, K. Krügel, A. Liebschner, I. Narberhaus, and U. K. Verfuß. Baltic Sea harbour porpoise populations: status and conservation needs derived from recent survey results. Marine Ecology Progress Series, 495:275–290, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m495p275.pdf.

Boman:2019:DGJ

[BBdG^{+}19] E. M. Boman, T. Bervoets, M. de Graaf, J. Dewenter, A. Maitz, M. P. Meijer Zu Schlochtern, J. Stapel, A. C. Smaal, and L. A. J. Nagelkerke. Diet and growth of juvenile queen conch Lobatus gigas (Gastropoda: Strombidae) in native, mixed and invasive seagrass habitats. Marine Ecology Progress Series, 621:143–154, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/621/m621p143.pdf.

Barahona:2019:EDF

[BBF^{+}19] M. Barahona, B. R. Broitman, S. Faugeron, L. Jaugeon, A. Ospina-Alvarez, D. Véliz, and S. A. Navarrete. Environmental and demographic factors influence the spatial genetic structure of an intertidal barnacle in central-northern Chile.

Marine Ecology Progress Series, 612:151–165, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/612/
m612p151.pdf.

Bo:2011:FDA

[BBG^{+}11] M. Bo, A. C. Baker, E. Gaino, H. H. Wirshing, F. Scoccia, and G. Bavestrello. First description of algal mutualistic endosymbiosis in a black coral (Anthozoa: Antipatharia). Marine Ecology Progress Series, 435:1–11, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Benoit-Bird:2012:CNB

[BBG12] K. J. Benoit-Bird and W. F. Gilly. Coordinated nocturnal behavior of foraging jumbo squid Dosidicus gigas. Marine Ecology Progress Series, 455:211–228, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m455p211.pdf.

Bonaglia:2013:ERM

[BBG^{+}13] S. Bonaglia, M. Bartoli, J. S. Gunnarsson, L. Rahm, C. Raymond, O. Svensson, S. Shakeri Yekta, and V. Brüchert. Effect of reoxygenation and Marenzelleria spp. bioturbation on Baltic Sea sediment metabolism. Marine Ecology Progress Series, 482:43–55, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m482p043.pdf.

Brown:2016:QES

[BBG^{+}16] T. A. Brown, S. T. Belt, M. Gosselin, M. Levasseur, M. Poulin, and C. J. Mundy. Quantitative estimates of sinking sea ice particulate organic carbon based on the biomarker IP 25. Marine Ecology Progress Series, 546:17–29, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/546/
m546p017.pdf.

Buchheister:2013:PDD

[BBGL13] A. Buchheister, C. F. Bonzek, J. Gartland, and R. J. Latour. Patterns and drivers of the demersal fish community of Chesapeake Bay. Marine Ecology Progress Series, 481:161–180, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/481/m481p161.pdf.

Balke:2011:WOT

[BBH^{+}11] T. Balke, T. J. Bouma, E. M. Horstman, E. L. Webb, P. L. A. Erftemeijer, and P. M. J. Herman. Windows of opportunity: thresholds to mangrove seedling establishment on tidal flats. Marine Ecology Progress Series, 440:1–9, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Bundy:2012:CPC

[BBH^{+}12] A. Bundy, E. C. Bohaboy, D. O. Hjermann, F. J. Mueter, C. Fu, and J. S. Link. Common patterns, common drivers: comparative analysis of aggregate surplus production across ecosystems. Marine Ecology Progress Series, 459:203–218, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
theme/m459p203.pdf.

Blanchet:2013:SBT

[BBH^{+}13] M. A. Blanchet, M. Biuw, G. J. G. Hofmeyr, P. J. N. de Bruyn, C. Lydersen, and K. M. Kovacs. At-sea behaviour of three krill predators breeding at Bouvetøya — Antarctic fur seals, macaroni penguins and chinstrap penguins. Marine Ecology Progress Series, 477:285–302, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/477/m477p285.pdf.

Bell:2016:ITO

[BBI16] L. E. Bell, B. A. Bluhm, and K. Iken. Influence of terrestrial organic matter in marine food webs of the Beaufort Sea shelf and slope. Marine Ecology Progress Series, 550:1–24, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Benoit-Bird:2011:AAE

[BBKH^{+}11] K. J. Benoit-Bird, K. Kuletz, S. Heppell, N. Jones, and B. Hoover. Active acoustic examination of the diving behavior of murres foraging on patchy prey. Marine Ecology Progress Series, 443:217–235, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m443p217.pdf.

Brown:2010:MPH

[BBKP10] H. Brown, M. C. Benfield, S. F. Keenan, and S. P. Powers. Movement patterns and home ranges of a pelagic carangid fish, Caranx crysos, around a petroleum platform complex. Marine Ecology Progress Series, 403:205–218, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/403/
m403p205.pdf.

Breed:2011:DFS

[BBL11] G. A. Breed, W. D. Bowen, and M. L. Leonard. Development of foraging strategies with age in a long-lived marine predator. Marine Ecology Progress Series, 431:267–279, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/431/
m431p267.pdf.

Ballachey:2013:QLT

[BBM13] B. E. Ballachey, J. L. Bodkin, and D. H. Monson. Quantifying long-term risks to sea otters from the 1989 ‘Exxon Valdez’ oil spill: Reply to Harwell & Gentile (2013). Marine Ecology Progress Series, 488:297–301, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m488p297.pdf.

Bacheler:2014:ECH

[BBM^{+}14] N. M. Bacheler, D. J. Berrane, W. A. Mitchell, C. M. Schobernd, Z. H. Schobernd, B. Z. Teer, and J. C. Ballenger. Environmental conditions and habitat characteristics influence trap and video detection probabilities for reef fish species. Marine Ecology Progress Series, 517:1–14, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Benoit-Bird:2013:NSS

[BBMH13] K. J. Benoit-Bird, N. E. McIntosh, and S. A. Heppell. Nested scales of spatial heterogeneity in juvenile walleye pollock Theragra chalcogramma in the southeastern Bering Sea. Marine Ecology Progress Series, 484:219–238, 2013. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m484p219.pdf.

Butcher:2014:LTE

[BBMM14] P. A. Butcher, A. J. Boulton, W. G. Macbeth, and H. A. Malcolm. Long-term effects of marine park zoning on giant mud crab Scylla serrata populations in three Australian estuaries. Marine Ecology Progress Series, 508:163–176, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/508/
m508p163.pdf.

Bockelmann:2012:WOE

[BBR12] A. C. Bockelmann, K. Beining, and T. B. H. Reusch. Widespread occurrence of endophytic Labyrinthula spp. in northern European eelgrass Zostera marina beds. Marine Ecology Progress Series, 445:109–116, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/445/
m445p109.pdf.

Brodeur:2014:SOB

[BBR^{+}14] R. D. Brodeur, C. Barceló, K. L. Robinson, E. A. Daly, and J. J. Ruzicka. Spatial overlap between forage fishes and the large medusa Chrysaora fuscescens in the northern California Current region. Marine Ecology Progress Series, 510:167–181, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/510/m510p167.pdf.

Bochdansky:2010:EHD

[BBRBG10] A. B. Bochdansky, S. M. Bollens, G. C. Rollwagen-Bollens, and A. H. Gibson. Effect of the heterotrophic dinoflagellate Oxyrrhis marina and the copepod Acartia tonsa on vertical carbon flux in and around thin layers of the phytoflagellate Isochrysis galbana. Marine Ecology Progress Series, 402:179–196, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/402/m402p179.pdf.

Baez:2014:ECL

[BBRM14] J. C. Báez, S. García Barcelona, R. Real, and D. Macías. Estimating by-catch of loggerhead turtles in the Mediterranean: Comment on Álvarez de Quevedo et al. (2013). Marine Ecology Progress Series, 504:301–302, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/504/m504p301.pdf. See

[dQSC13] and reply [dQSC14].

Bjornsson:2017:ECH

[BBS^{+}17] B. Björnsson, J. M. Burgos, J. Sólmundsson, S. Á. Ragnarsson, I. G. Jónsdóttir, and U. Skúladóttir. Effects of cod and haddock abundance on the distribution and abundance of northern shrimp. Marine Ecology Progress Series, 572:209–221, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/572/m572p209.pdf.

Benoit-Bird:2019:DFR

[BBSM19] K. J. Benoit-Bird, B. L. Southall, and M. A. Moline. Dynamic foraging by Risso’s dolphins revealed in four dimensions. Marine Ecology Progress Series, 632:221–234, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m632p221.pdf.

Bell:2014:FCM

[BBTW14] J. E. Bell, M. J. Bishop, R. B. Taylor, and J. E. Williamson. Facilitation cascade maintains a kelp community. Marine Ecology Progress Series, 501:1–10, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Bogdanova:2017:MCT

[BBW^{+}17] M. I. Bogdanova, A. Butler, S. Wanless, B. Moe, T. Anker-Nilssen, M. Frederiksen, T. Boulinier, L. S. Chivers, S. Christensen-Dalsgaard, S. Descamps, M. P. Harris, M. Newell, B. Olsen, R. A. Phillips, D. Shaw, H. Steen, H. Strøm, T. L. Thórarinsson, and F. Daunt. Multi-colony tracking reveals spatio-temporal variation in carry-over effects between breeding success and winter movements in a pelagic seabird. Marine Ecology Progress Series, 578:167–181, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m578p167.pdf.

Budden:2019:ESP

[BBW^{+}19] C. Budden, I. Butler, K. Wolfe, D. Deaker, H. Sweatman, and M. Byrne. Effect of sublethal predation on reproductive output of the crown-of-thorns starfish Acanthaster sp., with an overview of arm damage. Marine Ecology Progress Series, 629:103–116, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/629/m629p103.pdf.

Buckel:2012:WMH

[BBWG12] C. A. Buckel, C. A. Blanchette, R. R. Warner, and S. D. Gaines. Where a male is hard to find: consequences of male rarity in the surfgrass Phyllospadix torreyi. Marine Ecology Progress Series, 449:121–132, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m449p121.pdf.

Babarro:2011:BSM

[BC11] J. M. F. Babarro and E. Carrington. Byssus secretion of Mytilus galloprovincialis: effect of site at macro- and microgeographical scales within Ría de Vigo (NW Spain). Marine Ecology Progress Series, 435:125–140, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/435/
m435p125.pdf.

Brown:2013:WFM

[BC13] A. L. Brown and R. C. Carpenter. Water-flow mediated oxygen dynamics within massive Porites-algal turf interactions. Marine Ecology Progress Series, 490:1–10, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Browne:2014:MAL

[BC14] M. A. Browne and M. G. Chapman. Mitigating against the loss of species by adding artificial intertidal pools to existing seawalls. Marine Ecology Progress Series, 497:119– 129, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/497/m497p119.pdf.

Bochdansky:2015:NTE

[BC15] A. B. Bochdansky and M. A. Clouse. New tracer to estimate community predation rates of phagotrophic protists. Marine Ecology Progress Series, 524:55–69, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m524p055.pdf.

Banas:2016:TCB

[BC16] N. S. Banas and R. G. Campbell. Traits controlling body size in copepods: separating general constraints from speciesspecific strategies. Marine Ecology Progress Series, 558:21– 33, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m558p021.pdf.

Burford:2011:ASG

[BCB11] M. O. Burford, M. H. Carr, and G. Bernardi. Agestructured genetic analysis reveals temporal and geographic variation within and between two cryptic rockfish species. Marine Ecology Progress Series, 442:201–215, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/442/
m442p201.pdf.

Breton:2017:FTV

[BCB^{+}17] E. Breton, U. Christaki, S. Bonato, M. Didry, and L. F. Artigas. Functional trait variation and nitrogen use efficiency in temperate coastal phytoplankton. Marine Ecology Progress Series, 563:35–49, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/563/m563p035.pdf.

Bradbury:2011:CNM

[BCC^{+}11] I. R. Bradbury, M. W. Coulson, S. E. Campana, I. G. Paterson, and P. Bentzen. Contemporary nuclear and mitochondrial genetic clines in a north temperate estuarine fish reflect Pleistocene vicariance. Marine Ecology Progress Series, 438:207–218, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/438/m438p207.pdf.

Becking:2013:SSC

[BCdV13] L. E. Becking, D. F. R. Cleary, and N. J. de Voogd. Sponge species composition, abundance, and cover in marine lakes and coastal mangroves in Berau, Indonesia. Marine Ecology Progress Series, 481:105–120, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m481p105.pdf.

Bailleul:2010:MEF

[BCG10] F. Bailleul, C. Cotté, and C. Guinet. Mesoscale eddies as foraging area of a deep-diving predator, the southern elephant seal. Marine Ecology Progress Series, 408:251–264, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/408/
m408p251.pdf.

Burford:2019:GRM

[BCGG19] B. P. Burford, N. Carey, W. F. Gilly, and J. A. Goldbogen. Grouping reduces the metabolic demand of a social squid. Marine Ecology Progress Series, 612:141–150, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/612/
m612p141.pdf.

Beca-Carretero:2019:DIA

[BCGWS19] P. Beca-Carretero, F. Guihéneuf, G. Winters, and D. B. Stengel. Depth-induced adjustment of fatty acid and pigment composition suggests high biochemical plasticity in the tropical seagrass Halophila stipulacea. Marine Ecology Progress Series, 608:105–117, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/608/m608p105.pdf.

Borstad:2011:ECB

[BCH^{+}11] G. Borstad, W. Crawford, J. M. Hipfner, R. Thomson, and K. Hyatt. Environmental control of the breeding success of rhinoceros auklets at Triangle Island, British Columbia. Marine Ecology Progress Series, 424:285–302, 2011. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m424p285.pdf.

Breed:2018:SSI

[BCH^{+}18] G. A. Breed, M. F. Cameron, J. M. Ver Hoef, P. L. Boveng, A. Whiting, and K. J. Frost. Seasonal sea ice dynamics drive movement and migration of juvenile bearded seals Erignathus barbatus. Marine Ecology Progress Series, 600:223– 237, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/600/m600p223.pdf.

Brooks:2018:FCS

[BCHG18] M. T. Brooks, V. J. Coles, R. R. Hood, and J. F. R. Gower. Factors controlling the seasonal distribution of pelagic Sargassum. Marine Ecology Progress Series, 599:1–18, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Bosch:2015:STE

[BCK15] J. A. Bosch, J. C. Cornwell, and W. M. Kemp. Shortterm effects of nereid polychaete size and density on sediment inorganic nitrogen cycling under varying oxygen conditions. Marine Ecology Progress Series, 524:155–169, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m524p155.pdf.

Bekkevold:2011:GMS

[BCM^{+}11a] D. Bekkevold, L. A. W. Clausen, S. Mariani, C. André, E. M. C. Hatfield, E. Torstensen, N. Ryman, G. R. Carvalho, and D. E. Ruzzante. Genetic mixed-stock analysis of Atlantic herring populations in a mixed feeding area. Marine Ecology Progress Series, 442:187–199, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/442/m442p187.pdf.

Bryars:2011:NEC

[BCM11b] S. Bryars, G. Collings, and D. Miller. Nutrient exposure causes epiphytic changes and coincident declines in two temperate Australian seagrasses. Marine Ecology Progress Series, 441:89–103, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/441/m441p089.pdf.

Becker:2014:SPT

[BCM^{+}14] E. L. Becker, E. E. Cordes, S. A. Macko, R. W. Lee, and C. R. Fisher. Spatial patterns of tissue stable isotope contents give insight into the nutritional sources for seep communities on the Gulf of Mexico lower slope. Marine Ecology Progress Series, 498:133–145, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2014/498/m498p133.pdf.

Busch:2019:GGC

[BCM19] M. Busch, D. Caron, and S. Moorthi. Growth and grazing control of the dinoflagellate Lingulodinium polyedrum in a natural plankton community. Marine Ecology Progress Series, 611:45–58, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/611/m611p045.pdf.

Barria:2015:UER

[BCN15] C. Barría, M. Coll, and J. Navarro. Unravelling the ecological role and trophic relationships of uncommon and threatened elasmobranchs in the western Mediterranean Sea. Marine Ecology Progress Series, 539:225–240, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/539/
m539p225.pdf.

Beca-Carretero:2018:REW

[BCOMKJ18] P. Beca-Carretero, B. Olesen, N. Marbà, and D. Krause-Jensen. Response to experimental warming in northern eelgrass populations: comparison across a range of temperature adaptations. Marine Ecology Progress Series, 589:59– 72, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/589/m589p059.pdf.

Bezio:2018:ECS

[BCPC18] N. Bezio, J. H. Costello, E. Perry, and S. P. Colin. Effects of capture surface morphology on feeding success of scyphomedusae: a comparative study. Marine Ecology Progress Series, 596:83–93, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m596p083.pdf.

Bylenga:2015:FLD

[BCR15] C. H. Bylenga, V. J. Cummings, and K. G. Ryan. Fertilisation and larval development in an Antarctic bivalve, Laternula elliptica, under reduced pH and elevated temperatures. Marine Ecology Progress Series, 536:187–201, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/536/
m536p187.pdf.

Bartolino:2012:SDD

[BCS^{+}12] V. Bartolino, L. Ciannelli, P. Spencer, T. K. Wilderbuer, and K. S. Chan. Scale-dependent detection of the effects of harvesting a marine fish population. Marine Ecology Progress Series, 444:251–261, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/444/m444p251.pdf.

Boucher:2013:EIE

[BCSB13] J. M. Boucher, C. Chen, Y. Sun, and R. C. Beardsley. Effects of interannual environmental variability on the transportretention dynamics in haddock Melanogrammus aeglefinus larvae on Georges Bank. Marine Ecology Progress Series, 487: 201–215, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m487p201.pdf.

Beatty:2018:IEM

[BCSH18] D. S. Beatty, C. S. Clements, F. J. Stewart, and M. E. Hay. Intergenerational effects of macroalgae on a reef coral: major declines in larval survival but subtle changes in microbiomes. Marine Ecology Progress Series, 589:97–114, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/589/
m589p097.pdf.

Blicher:2010:SSV

[BCSR10] M. E. Blicher, C. Clemmesen, M. K. Sejr, and S. Rysgaard. Seasonal and spatial variations in the RNA: DNA ratio and its relation to growth in sub-Arctic scallops. Marine Ecology Progress Series, 407:87–98, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/407/m407p087.pdf.

Bodey:2018:FCI

[BCV^{+}18] T. W. Bodey, I. R. Cleasby, S. C. Votier, K. C. Hamer, J. Newton, S. C. Patrick, E. D. Wakefield, and S. Bearhop. Frequency and consequences of individual dietary specialisation in a wideranging marine predator, the northern gannet. Marine Ecology Progress Series, 604:251–262, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/604/m604p251.pdf.

Bednarz:2015:SVD

[BCvH^{+}15] V. N. Bednarz, U. Cardini, N. van Hoytema, M. M. D. Al-Rshaidat, and C. Wild. Seasonal variation in dinitrogen fixation and oxygen fluxes associated with two dominant zooxanthellate soft corals from the northern Red Sea. Marine Ecology Progress Series, 519:141–152, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/519/m519p141.pdf.

Becker:2012:ITA

[BCW12] A. Becker, C. Coppinger, and A. K. Whitfield. Influence of tides on assemblages and behaviour of fishes associated with shallow seagrass edges and bare sand. Marine Ecology Progress Series, 456:187–199, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/456/m456p187.pdf.

Bayer:2019:DET

[BCW19] S. R. Bayer, P. D. Countway, and R. A. Wahle. Developing an eDNA toolkit to quantify broadcast spawning events of the sea scallop Placopecten magellanicus: moving beyond fertilization assays. Marine Ecology Progress Series, 621:127–141, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m621p127.pdf.

Burdett-Coutts:2014:SLB

[BCWSR14] V. H. M. Burdett-Coutts, R. A. Wahle, P. V. R. Snelgrove, and R. Rochette. Spatial linkages between settling youngof-year and older juvenile lobsters. Marine Ecology Progress Series, 499:143–155, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m499p143.pdf.

Beukema:2011:BSD

[BD11a] J. J. Beukema and R. Dekker. Benthic species diversity and sediment composition: Comment on Compton et al. (2008). Marine Ecology Progress Series, 440:281–283, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/440/

m440p281.pdf. See [CTvdM^{+}08].

Bosley:2011:UEL

[BD11b] K. M. Bosley and B. R. Dumbauld. Use of extractable lipofuscin to estimate age structure of ghost shrimp populations in west coast estuaries of the USA. Marine Ecology Progress Series, 428:161–176, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/428/m428p161.pdf.

Blake:2012:CBE

[BD12] R. E. Blake and J. E. Duffy. Changes in biodiversity and environmental stressors influence community structure of an experimental eelgrass Zostera marina system. Marine Ecology Progress Series, 470:41–54, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/470/m470p041.pdf.

Beukema:2013:EBP

[BD13] J. J. Beukema and R. Dekker. Evaluation of Brey’s production/biomass model on the basis of a long-term data set on a clam population. Marine Ecology Progress Series, 489: 163–175, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/489/m489p163.pdf.

Beukema:2014:VPA

[BD14] J. J. Beukema and R. Dekker. Variability in predator abundance links winter temperatures and bivalve recruitment: correlative evidence from long-term data in a tidal flat. Marine Ecology Progress Series, 513:1–15, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Beukema:2015:DDG

[BD15] J. J. Beukema and R. Dekker. Density dependence of growth and production in a Wadden Sea population of the cockle Cerastoderma edule. Marine Ecology Progress Series, 538: 157–167, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/538/m538p157.pdf.

Beukema:2019:CCT

[BD19] J. J. Beukema and R. Dekker. The carrying capacity of a tidal flat area for suspension-feeding bivalves. Marine Ecology Progress Series, 629:55–65, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/629/m629p055.pdf.

Berge:2010:ELP

[BDAH10] T. Berge, N. Daugbjerg, B. Balling Andersen, and P. J. Hansen. Effect of lowered pH on marine phytoplankton growth rates. Marine Ecology Progress Series, 416:79–91, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2010/416/m416p079.pdf.

Burthe:2012:PTT

[BDB^{+}12] S. Burthe, F. Daunt, A. Butler, D. A. Elston, M. Frederiksen, D. Johns, M. Newell, S. J. Thackeray, and S. Wanless. Phenological trends and trophic mismatch across multiple levels of a North Sea pelagic food web. Marine Ecology Progress Series, 454:119–133, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m454p119.pdf.

Bijoux:2013:TDR

[BDB^{+}13] J. P. Bijoux, L. Dagorn, G. Berke, P. D. Cowley, M. Soria, J. C. Gaertner, and J. Robinson. Temporal dynamics, residency and site fidelity of spawning aggregations of a herbivorous tropical reef fish Siganus sutor. Marine Ecology Progress Series, 475:233–247, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/475/m475p233.pdf.

Brandt:2011:RHP

[BDBN11] M. J. Brandt, A. Diederichs, K. Betke, and G. Nehls. Responses of harbour porpoises to pile driving at the Horns Rev II offshore wind farm in the Danish North Sea. Marine Ecology Progress Series, 421:205–216, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/421/m421p205.pdf.

Brunel:2010:ETP

[BDC10] T. Brunel and M. Dickey-Collas. Effects of temperature and population density on von Bertalanffy growth parameters in Atlantic herring: a macro-ecological analysis. Marine Ecology Progress Series, 405:15–28, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/405/m405p015.pdf.

Beukema:2017:PCL

[BDD17] J. J. Beukema, R. Dekker, and J. Drent. Parallel changes of Limecola (Macoma) balthica populations in the Dutch Wadden Sea. Marine Ecology Progress Series, 585:71–79, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2017/585/m585p071.pdf.

Brandt:2018:DHP

[BDD^{+}18] M. J. Brandt, A. C. Dragon, A. Diederichs, M. A. Bellmann, V. Wahl, W. Piper, J. Nabe-Nielsen, and G. Nehls. Disturbance of harbour porpoises during construction of the first seven offshore wind farms in Germany. Marine Ecology Progress Series, 596:213–232, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/596/m596p213.pdf.

Beukema:2017:LTC

[BDDvdM17] J. J. Beukema, R. Dekker, J. Drent, and J. van der Meer. Long-term changes in annual growth of bivalves in the Wadden Sea: influences of temperature, food, and abundance. Marine Ecology Progress Series, 573:143–156, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/573/
m573p143.pdf.

Bridge:2011:VMC

[BDF^{+}11] T. C. L. Bridge, T. J. Done, A. Friedman, R. J. Beaman, S. B. Williams, O. Pizarro, and J. M. Webster. Variability in mesophotic coral reef communities along the Great Barrier Reef, Australia. Marine Ecology Progress Series, 428: 63–75, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/428/m428p063.pdf.

Bouchoucha:2016:PUM

[BDG^{+}16] M. Bouchoucha, A. M. Darnaude, A. Gudefin, R. Neveu, M. Verdoit-Jarraya, P. Boissery, and P. Lenfant. Potential use of marinas as nursery grounds by rocky fishes: insights from four Diplodus species in the Mediterranean. Marine Ecology Progress Series, 547:193–209, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/547/m547p193.pdf.

Burkholz:2019:TDS

[BDGB19] C. Burkholz, C. M. Duarte, and N. Garcias-Bonet. Thermal dependence of seagrass ecosystem metabolism in the Red Sea. Marine Ecology Progress Series, 614:79–90, 2019. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m614p079.pdf.

Bortolotto:2017:WDB

[BDH^{+}17] G. A. Bortolotto, D. Danilewicz, P. S. Hammond, L. Thomas, and A. N. Zerbini. Whale distribution in a breeding area: spatial models of habitat use and abundance of western South Atlantic humpback whales. Marine Ecology Progress Series, 585:213–227, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/585/m585p213.pdf.

Bendell:2010:CGF

[BDLW10] L. I. Bendell, C. Duckham, T. L’Esperance, and J. Whiteley. Changes in geochemical foreshore attributes as a consequence of intertidal shellfish aquaculture: a case study. Marine Ecology Progress Series, 404:91–108, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/404/
m404p091.pdf.

Barcelo:2013:HUA

[BDM^{+}13] C. Barceló, A. Domingo, P. Miller, L. Ortega, B. Giffoni, G. Sales, L. McNaughton, M. Marcovaldi, S. S. Heppell, and Y. Swimmer. High-use areas, seasonal movements and dive patterns of juvenile loggerhead sea turtles in the southwestern Atlantic Ocean. Marine Ecology Progress Series, 479:235–250, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/479/m479p235.pdf.

Bernard:2014:CSS

[BDM^{+}14] G. Bernard, M. L. Delgard, O. Maire, A. Ciutat, P. Lecroart, B. Deflandre, J. C. Duchêne, and A. Grémare. Comparative study of sediment particle mixing in a Zostera noltei meadow and a bare sediment mudflat. Marine Ecology Progress Series, 514:71–86, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/514/m514p071.pdf.

Ballard:2010:FSA

[BDNA10] G. Ballard, K. M. Dugger, N. Nur, and D. G. Ainley. Foraging strategies of Adélie penguins: adjusting body condition to cope with environmental variability. Marine Ecology Progress Series, 405:287–302, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m405p287.pdf.

Beukema:2010:LTV

[BDP10] J. J. Beukema, R. Dekker, and C. J. M. Philippart. Long-term variability in bivalve recruitment, mortality, and growth and their contribution to fluctuations in food stocks of shellfisheating birds. Marine Ecology Progress Series, 414:117–130, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps_oa/m414p117.pdf.

Bender:2014:WAP

[BDPD14] D. Bender, G. Diaz-Pulido, and S. Dove. Warming and acidification promote cyanobacterial dominance in turf algal assemblages. Marine Ecology Progress Series, 517:271– 284, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/517/m517p271.pdf.

Beukhof:2019:STV

[BDPL19] E. Beukhof, T. S. Dencker, L. Pecuchet, and M. Lindegren. Spatio-temporal variation in marine fish traits reveals community-wide responses to environmental change. Marine Ecology Progress Series, 610:205–222, 2019. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m610p205.pdf.

Boucher:2019:VPB

[BDR19] N. Boucher, A. E. Derocher, and E. S. Richardson. Variability in polar bear Ursus maritimus stable isotopes in relation to environmental change in the Canadian Beaufort Sea. Marine Ecology Progress Series, 630:215–225, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/630/
m630p215.pdf.

Baudry:2018:TMP

[BDS18a] J. Baudry, D. Dumont, and I. R. Schloss. Turbulent mixing and phytoplankton life history: a Lagrangian versus Eulerian model comparison. Marine Ecology Progress Series, 600: 55–70, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/600/m600p055.pdf.

Bourque:2018:FHN

[BDS^{+}18b] J. Bourque, R. Dietz, C. Sonne, J. St Leger, S. Iverson, A. Rosing-Asvid, M. Hansen, and M. A. McKinney. Feeding habits of a new Arctic predator: insight from full-depth blubber fatty acid signatures of Greenland, Faroe Islands, Denmark, and managed-care killer whales Orcinus orca. Marine Ecology Progress Series, 603:1–12, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Bartilotti:2014:DLR

[BdSC^{+}14] C. Bartilotti, A. dos Santos, M. Castro, Á. Peliz, and A. M. P. Santos. Decapod larval retention within distributional bands in a coastal upwelling ecosystem. Marine Ecology Progress Series, 507:233–247, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2014/507/m507p233.pdf.

Bradbury:2011:RNT

[BDT^{+}11] I. R. Bradbury, C. DiBacco, S. R. Thorrold, P. V. R. Snelgrove, and S. E. Campana. Resolving natal tags using otolith geochemistry in an estuarine fish, rainbow smelt Osmerus mordax. Marine Ecology Progress Series, 433:195– 204, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/433/m433p195.pdf.

Bergman:2016:NIH

[BDT^{+}16] J. L. Bergman, B. N. Dang, M. M. Tabatabaee, M. M. Mc-Gowan, C. R. Fong, S. J. Bittick, and P. Fong. Nutrients induce and herbivores maintain thallus toughness, a structural anti-herbivory defense in Turbinaria ornata. Marine Ecology Progress Series, 559:35–43, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/559/m559p035.pdf.

Benjamins:2016:RTU

[BDvGW16] S. Benjamins, A. Dale, N. van Geel, and B. Wilson. Riding the tide: use of a moving tidal-stream habitat by harbour porpoises. Marine Ecology Progress Series, 549:275–288, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m549p275.pdf.

Brosset:2019:FPG

[BDVP19] P. Brosset, J. M. Durant, E. Van Beveren, and S. Plourde. Fish population growth in the Gulf of St Lawrence: effects of climate, fishing and predator abundance. Marine Ecology Progress Series, 624:167–181, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/624/m624p167.pdf.

Beauchamp:2011:FAM

[Bea11a] G. Beauchamp. Fit of aggregation models to the distribution of group sizes in Northwest Atlantic seabirds. Marine Ecology Progress Series, 425:261–268, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/425/m425p261.pdf.

Beauchamp:2011:FRB

[Bea11b] G. Beauchamp. Functional relationship between group size and population density in Northwest Atlantic seabirds. Marine Ecology Progress Series, 435:225–233, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/435/
m435p225.pdf.

Beaugrand:2012:UBC

[Bea12] G. Beaugrand. Unanticipated biological changes and global warming. Marine Ecology Progress Series, 445:293–301, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2011/445/m445p293.pdf.

Blomqvist:2015:LOI

[BEEH15] S. Blomqvist, N. Ekeroth, R. Elmgren, and P. O. J. Hall. Long overdue improvement of box corer sampling. Marine Ecology Progress Series, 538:13–21, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m538p013.pdf.

Boissin:2015:CPG

[BEFC15] E. Boissin, E. Egea, J. P. Féral, and A. Chenuil. Contrasting population genetic structures in Amphipholis squamata, a complex of brooding, self-reproducing sister species sharing life history traits. Marine Ecology Progress Series, 539:165–177, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/539/m539p165.pdf.

Booth:2013:HPD

[BEG^{+}13] C. G. Booth, C. Embling, J. Gordon, S. V. Calderan, and P. S. Hammond. Habitat preferences and distribution of the harbour porpoise Phocoena phocoena west of Scotland. Marine Ecology Progress Series, 478:273–285, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/478/
m478p273.pdf.

Beninger:2011:DTS

[BEMD11] P. G. Beninger, R. W. Elner, M. Morançais, and P. Decottignies. Downward trophic shift during breeding migration in the shorebird Calidris mauri (western sandpiper). Marine Ecology Progress Series, 428:259–269, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/428/m428p259.pdf.

Benkwitt:2016:ILI

[Ben16] C. E. Benkwitt. Invasive lionfish increase activity and foraging movements at greater local densities. Marine Ecology Progress Series, 558:255–266, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m558p255.pdf.

Baden:2012:SSF

[BEP^{+}12] S. Baden, A. Emanuelsson, L. Pihl, C. J. Svensson, and P. Åberg. Shift in seagrass food web structure over decades is linked to overfishing. Marine Ecology Progress Series, 451: 61–73, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m451p061.pdf.

Bergsma:2012:CME

[Ber12a] G. S. Bergsma. Coral mutualists enhance fish abundance and diversity through a morphology-mediated facilitation cascade. Marine Ecology Progress Series, 451:151–161, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/451/
m451p151.pdf.

Berke:2012:BVE

[Ber12b] S. K. Berke. Biogeographic variability in ecosystem engineering: patterns in the abundance and behavior of the tubebuilding polychaete Diopatra cuprea. Marine Ecology Progress Series, 447:1–13, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Barrett:2017:OCE

[BER17] R. T. Barrett, K. E. Erikstad, and T. K. Reiertsen. Ocean climate and egg investment in the black-legged kittiwake Rissa tridactyla. Marine Ecology Progress Series, 579:129– 137, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/579/m579p129.pdf.

Bett:2013:CBS

[Bet13] Brian J. Bett. Characteristic benthic size spectra: potential sampling artefacts. Marine Ecology Progress Series, 487: 1–6, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/feature/m487p001.pdf. See comment [War14] and

reply [Bet14].

Bett:2014:MMR

[Bet14] B. J. Bett. Macroecology and meiobenthos: Reply to Warwick (2014). Marine Ecology Progress Series, 505:299–302, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_

oa/m505p299.pdf. See [Bet13, War14].

Bell:2014:SMF

[BFC14] R. J. Bell, M. J. Fogarty, and J. S. Collie. Stability in marine fish communities. Marine Ecology Progress Series, 504:221–239, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/504/m504p221.pdf.

Becker:2016:IDH

[BFD^{+}16] S. L. Becker, J. T. Finn, A. J. Danylchuk, C. G. Pollock, Z. Hillis-Starr, I. Lundgren, and A. Jordaan. Influence of detection history and analytic tools on quantifying spatial ecology of a predatory fish in a marine protected area. Marine Ecology Progress Series, 562:147–161, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/562/
m562p147.pdf.

Barrow:2018:EDG

[BFDM18] J. Barrow, J. Ford, R. Day, and J. Morrongiello. Environmental drivers of growth and predicted effects of climate change on a commercially important fish, Platycephalus laevigatus. Marine Ecology Progress Series, 598:201–212, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m598p201.pdf.

Briones-Fourzan:2019:AFA

[BFEMLÁJ19] P. Briones-Fourzán, A. F. Espinosa-Magaña, E. Lozano-Álvarez, and A. Jeffs. Analysis of fatty acids to examine larval and settlement biology of the Caribbean spiny lobster Panulirus argus. Marine Ecology Progress Series, 630: 137–148, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/630/m630p137.pdf.

Becker:2010:CCC

[BFF^{+}10] E. A. Becker, K. A. Forney, M. C. Ferguson, D. G. Foley, R. C. Smith, J. Barlow, and J. V. Redfern. Comparing California Current cetacean-habitat models developed using in situ and remotely sensed sea surface temperature data. Marine Ecology Progress Series, 413:163–183, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/413/m413p163.pdf.

Buscaino:2011:ABE

[BFG^{+}11a] G. Buscaino, F. Filiciotto, M. Gristina, A. Bellante, G. Buffa, V. Di Stefano, V. Maccarrone, G. Tranchida, C. Buscaino, and S. Mazzola. Acoustic behaviour of the European spiny lobster Palinurus elephas. Marine Ecology Progress Series, 441:177–184, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m441p177.pdf.

Buscaino:2011:DSE

[BFG^{+}11b] G. Buscaino, F. Filiciotto, M. Gristina, G. Buffa, A. Bellante, V. Maccarrone, B. Patti, and S. Mazzola. Defensive strategies of European spiny lobster Palinurus elephas during predator attack. Marine Ecology Progress Series, 423:143–154, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m423p143.pdf.

Battini:2019:SAI

[BFG^{+}19] N. Battini, N. Farías, C. B. Giachetti, E. Schwindt, and A. Bortolus. Staying ahead of invaders: using species distribution modeling to predict alien species’ potential niche shifts. Marine Ecology Progress Series, 612:127–140, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/612/
m612p127.pdf.

Bolte:2013:PGI

[BFH^{+}13a] S. Bolte, V. Fuentes, H. Haslob, B. Huwer, D. Thibault-Botha, D. Angel, B. Galil, J. Javidpour, A. G. Moss, and T. B. H. Reusch. Population genetics of the invasive ctenophore Mnemiopsis leidyi in Europe reveal source-sink dynamics and secondary dispersal to the Mediterranean Sea. Marine Ecology Progress Series, 485:25–36, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/485/m485p025.pdf.

Burkholder:2013:SPS

[BFH13b] D. A. Burkholder, J. W. Fourqurean, and M. R. Heithaus. Spatial pattern in seagrass stoichiometry indicates both n-limited and p-limited regions of an iconic p-limited subtropical bay. Marine Ecology Progress Series, 472:101–115, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/472/
m472p101.pdf.

Burmester:2017:TSA

[BFKR17] E. M. Burmester, J. R. Finnerty, L. Kaufman, and R. D. Rotjan. Temperature and symbiosis affect lesion recovery in experimentally wounded, facultatively symbiotic temperate corals. Marine Ecology Progress Series, 570:87–99, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/570/
m570p087.pdf.

Bricelj:2011:ETA

[BFL^{+}11] V. M. Bricelj, S. E. Ford, C. Lambert, A. Barbou, and C. Paillard. Effects of toxic Alexandrium tamarense on behavior, hemocyte responses and development of brown ring disease in Manila clams. Marine Ecology Progress Series, 430: 35–48, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/430/m430p035.pdf.

Brown:2019:HPC

[BFL^{+}19] A. L. Brown, T. K. Frazer, G. Li, J. R. Hilsenroth, J. S. Shima, and C. W. Osenberg. Hidden predators on coral reefs: muricid consumption of vermetids. Marine Ecology Progress Series, 615:121–131, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/615/m615p121.pdf.

Bradshaw:2013:MAB

[BFM^{+}13] C. J. A. Bradshaw, I. C. Field, C. R. McMahon, G. J. Johnson, M. G. Meekan, and R. C. Buckworth. More analytical bite in estimating targets for shark harvest. Marine Ecology Progress Series, 488:221–232, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/488/m488p221.pdf.

Bryan:2019:HRS

[BFN^{+}19] D. R. Bryan, M. W. Feeley, R. S. Nemeth, C. Pollock, and J. S. Ault. Home range and spawning migration patterns of queen triggerfish Balistes vetula in St. Croix, US Virgin Islands. Marine Ecology Progress Series, 616:123–139, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m616p123.pdf.

Beck:2016:WSE

[BFNB16] H. J. Beck, D. A. Feary, Y. Nakamura, and D. J. Booth. Wave-sheltered embayments are recruitment hotspots for tropical fishes on temperate reefs. Marine Ecology Progress Series, 546:197–212, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/546/m546p197.pdf.

Briones-Fourzan:2012:ETC

[BFPONS^{+}12] P. Briones-Fourzán, M. Pérez-Ortiz, F. Negrete-Soto, C. Barradas-Ortiz, and E. Lozano-Álvarez. Ecological traits of Caribbean sea anemones and symbiotic crustaceans. Marine Ecology Progress Series, 470:55–68, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m470p055.pdf.

Baker:2013:HRS

[BFRM13] R. Baker, B. Fry, L. P. Rozas, and T. J. Minello. Hydrodynamic regulation of salt marsh contributions to aquatic food webs. Marine Ecology Progress Series, 490:37–52, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/490/m490p037.pdf.

Bach:2010:LAC

[BFS10] L. Bach, A. Fischer, and J. Strand. Local anthropogenic contamination affects the fecundity and reproductive success of an Arctic amphipod. Marine Ecology Progress Series, 419:121–128, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/419/m419p121.pdf.

Bernard:2012:BIP

[BG12] A. T. F. Bernard and A. Götz. Bait increases the precision in count data from remote underwater video for most subtidal reef fish in the warm-temperate Agulhas bioregion. Marine Ecology Progress Series, 471:235–252, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/471/
m471p235.pdf.

Baillie:2019:FAR

[BG19] C. J. Baillie and J. H. Grabowski. Factors affecting recruitment, growth and survival of the eastern oyster Crassostrea virginica across an intertidal elevation gradient in southern New England. Marine Ecology Progress Series, 609:119–132, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/609/m609p119.pdf.

Bacheler:2017:CRA

[BGB^{+}17] N. M. Bacheler, N. R. Geraldi, M. L. Burton, R. C. Muñoz, and G. T. Kellison. Comparing relative abundance, lengths, and habitat of temperate reef fishes using simultaneous underwater visual census, video, and trap sampling. Marine Ecology Progress Series, 574:141–155, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m574p141.pdf.

Byrne:2013:EOW

[BGBD^{+}13] M. Byrne, M. Gonzalez-Bernat, S. Doo, S. Foo, N. Soars, and M. Lamare. Effects of ocean warming and acidification on embryos and non-calcifying larvae of the invasive sea star Patiriella regularis. Marine Ecology Progress Series, 473:235–246, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/473/m473p235.pdf.

Brownscombe:2014:FBA

[BGDC14] J. W. Brownscombe, L. F. G. Gutowsky, A. J. Danylchuk, and S. J. Cooke. Foraging behaviour and activity of a marine benthivorous fish estimated using tri-axial accelerometer biologgers. Marine Ecology Progress Series, 505:241– 251, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/505/m505p241.pdf.

Botto:2019:TNC

[BGIA19] F. Botto, E. Gaitán, O. O. Iribarne, and E. M. Acha. Trophic niche changes during settlement in the Argentine hake Merluccius hubbsi reveal the importance of pelagic food post metamorphosis. Marine Ecology Progress Series, 619:125– 136, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/619/m619p125.pdf.

Barclay:2019:VEO

[BGJ^{+}19] K. M. Barclay, B. Gaylord, B. M. Jellison, P. Shukla, E. Sanford, and L. R. Leighton. Variation in the effects of ocean acidification on shell growth and strength in two intertidal gastropods. Marine Ecology Progress Series, 626:109–121, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2019/626/m626p109.pdf.

Bronstein:2017:DIL

[BGK17] O. Bronstein, E. Georgopoulou, and A. Kroh. On the distribution of the invasive long-spined echinoid Diadema setosum and its expansion in the Mediterranean Sea. Marine Ecology Progress Series, 583:163–178, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/583/m583p163.pdf.

Barreiro:2011:ACW

[BGL^{+}11] F. Barreiro, M. Gómez, M. Lastra, J. López, and R. de la Huz. Annual cycle of wrack supply to sandy beaches: effect of the physical environment. Marine Ecology Progress Series, 433:65–74, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/433/m433p065.pdf.

Buxton:2017:VNN

[BGM^{+}17] R. T. Buxton, R. Galvan, M. F. McKenna, C. L. White, and V. Seher. Visitor noise at a nesting colony alters the behavior of a coastal seabird. Marine Ecology Progress Series, 570:233–246, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m570p233.pdf.

Bologna:2018:SEA

[BGM^{+}18] P. Bologna, J. J. Gaynor, R. Meredith, D. Restaino, and C. Barry. Stochastic event alters gelatinous zooplankton community structure: impacts of Hurricane Sandy in a Mid-Atlantic estuary. Marine Ecology Progress Series, 591:217– 227, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/591/m591p217.pdf.

Brook:2018:ESA

[BGO^{+}18] T. W. Brook, B. L. Gilby, A. D. Olds, R. M. Connolly, C. J. Henderson, and T. A. Schlacher. The effects of shoreline armouring on estuarine fish are contingent upon the broader urbanisation context. Marine Ecology Progress Series, 605:195–206, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/605/m605p195.pdf.

Barbosa:2016:LBD

[BGP^{+}16] A. C. C. Barbosa, C. C. Gomes, G. C. Pereira, M. Bueno, and A. A. V. Flores. Local biological drivers, not remote forcing, predict settlement rate to a subtropical barnacle population. Marine Ecology Progress Series, 543:201–208, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/543/
m543p201.pdf.

Brewster:2016:ENC

[BGS^{+}16] J. D. Brewster, C. Giraldo, H. Swanson, W. Walkusz, T. N. Loewen, J. D. Reist, G. A. Stern, and L. L. Loseto. Ecological niche of coastal Beaufort Sea fishes defined by stable isotopes and fatty acids. Marine Ecology Progress Series, 559:159–173, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/559/m559p159.pdf.

Bejarano:2013:ERE

[BGSM13] S. Bejarano, Y. Golbuu, T. Sapolu, and P. J. Mumby. Ecological risk and the exploitation of herbivorous reef fish across Micronesia. Marine Ecology Progress Series, 482:197–215, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/482/m482p197.pdf.

Brannock:2010:HRH

[BH10] P. M. Brannock and T. J. Hilbish. Hybridization results in high levels of sterility and restricted introgression between invasive and endemic marine blue mussels. Marine Ecology Progress Series, 406:161–171, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/406/m406p161.pdf.

Boudreau:2012:SVE

[BH12a] M. R. Boudreau and D. J. Hamilton. Seasonal variation in effects of multiple predators on an intertidal mussel bed: implications for interpretation of manipulative experiments. Marine Ecology Progress Series, 465:137–153, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/465/
m465p137.pdf.

Boyd:2012:TSB

[BH12b] Philip W. Boyd and David A. Hutchins. Theme section: Biological responses in an anthropogenically modified ocean. Marine Ecology Progress Series, 470:123–290, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m470_ThemeSection.pdf.

Bessey:2015:ENA

[BH15] C. Bessey and M. R. Heithaus. Ecological niche of an abundant teleost Pelates octolineatus in a subtropical seagrass ecosystem. Marine Ecology Progress Series, 541:195– 204, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/541/m541p195.pdf.

Barnes:2016:VEF

[BH16a] R. S. K. Barnes and S. Hamylton. On the very edge: faunal and functional responses to the interface between benthic seagrass and unvegetated sand assemblages. Marine Ecology Progress Series, 553:33–48, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/553/m553p033.pdf.

Berge:2016:RCP

[BH16b] T. Berge and P. J. Hansen. Role of the chloroplasts in the predatory dinoflagellate Karlodinium armiger. Marine Ecology Progress Series, 549:41–54, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/549/
m549p041.pdf.

Bannister:2011:IBD

[BHA^{+}11] R. J. Bannister, M. O. Hoogenboom, K. R. N. Anthony, C. N. Battershill, S. Whalan, N. S. Webster, and R. de Nys. Incongruence between the distribution of a common coral reef sponge and photosynthesis. Marine Ecology Progress Series, 423:95–100, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/423/m423p095.pdf.

Beltran-Heredia:2017:PSS

[BHAC17] E. Beltrán-Heredia, D. L. Aksnes, and F. J. Cao. Phytoplankton size scaling with nutrient concentration. Marine Ecology Progress Series, 571:59–64, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/571/m571p059.pdf.

Ben-Horin:2018:GVA

[BHASP18] T. Ben-Horin, S. K. Allen, Jr., J. M. Small, and D. A. Proestou. Genetic variation in anti-parasite behavior in oysters. Marine Ecology Progress Series, 594:107–117, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2018/594/m594p107.pdf.

Brandt:2013:SST

[BHD^{+}13] M. J. Brandt, C. Höschle, A. Diederichs, K. Betke, R. Matuschek, and G. Nehls. Seal scarers as a tool to deter harbour porpoises from offshore construction sites. Marine Ecology Progress Series, 475:291–302, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/475/m475p291.pdf.

Borges:2015:ESH

[BHDC15] C. D. G. Borges, S. J. Hawkins, C. P. Doncaster, and T. P. Crowe. Effects of simulated human exploitation of a key grazer, Patella vulgata, on rocky shore assemblages. Marine Ecology Progress Series, 533:163–176, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/533/m533p163.pdf.

Beaugrand:2014:DPS

[BHE14] Gregory Beaugrand, Xavier Harlay, and Martin Edwards. Detecting plankton shifts in the North Sea: a new abrupt ecosystem shift between 1996 and 2003. Marine Ecology Progress Series, 502:85–104, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/502/m502p085.pdf. See comment

[AF14].

Barnett:2017:GBF

[BHE^{+}17] L. A. K. Barnett, S. M. Hennessey, T. E. Essington, A. O. Shelton, B. E. Feist, T. A. Branch, and M. M. McClure. Getting to the bottom of fishery interactions with living habitats: spatiotemporal trends in disturbance of corals and sponges on the US west coast. Marine Ecology Progress Series, 574:29–47, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/574/m574p029.pdf.

Baggett:2010:NEG

[BHF^{+}10] L. P. Baggett, K. L. Heck, Jr., T. A. Frankovich, A. R. Armitage, and J. W. Fourqurean. Nutrient enrichment, grazer identity, and their effects on epiphytic algal assemblages: field experiments in subtropical turtlegrass Thalassia testudinum meadows. Marine Ecology Progress Series, 406:33–45, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m406p033.pdf.

Bessey:2016:ITM

[BHF^{+}16] C. Bessey, M. R. Heithaus, J. W. Fourqurean, K. R. Gastrich, and D. A. Burkholder. Importance of teleost macrograzers to seagrass composition in a subtropical ecosystem with abundant populations of megagrazers and predators. Marine Ecology Progress Series, 553:81–92, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/553/
m553p081.pdf.

Bruce:2019:BSM

[BHL^{+}19] B. D. Bruce, D. Harasti, K. Lee, C. Gallen, and A. R. Bradford. Broad-scale movements of juvenile white sharks Carcharodon carcharias in eastern Australia from acoustic and satellite telemetry. Marine Ecology Progress Series, 619: 1–15, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/feature/m619p001.pdf.

Burdorf:2016:LDE

[BHMCM16] L. D. W. Burdorf, S. Hidalgo-Martinez, P. L. M. Cook, and F. J. R. Meysman. Long-distance electron transport by cable bacteria in mangrove sediments. Marine Ecology Progress Series, 545:1–8, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Burrows:2012:MMR

[BHN^{+}12] J. A. Burrows, J. T. Harvey, K. M. Newton, D. A. Croll, and S. R. Benson. Marine mammal response to interannual variability in Monterey Bay, California. Marine Ecology Progress Series, 461:257–271, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/461/m461p257.pdf.

Bentley:2014:LGP

[BHN^{+}14] B. P. Bentley, E. S. Harvey, S. J. Newman, D. J. Welch, A. K. Smith, and W. J. Kennington. Local genetic patchiness but no regional differences between Indo–West Pacific populations of the dogtooth tuna Gymnosarda unicolor. Marine Ecology Progress Series, 506:267–277, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/506/m506p267.pdf.

Bailie:2011:GPS

[BHP11] D. A. Bailie, R. Hynes, and P. A. Prodöhl. Genetic parentage in the squat lobsters Munida rugosa and M. sarsi (Crustacea, Anomura, Galatheidae). Marine Ecology Progress Series, 421:173–182, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/421/m421p173.pdf.

Baker:2012:RIC

[BHP12] J. D. Baker, E. A. Howell, and J. J. Polovina. Relative influence of climate variability and direct anthropogenic impact on a sub-tropical Pacific top predator, the Hawaiian monk seal. Marine Ecology Progress Series, 469:175–189, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/469/
m469p175.pdf.

Burghart:2010:PFR

[BHT10] S. E. Burghart, T. L. Hopkins, and J. J. Torres. Partitioning of food resources in bathypelagic micronekton in the eastern Gulf of Mexico. Marine Ecology Progress Series, 399:131–140, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/399/m399p131.pdf.

Berthelsen:2015:BTT

[BHT15] A. K. Berthelsen, J. E. Hewitt, and R. B. Taylor. Biological traits and taxonomic composition of invertebrate assemblages associated with coralline turf along an environmental gradient. Marine Ecology Progress Series, 530:15–27, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/530/
m530p015.pdf.

Burkholder:2011:DTI

[BHTF11] D. A. Burkholder, M. R. Heithaus, J. A. Thomson, and J. W. Fourqurean. Diversity in trophic interactions of green sea turtles Chelonia mydas on a relatively pristine coastal foraging ground. Marine Ecology Progress Series, 439:277– 293, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/439/m439p277.pdf.

Baeza:2019:PGB

[BHUCMO19] J. A. Baeza, D. Holstein, R. Umaña-Castro, and L. M. Mejía-Ortíz. Population genetics and biophysical modeling inform metapopulation connectivity of the Caribbean king crab Maguimithrax spinosissimus. Marine Ecology Progress Series, 610:83–97, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/610/m610p083.pdf.

Bachiller:2015:TDO

[BI15] E. Bachiller and X. Irigoien. Trophodynamics and diet overlap of small pelagic fish species in the Bay of Biscay. Marine Ecology Progress Series, 534:179–198, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/534/
m534p179.pdf.

Beniash:2010:ELC

[BIL^{+}10] E. Beniash, A. Ivanina, N. S. Lieb, I. Kurochkin, and I. M. Sokolova. Elevated level of carbon dioxide affects metabolism and shell formation in oysters Crassostrea virginica. Marine Ecology Progress Series, 419:95–108, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/419/
m419p095.pdf.

Bergmark:2014:LPC

[BJ14] P. Bergmark and D. Jørgensen. Lophelia pertusa conservation in the North Sea using obsolete offshore structures as artificial reefs. Marine Ecology Progress Series, 516:275– 280, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/516/m516p275.pdf.

Beldade:2014:GSA

[BJCB^{+}14] R. Beldade, A. M. Jackson, R. Cudney-Bueno, P. T. Raimondi, and G. Bernardi. Genetic structure among spawning aggregations of the gulf coney Hyporthodus acanthistius. Marine Ecology Progress Series, 499:193–201, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/499/
m499p193.pdf.

Barry:2017:EIG

[BJF17] S. C. Barry, C. A. Jacoby, and T. K. Frazer. Environmental influences on growth and morphology of Thalassia testudinum. Marine Ecology Progress Series, 570:57–70, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/570/
m570p057.pdf.

Basterretxea:2011:SST

[BJG^{+}11] G. Basterretxea, A. Jordi, E. Garcés, S. Anglès, and A. Reñé. Seiches stimulate transient biogeochemical changes in a microtidal coastal ecosystem. Marine Ecology Progress Series, 423:15–28, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/423/m423p015.pdf.

Berkstrom:2013:ECN

[BJH13] C. Berkström, T. L. Jörgensen, and M. Hellström. Ecological connectivity and niche differentiation between two closely related fish species in the mangrove-seagrass-coral reef continuum. Marine Ecology Progress Series, 477:201– 215, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/477/m477p201.pdf.

Brown:2017:HCM

[BJL^{+}17] C. J. Brown, S. D. Jupiter, H. Y. Lin, S. Albert, C. Klein, J. M. Maina, V. J. D. Tulloch, A. S. Wenger, and P. J. Mumby. Habitat change mediates the response of coral reef fish populations to terrestrial run-off. Marine Ecology Progress Series, 576:55–68, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/576/m576p055.pdf.

Bangi:2019:RAT

[BJM19] H. G. P. Bangi and M. A. Juinio-Meñez. Resource allocation trade-offs in the sea urchin Tripneustes gratilla under relative storminess and wave exposure. Marine Ecology Progress Series, 608:165–182, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/608/m608p165.pdf.

Berkstrom:2012:EVI

[BJMS12] C. Berkström, G. P. Jones, M. I. McCormick, and M. Srinivasan. Ecological versatility and its importance for the distribution and abundance of coral reef wrasses. Marine Ecology Progress Series, 461:151–163, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/461/m461p151.pdf.

Burrows:2010:SVS

[BJRH10] M. T. Burrows, S. R. Jenkins, L. Robb, and R. Harvey. Spatial variation in size and density of adult and postsettlement Semibalanus balanoides: effects of oceanographic and local conditions. Marine Ecology Progress Series, 398: 207–219, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/398/m398p207.pdf.

Bond:2011:RSP

[BJS^{+}11] A. L. Bond, I. L. Jones, W. J. Sydeman, H. L. Major, S. Minobe, J. C. Williams, and G. V. Byrd. Reproductive success of planktivorous seabirds in the North Pacific is related to ocean climate on decadal scales. Marine Ecology Progress Series, 424:205–218, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/424/m424p205.pdf.

Burrows:2016:PDD

[BJS^{+}16] J. A. Burrows, D. W. Johnston, J. M. Straley, E. M. Chenoweth, C. Ware, C. Curtice, S. L. DeRuiter, and A. S. Friedlaender. Prey density and depth affect the fine-scale foraging behavior of humpback whales Megaptera novaeangliae in Sitka Sound, Alaska, USA. Marine Ecology Progress Series, 561:245–260, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/561/m561p245.pdf.

Bunt:2014:MHU

[BK14] C. M. Bunt and M. J. Kingsford. Movement, habitat utilization and behaviour of coral trout Plectropomus leopardus during and after the reproductive period on the southern Great Barrier Reef. Marine Ecology Progress Series, 496: 33–45, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/496/m496p033.pdf.

Buren:2012:LPD

[BKAM12] A. D. Buren, M. Koen-Alonso, and W. A. Montevecchi. Linking predator diet and prey availability: common murres and capelin in the Northwest Atlantic. Marine Ecology Progress Series, 445:25–35, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/445/m445p025.pdf.

Buren:2014:RHS

[BKAS14] A. D. Buren, M. Koen-Alonso, and G. B. Stenson. The role of harp seals, fisheries and food availability in driving the dynamics of northern cod. Marine Ecology Progress Series, 511:265–284, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m511p265.pdf.

Bonaldo:2011:RIP

[BKB11] R. M. Bonaldo, J. P. Krajewski, and D. R. Bellwood. Relative impact of parrotfish grazing scars on massive Porites corals at Lizard Island, Great Barrier Reef. Marine Ecology Progress Series, 423:223–233, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/423/m423p223.pdf.

Bourke:2014:IBU

[BKC14] M. F. Bourke, A. J. Kessler, and P. L. M. Cook. Influence of buried Ulva lactuca on denitrification in permeable sediments. Marine Ecology Progress Series, 498:85–94, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/498/
m498p085.pdf.

Brault:2019:TPF

[BKC^{+}19] E. K. Brault, P. L. Koch, D. P. Costa, M. D. McCarthy, L. A. Hückstädt, K. T. Goetz, K. W. McMahon, M. E. Goebel, O. Karlsson, J. Teilmann, T. Harkonen, and K. C. Harding. Trophic position and foraging ecology of Ross, Weddell, and crabeater seals revealed by compound-specific isotope analysis. Marine Ecology Progress Series, 611:1–18, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Bourque:2015:IPD

[BKF15] A. S. Bourque, W. J. Kenworthy, and J. W. Fourqurean. Impacts of physical disturbance on ecosystem structure in subtropical seagrass meadows. Marine Ecology Progress Series, 540:27–41, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/540/m540p027.pdf.

Bowen:2013:AAC

[BKHW13] J. L. Bowen, P. J. Kearns, M. Holcomb, and B. B. L. Ward. Acidification alters the composition of ammonia-oxidizing microbial assemblages in marine mesocosms. Marine Ecology Progress Series, 492:1–8, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Brault:2018:CNZ

[BKM^{+}18] E. K. Brault, P. L. Koch, K. W. McMahon, K. H. Broach, A. P. Rosenfield, W. Sauthoff, V. J. Loeb, K. R. Arrigo, and W. O. Smith, Jr. Carbon and nitrogen zooplankton isoscapes in West Antarctica reflect oceanographic transitions. Marine Ecology Progress Series, 593:29–45, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/593/m593p029.pdf.

Berriman:2015:SAB

[BKR^{+}15] J. S. Berriman, M. C. Kay, D. C. Reed, A. Rassweiler, D. A. Goldstein, and W. G. Wright. Shifts in attack behavior of an important kelp forest predator within marine reserves. Marine Ecology Progress Series, 522:193–201, 2015. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m522p193.pdf.

Broad:2010:ESR

[BKTD10] A. Broad, N. Knott, X. Turon, and A. R. Davis. Effects of a shark repulsion device on rocky reef fishes: no shocking outcomes. Marine Ecology Progress Series, 408:295– 298, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/408/m408p295.pdf.

Byron:2010:SFE

[BL10] C. J. Byron and J. S. Link. Stability in the feeding ecology of four demersal fish predators in the US northeast shelf large marine ecosystem. Marine Ecology Progress Series, 406:239–250, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m406p239.pdf.

Brischoux:2013:TCP

[BL13] F. Brischoux and H. B. Lillywhite. Trophic consequences of pelagic life-style in yellow-bellied sea snakes. Marine Ecology Progress Series, 478:231–238, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/478/m478p231.pdf.

Balestri:2014:SRB

[BL14] E. Balestri and C. Lardicci. Seagrass response to burial and breakage of expanding horizontal rhizomes: implications for clone spread. Marine Ecology Progress Series, 504:133– 145, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/504/m504p133.pdf.

Bromilow:2017:MGO

[BL17] A. M. Bromilow and R. N. Lipcius. Mechanisms governing ontogenetic habitat shifts: role of trade-offs, predation, and cannibalism for the blue crab. Marine Ecology Progress Series, 584:145–159, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m584p145.pdf.

Brosset:2016:LSP

[BLC^{+}16] Pablo Brosset, Baptiste Le Bourg, David Costalago, Daniela Bănaru, Elisabeth Van Beveren, Jean-Hervé Bourdeix, Jean-Marc Fromentin, Frédéric Ménard, and Claire Saraux. Linking small pelagic dietary shifts with ecosystem changes in the Gulf of Lions. Marine Ecology Progress Series, 554:157– 171, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/554/m554p157.pdf.

Barco:2016:LTK

[BLD^{+}16a] S. Barco, M. Law, B. Drummond, H. Koopman, C. Trapani, S. Reinheimer, S. Rose, W. M. Swingle, and A. Williard. Loggerhead turtles killed by vessel and fishery interaction in Virginia, USA, are healthy prior to death. Marine Ecology Progress Series, 555:221–234, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/555/m555p221.pdf.

Bashevkin:2016:PEL

[BLD^{+}16b] S. M. Bashevkin, D. Lee, P. Driver, E. Carrington, and S. B. George. Prior exposure to low salinity affects the vertical distribution of Pisaster ochraceus (Echinodermata: Asteroidea) larvae in haloclines. Marine Ecology Progress Series, 542:123–140, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/542/m542p123.pdf.

Baumgartner:2013:ABN

[BLE^{+}13] M. F. Baumgartner, N. S. J. Lysiak, H. C. Esch, A. N. Zerbini, C. L. Berchok, and P. J. Clapham. Associations between North Pacific right whales and their zooplanktonic prey in the southeastern Bering Sea. Marine Ecology Progress Series, 490:267–284, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m490p267.pdf.

Biermann:2015:OMP

[BLER15] A. Biermann, A. M. Lewandowska, A. Engel, and U. Riebesell. Organic matter partitioning and stoichiometry in response to rising water temperature and copepod grazing. Marine Ecology Progress Series, 522:49–65, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m522p049.pdf.

Blackett:2015:PEM

[BLHL15] M. Blackett, C. H. Lucas, R. A. Harmer, and P. Licandro. Population ecology of Muggiaea atlantica (Cnidaria, Siphonophora) in the Western English Channel. Marine Ecology Progress Series, 535:129–144, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m535p129.pdf.

Barber-Lluch:2019:RVB

[BLHRP^{+}19] E. Barber-Lluch, M. Hernández-Ruiz, A. Prieto, E. Fernández, and E. Teira. Role of vitamin B_{12} in the microbial plankton response to nutrient enrichment. Marine Ecology Progress Series, 626:29–42, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/626/m626p029.pdf.

Beaugrand:2011:NMA

[BLIM11] G. Beaugrand, S. Lenoir, F. Ibañez, and C. Manté. A new model to assess the probability of occurrence of a species, based on presence-only data. Marine Ecology Progress Series, 424:175–190, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/424/m424p175.pdf.

Brasso:2014:ONE

[BLJP14] R. L. Brasso, J. Lang, C. D. Jones, and M. J. Polito. Ontogenetic niche expansion influences mercury exposure in the Antarctic silverfish Pleuragramma antarcticum. Marine Ecology Progress Series, 504:253–263, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/504/
m504p253.pdf.

Buckley:2010:STM

[BLM10] L. J. Buckley, R. G. Lough, and D. Mountain. Seasonal trends in mortality and growth of cod and haddock larvae result in an optimal window for survival. Marine Ecology Progress Series, 405:57–69, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/405/m405p057.pdf.

Bundy:2012:TSC

[BLM^{+}12] A. Bundy, J. Link, T. Miller, E. Moksness, and K. Stergiou. Theme section: Comparative analysis of marine fisheries production. Marine Ecology Progress Series, 459:157–302, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m459_ThemeSection.pdf.

Brannock:2018:MCC

[BLM^{+}18] P. M. Brannock, D. R. Learman, A. R. Mahon, S. R. Santos, and K. M. Halanych. Meiobenthic community composition and biodiversity along a 5500 km transect of Western Antarctica: a metabarcoding analysis. Marine Ecology Progress Series, 603:47–60, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/603/m603p047.pdf.

Barrett-Lennard:2011:PGW

[BLMD^{+}11] L. G. Barrett-Lennard, C. O. Matkin, J. W. Durban, E. L. Saulitis, and D. Ellifrit. Predation on gray whales and prolonged feeding on submerged carcasses by transient killer whales at Unimak Island, Alaska. Marine Ecology Progress Series, 421:229–241, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/421/m421p229.pdf.

Bradley:2016:CSI

[BLPP16] C. J. Bradley, K. Longenecker, R. L. Pyle, and B. N. Popp. Compound-specific isotopic analysis of amino acids reveals dietary changes in mesophotic coral-reef fish. Marine Ecology Progress Series, 558:65–79, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/558/m558p065.pdf.

Baumgartner:2011:DVM

[BLS^{+}11] M. F. Baumgartner, N. S. J. Lysiak, C. Schuman, J. Urban-Rich, and F. W. Wenzel. Diel vertical migration behavior of Calanus finmarchicus and its influence on right and sei whale occurrence. Marine Ecology Progress Series, 423: 167–184, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/423/m423p167.pdf.

Boehme:2014:CGS

[BLT14] L. Boehme, M. Lonergan, and C. D. Todd. Comparison of gridded sea surface temperature datasets for marine ecosystem studies. Marine Ecology Progress Series, 516:7–22, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m516p007.pdf.

Barton:2019:ITD

[BLV^{+}19] M. B. Barton, S. Y. Litvin, J. J. Vollenweider, R. A. Heintz, B. L. Norcross, and K. M. Boswell. Implications of trophic discrimination factor selection for stable isotope food web models of low trophic levels in the Arctic nearshore. Marine Ecology Progress Series, 613:211–216, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/613/m613p211.pdf.

Baker:2010:GMJ

[BM10a] R. Baker and T. J. Minello. Growth and mortality of juvenile white shrimp Litopenaeus setiferus in a marsh pond. Marine Ecology Progress Series, 413:95–104, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m413p095.pdf.

Brunnschweiler:2010:BSM

[BM10b] Jürg M. Brunnschweiler and Jonathan McKenzie. Baiting sharks for marine tourism: Comment on Clua et al. (2010). Marine Ecology Progress Series, 420:283–284, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/420/

m420p283.pdf. See [CBL^{+}10b] and reply [CBL^{+}10a].

Burgess:2011:FEP

[BM11] S. C. Burgess and D. J. Marshall. Field estimates of planktonic larval duration in a marine invertebrate. Marine Ecology Progress Series, 440:151–161, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/440/m440p151.pdf.

Booth:2013:NRP

[BM13] J. M. Booth and C. D. McQuaid. Northern rockhopper penguins prioritise future reproduction over chick provisioning. Marine Ecology Progress Series, 486:289–304, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/486/
m486p289.pdf.

Byron:2016:PRS

[BM16] C. Byron and A. Morgan. Potential role of spiny dogfish in gray and harbor seal diets in the Gulf of Maine. Marine Ecology Progress Series, 550:249–270, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m550p249.pdf.

Barbour:2010:MUI

[BMA^{+}10] A. B. Barbour, M. L. Montgomery, A. A. Adamson, E. Díaz-Ferguson, and B. R. Silliman. Mangrove use by the invasive lionfish Pterois volitans. Marine Ecology Progress Series, 401:291–294, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/401/m401p291.pdf.

Buren:2019:CCL

[BMA^{+}19] A. D. Buren, H. M. Murphy, A. T. Adamack, G. K. Davoren, M. Koen-Alonso, W. A. Montevecchi, F. K. Mowbray, P. Pepin, P. M. Regular, D. Robert, G. A. Rose, G. B. Stenson, and D. Varkey. The collapse and continued low productivity of a keystone forage fish species. Marine Ecology Progress Series, 616:155–170, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m616p155.pdf.

Baird:2018:DBS

[BMÁN^{+}18] A. H. Baird, J. S. Madin, M. Álvarez-Noriega, L. Fontoura, J. T. Kerry, C. Y. Kuo, K. Precoda, D. Torres-Pulliza, R. M. Woods, K. J. A. Zawada, and T. P. Hughes. A decline in bleaching suggests that depth can provide a refuge from global warming in most coral taxa. Marine Ecology Progress Series, 603:257–264, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/603/m603p257.pdf.

Baguley:2015:CRD

[BMC^{+}15] J. G. Baguley, P. A. Montagna, C. Cooksey, J. L. Hyland, H. W. Bang, C. Morrison, A. Kamikawa, P. Bennetts, G. Saiyo, E. Parsons, M. Herdener, and M. Ricci. Community response of deep-sea soft-sediment metazoan meiofauna to the Deepwater Horizon blowout and oil spill. Marine Ecology Progress Series, 528:127–140, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m528p127.pdf.

Bombar:2011:DDM

[BMD^{+}11] D. Bombar, P. H. Moisander, J. W. Dippner, R. A. Foster, M. Voss, B. Karfeld, and J. P. Zehr. Distribution of diazotrophic microorganisms and nifH gene expression in the Mekong River plume during intermonsoon. Marine Ecology Progress Series, 424:39–52, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/424/m424p039.pdf.

Bellchambers:2011:MHA

[BMEL11] L. M. Bellchambers, J. J. Meeuwig, S. N. Evans, and P. Legendre. Modelling habitat associations of the common spider conch in the Cocos (Keeling) Islands. Marine Ecology Progress Series, 432:83–90, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/432/m432p083.pdf.

Bracken:2012:MSH

[BMF^{+}12] M. E. S. Bracken, B. A. Menge, M. M. Foley, C. J. B. Sorte, J. Lubchenco, and D. R. Schiel. Mussel selectivity for high-quality food drives carbon inputs into open-coast intertidal ecosystems. Marine Ecology Progress Series, 459: 53–62, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/459/m459p053.pdf.

Brosset:2015:IEV

[BMF^{+}15] P. Brosset, F. Ménard, J. M. Fromentin, S. Bonhommeau, C. Ulses, J. H. Bourdeix, J. L. Bigot, E. Van Beveren, D. Roos, and C. Saraux. Influence of environmental variability and age on the body condition of small pelagic fish in the Gulf of Lions. Marine Ecology Progress Series, 529:219–231, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/529/
m529p219.pdf.

Berejikian:2016:PPI

[BMJ16] B. A. Berejikian, M. E. Moore, and S. J. Jeffries. Predator– prey interactions between harbor seals and migrating steelhead trout smolts revealed by acoustic telemetry. Marine Ecology Progress Series, 543:21–35, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m543p021.pdf.

Belicka:2012:FAS

[BMJH12] L. L. Belicka, P. Matich, R. Jaffé, and M. R. Heithaus. Fatty acids and stable isotopes as indicators of early-life feeding and potential maternal resource dependency in the bull shark Carcharhinus leucas. Marine Ecology Progress Series, 455:245–256, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/455/m455p245.pdf.

Barley:2017:SDA

[BMM17] S. C. Barley, M. G. Meekan, and J. J. Meeuwig. Species diversity, abundance, biomass, size and trophic structure of fish on coral reefs in relation to shark abundance. Marine Ecology Progress Series, 565:163–179, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/565/m565p163.pdf.

Bellgrove:2010:RHF

[BMMS10] A. Bellgrove, P. F. McKenzie, J. L. McKenzie, and B. J. Sfiligoj. Restoration of the habitat-forming fucoid alga Hormosira banksii at effluent-affected sites: competitive exclusion by coralline turfs. Marine Ecology Progress Series, 419: 47–56, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/419/m419p047.pdf.

Brena:2015:SRP

[BMPC15] P. F. Brena, J. Mourier, S. Planes, and E. Clua. Shark and ray provisioning: functional insights into behavioral, ecological and physiological responses across multiple scales. Marine Ecology Progress Series, 538:273–283, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/538/
m538p273.pdf.

Brame:2014:SFC

[BMPH14] A. B. Brame, C. C. McIvor, E. B. Peebles, and D. J. Hollander. Site fidelity and condition metrics suggest sequential habitat use by juvenile common snook. Marine Ecology Progress Series, 509:255–269, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/509/m509p255.pdf.

Bishop:2017:EMS

[BMR17] N. Bishop, D. L. Martin, and C. Ross. Effects of multistress exposure on the infection dynamics of a Labyrinthula sp.–turtle grass pathosystem. Marine Ecology Progress Series, 581:119–133, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/581/m581p119.pdf.

Baker:2013:LTM

[BMSB13] A. C. Baker, T. R. McClanahan, C. J. Starger, and R. K. Boonstra. Long-term monitoring of algal symbiont communities in corals reveals stability is taxon dependent and driven by site-specific thermal regime. Marine Ecology Progress Series, 479:85–97, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m479p085.pdf.

Braun:2014:SFL

[BMSS14] C. Braun, N. K. Michiels, U. E. Siebeck, and D. Sprenger. Signalling function of long wavelength colours during agonistic male-male interactions in the wrasse Coris julis. Marine Ecology Progress Series, 504:277–286, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m504p277.pdf.

Broell:2017:TTE

[BMT17] F. Broell, J. S. P. McCain, and C. T. Taggart. Thermal time explains size-at-age variation in molluscs. Marine Ecology Progress Series, 573:157–165, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/573/m573p157.pdf.

Bedford:2015:PFU

[BMTC^{+}15] M. Bedford, J. Melbourne-Thomas, S. Corney, T. Jarvis, N. Kelly, and A. Constable. Prey-field use by a Southern Ocean top predator: enhanced understanding using integrated datasets. Marine Ecology Progress Series, 526:169– 181, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/526/m526p169.pdf.

Bueno:2010:RFT

[BMTF10] M. Bueno, G. A. O. Moser, B. R. C. Tocci, and A. A. V. Flores. Retention-favorable timing of propagule release in barnacles and periwinkles. Marine Ecology Progress Series, 414:155–165, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/414/m414p155.pdf.

Bentaleb:2011:FEM

[BMV^{+}11] I. Bentaleb, C. Martin, M. Vrac, B. Mate, P. Mayzaud, D. Siret, R. de Stephanis, and C. Guinet. Foraging ecology of Mediterranean fin whales in a changing environment elucidated by satellite tracking and baleen plate stable isotopes. Marine Ecology Progress Series, 438:285–302, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m438p285.pdf.

Barbera:2017:MBI

[BMV^{+}17] C. Barberá, S. Mallol, A. Vergés, M. Cabanellas-Reboredo, D. Díaz, and R. Goñi. Maerl beds inside and outside a 25year-old no-take area. Marine Ecology Progress Series, 572: 77–90, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/572/m572p077.pdf.

Batsleer:2018:EHC

[BMV^{+}18] J. Batsleer, P. Marchal, S. Vaz, V. Vermard, A. D. Rijnsdorp, and J. J. Poos. Exploring habitat credits to manage the benthic impact in a mixed fishery. Marine Ecology Progress Series, 586:167–179, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/586/m586p167.pdf.

Berezina:2019:IBI

[BMV19] N. A. Berezina, A. A. Maximov, and O. M. Vladimirova. Influence of benthic invertebrates on phosphorus flux at the sediment–water interface in the easternmost Baltic Sea. Marine Ecology Progress Series, 608:33–43, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/608/
m608p033.pdf.

Brito-Millan:2019:CRA

[BMVAS19] M. Brito-Millán, M. J. A. Vermeij, E. A. Alcantar, and S. A. Sandin. Coral reef assessments based on cover alone mask active dynamics of coral communities. Marine Ecology Progress Series, 630:55–68, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/630/m630p055.pdf.

Bertocci:2010:RRS

[BMVBC10] I. Bertocci, E. Maggi, S. Vaselli, and L. Benedetti-Cecchi. Resistance of rocky shore assemblages of algae and invertebrates to changes in intensity and temporal variability of aerial exposure. Marine Ecology Progress Series, 400:75–86, 2010. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m400p075.pdf.

Bernard:2014:BNF

[BMW^{+}14] R. J. Bernard, B. Mortazavi, L. Wang, A. C. Ortmann, H. MacIntyre, and W. C. Burnett. Benthic nutrient fluxes and limited denitrification in a sub-tropical groundwaterinfluenced coastal lagoon. Marine Ecology Progress Series, 504:13–26, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/504/m504p013.pdf.

Biggs:2016:STM

[BN16] C. R. Biggs and R. S. Nemeth. Spatial and temporal movement patterns of two snapper species at a multi-species spawning aggregation. Marine Ecology Progress Series, 558:129– 142, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/558/m558p129.pdf.

Bernasconi:2011:AOD

[BNAK11] M. Bernasconi, L. Nøttestad, B. E. Axelsen, and J. O. Krakstad. Acoustic observations of dusky dolphins Lagenorhynchus obscurus hunting Cape horse mackerel Trachurus capensis off Namibia. Marine Ecology Progress Series, 429:209–218, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m429p209.pdf.

Barrett:2012:LTD

[BNAN12] R. T. Barrett, E. B. Nilsen, and T. Anker-Nilssen. Longterm decline in egg size of Atlantic puffins Fratercula arctica is related to changes in forage fish stocks and climate conditions. Marine Ecology Progress Series, 457:1–10, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Boissonnot:2019:LFA

[BNE^{+}19] L. Boissonnot, B. Niehoff, B. Ehrenfels, J. E. Søreide, W. Hagen, and M. Graeve. Lipid and fatty acid turnover of the pteropods Limacina helicina, L. retroversa and Clione limacina from Svalbard waters. Marine Ecology Progress Series, 609:133–149, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/609/m609p133.pdf.

Boje:2014:SMV

[BNS^{+}14] J. Boje, S. Neuenfeldt, C. R. Sparrevohn, O. Eigaard, and J. W. Behrens. Seasonal migration, vertical activity, and winter temperature experience of Greenland halibut Reinhardtius hippoglossoides in West Greenland waters. Marine Ecology Progress Series, 508:211–222, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/508/m508p211.pdf.

Burnaford:2014:CMA

[BNW14] J. L. Burnaford, K. J. Nielsen, and S. L. Williams. Celestial mechanics affects emersion time and cover patterns of an ecosystem engineer, the intertidal kelp Saccharina sessilis. Marine Ecology Progress Series, 509:127–136, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/509/
m509p127.pdf.

Blackburn:2013:SBE

[BO13] N. J. Blackburn and R. J. Orth. Seed burial in eelgrass Zostera marina: the role of infauna. Marine Ecology Progress Series, 474:135–145, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/474/m474p135.pdf.

Bednarsek:2015:CPD

[BO15] N. Bednaršek and M. D. Ohman. Changes in pteropod distributions and shell dissolution across a frontal system in the California Current system. Marine Ecology Progress Series, 523:93–103, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/523/m523p093.pdf.

Blanco:2019:ILT

[BOÁNF19] M. Blanco, A. Ospina-Álvarez, S. A. Navarrete, and M. Fernández. Influence of larval traits on dispersal and connectivity patterns of two exploited marine invertebrates in central Chile. Marine Ecology Progress Series, 612:43–64, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2019/612/m612p043.pdf.

Baylis:2016:SSH

[BOC^{+}16] A. M. M. Baylis, R. A. Orben, D. P. Costa, J. P. Y. Arnould, and I. J. Staniland. Sexual segregation in habitat use is smaller than expected in a highly dimorphic marine predator, the southern sea lion. Marine Ecology Progress Series, 554:201–211, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/554/m554p201.pdf.

Bui:2010:ICS

[BOCB10] A. O. V. Bui, P. Ouellet, M. Castonguay, and J. C. Brêthes. Ichthyoplankton community structure in the northwest Gulf of St. Lawrence (Canada): past and present. Marine Ecology Progress Series, 412:189–205, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/412/m412p189.pdf.

Blok:2018:LHE

[BOKJ18] S. E. Blok, B. Olesen, and D. Krause-Jensen. Life history events of eelgrass Zostera marina l. populations across gradients of latitude and temperature. Marine Ecology Progress Series, 590:79–93, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m590p079.pdf.

Boltovskoy:2017:SVF

[Bol17] D. Boltovskoy. Seasonality in the vertical flux and species composition of Radiolaria polycystina (Protista): patterns and drivers. Marine Ecology Progress Series, 578:51–72, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2017/578/m578p051.pdf.

Brownlee:2016:MCS

[BOS16] E. F. Brownlee, R. J. Olson, and H. M. Sosik. Microzooplankton community structure investigated with imaging flow cytometry and automated live-cell staining. Marine Ecology Progress Series, 550:65–81, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m550p065.pdf.

Burek:2018:WER

[BOS18] K. E. Burek, J. M. O’Brien, and R. E. Scheibling. Wasted effort: recruitment and persistence of kelp on algal turf. Marine Ecology Progress Series, 600:3–19, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Burfeind:2013:WTB

[BOU13] D. D. Burfeind, K. R. O’Brien, and J. W. Udy. Water temperature and benthic light levels drive horizontal expansion of Caulerpa taxifolia in native and invasive locations. Marine Ecology Progress Series, 472:61–72, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/472/
m472p061.pdf.

Benner:2010:UON

[BP10] I. Benner and U. Passow. Utilization of organic nutrients by coccolithophores. Marine Ecology Progress Series, 404: 21–29, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/404/m404p021.pdf.

Bell:2012:SGR

[BP12a] I. Bell and D. A. Pike. Somatic growth rates of hawksbill turtles Eretmochelys imbricata in a northern Great Barrier Reef foraging area. Marine Ecology Progress Series, 446:275–283, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/446/m446p275.pdf.

Bownes:2012:CSC

[BP12b] S. J. Bownes and R. Perissinotto. Community structure and composition of meiofauna over a sea-induced mouth-breaching event in St Lucia Estuary, South Africa. Marine Ecology Progress Series, 463:105–126, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/463/m463p105.pdf.

Brewin:2011:RIP

[BPB11] P. E. Brewin, P. K. Probert, and M. F. Barker. Relative influence of processes structuring fjord deep-water macrofaunal communities across multiple spatial scales. Marine Ecology Progress Series, 425:175–191, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/425/m425p175.pdf.

Blower:2012:PGA

[BPB^{+}12] Dean C. Blower, John M. Pandolfi, Barry D. Bruce, Maria del C.Gomez-Cabrera, and Jennifer R. Ovenden. Population genetics of Australian white sharks reveals fine-scale spatial structure, transoceanic dispersal events and low effective population sizes. Marine Ecology Progress Series, 455: 229–244, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/455/m455p229.pdf.

Bernik:2018:SED

[BPB18] B. M. Bernik, J. H. Pardue, and M. J. Blum. Soil erodibility differs according to heritable trait variation and nutrientinduced plasticity in the salt marsh engineer Spartina alterniflora. Marine Ecology Progress Series, 601:1–14, 2018. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Bradley:2017:NPB

[BPC17] D. Bradley, Y. P. Papastamatiou, and J. E. Caselle. No persistent behavioural effects of SCUBA diving on reef sharks. Marine Ecology Progress Series, 567:173–184, 2017. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m567p173.pdf.

Brodeur:2015:CMH

[BPF15] M. C. Brodeur, M. F. Piehler, and F. J. Fodrie. Consumers mitigate heat stress and nutrient enrichment effects on eelgrass Zostera marina communities at its southern range limit. Marine Ecology Progress Series, 525:53–64, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/525/
m525p053.pdf.

Briceno:2018:CMR

[BPF^{+}18] F. A. Briceño, E. T. Polymeropoulos, Q. P. Fitzgibbon, J. M. Dambacher, and G. T. Pecl. Changes in metabolic rate of spiny lobster under predation risk. Marine Ecology Progress Series, 598:71–84, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/598/m598p071.pdf.

Braeckman:2010:RMF

[BPG^{+}10] U. Braeckman, P. Provoost, B. Gribsholt, D. Van Gansbeke, J. J. Middelburg, K. Soetaert, M. Vincx, and J. Vanaverbeke. Role of macrofauna functional traits and density in biogeochemical fluxes and bioturbation. Marine Ecology Progress Series, 399:173–186, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/399/m399p173.pdf.

Butler:2011:BCD

[BPG^{+}11] M. J. Butler IV, C. B. Paris, J. S. Goldstein, H. Matsuda, and R. K. Cowen. Behavior constrains the dispersal of longlived spiny lobster larvae. Marine Ecology Progress Series, 422:223–237, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m422p223.pdf.

Bogdanowicz:2013:GDM

[BPG^{+}13] W. Bogdanowicz, M. Pilot, M. Gajewska, E. Suchecka, and M. Golachowski. Genetic diversity in a moulting colony of southern elephant seals in comparison with breeding colonies. Marine Ecology Progress Series, 478:287–300, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/478/
m478p287.pdf.

Bierwagen:2019:TND

[BPH^{+}19] S. L. Bierwagen, H. Pethybridge, M. R. Heupel, A. Chin, and C. A. Simpfendorfer. Trophic niches determined from fatty acid profiles of sympatric coral reef mesopredators. Marine Ecology Progress Series, 632:159–174, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/632/
m632p159.pdf.

Brunner:2016:LSP

[BPHW16] E. L. Brunner, F. G. Prahl, B. Hales, and G. G. Waldbusser. A longitudinal study of Pacific oyster (Crassostrea gigas) larval development: isotope shifts during early shell formation reveal sub-lethal energetic stress. Marine Ecology Progress Series, 555:109–123, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/555/m555p109.pdf.

Bjorkland:2015:SIM

[BPJ^{+}15] R. H. Bjorkland, S. F. Pearson, S. J. Jeffries, M. M. Lance, A. Acevedo-Gutiérrez, and E. J. Ward. Stable isotope mixing models elucidate sex and size effects on the diet of a generalist marine predator. Marine Ecology Progress Series, 526:213–225, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/526/m526p213.pdf.

Browne:2014:PPC

[BPLT14] N. K. Browne, E. Precht, K. S. Last, and P. A. Todd. Photophysiological costs associated with acute sediment stress events in three near-shore turbid water corals. Marine Ecology Progress Series, 502:129–143, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/502/m502p129.pdf.

Batsleer:2013:MFM

[BPM^{+}13] J. Batsleer, J. J. Poos, P. Marchal, Y. Vermard, and A. D. Rijnsdorp. Mixed fisheries management: protecting the weakest link. Marine Ecology Progress Series, 479:177–190, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/479/m479p177.pdf.

Bernreuther:2018:TDM

[BPM^{+}18a] M. Bernreuther, J. Peters, C. Möllmann, J. Renz, J. Dutz, J. P. Herrmann, and A. Temming. Trophic decoupling of mesozooplankton production and the pelagic planktivores sprat Sprattus sprattus and herring Clupea harengus in the Central Baltic Sea. Marine Ecology Progress Series, 592:181– 196, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/592/m592p181.pdf.

Burdett:2018:CLS

[BPM^{+}18b] H. L. Burdett, G. Perna, L. McKay, G. Broomhead, and N. A. Kamenos. Community-level sensitivity of a calcifying ecosystem to acute in situ CO_{2} enrichment. Marine Ecology Progress Series, 587:73–80, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/587/m587p073.pdf.

Berkenbusch:2011:CBS

[BPN11] K. Berkenbusch, P. K. Probert, and S. D. Nodder. Comparative biomass of sediment benthos across a depth transect, Chatham Rise, Southwest Pacific Ocean. Marine Ecology Progress Series, 425:79–90, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/425/m425p079.pdf.

Beguer-Pon:2016:MMA

[BPOS^{+}16] M. Béguer-Pon, K. Ohashi, J. Sheng, M. Castonguay, and J. J. Dodson. Modeling the migration of the American eel in the Gulf of St. Lawrence. Marine Ecology Progress Series, 549:183–198, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/549/m549p183.pdf.

Barash:2018:SAS

[BPP^{+}18] A. Barash, R. Pickholtz, E. Pickholtz, L. Blaustein, and G. Rilov. Seasonal aggregations of sharks near coastal power plants in Israel: an emerging phenomenon. Marine Ecology Progress Series, 590:145–154, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/590/m590p145.pdf.

Beguer-Pon:2017:BVV

[BPSCD17] M. Béguer-Pon, S. Shan, M. Castonguay, and J. J. Dodson. Behavioural variability in the vertical and horizontal oceanic migrations of silver American eels. Marine Ecology Progress Series, 585:123–142, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/585/m585p123.pdf.

Bleich:2011:DMS

[BPSG11] S. Bleich, M. Powilleit, T. Seifert, and G. Graf. β-diversity as a measure of species turnover along the salinity gradient in the Baltic Sea, and its consistency with the Venice system. Marine Ecology Progress Series, 436:101–118, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/436/
m436p101.pdf.

Bostrom:2011:SEC

[BPSK11] C. Boström, S. J. Pittman, C. Simenstad, and R. T. Kneib. Seascape ecology of coastal biogenic habitats: advances, gaps, and challenges. Marine Ecology Progress Series, 427:191– 217, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m427p191.pdf.

Bishop:2015:BMG

[BPT15] A. Bishop, P. Pomeroy, and S. D. Twiss. Breeding male grey seals exhibit similar activity budgets across varying exposures to human activity. Marine Ecology Progress Series, 527:247–259, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/527/m527p247.pdf.

Booth:2018:IPI

[BPW^{+}18] H. S. Booth, T. J. Pusack, J. W. White, C. D. Stallings, and D. L. Kimbro. Intraspecific predator inhibition, not a prey size refuge, enables oyster population persistence during predator outbreaks. Marine Ecology Progress Series, 602:155–167, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/602/m602p155.pdf.

Beirao:2014:SPS

[BPWF14] J. Beirão, C. F. Purchase, B. F. Wringe, and I. A. Fleming. Sperm plasticity to seawater temperatures in Atlantic cod Gadus morhua is affected more by population origin than individual environmental exposure. Marine Ecology Progress Series, 495:263–274, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/495/m495p263.pdf.

Bilgmann:2014:MMU

[BPZ^{+}14] K. Bilgmann, G. J. Parra, N. Zanardo, L. B. Beheregaray, and L. M. Möller. Multiple management units of short-beaked common dolphins subject to fisheries bycatch off southern and southeastern Australia. Marine Ecology Progress Series, 500:265–279, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/500/m500p265.pdf.

Buchan:2016:FIO

[BQ16] S. J. Buchan and R. A. Quiñones. First insights into the oceanographic characteristics of a blue whale feeding ground in northern Patagonia, Chile. Marine Ecology Progress Series, 554:183–199, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/554/m554p183.pdf.

Bollens:2012:PED

[BQRB12] S. M. Bollens, J. A. Quenette, and G. Rollwagen-Bollens. Predator-enhanced diel vertical migration in a planktonic dinoflagellate. Marine Ecology Progress Series, 447:49–54, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2011/447/m447p049.pdf.

Borg:2010:SVC

[BRA^{+}10] J. A. Borg, A. A. Rowden, M. J. Attrill, P. J. Schembri, and M. B. Jones. Spatial variation in the composition of motile macroinvertebrate assemblages associated with two bed types of the seagrass Posidonia oceanica. Marine Ecology Progress Series, 406:91–104, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/406/m406p091.pdf.

Bi:2011:ECS

[BRB11a] H. Bi, K. A. Rose, and M. C. Benfield. Estimating copepod stage-specific mortality rates in open ocean waters: a case study from the northern Gulf of Mexico, USA. Marine Ecology Progress Series, 427:145–159, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/427/m427p145.pdf.

Button:2011:VRP

[BRB11b] C. A. Button and L. Rogers-Bennett. Vital rates of pink abalone Haliotis corrugata estimated from mark-recapture data to inform recovery. Marine Ecology Progress Series, 431:151–161, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/431/m431p151.pdf.

Brakel:2017:MVC

[BRB17] J. Brakel, T. B. H. Reusch, and A. C. Bockelmann. Moderate virulence caused by the protist Labyrinthula zosterae in ecosystem foundation species Zostera marina under nutrient limitation. Marine Ecology Progress Series, 571:97– 108, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/571/m571p097.pdf.

Bennice:2019:FSH

[BRBH19] C. O. Bennice, A. P. Rayburn, W. R. Brooks, and R. T. Hanlon. Fine-scale habitat partitioning facilitates sympatry between two octopus species in a shallow Florida lagoon. Marine Ecology Progress Series, 609:151–161, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/609/
m609p151.pdf.

Boschen:2015:MAS

[BRC^{+}15] R. E. Boschen, A. A. Rowden, M. R. Clark, S. J. Barton, A. Pallentin, and J. P. A. Gardner. Megabenthic assemblage structure on three New Zealand seamounts: implications for seafloor massive sulfide mining. Marine Ecology Progress Series, 523:1–14, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Brett:2014:RPG

[Bre14] M. T. Brett. Resource polygon geometry predicts Bayesian stable isotope mixing model bias. Marine Ecology Progress Series, 514:1–12, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Budenbender:2011:CAC

[BRF11] J. Büdenbender, U. Riebesell, and A. Form. Calcification of the Arctic coralline red algae Lithothamnion glaciale in response to elevated CO_{2}. Marine Ecology Progress Series, 441:79–87, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m441p079.pdf.

Bekkby:2014:LSW

[BRG^{+}14] T. Bekkby, E. Rinde, H. Gundersen, K. M. Norderhaug, J. K. Gitmark, and H. Christie. Length, strength and water flow: relative importance of wave and current exposure on morphology in kelp Laminaria hyperborea. Marine Ecology Progress Series, 506:61–70, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m506p061.pdf.

Brannock:2013:UHM

[BRH13] P. M. Brannock, M. A. Roberts, and T. J. Hilbish. Ubiquitous heteroplasmy in Mytilus spp. resulting from disruption in doubly uniparental inheritance regulation. Marine Ecology Progress Series, 480:131–143, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/480/m480p131.pdf.

Briggs:2012:SIM

[Bri12] John C. Briggs. As I see it: Marine species invasions in estuaries and harbors. Marine Ecology Progress Series, 449: 297–302, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/449/m449p297.pdf.

Burnell:2013:EOI

[BRIC13] O. W. Burnell, B. D. Russell, A. D. Irving, and S. D. Connell. Eutrophication offsets increased sea urchin grazing on seagrass caused by ocean warming and acidification. Marine Ecology Progress Series, 485:37–46, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/485/m485p037.pdf.

Barbraud:2012:ECC

[BRJ^{+}12] C. Barbraud, V. Rolland, S. Jenouvrier, M. Nevoux, K. Delord, and H. Weimerskirch. Effects of climate change and fisheries bycatch on Southern Ocean seabirds: a review. Marine Ecology Progress Series, 454:285–307, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m454p285.pdf.

Banks:2012:ISS

[BRK^{+}12] J. L. Banks, D. J. Ross, M. J. Keough, C. K. Macleod, and B. D. Eyre. Influence of small-scale patchiness on resilience of nutrient cycling to extended hypoxia in estuarine sediments. Marine Ecology Progress Series, 453:49–62, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m453p049.pdf.

Braccini:2017:AMR

[BRLM17] M. Braccini, K. Rensing, T. Langlois, and R. McAuley. Acoustic monitoring reveals the broad-scale movements of commercially important sharks. Marine Ecology Progress Series, 577:121–129, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/577/m577p121.pdf.

Baustian:2011:SMH

[BRMT11] M. M. Baustian, N. N. Rabalais, W. L. Morrison, and R. E. Turner. Seasonal microphytobenthos on the hypoxic northern Gulf of Mexico continental shelf. Marine Ecology Progress Series, 436:51–66, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/436/m436p051.pdf.

Blicher:2010:SGV

[BRS10] M. E. Blicher, S. Rysgaard, and M. K. Sejr. Seasonal growth variation in Chlamys islandica (Bivalvia) from sub-Arctic Greenland is linked to food availability and temperature. Marine Ecology Progress Series, 407:71–86, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/407/
m407p071.pdf.

Botha:2017:SSF

[BRT^{+}17] J. A. Botha, G. M. Rishworth, A. Thiebault, D. B. Green, and P. A. Pistorius. Sex-specific foraging over space and time in Cape gannets during chick rearing. Marine Ecology Progress Series, 579:157–167, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/579/m579p157.pdf.

Bruckner:2010:QDC

[Bru10] A. W. Bruckner. Quantifying the decline in Corallium rubrum populations: Reply to santangelo & bramanti (2010). Marine Ecology Progress Series, 418:299–303, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/418/
m418p299.pdf.

Blicher:2011:AOC

[BS11] M. E. Blicher and M. K. Sejr. Abundance, oxygen consumption and carbon demand of brittle stars in Young Sound and the NE Greenland shelf. Marine Ecology Progress Series, 422:139–144, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/422/m422p139.pdf.

Best:2012:TCS

[BS12] R. J. Best and J. J. Stachowicz. Trophic cascades in seagrass meadows depend on mesograzer variation in feeding rates, predation susceptibility, and abundance. Marine Ecology Progress Series, 456:29–42, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/456/m456p029.pdf.

Baird:2014:STH

[BS14a] H. P. Baird and J. S. Stark. Spatial and temporal heterogeneity in the distribution of an Antarctic amphipod and relationship with the sediment. Marine Ecology Progress Series, 502:169–183, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/502/m502p169.pdf.

Block:2014:SVS

[BS14b] H. E. Block and M. A. Steele. Spatial variation in selective mortality on larval traits in the coral reef fish Chromis viridis. Marine Ecology Progress Series, 509:303–308, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/509/
m509p303.pdf.

Bailey:2015:DFF

[BS15] M. Bailey and U. R. Sumaila. Destructive fishing and fisheries enforcement in eastern Indonesia. Marine Ecology Progress Series, 530:195–211, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m530p195.pdf.

Bellquist:2016:TSD

[BS16] L. Bellquist and B. X. Semmens. Temporal and spatial dynamics of ‘trophy’-sized demersal fishes off the California (USA) coast, 1966 to 2013. Marine Ecology Progress Series, 547:1–18, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Bringloe:2018:MDS

[BS18a] T. T. Bringloe and G. W. Saunders. Mitochondrial DNA sequence data reveal the origins of postglacial marine macroalgal flora in the Northwest Atlantic. Marine Ecology Progress Series, 589:45–58, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/589/m589p045.pdf. See correction

[BS18b].

Bringloe:2018:CMD

[BS18b] Trevor T. Bringloe and Gary W. Saunders. Correction: Mitochondrial DNA sequence data reveal the origins of postglacial marine macroalgal flora in the Northwest Atlantic, MEPS 589:45–58 (2017). Marine Ecology Progress Series, 605:241, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/abstracts/
meps/v589/c_p45-58/. See [BS18a].

Bethoney:2019:IEH

[BS19a] N. D. Bethoney and K. D. E. Stokesbury. Implications of extremely high recruitment: crowding and reduced growth within spatial closures. Marine Ecology Progress Series, 611: 157–165, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m611p157.pdf.

Byrne:2019:EML

[BS19b] M. Byrne and M. A. Sewell. Evolution of maternal lipid provisioning strategies in echinoids with non-feeding larvae: selection for high-quality juveniles. Marine Ecology Progress Series, 616:95–106, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/616/m616p095.pdf.

Brothers:2016:CBD

[BSA^{+}16] C. J. Brothers, K. E. Smith, M. O. Amsler, R. B. Aronson, H. Singh, and J. B. McClintock. Covering behavior of deep-water echinoids in Antarctica: possible response to predatory king crabs. Marine Ecology Progress Series, 553: 155–162, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m553p155.pdf.

Benoit:2011:EPG

[BSB^{+}11a] H. P. Benoît, D. P. Swain, W. D. Bowen, G. A. Breed, M. O. Hammill, and V. Harvey. Evaluating the potential for grey seal predation to explain elevated natural mortality in three fish species in the southern Gulf of St. Lawrence. Marine Ecology Progress Series, 442:149–167, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/442/m442p149.pdf.

Bongaerts:2011:SDM

[BSB^{+}11b] P. Bongaerts, E. M. Sampayo, T. C. L. Bridge, T. Ridgway, F. Vermeulen, N. Englebert, J. M. Webster, and O. Hoegh-Guldberg. Symbiodinium diversity in mesophotic coral communities on the Great Barrier Reef: a first assessment. Marine Ecology Progress Series, 439:117–126, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/439/
m439p117.pdf.

Borgianini:2012:SMM

[BSB12] S. A. Borgianini, R. Styles, and R. J. Brodie. Simple model of megalopal transport in narrow river-dominated estuaries. Marine Ecology Progress Series, 452:179–191, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/452/
m452p179.pdf.

Bergstrom:2013:EOW

[BSB13] L. Bergström, F. Sundqvist, and U. Bergström. Effects of an offshore wind farm on temporal and spatial patterns in the demersal fish community. Marine Ecology Progress Series, 485:199–210, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m485p199.pdf.

Bittler:2014:FIB

[BSB14] K. M. Bittler, L. P. Scheef, and E. J. Buskey. Freshwater inflows and blue crabs: the influence of salinity on selective tidal stream transport. Marine Ecology Progress Series, 514:137–148, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/514/m514p137.pdf.

Bacheler:2018:BGT

[BSB^{+}18] N. M. Bacheler, K. W. Shertzer, J. A. Buckel, P. J. Rudershausen, and B. J. Runde. Behavior of gray triggerfish Balistes capriscus around baited fish traps determined from fine-scale acoustic tracking. Marine Ecology Progress Series, 606:133–150, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/606/m606p133.pdf.

Boersch-Supan:2012:ESF

[BSBR^{+}12] P. H. Boersch-Supan, L. Boehme, J. F. Read, A. D. Rogers, and A. S. Brierley. Elephant seal foraging dives track prey distribution, not temperature: Comment on McIntyre et al. (2011). Marine Ecology Progress Series, 461:293–298, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2012/461/m461p293.pdf. See [MTP^{+}10].

Baker:2013:SMF

[BSC^{+}13] C. S. Baker, D. Steel, J. Calambokidis, E. Falcone, U. González-Peral, J. Barlow, A. M. Burdin, P. J. Clapham, J. K. B. Ford, C. M. Gabriele, D. Mattila, L. Rojas-Bracho, J. M. Straley, B. L. Taylor, J. Urbán, P. R. Wade, D. Weller, B. H. Witteveen, and M. Yamaguchi. Strong maternal fidelity and natal philopatry shape genetic structure in North Pacific humpback whales. Marine Ecology Progress Series, 494:291–306, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m494p291.pdf.

Beltran:2015:FSW

[BSC^{+}15] R. S. Beltran, M. Connolly Sadou, R. Condit, S. Peterson, C. Reichmuth, and D. P. Costa. Fine-scale whisker growth measurements can reveal temporal foraging patterns from stable isotope signatures. Marine Ecology Progress Series, 523:243–253, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/523/m523p243.pdf.

Bernardello:2016:CRS

[BSC^{+}16] R. Bernardello, E. Serrano, R. Coma, M. Ribes, and N. Bahamon. A comparison of remote-sensing SST and in situ seawater temperature in near-shore habitats in the western Mediterranean Sea. Marine Ecology Progress Series, 559: 21–34, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/559/m559p021.pdf.

Blewett:2017:EER

[BSC17] D. A. Blewett, P. W. Stevens, and J. Carter. Ecological effects of river flooding on abundance and body condition of a large, euryhaline fish. Marine Ecology Progress Series, 563:211–218, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/563/m563p211.pdf.

Burford:2014:LAM

[BSCH14] M. O. Burford, J. Scarpa, B. J. Cook, and M. P. Hare. Local adaptation of a marine invertebrate with a high dispersal potential: evidence from a reciprocal transplant experiment of the eastern oyster Crassostrea virginica. Marine Ecology Progress Series, 505:161–175, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/505/m505p161.pdf.

Bertocci:2014:ELA

[BSD^{+}14] I. Bertocci, M. I. Seabra, R. Dominguez, D. Jacinto, R. Ramírez, J. Coca, and F. Tuya. Effects of loss of algal canopies along temperature and irradiation gradients in continental Portugal and the Canary Islands. Marine Ecology Progress Series, 506:47–60, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/506/m506p047.pdf.

Brooker:2017:CSA

[BSD17] R. M. Brooker, T. L. Sih, and D. L. Dixson. Contact with seaweed alters prey selectivity in a coral-feeding reef fish. Marine Ecology Progress Series, 580:239–244, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/580/
m580p239.pdf.

Baker:2017:IEV

[BSF17] M. C. Baker, M. A. Steinhoff, and G. F. Fricano. Integrated effects of the Deepwater Horizon oil spill on nearshore ecosystems. Marine Ecology Progress Series, 576:219–234, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m576p219.pdf.

Bargu:2010:CDA

[BSG^{+}10] S. Bargu, M. Silver, T. Goldstein, K. Roberts, and F. Gulland. Complexity of domoic acid-related sea lion strandings in Monterey Bay, California: foraging patterns, climate events, and toxic blooms. Marine Ecology Progress Series, 418:213–222, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/418/m418p213.pdf.

Borland:2017:HTB

[BSG^{+}17] H. P. Borland, T. A. Schlacher, B. L. Gilby, R. M. Connolly, N. A. Yabsley, and A. D. Olds. Habitat type and beach exposure shape fish assemblages in the surf zones of ocean beaches. Marine Ecology Progress Series, 570:203–211, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/570/
m570p203.pdf.

Blicher:2013:PSM

[BSH13] M. E. Blicher, M. K. Sejr, and S. Høgslund. Population structure of Mytilus edulis in the intertidal zone in a sub-Arctic fjord, SW Greenland. Marine Ecology Progress Series, 487:89–100, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/487/m487p089.pdf.

Burden:2014:GSC

[BSH^{+}14] C. T. Burden, A. J. Stow, S. J. Hoggard, M. A. Coleman, and M. J. Bishop. Genetic structure of Carcinus maenas in southeast Australia. Marine Ecology Progress Series, 500: 139–147, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/500/m500p139.pdf.

Benoit:2018:ERS

[BSH^{+}18] H. P. Benoît, D. P. Swain, J. A. Hutchings, D. Knox, T. Doniol-Valcroze, and C. M. Bourne. Evidence for reproductive senescence in a broadly distributed harvested marine fish. Marine Ecology Progress Series, 592:207–224, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/592/
m592p207.pdf.

Barshis:2011:CUL

[BSK^{+}11] D. J. Barshis, E. E. Sotka, R. P. Kelly, A. Sivasundar, B. A. Menge, J. A. Barth, and S. R. Palumbi. Coastal upwelling is linked to temporal genetic variability in the acorn barnacle Balanus glandula. Marine Ecology Progress Series, 439:139–150, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/439/m439p139.pdf.

Behrens:2018:IVA

[BSN^{+}18] J. W. Behrens, J. C. Svendsen, S. Neuenfeldt, N. G. Andersen, and M. van Deurs. Individual variation in aerobic scope affects modeled vertical foraging migration in Atlantic cod Gadus morhua, but only in moderate hypoxia. Marine Ecology Progress Series, 599:201–208, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m599p201.pdf.

Billard:2010:FVS

[BSP^{+}10] E. Billard, E. Serrão, G. Pearson, C. Destombe, and M. Valero. Fucus vesiculosus and spiralis species complex: a nested model of local adaptation at the shore level. Marine Ecology Progress Series, 405:163–174, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m405p163.pdf.

Ben-Shlomo:2010:PGI

[BSRDR10] R. Ben-Shlomo, E. Reem, J. Douek, and B. Rinkevich. Population genetics of the invasive ascidian Botryllus schlosseri from South American coasts. Marine Ecology Progress Series, 412:85–92, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/412/m412p085.pdf.

Buchwald:2018:DQK

[BSS18] R. Buchwald, R. E. Scheibling, and A. G. B. Simpson. Detection and quantification of a keystone pathogen in a coastal marine ecosystem. Marine Ecology Progress Series, 606: 79–90, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/606/m606p079.pdf.

Bernreuther:2013:VRP

[BSST13] M. Bernreuther, J. Schmidt, D. Stepputtis, and A. Temming. Vertically resolved prey selectivity and competition of Baltic herring Clupea harengus and sprat Sprattus sprattus. Marine Ecology Progress Series, 489:177–195, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/489/
m489p177.pdf.

Bertics:2010:BDB

[BST^{+}10] V. J. Bertics, J. A. Sohm, T. Treude, C. E. T. Chow, D. G. Capone, J. A. Fuhrman, and W. Ziebis. Burrowing deeper into benthic nitrogen cycling: the impact of bioturbation on nitrogen fixation coupled to sulfate reduction. Marine Ecology Progress Series, 409:1–15, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ballard:2019:FSO

[BST^{+}19a] G. Ballard, A. E. Schmidt, V. Toniolo, S. Veloz, D. Jongsomjit, K. R. Arrigo, and D. G. Ainley. Fine-scale oceanographic features characterizing successful Adélie penguin foraging in the SW Ross Sea. Marine Ecology Progress Series, 608:263–277, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m608p263.pdf.

Belcher:2019:RRA

[BST19b] A. Belcher, R. A. Saunders, and G. A. Tarling. Respiration rates and active carbon flux of mesopelagic fishes (family Myctophidae) in the Scotia Sea, Southern Ocean. Marine Ecology Progress Series, 610:149–162, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/610/m610p149.pdf.

Bustnes:2013:CDF

[BSY13] J. O. Bustnes, G. H. Systad, and R. C. Ydenberg. Changing distribution of flocking sea ducks as non-regenerating food resources are depleted. Marine Ecology Progress Series, 484:249–257, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/484/m484p249.pdf.

Bailey:2010:EOF

[BT10] H. Bailey and P. Thompson. Effect of oceanographic features on fine-scale foraging movements of bottlenose dolphins. Marine Ecology Progress Series, 418:223–233, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/418/
m418p223.pdf.

Brady:2013:MJW

[BT13] D. C. Brady and T. E. Targett. Movement of juvenile weakfish Cynoscion regalis and spot Leiostomus xanthurus in relation to diel-cycling hypoxia in an estuarine tidal tributary. Marine Ecology Progress Series, 491:199–219, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/491/
m491p199.pdf.

Berthelsen:2014:AMR

[BT14] A. K. Berthelsen and R. B. Taylor. Arthropod mesograzers reduce epiphytic overgrowth of subtidal coralline turf. Marine Ecology Progress Series, 515:123–132, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/515/
m515p123.pdf.

Bevilacqua:2016:SSE

[BT16] S. Bevilacqua and A. Terlizzi. Species surrogacy in environmental impact assessment and monitoring: extending the BestAgg approach to asymmetrical designs. Marine Ecology Progress Series, 547:19–32, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/547/m547p019.pdf.

Blake:2017:GFP

[BTLF17] C. Blake, M. Thiel, B. A. López, and C. I. Fraser. Gall-forming protistan parasites infect southern bull kelp across the Southern Ocean, with prevalence increasing to the south. Marine Ecology Progress Series, 583:95–106, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m583p095.pdf.

Bulleri:2018:EMI

[BTM^{+}18] F. Bulleri, L. Thiault, S. C. Mills, M. M. Nugues, E. M. Eckert, G. Corno, and J. Claudet. Erect macroalgae influence epilithic bacterial assemblages and reduce coral recruitment. Marine Ecology Progress Series, 597:65–77, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/597/
m597p065.pdf.

Baylis:2018:GVF

[BTO^{+}18] A. M. M. Baylis, M. Tierney, R. A. Orben, I. J. Staniland, and P. Brickle. Geographic variation in the foraging behaviour of South American fur seals. Marine Ecology Progress Series, 596:233–245, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/596/m596p233.pdf.

Basedow:2010:HSM

[BTS10] S. L. Basedow, K. S. Tande, and L. C. Stige. Habitat selection by a marine copepod during the productive season in the subarctic. Marine Ecology Progress Series, 416:165–178, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m416p165.pdf.

Barnes:2014:ABA

[BTS^{+}14] M. K. Barnes, G. H. Tilstone, T. J. Smyth, D. J. Suggett, R. Astoreca, C. Lancelot, and J. C. Kromkamp. Absorptionbased algorithm of primary production for total and sizefractionated phytoplankton in coastal waters. Marine Ecology Progress Series, 504:73–89, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m504p073.pdf.

Ben-Tzvi:2010:RCD

[BTTK10] O. Ben-Tzvi, D. Tchernov, and M. Kiflawi. Role of coralderived chemical cues in microhabitat selection by settling Chromis viridis. Marine Ecology Progress Series, 409:181– 187, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/409/m409p181.pdf.

Burrows:2012:IWF

[Bur12] M. T. Burrows. Influences of wave fetch, tidal flow and ocean colour on subtidal rocky communities. Marine Ecology Progress Series, 445:193–207, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/445/m445p193.pdf.

Bush:2012:EAA

[Bus12] S. L. Bush. Economy of arm autotomy in the mesopelagic squid Octopoteuthis deletron. Marine Ecology Progress Series, 458:133–140, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/458/m458p133.pdf.

Bond:2017:ASS

[BVAB^{+}17] M. E. Bond, J. Valentin-Albanese, E. A. Babcock, D. Abercrombie, N. F. Lamb, A. Miranda, E. K. Pikitch, and D. D. Chapman. Abundance and size structure of a reef shark population within a marine reserve has remained stable for more than a decade. Marine Ecology Progress Series, 576:1–10, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Burford:2016:ISS

[BVC^{+}16] M. A. Burford, D. Valdez, G. Curwen, S. J. Faggotter, D. P. Ward, and K. R. O’Brien. Inundation of saline supratidal mudflats provides an important source of carbon and nutrients in an aquatic system. Marine Ecology Progress Series, 545:21–33, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/545/m545p021.pdf.

Burghart:2014:DBP

[BVD^{+}14] S. E. Burghart, L. Van Woudenberg, C. A. Daniels, S. D. Meyers, E. B. Peebles, and M. Breitbart. Disparity between planktonic fish egg and larval communities as indicated by DNA barcoding. Marine Ecology Progress Series, 503:195–204, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m503p195.pdf.

Bearham:2013:TLE

[BVG13] D. Bearham, M. A. Vanderklift, and J. R. Gunson. Temperature and light explain spatial variation in growth and productivity of the kelp Ecklonia radiata. Marine Ecology Progress Series, 476:59–70, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/476/m476p059.pdf.

Brin:2010:NIE

[BVGH10] L. D. Brin, I. Valiela, D. Goehringer, and B. Howes. Nitrogen interception and export by experimental salt marsh plots exposed to chronic nutrient addition. Marine Ecology Progress Series, 400:3–17, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Bednarz:2015:DFP

[BvHC^{+}15] V. N. Bednarz, N. van Hoytema, U. Cardini, M. S. Naumann, M. M. D. Al-Rshaidat, and C. Wild. Dinitrogen fixation and primary productivity by carbonate and silicate reef sand communities of the Northern Red Sea. Marine Ecology Progress Series, 527:47–57, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/527/m527p047.pdf.

Balestri:2010:ESD

[BVL10] E. Balestri, F. Vallerini, and C. Lardicci. Effect of seed density and sediment nutrient heterogeneity on recruitment and early patch growth in the seagrass Cymodocea nodosa. Marine Ecology Progress Series, 417:63–72, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/417/
m417p063.pdf.

Barbosa:2019:ISC

[BVM^{+}19] A. C. C. Barbosa, C. Vinagre, D. Mizrahi, R. C. Duarte, and A. A. V. Flores. Invasive sun corals and warming pose independent threats to the brain coral Mussismilia hispida in the southwestern Atlantic. Marine Ecology Progress Series, 629:43–54, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/629/m629p043.pdf.

Babarro:2016:IPP

[BVO16] J. M. F. Babarro, E. Vázquez, and C. Olabarria. Importance of phenotypic plastic traits on invasion success: response of Xenostrobus securis to the predatory dogwhelk Nucella lapillus. Marine Ecology Progress Series, 560:185– 198, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/560/m560p185.pdf.

Bode:2011:CMS

[BVP11] A. Bode, M. Varela, and R. Prego. Continental and marine sources of organic matter and nitrogen for rías of northern Galicia (Spain). Marine Ecology Progress Series, 437: 13–26, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m437p013.pdf.

Brante:2012:RPT

[BVQ12] A. Brante, C. Vilches, and A. Quiñones. Role of population traits and local social interactions in sex-change plasticity of the protandric marine gastropod Crepidula coquimbensis. Marine Ecology Progress Series, 458:123–132, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/458/
m458p123.pdf.

Braeckman:2011:CMA

[BVS^{+}11] U. Braeckman, C. Van Colen, K. Soetaert, M. Vincx, and J. Vanaverbeke. Contrasting macrobenthic activities differentially affect nematode density and diversity in a shallow subtidal marine sediment. Marine Ecology Progress Series, 422:179–191, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/422/m422p179.pdf.

Bandara:2016:SVS

[BVS^{+}16] K. Bandara, Ø. Varpe, J. E. Søreide, J. Wallenschus, J. Berge, and K. Eiane. Seasonal vertical strategies in a high-Arctic coastal zooplankton community. Marine Ecology Progress Series, 555:49–64, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m555p049.pdf.

Boudreau:2010:TCL

[BW10] S. A. Boudreau and B. Worm. Top-down control of lobster in the Gulf of Maine: insights from local ecological knowledge and research surveys. Marine Ecology Progress Series, 403:181–191, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m403p181.pdf.

Boudreau:2012:ERL

[BW12] S. A. Boudreau and B. Worm. Ecological role of large benthic decapods in marine ecosystems: a review. Marine Ecology Progress Series, 469:195–213, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m469p195.pdf.

Bauer:2013:FAC

[BW13] A. Bauer and J. J. Waniek. Factors affecting chlorophyll-a concentration in the central Beibu Gulf, South China Sea. Marine Ecology Progress Series, 474:67–88, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/474/
m474p067.pdf.

Boyce:2015:PEI

[BW15] D. G. Boyce and B. Worm. Patterns and ecological implications of historical marine phytoplankton change. Marine Ecology Progress Series, 534:251–272, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/534/
m534p251.pdf.

Baamstedt:2016:MDA

[BW16a] U. Båmstedt and J. Wikner. Mixing depth and allochthonous dissolved organic carbon: controlling factors of coastal trophic balance. Marine Ecology Progress Series, 561:17–29, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/561/m561p017.pdf. See erratum [BW16b].

Baamstedt:2016:EMD

[BW16b] Ulf Båmstedt and Johan Wikner. Erratum: Mixing depth and allochthonous dissolved organic carbon: controlling factors of coastal trophic balance, Mar Ecol Prog Ser 561:17–29, 2016. Marine Ecology Progress Series, 562:263, 2016. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m562p263.pdf. See [BW16a].

Bricker:2011:HCA

[BWCZ11] E. Bricker, M. Waycott, A. Calladine, and J. C. Zieman. High connectivity across environmental gradients and implications for phenotypic plasticity in a marine plant. Marine Ecology Progress Series, 423:57–67, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m423p057.pdf.

Beer:2011:PVB

[BWH11] N. A. Beer, S. R. Wing, and Y. Hu. Physical versus biological control of element incorporation into biogenic carbonate: an in situ experiment in a New Zealand fjord. Marine Ecology Progress Series, 433:289–301, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/433/m433p289.pdf.

Bodmer:2017:IET

[BWH^{+}17] M. D. V. Bodmer, P. M. Wheeler, A. M. Hendrix, D. N. Cesarano, A. S. East, and D. A. Exton. Interacting effects of temperature, habitat and phenotype on predator avoidance behaviour in Diadema antillarum: implications for restorative conservation. Marine Ecology Progress Series, 566:105– 115, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/566/m566p105.pdf.

Blake:2019:SCS

[BWH^{+}19] R. E. Blake, C. L. Ward, M. E. Hunsicker, A. O. Shelton, and A. B. Hollowed. Spatial community structure of groundfish is conserved across the Gulf of Alaska. Marine Ecology Progress Series, 626:145–160, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m626p145.pdf.

Bayer:2016:MSF

[BWJB16] S. R. Bayer, R. A. Wahle, P. A. Jumars, and D. C. Brady. Measuring scallop fertilization success in the field: chamber design and tests. Marine Ecology Progress Series, 551: 141–154, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/551/m551p141.pdf.

Baumgartner:2017:NAR

[BWLP17] M. F. Baumgartner, F. W. Wenzel, N. S. J. Lysiak, and M. R. Patrician. North Atlantic right whale foraging ecology and its role in human-caused mortality. Marine Ecology Progress Series, 581:165–181, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m581p165.pdf.

Bacheler:2015:MIA

[BWM^{+}15] N. M. Bacheler, P. E. Whitfield, R. C. Muñoz, B. B. Harrison, C. A. Harms, and C. A. Buckel. Movement of invasive adult lionfish Pterois volitans using telemetry: importance of controls to estimate and explain variable detection probabilities. Marine Ecology Progress Series, 527:205–220, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/527/
m527p205.pdf.

Burthe:2014:AVM

[BWN^{+}14] S. J. Burthe, S. Wanless, M. A. Newell, A. Butler, and F. Daunt. Assessing the vulnerability of the marine bird community in the western North Sea to climate change and other anthropogenic impacts. Marine Ecology Progress Series, 507:277–295, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/507/m507p277.pdf.

Blundell:2011:UGT

[BWP^{+}11] G. M. Blundell, J. N. Womble, G. W. Pendleton, S. A. Karpovich, S. M. Gende, and J. K. Herreman. Use of glacial and terrestrial habitats by harbor seals in Glacier Bay, Alaska: costs and benefits. Marine Ecology Progress Series, 429:277– 290, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m429p277.pdf.

Brosnan:2014:EIE

[BWRP14] I. G. Brosnan, D. W. Welch, E. L. Rechisky, and A. D. Porter. Evaluating the influence of environmental factors on yearling Chinook salmon survival in the Columbia River plume (USA). Marine Ecology Progress Series, 496:181–196, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m496p181.pdf.

Beer:2011:OEE

[BWS11] N. A. Beer, S. R. Wing, and S. E. Swearer. Otolith elemental evidence for spatial structuring in a temperate reef fish population. Marine Ecology Progress Series, 442:217– 227, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/442/m442p217.pdf.

Baker:2012:DPD

[BWS^{+}12] K. D. Baker, V. E. Wareham, P. V. R. Snelgrove, R. L. Haedrich, D. A. Fifield, E. N. Edinger, and K. D. Gilkinson. Distributional patterns of deep-sea coral assemblages in three submarine canyons off Newfoundland, Canada. Marine Ecology Progress Series, 445:235–249, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m445p235.pdf.

Barott:2012:NHC

[BWV^{+}12] K. L. Barott, G. J. Williams, M. J. A. Vermeij, J. Harris, J. E. Smith, F. L. Rohwer, and S. A. Sandin. Natural history of coral–algae competition across a gradient of human activity in the Line Islands. Marine Ecology Progress Series, 460:1–12, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Barnett:2013:TEA

[BYAA13] A. Barnett, J. L. Yick, K. G. Abrantes, and C. A. Awruch. Trophic ecology of an abundant predator and its relationship with fisheries. Marine Ecology Progress Series, 494: 241–248, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/494/m494p241.pdf.

Bar-Zeev:2011:TEP

[BZBR^{+}11] E. Bar-Zeev, T. Berman, E. Rahav, G. Dishon, B. Herut, and I. Berman-Frank. Transparent exopolymer particle (TEP) dynamics in the eastern Mediterranean Sea. Marine Ecology Progress Series, 431:107–118, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/431/m431p107.pdf.

Bailey:2012:NRT

[BZC^{+}12] K. M. Bailey, T. Zhang, K. S. Chan, S. M. Porter, and A. B. Dougherty. Near real-time forecasting of recruitment from larval surveys: application to Alaska pollock. Marine Ecology Progress Series, 452:205–217, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/452/m452p205.pdf.

Castellani:2014:SCA

[CA14] C. Castellani and Y. Altunba. Seasonal change in acclimatised respiration rate of Temora longicornis. Marine Ecology Progress Series, 500:83–101, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/500/m500p083.pdf.

Clausing:2014:ESD

[CAB^{+}14] R. J. Clausing, C. Annunziata, G. Baker, C. Lee, S. J. Bittick, and P. Fong. Effects of sediment depth on algal turf height are mediated by interactions with fish herbivory on a fringing reef. Marine Ecology Progress Series, 517:121–129, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/517/
m517p121.pdf.

Culloch:2016:ECR

[CAB^{+}16] R. M. Culloch, P. Anderwald, A. Brandecker, D. Haberlin, B. McGovern, R. Pinfield, F. Visser, M. Jessopp, and M. Cronin. Effect of construction-related activities and vessel traffic on marine mammals. Marine Ecology Progress Series, 549:231–242, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m549p231.pdf.

Chappell:2019:PDD

[CAB^{+}19a] P. D. Chappell, E. V. Armbrust, K. A. Barbeau, R. M. Bundy, J. W. Moffett, J. Vedamati, and B. D. Jenkins. Patterns of diatom diversity correlate with dissolved trace metal concentrations and longitudinal position in the northeast Pacific coastal-offshore transition zone. Marine Ecology Progress Series, 609:69–86, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m609p069.pdf.

Cleeland:2019:FIH

[CAB^{+}19b] J. B. Cleeland, R. Alderman, A. Bindoff, M. A. Lea, C. R. McMahon, R. A. Phillips, B. Raymond, M. D. Sumner, A. Terauds, S. J. Wotherspoon, and M. A. Hindell. Factors influencing the habitat use of sympatric albatrosses from Macquarie Island, Australia. Marine Ecology Progress Series, 609:221–237, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/609/m609p221.pdf.

Christofoletti:2011:EGI

[CAC11] R. A. Christofoletti, T. V. V. Almeida, and Á. M. Ciotti. Environmental and grazing influence on spatial variability of intertidal biofilm on subtropical rocky shores. Marine Ecology Progress Series, 424:15–23, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/424/m424p015.pdf.

Cama:2012:MSC

[CAC^{+}12] A. Cama, R. Abellana, I. Christel, X. Ferrer, and D. R. Vieites. Moving to the sea: a challenge for an inshore species, the slender-billed gull. Marine Ecology Progress Series, 463:285–295, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/463/m463p285.pdf.

Canepuccia:2010:EEB

[CAD^{+}10] A. D. Canepuccia, J. Alberti, P. Daleo, J. Pascual, J. L. Farina, and O. O. Iribarne. Ecosystem engineering by burrowing crabs increases cordgrass mortality caused by stem-boring insects. Marine Ecology Progress Series, 404:151–159, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/404/
m404p151.pdf.

Comeau:2012:EOA

[CAG12] S. Comeau, S. Alliouane, and J. P. Gattuso. Effects of ocean acidification on overwintering juvenile Arctic pteropods Limacina helicina. Marine Ecology Progress Series, 456:279– 284, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m456p279.pdf.

Claudino:2013:SIR

[CAG13] M. C. Claudino, P. C. Abreu, and A. M. Garcia. Stable isotopes reveal temporal and between-habitat changes in trophic pathways in a southwestern Atlantic estuary. Marine Ecology Progress Series, 489:29–42, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/489/m489p029.pdf.

Cortes:2017:SDL

[CAGS17] V. Cortés, J. M. Arcos, and J. González-Solís. Seabirds and demersal longliners in the northwestern Mediterranean: factors driving their interactions and bycatch rates. Marine Ecology Progress Series, 565:1–16, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Clark:2014:OCD

[CAKD14] C. D. Clark, P. Aiona, J. K. Keller, and W. J. De Bruyn. Optical characterization and distribution of chromophoric dissolved organic matter (CDOM) in soil porewater from a salt marsh ecosystem. Marine Ecology Progress Series, 516: 71–83, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/516/m516p071.pdf.

Colmenero:2010:SCT

[CALB10] A. I. Colmenero, J. Aguzzi, A. Lombarte, and A. Bozzano. Sensory constraints in temporal segregation in two species of anglerfish, Lophius budegassa and L. piscatorius. Marine Ecology Progress Series, 416:255–265, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/416/
m416p255.pdf.

Carlson:2012:FAB

[Car12] B. A. Carlson. Feeding activity by the blenny Exallias brevis causes multifocal bleaching in corals: Comment on Zvuloni et al. (2011). Marine Ecology Progress Series, 463:297–299, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_

oa/m463p297.pdf. See [ZAZS11, ZAZS12].

Caccavo:2019:SSL

[CAR^{+}19] J. A. Caccavo, J. R. Ashford, S. Ryan, C. Papetti, M. Schröder, and L. Zane. Spatial structuring and life history connectivity of Antarctic silverfish along the southern continental shelf of the Weddell Sea. Marine Ecology Progress Series, 624:195–212, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m624p195.pdf.

Clay:2019:PRA

[CASG^{+}19] T. A. Clay, J. Alfaro-Shigueto, B. J. Godley, N. Tregenza, and J. C. Mangel. Pingers reduce the activity of Burmeister’s porpoise around small-scale gillnet vessels. Marine Ecology Progress Series, 626:197–208, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/626/m626p197.pdf.

Cruz-Acevedo:2018:DSF

[CATAV18] E. Cruz-Acevedo, N. Tolimieri, and H. Aguirre-Villaseñor. Deep-sea fish assemblages (300−2100 m) in the eastern Pacific off northern Mexico. Marine Ecology Progress Series, 592:225–242, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/592/m592p225.pdf.

Carnegie:2011:DII

[CB11] R. B. Carnegie and E. M. Burreson. Declining impact of an introduced pathogen: Haplosporidium nelsoni in the oyster Crassostrea virginica in Chesapeake Bay. Marine Ecology Progress Series, 432:1–15, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Cook:2012:UFS

[CB12] A. M. Cook and A. Bundy. Use of fishes as sampling tools for understanding biodiversity and ecosystem functioning in the ocean. Marine Ecology Progress Series, 454:1–18, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Carr:2014:VMT

[CB14] L. A. Carr and K. E. Boyer. Variation at multiple trophic levels mediates a novel seagrass–grazer interaction. Marine Ecology Progress Series, 508:117–128, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/508/m508p117.pdf.

Chiswell:2017:ELL

[CB17] S. M. Chiswell and J. D. Booth. Evolution of long larval life in the Australasian rock lobster Jasus edwardsii. Marine Ecology Progress Series, 576:69–87, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/576/m576p069.pdf.

Campos:2010:FBS

[CBC^{+}10] J. Campos, A. Bio, J. F. M. F. Cardoso, R. Dapper, J. I. J. Witte, and H. W. van der Veer. Fluctuations of brown shrimp Crangon crangon abundance in the western Dutch Wadden Sea. Marine Ecology Progress Series, 405:203–219, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2010/405/m405p203.pdf.

Cama:2013:FMS

[CBC^{+}13] A. Cama, J. Bort, I. Christel, D. R. Vieites, and X. Ferrer. Fishery management has a strong effect on the distribution of Audouin’s gull. Marine Ecology Progress Series, 484:279–286, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/484/m484p279.pdf.

Cross:2016:GES

[CBC^{+}16] M. E. Cross, C. R. Bradley, T. F. Cross, S. Culloty, S. Lynch, P. McGinnity, R. M. O’Riordan, S. Vartia, and P. A. Prodöhl. Genetic evidence supports recolonisation by Mya arenaria of western Europe from North America. Marine Ecology Progress Series, 549:99–112, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/549/m549p099.pdf.

Coleman:2014:MAG

[CBD14] D. W. Coleman, M. Byrne, and A. R. Davis. Molluscs on acid: gastropod shell repair and strength in acidifying oceans. Marine Ecology Progress Series, 509:203–211, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/509/
m509p203.pdf.

Comerford:2013:TEG

[CBF^{+}13] S. Comerford, D. Brophy, C. J. Fox, N. Taylor, H. W. van der Veer, R. D. M. Nash, and A. J. Geffen. Temperature effect on growth and larval duration of plaice Pleuronectes platessa in three regions of the Northeast Atlantic. Marine Ecology Progress Series, 476:215–226, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/476/m476p215.pdf.

Cleary:2019:PDD

[CBF^{+}19] A. C. Cleary, M. Bester, J. Forcada, M. Goebel, S. D. Goldsworthy, C. Guinet, J. I. Hoffman, K. M. Kovacs, C. Lydersen, and A. D. Lowther. Prey differences drive local genetic adaptation in Antarctic fur seals. Marine Ecology Progress Series, 628:195–209, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/628/m628p195.pdf.

Clements:2012:ESF

[CBGDS12] C. Clements, V. Bonito, R. Grober-Dunsmore, and M. Sobey. Effects of small, Fijian community-based marine protected areas on exploited reef fishes. Marine Ecology Progress Series, 449:233–243, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m449p233.pdf.

Crosbie:2019:RRC

[CBJB19] A. J. Crosbie, T. C. L. Bridge, G. Jones, and A. H. Baird. Response of reef corals and fish at Osprey Reef to a thermal anomaly across a 30 m depth gradient. Marine Ecology Progress Series, 622:93–102, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/622/m622p093.pdf.

Cure:2012:CBR

[CBK^{+}12] K. Cure, C. E. Benkwitt, T. L. Kindinger, E. A. Pickering, T. J. Pusack, J. L. McIlwain, and M. A. Hixon. Comparative behavior of red lionfish Pterois volitans on native Pacific versus invaded Atlantic coral reefs. Marine Ecology Progress Series, 467:181–192, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/467/m467p181.pdf.

Carrea:2016:PSL

[CBKM16] C. Carrea, C. P. Burridge, C. K. King, and K. J. Miller. Population structure and long-term decline in three species of heart urchins Abatus spp. near-shore in the vestfold hills region, East Antarctica. Marine Ecology Progress Series, 545:227–238, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m545p227.pdf.

Clua:2010:EPS

[CBL^{+}10a] E. Clua, N. Buray, P. Legendre, J. Mourier, and S. Planes. Effects of provisioning on shark behaviour: Reply to Brunnschweiler & McKenzie (2010). Marine Ecology Progress Series, 420:285–288, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/420/m420p285.pdf. See

[BM10b].

Clua:2010:BRS

[CBL^{+}10b] Eric Clua, Nicolas Buray, Pierre Legendre, Johann Mourier, and Serge Planes. Behavioural response of sicklefin lemon sharks Negaprion acutidens to underwater feeding for ecotourism purposes. Marine Ecology Progress Series, 414: 257–266, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/414/m414p257.pdf. See comment

[BM10b].

Cummings:2010:ODS

[CBL^{+}10c] D. O. Cummings, D. J. Booth, R. W. Lee, S. J. Simpson, and A. J. Pile. Ontogenetic diet shifts in the reef fish Pseudanthias rubrizonatus from isolated populations on the North-West shelf of Australia. Marine Ecology Progress Series, 419:211–222, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/419/m419p211.pdf.

Carneiro:2017:MDQ

[CBLM^{+}17] A. P. B. Carneiro, A. S. Bonnet-Lebrun, A. Manica, I. J. Staniland, and R. A. Phillips. Methods for detecting and quantifying individual specialisation in movement and foraging strategies of marine predators. Marine Ecology Progress Series, 578:151–166, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m578p151.pdf.

Casini:2010:LFT

[CBMK10] M. Casini, V. Bartolino, J. C. Molinero, and G. Kornilovs. Linking fisheries, trophic interactions and climate: threshold dynamics drive herring Clupea harengus growth in the central Baltic Sea. Marine Ecology Progress Series, 413:241–252, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m413p241.pdf.

Carlucci:2018:EST

[CBR^{+}18] R. Carlucci, V. Bandelj, P. Ricci, F. Capezzuto, L. Sion, P. Maiorano, A. Tursi, C. Solidoro, and S. Libralato. Exploring spatio-temporal changes in the demersal and benthopelagic assemblages of the north-western Ionian Sea (central Mediterranean Sea). Marine Ecology Progress Series, 598:1–19, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Cherel:2019:AAS

[CBR19] Y. Cherel, P. Bustamante, and P. Richard. Amino acid δ ^{13}C and δ ^{15}N from sclerotized beaks: a new tool to investigate the foraging ecology of cephalopods, including giant and colossal squids. Marine Ecology Progress Series, 624:89–102, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m624p089.pdf.

Chargulaf:2013:SGR

[CBT13] C. A. Chargulaf, D. D. Burfeind, and I. R. Tibbetts. A sand goby realizes its niche both at high population densities and in the presence of the half bridled goby. Marine Ecology Progress Series, 488:247–254, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/488/m488p247.pdf.

Cozzoli:2013:ANL

[CBYH13] F. Cozzoli, T. J. Bouma, T. Ysebaert, and P. M. J. Herman. Application of non-linear quantile regression to macrozoobenthic species distribution modelling: comparing two contrasting basins. Marine Ecology Progress Series, 475:119– 133, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/475/m475p119.pdf.

Conway-Cranos:2012:GVR

[CC12] L. L. Conway-Cranos. Geographic variation in resilience: an experimental evaluation of four rocky intertidal assemblages. Marine Ecology Progress Series, 457:67–83, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/457/
m457p067.pdf.

Campos:2019:IVS

[CC19] P. Campos and L. Cardona. Individual variability in the settlement of juvenile green turtles in the western South Atlantic Ocean: relevance of currents and somatic growth rate. Marine Ecology Progress Series, 614:173–182, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/614/
m614p173.pdf.

Clavier:2014:BMS

[CCA^{+}14] J. Clavier, L. Chauvaud, E. Amice, P. Lazure, M. van der Geest, P. Labrosse, A. Diagne, A. Carlier, and S. Chauvaud. Benthic metabolism in shallow coastal ecosystems of the Banc d’Arguin, Mauritania. Marine Ecology Progress Series, 501:11–23, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/501/m501p011.pdf.

Camprasse:2017:MSF

[CCA^{+}17] E. C. M. Camprasse, Y. Cherel, J. P. Y. Arnould, A. J. Hoskins, P. Bustamante, and C. A. Bost. Mate similarity in foraging Kerguelen shags: a combined bio-logging and stable isotope investigation. Marine Ecology Progress Series, 578:183–196, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m578p183.pdf.

Camprasse:2017:III

[CCB^{+}17] E. C. M. Camprasse, Y. Cherel, P. Bustamante, J. P. Y. Arnould, and C. A. Bost. Intra- and inter-individual variation in the foraging ecology of a generalist subantarctic seabird, the gentoo penguin. Marine Ecology Progress Series, 578:227–242, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m578p227.pdf.

Clusa:2018:PLT

[CCC^{+}18] M. Clusa, C. Carreras, L. Cardona, A. Demetropoulos, D. Margaritoulis, A. F. Rees, A. A. Hamza, M. Khalil, Y. Levy, O. Turkozan, A. Aguilar, and M. Pascual. Philopatry in loggerhead turtles Caretta caretta: beyond the gender paradigm. Marine Ecology Progress Series, 588:201–213, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/588/
m588p201.pdf.

Claydon:2015:SDC

[CCDP15] J. A. B. Claydon, M. C. Calosso, G. A. De Leo, and R. B. J. Peachey. Spatial and demographic consequences of nurserydependence in reef fishes: an empirical and simulation study. Marine Ecology Progress Series, 525:171–183, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/525/
m525p171.pdf.

Cardona:2014:DPF

[CCE^{+}14] L. Cardona, M. Clusa, E. Eder, A. Demetropoulos, D. Margaritoulis, A. F. Rees, A. A. Hamza, M. Khalil, Y. Levy, O. Türkozan, I. Marín, and A. Aguilar. Distribution patterns and foraging ground productivity determine clutch size in Mediterranean loggerhead turtles. Marine Ecology Progress Series, 497:229–241, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2014/497/m497p229.pdf.

Castro:2010:FAE

[CCG^{+}10] L. R. Castro, G. Claramunt, H. E. González, M. C. Krautz, A. Llanos-Rivera, J. Méndez, W. Schneider, and S. Soto. Fatty acids in eggs of anchoveta Engraulis ringens during two contrasting winter spawning seasons. Marine Ecology Progress Series, 420:193–205, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/420/m420p193.pdf.

Criales:2019:BCL

[CCG^{+}19] M. M. Criales, L. Chérubin, R. Gandy, L. Garavelli, M. Ali Ghannami, and C. Crowley. Blue crab larval dispersal highlights population connectivity and implications for fishery management. Marine Ecology Progress Series, 625:53–70, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps_oa/m625p053.pdf.

Crossin:2014:TFM

[CCGP14] G. T. Crossin, S. J. Cooke, J. A. Goldbogen, and R. A. Phillips. Tracking fitness in marine vertebrates: current knowledge and opportunities for future research. Marine Ecology Progress Series, 496:1–17, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/496/m496p001.pdf.

Chang:2010:ESD

[CCHG10] J. H. Chang, Y. Chen, D. Holland, and J. Grabowski. Estimating spatial distribution of American lobster Homarus americanus using habitat variables. Marine Ecology Progress Series, 420:145–156, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/420/m420p145.pdf.

Cherel:2014:SYR

[CCJR14] Y. Cherel, M. Connan, A. Jaeger, and P. Richard. Seabird year-round and historical feeding ecology: blood and feather δ ^{13}C and δ ^{15}N values document foraging plasticity of small sympatric petrels. Marine Ecology Progress Series, 505:267– 280, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m505p267.pdf.

Conway-Cranos:2015:SIO

[CCKB^{+}15] L. Conway-Cranos, P. Kiffney, N. Banas, M. Plummer, S. Naman, P. MacCready, J. Bucci, and M. Ruckelshaus. Stable isotopes and oceanographic modeling reveal spatial and trophic connectivity among terrestrial, estuarine, and marine environments. Marine Ecology Progress Series, 533: 15–28, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/533/m533p015.pdf.

Cupido:2012:SSH

[CCM^{+}12] R. Cupido, S. Cocito, V. Manno, S. Ferrando, A. Peirano, M. Iannelli, L. Bramanti, and G. Santangelo. Sexual structure of a highly reproductive, recovering gorgonian population: quantifying reproductive output. Marine Ecology Progress Series, 469:25–36, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/469/m469p025.pdf.

Campbell:2014:RCA

[CCM^{+}14] J. E. Campbell, J. D. Craft, N. Muehllehner, C. Langdon, and V. J. Paul. Responses of calcifying algae (Halimeda spp.) to ocean acidification: implications for herbivores.

Marine Ecology Progress Series, 514:43–56, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/514/
m514p043.pdf.

Carter:2015:MSA

[CCM^{+}15] A. B. Carter, A. G. Carton, M. I. McCormick, A. J. Tobin, and A. J. Williams. Maternal size, not age, influences egg quality of a wild, protogynous coral reef fish Plectropomus leopardus. Marine Ecology Progress Series, 529:249–263, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/529/
m529p249.pdf.

Childs:2015:HCI

[CCNB15] A. R. Childs, P. D. Cowley, T. F. Næsje, and R. H. Bennett. Habitat connectivity and intra-population structure of an estuary-dependent fishery species. Marine Ecology Progress Series, 537:233–245, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2015/537/m537p233.pdf.

Christensen:2014:CFB

[CCP^{+}14] V. Christensen, M. Coll, C. Piroddi, J. Steenbeek, J. Buszowski, and D. Pauly. A century of fish biomass decline in the ocean. Marine Ecology Progress Series, 512:155–166, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m512p155.pdf.

Cimino:2018:OAR

[CCS^{+}18] M. Cimino, P. Colin, T. Schramek, S. Lindfield, M. Domeier, and E. Terrill. Oceanographic, acoustic, and remote approaches reveal the spatio-temporal dynamics of blackfin snapper at an aggregation site in Palau. Marine Ecology Progress Series, 601:185–201, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/601/m601p185.pdf.

Claydon:2012:PIL

[CCT12] J. A. B. Claydon, M. C. Calosso, and S. B. Traiger. Progression of invasive lionfish in seagrass, mangrove and reef habitats. Marine Ecology Progress Series, 448:119–129, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/448/
m448p119.pdf.

Choy:2013:PDO

[CD13] C. A. Choy and J. C. Drazen. Plastic for dinner? Observations of frequent debris ingestion by pelagic predatory fishes from the central North Pacific. Marine Ecology Progress Series, 485:155–163, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m485p155.pdf.

Crook:2014:UBC

[CD14] K. A. Crook and G. K. Davoren. Underwater behaviour of common murres foraging on capelin: influences of prey density and antipredator behaviour. Marine Ecology Progress Series, 501:279–290, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/501/m501p279.pdf.

Celentano:2016:ECM

[CD16a] E. Celentano and O. Defeo. Effects of climate on the mole crab Emerita brasiliensis on a dissipative beach in Uruguay. Marine Ecology Progress Series, 552:211–222, 2016. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m552p211.pdf.

Crook:2016:ISC

[CD16b] K. A. Crook and G. K. Davoren. Influence of spawning capelin Mallotus villosus on the distribution of green sea urchins Strongylocentrotus droebachiensis on the northeast Newfoundland coast. Marine Ecology Progress Series, 549: 125–135, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/549/m549p125.pdf.

Carvalho:2019:ANB

[CD19] P. C. Carvalho and G. K. Davoren. Associations of nonbreeding shearwater species on the northeastern Newfoundland coast. Marine Ecology Progress Series, 627:1–12, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Cleary:2018:FAK

[CDC18] A. C. Cleary, E. G. Durbin, and M. C. Casas. Feeding by Antarctic krill Euphausia superba in the West Antarctic Peninsula: differences between fjords and open waters. Marine Ecology Progress Series, 595:39–54, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/595/
m595p039.pdf.

Cole:2012:FHA

[CDCS12] R. G. Cole, N. K. Davey, G. D. Carbines, and R. Stewart. Fishhabitat associations in New Zealand: geographical contrasts. Marine Ecology Progress Series, 450:131–145, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/450/
m450p131.pdf.

Cleary:2016:WDS

[CDCZ16] A. C. Cleary, E. G. Durbin, M. C. Casas, and M. Zhou. Winter distribution and size structure of Antarctic krill Euphausia superba populations in-shore along the West Antarctic Peninsula. Marine Ecology Progress Series, 552:115–129, 2016. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m552p115.pdf.

Calo:2016:PDL

[CDD^{+}16] A. Calò, A. Di Franco, G. E. De Benedetto, A. Pennetta, Á. Pérez-Ruzafa, and J. A. García-Charton. Propagule dispersal and larval patch cohesiveness in a Mediterranean coastal fish. Marine Ecology Progress Series, 544:213–224, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/544/m544p213.pdf.

Christensen-Dalsgaard:2018:PWC

[CDMB^{+}18] S. Christensen-Dalsgaard, R. F. May, R. T. Barrett, M. Langset, B. K. Sandercock, and S. H. Lorentsen. Prevailing weather conditions and diet composition affect chick growth and survival in the black-legged kittiwake. Marine Ecology Progress Series, 604:237–249, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m604p237.pdf.

Carr:2012:MEC

[CDMW12] J. A. Carr, P. D’Odorico, K. J. McGlathery, and P. L. Wiberg. Modeling the effects of climate change on eelgrass stability and resilience: future scenarios and leading indicators of collapse. Marine Ecology Progress Series, 448:289–301, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m448p289.pdf.

Cleary:2012:KFS

[CDR12] A. C. Cleary, E. G. Durbin, and T. A. Rynearson. Krill feeding on sediment in the Gulf of Maine (North Atlantic). Marine Ecology Progress Series, 455:157–172, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m455p157.pdf.

Chambault:2016:INB

[CdTK^{+}16] P. Chambault, B. de Thoisy, L. Kelle, R. Berzins, M. Bonola, H. Delvaux, Y. Le Maho, and D. Chevallier. Inter-nesting behavioural adjustments of green turtles to an estuarine habitat in French Guiana. Marine Ecology Progress Series, 555: 235–248, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/555/m555p235.pdf.

Carlson:2017:FSS

[CDWC17] P. M. Carlson, K. Davis, R. R. Warner, and J. E. Caselle. Fine-scale spatial patterns of parrotfish herbivory are shaped by resource availability. Marine Ecology Progress Series, 577:165–176, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/577/m577p165.pdf.

Carvalho:2011:CSI

[CE11] M. C. Carvalho and B. D. Eyre. Carbon stable isotope discrimination during respiration in three seaweed species. Marine Ecology Progress Series, 437:41–49, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/437/
m437p041.pdf.

Cros:2013:HHL

[CE13] L. Cros and M. Estrada. Holo-heterococcolithophore life cycles: ecological implications. Marine Ecology Progress Series, 492:57–68, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/492/m492p057.pdf.

Cameron:2014:ESF

[CE14] C. M. Cameron and P. J. Edmunds. Effects of simulated fish predation on small colonies of massive Porites spp. and Pocillopora meandrina. Marine Ecology Progress Series, 508: 139–148, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m508p139.pdf.

Cunningham:2018:IFL

[CEC^{+}18] J. T. Cunningham, K. H. Elliott, K. Cottenie, S. A. Hatch, and S. R. Jacobs. Individual foraging location, but not dietary, specialization: implications for rhinoceros auklets as samplers of forage fish. Marine Ecology Progress Series, 605:225–240, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m605p225.pdf.

Camp:2019:MLG

[CED^{+}19] E. F. Camp, J. Edmondson, A. Doheny, J. Rumney, A. J. Grima, A. Huete, and D. J. Suggett. Mangrove lagoons of the Great Barrier Reef support coral populations persisting under extreme environmental conditions. Marine Ecology Progress Series, 625:1–14, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Calich:2018:OBH

[CEH18] H. Calich, M. Estevanez, and N. Hammerschlag. Overlap between highly suitable habitats and longline gear management areas reveals vulnerable and protected regions for highly migratory sharks. Marine Ecology Progress Series, 602:183– 195, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/602/m602p183.pdf.

Cormon:2016:NSS

[CEK^{+}16] X. Cormon, B. Ernande, A. Kempf, Y. Vermard, and P. Marchal. North Sea saithe Pollachius virens growth in relation to food availability, density dependence and temperature. Marine Ecology Progress Series, 542:141–151, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/542/
m542p141.pdf.

Celdran:2012:EEB

[CESAM12] D. Celdrán, E. Espinosa, A. Sánchez-Amat, and A. Marín. Effects of epibiotic bacteria on leaf growth and epiphytes of the seagrass Posidonia oceanica. Marine Ecology Progress Series, 456:21–27, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/456/m456p021.pdf.

Comeau:2014:DPO

[CESC14] S. Comeau, P. J. Edmunds, N. B. Spindel, and R. C. Carpenter. Diel pCO 2 oscillations modulate the response of the coral Acropora hyacinthus to ocean acidification. Marine Ecology Progress Series, 501:99–111, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m501p099.pdf.

Cook:2015:FCN

[CEW15] P. L. M. Cook, V. Evrard, and R. J. Woodland. Factors controlling nitrogen fixation in temperate seagrass beds. Marine Ecology Progress Series, 525:41–51, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/525/
m525p041.pdf.

Cesar:2012:BDA

[CF12] C. P. Cesar and C. L. J. Frid. Benthic disturbance affects intertidal food web dynamics: implications for investigations of ecosystem functioning. Marine Ecology Progress Series, 466:35–41, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m466p035.pdf.

Connell:2014:WSI

[CFA14] S. D. Connell, M. S. Foster, and L. Airoldi. As we see it: What are algal turfs? Towards a better description of turfs. Marine Ecology Progress Series, 495:299–307, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m495p299.pdf.

Carreja:2016:JAE

[CFA16] B. Carreja, M. Fernández, and S. Agustí. Joint additive effects of temperature and UVB radiation on zoeae of the crab Taliepus dentatus. Marine Ecology Progress Series, 550: 135–145, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/550/m550p135.pdf.

Crum:2014:MDC

[CFBG14] K. P. Crum, H. L. Fuchs, P. A. X. Bologna, and J. J. Gaynor. Model-to-data comparisons reveal influence of jellyfish interactions on plankton community dynamics. Marine Ecology Progress Series, 517:105–119, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/517/m517p105.pdf.

Cibic:2012:BEF

[CFC^{+}12] T. Cibic, A. Franzo, M. Celussi, C. Fabbro, and P. Del Negro. Benthic ecosystem functioning in hydrocarbon and heavy-metal contaminated sediments of an Adriatic lagoon. Marine Ecology Progress Series, 458:69–87, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/458/
m458p069.pdf.

Canepuccia:2018:DFB

[CFF^{+}18] A. D. Canepuccia, J. L. Farina, E. Fanjul, F. Botto, J. Pascual, and O. Iribarne. Driving forces behind latitudinal variations in plant-herbivore interactions in SW Atlantic salt marshes. Marine Ecology Progress Series, 603:93–103, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/603/
m603p093.pdf.

Crowe:2011:IEL

[CFH11] T. P. Crowe, N. J. Frost, and S. J. Hawkins. Interactive effects of losing key grazers and ecosystem engineers vary with environmental context. Marine Ecology Progress Series, 430:223–234, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m430p223.pdf.

Cardozo-Ferreira:2018:ILF

[CFMFFJ18] G. C. Cardozo-Ferreira, R. M. Macieira, R. B. Francini-Filho, and J. C. Joyeux. Inferring labrid functional roles through morphological and ecological traits. Marine Ecology Progress Series, 588:135–145, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/588/m588p135.pdf.

Cornwell:2016:CGP

[CFMN16] B. H. Cornwell, J. L. Fisher, S. G. Morgan, and J. E. Neigel. Chaotic genetic patchiness without sweepstakes reproduction in the shore crab Hemigrapsus oregonensis. Marine Ecology Progress Series, 548:139–152, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m548p139.pdf.

Cocito:2013:NAF

[CFPC^{+}13] S. Cocito, C. Ferrier-Pagès, R. Cupido, C. Rottier, W. Meier-Augenstein, H. Kemp, S. Reynaud, and A. Peirano. Nutrient acquisition in four Mediterranean gorgonian species. Marine Ecology Progress Series, 473:179–188, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/473/
m473p179.pdf.

Cimino:2014:LSC

[CFPF^{+}14] M. A. Cimino, W. R. Fraser, D. L. Patterson-Fraser, V. S. Saba, and M. J. Oliver. Large-scale climate and local weather drive interannual variability in Adélie penguin chick fledging mass. Marine Ecology Progress Series, 513:253–268, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m513p253.pdf.

Ciannelli:2013:TCE

[CFSM^{+}13] L. Ciannelli, J. A. D. Fisher, M. Skern-Mauritzen, M. E. Hunsicker, M. Hidalgo, K. T. Frank, and K. M. Bailey. Theory, consequences and evidence of eroding population spatial structure in harvested marine fishes: a review. Marine Ecology Progress Series, 480:227–243, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m480p227.pdf.

Chan:2010:TDM

[CG10] K. Y. K. Chan and D. Grünbaum. Temperature and diet modified swimming behaviors of larval sand dollar. Marine Ecology Progress Series, 415:49–59, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/415/m415p049.pdf.

Clay:2011:SPC

[CG11] T. W. Clay and D. Grünbaum. Swimming performance as a constraint on larval morphology in plutei. Marine Ecology Progress Series, 423:185–196, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/423/m423p185.pdf.

Clark:2016:DFC

[CG16] H. R. Clark and C. J. Gobler. Diurnal fluctuations in CO_{2} and dissolved oxygen concentrations do not provide a refuge from hypoxia and acidification for early-life-stage bivalves. Marine Ecology Progress Series, 558:1–14, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Chiaverano:2017:MPA

[CG17] L. M. Chiaverano and W. M. Graham. Morphological plasticity in Aurelia polyps, with subsequent effects on asexual fecundity and morphology of young medusae. Marine Ecology Progress Series, 582:79–92, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/582/m582p079.pdf.

Cortes:2018:SAB

[CGBGS18] V. Cortés, S. García-Barcelona, and J. González-Solís. Sexand age-biased mortality of three shearwater species in longline fisheries of the Mediterranean. Marine Ecology Progress Series, 588:229–241, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/588/m588p229.pdf.

Chiaverano:2015:PAB

[CGC15] L. M. Chiaverano, W. M. Graham, and J. H. Costello. Parasites alter behavior, reproductive output, and growth patterns of Aurelia medusae in a marine lake. Marine Ecology Progress Series, 540:87–98, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/540/m540p087.pdf.

Celentano:2010:EME

[CGD10] E. Celentano, N. L. Gutiérrez, and O. Defeo. Effects of morphodynamic and estuarine gradients on the demography and distribution of a sandy beach mole crab: implications for source-sink habitat dynamics. Marine Ecology Progress Series, 398:193–205, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/398/m398p193.pdf.

Compaire:2019:MMF

[CGEGC^{+}19] J. C. Compaire, J. Gómez-Enri, C. Gómez-Cama, P. Casademont, O. V. Sáez, F. Pastoriza-Martin, C. Díaz-Gil, R. Cabrera, A. Cosin, and M. C. Soriguer. Micro- and macroscale factors affecting fish assemblage structure in the rocky intertidal zone. Marine Ecology Progress Series, 610:175– 189, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/610/m610p175.pdf.

Carter:2010:PVM

[CGG10a] M. C. Carter, D. P. Gordon, and J. P. A. Gardner. Polymorphism and variation in modular animals: morphometric and density analyses of bryozoan avicularia. Marine Ecology Progress Series, 399:117–130, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/399/m399p117.pdf.

Cuveliers:2010:MVJ

[CGG^{+}10b] E. L. Cuveliers, A. J. Geffen, J. Guelinckx, J. A. M. Raeymaekers, J. Skadal, F. A. M. Volckaert, and G. E. Maes. Microchemical variation in juvenile Solea solea otoliths as a powerful tool for studying connectivity in the North Sea. Marine Ecology Progress Series, 401:211–220, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/401/m401p211.pdf.

Chenoweth:2011:TIH

[CGH11] E. M. Chenoweth, C. M. Gabriele, and D. F. Hill. Tidal influences on humpback whale habitat selection near headlands. Marine Ecology Progress Series, 423:279–289, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/423/
m423p279.pdf.

Chung:2012:ETM

[CGH12] C. C. Chung, G. C. Gong, and C. C. Hung. Effect of Typhoon Morakot on microphytoplankton population dynamics in the subtropical Northwest Pacific. Marine Ecology Progress Series, 448:39–49, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m448p039.pdf.

Commito:2016:SAM

[CGH^{+}16] J. A. Commito, N. J. Gownaris, D. E. Haulsee, S. E. Coleman, and B. F. Beal. Separation anxiety: mussels self-organize into similar power-law clusters regardless of predation threat cues. Marine Ecology Progress Series, 547:107–119, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/547/
m547p107.pdf.

Castellanos-Galindo:2013:MFA

[CGK13a] G. A. Castellanos-Galindo and U. Krumme. Mangrove fish assemblages from data-sparse regions and the measurement of ecological equivalence: Comment on Sheaves (2012). Marine Ecology Progress Series, 474:299–302, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/474/

m474p299.pdf. See [She12] and reply [She13].

Castellanos-Galindo:2013:TDS

[CGK13b] G. A. Castellanos-Galindo and U. Krumme. Tidal, diel and seasonal effects on intertidal mangrove fish in a high-rainfall area of the Tropical Eastern Pacific. Marine Ecology Progress Series, 494:249–265, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/494/m494p249.pdf.

Castellanos-Galindo:2010:TIF

[CGKW10] G. A. Castellanos-Galindo, U. Krumme, and T. J. Willis. Tidal influences on fish distributions on tropical eastern Pacific rocky shores (Colombia). Marine Ecology Progress Series, 416:241–254, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/416/m416p241.pdf.

Cote:2013:DRI

[CGM^{+}13] I. M. Côté, S. J. Green, J. A. Morris, Jr., J. L. Akins, and D. Steinke. Diet richness of invasive Indo–Pacific lionfish revealed by DNA barcoding. Marine Ecology Progress Series, 472:249–256, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/472/m472p249.pdf.

Correia:2019:TIB

[CGM^{+}19] E. Correia, J. P. Granadeiro, V. A. Mata, A. Regalla, and P. Catry. Trophic interactions between migratory seabirds, predatory fishes and small pelagics in coastal West Africa. Marine Ecology Progress Series, 622:177–189, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/622/
m622p177.pdf.

Caro:2011:CBN

[CGOC11] A. U. Caro, R. Guiñez, V. Ortiz, and J. C. Castilla. Competition between a native mussel and a non-indigenous invader for primary space on intertidal rocky shores in Chile. Marine Ecology Progress Series, 428:177–185, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/428/
m428p177.pdf.

Compte:2010:TPE

[CGQB10] J. Compte, S. Gascón, X. D. Quintana, and D. Boix. Toppredator effects of jellyfish Odessia maeotica in Mediterranean salt marshes. Marine Ecology Progress Series, 402: 147–159, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/402/m402p147.pdf.

Cardich:2015:CBF

[CGR^{+}15] J. Cardich, D. Gutiérrez, D. Romero, A. Pérez, L. Quipúzcoa, R. Marquina, W. Yupanqui, J. Solís, W. Carhuapoma, A. Sifeddine, and A. Rathburn. Calcareous benthic Foraminifera from the upper central Peruvian margin: control of the assemblage by pore water redox and sedimentary organic matter. Marine Ecology Progress Series, 535: 63–87, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/535/m535p063.pdf.

Carvalho:2019:TSI

[CGS^{+}19] A. Carvalho, I. Gomes, S. E. Swearer, H. Queiroga, and L. G. Peteiro. Temperature and salinity influence on element incorporation into Mytilus galloprovincialis larvae shells: discerning physiological from environmental control. Marine Ecology Progress Series, 626:83–96, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/626/m626p083.pdf.

Cope:2012:TSS

[CH12] J. M. Cope and M. A. Haltuch. Temporal and spatial summer groundfish assemblages in trawlable habitat off the west coast of the USA, 1977 to 2009. Marine Ecology Progress Series, 451:187–200, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/451/m451p187.pdf.

Clements:2015:MAB

[CH15] J. C. Clements and H. L. Hunt. Marine animal behaviour in a high CO_{2} ocean. Marine Ecology Progress Series, 536: 259–279, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/536/m536p259.pdf.

Chapman:2013:CRH

[Cha13] M. G. Chapman. Constructing replacement habitat for specialist and generalist molluscs—the effect of patch size. Marine Ecology Progress Series, 473:201–214, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/473/
m473p201.pdf.

Clemente:2011:CDE

[CHB11] S. Clemente, J. C. Hernández, and A. Brito. Contextdependent effects of marine protected areas on predatory interactions. Marine Ecology Progress Series, 437:119–133, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2011/437/m437p119.pdf.

Coates:2013:MHR

[CHB^{+}13] J. H. Coates, K. A. Hovel, J. L. Butler, A. P. Klimley, and S. G. Morgan. Movement and home range of pink abalone Haliotis corrugata: implications for restoration and population recovery. Marine Ecology Progress Series, 486:189–201, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m486p189.pdf.

Chalifour:2019:DAS

[CHDH^{+}19] B. Chalifour, J. R. H. Hoogveld, M. Derksen-Hooijberg, K. L. Harris, J. M. Urueña, W. G. Sawyer, T. van der Heide, and C. Angelini. Drought alters the spatial distribution, grazing patterns, and radula morphology of a fungal-farming salt marsh snail. Marine Ecology Progress Series, 620:1–13, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m620p001.pdf.

Clarke:2017:IIH

[CHE^{+}17] L. J. Clarke, K. M. Hughes, L. S. Esteves, R. J. H. Herbert, and R. A. Stillman. Intertidal invertebrate harvesting: a meta-analysis of impacts and recovery in an important waterbird prey resource. Marine Ecology Progress Series, 584:229–244, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m584p229.pdf.

Condie:2012:MEC

[CHFS12] S. A. Condie, D. Hayes, E. A. Fulton, and M. Savina. Modelling ecological change over half a century in a subtropical estuary: impacts of climate change, land-use, urbanization and freshwater extraction. Marine Ecology Progress Series, 457:43–66, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m457p043.pdf.

Corry:2018:FRS

[CHG^{+}18] M. Corry, D. Harasti, T. Gaston, D. Mazumder, T. Cresswell, and N. Moltschaniwskyj. Functional role of the soft coral Dendronephthya australis in the benthic food web of temperate estuaries. Marine Ecology Progress Series, 593: 61–72, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/593/m593p061.pdf.

Chiswell:2011:ACS

[Chi11] S. M. Chiswell. Annual cycles and spring blooms in phytoplankton: don’t abandon sverdrup completely. Marine Ecology Progress Series, 443:39–50, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m443p039.pdf.

Czerny:2016:DPR

[CHLR16] J. M. S. Czerny, H. Hauss, C. R. Löscher, and U. Riebesell. Dissolved N: P ratio changes in the eastern tropical North Atlantic: effect on phytoplankton growth and community structure. Marine Ecology Progress Series, 545:49–62, 2016. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m545p049.pdf.

Chapman:2011:MTF

[CHP^{+}11] E. W. Chapman, E. E. Hofmann, D. L. Patterson, C. A. Ribic, and W. R. Fraser. Marine and terrestrial factors affecting Adélie penguin Pygoscelis adeliae chick growth and recruitment off the western Antarctic Peninsula. Marine Ecology Progress Series, 436:273–289, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/436/m436p273.pdf.

Christiansen:2017:OPN

[Chr17] J. S. Christiansen. Opinion piece: No future for Euro– Arctic ocean fishes? Marine Ecology Progress Series, 575: 217–227, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/575/m575p217.pdf.

Clemente:2010:IKP

[CHRB10] S. Clemente, J. C. Hernández, A. Rodríguez, and A. Brito. Identifying keystone predators and the importance of preserving functional diversity in sublittoral rocky-bottom areas. Marine Ecology Progress Series, 413:55–67, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/413/
m413p055.pdf.

Chin:2016:POR

[CHST16] A. Chin, M. R. Heupel, C. A. Simpfendorfer, and A. J. Tobin. Population organisation in reef sharks: new variations in coastal habitat use by mobile marine predators. Marine Ecology Progress Series, 544:197–211, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/544/
m544p197.pdf.

Chen:2011:BDD

[CIF11] B. Chen, A. J. Irwin, and Z. V. Finkel. Biogeographic distribution of diversity and size-structure of organic-walled dinoflagellate cysts. Marine Ecology Progress Series, 425: 35–45, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/425/m425p035.pdf.

Calbet:2013:AFF

[CIM^{+}13] A. Calbet, S. Isari, R. A. Martínez, E. Saiz, S. Garrido, J. Peters, R. M. Borrat, and M. Alcaraz. Adaptations to feast and famine in different strains of the marine heterotrophic dinoflagellates Gyrodinium dominans and Oxyrrhis marina. Marine Ecology Progress Series, 483:67–84, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/483/
m483p067.pdf.

Caut:2013:DTS

[CJM^{+}13] S. Caut, M. J. Jowers, L. Michel, G. Lepoint, and A. T. Fisk. Diet- and tissue-specific incorporation of isotopes in the shark Scyliorhinus stellaris, a North Sea mesopredator.

Marine Ecology Progress Series, 492:185–198, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/492/
m492p185.pdf.

Constantine:2012:AHW

[CJS^{+}12] R. Constantine, J. A. Jackson, D. Steel, C. S. Baker, L. Brooks, D. Burns, P. Clapham, N. Hauser, B. Madon, D. Mattila, M. Oremus, M. Poole, J. Robbins, K. Thompson, and C. Garrigue. Abundance of humpback whales in Oceania using photo-identification and microsatellite genotyping. Marine Ecology Progress Series, 453:249–261, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/453/
m453p249.pdf.

Concepcion:2010:RNR

[CKC^{+}10] G. T. Concepcion, S. E. Kahng, M. W. Crepeau, E. C. Franklin, S. L. Coles, and R. J. Toonen. Resolving natural ranges and marine invasions in a globally distributed octocoral (genus Carijoa). Marine Ecology Progress Series, 401:113–127, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m401p113.pdf.

Clausen:2014:SEB

[CKJOM14] K. K. Clausen, D. Krause-Jensen, B. Olesen, and N. Marbà. Seasonality of eelgrass biomass across gradients in temperature and latitude. Marine Ecology Progress Series, 506: 71–85, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/506/m506p071.pdf.

Chow:2010:JEA

[CKK^{+}10] S. Chow, H. Kurogi, S. Katayama, D. Ambe, M. Okazaki, T. Watanabe, T. Ichikawa, M. Kodama, J. Aoyama, A. Shinoda, S. Watanabe, K. Tsukamoto, S. Miyazaki, S. Kimura, Y. Yamada, K. Nomura, H. Tanaka, Y. Kazeto, K. Hata, T. Handa, A. Tawa, and N. Mochioka. Japanese eel Anguilla japonica do not assimilate nutrition during the oceanic spawning migration: evidence from stable isotope analysis. Marine Ecology Progress Series, 402:233–238, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/402/
m402p233.pdf.

Cambie:2016:SIS

[CKM^{+}16] G. Cambiè, M. J. Kaiser, A. L. Marriott, J. Fox, G. Lambert, J. G. Hiddink, T. Overy, S. A. Bennet, M. J. Leng, and I. D. McCarthy. Stable isotope signatures reveal smallscale spatial separation in populations of European sea bass. Marine Ecology Progress Series, 546:213–223, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/546/
m546p213.pdf.

Compton:2011:SBT

[CKR^{+}11] T. J. Compton, C. Kraan, M. J. A. Rijkenberg, T. A. Troost, D. I. Rogers, J. Leyrer, G. Pearson, and T. Piersma. Sampling bivalves on tidal flats: possibility of missing rare species, versus smoothing of environmental variation. Marine Ecology Progress Series, 440:285–288, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/440/m440p285.pdf.

Collins:2012:BSM

[CKV12] P. C. Collins, R. Kennedy, and C. L. Van Dover. A biological survey method applied to seafloor massive sulphides (SMS) with contagiously distributed hydrothermal-vent fauna. Marine Ecology Progress Series, 452:89–107, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/452/
m452p089.pdf.

Chen:2010:TLB

[CL10a] B. Chen and H. Liu. Trophic linkages between grazers and ultraplankton within the microbial food web in subtropical coastal waters. Marine Ecology Progress Series, 407: 43–53, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/407/m407p043.pdf.

Chu:2010:HRM

[CL10b] J. W. F. Chu and S. P. Leys. High resolution mapping of community structure in three glass sponge reefs (Porifera, Hexactinellida). Marine Ecology Progress Series, 417:97– 113, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/417/m417p097.pdf.

Copeman:2010:EEF

[CL10c] L. A. Copeman and B. J. Laurel. Experimental evidence of fatty acid limited growth and survival in Pacific cod larvae. Marine Ecology Progress Series, 412:259–272, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/412/
m412p259.pdf.

Chakraborty:2015:PCS

[CL15a] S. Chakraborty and S. E. Lohrenz. Phytoplankton community structure in the river-influenced continental margin of the northern Gulf of Mexico. Marine Ecology Progress Series, 521:31–47, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m521p031.pdf.

Chen:2015:RAH

[CL15b] B. Chen and K. Liu. Responses of autotrophic and heterotrophic rates of plankton from a subtropical coastal site to short-term temperature modulations. Marine Ecology Progress Series, 527:59–71, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/527/m527p059.pdf.

Colden:2015:LSE

[CL15c] A. M. Colden and R. N. Lipcius. Lethal and sublethal effects of sediment burial on the eastern oyster Crassostrea virginica. Marine Ecology Progress Series, 527:105–117, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m527p105.pdf.

Coelho:2016:LBS

[CL16] M. A. G. Coelho and H. R. Lasker. Larval behavior and settlement dynamics of a ubiquitous Caribbean octocoral and its implications for dispersal. Marine Ecology Progress Series, 561:109–121, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/561/m561p109.pdf.

Chen:2017:DTJ

[CL17] L. Chen and C. Li. Different tolerances of jellyfish ephyrae (Aurelia sp. 1) and fish larvae (Paralichthys olivaceus) to nutrient limitations. Marine Ecology Progress Series, 569:1–13, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Cameron:2014:UCH

[CLB^{+}14] M. J. Cameron, V. Lucieer, N. S. Barrett, C. R. Johnson, and G. J. Edgar. Understanding community-habitat associations of temperate reef fishes using fine-resolution bathymetric measures of physical structure. Marine Ecology Progress Series, 506:213–229, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/506/m506p213.pdf.

Chin:2018:BDC

[CLB^{+}18] A. T. M. Chin, J. Linke, M. Boudreau, M. H. Thériault, S. C. Courtenay, R. Cormier, and M. J. Fortin. Beta diversity changes in estuarine fish communities due to environmental change. Marine Ecology Progress Series, 603:161– 173, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/603/m603p161.pdf.

Carroll:2019:IDG

[CLB19] J. Carroll and S. K. Lowerre-Barbieri. Interactions of dimorphic growth, reproductive behavior, and a size-regulated fishery: a case study using spotted seatrout Cynoscion nebulosus. Marine Ecology Progress Series, 608:233–245, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/608/
m608p233.pdf.

Chen:2012:EDE

[CLC12a] M. Chen, H. Liu, and B. Chen. Effects of dietary essential fatty acids on reproduction rates of a subtropical calanoid copepod, Acartia erythraea. Marine Ecology Progress Series, 455:95–110, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/455/m455p095.pdf.

Chivers:2012:FTT

[CLC^{+}12b] L. S. Chivers, M. G. Lundy, K. Colhoun, S. F. Newton, J. D. R. Houghton, and N. Reid. Foraging trip time-activity budgets and reproductive success in the black-legged kittiwake. Marine Ecology Progress Series, 456:269–277, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/456/
m456p269.pdf.

Callier:2013:SBA

[CLD^{+}13] M. D. Callier, S. Lefebvre, M. K. Dunagan, M. P. Bataille, J. Coughlan, and T. P. Crowe. Shift in benthic assemblages and organisms’ diet at salmon farms: community structure and stable isotope analyses. Marine Ecology Progress Series, 483:153–167, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/483/m483p153.pdf.

Carson:2013:TSI

[CLDC^{+}13] H. S. Carson, P. C. López-Duarte, G. S. Cook, F. J. Fodrie, B. J. Becker, C. DiBacco, and L. A. Levin. Temporal, spatial, and interspecific variation in geochemical signatures within fish otoliths, bivalve larval shells, and crustacean larvae. Marine Ecology Progress Series, 473:133–148, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/473/
m473p133.pdf.

Camillini:2019:BPS

[CLG19a] N. Camillini, M. Larsen, and R. N. Glud. Behavioural patterns of the soft-shell clam Mya arenaria: implications for benthic oxygen and nitrogen dynamics. Marine Ecology Progress Series, 622:103–119, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m622p103.pdf.

Cheney:2019:BPS

[CLG^{+}19b] D. Cheney, J. M. Logan, K. Gardner, E. Sly, B. Wysor, and S. Greenwood. Bioaccumulation of PCBs by a seaweed bloom (Ulva rigida) and transfer to higher trophic levels in an estuarine food web. Marine Ecology Progress Series, 611:75–93, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/611/m611p075.pdf.

Christiaen:2016:FIC

[CLGC16] B. Christiaen, J. C. Lehrter, J. Goff, and J. Cebrian. Functional implications of changes in seagrass species composition in two shallow coastal lagoons. Marine Ecology Progress Series, 557:111–121, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m557p111.pdf.

Chaabani:2019:ENM

[CLGCA^{+}19] S. Chaabani, P. J. López-González, P. Casado-Amezua, H. Pehlke, L. Weber, I. Martínez-Baraldés, and K. Jerosch. Ecological niche modelling of cold-water corals in the Southern Ocean (N Antarctic), present distribution and future projections due to temperature changes. Marine Ecology Progress Series, 628:73–93, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/628/m628p073.pdf.

Carlisle:2015:RHU

[CLH^{+}15] A. B. Carlisle, S. Y. Litvin, E. L. Hazen, D. J. Madigan, K. J. Goldman, R. N. Lea, and B. A. Block. Reconstructing habitat use by juvenile salmon sharks links upwelling to strandings in the California Current. Marine Ecology Progress Series, 525:217–228, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/525/m525p217.pdf.

Chen:2014:TEG

[CLHW14] B. Chen, H. Liu, B. Huang, and J. Wang. Temperature effects on the growth rate of marine picoplankton. Marine Ecology Progress Series, 505:37–47, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/505/m505p037.pdf.

Chan:2012:NAI

[CLJF12] A. Y. Chan, K. Lubarsky, K. N. Judy, and P. Fong. Nutrient addition increases consumption rates of tropical algae with different initial palatabilities. Marine Ecology Progress Series, 465:25–31, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/465/m465p025.pdf.

Chen:2013:EMF

[CLL13] M. Chen, H. Liu, and H. Li. Effect of mesozooplankton feeding selectivity on the dynamics of algae in the presence of intermediate grazers — a laboratory simulation. Marine Ecology Progress Series, 486:47–58, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/486/m486p047.pdf.

Casas:2015:ROR

[CLL15] S. M. Casas, J. La Peyre, and M. K. La Peyre. Restoration of oyster reefs in an estuarine lake: population dynamics and shell accretion. Marine Ecology Progress Series, 524:171–184, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/524/m524p171.pdf.

Colden:2017:RHD

[CLL17] A. M. Colden, R. J. Latour, and R. N. Lipcius. Reef height drives threshold dynamics of restored oyster reefs. Marine Ecology Progress Series, 582:1–13, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Calves:2013:GSE

[CLM^{+}13] I. Calvès, E. Lavergne, A. L. Meistertzheim, G. Charrier, H. Cabral, B. Guinand, L. Quiniou, and J. Laroche. Genetic structure of European flounder Platichthys flesus: effects of both the southern limit of the species’ range and chemical stress. Marine Ecology Progress Series, 472:257– 273, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/472/m472p257.pdf.

Chua:2013:NFR

[CLMB13a] C. M. Chua, W. Leggat, A. Moya, and A. H. Baird. Nearfuture reductions in pH will have no consistent ecological effects on the early life-history stages of reef corals. Marine Ecology Progress Series, 486:143–151, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/486/
m486p143.pdf.

Chua:2013:TAE

[CLMB13b] C. M. Chua, W. Leggat, A. Moya, and A. H. Baird. Temperature affects the early life history stages of corals more than near future ocean acidification. Marine Ecology Progress Series, 475:85–92, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/475/m475p085.pdf.

Clemente:2014:SUD

[CLMM^{+}14] S. Clemente, J. Lorenzo-Morales, J. C. Mendoza, C. López, C. Sangil, F. Alves, M. Kaufmann, and J. C. Hernández. Sea urchin Diadema africanum mass mortality in the subtropical eastern Atlantic: role of waterborne bacteria in a warming ocean. Marine Ecology Progress Series, 506:1–14, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Cummings:2013:TSC

[CLN^{+}13] D. O. Cummings, R. W. Lee, S. D. Nodder, S. J. Simpson, and S. P. Holmes. Trophic status and condition of Hyalinoecia longibranchiata from two regions of contrasting oceanic productivity. Marine Ecology Progress Series, 477: 147–159, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/477/m477p147.pdf.

Ceccarelli:2018:AOH

[CLP18] D. M. Ceccarelli, M. Logan, and S. W. Purcell. Analysis of optimal habitat for captive release of the sea cucumber Holothuria scabra. Marine Ecology Progress Series, 588: 85–100, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m588p085.pdf.

Condon:2014:TSJ

[CLPiU14] Robert H. Condon, Cathy H. Lucas, Kylie A. Pitt, and Shin ichi Uye. Theme section: Jellyfish blooms and ecological interaction. Marine Ecology Progress Series, 510:107, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m510p107.pdf.

Condon:2014:JBE

[CLPU14] R. H. Condon, C. H. Lucas, K. A. Pitt, and S. I. Uye. Jellyfish blooms and ecological interactions. Marine Ecology Progress Series, 510:109–110, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m510p109.pdf.

Crear:2016:SSM

[CLS^{+}16] D. P. Crear, D. D. Lawson, J. A. Seminoff, T. Eguchi, R. A. LeRoux, and C. G. Lowe. Seasonal shifts in the movement and distribution of green sea turtles Chelonia mydas in response to anthropogenically altered water temperatures. Marine Ecology Progress Series, 548:219–232, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/548/m548p219.pdf.

Campanya-Llovet:2019:IPQ

[CLS19] N. Campanyà-Llovet and P. V. R. Snelgrove. Influence of phytodetrital quality on macroinfaunal community structure and epifaunal response. Marine Ecology Progress Series, 608: 45–60, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m608p045.pdf.

Copeman:2017:TIL

[CLSS17] L. A. Copeman, B. J. Laurel, M. Spencer, and A. Sremba. Temperature impacts on lipid allocation among juvenile gadid species at the Pacific Arctic–Boreal interface: an experimental laboratory approach. Marine Ecology Progress Series, 566:183–198, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/566/m566p183.pdf.

Cermeno:2011:CDT

[CLW^{+}11] P. Cermeño, J. B. Lee, K. Wyman, O. Schofield, and P. G. Falkowski. Competitive dynamics in two species of marine phytoplankton under non-equilibrium conditions. Marine Ecology Progress Series, 429:19–28, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/429/
m429p019.pdf.

Cote:2010:PRI

[CM10] I. M. Côté and A. Maljković. Predation rates of Indo–Pacific lionfish on Bahamian coral reefs. Marine Ecology Progress Series, 404:219–225, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/404/m404p219.pdf.

Cole:2011:BSS

[CM11] V. J. Cole and C. D. McQuaid. Broad-scale spatial factors outweigh the influence of habitat structure on the fauna associated with a bioengineer. Marine Ecology Progress Series, 442:101–109, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/442/m442p101.pdf.

Cole:2012:NFR

[CM12] L. W. Cole and K. J. McGlathery. Nitrogen fixation in restored eelgrass meadows. Marine Ecology Progress Series, 448:235–246, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m448p235.pdf.

Casale:2014:FLY

[CM14] P. Casale and P. Mariani. The first ‘lost year’ of Mediterranean sea turtles: dispersal patterns indicate subregional management units for conservation. Marine Ecology Progress Series, 498:263–274, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/498/m498p263.pdf.

Clarkin:2012:ECI

[CMAJ12] E. Clarkin, C. A. Maggs, A. L. Allcock, and M. P. Johnson. Environment, not characteristics of individual algal rafts, affects composition of rafting invertebrate assemblages in Irish coastal waters. Marine Ecology Progress Series, 470: 31–40, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/470/m470p031.pdf.

Connan:2014:CSC

[CMB^{+}14] M. Connan, C. D. McQuaid, B. T. Bonnevie, M. J. Smale, and Y. Cherel. Combined stomach content, lipid and stable isotope analyses reveal spatial and trophic partitioning among three sympatric albatrosses from the Southern Ocean. Marine Ecology Progress Series, 497:259–272, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m497p259.pdf.

Castilla:2010:ERS

[CMC10a] J. C. Castilla, P. H. Manríquez, and A. Camaño. Effects of rocky shore coseismic uplift and the 2010 Chilean megaearthquake on intertidal biomarker species. Marine Ecology Progress Series, 418:17–23, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/418/m418p017.pdf.

Chollett:2010:UAD

[CMC10b] I. Chollett, P. J. Mumby, and J. Cortés. Upwelling areas do not guarantee refuge for coral reefs in a warming ocean. Marine Ecology Progress Series, 416:47–56, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/416/
m416p047.pdf.

Connelly:2015:SCQ

[CMC^{+}15] T. L. Connelly, J. W. McClelland, B. C. Crump, C. T. E. Kellogg, and K. H. Dunton. Seasonal changes in quantity and composition of suspended particulate organic matter in lagoons of the Alaskan Beaufort Sea. Marine Ecology Progress Series, 527:31–45, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/527/m527p031.pdf.

Carneiro:2016:CMS

[CMC^{+}16] A. P. B. Carneiro, A. Manica, T. A. Clay, J. R. D. Silk, M. King, and R. A. Phillips. Consistency in migration strategies and habitat preferences of brown skuas over two winters, a decade apart. Marine Ecology Progress Series, 553: 267–281, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/553/m553p267.pdf.

Corell:2012:DDL

[CME^{+}12] H. Corell, P. O. Moksnes, A. Engqvist, K. Döös, and P. R. Jonsson. Depth distribution of larvae critically affects their dispersal and the efficiency of marine protected areas. Marine Ecology Progress Series, 467:29–46, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m467p029.pdf.

Cannicci:2019:RLC

[CMF^{+}19] S. Cannicci, B. Mostert, S. Fratini, C. D. McQuaid, and F. Porri. Recruitment limitation and competent settlement of sesarmid crab larvae within East African mangrove forests. Marine Ecology Progress Series, 626:123–133, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/626/
m626p123.pdf.

Cowburn:2019:ECB

[CMGS19] B. Cowburn, C. Moritz, G. Grimsditch, and J. L. Solandt. Evidence of coral bleaching avoidance, resistance and recovery in the Maldives during the 2016 mass-bleaching event. Marine Ecology Progress Series, 626:53–67, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/626/
m626p053.pdf.

Cure:2014:HPN

[CMH14] K. Cure, J. L. McIlwain, and M. A. Hixon. Habitat plasticity in native Pacific red lionfish Pterois volitans facilitates successful invasion of the Atlantic. Marine Ecology Progress Series, 506:243–253, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/506/m506p243.pdf.

Clausius:2017:ECV

[CMHH17] E. Clausius, C. R. McMahon, R. Harcourt, and M. A. Hindell. Effect of climate variability on weaning mass in a declining population of southern elephant seals Mirounga leonina. Marine Ecology Progress Series, 568:249–260, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/568/
m568p249.pdf.

Cardona:2015:RSP

[CMIMGS15] L. Cardona, L. Martínez-Iñigo, R. Mateo, and J. González-Solís. The role of sardine as prey for pelagic predators in the western Mediterranean Sea assessed using stable isotopes and fatty acids. Marine Ecology Progress Series, 531:1–14, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Cathey:2012:MJM

[CMK12] A. M. Cathey, N. R. Miller, and D. G. Kimmel. Microchemistry of juvenile Mercenaria mercenaria shell: implications for modeling larval dispersal. Marine Ecology Progress Series, 465:155–168, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/465/m465p155.pdf.

Cottin:2014:CAL

[CMK^{+}14] M. Cottin, A. J. J. MacIntosh, A. Kato, A. Takahashi, M. Debin, T. Raclot, and Y. Ropert-Coudert. Corticosterone administration leads to a transient alteration of foraging behaviour and complexity in a diving seabird. Marine Ecology Progress Series, 496:249–262, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m496p249.pdf.

Cossu:2012:GSH

[CML^{+}12] P. Cossu, F. Maltagliati, T. Lai, M. Casu, M. Curini-Galletti, and A. Castelli. Genetic structure of Hediste diversicolor (Polychaeta, Nereididae) from the northwestern Mediterranean as revealed by DNA inter-simple sequence repeat (ISSR) markers. Marine Ecology Progress Series, 452: 171–178, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/452/m452p171.pdf.

Compton:2012:OHA

[CMLC12] T. J. Compton, M. A. Morrison, J. R. Leathwick, and G. D. Carbines. Ontogenetic habitat associations of a demersal fish species, Pagrus auratus, identified using boosted regression trees. Marine Ecology Progress Series, 462:219– 230, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/462/m462p219.pdf.

Connell:2018:PBD

[CMM^{+}18] P. E. Connell, C. Michel, G. Meisterhans, K. R. Arrigo, and D. A. Caron. Phytoplankton and bacterial dynamics on the Chukchi Sea Shelf during the spring-summer transition. Marine Ecology Progress Series, 602:49–62, 2018. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m602p049.pdf.

Christofoletti:2010:FOC

[CMO^{+}10] R. A. Christofoletti, V. A. Murakami, D. N. Oliveira, R. E. Barreto, and A. A. V. Flores. Foraging by the omnivorous crab Pachygrapsus transversus affects the structure of assemblages on sub-tropical rocky shores. Marine Ecology Progress Series, 420:125–134, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/420/m420p125.pdf.

Canepuccia:2011:SBM

[CMP^{+}11] A. D. Canepuccia, D. Montemayor, J. Pascual, J. L. Farina, and O. O. Iribarne. A stem-boring moth drives detritus production in SW Atlantic marshes. Marine Ecology Progress Series, 442:1–9, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Chaparro:2013:PSR

[CMP^{+}13] O. R. Chaparro, J. A. Montory, S. V. Pereda, R. J. Thompson, G. Rivera, S. J. A. Osores, and C. J. Segura. Particle selection and regulation of particle uptake by the slipper limpet Crepipatella fecunda. Marine Ecology Progress Series, 474:167–177, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/474/m474p167.pdf.

Caiger:2012:CLI

[CMR12] P. E. Caiger, J. C. Montgomery, and C. A. Radford. Chronic low-intensity noise exposure affects the hearing thresholds of juvenile snapper. Marine Ecology Progress Series, 466:225– 232, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/466/m466p225.pdf.

Certain:2011:ICB

[CMV^{+}11] G. Certain, J. Masse, O. Van Canneyt, P. Petitgas, G. Doremus, M. B. Santos, and V. Ridoux. Investigating the coupling between small pelagic fish and marine top predators using data collected from ecosystem-based surveys. Marine Ecology Progress Series, 422:23–39, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/422/m422p023.pdf.

Claiborne:2014:ESM

[CMW^{+}14] A. M. Claiborne, J. A. Miller, L. A. Weitkamp, D. J. Teel, and R. L. Emmett. Evidence for selective mortality in marine environments: the role of fish migration size, timing, and production type. Marine Ecology Progress Series, 515: 187–202, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/515/m515p187.pdf.

Cnudde:2015:TEI

[CMW^{+}15] C. Cnudde, T. Moens, E. Werbrouck, G. Lepoint, D. Van Gansbeke, and M. De Troch. Trophodynamics of estuarine intertidal harpacticoid copepods based on stable isotope composition and fatty acid profiles. Marine Ecology Progress Series, 524:225–239, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/524/m524p225.pdf.

Chu:2011:GSR

[CMYL11] J. W. F. Chu, M. Maldonado, G. Yahel, and S. P. Leys. Glass sponge reefs as a silicon sink. Marine Ecology Progress Series, 441:1–14, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Costalago:2012:OSC

[CNÁCP12] D. Costalago, J. Navarro, I. Álvarez-Calleja, and I. Palomera. Ontogenetic and seasonal changes in the feeding habits and trophic levels of two small pelagic fish species. Marine Ecology Progress Series, 460:169–181, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/460/m460p169.pdf.

Couturier:2018:VOH

[CNJ^{+}18] L. I. E. Couturier, P. Newman, F. R. A. Jaine, M. B. Bennett, W. N. Venables, E. F. Cagua, K. A. Townsend, S. J. Weeks, and A. J. Richardson. Variation in occupancy and habitat use of Mobula alfredi at a major aggregation site. Marine Ecology Progress Series, 599:125–145, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m599p125.pdf.

Cumming:2016:TTG

[CNSW16] R. A. Cumming, R. Nikula, H. G. Spencer, and J. M. Waters. Trans-tasman genetic connectivity in the intertidal airbreathing slug Onchidella nigricans. Marine Ecology Progress Series, 562:93–100, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/562/m562p093.pdf.

Croll:2012:MHU

[CNW^{+}12] D. A. Croll, K. M. Newton, K. Weng, F. Galván-Magaña, J. O’Sullivan, and H. Dewar. Movement and habitat use by the spine-tail devil ray in the Eastern Pacific Ocean. Marine Ecology Progress Series, 465:193–200, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/465/
m465p193.pdf.

Collin:2016:ISE

[CO16] R. Collin and I. Ochoa. Influence of seasonal environmental variation on the reproduction of four tropical marine gastropods. Marine Ecology Progress Series, 555:125– 139, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/555/m555p125.pdf.

Costa:2010:MTD

[COK10] T. L. Costa, T. D. O’Hara, and M. J. Keough. Measures of taxonomic distinctness do not reliably assess anthropogenic impacts on intertidal mollusc communities. Marine Ecology Progress Series, 413:81–93, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/413/m413p081.pdf.

Cole:2010:ACB

[Col10a] V. J. Cole. Alteration of the configuration of bioengineers affects associated taxa. Marine Ecology Progress Series, 416: 127–136, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/416/m416p127.pdf.

Collin:2010:RES

[Col10b] R. Collin. Repeatability of egg size in two marine gastropods: brood order and female size do not contribute to intraspecific variation. Marine Ecology Progress Series, 410: 89–96, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/410/m410p089.pdf.

Cook:2011:COM

[Coo11] G. S. Cook. Changes in otolith microchemistry over a protracted spawning season influence assignment of natal origin. Marine Ecology Progress Series, 423:197–209, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/423/
m423p197.pdf.

Carrasco:2010:STV

[CP10] N. K. Carrasco and R. Perissinotto. Spatial and temporal variations in the diet of the mysid Mesopodopsis africana in the St. Lucia Estuary (South Africa). Marine Ecology Progress Series, 417:127–138, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/417/m417p127.pdf.

Cole:2011:ISV

[CP11] A. J. Cole and M. S. Pratchett. Inter-specific variation in susceptibility to grazing among common reef corals. Marine Ecology Progress Series, 422:155–164, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/422/
m422p155.pdf.

Carroll:2011:PSI

[CPA^{+}11] E. Carroll, N. Patenaude, A. Alexander, D. Steel, R. Harcourt, S. Childerhouse, S. Smith, J. Bannister, R. Constantine, and C. S. Baker. Population structure and individual movement of southern right whales around New Zealand and Australia. Marine Ecology Progress Series, 432:257–268, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m432p257.pdf.

Canali:2011:EAO

[CPB^{+}11] E. Canali, G. Ponte, P. Belcari, F. Rocha, and G. Fiorito. Evaluating age in Octopus vulgaris: estimation, validation and seasonal differences. Marine Ecology Progress Series, 441:141–149, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/441/m441p141.pdf.

Carlson-Perret:2018:DFR

[CPEE18] N. L. Carlson-Perret, D. V. Erler, and B. D. Eyre. Dinitrogen (N_{2}) fixation rates in a subtropical seagrass meadow measured with a direct ^{1}5N–N_{2} tracer method. Marine Ecology Progress Series, 605:87–101, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/605/m605p087.pdf.

Canepuccia:2013:DPS

[CPF^{+}13] A. D. Canepuccia, C. F. Pérez, J. L. Farina, D. Alemany, and O. O. Iribarne. Dissimilarity in plant species diversity between salt marsh and neighboring environments decreases as environmental harshness increases. Marine Ecology Progress Series, 494:135–148, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/494/m494p135.pdf.

Ceia:2014:ASC

[CPF^{+}14] F. R. Ceia, V. H. Paiva, V. Fidalgo, L. Morais, A. Baeta, P. Crisóstomo, E. Mourato, S. Garthe, J. C. Marques, and J. A. Ramos. Annual and seasonal consistency in the feeding ecology of an opportunistic species, the yellow-legged gull Larus michahellis. Marine Ecology Progress Series, 497:273– 284, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/497/m497p273.pdf.

Christiansen:2016:SVD

[CPF^{+}16] F. Christiansen, N. F. Putman, R. Farman, D. M. Parker, M. R. Rice, J. J. Polovina, G. H. Balazs, and G. C. Hays. Spatial variation in directional swimming enables juvenile sea turtles to reach and remain in productive waters. Marine Ecology Progress Series, 557:247–259, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m557p247.pdf.

Cohen:2014:CPA

[CPG^{+}14] L. A. Cohen, L. Pichegru, D. Grémillet, J. Coetzee, L. Upfold, and P. G. Ryan. Changes in prey availability impact the foraging behaviour and fitness of Cape gannets over a decade. Marine Ecology Progress Series, 505:281–293, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/505/
m505p281.pdf.

Cornwall:2015:CMI

[CPHH15] C. E. Cornwall, C. A. Pilditch, C. D. Hepburn, and C. L. Hurd. Canopy macroalgae influence understorey corallines’ metabolic control of near-surface pH and oxygen concentration. Marine Ecology Progress Series, 525:81–95, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2015/525/m525p081.pdf.

Choy:2013:DFI

[CPID13] C. A. Choy, E. Portner, M. Iwane, and J. C. Drazen. Diets of five important predatory mesopelagic fishes of the central North Pacific. Marine Ecology Progress Series, 492:169– 184, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m492p169.pdf.

Clay:2017:EOG

[CPM^{+}17] Thomas A. Clay, Richard A. Phillips, Andrea Manica, Hazel A. Jackson, and M. de L.Brooke. Escaping the oligotrophic gyre? The year-round movements, foraging behaviour and habitat preferences of Murphy’s petrels. Marine Ecology Progress Series, 579:139–155, 2017. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m579p139.pdf.

Cook:2011:RPE

[CPMR11] A. S. C. P. Cook, M. Parsons, I. Mitchell, and R. A. Robinson. Reconciling policy with ecological requirements in biodiversity monitoring. Marine Ecology Progress Series, 434: 267–277, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m434p267.pdf.

Carrasco:2010:ZSL

[CPP10] N. K. Carrasco, R. Perissinotto, and D. Pillay. Zooplankton of the St. Lucia Estuary during the current drought cycle: a comparison between open- and closed-mouth conditions. Marine Ecology Progress Series, 399:157–171, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/399/
m399p157.pdf.

Carrasco:2012:OSM

[CPPM12] S. A. Carrasco, N. E. Phillips, and A. Pérez-Matus. Offspring size and maternal environments mediate the early juvenile performance of two congeneric whelks. Marine Ecology Progress Series, 459:73–83, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/459/m459p073.pdf.

Cleary:2014:CRN

[CPR^{+}14] D. F. R. Cleary, A. R. M. Polónia, W. Renema, B. W. Hoeksema, J. Wolstenholme, Y. Tuti, and N. J. de Voogd. Coral reefs next to a major conurbation: a study of temporal change (1985–2011) in coral cover and composition in the reefs of Jakarta, Indonesia. Marine Ecology Progress Series, 501:89–98, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m501p089.pdf.

Chevillot:2016:ASG

[CPR^{+}16] X. Chevillot, M. Pierre, A. Rigaud, H. Drouineau, A. Chaalali, B. Sautour, and J. Lobry. Abrupt shifts in the Gironde fish community: an indicator of ecological changes in an estuarine ecosystem. Marine Ecology Progress Series, 549:137– 151, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m549p137.pdf.

Corbisier:2014:TSB

[CPS^{+}14] T. N. Corbisier, M. A. V. Petti, L. S. H. Soares, E. Y. Muto, S. Bromberg, and I. Valiela. Trophic structure of benthic communities in the Cabo Frio upwelling system (southeastern Brazilian shelf): a temporal study using stable isotope analysis. Marine Ecology Progress Series, 512:23–38, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m512p023.pdf.

Conroy:2015:CEE

[CPS15] C. W. Conroy, P. M. Piccoli, and D. H. Secor. Carryover effects of early growth and river flow on partial migration in striped bass Morone saxatilis. Marine Ecology Progress Series, 541:179–194, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/541/m541p179.pdf.

Calder-Potts:2015:MSI

[CPSC^{+}15] R. Calder-Potts, J. I. Spicer, P. Calosi, H. S. Findlay, and S. Widdicombe. A mesocosm study investigating the effects of hypoxia and population density on respiration and reproductive biology in the brittlestar Amphiura filiformis. Marine Ecology Progress Series, 534:135–147, 2015. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m534p135.pdf.

Calder-Potts:2018:DDR

[CPSC^{+}18] R. N. Calder-Potts, J. I. Spicer, P. Calosi, H. S. Findlay, A. M. Queirós, and S. Widdicombe. Density-dependent responses of the brittlestar Amphiura filiformis to moderate hypoxia and consequences for nutrient fluxes. Marine Ecology Progress Series, 594:175–191, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/594/m594p175.pdf.

Charalampopoulou:2011:IPA

[CPTL11] A. Charalampopoulou, A. J. Poulton, T. Tyrrell, and M. I. Lucas. Irradiance and pH affect coccolithophore community composition on a transect between the North Sea and the Arctic Ocean. Marine Ecology Progress Series, 431: 25–43, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/431/m431p025.pdf.

Collet:2015:ARF

[CPW15] J. Collet, S. C. Patrick, and H. Weimerskirch. Albatrosses redirect flight towards vessels at the limit of their visual range. Marine Ecology Progress Series, 526:199–205, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/526/
m526p199.pdf.

Couto:2017:OBS

[CQR^{+}17] A. Couto, N. Queiroz, P. Relvas, M. Baptista, M. Furtado, J. Castro, M. Nunes, H. Morikawa, and R. Rosa. Occurrence of basking shark Cetorhinus maximus in southern Portuguese waters: a two-decade survey. Marine Ecology Progress Series, 564:77–86, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/564/m564p077.pdf.

Corman:2011:CSM

[CR11] S. S. Corman and C. T. Roman. Comparison of salt marsh creeks and ditches as habitat for nekton. Marine Ecology Progress Series, 434:57–66, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/434/m434p057.pdf.

Campbell:2014:EHI

[CR14] L. A. Campbell and J. A. Rice. Effects of hypoxia-induced habitat compression on growth of juvenile fish in the Neuse River Estuary, North Carolina, USA. Marine Ecology Progress Series, 497:199–213, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2014/497/m497p199.pdf.

Canesi:2016:TVS

[CR16] K. L. Canesi and T. A. Rynearson. Temporal variation of Skeletonema community composition from a long-term time series in Narragansett Bay identified using high-throughput DNA sequencing. Marine Ecology Progress Series, 556:1–16, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Carrier:2019:BCD

[CR19] T. J. Carrier and A. M. Reitzel. Bacterial community dynamics during embryonic and larval development of three confamilial echinoids. Marine Ecology Progress Series, 611: 179–188, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/611/m611p179.pdf.

Craig:2012:AEE

[Cra12] J. K. Craig. Aggregation on the edge: effects of hypoxia avoidance on the spatial distribution of brown shrimp and demersal fishes in the Northern Gulf of Mexico. Marine Ecology Progress Series, 445:75–95, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/445/m445p075.pdf.

Cabanellas-Reboredo:2012:MPE

[CRAP^{+}12] M. Cabanellas-Reboredo, J. Alós, M. Palmer, D. March, and R. O’Dor. Movement patterns of the European squid Loligo vulgaris during the inshore spawning season. Marine Ecology Progress Series, 466:133–144, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/466/m466p133.pdf.

Cruz-Rivera:2013:EAP

[CRF13] E. Cruz-Rivera and M. Friedlander. Effects of algal phenotype on mesograzer feeding. Marine Ecology Progress Series, 490:69–78, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/490/m490p069.pdf.

Cartaxana:2016:LMT

[CRG^{+}16] P. Cartaxana, L. Ribeiro, J. W. Goessling, S. Cruz, and M. Kühl. Light and O 2 microenvironments in two contrasting diatom-dominated coastal sediments. Marine Ecology Progress Series, 545:35–47, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/545/m545p035.pdf.

Collin:2018:DTS

[CRGM18] R. Collin, F. Rendina, V. Goodwin, and S. McCabe. Do tropical specialist sea urchins have higher thermal tolerances and optimal temperatures than their more widely distributed relatives? Marine Ecology Progress Series, 589: 153–166, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/589/m589p153.pdf.

Calliari:2019:DDG

[CRGT19] D. Calliari, L. Rodríguez-Graña, and P. Tiselius. Density dependent grazing rates in a natural microzooplankton community. Marine Ecology Progress Series, 622:83– 92, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/622/m622p083.pdf.

Cresson:2014:ARD

[CRHV14] P. Cresson, S. Ruitton, and M. Harmelin-Vivien. Artificial reefs do increase secondary biomass production: mechanisms evidenced by stable isotopes. Marine Ecology Progress Series, 509:15–26, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/509/m509p015.pdf.

Chandarana:2013:GSP

[CRJ13] J. Chandarana, D. J. Reish, and R. L. Johnson. Genetic structure of the polychaete Nereis grubei in the context of current patterns and life history. Marine Ecology Progress Series, 473:215–224, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/473/m473p215.pdf.

Collet:2018:IDC

[CRJ^{+}18] J. Collet, G. Richard, A. Janc, C. Guinet, and H. Weimerskirch. Influence of depredating cetaceans on albatross attraction and attendance patterns at fishing boats. Marine Ecology Progress Series, 605:49–59, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/605/m605p049.pdf.

Christiansen:2013:WWD

[CRL13] F. Christiansen, M. Rasmussen, and D. Lusseau. Whale watching disrupts feeding activities of minke whales on a feeding ground. Marine Ecology Progress Series, 478:239– 251, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/478/m478p239.pdf.

Cartaxana:2019:NSD

[CRR^{+}19] P. Cartaxana, F. Rey, M. Ribeiro, A. S. P. Moreira, M. R. M. Domingues, R. Calado, and S. Cruz. Nutritional state determines reproductive investment in the mixotrophic sea slug Elysia viridis. Marine Ecology Progress Series, 611:167– 177, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/611/m611p167.pdf.

Cermeno:2013:SRM

[CRRD^{+}13] P. Cermeño, T. Rodríguez-Ramos, M. Dornelas, F. G. Figueiras, E. Marañón, I. G. Teixeira, and S. Vallina. Species richness in marine phytoplankton communities is not correlated to ecosystem productivity. Marine Ecology Progress Series, 488:1–9, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Choy:2017:IAV

[CRRL17] E. S. Choy, B. Rosenberg, J. D. Roth, and L. L. Loseto. Interannual variation in environmental factors affect the prey and body condition of beluga whales in the eastern Beaufort Sea. Marine Ecology Progress Series, 579:213–225, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/579/
m579p213.pdf.

Calbet:2011:PGM

[CRS^{+}11] A. Calbet, K. Riisgaard, E. Saiz, S. Zamora, C. Stedmon, and T. G. Nielsen. Phytoplankton growth and microzooplankton grazing along a sub-Arctic fjord (Godthåbsfjord, west Greenland). Marine Ecology Progress Series, 442:11– 22, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/442/m442p011.pdf.

Copeman:2018:BED

[CRS^{+}18] L. Copeman, C. Ryer, M. Spencer, M. Ottmar, P. Iseri, A. Sremba, J. Wells, and C. Parrish. Benthic enrichment by diatom-sourced lipid promotes growth and condition in juvenile Tanner crabs around Kodiak Island, Alaska. Marine Ecology Progress Series, 597:161–178, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/597/
m597p161.pdf.

Cunning:2016:PBR

[CRWG16] Ross Cunning, Raphael Ritson-Williams, and Ruth D. Gates. Patterns of bleaching and recovery of Montipora capitata in Kāne’ohe Bay, Hawai’i, USA. Marine Ecology Progress Series, 551:131–139, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/551/m551p131.pdf.

Colton:2010:CTS

[CS10a] M. A. Colton and S. E. Swearer. A comparison of two survey methods: differences between underwater visual census and baited remote underwater video. Marine Ecology Progress Series, 400:19–36, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/400/m400p019.pdf.

Conti:2010:FYP

[CS10b] L. Conti and M. Scardi. Fisheries yield and primary productivity in large marine ecosystems. Marine Ecology Progress Series, 410:233–244, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/410/m410p233.pdf.

Cooper:2011:LCS

[CS11a] E. E. Cooper and A. L. Shanks. Latitude and coastline shape correlate with age-structure of Chlorostoma (Tegula) funebralis populations. Marine Ecology Progress Series, 424: 133–143, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/424/m424p133.pdf.

Currie:2011:SSL

[CS11b] J. J. Currie and D. C. Schneider. Spatial scaling from latitudinal gradients: size-specific fecundity in the American lobster Homarus americanus. Marine Ecology Progress Series, 439:193–201, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m439p193.pdf.

Connolly:2013:SAS

[CS13] R. M. Connolly and T. A. Schlacher. Sample acidification significantly alters stable isotope ratios of sulfur in aquatic plants and animals. Marine Ecology Progress Series, 493:1–8, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Cheung:2015:EIO

[CS15] W. W. L. Cheung and U. R. Sumaila. Economic incentives and overfishing: a bioeconomic vulnerability index. Marine Ecology Progress Series, 530:223–232, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m530p223.pdf.

Carreiro-Silva:2013:VGR

[CSABH^{+}13] M. Carreiro-Silva, A. H. Andrews, A. Braga-Henriques, V. de Matos, F. M. Porteiro, and R. S. Santos. Variability in growth rates of long-lived black coral Leiopathes sp. from the Azores. Marine Ecology Progress Series, 473:189–199, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m473p189.pdf.

Catano:2014:PRC

[CSB14] L. B. Catano, A. A. Shantz, and D. E. Burkepile. Predation risk, competition, and territorial damselfishes as drivers of herbivore foraging on Caribbean coral reefs. Marine Ecology Progress Series, 511:193–207, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/511/m511p193.pdf.

Chouvelon:2011:ISO

[CSC^{+}11] T. Chouvelon, J. Spitz, Y. Cherel, F. Caurant, R. Sirmel, P. Mèndez-Fernandez, and P. Bustamante. Inter-specific and ontogenic differences in δ ^{13}C and δ ^{15}N values and Hg and Cd concentrations in cephalopods. Marine Ecology Progress Series, 433:107–120, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m433p107.pdf.

Cebrian:2013:EGF

[CSC13a] J. Cebrian, J. P. Stutes, and B. Christiaen. Effects of grazing and fertilization on epiphyte growth dynamics under moderately eutrophic conditions: implications for grazing rate estimates. Marine Ecology Progress Series, 474:121–133, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m474p121.pdf.

Cox:2013:CST

[CSC13b] S. L. Cox, B. E. Scott, and C. J. Camphuysen. Combined spatial and tidal processes identify links between pelagic prey species and seabirds. Marine Ecology Progress Series, 479: 203–221, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/479/m479p203.pdf.

Chambellant:2013:TVW

[CSF13] M. Chambellant, I. Stirling, and S. H. Ferguson. Temporal variation in western Hudson Bay ringed seal Phoca hispida diet in relation to environment. Marine Ecology Progress Series, 481:269–287, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/481/m481p269.pdf.

Cannicci:2011:REC

[CSG11] S. Cannicci, R. Simoni, and F. Giomi. Role of the embryo in crab terrestrialisation: an ontogenetic approach. Marine Ecology Progress Series, 430:121–131, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m430p121.pdf.

Campbell:2017:EOA

[CSJP17] J. E. Campbell, J. M. Sneed, L. Johnston, and V. J. Paul. Effects of ocean acidification and contact with the brown alga Stypopodium zonale on the settlement and early survival of the coral Porites astreoides. Marine Ecology Progress Series, 577:67–77, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/577/m577p067.pdf.

Cavanaugh:2010:SGK

[CSKR10] K. C. Cavanaugh, D. A. Siegel, B. P. Kinlan, and D. C. Reed. Scaling giant kelp field measurements to regional scales using satellite observations. Marine Ecology Progress Series, 403:13–27, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m403p013.pdf.

Carreiro-Silva:2012:MDB

[CSM12] M. Carreiro-Silva and T. R. McClanahan. Macrobioerosion of dead branching Porites, 4 and 6 years after coral mass mortality. Marine Ecology Progress Series, 458:103–122, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m458p103.pdf.

Chalifour:2019:HUJ

[CSM^{+}19] L. Chalifour, D. C. Scott, M. MacDuffee, J. C. Iacarella, T. G. Martin, and J. K. Baum. Habitat use by juvenile salmon, other migratory fish, and resident fish species underscores the importance of estuarine habitat mosaics. Marine Ecology Progress Series, 625:145–162, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2019/625/m625p145.pdf.

Cutajar:2012:IIG

[CSN12] J. Cutajar, J. Shimeta, and D. Nugegoda. Impacts of the invasive grass Spartina anglica on benthic macrofaunal assemblages in a temperate Australian saltmarsh. Marine Ecology Progress Series, 464:107–120, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/464/m464p107.pdf.

Coll:2010:EAN

[CSPA10] M. Coll, A. Santojanni, I. Palomera, and E. Arneri. Ecosystem assessment of the North-Central Adriatic Sea: towards a multivariate reference framework. Marine Ecology Progress Series, 417:193–210, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/417/m417p193.pdf.

Cuna:2011:EMG

[CSQRA11] V. Cuña, M. Saura, H. Quesada, and E. Rolán-Alvarez. Extensive micro-geographical shell polymorphism in a planktotrophic marine intertidal snail. Marine Ecology Progress Series, 427:133–143, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/427/m427p133.pdf.

Caldeira:2014:EPC

[CSR^{+}14] C. Caldeira, A. M. P. Santos, P. Ré, M. A. Peck, E. Saiz, and S. Garrido. Effects of prey concentration on ingestion rates of European sardine Sardina pilchardus larvae in the laboratory. Marine Ecology Progress Series, 517:217–228, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/517/
m517p217.pdf.

Cavanaugh:2011:ECG

[CSRD11] K. C. Cavanaugh, D. A. Siegel, D. C. Reed, and P. E. Dennison. Environmental controls of giant-kelp biomass in the Santa Barbara Channel, California. Marine Ecology Progress Series, 429:1–17, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Clare:2016:SDB

[CSRF16] D. S. Clare, M. Spencer, L. A. Robinson, and C. L. J. Frid. Species densities, biological interactions and benthic ecosystem functioning: an in situ experiment. Marine Ecology Progress Series, 547:149–161, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/547/m547p149.pdf.

Cook:2015:BPB

[CSRS15] C. A. Cook, P. C. Sikkel, L. P. Renoux, and N. J. Smit. Blood parasite biodiversity of reef-associated fishes of the eastern Caribbean. Marine Ecology Progress Series, 533:1–13, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Croxall:2013:ASD

[CSS^{+}13] J. Croxall, C. Small, B. Sullivan, R. Wanless, E. Frere, B. Lascelles, I. Ramirez, M. Sato, and O. Yates. Appropriate scales and data to manage seabird–fishery interactions: Comment on Torres et al. (2013). Marine Ecology Progress Series, 493: 297–300, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/493/m493p297.pdf. See [TSTP13c].

Champion:2015:ZKP

[CSS15] C. Champion, I. M. Suthers, and J. A. Smith. Zooplanktivory is a key process for fish production on a coastal artificial reef. Marine Ecology Progress Series, 541:1–14, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Colomer:2017:IAC

[CSS^{+}17] J. Colomer, M. Soler, T. Serra, X. Casamitjana, and C. Oldham. Impact of anthropogenically created canopy gaps on wave attenuation in a Posidonia oceanica seagrass meadow. Marine Ecology Progress Series, 569:103–116, 2017. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m569p103.pdf.

Cure:2015:PSP

[CSV^{+}15] C. Curé, L. D. Sivle, F. Visser, P. J. Wensveen, S. Isojunno, C. M. Harris, P. H. Kvadsheim, F. P. A. Lam, and P. J. O. Miller. Predator sound playbacks reveal strong avoidance responses in a fight strategist baleen whale. Marine Ecology Progress Series, 526:267–282, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/526/m526p267.pdf.

Calderon:2010:TGV

[CT10] I. Calderón and X. Turon. Temporal genetic variability in the Mediterranean common sea urchin Paracentrotus lividus. Marine Ecology Progress Series, 408:149–159, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/408/
m408p149.pdf.

Ciotti:2013:SVG

[CTB13] B. J. Ciotti, T. E. Targett, and M. T. Burrows. Spatial variation in growth rate of early juvenile European plaice Pleuronectes platessa. Marine Ecology Progress Series, 475: 213–232, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/475/m475p213.pdf.

Chapman:2010:CIP

[CTMU10] M. G. Chapman, T. J. Tolhurst, R. J. Murphy, and A. J. Underwood. Complex and inconsistent patterns of variation in benthos, micro-algae and sediment over multiple spatial scales. Marine Ecology Progress Series, 398:33–47, 2010. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m398p033.pdf.

Cabral-Tena:2013:DCR

[CTRBLC^{+}13] R. A. Cabral-Tena, H. Reyes-Bonilla, S. Lluch-Cota, D. A. Paz-García, L. E. Calderón-Aguilera, O. Norzagaray-López, and E. F. Balart. Different calcification rates in males and females of the coral Porites panamensis in the Gulf of California. Marine Ecology Progress Series, 476:1–8, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Crossin:2015:HSF

[CTS^{+}15] G. T. Crossin, A. Takahashi, K. Q. Sakamoto, P. N. Trathan, and T. D. Williams. Habitat selection by foraging macaroni penguins correlates with hematocrit, an index of aerobic condition. Marine Ecology Progress Series, 530:163– 176, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/530/m530p163.pdf.

Chin:2012:RSI

[CTSH12] A. Chin, A. Tobin, C. Simpfendorfer, and M. Heupel. Reef sharks and inshore habitats: patterns of occurrence and implications for vulnerability. Marine Ecology Progress Series, 460:115–125, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/460/m460p115.pdf.

Compton:2008:DOR

[CTvdM^{+}08] Tanya J. Compton, Tineke A. Troost, Jaap van der Meer, Casper Kraan, Pieter J. C. Honkoop, Danny I. Rogers, Grant B. Pearson, Petra de Goeij, Pierrick Bocher, Marc S. S. Lavaleye, Jutta Leyrer, Mick G. Yates, Anne Dekinga, and Theunis Piersma. Distributional overlap rather than habitat differentiation characterizes co-occurrence of bivalves in intertidal soft sediment systems. Marine Ecology Progress Series, 373:25–35, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/373/m373p025.pdf. See comment

[BD11a].

Cerchio:2016:STH

[CTZ^{+}16] S. Cerchio, L. Trudelle, A. N. Zerbini, J. B. Charrassin, Y. Geyer, F. X. Mayer, N. Andrianarivelo, J. L. Jung, O. Adam, and H. C. Rosenbaum. Satellite telemetry of humpback whales off Madagascar reveals insights on breeding behavior and long-range movements within the southwest Indian Ocean. Marine Ecology Progress Series, 562: 193–209, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/562/m562p193.pdf.

Corrales-Ugalde:2017:NDC

[CUCS17] M. Corrales-Ugalde, S. P. Colin, and K. R. Sutherland. Nematocyst distribution corresponds to prey capture location in hydromedusae with different predation modes. Marine Ecology Progress Series, 568:101–110, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/568/m568p101.pdf.

Cook:2011:DPF

[CV11] K. Cook and M. A. Vanderklift. Depletion of predatory fish by fishing in a temperate reef ecosystem leads to indirect effects on prey, but not to lower trophic levels. Marine Ecology Progress Series, 432:195–205, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/432/m432p195.pdf.

Christiansen:2018:MBS

[CVC^{+}18] F. Christiansen, F. Vivier, C. Charlton, R. Ward, A. Amerson, S. Burnell, and L. Bejder. Maternal body size and condition determine calf growth rates in southern right whales. Marine Ecology Progress Series, 592:267–281, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/592/
m592p267.pdf.

Cox:2017:EMP

[CVM^{+}17] C. Cox, A. Valdivia, M. McField, K. Castillo, and J. F. Bruno. Establishment of marine protected areas alone does not restore coral reef communities in Belize. Marine Ecology Progress Series, 563:65–79, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/563/m563p065.pdf.

Cumbo:2018:TSP

[CvOB18] V. R. Cumbo, M. J. H. van Oppen, and A. H. Baird. Temperature and Symbiodinium physiology affect the establishment and development of symbiosis in corals. Marine Ecology Progress Series, 587:117–127, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/587/m587p117.pdf.

Cook:2011:SEH

[CVP11] K. Cook, M. A. Vanderklift, and A. G. B. Poore. Strong effects of herbivorous amphipods on epiphyte biomass in a temperate seagrass meadow. Marine Ecology Progress Series, 442:263–269, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/442/m442p263.pdf.

Capuska:2011:DSF

[CVW^{+}11] G. E. Machovsky Capuska, R. L. Vaughn, B. Würsig, G. Katzir, and D. Raubenheimer. Dive strategies and foraging effort in the Australasian gannet Morus serrator revealed by underwater videography. Marine Ecology Progress Series, 442:255–261, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2011/442/m442p255.pdf.

Carilli:2012:BFA

[CW12] J. Carilli and S. Walsh. Benthic Foraminiferal assemblages from Kiritimati (Christmas) Island indicate human-mediated nutrification has occurred over the scale of decades. Marine Ecology Progress Series, 456:87–99, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/456/
m456p087.pdf.

Curtis:2017:DSN

[CWAS17] J. S. Curtis, K. R. Wall, M. A. Albins, and C. D. Stallings. Diet shifts in a native mesopredator across a range of invasive lionfish biomass. Marine Ecology Progress Series, 573: 215–228, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/573/m573p215.pdf.

Cleasby:2015:SSW

[CWB^{+}15] I. R. Cleasby, E. D. Wakefield, T. W. Bodey, R. D. Davies, S. C. Patrick, J. Newton, S. C. Votier, S. Bearhop, and K. C. Hamer. Sexual segregation in a wide-ranging marine predator is a consequence of habitat selection. Marine Ecology Progress Series, 518:1–12, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Cass:2011:LCT

[CWD11] C. J. Cass, S. G. Wakeham, and K. L. Daly. Lipid composition of tropical and subtropical copepod species of the genus Rhincalanus (Copepoda: Eucalanidae): a novel fatty acid and alcohol signature. Marine Ecology Progress Series, 439:127–138, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/439/m439p127.pdf.

Chen:2019:IHD

[CWEE19] J. J. Chen, N. S. Wells, D. V. Erler, and B. D. Eyre. Importance of habitat diversity to changes in benthic metabolism over land-use gradients: evidence from three subtropical estuaries. Marine Ecology Progress Series, 631:31–47, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2019/631/m631p031.pdf.

Colleter:2015:IFA

[CWGG15] M. Colléter, C. J. Walters, P. Gatti, and D. Gascuel. Including foraging arena and top-down controls improves the modeling of trophic flows and fishing impacts in aquatic food webs. Marine Ecology Progress Series, 534:17–37, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/534/
m534p017.pdf.

Cartamil:2010:MPN

[CWK^{+}10] D. Cartamil, N. C. Wegner, D. Kacev, N. Ben-aderet, S. Kohin, and J. B. Graham. Movement patterns and nursery habitat of juvenile thresher sharks Alopias vulpinus in the Southern California Bight. Marine Ecology Progress Series, 404:249–258, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/404/m404p249.pdf.

Coker:2013:SGE

[CWMP13] D. J. Coker, S. P. W. Walker, P. L. Munday, and M. S. Pratchett. Social group entry rules may limit population resilience to patchy habitat disturbance. Marine Ecology Progress Series, 493:237–242, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/493/m493p237.pdf.

Calmanovici:2018:PMR

[CWR^{+}18] B. Calmanovici, D. Waayers, J. Reisser, J. Clifton, and M. Proietti. I^{3}S pattern as a mark–recapture tool to identify captured and free-swimming sea turtles: an assessment. Marine Ecology Progress Series, 589:263–268, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/589/
m589p263.pdf.

Carey:2013:SSJ

[CWS13] J. D. Carey, R. A. Wahle, and K. D. E. Stokesbury. Spatial scaling of juvenile-adult associations in northwest Atlantic sea scallop Placopecten magellanicus populations. Marine Ecology Progress Series, 493:185–194, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/493/m493p185.pdf.

Cartamil:2016:MDJ

[CWW^{+}16a] D. Cartamil, J. Wraith, N. C. Wegner, D. Kacev, C. H. Lam, O. Santana-Morales, O. Sosa-Nishizaki, M. Escobedo-Olvera, S. Kohin, J. B. Graham, and P. Hastings. Movements and distribution of juvenile common thresher sharks Alopias vulpinus in Pacific coast waters of the USA and Mexico. Marine Ecology Progress Series, 548:153–163, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/548/m548p153.pdf.

Choy:2016:FWT

[CWW^{+}16b] C. A. Choy, C. C. C. Wabnitz, M. Weijerman, P. A. Woodworth-Jefcoats, and J. J. Polovina. Finding the way to the top: how the composition of oceanic mid-trophic micronekton groups determines apex predator biomass in the central North Pacific. Marine Ecology Progress Series, 549:9–25, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m549p009.pdf.

Cherel:2017:FEI

[CXdG^{+}17] Y. Cherel, J. C. Xavier, S. de Grissac, C. Trouvé, and H. Weimerskirch. Feeding ecology, isotopic niche, and ingestion of fishery-related items of the wandering albatross Diomedea exulans at Kerguelen and Crozet Islands. Marine Ecology Progress Series, 565:197–215, 2017. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m565p197.pdf.

Chen:2010:CPM

[CY10] G. C. Chen and Y. Ye. Changes in properties of mangrove sediment due to foraging on Kandelia obovata leaves by crabs Parasesarmea plicatum (Grapsidae: Sesarminae). Marine Ecology Progress Series, 419:137–145, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/419/
m419p137.pdf.

Chen:2013:GPR

[CY13] Y. P. Chen and Y. Ye. Growth and physiological responses of saplings of two mangrove species to intertidal elevation.

Marine Ecology Progress Series, 482:107–118, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/482/
m482p107.pdf.

Chan:2010:OBR

[CZB10] K. S. Chan, T. Zhang, and K. M. Bailey. Otolith biochronology reveals factors underlying dynamics in marine fish larvae. Marine Ecology Progress Series, 412:1–10, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Criales:2013:ETB

[CZH^{+}13] M. M. Criales, I. C. Zink, B. K. Haus, J. Wylie, and J. A. Browder. Effect of turbulence on the behavior of pink shrimp postlarvae and implications for selective tidal stream transport behavior. Marine Ecology Progress Series, 477:161– 176, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/477/m477p161.pdf.

Deza:2010:HFP

[DA10] A. A. Deza and T. W. Anderson. Habitat fragmentation, patch size, and the recruitment and abundance of kelp forest fishes. Marine Ecology Progress Series, 416:229–240, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/416/
m416p229.pdf.

Davison:2011:PIM

[DA11] P. Davison and R. G. Asch. Plastic ingestion by mesopelagic fishes in the North Pacific subtropical gyre. Marine Ecology Progress Series, 432:173–180, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/432/m432p173.pdf.

Dzaugis:2017:IST

[DAB17] M. P. Dzaugis, R. J. Allman, and B. A. Black. Importance of the spring transition in the northern Gulf of Mexico as inferred from marine fish biochronologies. Marine Ecology Progress Series, 565:149–162, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/565/m565p149.pdf.

Dudgeon:2010:PSS

[DABP10] S. R. Dudgeon, R. B. Aronson, J. F. Bruno, and W. F. Precht. Phase shifts and stable states on coral reefs. Marine Ecology Progress Series, 413:201–216, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m413p201.pdf.

Daly:2013:WIB

[DABP13] E. A. Daly, T. D. Auth, R. D. Brodeur, and W. T. Peterson. Winter ichthyoplankton biomass as a predictor of early summer prey fields and survival of juvenile salmon in the northern California Current. Marine Ecology Progress Series, 484:203–217, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m484p203.pdf.

Distiller:2012:FAA

[DAC^{+}12] G. Distiller, R. Altwegg, R. J. M. Crawford, N. T. W. Klages, and B. Barham. Factors affecting adult survival and inter-colony movement at the three South African colonies of Cape gannet. Marine Ecology Progress Series, 461:245– 255, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/461/m461p245.pdf.

Domingues:2011:SCL

[DAD^{+}11] C. P. Domingues, M. J. Almeida, J. Dubert, R. Nolasco, N. Cordeiro, S. Waap, A. Sequeira, S. Tavares, and H. Queiroga. Supply of crab larvae to an estuary in the eastern Atlantic upwelling system exhibits predictable and haphazard variation at different temporal scales. Marine Ecology Progress Series, 425:113–124, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/425/m425p113.pdf.

Dahlgren:2010:PPC

[DAL^{+}10] K. Dahlgren, A. Andersson, U. Larsson, S. Hajdu, and U. Båmstedt. Planktonic production and carbon transfer efficiency along a north-south gradient in the Baltic Sea. Marine Ecology Progress Series, 409:77–94, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m409p077.pdf.

Dunn:2017:CIS

[DAM^{+}17] R. P. Dunn, A. H. Altieri, K. Miller, M. Yeager, and K. A. Hovel. Coral identity and structural complexity drive habitat associations and demographic processes for an increasingly important Caribbean herbivore. Marine Ecology Progress Series, 577:33–47, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/577/m577p033.pdf.

Dunn:2018:CBR

[DAM^{+}18] R. P. Dunn, A. H. Altieri, K. Miller, M. E. Yeager, and K. A. Hovel. Contrasting behavioral responses to predatory risk cues reflect different foraging strategies in two Caribbean sea urchins. Marine Ecology Progress Series, 604: 187–198, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/604/m604p187.pdf.

Dando:2011:SFH

[Dan11] P. R. Dando. Site fidelity, homing and spawning migrations of flounder Platichthys flesus in the Tamar estuary, South West England. Marine Ecology Progress Series, 430:183–196, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m430p183.pdf.

Dao:2013:RCS

[Dao13] M. H. Dao. Reassessment of the cell surface area limitation to nutrient uptake in phytoplankton. Marine Ecology Progress Series, 489:87–92, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m489p087.pdf.

Davis:2019:IRS

[Dav19] A. C. D. Davis. Integrating remote sensing and diver observations to predict the distribution of invasive lionfish on Bahamian coral reefs. Marine Ecology Progress Series, 623:1–11, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Derocher:2011:SIP

[DAW^{+}11] A. E. Derocher, M. Andersen, Ø. Wiig, J. Aars, E. Hansen, and M. Biuw. Sea ice and polar bear den ecology at Hopen Island, Svalbard. Marine Ecology Progress Series, 441:273–279, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m441p273.pdf.

DuBois:2019:RPE

[DAWS19] K. DuBois, J. M. Abbott, S. L. Williams, and J. J. Stachowicz. Relative performance of eelgrass genotypes shifts during an extreme warming event: disentangling the roles of multiple traits. Marine Ecology Progress Series, 615: 67–77, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/615/m615p067.pdf.

deBoer:2010:SDD

[dB10] M. N. de Boer. Spring distribution and density of minke whale Balaenoptera acutorostrata along an offshore bank in the central North Sea. Marine Ecology Progress Series, 408: 265–274, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/408/m408p265.pdf.

Dekaezemacker:2011:SFC

[DB11a] J. Dekaezemacker and S. Bonnet. Sensitivity of N_{2} fixation to combined nitrogen forms (NO_{3}^{−} and NH_{4}^{+}) in two strains of the marine diazotroph Crocosphaera watsonii (Cyanobacteria). Marine Ecology Progress Series, 438:33–46, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/438/
m438p033.pdf.

Drolet:2011:MPD

[DB11b] D. Drolet and M. A. Barbeau. Movement patterns drive within-mudflat distribution of an intertidal amphipod. Marine Ecology Progress Series, 431:255–265, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/431/
m431p255.pdf.

Dekker:2014:PAB

[DB14] R. Dekker and J. J. Beukema. Phenology of abundance of bivalve spat and of their epibenthic predators: limited evidence for mismatches after cold winters. Marine Ecology Progress Series, 513:17–27, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/513/m513p017.pdf.

Daly:2017:AOC

[DBA17] E. A. Daly, R. D. Brodeur, and T. D. Auth. Anomalous ocean conditions in 2015: impacts on spring Chinook salmon and their prey field. Marine Ecology Progress Series, 566:169–182, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/566/m566p169.pdf.

Deakos:2011:CMR

[DBB11] M. H. Deakos, J. D. Baker, and L. Bejder. Characteristics of a manta ray Manta alfredi population off Maui, Hawaii, and implications for management. Marine Ecology Progress Series, 429:245–260, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2011/429/m429p245.pdf.

DeTroch:2012:BFA

[DBC^{+}12a] M. De Troch, P. Boeckx, C. Cnudde, D. Van Gansbeke, A. Vanreusel, M. Vincx, and M. J. Caramujo. Bioconversion of fatty acids at the basis of marine food webs: insights from a compound-specific stable isotope analysis. Marine Ecology Progress Series, 465:53–67, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/465/m465p053.pdf.

Drolet:2012:PBM

[DBC^{+}12b] D. Drolet, T. T. Bringloe, M. R. S. Coffin, M. A. Barbeau, and D. J. Hamilton. Potential for between-mudflat movement and metapopulation dynamics in an intertidal burrowing amphipod. Marine Ecology Progress Series, 449:197– 209, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/449/m449p197.pdf.

DiCarlo:2011:RPO

[DBCB11] G. Di Carlo, L. Benedetti-Cecchi, and F. Badalamenti. Response of Posidonia oceanica growth to dredging effects of different magnitude. Marine Ecology Progress Series, 423: 39–45, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/423/m423p039.pdf.

Diederich:2015:DTL

[DBCP15] C. M. Diederich, S. M. Bashevkin, O. R. Chaparro, and J. A. Pechenik. Desiccation tolerance and lifting behavior in Crepidula fornicata (Gastropoda). Marine Ecology Progress Series, 528:235–243, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2015/528/m528p235.pdf.

Demopoulos:2016:IVD

[DBCS16] A. W. J. Demopoulos, J. R. Bourque, E. Cordes, and K. M. Stamler. Impacts of the Deepwater Horizon oil spill on deepsea coral-associated sediment communities. Marine Ecology Progress Series, 561:51–68, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/561/m561p051.pdf.

Delefosse:2012:MCR

[DBCV^{+}12] M. Delefosse, G. T. Banta, P. Canal-Vergés, G. Penha-Lopes, C. O. Quintana, T. Valdemarsen, and E. Kristensen. Macrobenthic community response to the Marenzelleria viridis (Polychaeta) invasion of a Danish estuary. Marine Ecology Progress Series, 461:83–94, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/461/m461p083.pdf.

Duhamel:2014:MRE

[DBDK14] S. Duhamel, K. M. Björkman, J. K. Doggett, and D. M. Karl. Microbial response to enhanced phosphorus cycling in the North Pacific subtropical gyre. Marine Ecology Progress Series, 504:43–58, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m504p043.pdf.

DeBoer:2012:PSD

[DBE^{+}12] T. S. DeBoer, A. C. Baker, M. V. Erdmann, Ambariyanto, P. R. Jones, and P. H. Barber. Patterns of Symbiodinium distribution in three giant clam species across the biodiverse Bird’s head region of Indonesia. Marine Ecology Progress Series, 444:117–132, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/444/m444p117.pdf.

Durrant:2018:SHP

[DBE^{+}18] H. M. S. Durrant, N. S. Barrett, G. J. Edgar, M. A. Coleman, and C. P. Burridge. Seascape habitat patchiness and hydrodynamics explain genetic structuring of kelp populations. Marine Ecology Progress Series, 587:81–92, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/587/
m587p081.pdf.

DePasquale:2015:VEL

[DBG15] E. DePasquale, H. Baumann, and C. J. Gobler. Vulnerability of early life stage Northwest Atlantic forage fish to ocean acidification and low oxygen. Marine Ecology Progress Series, 523:145–156, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/523/m523p145.pdf.

Durden:2016:IED

[DBH^{+}16] J. M. Durden, B. J. Bett, T. Horton, A. Serpell-Stevens, K. J. Morris, D. S. M. Billett, and H. A. Ruhl. Improving the estimation of deep-sea megabenthos biomass: dimension to wet weight conversions for abyssal invertebrates. Marine Ecology Progress Series, 552:71–79, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m552p071.pdf.

Dragon:2012:CAM

[DBHMG12a] A. C. Dragon, A. Bar-Hen, P. Monestiez, and C. Guinet. Comparative analysis of methods for inferring successful foraging areas from Argos and GPS tracking data. Marine Ecology Progress Series, 452:253–267, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m452p253.pdf.

Dragon:2012:HVM

[DBHMG12b] A. C. Dragon, A. Bar-Hen, P. Monestiez, and C. Guinet. Horizontal and vertical movements as predictors of foraging success in a marine predator. Marine Ecology Progress Series, 447:243–257, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/447/m447p243.pdf.

Durand:2013:PGS

[DBS^{+}13] J. D. Durand, H. Blel, K. N. Shen, E. T. Koutrakis, and B. Guinand. Population genetic structure of Mugil cephalus in the Mediterranean and Black Seas: a single mitochondrial clade and many nuclear barriers. Marine Ecology Progress Series, 474:243–261, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m474p243.pdf.

Davis:2014:SMP

[DBS14] B. Davis, R. Baker, and M. Sheaves. Seascape and metacommunity processes regulate fish assemblage structure in coastal wetlands. Marine Ecology Progress Series, 500:187– 202, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/500/m500p187.pdf.

Durden:2016:CIA

[DBS^{+}16] J. M. Durden, B. J. Bett, T. Schoening, K. J. Morris, T. W. Nattkemper, and H. A. Ruhl. Comparison of image annotation data generated by multiple investigators for benthic ecology. Marine Ecology Progress Series, 552:61–70, 2016. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m552p061.pdf.

Dyer:2019:VSI

[DBS^{+}19] D. C. Dyer, M. J. Butler, A. J. Smit, R. J. Anderson, and J. J. Bolton. Variability in stable isotope values of South African Laminariales, Ecklonia maxima and Laminaria pallida, over different spatial and temporal scales. Marine Ecology Progress Series, 628:55–71, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/628/m628p055.pdf.

Downie:2013:DHF

[DBTV13] R. A. Downie, R. C. Babcock, D. P. Thomson, and M. A. Vanderklift. Density of herbivorous fish and intensity of herbivory are influenced by proximity to coral reefs. Marine Ecology Progress Series, 482:217–225, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/482/m482p217.pdf.

Dalrymple:2015:EAC

[DC15] D. J. Dalrymple and R. H. Carmichael. Effects of age class on N removal capacity of oysters and implications for bioremediation. Marine Ecology Progress Series, 528:205– 220, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/528/m528p205.pdf.

Duffy:2017:OPE

[DC17] G. A. Duffy and S. L. Chown. Opinion piece: Explicitly integrating a third dimension in marine species distribution modelling. Marine Ecology Progress Series, 564:1–8, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Derville:2016:ECN

[DCB^{+}16] Solène Derville, Rochelle Constantine, C. Scott Baker, Marc Oremus, and Leight G. Torres. Environmental correlates of nearshore habitat distribution by the critically endangered Māui dolphin. Marine Ecology Progress Series, 551:261– 275, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/551/m551p261.pdf.

Drago:2010:RSS

[DCC^{+}10] M. Drago, L. Cardona, E. A. Crespo, M. F. Grandi, and A. Aguilar. Reduction of skull size in South American sea lions reveals density-dependent growth during population recovery. Marine Ecology Progress Series, 420:253–261, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/420/
m420p253.pdf.

Driggers:2012:FCS

[DCHI12] W. B. Driggers III, M. D. Campbell, E. R. Hoffmayer, and G. W. Ingram, Jr. Feeding chronology of six species of carcharhinid sharks in the western North Atlantic Ocean as inferred from longline capture data. Marine Ecology Progress Series, 465:185–192, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/465/m465p185.pdf.

Doubleday:2018:UPS

[DCIG18] Z. A. Doubleday, J. Cliff, C. Izzo, and B. M. Gillanders. Untapping the potential of sulfur isotope analysis in biominerals. Marine Ecology Progress Series, 598:159–166, 2018. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m598p159.pdf.

Davis:2017:PMP

[DCL^{+}17] K. Davis, P. M. Carlson, C. G. Lowe, R. R. Warner, and J. E. Caselle. Parrotfish movement patterns vary with spatiotemporal scale. Marine Ecology Progress Series, 577:149– 164, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/577/m577p149.pdf.

Domingues:2010:GSC

[DCT^{+}10] C. P. Domingues, S. Creer, M. I. Taylor, H. Queiroga, and G. R. Carvalho. Genetic structure of Carcinus maenas within its native range: larval dispersal and oceanographic variability. Marine Ecology Progress Series, 410:111–123, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/410/
m410p111.pdf.

Durant:2010:IFC

[DCW^{+}10] J. M. Durant, R. J. M. Crawford, A. C. Wolfaardt, K. Agenbag, J. Visagie, L. Upfold, and N. C. Stenseth. Influence of feeding conditions on breeding of African penguins— importance of adequate local food supplies. Marine Ecology Progress Series, 420:263–271, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/420/m420p263.pdf.

Duggan:2014:EFF

[DCW^{+}14] M. Duggan, R. M. Connolly, M. Whittle, G. Curwen, and M. A. Burford. Effects of freshwater flow extremes on intertidal biota of a wet-dry tropical estuary. Marine Ecology Progress Series, 502:11–23, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/502/m502p011.pdf.

Darnell:2015:CTG

[DD15] K. M. Darnell and K. H. Dunton. Consumption of turtle grass seeds and seedlings by crabs in the western Gulf of Mexico. Marine Ecology Progress Series, 520:153–163, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/520/
m520p153.pdf.

DiLorenzo:2014:FSN

[DDB^{+}14] M. Di Lorenzo, G. D’Anna, F. Badalamenti, V. M. Giacalone, R. M. Starr, and P. Guidetti. Fitting the size of no-take zones to species movement patterns: a case study on a Mediterranean seabream. Marine Ecology Progress Series, 502:245–255, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/502/m502p245.pdf.

Delgard:2016:BOE

[DDB^{+}16] M. L. Delgard, B. Deflandre, G. Bernard, M. Richard, E. Kochoni, C. Charbonnier, F. Cesbron, E. Metzger, A. Grémare, and P. Anschutz. Benthic oxygen exchange over a heterogeneous Zostera noltei meadow in a temperate coastal ecosystem. Marine Ecology Progress Series, 543:55–71, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/543/m543p055.pdf.

Dubois:2015:RPD

[DDC^{+}15] S. F. Dubois, F. Dérian, X. Caisey, C. Rigolet, J. C. Caprais, and E. Thiébaut. Role of pockmarks in diversity and species assemblages of coastal macrobenthic communities. Marine Ecology Progress Series, 529:91–105, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/529/
m529p091.pdf.

Denis:2011:LRC

[DDD^{+}11] V. Denis, J. Debreuil, S. De Palmas, J. Richard, M. M. M. Guillaume, and J. H. Bruggemann. Lesion regeneration capacities in populations of the massive coral Porites lutea at Réunion Island: environmental correlates. Marine Ecology Progress Series, 428:105–117, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/428/m428p105.pdf.

DeBruyn:2010:AET

[DDDR10] C. De Bruyn, B. David, C. De Ridder, and T. Rigaud. Asymmetric exploitation of two echinoid host species by a parasitic pea crab and its consequences for the parasitic life cycle. Marine Ecology Progress Series, 398:183–191, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/398/
m398p183.pdf.

Delaval:2018:NFC

[DDK^{+}18] A. Delaval, G. Dahle, H. Knutsen, J. Devine, and A. G. V. Salvanes. Norwegian fjords contain sub-populations of roundnose grenadier Coryphaenoides rupestris, a deep-water fish. Marine Ecology Progress Series, 586:181–192, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/586/
m586p181.pdf.

DeBruyn:2016:SSS

[DDM^{+}16] C. De Bruyn, B. David, S. Motreuil, G. Caulier, Q. Jossart, T. Rigaud, and C. De Ridder. Should I stay or should I go? Causes and dynamics of host desertion by a parasitic crab living on echinoids. Marine Ecology Progress Series, 546:163–171, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/546/m546p163.pdf.

Darrenougue:2013:GCR

[DDP^{+}13a] N. Darrenougue, P. De Deckker, C. Payri, S. Eggins, and S. Fallon. Growth and chronology of the rhodolith-forming, coralline red alga Sporolithon durum. Marine Ecology Progress Series, 474:105–119, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/474/m474p105.pdf.

Doropoulos:2013:HCR

[DDP13b] C. Doropoulos and G. Diaz-Pulido. High CO_{2} reduces the settlement of a spawning coral on three common species of crustose coralline algae. Marine Ecology Progress Series, 475: 93–99, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/475/m475p093.pdf.

Dolbeth:2015:RBF

[DDP15] M. Dolbeth, S. Dolédec, and M. Â. Pardal. Relationship between functional diversity and benthic secondary production in a disturbed estuary. Marine Ecology Progress Series, 539:33–46, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/539/m539p033.pdf.

Dixon:2019:STT

[DDP19] H. J. Dixon, J. B. Dempson, and M. Power. Short-term temporal variation in inshore/offshore feeding and trophic niche of Atlantic salmon Salmo salar off West Greenland. Marine Ecology Progress Series, 610:191–203, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/610/
m610p191.pdf.

Davidson:2010:AES

[DdR10] T. M. Davidson and C. E. de Rivera. Accelerated erosion of saltmarshes infested by the non-native burrowing crustacean Sphaeroma quoianum. Marine Ecology Progress Series, 419:129–136, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/419/m419p129.pdf.

Davidson:2014:DAM

[DdRH14] T. M. Davidson, C. E. de Rivera, and H. L. Hsieh. Damage and alteration of mangroves inhabited by a marine wood-borer. Marine Ecology Progress Series, 516:177–185, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/516/
m516p177.pdf.

duDot:2018:TOB

[dDTA^{+}18] T. Jeanniard du Dot, A. W. Trites, J. P. Y. Arnould, J. R. Speakman, and C. Guinet. Trade-offs between foraging efficiency and pup feeding rate of lactating northern fur seals in a declining population. Marine Ecology Progress Series, 600:207–222, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m600p207.pdf.

Jeanniard-du-Dot:2017:CHP

[dDTC^{+}17] T. Jeanniard du Dot, A. C. Thomas, Y. Cherel, A. W. Trites, and C. Guinet. Combining hard-part and DNA analyses of scats with biologging and stable isotopes can reveal different diet compositions and feeding strategies within a fur seal population. Marine Ecology Progress Series, 584:1–16, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Doring:2017:UOE

[DE17] J. Döring and W. Ekau. Using oocyte essential fatty acid composition to assess spawner reproductive potential under hypersaline conditions. Marine Ecology Progress Series, 584: 199–212, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m584p199.pdf.

Dehnhard:2015:LIP

[DED^{+}15] N. Dehnhard, M. Eens, L. Demongin, P. Quillfeldt, D. Suri, and M. Poisbleau. Limited individual phenotypic plasticity in the timing of and investment into egg laying in southern rockhopper penguins under climate change. Marine Ecology Progress Series, 524:269–281, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/524/m524p269.pdf.

Daly-Engel:2010:MMB

[DEGF^{+}10] T. S. Daly-Engel, R. D. Grubbs, K. A. Feldheim, B. W. Bowen, and R. J. Toonen. Is multiple mating beneficial or unavoidable? Low multiple paternity and genetic diversity in the shortspine spurdog Squalus mitsukurii. Marine Ecology Progress Series, 403:255–267, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/403/m403p255.pdf.

Dahnke:2010:SIC

[DEJN10] K. Dähnke, K. Emeis, A. Johannsen, and B. Nagel. Stable isotope composition and turnover of nitrate in the German Bight. Marine Ecology Progress Series, 408:7–18, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/408/
m408p007.pdf.

Drake:2015:RBL

[DEM15] P. T. Drake, C. A. Edwards, and S. G. Morgan. Relationship between larval settlement, alongshore wind stress and surface temperature in a numerical model of the central California coastal circulation. Marine Ecology Progress Series, 537:71–87, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m537p071.pdf.

Dahl:2019:DDC

[DEP19] K. A. Dahl, M. A. Edwards, and W. F. Patterson III. Densitydependent condition and growth of invasive lionfish in the northern Gulf of Mexico. Marine Ecology Progress Series, 623:145–159, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m623p145.pdf.

DeRock:2019:PLS

[DER^{+}19] P. De Rock, S. H. Elwen, J. P. Roux, R. H. Leeney, B. S. James, V. Visser, M. J. Martin, and T. Gridley. Predicting large-scale habitat suitability for cetaceans off Namibia using MinxEnt. Marine Ecology Progress Series, 619:149–167, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/619/
m619p149.pdf.

Dethier:2010:VRD

[Det10] M. N. Dethier. Variation in recruitment does not drive the cline in diversity along an estuarine gradient. Marine Ecology Progress Series, 410:43–54, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/410/m410p043.pdf.

Diaz:2015:DRS

[DEWK15] E. R. Díaz, J. Erlandsson, M. Westerbom, and P. Kraufvelin. Depth-related spatial patterns of sublittoral blue mussel beds and their associated macrofaunal diversity revealed by geostatistical analyses. Marine Ecology Progress Series, 540: 121–134, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/540/m540p121.pdf.

Dernbach:2015:FPW

[DF15] E. M. Dernbach and A. S. Freeman. Foraging preference of whelks Nucella lapillus is robust to influences of wave exposure and predator cues. Marine Ecology Progress Series, 540:135–144, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/540/m540p135.pdf.

Druon:2011:PFS

[DFAH11] J. N. Druon, J. M. Fromentin, F. Aulanier, and J. Heikkonen. Potential feeding and spawning habitats of Atlantic bluefin tuna in the Mediterranean Sea. Marine Ecology Progress Series, 439:223–240, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m439p223.pdf.

Dionisio:2017:ICC

[DFB^{+}17] G. Dionísio, F. Faleiro, M. Bilan, I. C. Rosa, M. Pimentel, J. Serôdio, R. Calado, and R. Rosa. Impact of climate change on the ontogenetic development of ‘solar-powered’ sea slugs. Marine Ecology Progress Series, 578:87–97, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/578/
m578p087.pdf.

Daly:2019:ATR

[DFD^{+}19] R. Daly, J. D. Filmalter, C. A. K. Daly, R. H. Bennett, M. A. M. Pereira, B. Q. Mann, S. W. Dunlop, and P. D. Cowley. Acoustic telemetry reveals multi-seasonal spatiotemporal dynamics of a giant trevally Caranx ignobilis aggregation. Marine Ecology Progress Series, 621:185–197, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/621/
m621p185.pdf.

Dissanayake:2018:EFM

[DFDC18] N. G. Dissanayake, C. L. J. Frid, T. P. Drylie, and B. A. Caswell. Ecological functioning of mudflats: global analysis reveals both regional differences and widespread conservation of functioning. Marine Ecology Progress Series, 604:1–20, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

DeSilva-Davila:2015:CPA

[DFGH^{+}15] R. De Silva-Dávila, C. Franco-Gordo, F. G. Hochberg, E. Godínez-Domínguez, R. Avendaño-Ibarra, J. Gómez-Gutiérrez, and C. J. Robinson. Cephalopod paralarval assemblages in the Gulf of California during 2004–2007. Marine Ecology Progress Series, 520:123–141, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/520/
m520p123.pdf.

DePaula:2017:EEE

[DFLC17] A. F. De Paula, B. G. Fleury, B. G. Lages, and J. C. Creed. Experimental evaluation of the effects of management of invasive corals on native communities. Marine Ecology Progress Series, 572:141–154, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/572/m572p141.pdf.

DiBacco:2011:SBV

[DFPH11] C. DiBacco, H. L. Fuchs, J. Pineda, and K. Helfrich. Swimming behavior and velocities of barnacle cyprids in a downwelling flume. Marine Ecology Progress Series, 433:131– 148, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/433/m433p131.pdf.

DeLecea:2013:PCB

[DFS13] A. M. De Lecea, S. T. Fennessy, and A. J. Smit. Processes controlling the benthic food web of a mesotrophic bight (KwaZulu-Natal, South Africa) revealed by stable isotope analysis. Marine Ecology Progress Series, 484:97–114, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/484/m484p097.pdf.

Drago:2015:SFS

[DFTZ^{+}15] M. Drago, V. Franco-Trecu, L. Zenteno, D. Szteren, E. A. Crespo, F. G. Riet Sapriza, L. de Oliveira, R. Machado, P. Inchausti, and L. Cardona. Sexual foraging segregation in South American sea lions increases during the pre-breeding period in the Río de la plata plume. Marine Ecology Progress Series, 525:261–272, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/525/m525p261.pdf.

deBruyn:2014:POC

[dG14] R. A. J. de Bruyn and L. A. Gosselin. Prevalence of ontogenetic changes in colour brightness among benthic invertebrates and their association with microhabitat shifts. Marine Ecology Progress Series, 498:147–159, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/498/m498p147.pdf.

Davidson:2015:NCP

[DGA^{+}15] A. Davidson, J. N. Griffin, C. Angelini, F. Coleman, R. L. Atkins, and B. R. Silliman. Non-consumptive predator effects intensify grazer-plant interactions by driving vertical habitat shifts. Marine Ecology Progress Series, 537:49–58, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2015/537/m537p049.pdf.

Dias:2012:WDN

[DGC12] M. P. Dias, J. P. Granadeiro, and P. Catry. Working the day or the night shift? Foraging schedules of Cory’s shearwaters vary according to marine habitat. Marine Ecology Progress Series, 467:245–252, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/467/m467p245.pdf.

Dalponti:2018:HHL

[DGC18] G. Dalponti, R. D. Guariento, and A. Caliman. Hunting high or low: body size drives trophic position among and within marine predators. Marine Ecology Progress Series, 597:39–46, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/597/m597p039.pdf.

Regaudie-de-Gioux:2017:IMU

[dGCF^{+}17] A. Regaudie de Gioux, A. Castagna, A. Ferreira, M. Abbrecht, E. S. Braga, and A. M. Ciotti. Influence of mixed upwelled waters on metabolic balance in a subtropical coastal ecosystem: São Sebastião Channel, southern Brazil. Marine Ecology Progress Series, 573:61–72, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/573/m573p061.pdf.

DAmbra:2018:DOB

[DGCH18] I. D’Ambra, W. M. Graham, R. H. Carmichael, and F. J. Hernandez, Jr. Dietary overlap between jellyfish and forage fish in the northern Gulf of Mexico. Marine Ecology Progress Series, 587:31–40, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/587/m587p031.pdf.

Davies:2014:SAS

[DGDJ14] K. T. A. Davies, W. C. Gentleman, C. DiBacco, and C. L. Johnson. Semi-annual spawning in marine scallops strengthens larval recruitment and connectivity on Georges Bank: a model study. Marine Ecology Progress Series, 516:209– 227, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/516/m516p209.pdf.

Dunn:2010:FHN

[DGFH10] M. R. Dunn, L. Griggs, J. Forman, and P. Horn. Feeding habits and niche separation among the deep-sea chimaeroid fishes Harriotta raleighana, Hydrolagus bemisi and Hydrolagus novaezealandiae. Marine Ecology Progress Series, 407: 209–225, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/407/m407p209.pdf.

Davoren:2010:IPB

[DGMB10] G. K. Davoren, S. Garthe, W. A. Montevecchi, and S. Benvenuti. Influence of prey behaviour and other predators on the foraging activities of a marine avian predator in a low Arctic ecosystem. Marine Ecology Progress Series, 404:275– 287, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/404/m404p275.pdf.

Desidera:2019:AFC

[DGP^{+}19] E. Desiderà, P. Guidetti, P. Panzalis, A. Navone, C. A. Valentini-Poirrier, P. Boissery, C. Gervaise, and L. Di Iorio. Acoustic fish communities: sound diversity of rocky habitats reflects fish species diversity. Marine Ecology Progress Series, 608:183–197, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/608/m608p183.pdf.

Dobretsov:2010:ESP

[DGQ10] S. V. Dobretsov, L. Gosselin, and P. Y. Qian. Effects of solar PAR and UV radiation on tropical biofouling communities. Marine Ecology Progress Series, 402:31–43, 2010. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m402p031.pdf.

deRivera:2011:MLT

[dGR11] C. E. deRivera, E. D. Grosholz, and G. M. Ruiz. Multiple and long-term effects of an introduced predatory crab. Marine Ecology Progress Series, 429:145–155, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/429/
m429p145.pdf.

Dell:2016:IDG

[DH16] C. Dell and M. E. Hay. Induced defence to grazing by vertebrate herbivores: uncommon or under-investigated? Marine Ecology Progress Series, 561:137–145, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/561/
m561p137.pdf.

Durant:2017:ASH

[DH17] J. M. Durant and D. Ø. Hjermann. Age-structure, harvesting and climate effects on population growth of arcto-boreal fish stocks. Marine Ecology Progress Series, 577:177–188, 2017.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m577p177.pdf.

Darnaude:2018:VOV

[DH18a] A. M. Darnaude and E. Hunter. Validation of otolith δ 18O values as effective natural tags for shelf-scale geolocation of migrating fish. Marine Ecology Progress Series, 598: 167–185, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m598p167.pdf.

Doerr:2018:SDD

[DH18b] J. C. Doerr and R. L. Hill. Spatial distribution, density, and habitat associations of queen conch Strombus gigas in St. Croix, US Virgin Islands. Marine Ecology Progress Series, 594:119–133, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/594/m594p119.pdf.

Doropoulos:2013:HSI

[DHAV13] C. Doropoulos, G. A. Hyndes, D. Abecasis, and A. Vergés. Herbivores strongly influence algal recruitment in both coraland algal-dominated coral reef habitats. Marine Ecology Progress Series, 486:153–164, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/486/m486p153.pdf.

Duke:2018:LIV

[DHB18] E. M. Duke, A. E. Harada, and R. S. Burton. Large interannual variation in spawning in San Diego Marine Protected Areas captured by molecular identification of fish eggs. Marine Ecology Progress Series, 604:199–210, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/604/
m604p199.pdf.

Durant:2013:EMM

[DHF^{+}13] J. M. Durant, D. Ø. Hjermann, T. Falkenhaug, D. J. Gifford, L. J. Naustvoll, B. K. Sullivan, G. Beaugrand, and N. C. Stenseth. Extension of the match-mismatch hypothesis to predator-controlled systems. Marine Ecology Progress Series, 474:43–52, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/474/m474p043.pdf.

Dumont:2011:ASS

[DHG11] C. P. Dumont, L. G. Harris, and C. F. Gaymer. Anthropogenic structures as a spatial refuge from predation for the invasive bryozoan Bugula neritina. Marine Ecology Progress Series, 427:95–103, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/427/m427p095.pdf.

Danielsen:2016:SCD

[DHG16] N. S. T. Danielsen, R. B. Hedeholm, and P. Grønkjær. Seasonal changes in diet and lipid content of northern sand lance Ammodytes dubius on Fyllas Bank, West Greenland. Marine Ecology Progress Series, 558:97–113, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/558/
m558p097.pdf.

Dong:2015:LVP

[DHGW15] Y. W. Dong, G. D. Han, M. Ganmanee, and J. Wang. Latitudinal variability of physiological responses to heat stress of the intertidal limpet Cellana toreuma along the Asian coast. Marine Ecology Progress Series, 529:107–119, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/529/
m529p107.pdf.

Drif:2010:SAE

[DHH10] K. Drif, A. G. Hirst, and S. Hay. Seasonal abundance and egg production rates of Oithona similis and Pseudocalanus elongatus in the northern North Sea: a first comparison of egg-ratio and incubation methods. Marine Ecology Progress Series, 415:159–175, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/415/m415p159.pdf.

Dupont:2010:EIR

[DHJ10] J. M. Dupont, P. Hallock, and W. C. Jaap. Ecological impacts of the 2005 red tide on artificial reef epibenthic macroinvertebrate and fish communities in the eastern Gulf of Mexico. Marine Ecology Progress Series, 415:189–200, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/415/
m415p189.pdf.

Dauble:2012:SPY

[DHJ12] A. D. Dauble, S. A. Heppell, and M. L. Johansson. Settlement patterns of young-of-the-year rockfish among six Oregon estuaries experiencing different levels of human development. Marine Ecology Progress Series, 448:143–154, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/448/
m448p143.pdf.

Decker:2017:MDR

[DHJ^{+}17] C. Decker, S. Hassani, M. D. Jezequel, C. Rault, C. Dumas, E. Méheust, E. Alfonsi, and J. L. Jung. Mitochondrial DNA reveals historical maternal lineages and a postglacial expansion of the grey seal in European waters. Marine Ecology Progress Series, 566:217–227, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/566/m566p217.pdf.

Davis:2017:DKA

[DHK^{+}17] L. B. Davis, E. E. Hofmann, J. M. Klinck, A. Piñones, and M. S. Dinniman. Distributions of krill and Antarctic silverfish and correlations with environmental variables in the western Ross Sea, Antarctica. Marine Ecology Progress Series, 584:45–65, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m584p045.pdf.

Drenner:2014:VTE

[DHM^{+}14] S. M. Drenner, S. G. Hinch, E. G. Martins, D. Robichaud, T. D. Clark, L. A. Thompson, D. A. Patterson, S. J. Cooke, and R. E. Thomson. Variable thermal experience and diel thermal patterns of homing sockeye salmon in coastal marine waters. Marine Ecology Progress Series, 496:109–124, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/496/m496p109.pdf.

Durant:2013:PGA

[DHR^{+}13] J. M. Durant, M. Hidalgo, T. Rouyer, D. Ø. Hjermann, L. Ciannelli, A. M. Eikeset, N. Yaragina, and N. C. Stenseth. Population growth across heterogeneous environments: effects of harvesting and age structure. Marine Ecology Progress Series, 480:277–287, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m480p277.pdf.

Deakin:2019:SDM

[DHS^{+}19] Z. Deakin, K. C. Hamer, R. B. Sherley, S. Bearhop, T. W. Bodey, B. L. Clark, W. J. Grecian, M. Gummery, J. Lane, G. Morgan, L. Morgan, R. A. Phillips, E. D. Wakefield, and S. C. Votier. Sex differences in migration and demography of a wide-ranging seabird, the northern gannet. Marine Ecology Progress Series, 622:191–201, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m622p191.pdf.

Davenport:2011:DRA

[DHSL11] J. Davenport, R. N. Hughes, M. Shorten, and P. S. Larsen. Drag reduction by air release promotes fast ascent in jumping emperor penguins — a novel hypothesis. Marine Ecology Progress Series, 430:171–182, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m430p171.pdf.

Divine:2015:RBF

[DIB15] L. M. Divine, K. Iken, and B. A. Bluhm. Regional benthic food web structure on the Alaska Beaufort Sea shelf. Marine Ecology Progress Series, 531:15–32, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m531p015.pdf.

DOnghia:2011:DBD

[DIG^{+}11] G. D’Onghia, A. Indennidate, A. Giove, A. Savini, F. Capezzuto, L. Sion, A. Vertino, and P. Maiorano. Distribution and behaviour of deep-sea benthopelagic fauna observed using towed cameras in the Santa Maria di Leuca cold-water coral province. Marine Ecology Progress Series, 443:95–110, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m443p095.pdf.

deJonge:2012:LTM

[dJdBdJB12] V. N. de Jonge, W. F. de Boer, D. J. de Jong, and V. S. Brauer. Long-term mean annual microphytobenthos chlorophyll-a variation correlates with air temperature. Marine Ecology Progress Series, 468:43–56, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m468p043.pdf.

deJuan:2011:RIL

[dJH11] S. de Juan and J. Hewitt. Relative importance of local biotic and environmental factors versus regional factors in driving macrobenthic species richness in intertidal areas. Marine Ecology Progress Series, 423:117–129, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/423/
m423p117.pdf.

Dmitrieva:2016:IVS

[DJJ^{+}16] L. Dmitrieva, M. Jüssi, I. Jüssi, Y. Kasymbekov, M. Verevkin, M. Baimukanov, S. Wilson, and S. J. Goodman. Individual variation in seasonal movements and foraging strategies of a land-locked, ice-breeding pinniped. Marine Ecology Progress Series, 554:241–256, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m554p241.pdf.

Duineveld:2012:STV

[DJL^{+}12] G. C. A. Duineveld, R. M. Jeffreys, M. S. S. Lavaleye, A. J. Davies, M. J. N. Bergman, T. Watmough, and R. Witbaard. Spatial and tidal variation in food supply to shallow coldwater coral reefs of the Mingulay Reef complex (Outer Hebrides, Scotland). Marine Ecology Progress Series, 444:97– 115, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/444/m444p097.pdf.

Degen:2016:PDM

[DJL^{+}16] R. Degen, L. L. Jørgensen, P. Ljubin, I. H. Ellingsen, H. Pehlke, and T. Brey. Patterns and drivers of megabenthic secondary production on the Barents Sea shelf. Marine Ecology Progress Series, 546:1–16, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

deJonge:2019:SFW

[dJSB19] V. N. de Jonge, U. Schückel, and D. Baird. Subsets of food webs cannot be used as a substitute to assess the functioning of entire ecosystems. Marine Ecology Progress Series, 613:49–66, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/613/m613p049.pdf.

deJuan:2014:CDE

[dJTH^{+}14] S. de Juan, S. F. Thrush, J. E. Hewitt, J. Halliday, and A. M. Lohrer. Cumulative degradation in estuaries: contribution of individual species to community recovery. Marine Ecology Progress Series, 510:25–38, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/510/m510p025.pdf.

Dawe:2012:TIB

[DKAC^{+}12] E. G. Dawe, M. Koen-Alonso, D. Chabot, D. Stansbury, and D. Mullowney. Trophic interactions between key predatory fishes and crustaceans: comparison of two Northwest Atlantic systems during a period of ecosystem change. Marine Ecology Progress Series, 469:233–248, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m469p233.pdf.

Dorresteijn:2012:CAF

[DKB^{+}12] I. Dorresteijn, A. S. Kitaysky, C. Barger, Z. M. Benowitz-Fredericks, G. V. Byrd, M. Shultz, and R. Young. Climate affects food availability to planktivorous least auklets Aethia pusilla through physical processes in the southeastern Bering Sea. Marine Ecology Progress Series, 454:207–220, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m454p207.pdf.

Dean:2015:SMC

[DKF^{+}15] B. Dean, H. Kirk, A. Fayet, A. Shoji, R. Freeman, K. Leonard, C. M. Perrins, and T. Guilford. Simultaneous multi-colony tracking of a pelagic seabird reveals cross-colony utilization of a shared foraging area. Marine Ecology Progress Series, 538:239–248, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m538p239.pdf.

Daase:2018:NIB

[DKL^{+}18] M. Daase, K. Kosobokova, K. S. Last, J. H. Cohen, M. Choquet, M. Hatlebakk, and J. E. Søreide. New insights into the biology of Calanus spp. (Copepoda) males in the Arctic. Marine Ecology Progress Series, 607:53–69, 2018. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m607p053.pdf.

Dimech:2012:EEF

[DKRS12] M. Dimech, M. J. Kaiser, S. Ragonese, and P. J. Schembri. Ecosystem effects of fishing on the continental slope in the Central Mediterranean Sea. Marine Ecology Progress Series, 449:41–54, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/449/m449p041.pdf.

DeRyck:2016:DLM

[DKV^{+}16] D. J. R. De Ryck, N. Koedam, T. Van der Stocken, R. M. van der Ven, J. Adams, and L. Triest. Dispersal limitation of the mangrove Avicennia marina at its South African range limit in strong contrast to connectivity in its core East African region. Marine Ecology Progress Series, 545:123– 134, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/545/m545p123.pdf.

Dawson:2013:PRC

[DL13] S. M. Dawson and D. M. Lusseau. Pseudo-replication confounds the assessment of long-distance detection of gillnets by porpoises: Comment on Nielsen et al. (2012). Marine Ecology Progress Series, 478:301–302, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/478/m478p301.pdf. See

[NWH^{+}12].

Dolecal:2014:CMF

[DL14] R. E. Dolecal and J. D. Long. Chemically mediated foraging by subtidal marine predators: a field test of tritrophic cues. Marine Ecology Progress Series, 498:161–171, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/498/
m498p161.pdf.

Gallardo:2019:LEI

[dlÁGRB^{+}19] M. d. l. Á. Gallardo, I. Rojas, K. Brokordt, G. Lovrich, V. Nuñez, K. Paschke, M. Thiel, and B. Yannicelli. Life on the edge: incubation behaviour and physiological performance of squat lobsters in oxygen-minimum conditions. Marine Ecology Progress Series, 623:51–70, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/623/m623p051.pdf.

Decker:2013:LCE

[DLC^{+}13] M. B. Decker, H. Liu, L. Ciannelli, C. Ladd, W. Cheng, and K. S. Chan. Linking changes in eastern Bering Sea jellyfish populations to environmental factors via nonlinear time series models. Marine Ecology Progress Series, 494: 179–189, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/494/m494p179.pdf.

Dairain:2019:IPB

[DLdM19] A. Dairain, A. Legeay, and X. de Montaudouin. Influence of parasitism on bioturbation: from host to ecosystem functioning. Marine Ecology Progress Series, 619:201– 214, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/619/m619p201.pdf.

delaGuardia:2017:SIC

[dlGMD^{+}17] L. Castro de la Guardia, P. G. Myers, A. E. Derocher, N. J. Lunn, and A. D. Terwisscha van Scheltinga. Sea ice cycle in western Hudson Bay, Canada, from a polar bear perspective. Marine Ecology Progress Series, 564:225–233, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/564/
m564p225.pdf.

Deng:2018:EUP

[DLL^{+}18a] X. Deng, T. Liu, C. Y. Liu, S. K. Liang, Y. B. Hu, Y. M. Jin, and X. C. Wang. Effects of Ulva prolifera blooms on the carbonate system in the coastal waters of Qingdao.

Marine Ecology Progress Series, 605:73–86, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/605/
m605p073.pdf.

Dong:2018:SDP

[DLL^{+}18b] Y. Dong, Q. P. Li, Z. Liu, Z. Wu, and W. Zhou. Sizedependent phytoplankton growth and grazing in the northern South China Sea. Marine Ecology Progress Series, 599: 35–47, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/599/m599p035.pdf.

David:2015:CSU

[DLRF15] C. David, B. Lange, B. Rabe, and H. Flores. Community structure of under-ice fauna in the Eurasian central Arctic Ocean in relation to environmental properties of sea-ice habitats. Marine Ecology Progress Series, 522:15–32, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/522/
m522p015.pdf.

delosSantos:2010:ASZ

[dlSBB^{+}10] C. B. de los Santos, F. G. Brun, T. J. Bouma, J. J. Vergara, and J. L. Pérez-Lloréns. Acclimation of seagrass Zostera noltii to co-occurring hydrodynamic and light stresses. Marine Ecology Progress Series, 398:127–135, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/398/
m398p127.pdf.

delosSantos:2012:LFP

[dlSBO^{+}12] C. B. de los Santos, F. G. Brun, Y. Onoda, M. L. Cambridge, T. J. Bouma, J. J. Vergara, and J. L. Pérez-Lloréns. Leaffracture properties correlated with nutritional traits in nine Australian seagrass species: implications for susceptibility to herbivory. Marine Ecology Progress Series, 458:89–102, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m458p089.pdf.

delosSantos:2017:STG

[dlSGS17] C. B. de los Santos, J. A. Godbold, and M. Solan. Shortterm growth and biomechanical responses of the temperate seagrass Cymodocea nodosa to CO_{2} enrichment. Marine Ecology Progress Series, 572:91–102, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m572p091.pdf.

delosSantos:2016:CAM

[dlSOV^{+}16] C. B. de los Santos, Y. Onoda, J. J. Vergara, J. L. Pérez-Lloréns, T. J. Bouma, Y. A. La Nafie, M. L. Cambridge, and F. G. Brun. A comprehensive analysis of mechanical and morphological traits in temperate and tropical seagrass species. Marine Ecology Progress Series, 551:81–94, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/551/
m551p081.pdf.

Defeo:2011:CBM

[DM11] O. Defeo and A. McLachlan. Coupling between macrofauna community structure and beach type: a deconstructive metaanalysis. Marine Ecology Progress Series, 433:29–41, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m433p029.pdf.

Denley:2017:LSS

[DM17] D. Denley and A. Metaxas. Lack of substrate specificity contributes to invasion success and persistence of Membranipora membranacea in the northwest Atlantic. Marine Ecology Progress Series, 580:117–129, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/580/m580p117.pdf.

deMontaudouin:2018:ITO

[dMA18] X. de Montaudouin and C. Accolla. Intraspecific trade-offs between facilitation and competition in the invasive mollusc Crepidula fornicata. Marine Ecology Progress Series, 604:163–171, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/604/m604p163.pdf.

deMontaudouin:2012:ETP

[dMBC12] X. de Montaudouin, H. Bazairi, and S. Culloty. Effect of trematode parasites on cockle Cerastoderma edule growth and condition index: a transplant experiment. Marine Ecology Progress Series, 471:111–121, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/471/m471p111.pdf.

Dias:2016:ECU

[DMC^{+}16] E. Dias, P. Morais, A. M. Cotter, C. Antunes, and J. C. Hoffman. Estuarine consumers utilize marine, estuarine and terrestrial organic matter and provide connectivity among these food webs. Marine Ecology Progress Series, 554:21– 34, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/554/m554p021.pdf.

DeCesare:2017:DPB

[DMC^{+}17] S. De Cesare, T. Meziane, L. Chauvaud, J. Richard, M. K. Sejr, J. Thébault, G. Winkler, and F. Olivier. Dietary plasticity in the bivalve astarte moerchi revealed by a multimarker study in two Arctic fjords. Marine Ecology Progress Series, 567:157–172, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/567/m567p157.pdf.

Daverat:2012:PEF

[DMD^{+}12] F. Daverat, P. Morais, E. Dias, J. Babaluk, J. Martin, M. Eon, R. Fablet, C. Pécheyran, and C. Antunes. Plasticity of European flounder life history patterns discloses alternatives to catadromy. Marine Ecology Progress Series, 465:267–280, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/465/m465p267.pdf.

Daigle:2014:BSP

[DMd14] R. M. Daigle, A. Metaxas, and B. deYoung. Bay-scale patterns in the distribution, aggregation and spatial variability of larvae of benthic invertebrates. Marine Ecology Progress Series, 503:139–156, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/503/m503p139.pdf.

Dell:2015:EMP

[DMH15] C. Dell, J. P. Montoya, and M. E. Hay. Effect of marine protected areas (MPAs) on consumer diet: MPA fish feed higher in the food chain. Marine Ecology Progress Series, 540:227–234, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/540/m540p227.pdf.

Drake:2014:HUP

[DMHS^{+}14] P. Drake, E. Marco-Herrero, M. D. Subida, A. M. Arias, and J. A. Cuesta. Host use pattern of the pea crab Afropinnotheres monodi: potential effects on its reproductive success and geographical expansion. Marine Ecology Progress Series, 498:203–215, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/498/m498p203.pdf.

Drouin:2011:HAD

[DMJ11] A. Drouin, C. W. McKindsey, and L. E. Johnson. Higher abundance and diversity in faunal assemblages with the invasion of Codium fragile ssp. fragile in eelgrass meadows. Marine Ecology Progress Series, 424:105–117, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/424/
m424p105.pdf.

Davenport:2011:CSA

[DMK11] J. Davenport, T. V. Moloney, and J. Kelly. Common sea anemones Actinia equina are predominantly sessile intertidal scavengers. Marine Ecology Progress Series, 430:147– 155, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m430p147.pdf.

deMontaudouin:2011:DPT

[dML11] X. de Montaudouin and L. Lanceleur. Distribution of parasites in their second intermediate host, the cockle Cerastoderma edule: community heterogeneity and spatial scale. Marine Ecology Progress Series, 428:187–199, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/428/
m428p187.pdf.

Donelson:2010:EEW

[DMM^{+}10] J. M. Donelson, P. L. Munday, M. I. McCormick, N. W. Pankhurst, and P. M. Pankhurst. Effects of elevated water temperature and food availability on the reproductive performance of a coral reef fish. Marine Ecology Progress Series, 401:233–243, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/401/m401p233.pdf.

Dawe:2012:ETS

[DMMW12] E. G. Dawe, D. R. Mullowney, M. Moriyasu, and E. Wade. Effects of temperature on size-at-terminal molt and molting frequency in snow crab Chionoecetes opilio from two Canadian Atlantic ecosystems. Marine Ecology Progress Series, 469:279–296, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m469p279.pdf.

Dawe:2012:TSE

[DMP12] Earl G. Dawe, Franz Mueter, and Ólafu K. Pálsson. Theme section: Effects of climate and predation on subarctic crustacean populations. Marine Ecology Progress Series, 469: 191–306, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m469_ThemeSection.pdf.

Diaz:2011:DTL

[DMPG11] D. Díaz, S. Mallol, A. M. Parma, and R. Goñi. Decadal trend in lobster reproductive output from a temperate marine protected area. Marine Ecology Progress Series, 433: 149–157, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/433/m433p149.pdf.

Doblin:2010:TTS

[DMR10] M. A. Doblin, K. R. Murphy, and G. M. Ruiz. Thresholds for tracing ships’ ballast water: an Australian case study. Marine Ecology Progress Series, 408:19–32, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/408/
m408p019.pdf.

Davis:2014:BCR

[DMS14] B. Davis, C. Mattone, and M. Sheaves. Bottom-up control regulates patterns of fish connectivity and assemblage structure in coastal wetlands. Marine Ecology Progress Series, 500:175–186, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2014/500/m500p175.pdf. See erra-

tum [Ano14c].

Deppe:2014:DPF

[DMT^{+}14] L. Deppe, K. F. McGregor, F. Tomasetto, J. V. Briskie, and R. P. Scofield. Distribution and predictability of foraging areas in breeding Chatham albatrosses Thalassarche eremita in relation to environmental characteristics. Marine Ecology Progress Series, 498:287–301, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2014/498/m498p287.pdf.

Dutz:2010:LCS

[DMvB10] J. Dutz, V. Mohrholz, and J. E. E. van Beusekom. Life cycle and spring phenology of Temora longicornis in the Baltic Sea. Marine Ecology Progress Series, 406:223–238, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/406/
m406p223.pdf.

Davidson:2016:ECC

[DMW^{+}16] A. T. Davidson, J. McKinlay, K. Westwood, P. G. Thomson, R. van den Enden, M. de Salas, S. Wright, R. Johnson, and K. Berry. Enhanced CO_{2} concentrations change the structure of Antarctic marine microbial communities. Marine Ecology Progress Series, 552:93–113, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/552/m552p093.pdf.

Dillon:2017:DDT

[DNO17] E. M. Dillon, R. D. Norris, and A. O’Dea. Dermal denticles as a tool to reconstruct shark communities. Marine Ecology Progress Series, 566:117–134, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/566/m566p117.pdf.

Delacy:2017:AAS

[DOH^{+}17] C. R. Delacy, A. Olsen, L. A. Howey, D. D. Chapman, E. J. Brooks, and M. E. Bond. Affordable and accurate stereovideo system for measuring dimensions underwater: a case study using oceanic whitetip sharks Carcharhinus longimanus. Marine Ecology Progress Series, 574:75–84, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/574/
m574p075.pdf.

Dawson:2011:ISG

[DP11] J. Dawson and D. Pillay. Influence of starfish grazing on lagoonal microalgal communities: non-competitive mechanisms for unimodal effects on diversity. Marine Ecology Progress Series, 435:75–82, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/435/m435p075.pdf.

Diederich:2013:TTC

[DP13] C. M. Diederich and J. A. Pechenik. Thermal tolerance of Crepidula fornicata (Gastropoda) life history stages from intertidal and subtidal subpopulations. Marine Ecology Progress Series, 486:173–187, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/486/m486p173.pdf.

DeBose:2014:CSM

[DP14a] J. L. DeBose and V. J. Paul. Chemical signatures of multispecies foraging aggregations are attractive to fish. Marine Ecology Progress Series, 498:243–248, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/498/
m498p243.pdf.

Dudley:2014:UEM

[DP14b] P. N. Dudley and W. P. Porter. Using empirical and mechanistic models to assess global warming threats to leatherback sea turtles. Marine Ecology Progress Series, 501:265–278, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2014/501/m501p265.pdf.

Diogo:2016:FDI

[DP16] H. Diogo and J. G. Pereira. Fishing in the dark: the importance of integrating a nocturnal component into recreational fishing surveys. Marine Ecology Progress Series, 542: 187–193, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/542/m542p187.pdf.

Dyer:2013:PFD

[DPC13] D. C. Dyer, R. Perissinotto, and N. K. Carrasco. Postflood dietary variation in the Mozambique tilapia Oreochromis mossambicus in the St Lucia Estuary, South Africa. Marine Ecology Progress Series, 476:199–214, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/476/
m476p199.pdf.

Druon:2012:PFH

[DPD^{+}12] J. N. Druon, S. Panigada, L. David, A. Gannier, P. Mayol, A. Arcangeli, A. Cañadas, S. Laran, N. Di Méglio, and P. Gauffier. Potential feeding habitat of fin whales in the western Mediterranean Sea: an environmental niche model. Marine Ecology Progress Series, 464:289–306, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m464p289.pdf.

Drew:2013:ECT

[DPH13] G. S. Drew, J. F. Piatt, and D. F. Hill. Effects of currents and tides on fine-scale use of marine bird habitats in a Southeast Alaska hotspot. Marine Ecology Progress Series, 487:275–286, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m487p275.pdf.

Downs:2014:NIF

[DPH14] E. E. Downs, B. N. Popp, and C. M. Holl. Nitrogen isotope fractionation and amino acid turnover rates in the Pacific white shrimp Litopenaeus vannamei. Marine Ecology Progress Series, 516:239–250, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2014/516/m516p239.pdf.

Dillon:2015:FEC

[DPM15] K. S. Dillon, M. S. Peterson, and C. A. May. Functional equivalence of constructed and natural intertidal eastern oyster reef habitats in a northern Gulf of Mexico estuary. Marine Ecology Progress Series, 528:187–203, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/528/m528p187.pdf.

Du:2015:IVP

[DPO15] X. Du, W. Peterson, and L. O’Higgins. Interannual variations in phytoplankton community structure in the northern California Current during the upwelling seasons of 2001–2010. Marine Ecology Progress Series, 519:75–87, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/519/
m519p075.pdf.

Davis:2017:SCT

[DPO^{+}17] J. P. Davis, K. A. Pitt, A. D. Olds, A. R. Harborne, and R. M. Connolly. Seagrass corridors and tidal state modify how fish use habitats on intertidal coral reef flats. Marine Ecology Progress Series, 581:135–147, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/581/m581p135.pdf.

Dahl:2016:EAL

[DPS16] K. A. Dahl, W. F. Patterson III, and R. A. Snyder. Experimental assessment of lionfish removals to mitigate reef fish community shifts on northern Gulf of Mexico artificial reefs. Marine Ecology Progress Series, 558:207–221, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m558p207.pdf.

Daniels:2016:SSC

[DPY^{+}16] C. J. Daniels, A. J. Poulton, J. R. Young, M. Esposito, M. P. Humphreys, M. Ribas-Ribas, E. Tynan, and T. Tyrrell. Species-specific calcite production reveals Coccolithus pelagicus as the key calcifier in the Arctic Ocean. Marine Ecology Progress Series, 555:29–47, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m555p029.pdf.

DiFranco:2013:PVE

[DQC^{+}13] A. Di Franco, K. Qian, A. Calò, M. Di Lorenzo, S. Planes, and P. Guidetti. Patterns of variability in early life traits of a Mediterranean coastal fish. Marine Ecology Progress Series, 476:227–235, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/476/m476p227.pdf.

DiFiore:2019:GHC

[DQM^{+}19] B. P. DiFiore, S. A. Queenborough, E. M. P. Madin, V. J. Paul, M. B. Decker, and A. C. Stier. Grazing halos on coral reefs: predation risk, herbivore density, and habitat size influence grazing patterns that are visible from space. Marine Ecology Progress Series, 627:71–81, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/627/m627p071.pdf.

deQuevedo:2013:MRC

[dQSC13] I. Álvarez de Quevedo, M. San Félix, and L. Cardona. Mortality rates in by-caught loggerhead turtle Caretta caretta in the Mediterranean Sea and implications for the Atlantic populations. Marine Ecology Progress Series, 489:225–234, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/489/m489p225.pdf. See comment [BBRM14] and

reply [dQSC14].

deQuevedo:2014:TTC

[dQSC14] I. Álvarez de Quevedo, M. San Félix, and L. Cardona. Temporal trends in the by-catch of loggerhead turtles Caretta caretta in the Mediterranean Sea: Reply to Báez et al. (2014). Marine Ecology Progress Series, 504:303–304, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/504/

m504p303.pdf. See [dQSC13, BBRM14].

Damare:2010:ASP

[DR10] V. Damare and S. Raghukumar. Association of the stramenopilan protists, the aplanochytrids, with zooplankton of the equatorial Indian Ocean. Marine Ecology Progress Series, 399:53–68, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/399/m399p053.pdf.

Dance:2016:SSV

[DR16] M. A. Dance and J. R. Rooker. Stage-specific variability in habitat associations of juvenile red drum across a latitudinal gradient. Marine Ecology Progress Series, 557:221– 235, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/557/m557p221.pdf.

Decker:2018:JFF

[DRD^{+}18] M. B. Decker, K. L. Robinson, S. Dorji, K. D. Cieciel, C. Barceló, J. J. Ruzicka, and R. D. Brodeur. Jellyfish and forage fish spatial overlap on the eastern Bering Sea shelf during periods of high and low jellyfish biomass. Marine Ecology Progress Series, 591:57–69, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m591p057.pdf.

Duarte:2015:EAF

[DRDV15] L. Duarte, F. Rossi, C. Docal, and R. M. Viejo. Effects of alga Fucus serratus decline on benthic assemblages and trophic linkages at its retreating southern range edge. Marine Ecology Progress Series, 527:87–103, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/527/m527p087.pdf.

Degen:2012:CST

[DRG^{+}12] R. Degen, L. Riavitz, S. Gollner, A. Vanreusel, C. Plum, and M. Bright. Community study of tubeworm-associated epizooic meiobenthos from deep-sea cold seeps and hot vents. Marine Ecology Progress Series, 468:135–148, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/468/
m468p135.pdf.

Driscoll:2015:TDG

[DRH15] R. M. Driscoll, C. S. Reiss, and B. T. Hentschel. Temperaturedependent growth of Thysanoessa macrura: inter-annual and spatial variability around Elephant Island, Antarctica. Marine Ecology Progress Series, 529:49–61, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/529/
m529p049.pdf.

Davies:2013:TSC

[DRT13] K. T. A. Davies, T. Ross, and C. T. Taggart. Tidal and subtidal currents affect deep aggregations of right whale prey, Calanus spp., along a shelf-basin margin. Marine Ecology Progress Series, 479:263–282, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/479/m479p263.pdf.

Dorado:2012:IBS

[DRWQ12] S. Dorado, J. R. Rooker, B. Wissel, and A. Quigg. Isotope baseline shifts in pelagic food webs of the Gulf of Mexico. Marine Ecology Progress Series, 464:37–49, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m464p037.pdf.

Demopoulos:2010:IMA

[DS10] A. W. J. Demopoulos and C. R. Smith. Invasive mangroves alter macrofaunal community structure and facilitate opportunistic exotics. Marine Ecology Progress Series, 404: 51–67, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/404/m404p051.pdf.

Durant:2013:TSH

[DS13] J. M. Durant and N. C. Stenseth. Theme section: Harvested fish stocks in a changing environment. Marine Ecology Progress Series, 480:199–301, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m480_ThemeSection.pdf.

Dijkstra:2016:IRC

[DS16] J. A. Dijkstra and C. Simkanin. Intraspecific response of colonial ascidians to variable salinity stress in an era of global change. Marine Ecology Progress Series, 551:215– 225, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/551/m551p215.pdf.

DAlessandro:2013:SMD

[DSC13] E. K. D’Alessandro, S. Sponaugle, and R. K. Cowen. Selective mortality during the larval and juvenile stages of snappers (Lutjanidae) and great barracuda Sphyraena barracuda. Marine Ecology Progress Series, 474:227–242, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/474/
m474p227.pdf.

daSilva:2011:STR

[dSdSdCO^{+}11] A. C. C. D. da Silva, E. A. P. dos Santos, F. L. d. C. Oliveira, M. I. Weber, J. A. F. Batista, T. Z. Serafini, and J. C. de Castilhos. Satellite-tracking reveals multiple foraging strategies and threats for olive ridley turtles in Brazil. Marine Ecology Progress Series, 443:237–247, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/443/
m443p237.pdf.

Dethier:2013:AAV

[DSG^{+}13] M. N. Dethier, E. Sosik, A. W. E. Galloway, D. O. Duggins, and C. A. Simenstad. Addressing assumptions: variation in stable isotopes and fatty acids of marine macrophytes can confound conclusions of food web studies. Marine Ecology Progress Series, 478:1–14, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Daewel:2015:PPE

[DSG15] U. Daewel, C. Schrum, and A. K. Gupta. The predictive potential of early life stage individual-based models (IBMs): an example for Atlantic cod Gadus morhua in the North Sea. Marine Ecology Progress Series, 534:199–219, 2015. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m534p199.pdf.

Dorman:2015:MKA

[DSGR^{+}15] J. G. Dorman, W. J. Sydeman, M. García-Reyes, R. A. Zeno, and J. A. Santora. Modeling krill aggregations in the centralnorthern California Current. Marine Ecology Progress Series, 528:87–99, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/528/m528p087.pdf.

Damerau:2014:PGS

[DSH14] M. Damerau, W. Salzburger, and R. Hanel. Population genetic structure of Lepidonotothen larseni revisited: cyb and microsatellites suggest limited connectivity in the Southern Ocean. Marine Ecology Progress Series, 517:251–263, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2014/517/m517p251.pdf.

Desmond:2018:ECS

[DSJNH18] M. J. Desmond, R. Suárez-Jiménez, W. A. Nelson, and C. D. Hepburn. Epifaunal community structure within southern New Zealand kelp forests. Marine Ecology Progress Series, 596:71–81, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/596/m596p071.pdf.

Dunweber:2010:SFS

[DSK^{+}10] M. Dünweber, R. Swalethorp, S. Kjellerup, T. G. Nielsen, K. E. Arendt, M. Hjorth, K. Tönnesson, and E. F. Møller. Succession and fate of the spring diatom bloom in Disko Bay, western Greenland. Marine Ecology Progress Series, 419: 11–29, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/419/m419p011.pdf.

Flores-de-Santiago:2012:SCL

[dSKFV12] F. Flores de Santiago, J. M. Kovacs, and F. Flores-Verdugo. Seasonal changes in leaf chlorophyll-a content and morphology in a sub-tropical mangrove forest of the Mexican Pacific. Marine Ecology Progress Series, 444:57–68, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m444p057.pdf.

Daase:2011:EFQ

[DSM11a] M. Daase, J. E. Søreide, and D. Martynova. Effects of food quality on naupliar development in Calanus glacialis at subzero temperatures. Marine Ecology Progress Series, 429:111–124, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/429/m429p111.pdf.

dosSantos:2011:PTP

[dSM11b] G. A. P. dos Santos and T. Moens. Populations of two prey nematodes and their interaction are controlled by a predatory nematode. Marine Ecology Progress Series, 427:117– 131, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/427/m427p117.pdf.

DAlessandro:2010:LES

[DSS10] E. D’Alessandro, S. Sponaugle, and J. E. Serafy. Larval ecology of a suite of snappers (family: Lutjanidae) in the Straits of Florida, western Atlantic Ocean. Marine Ecology Progress Series, 410:159–175, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/410/m410p159.pdf.

DeSalvo:2010:TRH

[DSVM10] M. K. DeSalvo, S. Sunagawa, C. R. Voolstra, and M. Medina. Transcriptomic responses to heat stress and bleaching in the elkhorn coral Acropora palmata. Marine Ecology Progress Series, 402:97–113, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/402/m402p097.pdf.

DuPreez:2011:STR

[DT11] C. Du Preez and V. Tunnicliffe. Shortspine thornyhead and rockfish (Scorpaenidae) distribution in response to substratum, biogenic structures and trawling. Marine Ecology Progress Series, 425:217–231, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m425p217.pdf.

DiBacco:2015:RPL

[DT15] C. DiBacco and T. W. Therriault. Reproductive periodicity and larval vertical migration behavior of European green crab Carcinus maenas in a non-native habitat. Marine Ecology Progress Series, 536:123–134, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/536/m536p123.pdf.

Dahne:2017:BCA

[DTC^{+}17] M. Dähne, J. Tougaard, J. Carstensen, A. Rose, and J. Nabe-Nielsen. Bubble curtains attenuate noise from offshore wind farm construction and reduce temporary habitat loss for harbour porpoises. Marine Ecology Progress Series, 580:221– 237, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m580p221.pdf.

Davidson:2016:EED

[DTG16] M. I. Davidson, T. E. Targett, and P. A. Grecay. Evaluating the effects of diel-cycling hypoxia and pH on growth and survival of juvenile summer flounder Paralichthys dentatus. Marine Ecology Progress Series, 556:223–235, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/556/
m556p223.pdf.

Dunphy:2015:CSD

[DTL^{+}15] B. J. Dunphy, G. A. Taylor, T. J. Landers, R. L. Sagar, B. L. Chilvers, L. Ranjard, and M. J. Rayner. Comparative seabird diving physiology: first measures of haematological parameters and oxygen stores in three New Zealand procellariiformes. Marine Ecology Progress Series, 523:187–198, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/523/
m523p187.pdf.

Dinasquet:2012:CEC

[DTM^{+}12] J. Dinasquet, J. Titelman, L. F. Møller, O. Setälä, L. Granhag, T. Andersen, U. Båmstedt, M. Haraldsson, A. Hosia, T. Katajisto, T. Kragh, J. Kuparinen, M. L. Schrøter, M. Søndergaard, P. Tiselius, and L. Riemann. Cascading effects of the ctenophore Mnemiopsis leidyi on the planktonic food web in a nutrient-limited estuarine system. Marine Ecology Progress Series, 460:49–61, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/460/m460p049.pdf.

Davies:2014:WMS

[DTS14] K. T. A. Davies, C. T. Taggart, and R. K. Smedbol. Water mass structure defines the diapausing copepod distribution in a right whale habitat on the Scotian Shelf. Marine Ecology Progress Series, 497:69–85, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/497/m497p069.pdf.

DeLaRocha:2010:RRN

[DTVH10] C. L. De La Rocha, A. Terbrüggen, C. Völker, and S. Hohn. Response to and recovery from nitrogen and silicon starvation in Thalassiosira weissflogii: growth rates, nutrient uptake and C, Si and N content per cell. Marine Ecology Progress Series, 412:57–68, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/412/m412p057.pdf.

Denis:2018:GFT

[DVA^{+}18] J. Denis, D. Vincent, E. Antajan, C. Vallet, J. Mestre, V. Lefebvre, J. Caboche, R. Cordier, P. Marchal, and C. Loots. Gut fluorescence technique to quantify pigment feeding in downs herring larvae. Marine Ecology Progress Series, 607: 129–142, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/607/m607p129.pdf.

Dolbeth:2016:EEF

[DVB^{+}16] M. Dolbeth, A. L. Vendel, A. Baeta, A. Pessanha, and J. Patrício. Exploring ecosystem functioning in two Brazilian estuaries integrating fish diversity, species traits and food webs. Marine Ecology Progress Series, 560:41–55, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/560/m560p041.pdf.

Dutz:2012:SDF

[DvBH12] J. Dutz, J. E. E. van Beusekom, and R. Hinrichs. Seasonal dynamics of fecundity and recruitment of Temora longicornis in the Baltic Sea. Marine Ecology Progress Series, 462: 51–66, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/462/m462p051.pdf.

Diez:2013:WNS

[DVS13] B. Díez, L. Van Nieuwerburgh, and P. Snoeijs. Water nutrient stoichiometry modifies the nutritional quality of phytoplankton and somatic growth of crustacean mesozooplankton. Marine Ecology Progress Series, 489:93–105, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/489/
m489p093.pdf.

DeLaender:2010:CTH

[DVSM10] F. De Laender, D. Van Oevelen, K. Soetaert, and J. J. Middelburg. Carbon transfer in herbivore- and microbial loopdominated pelagic food webs in the southern Barents Sea during spring and summer. Marine Ecology Progress Series, 398:93–107, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/398/m398p093.pdf.

DeSmet:2016:LCS

[DvV^{+}16] B. De Smet, D. van Oevelen, M. Vincx, J. Vanaverbeke, and K. Soetaert. Lanice conchilega structures carbon flows in soft-bottom intertidal areas. Marine Ecology Progress Series, 552:47–60, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m552p047.pdf.

DeMeester:2018:ASS

[DVV^{+}18] N. De Meester, T. Van Daele, J. Van Malderen, L. Monteiro, C. Van Colen, S. Derycke, and T. Moens. Active and species-specific dispersal behaviour in a marine nematode cryptic species complex. Marine Ecology Progress Series, 600:71–83, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/600/m600p071.pdf.

Davis:2012:NPF

[DW12] J. P. Davis and S. R. Wing. Niche partitioning in the fiordland wrasse guild. Marine Ecology Progress Series, 446:207– 220, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/446/m446p207.pdf.

Dong:2018:IOO

[DWD^{+}18] J. Dong, B. Wang, Y. Duan, W. D. Yoon, A. Wang, X. Liu, Y. Li, M. Sun, and Y. Chai. Initial occurrence, ontogenic distribution-shifts and advection of Nemopilema nomurai (Scyphozoa: Rhizostomeae) in Liaodong Bay, China, from 2005–2015. Marine Ecology Progress Series, 591:185– 197, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/591/m591p185.pdf.

DeMartini:2013:PSP

[DWK13] E. E. DeMartini, J. L. K. Wren, and D. R. Kobayashi. Persistent spatial patterns of recruitment in a guild of Hawaiian coral reef fishes. Marine Ecology Progress Series, 485: 165–179, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/485/m485p165.pdf.

Durland:2019:CLD

[DWL19] E. Durland, G. Waldbusser, and C. Langdon. Comparison of larval development in domesticated and naturalized stocks of the Pacific oyster Crassostrea gigas exposed to high pCO_{2} conditions. Marine Ecology Progress Series, 621:107–125, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m621p107.pdf.

Davies:2013:DDF

[DWLH13] R. D. Davies, S. Wanless, S. Lewis, and K. C. Hamer. Density-dependent foraging and colony growth in a pelagic seabird species under varying environmental conditions. Marine Ecology Progress Series, 485:287–294, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/485/
m485p287.pdf.

Dornhoffer:2015:MLI

[DWM15] T. M. Dornhoffer, G. G. Waldbusser, and C. Meile. Modeling lugworm irrigation behavior effects on sediment nitrogen cycling. Marine Ecology Progress Series, 534:121– 134, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/534/m534p121.pdf.

Dale:2011:NHU

[DWPH11] J. J. Dale, N. J. Wallsgrove, B. N. Popp, and K. N. Holland. Nursery habitat use and foraging ecology of the brown stingray Dasyatis lata determined from stomach contents, bulk and amino acid stable isotopes. Marine Ecology Progress Series, 433:221–236, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/433/m433p221.pdf.

Duckworth:2012:EWT

[DWV^{+}12] A. R. Duckworth, L. West, T. Vansach, A. Stubler, and M. Hardt. Effects of water temperature and pH on growth and metabolite biosynthesis of coral reef sponges. Marine Ecology Progress Series, 462:67–77, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/462/m462p067.pdf.

DAloia:2018:LDE

[DXF^{+}18] C. C. D’Aloia, A. Xuereb, M. J. Fortin, S. M. Bogdanowicz, and P. M. Buston. Limited dispersal explains the spatial distribution of siblings in a reef fish population. Marine Ecology Progress Series, 607:143–154, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/607/m607p143.pdf.

Diop:2018:FET

[DZB^{+}18] N. Diop, L. Zango, A. Beard, C. T. Ba, P. I. Ndiaye, L. Henry, E. Clingham, S. Oppel, and J. González-Solís. Foraging ecology of tropicbirds breeding in two contrasting marine environments in the tropical Atlantic. Marine Ecology Progress Series, 607:221–236, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/607/m607p221.pdf.

Dierssen:2010:BES

[DZDB10] H. M. Dierssen, R. C. Zimmerman, L. A. Drake, and D. Burdige. Benthic ecology from space: optics and net primary production in seagrass and benthic algae across the Great Bahama Bank. Marine Ecology Progress Series, 411:1–15, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Daniels:2011:SHB

[DZH^{+}11] C. A. Daniels, A. Zeifman, K. Heym, K. B. Ritchie, C. A. Watson, I. Berzins, and M. Breitbart. Spatial heterogeneity of bacterial communities in the mucus of Montastraea annularis.

Marine Ecology Progress Series, 426:29–40, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/426/
m426p029.pdf.

Dunn:2016:BBE

[DZL16] P. H. Dunn, V. Zarulli, and D. A. Levitis. Beyond being eaten or swept away: ontogenetic transitions drive developmental mortality in marine barnacle larvae. Marine Ecology Progress Series, 559:103–116, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/559/m559p103.pdf.

Edmunds:2019:WMC

[EAB^{+}19] P. J. Edmunds, T. C. Adam, A. C. Baker, S. S. Doo, P. W. Glynn, D. P. Manzello, N. J. Silbiger, T. B. Smith, and P. Fong. Why more comparative approaches are required in time-series analyses of coral reef ecosystems. Marine Ecology Progress Series, 608:297–306, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m608p297.pdf.

Elsdon:2010:EES

[EAMT10] T. S. Elsdon, S. Ayvazian, K. W. McMahon, and S. R. Thorrold. Experimental evaluation of stable isotope fractionation in fish muscle and otoliths. Marine Ecology Progress Series, 408:195–205, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/408/m408p195.pdf.

Eriksson:2010:PTS

[EBB10] A. C. Eriksson, D. J. Booth, and P. A. Biro. ‘Personality’ in two species of temperate damselfish. Marine Ecology Progress Series, 420:273–276, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/420/m420p273.pdf.

Eriksen:2018:CDI

[EBDB18] E. Eriksen, B. Bogstad, A. Dolgov, and I. M. Beck. Cod diet as an indicator of Ctenophora abundance dynamics in the Barents Sea. Marine Ecology Progress Series, 591:87– 100, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/591/m591p087.pdf.

Eriksson:2012:SCA

[EBdlTC12] H. Eriksson, M. Byrne, and M. de la Torre-Castro. Sea cucumber (Aspidochirotida) community, distribution and habitat utilization on the reefs of Mayotte, Western Indian Ocean. Marine Ecology Progress Series, 452:159–170, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/452/
m452p159.pdf.

Ebert:2010:DPP

[Ebe10] T. A. Ebert. Demographic patterns of the purple sea urchin Strongylocentrotus purpuratus along a latitudinal gradient, 1985–1987. Marine Ecology Progress Series, 406:105–120, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2010/406/m406p105.pdf.

Ekeroth:2016:NFR

[EBH16] N. Ekeroth, S. Blomqvist, and P. O. J. Hall. Nutrient fluxes from reduced Baltic Sea sediment: effects of oxygenation and macrobenthos. Marine Ecology Progress Series, 544:77–92, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m544p077.pdf.

Ellis:2017:WSI

[EBH17] T. A. Ellis, J. A. Buckel, and J. E. Hightower. Winter severity influences spotted seatrout mortality in a southeast US estuarine system. Marine Ecology Progress Series, 564:145– 161, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/564/m564p145.pdf.

Engel:2012:EBL

[EBK^{+}12] M. Engel, A. Behnke, J. Klier, C. Buschbaum, N. Volkenborn, and T. Stoeck. Effects of the bioturbating lugworm Arenicola marina on the structure of benthic protistan communities. Marine Ecology Progress Series, 471:87–99, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/471/
m471p087.pdf.

Eastman:2011:ARL

[EBOM11] J. T. Eastman, E. Barrera-Oro, and E. Moreira. Adaptive radiation at a low taxonomic level: divergence in buoyancy of the ecologically similar Antarctic fish Notothenia coriiceps and N. rossii. Marine Ecology Progress Series, 438: 195–206, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/438/m438p195.pdf.

Elschot:2015:EEL

[EBT^{+}15] K. Elschot, J. P. Bakker, S. Temmerman, J. van de Koppel, and T. J. Bouma. Ecosystem engineering by large grazers enhances carbon stocks in a tidal salt marsh. Marine Ecology Progress Series, 537:9–21, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/537/m537p009.pdf.

Eierman:2014:FBP

[EC14] L. E. Eierman and R. C. Connor. Foraging behavior, prey distribution, and microhabitat use by bottlenose dolphins Tursiops truncatus in a tropical atoll. Marine Ecology Progress Series, 503:279–288, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2014/503/m503p279.pdf.

Escalle:2015:CTP

[ECC^{+}15] L. Escalle, A. Capietto, P. Chavance, L. Dubroca, A. Delgado De Molina, H. Murua, D. Gaertner, E. Romanov, J. Spitz, J. J. Kiszka, L. Floch, A. Damiano, and B. Merigot. Cetaceans and tuna purse seine fisheries in the Atlantic and Indian Oceans: interactions but few mortalities. Marine Ecology Progress Series, 522:255–268, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/522/m522p255.pdf.

Eder:2012:FDL

[ECM^{+}12] E. Eder, A. Ceballos, S. Martins, H. Pérez-García, I. Marín, A. Marco, and L. Cardona. Foraging dichotomy in loggerhead sea turtles Caretta caretta off northwestern Africa. Marine Ecology Progress Series, 470:113–122, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/470/
m470p113.pdf.

Espinasse:2014:DZH

[ECZD14] B. Espinasse, F. Carlotti, M. Zhou, and J. L. Devenon. Defining zooplankton habitats in the Gulf of Lion (NW Mediterranean Sea) using size structure and environmental conditions. Marine Ecology Progress Series, 506:31–46, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/506/
m506p031.pdf.

Edgell:2012:CSS

[ED12] T. C. Edgell and M. W. Demarchi. California and Steller sea lion use of a major winter haulout in the Salish Sea over 45 years. Marine Ecology Progress Series, 467:253– 262, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/467/m467p253.pdf.

Edmunds:2010:PBP

[Edm10] P. J. Edmunds. Population biology of Porites astreoides and Diploria strigosa on a shallow Caribbean reef. Marine Ecology Progress Series, 418:87–104, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m418p087.pdf.

Edmunds:2013:DSC

[Edm13] P. J. Edmunds. Decadal-scale changes in the community structure of coral reefs of St. John, US Virgin Islands. Marine Ecology Progress Series, 489:107–123, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m489p107.pdf.

Edmunds:2014:LSV

[Edm14] P. J. Edmunds. Landscape-scale variation in coral reef community structure in the United States Virgin Islands. Marine Ecology Progress Series, 509:137–152, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m509p137.pdf.

Elahi:2014:MVO

[EDS14] R. Elahi, T. R. Dwyer, and K. P. Sebens. Mesoscale variability in oceanographic retention sets the abiotic stage for subtidal benthic diversity. Marine Ecology Progress Series, 498:117–132, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/498/m498p117.pdf.

Evensen:2015:EPS

[EES15] N. R. Evensen, P. J. Edmunds, and K. Sakai. Effects of pCO_{2} on spatial competition between the corals Montipora aequituberculata and Porites lutea. Marine Ecology Progress Series, 541:123–134, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m541p123.pdf.

Ellis:2016:NGI

[EF16] R. D. Ellis and M. E. Faletti. Native grouper indirectly ameliorates the negative effects of invasive lionfish. Marine Ecology Progress Series, 558:267–279, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m558p267.pdf.

Escobar-Flores:2013:ACP

[EFOM13] P. Escobar-Flores, R. L. O’Driscoll, and J. C. Montgomery. Acoustic characterization of pelagic fish distribution across the South Pacific Ocean. Marine Ecology Progress Series, 490:169–183, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/490/m490p169.pdf.

Escobar-Flores:2018:PDR

[EFOM18] P. C. Escobar-Flores, R. L. O’Driscoll, and J. C. Montgomery. Predicting distribution and relative abundance of mid-trophic level organisms using oceanographic parameters and acoustic backscatter. Marine Ecology Progress Series, 592: 37–56, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/592/m592p037.pdf.

Ellis:2012:DDW

[EFR12] J. I. Ellis, G. Fraser, and J. Russell. Discharged drilling waste from oil and gas platforms and its effects on benthic communities. Marine Ecology Progress Series, 456:285– 302, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/456/m456p285.pdf.

Edwards:2015:DSM

[EGH^{+}15] A. J. Edwards, J. R. Guest, A. J. Heyward, R. D. Villanueva, M. V. Baria, I. S. F. Bollozos, and Y. Golbuu. Direct seeding of mass-cultured coral larvae is not an effective option for reef rehabilitation. Marine Ecology Progress Series, 525:105–116, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/525/m525p105.pdf.

Eaglesham:2013:WDC

[EH13] J. B. Eaglesham and I. Hewson. Widespread detection of circular replication initiator protein (rep)-encoding ssDNA viral genomes in estuarine, coastal and open ocean net plankton. Marine Ecology Progress Series, 494:65–72, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/494/
m494p065.pdf.

Ellis:2017:PGS

[EHD^{+}17] C. D. Ellis, D. J. Hodgson, C. L. Daniels, M. Collins, and A. G. F. Griffiths. Population genetic structure in European lobsters: implications for connectivity, diversity and hatchery stocking. Marine Ecology Progress Series, 563:123–137, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m563p123.pdf.

Eggertsen:2016:ETC

[EHG16] L. Eggertsen, L. Hammar, and M. Gullström. Effects of tidal current-induced flow on reef fish behaviour and function on a subtropical rocky reef. Marine Ecology Progress Series, 559:175–192, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/559/m559p175.pdf.

Ebert:2012:OCB

[EHR12] T. A. Ebert, J. C. Hernández, and M. P. Russell. Ocean conditions and bottom-up modifications of gonad development in the sea urchin Strongylocentrotus purpuratus over space and time. Marine Ecology Progress Series, 467:147– 166, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/467/m467p147.pdf.

Eads:2016:IEO

[EKE16] A. R. Eads, W. J. Kennington, and J. P. Evans. Interactive effects of ocean warming and acidification on sperm motility and fertilization in the mussel Mytilus galloprovincialis. Marine Ecology Progress Series, 562:101–111, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/562/
m562p101.pdf.

Evans:2013:FBC

[EKOÅ13] T. J. Evans, M. Kadin, O. Olsson, and S. Åkesson. Foraging behaviour of common murres in the Baltic Sea, recorded by simultaneous attachment of GPS and time-depth recorder devices. Marine Ecology Progress Series, 475:277–289, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/475/m475p277.pdf.

Engelsma:2010:EBO

[EKR^{+}10] M. Y. Engelsma, S. Kerkhoff, I. Roozenburg, O. L. M. Haenen, A. van Gool, W. Sistermans, S. Wijnhoven, and H. Hummel. Epidemiology of Bonamia ostreae infecting European flat oysters Ostrea edulis from Lake Grevelingen, The Netherlands. Marine Ecology Progress Series, 409:131–142, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/409/
m409p131.pdf.

Edmunds:2016:CRS

[EL16] P. J. Edmunds and H. R. Lasker. Cryptic regime shift in benthic community structure on shallow reefs in St. John, US Virgin Islands. Marine Ecology Progress Series, 559:1–12, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Edmunds:2019:RPS

[EL19] P. J. Edmunds and H. R. Lasker. Regulation of population size of arborescent octocorals on shallow Caribbean reefs. Marine Ecology Progress Series, 615:1–14, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Edmunds:2010:LSV

[ELA10] P. J. Edmunds, J. J. Leichter, and M. Adjeroud. Landscapescale variation in coral recruitment in Moorea, French Polynesia. Marine Ecology Progress Series, 414:75–89, 2010. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m414p075.pdf.

Ezzat:2019:SFI

[ELM^{+}19] L. Ezzat, T. Lamy, R. L. Maher, K. S. Munsterman, K. Landfield, E. R. Schmeltzer, C. A. Gaulke, D. E. Burkepile, and R. Vega Thurber. Surgeonfish feces increase microbial opportunism in reef-building corals. Marine Ecology Progress Series, 631:81–97, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/631/m631p081.pdf.

Eero:2011:ETS

[EM11] M. Eero and B. R. MacKenzie. Extending time series of fish biomasses using a simple surplus production-based approach. Marine Ecology Progress Series, 440:191–202, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/440/
m440p191.pdf.

Eggers:2013:IDC

[EM13] S. L. Eggers and B. Matthiessen. Initial dominance in coccolithophore communities affects community structure but does not translate into altered community functioning. Marine Ecology Progress Series, 473:67–77, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m473p067.pdf.

Elliott:2015:SPM

[EMG^{+}15] A. Elliott, C. J. Mundy, M. Gosselin, M. Poulin, K. Campbell, and F. Wang. Spring production of mycosporine-like amino acids and other UV-absorbing compounds in sea ice-associated algae communities in the Canadian Arctic. Marine Ecology Progress Series, 541:91–104, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m541p091.pdf.

Espinoza:2019:MTS

[EMH^{+}19] M. Espinoza, J. Matley, M. R. Heupel, A. J. Tobin, A. T. Fisk, and C. A. Simpfendorfer. Multi-tissue stable isotope analysis reveals resource partitioning and trophic relationships of large reef-associated predators. Marine Ecology Progress Series, 615:159–176, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2019/615/m615p159.pdf.

Eurich:2018:HSA

[EMJ18] J. G. Eurich, M. I. McCormick, and G. P. Jones. Habitat selection and aggression as determinants of fine-scale partitioning of coral reef zones in a guild of territorial damselfishes. Marine Ecology Progress Series, 587:201–215, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/587/
m587p201.pdf.

Eerkes-Medrano:2013:CEH

[EMMSL13] D. Eerkes-Medrano, B. A. Menge, C. Sislak, and C. J. Langdon. Contrasting effects of hypoxic conditions on survivorship of planktonic larvae of rocky intertidal invertebrates. Marine Ecology Progress Series, 478:139–151, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/478/
m478p139.pdf.

Elliott:2012:YLD

[EMTB^{+}12] K. H. Elliott, L. McFarlane-Tranquilla, C. M. Burke, A. Hedd, W. A. Montevecchi, and W. G. Anderson. Year-long deployments of small geolocators increase corticosterone levels in murres. Marine Ecology Progress Series, 466:1–7, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Escobar:2011:RRV

[EN11] J. B. Escobar and S. A. Navarrete. Risk recognition and variability in escape responses among intertidal molluskan grazers to the sun star Heliaster helianthus. Marine Ecology Progress Series, 421:151–161, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/421/m421p151.pdf.

Eleftheriou:2013:MHR

[EN13] Anastasios Eleftheriou and Thomas Noji. In memoriam: Heye Rumohr (1947–2013). Marine Ecology Progress Series, ??:??, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m494i.pdf.

Enge:2011:RBF

[ENO^{+}11] A. J. Enge, H. Nomaki, N. O. Ogawa, U. Witte, M. M. Moeseneder, G. Lavik, N. Ohkouchi, H. Kitazato, M. Kučera, and P. Heinz. Response of the benthic Foraminiferal community to a simulated short-term phytodetritus pulse in the abyssal North Pacific. Marine Ecology Progress Series, 438: 129–142, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m438p129.pdf.

Engel:2014:NDE

[ENW^{+}14] A. Engel, C. Cisternas Novoa, M. Wurst, S. Endres, T. Tang, M. Schartau, and C. Lee. No detectable effect of CO_{2} on elemental stoichiometry of Emiliania huxleyi in nutrient-limited, acclimated continuous cultures. Marine Ecology Progress Series, 507:15–30, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m507p015.pdf.

Erga:2015:GDV

[EOA15] S. R. Erga, C. D. Olseng, and L. H. Aarø. Growth and diel vertical migration patterns of the toxic dinoflagellate Protoceratium reticulatum in a water column with salinity stratification: the role of bioconvection and light. Marine Ecology Progress Series, 539:47–64, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m539p047.pdf.

Ebrahim:2014:HSS

[EOM^{+}14] A. Ebrahim, A. D. Olds, P. S. Maxwell, K. A. Pitt, D. D. Burfeind, and R. M. Connolly. Herbivory in a subtropical seagrass ecosystem: separating the functional role of different grazers. Marine Ecology Progress Series, 511:83–91, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/511/
m511p083.pdf.

Eggers:2015:IHT

[EOMS15] F. Eggers, E. Moland Olsen, E. Moland, and A. Slotte. Individual habitat transitions of Atlantic herring Clupea harengus in a human-modified coastal system. Marine Ecology Progress Series, 520:245–256, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m520p245.pdf.

Eguiguren:2019:HUC

[EPC^{+}19] A. Eguiguren, E. Pirotta, M. Cantor, L. Rendell, and H. Whitehead. Habitat use of culturally distinct Galápagos sperm whale Physeter macrocephalus clans. Marine Ecology Progress Series, 609:257–270, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/609/m609p257.pdf.

Every:2017:NMS

[EPF^{+}17] S. L. Every, H. R. Pethybridge, C. J. Fulton, P. M. Kyne, and D. A. Crook. Niche metrics suggest euryhaline and coastal elasmobranchs provide trophic connections among marine and freshwater biomes in northern Australia. Marine Ecology Progress Series, 565:181–196, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/565/m565p181.pdf.

Edmunds:2014:LTC

[EPL^{+}14] P. J. Edmunds, X. Pochon, D. R. Levitan, D. M. Yost, M. Belcaid, H. M. Putnam, and R. D. Gates. Long-term changes in Symbiodinium communities in Orbicella annularis in St. John, US Virgin Islands. Marine Ecology Progress Series, 506:129–144, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m506p129.pdf.

Eddy:2016:DIL

[EPM^{+}16] C. Eddy, J. Pitt, J. A. Morris, Jr., S. Smith, G. Goodbody-Gringley, and D. Bernal. Diet of invasive lionfish (Pterois volitans and P. miles) in Bermuda. Marine Ecology Progress Series, 558:193–206, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m558p193.pdf.

Elias-Piera:2013:TES

[EPRGO13] F. Elias-Piera, S. Rossi, J. M. Gili, and C. Orejas. Trophic ecology of seven Antarctic gorgonian species. Marine Ecology Progress Series, 477:93–106, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/477/m477p093.pdf.

Emmerson:2011:RCE

[EPS11] L. Emmerson, R. Pike, and C. Southwell. Reproductive consequences of environment-driven variation in Adélie penguin breeding phenology. Marine Ecology Progress Series, 440:203–216, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/440/m440p203.pdf.

Erikstad:2013:SFI

[ERB^{+}13a] K. E. Erikstad, T. K. Reiertsen, R. T. Barrett, F. Vikebø, and H. Sandvik. Seabird–fish interactions: the fall and rise of a common guillemot Uria aalge population. Marine Ecology Progress Series, 475:267–276, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/475/m475p267.pdf. See

erratum [ERB^{+}13b].

Erikstad:2013:ESF

[ERB^{+}13b] Kjell Einar Erikstad, Tone Kristin Reiertsen, Robert T. Barrett, Frode Vikebø, and Hanno Sandvik. Erratum: Seabird– fish interactions: the fall and rise of a common guillemot Uria aalge population. Marine Ecology Progress Series, ??: ??, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/481/m481p305.pdf. See [ERB^{+}13a].

Eliasen:2011:SLB

[ERG^{+}11] K. Eliasen, J. Reinert, E. Gaard, B. Hansen, J. A. Jacobsen, P. Grønkjær, and J. T. Christensen. Sandeel as a link between primary production and higher trophic levels on the Faroe shelf. Marine Ecology Progress Series, 438: 185–194, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/438/m438p185.pdf.

Edwards:2012:TVL

[ES12a] K. F. Edwards and J. J. Stachowicz. Temporally varying larval settlement, competition, and coexistence in a sessile invertebrate community. Marine Ecology Progress Series, 462:93–102, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/462/m462p093.pdf.

Elahi:2012:CMN

[ES12b] R. Elahi and K. P. Sebens. Consumers mediate natural variation between prey richness and resource use in a benthic marine community. Marine Ecology Progress Series, 452:131–143, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/452/m452p131.pdf.

Easson:2014:CEA

[ESBG14] C. G. Easson, M. Slattery, D. M. Baker, and D. J. Gochfeld. Complex ecological associations: competition and facilitation in a sponge-algal interaction. Marine Ecology Progress Series, 507:153–167, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/507/m507p153.pdf.

Elahi:2016:OWD

[ESD16] R. Elahi, K. P. Sebens, and G. A. De Leo. Ocean warming and the demography of declines in coral body size. Marine Ecology Progress Series, 560:147–158, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/560/m560p147.pdf.

Evrard:2010:CNF

[ESH^{+}10] V. Evrard, K. Soetaert, C. H. R. Heip, M. Huettel, M. A. Xenopoulos, and J. J. Middelburg. Carbon and nitrogen flows through the benthic food web of a photic subtidal sandy sediment. Marine Ecology Progress Series, 416:1–16, 2010. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Eurich:2014:CSL

[ESW14] J. G. Eurich, R. L. Selden, and R. R. Warner. California spiny lobster preference for urchins from kelp forests: implications for urchin barren persistence. Marine Ecology Progress Series, 498:217–225, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/498/m498p217.pdf.

Elliott:2011:ICA

[ET11] D. T. Elliott and K. W. Tang. Influence of carcass abundance on estimates of mortality and assessment of population dynamics in Acartia tonsa. Marine Ecology Progress Series, 427:1–12, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Erler:2013:DAN

[ETAE13] D. V. Erler, L. A. Trott, D. M. Alongi, and B. D. Eyre. Denitrification, anammox and nitrate reduction in sediments of the southern Great Barrier Reef lagoon. Marine Ecology Progress Series, 478:57–70, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/478/m478p057.pdf.

Enochs:2011:EDM

[ETB^{+}11] I. C. Enochs, L. T. Toth, V. W. Brandtneris, J. C. Afflerbach, and D. P. Manzello. Environmental determinants of motile cryptofauna on an eastern Pacific coral reef. Marine Ecology Progress Series, 438:105–118, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/438/m438p105.pdf.

Eriksson:2013:PMP

[ETB13] H. Eriksson, B. V. Thorne, and M. Byrne. Population metrics in protected commercial sea cucumber populations (curryfish: Stichopus herrmanni) on One Tree Reef, Great Barrier Reef. Marine Ecology Progress Series, 473:225–234, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/473/
m473p225.pdf.

Eriksen:2018:SSP

[ETD^{+}18] R. Eriksen, T. W. Trull, D. Davies, P. Jansen, A. T. Davidson, K. Westwood, and R. van den Enden. Seasonal succession of phytoplankton community structure from autonomous sampling at the Australian Southern Ocean time series (SOTS) observatory. Marine Ecology Progress Series, 589:13–31, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/589/m589p013.pdf.

Elliott:2017:JGH

[ETHB17] S. A. M. Elliott, W. R. Turrell, M. R. Heath, and D. M. Bailey. Juvenile gadoid habitat and ontogenetic shift observations using stereo-video baited cameras. Marine Ecology Progress Series, 568:123–135, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m568p123.pdf.

Espeland:2010:DVM

[ETO^{+}10] S. H. Espeland, A. G. Thoresen, E. M. Olsen, L. C. Stige, H. Knutsen, J. Gjøsæter, and N. C. Stenseth. Diel vertical migration patterns in juvenile cod from the Skagerrak coast. Marine Ecology Progress Series, 405:29–37, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/405/
m405p029.pdf.

Eble:2011:EPL

[ETS^{+}11] J. A. Eble, R. J. Toonen, L. Sorenson, L. V. Basch, Y. P. Papastamatiou, and B. W. Bowen. Escaping paradise: larval export from Hawaii in an Indo-Pacific reef fish, the yellow tang Zebrasoma flavescens. Marine Ecology Progress Series, 428:245–258, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/428/m428p245.pdf.

Eberlein:2016:IEO

[EVB^{+}16] T. Eberlein, D. B. Van de Waal, K. M. Brandenburg, U. John, M. Voss, E. P. Achterberg, and B. Rost. Interactive effects of ocean acidification and nitrogen limitation on two bloomforming dinoflagellate species. Marine Ecology Progress Series, 543:127–140, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m543p127.pdf.

Eigaard:2014:PPE

[EvDB^{+}14] O. R. Eigaard, M. van Deurs, J. W. Behrens, D. Bekkevold, K. Brander, M. Plambech, K. Schreiber Plet-Hansen, and H. Mosegaard. Prey or predator—expanding the food web role of sandeel Ammodytes marinus. Marine Ecology Progress Series, 516:267–273, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m516p267.pdf.

Espinel-Velasco:2018:EOA

[EVHA^{+}18] N. Espinel-Velasco, L. Hoffmann, A. Agüera, M. Byrne, S. Dupont, S. Uthicke, N. S. Webster, and M. Lamare. Effects of ocean acidification on the settlement and metamorphosis of marine invertebrate and fish larvae: a review. Marine Ecology Progress Series, 606:237–257, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/606/m606p237.pdf.

Easter:2016:SMP

[EW16] E. E. Easter and J. W. White. Spatial management for protogynous sex-changing fishes: a general framework for coastal systems. Marine Ecology Progress Series, 543:223–240, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m543p223.pdf.

Espinasse:2012:AFW

[EZZ^{+}12] B. Espinasse, M. Zhou, Y. Zhu, E. L. Hazen, A. S. Friedlaender, D. P. Nowacek, D. Chu, and F. Carlotti. Austral fall– winter transition of mesozooplankton assemblages and krill aggregations in an embayment west of the Antarctic Peninsula. Marine Ecology Progress Series, 452:63–80, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/452/
m452p063.pdf.

Floyd:2010:IEN

[FA10a] E. Y. Floyd and T. W. Anderson. Interactive effects of nutritional condition and refuge availability on survival of a temperate reef goby. Marine Ecology Progress Series, 407: 257–269, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/407/m407p257.pdf.

Fuentes:2010:EPC

[FA10b] M. M. P. B. Fuentes and D. Abbs. Effects of projected changes in tropical cyclone frequency on sea turtles. Marine Ecology Progress Series, 412:283–292, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/412/m412p283.pdf.

Farmer:2011:GSM

[FA11] N. A. Farmer and J. S. Ault. Grouper and snapper movements and habitat use in Dry Tortugas, Florida. Marine Ecology Progress Series, 433:169–184, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/433/m433p169.pdf.

Flowers:2016:RBP

[FABH^{+}16] K. I. Flowers, M. J. Ajemian, K. Bassos-Hull, K. A. Feldheim, R. E. Hueter, Y. P. Papastamatiou, and D. D. Chapman. A review of batoid philopatry, with implications for future research and population management. Marine Ecology Progress Series, 562:251–261, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/562/m562p251.pdf.

Flores:2012:ICC

[FAK^{+}12] H. Flores, A. Atkinson, S. Kawaguchi, B. A. Krafft, G. Milinevsky, S. Nicol, C. Reiss, G. A. Tarling, R. Werner, E. Bravo Rebolledo, V. Cirelli, J. Cuzin-Roudy, S. Fielding, J. A. van Franeker, J. J. Groeneveld, M. Haraldsson, A. Lombana, E. Marschoff, B. Meyer, E. A. Pakhomov, A. P. Van de Putte, E. Rombolá, K. Schmidt, V. Siegel, M. Teschke, H. Tonkes, J. Y. Toullec, P. N. Trathan, N. Tremblay, and T. Werner. Impact of climate change on Antarctic krill. Marine Ecology Progress Series, 458:1–19, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Faust:2017:ORE

[FAM^{+}17] E. Faust, C. André, S. Meurling, J. Kochmann, H. Christiansen, L. F. Jensen, G. Charrier, A. T. Laugen, and Å. Strand. Origin and route of establishment of the invasive Pacific oyster Crassostrea gigas in Scandinavia. Marine Ecology Progress Series, 575:95–105, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m575p095.pdf.

Fernandez:2015:AEE

[FAT^{+}15] Á. Fernández, F. Arenas, A. Trilla, S. Rodríguez, L. Rueda, and B. Martínez. Additive effects of emersion stressors on the ecophysiological performance of two intertidal seaweeds. Marine Ecology Progress Series, 536:135–147, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/536/
m536p135.pdf.

Fukunaga:2010:ICE

[FAWBF10] A. Fukunaga, M. J. Anderson, J. G. Webster-Brown, and R. B. Ford. Individual and combined effects of heavy metals on estuarine infaunal communities. Marine Ecology Progress Series, 402:123–136, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/402/m402p123.pdf.

Frojan:2014:MPC

[FAZ^{+}14] M. Froján, B. Arbones, D. Zúñiga, C. G. Castro, and F. G. Figueiras. Microbial plankton community in the Ría de Vigo (NW Iberian upwelling system): impact of the culture of Mytilus galloprovincialis. Marine Ecology Progress Series, 498:43–54, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/498/m498p043.pdf.

Fuentes:2014:MER

[FB14] K. Fuentes and A. Brante. Megadisturbances and ecological recovery in the intertidal zone: the role of interspecific competition between bioengineer mussel species. Marine Ecology Progress Series, 515:161–172, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/515/m515p161.pdf.

Felix:2011:DBH

[FBA11] F. Félix and N. Botero-Acosta. Distribution and behaviour of humpback whale mother-calf pairs during the breeding season off Ecuador. Marine Ecology Progress Series, 426:277– 287, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/426/m426p277.pdf.

Filgueira:2015:WSI

[FBC^{+}15] R. Filgueira, C. J. Byron, L. A. Comeau, B. Costa-Pierce, P. J. Cranford, J. G. Ferreira, J. Grant, T. Guyondet, H. M. Jansen, T. Landry, C. W. McKindsey, J. K. Petersen, G. K. Reid, S. M. C. Robinson, A. Smaal, R. Sonier, Ø. Strand, and T. Strohmeier. As we see it: An integrated ecosystem approach for assessing the potential role of cultivated bivalve shells as part of the carbon trading system. Marine Ecology Progress Series, 518:281–287, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m518p281.pdf.

Freeman:2015:SSM

[FBET15] C. J. Freeman, D. M. Baker, C. G. Easson, and R. W. Thacker. Shifts in sponge-microbe mutualisms across an experimental irradiance gradient. Marine Ecology Progress Series, 526:41–53, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/526/m526p041.pdf.

Feary:2018:TPO

[FBGH18] D. A. Feary, A. G. Bauman, J. Guest, and A. S. Hoey. Trophic plasticity in an obligate corallivorous butterflyfish. Marine Ecology Progress Series, 605:165–171, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/605/
m605p165.pdf.

Fraschetti:2012:IEP

[FBGT12] S. Fraschetti, S. Bevilacqua, G. Guarnieri, and A. Terlizzi. Idiosyncratic effects of protection in a remote marine reserve. Marine Ecology Progress Series, 466:21–34, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m466p021.pdf.

Fulweiler:2013:ECM

[FBNJ13] R. W. Fulweiler, S. M. Brown, S. W. Nixon, and B. D. Jenkins. Evidence and a conceptual model for the co-occurrence of nitrogen fixation and denitrification in heterotrophic marine sediments. Marine Ecology Progress Series, 482:57– 68, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/482/m482p057.pdf.

Follana-Berna:2019:EDR

[FBPCC^{+}19] G. Follana-Berná, M. Palmer, A. Campos-Candela, P. Arechavala-Lopez, C. Diaz-Gil, J. Alós, I. A. Catalan, S. Balle, J. Coll, G. Morey, F. Verger, and A. Grau. Estimating the density of resident coastal fish using underwater cameras: accounting for individual detectability. Marine Ecology Progress Series, 615:177–188, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/615/m615p177.pdf.

Feagans-Bartow:2014:EOR

[FBS14] J. N. Feagans-Bartow and T. T. Sutton. Ecology of the oceanic rim: pelagic eels as key ecosystem components. Marine Ecology Progress Series, 502:257–266, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/502/
m502p257.pdf.

Fanelli:2010:TVF

[FC10] E. Fanelli and J. E. Cartes. Temporal variations in the feeding habits and trophic levels of three deep-sea demersal fishes from the western Mediterranean Sea, based on stomach contents and stable isotope analyses. Marine Ecology Progress Series, 402:213–232, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/402/m402p213.pdf.

Fitch:2012:CEI

[FC12] J. E. Fitch and T. P. Crowe. Combined effects of inorganic nutrients and organic enrichment on intertidal benthic macrofauna: an experimental approach. Marine Ecology Progress Series, 461:59–70, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/461/m461p059.pdf.

Fontanez:2013:PRH

[FC13] K. M. Fontanez and C. M. Cavanaugh. Phylogenetic relationships of hydrothermal vent mussels (Bathymodiolinae) and their symbionts. Marine Ecology Progress Series, 474: 147–154, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/474/m474p147.pdf.

Fraser:2019:THS

[FC19] C. M. L. Fraser and B. K. K. Chan. Too hot for sex: mating behaviour and fitness in the intertidal barnacle Fistulobalanus albicostatus under extreme heat stress. Marine Ecology Progress Series, 610:99–108, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/610/m610p099.pdf.

Forrestal:2012:EEB

[FCDC12] F. Forrestal, M. Coll, D. J. Die, and V. Christensen. Ecosystem effects of bluefin tuna Thunnus thynnus thynnus aquaculture in the NW Mediterranean Sea. Marine Ecology Progress Series, 456:215–231, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/456/m456p215.pdf.

Filmalter:2015:FSD

[FCFD15] J. Filmalter, P. Cowley, F. Forget, and L. Dagorn. Finescale 3-dimensional movement behaviour of silky sharks Carcharhinus falciformis associated with fish aggregating devices (FADs). Marine Ecology Progress Series, 539:207–223, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m539p207.pdf.

Fort:2010:GSV

[FCH^{+}10] J. Fort, Y. Cherel, A. M. A. Harding, J. Welcker, D. Jakubas, H. Steen, N. J. Karnovsky, and D. Grémillet. Geographic and seasonal variability in the isotopic niche of little auks. Marine Ecology Progress Series, 414:293–302, 2010. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m414p293.pdf.

Fassatoui:2012:RBH

[FCR12] C. Fassatoui, A. Chenuil, and M. S. Romdhane. Relationships between heterozygosity, growth parameters and age in the common pandora Pagellus erythrinus (Sparidae) in the Gabès Gulf (Tunisia). Marine Ecology Progress Series, 445: 251–261, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/445/m445p251.pdf.

Falkenberg:2014:HME

[FCR14] L. J. Falkenberg, S. D. Connell, and B. D. Russell. Herbivory mediates the expansion of an algal habitat under nutrient and CO_{2} enrichment. Marine Ecology Progress Series, 497:87–92, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/497/m497p087.pdf.

Fleeger:2015:RSM

[FCR^{+}15] J. W. Fleeger, K. R. Carman, M. R. Riggio, I. A. Mendelssohn, Q. X. Lin, A. Hou, D. R. Deis, and S. Zengel. Recovery of salt marsh benthic microalgae and meiofauna following the Deepwater Horizon oil spill linked to recovery of Spartina alterniflora. Marine Ecology Progress Series, 536: 39–54, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/536/m536p039.pdf.

Freitas:2018:FBJ

[FCRD18] C. Freitas, R. Caldeira, J. Reis, and T. Dellinger. Foraging behavior of juvenile loggerhead sea turtles in the open ocean: from Lévy exploration to area-restricted search. Marine Ecology Progress Series, 595:203–215, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/595/
m595p203.pdf.

Fraser:2015:OKG

[FCS15] C. M. L. Fraser, R. A. Coleman, and F. Seebacher. Orientation in a keystone grazer: interactions between habitat and individual identity drive patterns of resting behaviour. Marine Ecology Progress Series, 522:145–156, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/522/
m522p145.pdf.

Frodella:2016:PCP

[FCV^{+}16] N. Frodella, R. Cannas, A. Velonà, P. Carbonara, E. D. Farrell, F. Fiorentino, M. C. Follesa, G. Garofalo, F. Hemida, C. Mancusi, M. Stagioni, N. Ungaro, F. Serena, F. Tinti, and A. Cariani. Population connectivity and phylogeography of the Mediterranean endemic skate Raja polystigma and evidence of its hybridization with the parapatric sibling R. montagui. Marine Ecology Progress Series, 554:99–113, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/554/
m554p099.pdf.

Foveau:2013:DPB

[FDDD13] A. Foveau, N. Desroy, J. C. Dauvin, and J. M. Dewarumez. Distribution patterns in the benthic diversity of the eastern English Channel. Marine Ecology Progress Series, 479: 115–126, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/479/m479p115.pdf.

Filbee-Dexter:2016:LSD

[FDFS16] K. Filbee-Dexter, C. J. Feehan, and R. E. Scheibling. Largescale degradation of a kelp ecosystem in an ocean warming hotspot. Marine Ecology Progress Series, 543:141–152, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/543/m543p141.pdf.

Ferguson:2010:BWB

[FDLL10] S. H. Ferguson, L. Dueck, L. L. Loseto, and S. P. Luque. Bowhead whale Balaena mysticetus seasonal selection of sea ice. Marine Ecology Progress Series, 411:285–297, 2010. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m411p285.pdf.

Foo:2016:ACS

[FDP^{+}16] S. A. Foo, S. A. Dworjanyn, A. G. B. Poore, J. Harianto, and M. Byrne. Adaptive capacity of the sea urchin Heliocidaris erythrogramma to ocean change stressors: responses from gamete performance to the juvenile. Marine Ecology Progress Series, 556:161–172, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/556/m556p161.pdf.

Filbee-Dexter:2012:HMD

[FDS12] K. Filbee-Dexter and R. E. Scheibling. Hurricane-mediated defoliation of kelp beds and pulsed delivery of kelp detritus to offshore sedimentary habitats. Marine Ecology Progress Series, 455:51–64, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/455/m455p051.pdf.

Filbee-Dexter:2014:SUB

[FDS14] K. Filbee-Dexter and R. E. Scheibling. Sea urchin barrens as alternative stable states of collapsed kelp ecosystems. Marine Ecology Progress Series, 495:1–25, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Farber:2018:IEO

[FDVL18] L. Färber, J. M. Durant, Y. Vindenes, and Ø. Langangen. Increased early offspring growth can offset the costs of longdistance spawning migration in fish. Marine Ecology Progress Series, 600:141–150, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m600p141.pdf.

Fisher:2015:HWE

[FEA^{+}15] B. Fisher, A. M. Ellis, D. K. Adams, H. E. Fox, and E. R. Selig. Health, wealth, and education: the socioeconomic backdrop for marine conservation in the developing world. Marine Ecology Progress Series, 530:233–242, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m530p233.pdf.

Frederiksen:2011:MLT

[FEE^{+}11] M. Frederiksen, D. A. Elston, M. Edwards, A. D. Mann, and S. Wanless. Mechanisms of long-term decline in size of lesser sandeels in the North Sea explored using a growth and phenology model. Marine Ecology Progress Series, 432: 137–147, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/432/m432p137.pdf.

Fox:2014:PHC

[FESPR14] C. H. Fox, R. El-Sabaawi, P. C. Paquet, and T. E. Reimchen. Pacific herring Clupea pallasii and wrack macrophytes subsidize semi-terrestrial detritivores. Marine Ecology Progress Series, 495:49–64, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/495/m495p049.pdf.

Fuchs:2010:PCP

[FF10] H. L. Fuchs and P. J. S. Franks. Plankton community properties determined by nutrients and size-selective feeding. Marine Ecology Progress Series, 413:1–15, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Fernandez:2012:ARI

[FF12] C. Fernandez and L. Farías. Assimilation and regeneration of inorganic nitrogen in a coastal upwelling system: ammonium and nitrate utilization. Marine Ecology Progress Series, 451: 1–14, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Freeman:2016:ROE

[FF16] L. A. Freeman and S. E. Freeman. Rapidly obtained ecosystem indicators from coral reef soundscapes. Marine Ecology Progress Series, 561:69–82, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m561p069.pdf.

Fisher:2010:GVM

[FFL10] J. A. D. Fisher, K. T. Frank, and W. C. Leggett. Global variation in marine fish body size and its role in biodiversityecosystem functioning. Marine Ecology Progress Series, 405: 1–13, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Figurski:2016:SSD

[FFLS16] J. D. Figurski, J. Freiwald, S. I. Lonhart, and C. D. Storlazzi. Seasonal sediment dynamics shape temperate bedrock reef communities. Marine Ecology Progress Series, 552:19– 29, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m552p019.pdf.

Fung:2013:WSS

[FFS^{+}13] T. Fung, K. D. Farnsworth, S. Shephard, D. G. Reid, and A. G. Rossberg. Why the size structure of marine communities can require decades to recover from fishing. Marine Ecology Progress Series, 484:155–171, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/484/m484p155.pdf.

Feehan:2014:HEK

[FFS14] C. J. Feehan, F. T. Y. Francis, and R. E. Scheibling. Harbouring the enemy: kelp holdfasts protect juvenile sea urchins from predatory crabs. Marine Ecology Progress Series, 514: 149–161, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/514/m514p149.pdf.

Fire:2015:BAM

[FFS^{+}15] S. E. Fire, L. J. Flewelling, M. Stolen, W. Noke Durden, M. de Wit, A. C. Spellman, and Z. Wang. Brevetoxinassociated mass mortality event of bottlenose dolphins and manatees along the east coast of Florida, USA. Marine Ecology Progress Series, 526:241–251, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/526/m526p241.pdf.

Faria:2010:OSB

[FG10] A. M. Faria and E. J. Gonçalves. Ontogeny of swimming behaviour of two temperate clingfishes, Lepadogaster lepadogaster and L. purpurea (Gobiesocidae). Marine Ecology Progress Series, 414:237–248, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/414/m414p237.pdf.

Foster:2016:SRC

[FG16a] T. Foster and J. P. Gilmour. Seeing red: Coral larvae are attracted to healthy-looking reefs. Marine Ecology Progress Series, 559:65–71, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/559/m559p065.pdf.

Frey:2016:SDG

[FG16b] D. L. Frey and P. Gagnon. Spatial dynamics of the green sea urchin Strongylocentrotus droebachiensis in food-depleted habitats. Marine Ecology Progress Series, 552:223–240, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/552/m552p223.pdf.

Fuller:2013:IST

[FGH^{+}13] W. J. Fuller, B. J. Godley, D. J. Hodgson, S. E. Reece, M. J. Witt, and A. C. Broderick. Importance of spatiotemporal data for predicting the effects of climate change on marine turtle sex ratios. Marine Ecology Progress Series, 488:267–274, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/488/m488p267.pdf.

Fu:2012:RIF

[FGL^{+}12] C. Fu, S. Gaichas, J. S. Link, A. Bundy, J. L. Boldt, A. M. Cook, R. Gamble, K. Rong Utne, H. Liu, and K. D. Friedland. Relative importance of fisheries, trophodynamic and environmental drivers in a series of marine ecosystems. Marine Ecology Progress Series, 459:169–184, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m459p169.pdf.

Ferguson:2016:MPZ

[FGOS16] A. J. P. Ferguson, R. K. Gruber, M. Orr, and P. Scanes. Morphological plasticity in Zostera muelleri across light, sediment, and nutrient gradients in Australian temperate coastal lakes. Marine Ecology Progress Series, 556:91–104, 2016. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m556p091.pdf.

Fischer:2017:ILN

[FGP17] R. Fischer, H. A. Giebel, and R. Ptacnik. Identity of the limiting nutrient (N vs. P) affects the competitive success of mixotrophs. Marine Ecology Progress Series, 563: 51–63, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/563/m563p051.pdf.

Ferrari:2012:SMC

[FGRM12] R. Ferrari, M. Gonzalez-Rivero, and P. J. Mumby. Size matters in competition between corals and macroalgae. Marine Ecology Progress Series, 467:77–88, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/467/m467p077.pdf.

Farrugia:2016:FUS

[FGTS16] T. J. Farrugia, K. J. Goldman, C. Tribuzio, and A. C. Seitz. First use of satellite tags to examine movement and habitat use of big skates Beringraja binoculata in the Gulf of Alaska. Marine Ecology Progress Series, 556:209–221, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/556/
m556p209.pdf.

Flanders:2015:KSA

[FGW^{+}15] N. P. Flanders, B. Gardner, K. J. Winiarski, P. W. C. Paton, T. Allison, and A. F. O’Connell. Key seabird areas in southern New England identified using a community occupancy model. Marine Ecology Progress Series, 533:277–290, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/533/
m533p277.pdf.

First:2010:EFS

[FH10] M. R. First and J. T. Hollibaugh. Environmental factors shaping microbial community structure in salt marsh sediments. Marine Ecology Progress Series, 399:15–26, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/399/
m399p015.pdf.

Fisher:2018:DIT

[FH18] M. Fisher and E. Hunter. Digital imaging techniques in otolith data capture, analysis and interpretation. Marine Ecology Progress Series, 598:213–231, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m598p213.pdf.

Fox:2012:IOM

[FHB^{+}12] H. E. Fox, K. M. Haisfield, M. S. Brown, T. C. Stevenson, B. N. Tissot, W. J. Walsh, and I. D. Williams. Influences of oceanographic and meteorological features on reef fish recruitment in Hawai’i. Marine Ecology Progress Series, 463:259–272, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/463/m463p259.pdf.

Froehlich:2015:STV

[FHE^{+}15] H. E. Froehlich, S. M. Hennessey, T. E. Essington, A. H. Beaudreau, and P. S. Levin. Spatial and temporal variation in nearshore macrofaunal community structure in a seasonally hypoxic estuary. Marine Ecology Progress Series, 520: 67–83, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/520/m520p067.pdf.

Fleming:2014:SJP

[FHG^{+}14] N. E. C. Fleming, C. Harrod, D. C. Griffin, J. Newton, and J. D. R. Houghton. Scyphozoan jellyfish provide short-term reproductive habitat for hyperiid amphipods in a temperate near-shore environment. Marine Ecology Progress Series, 510:229–240, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m510p229.pdf.

Fox:2017:PML

[FHH^{+}17] C. H. Fox, F. H. Huettmann, G. K. A. Harvey, K. H. Morgan, J. Robinson, R. Williams, and P. C. Paquet. Predictions from machine learning ensembles: marine bird distribution and density on Canada’s Pacific coast. Marine Ecology Progress Series, 566:199–216, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m566p199.pdf.

Ferreira:2015:PBC

[FHJ15] J. G. Ferreira, S. J. Hawkins, and S. R. Jenkins. Physical and biological control of fucoid recruitment in range edge and range centre populations. Marine Ecology Progress Series, 518:85–94, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/518/m518p085.pdf.

Fitzpatrick:2015:EFF

[FHL^{+}15] B. M. Fitzpatrick, E. S. Harvey, T. J. Langlois, R. Babcock, and E. Twiggs. Effects of fishing on fish assemblages at the reefscape scale. Marine Ecology Progress Series, 524: 241–253, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/524/m524p241.pdf.

Fan:2017:PLT

[FHL^{+}17] T. Y. Fan, Y. C. Hsieh, K. H. Lin, F. W. Kuo, K. Soong, C. J. McRae, P. J. Edmunds, and L. S. Fang. Plasticity in lunar timing of larval release of two brooding pocilloporid corals in an internal tide-induced upwelling reef. Marine Ecology Progress Series, 569:117–127, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m569p117.pdf.

Fernandez:2011:GSB

[FHM^{+}11] M. V. Fernández, S. Heras, F. Maltagliati, A. Turco, and M. I. Roldán. Genetic structure in the blue and red shrimp Aristeus antennatus and the role played by hydrographical and oceanographical barriers. Marine Ecology Progress Series, 421:163–171, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/421/m421p163.pdf.

Foo:2019:IFH

[FHMG19] D. Foo, M. Hindell, C. McMahon, and S. Goldsworthy. Identifying foraging habitats of adult female long-nosed fur seal Arctocephalus forsteri based on vibrissa stable isotopes. Marine Ecology Progress Series, 628:223–234, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/628/
m628p223.pdf.

Fiechter:2016:FCE

[FHRC16] J. Fiechter, L. A. Huckstadt, K. A. Rose, and D. P. Costa. A fully coupled ecosystem model to predict the foraging ecology of apex predators in the California Current. Marine Ecology Progress Series, 556:273–285, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/556/m556p273.pdf.

Friedrichs:2013:SBP

[FHS^{+}13] L. Friedrichs, M. Hörnig, L. Schulze, A. Bertram, S. Jansen, and C. Hamm. Size and biomechanic properties of diatom frustules influence food uptake by copepods. Marine Ecology Progress Series, 481:41–51, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/481/m481p041.pdf.

Farchadi:2019:MDH

[FHTY19] N. Farchadi, M. G. Hinton, A. R. Thompson, and Z. Y. Yin. Modeling the dynamic habitats of mobile pelagic predators (Makaira nigricans and Istiompax indica) in the eastern Pacific Ocean. Marine Ecology Progress Series, 622:157– 176, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/622/m622p157.pdf.

Fielding:2014:PCR

[Fie14] S. R. Fielding. Predicting coccolithophore rain ratio responses to calcite saturation state. Marine Ecology Progress Series, 500:57–65, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/500/m500p057.pdf.

Fisher:2010:PLA

[Fis10] J. A. D. Fisher. Parasite-like associations in rocky intertidal assemblages: implications for escalated gastropod defenses. Marine Ecology Progress Series, 399:199–209, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/399/
m399p199.pdf.

Fiksen:2011:MOB

[FJ11] Ø. Fiksen and C. Jørgensen. Model of optimal behaviour in fish larvae predicts that food availability determines survival, but not growth. Marine Ecology Progress Series, 432: 207–219, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m432p207.pdf.

Finstad:2012:EIT

[FJ12] A. G. Finstad and B. Jonsson. Effect of incubation temperature on growth performance in Atlantic salmon. Marine Ecology Progress Series, 454:75–82, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/454/
m454p075.pdf.

Fagerberg:2010:MSR

[FJC10] T. Fagerberg, T. Jephson, and P. Carlsson. Molecular size of riverine dissolved organic matter influences coastal phytoplankton communities. Marine Ecology Progress Series, 409: 17–25, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/409/m409p017.pdf.

Franklin:2013:PMC

[FJD13] E. C. Franklin, P. L. Jokiel, and M. J. Donahue. Predictive modeling of coral distribution and abundance in the Hawaiian Islands. Marine Ecology Progress Series, 481:121– 132, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/481/m481p121.pdf.

Flukes:2012:FSU

[FJL12] E. B. Flukes, C. R. Johnson, and S. D. Ling. Forming sea urchin barrens from the inside out: an alternative pattern of overgrazing. Marine Ecology Progress Series, 464: 179–194, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m464p179.pdf.

Flukes:2014:TKC

[FJW14] E. B. Flukes, C. R. Johnson, and J. T. Wright. Thinning of kelp canopy modifies understory assemblages: the importance of canopy density. Marine Ecology Progress Series, 514:57–70, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/514/m514p057.pdf.

Foley:2010:CBA

[FK10] M. M. Foley and P. L. Koch. Correlation between allochthonous subsidy input and isotopic variability in the giant kelp Macrocystis pyrifera in central California, USA. Marine Ecology Progress Series, 409:41–50, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/409/
m409p041.pdf.

Freitas:2012:IFI

[FKA^{+}12] C. Freitas, K. M. Kovacs, M. Andersen, J. Aars, S. Sandven, M. Skern-Mauritzen, O. Pavlova, and C. Lydersen. Importance of fast ice and glacier fronts for female polar bears and their cubs during spring in Svalbard, Norway. Marine Ecology Progress Series, 447:289–304, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/447/m447p289.pdf.

Feutry:2015:WMS

[FKP^{+}15] P. Feutry, P. M. Kyne, R. D. Pillans, X. Chen, J. Marthick, D. L. Morgan, and P. M. Grewe. Whole mitogenome sequencing refines population structure of the critically endangered sawfish Pristis pristis. Marine Ecology Progress Series, 533:237–244, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m533p237.pdf.

Fujii:2013:DSA

[FKRJ13] T. Fujii, N. M. Kilgallen, A. A. Rowden, and A. J. Jamieson. Deep-sea amphipod community structure across abyssal to hadal depths in the Peru–Chile and Kermadec trenches. Marine Ecology Progress Series, 492:125–138, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/492/
m492p125.pdf.

Figuerola:2015:DPA

[FKT15] B. Figuerola, P. Kuklinski, and P. D. Taylor. Depth patterns in Antarctic bryozoan skeletal Mg-calcite: Can they provide an analogue for future environmental changes? Marine Ecology Progress Series, 540:109–120, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/540/m540p109.pdf.

Freitas:2012:LTH

[FKvdV12] V. Freitas, S. A. L. M. Kooijman, and H. W. van der Veer. Latitudinal trends in habitat quality of shallow-water flatfish nurseries. Marine Ecology Progress Series, 471:203– 214, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/471/m471p203.pdf.

Favero:2018:MLT

[FKZTT18] J. M. Favero, M. Katsuragawa, M. L. Zani-Teixeira, and J. T. Turner. Modeling long-term fluctuations in the distribution and abundance of Engraulis anchoita eggs and larvae in the southeastern Brazilian bight. Marine Ecology Progress Series, 587:159–173, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/587/m587p159.pdf.

Fairweather:2010:PFE

[FL10] P. G. Fairweather and R. E. Lester. Predicting future ecological degradation based on modelled thresholds. Marine Ecology Progress Series, 413:291–304, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m413p291.pdf.

Franze:2017:MFW

[FL17] G. Franzè and P. J. Lavrentyev. Microbial food web structure and dynamics across a natural temperature gradient in a productive polar shelf system. Marine Ecology Progress Series, 569:89–102, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/569/m569p089.pdf.

Farlin:2010:FDA

[FLAL10] J. P. Farlin, L. S. Lewis, T. W. Anderson, and C. T. Lai. Functional diversity in amphipods revealed by stable isotopes in an eelgrass ecosystem. Marine Ecology Progress Series, 420:277–281, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/420/m420p277.pdf.

Field:2014:SIA

[FLC^{+}14] John C. Field, Steven Y. Litvin, Aaron Carlisle, Julia S. Stewart, Wiliam F. Gilly, and Rocio I. Ruiz-Cooley. Stable isotope analysis of Humboldt squid prey: Comment on Miller et al. (2013). Marine Ecology Progress Series, 500: 281–285, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/500/m500p281.pdf. See [MBS^{+}13] and

reply [MBS^{+}14].

Fripiat:2011:ESR

[FLE^{+}11] F. Fripiat, K. Leblanc, M. Elskens, A. J. Cavagna, L. Armand, L. André, F. Dehairs, and D. Cardinal. Efficient silicon recycling in summer in both the polar frontal and subantarctic zones of the Southern Ocean. Marine Ecology Progress Series, 435:47–61, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/435/m435p047.pdf.

Fisk:2012:APT

[FLK12] A. T. Fisk, C. Lydersen, and K. M. Kovacs. Archival pop-off tag tracking of Greenland sharks Somniosus microcephalus in the high Arctic waters of Svalbard, Norway. Marine Ecology Progress Series, 468:255–265, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/468/m468p255.pdf.

Feng:2018:HIM

[FLS^{+}18] S. Feng, J. Lin, S. Sun, F. Zhang, and C. Li. Hyposalinity and incremental micro-zooplankton supply in earlydeveloped Nemopilema nomurai polyp survival, growth, and podocyst reproduction. Marine Ecology Progress Series, 591: 117–128, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/591/m591p117.pdf.

Fuentes-Lema:2015:ESU

[FLSG^{+}15] A. Fuentes-Lema, C. Sobrino, N. González, M. Estrada, and P. J. Neale. Effect of solar UVR on the production of particulate and dissolved organic carbon from phytoplankton assemblages in the Indian Ocean. Marine Ecology Progress Series, 535:47–61, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/535/m535p047.pdf.

Fraser:2014:GSB

[FM14] M. R. Fraser and M. I. McCormick. Gender-specific benefits of eating eggs at resident reef fish spawning aggregation sites. Marine Ecology Progress Series, 517:209–216, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/517/
m517p209.pdf.

Fourqurean:2010:EES

[FMC^{+}10] J. W. Fourqurean, S. Manuel, K. A. Coates, W. J. Kenworthy, and S. R. Smith. Effects of excluding sea turtle herbivores from a seagrass bed: Overgrazing may have led to loss of seagrass meadows in Bermuda. Marine Ecology Progress Series, 419:223–232, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/419/m419p223.pdf.

Forcioli:2011:SDI

[FMC^{+}11] D. Forcioli, P. L. Merle, C. Caligara, M. Ciosi, C. Muti, P. Francour, C. Cerrano, and D. Allemand. Symbiont diversity is not involved in depth acclimation in the Mediterranean sea whip Eunicella singularis. Marine Ecology Progress Series, 439:57–71, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/439/m439p057.pdf.

Fournet:2018:HWM

[FMG^{+}18] M. E. H. Fournet, L. P. Matthews, C. M. Gabriele, S. Haver, D. K. Mellinger, and H. Klinck. Humpback whales Megaptera novaeangliae alter calling behavior in response to natural sounds and vessel noise. Marine Ecology Progress Series, 607:251–268, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m607p251.pdf.

Fields:2015:CSC

[FMH^{+}15] L. Fields, J. Mercer, K. J. W. Hyde, M. Brush, S. W. Nixon, C. Oviatt, M. L. Schwartz, D. Ullman, and D. Codiga. Comparison of surface chlorophyll, primary production, and satellite imagery in hydrographically different sounds off southern New England. Marine Ecology Progress Series, 535: 29–45, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/535/m535p029.pdf.

Foo:2019:ISO

[FMH^{+}19] D. Foo, C. McMahon, M. Hindell, S. Goldsworthy, and F. Bailleul. Influence of shelf oceanographic variability on alternate foraging strategies in long-nosed fur seals. Marine Ecology Progress Series, 615:189–204, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/615/
m615p189.pdf.

Font-Munoz:2018:APC

[FMJA^{+}18] J. S. Font-Muñoz, A. Jordi, S. Anglès, P. Ferriol, E. Garcés, and G. Basterretxea. Assessing phytoplankton community composition using combined pigment and particle size distribution analysis. Marine Ecology Progress Series, 594: 51–63, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/594/m594p051.pdf.

Fabrizio:2013:HAD

[FMP13] M. C. Fabrizio, J. P. Manderson, and J. P. Pessutti. Habitat associations and dispersal of black sea bass from a mid-Atlantic Bight reef. Marine Ecology Progress Series, 482: 241–253, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m482p241.pdf.

Firstater:2016:CCE

[FNA^{+}16] F. N. Firstater, M. Narvarte, M. F. Alvarez, M. E. Fanjul, and O. O. Iribarne. Cordgrass canopy elicits weak effects on sediment properties and microphytobenthic abundance in a harsh environment. Marine Ecology Progress Series, 550:101–110, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/550/m550p101.pdf.

Fagerli:2013:LSU

[FNC13] C. W. Fagerli, K. M. Norderhaug, and H. C. Christie. Lack of sea urchin settlement may explain kelp forest recovery in overgrazed areas in Norway. Marine Ecology Progress Series, 488:119–132, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m488p119.pdf.

Fagerli:2014:PDS

[FNC^{+}14] C. W. Fagerli, K. M. Norderhaug, H. Christie, M. F. Pedersen, and S. Fredriksen. Predators of the destructive sea urchin Strongylocentrotus droebachiensis on the Norwegian coast. Marine Ecology Progress Series, 502:207–218, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m502p207.pdf.

Farmer:2018:RES

[FNF^{+}18] N. A. Farmer, D. P. Noren, E. M. Fougères, A. Machernis, and K. Baker. Resilience of the endangered sperm whale Physeter macrocephalus to foraging disturbance in the Gulf of Mexico, USA: a bioenergetic approach. Marine Ecology Progress Series, 589:241–261, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/589/m589p241.pdf.

Fogarty:2012:CAC

[Fog12] N. D. Fogarty. Caribbean acroporid coral hybrids are viable across life history stages. Marine Ecology Progress Series, 446:145–159, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/446/m446p145.pdf.

Fouzai:2015:ETF

[FOJF15] N. Fouzai, A. F. Opdal, C. Jørgensen, and Ø. Fiksen. Effects of temperature and food availability on larval cod survival: a model for behaviour in vertical gradients. Marine Ecology Progress Series, 529:199–212, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m529p199.pdf.

Fogarty:2012:ASP

[FOL12] M. J. Fogarty, W. J. Overholtz, and J. S. Link. Aggregate surplus production models for demersal fishery resources of the Gulf of Maine. Marine Ecology Progress Series, 459: 247–258, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m459p247.pdf.

Fox:2016:BLR

[Fox16] M. D. Fox. Biomass loss reduces growth and resource translocation in giant kelp Macrocystis pyrifera. Marine Ecology Progress Series, 562:65–77, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/562/m562p065.pdf.

Frank:2016:AED

[FPBL16] K. T. Frank, B. Petrie, D. Boyce, and W. C. Leggett. Anomalous ecosystem dynamics following the apparent collapse of a keystone forage species. Marine Ecology Progress Series, 553:185–202, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/553/m553p185.pdf.

Fire:2012:SEE

[FPC^{+}12] S. E. Fire, J. Pruden, D. Couture, Z. Wang, M. Y. Dechraoui Bottein, B. L. Haynes, T. Knott, D. Bouchard, A. Lichtenwalner, and G. Wippelhauser. Saxitoxin exposure in an endangered fish: association of a shortnose sturgeon mortality event with a harmful algal bloom. Marine Ecology Progress Series, 460:145–153, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/460/m460p145.pdf.

Fanelli:2013:TWD

[FPC^{+}13] E. Fanelli, V. Papiol, J. E. Cartes, P. Rumolo, and C. López-Pérez. Trophic webs of deep-sea megafauna on mainland and insular slopes of the NW Mediterranean: a comparison by stable isotope analysis. Marine Ecology Progress Series, 490:199–221, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/490/m490p199.pdf.

Fuentes:2018:JBA

[FPC^{+}18] V. L. Fuentes, J. E. Purcell, R. H. Condon, F. Lombard, and C. H. Lucas. Jellyfish blooms: advances and challenges. Marine Ecology Progress Series, 591:3–5, 2018. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m591p003.pdf.

Fossette:2012:BGA

[FPL^{+}12] S. Fossette, N. F. Putman, K. J. Lohmann, R. Marsh, and G. C. Hays. A biologist’s guide to assessing ocean currents: a review. Marine Ecology Progress Series, 457:285–301, 2012.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m457p285.pdf.

Fisher:2014:DLA

[FPM14] J. L. Fisher, W. T. Peterson, and S. G. Morgan. Does larval advection explain latitudinal differences in recruitment across upwelling regimes? Marine Ecology Progress Series, 503:123–137, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m503p123.pdf.

Fuhrmann:2017:TNI

[FPN17] M. M. Fuhrmann, T. Pedersen, and E. M. Nilssen. Trophic niche of the invasive red king crab Paralithodes camtschaticus in a benthic food web. Marine Ecology Progress Series, 565:113–129, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/565/m565p113.pdf.

Fox:2018:PHS

[FPR18] C. H. Fox, P. C. Paquet, and T. E. Reimchen. Pacific herring spawn events influence nearshore subtidal and intertidal species. Marine Ecology Progress Series, 595:157–169, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m595p157.pdf.

Froese:2011:WSI

[FQ11] R. Froese and M. Quaas. AS WE SEE IT: Three options for rebuilding the cod stock in the eastern Baltic Sea. Marine Ecology Progress Series, 434:197–200, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/434/
m434p197.pdf.

Falkenberg:2013:FHI

[FRC13a] L. J. Falkenberg, B. D. Russell, and S. D. Connell. Future herbivory: the indirect effects of enriched CO_{2} may rival its direct effects. Marine Ecology Progress Series, 492: 85–95, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/492/m492p085.pdf.

Fratini:2013:PGI

[FRC^{+}13b] S. Fratini, L. Ragionieri, G. Cutuli, M. Vannini, and S. Cannicci. Pattern of genetic isolation in the crab Pachygrapsus marmoratus within the Tuscan Archipelago (Mediterranean Sea). Marine Ecology Progress Series, 478:173–183, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/478/m478p173.pdf.

Fernandez-Reiriz:2015:FAC

[FRGI15] M. J. Fernández-Reiriz, J. L. Garrido, and J. Irisarri. Fatty acid composition in Mytilus galloprovincialis organs: trophic interactions, sexual differences and differential anatomical distribution. Marine Ecology Progress Series, 528:221–234, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2015/528/m528p221.pdf.

Foden:2010:RUS

[FRJ10] J. Foden, S. I. Rogers, and A. P. Jones. Recovery of UK seabed habitats from benthic fishing and aggregate extraction— towards a cumulative impact assessment. Marine Ecology Progress Series, 411:259–270, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/411/m411p259.pdf.

Foden:2011:HPU

[FRJ11] J. Foden, S. I. Rogers, and A. P. Jones. Human pressures on UK seabed habitats: a cumulative impact assessment. Marine Ecology Progress Series, 428:33–47, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/428/
m428p033.pdf.

Fitzsimmons:2017:HGL

[FRMS17] M. G. Fitzsimmons, M. E. Rector, D. W. McKay, and A. E. Storey. High growth and low corticosterone in foodsupplemented Atlantic puffin Fratercula arctica chicks under poor foraging conditions. Marine Ecology Progress Series, 565:217–226, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/565/m565p217.pdf.

Fujimura:2017:NSO

[FRP^{+}17] A. G. Fujimura, A. J. H. M. Reniers, C. B. Paris, A. L. Shanks, J. H. MacMahan, and S. G. Morgan. Numerical simulations of onshore transport of larvae and detritus to a steep pocket beach. Marine Ecology Progress Series, 582:33–43, 2017. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m582p033.pdf.

Fernandez-Reiriz:2012:TJM

[FRRÁS^{+}12] M. J. Fernández-Reiriz, P. Range, X. A. Álvarez-Salgado, J. Espinosa, and U. Labarta. Tolerance of juvenile Mytilus galloprovincialis to experimental seawater acidification. Marine Ecology Progress Series, 454:65–74, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/454/
m454p065.pdf.

Fernandez-Reiriz:2011:PEJ

[FRRÁSL11] M. J. Fernández-Reiriz, P. Range, X. A. Álvarez-Salgado, and U. Labarta. Physiological energetics of juvenile clams Ruditapes decussatus in a high CO_{2} coastal ocean. Marine Ecology Progress Series, 433:97–105, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/433/m433p097.pdf.

Fitzgibbon:2014:TTN

[FRTB14] Q. P. Fitzgibbon, N. Ruff, S. R. Tracey, and S. C. Battaglene. Thermal tolerance of the nektonic puerulus stage of spiny lobsters and implications of ocean warming. Marine Ecology Progress Series, 515:173–186, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/515/m515p173.pdf.

Fiedler:2013:ETC

[FRV^{+}13] P. C. Fiedler, J. V. Redfern, J. Van Noord, C. Hall, R. L. Pitman, and L. T. Ballance. Effects of a tropical cyclone on a pelagic ecosystem from the physical environment to top predators. Marine Ecology Progress Series, 484:1–16, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Fry:2013:AAS

[Fry13a] Brian Fry. Alternative approaches for solving underdetermined isotope mixing problems. Marine Ecology Progress Series, 472:1–13, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/feature/m472p001.pdf. See comment

[SWP^{+}13] and reply [Fry13b].

Fry:2013:RCM

[Fry13b] Brian Fry. Reply comment: Minmax solutions for underdetermined isotope mixing problems: Reply to Semmens et al. (2013). Marine Ecology Progress Series, 490:291–294, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_

oa/m490p291.pdf. See [Fry13a, SWP^{+}13].

Feehan:2014:DCS

[FS14] C. J. Feehan and R. E. Scheibling. Disease as a control of sea urchin populations in Nova Scotian kelp beds. Marine Ecology Progress Series, 500:149–158, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/500/m500p149.pdf.

Fuchs:2018:EDR

[FS18] H. L. Fuchs and J. A. Specht. Evidence for diverse responses to viscosity in suspension-feeding bivalves: Reply to Riisgård & Larsen (2018). Marine Ecology Progress Series, 596:267–270, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/596/m596p267.pdf.

Falkner:2015:EMP

[FSB15] I. Falkner, M. A. Sewell, and M. Byrne. Evolution of maternal provisioning in ophiuroid echinoderms: characterisation of egg composition in planktotrophic and lecithotrophic developers. Marine Ecology Progress Series, 525:1–13, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Feng:2014:LSS

[FSD^{+}14] J. Feng, L. C. Stige, J. M. Durant, D. O. Hessen, L. Zhu, D. Ø. Hjermann, M. Llope, and N. C. Stenseth. Large-scale season-dependent effects of temperature and zooplankton on phytoplankton in the North Atlantic. Marine Ecology Progress Series, 502:25–37, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/502/m502p025.pdf.

Friedlander:2010:SPS

[FSDS10] A. M. Friedlander, S. A. Sandin, E. E. DeMartini, and E. Sala. Spatial patterns of the structure of reef fish assemblages at a pristine atoll in the central Pacific. Marine Ecology Progress Series, 410:219–231, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/410/m410p219.pdf.

Freeman:2016:SCN

[FSE^{+}16] C. J. Freeman, E. W. Stoner, C. G. Easson, K. O. Matterson, and D. M. Baker. Symbiont carbon and nitrogen assimilation in the Cassiopea–Symbiodinium mutualism. Marine Ecology Progress Series, 544:281–286, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/544/m544p281.pdf.

Freeman:2017:VVS

[FSE^{+}17] C. J. Freeman, E. W. Stoner, C. G. Easson, K. O. Matterson, and D. M. Baker. Variation in δ ^{13}C and δ ^{15}N values suggests a coupling of host and symbiont metabolism in the Symbiodinium−Cassiopea mutualism. Marine Ecology Progress Series, 571:245–251, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/571/m571p245.pdf.

Fayet:2017:WPS

[FSF^{+}17] A. L. Fayet, A. Shoji, R. Freeman, C. M. Perrins, and T. Guilford. Within-pair similarity in migration route and female winter foraging effort predict pair breeding performance in a monogamous seabird. Marine Ecology Progress Series, 569: 243–252, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m569p243.pdf.

Frankenbach:2018:PRM

[FSFS18] S. Frankenbach, W. Schmidt, J. C. Frommlet, and J. Serôdio. Photoinactivation, repair and the motility-physiology trade-off in microphytobenthos. Marine Ecology Progress Series, 601:

41–57, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/601/m601p041.pdf.

Freitas:2016:CCE

[FSH^{+}16] R. F. Freitas, E. C. Schrack, Q. He, B. R. Silliman, E. B. Furlong, A. C. Telles, and C. S. B. Costa. Consumer control of the establishment of marsh foundation plants in intertidal mudflats. Marine Ecology Progress Series, 547:79– 89, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/547/m547p079.pdf.

Feehan:2012:AFB

[FSLG12] C. Feehan, R. E. Scheibling, and J. S. Lauzon-Guay. Aggregative feeding behavior in sea urchins leads to destructive grazing in a Nova Scotian kelp bed. Marine Ecology Progress Series, 444:69–83, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/444/m444p069.pdf.

Fu:2014:BSR

[FSM^{+}14] Z. Fu, M. Shibata, R. Makabe, H. Ikeda, and S. Uye. Body size reduction under starvation, and the point of no return, in ephyrae of the moon jellyfish Aurelia aurita. Marine Ecology Progress Series, 510:255–263, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m510p255.pdf.

Fluhr:2017:WSG

[FSP^{+}17] J. Fluhr, H. Strøm, R. Pradel, O. Duriez, G. Beaugrand, and S. Descamps. Weakening of the subpolar gyre as a key driver of North Atlantic seabird demography: a case study with Brünnich’s guillemots in Svalbard. Marine Ecology Progress Series, 563:1–11, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Flight:2010:PSF

[FSR10] P. A. Flight, S. D. Schoepfer, and D. M. Rand. Physiological stress and the fitness effects of Mpi genotypes in the acorn barnacle Semibalanus balanoides. Marine Ecology Progress Series, 404:139–149, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/404/m404p139.pdf.

Ford:2017:SPH

[FSR17] B. M. Ford, B. A. Stewart, and J. D. Roberts. Species pools and habitat complexity define Western Australian marine fish community composition. Marine Ecology Progress Series, 574:157–166, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/574/m574p157.pdf.

Friedman:2018:EGR

[FSS^{+}18] W. R. Friedman, J. A. Santora, I. D. Schroeder, D. D. Huff, R. D. Brodeur, J. C. Field, and B. K. Wells. Environmental and geographic relationships among salmon forage assemblages along the continental shelf of the California Current. Marine Ecology Progress Series, 596:181–198, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/596/
m596p181.pdf.

Froeschke:2010:EIO

[FSW10] J. Froeschke, G. W. Stunz, and M. L. Wildhaber. Environmental influences on the occurrence of coastal sharks in estuarine waters. Marine Ecology Progress Series, 407:279– 292, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/407/m407p279.pdf.

Franklin:2010:DDD

[FSY^{+}10a] D. J. Franklin, M. Steinke, J. Young, I. Probert, and G. Malin. Dimethylsulphoniopropionate (DMSP), DMSP-lyase activity (DLA) and dimethylsulphide (DMS) in 10 species of coccolithophore. Marine Ecology Progress Series, 410:13– 23, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/410/m410p013.pdf.

Fukuwaka:2010:BMC

[FSY^{+}10b] M. Fukuwaka, S. Sato, O. Yamamura, O. Sakai, T. Nagasawa, A. Nishimura, and T. Azumaya. Biomass and mortality of chum salmon in the pelagic Bering Sea. Marine Ecology Progress Series, 403:219–230, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/403/m403p219.pdf.

Forehead:2010:MCS

[FT10] H. I. Forehead and P. A. Thompson. Microbial communities of subtidal shallow sandy sediments change with depth and wave disturbance, but nutrient exchanges remain similar. Marine Ecology Progress Series, 414:11–26, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/414/
m414p011.pdf.

Ferriss:2014:RIA

[FTB14] B. E. Ferriss, M. Trudel, and B. R. Beckman. Regional and inter-annual trends in marine growth of juvenile salmon in coastal pelagic ecosystems of British Columbia, Canada. Marine Ecology Progress Series, 503:247–261, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m503p247.pdf.

Freshwater:2017:EDD

[FTB^{+}17] C. Freshwater, M. Trudel, T. D. Beacham, S. C. H. Grant, S. C. Johnson, C. E. M. Neville, S. Tucker, and F. Juanes. Effects of density during freshwater and early marine rearing on juvenile sockeye salmon size, growth, and migration. Marine Ecology Progress Series, 579:97–110, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/579/
m579p097.pdf.

Freuchet:2015:IES

[FTF15] F. Freuchet, R. Tremblay, and A. A. V. Flores. Interacting environmental stressors modulate reproductive output and larval performance in a tropical intertidal barnacle. Marine Ecology Progress Series, 532:161–175, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/532/m532p161.pdf.

Fu:2012:GCF

[FTH12] F. X. Fu, A. O. Tatters, and D. A. Hutchins. Global change and the future of harmful algal blooms in the ocean.

Marine Ecology Progress Series, 470:207–233, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m470p207.pdf.

Fedosov:2014:TDE

[FTKK14] A. E. Fedosov, A. V. Tiunov, S. I. Kiyashko, and Y. I. Kantor. Trophic diversification in the evolution of predatory marine gastropods of the family terebridae as inferred from stable isotope data. Marine Ecology Progress Series, 497:143–156, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/497/m497p143.pdf.

Fortune:2013:ERN

[FTM^{+}13] S. M. E. Fortune, A. W. Trites, C. A. Mayo, D. A. S. Rosen, and P. K. Hamilton. Energetic requirements of North Atlantic right whales and the implications for species recovery. Marine Ecology Progress Series, 478:253–272, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/478/
m478p253.pdf.

Feis:2015:MVH

[FTO^{+}15] M. E. Feis, D. W. Thieltges, J. L. Olsen, X. de Montaudouin, K. T. Jensen, H. Bazaïri, S. C. Culloty, and P. C. Luttikhuizen. The most vagile host as the main determinant of population connectivity in marine macroparasites. Marine Ecology Progress Series, 520:85–99, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/520/m520p085.pdf.

Friedlaender:2013:EDV

[FTS^{+}13] A. S. Friedlaender, R. B. Tyson, A. K. Stimpert, A. J. Read, and D. P. Nowacek. Extreme diel variation in the feeding behavior of humpback whales along the western Antarctic Peninsula during autumn. Marine Ecology Progress Series, 494:281–289, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/494/m494p281.pdf.

Fauchald:2017:SPS

[FTT^{+}17] P. Fauchald, A. Tarroux, T. Tveraa, Y. Cherel, Y. Ropert-Coudert, A. Kato, O. P. Love, Ø. Varpe, and S. Descamps. Spring phenology shapes the spatial foraging behavior of Antarctic petrels. Marine Ecology Progress Series, 568:203– 215, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m568p203.pdf.

Fox:2012:IMF

[FTvdK^{+}12] C. J. Fox, M. I. Taylor, J. van der Kooij, N. Taylor, S. P. Milligan, A. Albaina, S. Pascoal, D. Lallias, M. Maillard, and E. Hunter. Identification of marine fish egg predators using molecular probes. Marine Ecology Progress Series, 462:205–218, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m462p205.pdf.

Fidler:2017:RWB

[FTWA17] R. Y. Fidler, R. G. Turingan, A. T. White, and M. N. R. Alava. Reef-wide beneficial shifts in fish population structure following establishment of marine protected areas in Philippine coral reefs. Marine Ecology Progress Series, 570:187– 202, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/570/m570p187.pdf.

Fujii:2016:PIO

[Fuj16] T. Fujii. Potential influence of offshore oil and gas platforms on the feeding ecology of fish assemblages in the North Sea. Marine Ecology Progress Series, 542:167–186, 2016. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m542p167.pdf.

Fernandez-Urruzola:2016:BMA

[FUOG^{+}16] I. Fernández-Urruzola, N. Osma, M. Gómez, S. Montesdeoca-Esponda, and T. T. Packard. Building a model of ammonium excretion in two species of marine zooplankton based on glutamate dehydrogenase kinetics. Marine Ecology Progress Series, 550:83–99, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/550/m550p083.pdf.

Fawcett:2011:PSN

[FW11] S. E. Fawcett and B. B. Ward. Phytoplankton succession and nitrogen utilization during the development of an upwelling bloom. Marine Ecology Progress Series, 428:13–31, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2011/428/m428p013.pdf.

Franco:2015:HDK

[FWB^{+}15] J. N. Franco, T. Wernberg, I. Bertocci, P. Duarte, D. Jacinto, N. Vasco-Rodrigues, and F. Tuya. Herbivory drives kelp recruits into ’hiding’ in a warm ocean climate. Marine Ecology Progress Series, 536:1–9, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ford:2010:ICE

[FWF^{+}10] J. R. Ford, R. J. Williams, A. M. Fowler, D. R. Cox, and I. M. Suthers. Identifying critical estuarine seagrass habitat for settlement of coastally spawned fish. Marine Ecology Progress Series, 408:181–193, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/408/m408p181.pdf.

Freedman:2015:CMJ

[FWL15] R. Freedman, C. R. Whitcraft, and C. G. Lowe. Connectivity and movements of juvenile predatory fishes between discrete restored estuaries in southern California. Marine Ecology Progress Series, 520:191–201, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/520/m520p191.pdf.

Fevolden:2012:SDV

[FWPP12] S. E. Fevolden, J. I. Westgaard, T. Pedersen, and K. Præbel. Settling-depth vs. genotype and size vs. genotype correlations at the Pan I locus in 0-group Atlantic cod Gadus morhua. Marine Ecology Progress Series, 468:267–278, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m468p267.pdf.

Feng:2018:STS

[FWS^{+}18] S. Feng, S. W. Wang, S. Sun, F. Zhang, G. T. Zhang, M. T. Liu, and S. Uye. Strobilation of three scyphozoans (Aurelia coelurea, Nemopilema nomurai, and Rhopilema esculentum) in the field at Jiaozhou Bay, China. Marine Ecology Progress Series, 591:141–153, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/591/m591p141.pdf.

Feng:2016:LBF

[FWW^{+}16] D. Q. Feng, W. Wang, X. Wang, Y. Qiu, and C. H. Ke. Low barnacle fouling on leaves of the mangrove plant Sonneratia apetala and possible anti-barnacle defense strategies. Marine Ecology Progress Series, 544:169–182, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/544/
m544p169.pdf.

Fernandez:2018:MTH

[FYG^{+}18] M. Fernandez, C. Yesson, A. Gannier, P. I. Miller, and J. M. N. Azevedo. A matter of timing: how temporal scale selection influences cetacean ecological niche modelling. Marine Ecology Progress Series, 595:217–231, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/595/m595p217.pdf.

Fujita:2016:CPT

[FZI^{+}16] J. Fujita, K. Zenimoto, A. Iguchi, Y. Kai, M. Ueno, and Y. Yamashita. Comparative phylogeography to test for predictions of marine larval dispersal in three amphidromous shrimps. Marine Ecology Progress Series, 560:105–120, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/560/
m560p105.pdf.

Goldstein:2011:EAF

[GA11] S. T. Goldstein and E. Alve. Experimental assembly of Foraminiferal communities from coastal propagule banks. Marine Ecology Progress Series, 437:1–11, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Gedan:2011:UFN

[GAB11] K. B. Gedan, A. H. Altieri, and M. D. Bertness. Uncertain future of New England salt marshes. Marine Ecology Progress Series, 434:229–237, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m434p229.pdf.

Goutte:2014:AVT

[GAB^{+}14] A. Goutte, F. Angelier, C. Bech, C. Clément-Chastel, G. Dell’Omo, G. W. Gabrielsen, Á. Z. Lendvai, B. Moe, E. Noreen, D. Pinaud, S. Tartu, and O. Chastel. Annual variation in the timing of breeding, pre-breeding foraging areas and corticosterone levels in an Arctic population of black-legged kittiwakes. Marine Ecology Progress Series, 496:233–247, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/496/m496p233.pdf.

Green:2011:FBP

[GAC11] S. J. Green, J. L. Akins, and I. M. Côté. Foraging behaviour and prey consumption in the Indo–Pacific lionfish on Bahamian coral reefs. Marine Ecology Progress Series, 433:159–167, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/433/m433p159.pdf.

Gilg:2016:DGE

[GAF^{+}16] I. C. Gilg, S. D. Archer, S. A. Floge, D. M. Fields, A. I. Vermont, A. H. Leavitt, W. H. Wilson, and J. Martínez Martínez. Differential gene expression is tied to photochemical efficiency reduction in virally infected Emiliania huxleyi. Marine Ecology Progress Series, 555:13–27, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m555p013.pdf.

Gallegos:2014:LTVa

[Gal14a] C. L. Gallegos. Long-term variations in primary production in a eutrophic sub-estuary. I. Seasonal and spatial patterns. Marine Ecology Progress Series, 502:53–67, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m502p053.pdf.

Gallegos:2014:LTVb

[Gal14b] C. L. Gallegos. Long-term variations in primary production in a eutrophic sub-estuary. II. Interannual variations and modeling. Marine Ecology Progress Series, 502:69–83, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m502p069.pdf.

Georgiou:2016:TLF

[GALS16] D. Georgiou, A. Alexandre, J. Luis, and R. Santos. Temperature is not a limiting factor for the expansion of Halophila stipulacea throughout the Mediterranean Sea. Marine Ecology Progress Series, 544:159–167, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/544/m544p159.pdf.

Gestoso:2014:BRF

[GAO14] I. Gestoso, F. Arenas, and C. Olabarria. Biotic resistance and facilitation of a non-indigenous mussel vary with environmental context. Marine Ecology Progress Series, 506: 163–173, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/506/m506p163.pdf.

Gardner:2013:BCT

[Gar13] J. P. A. Gardner. Bottom-up control of temperate rocky intertidal community structure: evidence from a transplant experiment. Marine Ecology Progress Series, 491:137–151, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/491/m491p137.pdf.

Goatley:2010:BMS

[GB10a] C. H. R. Goatley and D. R. Bellwood. Biologically mediated sediment fluxes on coral reefs: sediment removal and off-reef transportation by the surgeonfish Ctenochaetus striatus. Marine Ecology Progress Series, 415:237–245, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/415/
m415p237.pdf.

Green:2010:RSE

[GB10b] T. J. Green and A. C. Barnes. Reduced salinity, but not estuarine acidification, is a cause of immune-suppression in the Sydney rock oyster Saccostrea glomerata. Marine Ecology Progress Series, 402:161–170, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/402/m402p161.pdf.

Golani:2012:DIP

[GB12] D. Golani and G. Bernardi. Differential invading potential among cryptic species of a Lessepsian bioinvader, the blotchfin dragonet Callionymus filamentosus. Marine Ecology Progress Series, 450:159–166, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/450/m450p159.pdf.

Gleason:2016:RPT

[GB16] L. U. Gleason and R. S. Burton. Regional patterns of thermal stress and constitutive gene expression in the marine snail Chlorostoma funebralis in northern and southern California. Marine Ecology Progress Series, 556:143–159, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/556/
m556p143.pdf.

Griffen:2016:WSI

[GBC^{+}16] B. D. Griffen, B. A. Belgrad, Z. J. Cannizzo, E. R. Knotts, and E. R. Hancock. As we see it: Rethinking our approach to multiple stressor studies in marine environments. Marine Ecology Progress Series, 543:273–281, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m543p273.pdf.

Gaspar:2012:ODJ

[GBD^{+}12] P. Gaspar, S. R. Benson, P. H. Dutton, A. Réveillère, G. Jacob, C. Meetoo, A. Dehecq, and S. Fossette. Oceanic dispersal of juvenile leatherback turtles: going beyond passive drift modeling. Marine Ecology Progress Series, 457:265–284, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m457p265.pdf.

Gutt:2013:ASS

[GBD13] J. Gutt, A. Böhmer, and W. Dimmler. Antarctic sponge spicule mats shape macrobenthic diversity and act as a silicon trap. Marine Ecology Progress Series, 480:57–71, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/480/m480p057.pdf.

Geraldi:2017:ABS

[GBE^{+}17] N. R. Geraldi, C. Bertolini, M. C. Emmerson, D. Roberts, J. D. Sigwart, and N. E. O’Connor. Aggregations of brittle stars can perform similar ecological roles as mussel reefs. Marine Ecology Progress Series, 563:157–167, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/563/m563p157.pdf.

Glud:2010:BEA

[GBH^{+}10] R. N. Glud, P. Berg, A. Hume, P. Batty, M. E. Blicher, K. Lennert, and S. Rysgaard. Benthic O 2 exchange across hard-bottom substrates quantified by eddy correlation in a sub-Arctic fjord. Marine Ecology Progress Series, 417:1–12, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Gerlach:2012:PSS

[GBH12] Gabriele Gerlach, Kendra C. Buresch, and Roger T. Hanlon. Population structure of the squid Doryteuthis pealeii on the eastern coast of the USA: Comment on Shaw et al. (2010). Marine Ecology Progress Series, 450:281–283, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/450/

m450p281.pdf. See [SHM^{+}10a].

Gillis:2014:PLS

[GBJ^{+}14] L. G. Gillis, T. J. Bouma, C. G. Jones, M. M. van Katwijk, I. Nagelkerken, C. J. L. Jeuken, P. M. J. Herman, and A. D. Ziegler. Potential for landscape-scale positive interactions among tropical marine ecosystems. Marine Ecology Progress Series, 503:289–303, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m503p289.pdf.

Gillis:2014:LTM

[GBK^{+}14] L. G. Gillis, T. J. Bouma, W. Kiswara, A. D. Ziegler, and P. M. J. Herman. Leaf transport in mimic mangrove forests and seagrass beds. Marine Ecology Progress Series, 498:95– 102, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/498/m498p095.pdf.

Gharouni:2015:SIS

[GBL^{+}15] A. Gharouni, M. A. Barbeau, A. Locke, L. Wang, and J. Watmough. Sensitivity of invasion speed to dispersal and demography: an application of spreading speed theory to the green crab invasion on the northwest Atlantic coast. Marine Ecology Progress Series, 541:135–150, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/541/m541p135.pdf.

Gaichas:2012:WDM

[GBM^{+}12] S. K. Gaichas, A. Bundy, T. J. Miller, E. Moksness, and Konstantinos I. Stergiou. What drives marine fisheries production? Marine Ecology Progress Series, 459:159–163, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m459p159.pdf.

Gardner:2012:ADO

[GBMW12] J. P. A. Gardner, C. Boesche, J. M. Meyer, and A. R. Wood. Analyses of DNA obtained from shells and brinepreserved meat of the giant clam Tridacna maxima from the central Pacific Ocean. Marine Ecology Progress Series, 453:297–301, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/453/m453p297.pdf.

Gibson:2011:FED

[GBR11] R. N. Gibson, M. T. Burrows, and L. Robb. Field experiments on depth selection by juvenile plaice Pleuronectes platessa. Marine Ecology Progress Series, 430:197–205, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m430p197.pdf.

Goberville:2010:CDC

[GBS^{+}10] E. Goberville, G. Beaugrand, B. Sautour, P. Tréguer, and SOMLIT Team. Climate-driven changes in coastal marine systems of western Europe. Marine Ecology Progress Series, 408:129–148, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/408/m408p129.pdf.

Garces:2011:CMC

[GBTS11] E. Garcés, G. Basterretxea, and A. Tovar-Sánchez. Changes in microbial communities in response to submarine groundwater input. Marine Ecology Progress Series, 438:47–58, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m438p047.pdf.

Garate:2015:CSP

[GBU15] L. Garate, A. Blanquer, and M. J. Uriz. Calcareous spherules produced by intracellular symbiotic bacteria protect the sponge Hemimycale columella from predation better than secondary metabolites. Marine Ecology Progress Series, 523:81–92, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m523p081.pdf.

Grubisic:2012:ESD

[GBW^{+}12] L. M. Grubisic, A. Brutemark, G. A. Weyhenmeyer, J. Wikner, U. Båmstedt, and S. Bertilsson. Effects of stratification depth and dissolved organic matter on brackish bacterioplankton communities. Marine Ecology Progress Series, 453: 37–48, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m453p037.pdf.

Gollety:2013:CBE

[GC13a] C. Golléty and T. P. Crowe. Contribution of biofilm to ecosystem functioning in rock pools with different macroalgal assemblages. Marine Ecology Progress Series, 482:69– 79, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/482/m482p069.pdf.

Green:2013:PBE

[GC13b] D. S. Green and T. P. Crowe. Physical and biological effects of introduced oysters on biodiversity in an intertidal boulder field. Marine Ecology Progress Series, 482:119– 132, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/482/m482p119.pdf.

Greco:2013:ELD

[GCBB13] M. Greco, A. Chiappetta, L. Bruno, and M. B. Bitonti. Effects of light deficiency on genome methylation in Posidonia oceanica. Marine Ecology Progress Series, 473:103– 114, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/473/m473p103.pdf.

Goss-Custard:2019:IHM

[GCBS19] J. D. Goss-Custard, K. M. Bowgen, and R. A. Stillman. Increasing the harvest for mussels Mytilus edulis without harming oystercatchers Haematopus ostralegus. Marine Ecology Progress Series, 612:101–110, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m612p101.pdf.

Gonzalez:2010:PPP

[GCC^{+}10] H. E. González, M. J. Calderón, L. Castro, A. Clement, L. A. Cuevas, G. Daneri, J. L. Iriarte, L. Lizárraga, R. Martínez, E. Menschel, N. Silva, C. Carrasco, C. Valenzuela, C. A. Vargas, and C. Molinet. Primary production and plankton dynamics in the Reloncaví Fjord and the Interior Sea of Chiloé, Northern Patagonia, Chile. Marine Ecology Progress Series, 402:13–30, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/402/m402p013.pdf.

Goutte:2014:IIA

[GCC^{+}14] A. Goutte, J. B. Charrassin, Y. Cherel, A. Carravieri, S. De Grissac, and G. Massé. Importance of ice algal production for top predators: new insights using sea-ice biomarkers. Marine Ecology Progress Series, 513:269–275, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m513p269.pdf.

Garrido:2016:ETG

[GCC^{+}16] S. Garrido, A. Cristóvão, C. Caldeira, R. Ben-Hamadou, N. Baylina, H. Batista, E. Saiz, M. A. Peck, P. Ré, and A. M. P. Santos. Effect of temperature on the growth, survival, development and foraging behaviour of Sardina pilchardus larvae. Marine Ecology Progress Series, 559:131–145, 2016. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m559p131.pdf.

Guerra-Castro:2016:SSV

[GCCCM16] E. J. Guerra-Castro, J. E. Conde, and J. J. Cruz-Motta. Scales of spatial variation in tropical benthic assemblages and their ecological relevance: epibionts on Caribbean mangrove roots as a model system. Marine Ecology Progress Series, 548:97–110, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/548/m548p097.pdf.

Guerra-Castro:2018:CSD

[GCCM18] E. J. Guerra-Castro and J. J. Cruz-Motta. Colonization and succession as drivers of small-scale spatial variability in epibionts on mangrove roots in the Southern Caribbean. Marine Ecology Progress Series, 588:15–27, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/588/
m588p015.pdf.

Greer:2019:LSO

[GCDH19] A. T. Greer, L. M. Chiaverano, J. G. Ditty, and F. J. Hernandez. In situ observations of fish larvae encased within a pelagic gelatinous matrix. Marine Ecology Progress Series, 614:209–214, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/614/m614p209.pdf.

Garden:2014:RDC

[GCFW14] C. J. Garden, K. Currie, C. I. Fraser, and J. M. Waters. Rafting dispersal constrained by an oceanographic boundary. Marine Ecology Progress Series, 501:297–302, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/501/
m501p297.pdf.

Guinand:2016:CGV

[GCL^{+}16] B. Guinand, C. Chauvel, M. Lechene, J. Tournois, C. S. Tsigenopoulos, A. M. Darnaude, D. J. McKenzie, and P. A. Gagnaire. Candidate gene variation in gilthead sea bream reveals complex spatiotemporal selection patterns between marine and lagoon habitats. Marine Ecology Progress Series, 558:115–127, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m558p115.pdf.

Gertseva:2010:GVS

[GCM10] V. V. Gertseva, J. M. Cope, and S. E. Matson. Growth variability in the splitnose rockfish Sebastes diploproa of the northeast Pacific Ocean: pattern revisited. Marine Ecology Progress Series, 413:125–136, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/413/m413p125.pdf.

Green:2017:CIM

[GCP^{+}17] D. S. Green, H. Christie, N. Pratt, B. Boots, J. A. Godbold, M. Solan, and C. Hauton. Competitive interactions moderate the effects of elevated temperature and atmospheric CO_{2} on the health and functioning of oysters. Marine Ecology Progress Series, 582:93–103, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m582p093.pdf.

Grieve:2016:REI

[GCR16] B. D. Grieve, E. N. Curchitser, and R. R. Rykaczewski. Range expansion of the invasive lionfish in the Northwest Atlantic with climate change. Marine Ecology Progress Series, 546:225–237, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/546/m546p225.pdf.

Gimenez:2017:IIN

[GCR^{+}17] J. Giménez, A. Cañadas, F. Ramírez, I. Afán, S. García-Tiscar, C. Fernández-Maldonado, J. J. Castillo, and R. de Stephanis. Intra- and interspecific niche partitioning in striped and common dolphins inhabiting the southwestern Mediterranean Sea. Marine Ecology Progress Series, 567:199–210, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2017/567/m567p199.pdf.

Green:2015:FDC

[GCRP15] D. B. Green, J. C. Coetzee, G. M. Rishworth, and P. A. Pistorius. Foraging distribution of Cape gannets in relation to oceanographic features, prey availability and marine protected areas. Marine Ecology Progress Series, 537:277– 288, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/537/m537p277.pdf.

Golikov:2018:OCS

[GCS^{+}18] A. V. Golikov, F. R. Ceia, R. M. Sabirov, Z. I. Zaripova, M. E. Blicher, D. V. Zakharov, and J. C. Xavier. Ontogenetic changes in stable isotope (δ ^{13}C and δ ^{1}5N) values in squid Gonatus fabricii (Cephalopoda) reveal its important ecological role in the Arctic. Marine Ecology Progress Series, 606:65–78, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/606/m606p065.pdf.

Golikov:2019:FST

[GCS^{+}19] A. V. Golikov, F. R. Ceia, R. M. Sabirov, A. N. Belyaev, M. E. Blicher, N. H. Arboe, D. V. Zakharov, and J. C. Xavier. Food spectrum and trophic position of an Arctic cephalopod, Rossia palpebrosa (Sepiolida), inferred by stomach contents and stable isotope (δ ^{13}C and δ ^{1}5N) analyses. Marine Ecology Progress Series, 632:131–144, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2019/632/m632p131.pdf.

Gillis:2018:FED

[GCSF18] A. J. Gillis, S. A. Ceriani, J. A. Seminoff, and M. M. P. B. Fuentes. Foraging ecology and diet selection of juvenile green turtles in the Bahamas: insights from stable isotope analysis and prey mapping. Marine Ecology Progress Series, 599:225–238, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/599/m599p225.pdf.

Genazzio:2015:PET

[GD15] M. A. Genazzio and M. J. Durako. Photochemical efficiency of Thalassia testudinum varies in response to repeated shading events and unpredictable weather. Marine Ecology Progress Series, 539:127–137, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2015/539/m539p127.pdf.

Gruss:2018:ISA

[GDA^{+}18] A. Grüss, M. D. Drexler, C. H. Ainsworth, J. J. Roberts, R. H. Carmichael, N. F. Putman, P. M. Richards, E. Chancellor, E. A. Babcock, and M. S. Love. Improving the spatial allocation of marine mammal and sea turtle biomasses in spatially explicit ecosystem models. Marine Ecology Progress Series, 602:255–274, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/602/m602p255.pdf.

Gama:2016:GTC

[GDB^{+}16] L. R. Gama, C. Domit, M. K. Broadhurst, M. M. P. B. Fuentes, and R. B. Millar. Green turtle Chelonia mydas foraging ecology at 25^{o}S in the western Atlantic: evidence to support a feeding model driven by intrinsic and extrinsic variability. Marine Ecology Progress Series, 542:209–219, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/542/
m542p209.pdf.

Gullage:2017:PDM

[GDE17] L. Gullage, R. Devillers, and E. Edinger. Predictive distribution modelling of cold-water corals in the Newfoundland and Labrador region. Marine Ecology Progress Series, 582: 57–77, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/582/m582p057.pdf.

Guery:2019:WEC

[GDH^{+}19] L. Guéry, S. Descamps, K. I. Hodges, R. Pradel, B. Moe, S. A. Hanssen, K. E. Erikstad, G. W. Gabrielsen, H. G. Gilchrist, S. Jenouvrier, and J. Bêty. Winter extratropical cyclone influence on seabird survival: variation between and within common eider Somateria mollissima populations. Marine Ecology Progress Series, 627:155–170, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/627/m627p155.pdf.

Grossman:2019:PPG

[GDJdBSL19] A. Grossman, F. G. Daura-Jorge, M. de Brito Silva, and G. O. Longo. Population parameters of green turtle adult males in the mixed ground of Atol das Rocas, Brazil. Marine Ecology Progress Series, 609:197–207, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/609/m609p197.pdf.

Garaffo:2011:MHU

[GDP^{+}11] G. V. Garaffo, S. L. Dans, S. N. Pedraza, M. Degrati, A. Schiavini, R. González, and E. A. Crespo. Modeling habitat use for dusky dolphin and Commerson’s dolphin in Patagonia. Marine Ecology Progress Series, 421:217–227, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/421/
m421p217.pdf.

Gross:2018:HUP

[GDPR18] C. Gross, C. Donoghue, C. Pruitt, and J. L. Ruesink. Habitat use patterns and edge effects across a seagrassunvegetated ecotone depend on species-specific behaviors and sampling methods. Marine Ecology Progress Series, 598: 21–33, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/598/m598p021.pdf.

Garcia:2013:PMD

[GDSPRA13] S. D. García, A. P. Diz, A. Sá-Pinto, and E. Rolán-Alvarez. Proteomic and morphological divergence in micro-allopatric morphotypes of Melarhaphe neritoides in the absence of genetic differentiation. Marine Ecology Progress Series, 475: 145–153, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/475/m475p145.pdf.

Gaston:2014:SDC

[GE14] A. J. Gaston and K. H. Elliott. Seabird diet changes in northern Hudson Bay, 1981–2013, reflect the availability of schooling prey. Marine Ecology Progress Series, 513:211–223, 2014.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m513p211.pdf.

Glazier:2017:GDA

[GE17] A. E. Glazier and R. J. Etter. Genetic divergence across an oxygen minimum zone. Marine Ecology Progress Series, 577:79–91, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/577/m577p079.pdf.

Geange:2010:ELH

[Gea10] S. W. Geange. Effects of larger heterospecifics and structural refuge on the survival of a coral reef fish, Thalassoma hardwicke. Marine Ecology Progress Series, 407:197– 207, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/407/m407p197.pdf.

Glynn:2014:EPR

[GEA^{+}14] P. W. Glynn, I. C. Enochs, J. A. Afflerbach, V. W. Brandtneris, and J. E. Serafy. Eastern Pacific reef fish responses to coral recovery following El Niño disturbances. Marine Ecology Progress Series, 495:233–247, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/495/m495p233.pdf.

Galland:2017:CCF

[GEAOH17] G. R. Galland, B. Erisman, O. Aburto-Oropeza, and P. A. Hastings. Contribution of cryptobenthic fishes to estimating community dynamics of sub-tropical reefs. Marine Ecology Progress Series, 584:175–184, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/584/m584p175.pdf.

Galloway:2014:QEI

[GED^{+}14] A. W. E. Galloway, M. E. Eisenlord, M. N. Dethier, G. W. Holtgrieve, and M. T. Brett. Quantitative estimates of isopod resource utilization using a Bayesian fatty acid mixing model. Marine Ecology Progress Series, 507:219–232, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/507/
m507p219.pdf.

Gochfeld:2012:DNE

[GEF^{+}12] D. J. Gochfeld, C. G. Easson, C. J. Freeman, R. W. Thacker, and J. B. Olson. Disease and nutrient enrichment as potential stressors on the Caribbean sponge Aplysina cauliformis and its bacterial symbionts. Marine Ecology Progress Series, 456:101–111, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/456/m456p101.pdf.

Garcia-Echauri:2019:RSS

[GELJ19] L. L. García-Echauri, G. Liggins, and A. Jeffs. Resistance to starvation in settling spiny lobsters in warming coastal waters. Marine Ecology Progress Series, 623:39–50, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/623/
m623p039.pdf.

Gelpi:2013:UDM

[GFC^{+}13] C. G. Gelpi, Jr., B. Fry, R. E. Condrey, J. W. Fleeger, and S. F. Dubois. Using δ ^{13}C and δ ^{15}N to determine the migratory history of offshore Louisiana blue crab spawning stocks. Marine Ecology Progress Series, 494:205–218, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/494/
m494p205.pdf.

Giddens:2014:ERI

[GFC^{+}14] J. Giddens, A. M. Friedlander, E. Conklin, C. Wiggins, K. Stamoulis, and M. K. Donovan. Experimental removal of the invasive peacock hind (roi) Cephalopholis argus, in Puakō, Hawai’i: methods for assessing and managing marine invasive species. Marine Ecology Progress Series, 511:209–221, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/511/
m511p209.pdf.

Girard:2016:MSO

[GFF16] F. Girard, B. Fu, and C. R. Fisher. Mutualistic symbiosis with ophiuroids limited the impact of the Deepwater Horizon oil spill on deep-sea octocorals. Marine Ecology Progress Series, 549:89–98, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m549p089.pdf.

Gasca:2012:HAC

[GFGGDSM12] R. Gasca, C. Franco-Gordo, E. Godínez-Domínguez, and E. Suárez-Morales. Hyperiid amphipod community in the Eastern Tropical Pacific before, during, and after El Niño 1997–1998. Marine Ecology Progress Series, 455:123–139, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps_oa/m455p123.pdf.

Galaiduk:2013:FDB

[GFKC13] R. Galaiduk, W. F. Figueira, M. J. Kingsford, and B. G. Curley. Factors driving the biogeographic distribution of two temperate Australian damselfishes and ramifications for range shifts. Marine Ecology Progress Series, 484:189–202, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/484/m484p189.pdf.

Galimany:2017:FCB

[GFL^{+}17] E. Galimany, C. J. Freeman, J. Lunt, A. Domingos, P. Sacks, and L. Walters. Feeding competition between the native oyster Crassostrea virginica and the invasive mussel Mytella charruana. Marine Ecology Progress Series, 564:57–66, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/564/m564p057.pdf.

Glockner-Fagetti:2019:LSS

[GFP19] A. Glockner-Fagetti and N. E. Phillips. Low salinity and sediment stress on sea urchin Evechinus chloroticus larvae has latent effects on juvenile performance. Marine Ecology Progress Series, 619:85–96, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/619/m619p085.pdf.

Goessling:2018:MLF

[GFR^{+}18] J. W. Goessling, S. Frankenbach, L. Ribeiro, J. Serôdio, and M. Kühl. Modulation of the light field related to valve optical properties of raphid diatoms: implications for niche differentiation in the microphytobenthos. Marine Ecology Progress Series, 588:29–42, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/588/m588p029.pdf.

Gahagan:2015:PMS

[GFS15] B. I. Gahagan, D. A. Fox, and D. H. Secor. Partial migration of striped bass: revisiting the contingent hypothesis. Marine Ecology Progress Series, 525:185–197, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/525/
m525p185.pdf.

Guo:2014:SVP

[GFW^{+}14] S. Guo, Y. Feng, L. Wang, M. Dai, Z. Liu, Y. Bai, and J. Sun. Seasonal variation in the phytoplankton community of a continental-shelf sea: the East China Sea. Marine Ecology Progress Series, 516:103–126, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m516p103.pdf.

Gosnell:2012:KII

[GG12] J. S. Gosnell and S. D. Gaines. Keystone intimidators in the intertidal: non-consumptive effects of a keystone sea star regulate feeding and growth in whelks. Marine Ecology Progress Series, 450:107–114, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/450/m450p107.pdf.

Griffith:2016:TCT

[GG16] A. W. Griffith and C. J. Gobler. Temperature controls the toxicity of the ichthyotoxic dinoflagellate Cochlodinium polykrikoides. Marine Ecology Progress Series, 545:63–76, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/545/m545p063.pdf.

Guillemette:2018:BFS

[GGB^{+}18] M. Guillemette, F. Grégoire, D. Bouillet, J. F. Rail, F. Bolduc, A. Caron, and D. Pelletier. Breeding failure of seabirds in relation to fish depletion: Is there one universal threshold of food abundance? Marine Ecology Progress Series, 587:235–245, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m587p235.pdf.

Goatley:2016:RSP

[GGCB16] C. H. R. Goatley, A. González-Cabello, and D. R. Bellwood. Reef-scale partitioning of cryptobenthic fish assemblages across the Great Barrier Reef, Australia. Marine Ecology Progress Series, 544:271–280, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/544/
m544p271.pdf.

Guttridge:2011:AIL

[GGD^{+}11] T. L. Guttridge, S. H. Gruber, J. D. DiBattista, K. A. Feldheim, D. P. Croft, S. Krause, and J. Krause. Assortative interactions and leadership in a free-ranging population of juvenile lemon shark Negaprion brevirostris. Marine Ecology Progress Series, 423:235–245, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/423/m423p235.pdf.

Gaichas:2012:ARA

[GGF^{+}12a] S. Gaichas, R. Gamble, M. Fogarty, H. Benoît, T. Essington, C. Fu, M. Koen-Alonso, and J. Link. Assembly rules for aggregate-species production models: simulations in support of management strategy evaluation. Marine Ecology Progress Series, 459:275–292, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m459p275.pdf.

Guttridge:2012:DDI

[GGF^{+}12b] T. L. Guttridge, S. H. Gruber, B. R. Franks, S. T. Kessel, K. S. Gledhill, J. Uphill, J. Krause, and D. W. Sims. Deep danger: intra-specific predation risk influences habitat use and aggregation formation of juvenile lemon sharks Negaprion brevirostris. Marine Ecology Progress Series, 445:279–291, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2011/445/m445p279.pdf.

Gonzalez:2014:NNM

[GGF^{+}14] D. J. Gonzalez, R. A. Gonzalez, B. A. Froelich, J. D. Oliver, R. T. Noble, and K. J. McGlathery. Non-native macroalga may increase concentrations of Vibrio bacteria on intertidal mudflats. Marine Ecology Progress Series, 505:29–36, 2014.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m505p029.pdf.

Gollner:2015:SMD

[GGFB15] S. Gollner, B. Govenar, C. R. Fisher, and M. Bright. Size matters at deep-sea hydrothermal vents: different diversity and habitat fidelity patterns of meio- and macrofauna. Marine Ecology Progress Series, 520:57–66, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m520p057.pdf.

Gomez:2018:SDP

[GGG18] C. G. Gomez, H. M. Guzman, and A. Gonzalez. Stability and dynamic properties of octocoral communities in the Tropical Eastern Pacific. Marine Ecology Progress Series, 588: 71–84, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/588/m588p071.pdf.

Gonzalez:2016:CFT

[GGK^{+}16] H. E. González, M. Graeve, G. Kattner, N. Silva, L. Castro, J. L. Iriarte, L. Osmán, G. Daneri, and C. A. Vargas. Carbon flow through the pelagic food web in southern Chilean Patagonia: relevance of Euphausia vallentini as a key species. Marine Ecology Progress Series, 557:91–110, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/557/
m557p091.pdf.

Galindo:2017:PCP

[GGL^{+}17] V. Galindo, M. Gosselin, J. Lavaud, C. J. Mundy, B. Else, J. Ehn, M. Babin, and S. Rysgaard. Pigment composition and photoprotection of Arctic sea ice algae during spring. Marine Ecology Progress Series, 585:49–69, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/585/
m585p049.pdf.

Gledhill:2015:HBQ

[GGLT15] M. Gledhill, L. J. A. Gerringa, P. Laan, and K. R. Timmermans. Heme-b quotas are low in Southern Ocean phytoplankton. Marine Ecology Progress Series, 532:29–40, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2015/532/m532p029.pdf.

Gomez-Gutierrez:2012:SGM

[GGMGR12] J. Gómez-Gutiérrez, S. Martínez-Gómez, and C. J. Robinson. Seasonal growth, molt, and egg production rates of Nyctiphanes simplex (Crustacea: Euphausiacea) juveniles and adults in the Gulf of California. Marine Ecology Progress Series, 455:173–194, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/455/m455p173.pdf.

Green-Gavrielidis:2018:BFM

[GGMTGC18] L. A. Green-Gavrielidis, F. MacKechnie, C. S. Thornber, and M. Gomez-Chiarri. Bloom-forming macroalgae (Ulva spp.) inhibit the growth of co-occurring macroalgae and decrease eastern oyster larval survival. Marine Ecology Progress Series, 595:27–37, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/595/m595p027.pdf.

Gladstone-Gallagher:2017:SDE

[GGNL^{+}17] R. V. Gladstone-Gallagher, H. R. Needham, A. M. Lohrer, C. J. Lundquist, and C. A. Pilditch. Site-dependent effects of bioturbator–detritus interactions alter soft-sediment ecosystem function. Marine Ecology Progress Series, 569: 145–161, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/569/m569p145.pdf.

Gosselin:2019:FEI

[GGPDS19] L. A. Gosselin, R. Gallego, J. Peters-Didier, and M. A. Sewell. Field evidence of interpopulation variation in oocyte size of a marine invertebrate under contrasting temperature and food availability. Marine Ecology Progress Series, 619: 69–84, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/619/m619p069.pdf.

Gouezo:2015:ITS

[GGvW^{+}15] M. Gouezo, Y. Golbuu, R. van Woesik, L. Rehm, S. Koshiba, and C. Doropoulos. Impact of two sequential super typhoons on coral reef communities in Palau. Marine Ecology Progress Series, 540:73–85, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/540/m540p073.pdf.

Gomez:2011:MFP

[GH11] I. Gómez and P. Huovinen. Morpho-functional patterns and zonation of South Chilean seaweeds: the importance of photosynthetic and bio-optical traits. Marine Ecology Progress Series, 422:77–91, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/422/m422p077.pdf.

Gilman:2015:BTI

[GHCC15] S. Gilman, H. Hayford, C. Craig, and E. Carrington. Body temperatures of an intertidal barnacle and two whelk predators in relation to shore height, solar aspect, and microhabitat. Marine Ecology Progress Series, 536:77–88, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/536/
m536p077.pdf.

Goni:2010:NCS

[GHD^{+}10] R. Goñi, R. Hilborn, D. Díaz, S. Mallol, and S. Adlerstein. Net contribution of spillover from a marine reserve to fishery catches. Marine Ecology Progress Series, 400:233– 243, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/400/m400p233.pdf.

Gabara:2018:RSL

[GHES18] S. S. Gabara, S. L. Hamilton, M. S. Edwards, and D. L. Steller. Rhodolith structural loss decreases abundance, diversity, and stability of benthic communities at Santa Catalina Island, CA. Marine Ecology Progress Series, 595:71–88, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/595/
m595p071.pdf.

Gallucci:2012:HAC

[GHGF12] F. Gallucci, P. Hutchings, P. Gribben, and G. Fonseca. Habitat alteration and community-level effects of an invasive ecosystem engineer: a case study along the coast of NSW, Australia.

Marine Ecology Progress Series, 449:95–108, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/449/
m449p095.pdf.

Grear:2011:MPR

[GHGG11] J. S. Grear, D. Borsay Horowitz, and R. Gutjahr-Gobell. Mysid population responses to resource limitation differ from those predicted by cohort studies. Marine Ecology Progress Series, 432:115–123, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/432/m432p115.pdf.

Gao:2012:RMP

[GHHH12] K. Gao, E. W. Helbling, D. P. Häder, and D. A. Hutchins. Responses of marine primary producers to interactions between ocean acidification, solar radiation, and warming. Marine Ecology Progress Series, 470:167–189, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m470p167.pdf.

Guarderas:2011:EEM

[GHL11] A. P. Guarderas, S. D. Hacker, and J. Lubchenco. Ecological effects of marine reserves in Latin America and the Caribbean. Marine Ecology Progress Series, 429:219–225, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/429/
m429p219.pdf.

Guest:2010:MSS

[GHNF10] M. A. Guest, A. J. Hirst, P. D. Nichols, and S. D. Frusher. Multi-scale spatial variation in stable isotope and fatty acid profiles amongst temperate reef species: implications for design and interpretation of trophic studies. Marine Ecology Progress Series, 410:25–41, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/410/m410p025.pdf.

Guannel:2015:TBD

[GHT^{+}15a] M. L. Guannel, D. Haring, M. J. Twiner, Z. Wang, A. E. Noble, P. A. Lee, M. A. Saito, and G. Rocap. Toxigenicity and biogeography of the diatom Pseudo-nitzschia across distinct environmental regimes in the South Atlantic Ocean. Marine Ecology Progress Series, 526:67–87, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/526/
m526p067.pdf.

Gunther:2015:LCO

[GHT15b] C. C. Günther, J. P. Herrmann, and A. Temming. Laboratory calibration of optimal growth to deduce in situ feeding conditions of early juvenile sprat Sprattus sprattus from otoliths. Marine Ecology Progress Series, 525:199–215, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/525/
m525p199.pdf.

Gosselin:2010:ESR

[GJ10] L. A. Gosselin and L. A. Jones. Effects of solar radiation on barnacle settlement, early post-settlement mortality and community development in the intertidal zone. Marine Ecology Progress Series, 407:149–158, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/407/m407p149.pdf.

Gotz:2013:ADD

[GJ13] T. Götz and V. M. Janik. Acoustic deterrent devices to prevent pinniped depredation: efficiency, conservation concerns and possible solutions. Marine Ecology Progress Series, 492:285–302, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/492/m492p285.pdf.

Gardiner:2016:HSS

[GJ16] N. M. Gardiner and G. P. Jones. Habitat specialisation, site fidelity and sociality predict homing success in coral reef cardinalfish. Marine Ecology Progress Series, 558:81– 96, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/558/m558p081.pdf.

Goldstein:2018:DDW

[GJSR18] J. Goldstein, C. Jürgensen, U. K. Steiner, and H. U. Riisgård. Density-driven water exchange controls seasonal declines in jellyfish (Aurelia aurita) abundance in a shallow fjord system.

Marine Ecology Progress Series, 597:137–145, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/597/
m597p137.pdf.

Gould:2010:DGR

[GK10] A. L. Gould and W. J. Kimmerer. Development, growth, and reproduction of the cyclopoid copepod Limnoithona tetraspina in the upper San Francisco Estuary. Marine Ecology Progress Series, 412:163–177, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/412/m412p163.pdf.

Gilad:2018:ULS

[GKBEF18] E. Gilad, S. M. Kidwell, Y. Benayahu, and Y. Edelman-Furstenberg. Unrecognized loss of seagrass communities based on molluscan death assemblages: historic baseline shift in tropical Gulf of Aqaba, Red Sea. Marine Ecology Progress Series, 589:73–83, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/589/m589p073.pdf.

Ghedini:2011:PES

[GKC11] G. Ghedini, J. C. Klein, and R. A. Coleman. Potential effects of storm-water run-off on assemblages of mobile invertebrates. Marine Ecology Progress Series, 439:169–180, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/439/
m439p169.pdf.

Ghazilou:2019:IBP

[GKE^{+}19] A. Ghazilou, E. Koochaknejad, H. Ershadifar, H. Negarestan, K. Kor, and G. Baskaleh. Infestation biology of Phallusia nigra (Tunicata, Phlebobranchia) on hard corals in a subtropical bay. Marine Ecology Progress Series, 626:135– 143, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/626/m626p135.pdf.

Germain:2013:NIF

[GKHM13] L. R. Germain, P. L. Koch, J. Harvey, and M. D. Mc-Carthy. Nitrogen isotope fractionation in amino acids from harbor seals: implications for compound-specific trophic position calculations. Marine Ecology Progress Series, 482:265– 277, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/482/m482p265.pdf.

Gerstenmaier:2016:GDN

[GKHS16] C. E. Gerstenmaier, S. A. Krueger-Hadfield, and E. E. Sotka. Genotypic diversity in a non-native ecosystem engineer has variable impacts on productivity. Marine Ecology Progress Series, 556:79–89, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/556/m556p079.pdf.

Greenfield:2016:MFG

[GKPT16] B. L. Greenfield, C. Kraan, C. A. Pilditch, and S. F. Thrush. Mapping functional groups can provide insight into ecosystem functioning and potential resilience of intertidal sandflats. Marine Ecology Progress Series, 548:1–10, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Greene:2015:FYC

[GKR^{+}15] C. Greene, L. Kuehne, C. Rice, K. Fresh, and D. Penttila. Forty years of change in forage fish and jellyfish abundance across greater Puget Sound, Washington (USA): anthropogenic and climate associations. Marine Ecology Progress Series, 525:153–170, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m525p153.pdf.

Griffiths:2019:EAS

[GKRG^{+}19] S. P. Griffiths, K. Kesner-Reyes, C. Garilao, L. M. Duffy, and M. H. Román. Ecological assessment of the sustainable impacts of fisheries (EASI-Fish): a flexible vulnerability assessment approach to quantify the cumulative impacts of fishing in data-limited settings. Marine Ecology Progress Series, 625:89–113, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/625/m625p089.pdf.

Giarrizzo:2010:SSM

[GKW10] T. Giarrizzo, U. Krumme, and W. Wosniok. Size-structured migration and feeding patterns in the banded puffer fish Colomesus psittacus (Tetraodontidae) from north Brazilian mangrove creeks. Marine Ecology Progress Series, 419:157– 170, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/419/m419p157.pdf.

Gamble:2012:UAP

[GL12] R. J. Gamble and J. S. Link. Using an aggregate production simulation model with ecological interactions to explore effects of fishing and climate on a fish community. Marine Ecology Progress Series, 459:259–274, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m459p259.pdf.

Gambrel:2016:ICA

[GL16] B. Gambrel and H. R. Lasker. Interactions in the canopy among Caribbean reef octocorals. Marine Ecology Progress Series, 546:85–95, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/546/m546p085.pdf.

Glaser:2010:IVP

[Gla10] S. M. Glaser. Interdecadal variability in predator–prey interactions of juvenile North Pacific albacore in the California Current system. Marine Ecology Progress Series, 414: 209–221, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/414/m414p209.pdf.

Grangere:2010:MSH

[GLB^{+}10] K. Grangeré, S. Lefebvre, C. Bacher, P. Cugier, and A. Ménesguen. Modelling the spatial heterogeneity of ecological processes in an intertidal estuarine bay: dynamic interactions between bivalves and phytoplankton. Marine Ecology Progress Series, 415:141–158, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/415/m415p141.pdf.

Galimany:2017:FBE

[GLF^{+}17] E. Galimany, J. Lunt, C. J. Freeman, S. Reed, I. Segura-García, and V. J. Paul. Feeding behavior of eastern oysters Crassostrea virginica and hard clams Mercenaria mercenaria in shallow estuaries. Marine Ecology Progress Series, 567:125–137, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/567/m567p125.pdf.

Graba-Landry:2018:OWG

[GLHM^{+}18] A. Graba-Landry, A. S. Hoey, J. K. Matley, H. Sheppard-Brennand, A. G. B. Poore, M. Byrne, and S. A. Dworjanyn. Ocean warming has greater and more consistent negative effects than ocean acidification on the growth and health of subtropical macroalgae. Marine Ecology Progress Series, 595: 55–69, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/595/m595p055.pdf.

Gaos:2018:RCM

[GLJ^{+}18] A. R. Gaos, R. L. Lewison, M. P. Jensen, M. J. Liles, A. Henriquez, S. Chavarria, C. M. Pacheco, M. Valle, D. Melero, V. Gadea, E. Altamirano, P. Torres, F. Vallejo, C. Miranda, C. LeMarie, J. Lucero, K. Oceguera, D. Chácon, L. Fonseca, M. Abrego, J. A. Seminoff, E. E. Flores, I. Llamas, R. Donadi, B. Peña, J. P. Muñoz, D. A. Ruales, J. A. Chaves, S. Otterstrom, A. Zavala, C. E. Hart, R. Brittain, J. Alfaro-Shigueto, J. Mangel, I. L. Yañez, and P. H. Dutton. Rookery contributions, movements and conservation needs of hawksbill turtles at foraging grounds in the eastern Pacific Ocean. Marine Ecology Progress Series, 586:203–216, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/586/
m586p203.pdf.

Gartner:2010:LRD

[GLM^{+}10] A. Gartner, P. S. Lavery, K. McMahon, A. Brearley, and H. Barwick. Light reductions drive macroinvertebrate changes in Amphibolis griffithii seagrass habitat. Marine Ecology Progress Series, 401:87–100, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/401/m401p087.pdf.

Girard:2016:CBI

[GLM16] F. Girard, M. Lacharité, and A. Metaxas. Colonization of benthic invertebrates in a submarine canyon in the NW Atlantic. Marine Ecology Progress Series, 544:53–64, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/544/
m544p053.pdf.

Guilpin:2019:FEP

[GLM^{+}19] M. Guilpin, V. Lesage, I. McQuinn, J. A. Goldbogen, J. Potvin, T. Jeanniard du Dot, T. Doniol-Valcroze, R. Michaud, M. Moisan, and G. Winkler. Foraging energetics and prey density requirements of western North Atlantic blue whales in the Estuary and Gulf of St. Lawrence, Canada. Marine Ecology Progress Series, 625:205–223, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/625/
m625p205.pdf.

Gavrilchuk:2014:TNP

[GLR^{+}14] K. Gavrilchuk, V. Lesage, C. Ramp, R. Sears, M. Bérubé, S. Bearhop, and G. Beauplet. Trophic niche partitioning among sympatric baleen whale species following the collapse of groundfish stocks in the Northwest Atlantic. Marine Ecology Progress Series, 497:285–301, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m497p285.pdf.

Galindo:2015:UIM

[GLS^{+}15] V. Galindo, M. Levasseur, M. Scarratt, C. J. Mundy, M. Gosselin, R. P. Kiene, M. Gourdal, and M. Lizotte. Under-ice microbial dimethylsulfoniopropionate metabolism during the melt period in the Canadian Arctic Archipelago. Marine Ecology Progress Series, 524:39–53, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m524p039.pdf.

Gaos:2012:SEC

[GLW^{+}12] A. R. Gaos, R. L. Lewison, B. P. Wallace, I. L. Yañez, M. J. Liles, W. J. Nichols, A. Baquero, C. R. Hasbún, M. Vasquez, J. Urteaga, and J. A. Seminoff. Spatial ecology of critically endangered hawksbill turtles Eretmochelys imbricata: implications for management and conservation. Marine Ecology Progress Series, 450:181–194, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/450/m450p181.pdf.

Gray:2017:MUO

[GLW^{+}17] M. W. Gray, C. J. Langdon, G. G. Waldbusser, B. Hales, and S. Kramer. Mechanistic understanding of ocean acidification impacts on larval feeding physiology and energy budgets of the mussel Mytilus californianus. Marine Ecology Progress Series, 563:81–94, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/563/m563p081.pdf.

Goecke:2010:CIB

[GLWI10] F. Goecke, A. Labes, J. Wiese, and J. F. Imhoff. Chemical interactions between marine macroalgae and bacteria. Marine Ecology Progress Series, 409:267–299, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/409/
m409p267.pdf.

Guo:2014:RMG

[GLY^{+}14] C. Guo, H. Liu, J. Yu, S. Zhang, and C. J. Wu. Role of microzooplankton grazing in regulating phytoplankton biomass and community structure in response to atmospheric aerosol input. Marine Ecology Progress Series, 507:69–79, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/507/
m507p069.pdf.

Graham:2010:MLN

[GM10] S. A. Graham and I. A. Mendelssohn. Multiple levels of nitrogen applied to an oligohaline marsh identify a plant community response sequence to eutrophication. Marine Ecology Progress Series, 417:73–82, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/417/m417p073.pdf.

Gao:2013:GMV

[GM13a] J. Gao and S. B. Munch. Genetic and maternal variation in early growth in the Atlantic silverside Menidia menidia. Marine Ecology Progress Series, 485:211–222, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/485/
m485p211.pdf.

Gulbransen:2013:NTM

[GM13b] D. Gulbransen and K. McGlathery. Nitrogen transfers mediated by a perennial, non-native macroalga: a ^{15}N tracer study. Marine Ecology Progress Series, 482:299–304, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/482/
m482p299.pdf.

Genitsaris:2016:SSV

[GMB^{+}16] S. Genitsaris, S. Monchy, E. Breton, E. Lecuyer, and U. Christaki. Small-scale variability of protistan planktonic communities relative to environmental pressures and biotic interactions at two adjacent coastal stations. Marine Ecology Progress Series, 548:61–75, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m548p061.pdf.

Gonzalez-Mirelis:2011:IBC

[GMBL11] G. Gonzalez-Mirelis, P. Bergström, and M. Lindegarth. Interaction between classification detail and prediction of community types: implications for predictive modelling of benthic biotopes. Marine Ecology Progress Series, 432:31–44, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2011/432/m432p031.pdf.

Gacia:2012:TIS

[GMC^{+}12] E. Gacia, N. Marbà, J. Cebrián, R. Vaquer-Sunyer, N. Garcias-Bonet, and C. M. Duarte. Thresholds of irradiance for seagrass Posidonia oceanica meadow metabolism. Marine Ecology Progress Series, 466:69–79, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m466p069.pdf.

Graff:2016:POP

[GMD16] J. R. Graff and S. Menden-Deuer. Physical and optical properties of phytoplankton-rich layers in a coastal fjord: a step toward prediction and strategic sampling of plankton patchiness. Marine Ecology Progress Series, 544:1–14, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Giorgini:2019:RBI

[GME^{+}19] M. Giorgini, A. Miguez, K. S. Esquius, C. Diaz de Astarloa, O. Iribarne, E. Fanjul, and M. Escapa. Regenerative bioturbation by intertidal burrowing crabs modifies microphytobenthic composition and enhances primary production in SW Atlantic mudflats. Marine Ecology Progress Series, 632:43–57, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/632/m632p043.pdf.

Guillaumot:2018:MIS

[GMES18] C. Guillaumot, A. Martin, M. Eléaume, and T. Saucède. Methods for improving species distribution models in datapoor areas: example of sub-Antarctic benthic species on the Kerguelen Plateau. Marine Ecology Progress Series, 594:149– 164, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/594/m594p149.pdf.

Gambi:2016:DFT

[GMG^{+}16] M. C. Gambi, L. Musco, A. Giangrande, F. Badalamenti, F. Micheli, and K. J. Kroeker. Distribution and functional traits of polychaetes in a CO_{2} vent system: winners and losers among closely related species. Marine Ecology Progress Series, 550:121–134, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/550/m550p121.pdf.

George:2019:FSM

[GMHS19] L. W. George, A. P. B. Martins, M. R. Heupel, and C. A. Simpfendorfer. Fine-scale movements of juvenile blacktip reef sharks Carcharhinus melanopterus in a shallow nearshore nursery. Marine Ecology Progress Series, 623:85–97, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2019/623/m623p085.pdf.

Goetz:2018:HPD

[GMK18] K. T. Goetz, B. I. McDonald, and G. L. Kooyman. Habitat preference and dive behavior of non-breeding emperor penguins in the eastern Ross Sea, Antarctica. Marine Ecology Progress Series, 593:155–171, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/593/m593p155.pdf.

Grant:2019:ESS

[GMK^{+}19] N. Grant, E. Matveev, A. S. Kahn, S. K. Archer, A. Dunham, R. J. Bannister, D. Eerkes-Medrano, and S. P. Leys. Effect of suspended sediments on the pumping rates of three species of glass sponge in situ. Marine Ecology Progress Series, 615:79–100, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m615p079.pdf.

Gooley:2010:RCD

[GMM10] T. A. Gooley, D. J. Marshall, and K. Monro. Responses to conspecific density in an arborescent bryozoan. Marine Ecology Progress Series, 415:83–90, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/415/m415p083.pdf.

Gill:2011:BWH

[GMP^{+}11] P. C. Gill, M. G. Morrice, B. Page, R. Pirzl, A. H. Levings, and M. Coyne. Blue whale habitat selection and within-season distribution in a regional upwelling system off southern Australia. Marine Ecology Progress Series, 421:243–263, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m421p243.pdf.

Gusmao:2013:FBS

[GMR13a] L. F. M. Gusmão, A. D. McKinnon, and A. J. Richardson. Female-biased sex ratios in marine pelagic copepods: Response to Hirst et al. (2013). Marine Ecology Progress Series, 489:299–301, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m489p299.pdf.

Gusmao:2013:NEP

[GMR13b] Luiz Felipe Mendes Gusmão, A. David McKinnon, and Anthony J. Richardson. No evidence of predation causing femalebiased sex ratios in marine pelagic copepods. Marine Ecology Progress Series, 482:265–277, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/482/m482p265.pdf. See

comment [HBCK13].

Gouraguine:2019:IUD

[GMR^{+}19] A. Gouraguine, C. J. Melián, O. Reñones, H. Hinz, H. Baxter, L. Cardona, and J. Moranta. Implications of using different metrics for niche analysis in ecological communities. Marine Ecology Progress Series, 630:1–12, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Gosnell:2014:DOF

[GMSC14] J. S. Gosnell, R. J. A. Macfarlan, N. T. Shears, and J. E. Caselle. A dynamic oceanographic front drives biogeographical structure in invertebrate settlement along Santa Cruz Island, California, USA. Marine Ecology Progress Series, 507:181–196, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m507p181.pdf.

Garcia-Munoz:2014:FCP

[GMSL^{+}14] C. García-Muñoz, C. Sobrino, L. M. Lubián, C. M. García, S. Martínez-García, and P. Sangrà. Factors controlling phytoplankton physiological state around the South Shetland Islands (Antarctica). Marine Ecology Progress Series, 498: 55–71, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/498/m498p055.pdf.

Gontikaki:2011:PLD

[GMT^{+}11] E. Gontikaki, D. J. Mayor, B. Thornton, K. Black, and U. Witte. Processing of ^{13}C-labelled diatoms by a bathyal community at sub-zero temperatures. Marine Ecology Progress Series, 421:39–50, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/421/m421p039.pdf.

Gaillard:2015:DTB

[GMT^{+}15] B. Gaillard, T. Meziane, R. Tremblay, P. Archambault, K. K. S. Layton, A. L. Martel, and F. Olivier. Dietary tracers in Bathyarca glacialis from contrasting trophic regions in the Canadian Arctic. Marine Ecology Progress Series, 536:175–186, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/536/m536p175.pdf.

Gaillard:2017:FRB

[GMT^{+}17] B. Gaillard, T. Meziane, R. Tremblay, P. Archambault, M. E. Blicher, L. Chauvaud, S. Rysgaard, and F. Olivier. Food resources of the bivalve Astarte elliptica in a sub-Arctic fjord: a multi-biomarker approach. Marine Ecology Progress Series, 567:139–156, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/567/m567p139.pdf.

Ganias:2011:BSS

[GN11] K. Ganias and C. Nunes. Bathymetric segregation of spawning stages in the Atlantic sardine Sardina pilchardus. Marine Ecology Progress Series, 428:235–244, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/428/
m428p235.pdf.

Grol:2011:PEJ

[GNBM11] M. G. G. Grol, I. Nagelkerken, N. Bosch, and E. H. Meesters. Preference of early juveniles of a coral reef fish for distinct lagoonal microhabitats is not related to common measures of structural complexity. Marine Ecology Progress Series, 432:221–233, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m432p221.pdf.

Griffin:2010:CEE

[GNC^{+}10] J. N. Griffin, L. M. L. J. Noël, T. P. Crowe, M. T. Burrows, S. J. Hawkins, R. C. Thompson, and S. R. Jenkins. Consumer effects on ecosystem functioning in rock pools: roles of species richness and composition. Marine Ecology Progress Series, 420:45–56, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/420/m420p045.pdf.

Grahl-Nielsen:2011:FAH

[GNHLN11] O. Grahl-Nielsen, T. Haug, U. Lindstrøm, and K. T. Nilssen. Fatty acids in harp seal blubber do not necessarily reflect their diet. Marine Ecology Progress Series, 426:263–276, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2011/426/m426p263.pdf.

Giron-Nava:2017:QAL

[GNJJ^{+}17] A. Giron-Nava, C. C. James, A. F. Johnson, D. Dannecker, B. Kolody, A. Lee, M. Nagarkar, G. M. Pao, H. Ye, D. G. Johns, and G. Sugihara. Quantitative argument for longterm ecological monitoring. Marine Ecology Progress Series, 572:269–274, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m572p269.pdf.

Gochfeld:2010:TDF

[Goc10] D. J. Gochfeld. Territorial damselfishes facilitate survival of corals by providing an associational defense against predators. Marine Ecology Progress Series, 398:137–148, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/398/
m398p137.pdf.

Gilby:2018:SMS

[GOC^{+}18] B. L. Gilby, A. D. Olds, R. M. Connolly, P. S. Maxwell, C. J. Henderson, and T. A. Schlacher. Seagrass meadows shape fish assemblages across estuarine seascapes. Marine Ecology Progress Series, 588:179–189, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/588/m588p179.pdf.

Gonzalez-Ortegon:2014:EMP

[GOG14] E. González-Ortegón and L. Giménez. Environmentally mediated phenotypic links and performance in larvae of a marine invertebrate. Marine Ecology Progress Series, 502:185– 195, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/502/m502p185.pdf.

Gianelli:2019:EOW

[GOM^{+}19] I. Gianelli, L. Ortega, Y. Marín, A. R. Piola, and O. Defeo. Evidence of ocean warming in Uruguay’s fisheries landings: the mean temperature of the catch approach. Marine Ecology Progress Series, 625:115–125, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m625p115.pdf.

Gauthier:2014:SCA

[GOO14] S. Gauthier, J. Oeffner, and R. L. O’Driscoll. Species composition and acoustic signatures of mesopelagic organisms in a subtropical convergence zone, the New Zealand Chatham Rise. Marine Ecology Progress Series, 503:23–40, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/503/
m503p023.pdf.

Goyert:2014:RAP

[Goy14] H. F. Goyert. Relationship among prey availability, habitat, and the foraging behavior, distribution, and abundance of common terns Sterna hirundo and roseate terns S. dougallii. Marine Ecology Progress Series, 506:291–302, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/506/
m506p291.pdf.

Gennaro:2011:SBT

[GP11a] P. Gennaro and L. Piazzi. Synergism between two anthropic impacts: Caulerpa racemosa var. cylindracea invasion and seawater nutrient enrichment. Marine Ecology Progress Series, 427:59–70, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/427/m427p059.pdf.

Geraldi:2011:SCP

[GP11b] N. R. Geraldi and S. P. Powers. Subtle changes in prey foraging behavior have cascading effects in a shallow estuary. Marine Ecology Progress Series, 427:51–58, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/427/
m427p051.pdf.

Goshima:2012:BBA

[GP12] S. Goshima and C. H. Peterson. Both below- and aboveground shoalgrass structure influence whelk predation on hard clams. Marine Ecology Progress Series, 451:75–92, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m451p075.pdf.

Gomes:2016:WMU

[GPA^{+}16] I. Gomes, L. G. Peteiro, R. Albuquerque, R. Nolasco, J. Dubert, S. E. Swearer, and H. Queiroga. Wandering mussels: using natural tags to identify connectivity patterns among Marine Protected Areas. Marine Ecology Progress Series, 552:159–176, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/552/m552p159.pdf.

Gutierrez:2015:WAL

[GPB^{+}15] J. L. Gutiérrez, M. G. Palomo, M. Bagur, L. P. Arribas, and S. A. Soria. Wave action limits crowding in an intertidal mussel. Marine Ecology Progress Series, 518:153– 163, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/518/m518p153.pdf.

Guerreiro:2015:HTE

[GPC^{+}15] M. Guerreiro, R. A. Phillips, Y. Cherel, F. R. Ceia, P. Alvito, R. Rosa, and J. C. Xavier. Habitat and trophic ecology of Southern Ocean cephalopods from stable isotope analyses. Marine Ecology Progress Series, 530:119–134, 2015. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m530p119.pdf.

Gislason:2013:GVN

[GPM^{+}13] Ó. S. Gíslason, S. Pálsson, N. J. McKeown, H. P. Halldórsson, P. W. Shaw, and J. Svavarsson. Genetic variation in a newly established population of the Atlantic rock crab Cancer irroratus in Iceland. Marine Ecology Progress Series, 494:219–230, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/494/m494p219.pdf.

Grabowski:2017:PIV

[GPRR17] J. H. Grabowski, S. P. Powers, H. Roman, and S. Rouhani. Potential impacts of the 2010 Deepwater Horizon oil spill on subtidal oysters in the Gulf of Mexico. Marine Ecology Progress Series, 576:163–174, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m576p163.pdf.

Gyory:2013:TTL

[GPS13] J. Gyory, J. Pineda, and A. Solow. Turbidity triggers larval release by the intertidal barnacle Semibalanus balanoides. Marine Ecology Progress Series, 476:141–151, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/476/
m476p141.pdf.

Gihwala:2017:DIF

[GPV17] K. N. Gihwala, D. Pillay, and M. Varughese. Differential impacts of foraging plasticity by greater flamingo Phoenicopterus roseus on intertidal soft sediments. Marine Ecology Progress Series, 569:227–242, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/569/m569p227.pdf.

Goldsworthy:2010:SDS

[GPW^{+}10] S. D. Goldsworthy, B. Page, A. Welling, M. Chambellant, and C. J. A. Bradshaw. Selection of diving strategy by Antarctic fur seals depends on where and when foraging takes place. Marine Ecology Progress Series, 409:255–266, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/409/
m409p255.pdf.

Givens:2015:CGM

[GRBH15] C. E. Givens, B. Ransom, N. Bano, and J. T. Hollibaugh. Comparison of the gut microbiomes of 12 bony fish and 3 shark species. Marine Ecology Progress Series, 518:209– 223, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/518/m518p209.pdf.

Gong:2018:SDF

[GRCH^{+}18] Y. Gong, R. I. Ruiz-Cooley, M. E. Hunsicker, Y. Li, and X. Chen. Sexual dimorphism in feeding apparatus and niche partitioning in juvenile jumbo squid Dosidicus gigas. Marine Ecology Progress Series, 607:99–112, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/607/
m607p099.pdf.

Garcia-Robledo:2010:FLI

[GRCP^{+}10] E. García-Robledo, A. Corzo, S. Papaspyrou, J. L. Jiménez-Arias, and D. Villahermosa. Freeze-lysable inorganic nutrients in intertidal sediments: dependence on microphytobenthos abundance. Marine Ecology Progress Series, 403:155– 163, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m403p155.pdf.

Graff:2011:BAP

[GRD11] J. R. Graff, J. E. B. Rines, and P. L. Donaghay. Bacterial attachment to phytoplankton in the pelagic marine environment. Marine Ecology Progress Series, 441:15–24, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/441/
m441p015.pdf.

Grey:2010:ELE

[Gre10] E. K. Grey. Effects of large enemies on success of exotic species in marine fouling communities of Washington, USA. Marine Ecology Progress Series, 411:89–100, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/411/
m411p089.pdf.

Gonzalez-Rivero:2012:IMC

[GRFSM12] M. González-Rivero, R. Ferrari, C. H. L. Schönberg, and P. J. Mumby. Impacts of macroalgal competition and parrotfish predation on the growth of a common bioeroding sponge. Marine Ecology Progress Series, 444:133–142, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/444/
m444p133.pdf.

Gardes:2012:EMA

[GRG^{+}12] A. Gärdes, Y. Ramaye, H. P. Grossart, U. Passow, and M. S. Ullrich. Effects of Marinobacter adhaerens HP15 on polymer exudation by Thalassiosira weissflogii at different N: P ratios. Marine Ecology Progress Series, 461:1–14, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Garcia:2014:MPF

[GRL^{+}14] J. Garcia, Y. Rousseau, H. Legrand, G. Saragoni, and P. Lenfant. Movement patterns of fish in a Martinique MPA: implications for marine reserve design. Marine Ecology Progress Series, 513:171–185, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/513/m513p171.pdf.

Golet:2015:PPW

[GRL^{+}15] W. J. Golet, N. R. Record, S. Lehuta, M. Lutcavage, B. Galuardi, A. B. Cooper, and A. J. Pershing. The paradox of the pelagics: why bluefin tuna can go hungry in a sea of plenty. Marine Ecology Progress Series, 527:181–192, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m527p181.pdf.

Grol:2014:GPP

[GRN14] M. G. G. Grol, A. L. Rypel, and I. Nagelkerken. Growth potential and predation risk drive ontogenetic shifts among nursery habitats in a coral reef fish. Marine Ecology Progress Series, 502:229–244, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m502p229.pdf.

Giles:2016:GPD

[GRN^{+}16] J. L. Giles, C. Riginos, G. J. P. Naylor, Dharmadi, and J. R. Ovenden. Genetic and phenotypic diversity in the wedgefish Rhynchobatus australiae, a threatened ray of high value in the shark fin trade. Marine Ecology Progress Series, 548:165–180, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/548/m548p165.pdf.

Green:2011:MSA

[GS11] K. M. Green and R. M. Starr. Movements of small adult black rockfish: implications for the design of MPAs. Marine Ecology Progress Series, 436:219–230, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m436p219.pdf.

Gravestock:2015:VES

[GS15] P. Gravestock and C. Sheppard. Valuing the ecosystem services of the Chagos: a review of challenges and estimates. Marine Ecology Progress Series, 530:255–270, 2015. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m530p255.pdf.

Gongol:2016:SVR

[GS16] C. Gongol and C. Savage. Spatial variation in rates of benthic denitrification and environmental controls in four New Zealand estuaries. Marine Ecology Progress Series, 556:59– 77, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/556/m556p059.pdf.

Glaspie:2017:RHP

[GS17] C. N. Glaspie and R. D. Seitz. Role of habitat and predators in maintaining functional diversity of estuarine bivalves. Marine Ecology Progress Series, 570:113–125, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/570/
m570p113.pdf.

Ginther:2018:LRE

[GS18] S. C. Ginther and M. A. Steele. Limited recruitment of an ecologically and economically important fish, Paralabrax clathratus, to an invasive alga. Marine Ecology Progress Series, 602:213–224, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/602/m602p213.pdf.

Gallo:2018:RDA

[GSA^{+}18] F. Gallo, K. G. Schulz, E. B. Azevedo, J. Madruga, and J. Barcelos e Ramos. Responses of the diatom Asterionellopsis glacialis to increasing sea water CO_{2} concentrations and turbulence. Marine Ecology Progress Series, 589:33– 44, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/589/m589p033.pdf.

Garcia-Seoane:2016:ADL

[GSÁCM^{+}16] E. García-Seoane, G. Álvarez-Colombo, J. Miquel, J. M. Rodríguez, C. Guevara-Fletcher, P. Álvarez, and F. Saborido-Rey. Acoustic detection of larval fish aggregations in Galician waters (NW Spain). Marine Ecology Progress Series, 551: 31–44, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/551/m551p031.pdf.

Georgian:2014:HRE

[GSC14] S. E. Georgian, W. Shedd, and E. E. Cordes. High-resolution ecological niche modelling of the cold-water coral Lophelia pertusa in the Gulf of Mexico. Marine Ecology Progress Series, 506:145–161, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/506/m506p145.pdf.

Gilman:2017:DGT

[GSC17] E. Gilman, P. Suuronen, and M. Chaloupka. Discards in global tuna fisheries. Marine Ecology Progress Series, 582: 231–252, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/582/m582p231.pdf.

Gallagher:2014:PSR

[GSCH14] A. J. Gallagher, J. E. Serafy, S. J. Cooke, and N. Hammerschlag. Physiological stress response, reflex impairment, and survival of five sympatric shark species following experimental capture and release. Marine Ecology Progress Series, 496:207–218, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/496/m496p207.pdf.

Gauthier:2010:MMM

[GSD10] O. Gauthier, J. Sarrazin, and D. Desbruyères. Measure and mis-measure of species diversity in deep-sea chemosynthetic communities. Marine Ecology Progress Series, 402: 285–302, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/402/m402p285.pdf.

Gleiber:2012:TSV

[GSD12] M. R. Gleiber, D. K. Steinberg, and H. W. Ducklow. Time series of vertical flux of zooplankton fecal pellets on the continental shelf of the western Antarctic Peninsula. Marine Ecology Progress Series, 471:23–36, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/471/m471p023.pdf.

Geraldi:2013:AJO

[GSF^{+}13] N. R. Geraldi, M. Simpson, S. R. Fegley, P. Holmlund, and C. H. Peterson. Addition of juvenile oysters fails to enhance oyster reef development in Pamlico Sound. Marine Ecology Progress Series, 480:119–129, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m480p119.pdf.

Goodell:2018:CRS

[GSF18] W. Goodell, K. A. Stamoulis, and A. M. Friedlander. Coupling remote sensing with in situ surveys to determine reef fish habitat associations for the design of marine protected areas. Marine Ecology Progress Series, 588:121–134, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/588/
m588p121.pdf.

Gillies:2013:EFW

[GSJS13] C. L. Gillies, J. S. Stark, G. J. Johnstone, and S. D. A. Smith. Establishing a food web model for coastal Antarctic benthic communities: a case study from the vestfold hills. Marine Ecology Progress Series, 478:27–41, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m478p027.pdf.

Guazzo:2019:GWM

[GSJS^{+}19] R. A. Guazzo, A. Schulman-Janiger, M. H. Smith, J. Barlow, G. L. D’Spain, D. B. Rimington, and J. A. Hildebrand. Gray whale migration patterns through the Southern California Bight from multi-year visual and acoustic monitoring. Marine Ecology Progress Series, 625:181–203, 2019. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m625p181.pdf.

Gosnell:2017:COS

[GSL17] J. S. Gosnell, K. Spurgin, and E. A. Levine. Caged oysters still get scared: Predator presence and density influence growth in oysters, but only at very close ranges. Marine Ecology Progress Series, 568:111–122, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/568/m568p111.pdf.

Glover:2011:PGS

[GSM^{+}11a] K. A. Glover, Å. B. Stølen, A. Messmer, B. F. Koop, O. Torrissen, and F. Nilsen. Population genetic structure of the parasitic copepod Lepeophtheirus salmonis throughout the Atlantic. Marine Ecology Progress Series, 427:161–172, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m427p161.pdf.

Goodkin:2011:CCH

[GSM^{+}11b] N. F. Goodkin, A. D. Switzer, D. McCorry, L. DeVantier, J. D. True, K. A. Hughen, N. Angeline, and T. T. Yang. Coral communities of Hong Kong: long-lived corals in a marginal reef environment. Marine Ecology Progress Series, 426:185–196, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/426/m426p185.pdf.

Grattepanche:2015:DAP

[GSMK15] J. D. Grattepanche, L. F. Santoferrara, G. B. McManus, and L. A. Katz. Distinct assemblage of planktonic ciliates dominates both photic and deep waters on the New England shelf. Marine Ecology Progress Series, 526:1–9, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Glaspie:2018:IHP

[GSO^{+}18] C. N. Glaspie, R. D. Seitz, M. B. Ogburn, C. F. Dungan, and A. H. Hines. Impacts of habitat, predators, recruitment, and disease on soft-shell clams Mya arenaria and stout razor clams Tagelus plebeius in Chesapeake Bay. Marine Ecology Progress Series, 603:117–133, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/603/m603p117.pdf.

Garrido:2015:TEP

[GSP^{+}15] S. Garrido, A. Silva, J. Pastor, R. Dominguez, A. V. Silva, and A. M. Santos. Trophic ecology of pelagic fish species off the Iberian coast: diet overlap, cannibalism and intraguild predation. Marine Ecology Progress Series, 539:271– 285, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/539/m539p271.pdf.

Galvan:2010:PSI

[GSR10] D. E. Galván, C. J. Sweeting, and W. D. K. Reid. Power of stable isotope techniques to detect size-based feeding in marine fishes. Marine Ecology Progress Series, 407:271– 278, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/407/m407p271.pdf.

Geange:2013:CHA

[GSS13] S. W. Geange, A. C. Stier, and J. S. Shima. Competitive hierarchies among three species of juvenile coral reef fishes. Marine Ecology Progress Series, 472:239–248, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/472/
m472p239.pdf.

Gonzalez-Solis:2011:CSI

[GSSM^{+}11] J. González-Solís, M. Smyrli, T. Militão, D. Gremillet, T. Tveraa, R. A. Phillips, and T. Boulinier. Combining stable isotope analyses and geolocation to reveal kittiwake migration. Marine Ecology Progress Series, 435:251–261, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/435/
m435p251.pdf.

Greene:2011:GII

[GSTK11] V. E. Greene, L. J. Sullivan, J. K. Thompson, and W. J. Kimmerer. Grazing impact of the invasive clam Corbula amurensis on the microplankton assemblage of the northern San Francisco Estuary. Marine Ecology Progress Series, 431:183–193, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m431p183.pdf.

Green:2012:HCS

[GSU12] B. C. Green, D. J. Smith, and G. J. C. Underwood. Habitat connectivity and spatial complexity differentially affect mangrove and salt marsh fish assemblages. Marine Ecology Progress Series, 466:177–192, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/466/m466p177.pdf.

Grigor:2014:SEL

[GSV14] J. J. Grigor, J. E. Søreide, and Ø. Varpe. Seasonal ecology and life-history strategy of the high-latitude predatory zooplankter Parasagitta elegans. Marine Ecology Progress Series, 499:77–88, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m499p077.pdf.

Guidone:2010:SGF

[GTF10] M. Guidone, C. S. Thornber, and E. Field. Snail grazing facilitates growth of a bloom-forming alga. Marine Ecology Progress Series, 420:83–89, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/420/m420p083.pdf.

Grecian:2016:CMR

[GTL^{+}16] W. J. Grecian, G. A. Taylor, G. Loh, R. A. R. McGill, C. M. Miskelly, R. A. Phillips, D. R. Thompson, and R. W. Furness. Contrasting migratory responses of two closely related seabirds to long-term climate change. Marine Ecology Progress Series, 559:231–242, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m559p231.pdf.

Grabowski:2014:SBA

[GTM14] T. B. Grabowski, V. Thorsteinsson, and G. Marteinsdóttir. Spawning behavior in Atlantic cod: analysis by use of data storage tags. Marine Ecology Progress Series, 506:279–290, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2014/506/m506p279.pdf.

Gimenez:2017:SDN

[GTP^{+}17] L. Giménez, G. Torres, A. Pettersen, M. T. Burrows, A. Estevez, and S. R. Jenkins. Scale-dependent natural variation in larval nutritional reserves in a marine invertebrate: implications for recruitment and cross-ecosystem coupling. Marine Ecology Progress Series, 570:141–155, 2017. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m570p141.pdf.

Gazeau:2015:CAC

[GUC^{+}15] F. Gazeau, L. Urbini, T. E. Cox, S. Alliouane, and J. P. Gattuso. Comparison of the alkalinity and calcium anomaly techniques to estimate rates of net calcification. Marine Ecology Progress Series, 527:1–12, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Gauffier:2018:CMF

[GVG^{+}18] P. Gauffier, P. Verborgh, J. Giménez, R. Esteban, J. M. Salazar Sierra, and R. de Stephanis. Contemporary migration of fin whales through the Strait of Gibraltar. Marine Ecology Progress Series, 588:215–228, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/588/m588p215.pdf.

Griffen:2015:EED

[GVGH15] B. D. Griffen, M. Vogel, L. Goulding, and R. Hartman. Energetic effects of diet choice by invasive Asian shore crabs: implications for persistence when prey are scarce. Marine Ecology Progress Series, 522:181–192, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/522/m522p181.pdf.

Guinet:2014:SES

[GVGP^{+}14] C. Guinet, J. Vacquié-Garcia, B. Picard, G. Bessigneul, Y. Lebras, A. C. Dragon, M. Viviant, J. P. Y. Arnould, and F. Bailleul. Southern elephant seal foraging success in relation to temperature and light conditions: insight into prey distribution. Marine Ecology Progress Series, 499:285–301, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m499p285.pdf.

Gatune:2014:GSP

[GVR^{+}14] C. Gatune, A. Vanreusel, R. Ruwa, P. Bossier, and M. De Troch. Growth and survival of post-larval giant tiger shrimp Penaeus monodon feeding on mangrove leaf litter biofilms. Marine Ecology Progress Series, 511:117–128, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/511/
m511p117.pdf.

Goncalves:2014:MPS

[GvSGCK14] R. J. Gonçalves, H. van Someren Gréve, D. Couespel, and T. Kiørboe. Mechanisms of prey size selection in a suspension-feeding copepod, Temora longicornis. Marine Ecology Progress Series, 517:61–74, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/517/
m517p061.pdf.

Garcia:2012:FSS

[GVW^{+}12] A. M. Garcia, J. P. Vieira, K. O. Winemiller, L. E. Moraes, and E. T. Paes. Factoring scales of spatial and temporal variation in fish abundance in a subtropical estuary. Marine Ecology Progress Series, 461:121–135, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/461/m461p121.pdf.

Goldstein:2015:SMA

[GW15] J. S. Goldstein and W. H. Watson III. Seasonal movements of American lobsters in southern Gulf of Maine coastal waters: patterns, environmental triggers, and implications for larval release. Marine Ecology Progress Series, 524:197– 211, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/524/m524p197.pdf.

Greer:2016:LFU

[GWGC16] A. T. Greer, C. B. Woodson, C. M. Guigand, and R. K. Cowen. Larval fishes utilize Batesian mimicry as a survival strategy in the plankton. Marine Ecology Progress Series, 551:1–12, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Glew:2018:MLD

[GWH^{+}18] K. St John Glew, S. Wanless, M. P. Harris, F. Daunt, K. E. Erikstad, H. Strøm, and C. N. Trueman. Moult location and diet of auks in the North Sea inferred from coupled light-based and isotope-based geolocation. Marine Ecology Progress Series, 599:239–251, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m599p239.pdf.

Gremillet:2012:LAB

[GWK^{+}12] D. Grémillet, J. Welcker, N. J. Karnovsky, W. Walkusz, M. E. Hall, J. Fort, Z. W. Brown, J. R. Speakman, and A. M. A. Harding. Little auks buffer the impact of current Arctic climate change. Marine Ecology Progress Series, 454:197–206, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m454p197.pdf.

Gryba:2019:IFL

[GWK^{+}19] R. D. Gryba, F. K. Wiese, B. P. Kelly, A. L. Von Duyke, R. S. Pickart, and D. A. Stockwell. Inferring foraging locations and water masses preferred by spotted seals Phoca largha and bearded seals Erignathus barbatus. Marine Ecology Progress Series, 631:209–224, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2019/631/m631p209.pdf.

Gove:2015:CRB

[GWM^{+}15] J. M. Gove, G. J. Williams, M. A. McManus, S. J. Clark, J. S. Ehses, and L. M. Wedding. Coral reef benthic regimes exhibit non-linear threshold responses to natural physical drivers. Marine Ecology Progress Series, 522:33–48, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/522/
m522p033.pdf.

Graff:2016:PNP

[GWM^{+}16] J. R. Graff, T. K. Westberry, A. J. Milligan, M. B. Brown, G. Dall’Olmo, K. M. Reifel, and M. J. Behrenfeld. Photoacclimation of natural phytoplankton communities. Marine Ecology Progress Series, 542:51–62, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m542p051.pdf.

Greiner:2016:SSC

[GWME16] J. T. Greiner, G. M. Wilkinson, K. J. McGlathery, and K. A. Emery. Sources of sediment carbon sequestered in restored seagrass meadows. Marine Ecology Progress Series, 551:95–105, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m551p095.pdf.

Gasbarro:2018:CFD

[GWT18] R. Gasbarro, D. Wan, and V. Tunnicliffe. Composition and functional diversity of macrofaunal assemblages on vertical walls of a deep northeast Pacific fjord. Marine Ecology Progress Series, 597:47–64, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/597/m597p047.pdf.

Gagnon:2015:CIS

[GYRJ15] K. Gagnon, J. Yli-Rosti, and V. Jormalainen. Cormorantinduced shifts in littoral communities. Marine Ecology Progress Series, 541:15–30, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/541/m541p015.pdf.

Gillis:2016:MLT

[GZB16] L. G. Gillis, M. Zimmer, and T. J. Bouma. Mangrove leaf transportation: Do mimic Avicennia and Rhizophora roots retain or donate leaves? Marine Ecology Progress Series, 551:107–115, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/551/m551p107.pdf.

Gonzalez-Zevallos:2011:SAI

[GZYS11] D. González-Zevallos, P. Yorio, and W. S. Svagelj. Seabird attendance and incidental mortality at shrimp fisheries in Golfo San Jorge, Argentina. Marine Ecology Progress Series, 432:125–135, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/432/m432p125.pdf.

Hoskins:2014:RBL

[HA14] A. J. Hoskins and J. P. Y. Arnould. Relationship between long-term environmental fluctuations and diving effort of female Australian fur seals. Marine Ecology Progress Series, 511:285–295, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/511/m511p285.pdf.

Holsman:2015:CME

[HA15] K. K. Holsman and K. Aydin. Comparative methods for evaluating climate change impacts on the foraging ecology of Alaskan groundfish. Marine Ecology Progress Series, 521: 217–235, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/521/m521p217.pdf.

Harmon:2018:PIP

[HA18] E. A. Harmon and J. D. Allen. Predator-induced plasticity in egg capsule deposition in the mud snail Tritia obsoleta. Marine Ecology Progress Series, 586:113–125, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/586/
m586p113.pdf.

HortaeCosta:2014:TFA

[HAF^{+}14] B. Horta e Costa, J. Assis, G. Franco, K. Erzini, M. Henriques, E. J. Gonçalves, and J. E. Caselle. Tropicalization of fish assemblages in temperate biogeographic transition zones. Marine Ecology Progress Series, 504:241–252, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/504/
m504p241.pdf.

Hill:2017:ABG

[HAF^{+}17] R. W. Hill, E. J. Armstrong, A. M. Florn, C. Li, R. W. Walquist, and A. Edward. Abundant betaines in giant clams (Tridacnidae) and western Pacific reef corals, including study of coral betaine acclimatization. Marine Ecology Progress Series, 576:27–41, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m576p027.pdf.

Hutchinson:2016:GRT

[HAHM16] E. Hutchinson, S. Atkinson, and A. Hoover-Miller. Growth and reproductive tracts from fetal to adult harbor seals in the Gulf of Alaska. Marine Ecology Progress Series, 557: 277–288, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/557/m557p277.pdf.

Halvorsen:2015:SLS

[Hal15] E. Halvorsen. Significance of lipid storage levels for reproductive output in the Arctic copepod Calanus hyperboreus. Marine Ecology Progress Series, 540:259–265, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/540/
m540p259.pdf.

Hall:2015:ILT

[HAM^{+}15] A. Goodman Hall, L. Avens, J. Braun McNeill, B. Wallace, and L. R. Goshe. Inferring long-term foraging trends of individual juvenile loggerhead sea turtles using stable isotopes. Marine Ecology Progress Series, 537:265–276, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/537/
m537p265.pdf.

Harasti:2016:DSP

[Har16] D. Harasti. Declining seahorse populations linked to loss of essential marine habitats. Marine Ecology Progress Series, 546:173–181, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/546/m546p173.pdf.

Hatch:2013:KDC

[Hat13] S. A. Hatch. Kittiwake diets and chick production signal a 2008 regime shift in the Northeast Pacific. Marine Ecology Progress Series, 477:271–284, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m477p271.pdf.

Hobson:2012:EIY

[HB12] K. A. Hobson and A. L. Bond. Extending an indicator: yearround information on seabird trophic ecology from multipletissue stable-isotope analyses. Marine Ecology Progress Series, 461:233–243, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/461/m461p233.pdf.

Hansen:2013:POD

[HB13] J. L. S. Hansen and J. Bendtsen. Parameterisation of oxygen dynamics in the bottom water of the Baltic Sea-North Sea transition zone. Marine Ecology Progress Series, 481: 25–39, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/481/m481p025.pdf.

Huenerlage:2015:TLK

[HB15a] K. Huenerlage and F. Buchholz. Thermal limits of krill species from the high-Arctic Kongsfjord (Spitsbergen). Marine Ecology Progress Series, 535:89–98, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/535/
m535p089.pdf.

Humanes:2015:LSS

[HB15b] A. Humanes and C. Bastidas. In situ settlement rates and early survivorship of hard corals: a good year for a Caribbean reef. Marine Ecology Progress Series, 539:139– 151, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/539/m539p139.pdf.

Han:2012:RZN

[HBB^{+}12] Q. Han, T. J. Bouma, F. G. Brun, W. Suykerbuyk, and M. M. van Katwijk. Resilience of Zostera noltii to burial or erosion disturbances. Marine Ecology Progress Series, 449:133–143, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m449p133.pdf.

Haulsee:2018:SEC

[HBB^{+}18] D. E. Haulsee, M. W. Breece, L. M. Brown, B. M. Wetherbee, D. A. Fox, and M. J. Oliver. Spatial ecology of Carcharias taurus in the northwestern Mid-Atlantic coastal ocean. Marine Ecology Progress Series, 597:191–206, 2018. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m597p191.pdf.

Hoffmann:2012:IOW

[HBBH12] L. J. Hoffmann, E. Breitbarth, P. W. Boyd, and K. A. Hunter. Influence of ocean warming and acidification on trace metal biogeochemistry. Marine Ecology Progress Series, 470:191–205, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m470p191.pdf.

Heldt:2015:BRH

[HBC15] K. A. Heldt, W. C. Bridges, Jr., and M. J. Childress. Behavioral response to habitat loss in juvenile spiny lobsters. Marine Ecology Progress Series, 521:117–128, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/521/
m521p117.pdf.

Hirst:2013:FBS

[HBCK13] A. G. Hirst, D. Bonnet, D. V. P. Conway, and T. Kiørboe. Female-biased sex ratios in marine pelagic copepods: Comment on Gusmão et al. (2013). Marine Ecology Progress Series, 489:297–298, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m489p297.pdf. See [GMR13b].

Henschke:2013:SFT

[HBE^{+}13] N. Henschke, D. A. Bowden, J. D. Everett, S. P. Holmes, R. J. Kloser, R. W. Lee, and I. M. Suthers. Salp-falls in the Tasman Sea: a major food input to deep-sea benthos. Marine Ecology Progress Series, 491:165–175, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/491/m491p165.pdf.

Halsey:2010:CFD

[HBF^{+}10] L. G. Halsey, P. J. Butler, A. Fahlman, C. A. Bost, and Y. Handrich. Changes in the foraging dive behaviour and energetics of king penguins through summer and autumn: a month by month analysis. Marine Ecology Progress Series, 401:279–289, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/401/m401p279.pdf.

Howells:2017:DDS

[HBG^{+}17] R. J. Howells, S. J. Burthe, J. A. Green, M. P. Harris, M. A. Newell, A. Butler, D. G. Johns, E. J. Carnell, S. Wanless, and F. Daunt. From days to decades: short- and long-term variation in environmental conditions affect offspring diet composition of a marine top predator. Marine Ecology Progress Series, 583:227–242, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m583p227.pdf.

Hardege:2011:IFS

[HBHF^{+}11] J. D. Hardege, H. D. Bartels-Hardege, N. Fletcher, J. A. Terschak, M. Harley, M. A. Smith, L. Davidson, D. Hayden, C. T. Müller, M. Lorch, K. Welham, T. Walther, and R. Bublitz. Identification of a female sex pheromone in Carcinus maenas. Marine Ecology Progress Series, 436:177–189, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/436/
m436p177.pdf.

Holdredge:2010:FCC

[HBHG10] C. Holdredge, M. D. Bertness, N. C. Herrmann, and K. B. Gedan. Fiddler crab control of cordgrass primary production in sandy sediments. Marine Ecology Progress Series, 399:253–259, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/399/m399p253.pdf.

Hammer:2018:HTC

[HBHS^{+}18] K. J. Hammer, J. Borum, H. Hasler-Sheetal, E. C. Shields, K. Sand-Jensen, and K. A. Moore. High temperatures cause reduced growth, plant death and metabolic changes in eelgrass Zostera marina. Marine Ecology Progress Series, 604: 121–132, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/604/m604p121.pdf.

Hammerschlag:2018:PDM

[HBI^{+}18] N. Hammerschlag, S. C. Barley, D. J. Irschick, J. J. Meeuwig, E. R. Nelson, and M. G. Meekan. Predator declines and morphological changes in prey: evidence from coral reefs depleted of sharks. Marine Ecology Progress Series, 586:127– 139, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/586/m586p127.pdf.

Huckstadt:2012:DSC

[HBK^{+}12] L. A. Hückstädt, J. M. Burns, P. L. Koch, B. I. McDonald, D. E. Crocker, and D. P. Costa. Diet of a specialist in a changing environment: the crabeater seal along the western Antarctic Peninsula. Marine Ecology Progress Series, 455:287–301, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/455/m455p287.pdf.

Hardy:2017:ATE

[HBK^{+}17] N. Hardy, T. Berry, B. P. Kelaher, S. D. Goldsworthy, M. Bunce, M. A. Coleman, B. M. Gillanders, S. D. Connell, M. Blewitt, and W. Figueira. Assessing the trophic ecology of top predators across a recolonisation frontier using DNA metabarcoding of diets. Marine Ecology Progress Series, 573:237–254, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/573/m573p237.pdf.

Holliday:2012:LFA

[HBM^{+}12] D. Holliday, L. E. Beckley, N. Millar, M. P. Olivar, D. Slawinski, M. Feng, and P. A. Thompson. Larval fish assemblages and particle back-tracking define latitudinal and cross-shelf variability in an eastern Indian Ocean boundary current. Marine Ecology Progress Series, 460:127–144, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/460/
m460p127.pdf.

Haulsee:2015:HSC

[HBM^{+}15] D. E. Haulsee, M. W. Breece, D. C. Miller, B. M. Wetherbee, D. A. Fox, and M. J. Oliver. Habitat selection of a coastal shark species estimated from an autonomous underwater vehicle. Marine Ecology Progress Series, 528:277–288, 2015. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m528p277.pdf.

Horne:2016:CPL

[HBP^{+}16] J. B. Horne, I. R. Bradbury, I. G. Paterson, D. Hardie, J. Hutchings, B. J. Laurel, P. V. R. Snelgrove, C. J. Morris, R. S. Gregory, and P. Bentzen. Complex post-larval dispersal processes in Atlantic cod revealed by age-based genetics and relatedness analysis. Marine Ecology Progress Series, 556:237–250, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m556p237.pdf.

Hughes:2010:GDG

[HBS10] A. R. Hughes, R. J. Best, and J. J. Stachowicz. Genotypic diversity and grazer identity interactively influence seagrass and grazer biomass. Marine Ecology Progress Series, 403: 43–51, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/403/m403p043.pdf.

Hazen:2013:TSB

[HBS^{+}13] E. L. Hazen, S. J. Bograd, R. M. Suryan, Y. Watanuki, and R. P. Wilson. Theme section: Biophysical coupling of marine hotspots. Marine Ecology Progress Series, 487:176–304, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m487_ThemeSection.pdf.

Hargreaves:2017:TGC

[HBS17] D. Hargreaves, A. Buckland, and M. Sheaves. Trophic guild concept: factors affecting within-guild consistency for tropical estuarine fish. Marine Ecology Progress Series, 564:175– 186, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/564/m564p175.pdf.

Hewitt:2013:CDS

[HC13] J. E. Hewitt and V. J. Cummings. Context-dependent success of restoration of a key species, biodiversity and community composition. Marine Ecology Progress Series, 479:63–73, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m479p063.pdf.

Hoogenboom:2011:BAC

[HCA11] M. O. Hoogenboom, S. R. Connolly, and K. R. N. Anthony. Biotic and abiotic correlates of tissue quality for common scleractinian corals. Marine Ecology Progress Series, 438:119–128, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/438/m438p119.pdf.

Hardison:2010:FMB

[HCAV10] A. K. Hardison, E. A. Canuel, I. C. Anderson, and B. Veuger. Fate of macroalgae in benthic systems: carbon and nitrogen cycling within the microbial community. Marine Ecology Progress Series, 414:41–55, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/414/m414p041.pdf.

Hernaez:2011:PDS

[HCBV^{+}11] P. Hernáez, T. McLellan Clarke, C. Benavides-Varela, F. Villalobos-Rojas, J. Nívia-Ruiz, and I. S. Wehrtmann. Population demography and spatial distribution of the mantis shrimp Squilla biformis (Stomatopoda, Squillidae) from Pacific Costa Rica. Marine Ecology Progress Series, 424:157– 168, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/424/m424p157.pdf.

Heggoy:2015:GLI

[HCDR^{+}15] O. Heggøy, S. Christensen-Dalsgaard, P. S. Ranke, O. Chastel, and C. Bech. GPS-loggers influence behaviour and physiology in the black-legged kittiwake Rissa tridactyla. Marine Ecology Progress Series, 521:237–248, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/521/m521p237.pdf.

Hoang:2016:SCW

[HCF^{+}16] T. C. Hoang, A. J. Cole, R. K. Fotedar, M. J. O’Leary, M. W. Lomas, and S. Roy. Seasonal changes in water quality and Sargassum biomass in southwest Australia. Marine Ecology Progress Series, 551:63–79, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/551/m551p063.pdf.

Hernandez:2010:ETS

[HCG^{+}10] J. C. Hernández, S. Clemente, D. Girard, Á. Pérez-Ruzafa, and A. Brito. Effect of temperature on settlement and postsettlement survival in a barrens-forming sea urchin. Marine Ecology Progress Series, 413:69–80, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/413/
m413p069.pdf.

Hendrickx:2015:INN

[HCG15] J. P. Hendrickx, R. G. Creese, and P. E. Gribben. Impacts of a non-native gastropod with a limited distribution; less conspicuous invaders matter too. Marine Ecology Progress Series, 537:151–162, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m537p151.pdf.

Hernandez:2013:ETS

[HCGP13] F. J. Hernandez, Jr., L. Carassou, W. M. Graham, and S. P. Powers. Evaluation of the taxonomic sufficiency approach for ichthyoplankton community analysis. Marine Ecology Progress Series, 491:77–90, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/491/m491p077.pdf.

Hu:2011:ESF

[HCJ^{+}11] S. Hu, C. Chen, R. Ji, D. W. Townsend, R. Tian, R. C. Beardsley, and C. S. Davis. Effects of surface forcing on interannual variability of the fall phytoplankton bloom in the Gulf of Maine revealed using a process-oriented model. Marine Ecology Progress Series, 427:29–49, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/427/
m427p029.pdf.

Henriksen:2018:OFR

[HCJ^{+}18] O. Henriksen, A. Christensen, S. Jonasdottir, B. R. MacKenzie, K. E. Nielsen, H. Mosegård, and M. van Deurs. Oceanographic flow regime and fish recruitment: reversed circulation in the North Sea coincides with unusually strong sandeel recruitment. Marine Ecology Progress Series, 607:187–205, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m607p187.pdf.

Hovey:2012:SSN

[HCK12] R. K. Hovey, M. L. Cambridge, and G. A. Kendrick. Season and sediment nutrient additions affect root architecture in the temperate seagrasses Posidonia australis and P. sinuosa. Marine Ecology Progress Series, 446:23–30, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/446/
m446p023.pdf.

Heyman:2010:DED

[HCL10] W. D. Heyman, L. M. Carr, and P. S. Lobel. Diver ecotourism and disturbance to reef fish spawning aggregations: It is better to be disturbed than to be dead. Marine Ecology Progress Series, 419:201–210, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/419/m419p201.pdf.

Hamilton:2011:EGO

[HCL^{+}11a] S. L. Hamilton, J. E. Caselle, C. A. Lantz, T. L. Egloff, E. Kondo, S. D. Newsome, K. Loke-Smith, D. J. Pondella II, K. A. Young, and C. G. Lowe. Extensive geographic and ontogenetic variation characterizes the trophic ecology of a temperate reef fish on southern California (USA) rocky reefs. Marine Ecology Progress Series, 429:227–244, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/429/
m429p227.pdf.

Hinz:2011:TDA

[HCL^{+}11b] H. Hinz, E. Capasso, M. Lilley, M. Frost, and S. R. Jenkins. Temporal differences across a bio-geographical boundary reveal slow response of sub-littoral benthos to climate change. Marine Ecology Progress Series, 423:69–82, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/423/
m423p069.pdf.

Hobbs:2018:PAD

[HCLB18] L. Hobbs, F. R. Cottier, K. S. Last, and J. Berge. Pan-Arctic diel vertical migration during the polar night. Marine Ecology Progress Series, 605:61–72, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m605p061.pdf.

Haber:2011:CDA

[HCM^{+}11] M. Haber, M. Carbone, E. Mollo, M. Gavagnin, and M. Ilan. Chemical defense against predators and bacterial fouling in the Mediterranean sponges Axinella polypoides and A. verrucosa. Marine Ecology Progress Series, 422:113–122, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/422/
m422p113.pdf.

Hoskins:2015:IIV

[HCW^{+}15] A. J. Hoskins, D. P. Costa, K. E. Wheatley, J. R. Gibbens, and J. P. Y. Arnould. Influence of intrinsic variation on foraging behaviour of adult female Australian fur seals. Marine Ecology Progress Series, 526:227–239, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/526/m526p227.pdf.

Hultgren:2010:SHC

[HD10] K. M. Hultgren and J. E. Duffy. Sponge host characteristics shape the community structure of their shrimp associates. Marine Ecology Progress Series, 407:1–12, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Higgins:2010:MDF

[HDB^{+}10] R. M. Higgins, B. S. Danilowicz, J. A. Balbuena, A. K. Daníelsdóttir, A. J. Geffen, W. G. Meijer, J. Modin, F. E. Montero, C. Pampoulie, D. Perdiguero-Alonso, A. Schreiber, M. Ö. Stefánsson, and B. Wilson. Multi-disciplinary fingerprints reveal the harvest location of cod Gadus morhua in the northeast Atlantic. Marine Ecology Progress Series, 404:197–206, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/404/m404p197.pdf.

Hubert:2017:GLH

[HDP^{+}17] N. Hubert, A. Dettai, P. Pruvost, C. Cruaud, M. Kulbicki, R. F. Myers, and P. Borsa. Geography and life history traits account for the accumulation of cryptic diversity among Indo– West Pacific coral reef fishes. Marine Ecology Progress Series, 583:179–193, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m583p179.pdf.

Huebner:2012:HPH

[HDT^{+}12] L. K. Huebner, B. Dailey, B. M. Titus, M. Khalaf, and N. E. Chadwick. Host preference and habitat segregation among Red Sea anemonefish: effects of sea anemone traits and fish life stages. Marine Ecology Progress Series, 464:1–15, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Hickman:2010:MEC

[HDWF10] A. E. Hickman, S. Dutkiewicz, R. G. Williams, and M. J. Follows. Modelling the effects of chromatic adaptation on phytoplankton community structure in the oligotrophic ocean. Marine Ecology Progress Series, 406:1–17, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Hunsicker:2010:PRC

[HEAI10] M. E. Hunsicker, T. E. Essington, K. Y. Aydin, and B. Ishida. Predatory role of the commander squid Berryteuthis magister in the eastern Bering Sea: insights from stable isotopes and food habits. Marine Ecology Progress Series, 415:91– 108, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/415/m415p091.pdf.

Henschke:2011:DLH

[HEB^{+}11] N. Henschke, J. D. Everett, M. E. Baird, M. D. Taylor, and I. M. Suthers. Distribution of life-history stages of the salp Thalia democratica in shelf waters during a spring bloom. Marine Ecology Progress Series, 430:49–62, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/430/
m430p049.pdf.

Hendy:2013:DUT

[HED^{+}13] I. W. Hendy, J. Eme, T. F. Dabruzzi, R. V. Nembhard, S. M. Cragg, and W. A. Bennett. Dartfish use teredinid tunnels in fallen mangrove wood as a low-tide refuge. Marine Ecology Progress Series, 486:237–245, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/486/m486p237.pdf.

Holsman:2012:CAC

[HEM^{+}12] K. K. Holsman, T. Essington, T. J. Miller, M. Koen-Alonso, and W. J. Stockhausen. Comparative analysis of cod and herring production dynamics across 13 northern hemisphere marine ecosystems. Marine Ecology Progress Series, 459: 231–246, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m459p231.pdf.

Hameed:2018:IVS

[HEMJ18] S. O. Hameed, M. L. Elliott, S. G. Morgan, and J. Jahncke. Interannual variation and spatial distribution of decapod larvae in a region of persistent coastal upwelling. Marine Ecology Progress Series, 587:55–71, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m587p055.pdf.

Hamilton:2019:SVR

[HER^{+}19] V. Hamilton, K. Evans, B. Raymond, E. Betty, and M. A. Hindell. Spatial variability in responses to environmental conditions in southern hemisphere long-finned pilot whales. Marine Ecology Progress Series, 629:207–218, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/629/
m629p207.pdf.

Haywood:2019:GRI

[HFG^{+}19] J. C. Haywood, W. J. Fuller, B. J. Godley, J. D. Shutler, S. Widdicombe, and A. C. Broderick. Global review and inventory: how stable isotopes are helping us understand ecology and inform conservation of marine turtles. Marine Ecology Progress Series, 613:217–245, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m613p217.pdf.

Huse:2010:ENW

[HFH10] G. Huse, A. Fernö, and J. C. Holst. Establishment of new wintering areas in herring co-occurs with peaks in the ‘first time/repeat spawner’ ratio. Marine Ecology Progress Series, 409:189–198, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/409/m409p189.pdf.

Hunting:2013:SDS

[HFK^{+}13] E. R. Hunting, O. Franken, F. Knopperts, M. H. S. Kraak, R. Vargas, W. F. M. Röling, and H. G. van der Geest. Substrate as a driver of sponge distributions in mangrove ecosystems. Marine Ecology Progress Series, 486:133–141, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m486p133.pdf.

Hidalgo:2011:EPS

[HFLI11] F. J. Hidalgo, F. N. Firstater, B. J. Lomovasky, and O. O. Iribarne. Effects of a predatory starfish on substrate colonization by a dominant mussel. Marine Ecology Progress Series, 432:103–114, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/432/m432p103.pdf.

Hamilton:2012:OSS

[HFM12] B. M. Hamilton, P. G. Fairweather, and B. McDonald. One species of seagrass cannot act as a surrogate for others in relation to providing habitat for other taxa. Marine Ecology Progress Series, 456:43–51, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/456/m456p043.pdf.

Heck:2015:SCN

[HFM^{+}15] K. L. Heck, Jr., F. J. Fodrie, S. Madsen, C. J. Baillie, and D. A. Byron. Seagrass consumption by native and a tropically associated fish species: potential impacts of the tropicalization of the northern Gulf of Mexico. Marine Ecology Progress Series, 520:165–173, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m520p165.pdf.

Hall:2016:RTT

[HFMD16] M. O. Hall, B. T. Furman, M. Merello, and M. J. Durako. Recurrence of Thalassia testudinum seagrass die-off in Florida Bay, USA: initial observations. Marine Ecology Progress Series, 560:243–249, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m560p243.pdf.

Hosia:2014:TAP

[HFN14] A. Hosia, T. Falkenhaug, and L. J. Naustvoll. Trends in abundance and phenology of Aurelia aurita and Cyanea spp. at a Skagerrak location, 1992–2011. Marine Ecology Progress Series, 498:103–115, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/498/m498p103.pdf.

Hjermann:2013:SAN

[HFR^{+}13] D. Ø. Hjermann, J. A. D. Fisher, T. Rouyer, K. T. Frank, and N. C. Stenseth. Spatial analysis of North Sea cod recruitment: concurrent effects of changes in spawning stock biomass, temperature and herring abundance. Marine Ecology Progress Series, 480:263–275, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m480p263.pdf.

Harvey:2018:CZA

[HFR^{+}18] J. B. J. Harvey, J. L. Fisher, J. P. Ryan, S. B. Johnson, W. T. Peterson, and R. C. Vrijenhoek. Changes in zooplankton assemblages in northern Monterey Bay, California, during a fall transition. Marine Ecology Progress Series, 604: 99–120, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/604/m604p099.pdf.

Heath:2010:WFB

[HG10] J. P. Heath and H. G. Gilchrist. When foraging becomes unprofitable: energetics of diving in tidal currents by common eiders wintering in the Arctic. Marine Ecology Progress Series, 403:279–290, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/403/m403p279.pdf.

Harwell:2013:QLT

[HG13] Mark A. Harwell and John H. Gentile. Quantifying long-term risks to sea otters from the 1989 ‘Exxon Valdez’ oil spill: Comment on Bodkin et al. (2012). Marine Ecology Progress Series, 488:291–296, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m488p291.pdf. See [BBC^{+}12a].

Hancock:2017:ECT

[HG17] E. R. Hancock and B. D. Griffen. Energetic consequences of temperature and sequential autotomization for the stone crab, Menippe spp. Marine Ecology Progress Series, 582: 133–146, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/582/m582p133.pdf.

Hixon:2016:LMM

[HGA^{+}16a] M. A. Hixon, S. J. Green, M. A. Albins, J. L. Akins, and J. A. Morris, Jr. Lionfish: a major marine invasion. Marine Ecology Progress Series, 558:161–165, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m558p161.pdf.

Hixon:2016:IAC

[HGA^{+}16b] Mark Hixon, Stephanie Green, Mark Albins, Ivan Nagelkerken, Tim McClanahan, and Charles Birkeland. Invasion of Atlantic coastal ecosystems by Pacific lionfish. Marine Ecology Progress Series, 558:159–160, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m558p159.pdf.

Hixon:2016:TSI

[HGA^{+}16c] Mark Hixon, Stephanie Green, Mark Albins, Ivan Nagelkerken, Tim McClanahan, and Charles Birkeland. Theme section: Invasion of Atlantic coastal ecosystems by Pacific lionfish. Marine Ecology Progress Series, 558:159–279, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m558_ThemeSection.pdf.

Honey:2013:HBM

[HGB^{+}13] D. J. Honey, M. Gledhill, T. S. Bibby, F. E. Legiret, N. J. Pratt, A. E. Hickman, T. Lawson, and E. P. Achterberg. Heme-b in marine phytoplankton and particulate material from the North Atlantic Ocean. Marine Ecology Progress Series, 483:1–17, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Hayford:2015:FBM

[HGC15] H. A. Hayford, S. E. Gilman, and E. Carrington. Foraging behavior minimizes heat exposure in a complex thermal landscape. Marine Ecology Progress Series, 518:165– 175, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/518/m518p165.pdf.

Hemery:2011:PHP

[HGFA^{+}11] L. G. Hemery, B. Galton-Fenzi, N. Améziane, M. J. Riddle, S. R. Rintoul, R. J. Beaman, A. L. Post, and M. Eléaume. Predicting habitat preferences for Anthometrina adriani (Echinodermata) on the East Antarctic continental shelf. Marine Ecology Progress Series, 441:105–116, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/441/
m441p105.pdf.

Hershberger:2010:KVS

[HGG^{+}10] P. Hershberger, J. Gregg, C. Grady, R. Collins, and J. Winton. Kinetics of viral shedding provide insights into the epidemiology of viral hemorrhagic septicemia in Pacific herring. Marine Ecology Progress Series, 400:187–193, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/400/
m400p187.pdf.

Halttunen:2018:MAS

[HGG^{+}18] E. Halttunen, K. Ø. Gjelland, K. A. Glover, I. Askeland Johnsen, R. M. Serra-Llinares, Ø. Skaala, R. Nilsen, P. A. Bjørn, Ø. Karlsen, B. Finstad, and O. T. Skilbrei. Migration of Atlantic salmon post-smolts in a fjord with high infestation pressure of salmon lice. Marine Ecology Progress Series, 592:243–256, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m592p243.pdf.

Hughes:2014:ASS

[HGKB14] A. R. Hughes, P. E. Gribben, D. L. Kimbro, and M. J. Bishop. Additive and site-specific effects of two foundation species on invertebrate community structure. Marine Ecology Progress Series, 508:129–138, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2014/508/m508p129.pdf.

Hajdu:2015:DAA

[HGL15] S. Hajdu, E. Gorokhova, and U. Larsson. In-depth analysis of an alternate-stage Prymnesium polylepis (Haptophyta) bloom and long-term trends in abundance of prymnesiales species in the Baltic Sea. Marine Ecology Progress Series, 526:55–66, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/526/m526p055.pdf.

Hughes:2010:AAC

[HGPM10] A. D. Hughes, A. G. Grottoli, T. K. Pease, and Y. Matsui. Acquisition and assimilation of carbon in non-bleached and bleached corals. Marine Ecology Progress Series, 420: 91–101, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/420/m420p091.pdf.

Haney:2014:BMVa

[HGS14a] J. C. Haney, H. J. Geiger, and J. W. Short. Bird mortality from the Deepwater Horizon oil spill. I. Exposure probability in the offshore Gulf of Mexico. Marine Ecology Progress Series, 513:225–237, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m513p225.pdf. See comment

[SB15a] and reply [HGS15].

Haney:2014:BMVb

[HGS14b] J. C. Haney, H. J. Geiger, and J. W. Short. Bird mortality from the Deepwater Horizon oil spill. II. Carcass sampling and exposure probability in the coastal Gulf of Mexico. Marine Ecology Progress Series, 513:239–252, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m513p239.pdf. See

comment [SB15a] and reply [HGS15].

Haney:2015:BMD

[HGS15] J. C. Haney, H. J. Geiger, and J. W. Short. Bird mortality due to the Deepwater Horizon oil spill: Reply to Sackmann & Becker (2015). Marine Ecology Progress Series, 534:279–283, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/

meps_oa/m534p279.pdf. See [HGS14a, HGS14b, SB15a].

Holt:2011:MBL

[HGW^{+}11] A. R. Holt, J. A. Godbold, P. C. L. White, A. M. Slater, E. G. Pereira, and M. Solan. Mismatches between legislative frameworks and benefits restrict the implementation of the ecosystem approach in coastal environments. Marine Ecology Progress Series, 434:213–228, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m434p213.pdf.

Henderson:2015:BAI

[HH15] J. Henderson and S. D. Hacker. Buried alive: an invasive seagrass (Zostera japonica) changes its reproductive allocation in response to sediment disturbance. Marine Ecology Progress Series, 532:123–136, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2015/532/m532p123.pdf.

Huwer:2014:CJS

[HHB^{+}14] B. Huwer, H. H. Hinrichsen, U. Böttcher, R. Voss, and F. Köster. Characteristics of juvenile survivors reveal spatiotemporal differences in early life stage survival of Baltic cod. Marine Ecology Progress Series, 511:165–180, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/511/
m511p165.pdf.

Henzler:2010:FCR

[HHG10] C. M. Henzler, E. A. Hoaglund, and S. D. Gaines. FISH-CS — a rapid method for counting and sorting species of marine zooplankton. Marine Ecology Progress Series, 410:1–11, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Hansen:2011:POF

[HHH11a] A. T. Hansen, M. Hondzo, and C. L. Hurd. Photosynthetic oxygen flux by Macrocystis pyrifera: a mass transfer model with experimental validation. Marine Ecology Progress Series, 434:45–55, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m434p045.pdf.

Heithaus:2011:TDR

[HHH^{+}11b] E. R. Heithaus, P. A. Heithaus, M. R. Heithaus, D. Burkholder, and C. A. Layman. Trophic dynamics in a relatively pristine subtropical fringing mangrove community. Marine Ecology Progress Series, 428:49–61, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/428/m428p049.pdf.

Huang:2011:NEC

[HHH^{+}11c] Y. C. A. Huang, H. J. Hsieh, S. C. Huang, P. J. Meng, Y. S. Chen, S. Keshavmurthy, Y. Nozawa, and C. A. Chen. Nutrient enrichment caused by marine cage culture and its influence on subtropical coral communities in turbid waters. Marine Ecology Progress Series, 423:83–93, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m423p083.pdf.

Hensgen:2014:LPI

[HHH^{+}14] G. M. Hensgen, G. J. Holt, S. A. Holt, J. A. Williams, and G. W. Stunz. Landscape pattern influences nekton diversity and abundance in seagrass meadows. Marine Ecology Progress Series, 507:139–152, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2014/507/m507p139.pdf.

Harii:2014:IMD

[HHI^{+}14] S. Harii, C. Hongo, M. Ishihara, Y. Ide, and H. Kayanne. Impacts of multiple disturbances on coral communities at Ishigaki Island, Okinawa, Japan, during a 15 year survey. Marine Ecology Progress Series, 509:171–180, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/509/
m509p171.pdf.

Han:2010:PGS

[HHLT10] Y. S. Han, C. L. Hung, Y. F. Liao, and W. N. Tzeng. Population genetic structure of the Japanese eel Anguilla japonica: panmixia at spatial and temporal scales. Marine Ecology Progress Series, 401:221–232, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/401/m401p221.pdf.

Hirai:2017:MBD

[HHNI17] J. Hirai, K. Hidaka, S. Nagai, and T. Ichikawa. Molecularbased diet analysis of the early post-larvae of Japanese sardine Sardinops melanostictus and Pacific round herring Etrumeus teres. Marine Ecology Progress Series, 564:99– 113, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/564/m564p099.pdf.

Healy:2017:RSM

[HHP^{+}17] S. J. Healy, S. G. Hinch, A. D. Porter, E. L. Rechisky, D. W. Welch, E. J. Eliason, A. G. Lotto, and N. B. Furey. Routespecific movements and survival during early marine migration of hatchery steelhead Oncorhynchus mykiss smolts in coastal British Columbia. Marine Ecology Progress Series, 577:131–147, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/577/m577p131.pdf.

Hammerschlag:2010:IPR

[HHS10] N. Hammerschlag, M. R. Heithaus, and J. E. Serafy. Influence of predation risk and food supply on nocturnal fish foraging distributions along a mangrove-seagrass ecotone. Marine Ecology Progress Series, 414:223–235, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/414/
m414p223.pdf.

Hansen:2012:SVC

[HHS^{+}12a] J. H. Hansen, R. B. Hedeholm, K. Sünksen, J. Tang Christensen, and P. Grønkjær. Spatial variability of carbon (δ 13C) and nitrogen (δ ^{1}5N) stable isotope ratios in an Arctic marine food web. Marine Ecology Progress Series, 467: 47–59, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/467/m467p047.pdf.

Harvey:2012:WFT

[HHS^{+}12b] V. Harvey, M. O. Hammill, D. P. Swain, G. A. Breed, C. Lydersen, and K. M. Kovacs. Winter foraging by a top predator, the grey seal Halichoerus grypus, in relation to the distribution of prey. Marine Ecology Progress Series, 462:273–286, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m462p273.pdf.

Hass:2012:CCH

[HHS12c] T. Hass, J. Hyman, and B. X. Semmens. Climate change, heightened hurricane activity, and extinction risk for an endangered tropical seabird, the black-capped petrel Pterodroma hasitata. Marine Ecology Progress Series, 454:251–261, 2012.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m454p251.pdf.

Hughes:2011:IFI

[HHWW11] B. B. Hughes, J. C. Haskins, K. Wasson, and E. Watson. Identifying factors that influence expression of eutrophication in a central California estuary. Marine Ecology Progress Series, 439:31–43, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/439/m439p031.pdf.

Hirase:2014:LTV

[HI14] S. Hirase and M. Ikeda. Long-term vicariance and post-glacial expansion in the Japanese rocky intertidal goby Chaenogobius annularis. Marine Ecology Progress Series, 499:217– 231, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/499/m499p217.pdf.

Hornby:2017:BWD

[HIH^{+}17] C. A. Hornby, J. Iacozza, C. Hoover, D. G. Barber, and L. L. Loseto. Beluga whale Delphinapterus leucas late summer habitat use and support for foraging areas in the Canadian Beaufort Sea. Marine Ecology Progress Series, 574: 243–257, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/574/m574p243.pdf.

Hirase:2012:PIG

[HIKK12] S. Hirase, M. Ikeda, M. Kanno, and A. Kijima. Phylogeography of the intertidal goby Chaenogobius annularis associated with paleoenvironmental changes around the Japanese Archipelago. Marine Ecology Progress Series, 450:167–179, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2012/450/m450p167.pdf.

Hutchinson:2015:PRS

[HIMH15] M. R. Hutchinson, D. G. Itano, J. A. Muir, and K. N. Holland. Post-release survival of juvenile silky sharks captured in a tropical tuna purse seine fishery. Marine Ecology Progress Series, 521:143–154, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m521p143.pdf.

Hoover:2013:ELA

[HJ13] R. R. Hoover and C. M. Jones. Effect of laser ablation depth in otolith life history scans. Marine Ecology Progress Series, 486:247–256, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/486/m486p247.pdf.

Hein:2016:DBB

[HJ16] S. R. Hein and M. W. Jacobs. Decorating behavior begins immediately after metamorphosis in the decorator crab Oregonia gracilis. Marine Ecology Progress Series, 555:141– 150, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/555/m555p141.pdf.

Henriksen:2012:ETF

[HJMN^{+}12] M. V. Henriksen, S. Jung-Madsen, T. G. Nielsen, E. F. Møller, K. V. Henriksen, S. Markager, and B. W. Hansen. Effects of temperature and food availability on feeding and egg production of Calanus hyperboreus from Disko Bay, western Greenland. Marine Ecology Progress Series, 447:109– 126, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/447/m447p109.pdf.

Haas:2010:OMR

[HJN^{+}10] A. F. Haas, C. Jantzen, M. S. Naumann, R. Iglesias-Prieto, and C. Wild. Organic matter release by the dominant primary producers in a Caribbean reef lagoon: implication for in situ O_{2} availability. Marine Ecology Progress Series, 409: 27–39, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/409/m409p027.pdf.

Hawkes:2019:GSG

[HKB^{+}19a] N. Hawkes, M. Korabik, L. Beazley, H. T. Rapp, J. R. Xavier, and E. Kenchington. Glass sponge grounds on the Scotian Shelf and their associated biodiversity. Marine Ecology Progress Series, 614:91–109, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m614p091.pdf. See corri-

gendum [Ano19b].

Hetherington:2019:RET

[HKB^{+}19b] E. D. Hetherington, C. M. Kurle, S. R. Benson, T. T. Jones, and J. A. Seminoff. Re-examining trophic dead ends: stable isotope values link gelatinous zooplankton to leatherback turtles in the California Current. Marine Ecology Progress Series, 632:205–219, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m632p205.pdf.

Hamilton:2017:SOA

[HKI^{+}17] C. D. Hamilton, K. M. Kovacs, R. A. Ims, J. Aars, H. Strøm, and C. Lydersen. Spatial overlap among an Arctic predator, prey and scavenger in the marginal ice zone. Marine Ecology Progress Series, 573:45–59, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/573/m573p045.pdf.

Haavisto:2017:DRM

[HKJ17] F. Haavisto, R. Koivikko, and V. Jormalainen. Defensive role of macroalgal phlorotannins: benefits and trade-offs under natural herbivory. Marine Ecology Progress Series, 566: 79–90, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/566/m566p079.pdf.

Hallenbeck:2012:RSD

[HKL12] T. R. Hallenbeck, R. G. Kvitek, and J. Lindholm. Rippled scour depressions add ecologically significant heterogeneity to soft-bottom habitats on the continental shelf. Marine Ecology Progress Series, 468:119–133, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/468/m468p119.pdf.

Hamilton:2015:YRH

[HKL15] C. D. Hamilton, K. M. Kovacs, and C. Lydersen. Yearround haul-out behaviour of male walruses Odobenus rosmarus in the Northern Barents Sea. Marine Ecology Progress Series, 519:251–263, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/519/m519p251.pdf.

Hamilton:2019:SSU

[HKL19] C. D. Hamilton, K. M. Kovacs, and C. Lydersen. Sympatric seals use different habitats in an Arctic glacial fjord. Marine Ecology Progress Series, 615:205–220, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/615/
m615p205.pdf.

Heisterkamp:2012:ICI

[HKS^{+}12] I. M. Heisterkamp, A. Kamp, A. T. Schramm, D. de Beer, and P. Stief. Indirect control of the intracellular nitrate pool of intertidal sediment by the polychaete Hediste diversicolor. Marine Ecology Progress Series, 445:181–192, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/445/
m445p181.pdf.

Heupel:2014:SER

[HKSD14] M. R. Heupel, D. M. Knip, C. A. Simpfendorfer, and N. K. Dulvy. Sizing up the ecological role of sharks as predators. Marine Ecology Progress Series, 495:291–298, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m495p291.pdf.

Hazelton:2010:GSP

[HKT10] E. L. G. Hazelton, T. J. Knight, and T. A. Theodose. Glutamine synthetase partitioning in native and introduced salt marsh grasses. Marine Ecology Progress Series, 414: 57–64, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/414/m414p057.pdf.

Hughes:2014:GDM

[HL14] A. R. Hughes and K. E. Lotterhos. Genotypic diversity at multiple spatial scales in the foundation marsh species, Spartina alterniflora. Marine Ecology Progress Series, 497: 105–117, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/497/m497p105.pdf.

Hirst:2016:LNS

[HLB^{+}16] A. J. Hirst, A. R. Longmore, D. Ball, P. L. M. Cook, and G. P. Jenkins. Linking nitrogen sources utilised by seagrass in a temperate marine embayment to patterns of seagrass change during drought. Marine Ecology Progress Series, 549:79–88, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/549/m549p079.pdf.

Hollyman:2018:SWB

[HLC^{+}18] P. R. Hollyman, M. J. Leng, S. R. N. Chenery, V. V. Laptikhovsky, and C. A. Richardson. Statoliths of the whelk Buccinum undatum: a novel age determination tool. Marine Ecology Progress Series, 598:261–272, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m598p261.pdf.

Hyndes:2012:DPC

[HLD12] G. A. Hyndes, P. S. Lavery, and C. Doropoulos. Dual processes for cross-boundary subsidies: incorporation of nutrients from reef-derived kelp into a seagrass ecosystem. Marine Ecology Progress Series, 445:97–107, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/445/m445p097.pdf.

Hamilton:2019:BOB

[HLF^{+}19] C. D. Hamilton, C. Lydersen, M. A. Fedak, C. Freitas, M. A. Hindell, and K. M. Kovacs. Behavioural ontogeny of bearded seals Erignathus barbatus through the first year of life. Marine Ecology Progress Series, 627:179–194, 2019. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m627p179.pdf.

Hunter:2018:IUS

[HLH18] E. Hunter, V. V. Laptikhovsky, and P. R. Hollyman. Innovative use of sclerochronology in marine resource management. Marine Ecology Progress Series, 598:155–158, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m598p155.pdf.

Hamilton:2016:CHU

[HLIK16] C. D. Hamilton, C. Lydersen, R. A. Ims, and K. M. Kovacs. Coastal habitat use by ringed seals Pusa hispida following a regional sea-ice collapse: importance of glacial refugia in a changing Arctic. Marine Ecology Progress Series, 545:261–277, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/545/m545p261.pdf.

Hong:2017:NNC

[HLL^{+}17] H. Hong, D. Li, W. Lin, W. Li, and D. Shi. Nitrogen nutritional condition affects the response of energy metabolism in diatoms to elevated carbon dioxide. Marine Ecology Progress Series, 567:41–56, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/567/m567p041.pdf.

Hundt:2019:SCS

[HLN^{+}19] P. J. Hundt, E. B. Liddle, S. V. Nielsen, B. J. Pinto, and T. Gamble. Sex chromosomes and sex-specific molecular markers in Indo–Pacific combtooth blennies (Blenniidae, Istiblennius). Marine Ecology Progress Series, 627:195– 200, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/627/m627p195.pdf.

Hauser:2015:RDB

[HLPS^{+}15] D. D. W. Hauser, K. L. Laidre, S. L. Parker-Stetter, J. K. Horne, R. S. Suydam, and P. R. Richard. Regional diving behavior of Pacific Arctic beluga whales Delphinapterus leucas and possible associations with prey. Marine Ecology Progress Series, 541:245–264, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m541p245.pdf.

Harris:2015:QSS

[HLS15a] J. L. Harris, L. S. Lewis, and J. E. Smith. Quantifying scales of spatial variability in algal turf assemblages on coral reefs. Marine Ecology Progress Series, 532:41–57, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/532/
m532p041.pdf.

Hines:2015:EES

[HLS^{+}15b] D. E. Hines, J. A. Lisa, B. Song, C. R. Tobias, and S. R. Borrett. Estimating the effects of seawater intrusion on an estuarine nitrogen cycle by comparative network analysis. Marine Ecology Progress Series, 524:137–154, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/524/
m524p137.pdf.

Haggarty:2017:YYR

[HLS17] D. R. Haggarty, K. E. Lotterhos, and J. B. Shurin. Youngof-the-year recruitment does not predict the abundance of older age classes in black rockfish in Barkley Sound, British Columbia, Canada. Marine Ecology Progress Series, 574: 113–126, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/574/m574p113.pdf.

Heupel:2018:TRS

[HLS18] M. R. Heupel, E. J. I. Lédée, and C. A. Simpfendorfer. Telemetry reveals spatial separation of co-occurring reef sharks. Marine Ecology Progress Series, 589:179–192, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/589/
m589p179.pdf.

Han:2016:BDE

[HLW^{+}16] Y. S. Han, Y. F. Lin, C. R. Wu, Y. Iizuka, T. R. Castillo, I. U. Yambot, M. D. Mamalangkap, and A. V. Yambot. Biogeographic distribution of the eel Anguilla luzonensis: dependence upon larval duration and oceanic currents. Marine Ecology Progress Series, 551:227–238, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/551/
m551p227.pdf.

Holt:2009:DST

[HM09] C. A. Holt and N. Mantua. Defining spring transition: regional indices for the California Current System. Marine Ecology Progress Series, 393:285–299, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m393p285.pdf. See erratum

[Ano10d].

Holmes:2010:SSP

[HM10] T. H. Holmes and M. I. McCormick. Size-selectivity of predatory reef fish on juvenile prey. Marine Ecology Progress Series, 399:273–283, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/399/m399p273.pdf.

Houk:2013:TIE

[HM13] P. Houk and C. Musburger. Trophic interactions and ecological stability across coral reefs in the Marshall Islands. Marine Ecology Progress Series, 488:23–34, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m488p023.pdf.

Harvey:2017:OWA

[HM17] B. P. Harvey and P. J. Moore. Ocean warming and acidification prevent compensatory response in a predator to reduced prey quality. Marine Ecology Progress Series, 563: 111–122, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/563/m563p111.pdf.

Hoover-Miller:2011:PDA

[HMAC^{+}11] A. Hoover-Miller, S. Atkinson, S. Conlon, J. Prewitt, and P. Armato. Persistent decline in abundance of harbor seals Phoca vitulina richardsi over three decades in Aialik Bay, an Alaskan tidewater glacial fjord. Marine Ecology Progress Series, 424:259–271, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m424p259.pdf.

Hazen:2012:OMT

[HMB^{+}12] E. L. Hazen, S. M. Maxwell, H. Bailey, S. J. Bograd, M. Hamann, P. Gaspar, B. J. Godley, and G. L. Shillinger. Ontogeny in marine tagging and tracking science: technologies and data gaps. Marine Ecology Progress Series, 457: 221–240, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m457p221.pdf.

Hocevar:2018:SFP

[HMBP18] S. Hočevar, A. Malej, B. Boldin, and J. E. Purcell. Seasonal fluctuations in population dynamics of Aurelia aurita polyps in situ with a modelling perspective. Marine Ecology Progress Series, 591:155–166, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/591/m591p155.pdf.

Hurst:2018:SVP

[HMF^{+}18] T. P. Hurst, J. A. Miller, N. Ferm, R. A. Heintz, and E. V. Farley. Spatial variation in potential and realized growth of juvenile Pacific cod in the southeastern Bering Sea. Marine Ecology Progress Series, 590:171–185, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/590/m590p171.pdf.

Heidemann:2012:SOM

[HMH^{+}12] F. Heidemann, L. Marohn, H. H. Hinrichsen, B. Huwer, K. Hüssy, A. Klügel, U. Böttcher, and R. Hanel. Suitability of otolith microchemistry for stock separation of Baltic cod. Marine Ecology Progress Series, 465:217–226, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/465/
m465p217.pdf.

Halsband:2018:JSD

[HMH^{+}18] C. Halsband, S. Majaneva, A. Hosia, P. A. Emaus, F. Gaardsted, Q. Zhou, O. A. Nøst, and P. E. Renaud. Jellyfish summer distribution, diversity and impact on fish farms in a Nordic fjord. Marine Ecology Progress Series, 591:267–279, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m591p267.pdf.

Horinouchi:2013:HPR

[HMJ^{+}13] M. Horinouchi, N. Mizuno, Y. Jo, M. Fujita, Y. Suzuki, F. Aranishi, and M. Sano. Habitat preference rather than predation risk determines the distribution patterns of filefish Rudarius ercodes in and around seagrass habitats. Marine Ecology Progress Series, 488:255–266, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/488/
m488p255.pdf.

Harborne:2011:BMS

[HMKF11] A. R. Harborne, P. J. Mumby, E. V. Kennedy, and R. Ferrari. Biotic and multi-scale abiotic controls of habitat quality: their effect on coral-reef fishes. Marine Ecology Progress Series, 437:201–214, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/437/m437p201.pdf.

Humphries:2014:DIC

[HMM14] A. T. Humphries, T. R. McClanahan, and C. D. McQuaid. Differential impacts of coral reef herbivores on algal succession in Kenya. Marine Ecology Progress Series, 504:119–132, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m504p119.pdf.

Hodapp:2015:SEM

[HMM^{+}15] D. Hodapp, S. Meier, F. Muijsers, T. H. Badewien, and H. Hillebrand. Structural equation modeling approach to the diversity–productivity relationship of Wadden Sea phytoplankton. Marine Ecology Progress Series, 523:31–40, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m523p031.pdf.

Hedd:2012:TEM

[HMO^{+}12] A. Hedd, W. A. Montevecchi, H. Otley, R. A. Phillips, and D. A. Fifield. Trans-equatorial migration and habitat use by sooty shearwaters Puffinus griseus from the South Atlantic during the nonbreeding season. Marine Ecology Progress Series, 449:277–290, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/449/m449p277.pdf.

Harland:2015:NLE

[HMP15] H. Harland, C. D. MacLeod, and R. Poulin. Non-linear effects of ocean acidification on the transmission of a marine intertidal parasite. Marine Ecology Progress Series, 536: 55–64, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/536/m536p055.pdf.

Horswill:2016:DIV

[HMR^{+}16] C. Horswill, J. Matthiopoulos, N. Ratcliffe, J. A. Green, P. N. Trathan, R. A. R. McGill, R. A. Phillips, and T. C. O’Connell. Drivers of intrapopulation variation in resource use in a generalist predator, the macaroni penguin. Marine Ecology Progress Series, 548:233–247, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m548p233.pdf.

Hammerschlag:2010:RPR

[HMS10] N. Hammerschlag, A. B. Morgan, and J. E. Serafy. Relative predation risk for fishes along a subtropical mangroveseagrass ecotone. Marine Ecology Progress Series, 401:259– 267, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/401/m401p259.pdf.

Hickman:2012:PPN

[HMS^{+}12] A. E. Hickman, C. M. Moore, J. Sharples, M. I. Lucas, G. H. Tilstone, V. Krivtsov, and P. M. Holligan. Primary production and nitrate uptake within the seasonal thermocline of a stratified shelf sea. Marine Ecology Progress Series, 463:39–57, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m463p039.pdf.

Horne:2012:SCT

[HMW^{+}12] J. B. Horne, P. Momigliano, D. J. Welch, S. J. Newman, and L. van Herwerden. Searching for common threads in threadfins: phylogeography of Australian polynemids in space and time. Marine Ecology Progress Series, 449:263– 276, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/449/m449p263.pdf.

Herbon:2013:EDS

[HN13] C. M. Herbon and I. Nordhaus. Experimental determination of stable carbon and nitrogen isotope fractionation between mangrove leaves and crabs. Marine Ecology Progress Series, 490:91–105, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m490p091.pdf.

Huijbers:2011:HTS

[HNG^{+}11] C. M. Huijbers, I. Nagelkerken, L. L. Govers, M. van de Kerk, J. J. Oldenburger, and J. H. F. de Brouwer. Habitat type and schooling interactively determine refuge-seeking behavior in a coral reef fish throughout ontogeny. Marine Ecology Progress Series, 437:241–251, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m437p241.pdf.

Hamabata:2018:SCG

[HNK^{+}18] T. Hamabata, H. Nishizawa, I. Kawazu, K. Kameda, N. Kamezaki, and T. Hikida. Stock composition of green turtles Chelonia mydas foraging in the Ryukyu Archipelago differs with size class. Marine Ecology Progress Series, 600: 151–163, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/600/m600p151.pdf.

Hosokawa:2015:SDS

[HNMK15] S. Hosokawa, M. Nakaoka, E. Miyoshi, and T. Kuwae. Seed dispersal in the seagrass Zostera marina is mostly within the parent bed in a protected bay. Marine Ecology Progress Series, 523:41–56, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m523p041.pdf.

Hubbard:2014:MCP

[HOA14] K. A. Hubbard, C. E. Olson, and E. V. Armbrust. Molecular characterization of Pseudo-nitzschia community structure and species ecology in a hydrographically complex estuarine system (Puget Sound, Washington, USA). Marine Ecology Progress Series, 507:39–55, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/507/m507p039.pdf.

Hoey:2010:SMM

[Hoe10] A. S. Hoey. Size matters: macroalgal height influences the feeding response of coral reef herbivores. Marine Ecology Progress Series, 411:299–302, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/411/m411p299.pdf.

Hunsicker:2012:PTC

[HOE^{+}12] M. E. Hunsicker, R. J. Olson, T. E. Essington, M. N. Maunder, L. M. Duffy, and J. F. Kitchell. Potential for topdown control on tropical tunas based on size structure of predator–prey interactions. Marine Ecology Progress Series, 445:263–277, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/445/m445p263.pdf.

Hoff:2010:ISN

[Hof10] G. R. Hoff. Identification of skate nursery habitat in the eastern Bering Sea. Marine Ecology Progress Series, 403: 243–254, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/403/m403p243.pdf.

Henderson:2017:MRS

[HOL^{+}17] C. J. Henderson, A. D. Olds, S. Y. Lee, B. L. Gilby, P. S. Maxwell, R. M. Connolly, and T. Stevens. Marine reserves and seascape context shape fish assemblages in seagrass ecosystems. Marine Ecology Progress Series, 566:135– 144, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/566/m566p135.pdf.

Huete-Ortega:2014:DPS

[HORRLS^{+}14] M. Huete-Ortega, T. Rodríguez-Ramos, D. C. López-Sandoval, P. Cermeño, J. M. Blanco, R. L. Palomino, J. Rodríguez, and E. Marañón. Distinct patterns in the sizescaling of abundance and metabolism in coastal and openocean phytoplankton communities. Marine Ecology Progress Series, 515:61–71, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/515/m515p061.pdf.

Hosono:2014:TER

[Hos14] T. Hosono. Temperature explains reproductive dynamics in caprellids at different latitudes. Marine Ecology Progress Series, 511:129–141, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/511/m511p129.pdf.

Hall:2011:VMP

[HP11a] N. S. Hall and H. W. Pearl. Vertical migration patterns of phytoflagellates in relation to light and nutrient availability in a shallow microtidal estuary. Marine Ecology Progress Series, 425:1–19, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Hart:2011:ATI

[HP11b] P. J. B. Hart and E. Pearson. An application of the theory of island biogeography to fish speciation on seamounts. Marine Ecology Progress Series, 430:281–288, 2011. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m430p281.pdf.

Hammer:2013:DWP

[HP13] K. M. Hammer and S. A. Pedersen. Deep-water prawn Pandalus borealis displays a relatively high pH regulatory capacity in response to CO_{2}-induced acidosis. Marine Ecology Progress Series, 492:139–151, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/492/m492p139.pdf.

Hondula:2014:MSC

[HP14] K. L. Hondula and M. L. Pace. Macroalgal support of cultured hard clams in a low nitrogen coastal lagoon. Marine Ecology Progress Series, 498:187–201, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/498/m498p187.pdf.

Hanke:2017:EIO

[HPA17a] M. H. Hanke, M. H. Posey, and T. D. Alphin. The effects of intertidal oyster reef habitat characteristics on faunal utilization. Marine Ecology Progress Series, 581:57–70, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/581/
m581p057.pdf.

Hanke:2017:IHC

[HPA17b] M. H. Hanke, M. H. Posey, and T. D. Alphin. The influence of habitat characteristics on intertidal oyster Crassostrea virginica populations. Marine Ecology Progress Series, 571:121–138, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/571/m571p121.pdf.

Halliday:2018:PMS

[HPB^{+}18] W. D. Halliday, M. K. Pine, A. P. H. Bose, S. Balshine, and F. Juanes. The plainfin midshipman’s soundscape at two sites around Vancouver Island, British Columbia. Marine Ecology Progress Series, 603:189–200, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/603/m603p189.pdf.

Harris:2015:BII

[HPH^{+}15] R. J. Harris, C. A. Pilditch, J. E. Hewitt, A. M. Lohrer, C. Van Colen, M. Townsend, and S. F. Thrush. Biotic interactions influence sediment erodibility on wave-exposed sandflats. Marine Ecology Progress Series, 523:15–30, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/523/
m523p015.pdf.

Huebert:2018:MLF

[HPH^{+}18] K. B. Huebert, J. Pätsch, M. Hufnagl, M. Kreus, and M. A. Peck. Modeled larval fish prey fields and growth rates help predict recruitment success of cod and anchovy in the North Sea. Marine Ecology Progress Series, 600:111–126, 2018. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m600p111.pdf.

Hammerschlag-Peyer:2010:IVH

[HPL10] C. M. Hammerschlag-Peyer and C. A. Layman. Intrapopulation variation in habitat use by two abundant coastal fish species. Marine Ecology Progress Series, 415:211–220, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2010/415/m415p211.pdf.

Ho:2012:PGS

[HPL12] A. L. F. C. Ho, C. L. Pruett, and J. Lin. Population genetic structure, coloration, and morphometrics of yellowhead jawfish Opistognathus aurifrons (Perciformes: Opistognathidae) in the Caribbean region. Marine Ecology Progress Series, 444:275–287, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/444/m444p275.pdf.

Han:2016:DSA

[HPL^{+}16] J. Han, J. Puthumana, M. C. Lee, S. Kim, and J. S. Lee. Different susceptibilities of the Antarctic and temperate copepods Tigriopus kingsejongensis and T. japonicus to ultraviolet (UV) radiation. Marine Ecology Progress Series, 561: 99–107, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/561/m561p099.pdf.

Hurrey:2013:MSR

[HPLS13] L. P. Hurrey, C. R. Pitcher, C. E. Lovelock, and S. Schmidt. Macroalgal species richness and assemblage composition of the Great Barrier Reef seabed. Marine Ecology Progress Series, 492:69–83, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/492/m492p069.pdf.

Holstein:2014:CIM

[HPM14] D. M. Holstein, C. B. Paris, and P. J. Mumby. Consistency and inconsistency in multispecies population network dynamics of coral reef ecosystems. Marine Ecology Progress Series, 499:1– 18, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Hammerschmidt:2011:DMA

[HPMB11] K. Hammerschmidt, A. J. Pemberton, N. K. Michiels, and J. D. D. Bishop. Differential maternal allocation following mixed insemination contributes to variation in oocyte size in a sea squirt. Marine Ecology Progress Series, 422:123– 128, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/422/m422p123.pdf.

Herfort:2011:PRG

[HPMZ11] L. Herfort, T. D. Peterson, L. A. McCue, and P. Zuber. Protist 18S rRNA gene sequence analysis reveals multiple sources of organic matter contributing to turbidity maxima of the Columbia River estuary. Marine Ecology Progress Series, 438:19–31, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2011/438/m438p019.pdf. See errata

[Ano14b].

Hitt:2011:DMF

[HPN11] S. Hitt, S. J. Pittman, and R. S. Nemeth. Diel movements of fishes linked to benthic seascape structure in a Caribbean coral reef ecosystem. Marine Ecology Progress Series, 427: 275–291, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m427p275.pdf.

Hardge:2017:SIO

[HPN^{+}17] K. Hardge, I. Peeken, S. Neuhaus, T. Krumpen, T. Stoeck, and K. Metfies. Sea ice origin and sea ice retreat as possible drivers of variability in Arctic marine protist composition. Marine Ecology Progress Series, 571:43–57, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/571/
m571p043.pdf.

Hill:2010:QWE

[HPP^{+}10] N. A. Hill, A. R. Pepper, M. L. Puotinen, M. G. Hughes, G. J. Edgar, N. S. Barrett, R. D. Stuart-Smith, and R. Leaper. Quantifying wave exposure in shallow temperate reef systems: applicability of fetch models for predicting algal biodiversity. Marine Ecology Progress Series, 417:83–95, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/417/
m417p083.pdf.

Hurst:2015:EMT

[HPQ15] T. A. Hurst, A. J. Pope, and G. P. Quinn. Exposure mediates transitions between bare and vegetated states in temperate mangrove ecosystems. Marine Ecology Progress Series, 533:121–134, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/533/m533p121.pdf.

Hinke:2012:FRT

[HPR^{+}12] J. T. Hinke, M. J. Polito, C. S. Reiss, S. G. Trivelpiece, and W. Z. Trivelpiece. Flexible reproductive timing can buffer reproductive success of Pygoscelis spp. penguins in the Antarctic Peninsula region. Marine Ecology Progress Series, 454:91–104, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/454/m454p091.pdf.

Hansen:2012:WTD

[HR12] J. C. R. Hansen and M. A. Reidenbach. Wave and tidally driven flows in eelgrass beds and their effect on sediment suspension. Marine Ecology Progress Series, 448:271–287, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m448p271.pdf.

Hubbard:2015:ELS

[HR15] A. B. Hubbard and M. A. Reidenbach. Effects of larval swimming behavior on the dispersal and settlement of the eastern oyster Crassostrea virginica. Marine Ecology Progress Series, 535:161–176, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/535/m535p161.pdf.

Hogan:2019:QMI

[HR19] S. Hogan and M. A. Reidenbach. Quantifying and mapping intertidal oyster reefs utilizing LiDAR-based remote sensing. Marine Ecology Progress Series, 630:83–99, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/630/
m630p083.pdf.

Hidalgo:2014:HBC

[HRÁB^{+}14] M. Hidalgo, P. Reglero, D. Álvarez-Berastegui, A. P. Torres, I. Álvarez, J. M. Rodriguez, A. Carbonell, N. Zaragoza, A. Tor, R. Goñi, S. Mallol, R. Balbín, and F. Alemany. Hydrographic and biological components of the seascape structure the meroplankton community in a frontal system. Marine Ecology Progress Series, 505:65–80, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/505/m505p065.pdf.

Hidas:2013:ATR

[HRAM13] E. Z. Hidas, K. G. Russell, D. J. Ayre, and T. E. Minchinton. Abundance of Tesseropora rosea at the margins of its biogeographic range is closely linked to recruitment, but not fecundity. Marine Ecology Progress Series, 483:199– 208, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/483/m483p199.pdf.

Hoeksema:2012:TAR

[HRC12] B. W. Hoeksema, P. J. Roos, and G. C. Cadée. Trans-Atlantic rafting by the brooding reef coral Favia fragum on man-made flotsam. Marine Ecology Progress Series, 445:209–218, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m445p209.pdf.

Hunt:2018:TSI

[HRK^{+}18] G. L. Hunt, Jr., M. Renner, K. J. Kuletz, S. Salo, L. Eisner, P. H. Ressler, C. Ladd, and J. A. Santora. Timing of seaice retreat affects the distribution of seabirds and their prey in the southeastern Bering Sea. Marine Ecology Progress Series, 593:209–230, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m593p209.pdf.

Hidalgo:2011:SEF

[HRM^{+}11] M. Hidalgo, T. Rouyer, J. C. Molinero, E. Massutí, J. Moranta, B. Guijarro, and N. C. Stenseth. Synergistic effects of fishing-induced demographic changes and climate variation on fish population dynamics. Marine Ecology Progress Series, 426:1–12, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Hagerty:2018:CNL

[HRP18] M. L. Hagerty, N. Reyns, and J. Pineda. Constrained nearshore larval distributions and thermal stratification. Marine Ecology Progress Series, 595:105–122, 2018. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_

oa/m595p105.pdf. See corrigendum [Ano19a].

Hernandez-Ruiz:2018:AAU

[HRPBLT18] M. Hernández-Ruiz, A. Prieto, E. Barber-Lluch, and E. Teira. Amino acid utilization by eukaryotic picophytoplankton in a coastal upwelling system. Marine Ecology Progress Series, 588:43–57, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/588/m588p043.pdf.

Howell:2016:OSD

[HRS^{+}16] L. N. Howell, K. J. Reich, D. J. Shaver, A. M. Landry, Jr., and C. C. Gorga. Ontogenetic shifts in diet and habitat of juvenile green sea turtles in the northwestern Gulf of Mexico. Marine Ecology Progress Series, 559:217–229, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/559/
m559p217.pdf.

Hedger:2017:DBA

[HRS^{+}17] R. D. Hedger, A. H. Rikardsen, J. F. Strøm, D. A. Righton, E. B. Thorstad, and T. F. Næsje. Diving behaviour of Atlantic salmon at sea: effects of light regimes and temperature stratification. Marine Ecology Progress Series, 574: 127–140, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/574/m574p127.pdf.

Hitchman:2012:LDS

[HRT12] S. M. Hitchman, N. B. Reyns, and A. R. Thompson. Larvae define spawning habitat of bocaccio rockfish Sebastes paucispinis within and around a large southern California marine reserve. Marine Ecology Progress Series, 465:227–242, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/465/
m465p227.pdf.

Hertz:2014:EPP

[HRT^{+}14] E. Hertz, J. P. W. Robinson, M. Trudel, A. Mazumder, and J. K. Baum. Estimation of predator–prey mass ratios using stable isotopes: sources of errors. Marine Ecology Progress Series, 516:1–6, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Hart:2011:MSS

[HS11a] D. R. Hart and B. V. Shank. Mortality of sea scallops Placopecten magellanicus in the Mid-Atlantic Bight: Comment on Stokesbury et al. (2011). Marine Ecology Progress Series, 443:293–297, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/443/m443p293.pdf. See

[SCHO11b] and reply [SCHO11a].

Heupel:2011:ENA

[HS11b] M. R. Heupel and C. A. Simpfendorfer. Estuarine nursery areas provide a low-mortality environment for young bull sharks Carcharhinus leucas. Marine Ecology Progress Series, 433:237–244, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/433/m433p237.pdf.

Haraldsson:2014:SDL

[HS14a] M. Haraldsson and V. Siegel. Seasonal distribution and life history of Thysanoessa macrura (Euphausiacea, Crustacea) in high latitude waters of the Lazarev Sea, Antarctica. Marine Ecology Progress Series, 495:105–118, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/495/
m495p105.pdf.

Heupel:2014:IEB

[HS14b] M. R. Heupel and C. A. Simpfendorfer. Importance of environmental and biological drivers in the presence and space use of a reef-associated shark. Marine Ecology Progress Series, 496:47–57, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/496/m496p047.pdf.

Hamza:2015:HFH

[HS15a] F. Hamza and S. Selmi. Habitat features and human presence as predictors of the abundance of shorebirds and wading birds wintering in the Gulf of Gabès, Tunisia. Marine Ecology Progress Series, 540:251–258, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/540/m540p251.pdf.

Howe:2015:USI

[HS15b] E. R. Howe and C. A. Simenstad. Using stable isotopes to discern mechanisms of connectivity in estuarine detritusbased food webs. Marine Ecology Progress Series, 518:13– 29, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/518/m518p013.pdf.

Hickford:2016:OMA

[HS16] M. J. H. Hickford and D. R. Schiel. Otolith microchemistry of the amphidromous Galaxias maculatus shows recruitment to coastal rivers from unstructured larval pools. Marine Ecology Progress Series, 548:197–207, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/548/m548p197.pdf.

Helyer:2017:FEI

[HS17] J. Helyer and J. F. Samhouri. Fishing and environmental influences on estimates of unfished herbivorous fish biomass across the Hawaiian Archipelago. Marine Ecology Progress Series, 575:1–15, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Howes:2015:PTS

[HSA^{+}15] E. L. Howes, L. Stemmann, C. Assailly, J. O. Irisson, M. Dima, J. Bijma, and J. P. Gattuso. Pteropod time series from the North Western Mediterranean (1967–2003): impacts of pH and climate variability. Marine Ecology Progress Series, 531:193–206, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m531p193.pdf.

Hofmann:2012:CBC

[HSB12] L. C. Hofmann, S. Straub, and K. Bischof. Competition between calcifying and noncalcifying temperate marine macroalgae under elevated CO_{2} levels. Marine Ecology Progress Series, 464:89–105, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/464/m464p089.pdf.

Hartmann:2015:COC

[HSC^{+}15] A. C. Hartmann, S. A. Sandin, V. F. Chamberland, K. L. Marhaver, J. M. de Goeij, and M. J. A. Vermeij. Crude oil contamination interrupts settlement of coral larvae after direct exposure ends. Marine Ecology Progress Series, 536:163–173, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m536p163.pdf.

Hoskins:2017:FNS

[HSCA17] A. J. Hoskins, N. Schumann, D. P. Costa, and J. P. Y. Arnould. Foraging niche separation in sympatric temperatelatitude fur seal species. Marine Ecology Progress Series, 566:229–241, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/566/m566p229.pdf.

Higgins:2013:CPS

[HSDB13] K. L. Higgins, J. M. Semmens, Z. A. Doubleday, and C. P. Burridge. Comparisons of population structuring in sympatric octopus species with and without a pelagic larval stage. Marine Ecology Progress Series, 486:203–212, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/486/
m486p203.pdf.

Heisterkamp:2010:NOP

[HSdBS10] I. M. Heisterkamp, A. Schramm, D. de Beer, and P. Stief. Nitrous oxide production associated with coastal marine invertebrates. Marine Ecology Progress Series, 415:1–9, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Hart:2012:HRH

[HSF^{+}12] K. M. Hart, A. R. Sartain, I. Fujisaki, H. L. Pratt, Jr., D. Morley, and M. W. Feeley. Home range, habitat use, and migrations of hawksbill turtles tracked from Dry Tortugas National Park, Florida, USA. Marine Ecology Progress Series, 457:193–207, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m457p193.pdf.

Hamylton:2012:SMB

[HSH12] S. Hamylton, T. Spencer, and A. B. Hagan. Spatial modelling of benthic cover using remote sensing data in the Aldabra lagoon, western Indian Ocean. Marine Ecology Progress Series, 460:35–47, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/460/m460p035.pdf.

Heins:2018:AIT

[HSH18] A. Heins, I. Sötje, and S. Holst. Assessment of investigation techniques for scyphozoan statoliths, with focus on early development of the jellyfish Sanderia malayensis. Marine Ecology Progress Series, 591:37–56, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m591p037.pdf.

Harrison:2013:FEE

[HSM13] C. S. Harrison, D. A. Siegel, and S. Mitarai. Filamentation and eddy–eddy interactions in marine larval accumulation and transport. Marine Ecology Progress Series, 472: 27–44, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/472/m472p027.pdf.

Hof:2017:FCS

[HSMB17] C. A. M. Hof, E. Smallwood, J. Meager, and I. P. Bell. First citizen-science population abundance and growth rate estimates for green sea turtles Chelonia mydas foraging in the northern Great Barrier Reef, Australia. Marine Ecology Progress Series, 574:181–191, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/574/m574p181.pdf.

Hill-Spanik:2015:PPD

[HSMS^{+}15] K. M. Hill-Spanik, J. R. McDowell, N. A. Stokes, K. S. Reece, E. M. Burreson, and R. B. Carnegie. Phylogeographic perspective on the distribution and dispersal of a marine pathogen, the oyster parasite Bonamia exitiosa. Marine Ecology Progress Series, 536:65–76, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/536/m536p065.pdf.

Hamilton:2014:QPF

[HSPS14] S. L. Hamilton, J. E. Smith, N. N. Price, and S. A. Sandin. Quantifying patterns of fish herbivory on Palmyra Atoll (USA), an uninhabited predator-dominated central Pacific coral reef. Marine Ecology Progress Series, 501:141–155, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m501p141.pdf.

Hubas:2010:BNS

[HSR^{+}10] C. Hubas, C. Sachidhanandam, H. Rybarczyk, H. V. Lubarsky, A. Rigaux, T. Moens, and D. M. Paterson. Bacterivorous nematodes stimulate microbial growth and exopolymer production in marine sediment microcosms. Marine Ecology Progress Series, 419:85–94, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/419/m419p085.pdf.

Holles:2013:BND

[HSR^{+}13] S. Holles, S. D. Simpson, A. N. Radford, L. Berten, and D. Lecchini. Boat noise disrupts orientation behaviour in a coral reef fish. Marine Ecology Progress Series, 485:295– 300, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/485/m485p295.pdf.

Haynert:2011:BDF

[HSRP11] K. Haynert, J. Schönfeld, U. Riebesell, and I. Polovodova. Biometry and dissolution features of the benthic foraminifer Ammonia aomoriensis at high pCO 2. Marine Ecology Progress Series, 432:53–67, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/432/m432p053.pdf.

Hoque:2010:SVS

[HSS^{+}10] A. T. M. R. Hoque, S. Sharma, R. Suwa, S. Mori, and A. Hagihara. Seasonal variation in the size-dependent respiration of mangroves Kandelia obovata. Marine Ecology Progress Series, 404:31–37, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/404/m404p031.pdf. See corrigen-

dum [Ano12c].

Humphries:2017:DVM

[HSS17] N. E. Humphries, S. J. Simpson, and D. W. Sims. Diel vertical migration and central place foraging in benthic predators. Marine Ecology Progress Series, 582:163–180, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/582/
m582p163.pdf.

Henschke:2018:MPD

[HSS18] N. Henschke, C. A. Stock, and J. L. Sarmiento. Modeling population dynamics of scyphozoan jellyfish (Aurelia spp.) in the Gulf of Mexico. Marine Ecology Progress Series, 591:167–183, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/591/m591p167.pdf.

Huebner:2019:CSC

[HSS^{+}19] L. K. Huebner, C. P. Shea, P. M. Schueller, A. D. Terrell, S. G. Ratchford, and N. E. Chadwick. Crustacean symbiosis with Caribbean sea anemones Bartholomea annulata: occupancy modeling, habitat partitioning, and persistence. Marine Ecology Progress Series, 631:99–116, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/631/
m631p099.pdf.

Humphries:2016:TCT

[HSWS16] N. E. Humphries, S. J. Simpson, V. J. Wearmouth, and D. W. Sims. Two’s company, three’s a crowd: fine-scale habitat partitioning by depth among sympatric species of marine mesopredator. Marine Ecology Progress Series, 561: 173–187, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/561/m561p173.pdf.

Haselmair:2010:BMB

[HSZR10] A. Haselmair, M. Stachowitsch, M. Zuschin, and B. Riedel. Behaviour and mortality of benthic crustaceans in response to experimentally induced hypoxia and anoxia in situ. Marine Ecology Progress Series, 414:195–208, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m414p195.pdf.

Hewitt:2010:EEA

[HT10] J. E. Hewitt and S. F. Thrush. Empirical evidence of an approaching alternate state produced by intrinsic community dynamics, climatic variability and management actions. Marine Ecology Progress Series, 413:267–276, 2010. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m413p267.pdf.

Hufnagl:2011:GBSa

[HT11a] M. Hufnagl and A. Temming. Growth in the brown shrimp Crangon crangon. I. Effects of food, temperature, size, gender, moulting, and cohort. Marine Ecology Progress Series, 435:141–154, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/435/m435p141.pdf.

Hufnagl:2011:GBSb

[HT11b] M. Hufnagl and A. Temming. Growth in the brown shrimp Crangon crangon. II. Meta-analysis and modelling. Marine Ecology Progress Series, 435:155–172, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/435/
m435p155.pdf.

Hirai:2015:MCA

[HT15] J. Hirai and A. Tsuda. Metagenetic community analysis of epipelagic planktonic copepods in the tropical and subtropical Pacific. Marine Ecology Progress Series, 534:65– 78, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/534/m534p065.pdf.

Hertz:2015:CSV

[HTB^{+}15] E. Hertz, M. Trudel, R. D. Brodeur, E. A. Daly, L. Eisner, E. V. Farley, Jr., J. A. Harding, R. B. MacFarlane, S. Mazumder, J. H. Moss, J. M. Murphy, and A. Mazumder. Continental-scale variability in the feeding ecology of juvenile Chinook salmon along the coastal Northeast Pacific Ocean. Marine Ecology Progress Series, 537:247–263, 2015. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m537p247.pdf.

Hellemann:2017:DOE

[HTB^{+}17] D. Hellemann, P. Tallberg, I. Bartl, M. Voss, and S. Hietanen. Denitrification in an oligotrophic estuary: a delayed sink for riverine nitrate. Marine Ecology Progress Series, 583:63–80, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m583p063.pdf.

Howard:2017:CEI

[HTC17] B. R. Howard, T. W. Therriault, and I. M. Côté. Contrasting ecological impacts of native and non-native marine crabs: a global meta-analysis. Marine Ecology Progress Series, 577:93–103, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/577/m577p093.pdf.

Hertz:2018:OCS

[HTCH^{+}18] E. Hertz, M. Trudel, M. Carrasquilla-Henao, L. Eisner, E. V. Farley, Jr., J. H. Moss, J. M. Murphy, and A. Mazumder. Oceanography and community structure drive zooplankton carbon and nitrogen dynamics in the eastern Bering Sea. Marine Ecology Progress Series, 601:97–108, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/601/
m601p097.pdf.

Hogan:2010:VCI

[HTH10] J. D. Hogan, R. J. Thiessen, and D. D. Heath. Variability in connectivity indicated by chaotic genetic patchiness within and among populations of a marine fish. Marine Ecology Progress Series, 417:263–275, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/417/m417p263.pdf.

Hosia:2011:IBN

[HTHH11] A. Hosia, J. Titelman, L. J. Hansson, and M. Haraldsson. Interactions between native and alien ctenophores: Beroe gracilis and Mnemiopsis leidyi in gullmarsfjorden. Marine Ecology Progress Series, 422:129–138, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/422/m422p129.pdf.

Hafker:2018:CFD

[HTHM18] N. S. Häfker, M. Teschke, L. Hüppe, and B. Meyer. Calanus finmarchicus diel and seasonal rhythmicity in relation to endogenous timing under extreme polar photoperiods. Marine Ecology Progress Series, 603:79–92, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m603p079.pdf.

Hasegawa:2017:CBO

[HTK^{+}17] T. Hasegawa, N. Takatsuki, Y. Kawabata, R. Kawabe, G. N. Nishihara, A. Ishimatsu, K. Soyano, K. Okamura, S. Furukawa, M. Yamada, M. Shimoda, T. Kinoshita, N. Yamawaki, Y. Morii, and Y. Sakakura. Continuous behavioral observation reveals the function of drifting seaweeds for Seriola spp. juveniles. Marine Ecology Progress Series, 573:101–115, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/573/m573p101.pdf.

Hinz:2011:ESD

[HTR^{+}11] H. Hinz, D. Tarrant, A. Ridgeway, M. J. Kaiser, and J. G. Hiddink. Effects of scallop dredging on temperate reef fauna. Marine Ecology Progress Series, 432:91–102, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/432/
m432p091.pdf.

Hawkes:2012:MPH

[HTR^{+}12] L. A. Hawkes, J. Tomás, O. Revuelta, Y. M. León, J. M. Blumenthal, A. C. Broderick, M. Fish, J. A. Raga, M. J. Witt, and B. J. Godley. Migratory patterns in hawksbill turtles described by satellite tracking. Marine Ecology Progress Series, 461:223–232, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/461/m461p223.pdf.

Hayashi:2019:SSD

[HTR19a] K. Hayashi, K. Tachihara, and J. D. Reimer. Species and sexual differences in human-oriented behavior of anemonefish at Okinawa Island, Japan. Marine Ecology Progress Series, 616:219–224, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/616/m616p219.pdf.

Hovinen:2019:RBI

[HTR^{+}19b] J. E. H. Hovinen, A. Tarroux, F. Ramírez, M. Forero, and S. Descamps. Relationships between isotopic ratios, body condition and breeding success in a high Arctic seabird community. Marine Ecology Progress Series, 613:183–195, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2019/613/m613p183.pdf.

Harms-Tuohy:2016:UDM

[HTSA16] C. A. Harms-Tuohy, N. V. Schizas, and R. S. Appeldoorn. Use of DNA metabarcoding for stomach content analysis in the invasive lionfish Pterois volitans in Puerto Rico. Marine Ecology Progress Series, 558:181–191, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m558p181.pdf.

Haraldsson:2012:RBF

[HTT^{+}12] M. Haraldsson, K. Tönnesson, P. Tiselius, T. F. Thingstad, and D. L. Aksnes. Relationship between fish and jellyfish as a function of eutrophication and water clarity. Marine Ecology Progress Series, 471:73–85, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m471p073.pdf.

Herbeck:2013:PAE

[HU13] L. S. Herbeck and D. Unger. Pond aquaculture effluents traced along back-reef waters by standard water quality parameters, δ ^{15}N in suspended matter and phytoplankton bioassays. Marine Ecology Progress Series, 478:71–86, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/478/
m478p071.pdf.

Hughes:2011:WRF

[Hug11] D. J. Hughes. Where’s the ‘reef’? A five year study of serpulid tube bioerosion in a Scottish sea loch. Marine Ecology Progress Series, 430:273–280, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m430p273.pdf.

Hernandez:2014:PDS

[Hv14] A. L. Molina Hernández and B. I. van Tussenbroek. Patch dynamics and species shifts in seagrass communities under moderate and high grazing pressure by green sea turtles. Marine Ecology Progress Series, 517:143–157, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/517/
m517p143.pdf.

Heyns-Veale:2019:CWE

[HVBG^{+}19] E. R. Heyns-Veale, A. T. F. Bernard, A. Götz, B. Q. Mann, J. Q. Maggs, and M. K. S. Smith. Community-wide effects of protection reveal insights into marine protected area effectiveness for reef fish. Marine Ecology Progress Series, 620: 99–117, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m620p099.pdf.

Huisman:2016:HSN

[HVD^{+}16] J. Huisman, P. Verhelst, K. Deneudt, P. Goethals, T. Moens, L. A. J. Nagelkerke, C. Nolting, J. Reubens, P. P. Schollema, H. V. Winter, and A. Mouton. Heading south or north: novel insights on European silver eel Anguilla anguilla migration in the North Sea. Marine Ecology Progress Series, 554:257–262, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/554/m554p257.pdf.

Holland:2012:ECN

[HvEBC12] D. P. Holland, I. van Erp, J. Beardall, and P. L. M. Cook. Environmental controls on the nitrogen-fixing cyanobacterium Nodularia spumigena in a temperate lagoon system in SE Australia. Marine Ecology Progress Series, 461:47–57, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2012/461/m461p047.pdf.

Hancock:2018:SIR

[HVJ^{+}18] J. M. Hancock, S. Vieira, V. Jimenez, J. Carvalho Rio, and R. Rebelo. Stable isotopes reveal dietary differences and site fidelity in juvenile green turtles foraging around São Tomé Island, West Central Africa. Marine Ecology Progress Series, 600:165–177, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/600/m600p165.pdf.

Heithaus:2013:ARP

[HVK^{+}13] M. R. Heithaus, J. J. Vaudo, S. Kreicker, C. A. Layman, M. Krützen, D. A. Burkholder, K. Gastrich, C. Bessey, R. Sarabia, K. Cameron, A. Wirsing, J. A. Thomson, and M. M. Dunphy-Daly. Apparent resource partitioning and trophic structure of large-bodied marine predators in a relatively pristine seagrass ecosystem. Marine Ecology Progress Series, 481:225–237, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/481/m481p225.pdf.

Haye:2012:GSR

[HVT12] P. A. Haye, A. I. Varela, and M. Thiel. Genetic signatures of rafting dispersal in algal-dwelling brooders Limnoria spp. (Isopoda) along the SE Pacific (Chile). Marine Ecology Progress Series, 455:111–122, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/455/m455p111.pdf.

Haydar:2011:PIP

[HW11a] D. Haydar and W. J. Wolff. Predicting invasion patterns in coastal ecosystems: relationship between vector strength and vector tempo. Marine Ecology Progress Series, 431:1–10, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Hemond:2011:MLI

[HW11b] E. M. Hemond and A. E. Wilbur. Microsatellite loci indicate population structure and selection between Atlantic and Gulf of Mexico populations of the bay scallop Argopecten irradians. Marine Ecology Progress Series, 423:131–142, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/423/
m423p131.pdf.

Harden:2012:USB

[HW12a] L. A. Harden and A. S. Williard. Using spatial and behavioral data to evaluate the seasonal bycatch risk of diamondback terrapins Malaclemys terrapin in crab pots. Marine Ecology Progress Series, 467:207–217, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/467/m467p207.pdf.

Hunt:2012:TIMb

[HW12b] G. L. Hunt, Jr. and R. P. Wilson. Technological innovation in marine ornithology. Marine Ecology Progress Series, 451:227–229, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m451p227.pdf.

Hunt:2012:TIMa

[HW12c] George L. Hunt, Jr. and Rory P. Wilson. Technological innovation in marine ornithology. Marine Ecology Progress Series, 451:227–304, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m451_ThemeSection.pdf.

Hill:2013:HCP

[HW13] J. M. Hill and M. J. Weissburg. Habitat complexity and predator size mediate interactions between intraguild blue crab predators and mud crab prey in oyster reefs. Marine Ecology Progress Series, 488:209–219, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/488/m488p209.pdf.

Hennicke:2014:CVO

[HW14] J. C. Hennicke and H. Weimerskirch. Coping with variable and oligotrophic tropical waters: foraging behaviour and flexibility of the Abbott’s booby Papasula abbotti. Marine Ecology Progress Series, 499:259–273, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/499/m499p259.pdf.

Howarth:2018:EBT

[HWB^{+}18] L. M. Howarth, J. J. Waggitt, S. G. Bolam, J. Eggleton, P. J. Somerfield, and J. G. Hiddink. Effects of bottom trawling and primary production on the composition of biological traits in benthic assemblages. Marine Ecology Progress Series, 602:31–48, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/602/m602p031.pdf.

Henderiks:2012:ECE

[HWE^{+}12] J. Henderiks, A. Winter, M. Elbrächter, R. Feistel, A. van der Plas, G. Nausch, and R. Barlow. Environmental controls on Emiliania huxleyi morphotypes in the Benguela coastal upwelling system (SE Atlantic). Marine Ecology Progress Series, 448:51–66, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m448p051.pdf.

Hanson:2013:RML

[HWET13] N. N. Hanson, C. M. Wurster, Edinburgh Ion Microprobe Facility (EIMF), and C. D. Todd. Reconstructing marine lifehistory strategies of wild Atlantic salmon from the stable isotope composition of otoliths. Marine Ecology Progress Series, 475:249–266, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/475/m475p249.pdf.

Hammann:2013:ISS

[HWR^{+}13] M. Hammann, G. Wang, E. Rickert, S. M. Boo, and F. Weinberger. Invasion success of the seaweed Gracilaria vermiculophylla correlates with low palatibility. Marine Ecology Progress Series, 486:93–103, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m486p093.pdf.

Hovinen:2014:SDF

[HWR^{+}14] J. E. H. Hovinen, J. Welcker, A. Rabindranath, Z. W. Brown, H. Hop, J. Berge, and H. Steen. At-sea distribution of foraging little auks relative to physical factors and food supply. Marine Ecology Progress Series, 503:263–277, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/503/
m503p263.pdf.

Hanley:2019:EGS

[HWSK19] T. C. Hanley, J. W. White, C. D. Stallings, and D. L. Kimbro. Environmental gradients shape the combined effects of multiple parasites on oyster hosts in the northern Gulf of Mexico. Marine Ecology Progress Series, 612:111–125, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/612/
m612p111.pdf.

Hoffle:2012:DAI

[HWTH12] H. Höffle, T. Wernberg, M. S. Thomsen, and M. Holmer. Drift algae, an invasive snail and elevated temperature reduce ecological performance of a warm-temperate seagrass, through additive effects. Marine Ecology Progress Series, 450: 67–80, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/450/m450p067.pdf.

Herrera:2019:CCI

[HWVM19] M. Herrera, D. S. Wethey, E. Vázquez, and G. Macho. Climate change implications for reproductive success: temperature effect on penis development in the barnacle Semibalanus balanoides. Marine Ecology Progress Series, 610:109– 123, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/610/m610p109.pdf.

Huang:2014:PNR

[HYC^{+}14] H. Huang, X. C. Yuan, W. J. Cai, C. L. Zhang, X. Li, and S. Liu. Positive and negative responses of coral calcification to elevated pCO_{2}: case studies of two coral species and the implications of their responses. Marine Ecology Progress Series, 502:145–156, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/502/m502p145.pdf.

Huff:2013:EFI

[HYL^{+}13] D. D. Huff, M. M. Yoklavich, M. S. Love, D. L. Watters, F. Chai, and S. T. Lindley. Environmental factors that influence the distribution, size, and biotic relationships of the Christmas tree coral Antipathes dendrochristos in the Southern California Bight. Marine Ecology Progress Series, 494: 159–177, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m494p159.pdf.

Iriarte:2010:DOC

[IAV^{+}10] A. Iriarte, G. Aravena, F. Villate, I. Uriarte, B. Ibáñez, M. Llope, and N. C. Stenseth. Dissolved oxygen in contrasting estuaries of the Bay of Biscay: effects of temperature, river discharge and chlorophyll-a. Marine Ecology Progress Series, 418:57–71, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/418/m418p057.pdf.

Ibanez:2011:GDD

[ICT^{+}11] C. M. Ibáñez, L. A. Cubillos, R. Tafur, J. Argüelles, C. Yamashiro, and E. Poulin. Genetic diversity and demographic history of Dosidicus gigas (Cephalopoda: Ommastrephidae) in the Humboldt Current System. Marine Ecology Progress Series, 431:163–171, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2011/431/m431p163.pdf.

Ivesa:2010:DPD

[ICUM10] L. Iveša, M. G. Chapman, A. J. Underwood, and R. J. Murphy. Differential patterns of distribution of limpets on intertidal seawalls: experimental investigation of the roles of recruitment, survival and competition. Marine Ecology Progress Series, 407:55–69, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m407p055.pdf.

Irigoyen:2018:HUS

[IDT^{+}18] A. J. Irigoyen, A. M. De Wysiecki, G. Trobbiani, N. Bovcon, C. A. Awruch, F. Argemi, and A. J. Jaureguizar. Habitat use, seasonality and demography of an apex predator: sevengill shark Notorynchus cepedianus in northern Patagonia. Marine Ecology Progress Series, 603:147–160, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/603/
m603p147.pdf.

Infantes:2016:EZM

[IEM16] E. Infantes, L. Eriander, and P. O. Moksnes. Eelgrass (Zostera marina) restoration on the west coast of Sweden using seeds. Marine Ecology Progress Series, 546:31–45, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/546/
m546p031.pdf.

Izzo:2016:VCD

[IHD^{+}16] C. Izzo, C. Huveneers, M. Drew, C. J. A. Bradshaw, S. C. Donnellan, and B. M. Gillanders. Vertebral chemistry demonstrates movement and population structure of bronze whaler. Marine Ecology Progress Series, 556:195–207, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/556/
m556p195.pdf.

Idjadi:2010:RSU

[IHP10] J. A. Idjadi, R. N. Haring, and W. F. Precht. Recovery of the sea urchin Diadema antillarum promotes scleractinian coral growth and survivorship on shallow Jamaican reefs. Marine Ecology Progress Series, 403:91–100, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m403p091.pdf.

Issaris:2012:OEM

[IKS^{+}12] Y. Issaris, S. Katsanevakis, M. Salomidi, K. Tsiamis, N. Katsiaras, and G. Verriopoulos. Occupancy estimation of marine species: dealing with imperfect detectability. Marine Ecology Progress Series, 453:95–106, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/453/m453p095.pdf.

Imbs:2010:DLF

[ILDL10] A. B. Imbs, N. A. Latyshev, T. N. Dautova, and Y. Y. Latypov. Distribution of lipids and fatty acids in corals by their taxonomic position and presence of zooxanthellae. Marine Ecology Progress Series, 409:65–75, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/409/
m409p065.pdf.

Ikeda:2011:MCC

[IMD11] T. Ikeda, A. D. McKinnon, and P. J. Doherty. Metabolism and chemical composition of small teleost fishes from tropical inshore waters. Marine Ecology Progress Series, 435: 197–207, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/435/m435p197.pdf.

Isojunno:2012:HPH

[IME12] S. Isojunno, J. Matthiopoulos, and P. G. H. Evans. Harbour porpoise habitat preferences: robust spatio-temporal inferences from opportunistic data. Marine Ecology Progress Series, 448:155–170, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/448/m448p155.pdf.

Ichii:2011:CAN

[IMS^{+}11] T. Ichii, K. Mahapatra, M. Sakai, T. Wakabayashi, H. Okamura, H. Igarashi, D. Inagake, and Y. Okada. Changes in abundance of the neon flying squid Ommastrephes bartramii in relation to climate change in the central North Pacific Ocean.

Marine Ecology Progress Series, 441:151–164, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/441/

m441p151.pdf. See erratum [Ano12d].

Ikeda:2017:BCA

[IMU17] H. Ikeda, C. Mizota, and S. Uye. Bioenergetic characterization in Aurelia aurita (Cnidaria: Scyphozoa) polyps and application to natural polyp populations. Marine Ecology Progress Series, 568:87–100, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/568/m568p087.pdf.

Ingeman:2016:LCI

[Ing16] K. E. Ingeman. Lionfish cause increased mortality rates and drive local extirpation of native prey. Marine Ecology Progress Series, 558:235–245, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m558p235.pdf.

Irion:2017:PAW

[INK^{+}17] D. T. Irion, L. R. Noble, A. A. Kock, E. Gennari, M. L. Dicken, A. M. Hewitt, A. V. Towner, A. J. Booth, M. J. Smale, and G. Cliff. Pessimistic assessment of white shark population status in South Africa: Comment on Andreotti et al. (2016). Marine Ecology Progress Series, 577:251–255, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/577/

m577p251.pdf. See [ARvdW^{+}16, AvdHH^{+}17].

Igulu:2013:OOW

[INvdB^{+}13] M. M. Igulu, I. Nagelkerken, M. van der Beek, M. Schippers, R. van Eck, and Y. D. Mgaya. Orientation from open water to settlement habitats by coral reef fish: behavioral flexibility in the use of multiple reliable cues. Marine Ecology Progress Series, 493:243–257, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m493p243.pdf.

Infantes:2012:ESP

[IOS^{+}12] E. Infantes, A. Orfila, G. Simarro, J. Terrados, M. Luhar, and H. Nepf. Effect of a seagrass (Posidonia oceanica) meadow on wave propagation. Marine Ecology Progress Series, 456:63–72, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/456/m456p063.pdf.

Itoh:2016:OOF

[IS16] T. Itoh and O. Sakai. Open-ocean foraging ecology of southern bluefin tuna Thunnus maccoyii based on stomach contents. Marine Ecology Progress Series, 555:203–219, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m555p203.pdf.

Ivory:2014:UNR

[ITT14] J. A. Ivory, K. W. Tang, and K. Takahashi. Use of neutral red in short-term sediment traps to distinguish between zooplankton swimmers and carcasses. Marine Ecology Progress Series, 505:107–117, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2014/505/m505p107.pdf.

Ingels:2010:PUB

[IVD^{+}10] J. Ingels, P. Van den Driessche, I. De Mesel, S. Vanhove, T. Moens, and A. Vanreusel. Preferred use of bacteria over phytoplankton by deep-sea nematodes in polar regions. Marine Ecology Progress Series, 406:121–133, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/406/
m406p121.pdf.

Ingeman:2015:NPM

[IW15] K. E. Ingeman and M. S. Webster. Native prey mortality increases but remains density-dependent following lionfish invasion. Marine Ecology Progress Series, 531:241–252, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2015/531/m531p241.pdf.

Johnson:2016:ZMR

[JAPK16] M. G. Johnson, C. P. Andersen, D. L. Phillips, and J. E. Kaldy. Zostera marina root demography in an intertidal estuarine environment measured using minirhizotron technology. Marine Ecology Progress Series, 557:123–132, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/557/
m557p123.pdf.

Jezequel:2019:RBE

[JBCGC19] Y. Jézéquel, J. Bonnel, J. Coston-Guarini, and L. Chauvaud. Revisiting the bioacoustics of European spiny lobsters Palinurus elephas: comparison of antennal rasps in tanks and in situ. Marine Ecology Progress Series, 615:143–157, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/615/
m615p143.pdf.

Johansson:2010:SUM

[JBD10] C. L. Johansson, D. R. Bellwood, and M. Depczynski. Sea urchins, macroalgae and coral reef decline: a functional evaluation of an intact reef system, Ningaloo, Western Australia. Marine Ecology Progress Series, 414:65–74, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/414/
m414p065.pdf.

Jensen:2016:STG

[JBL^{+}16] M. P. Jensen, I. Bell, C. J. Limpus, M. Hamann, S. Ambar, T. Whap, C. David, and N. N. FitzSimmons. Spatial and temporal genetic variation among size classes of green turtles (Chelonia mydas) provides information on oceanic dispersal and population dynamics. Marine Ecology Progress Series, 543:241–256, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/543/m543p241.pdf.

Johansen:2013:HPV

[JBM^{+}13] R. Johansen, Ø. Bergh, I. Modahl, G. Dahle, B. Gjerset, J. C. Holst, and N. Sandlund. High prevalence of viral haemorrhagic septicaemia virus (VHSV) in Norwegian springspawning herring. Marine Ecology Progress Series, 478:223– 230, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/478/m478p223.pdf.

Jaschinski:2011:SVC

[JBS11] S. Jaschinski, D. C. Brepohl, and U. Sommer. Seasonal variation in carbon sources of mesograzers and small predators in an eelgrass community: stable isotope and fatty acid analyses. Marine Ecology Progress Series, 431:69–82, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/431/
m431p069.pdf.

Jonsdottir:2012:PAC

[JBS12] I. G. Jónsdóttir, H. Björnsson, and U. Skúladóttir. Predation by Atlantic cod Gadus morhua on northern shrimp Pandalus borealis in inshore and offshore areas of Iceland. Marine Ecology Progress Series, 469:223–232, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m469p223.pdf.

Janousek:2016:PES

[JBT^{+}16] C. N. Janousek, K. J. Buffington, K. M. Thorne, G. R. Guntenspergen, J. Y. Takekawa, and B. D. Dugger. Potential effects of sea-level rise on plant productivity: species-specific responses in northeast Pacific tidal marshes. Marine Ecology Progress Series, 548:111–125, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m548p111.pdf.

Joseph:2015:RBI

[JC15] L. Joseph and M. Cusson. Resistance of benthic intertidal communities to multiple disturbances and stresses. Marine Ecology Progress Series, 534:49–64, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/534/m534p049.pdf.

Johnston:2017:MET

[JC17] C. A. Johnston and O. N. Caretti. Mangrove expansion into temperate marshes alters habitat quality for recruiting Callinectes spp. Marine Ecology Progress Series, 573:1–14, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Janssen:2013:MCF

[JCA13] A. Janssen, P. Chevaldonné, and P. Martínez Arbizu. Meiobenthic copepod fauna of a marine cave (NW Mediterranean) closely resembles that of deep-sea communities. Marine Ecology Progress Series, 479:99–113, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/479/
m479p099.pdf.

Johansson:2012:MSC

[JCF^{+}12] M. L. Johansson, K. Clifford, B. Fodness, N. A. Vazquez, and M. A. Banks. Mate selection in captive-breeding rockfishes Sebastes spp.: inference from parentage analysis and the major histocompatibility complex (MHC). Marine Ecology Progress Series, 460:195–206, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/460/m460p195.pdf.

Jenkins:2010:HVR

[JCM10] G. P. Jenkins, S. D. Conron, and A. K. Morison. Highly variable recruitment in an estuarine fish is determined by salinity stratification and freshwater flow: implications of a changing climate. Marine Ecology Progress Series, 417:249– 261, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/417/m417p249.pdf.

Juanes:2013:FWE

[JCM13] F. Juanes, P. J. Clarke, and J. Murt. Fall and winter estuarine recruitment of bluefish Pomatomus saltatrix: selectivity for large lipid-rich prey increases depleted energy levels. Marine Ecology Progress Series, 492:235–252, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/492/
m492p235.pdf.

Jaeger:2010:USI

[JCRC10] A. Jaeger, M. Connan, P. Richard, and Y. Cherel. Use of stable isotopes to quantify seasonal changes of trophic niche and levels of population and individual specialisation in seabirds. Marine Ecology Progress Series, 401:269–277, 2010. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m401p269.pdf.

Jiang:2012:PCD

[JCZ^{+}12] Z. B. Jiang, Q. Z. Chen, J. N. Zeng, Y. B. Liao, L. Shou, and J. Liu. Phytoplankton community distribution in relation to environmental parameters in three aquaculture systems in a Chinese subtropical eutrophic bay. Marine Ecology Progress Series, 446:73–89, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/446/m446p073.pdf.

Johnson:2018:LSS

[JD18] D. W. Johnson and A. Dunk. A large-scale study of competition of two temperate reef fishes: temperature, functional diversity, and regional differences in dynamics. Marine Ecology Progress Series, 593:97–109, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/593/m593p097.pdf.

Jimenez:2010:TRN

[JDB^{+}10] H. Jimenez, P. Dumas, L. Bigot, J. M. Amouroux, and J. Ferraris. Taxonomic resolution needed to describe invertebrate assemblages and to detect harvesting effects on coral reef ecosystems. Marine Ecology Progress Series, 406:211– 222, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/406/m406p211.pdf.

Jossart:2013:NEH

[JDD^{+}13] Q. Jossart, B. David, C. De Bruyn, C. De Ridder, T. Rigaud, and R. A. Wattier. No evidence of host specialization in a parasitic pea-crab exploiting two echinoid hosts. Marine Ecology Progress Series, 475:167–176, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/475/m475p167.pdf.

Jung:2017:LTS

[JDG^{+}17] A. S. Jung, R. Dekker, M. Germain, C. J. M. Philippart, J. I. J. Witte, and H. W. van der Veer. Long-term shifts in intertidal predator and prey communities in the Wadden Sea and consequences for food requirements and supply. Marine Ecology Progress Series, 579:37–53, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/579/m579p037.pdf.

Jacob:2011:CMP

[JDQS11] B. Jacob, G. Daneri, R. A. Quiñones, and M. Sobarzo. Community metabolism, phytoplankton size structure and heterotrophic prokaryote production in a highly productive upwelling zone off northern Chile. Marine Ecology Progress Series, 430:23–34, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m430p023.pdf.

Jin:2017:HLS

[JDX^{+}17] P. Jin, J. Ding, T. Xing, U. Riebesell, and K. Gao. High levels of solar radiation offset impacts of ocean acidification on calcifying and non-calcifying strains of Emiliania huxleyi. Marine Ecology Progress Series, 568:47–58, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m568p047.pdf.

Johnson:2010:PDS

[JE10] E. G. Johnson and D. B. Eggleston. Population density, survival and movement of blue crabs in estuarine salt marsh nurseries. Marine Ecology Progress Series, 407:135– 147, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/407/m407p135.pdf.

Jay:2012:WAU

[JFK12] C. V. Jay, A. S. Fischbach, and A. A. Kochnev. Walrus areas of use in the Chukchi Sea during sparse sea ice cover. Marine Ecology Progress Series, 468:1–13, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Jenewein:2013:OSS

[JG13] B. T. Jenewein and L. A. Gosselin. Ontogenetic shift in stress tolerance thresholds of Mytilus trossulus effects of desiccation and heat on juvenile mortality. Marine Ecology Progress Series, 481:147–159, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/481/m481p147.pdf.

Johnson:2012:MSM

[JGCRS12] D. W. Johnson, K. Grorud-Colvert, T. L. Rankin, and S. Sponaugle. Measuring selective mortality from otoliths and similar structures: a practical guide for describing multivariate selection from cross-sectional data. Marine Ecology Progress Series, 471:151–163, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/471/m471p151.pdf.

Jones:2015:TBC

[JGG^{+}15] B. T. Jones, J. Gyory, E. K. Grey, M. Bartlein, D. S. Ko, R. W. Nero, and C. M. Taylor. Transport of blue crab larvae in the northern Gulf of Mexico during the Deepwater Horizon oil spill. Marine Ecology Progress Series, 527: 143–156, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/527/m527p143.pdf.

Jankowski:2015:DGD

[JGJ15] M. W. Jankowski, N. A. J. Graham, and G. P. Jones. Depth gradients in diversity, distribution and habitat specialisation in coral reef fishes: implications for the depth–refuge hypothesis. Marine Ecology Progress Series, 540:203–215, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/540/
m540p203.pdf.

Jones:2012:RDW

[JGL12] D. O. B. Jones, A. R. Gates, and B. Lausen. Recovery of deep-water megafaunal assemblages from hydrocarbon drilling disturbance in the Faroe–Shetland Channel. Marine Ecology Progress Series, 461:71–82, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m461p071.pdf.

Johnson:2016:IGR

[JGM^{+}16] R. C. Johnson, J. C. Garza, R. B. MacFarlane, C. B. Grimes, C. C. Phillis, P. L. Koch, P. K. Weber, and M. H. Carr. Isotopes and genes reveal freshwater origins of Chinook salmon Oncorhynchus tshawytscha aggregations in California’s coastal ocean. Marine Ecology Progress Series, 548:181–196, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m548p181.pdf.

Johnson:2014:ODT

[JGS14] K. D. Johnson, J. H. Grabowski, and D. L. Smee. Omnivory dampens trophic cascades in estuarine communities. Marine Ecology Progress Series, 507:197–206, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/507/
m507p197.pdf.

Jakubas:2011:FED

[JGWJ^{+}11] D. Jakubas, M. Głuchowska, K. Wojczulanis-Jakubas, N. J. Karnovsky, L. Keslinka, D. Kidawa, W. Walkusz, R. Boehnke, M. Cisek, S. Kwaśniewski, and L. Stempniewicz. Foraging effort does not influence body condition and stress level in little auks. Marine Ecology Progress Series, 432:277–290, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m432p277.pdf.

Jennings:2011:SMI

[JH11] L. B. Jennings and H. L. Hunt. Small macrobenthic invertebrates affect the mortality and growth of early post-settlement sea urchins and sea stars in subtidal cobble habitat. Marine Ecology Progress Series, 431:173–182, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/431/
m431p173.pdf.

Jennings:2014:SPE

[JH14] L. B. Jennings and H. L. Hunt. Spatial patterns in early post-settlement processes of the green sea urchin Strongylocentrotus droebachiensis. Marine Ecology Progress Series, 502:219–228, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/502/m502p219.pdf.

Johnson:2019:TIP

[JHL^{+}19] A. C. Johnson, K. A. Hobson, N. J. Lunn, D. McGeachy, E. S. Richardson, and A. E. Derocher. Temporal and intrapopulation patterns in polar bear foraging ecology in western Hudson Bay. Marine Ecology Progress Series, 619:187– 199, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m619p187.pdf.

Januchowski-Hartley:2012:ISG

[JHNL12] F. A. Januchowski-Hartley, K. L. Nash, and R. J. Lawton. Influence of spear guns, dive gear and observers on estimating fish flight initiation distance on coral reefs. Marine Ecology Progress Series, 469:113–119, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/469/m469p113.pdf.

Ji:2011:CFD

[Ji11] R. Ji. Calanus finmarchicus diapause initiation: new view from traditional life history-based model. Marine Ecology Progress Series, 440:105–114, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/440/m440p105.pdf.

Jonsson:2014:LET

[JJF14] B. Jonsson, N. Jonsson, and A. G. Finstad. Linking embryonic temperature with adult reproductive investment in Atlantic salmon Salmo salar. Marine Ecology Progress Series, 515:217–226, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m515p217.pdf.

Jonsson:2016:OSS

[JJJ16] B. Jonsson, M. Jonsson, and N. Jonsson. Optimal size at seaward migration in an anadromous salmonid. Marine Ecology Progress Series, 559:193–200, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/559/m559p193.pdf.

Jones:2012:STP

[JJNP12] R. Jones, R. Johnson, T. Noyes, and R. Parsons. Spatial and temporal patterns of coral black band disease in relation to a major sewage outfall. Marine Ecology Progress Series, 462:79–92, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/462/m462p079.pdf.

Jaeger:2013:SID

[JJP^{+}13] A. Jaeger, S. Jaquemet, R. A. Phillips, R. M. Wanless, P. Richard, and Y. Cherel. Stable isotopes document interand intra-specific variation in feeding ecology of nine large southern procellariiformes. Marine Ecology Progress Series, 490:255–266, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m490p255.pdf.

Juhl:2011:SDD

[JKM11] A. R. Juhl, C. Krembs, and K. M. Meiners. Seasonal development and differential retention of ice algae and other organic fractions in first-year Arctic sea ice. Marine Ecology Progress Series, 436:1–16, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Johnston:2012:EMG

[JL12] C. A. Johnston and R. N. Lipcius. Exotic macroalga Gracilaria vermiculophylla provides superior nursery habitat for native blue crab in Chesapeake Bay. Marine Ecology Progress Series, 467:137–146, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m467p137.pdf.

Jones:2017:RSH

[JL17] E. Jones and J. Long. The relative strength of an herbivoreinduced seaweed defense varies with herbivore species. Marine Ecology Progress Series, 581:33–44, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/581/
m581p033.pdf.

Jeffreys:2012:LES

[JLL^{+}12] R. M. Jeffreys, L. A. Levin, P. A. Lamont, C. Woulds, C. R. Whitcraft, G. F. Mendoza, G. A. Wolff, and G. L. Cowie. Living on the edge: single-species dominance at the Pakistan oxygen minimum zone boundary. Marine Ecology Progress Series, 470:79–99, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/470/m470p079.pdf.

Johannesen:2012:FHE

[JLM^{+}12] E. Johannesen, U. Lindstrøm, K. Michalsen, M. Skern-Mauritzen, P. Fauchald, B. Bogstad, and A. Dolgov. Feeding in a heterogeneous environment: spatial dynamics in summer foraging Barents Sea cod. Marine Ecology Progress Series, 458:181–197, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m458p181.pdf.

Joumaa:2017:TDA

[JLPG17] J. Jouma’a, Y. Le Bras, B. Picard, and C. Guinet. Threedimensional assessment of hunting strategies in a deep diving predator, southern elephant seal Mirounga leonina. Marine Ecology Progress Series, 573:255–268, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/573/
m573p255.pdf.

Jovanovic:2014:ODP

[JLQ^{+}14] Z. Jovanovic, M. Larsen, C. Organo Quintana, E. Kristensen, and R. N. Glud. Oxygen dynamics and porewater transport in sediments inhabited by the invasive polychaete Marenzelleria viridis. Marine Ecology Progress Series, 504:181–192, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m504p181.pdf.

Jones:2017:CSC

[JM17] D. O. B. Jones and J. W. Murray. Controls on the standing crop of benthic Foraminifera at an oceanic scale. Marine Ecology Progress Series, 581:71–83, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m581p071.pdf.

Juinio-Menez:2010:EIF

[JMB10] M. A. Juinio-Meñez and H. G. P. Bangi. Extrinsic and intrinsic factors affecting the metamorphic rate of Tripneustes gratilla (Echinodermata: Echinoidea). Marine Ecology Progress Series, 402:137–145, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/402/m402p137.pdf.

James:2017:CDR

[JMB^{+}17] L. C. James, M. P. Marzloff, N. Barrett, A. Friedman, and C. R. Johnson. Changes in deep reef benthic community composition across a latitudinal and environmental gradient in temperate Eastern Australia. Marine Ecology Progress Series, 565:35–52, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m565p035.pdf.

Jungblut:2018:SLS

[JMB^{+}18] S. Jungblut, M. L. McCarthy, K. Boos, R. Saborowski, and W. Hagen. Seasonal lipid storage and dietary preferences of native European versus invasive Asian shore crabs. Marine Ecology Progress Series, 602:169–181, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/602/
m602p169.pdf.

Jimenez-Melero:2013:SPA

[JMGG13] R. Jiménez-Melero, J. D. Gilbert, and F. Guerrero. Secondary production of Arctodiaptomus salinus in a shallow saline pond: comparison of methods. Marine Ecology Progress Series, 483:103–116, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/483/m483p103.pdf.

Jarvis:2012:CEI

[JMK12] J. C. Jarvis, K. A. Moore, and W. J. Kenworthy. Characterization and ecological implication of eelgrass life history strategies near the species’ southern limit in the western North Atlantic. Marine Ecology Progress Series, 444:43–56, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/444/
m444p043.pdf.

Jones:2018:IFC

[JMK^{+}18] K. R. Jones, J. M. Maina, S. Kark, T. R. McClanahan, C. J. Klein, and M. Beger. Incorporating feasibility and collaboration into large-scale planning for regional recovery of coral reef fisheries. Marine Ecology Progress Series, 604:211–222, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m604p211.pdf.

Jensen:2019:SED

[JMO^{+}19] O. P. Jensen, C. W. Martin, K. L. Oken, F. J. Fodrie, P. C. López-Duarte, K. W. Able, and B. J. Roberts. Simultaneous estimation of dispersal and survival of the gulf killifish Fundulus grandis from a batch-tagging experiment. Marine Ecology Progress Series, 624:183–194, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/624/m624p183.pdf.

Juttner:2010:OCD

[JMPZ10] F. Jüttner, P. Messina, C. Patalano, and V. Zupo. Odour compounds of the diatom Cocconeis scutellum: effects on benthic herbivores living on Posidonia oceanica. Marine Ecology Progress Series, 400:63–73, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/400/m400p063.pdf.

Jones:2015:PSU

[JMS^{+}15] E. L. Jones, B. J. McConnell, S. Smout, P. S. Hammond, C. D. Duck, C. D. Morris, D. Thompson, D. J. F. Russell, C. Vincent, M. Cronin, R. J. Sharples, and J. Matthiopoulos. Patterns of space use in sympatric marine colonial predators reveal scales of spatial partitioning. Marine Ecology Progress Series, 534:235–249, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/534/m534p235.pdf.

Josefson:2012:BDP

[JNN12] A. B. Josefson, J. Norkko, and A. Norkko. Burial and decomposition of plant pigments in surface sediments of the Baltic Sea: role of oxygen and benthic fauna. Marine Ecology Progress Series, 455:33–49, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/455/m455p033.pdf.

Joh:2013:IGD

[JNYT13] M. Joh, M. Nakaya, N. Yoshida, and T. Takatsu. Interannual growth differences and growth-selective survival in larvae and juveniles of marbled sole Pseudopleuronectes yokohamae. Marine Ecology Progress Series, 494:267–279, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/494/
m494p267.pdf.

Janiak:2018:ASV

[JOFP18] D. S. Janiak, R. W. Osman, C. J. Freeman, and V. J. Paul. Artificial structures versus mangrove prop roots: a general comparison of epifaunal communities within the Indian River Lagoon, Florida, USA. Marine Ecology Progress Series, 607:85–98, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/607/m607p085.pdf.

Johnson:2010:RBC

[Joh10] A. E. Johnson. Reducing bycatch in coral reef trap fisheries: escape gaps as a step towards sustainability. Marine Ecology Progress Series, 415:201–209, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/415/m415p201.pdf.

Johnson:2011:HMI

[Joh11] D. S. Johnson. High-marsh invertebrates are susceptible to eutrophication. Marine Ecology Progress Series, 438:143– 152, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/438/m438p143.pdf.

Jones:2011:CMM

[Jon11] I. S. F. Jones. Contrasting micro- and macro-nutrient nourishment of the ocean. Marine Ecology Progress Series, 425: 281–296, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m425p281.pdf.

Johnston:2015:CSI

[JP15] M. W. Johnston and S. J. Purkis. A coordinated and sustained international strategy is required to turn the tide on the Atlantic lionfish invasion. Marine Ecology Progress Series, 533:219–235, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/533/m533p219.pdf.

Juul-Pedersen:2015:SIP

[JPAM^{+}15] T. Juul-Pedersen, K. E. Arendt, J. Mortensen, M. E. Blicher, D. H. Søgaard, and S. Rysgaard. Seasonal and interannual phytoplankton production in a sub-Arctic tidewater outlet glacier fjord, SW Greenland. Marine Ecology Progress Series, 524:27–38, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m524p027.pdf.

Jost:2012:ETT

[JPF12] J. A. Jost, S. M. Podolski, and M. Frederich. Enhancing thermal tolerance by eliminating the pejus range: a comparative study with three decapod crustaceans. Marine Ecology Progress Series, 444:263–274, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/444/m444p263.pdf.

Jorgensen:2019:IMS

[JPI^{+}19] L. L. Jørgensen, R. Primicerio, R. B. Ingvaldsen, M. Fossheim, N. Strelkova, T. H. Thangstad, I. Manushin, and D. Zakharov. Impact of multiple stressors on sea bed fauna in a warming Arctic. Marine Ecology Progress Series, 608:1–12, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Jovanovic:2015:RDY

[JPL^{+}15] Z. Jovanovic, M. Ø. Pedersen, M. Larsen, E. Kristensen, and R. N. Glud. Rhizosphere O 2 dynamics in young Zostera marina and Ruppia maritima. Marine Ecology Progress Series, 518:95–105, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m518p095.pdf.

Juul-Pedersen:2010:SEP

[JPMG10] T. Juul-Pedersen, C. Michel, and M. Gosselin. Sinking export of particulate organic material from the euphotic zone in the eastern Beaufort Sea. Marine Ecology Progress Series, 410:55–70, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m410p055.pdf.

Jones:2017:MMM

[JPP^{+}17] T. Jones, J. K. Parrish, A. E. Punt, V. L. Trainer, R. Kudela, J. Lang, M. S. Brancato, A. Odell, and B. Hickey. Mass mortality of marine birds in the Northeast Pacific caused by Akashiwo sanguinea. Marine Ecology Progress Series, 579:111–127, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m579p111.pdf.

Jager:2015:PGM

[JR15] T. Jager and E. Ravagnan. Parameterising a generic model for the dynamic energy budget of Antarctic krill Euphausia superba. Marine Ecology Progress Series, 519:115–128, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2014/519/m519p115.pdf.

Jimenez-Ramos:2018:NLE

[JREVB18] R. Jiménez-Ramos, L. G. Egea, J. J. Vergara, and F. G. Brun. Nutrient load and epiphytes are drivers of increased herbivory in seagrass communities. Marine Ecology Progress Series, 599:49–64, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/599/m599p049.pdf.

Jokiel:2013:CGV

[JRK13] P. L. Jokiel, K. S. Rodgers, and S. A. Karl. Comparison of grafting versus a molecular genetic technique in identifying clone mates of the reef coral Porites rus. Marine Ecology Progress Series, 477:87–92, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/477/m477p087.pdf.

Johnsen:2017:CSA

[JROS17] E. Johnsen, G. Rieucau, E. Ona, and G. Skaret. Collective structures anchor massive schools of lesser sandeel to the seabed, increasing vulnerability to fishery. Marine Ecology Progress Series, 573:229–236, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/573/m573p229.pdf.

Jacobs:2015:IBM

[JRvdM15] P. Jacobs, R. Riegman, and J. van der Meer. Impact of the blue mussel Mytilus edulis on the microbial food web in the western Wadden Sea, the Netherlands. Marine Ecology Progress Series, 527:119–131, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/527/m527p119.pdf.

Jaine:2014:MHU

[JRW^{+}14] F. R. A. Jaine, C. A. Rohner, S. J. Weeks, L. I. E. Couturier, M. B. Bennett, K. A. Townsend, and A. J. Richardson. Movements and habitat use of reef manta rays off eastern Australia: offshore excursions, deep diving and eddy affinity revealed by satellite telemetry. Marine Ecology Progress Series, 510:73–86, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m510p073.pdf.

Jaschinski:2010:PEM

[JS10] S. Jaschinski and U. Sommer. Positive effects of mesograzers on epiphytes in an eelgrass system. Marine Ecology Progress Series, 401:77–85, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/401/m401p077.pdf.

Johnson:2012:SMR

[JS12a] K. D. Johnson and D. L. Smee. Size matters for risk assessment and resource allocation in bivalves. Marine Ecology Progress Series, 462:103–110, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/462/m462p103.pdf.

Johnson:2012:FDF

[JS12b] S. P. Johnson and D. E. Schindler. Four decades of foraging history: stock-specific variation in the carbon and nitrogen stable isotope signatures of Alaskan sockeye salmon. Marine Ecology Progress Series, 460:155–167, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m460p155.pdf.

Jahnke:2015:IGM

[JSBP15] M. Jahnke, I. A. Serra, G. Bernard, and G. Procaccini. The importance of genetic make-up in seagrass restoration: a case study of the seagrass Zostera noltei. Marine Ecology Progress Series, 532:111–122, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2015/532/m532p111.pdf.

Jenkins:2015:DIS

[JSCM15] G. P. Jenkins, D. Spooner, S. Conron, and J. R. Morrongiello. Differing importance of salinity stratification and freshwater flow for the recruitment of apex species of estuarine fish. Marine Ecology Progress Series, 523:125–144, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/523/
m523p125.pdf.

Jusufovski:2019:CBM

[JSK19] D. Jusufovski, C. Saavedra, and A. Kuparinen. Competition between marine mammals and fisheries in contemporary harvested marine ecosystems. Marine Ecology Progress Series, 627:207–232, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/627/m627p207.pdf. See corrigen-

dum [Ano21].

Jungbluth:2017:SSG

[JSLG17] M. J. Jungbluth, K. E. Selph, P. H. Lenz, and E. Goetze. Species-specific grazing and significant trophic impacts by two species of copepod nauplii, Parvocalanus crassirostris and Bestiolina similis. Marine Ecology Progress Series, 572: 57–76, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/572/m572p057.pdf.

Jordan:2014:SAI

[JSMO14] F. Jordán, M. Scotti, Á. Mike, and M. Ortiz. Strong asymmetrical inter-specific relationships in food web simulations. Marine Ecology Progress Series, 512:89–98, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m512p089.pdf.

Johnson:2018:SBB

[JSOK18] A. J. Johnson, J. Statton, R. J. Orth, and G. A. Kendrick. A sediment bioturbator bottleneck to seedling recruitment for the seagrass Posidonia australis. Marine Ecology Progress Series, 595:89–103, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/595/m595p089.pdf.

Jansen:2011:SVN

[JSS^{+}11] H. M. Jansen, Ø. Strand, T. Strohmeier, C. Krogness, M. Verdegem, and A. Smaal. Seasonal variability in nutrient regeneration by mussel Mytilus edulis rope culture in oligotrophic systems. Marine Ecology Progress Series, 431:

137–149, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/431/m431p137.pdf.

Jellyman:2010:VMM

[JT10a] D. Jellyman and K. Tsukamoto. Vertical migrations may control maturation in migrating female Anguilla dieffenbachii. Marine Ecology Progress Series, 404:241–247, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/404/
m404p241.pdf.

Jones:2010:EHM

[JT10b] E. Jones and C. S. Thornber. Effects of habitat-modifying invasive macroalgae on epiphytic algal communities. Marine Ecology Progress Series, 400:87–100, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/400/
m400p087.pdf.

Jewell:2012:GAC

[JTH^{+}12] R. Jewell, L. Thomas, C. M. Harris, K. Kaschner, R. Wiff, P. S. Hammond, and N. J. Quick. Global analysis of cetacean line-transect surveys: detecting trends in cetacean density. Marine Ecology Progress Series, 453:227–240, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m453p227.pdf.

Jakobsson-Thor:2018:SWD

[JTTB^{+}18] S. Jakobsson-Thor, G. B. Toth, J. Brakel, A. C. Bockelmann, and H. Pavia. Seagrass wasting disease varies with salinity and depth in natural Zostera marina populations. Marine Ecology Progress Series, 587:105–115, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m587p105.pdf.

Jakobsson-Thor:2019:SWD

[JTTP19] S. Jakobsson-Thor, G. B. Toth, and H. Pavia. Seagrass wasting disease along a naturally occurring salinity gradient. Marine Ecology Progress Series, 614:67–77, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m614p067.pdf.

Jakubas:2013:FCC

[JTWJ^{+}13] D. Jakubas, E. Trudnowska, K. Wojczulanis-Jakubas, L. Iliszko, D. Kidawa, M. Darecki, K. Błachowiak-Samołyk, and L. Stempniewicz. Foraging closer to the colony leads to faster growth in little auks. Marine Ecology Progress Series, 489:263–278, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/489/m489p263.pdf.

Jenkins:2016:TSF

[JU16] S. R. Jenkins and M. Uyà. Temporal scale of field experiments in benthic ecology. Marine Ecology Progress Series, 547:273–286, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/547/m547p273.pdf.

Jay:2010:DMW

[JUK^{+}10] C. V. Jay, M. S. Udevitz, R. Kwok, A. S. Fischbach, and D. C. Douglas. Divergent movements of walrus and sea ice in the northern Bering Sea. Marine Ecology Progress Series, 407:293–302, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/407/m407p293.pdf.

Jackson:2010:LEP

[JUMS10] A. C. Jackson, A. J. Underwood, R. J. Murphy, and G. A. Skilleter. Latitudinal and environmental patterns in abundance and composition of epilithic microphytobenthos. Marine Ecology Progress Series, 417:27–38, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m417p027.pdf.

Johansen:2010:KEB

[JVSD10] J. L. Johansen, R. Vaknin, J. F. Steffensen, and P. Domenici. Kinematics and energetic benefits of schooling in the labriform fish, striped surfperch Embiotoca lateralis. Marine Ecology Progress Series, 420:221–229, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/420/m420p221.pdf.

Jack:2010:MOG

[JW10] L. Jack and S. R. Wing. Maintenance of old-growth size structure and fecundity of the red rock lobster Jasus edwardsii among marine protected areas in Fiordland, New Zealand. Marine Ecology Progress Series, 404:161–172, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/404/
m404p161.pdf.

Jack:2011:IVT

[JW11] L. Jack and S. R. Wing. Individual variability in trophic position and diet of a marine omnivore is linked to kelp bed habitat. Marine Ecology Progress Series, 443:129–139, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2011/443/m443p129.pdf.

Johns:2017:ATV

[JWB^{+}17] M. E. Johns, P. Warzybok, R. W. Bradley, J. Jahncke, M. Lindberg, and G. A. Breed. Age, timing, and a variable environment affect double brooding of a long-lived seabird. Marine Ecology Progress Series, 564:187–197, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/564/
m564p187.pdf.

Jones:2010:CTO

[JWBS10] D. L. Jones, J. F. Walter, E. N. Brooks, and J. E. Serafy. Connectivity through ontogeny: fish population linkages among mangrove and coral reef habitats. Marine Ecology Progress Series, 401:245–258, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m401p245.pdf.

Jack:2011:NTE

[JWHR11] L. Jack, S. R. Wing, Y. Hu, and M. Roberts. Natural trace elemental markers for adult red rock lobsters Jasus edwardsii vary among replicate distinct water masses. Marine Ecology Progress Series, 443:141–151, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/443/m443p141.pdf.

Jantzen:2010:EPW

[JWR^{+}10] C. Jantzen, C. Wild, M. Rasheed, M. El-Zibdah, and C. Richter. Enhanced pore-water nutrient fluxes by the upsidedown jellyfish Cassiopea sp. in a Red Sea coral reef. Marine Ecology Progress Series, 411:117–125, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/411/
m411p117.pdf.

Jeglinski:2012:ABM

[JWR^{+}12] J. W. E. Jeglinski, C. Werner, P. W. Robinson, D. P. Costa, and F. Trillmich. Age, body mass and environmental variation shape the foraging ontogeny of Galapagos sea lions. Marine Ecology Progress Series, 453:279–296, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m453p279.pdf.

Jewell:2014:CHU

[JWT^{+}14] O. J. D. Jewell, M. A. Wcisel, A. V. Towner, W. Chivell, L. van der Merwe, and M. N. Bester. Core habitat use of an apex predator in a complex marine landscape. Marine Ecology Progress Series, 506:231–242, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m506p231.pdf.

Kindinger:2016:NPI

[KA16] T. L. Kindinger and E. R. Anderson. NOTE: Preferences of invasive lionfish and native grouper between congeneric prey fishes. Marine Ecology Progress Series, 558:247–253, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m558p247.pdf.

Kaminsky:2015:BPF

[KAAI15] J. Kaminsky, J. Alberti, M. Aguiar, and O. Iribarne. Biological and physical factors affecting the colonization of vegetationfree patches in a SW Atlantic salt marsh. Marine Ecology Progress Series, 531:33–41, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/531/m531p033.pdf.

Kavanaugh:2015:ECS

[KAD^{+}15] M. T. Kavanaugh, F. N. Abdala, H. Ducklow, D. Glover, W. Fraser, D. Martinson, S. Stammerjohn, O. Schofield, and S. C. Doney. Effect of continental shelf canyons on phytoplankton biomass and community composition along the western Antarctic Peninsula. Marine Ecology Progress Series, 524:11–26, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m524p011.pdf.

Krawczyk:2015:STD

[KAJP^{+}15] D. W. Krawczyk, K. E. Arendt, T. Juul-Pedersen, M. K. Sejr, M. E. Blicher, and H. H. Jakobsen. Spatial and temporal distribution of planktonic protists in the East Greenland fjord and offshore waters. Marine Ecology Progress Series, 538:99–116, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m538p099.pdf.

Kelkar:2013:GTH

[KAMA13] N. Kelkar, R. Arthur, N. Marba, and T. Alcoverro. Green turtle herbivory dominates the fate of seagrass primary production in the Lakshadweep islands (Indian Ocean). Marine Ecology Progress Series, 485:235–243, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/485/
m485p235.pdf.

Katsikatsou:2012:FSP

[KAP^{+}12] M. Katsikatsou, A. Anestis, H. O. Pörtner, A. Vratsistas, K. Aligizaki, and B. Michaelidis. Field studies and projections of climate change effects on the bearded horse mussel Modiolus barbatus in the Gulf of Thermaikos, Greece. Marine Ecology Progress Series, 449:183–196, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/449/m449p183.pdf.

Kimmance:2014:RPI

[KAP^{+}14] S. A. Kimmance, M. J. Allen, A. Pagarete, J. Martínez Martínez, and W. H. Wilson. Reduction in photosystem II efficiency during a virus-controlled Emiliania huxleyi bloom. Marine Ecology Progress Series, 495:65–76, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/495/
m495p065.pdf.

Katsanevakis:2016:TCA

[Kat16] S. Katsanevakis. Transplantation as a conservation action to protect the Mediterranean fan mussel Pinna nobilis. Marine Ecology Progress Series, 546:113–122, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/546/
m546p113.pdf.

Kuklinski:2010:FBB

[KB10] P. Kuklinski and D. K. A. Barnes. First bipolar benthic brooder. Marine Ecology Progress Series, 401:15–20, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2009/401/m401p015.pdf.

Karnauskas:2012:CBA

[KB12] M. Karnauskas and E. A. Babcock. Comparisons between abundance estimates from underwater visual census and catchper-unit-effort in a patch reef system. Marine Ecology Progress Series, 468:217–230, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/468/m468p217.pdf.

Kintzing:2014:EPU

[KB14] M. D. Kintzing and M. J. Butler IV. Effects of predation upon the long-spined sea urchin Diadema antillarum by the spotted spiny lobster Panulirus guttatus. Marine Ecology Progress Series, 495:185–191, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/495/m495p185.pdf.

Kaldy:2013:LST

[KBA13] J. E. Kaldy, C. A. Brown, and C. P. Andersen. In situ ^{13}C tracer experiments elucidate carbon translocation rates and allocation patterns in eelgrass Zostera marina. Marine Ecology Progress Series, 487:27–39, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/487/m487p027.pdf.

Knox:2017:HUD

[KBA17] T. C. Knox, A. M. M. Baylis, and J. P. Y. Arnould. Habitat use and diving behaviour of male Australian fur seals. Marine Ecology Progress Series, 566:243–256, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/566/
m566p243.pdf.

Kovach:2010:FSS

[KBB^{+}10a] A. I. Kovach, T. S. Breton, D. L. Berlinsky, L. Maceda, and I. Wirgin. Fine-scale spatial and temporal genetic structure of Atlantic cod off the Atlantic coast of the USA. Marine Ecology Progress Series, 410:177–195, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/410/m410p177.pdf.

Kruse:2010:RMC

[KBB10b] S. Kruse, T. Brey, and U. Bathmann. Role of midwater chaetognaths in Southern Ocean pelagic energy flow. Marine Ecology Progress Series, 416:105–113, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/416/
m416p105.pdf.

Kramer:2014:BCC

[KBB14] M. J. Kramer, D. R. Bellwood, and O. Bellwood. Benthic crustacea on coral reefs: a quantitative survey. Marine Ecology Progress Series, 511:105–116, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/511/m511p105.pdf.

Kramer:2016:FMU

[KBB16] M. J. Kramer, O. Bellwood, and D. R. Bellwood. Foraging and microhabitat use by crustacean-feeding wrasses on coral reefs. Marine Ecology Progress Series, 548:277–282, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/548/
m548p277.pdf.

Kulbicki:2012:DIP

[KBC^{+}12] M. Kulbicki, J. Beets, P. Chabanet, K. Cure, E. Darling, S. R. Floeter, R. Galzin, A. Green, M. Harmelin-Vivien, M. Hixon, Y. Letourneur, T. L. de Loma, T. McClanahan, J. McIlwain, G. MouTham, R. Myers, J. K. O’Leary, S. Planes, L. Vigliola, and L. Wantiez. Distributions of Indo–Pacific lionfishes Pterois spp. in their native ranges: implications for the Atlantic invasion. Marine Ecology Progress Series, 446:189–205, 2012.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m446p189.pdf.

Karpov:2012:MFC

[KBG12] K. A. Karpov, M. Bergen, and J. J. Geibel. Monitoring fish in California Channel Islands marine protected areas with a remotely operated vehicle: the first five years. Marine Ecology Progress Series, 453:159–172, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m453p159.pdf.

Kennington:2013:SSG

[KBG^{+}13] W. J. Kennington, O. Berry, D. M. Groth, M. S. Johnson, and R. Melville-Smith. Spatial scales of genetic patchiness in the western rock lobster Panulirus cygnus. Marine Ecology Progress Series, 486:213–221, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/486/m486p213.pdf.

Kristensen:2019:BVS

[KBGA19] M. L. Kristensen, K. Birnie-Gauvin, and K. Aarestrup. Behaviour of veteran sea trout Salmo trutta in a dangerous fjord system. Marine Ecology Progress Series, 616:141–153, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2019/616/m616p141.pdf.

Kaiser:2011:I

[KBH11a] M. J. Kaiser, M. T. Burrows, and H. Hughes. Introduction. Marine Ecology Progress Series, 430:99–101, 2011. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m430p099.pdf.

Kaiser:2011:TSE

[KBH11b] Michel J. Kaiser, Michael T. Burrows, and Helen Hughes. Theme section: Evolution and ecology of marine biodiversity: mechanisms and dynamics. Marine Ecology Progress Series, 430:99–288, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m430_ThemeSection.pdf.

Klimpel:2010:ASC

[KBK^{+}10] S. Klimpel, M. W. Busch, T. Kuhn, A. Rohde, and H. W. Palm. The Anisakis simplex complex off the South Shetland Islands (Antarctica): endemic populations versus introduction through migratory hosts. Marine Ecology Progress Series, 403: 1–11, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Klepac:2015:SSC

[KBK^{+}15] C. N. Klepac, J. Beal, C. D. Kenkel, A. Sproles, J. M. Polinski, M. A. Williams, M. V. Matz, and J. D. Voss. Seasonal stability of coral–Symbiodinium associations in the subtropical coral habitat of St. Lucie Reef, Florida. Marine Ecology Progress Series, 532:137–151, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2015/532/m532p137.pdf.

Kahng:2011:SRO

[KBL11] S. E. Kahng, Y. Benayahu, and H. R. Lasker. Sexual reproduction in octocorals. Marine Ecology Progress Series, 443:265–283, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/443/m443p265.pdf.

Koussoroplis:2010:NIM

[KBP^{+}10] A. M. Koussoroplis, A. Bec, M. E. Perga, E. Koutrakis, C. Desvilettes, and G. Bourdier. Nutritional importance of minor dietary sources for leaping grey mullet Liza saliens (Mugilidae) during settlement: insights from fatty acid δ 13C analysis. Marine Ecology Progress Series, 404:207–217, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2010/404/m404p207.pdf.

Kolesar:2010:EHM

[KBPD10] S. E. Kolesar, D. L. Breitburg, J. E. Purcell, and M. B. Decker. Effects of hypoxia on Mnemiopsis leidyi, ichthyoplankton and copepods: clearance rates and vertical habitat overlap. Marine Ecology Progress Series, 411:173–188, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/411/
m411p173.pdf.

Katija:2011:QFG

[KBR^{+}11] K. Katija, W. T. Beaulieu, C. Regula, S. P. Colin, J. H. Costello, and J. O. Dabiri. Quantification of flows generated by the hydromedusa Aequorea victoria: a Lagrangian coherent structure analysis. Marine Ecology Progress Series, 435:111–123, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/435/m435p111.pdf.

Kollars:2016:IDA

[KBS16] N. M. Kollars, J. E. Byers, and E. E. Sotka. Invasive décor: an association between a native decorator worm and a nonnative seaweed can be mutualistic. Marine Ecology Progress Series, 545:135–145, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/545/m545p135.pdf.

Karnovsky:2011:ICC

[KBW^{+}11] N. J. Karnovsky, Z. W. Brown, J. Welcker, A. M. A. Harding, W. Walkusz, A. Cavalcanti, J. Hardin, A. Kitaysky, G. Gabrielsen, and D. Grémillet. Inter-colony comparison of diving behavior of an Arctic top predator: implications for warming in the Greenland Sea. Marine Ecology Progress Series, 440:229–240, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m440p229.pdf.

Keppel:2015:ECV

[KBW^{+}15] A. G. Keppel, D. L. Breitburg, G. H. Wikfors, R. B. Burrell, and V. M. Clark. Effects of co-varying diel-cycling hypoxia and pH on disease susceptibility in the eastern oyster Crassostrea virginica. Marine Ecology Progress Series, 538:169–183, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/538/m538p169.pdf. See erratum

[Ano16].

Kelaher:2018:OWA

[KCB18] B. P. Kelaher, M. A. Coleman, and M. J. Bishop. Ocean warming, but not acidification, accelerates seagrass decomposition under near-future climate scenarios. Marine Ecology Progress Series, 605:103–110, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/605/m605p103.pdf.

Kerr:2012:RER

[KCCG12] K. A. Kerr, J. H. Christy, R. Collin, and F. Guichard. Reducing error in reproductive timing caused by temperature variation: interspecific differences in behavioural adjustment by fiddler crabs. Marine Ecology Progress Series, 459:1–16, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Koehl:2013:WCS

[KCD13] M. A. R. Koehl, J. P. Crimaldi, and D. E. Dombroski. Wind chop and ship wakes determine hydrodynamic stresses on larvae settling on different microhabitats in fouling communities. Marine Ecology Progress Series, 479:47–62, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/479/
m479p047.pdf.

Krumme:2014:IMF

[KCE14] U. Krumme, M. Audfroid Calderón, and A. Echterhoff. Intertidal migration of the four-eyed fish Anableps anableps in North Brazilian mangrove creeks. Marine Ecology Progress Series, 509:271–287, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/509/m509p271.pdf.

Kessel:2014:PTR

[KCF^{+}14] S. T. Kessel, D. D. Chapman, B. R. Franks, T. Gedamke, S. H. Gruber, J. M. Newman, E. R. White, and R. G. Perkins. Predictable temperature-regulated residency, movement and migration in a large, highly mobile marine predator (Negaprion brevirostris). Marine Ecology Progress Series, 514:175–190, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/514/m514p175.pdf.

Kerr:2014:PPR

[KCGC14] K. A. Kerr, A. Cornejo, F. Guichard, and R. Collin. Planktonic predation risk varies with prey life history stage and diurnal phase. Marine Ecology Progress Series, 503:99–109, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2014/503/m503p099.pdf.

Kohler:2011:GVT

[KCH^{+}11] S. A. Kohler, M. Connan, J. M. Hill, C. Mablouké, B. Bonnevie, K. Ludynia, J. Kemper, J. Huisamen, L. G. Underhill, Y. Cherel, C. D. McQuaid, and S. Jaquemet. Geographic variation in the trophic ecology of an avian rocky shore predator, the African black oystercatcher, along the southern African coastline. Marine Ecology Progress Series, 435:235–249, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/435/m435p235.pdf.

Krueger:2011:SGR

[KCL^{+}11] B. H. Krueger, M. Y. Chaloupka, P. A. Leighton, J. A. Dunn, and J. A. Horrocks. Somatic growth rates for a hawksbill turtle population in coral reef habitat around Barbados. Marine Ecology Progress Series, 432:269–276, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/432/
m432p269.pdf.

Kersting:2013:UPS

[KCLLL13] D. K. Kersting, C. Casado, S. López-Legentil, and C. Linares. Unexpected patterns in the sexual reproduction of the Mediterranean scleractinian coral Cladocora caespitosa. Marine Ecology Progress Series, 486:165–171, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/486/
m486p165.pdf.

Kough:2016:DAL

[KCLP16] A. S. Kough, R. Claro, K. C. Lindeman, and C. B. Paris. Decadal analysis of larval connectivity from Cuban snapper (Lutjanidae) spawning aggregations based on biophysical modeling. Marine Ecology Progress Series, 550:175– 190, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/550/m550p175.pdf.

Kenworthy:2017:SRF

[KCMH^{+}17] W. J. Kenworthy, N. Cosentino-Manning, L. Handley, M. Wild, and S. Rouhani. Seagrass response following exposure to Deepwater Horizon oil in the Chandeleur Islands, Louisiana (USA). Marine Ecology Progress Series, 576:145– 161, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m576p145.pdf.

Kellogg:2013:DNA

[KCOP13] M. L. Kellogg, J. C. Cornwell, M. S. Owens, and K. T. Paynter. Denitrification and nutrient assimilation on a restored oyster reef. Marine Ecology Progress Series, 480:1–19, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Klein:2011:DSE

[KCP^{+}11] C. Klein, P. Claquin, A. Pannard, C. Napoléon, B. Le Roy, and B. Véron. Dynamics of soluble extracellular polymeric substances and transparent exopolymer particle pools in coastal ecosystems. Marine Ecology Progress Series, 427: 13–27, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/427/m427p013.pdf.

Kim:2010:EFA

[KCR^{+}10] T. W. Kim, J. H. Christy, J. R. Rissanen, P. D. Ribeiro, and J. C. Choe. Effect of food addition on the reproductive intensity and timing of both sexes of an intertidal crab. Marine Ecology Progress Series, 401:183–194, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/401/
m401p183.pdf.

Kneebone:2012:SRH

[KCS12] J. Kneebone, J. Chisholm, and G. B. Skomal. Seasonal residency, habitat use, and site fidelity of juvenile sand tiger sharks Carcharias taurus in a Massachusetts estuary. Marine Ecology Progress Series, 471:165–181, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/471/m471p165.pdf.

Kough:2017:EEM

[KCS^{+}17] A. S. Kough, H. Cronin, R. Skubel, C. A. Belak, and A. W. Stoner. Efficacy of an established marine protected area at sustaining a queen conch Lobatus gigas population during three decades of monitoring. Marine Ecology Progress Series, 573:177–189, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m573p177.pdf.

Keller:2017:SSR

[KCW^{+}17] A. A. Keller, L. Ciannelli, W. W. Wakefield, V. Simon, J. A. Barth, and S. D. Pierce. Species-specific responses of demersal fishes to near-bottom oxygen levels within the California Current large marine ecosystem. Marine Ecology Progress Series, 568:151–173, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/568/m568p151.pdf.

Koski:2010:SCF

[KDB^{+}10] M. Koski, J. Dutz, W. Klein Breteler, S. Rampen, and A. Noordeloos. Seasonal changes in food quantity and quality of the common North Sea copepods Temora longicornis and Pseudocalanus elongatus: a bioassay approach. Marine Ecology Progress Series, 399:141–155, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/399/m399p141.pdf.

Kelly:2018:ECI

[KdEPW18] S. Kelly, E. de Eyto, R. Poole, and M. White. Ecological consequences of internal seiches in a semi-enclosed, anoxic coastal basin. Marine Ecology Progress Series, 603:265– 272, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/603/m603p265.pdf.

Kininmonth:2010:MML

[KDJP10] S. Kininmonth, M. Drechsler, K. Johst, and H. P. Possingham. Metapopulation mean life time within complex networks. Marine Ecology Progress Series, 417:139–149, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/417/
m417p139.pdf.

Kvile:2014:TEC

[KDO^{+}14] K. Ø. Kvile, P. Dalpadado, E. Orlova, N. C. Stenseth, and L. C. Stige. Temperature effects on Calanus finmarchicus vary in space, time and between developmental stages. Marine Ecology Progress Series, 517:85–104, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/517/
m517p085.pdf.

Kjellerup:2012:EFW

[KDS^{+}12] S. Kjellerup, M. Dünweber, R. Swalethorp, T. G. Nielsen, E. F. Møller, S. Markager, and B. W. Hansen. Effects of a future warmer ocean on the coexisting copepods Calanus finmarchicus and C. glacialis in Disko Bay, western Greenland. Marine Ecology Progress Series, 447:87–108, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/447/
m447p087.pdf.

Kulk:2012:TDG

[KdVvdP^{+}12] G. Kulk, P. de Vries, W. H. van de Poll, R. J. W. Visser, and A. G. J. Buma. Temperature-dependent growth and photophysiology of prokaryotic and eukaryotic oceanic picophytoplankton. Marine Ecology Progress Series, 466:43– 55, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/466/m466p043.pdf.

Kandratavicius:2018:REF

[KdWV^{+}18] N. Kandratavicius, C. Pastor de Ward, N. Venturini, L. Giménez, M. Rodriguez, and P. Muniz. Response of estuarine free-living nematode assemblages to organic enrichment: an experimental approach. Marine Ecology Progress Series, 602:117–133, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/602/m602p117.pdf.

Kaltenberg:2010:TFF

[KEBB10] A. M. Kaltenberg, R. L. Emmett, and K. J. Benoit-Bird. Timing of forage fish seasonal appearance in the Columbia River plume and link to ocean conditions. Marine Ecology Progress Series, 419:171–184, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/419/m419p171.pdf.

Kolb:2010:ESN

[KEH10] G. S. Kolb, J. Ekholm, and P. A. Hambäck. Effects of seabird nesting colonies on algae and aquatic invertebrates in coastal waters. Marine Ecology Progress Series, 417:287– 300, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/417/m417p287.pdf.

Kasten:2013:DET

[KF13a] P. Kasten and A. A. V. Flores. Disruption of endogenous tidal rhythms of larval release linked to food supply and heat stress in an intertidal barnacle. Marine Ecology Progress Series, 472:185–198, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/472/m472p185.pdf.

Kennison:2013:HAT

[KF13b] R. L. Kennison and P. Fong. High amplitude tides that result in floating mats decouple algal distribution from patterns of recruitment and nutrient sources. Marine Ecology Progress Series, 494:73–86, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/494/m494p073.pdf.

Kitagawa:2017:ROS

[KF17] T. Kitagawa and K. Fujioka. Rapid ontogenetic shift in juvenile Pacific bluefin tuna diet. Marine Ecology Progress Series, 571:253–257, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m571p253.pdf.

Kutti:2015:ISC

[KFB15] T. Kutti, J. H. Fosså, and O. A. Bergstad. Influence of structurally complex benthic habitats on fish distribution. Marine Ecology Progress Series, 520:175–190, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/520/
m520p175.pdf.

Kohlbach:2019:LSI

[KFBM19] D. Kohlbach, S. H. Ferguson, T. A. Brown, and C. Michel. Landfast sea ice-benthic coupling during spring and potential impacts of system changes on food web dynamics in Eclipse Sound, Canadian Arctic. Marine Ecology Progress Series, 627:33–48, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m627p033.pdf.

Kozak:2014:SIV

[KFGSMPG14] E. R. Kozak, C. Franco-Gordo, E. Suárez-Morales, and R. Palomares-García. Seasonal and interannual variability of the calanoid copepod community structure in shelf waters of the Eastern Tropical Pacific. Marine Ecology Progress Series, 507:95–110, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/507/m507p095.pdf.

Keller:2018:CRS

[KFW^{+}18] A. A. Keller, P. H. Frey, J. R. Wallace, M. A. Head, C. R. Wetzel, J. M. Cope, and J. H. Harms. Canary rockfishes Sebastes pinniger return from the brink: catch, distribution and life history along the US west coast (Washington to California). Marine Ecology Progress Series, 599:181–200, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/599/
m599p181.pdf.

Klein:2016:CPH

[KGJ^{+}16] E. S. Klein, S. M. Glaser, A. Jordaan, L. Kaufman, and A. A. Rosenberg. A complex past: historical and contemporary fisheries demonstrate nonlinear dynamics and a loss of determinism. Marine Ecology Progress Series, 557:237–246, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m557p237.pdf.

Koslow:2011:IDI

[KGLLW11] J. A. Koslow, R. Goericke, A. Lara-Lopez, and W. Watson. Impact of declining intermediate-water oxygen on deepwater fishes in the California Current. Marine Ecology Progress Series, 436:207–218, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/436/m436p207.pdf.

Kobara:2010:SBG

[KH10] S. Kobara and W. D. Heyman. Sea bottom geomorphology of multi-species spawning aggregation sites in Belize. Marine Ecology Progress Series, 405:243–254, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/405/
m405p243.pdf.

Kordas:2016:DRC

[KH16] R. L. Kordas and C. D. G. Harley. Demographic responses of coexisting species to in situ warming. Marine Ecology Progress Series, 546:147–161, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/546/m546p147.pdf.

Kitaysky:2018:ISR

[KH18] A. S. Kitaysky and G. L. Hunt, Jr. INTRODUCTION : Seabird responses to a changing Bering Sea. Marine Ecology Progress Series, 593:189–194, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m593p189.pdf.

Kahng:2012:ELH

[KHA^{+}12] S. E. Kahng, E. J. Hochberg, A. Apprill, D. Wagner, D. G. Luck, D. Perez, and R. R. Bidigare. Efficient light harvesting in deep-water zooxanthellate corals. Marine Ecology Progress Series, 455:65–77, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/455/m455p065.pdf.

Keller:2017:DCC

[KHÁB^{+}17] S. Keller, M. Hidalgo, D. Álvarez-Berastegui, I. Bitetto, L. Casciaro, D. Cuccu, A. Esteban, G. Garofalo, M. Gonzalez, B. Guijarro, M. Josephides, A. Jadaud, E. Lefkaditou, P. Maiorano, C. Manfredi, B. Marceta, R. Micallef, P. Peristeraki, G. Relini, P. Sartor, M. T. Spedicato, G. Tserpes, and A. Quetglas. Demersal cephalopod communities in the Mediterranean: a large-scale analysis. Marine Ecology Progress Series, 584:105–118, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/584/m584p105.pdf.

Karnauskas:2011:PES

[KHBL11] M. Karnauskas, B. E. Huntington, E. A. Babcock, and D. Lirman. Pre-existing spatial patterns in fish abundances influence species-specific responses in a Caribbean marine reserve. Marine Ecology Progress Series, 432:235–246, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/432/
m432p235.pdf.

Kristensen:2011:CEP

[KHD^{+}11] E. Kristensen, T. Hansen, M. Delefosse, G. T. Banta, and C. O. Quintana. Contrasting effects of the polychaetes Marenzelleria viridis and Nereis diversicolor on benthic metabolism and solute transport in sandy coastal sediment. Marine Ecology Progress Series, 425:125–139, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/425/
m425p125.pdf.

Kim:2019:ECR

[KHD19] S. Kim, K. Hammerstrom, and P. Dayton. Epifaunal community response to iceberg-mediated environmental change in McMurdo Sound, Antarctica. Marine Ecology Progress Series, 613:1–14, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Kinney:2011:CCS

[KHF^{+}11] M. J. Kinney, N. E. Hussey, A. T. Fisk, A. J. Tobin, and C. A. Simpfendorfer. Communal or competitive? Stable isotope analysis provides evidence of resource partitioning within a communal shark nursery. Marine Ecology Progress Series, 439:263–276, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/439/m439p263.pdf.

Kim:2019:CNA

[KHG^{+}19] H. G. Kim, L. E. Hawkins, J. A. Godbold, C. W. Oh, H. S. Rho, and S. J. Hawkins. Comparison of nematode assemblages associated with Sargassum muticum in its native range in South Korea and as an invasive species in the English Channel.

Marine Ecology Progress Series, 611:95–110, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/611/
m611p095.pdf.

Kamel:2012:FSG

[KHGS12] S. J. Kamel, A. R. Hughes, R. K. Grosberg, and J. J. Stachowicz. Fine-scale genetic structure and relatedness in the eelgrass Zostera marina. Marine Ecology Progress Series, 447:127–137, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/447/m447p127.pdf.

Keesing:2018:MRS

[KHH18] J. K. Keesing, A. R. Halford, and K. C. Hall. Mortality rates of small juvenile crown-of-thorns starfish Acanthaster planci on the Great Barrier Reef: implications for population size and larval settlement thresholds for outbreaks. Marine Ecology Progress Series, 597:179–190, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/597/m597p179.pdf.

Karnovsky:2012:LLE

[KHI12] N. J. Karnovsky, K. A. Hobson, and S. J. Iverson. From lavage to lipids: estimating diets of seabirds. Marine Ecology Progress Series, 451:263–284, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m451p263.pdf.

Krietsch:2017:CVI

[KHK^{+}17] J. Krietsch, S. Hahn, M. Kopp, R. A. Phillips, H. U. Peter, and S. Lisovski. Consistent variation in individual migration strategies of brown skuas. Marine Ecology Progress Series, 578:213–225, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m578p213.pdf.

Knowlton:2012:MNA

[KHM^{+}12] A. R. Knowlton, P. K. Hamilton, M. K. Marx, H. M. Pettis, and S. D. Kraus. Monitoring North Atlantic right whale Eubalaena glacialis entanglement rates: a 30 yr retrospective. Marine Ecology Progress Series, 466:293–302, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m466p293.pdf.

Kool:2015:SLC

[KHN15] J. T. Kool, Z. Huang, and S. L. Nichol. Simulated larval connectivity among Australia’s southwest submarine canyons. Marine Ecology Progress Series, 539:77–91, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/539/
m539p077.pdf.

Knip:2010:SNE

[KHS10a] D. M. Knip, M. R. Heupel, and C. A. Simpfendorfer. Sharks in nearshore environments: models, importance, and consequences. Marine Ecology Progress Series, 402:1–11, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Kreibich:2010:FUT

[KHS10b] T. Kreibich, W. Hagen, and R. Saborowski. Food utilization of two pelagic crustaceans in the Greenland Sea: Meganyctiphanes norvegica (Euphausiacea) and Hymenodora glacialis (Decapoda, Caridea). Marine Ecology Progress Series, 413: 105–115, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/413/m413p105.pdf.

Knip:2011:OSM

[KHS^{+}11] D. M. Knip, M. R. Heupel, C. A. Simpfendorfer, A. J. Tobin, and J. Moloney. Ontogenetic shifts in movement and habitat use of juvenile pigeye sharks Carcharhinus amboinensis in a tropical nearshore region. Marine Ecology Progress Series, 425:233–246, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/425/m425p233.pdf.

Kodama:2017:DCF

[KHT^{+}17] T. Kodama, J. Hirai, S. Tamura, T. Takahashi, Y. Tanaka, T. Ishihara, A. Tawa, H. Morimoto, and S. Ohshimo. Diet composition and feeding habits of larval Pacific bluefin tuna Thunnus orientalis in the Sea of Japan: integrated morphological and metagenetic analysis. Marine Ecology Progress Series, 583:211–226, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/583/m583p211.pdf.

Karnovsky:2010:FDL

[KHW^{+}10] N. Karnovsky, A. Harding, W. Walkusz, S. Kwaśniewski, I. Goszczko, J. Wiktor, Jr., H. Routti, A. Bailey, L. McFadden, Z. Brown, G. Beaugrand, and D. Grémillet. Foraging distributions of little auks Alle alle across the Greenland Sea: implications of present and future Arctic climate change. Marine Ecology Progress Series, 415:283–293, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m415p283.pdf.

King:2012:SSR

[KHW^{+}12] S. D. King, G. A. Harper, J. B. Wright, J. C. McInnes, J. E. van der Lubbe, M. L. Dobbins, and S. J. Murray. Site-specific reproductive failure and decline of a population of the endangered yellow-eyed penguin: a case for foraging habitat quality. Marine Ecology Progress Series, 467:233–244, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/467/
m467p233.pdf.

Kiszka:2015:BDE

[KHW15] J. J. Kiszka, M. R. Heithaus, and A. J. Wirsing. Behavioural drivers of the ecological roles and importance of marine mammals. Marine Ecology Progress Series, 523:267– 281, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/523/m523p267.pdf.

Keller:2019:CLL

[KHW^{+}19] A. A. Keller, J. H. Harms, J. R. Wallace, C. Jones, J. A. Benante, and A. Chappell. Changes in long-lived rockfishes after more than a decade of protection within California’s largest marine reserve. Marine Ecology Progress Series, 623: 175–193, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/623/m623p175.pdf.

Kinne:2010:VME

[Kin10a] O. Kinne. 400 volumes of marine ecology progress series (MEPS): a unique success story continues MEPS 400:1. Marine Ecology Progress Series, ??:??, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/misc/m400p001.pdf.

Kinne:2010:MDA

[Kin10b] Otto Kinne. In memoriam: Don Alderdice (1924–2010). Marine Ecology Progress Series, 408:1, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m408p000.pdf.

Kindinger:2016:SEI

[Kin16] T. L. Kindinger. Symmetrical effects of interspecific competition on congeneric coral-reef fishes. Marine Ecology Progress Series, 555:1–11, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Kiorboe:2016:FMP

[Kiø16] T. Kiørboe. Foraging mode and prey size spectra of suspension-feeding copepods and other zooplankton. Marine Ecology Progress Series, 558:15–20, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/558/
m558p015.pdf.

Kitagawa:2013:OPB

[KIU^{+}13] T. Kitagawa, T. Ishimura, R. Uozato, K. Shirai, Y. Amano, A. Shinoda, T. Otake, U. Tsunogai, and S. Kimura. Otolith δ 18O of Pacific bluefin tuna Thunnus orientalis as an indicator of ambient water temperature. Marine Ecology Progress Series, 481:199–209, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m481p199.pdf.

Knutsen:2012:PGS

[KJBS12] H. Knutsen, P. E. Jorde, O. A. Bergstad, and M. Skogen. Population genetic structure in a deepwater fish Coryphaenoides rupestris: patterns and processes. Marine Ecology Progress Series, 460:233–246, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/460/m460p233.pdf.

Kyhn:2015:PCT

[KJC^{+}15] L. A. Kyhn, P. B. Jørgensen, J. Carstensen, N. I. Bech, J. Tougaard, T. Dabelsteen, and J. Teilmann. Pingers cause temporary habitat displacement in the harbour porpoise Phocoena phocoena. Marine Ecology Progress Series, 526:253– 265, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/526/m526p253.pdf.

Krause-Jensen:2011:SBC

[KJCN^{+}11] D. Krause-Jensen, J. Carstensen, S. L. Nielsen, T. Dalsgaard, P. B. Christensen, H. Fossing, and M. B. Rasmussen. Sea bottom characteristics affect depth limits of eelgrass Zostera marina. Marine Ecology Progress Series, 425:91–102, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m425p091.pdf.

King:2015:ECG

[KJW^{+}15] A. L. King, B. D. Jenkins, J. R. Wallace, Y. Liu, G. H. Wikfors, L. M. Milke, and S. L. Meseck. Effects of CO_{2} on growth rate, C:N:P, and fatty acid composition of seven marine phytoplankton species. Marine Ecology Progress Series, 537:59–69, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m537p059.pdf.

Kayfetz:2017:ABC

[KK17] K. Kayfetz and W. Kimmerer. Abiotic and biotic controls on the copepod Pseudodiaptomus forbesi in the upper San Francisco Estuary. Marine Ecology Progress Series, 581:85– 101, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m581p085.pdf.

Kaartvedt:2012:IWM

[KKA12] S. Kaartvedt, T. A. Klevjer, and D. L. Aksnes. Internal wavemediated shading causes frequent vertical migrations in fishes. Marine Ecology Progress Series, 452:1–10, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Katajisto:2013:FYA

[KKL13] T. Katajisto, L. Karjala, and M. Lehtiniemi. Fifteen years after invasion: egg bank of the predatory cladoceran Cercopagis pengoi in the Baltic Sea. Marine Ecology Progress Series, 482:81–92, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/482/m482p081.pdf.

Kakela:2010:FAS

[KKMA^{+}10] R. Käkelä, A. Käkelä, A. Martínez-Abraín, B. Sarzo, M. Louzao, C. Gerique, E. Villuendas, U. Strandberg, R. W. Furness, and D. Oro. Fatty acid signature analysis confirms foraging resources of a globally endangered Mediterranean seabird species: calibration test and application to the wild. Marine Ecology Progress Series, 398:245–258, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/398/
m398p245.pdf.

Khalaman:2016:AVA

[KKMM16] V. V. Khalaman, A. Y. Komendantov, S. S. Malavenda, and T. A. Mikhaylova. Algae versus animals in early fouling communities of the White Sea. Marine Ecology Progress Series, 553:13–32, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/553/m553p013.pdf.

Kodama:2014:VDT

[KKO^{+}14] T. Kodama, H. Kurogi, M. Okazaki, T. Jinbo, S. Chow, T. Tomoda, T. Ichikawa, and T. Watanabe. Vertical distribution of transparent exopolymer particle (TEP) concentration in the oligotrophic western tropical North Pacific. Marine Ecology Progress Series, 513:29–37, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/513/m513p029.pdf.

Kim:2014:APL

[KKPL14] S. H. Kim, J. H. Kim, S. R. Park, and K. S. Lee. Annual and perennial life history strategies of Zostera marina populations under different light regimes. Marine Ecology Progress Series, 509:1–13, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Kelly:2012:DAS

[KKS12a] J. R. Kelly, K. A. Krumhansl, and R. E. Scheibling. Drift algal subsidies to sea urchins in low-productivity habitats. Marine Ecology Progress Series, 452:145–157, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/452/
m452p145.pdf.

Kirui:2012:ESR

[KKS^{+}12b] B. K. Y. Kirui, J. G. Kairo, M. W. Skov, M. Mencuccini, and M. Huxham. Effects of species richness, identity and environmental variables on growth in planted mangroves in Kenya. Marine Ecology Progress Series, 465:1–10, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Kiyashko:2014:TSA

[KKS^{+}14] S. I. Kiyashko, V. I. Kharlamenko, K. Sanamyan, I. L. Alalykina, and L. Würzberg. Trophic structure of the abyssal benthic community in the Sea of Japan inferred from stable isotope and fatty acid analyses. Marine Ecology Progress Series, 500:121–137, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/500/m500p121.pdf.

Kim:2015:USK

[KKY15] J. K. Kim, G. P. Kraemer, and C. Yarish. Use of sugar kelp aquaculture in Long Island Sound and the Bronx River Estuary for nutrient extraction. Marine Ecology Progress Series, 531:155–166, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/531/m531p155.pdf.

Kindt-Larsen:2016:IHR

[KLBT^{+}16] L. Kindt-Larsen, C. W. Berg, J. Tougaard, T. K. Sørensen, K. Geitner, S. Northridge, S. Sveegaard, and F. Larsen. Identification of high-risk areas for harbour porpoise Phocoena phocoena bycatch using remote electronic monitoring and satellite telemetry data. Marine Ecology Progress Series, 555:261–271, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/555/m555p261.pdf.

Kim:2018:MRD

[KLD18] H. Kim, D. E. Lee, and H. W. Ducklow. Mixing regimedependent causality between phytoplankton and bacteria in the subtropical North Atlantic Ocean ecosystem. Marine Ecology Progress Series, 600:41–53, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/600/
m600p041.pdf.

Kay:2012:EMR

[KLKM12] M. C. Kay, H. S. Lenihan, M. J. Kotchen, and C. J. Miller. Effects of marine reserves on California spiny lobster are robust and modified by fine-scale habitat features and distance from reserve borders. Marine Ecology Progress Series, 451:137–150, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/451/m451p137.pdf.

Kruzic:2014:IBC

[KLMR14] P. Kružić, L. Lipej, B. Mavrič, and P. Rodić. Impact of bleaching on the coral Cladocora caespitosa in the eastern Adriatic Sea. Marine Ecology Progress Series, 509:193– 202, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/509/m509p193.pdf.

Kolts:2013:EBS

[KLN^{+}13] J. M. Kolts, J. R. Lovvorn, C. A. North, J. M. Grebmeier, and L. W. Cooper. Effects of body size, gender, and prey availability on diets of snow crabs in the northern Bering Sea. Marine Ecology Progress Series, 483:209–220, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/483/
m483p209.pdf.

Kolts:2015:ODM

[KLNJ15] J. M. Kolts, J. R. Lovvorn, C. A. North, and M. A. Janout. Oceanographic and demographic mechanisms affecting population structure of snow crabs in the northern Bering Sea. Marine Ecology Progress Series, 518:193–208, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/518/
m518p193.pdf.

Kouris:2010:BMF

[KLSJ10] A. Kouris, H. Limén, C. J. Stevens, and S. K. Juniper. Blue mats: faunal composition and food web structure in colonial ciliate (Folliculinopsis sp.) mats at Northeast Pacific hydrothermal vents. Marine Ecology Progress Series, 412:93– 101, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/412/m412p093.pdf.

Kipson:2012:ETS

[KLT^{+}12] S. Kipson, C. Linares, N. Teixidó, T. Bakran-Petricioli, and J. Garrabou. Effects of thermal stress on early developmental stages of a gorgonian coral. Marine Ecology Progress Series, 470:69–78, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/470/m470p069.pdf.

Kelly:2010:UPD

[KM10a] N. Kelly and A. Metaxas. Understanding population dynamics of a numerically dominant species at hydrothermal vents: a matrix modeling approach. Marine Ecology Progress Series, 403:113–128, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/403/m403p113.pdf.

Kiehn:2010:VDD

[KM10b] W. M. Kiehn and J. T. Morris. Variability in dimethylsulfoniopropionate (DMSP) concentrations in Spartina alterniflora and the effect on Littoraria irrorata. Marine Ecology Progress Series, 406:47–55, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/406/m406p047.pdf.

Kuhlmann:2014:DSO

[KM14] M. L. Kuhlmann and B. M. McCabe. Diet specialization in Octopus vulgaris at San Salvador, Bahamas. Marine Ecology Progress Series, 516:229–237, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/516/m516p229.pdf.

Kurle:2017:STV

[KM17] C. M. Kurle and J. K. McWhorter. Spatial and temporal variability within marine isoscapes: implications for interpreting stable isotope data from marine systems. Marine Ecology Progress Series, 568:31–45, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/568/m568p031.pdf.

Kiszka:2016:UUA

[KMGH16] J. J. Kiszka, J. Mourier, K. Gastrich, and M. R. Heithaus. Using unmanned aerial vehicles (UAVs) to investigate shark and ray densities in a shallow coral lagoon. Marine Ecology Progress Series, 560:237–242, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/560/m560p237.pdf.

Kunze:2013:FTB

[KML13] H. B. Kunze, S. G. Morgan, and K. M. Lwiza. Field test of the behavioral regulation of larval transport. Marine Ecology Progress Series, 487:71–87, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m487p071.pdf.

Koslow:2015:DDC

[KMM15] J. A. Koslow, E. F. Miller, and J. A. McGowan. Dramatic declines in coastal and oceanic fish communities off California. Marine Ecology Progress Series, 538:221–227, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/538/
m538p221.pdf.

Knoester:2019:BHF

[KMO19] E. G. Knoester, A. J. Murk, and R. Osinga. Benefits of herbivorous fish outweigh costs of corallivory in coral nurseries placed close to a Kenyan patch reef. Marine Ecology Progress Series, 611:143–155, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/611/m611p143.pdf.

Kendall:2011:PSD

[KMP11] M. S. Kendall, T. J. Miller, and S. J. Pittman. Patterns of scale-dependency and the influence of map resolution on the seascape ecology of reef fish. Marine Ecology Progress Series, 427:259–274, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m427p259.pdf.

Kleisner:2014:RBM

[KMP14] K. Kleisner, H. Mansour, and D. Pauly. Region-based MTI: resolving geographic expansion in the marine trophic index. Marine Ecology Progress Series, 512:185–199, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m512p185.pdf.

Koenigs:2015:TPR

[KMP15] C. Koenigs, R. J. Miller, and H. M. Page. Top predators rely on carbon derived from giant kelp Macrocystis pyrifera. Marine Ecology Progress Series, 537:1–8, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Kaplan:2015:CRS

[KMPS15] M. B. Kaplan, T. A. Mooney, J. Partan, and A. R. Solow. Coral reef species assemblages are associated with ambient soundscapes. Marine Ecology Progress Series, 533:93– 107, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/533/m533p093.pdf.

Kuwae:2010:FMS

[KMSW10] T. Kuwae, E. Miyoshi, S. Sassa, and Y. Watabe. Foraging mode shift in varying environmental conditions by dunlin Calidris alpina. Marine Ecology Progress Series, 406: 281–289, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m406p281.pdf.

Kauffman:2018:HAE

[KMV18] T. C. Kauffman, C. W. Martin, and J. F. Valentine. Hydrological alteration exacerbates the negative impacts of invasive Eurasian milfoil Myriophyllum spicatum by creating hypoxic conditions in a northern Gulf of Mexico estuary. Marine Ecology Progress Series, 592:97–108, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/592/
m592p097.pdf.

Koslow:2017:ICB

[KMW17] J. A. Koslow, H. McMonagle, and W. Watson. Influence of climate on the biodiversity and community structure of fishes in the southern California Current. Marine Ecology Progress Series, 571:193–206, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/571/m571p193.pdf.

Kamenev:2012:BFD

[KN12] G. M. Kamenev and D. A. Nekrasov. Bivalve fauna and distribution in the Amur River estuary — a warm-water ecosystem in the cold-water Pacific region. Marine Ecology Progress Series, 455:195–210, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/455/m455p195.pdf.

Kraft:2013:FER

[KNB^{+}13] A. Kraft, E. M. Nöthig, E. Bauerfeind, D. J. Wildish, G. W. Pohle, U. V. Bathmann, A. Beszczynska-Möller, and M. Klages. First evidence of reproductive success in a southern invader indicates possible community shifts among Arctic zooplankton. Marine Ecology Progress Series, 493:291– 296, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/493/m493p291.pdf.

Kauppi:2017:SVE

[KNIN17] L. Kauppi, J. Norkko, J. Ikonen, and A. Norkko. Seasonal variability in ecosystem functions: quantifying the contribution of invasive species to nutrient cycling in coastal ecosystems. Marine Ecology Progress Series, 572:193–207, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/572/
m572p193.pdf.

Kimirei:2015:DFP

[KNS^{+}15] I. A. Kimirei, I. Nagelkerken, N. Slooter, E. T. Gonzalez, C. M. Huijbers, Y. D. Mgaya, and A. L. Rypel. Demography of fish populations reveals new challenges in appraising juvenile habitat values. Marine Ecology Progress Series, 518:225–237, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m518p225.pdf.

Kawabata:2015:DMR

[KNY^{+}15] Y. Kawabata, A. Nanami, K. Yamamoto, T. Sato, K. Kuwahara, M. Koga, K. Kawaguchi, T. Yamaguchi, I. Ohta, R. Kawabe, G. N. Nishihara, M. Yagi, and K. Soyano. Duration of migration and reproduction in males is dependent on energy reserve in a fish forming spawning aggregations. Marine Ecology Progress Series, 534:149–161, 2015. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_

oa/m534p149.pdf. See errata [Ano15a].

Karlson:2012:SCG

[KO12] R. H. Karlson and R. W. Osman. Species composition and geographic distribution of invertebrates in fouling communities along the east coast of the USA: a regional perspective. Marine Ecology Progress Series, 458:255–268, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m458p255.pdf.

Kadin:2012:CEF

[KÖHSO12] M. Kadin, H. Österblom, J. Hentati-Sundberg, and O. Olsson. Contrasting effects of food quality and quantity on a marine top predator. Marine Ecology Progress Series, 444: 239–249, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m444p239.pdf.

Kennedy:2017:CRT

[KOLDP17] E. V. Kennedy, A. Ordoñez, B. E. Lewis, and G. Diaz-Pulido. Comparison of recruitment tile materials for monitoring coralline algae responses to a changing climate. Marine Ecology Progress Series, 569:129–144, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/569/
m569p129.pdf.

Kumblad:2015:RRB

[KOPW15] L. Kumblad, H. Oskarsson, C. Palmer, and A. K. E. Wiklund. Response and recovery of Baltic Sea blue mussels from exposure to pharmaceuticals. Marine Ecology Progress Series, 526:89–100, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/526/m526p089.pdf.

Kosman:2011:IVL

[KP11] E. T. Kosman and B. Pernet. Intraspecific variation in larval size and its effects on juvenile lophophore size in four bryozoans. Marine Ecology Progress Series, 429:67– 73, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/429/m429p067.pdf.

Kacenas:2018:DDE

[KP18] S. E. Kacenas and R. D. Podolsky. Density-dependent expression of plasticity in larval morphology: effects of actual and apparent competitors. Marine Ecology Progress Series, 593:1–13, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Kraufvelin:2011:SDD

[KPB11] P. Kraufvelin, J. Perus, and E. Bonsdorff. Scale-dependent distribution of soft-bottom infauna and possible structuring forces in low diversity systems. Marine Ecology Progress Series, 426:13–28, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m426p013.pdf.

Kenworthy:2016:RBA

[KPB16] J. M. Kenworthy, D. M. Paterson, and M. J. Bishop. Response of benthic assemblages to multiple stressors: comparative effects of nutrient enrichment and physical disturbance. Marine Ecology Progress Series, 562:37–51, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/562/
m562p037.pdf.

Kruse:2015:UTR

[KPH^{+}15] S. Kruse, E. A. Pakhomov, B. P. V. Hunt, Y. Chikaraishi, N. O. Ogawa, and U. Bathmann. Uncovering the trophic relationship between Themisto gaudichaudii and Salpa thompsoni in the Antarctic polar frontal zone. Marine Ecology Progress Series, 529:63–74, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/529/m529p063.pdf.

Kenchington:2013:ADF

[KPKA13] E. Kenchington, D. Power, and M. Koen-Alonso. Associations of demersal fish with sponge grounds on the continental slopes of the northwest Atlantic. Marine Ecology Progress Series, 477:217–230, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/477/m477p217.pdf.

Kristensen:2012:WBN

[KPLD^{+}12] E. Kristensen, G. Penha-Lopes, M. Delefosse, T. Valdemarsen, C. O. Quintana, and G. T. Banta. What is bioturbation? The need for a precise definition for fauna in aquatic sciences. Marine Ecology Progress Series, 446:285–302, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m446p285.pdf.

Kopp:2011:SPS

[KPM^{+}11] M. Kopp, H. U. Peter, O. Mustafa, S. Lisovski, M. S. Ritz, R. A. Phillips, and S. Hahn. South polar skuas from a single breeding population overwinter in different oceans though show similar migration patterns. Marine Ecology Progress Series, 435:263–267, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m435p263.pdf.

Kroll:2016:EEE

[KPP^{+}16] I. R. Kroll, A. K. Poray, B. J. Puckett, D. B. Eggleston, and F. J. Fodrie. Environmental effects on elemental signatures in eastern oyster Crassostrea virginica shells: using geochemical tagging to assess population connectivity. Marine Ecology Progress Series, 543:173–186, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/543/m543p173.pdf.

Kruger:2018:HFC

[KPr^{+}18] L. Krüger, J. M. Pereira, I. Ramí rez, J. A. Ramos, and V. H. Paiva. How the future climate may modulate the non-breeding distribution of a vulnerable gadfly petrel. Marine Ecology Progress Series, 599:253–266, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/599/m599p253.pdf.

Kough:2014:LSS

[KPS14] A. S. Kough, C. B. Paris, and E. Staaterman. In situ swimming and orientation behavior of spiny lobster (Panulirus argus) postlarvae. Marine Ecology Progress Series, 504:207–219, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/504/m504p207.pdf.

Kendall:2013:CLH

[KPW^{+}13] M. S. Kendall, M. Poti, T. T. Wynne, B. P. Kinlan, and L. B. Bauer. Consequences of the life history traits of pelagic larvae on interisland connectivity during a changing climate. Marine Ecology Progress Series, 489:43–59, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m489p043.pdf.

Kerwath:2019:TDW

[KPW^{+}19] S. E. Kerwath, D. Parker, H. Winker, W. Potts, B. Mann, C. Wilke, and C. Attwood. Tracking the decline of the world’s largest seabream against policy adjustments. Marine Ecology Progress Series, 610:163–173, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m610p163.pdf.

Khan:2016:CBR

[KR16] B. Khan and A. H. Ringwood. Cellular biomarker responses to hypoxia in eastern oysters and Atlantic ribbed marsh mussels. Marine Ecology Progress Series, 546:123–133, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/546/
m546p123.pdf.

Kaartvedt:2017:CWC

[KRA17] S. Kaartvedt, A. Røstad, and D. L. Aksnes. Changing weather causes behavioral responses in the lower mesopelagic. Marine Ecology Progress Series, 574:259–263, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/574/
m574p259.pdf.

Kolasinski:2011:SPO

[KRC^{+}11] J. Kolasinski, K. Rogers, P. Cuet, B. Barry, and P. Frouin. Sources of particulate organic matter at the ecosystem scale: a stable isotope and trace element study in a tropical coral reef. Marine Ecology Progress Series, 443:77–93, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/443/
m443p077.pdf.

Kumala:2017:ODF

[KRC17] L. Kumala, H. U. Riisgård, and D. E. Canfield. Osculum dynamics and filtration activity in small single-osculum explants of the demosponge Halichondria panicea. Marine Ecology Progress Series, 572:117–128, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m572p117.pdf.

Kostecki:2012:TEJ

[KRD^{+}12] C. Kostecki, J. M. Roussel, N. Desroy, G. Roussel, J. Lanshere, H. Le Bris, and O. Le Pape. Trophic ecology of juvenile flatfish in a coastal nursery ground: contributions of intertidal primary production and freshwater particulate organic matter. Marine Ecology Progress Series, 449:221–232, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/449/
m449p221.pdf.

Kristensen:2018:COB

[KRDV18] E. Kristensen, H. Røy, K. Debrabant, and T. Valdemarsen. Carbon oxidation and bioirrigation in sediments along a Skagerrak–Kattegat–Belt Sea depth transect. Marine Ecology Progress Series, 604:33–50, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m604p033.pdf.

Kristensen:2018:TDP

[KRdVG^{+}18] M. L. Kristensen, D. Righton, D. del Villar-Guerra, H. Baktoft, and K. Aarestrup. Temperature and depth preferences of adult sea trout Salmo trutta during the marine migration phase. Marine Ecology Progress Series, 599:209– 224, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/599/m599p209.pdf.

Kaartvedt:2019:HMF

[KROA19] S. Kaartvedt, A. Røstad, A. F. Opdal, and D. L. Aksnes. Herding mesopelagic fish by light. Marine Ecology Progress Series, 625:225–231, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m625p225.pdf.

Kitchens:2018:DAY

[KRR^{+}18] L. L. Kitchens, J. R. Rooker, L. Reynal, B. J. Falterman, E. Saillant, and H. Murua. Discriminating among yellowfin tuna Thunnus albacares nursery areas in the Atlantic Ocean using otolith chemistry. Marine Ecology Progress Series, 603:201–213, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/603/m603p201.pdf.

Kelly:2018:END

[KRW18] J. P. Kelly, C. A. Rothenbach, and W. W. Weathers. Echoes of numerical dependence: responses of wintering waterbirds to Pacific herring spawns. Marine Ecology Progress Series, 597:243–257, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/597/m597p243.pdf.

Krumhansl:2011:DPN

[KS11] K. A. Krumhansl and R. E. Scheibling. Detrital production in Nova Scotian kelp beds: patterns and processes. Marine Ecology Progress Series, 421:67–82, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/421/
m421p067.pdf.

Kelly:2012:FAD

[KS12a] J. R. Kelly and R. E. Scheibling. Fatty acids as dietary tracers in benthic food webs. Marine Ecology Progress Series, 446:1– 22, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Krumhansl:2012:DSS

[KS12b] K. A. Krumhansl and R. E. Scheibling. Detrital subsidy from subtidal kelp beds is altered by the invasive green alga Codium fragile ssp. fragile. Marine Ecology Progress Series, 456:73–85, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/456/m456p073.pdf.

Krumhansl:2012:PFK

[KS12c] K. A. Krumhansl and R. E. Scheibling. Production and fate of kelp detritus. Marine Ecology Progress Series, 467: 281–302, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/467/m467p281.pdf.

Krug:2013:LES

[KS13] J. M. Krug and M. A. Steele. Larval exposure to shared oceanography does not cause spatially correlated recruitment in kelp forest fishes. Marine Ecology Progress Series, 477: 177–188, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/477/m477p177.pdf.

Kaartvedt:2012:ETA

[KSA12] S. Kaartvedt, A. Staby, and D. L. Aksnes. Efficient trawl avoidance by mesopelagic fishes causes large underestimation of their biomass. Marine Ecology Progress Series, 456:1–6, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Kaga:2013:LCC

[KSA^{+}13] T. Kaga, S. Sato, T. Azumaya, N. D. Davis, and M. Fukuwaka. Lipid content of chum salmon Oncorhynchus keta affected by pink salmon O. gorbuscha abundance in the central Bering Sea. Marine Ecology Progress Series, 478:211–221, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/478/m478p211.pdf.

Kelly:2011:IMS

[KSB11] J. R. Kelly, R. E. Scheibling, and T. Balch. Invasionmediated shifts in the macrobenthic assemblage of a rocky subtidal ecosystem. Marine Ecology Progress Series, 437: 69–78, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/437/m437p069.pdf.

Kiszka:2011:ENS

[KSBM^{+}11] J. Kiszka, B. Simon-Bouhet, L. Martinez, C. Pusineri, P. Richard, and V. Ridoux. Ecological niche segregation within a community of sympatric dolphins around a tropical island. Marine Ecology Progress Series, 433:273–288, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/433/
m433p273.pdf.

Karp:2018:FCR

[KSF18] M. A. Karp, R. D. Seitz, and M. C. Fabrizio. Faunal communities on restored oyster reefs: effects of habitat complexity and environmental conditions. Marine Ecology Progress Series, 590:35–51, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/590/m590p035.pdf.

Kashef:2014:OCS

[KSFL14] N. S. Kashef, S. M. Sogard, R. Fisher, and J. L. Largier. Ontogeny of critical swimming speeds for larval and pelagic juvenile rockfishes (Sebastes spp., family Scorpaenidae). Marine Ecology Progress Series, 500:231–243, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/500/
m500p231.pdf.

Kanno:2019:SVM

[KSHS19] S. Kanno, A. M. Schlaff, M. R. Heupel, and C. A. Simpfendorfer. Stationary video monitoring reveals habitat use of stingrays in mangroves. Marine Ecology Progress Series, 621:155–168, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/621/m621p155.pdf.

Kerr:2007:SIS

[KSK07] L. A. Kerr, D. H. Secor, and R. T. Kraus. Stable isotope (δ 13C and δ ^{18}O) and Sr/Ca composition of otoliths as proxies for environmental salinity experienced by an estuarine fish. Marine Ecology Progress Series, 349:245–253, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/349/

m349p245.pdf. See erratum [Ano10e].

Krafft:2012:AKS

[KSK^{+}12] B. A. Krafft, G. Skaret, T. Knutsen, W. Melle, T. Klevjer, and H. Søiland. Antarctic krill swarm characteristics in the Southeast Atlantic sector of the Southern Ocean. Marine Ecology Progress Series, 465:69–83, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m465p069.pdf.

Kearney:2013:AAI

[KSS13] K. A. Kearney, C. Stock, and J. L. Sarmiento. Amplification and attenuation of increased primary production in a marine food web. Marine Ecology Progress Series, 491:1–14, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Klevjer:2010:SCA

[KTF10] T. A. Klevjer, G. A. Tarling, and S. Fielding. Swarm characteristics of Antarctic krill Euphausia superba relative to the proximity of land during summer in the Scotia Sea. Marine Ecology Progress Series, 409:157–170, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/409/
m409p157.pdf.

Knott:2018:SVD

[KTH^{+}18] K. E. Knott, A. Thonig, S. Heiskanen, B. Winding Hansen, and G. T. Banta. Seasonal variation in diversity of marine benthic invertebrates leads to a positive species–genetic diversity correlation (SGDC). Marine Ecology Progress Series, 592:129–140, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/592/m592p129.pdf.

Komorita:2012:ONS

[KTK^{+}12] T. Komorita, H. Tsutsumi, R. Kajihara, N. Suga, S. Shibanuma, T. Yamada, and S. Montani. Oceanic nutrient supply and uptake by microphytobenthos of the Hichirippu Lagoon, Hokkaido, Japan. Marine Ecology Progress Series, 446:161– 171, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/446/m446p161.pdf.

Kogovsek:2014:JBC

[KTKM14] T. Kogovšek, T. Tinta, K. Klun, and A. Malej. Jellyfish biochemical composition: importance of standardised sample processing. Marine Ecology Progress Series, 510:275–288, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m510p275.pdf.

Kanamori:2019:CCS

[KTNO19] Y. Kanamori, A. Takasuka, S. Nishijima, and H. Okamura. Climate change shifts the spawning ground northward and extends the spawning period of chub mackerel in the western North Pacific. Marine Ecology Progress Series, 624:155– 166, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m624p155.pdf.

Kokubun:2018:IAC

[KTP^{+}18] N. Kokubun, A. Takahashi, R. Paredes, R. C. Young, N. N. Sato, T. Yamamoto, D. M. Kikuchi, E. V. Kitaiskaia, M. Ito, Y. Watanuki, A. Will, R. Lauth, M. D. Romano, and A. S. Kitaysky. Inter-annual climate variability affects foraging behavior and nutritional state of thick-billed murres breeding in the southeastern Bering Sea. Marine Ecology Progress Series, 593:195–208, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m593p195.pdf.

Kluger:2016:AEI

[KTR^{+}16] L. C. Kluger, M. H. Taylor, E. Barriga Rivera, E. Torres Silva, and M. Wolff. Assessing the ecosystem impact of scallop bottom culture through a community analysis and trophic modelling approach. Marine Ecology Progress Series, 547:121–135, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/547/m547p121.pdf.

Kinney:2013:CSM

[KV13] E. L. Kinney and I. Valiela. Changes in δ ^{15}N in salt marsh sediments in a long-term fertilization study. Marine Ecology Progress Series, 477:41–52, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/477/m477p041.pdf.

Karati:2018:RPF

[KVR^{+}18] K. K. Karati, G. Vineetha, T. V. Raveendran, K. R. Muraleedharan, H. Habeebrehman, K. P. Philson, and C. T. Achuthankutty. River plume fronts and their implications for the biological production of the Bay of Bengal, Indian Ocean. Marine Ecology Progress Series, 597:79–98, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/597/
m597p079.pdf.

Kelaher:2013:PRS

[KVY^{+}13] B. P. Kelaher, J. Van Den Broek, P. H. York, M. J. Bishop, and D. J. Booth. Positive responses of a seagrass ecosystem to experimental nutrient enrichment. Marine Ecology Progress Series, 487:15–25, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/487/m487p015.pdf.

Knights:2010:RRI

[KW10] A. M. Knights and K. Walters. Recruit-recruit interactions, density-dependent processes and population persistence in the eastern oyster Crassostrea virginica. Marine Ecology Progress Series, 404:79–90, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/404/m404p079.pdf.

Koivisto:2012:ICA

[KW12] M. Koivisto and M. Westerbom. Invertebrate communities associated with blue mussel beds in a patchy environment: a landscape ecology approach. Marine Ecology Progress Series, 471:101–110, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/471/m471p101.pdf.

Kainge:2017:FSE

[KW17] P. Kainge and K. Wieland. Fine-scale environmental effects on Cape hake survey catch rates in the northern Benguela, using data from a trawl-mounted instrument package. Marine Ecology Progress Series, 584:185–198, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/584/
m584p185.pdf.

Kilfoil:2017:BRU

[KWC^{+}17] J. P. Kilfoil, A. J. Wirsing, M. D. Campbell, J. J. Kiszka, K. R. Gastrich, M. R. Heithaus, Y. Zhang, and M. E. Bond. Baited remote underwater video surveys undercount sharks at high densities: insights from full-spherical camera technologies. Marine Ecology Progress Series, 585:113–121, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/585/
m585p113.pdf.

Kleisner:2010:MSA

[KWDD10] K. M. Kleisner, J. F. Walter III, S. L. Diamond, and D. J. Die. Modeling the spatial autocorrelation of pelagic fish abundance. Marine Ecology Progress Series, 411:203–213, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/411/
m411p203.pdf.

Kirchner:2018:HFM

[KWH^{+}18] T. Kirchner, D. N. Wiley, E. L. Hazen, S. E. Parks, L. G. Torres, and A. S. Friedlaender. Hierarchical foraging movement of humpback whales relative to the structure of their prey. Marine Ecology Progress Series, 607:237–250, 2018. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m607p237.pdf.

Kesselring:2012:IOS

[KWM12] H. Kesselring, R. Wheatley, and D. J. Marshall. Initial offspring size mediates trade-off between fecundity and longevity in the field. Marine Ecology Progress Series, 465:129–136, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2012/465/m465p129.pdf.

Keen:2017:WWS

[KWM^{+}17] E. M. Keen, J. Wray, H. Meuter, K. L. Thompson, J. P. Barlow, and C. R. Picard. ‘whale wave’: shifting strategies structure the complex use of critical fjord habitat by humpbacks. Marine Ecology Progress Series, 567:211–233, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/567/
m567p211.pdf.

Kappes:2011:VRP

[KWPL11] M. A. Kappes, H. Weimerskirch, D. Pinaud, and M. Le Corre. Variability of resource partitioning in sympatric tropical boobies. Marine Ecology Progress Series, 441:281–294, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2011/441/m441p281.pdf.

Kenitz:2013:PPS

[KWS^{+}13] K. Kenitz, R. G. Williams, J. Sharples, Ö. Selsil, and V. N. Biktashev. The paradox of the plankton: species competition and nutrient feedback sustain phytoplankton diversity. Marine Ecology Progress Series, 490:107–119, 2013. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m490p107.pdf.

Keller:2014:DDF

[KWW^{+}14] A. A. Keller, W. W. Wakefield, C. E. Whitmire, B. H. Horness, M. A. Bellman, and K. L. Bosley. Distribution of demersal fishes along the US west coast (Canada to Mexico) in relation to spatial fishing closures (2003–2011). Marine Ecology Progress Series, 501:169–190, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/501/m501p169.pdf.

Layman:2012:CTE

[LA12a] C. A. Layman and J. E. Allgeier. Characterizing trophic ecology of generalist consumers: a case study of the invasive lionfish in the Bahamas. Marine Ecology Progress Series, 448:131–141, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/448/m448p131.pdf.

Luo:2012:VMR

[LA12b] J. Luo and J. S. Ault. Vertical movement rates and habitat use of Atlantic tarpon. Marine Ecology Progress Series, 467:167–180, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m467p167.pdf.

Lough:2014:TRV

[LA14] R. G. Lough and A. Aretxabaleta. Transport and retention of vertically migrating adult mysid and decapod shrimp in the tidal front on Georges Bank. Marine Ecology Progress Series, 514:119–135, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2014/514/m514p119.pdf.

Llabres:2010:SPC

[LAALH10] M. Llabrés, S. Agustí, P. Alonso-Laita, and G. J. Herndl. Synechococcus and Prochlorococcus cell death induced by UV radiation and the penetration of lethal UVR in the Mediterranean Sea. Marine Ecology Progress Series, 399:27–37, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m399p027.pdf.

Lardies:2010:PCM

[LAB10] M. A. Lardies, M. B. Arias, and L. D. Bacigalupe. Phenotypic covariance matrix in life-history traits along a latitudinal gradient: a study case in a geographically widespread crab on the coast of Chile. Marine Ecology Progress Series, 412:179–187, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/412/m412p179.pdf.

Lozano-Alvarez:2010:ISA

[LÁBFÁF^{+}10] E. Lozano-Álvarez, P. Briones-Fourzán, L. Álvarez-Filip, H. M. Weiss, F. Negrete-Soto, and C. Barradas-Ortiz. Influence of shelter availability on interactions between Caribbean spiny lobsters and moray eels: implications for artificial lobster enhancement. Marine Ecology Progress Series, 400:175– 185, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m400p175.pdf.

Lindo-Atichati:2012:VMS

[LABG^{+}12] D. Lindo-Atichati, F. Bringas, G. Goni, B. Muhling, F. E. Muller-Karger, and S. Habtes. Varying mesoscale structures influence larval fish distribution in the northern Gulf of Mexico. Marine Ecology Progress Series, 463:245–257, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2012/463/m463p245.pdf.

Lenz:2017:AGJ

[LABM17] A. J. Lenz, L. Avens, and M. Borges-Martins. Age and growth of juvenile green turtles Chelonia mydas in the western South Atlantic Ocean. Marine Ecology Progress Series, 568:191–201, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/568/m568p191.pdf.

Lowther:2015:RVT

[LAH^{+}15] A. D. Lowther, H. Ahonen, G. Hofmeyr, W. C. Oosthuizen, P. J. N. De Bruyn, C. Lydersen, and K. M. Kovacs. Reliability of VHF telemetry data for measuring attendance patterns of marine predators: a comparison with time depth recorder data. Marine Ecology Progress Series, 538:249– 256, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/538/m538p249.pdf.

Langbehn:2019:LCZ

[LAK^{+}19] T. J. Langbehn, D. L. Aksnes, S. Kaartvedt, Ø. Fiksen, and C. Jørgensen. Light comfort zone in a mesopelagic fish emerges from adaptive behaviour along a latitudinal gradient. Marine Ecology Progress Series, 623:161–174, 2019. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m623p161.pdf.

Lavaud:2018:NIS

[LAL^{+}18] R. Lavaud, S. Artigaud, F. Le Grand, A. Donval, P. Soudant, J. Flye-Sainte-Marie, T. Strohmeier, Ø. Strand, A. Leynaert, B. Beker, A. Chatterjee, and F. Jean. New insights into the seasonal feeding ecology of Pecten maximus using pigments, fatty acids and sterols analyses. Marine Ecology Progress Series, 590:109–129, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/590/m590p109.pdf.

Lobon:2011:HML

[LALÁC11] C. M. Lobón, J. L. Acuña, M. López-Álvarez, and F. L. Capitanio. Heritability of morphological and life history traits in a pelagic tunicate. Marine Ecology Progress Series, 422: 145–154, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/422/m422p145.pdf.

Lopez-Acosta:2018:SUS

[LALCM18] M. López-Acosta, A. Leynaert, V. Coquille, and M. Maldonado. Silicon utilization by sponges: an assessment of seasonal changes. Marine Ecology Progress Series, 605:111–123, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m605p111.pdf.

Lathlean:2010:SSB

[LAM10] J. A. Lathlean, D. J. Ayre, and T. E. Minchinton. Supplyside biogeography: geographic patterns of settlement and early mortality for a barnacle approaching its range limit. Marine Ecology Progress Series, 412:141–150, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/412/
m412p141.pdf.

Lathlean:2011:RIT

[LAM11] J. A. Lathlean, D. J. Ayre, and T. E. Minchinton. Rocky intertidal temperature variability along the southeast coast of Australia: comparing data from in situ loggers, satellitederived SST and terrestrial weather stations. Marine Ecology Progress Series, 439:83–95, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/439/m439p083.pdf.

Lathlean:2013:TVL

[LAM13] J. A. Lathlean, D. J. Ayre, and T. E. Minchinton. Temperature variability at the larval scale affects early survival and growth of an intertidal barnacle. Marine Ecology Progress Series, 475:155–166, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/475/m475p155.pdf.

Lorentsen:2018:SGF

[LANE18] S. H. Lorentsen, T. Anker-Nilssen, and K. E. Erikstad. Seabirds as guides for fisheries management: European shag Phalacrocorax aristotelis diet as indicator of saithe Pollachius virens recruitment. Marine Ecology Progress Series, 586:193– 201, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m586p193.pdf.

Lorentsen:2015:FFA

[LANER15] S. H. Lorentsen, T. Anker-Nilssen, K. E. Erikstad, and N. Røv. Forage fish abundance is a predictor of timing of breeding and hatching brood size in a coastal seabird. Marine Ecology Progress Series, 519:209–220, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m519p209.pdf.

Laptikhovsky:2019:OPW

[Lap19] V. Laptikhovsky. Opinion piece: White blood of the Antarctic icefish: Why? Marine Ecology Progress Series, 626: 227–231, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/626/m626p227.pdf.

Lefevre:2011:TVC

[LB11] C. D. Lefèvre and D. R. Bellwood. Temporal variation in coral reef ecosystem processes: herbivory of macroalgae by fishes. Marine Ecology Progress Series, 422:239–251, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/422/
m422p239.pdf.

Lake:2015:MER

[LB15a] S. J. Lake and M. J. Brush. Modeling estuarine response to load reductions in a warmer climate: York River Estuary, Virginia, USA. Marine Ecology Progress Series, 538: 81–98, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/538/m538p081.pdf.

Lefevre:2015:DRS

[LB15b] C. D. Lefèvre and D. R. Bellwood. Disturbance and recolonisation by small reef fishes: the role of local movement versus recruitment. Marine Ecology Progress Series, 537:205–215, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/537/m537p205.pdf.

Lake:2013:IVE

[LBAK13] S. J. Lake, M. J. Brush, I. C. Anderson, and H. I. Kator. Internal versus external drivers of periodic hypoxia in a coastal plain tributary estuary: the York River, Virginia. Marine Ecology Progress Series, 492:21–39, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/492/
m492p021.pdf.

Luttikhuizen:2019:LGC

[LBC^{+}19] P. C. Luttikhuizen, J. Beermann, R. P. M. A. Crooijmans, R. G. Jak, and J. W. P. Coolen. Low genetic connectivity in a fouling amphipod among man-made structures in the southern North Sea. Marine Ecology Progress Series, 615:133–142, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/615/m615p133.pdf.

Lynch:2010:NRN

[LBCL10] P. D. Lynch, M. J. Brush, E. D. Condon, and R. J. Latour. Net removal of nitrogen through ingestion of phytoplankton by Atlantic menhaden Brevoortia tyrannus in Chesapeake Bay. Marine Ecology Progress Series, 401:195–209, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/401/
m401p195.pdf.

Lowerre-Barbieri:2015:ARR

[LBCS^{+}15] S. Lowerre-Barbieri, L. Crabtree, T. Switzer, S. Walters Burnsed, and C. Guenther. Assessing reproductive resilience: an example with South Atlantic red snapper Lutjanus campechanus. Marine Ecology Progress Series, 526: 125–141, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m526p125.pdf.

Laurel:2015:SPD

[LBD^{+}15] B. J. Laurel, A. J. Basilio, C. Danley, C. H. Ryer, and M. Spencer. Substrate preference and delayed settlement in northern rock sole larvae Lepidopsetta polyxystra. Marine Ecology Progress Series, 519:183–193, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/519/
m519p183.pdf.

Litz:2010:EVO

[LBE^{+}10] M. N. C. Litz, R. D. Brodeur, R. L. Emmett, S. S. Heppell, R. S. Rasmussen, L. O’Higgins, and M. S. Morris. Effects of variable oceanographic conditions on forage fish lipid content and fatty acid composition in the northern California Current. Marine Ecology Progress Series, 405:71–85, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/405/
m405p071.pdf.

Lord:2017:ICC

[LBG17] J. P. Lord, J. P. Barry, and D. Graves. Impact of climate change on direct and indirect species interactions. Marine Ecology Progress Series, 571:1–11, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Laptikhovsky:2018:CRW

[LBH18] V. V. Laptikhovsky, C. J. Barrett, and P. R. Hollyman. From coral reefs to whale teeth: estimating mortality from natural accumulations of skeletal materials. Marine Ecology Progress Series, 598:273–291, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m598p273.pdf.

Legrand:2016:CWH

[LBJ^{+}16] B. Legrand, A. Benneveau, A. Jaeger, P. Pinet, G. Potin, S. Jaquemet, and M. Le Corre. Current wintering habitat of an endemic seabird of Réunion Island, Barau’s petrel Pterodroma baraui, and predicted changes induced by global warming. Marine Ecology Progress Series, 550:235–248, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/550/
m550p235.pdf.

Lidgard:2014:PBA

[LBJI14] D. C. Lidgard, W. D. Bowen, I. D. Jonsen, and S. J. Iverson. Predator-borne acoustic transceivers and GPS tracking reveal spatiotemporal patterns of encounters with acoustically tagged fish in the open ocean. Marine Ecology Progress Series, 501:157–168, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/501/m501p157.pdf.

Levitan:2014:LTR

[LBJK14] D. R. Levitan, W. Boudreau, J. Jara, and N. Knowlton. Longterm reduced spawning in Orbicella coral species due to temperature stress. Marine Ecology Progress Series, 515:1–10, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Lind:2012:SFA

[LBL^{+}12] Y. Lind, B. M. Bäcklin, K. Lundström, S. M. Budge, M. Walton, and O. Karlsson. Stability of fatty acid composition in seal blubber during long-term storage. Marine Ecology Progress Series, 461:283–291, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/461/m461p283.pdf.

Leat:2013:IWA

[LBM^{+}13] E. H. K. Leat, S. Bourgeon, E. Magnusdottir, G. W. Gabrielsen, W. J. Grecian, S. A. Hanssen, K. Olafsdottir, A. Petersen, R. A. Phillips, H. Strøm, S. Ellis, A. T. Fisk, J. O. Bustnes, R. W. Furness, and K. Borgå. Influence of wintering area on persistent organic pollutants in a breeding migratory seabird. Marine Ecology Progress Series, 491: 277–293, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/491/m491p277.pdf.

Ladd:2018:EOV

[LBM^{+}18] T. M. Ladd, J. A. Bullington, P. G. Matson, R. M. Kudela, and M. D. Iglesias-Rodríguez. Exposure to oil from the 2015 Refugio spill alters the physiology of a common harmful algal bloom species, Pseudo-nitzschia australis, and the ubiquitous coccolithophore, Emiliania huxleyi. Marine Ecology Progress Series, 603:61–78, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/603/m603p061.pdf.

Long:2015:STS

[LBMZ15] M. H. Long, P. Berg, K. J. McGlathery, and J. C. Zieman. Sub-tropical seagrass ecosystem metabolism measured by eddy covariance. Marine Ecology Progress Series, 529:75–90, 2015.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m529p075.pdf.

Lalande:2011:DPO

[LBN11] C. Lalande, E. Bauerfeind, and E. M. Nöthig. Downward particulate organic carbon export at high temporal resolution in the eastern Fram Strait: influence of Atlantic water on flux composition. Marine Ecology Progress Series, 440:127–136, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/440/m440p127.pdf.

Lara:2016:HEM

[LBP^{+}16] M. Lara, T. J. Bouma, G. Peralta, J. van Soelen, and J. L. Pérez-Lloréns. Hydrodynamic effects of macrophyte microtopography: spatial consequences of interspecific benthic transitions. Marine Ecology Progress Series, 561:123– 136, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/561/m561p123.pdf.

Lassoued:2019:BEP

[LBP^{+}19] J. Lassoued, J. M. F. Babarro, X. A. Padín, L. A. Comeau, N. Bejaoui, and F. F. Pérez. Behavioural and eco-physiological responses of the mussel Mytilus galloprovincialis to acidification and distinct feeding regimes. Marine Ecology Progress Series, 626:97–108, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/626/m626p097.pdf.

Lewis:2019:FCM

[LBR^{+}19] K. P. Lewis, A. D. Buren, P. M. Regular, F. K. Mowbray, and H. M. Murphy. Forecasting capelin Mallotus villosus biomass on the Newfoundland shelf. Marine Ecology Progress Series, 616:171–183, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m616p171.pdf.

Loury:2015:TEG

[LBS^{+}15] E. K. Loury, S. M. Bros, R. M. Starr, D. A. Ebert, and G. M. Cailliet. Trophic ecology of the gopher rockfish Sebastes carnatus inside and outside of central California marine protected areas. Marine Ecology Progress Series, 536:229–241, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m536p229.pdf.

Linares:2010:MPA

[LBT^{+}10] C. Linares, O. Bianchimani, O. Torrents, C. Marschal, P. Drap, and J. Garrabou. Marine protected areas and the conservation of long-lived marine invertebrates: the Mediterranean red coral. Marine Ecology Progress Series, 402:69– 79, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/402/m402p069.pdf.

Lavers:2013:LSM

[LBVH13] J. L. Lavers, A. L. Bond, S. L. Van Wilgenburg, and K. A. Hobson. Linking at-sea mortality of a pelagic shearwater to breeding colonies of origin using biogeochemical markers. Marine Ecology Progress Series, 491:265–275, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/491/
m491p265.pdf.

Lowerre-Barbieri:2013:SFR

[LBWB^{+}13] S. K. Lowerre-Barbieri, S. Walters, J. Bickford, W. Cooper, and R. Muller. Site fidelity and reproductive timing at a spotted seatrout spawning aggregation site: individual versus population scale behavior. Marine Ecology Progress Series, 481:181–197, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/481/m481p181.pdf.

Lefebvre:2012:SEA

[LC12] F. Lefebvre and A. J. Crivelli. Salinity effects on anguillicolosis in Atlantic eels: a natural tool for disease control. Marine Ecology Progress Series, 471:193–202, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/471/
m471p193.pdf.

Li:2013:LEH

[LC13] A. Li and J. M. Y. Chiu. Latent effects of hypoxia on the gastropod Crepidula onyx. Marine Ecology Progress Series, 480:145–154, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/480/m480p145.pdf.

Liversage:2018:CEE

[LC18] K. Liversage and M. G. Chapman. Coastal ecological engineering and habitat restoration: incorporating biologically diverse boulder habitat. Marine Ecology Progress Series, 593:173–185, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/593/m593p173.pdf.

Lawson:2010:MDB

[LCB10] G. L. Lawson, M. R. Castleton, and B. A. Block. Movements and diving behavior of Atlantic bluefin tuna Thunnus thynnus in relation to water column structure in the northwestern Atlantic. Marine Ecology Progress Series, 400: 245–265, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/400/m400p245.pdf.

Lucey:2012:CAS

[LCB^{+}12] S. M. Lucey, A. M. Cook, J. L. Boldt, J. S. Link, T. E. Essington, and T. J. Miller. Comparative analyses of surplus production dynamics of functional feeding groups across 12 northern hemisphere marine ecosystems. Marine Ecology Progress Series, 459:219–229, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m459p219.pdf.

Lubcker:2016:VGP

[LCB^{+}16] N. Lübcker, R. Condit, R. S. Beltran, P. J. N. de Bruyn, and M. N. Bester. Vibrissal growth parameters of southern elephant seals Mirounga leonina: obtaining fine-scale, timebased stable isotope data. Marine Ecology Progress Series, 559:243–255, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/559/m559p243.pdf.

Lopez-Castro:2013:STP

[LCBK^{+}13] M. C. López-Castro, K. A. Bjorndal, G. D. Kamenov, R. Zenil-Ferguson, and A. B. Bolten. Sea turtle population structure and connections between oceanic and neritic foraging areas in the Atlantic revealed through trace elements. Marine Ecology Progress Series, 490:233–246, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/490/m490p233.pdf.

Liu:2011:TES

[LCC^{+}11] B. Liu, X. Chen, Y. Chen, H. Lu, and W. Qian. Trace elements in the statoliths of jumbo flying squid off the Exclusive Economic Zones of Chile and Peru. Marine Ecology Progress Series, 429:93–101, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/429/m429p093.pdf.

Linares:2012:ETA

[LCC12] C. Linares, E. Cebrian, and R. Coma. Effects of turf algae on recruitment and juvenile survival of gorgonian corals. Marine Ecology Progress Series, 452:81–88, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/452/
m452p081.pdf.

Lim:2019:APW

[LCCT19] A. Y. H. Lim, I. Z. W. Chan, L. R. Carrasco, and P. A. Todd. Aposematism in pink warty sea cucumbers: independent effects of chromatic and achromatic cues. Marine Ecology Progress Series, 631:157–164, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m631p157.pdf.

Long:2011:PDE

[LCD11] J. D. Long, E. Cochrane, and R. E. Dolecal. Previous disturbance enhances the negative effects of trampling on barnacles. Marine Ecology Progress Series, 437:165–173, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/437/
m437p165.pdf.

Levin:2014:BDR

[LCF^{+}14] N. Levin, M. Coll, S. Fraschetti, G. Gal, S. Giakoumi, C. Göke, J. J. Heymans, S. Katsanevakis, T. Mazor, B. Öztürk, G. Rilov, J. Gajewski, J. Steenbeek, and S. Kark. Biodiversity data requirements for systematic conservation planning in the Mediterranean Sea. Marine Ecology Progress Series, 508:261–281, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/508/m508p261.pdf.

Lopez:2010:PPI

[LCFR10] M. S. López, R. Coutinho, C. E. L. Ferreira, and G. Rilov. Predator–prey interactions in a bioinvasion scenario: differential predation by native predators on two exotic rocky intertidal bivalves. Marine Ecology Progress Series, 403:101– 112, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/403/m403p101.pdf.

Leopold:2017:FMW

[LCG^{+}17] M. Léopold, O. Chateau, H. Gabriault, J. Ham, S. Andréfouët, J. Raubani, and P. Dumas. Fish movements within community-managed fishery reserve networks: an acoustic survey of Lethrinus harak in Vanuatu. Marine Ecology Progress Series, 571:153–168, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/571/m571p153.pdf.

Laiz-Carrion:2019:SIA

[LCGS^{+}19] R. Laiz-Carrión, T. Gerard, J. J. Suca, E. Malca, A. Uriarte, J. M. Quintanilla, S. Privoznik, J. K. Llopiz, J. Lamkin, and A. García. Stable isotope analysis indicates resource partitioning and trophic niche overlap in larvae of four tuna species in the Gulf of Mexico. Marine Ecology Progress Series, 619:53–68, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m619p053.pdf.

Lance:2012:HSD

[LCJ^{+}12] M. M. Lance, W. Y. Chang, S. J. Jeffries, S. F. Pearson, and A. Acevedo-Gutiérrez. Harbor seal diet in northern Puget Sound: implications for the recovery of depressed fish stocks. Marine Ecology Progress Series, 464:257–271, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/464/
m464p257.pdf.

Liu:2011:PRT

[LCL^{+}11] C. C. Liu, J. M. Y. Chiu, L. Li, P. K. S. Shin, and S. G. Cheung. Physiological responses of two sublittoral nassariid gastropods to hypoxia. Marine Ecology Progress Series, 429: 75–85, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/429/m429p075.pdf.

Lourenco:2013:RDG

[LCL^{+}13] P. M. Lourenço, P. Catry, M. Lecoq, I. Ramírez, and J. P. Granadeiro. Role of disturbance, geology and other environmental factors in determining abundance and diversity in coastal avian communities during winter. Marine Ecology Progress Series, 479:223–234, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/479/m479p223.pdf.

Lourenco:2017:ITL

[LCL^{+}17] P. M. Lourenço, T. Catry, R. J. Lopes, T. Piersma, and J. P. Granadeiro. Invisible trophic links? Quantifying the importance of non-standard food sources for key intertidal avian predators in the Eastern Atlantic. Marine Ecology Progress Series, 563:219–232, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/563/m563p219.pdf.

Letessier:2016:EPB

[LCM^{+}16] T. B. Letessier, M. J. Cox, J. J. Meeuwig, P. H. Boersch-Supan, and A. S. Brierley. Enhanced pelagic biomass around coral atolls. Marine Ecology Progress Series, 546:271–276, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/546/m546p271.pdf.

Lukowiak:2018:HCC

[ŁCM^{+}18] M. Łukowiak, K. L. Cramer, D. Madzia, M. G. Hynes, R. D. Norris, and A. O’Dea. Historical change in a Caribbean reef sponge community and long-term loss of sponge predators. Marine Ecology Progress Series, 601:127–137, 2018. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m601p127.pdf.

Laiz-Carrion:2013:HPC

[LCQT^{+}13] R. Laiz-Carrión, J. M. Quintanilla, A. P. Torres, F. Alemany, and A. García. Hydrographic patterns conditioning variable trophic pathways and early life dynamics of bullet tuna Auxis rochei larvae in the Balearic Sea. Marine Ecology Progress Series, 475:203–212, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/475/m475p203.pdf.

Lindsay:2016:CEB

[LCR^{+}16] R. E. Lindsay, R. Constantine, J. Robbins, D. K. Mattila, A. Tagarino, and T. E. Dennis. Characterising essential breeding habitat for whales informs the development of large-scale Marine Protected Areas in the South Pacific. Marine Ecology Progress Series, 548:263–275, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/548/m548p263.pdf.

Liu:2010:MSF

[LCS^{+}10] H. Liu, M. Chen, K. Suzuki, C. K. Wong, and B. Chen. Mesozooplankton selective feeding in subtropical coastal waters as revealed by HPLC pigment analysis. Marine Ecology Progress Series, 407:111–123, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/407/m407p111.pdf.

Laurenceau-Cornec:2015:PMC

[LCTD^{+}15] E. C. Laurenceau-Cornec, T. W. Trull, D. M. Davies, C. L. De La Rocha, and S. Blain. Phytoplankton morphology controls on marine snow sinking velocity. Marine Ecology Progress Series, 520:35–56, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m520p035.pdf.

Linares:2010:FCE

[LD10] C. Linares and D. F. Doak. Forecasting the combined effects of disparate disturbances on the persistence of long-lived gorgonians: a case study of Paramuricea clavata. Marine Ecology Progress Series, 402:59–68, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/402/m402p059.pdf.

Lewis:2017:EGM

[LD17a] N. S. Lewis and T. H. DeWitt. Effect of green macroalgal blooms on the behavior, growth, and survival of cockles Clinocardium nuttallii in Pacific NW estuaries. Marine Ecology Progress Series, 582:105–120, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/582/
m582p105.pdf.

Lowen:2017:DCG

[LD17b] J. B. Lowen and C. DiBacco. Distributional changes in a guild of non-indigenous tunicates in the NW Atlantic under highresolution climate projections. Marine Ecology Progress Series, 570:173–186, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/570/m570p173.pdf.

Liu:2012:PND

[LDAE12] S. Y. V. Liu, C. F. Dai, G. R. Allen, and M. V. Erdmann. Phylogeography of the neon damselfish Pomacentrus coelestis indicates a cryptic species and different species origins in the West Pacific Ocean. Marine Ecology Progress Series, 458:155–167, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/458/m458p155.pdf.

Lindegren:2010:RSR

[LDM10] M. Lindegren, R. Diekmann, and C. Möllmann. Regime shifts, resilience and recovery of a cod stock. Marine Ecology Progress Series, 402:239–253, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/402/m402p239.pdf.

Lyon:2017:DSM

[LDP^{+}17] B. J. Lyon, R. G. Dwyer, R. D. Pillans, H. A. Campbell, and C. E. Franklin. Distribution, seasonal movements and habitat utilisation of an endangered shark, Glyphis glyphis, from northern Australia. Marine Ecology Progress Series, 573:203–213, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/573/m573p203.pdf.

Lecours:2015:SSG

[LDS^{+}15] V. Lecours, R. Devillers, D. C. Schneider, V. L. Lucieer, C. J. Brown, and E. N. Edinger. Spatial scale and geographic context in benthic habitat mapping: review and future directions. Marine Ecology Progress Series, 535:259–284, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m535p259.pdf.

Lovvorn:2013:NOT

[LDT^{+}13] J. R. Lovvorn, S. E. W. De La Cruz, J. Y. Takekawa, L. E. Shaskey, and S. E. Richman. Niche overlap, threshold food densities, and limits to prey depletion for a diving duck assemblage in an estuarine bay. Marine Ecology Progress Series, 476:251–268, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/476/m476p251.pdf.

Lagarde:2017:RPO

[LdU^{+}17] F. Lagarde, E. Roque d’orbcastel, M. Ubertini, S. Mortreux, I. Bernard, A. Fiandrino, C. Chiantella, B. Bec, C. Roques, D. Bonnet, G. Miron, M. Richard, S. Pouvreau, and C. Lett. Recruitment of the Pacific oyster Crassostrea gigas in a shellfish-exploited Mediterranean lagoon: discovery, driving factors and a favorable environmental window. Marine Ecology Progress Series, 578:1–17, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/feature/m578p001.pdf.

Lossent:2017:MDS

[LDVP^{+}17] J. Lossent, L. Di Iorio, C. A. Valentini-Poirier, P. Boissery, and C. Gervaise. Mapping the diversity of spectral shapes discriminates between adjacent benthic biophonies. Marine Ecology Progress Series, 585:31–48, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/585/m585p031.pdf.

Lenihan:2010:RPV

[LE10] H. S. Lenihan and P. J. Edmunds. Response of Pocillopora verrucosa to corallivory varies with environmental conditions. Marine Ecology Progress Series, 409:51–63, 2010. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m409p051.pdf.

Lindegren:2013:TDC

[LE13] M. Lindegren and M. Eero. Threshold-dependent climate effects and high mortality limit recruitment and recovery of the Kattegat cod. Marine Ecology Progress Series, 490:223– 232, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/490/m490p223.pdf.

Lillis:2014:ESD

[LEB14a] A. Lillis, D. B. Eggleston, and D. R. Bohnenstiehl. Estuarine soundscapes: distinct acoustic characteristics of oyster reefs compared to soft-bottom habitats. Marine Ecology Progress Series, 505:1–17, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Lillis:2014:SVL

[LEB14b] A. Lillis, D. B. Eggleston, and D. R. Bohnenstiehl. Soundscape variation from a larval perspective: the case for habitatassociated sound as a settlement cue for weakly swimming estuarine larvae. Marine Ecology Progress Series, 509:57– 70, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/509/m509p057.pdf.

Lyon:2019:FCS

[LEB^{+}19] R. P. Lyon, D. B. Eggleston, D. R. Bohnenstiehl, C. A. Layman, S. W. Ricci, and J. E. Allgeier. Fish community structure, habitat complexity, and soundscape characteristics of patch reefs in a tropical, back-reef system. Marine Ecology Progress Series, 609:33–48, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/609/m609p033.pdf.

Leclerc:2018:PSV

[Lec18] J. C. Leclerc. Patterns of spatial variability between contrasting substrata: a boulder-field study. Marine Ecology Progress Series, 597:23–38, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/597/m597p023.pdf.

Litchman:2012:PNT

[LEKT12] E. Litchman, K. F. Edwards, C. A. Klausmeier, and M. K. Thomas. Phytoplankton niches, traits and eco-evolutionary responses to global environmental change. Marine Ecology Progress Series, 470:235–248, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m470p235.pdf.

Longphuirt:2013:SSS

[LERS13] S. Ní Longphuirt, C. Eschmann, C. Russell, and D. B. Stengel. Seasonal and species-specific response of five brown macroalgae to high atmospheric CO_{2}. Marine Ecology Progress Series, 493:91–102, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/493/m493p091.pdf.

Lok:2012:SAB

[LET^{+}12] E. K. Lok, D. Esler, J. Y. Takekawa, S. W. De La Cruz, W. S. Boyd, D. R. Nysewander, J. R. Evenson, and D. H. Ward. Spatiotemporal associations between Pacific herring spawn and surf scoter spring migration: evaluating a ‘silver wave’ hypothesis. Marine Ecology Progress Series, 457: 139–150, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/457/m457p139.pdf.

Lonnstedt:2014:NHB

[LF14] O. M. Lönnstedt and A. J. Frisch. NOTE: Habitat bleaching disrupts threat responses and persistence in anemonefish. Marine Ecology Progress Series, 517:265–270, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/517/
m517p265.pdf.

Lockfield:2013:MFH

[LFD13] K. C. Lockfield, J. W. Fleeger, and L. A. Deegan. Mummichog Fundulus heteroclitus responses to long-term, wholeecosystem nutrient enrichment. Marine Ecology Progress Series, 492:211–222, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/492/m492p211.pdf.

Lilley:2014:CGJ

[LFE^{+}14] M. K. S. Lilley, M. Ferraris, A. Elineau, L. Berline, P. Cuvilliers, L. Gilletta, A. Thiéry, G. Gorsky, and F. Lombard. Culture and growth of the jellyfish Pelagia noctiluca in the laboratory. Marine Ecology Progress Series, 510:265–273, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m510p265.pdf.

Large:2015:CPE

[LFFL15] S. I. Large, G. Fay, K. D. Friedland, and J. S. Link. Critical points in ecosystem responses to fishing and environmental pressures. Marine Ecology Progress Series, 521:1–17, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Liu:2012:NDF

[LFG^{+}12] H. Liu, M. J. Fogarty, S. M. Glaser, I. Altman, C. h. Hsieh, L. Kaufman, A. A. Rosenberg, and G. Sugihara. Nonlinear dynamic features and co-predictability of the Georges Bank fish community. Marine Ecology Progress Series, 464:195– 207, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/464/m464p195.pdf.

Langaard:2014:SDS

[LFJ^{+}14] L. Langård, O. A. Fatnes, A. Johannessen, G. Skaret, B. E. Axelsen, L. Nøttestad, A. Slotte, K. H. Jensen, and A. Fernö. State-dependent spatial and intra-school dynamics in prespawning herring Clupea harengus in a semi-enclosed ecosystem. Marine Ecology Progress Series, 501:251–263, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2014/501/m501p251.pdf.

Li:2011:MGD

[LFL11] Q. P. Li, P. J. S. Franks, and M. R. Landry. Microzooplankton grazing dynamics: parameterizing grazing models with dilution experiment data from the California Current Ecosystem. Marine Ecology Progress Series, 438:59–69, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/438/
m438p059.pdf.

Lienart:2016:GVS

[LFM^{+}16] C. Liénart, E. Feunteun, M. J. Miller, J. Aoyama, J. M. Mortillaro, C. Hubas, M. Kuroki, S. Watanabe, C. Dupuy, A. Carpentier, T. Otake, K. Tsukamoto, and T. Meziane. Geographic variation in stable isotopic and fatty acid composition of anguilliform leptocephali and particulate organic matter in the South Pacific. Marine Ecology Progress Series, 544:225–241, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/544/m544p225.pdf.

Lages:2011:CTR

[LFMC11] B. G. Lages, B. G. Fleury, C. Menegola, and J. C. Creed. Change in tropical rocky shore communities due to an alien coral invasion. Marine Ecology Progress Series, 438:85–96, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2011/438/m438p085.pdf.

Lynch:2012:DAB

[LFN^{+}12] H. J. Lynch, W. F. Fagan, R. Naveen, S. G. Trivelpiece, and W. Z. Trivelpiece. Differential advancement of breeding phenology in response to climate may alter staggered breeding among sympatric pygoscelid penguins. Marine Ecology Progress Series, 454:135–145, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m454p135.pdf.

Lammers:2011:HWM

[LFPA^{+}11] M. O. Lammers, P. I. Fisher-Pool, W. W. L. Au, C. G. Meyer, K. B. Wong, and R. E. Brainard. Humpback whale Megaptera novaeangliae song reveals wintering activity in the Northwestern Hawaiian Islands. Marine Ecology Progress Series, 423:261–268, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m423p261.pdf.

Leal:2014:TEF

[LFPC^{+}14] M. C. Leal, C. Ferrier-Pagès, R. Calado, J. A. Brandes, M. E. Frischer, and J. C. Nejstgaard. Trophic ecology of the facultative symbiotic coral Oculina arbuscula. Marine Ecology Progress Series, 504:171–179, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/504/m504p171.pdf.

Luiz:2013:WSI

[LFRF13] O. J. Luiz, S. R. Floeter, L. A. Rocha, and C. E. L. Ferreira. As we see it: Perspectives for the lionfish invasion in the South Atlantic: Are Brazilian reefs protected by the currents? Marine Ecology Progress Series, 485:1–7, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Lagarde:2019:DTS

[LFU^{+}19] F. Lagarde, A. Fiandrino, M. Ubertini, E. Roque d’Orbcastel, S. Mortreux, C. Chiantella, B. Bec, D. Bonnet, C. Roques, I. Bernard, M. Richard, T. Guyondet, S. Pouvreau, and C. Lett. Duality of trophic supply and hydrodynamic connectivity drives spatial patterns of Pacific oyster recruitment. Marine Ecology Progress Series, 632:81–100, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/632/
m632p081.pdf.

Liu:2018:ELN

[LG18] D. Liu and P. M. Glibert. Ecophysiological linkage of nitrogen enrichment to heavily silicified diatoms in winter. Marine Ecology Progress Series, 604:51–63, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m604p051.pdf.

Lehtiniemi:2013:DRA

[LGB^{+}13] M. Lehtiniemi, E. Gorokhova, S. Bolte, H. Haslob, B. Huwer, T. Katajisto, L. Lennuk, S. Majaneva, A. Põllumäe, M. Schaber, O. Setälä, T. B. H. Reusch, S. Viitasalo-Frösén, I. Vuorinen, and P. Välipakka. Distribution and reproduction of the Arctic ctenophore Mertensia ovum in the Baltic Sea. Marine Ecology Progress Series, 491:111–124, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/491/
m491p111.pdf.

Latour:2017:STD

[LGB17] R. J. Latour, J. Gartland, and C. F. Bonzek. Spatiotemporal trends and drivers of fish condition in Chesapeake Bay. Marine Ecology Progress Series, 579:1–17, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Lazoski:2011:PPA

[LGBSC11] C. Lazoski, J. Gusmão, P. Boudry, and A. M. Solé-Cava. Phylogeny and phylogeography of Atlantic oyster species: evolutionary history, limited genetic connectivity and isolation by distance. Marine Ecology Progress Series, 426:197– 212, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/426/m426p197.pdf.

Liu:2010:LWE

[LGC10] H. C. Liu, G. C. Gong, and J. Chang. Lateral water exchange between shelf-margin upwelling and Kuroshio waters influences phosphorus stress in microphytoplankton. Marine Ecology Progress Series, 409:121–130, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m409p121.pdf.

Licursi:2010:EOT

[LGD10] M. Licursi, N. Gómez, and J. Donadelli. Ecological optima and tolerances of coastal benthic diatoms in the freshwatermixohaline zone of the Río de la plata estuary. Marine Ecology Progress Series, 418:105–117, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/418/m418p105.pdf.

Long:2012:DDI

[LGJH12] W. C. Long, E. F. Gamelin, E. G. Johnson, and A. H. Hines. Density-dependent indirect effects: apparent mutualism and apparent competition coexist in a two-prey system. Marine Ecology Progress Series, 456:139–148, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/456/
m456p139.pdf.

Lauzon-Guay:2011:RSE

[LGL11] Jean-Sébastian Lauzon-Guay and Devin A. Lyons. Reproducibility in simulation experiments: Comment on Munguia et al. (2010). Marine Ecology Progress Series, 432:299–302, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_

oa/m432p299.pdf. See [MOH^{+}10] and reply [MHO^{+}11].

Link:2012:SLL

[LGM^{+}12] J. S. Link, S. Gaichas, T. J. Miller, T. Essington, A. Bundy, J. Boldt, K. F. Drinkwater, and E. Moksness. Synthesizing lessons learned from comparing fisheries production in 13 northern hemisphere ecosystems: emergent fundamental features. Marine Ecology Progress Series, 459:293–302, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m459p293.pdf.

LaScala-Gruenewald:2016:QTE

[LGMB^{+}16] D. E. LaScala-Gruenewald, L. P. Miller, M. E. S. Bracken, B. J. Allen, and M. W. Denny. Quantifying the top-down effects of grazers on a rocky shore: selective grazing and the potential for competition. Marine Ecology Progress Series, 553:49–66, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m553p049.pdf.

Llaveria:2010:SST

[LGR^{+}10] G. Llaveria, E. Garcés, O. N. Ross, R. I. Figueroa, N. Sampedro, and E. Berdalet. Small-scale turbulence can reduce parasite infectivity to dinoflagellates. Marine Ecology Progress Series, 412:45–56, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m412p045.pdf.

Lauzon-Guay:2010:SDE

[LGS10] J. S. Lauzon-Guay and R. E. Scheibling. Spatial dynamics, ecological thresholds and phase shifts: modelling grazer aggregation and gap formation in kelp beds. Marine Ecology Progress Series, 403:29–41, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/403/m403p029.pdf.

Luo:2014:EDF

[LGT^{+}14] J. Y. Luo, B. Grassian, D. Tang, J. O. Irisson, A. T. Greer, C. M. Guigand, S. McClatchie, and R. K. Cowen. Environmental drivers of the fine-scale distribution of a gelatinous zooplankton community across a mesoscale front. Marine Ecology Progress Series, 510:129–149, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m510p129.pdf.

Levas:2015:OCF

[LGW^{+}15] S. Levas, A. G. Grottoli, M. E. Warner, W. J. Cai, J. Bauer, V. Schoepf, J. H. Baumann, Y. Matsui, C. Gearing, T. F. Melman, K. D. Hoadley, D. T. Pettay, X. Hu, Q. Li, H. Xu, and Y. Wang. Organic carbon fluxes mediated by corals at elevated pCO_{2} and temperature. Marine Ecology Progress Series, 519:153–164, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/519/m519p153.pdf.

Loflen:2010:BRV

[LH10] C. L. Loflen and K. A. Hovel. Behavioral responses to variable predation risk in the California spiny lobster Panulirus interruptus. Marine Ecology Progress Series, 420:135– 144, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/420/m420p135.pdf.

Lannin:2011:VPD

[LH11] R. Lannin and K. Hovel. Variable prey density modifies the effects of seagrass habitat structure on predator– prey interactions. Marine Ecology Progress Series, 442:59– 70, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/442/m442p059.pdf.

Langlois:2014:CST

[LH14] L. A. Langlois and M. O. Hoogenboom. Capacity for shortterm physiological acclimation to light does not control the lower depth distributions of branching corals. Marine Ecology Progress Series, 508:149–162, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/508/m508p149.pdf.

LeGault:2016:CAG

[LH16] K. N. LeGault and H. L. Hunt. Cannibalism among green sea urchins Strongylocentrotus droebachiensis in the laboratory and field. Marine Ecology Progress Series, 542:1–12, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Loffler:2019:MRE

[LH19] Z. Loffler and A. S. Hoey. Microtopographic refuges enhance recruitment and survival but inhibit growth of propagules of the tropical macroalga Sargassum swartzii. Marine Ecology Progress Series, 627:61–70, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/627/m627p061.pdf.

Longo:2015:RTM

[LHB^{+}15] C. Longo, S. Hornborg, V. Bartolino, M. T. Tomczak, L. Ciannelli, S. Libralato, and A. Belgrano. Role of trophic models and indicators in current marine fisheries management. Marine Ecology Progress Series, 538:257–272, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/538/
m538p257.pdf.

Lord:2019:OAM

[LHB19] J. P. Lord, E. M. Harper, and J. P. Barry. Ocean acidification may alter predator–prey relationships and weaken nonlethal interactions between gastropods and crabs. Marine Ecology Progress Series, 616:83–94, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m616p083.pdf.

Lee:2015:HMS

[LHC^{+}15] C. L. Lee, Y. H. Huang, C. Y. Chung, S. C. Hsiao, and H. J. Lin. Herbivory in multi-species, tropical seagrass beds. Marine Ecology Progress Series, 525:65–80, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/525/
m525p065.pdf.

Lee:2017:ODT

[LHDH17] K. A. Lee, C. Huveneers, T. Duong, and R. G. Harcourt. The ocean has depth: two- versus three-dimensional space use estimators in a demersal reef fish. Marine Ecology Progress Series, 572:223–241, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/572/m572p223.pdf.

Lowther:2011:CHF

[LHHG11] A. D. Lowther, R. G. Harcourt, D. J. Hamer, and S. D. Goldsworthy. Creatures of habit: foraging habitat fidelity of adult female Australian sea lions. Marine Ecology Progress Series, 443:249–263, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/443/m443p249.pdf.

Laidre:2011:LLE

[LHJ11] K. L. Laidre and M. P. Heide-Jørgensen. Life in the lead: extreme densities of narwhals Monodon monoceros in the offshore pack ice. Marine Ecology Progress Series, 423:269– 278, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/423/m423p269.pdf.

Laidre:2010:SAB

[LHJH^{+}10] K. L. Laidre, M. P. Heide-Jørgensen, P. Heagerty, A. Cossio, B. Bergström, and M. Simon. Spatial associations between large baleen whales and their prey in West Greenland. Marine Ecology Progress Series, 402:269–284, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/402/
m402p269.pdf.

Lee:2011:LSP

[LHL11] L. H. Lee, L. Y. Hsieh, and H. J. Lin. In situ production and respiration of the benthic community during emersion on subtropical intertidal sandflats. Marine Ecology Progress Series, 441:33–47, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/441/m441p033.pdf.

Lagrue:2016:SAB

[LHPP16] C. Lagrue, K. Heaphy, B. Presswell, and R. Poulin. Strong association between parasitism and phenotypic variation in a supralittoral amphipod. Marine Ecology Progress Series, 553:111–123, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/553/m553p111.pdf.

Latasa:2014:PDB

[LHS^{+}14] M. Latasa, J. Henjes, R. Scharek, P. Assmy, R. Röttgers, and V. Smetacek. Progressive decoupling between phytoplankton growth and microzooplankton grazing during an iron-induced phytoplankton bloom in the Southern Ocean (EIFEX). Marine Ecology Progress Series, 513:39–50, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m513p039.pdf.

Ledee:2016:MPT

[LHT^{+}16] E. J. I. Lédée, M. R. Heupel, A. J. Tobin, A. Mapleston, and C. A. Simpfendorfer. Movement patterns of two carangid species in inshore habitats characterised using network analysis. Marine Ecology Progress Series, 553:219– 232, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/553/m553p219.pdf.

Luo:2017:IDS

[LHT^{+}17] M. Luo, J. Huang, C. Tong, Y. Liu, X. Duan, and Y. Hu. Iron dynamics in a subtropical estuarine tidal marsh: effect of season and vegetation. Marine Ecology Progress Series, 577: 1–15, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Lamare:2018:LTW

[LHU^{+}18] M. Lamare, J. Harianto, S. Uthicke, A. Agüera, S. Karelitz, D. Pecorino, J. Chin, and M. Byrne. Larval thermal windows in native and hybrid Pseudoboletia progeny (Echinoidea) as potential drivers of the hybridization zone. Marine Ecology Progress Series, 598:99–112, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/598/m598p099.pdf.

Lyndby:2019:BOP

[LHW^{+}19] N. H. Lyndby, J. B. Holm, D. Wangpraseurt, C. Ferrier-Pagès, and M. Kühl. Bio-optical properties and radiative energy budgets in fed and unfed scleractinian corals (Pocillopora sp.) during thermal bleaching. Marine Ecology Progress Series, 629: 1–17, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Lamont:2018:SHU

[LI18] M. M. Lamont and A. R. Iverson. Shared habitat use by juveniles of three sea turtle species. Marine Ecology Progress Series, 606:187–200, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/606/m606p187.pdf.

Lowen:2013:PID

[LIT13] J. B. Lowen, D. J. Innes, and R. J. Thompson. Predatorinduced defenses differ between sympatric Mytilus edulis and M. trossulus. Marine Ecology Progress Series, 475:135–143, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/475/m475p135.pdf.

Lamb:2010:TRC

[LJ10] R. W. Lamb and D. W. Johnson. Trophic restructuring of coral reef fish communities in a large marine reserve. Marine Ecology Progress Series, 408:169–180, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/408/m408p169.pdf.

Liversage:2014:ABF

[LJB14] K. Liversage, N. Janetzki, and K. Benkendorff. Associations of benthic fauna with different rock types, and evidence of changing effects during succession. Marine Ecology Progress Series, 505:131–143, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m505p131.pdf.

Lambert:2011:QPI

[LJK^{+}11] G. I. Lambert, S. Jennings, M. J. Kaiser, H. Hinz, and J. G. Hiddink. Quantification and prediction of the impact of fishing on epifaunal communities. Marine Ecology Progress Series, 430:71–86, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/430/m430p071.pdf.

Lea:2010:DOL

[LJM^{+}10] M. A. Lea, D. Johnson, S. Melin, R. Ream, and T. Gelatt. Diving ontogeny and lunar responses in a highly migratory mammal, the northern fur seal Callorhinus ursinus. Marine Ecology Progress Series, 419:233–247, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/419/m419p233.pdf.

Lelievre:2012:IMI

[LJM^{+}12] S. Lelièvre, M. Jérôme, G. E. Maes, S. Vaz, S. Calaivany, and V. Verrez-Bagnis. Integrating molecular identification of pelagic eggs with geostatistical mapping to improve the delineation of North Sea fish spawning grounds. Marine Ecology Progress Series, 445:161–172, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/445/m445p161.pdf.

Lehtonen:2015:DEF

[LK15] T. K. Lehtonen and C. Kvarnemo. Density effects on fish egg survival and infections depend on salinity. Marine Ecology Progress Series, 540:183–191, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/540/m540p183.pdf.

Loxton:2014:VMC

[LKB^{+}14] J. Loxton, P. Kuklinski, D. K. A. Barnes, J. Najorka, M. Spencer Jones, and J. S. Porter. Variability of Mg-calcite in Antarctic bryozoan skeletons across spatial scales. Marine Ecology Progress Series, 507:169–180, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/507/
m507p169.pdf.

Lagasse:2015:SAR

[LKC^{+}15] C. R. Lagasse, A. Knudby, J. Curtis, J. L. Finney, and S. P. Cox. Spatial analyses reveal conservation benefits for cold-water corals and sponges from small changes in a trawl fishery footprint. Marine Ecology Progress Series, 528:161– 172, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/528/m528p161.pdf.

Lanari:2017:DED

[LKC^{+}17] M. Lanari, H. Kennedy, M. S. Copertino, M. Wallner-Kersanach, and M. C. Claudino. Dynamics of estuarine drift macroalgae: growth cycles and contributions to sediments in shallow areas. Marine Ecology Progress Series, 570:

41–55, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/570/m570p041.pdf.

Lee:2012:CME

[LKCH12] D. Y. Lee, D. P. Keller, B. C. Crump, and R. R. Hood. Community metabolism and energy transfer in the Chesapeake Bay estuarine turbidity maximum. Marine Ecology Progress Series, 449:65–82, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/449/m449p065.pdf.

Lewis:2017:SGC

[LKDP17] B. Lewis, E. V. Kennedy, and G. Diaz-Pulido. Seasonal growth and calcification of a reef-building crustose coralline alga on the Great Barrier Reef. Marine Ecology Progress Series, 568:73–86, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/568/m568p073.pdf.

Levy:2017:SHH

[LKL^{+}17] Y. Levy, T. Keren, N. Leader, G. Weil, D. Tchernov, and G. Rilov. Spatiotemporal hotspots of habitat use by loggerhead (Caretta caretta) and green (Chelonia mydas) sea turtles in the Levant basin as tools for conservation. Marine Ecology Progress Series, 575:165–179, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/575/m575p165.pdf.

Lee:2019:CBS

[LKL^{+}19] D. H. Lee, J. H. Kim, Y. M. Lee, Y. K. Jin, C. Paull, D. Kim, and K. H. Shin. Chemosynthetic bacterial signatures in frenulata tubeworm Oligobrachia sp. in an active mud volcano of the Canadian Beaufort Sea. Marine Ecology Progress Series, 628:95–104, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/628/m628p095.pdf.

Li:2011:SDP

[LKLC11] B. Li, D. M. Karl, R. M. Letelier, and M. J. Church. Size-dependent photosynthetic variability in the North Pacific subtropical gyre. Marine Ecology Progress Series, 440: 27–40, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m440p027.pdf.

Lien:2013:HGS

[LKN^{+}13] Y. T. Lien, S. Keshavmurthy, Y. Nakano, S. Plathong, H. Huang, C. M. Hsu, H. Fukami, Y. Yamashita, H. J. Hsieh, J. T. Wang, and C. A. Chen. Host genetics and Symbiodinium D diversity in a stress-tolerant scleractinian coral, Oulastrea crispata, in the West Pacific. Marine Ecology Progress Series, 473:163–177, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/473/m473p163.pdf.

Loxton:2014:VSM

[LKN^{+}14] J. Loxton, P. Kuklinski, J. Najorka, M. Spencer Jones, and J. S. Porter. Variability in the skeletal mineralogy of temperate bryozoans: the relative influence of environmental and biological factors. Marine Ecology Progress Series, 510: 45–57, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/510/m510p045.pdf.

Logan:2012:LDM

[LKS12] C. A. Logan, L. E. Kost, and G. N. Somero. Latitudinal differences in Mytilus californianus thermal physiology. Marine Ecology Progress Series, 450:93–105, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/450/
m450p093.pdf.

Logan:2018:RIR

[LL18] R. K. Logan and C. G. Lowe. Residency and inter-reef connectivity of three gamefishes between natural reefs and a large mitigation artificial reef. Marine Ecology Progress Series, 593:111–126, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/593/m593p111.pdf.

Lopez-Lopez:2015:RPA

[LLBP15] L. Lopez-Lopez, V. Bartolino, and I. Preciado. Role of prey abundance and geographical variables in a demersal top predator’s feeding habits (Merluccius merluccius). Marine Ecology Progress Series, 541:165–177, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/541/m541p165.pdf.

Lai:2018:BSS

[LLBT18] S. Lai, L. H. L. Loke, T. J. Bouma, and P. A. Todd. Biodiversity surveys and stable isotope analyses reveal key differences in intertidal assemblages between tropical seawalls and rocky shores. Marine Ecology Progress Series, 587:41–53, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m587p041.pdf.

Lin:2014:TSM

[LLC^{+}14] H. Y. Lin, P. Y. Lin, N. N. Chang, J. C. Shiao, and S. J. Kao. Trophic structure of megabenthic food webs along depth gradients in the South China Sea and off northeastern Taiwan. Marine Ecology Progress Series, 501:53–66, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/501/
m501p053.pdf.

Chen:2011:COT

[lLChTC11] Y. l. Lee Chen, S. h. Tuo, and H. Y. Chen. Co-occurrence and transfer of fixed nitrogen from Trichodesmium spp. to diatoms in the low-latitude Kuroshio Current in the NW Pacific. Marine Ecology Progress Series, 421:25–38, 2011. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m421p025.pdf.

Linacre:2012:TDC

[LLCM^{+}12] L. Linacre, M. R. Landry, R. Cajal-Medrano, J. R. Lara-Lara, J. M. Hernández-Ayón, R. R. Mouriño-Pérez, E. García-Mendoza, and C. Bazán-Guzmán. Temporal dynamics of carbon flow through the microbial plankton community in a coastal upwelling system off northern Baja California, Mexico. Marine Ecology Progress Series, 461:31–46, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/461/
m461p031.pdf.

Lord:2012:EWD

[LLD^{+}12] C. Lord, J. Lorion, A. Dettai, S. Watanabe, K. Tsukamoto, C. Cruaud, and P. Keith. From endemism to widespread distribution: phylogeography of three amphidromous Sicyopterus species (teleostei: Gobioidei: Sicydiinae). Marine Ecology Progress Series, 455:269–285, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/455/m455p269.pdf.

Lucey:2018:CLH

[LLF^{+}18] N. M. Lucey, C. Lombardi, M. Florio, S. D. Rundle, P. Calosi, and M. C. Gambi. A comparison of life-history traits in calcifying spirorbinae polychaetes living along natural pH gradients. Marine Ecology Progress Series, 589:141–152, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/589/
m589p141.pdf.

Li:2011:DDS

[LLH^{+}11] T. Li, S. Liu, L. Huang, H. Huang, J. Lian, Y. Yan, and S. Lin. Diatom to dinoflagellate shift in the summer phytoplankton community in a bay impacted by nuclear power plant thermal effluent. Marine Ecology Progress Series, 424: 75–85, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/424/m424p075.pdf.

Laneri:2010:TRI

[LLMA^{+}10] K. Laneri, M. Louzao, A. Martínez-Abraín, J. M. Arcos, E. J. Belda, J. Guallart, A. Sánchez, M. Giménez, R. Maestre, and D. Oro. Trawling regime influences longline seabird bycatch in the Mediterranean: new insights from a small-scale fishery. Marine Ecology Progress Series, 420:241–252, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/420/
m420p241.pdf.

Lombard:2010:PEN

[LLP^{+}10] F. Lombard, L. Legendre, M. Picheral, A. Sciandra, and G. Gorsky. Prediction of ecological niches and carbon export by appendicularians using a new multispecies ecophysiological model. Marine Ecology Progress Series, 398:109– 125, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/398/m398p109.pdf.

Lekunberri:2012:RBI

[LLRC^{+}12] I. Lekunberri, T. Lefort, C. Romera-Castillo, C. Cardelús, M. Coll-Lladó, C. Ruiz-González, C. Marrasé, and J. M. Gasol. Relationship between induced phytoplankton blooms and the structure and dynamics of the free-living heterotrophic bacterial community. Marine Ecology Progress Series, 448: 23–37, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/448/m448p023.pdf.

Larsson:2018:HEA

[LLS^{+}18] M. E. Larsson, O. F. Laczka, I. M. Suthers, P. A. Ajani, and M. A. Doblin. Hitchhiking in the East Australian Current: rafting as a dispersal mechanism for harmful epibenthic dinoflagellates. Marine Ecology Progress Series, 596: 49–60, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/596/m596p049.pdf.

Larsson:2013:SGR

[LLv13] A. I. Larsson, T. Lundälv, and D. van Oevelen. Skeletal growth, respiration rate and fatty acid composition in the coldwater coral Lophelia pertusa under varying food conditions. Marine Ecology Progress Series, 483:169–184, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/483/
m483p169.pdf.

Lathlean:2012:MTS

[LM12] J. A. Lathlean and T. E. Minchinton. Manipulating thermal stress on rocky shores to predict patterns of recruitment of marine invertebrates under a changing climate. Marine Ecology Progress Series, 467:121–136, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/467/m467p121.pdf.

Low:2018:LFT

[LM18] N. H. N. Low and F. Micheli. Lethal and functional thresholds of hypoxia in two key benthic grazers. Marine Ecology Progress Series, 594:165–173, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/594/m594p165.pdf.

Lellys:2019:PFM

[LMB^{+}19] N. T. Lellys, R. L. Moura, R. M. Bonaldo, R. B. Francini-Filho, and F. Z. Gibran. Parrotfish functional morphology and bioerosion on SW Atlantic reefs. Marine Ecology Progress Series, 629:149–163, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2019/629/m629p149.pdf.

Livore:2018:HMS

[LMBM18] J. P. Livore, M. M. Mendez, G. Bigatti, and F. Márquez. Habitat-modulated shell shape and spatial segregation in a Patagonian false limpet (Siphonaria lessonii). Marine Ecology Progress Series, 606:55–63, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/606/m606p055.pdf.

LePort:2014:BMS

[LMC14] A. Le Port, J. C. Montgomery, and A. E. Croucher. Biophysical modelling of snapper Pagrus auratus larval dispersal from a temperate MPA. Marine Ecology Progress Series, 515:203–215, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/515/m515p203.pdf.

Lorentsen:2019:RSE

[LMCD19] S. H. Lorentsen, J. Mattisson, and S. Christensen-Dalsgaard. Reproductive success in the European shag is linked to annual variation in diet and foraging trip metrics. Marine Ecology Progress Series, 619:137–147, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m619p137.pdf.

Lawerence:2012:DPG

[LMD12a] C. Lawerence and S. Menden-Deuer. Drivers of protistan grazing pressure: seasonal signals of plankton community composition and environmental conditions. Marine Ecology Progress Series, 459:39–52, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/459/m459p039.pdf.

Lloyd:2012:PVD

[LMd12b] M. J. Lloyd, A. Metaxas, and B. deYoung. Patterns in vertical distribution and their potential effects on transport of larval benthic invertebrates in a shallow embayment. Marine Ecology Progress Series, 469:37–52, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/469/m469p037.pdf.

Lloyd:2012:PBF

[LMd12c] M. J. Lloyd, A. Metaxas, and B. deYoung. Physical and biological factors affect the vertical distribution of larvae of benthic gastropods in a shallow embayment. Marine Ecology Progress Series, 464:135–151, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/464/m464p135.pdf.

Langer:2019:AAC

[LME^{+}19] J. A. F. Langer, C. L. Meunier, U. Ecker, H. G. Horn, K. Schwenk, and M. Boersma. Acclimation and adaptation of the coastal calanoid copepod Acartia tonsa to ocean acidification: a long-term laboratory investigation. Marine Ecology Progress Series, 619:35–51, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/619/m619p035.pdf.

Lloyd:2012:TAP

[LMG12] H. B. Lloyd, B. R. Murray, and P. E. Gribben. Trait and abundance patterns in two marine molluscs: the influence of abiotic conditions operating across multiple spatial scales. Marine Ecology Progress Series, 463:205–214, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/463/
m463p205.pdf.

LeCorre:2013:VRD

[LMGJ13] N. Le Corre, A. L. Martel, F. Guichard, and L. E. Johnson. Variation in recruitment: differentiating the roles of primary and secondary settlement of blue mussels Mytilus spp. Marine Ecology Progress Series, 481:133–146, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/481/
m481p133.pdf.

Losh:2012:MIC

[LMH12] J. L. Losh, F. M. M. Morel, and B. M. Hopkinson. Modest increase in the C:N ratio of N-limited phytoplankton in the California Current in response to high CO_{2}. Marine Ecology Progress Series, 468:31–42, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/468/m468p031.pdf.

Lefebvre:2012:IHR

[LMLB12] S. Lemieux Lefebvre, R. Michaud, V. Lesage, and D. Berteaux. Identifying high residency areas of the threatened St. Lawrence beluga whale from fine-scale movements of individuals and coarse-scale movements of herds. Marine Ecology Progress Series, 450:243–257, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m450p243.pdf.

Levine:2012:DPN

[LMP12] M. R. Levine and G. Muller-Parker. Distribution patterns and nutritional contributions of algal symbionts in the sea anemone Anthopleura xanthogrammica. Marine Ecology Progress Series, 453:79–94, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/453/m453p079.pdf.

Lesage:2010:SIT

[LMR^{+}10] V. Lesage, Y. Morin, È. Rioux, C. Pomerleau, S. H. Ferguson, and É. Pelletier. Stable isotopes and trace elements as indicators of diet and habitat use in cetaceans: predicting errors related to preservation, lipid extraction, and lipid normalization. Marine Ecology Progress Series, 419:249–265, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m419p249.pdf.

Langlois:2012:DBR

[LMR^{+}12] R. J. Langlois, M. M. Mills, C. Ridame, P. Croot, and J. LaRoche. Diazotrophic bacteria respond to Saharan dust additions. Marine Ecology Progress Series, 470:1–14, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Larkin:2017:MBA

[LMRRB17] P. D. Larkin, T. J. Maloney, S. Rubiano-Rincon, and M. M. Barrett. A map-based approach to assessing genetic diversity, structure, and connectivity in the seagrass Halodule wrightii. Marine Ecology Progress Series, 567:95–107, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/567/
m567p095.pdf.

Lopez-Mendilaharsu:2019:SDJ

[LMSC^{+}19] M. Lopez-Mendilaharsu, G. Sales, R. Coluchi, M. Â. Marcovaldi, and B. Giffoni. At-sea distribution of juvenile leatherback turtles: new insights from bycatch data in the Atlantic Ocean. Marine Ecology Progress Series, 621:199– 208, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/621/m621p199.pdf.

Lowe:2015:HES

[LMT15] M. L. Lowe, M. A. Morrison, and R. B. Taylor. Harmful effects of sediment-induced turbidity on juvenile fish in estuaries. Marine Ecology Progress Series, 539:241–254, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/539/
m539p241.pdf.

Lowen:2016:ESR

[LMTD16] J. B. Lowen, C. W. McKindsey, T. W. Therriault, and C. DiBacco. Effects of spatial resolution on predicting the distribution of aquatic invasive species in nearshore marine environments. Marine Ecology Progress Series, 556:17–30, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/556/m556p017.pdf.

Lemoine:2019:RUP

[LMV^{+}19] M. Lemoine, T. Moens, A. M. Vafeiadou, L. A. V. Bezerra, and P. Lana. Resource utilization of puffer fish in a subtropical bay as revealed by stable isotope analysis and food web modeling. Marine Ecology Progress Series, 626:161–175, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/626/
m626p161.pdf.

Lawson:2012:ESS

[LMW12] S. E. Lawson, K. J. McGlathery, and P. L. Wiberg. Enhancement of sediment suspension and nutrient flux by benthic macrophytes at low biomass. Marine Ecology Progress Series, 448:259–270, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m448p259.pdf.

Lavery:2013:RDO

[LMW^{+}13] P. S. Lavery, K. McMahon, J. Weyers, M. C. Boyce, and C. E. Oldham. Release of dissolved organic carbon from seagrass wrack and its implications for trophic connectivity. Marine Ecology Progress Series, 494:121–133, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/494/
m494p121.pdf.

Long:2016:ELO

[LMZ16] M. H. Long, T. A. Mooney, and C. Zakroff. Extreme low oxygen and decreased pH conditions naturally occur within developing squid egg capsules. Marine Ecology Progress Series, 550:111–119, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/550/m550p111.pdf.

Lucey:2010:SSA

[LN10] S. M. Lucey and J. A. Nye. Shifting species assemblages in the Northeast US continental shelf large marine ecosystem. Marine Ecology Progress Series, 415:23–33, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/415/
m415p023.pdf.

Laviale:2011:RBP

[LN11] M. Laviale and J. Neveux. Relationships between pigment ratios and growth irradiance in 11 marine phytoplankton species. Marine Ecology Progress Series, 425:63–77, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/425/
m425p063.pdf.

Louzao:2011:ECP

[LNF^{+}11] M. Louzao, J. Navarro, M. G. Forero, J. M. Igual, M. Genovart, K. A. Hobson, and D. Oro. Exploiting the closest productive area: geographical segregation of foraging grounds in a critically endangered seabird. Marine Ecology Progress Series, 429:291–301, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/429/m429p291.pdf.

LeGouvello:2017:IPP

[LNH^{+}17] D. Z. M. Le Gouvello, R. Nel, L. R. Harris, K. Bezuidenhout, and S. Woodborne. Identifying potential pathways for turtle-derived nutrients cycling through beach ecosystems. Marine Ecology Progress Series, 583:49–62, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/583/
m583p049.pdf.

Liu:2016:RPP

[LNIH16] Q. Liu, N. Nishibori, I. Imai, and J. T. Hollibaugh. Response of polyamine pools in marine phytoplankton to nutrient limitation and variation in temperature and salinity. Marine Ecology Progress Series, 544:93–105, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m544p093.pdf.

Lovvorn:2016:PEC

[LNK^{+}16] J. R. Lovvorn, C. A. North, J. M. Kolts, J. M. Grebmeier, L. W. Cooper, and X. Cui. Projecting the effects of climatedriven changes in organic matter supply on benthic food webs in the northern Bering Sea. Marine Ecology Progress Series, 548:11–30, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/548/m548p011.pdf.

Ljungberg:2012:PSJ

[LNP12] P. Ljungberg, P. A. Nilsson, and A. Persson. Prey selectivity by juvenile Atlantic cod Gadus morhua in three coastal habitat types. Marine Ecology Progress Series, 466:215– 223, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/466/m466p215.pdf.

Lam:2010:ISS

[LNS10] C. H. Lam, A. Nielsen, and J. R. Sibert. Incorporating sea-surface temperature to the light-based geolocation model TrackIt. Marine Ecology Progress Series, 419:71–84, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m419p071.pdf.

Lezama-Ochoa:2011:SPS

[LOBW^{+}11] A. Lezama-Ochoa, M. Ballón, M. Woillez, D. Grados, X. Irigoien, and A. Bertrand. Spatial patterns and scaledependent relationships between macrozooplankton and fish in the Bay of Biscay: an acoustic study. Marine Ecology Progress Series, 439:151–168, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/439/m439p151.pdf.

Lund:2011:LHF

[LOES11] M. B. Lund, E. M. Olsen, S. H. Espeland, and N. C. Stenseth. Life history of fjord cod from the Skagerrak in the mid-2000s compared to 1905. Marine Ecology Progress Series, 424:169–174, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/424/m424p169.pdf.

Lokkeborg:2011:BPM

[Løk11] S. Løkkeborg. Best practices to mitigate seabird bycatch in longline, trawl and gillnet fisheries—efficiency and practical applicability. Marine Ecology Progress Series, 435:285– 303, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m435p285.pdf.

Labrousse:2019:FOB

[LOS^{+}19] S. Labrousse, F. Orgeret, A. R. Solow, C. Barbraud, C. A. Bost, J. B. Sallée, H. Weimerskirch, and S. Jenouvrier. First odyssey beneath the sea ice of juvenile emperor penguins in East Antarctica. Marine Ecology Progress Series, 609:1–16, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Leong:2010:ERT

[LP10] W. Leong and J. R. Pawlik. Evidence of a resource trade-off between growth and chemical defenses among Caribbean coral reef sponges. Marine Ecology Progress Series, 406:71–78, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m406p071.pdf.

Loh:2012:FFN

[LP12] T. L. Loh and J. R. Pawlik. Friend or foe? No evidence that association with the sponge Mycale laevis provides a benefit to corals of the genus Montastraea. Marine Ecology Progress Series, 465:111–117, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m465p111.pdf.

Lynghammar:2016:WPM

[LPB^{+}16] A. Lynghammar, K. Præbel, S. Bhat, S. E. Fevolden, and J. S. Christiansen. Widespread physical mixing of starry ray from differentiated populations and life histories in the North Atlantic. Marine Ecology Progress Series, 562:123– 134, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/562/m562p123.pdf.

Lauriano:2011:ASA

[LPC^{+}11a] G. Lauriano, S. Panigada, P. Casale, N. Pierantonio, and G. P. Donovan. Aerial survey abundance estimates of the loggerhead sea turtle Caretta caretta in the Pelagos sanctuary, northwestern Mediterranean Sea. Marine Ecology Progress Series, 437:291–302, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/437/m437p291.pdf.

Linares:2011:RAH

[LPC11b] C. Linares, M. S. Pratchett, and D. J. Coker. Recolonisation of Acropora hyacinthus following climate-induced coral bleaching on the Great Barrier Reef. Marine Ecology Progress Series, 438:97–104, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/438/m438p097.pdf.

Lefebvre:2019:HRO

[LPC19] A. Lefebvre and E. Poisson-Caillault. High resolution overview of phytoplankton spectral groups and hydrological conditions in the eastern English Channel using unsupervised clustering. Marine Ecology Progress Series, 608:73–92, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/608/
m608p073.pdf.

Leis:2014:OFL

[LPI^{+}14] J. M. Leis, C. B. Paris, J. O. Irisson, M. N. Yerman, and U. E. Siebeck. Orientation of fish larvae in situ is consistent among locations, years and methods, but varies with time of day. Marine Ecology Progress Series, 505:193–208, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/505/
m505p193.pdf.

Long:2013:STL

[LPJ^{+}13] J. D. Long, L. Porturas, E. Jones, C. Kwan, and G. C. Trussell. Seaweed traits linked to wave exposure determine predator avoidance. Marine Ecology Progress Series, 483: 143–151, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/483/m483p143.pdf.

Liu:2016:PEA

[LPK^{+}16] D. Liu, Y. Peng, J. K. Keesing, Y. Wang, and P. Richard. Paleo-ecological analyses to assess long-term environmental effects of pearl farming in Western Australia. Marine Ecology Progress Series, 552:145–158, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/552/m552p145.pdf.

Leonard-Pingel:2019:GBR

[LPKT^{+}19] J. S. Leonard-Pingel, S. M. Kidwell, A. Tomašových, C. R. Alexander, and D. B. Cadien. Gauging benthic recovery from 20 th century pollution on the southern California continental shelf using bivalves from sediment cores. Marine Ecology Progress Series, 615:101–119, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/615/m615p101.pdf.

Lefebvre:2017:MGI

[LPSK17] L. W. Lefebvre, J. A. Provancha, D. H. Slone, and W. J. Kenworthy. Manatee grazing impacts on a mixed species seagrass bed. Marine Ecology Progress Series, 564:29–45, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/564/m564p029.pdf.

Libralato:2014:TME

[LPSL14a] S. Libralato, F. Pranovi, Konstantinos I. Stergiou, and J. S. Link. Trophodynamics in marine ecology: 70 years after Lindeman. Marine Ecology Progress Series, 512:1–7, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m512p001.pdf.

Libralato:2014:TST

[LPSL14b] Simone Libralato, Fabio Pranovi, Konstantinos Stergiou, and Jason Link. Theme section: Trophodynamics in marine ecology. Marine Ecology Progress Series, 512:1–237, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m512_ThemeSection.zip.

Legendre:2015:FBC

[LR15] L. Legendre and R. B. Rivkin. Flows of biogenic carbon within marine pelagic food webs: roles of microbial competition switches. Marine Ecology Progress Series, 521:19– 30, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/521/m521p019.pdf.

Ley:2018:UOM

[LR18] J. A. Ley and H. J. Rolls. Using otolith microchemistry to assess nursery habitat contribution and function at a fine spatial scale. Marine Ecology Progress Series, 606:151–173, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m606p151.pdf.

Leduc:2012:NBD

[LRB^{+}12a] D. Leduc, A. A. Rowden, D. A. Bowden, S. D. Nodder, P. K. Probert, C. A. Pilditch, G. C. A. Duineveld, and R. Witbaard.

Nematode beta diversity on the continental slope of New Zealand: spatial patterns and environmental drivers. Marine Ecology Progress Series, 454:37–52, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m454p037.pdf.

Leduc:2012:URB

[LRB^{+}12b] D. Leduc, A. A. Rowden, D. A. Bowden, P. K. Probert, C. A. Pilditch, and S. D. Nodder. Unimodal relationship between biomass and species richness of deep-sea nematodes: implications for the link between productivity and diversity. Marine Ecology Progress Series, 454:53–64, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/454/
m454p053.pdf.

Lagarde:2018:TEI

[LRB^{+}18] F. Lagarde, M. Richard, B. Bec, C. Roques, S. Mortreux, I. Bernard, C. Chiantella, G. Messiaen, J. B. Nadalini, M. Hori, M. Hamaguchi, S. Pouvreau, E. Roque d’Orbcastel, and R. Tremblay. Trophic environments influence size at metamorphosis and recruitment performance of Pacific oysters. Marine Ecology Progress Series, 602:135–153, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/602/
m602p135.pdf.

Loeffler:2015:EGN

[LRBS15] C. R. Loeffler, M. L. Richlen, M. E. Brandt, and T. B. Smith. Effects of grazing, nutrients, and depth on the ciguateracausing dinoflagellate Gambierdiscus in the US Virgin Islands. Marine Ecology Progress Series, 531:91–104, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/531/
m531p091.pdf.

Lasley-Rasher:2015:ITG

[LRBSJ15] R. S. Lasley-Rasher, D. C. Brady, B. E. Smith, and P. A. Jumars. It takes guts to locate elusive crustacean prey. Marine Ecology Progress Series, 538:1–12, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Lovvorn:2013:WEA

[LRC^{+}13] J. R. Lovvorn, M. F. Raisbeck, L. W. Cooper, G. A. Cutter, M. W. Miller, M. L. Brooks, J. M. Grebmeier, A. C. Matz, and C. M. Schaefer. Wintering eiders acquire exceptional Se and Cd burdens in the Bering Sea: physiological and oceanographic factors. Marine Ecology Progress Series, 489:245–261, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/489/m489p245.pdf.

Lisle:2014:ASS

[LRH14] John Lisle, Crhis Reich, and Robert Halley. Aragonite saturation states and nutrient fluxes in coral reef sediments in Biscayne National Park, FL, USA. Marine Ecology Progress Series, 509:71–85, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/509/m509p071.pdf.

Lee:2018:ECR

[LRHP18] K. A. Lee, M. Roughan, R. G. Harcourt, and V. M. Peddemors. Environmental correlates of relative abundance of potentially dangerous sharks in nearshore areas, southeastern Australia. Marine Ecology Progress Series, 599:157– 179, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/599/m599p157.pdf.

Lyons:2015:RMS

[LRK^{+}15] M. Lyons, C. Roelfsema, E. Kovacs, J. Samper-Villarreal, M. Saunders, P. Maxwell, and S. Phinn. Rapid monitoring of seagrass biomass using a simple linear modelling approach, in the field and from space. Marine Ecology Progress Series, 530:1–14, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Leclerc:2013:TVO

[LRL^{+}13] J. C. Leclerc, P. Riera, C. Leroux, L. Lévêque, and D. Davoult. Temporal variation in organic matter supply in kelp forests: linking structure to trophic functioning. Marine Ecology Progress Series, 494:87–105, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/494/m494p087.pdf.

Lyngsgaard:2014:DPP

[LRM^{+}14] M. M. Lyngsgaard, K. Richardson, S. Markager, M. H. Nielsen, M. Olesen, and J. P. A. Christensen. Deep primary production in coastal pelagic systems: importance for ecosystem functioning. Marine Ecology Progress Series, 517:15–33, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m517p015.pdf.

Lubcker:2017:LTL

[LRO^{+}17] N. Lübcker, R. R. Reisinger, W. C. Oosthuizen, P. J. N. de Bruyn, A. van Tonder, P. A. Pistorius, and M. N. Bester. Low trophic level diet of juvenile southern elephant seals Mirounga leonina from Marion Island: a stable isotope investigation using vibrissal regrowths. Marine Ecology Progress Series, 577:237–250, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/577/m577p237.pdf.

Lowe:2012:SLA

[LRQ^{+}12] A. T. Lowe, R. M. Ross, L. B. Quetin, M. Vernet, and C. H. Fritsen. Simulating larval Antarctic krill growth and condition factor during fall and winter in response to environmental variability. Marine Ecology Progress Series, 452: 27–43, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/452/m452p027.pdf.

Leandro:2010:ENA

[LRR^{+}10] L. F. Leandro, R. M. Rolland, P. B. Roth, N. Lundholm, Z. Wang, and G. J. Doucette. Exposure of the North Atlantic right whale Eubalaena glacialis to the marine algal biotoxin, domoic acid. Marine Ecology Progress Series, 398: 287–303, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/398/m398p287.pdf.

Lasley-Rasher:2011:PCH

[LRRM^{+}11] R. S. Lasley-Rasher, D. B. Rasher, Z. H. Marion, R. B. Taylor, and M. E. Hay. Predation constrains host choice for a marine mesograzer. Marine Ecology Progress Series, 434: 91–99, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/434/m434p091.pdf.

Laurel:2012:ENA

[LRS^{+}12] B. J. Laurel, C. H. Ryer, M. Spencer, P. Iseri, B. Knoth, and A. Stoner. Effects of natural and anthropogenic disturbance on polychaete worm tubes and age-0 flatfish distribution. Marine Ecology Progress Series, 466:193–203, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/466/
m466p193.pdf.

Lunt:2017:WEF

[LRS17] J. Lunt, J. Reustle, and D. L. Smee. Wave energy and flow reduce the abundance and size of benthic species on oyster reefs. Marine Ecology Progress Series, 569:25–36, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/569/
m569p025.pdf.

Luskow:2019:SCB

[LRSB19] F. Lüskow, H. U. Riisgård, V. Solovyeva, and J. R. Brewer. Seasonal changes in bacteria and phytoplankton biomass control the condition index of the demosponge Halichondria panicea in temperate Danish waters. Marine Ecology Progress Series, 608:119–132, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/608/m608p119.pdf.

Lionard:2012:CEI

[LRTLF12] M. Lionard, S. Roy, M. Tremblay-Létourneau, and G. A. Ferreyra. Combined effects of increased UV-b and temperature on the pigment-determined marine phytoplankton community of the St. Lawrence Estuary. Marine Ecology Progress Series, 445:219–234, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/445/m445p219.pdf.

Lewandowska:2010:CCS

[LS10a] A. Lewandowska and U. Sommer. Climate change and the spring bloom: a mesocosm study on the influence of light and temperature on phytoplankton and mesozooplankton. Marine Ecology Progress Series, 405:101–111, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/405/
m405p101.pdf.

Lillis:2010:NBH

[LS10b] A. Lillis and P. V. R. Snelgrove. Near-bottom hydrodynamic effects on postlarval settlement in the American lobster Homarus americanus. Marine Ecology Progress Series, 401:161–172, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m401p161.pdf.

Larson:2012:RMB

[LS12] F. Larson and K. Sundbäck. Recovery of microphytobenthos and benthic functions after sediment deposition. Marine Ecology Progress Series, 446:31–44, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/446/
m446p031.pdf.

Lathlean:2014:ITM

[LS14] J. Lathlean and L. Seuront. Infrared thermography in marine ecology: methods, previous applications and future challenges. Marine Ecology Progress Series, 514:263–277, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/514/
m514p263.pdf.

Lesser:2015:PCS

[LS15] M. P. Lesser and M. Slattery. Picoplankton consumption supports the ascidian Cnemidocarpa verrucosa in McMurdo Sound, Antarctica. Marine Ecology Progress Series, 525:117– 126, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/525/m525p117.pdf.

Longtin:2016:RCS

[LS16] C. M. Longtin and G. W. Saunders. The relative contribution of Saccharina nigripes (Phaeophyceae) to the Bay of Fundy laminariaceae: spatial and temporal variability. Marine Ecology Progress Series, 543:153–162, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/543/
m543p153.pdf.

Lewis:2019:FDA

[LS19] L. S. Lewis and J. E. Smith. Functional diversity among herbivorous sea urchins on a coral reef: grazing rate, dietary preference, and metabolism. Marine Ecology Progress Series, 625:71–87, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/625/m625p071.pdf.

Ladd:2017:TSR

[LSBB17] M. C. Ladd, A. A. Shantz, E. Bartels, and D. E. Burkepile. Thermal stress reveals a genotype-specific tradeoff between growth and tissue loss in restored Acropora cervicornis. Marine Ecology Progress Series, 572:129–139, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/572/
m572p129.pdf.

Liu:2013:PAP

[LSD^{+}13] D. Liu, X. Shen, B. Di, Y. Shi, J. K. Keesing, Y. Wang, and Y. Wang. Palaeoecological analysis of phytoplankton regime shifts in response to coastal eutrophication. Marine Ecology Progress Series, 475:1–14, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Levasseur:2019:EHN

[LSFQ19] K. E. Levasseur, S. P. Stapleton, M. C. Fuller, and J. M. Quattro. Exceptionally high natal homing precision in hawksbill sea turtles to insular rookeries of the Caribbean. Marine Ecology Progress Series, 620:155–171, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/620/m620p155.pdf.

Leis:2015:LSO

[LSH^{+}15] J. M. Leis, U. E. Siebeck, A. C. Hay, C. B. Paris, O. Chateau, and L. Wantiez. In situ orientation of fish larvae can vary among regions. Marine Ecology Progress Series, 537:191– 203, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/537/m537p191.pdf.

Lane:2019:EWT

[LSH19a] J. V. Lane, D. V. Spracklen, and K. C. Hamer. Effects of windscape on three-dimensional foraging behaviour in a wideranging marine predator, the northern gannet. Marine Ecology Progress Series, 628:183–193, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m628p183.pdf.

Lee:2019:EDA

[LSH^{+}19b] K. A. Lee, A. F. Smoothey, R. G. Harcourt, M. Roughan, P. A. Butcher, and V. M. Peddemors. Environmental drivers of abundance and residency of a large migratory shark, Carcharhinus leucas, inshore of a dynamic western boundary current. Marine Ecology Progress Series, 622:121–137, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2019/622/m622p121.pdf.

Langaard:2015:TIH

[LSJ^{+}15] L. Langård, G. Skaret, K. H. Jensen, A. Johannessen, A. Slotte, L. Nøttestad, and A. Fernö. Tracking individual herring within a semi-enclosed coastal marine ecosystem: 3-dimensional dynamics from pre- to post-spawning. Marine Ecology Progress Series, 518:267–279, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/518/
m518p267.pdf.

Lamb:2017:DCP

[LSJ17] J. S. Lamb, Y. G. Satgé, and P. G. R. Jodice. Diet composition and provisioning rates of nestlings determine reproductive success in a subtropical seabird. Marine Ecology Progress Series, 581:149–164, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m581p149.pdf.

LaRue:2019:PEF

[LSN^{+}19] M. A. LaRue, L. Salas, N. Nur, D. G. Ainley, S. Stammerjohn, L. Barrington, K. Stamatiou, J. Pennycook, M. Dozier, J. Saints, and H. Nakamura. Physical and ecological factors explain the distribution of Ross Sea Weddell seals during the breeding season. Marine Ecology Progress Series, 612:193–208, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/612/m612p193.pdf.

Leung:2013:PBR

[LSQ^{+}13] Y. S. Leung, P. K. S. Shin, J. W. Qiu, P. O. Ang, J. M. Y. Chiu, V. Thiyagarajan, and S. G. Cheung. Physiological and behavioural responses of different life stages of a serpulid polychaete to hypoxia. Marine Ecology Progress Series, 477:135–145, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/477/m477p135.pdf.

Lopez-Sandoval:2013:EOC

[LSRRCM13] D. C. López-Sandoval, T. Rodríguez-Ramos, P. Cermeño, and E. Marañón. Exudation of organic carbon by marine phytoplankton: dependence on taxon and cell size. Marine Ecology Progress Series, 477:53–60, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/477/m477p053.pdf.

Large:2011:ECI

[LST11] S. I. Large, D. L. Smee, and G. C. Trussell. Environmental conditions influence the frequency of prey responses to predation risk. Marine Ecology Progress Series, 422:41– 49, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/422/m422p041.pdf.

Latasa:2010:PPG

[LSV^{+}10a] M. Latasa, R. Scharek, M. Vidal, G. Vila-Reixach, A. Gutiérrez-Rodríguez, M. Emelianov, and J. M. Gasol. Preferences of phytoplankton groups for waters of different trophic status in the northwestern Mediterranean Sea. Marine Ecology Progress Series, 407:27–42, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/407/m407p027.pdf.

Lyons:2010:STV

[LSV10b] D. A. Lyons, R. E. Scheibling, and K. L. Van Alstyne. Spatial and temporal variation in DMSP content in the invasive seaweed Codium fragile ssp. fragile: effects of temperature, light and grazing. Marine Ecology Progress Series, 417:51–61, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/417/m417p051.pdf.

Landry:2011:DPG

[LSY11] M. R. Landry, K. E. Selph, and E. J. Yang. Decoupled phytoplankton growth and microzooplankton grazing in the deep euphotic zone of the eastern equatorial Pacific. Marine Ecology Progress Series, 421:13–24, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/421/m421p013.pdf.

La:2018:FED

[LSY^{+}18] H. S. La, K. Shimada, E. J. Yang, K. H. Cho, S. Y. Ha, J. Jung, J. O. Min, S. H. Kang, and H. K. Ha. Further evidence of diel vertical migration of copepods under Arctic sea ice during summer. Marine Ecology Progress Series, 592:283–289, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/592/m592p283.pdf.

Lundquist:2010:IBD

[LTCH10] C. J. Lundquist, S. F. Thrush, G. Coco, and J. E. Hewitt. Interactions between disturbance and dispersal reduce persistence thresholds in a benthic community. Marine Ecology Progress Series, 413:217–228, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m413p217.pdf.

Lifavi:2017:EDC

[LTG17] D. M. Lifavi, T. E. Targett, and P. A. Grecay. Effects of diel-cycling hypoxia and acidification on juvenile weakfish Cynoscion regalis growth, survival, and activity. Marine Ecology Progress Series, 564:163–174, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/564/
m564p163.pdf.

Longmore:2011:OGI

[LTN^{+}11] C. Longmore, C. N. Trueman, F. Neat, E. J. O’Gorman, J. A. Milton, and S. Mariani. Otolith geochemistry indicates life-long spatial population structuring in a deep-sea fish, Coryphaenoides rupestris. Marine Ecology Progress Series, 435:209–224, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/435/m435p209.pdf.

LaValle:2019:MBG

[LTN19] F. F. La Valle, F. I. Thomas, and C. E. Nelson. Macroalgal biomass, growth rates, and diversity are influenced by submarine groundwater discharge and local hydrodynamics in tropical reefs. Marine Ecology Progress Series, 621:51–67, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m621p051.pdf.

Long:2013:PUC

[LTP13] J. D. Long, G. Toth, and H. Pavia. Proximate and ultimate causes for transatlantic variation in seaweed defenses. Marine Ecology Progress Series, 493:83–89, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/493/
m493p083.pdf.

Lauerburg:2018:FFC

[LTP^{+}18] R. A. M. Lauerburg, A. Temming, J. K. Pinnegar, P. Kotterba, A. F. Sell, A. Kempf, and J. Floeter. Forage fish control population dynamics of North Sea whiting Merlangius merlangus. Marine Ecology Progress Series, 594:213–230, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/594/
m594p213.pdf.

Lehette:2012:KEE

[LTSDHL12] P. Lehette, A. Tovar-Sánchez, C. M. Duarte, and S. Hernández-León. Krill excretion and its effect on primary production. Marine Ecology Progress Series, 459:29–38, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/459/
m459p029.pdf.

Lenes:2013:SCD

[LWD13] J. M. Lenes, J. J. Walsh, and B. P. Darrow. Simulating cell death in the termination of Karenia brevis blooms: implications for predicting aerosol toxicity vectors to humans.

Marine Ecology Progress Series, 493:71–81, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/493/
m493p071.pdf.

Litvin:2014:HUP

[LWG14] S. Y. Litvin, M. P. Weinstein, and V. G. Guida. Habitat utilization patterns determine the physiological condition of Cynoscion regalis during estuarine residency. Marine Ecology Progress Series, 510:87–99, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/510/m510p087.pdf.

Lagos:2015:ALO

[LWM15] M. E. Lagos, C. R. White, and D. J. Marshall. Avoiding low-oxygen environments: oxytaxis as a mechanism of habitat selection in a marine invertebrate. Marine Ecology Progress Series, 540:99–107, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/540/m540p099.pdf.

Lecchini:2013:FLP

[LWP^{+}13] D. Lecchini, V. P. Waqalevu, E. Parmentier, C. A. Radford, and B. Banaigs. Fish larvae prefer coral over algal water cues: implications of coral reef degradation. Marine Ecology Progress Series, 475:303–307, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/475/m475p303.pdf.

Leu:2010:IIR

[LWS^{+}10] E. Leu, J. Wiktor, J. E. Søreide, J. Berge, and S. Falk-Petersen. Increased irradiance reduces food quality of sea ice algae. Marine Ecology Progress Series, 411:49–60, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m411p049.pdf.

Logan:2017:CUV

[LYH^{+}17] J. M. Logan, M. A. Young, E. S. Harvey, A. C. G. Schimel, and D. Ierodiaconou. Combining underwater video methods improves effectiveness of demersal fish assemblage surveys across habitats. Marine Ecology Progress Series, 582: 181–200, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/582/m582p181.pdf.

Lyons:2014:BDA

[Lyo14] P. J. Lyons. Behavioral differences among mutualist species in a shrimp–goby association. Marine Ecology Progress Series, 510:101–106, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/510/m510p101.pdf.

Landes:2012:AWA

[LZ12] A. Landes and M. Zimmer. Acidification and warming affect both a calcifying predator and prey, but not their interaction. Marine Ecology Progress Series, 450:1–10, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Li:2014:MDD

[LZF^{+}14] J. Li, H. Zhou, J. Fang, Y. Sun, and S. Dasgupta. Microbial distribution in different spatial positions within the walls of a black sulfide hydrothermal chimney. Marine Ecology Progress Series, 508:67–85, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/508/m508p067.pdf.

McFarlin:2013:FDD

[MA13a] C. R. McFarlin and M. Alber. Foliar DMSO: DMSP ratio and metal content as indicators of stress in Spartina alterniflora. Marine Ecology Progress Series, 474:1–13, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Morton:2013:SPI

[MA13b] D. N. Morton and T. W. Anderson. Spatial patterns of invertebrate settlement in giant kelp forests. Marine Ecology Progress Series, 485:75–89, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/485/m485p075.pdf.

McClanahan:2014:CRI

[MA14] T. R. McClanahan and C. A. Abunge. Catch rates and income are associated with fisheries management restrictions and not an environmental disturbance, in a heavily exploited tropical fishery. Marine Ecology Progress Series, 513:201–210, 2014.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m513p201.pdf.

Montagnes:2010:RCO

[MAB^{+}10] D. J. S. Montagnes, J. Allen, L. Brown, C. Bulit, R. Davidson, S. Fielding, M. Heath, N. P. Holliday, J. Rasmussen, R. Sanders, J. J. Waniek, and D. Wilson. Role of ciliates and other microzooplankton in the Irminger Sea (NW Atlantic Ocean). Marine Ecology Progress Series, 411:101– 115, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/411/m411p101.pdf.

McIntyre:2012:ESF

[MAB^{+}12a] T. McIntyre, I. J. Ansorge, H. Bornemann, J. Plötz, C. A. Tosh, and M. N. Bester. Elephant seal foraging dives do indeed track prey distribution, but temperature influences the distribution of prey: Reply to boersch-supan et al. (2012). Marine Ecology Progress Series, 461:299–303, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/461/
m461p299.pdf.

Milligan:2012:FNQ

[MAB12b] A. J. Milligan, U. A. Aparicio, and M. J. Behrenfeld. Fluorescence and nonphotochemical quenching responses to simulated vertical mixing in the marine diatom Thalassiosira weissflogii. Marine Ecology Progress Series, 448:67–78, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/448/
m448p067.pdf.

Martinez:2017:DIE

[MAD17] G. Martínez, M. Arim, and O. Defeo. Distribution of the isopod Excirolana braziliensis on sandy beaches of the Atlantic and Pacific Oceans. Marine Ecology Progress Series, 583:137–148, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m583p137.pdf.

Martinez:2018:GTS

[MAD18] G. Martínez, M. Arim, and O. Defeo. Genetics, taxonomy and species complex in sandy beach macrofauna: Reply to mattos et al. (2018). Marine Ecology Progress Series, 601:273–275, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m601p273.pdf.

McKnight:2013:SSP

[MADI13] A. McKnight, A. J. Allyn, D. C. Duffy, and D. B. Irons. ‘stepping stone’ pattern in Pacific Arctic tern migration reveals the importance of upwelling areas. Marine Ecology Progress Series, 491:253–264, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m491p253.pdf.

Moriarty:2012:VHM

[MAHK12] P. E. Moriarty, K. S. Andrews, C. J. Harvey, and M. Kawase. Vertical and horizontal movement patterns of scyphozoan jellyfish in a fjord-like estuary. Marine Ecology Progress Series, 455:1–12, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Miller:2015:WMD

[MAK^{+}15] L. P. Miller, B. J. Allen, F. A. King, D. R. Chilin, V. M. Reynoso, and M. W. Denny. Warm microhabitats drive both increased respiration and growth rates of intertidal consumers. Marine Ecology Progress Series, 522:127–143, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/522/
m522p127.pdf.

Murphy:2015:ENR

[MAL15] A. E. Murphy, I. C. Anderson, and M. W. Luckenbach. Enhanced nutrient regeneration at commercial hard clam (Mercenaria mercenaria) beds and the role of macroalgae. Marine Ecology Progress Series, 530:135–151, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/530/
m530p135.pdf.

McGinley:2012:SSC

[MAP^{+}12] M. P. McGinley, M. D. Aschaffenburg, D. T. Pettay, R. T. Smith, T. C. LaJeunesse, and M. E. Warner. Symbiodinium spp. in colonies of eastern Pacific Pocillopora spp. are highly stable despite the prevalence of low-abundance background populations. Marine Ecology Progress Series, 462:1–7, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Martin:2017:AOC

[Mar17] C. W. Martin. Avoidance of oil contaminated sediments by estuarine fishes. Marine Ecology Progress Series, 576:125– 134, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m576p125.pdf.

Martone:2010:BIC

[MAS10a] P. T. Martone, M. Alyono, and S. Stites. Bleaching of an intertidal coralline alga: untangling the effects of light, temperature, and desiccation. Marine Ecology Progress Series, 416:57–67, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/416/m416p057.pdf.

Meyer:2010:SVB

[MAS^{+}10b] B. Meyer, L. Auerswald, V. Siegel, S. Spahić, C. Pape, B. A. Fach, M. Teschke, A. L. Lopata, and V. Fuentes. Seasonal variation in body composition, metabolic activity, feeding, and growth of adult krill Euphausia superba in the Lazarev Sea. Marine Ecology Progress Series, 398:1–18, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Martinez-Alarcon:2019:SLD

[MASMH19] D. Martínez-Alarcón, R. Saborowski, E. Melis, and W. Hagen. Seasonal lipid dynamics of the shrimps Crangon crangon and Pandalus montagui in the German Bight (North Sea). Marine Ecology Progress Series, 625:41–52, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/625/
m625p041.pdf.

Mangel:2011:PCM

[MASW^{+}11] J. C. Mangel, J. Alfaro-Shigueto, M. J. Witt, P. H. Dutton, J. A. Seminoff, and B. J. Godley. Post-capture movements of loggerhead turtles in the southeastern Pacific Ocean assessed by satellite tracking. Marine Ecology Progress Series, 433:261–272, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/433/m433p261.pdf.

Matassa:2010:PSU

[Mat10a] C. M. Matassa. Purple sea urchins Strongylocentrotus purpuratus reduce grazing rates in response to risk cues from the spiny lobster Panulirus interruptus. Marine Ecology Progress Series, 400:283–288, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/400/m400p283.pdf.

Matsumoto:2010:ELS

[Mat10b] A. K. Matsumoto. Estimation of in situ distribution of carbonate produced from cold-water octocorals on a Japanese seamount in the NW Pacific. Marine Ecology Progress Series, 399:81–102, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m399p081.pdf.

Martin:2011:RSI

[MAT^{+}11] A. Martin, M. J. Anderson, C. Thorn, S. K. Davy, and K. G. Ryan. Response of sea-ice microbial communities to environmental disturbance: an in situ transplant experiment in the Antarctic. Marine Ecology Progress Series, 424: 25–37, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/424/m424p025.pdf.

Martinez-Abrain:2019:DPP

[MATM^{+}19] A. Martínez-Abraín, P. Santidrián Tomillo, J. Mouriño, S. Tenan, and D. Oro. Delayed predator–prey collapses: the case of black-legged kittiwakes and Iberian sardines. Marine Ecology Progress Series, 631:201–207, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/631/
m631p201.pdf.

Manzanera:2011:RPO

[MATR11] M. Manzanera, T. Alcoverro, F. Tomas, and J. Romero. Response of Posidonia oceanica to burial dynamics. Marine Ecology Progress Series, 423:47–56, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/423/
m423p047.pdf.

Manabe:2011:FOO

[MAW^{+}11] R. Manabe, J. Aoyama, K. Watanabe, M. Kawai, M. J. Miller, and K. Tsukamoto. First observations of the oceanic migration of Japanese eel, from pop-up archival transmitting tags. Marine Ecology Progress Series, 437:229–240, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/437/
m437p229.pdf.

Morishita:2011:BSU

[MB11] V. R. Morishita and R. E. Barreto. Black sea urchins evaluate predation risk using chemical signals from a predator and injured con- and heterospecific prey. Marine Ecology Progress Series, 435:173–181, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2011/435/m435p173.pdf.

Malek:2016:PEH

[MB16] J. C. Malek and J. E. Byers. Predator effects on host–parasite interactions in the eastern oyster Crassostrea virginica. Marine Ecology Progress Series, 556:131–141, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/556/
m556p131.pdf.

Minello:2018:ETF

[MB18] T. J. Minello and M. C. Benfield. Effects of turbidity on feeding of southern flounder on estuarine prey. Marine Ecology Progress Series, 594:203–212, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/594/m594p203.pdf.

Macreadie:2011:WSI

[MBB11] P. I. Macreadie, M. J. Bishop, and D. J. Booth. As we see it; implications of climate change for macrophytic rafts and their hitchhikers. Marine Ecology Progress Series, 443: 285–292, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/443/m443p285.pdf.

McCoy:2012:PGS

[MBB^{+}12a] K. D. McCoy, P. Beis, A. Barbosa, J. J. Cuervo, W. R. Fraser, J. González-Solís, E. Jourdain, M. Poisbleau, P. Quillfeldt, E. Léger, and M. Dietrich. Population genetic structure and colonisation of the western Antarctic Peninsula by the seabird tick Ixodes uriae. Marine Ecology Progress Series, 459:109–120, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m459p109.pdf.

Miles:2012:VTP

[MBB^{+}12b] A. K. Miles, L. Bowen, B. Ballachey, J. L. Bodkin, M. Murray, J. L. Estes, R. A. Keister, and J. Stott. Variations of transcript profiles between sea otters Enhydra lutris from Prince William Sound, Alaska, and clinically normal reference otters. Marine Ecology Progress Series, 451:201–212, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/451/
m451p201.pdf.

Madigan:2015:DSS

[MBB^{+}15] D. J. Madigan, E. J. Brooks, M. E. Bond, J. Gelsleichter, L. A. Howey, D. L. Abercrombie, A. Brooks, and D. D. Chapman. Diet shift and site-fidelity of oceanic whitetip sharks Carcharhinus longimanus along the Great Bahama Bank. Marine Ecology Progress Series, 529:185–197, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m529p185.pdf.

Madin:2018:CEC

[MBB^{+}18] J. S. Madin, A. H. Baird, T. C. L. Bridge, S. R. Connolly, K. J. A. Zawada, and M. Dornelas. Cumulative effects of cyclones and bleaching on coral cover and species richness at Lizard Island. Marine Ecology Progress Series, 604:263–268, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/604/m604p263.pdf.

Maggi:2012:CAM

[MBBBC12] E. Maggi, F. Bulleri, I. Bertocci, and L. Benedetti-Cecchi. Competitive ability of macroalgal canopies overwhelms the effects of variable regimes of disturbance. Marine Ecology Progress Series, 465:99–109, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m465p099.pdf.

Miller:2016:AIS

[MBBK16] S. H. Miller, D. L. Breitburg, R. B. Burrell, and A. G. Keppel. Acidification increases sensitivity to hypoxia in important forage fishes. Marine Ecology Progress Series, 549:1–8, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Mollmann:2011:BEB

[MBC^{+}11] C. Möllmann, T. Blenckner, M. Casini, A. Gårdmark, and M. Lindegren. Beauty is in the eye of the beholder: management of Baltic cod stock requires an ecosystem approach. Marine Ecology Progress Series, 431:293–297, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/431/
m431p293.pdf.

Mendez:2017:GVF

[MBC^{+}17] L. Mendez, P. Borsa, S. Cruz, S. de Grissac, J. Hennicke, J. Lallemand, A. Prudor, and H. Weimerskirch. Geographical variation in the foraging behaviour of the pantropical red-footed booby. Marine Ecology Progress Series, 568:217– 230, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m568p217.pdf.

Maggi:2018:TCS

[MBC18] E. Maggi and L. Benedetti-Cecchi. Trophic compensation stabilizes marine primary producers exposed to artificial light at night. Marine Ecology Progress Series, 606:1–5, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Mieszkowska:2019:OPM

[MBCB^{+}19] N. Mieszkowska, L. Benedetti-Cecchi, M. T. Burrows, M. C. Mangano, A. Queirós, L. Seuront, and G. Sarà. Opinion piece: Multinational, integrated approaches to forecasting and managing the impacts of climate change on intertidal species. Marine Ecology Progress Series, 613:247–252, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m613p247.pdf.

Morfin:2016:HTD

[MBF16] M. Morfin, N. Bez, and J. M. Fromentin. Habitats of ten demersal species in the Gulf of Lions and potential implications for spatial management. Marine Ecology Progress Series, 547:219–232, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/547/m547p219.pdf.

Moore:2015:MPA

[MBG^{+}15] M. E. Moore, B. A. Berejikian, F. A. Goetz, A. G. Berger, S. S. Hodgson, E. J. Connor, and T. P. Quinn. Multi-population analysis of Puget Sound steelhead survival and migration behavior. Marine Ecology Progress Series, 537:217–232, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m537p217.pdf.

Maire:2016:HDM

[MBG^{+}16] O. Maire, C. Barras, T. Gestin, M. P. Nardelli, A. Romero-Ramirez, J. C. Duchêne, and E. Geslin. How does macrofaunal bioturbation influence the vertical distribution of living benthic Foraminifera? Marine Ecology Progress Series, 561:83–97, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/561/m561p083.pdf.

MacDonald:2016:DBP

[MBJ16] C. MacDonald, T. C. L. Bridge, and G. P. Jones. Depth, bay position and habitat structure as determinants of coral reef fish distributions: Are deep reefs a potential refuge? Marine Ecology Progress Series, 561:217–231, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/561/
m561p217.pdf.

Monczak:2017:LTA

[MBKM17] A. Monczak, A. Berry, C. Kehrer, and E. W. Montie. Longterm acoustic monitoring of fish calling provides baseline estimates of reproductive timelines in the May River estuary, southeastern USA. Marine Ecology Progress Series, 581:1–19, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Macdonald:2014:VMM

[MBM^{+}14] R. G. Macdonald, D. G. Bowers, D. McKee, G. W. Graham, and W. A. M. Nimmo-Smith. Vertical migration maintains phytoplankton position in a tidal channel with residual flow. Marine Ecology Progress Series, 509:113–126, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/509/
m509p113.pdf.

Miller:2012:BSC

[MBP^{+}12] C. A. Miller, P. B. Best, W. L. Perryman, M. F. Baumgartner, and M. J. Moore. Body shape changes associated with reproductive status, nutritive condition and growth in right whales Eubalaena glacialis and E. australis. Marine Ecology Progress Series, 459:135–156, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m459p135.pdf.

Majaneva:2013:APP

[MBR^{+}13] S. Majaneva, J. Berge, P. E. Renaud, A. Vader, E. Stübner, A. M. Rao, Ø. Sparre, and M. Lehtiniemi. Aggregations of predators and prey affect predation impact of the Arctic ctenophore Mertensia ovum. Marine Ecology Progress Series, 476:87–100, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/476/m476p087.pdf.

McGinty:2018:TSC

[MBR^{+}18] N. McGinty, A. D. Barton, N. R. Record, Z. V. Finkel, and A. J. Irwin. Traits structure copepod niches in the North Atlantic and Southern Ocean. Marine Ecology Progress Series, 601:109–126, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/601/m601p109.pdf.

Miller:2010:PDS

[MBRO10] T. W. Miller, R. D. Brodeur, G. Rau, and K. Omori. Prey dominance shapes trophic structure of the northern California Current pelagic food web: evidence from stable isotopes and diet analysis. Marine Ecology Progress Series, 420:15–26, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m420p015.pdf.

Miller:2013:CPH

[MBS^{+}13] Todd W. Miller, Keith L. Bosley, Junya Shibata, Richard D. Brodeur, Koji Omori, and Robert Emmett. Contribution of prey to Humboldt squid Dosidicus gigas in the northern California Current, revealed by stable isotope analyses. Marine Ecology Progress Series, 477:123–134, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/477/

m477p123.pdf. See comment [FLC^{+}14] and reply [MBS^{+}14].

Miller:2014:UMM

[MBS^{+}14] T. W. Miller, K. L. Bosley, J. Shibata, R. D. Brodeur, K. Omori, and R. Emmett. Use of mixing models for Humboldt squid diet analysis: Reply to Field et al. (2014). Marine Ecology Progress Series, 500:287–290, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/500/

m500p287.pdf. See [MBS^{+}13, FLC^{+}14].

McGovern:2018:HFW

[MBS^{+}18] M. McGovern, J. Berge, B. Szymczycha, J. M. Wesławski, and P. E. Renaud. Hyperbenthic food-web structure in an Arctic fjord. Marine Ecology Progress Series, 603:29–46, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m603p029.pdf.

Martin:2011:SCN

[MBSG11] C. Martin, I. Bentaleb, S. Steelandt, and C. Guinet. Stable carbon and nitrogen isotope variations in canine dentine growth layers of Kerguelen southern elephant seals. Marine Ecology Progress Series, 439:295–305, 2011. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m439p295.pdf.

Morrow:2014:VGC

[MBT14] K. Morrow, S. S. Bell, and A. Tewfik. Variation in ghost crab trophic links on sandy beaches. Marine Ecology Progress Series, 502:197–206, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2014/502/m502p197.pdf.

Modica:2011:PFS

[MBV^{+}11] L. Modica, A. Bozzano, F. Velasco, G. Albertelli, and I. Olaso. Predation, feeding strategy and food daily ration in juvenile European hake. Marine Ecology Progress Series, 440: 177–189, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/440/m440p177.pdf.

Miller:2018:CGS

[MBvO^{+}18] K. J. Miller, H. P. Baird, J. van Oosterom, J. Mondon, and C. K. King. Complex genetic structure revealed in the circum-Antarctic broadcast spawning sea urchin Sterechinus neumayeri. Marine Ecology Progress Series, 601:153–166, 2018. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m601p153.pdf.

Macdonald:2012:FFC

[MBvR12] T. A. Macdonald, B. J. Burd, and A. van Roodselaar. Facultative feeding and consistency of trophic structure in marine soft-bottom macrobenthic communities. Marine Ecology Progress Series, 445:129–140, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/445/m445p129.pdf.

Matta:2010:CDS

[MBW10] E. M. Matta, B. A. Black, and T. K. Wilderbuer. Climatedriven synchrony in otolith growth-increment chronologies for three Bering Sea flatfish species. Marine Ecology Progress Series, 413:137–145, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/413/m413p137.pdf.

Mienis:2019:EAE

[MBW^{+}19] F. Mienis, T. J. Bouma, R. Witbaard, D. van Oevelen, and G. C. A. Duineveld. Experimental assessment of the effects of coldwater coral patches on water flow. Marine Ecology Progress Series, 609:101–117, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/609/m609p101.pdf.

Macdonald:2010:VSC

[MC10a] J. I. Macdonald and D. A. Crook. Variability in Sr: Ca and Ba:Ca ratios in water and fish otoliths across an estuarine salinity gradient. Marine Ecology Progress Series, 413:147–161, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/413/m413p147.pdf.

Manriquez:2010:FEG

[MC10b] P. H. Manríquez and J. C. Castilla. Fertilization efficiency and gamete viability in the ascidian Pyura praeputialis in Chile. Marine Ecology Progress Series, 409:107–119, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/409/
m409p107.pdf.

Manriquez:2011:BTC

[MC11] P. H. Manríquez and J. C. Castilla. Behavioural traits of competent Concholepas concholepas (loco) larvae. Marine Ecology Progress Series, 430:207–221, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m430p207.pdf.

Major:2012:ECB

[MC12a] H. L. Major and A. M. Chubaty. Estimating colony and breeding population size for nocturnal burrow-nesting seabirds. Marine Ecology Progress Series, 454:83–90, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/454/
m454p083.pdf.

Merot:2012:EFA

[MC12b] C. Mérot and R. Collin. Effects of food availability on sex change in two species of Crepidula (Gastropoda: Calyptraeidae). Marine Ecology Progress Series, 449:173–181, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2012/449/m449p173.pdf.

McDuie:2016:TEM

[MC16] F. McDuie and B. C. Congdon. Trans-equatorial migration and non-breeding habitat of tropical shearwaters: implications for modelling pelagic important bird areas. Marine Ecology Progress Series, 550:219–234, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/550/m550p219.pdf.

McAfee:2018:IDT

[MCBR18] D. McAfee, V. R. Cumbo, M. J. Bishop, and D. A. Raftos. Intraspecific differences in the transcriptional stress response of two populations of Sydney rock oyster increase with rising temperatures. Marine Ecology Progress Series, 589:115– 127, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/589/m589p115.pdf.

McConnaughey:2012:ZOC

[McC12] T. A. McConnaughey. Zooxanthellae that open calcium channels: implications for reef corals. Marine Ecology Progress Series, 460:277–287, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/460/m460p277.pdf.

McClanahan:2014:RFG

[McC14] T. R. McClanahan. Recovery of functional groups and trophic relationships in tropical fisheries closures. Marine Ecology Progress Series, 497:13–23, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m497p013.pdf.

McMahan:2016:GJA

[MCC^{+}16] M. D. McMahan, D. F. Cowan, Y. Chen, G. D. Sherwood, and J. H. Grabowski. Growth of juvenile American lobster Homarus americanus in a changing environment. Marine Ecology Progress Series, 557:177–187, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/557/
m557p177.pdf.

McClanahan:2017:CCS

[McC17] T. R. McClanahan. Changes in coral sensitivity to thermal anomalies. Marine Ecology Progress Series, 570:71–85, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m570p071.pdf.

McClanahan:2019:CRF

[McC19] T. R. McClanahan. Coral reef fish communities, diversity, and their fisheries and biodiversity status in East Africa. Marine Ecology Progress Series, 632:175–191, 2019. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m632p175.pdf.

McDonald:2013:GVM

[McD13] J. H. McDonald. Geographic variation in Megalorchestia californiana allele frequencies may be caused by winter rather than summer temperatures. Marine Ecology Progress Series, 488:201–207, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/488/m488p201.pdf.

Mourino-Carballido:2016:NSC

[MCHC^{+}16] B. Mouriño-Carballido, E. Hojas, P. Cermeño, P. Chouciño, B. Fernández-Castro, M. Latasa, E. Marañón, X. A. G. Morán, and M. Vidal. Nutrient supply controls picoplankton community structure during three contrasting seasons in the northwestern Mediterranean Sea. Marine Ecology Progress Series, 543:1–19, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Moody:2013:ESE

[MCK^{+}13] R. M. Moody, J. Cebrian, S. M. Kerner, K. L. Heck, Jr., S. P. Powers, and C. Ferraro. Effects of shoreline erosion on salt-marsh floral zonation. Marine Ecology Progress Series, 488:145–155, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m488p145.pdf.

Monteiro:2017:SGS

[MCL^{+}17] S. S. Monteiro, F. Caurant, A. López, J. Cedeira, M. Ferreira, J. V. Vingada, C. Eira, and P. Méndez-Fernandez. Sympatric Globicephala species: feeding ecology and contamination status based on stable isotopes and trace elements. Marine Ecology Progress Series, 563:233–247, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/563/
m563p233.pdf.

Miller:2018:ITT

[MCP^{+}18] M. G. R. Miller, N. Carlile, J. Scutt Phillips, F. McDuie, and B. C. Congdon. Importance of tropical tuna for seabird foraging over a marine productivity gradient. Marine Ecology Progress Series, 586:233–249, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/586/m586p233.pdf.

Maggs:2019:SSS

[MCPC19] J. Q. Maggs, P. D. Cowley, S. N. Porter, and A. R. Childs. Should I stay or should I go? Intra-population variability in movement behaviour of wide-ranging and resident coastal fishes. Marine Ecology Progress Series, 619:111– 124, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/619/m619p111.pdf.

Melatunan:2013:EOA

[MCR^{+}13] S. Melatunan, P. Calosi, S. D. Rundle, S. Widdicombe, and A. J. Moody. Effects of ocean acidification and elevated temperature on shell plasticity and its energetic basis in an intertidal gastropod. Marine Ecology Progress Series, 472: 155–168, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/472/m472p155.pdf.

Muhlin:2011:MRI

[MCRB11] J. F. Muhlin, M. A. Coleman, T. A. V. Rees, and S. H. Brawley. Modeling of reproduction in the intertidal macrophyte Fucus vesiculosus and implications for spatial subsidies in the nearshore environment. Marine Ecology Progress Series, 440:79–94, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/440/m440p079.pdf.

Meade:2015:SPA

[MCS^{+}15] J. Meade, M. B. Ciaglia, D. J. Slip, J. Negrete, M. E. I. Márquez, J. Mennucci, and T. L. Rogers. Spatial patterns in activity of leopard seals Hydrurga leptonyx in relation to sea ice. Marine Ecology Progress Series, 521:265– 275, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/521/m521p265.pdf.

Munsch:2015:ESA

[MCT15] S. H. Munsch, J. R. Cordell, and J. D. Toft. Effects of seawall armoring on juvenile Pacific salmon diets in an urban estuarine embayment. Marine Ecology Progress Series, 535:213–229, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/535/m535p213.pdf.

Munsch:2016:FSH

[MCT16] S. H. Munsch, J. R. Cordell, and J. D. Toft. Fine-scale habitat use and behavior of a nearshore fish community: nursery functions, predation avoidance, and spatiotemporal habitat partitioning. Marine Ecology Progress Series, 557:1–15, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Munoz-Colmenero:2012:ELA

[MCTHGV12] M. Muñoz-Colmenero, P. Turrero, J. L. Horreo, and E. Garcia-Vazquez. Evolution of limpet assemblages driven by environmental changes and harvesting in North Iberia. Marine Ecology Progress Series, 466:121–131, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/466/
m466p121.pdf.

Matias:2012:IDR

[MCUO12] M. G. Matias, M. G. Chapman, A. J. Underwood, and N. E. O’Connor. Increasing density of rare species of intertidal gastropods: tests of competitive ability compared with common species. Marine Ecology Progress Series, 453:107–116, 2012.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m453p107.pdf.

Munoz:2011:DIL

[MCW11] R. C. Muñoz, C. A. Currin, and P. E. Whitfield. Diet of invasive lionfish on hard bottom reefs of the Southeast USA: insights from stomach contents and stable isotopes. Marine Ecology Progress Series, 432:181–193, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/432/
m432p181.pdf.

Madigan:2016:ITR

[MCW^{+}16] D. J. Madigan, W. C. Chiang, N. J. Wallsgrove, B. N. Popp, T. Kitagawa, C. A. Choy, J. Tallmon, N. Ahmed, N. S. Fisher, and C. l. Sun. Intrinsic tracers reveal recent foraging ecology of giant Pacific bluefin tuna at their primary spawning grounds. Marine Ecology Progress Series, 553:253–266, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/553/
m553p253.pdf.

Muzaffar:2017:SDD

[MCW^{+}17] S. B. Muzaffar, C. Clarke, R. Whelan, R. Gubiani, and T. R. Cook. Short distance directional migration in the threatened socotra cormorant: link to primary productivity and implications for conservation. Marine Ecology Progress Series, 575:181–194, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/575/m575p181.pdf.

Meyer:2011:ESB

[MD11] A. L. Meyer and J. Dierking. Elevated size and body condition and altered feeding ecology of the grouper Cephalopholis argus in non-native habitats. Marine Ecology Progress Series, 439:203–212, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/439/m439p203.pdf.

Moore:2016:FSI

[MD16a] A. F. P. Moore and J. E. Duffy. Foundation species identity and trophic complexity affect experimental seagrass communities. Marine Ecology Progress Series, 556:105–121, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/556/m556p105.pdf.

Mos:2016:EMC

[MD16b] B. Mos and S. A. Dworjanyn. Early metamorphosis is costly and avoided by young, but physiologically competent, marine larvae. Marine Ecology Progress Series, 559:117– 129, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/559/m559p117.pdf.

Mateo:2010:RMN

[MDA^{+}10] I. Mateo, E. G. Durbin, R. S. Appeldoorn, A. J. Adams, F. Juanes, R. Kingsley, P. Swart, and D. Durant. Role of mangroves as nurseries for French grunt Haemulon flavolineatum and schoolmaster Lutjanus apodus assessed by otolith elemental fingerprints. Marine Ecology Progress Series, 402: 197–212, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/402/m402p197.pdf.

Mooney:2011:IGD

[MDAPH11] B. D. Mooney, J. J. Dorantes-Aranda, A. R. Place, and G. M. Hallegraeff. Ichthyotoxicity of gymnodinioid dinoflagellates: PUFA and superoxide effects in sheepshead minnow larvae and rainbow trout gill cells. Marine Ecology Progress Series, 426:213–224, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/426/m426p213.pdf.

Moran:2013:CFT

[MDE13] X. A. G. Morán, H. W. Ducklow, and M. Erickson. Carbon fluxes through estuarine bacteria reflect coupling with phytoplankton. Marine Ecology Progress Series, 489:75–85, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/489/m489p075.pdf.

Menden-Deuer:2010:SDP

[MDF10] S. Menden-Deuer and K. Fredrickson. Structure-dependent, protistan grazing and its implication for the formation, maintenance and decline of plankton patches. Marine Ecology Progress Series, 420:57–71, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m420p057.pdf.

Mueller:2019:OOC

[MDJN19] P. Mueller, H. T. Do, K. Jensen, and S. Nolte. Origin of organic carbon in the topsoil of Wadden Sea salt marshes. Marine Ecology Progress Series, 624:39–50, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/624/
m624p039.pdf.

Moore:2013:DMP

[MDKM13] C. H. Moore, J. C. Drazen, C. D. Kelley, and W. F. X. E. Misa. Deepwater marine protected areas of the main Hawaiian Islands: establishing baselines for commercially valuable bottomfish populations. Marine Ecology Progress Series, 476:167–183, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/476/m476p167.pdf.

Ma:2017:STD

[MDLM17] K. C. K. Ma, D. Deibel, J. B. Lowen, and C. H. McKenzie. Spatio-temporal dynamics of ascidian larval recruitment and colony abundance in a non-indigenous Newfoundland population. Marine Ecology Progress Series, 585:99– 112, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/585/m585p099.pdf.

McMahon:2018:FRD

[MDM18] S. J. McMahon, J. M. Donelson, and P. L. Munday. Food ration does not influence the effect of elevated CO_{2} on antipredator behaviour of a reef fish. Marine Ecology Progress Series, 586:155–165, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/586/m586p155.pdf.

Musyl:2011:RPP

[MDNL^{+}11] M. K. Musyl, M. L. Domeier, N. Nasby-Lucas, R. W. Brill, L. M. McNaughton, J. Y. Swimmer, M. S. Lutcavage, S. G. Wilson, B. Galuardi, and J. B. Liddle. Review: Performance of pop-up satellite archival tags. Marine Ecology Progress Series, 433:1–28, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

MacIntosh:2012:SMC

[MdNW12] H. MacIntosh, R. de Nys, and S. Whalan. Shipworms as a model for competition and coexistence in specialized habitats. Marine Ecology Progress Series, 461:95–105, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/461/
m461p095.pdf.

Miner:2013:WIE

[MDPG13] B. G. Miner, D. A. Donovan, L. M. Portis, and T. C. Goulding. Whelks induce an effective defense against sea stars. Marine Ecology Progress Series, 493:195–206, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/493/
m493p195.pdf.

Mollet:2016:REF

[MDR16] F. M. Mollet, U. Dieckmann, and A. D. Rijnsdorp. Reconstructing the effects of fishing on life-history evolution in North Sea plaice Pleuronectes platessa. Marine Ecology Progress Series, 542:195–208, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/542/m542p195.pdf.

Montresor:2013:DGP

[MDS^{+}13] M. Montresor, C. Di Prisco, D. Sarno, F. Margiotta, and A. Zingone. Diversity and germination patterns of diatom resting stages at a coastal Mediterranean site. Marine Ecology Progress Series, 484:79–95, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m484p079.pdf.

Murray:2014:SST

[MDS14] F. Murray, A. Douglas, and M. Solan. Species that share traits do not necessarily form distinct and universally applicable functional effect groups. Marine Ecology Progress Series, 516:23–34, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m516p023.pdf.

Maduna:2016:WTO

[MdSW^{+}16] S. N. Maduna, C. da Silva, S. P. Wintner, R. Roodt-Wilding, and A. E. Bester van der Merwe. When two oceans meet: regional population genetics of an exploited coastal shark, Mustelus mustelus. Marine Ecology Progress Series, 544:183– 196, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/544/m544p183.pdf.

Maher:2011:BCM

[ME11] D. Maher and B. D. Eyre. Benthic carbon metabolism in southeast Australian estuaries: habitat importance, driving forces, and application of artificial neural network models. Marine Ecology Progress Series, 439:97–115, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/439/
m439p097.pdf.

Maas:2011:MRA

[MEDS11] A. E. Maas, L. E. Elder, H. M. Dierssen, and B. A. Seibel. Metabolic response of Antarctic pteropods (Mollusca: Gastropoda) to food deprivation and regional productivity. Marine Ecology Progress Series, 441:129–139, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/441/
m441p129.pdf.

McGaw:2011:PDN

[MEK11] I. J. McGaw, T. C. Edgell, and M. J. Kaiser. Population demographics of native and newly invasive populations of the green crab Carcinus maenas. Marine Ecology Progress Series, 430:235–240, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m430p235.pdf.

Manzello:2015:RDG

[MEKC15] D. P. Manzello, I. C. Enochs, G. Kolodziej, and R. Carlton. Recent decade of growth and calcification of Orbicella faveolata in the Florida Keys: an inshore-offshore comparison. Marine Ecology Progress Series, 521:81–89, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m521p081.pdf.

Mousing:2014:GPP

[MER14] E. A. Mousing, M. Ellegaard, and K. Richardson. Global patterns in phytoplankton community size structure — evidence for a direct temperature effect. Marine Ecology Progress Series, 497:25–38, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m497p025.pdf.

Metaxas:2011:SPL

[Met11] A. Metaxas. Spatial patterns of larval abundance at hydrothermal vents on seamounts: evidence for recruitment limitation. Marine Ecology Progress Series, 437:103–117, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/437/
m437p103.pdf.

Morgan:2010:LBR

[MF10] S. G. Morgan and J. L. Fisher. Larval behavior regulates nearshore retention and offshore migration in an upwelling shadow and along the open coast. Marine Ecology Progress Series, 404:109–126, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/404/m404p109.pdf.

Minchinton:2013:SDA

[MF13] T. E. Minchinton and K. J. Fels. Sediment disturbance associated with trampling by humans alters species assemblages on a rocky intertidal seashore. Marine Ecology Progress Series, 472:129–140, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/472/m472p129.pdf.

Matthews:2015:SFB

[MF15] C. J. D. Matthews and S. H. Ferguson. Seasonal foraging behaviour of Eastern Canada–West Greenland bowhead whales: an assessment of isotopic cycles along baleen. Marine Ecology Progress Series, 522:269–286, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/522/m522p269.pdf.

Matley:2012:SPA

[MFD12] J. K. Matley, A. T. Fisk, and T. A. Dick. Seabird predation on Arctic cod during summer in the Canadian Arctic.

Marine Ecology Progress Series, 450:219–228, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/450/
m450p219.pdf.

Michel:2017:OSR

[MFDW17] J. Michel, S. R. Fegley, J. A. Dahlin, and C. Wood. Oil spill response-related injuries on sand beaches: when shoreline treatment extends the impacts beyond the oil. Marine Ecology Progress Series, 576:203–218, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m576p203.pdf.

Maneja:2013:EOA

[MFG^{+}13] R. H. Maneja, A. Y. Frommel, A. J. Geffen, A. Folkvord, U. Piatkowski, M. Y. Chang, and C. Clemmesen. Effects of ocean acidification on the calcification of otoliths of larval Atlantic cod Gadus morhua. Marine Ecology Progress Series, 477:251–258, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/477/m477p251.pdf.

Morys:2016:VBV

[MFG16] C. Morys, S. Forster, and G. Graf. Variability of bioturbation in various sediment types and on different spatial scales in the southwestern Baltic Sea. Marine Ecology Progress Series, 557:31–49, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/557/m557p031.pdf.

Mayr:2011:SIV

[MFH^{+}11] C. C. Mayr, G. Försterra, V. Häussermann, A. Wunderlich, J. Grau, M. Zieringer, and A. V. Altenbach. Stable isotope variability in a Chilean fjord food web: implications for n- and c-cycles. Marine Ecology Progress Series, 428:89–104, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/428/m428p089.pdf.

Mei:2011:PGA

[MFI11] Z. P. Mei, Z. V. Finkel, and A. J. Irwin. Phytoplankton growth allometry and size dependent C:N stoichiometry revealed by a variable quota model. Marine Ecology Progress Series, 434:29–43, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/434/m434p029.pdf.

Mittermayr:2014:TVS

[MFS14] A. Mittermayr, S. E. Fox, and U. Sommer. Temporal variation in stable isotope composition (δ ^{13}C, δ ^{15}N and δ ^{34}S) of a temperate Zostera marina food web. Marine Ecology Progress Series, 505:95–105, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/505/m505p095.pdf.

McConnico:2014:PBL

[MFSRR14] L. A. McConnico, M. S. Foster, D. L. Steller, and R. Riosmena-Rodríguez. Population biology of a long-lived rhodolith: the consequences of becoming old and large. Marine Ecology Progress Series, 504:109–118, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/504/m504p109.pdf.

Minutoli:2012:MCR

[MG12] R. Minutoli and L. Guglielmo. Mesozooplankton carbon requirement in the Tyrrhenian Sea: its vertical distribution, diel variability and relation to particle flux. Marine Ecology Progress Series, 446:91–105, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/446/m446p091.pdf.

McTee:2014:NDD

[MG14] S. A. McTee and J. R. Grubich. Native densities, distribution, and diurnal activity of Red Sea lionfishes (Scorpaenidae). Marine Ecology Progress Series, 508:223–232, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/508/
m508p223.pdf.

Millar:2018:MSR

[MG18] K. R. Millar and P. Gagnon. Mechanisms of stability of rhodolith beds: sedimentological aspects. Marine Ecology Progress Series, 594:65–83, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/594/m594p065.pdf.

Medina:2015:FPA

[MGAV15] A. Medina, N. Goñi, H. Arrizabalaga, and J. L. Varela. Feeding patterns of age-0 bluefin tuna in the western Mediterranean inferred from stomach-content and isotope analyses. Marine Ecology Progress Series, 527:193–204, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/527/
m527p193.pdf.

Minutoli:2014:MCR

[MGB^{+}14] R. Minutoli, A. Granata, C. Brugnano, G. Zagami, and L. Guglielmo. Mesozooplankton carbon requirement in the southern Adriatic Sea: vertical distribution, diel and seasonal variability, relation to particle flux. Marine Ecology Progress Series, 495:91–104, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/495/m495p091.pdf.

Montevecchi:2019:PDS

[MGB^{+}19] W. A. Montevecchi, K. Gerrow, A. D. Buren, G. K. Davoren, K. P. Lewis, M. W. Montevecchi, and P. M. Regular. Pursuitdiving seabird endures regime shift involving a three-decade decline in forage fish mass and abundance. Marine Ecology Progress Series, 627:171–178, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m627p171.pdf.

Moore:2016:AES

[MGCH^{+}16] J. W. Moore, J. Gordon, C. Carr-Harris, A. S. Gottesfeld, S. M. Wilson, and J. H. Russell. Assessing estuaries as stopover habitats for juvenile Pacific salmon. Marine Ecology Progress Series, 559:201–215, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/559/m559p201.pdf.

Munday:2011:OAD

[MGD^{+}11] P. L. Munday, M. Gagliano, J. M. Donelson, D. L. Dixson, and S. R. Thorrold. Ocean acidification does not affect the early life history development of a tropical marine fish. Marine Ecology Progress Series, 423:211–221, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/423/m423p211.pdf.

Martinez-Garcia:2010:DRP

[MGFÁS^{+}10] S. Martínez-García, E. Fernández, X. A. Álvarez-Salgado, J. González, C. Lønborg, E. Marañón, X. A. G. Morán, and E. Teira. Differential responses of phytoplankton and heterotrophic bacteria to organic and inorganic nutrient additions in coastal waters off the NW Iberian Peninsula. Marine Ecology Progress Series, 416:17–33, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/416/
m416p017.pdf.

Mabin:2013:VMR

[MGFW13] C. J. T. Mabin, P. E. Gribben, A. Fischer, and J. T. Wright. Variation in the morphology, reproduction and development of the habitat-forming kelp Ecklonia radiata with changing temperature and nutrients. Marine Ecology Progress Series, 483:117–131, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/483/m483p117.pdf.

Mundy:2014:REF

[MGG^{+}14] C. J. Mundy, M. Gosselin, Y. Gratton, K. Brown, V. Galindo, K. Campbell, M. Levasseur, D. Barber, T. Papakyriakou, and S. Bélanger. Role of environmental factors on phytoplankton bloom initiation under landfast sea ice in Resolute Passage, Canada. Marine Ecology Progress Series, 497:39–49, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m497p039.pdf.

Meyer-Gutbrod:2015:CAC

[MGGSP15] E. L. Meyer-Gutbrod, C. H. Greene, P. J. Sullivan, and A. J. Pershing. Climate-associated changes in prey availability drive reproductive dynamics of the North Atlantic right whale population. Marine Ecology Progress Series, 535:243– 258, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/535/m535p243.pdf.

Morgan:2016:TMI

[MGL^{+}16] S. G. Morgan, S. A. Gravem, A. C. Lipus, M. Grabiel, and B. G. Miner. Trait-mediated indirect interactions among residents of rocky shore tidepools. Marine Ecology Progress Series, 552:31–46, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m552p031.pdf.

Miller:2010:QCJ

[MGM10] J. A. Miller, A. Gray, and J. Merz. Quantifying the contribution of juvenile migratory phenotypes in a population of Chinook salmon Oncorhynchus tshawytscha. Marine Ecology Progress Series, 408:227–240, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/408/m408p227.pdf.

Malinka:2018:FLS

[MGM^{+}18] C. E. Malinka, D. M. Gillespie, J. D. J. Macaulay, R. Joy, and C. E. Sparling. First in situ passive acoustic monitoring for marine mammals during operation of a tidal turbine in Ramsey Sound, Wales. Marine Ecology Progress Series, 590:247–266, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/590/m590p247.pdf.

Manriquez:2012:GVT

[MGO^{+}12] P. H. Manríquez, S. P. Galaz, T. Opitz, S. Hamilton, G. Paradis, R. R. Warner, J. C. Castilla, F. A. Labra, and N. A. Lagos. Geographic variation in trace-element signatures in the statoliths of near-hatch larvae and recruits of Concholepas concholepas (loco). Marine Ecology Progress Series, 448:105–118, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/448/m448p105.pdf.

Morais:2019:MFT

[MGOL19] G. C. Morais, J. B. Gusmao, V. M. Oliveira, and P. Lana. Macrobenthic functional trait diversity at multiple scales along a subtropical estuarine gradient. Marine Ecology Progress Series, 624:23–37, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/624/m624p023.pdf.

Macreadie:2011:HSS

[MGP11] P. I. Macreadie, N. R. Geraldi, and C. H. Peterson. How smallscale variation in oyster reef patchiness influences predation on bivalves. Marine Ecology Progress Series, 429:87–91, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m429p087.pdf.

Macreadie:2012:PFH

[MGP12] P. I. Macreadie, N. R. Geraldi, and C. H. Peterson. Preference for feeding at habitat edges declines among juvenile blue crabs as oyster reef patchiness increases and predation risk grows. Marine Ecology Progress Series, 466:145–153, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m466p145.pdf.

McDuie:2013:DCD

[MGPC13] F. McDuie, W. Goulding, D. R. Peck, and B. C. Congdon. Divergence in chick developmental patterns among wedgetailed shearwater populations. Marine Ecology Progress Series, 485:275–285, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/485/m485p275.pdf.

Martino:2010:RSB

[MH10] E. J. Martino and E. D. Houde. Recruitment of striped bass in Chesapeake Bay: spatial and temporal environmental variability and availability of zooplankton prey. Marine Ecology Progress Series, 409:213–228, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/409/m409p213.pdf.

Matthews:2011:MLS

[MH11a] J. A. Matthews and B. T. Hentschel. Measuring in situ growth rates of an infaunal polychaete at different intertidal elevations. Marine Ecology Progress Series, 442:111– 121, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/442/m442p111.pdf.

McGrew:2011:BPI

[MH11b] M. McGrew and K. M. Hultgren. Bopyrid parasite infestation affects activity levels and morphology of the eusocial snapping shrimp Synalpheus elizabethae. Marine Ecology Progress Series, 431:195–204, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/431/m431p195.pdf.

Moulton:2011:CVS

[MH11c] O. M. Moulton and S. D. Hacker. Congeneric variation in surfgrasses and ocean conditions influence macroinvertebrate community structure. Marine Ecology Progress Series, 433:53–63, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/433/m433p053.pdf.

Matich:2012:EET

[MH12a] P. Matich and M. R. Heithaus. Effects of an extreme temperature event on the behavior and age structure of an estuarine top predator, Carcharhinus leucas. Marine Ecology Progress Series, 447:165–178, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2011/447/m447p165.pdf.

McClanahan:2012:DSL

[MH12b] T. R. McClanahan and A. T. Humphries. Differential and slow life-history responses of fishes to coral reef closures. Marine Ecology Progress Series, 469:121–131, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m469p121.pdf.

Myers:2013:ACE

[MH13] J. A. Myers and K. L. Heck, Jr. Amphipod control of epiphyte load and its concomitant effects on shoalgrass Halodule wrightii biomass. Marine Ecology Progress Series, 483:133–142, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m483p133.pdf.

Marshak:2019:CIA

[MH19] A. R. Marshak and K. L. Heck, Jr. Competitive interactions among juvenile and adult life stages of northern Gulf of Mexico red snapper Lutjanus campechanus and a tropical range-expanding congener. Marine Ecology Progress Series, 622:139–155, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/622/m622p139.pdf.

Meager:2018:LTT

[MHAP18] J. J. Meager, E. R. Hawkins, I. Ansmann, and G. J. Parra. Long-term trends in habitat use and site fidelity by Australian humpback dolphins Sousa sahulensis in a near-urban embayment. Marine Ecology Progress Series, 603:227–242, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2018/603/m603p227.pdf.

Mancini:2014:RBS

[MHB14] P. L. Mancini, K. A. Hobson, and L. Bugoni. Role of body size in shaping the trophic structure of tropical seabird communities. Marine Ecology Progress Series, 497:243–257, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2014/497/m497p243.pdf.

McClanahan:2011:CRM

[MHC11] T. R. McClanahan, B. E. Huntington, and B. Cokos. Coral responses to macroalgal reduction and fisheries closure on Caribbean patch reefs. Marine Ecology Progress Series, 437: 89–102, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m437p089.pdf.

McConnico:2018:NPR

[MHCRR18] L. A. McConnico, G. Hernández-Carmona, and R. Riosmena-Rodríguez. Nutrient production in rhodolith beds: impact of a foundation species and its associates. Marine Ecology Progress Series, 590:53–66, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/590/m590p053.pdf.

Martins:2018:BND

[MHCS18] A. P. B. Martins, M. R. Heupel, A. Chin, and C. A. Simpfendorfer. Batoid nurseries: definition, use and importance. Marine Ecology Progress Series, 595:253–267, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2018/595/m595p253.pdf.

Moore:2014:PRB

[MHF^{+}14] E. K. Moore, H. R. Harvey, J. F. Faux, D. R. Goodlett, and B. L. Nunn. Protein recycling in Bering Sea algal incubations. Marine Ecology Progress Series, 515:45–59, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m515p045.pdf.

Martins:2019:DIE

[MHF^{+}19] G. M. Martins, C. D. G. Harley, J. Faria, M. Vale, S. J. Hawkins, A. I. Neto, and F. Arenas. Direct and indirect effects of climate change squeeze the local distribution of a habitat-forming seaweed. Marine Ecology Progress Series, 626:43–52, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/626/m626p043.pdf.

McLean:2010:LDA

[MHFN10] D. L. McLean, E. S. Harvey, D. V. Fairclough, and S. J. Newman. Large decline in the abundance of a targeted tropical lethrinid in areas open and closed to fishing. Marine Ecology Progress Series, 418:189–199, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/418/m418p189.pdf.

Max:2013:BPO

[MHGW13] L. M. Max, S. L. Hamilton, S. D. Gaines, and R. R. Warner. Benthic processes and overlying fish assemblages drive the composition of benthic detritus on a central Pacific coral reef. Marine Ecology Progress Series, 482:181–195, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/482/
m482p181.pdf.

Martien:2017:UPG

[MHHB^{+}17] K. K. Martien, B. L. Hancock-Hanser, R. W. Baird, J. J. Kiszka, J. M. Aschettino, M. Oremus, and M. C. Hill. Unexpected patterns of global population structure in melonheaded whales Peponocephala electra. Marine Ecology Progress Series, 577:205–220, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m577p205.pdf.

McLeod:2013:USF

[MHHF13] R. J. McLeod, G. A. Hyndes, C. L. Hurd, and R. D. Frew. Unexpected shifts in fatty acid composition in response to diet in a common littoral amphipod. Marine Ecology Progress Series, 479:1–12, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Murray:2011:FRF

[MHK11] L. G. Murray, H. Hinz, and M. J. Kaiser. Functional response of fishers in the isle of man scallop fishery. Marine Ecology Progress Series, 430:157–169, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m430p157.pdf.

Mongillo:2012:PPD

[MHN^{+}12] T. M. Mongillo, E. E. Holmes, D. P. Noren, G. R. VanBlaricom, A. E. Punt, S. M. O’Neill, G. M. Ylitalo, M. B. Hanson, and P. S. Ross. Predicted polybrominated diphenyl ether (PBDE) and polychlorinated biphenyl (PCB) accumulation in southern resident killer whales. Marine Ecology Progress Series, 453:263–277, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/453/m453p263.pdf.

Miller:2014:MRF

[MHNL14] K. Miller, F. Huettmann, B. Norcross, and M. Lorenz. Multivariate random forest models of estuarine-associated fish and invertebrate communities. Marine Ecology Progress Series, 500:159–174, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/500/m500p159.pdf.

Munguia:2011:CMP

[MHO^{+}11] P. Munguia, J. Hamilton, R. W. Osman, R. B. Whitlatch, and R. N. Zajac. Clarification of model parameters: Reply to Lauzon-Guay & Lyons (2011) MEPS 432:302. Marine Ecology Progress Series, ??:??, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m432p299.pdf. See [MOH^{+}10,

?].

Mbongwa:2019:EBI

[MHPvdH19] N. A. Mbongwa, C. Hui, A. Pulfrich, and S. von der Heyden. Every beach an island — deep population divergence and possible loss of genetic diversity in Tylos granulatus, a sandy shore isopod. Marine Ecology Progress Series, 614: 111–123, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/614/m614p111.pdf.

Mills:2017:ARC

[MHQ17] K. A. Mills, P. A. Hamer, and G. P. Quinn. Artificial reefs create distinct fish assemblages. Marine Ecology Progress Series, 585:155–173, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/585/m585p155.pdf.

McGowan:2019:ETD

[MHR19a] D. W. McGowan, J. K. Horne, and L. A. Rogers. Effects of temperature on the distribution and density of capelin in the Gulf of Alaska. Marine Ecology Progress Series, 620: 119–138, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/620/m620p119.pdf.

Morgan:2019:IFB

[MHR^{+}19b] E. A. Morgan, C. Hassall, C. P. F. Redfern, R. M. Bevan, and K. C. Hamer. Individuality of foraging behaviour in a shortranging benthic marine predator: incidence and implications. Marine Ecology Progress Series, 609:209–219, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/609/
m609p209.pdf.

Mittermayr:2014:SAR

[MHS14] A. Mittermayr, T. Hansen, and U. Sommer. Simultaneous analysis of δ ^{13}C, δ ^{15}N and δ ^{34}S ratios uncovers food web relationships and the trophic importance of epiphytes in an eelgrass Zostera marina community. Marine Ecology Progress Series, 497:93–103, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/497/m497p093.pdf.

Matley:2015:DSU

[MHS15] J. K. Matley, M. R. Heupel, and C. A. Simpfendorfer. Depth and space use of leopard coralgrouper Plectropomus leopardus using passive acoustic tracking. Marine Ecology Progress Series, 521:201–216, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/521/m521p201.pdf.

McKeown:2017:MGB

[MHS17] N. J. McKeown, L. Hauser, and P. W. Shaw. Microsatellite genotyping of brown crab Cancer pagurus reveals fine scale selection and ‘non-chaotic’ genetic patchiness within a high gene flow system. Marine Ecology Progress Series, 566: 91–103, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/566/m566p091.pdf.

McMillan:2018:NTR

[MHSG18] M. N. McMillan, C. Huveneers, J. M. Semmens, and B. M. Gillanders. Natural tags reveal populations of conservation dependent school shark use different pupping areas. Marine Ecology Progress Series, 599:147–156, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/599/
m599p147.pdf.

Michael:2013:IBH

[MHVV13] P. J. Michael, G. A. Hyndes, M. A. Vanderklift, and A. Vergés. Identity and behaviour of herbivorous fish influence large-scale spatial patterns of macroalgal herbivory in a coral reef. Marine Ecology Progress Series, 482:227–240, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/482/
m482p227.pdf.

McKinney:2012:FHW

[MHW^{+}12] J. A. McKinney, E. R. Hoffmayer, W. Wu, R. Fulford, and J. Hendon. Feeding habitat of the whale shark Rhincodon typus in the northern Gulf of Mexico determined using species distribution modelling. Marine Ecology Progress Series, 458: 199–211, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m458p199.pdf.

McKnight:2011:WDA

[MIA^{+}11] A. McKnight, D. B. Irons, A. J. Allyn, K. M. Sullivan, and R. M. Suryan. Winter dispersal and activity patterns of post-breeding black-legged kittiwakes Rissa tridactyla from Prince William Sound, Alaska. Marine Ecology Progress Series, 442:241–253, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m442p241.pdf.

McDevitt-Irwin:2016:RNR

[MIIB16] J. M. McDevitt-Irwin, J. C. Iacarella, and J. K. Baum. Reassessing the nursery role of seagrass habitats from temperate to tropical regions: a meta-analysis. Marine Ecology Progress Series, 557:133–143, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/557/m557p133.pdf.

Montes:2012:WGS

[MIM^{+}12] I. Montes, M. Iriondo, C. Manzano, H. Arrizabalaga, E. Jiménez, M. Á. Pardo, N. Goñi, C. A. Davies, and A. Estonba. Worldwide genetic structure of albacore Thunnus alalunga revealed by microsatellite DNA markers. Marine Ecology Progress Series, 471:183–191, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m471p183.pdf.

Miura:2012:SOC

[Miu12] O. Miura. Social organization and caste formation in three additional parasitic flatworm species. Marine Ecology Progress Series, 465:119–127, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/465/m465p119.pdf.

Monk:2010:HSM

[MIV^{+}10] J. Monk, D. Ierodiaconou, V. L. Versace, A. Bellgrove, E. Harvey, A. Rattray, L. Laurenson, and G. P. Quinn. Habitat suitability for marine fishes using presence-only modelling and multibeam sonar. Marine Ecology Progress Series, 420:157–174, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/420/m420p157.pdf.

McKinley:2010:ICS

[MJ10] A. McKinley and E. L. Johnston. Impacts of contaminant sources on marine fish abundance and species richness: a review and meta-analysis of evidence from the field. Marine Ecology Progress Series, 420:175–191, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/420/
m420p175.pdf.

McClanahan:2017:MCR

[MJ17] T. R. McClanahan and C. Jadot. Managing coral reef fish community biomass is a priority for biodiversity conservation in Madagascar. Marine Ecology Progress Series, 580: 169–190, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m580p169.pdf.

Moller:2018:ZPM

[MJA^{+}18] E. F. Møller, K. Lambert Johansen, M. D. Agersted, F. Rigét, D. S. Clausen, J. Larsen, P. Lyngs, A. Middelbo, and A. Mosbech. Zooplankton phenology may explain the North water polynya’s importance as a breeding area for little auks. Marine Ecology Progress Series, 605:207–223, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/605/
m605p207.pdf.

Murphy:2011:DVM

[MJHS11] H. M. Murphy, G. P. Jenkins, P. A. Hamer, and S. E. Swearer. Diel vertical migration related to foraging success in snapper Chrysophrys auratus larvae. Marine Ecology Progress Series, 433:185–194, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/433/m433p185.pdf.

Murphy:2013:IVL

[MJHS13] H. M. Murphy, G. P. Jenkins, P. A. Hamer, and S. E. Swearer. Interannual variation in larval abundance and growth in snapper Chrysophrys auratus (Sparidae) is related to prey availability and temperature. Marine Ecology Progress Series, 487:151–162, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/487/m487p151.pdf.

Manriquez:2014:OAA

[MJM^{+}14] P. H. Manríquez, M. E. Jara, M. L. Mardones, R. Torres, J. M. Navarro, M. A. Lardies, C. A. Vargas, C. Duarte, and N. A. Lagos. Ocean acidification affects predator avoidance behaviour but not prey detection in the early ontogeny of a keystone species. Marine Ecology Progress Series, 502: 157–167, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/502/m502p157.pdf.

Manriquez:2013:EPR

[MJO^{+}13] P. H. Manríquez, M. E. Jara, T. Opitz, J. C. Castilla, and N. A. Lagos. Effects of predation risk on survival, behaviour and morphological traits of small juveniles of Concholepas concholepas (loco). Marine Ecology Progress Series, 472:169–183, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/472/m472p169.pdf.

Martins:2014:EPR

[MJR^{+}14] G. M. Martins, S. R. Jenkins, R. Ramírez, F. Tuya, A. I. Neto, and F. Arenas. Early patterns of recovery from disturbance in intertidal algal assemblages: consistency across regions within a marine province. Marine Ecology Progress Series, 517:131–142, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/517/m517p131.pdf.

Magnhagen:2017:EMN

[MJS17] C. Magnhagen, K. Johansson, and P. Sigray. Effects of motorboat noise on foraging behaviour in Eurasian perch and roach: a field experiment. Marine Ecology Progress Series, 564:115–125, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m564p115.pdf.

Moller:2012:MSZ

[MJT^{+}12] K. O. Möller, M. St John, A. Temming, J. Floeter, A. F. Sell, J. P. Herrmann, and C. Möllmann. Marine snow, zooplankton and thin layers: indications of a trophic link from small-scale sampling with the video plankton recorder. Marine Ecology Progress Series, 468:57–69, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m468p057.pdf.

Manriquez:2014:EOA

[MJT^{+}14] P. H. Manríquez, M. E. Jara, R. Torres, M. L. Mardones, N. A. Lagos, M. A. Lardies, C. A. Vargas, C. Duarte, and J. M. Navarro. Effects of ocean acidification on larval development and early post-hatching traits in Concholepas concholepas (loco). Marine Ecology Progress Series, 514:87– 103, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/514/m514p087.pdf.

Martinez-Jauregui:2012:PRM

[MJTC^{+}12] M. Martínez-Jauregui, G. Tavecchia, M. A. Cedenilla, T. Coulson, P. Fernández de Larrinoa, M. Muñoz, and L. M. González. Population resilience of the Mediterranean monk seal Monachus monachus at Cabo Blanco peninsula. Marine Ecology Progress Series, 461:273–281, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/461/
m461p273.pdf.

Moreno:2011:ITE

[MJV^{+}11] R. Moreno, L. Jover, A. Velando, I. Munilla, and C. Sanpera. Influence of trophic ecology and spatial variation on the isotopic fingerprints of seabirds. Marine Ecology Progress Series, 442:229–239, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/442/m442p229.pdf.

Mabin:2019:PRT

[MJW19] C. J. T. Mabin, C. R. Johnson, and J. T. Wright. Physiological response to temperature, light, and nitrates in the giant kelp Macrocystis pyrifera from Tasmania, Australia. Marine Ecology Progress Series, 614:1–19, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

MacKenzie:2013:IEM

[MK13] R. A. MacKenzie and C. L. Kryss. Impacts of exotic mangroves and chemical eradication of mangroves on tide pool fish assemblages. Marine Ecology Progress Series, 472:219– 237, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/472/m472p219.pdf.

Morissette:2010:EMC

[MKG10] L. Morissette, K. Kaschner, and L. R. Gerber. Ecosystem models clarify the trophic role of whales off Northwest Africa. Marine Ecology Progress Series, 404:289–302, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/404/
m404p289.pdf.

Miyake:2015:RVB

[MKI^{+}15] Y. Miyake, S. Kimura, S. Itoh, S. Chow, K. Murakami, S. Katayama, A. Takeshige, and H. Nakata. Roles of vertical behavior in the open-ocean migration of teleplanic larvae: a modeling approach to the larval transport of Japanese spiny lobster. Marine Ecology Progress Series, 539:93–109, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2015/539/m539p093.pdf.

Murase:2019:CDL

[MKII19] I. Murase, T. Kawakami, T. Irie, and K. Iguchi. Counterdirectional latitudinal clines of size at upstream migration between two adjacent water bodies in a Japanese amphidromous fish. Marine Ecology Progress Series, 624:143–154, 2019. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m624p143.pdf.

Miyake:2011:PCE

[MKK^{+}11] Y. Miyake, S. Kimura, T. Kawamura, T. Kitagawa, T. Takahashi, and H. Takami. Population connectivity of Ezo abalone on the northern Pacific coast of Japan in relation to the establishment of harvest refugia. Marine Ecology Progress Series, 440:137–150, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/440/m440p137.pdf.

Mae:2013:DHD

[MKK13] Y. Mae, M. Kanno, and A. Kijima. Detection of a highly divergent population structure and identification of a cryptic species in the East Asian dogwhelk Nucella heyseana. Marine Ecology Progress Series, 484:131–141, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/484/
m484p131.pdf.

McLaskey:2016:DEP

[MKM^{+}16] A. K. McLaskey, J. E. Keister, P. McElhany, M. B. Olson, D. S. Busch, M. Maher, and A. K. Winans. Development of Euphausia pacifica (krill) larvae is impaired under pCO_{2} levels currently observed in the Northeast Pacific. Marine Ecology Progress Series, 555:65–78, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m555p065.pdf.

Mozdzer:2011:NUS

[MKMZ11] T. J. Mozdzer, M. Kirwan, K. J. McGlathery, and J. C. Zieman. Nitrogen uptake by the shoots of smooth cordgrass Spartina alterniflora. Marine Ecology Progress Series, 433: 43–52, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/433/m433p043.pdf.

Mangano:2013:ETD

[MKPS13] M. C. Mangano, M. J. Kaiser, E. M. D. Porporato, and N. Spanò. Evidence of trawl disturbance on mega-epibenthic communities in the southern Tyrrhenian Sea. Marine Ecology Progress Series, 475:101–117, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/475/m475p101.pdf.

McLean:2018:DGI

[MKS18] E. L. McLean, N. V. Katenka, and B. A. Seibel. Decreased growth and increased shell disease in early benthic phase Homarus americanus in response to elevated CO_{2}. Marine Ecology Progress Series, 596:113–126, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/596/
m596p113.pdf.

Murillo:2018:SAP

[MKT^{+}18] F. J. Murillo, E. Kenchington, G. Tompkins, L. Beazley, E. Baker, A. Knudby, and W. Walkusz. Sponge assemblages and predicted archetypes in the eastern Canadian Arctic. Marine Ecology Progress Series, 597:115–135, 2018. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m597p115.pdf.

Maldonado:2011:SSE

[ML11] E. M. Maldonado and M. I. Latz. Species-specific effects of fluid shear on grazing by sea urchin larvae: comparison of experimental results with encounter-model predictions. Marine Ecology Progress Series, 436:119–130, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/436/
m436p119.pdf.

Maldonado:2013:RGG

[MLASTS13] M. Maldonado, M. López-Acosta, L. Sánchez-Tocino, and C. Sitjà. The rare, giant gorgonian Ellisella paraplexauroides: demographics and conservation concerns. Marine Ecology Progress Series, 479:127–141, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m479p127.pdf.

Midwood:2014:DCM

[MLB^{+}14] J. D. Midwood, M. H. Larsen, M. Boel, N. Jepsen, K. Aarestrup, and S. J. Cooke. Does cortisol manipulation influence outmigration behaviour, survival and growth of sea trout? A field test of carryover effects in wild fish. Marine Ecology Progress Series, 496:135–144, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/496/m496p135.pdf.

Mendez:2018:ETI

[MLB18] M. M. Mendez, J. P. Livore, and G. Bigatti. Effects of trampling on intertidal mussel beds: importance of disturbance intensity. Marine Ecology Progress Series, 606:231– 235, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/606/m606p231.pdf.

Mendez:2019:INA

[MLB19] M. M. Mendez, J. P. Livore, and G. Bigatti. Interaction of natural and anthropogenic stressors on rocky shores: community resistance to trampling. Marine Ecology Progress Series, 631:117–126, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/631/m631p117.pdf.

McNamara:2014:RES

[MLC14] M. E. McNamara, D. J. Lonsdale, and R. M. Cerrato. Role of eutrophication in structuring planktonic communities in the presence of the ctenophore Mnemiopsis leidyi. Marine Ecology Progress Series, 510:151–165, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/510/m510p151.pdf.

Mazur:2019:CCM

[MLCC19] M. D. Mazur, B. Li, J. H. Chang, and Y. Chen. Contributions of a conservation measure that protects the spawning stock to drastic increases in the Gulf of Maine American lobster fishery. Marine Ecology Progress Series, 631:127–139, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/631/
m631p127.pdf.

Moksness:2012:BMP

[MLDG12] E. Moksness, J. Link, K. Drinkwater, and S. Gaichas. Bernard megrey: pioneer of comparative marine ecosystem analyses. Marine Ecology Progress Series, 459:165–167, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m459p165.pdf.

Milewski:2018:SCA

[MLF^{+}18] I. Milewski, R. H. Loucks, B. Fisher, R. E. Smith, J. S. P. McCain, and H. K. Lotze. Sea-cage aquaculture impacts market and berried lobster (Homarus americanus) catches. Marine Ecology Progress Series, 598:85–97, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/598/

m598p085.pdf. See corrigendum [Ano19c].

MacDonald:2012:HRE

[MLM^{+}12] B. D. MacDonald, R. L. Lewison, S. V. Madrak, J. A. Seminoff, and T. Eguchi. Home ranges of East Pacific green turtles Chelonia mydas in a highly urbanized temperate foraging ground. Marine Ecology Progress Series, 461:211– 221, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/461/m461p211.pdf.

Monsinjon:2019:ETD

[MLML^{+}19] J. Monsinjon, M. Lopez-Mendilaharsu, P. Lara, A. Santos, M. A. G. dei Marcovaldi, M. Girondot, and M. M. P. B. Fuentes. Effects of temperature and demography on the phenology of loggerhead sea turtles in Brazil. Marine Ecology Progress Series, 623:209–219, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/623/m623p209.pdf.

Morrow:2013:BSA

[MLP^{+}13] K. M. Morrow, M. R. Liles, V. J. Paul, A. Moss, and N. E. Chadwick. Bacterial shifts associated with coral-macroalgal competition in the Caribbean Sea. Marine Ecology Progress Series, 488:103–117, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/488/m488p103.pdf.

Moriceau:2014:BSD

[MLP^{+}14] B. Moriceau, G. G. Laruelle, U. Passow, P. Van Cappellen, and O. Ragueneau. Biogenic silica dissolution in diatom aggregates: insights from reactive transport modelling. Marine Ecology Progress Series, 517:35–49, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/517/
m517p035.pdf.

McGeady:2019:TMF

[MLP19] R. McGeady, C. Lordan, and A. M. Power. Twilight migrators: factors determining larval vertical distribution in Nephrops norvegicus with implications for larval retention. Marine Ecology Progress Series, 631:141–155, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m631p141.pdf.

Maharaj:2018:RVS

[MLPC18] R. R. Maharaj, V. W. Y. Lam, D. Pauly, and W. W. L. Cheung. Regional variability in the sensitivity of Caribbean reef fish assemblages to ocean warming. Marine Ecology Progress Series, 590:201–209, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/590/m590p201.pdf.

Muhling:2012:DSR

[MLR12] B. A. Muhling, J. T. Lamkin, and W. J. Richards. Decadalscale responses of larval fish assemblages to multiple ecosystem processes in the northern Gulf of Mexico. Marine Ecology Progress Series, 450:37–53, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/450/m450p037.pdf.

Markel:2017:TVE

[MLR17] R. W. Markel, K. E. Lotterhos, and C. L. K. Robinson. Temporal variability in the environmental and geographic predictors of spatial-recruitment in nearshore rockfishes. Marine Ecology Progress Series, 574:97–111, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/574/
m574p097.pdf.

Moutopoulos:2014:ELD

[MLS^{+}14] D. K. Moutopoulos, S. Libralato, C. Solidoro, K. Erzini, and Konstantinos I. Stergiou. Effect of landings data disaggregation on ecological indicators. Marine Ecology Progress Series, 509:27–38, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m509p027.pdf.

Musetta-Lambert:2015:ICD

[MLSK^{+}15] J. L. Musetta-Lambert, R. A. Scrosati, E. A. Keppel, M. A. Barbeau, M. A. Skinner, and S. C. Courtenay. Intertidal communities differ between breakwaters and natural rocky areas on ice-scoured Northwest Atlantic coasts. Marine Ecology Progress Series, 539:19–31, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/539/m539p019.pdf.

McCormick:2010:IAI

[MM10] M. I. McCormick and M. G. Meekan. The importance of attitude: the influence of behaviour on survival at an ontogenetic boundary. Marine Ecology Progress Series, 407: 173–185, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/407/m407p173.pdf.

Madin:2011:WSI

[MM11] E. M. P. Madin and J. S. Madin. As we see it: Predators, facilitators, or both? Re-evaluating an apparent predator– prey relationship. Marine Ecology Progress Series, 431:299– 302, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/431/m431p299.pdf.

Martone:2012:GVD

[MM12] R. G. Martone and F. Micheli. Geographic variation in demography of a temperate reef snail: importance of multiple life-history traits. Marine Ecology Progress Series, 457: 85–99, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/457/m457p085.pdf.

Maxwell:2013:FSP

[MM13a] S. M. Maxwell and L. E. Morgan. Foraging of seabirds on pelagic fishes: implications for management of pelagic marine protected areas. Marine Ecology Progress Series, 481: 289–303, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/481/m481p289.pdf.

McAlister:2013:EVE

[MM13b] J. S. McAlister and A. L. Moran. Effects of variation in egg energy and exogenous food on larval development in congeneric sea urchins. Marine Ecology Progress Series, 490: 155–167, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m490p155.pdf.

Miller:2013:IDD

[MM13c] S. H. Miller and S. G. Morgan. Interspecific differences in depth preference: regulation of larval transport in an upwelling system. Marine Ecology Progress Series, 476:301–306, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m476p301.pdf.

McClanahan:2016:SIF

[MM16] T. R. McClanahan and N. A. Muthiga. Similar impacts of fishing and environmental stress on calcifying organisms in Indian Ocean coral reefs. Marine Ecology Progress Series, 560:87–103, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m560p087.pdf.

McLean:2018:ELH

[MMA18] M. McLean, D. Mouillot, and A. Auber. Ecological and life history traits explain a climate-induced shift in a temperate marine fish community. Marine Ecology Progress Series, 606:175–186, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/606/m606p175.pdf.

Mercier:2012:MEO

[MMB^{+}12a] L. Mercier, D. Mouillot, O. Bruguier, L. Vigliola, and A. M. Darnaude. Multi-element otolith fingerprints unravel sea–lagoon lifetime migrations of gilthead sea bream Sparus aurata. Marine Ecology Progress Series, 444:175– 194, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/444/m444p175.pdf.

Morley:2012:STV

[MMB^{+}12b] S. A. Morley, S. M. Martin, A. E. Bates, M. S. Clark, J. Ericson, M. Lamare, and L. S. Peck. Spatial and temporal variation in the heat tolerance limits of two abundant Southern Ocean invertebrates. Marine Ecology Progress Series, 450:81–92, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/450/m450p081.pdf.

Morse:2019:CSA

[MMB^{+}19] P. Morse, M. A. Mole, M. N. Bester, R. Johnson, U. Scacco, and E. Gennari. Cape fur seals (Arctocephalus pusillus pusillus) adjust traversing behaviour with lunar conditions in the high white shark (Carcharodon carcharias) density waters of Mossel Bay, South Africa. Marine Ecology Progress Series, 622:219–230, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/622/m622p219.pdf.

Mills:2013:PED

[MMBA13] S. W. Mills, L. S. Mullineaux, S. E. Beaulieu, and D. K. Adams. Persistent effects of disturbance on larval patterns in the plankton after an eruption on the East Pacific rise. Marine Ecology Progress Series, 491:67–76, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/491/
m491p067.pdf.

Mille:2016:DCO

[MMC^{+}16] T. Mille, K. Mahé, M. Cachera, M. C. Villanueva, H. de Pontual, and B. Ernande. Diet is correlated with otolith shape in marine fish. Marine Ecology Progress Series, 555:167– 184, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/555/m555p167.pdf.

Mitchell:2018:QSD

[MMC^{+}18] J. D. Mitchell, D. L. McLean, S. P. Collin, S. Taylor, G. Jackson, R. Fisher, and T. J. Langlois. Quantifying shark depredation in a recreational fishery in the Ningaloo Marine Park and Exmouth Gulf, Western Australia. Marine Ecology Progress Series, 587:141–157, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/587/m587p141.pdf.

Mitchell:2019:SDB

[MMCL19] J. D. Mitchell, D. L. McLean, S. P. Collin, and T. J. Langlois. Shark depredation and behavioural interactions with fishing gear in a recreational fishery in Western Australia. Marine Ecology Progress Series, 616:107–122, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/616/
m616p107.pdf.

Miller:2014:PCM

[MMCW14] M. W. Miller, C. Marmet, C. M. Cameron, and D. E. Williams. Prevalence, consequences, and mitigation of fireworm predation on endangered staghorn coral. Marine Ecology Progress Series, 516:187–194, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2014/516/m516p187.pdf.

Mikkelsen:2013:RES

[MMD^{+}13] L. Mikkelsen, K. N. Mouritsen, K. Dahl, J. Teilmann, and J. Tougaard. Re-established stony reef attracts harbour porpoises Phocoena phocoena. Marine Ecology Progress Series, 481:239–248, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m481p239.pdf.

Marcello:2012:ETG

[MMDM12] L. A. Marcello, F. J. Mueter, E. G. Dawe, and M. Moriyasu. Effects of temperature and gadid predation on snow crab recruitment: comparisons between the Bering Sea and Atlantic Canada. Marine Ecology Progress Series, 469:249–261, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m469p249.pdf.

Morse:2014:PND

[MMEM14] R. E. Morse, M. R. Mulholland, T. A. Egerton, and H. G. Marshall. Phytoplankton and nutrient dynamics in a tidally dominated eutrophic estuary: daily variability and controls on bloom formation. Marine Ecology Progress Series, 503: 59–74, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/503/m503p059.pdf.

Marcoux:2012:CTV

[MMFF12] M. Marcoux, B. C. McMeans, A. T. Fisk, and S. H. Ferguson. Composition and temporal variation in the diet of beluga whales, derived from stable isotopes. Marine Ecology Progress Series, 471:283–291, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/471/m471p283.pdf.

Monteiro:2015:LFP

[MMFP^{+}15] S. S. Monteiro, P. Méndez-Fernandez, S. Piertney, C. F. Moffat, M. Ferreira, J. V. Vingada, A. López, A. Brownlow, P. Jepson, B. Mikkelsen, M. Niemeyer, J. C. Carvalho, and G. J. Pierce. Long-finned pilot whale population diversity and structure in Atlantic waters assessed through biogeochemical and genetic markers. Marine Ecology Progress Series, 536:243–257, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/536/m536p243.pdf.

Maggi:2015:LLS

[MMG^{+}15] E. Maggi, M. Milazzo, M. Graziano, R. Chemello, and L. Benedetti-Cecchi. Latitudinal- and local-scale variations in a rocky intertidal interaction web. Marine Ecology Progress Series, 534:39–48, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m534p039.pdf.

Matheson:2016:LED

[MMG^{+}16] K. Matheson, C. H. McKenzie, R. S. Gregory, D. A. Robichaud, I. R. Bradbury, P. V. R. Snelgrove, and G. A. Rose. Linking eelgrass decline and impacts on associated fish communities to European green crab Carcinus maenas invasion. Marine Ecology Progress Series, 548:31–45, 2016. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m548p031.pdf.

Murray:2014:OSO

[MMGB14] C. S. Murray, A. Malvezzi, C. J. Gobler, and H. Baumann. Offspring sensitivity to ocean acidification changes seasonally in a coastal marine fish. Marine Ecology Progress Series, 504: 1–11, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Mahfouz:2017:MAA

[MMH^{+}17] C. Mahfouz, T. Meziane, F. Henry, C. Abi-Ghanem, J. Spitz, T. Jauniaux, T. Bouveroux, G. Khalaf, and R. Amara. Multiapproach analysis to assess diet of harbour porpoises Phocoena phocoena in the southern North Sea. Marine Ecology Progress Series, 563:249–259, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/563/m563p249.pdf.

McLeod:2015:LVL

[MMM^{+}15] I. M. McLeod, M. I. McCormick, P. L. Munday, T. D. Clark, A. S. Wenger, R. M. Brooker, M. Takahashi, and G. P. Jones. Latitudinal variation in larval development of coral reef fishes: implications of a warming ocean. Marine Ecology Progress Series, 521:129–141, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2015/521/m521p129.pdf.

Monczak:2019:SPS

[MMM^{+}19] A. Monczak, C. Mueller, M. E. Miller, Y. Ji, S. A. Borgianini, and E. W. Montie. Sound patterns of snapping shrimp, fish, and dolphins in an estuarine soundscape of the southeastern USA. Marine Ecology Progress Series, 609: 49–68, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/609/m609p049.pdf.

Morey:2012:ACG

[MMNR^{+}12] G. Morey, B. Morales-Nin, F. Riera, A. Grau, A. J. Geffen, S. Pérez-Mayol, M. Y. Chang, and A. M. Grau. Atlantic cod Gadus morhua in the Mediterranean: a surprising immigrant. Marine Ecology Progress Series, 467:277–280, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/467/
m467p277.pdf.

Marco-Mendez:2012:EMT

[MMPH^{+}12] C. Marco-Méndez, P. Prado, K. L. Heck, J. Cebrián, and J. L. Sánchez-Lizaso. Epiphytes mediate the trophic role of sea urchins in Thalassia testudinum seagrass beds. Marine Ecology Progress Series, 460:91–100, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m460p091.pdf.

Marriott:2016:DNG

[MMRC16] A. L. Marriott, I. D. McCarthy, A. L. Ramsay, and S. R. N. Chenery. Discriminating nursery grounds of juvenile plaice (Pleuronectes platessa) in the south-eastern Irish Sea using otolith microchemistry. Marine Ecology Progress Series, 546: 183–195, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/546/m546p183.pdf.

Mercado-Molina:2018:TLR

[MMRDS18] A. E. Mercado-Molina, C. P. Ruiz-Diaz, and A. M. Sabat. Tissue loss rather than colony size determines the demographic fate of the branching coral Acropora cervicornis. Marine Ecology Progress Series, 597:147–159, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m597p147.pdf.

McHuron:2017:EPR

[MMSC17] E. A. McHuron, M. Mangel, L. K. Schwarz, and D. P. Costa. Energy and prey requirements of California sea lions under variable environmental conditions. Marine Ecology Progress Series, 567:235–247, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/567/m567p235.pdf.

Miller:2013:IVA

[MMWG13a] S. H. Miller, S. G. Morgan, J. W. White, and P. G. Green. Interannual variability in an atlas of trace element signatures for determining population connectivity. Marine Ecology Progress Series, 474:179–190, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m474p179.pdf.

Miller:2013:TES

[MMWG13b] S. H. Miller, S. G. Morgan, J. W. White, and P. G. Green. Trace element signatures in larval soft tissues reveal transport, but not population connectivity. Marine Ecology Progress Series, 481:1–10, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Munroe:2010:PBF

[MN10] D. M. Munroe and T. Noda. Physical and biological factors contributing to changes in the relative importance of recruitment to population dynamics in open populations. Marine Ecology Progress Series, 412:151–162, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/412/
m412p151.pdf.

McCabe:2018:RIR

[MN18] M. K. McCabe and S. A. Navarrete. Reproductive investment in rocky intertidal mussels: spatiotemporal variability and environmental determinants. Marine Ecology Progress Series, 599:107–124, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/599/m599p107.pdf.

McClain:2011:LSF

[MNKB11] C. R. McClain, J. C. Nekola, L. Kuhnz, and J. P. Barry. Local-scale faunal turnover on the deep Pacific seafloor. Marine Ecology Progress Series, 422:193–200, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/422/
m422p193.pdf.

Marion:2012:SEE

[MO12] S. R. Marion and R. J. Orth. Seedling establishment in eelgrass: seed burial effects on winter losses of developing seedlings. Marine Ecology Progress Series, 448:197–207, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m448p197.pdf.

Murray:2013:ERL

[MO13] K. T. Murray and C. D. Orphanides. Estimating the risk of loggerhead turtle Caretta caretta bycatch in the US mid-Atlantic using fishery-independent and -dependent data. Marine Ecology Progress Series, 477:259–270, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/477/
m477p259.pdf.

Munguia:2010:MPE

[MOH^{+}10] Pablo Munguia, Richard W. Osman, John Hamilton, Robert B. Whitlatch, and Roman N. Zajac. Modeling of priority effects and species dominance in Long Island Sound benthic communities. Marine Ecology Progress Series, 413: 229–240, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m413p229.pdf. See comment [LGL11] and

reply [MHO^{+}11].

Moland:2011:APW

[MOK^{+}11] E. Moland, E. Moland Olsen, H. Knutsen, J. A. Knutsen, S. E. Enersen, C. André, and N. C. Stenseth. Activity patterns of wild European lobster Homarus gammarus in coastal marine reserves: implications for future reserve design. Marine Ecology Progress Series, 429:197–207, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/429/
m429p197.pdf.

Muko:2018:LTC

[MOK^{+}18] S. Muko, S. Ohshimo, H. Kurota, T. Yasuda, and M. A. Fukuwaka. Long-term change in the distribution of Japanese sardine in the Sea of Japan during population fluctuations. Marine Ecology Progress Series, 593:141–154, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/593/
m593p141.pdf.

Mathies:2014:EIF

[MOMF14] N. H. Mathies, M. B. Ogburn, G. McFall, and S. Fangman. Environmental interference factors affecting detection range in acoustic telemetry studies using fixed receiver arrays. Marine Ecology Progress Series, 495:27–38, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/495/
m495p027.pdf.

Martin:2015:EPF

[MOP^{+}15] T. S. H. Martin, A. D. Olds, K. A. Pitt, A. B. Johnston, I. R. Butler, P. S. Maxwell, and R. M. Connolly. Effective protection of fish on inshore coral reefs depends on the scale of mangrove–reef connectivity. Marine Ecology Progress Series, 527:157–165, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/527/m527p157.pdf.

Manley:2015:RDP

[MORM15] S. R. Manley, R. J. Orth, and L. Ruiz-Montoya. Roles of dispersal and predation in determining seedling recruitment patterns in a foundational marine angiosperm. Marine Ecology Progress Series, 533:109–120, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/533/m533p109.pdf.

Moland:2010:MIO

[MOS10] E. Moland, E. Moland Olsen, and N. C. Stenseth. Maternal influences on offspring size variation and viability in wild European lobster Homarus gammarus. Marine Ecology Progress Series, 400:165–173, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/400/m400p165.pdf.

Mouritsen:2017:PRP

[Mou17] K. N. Mouritsen. Periwinkle regulation: parasitism and epibiosis are linked. Marine Ecology Progress Series, 579: 227–231, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/579/m579p227.pdf.

Mikkelsen:2012:IRK

[MP12] N. Mikkelsen and T. Pedersen. Invasive red king crab affects lumpsucker recruitment by egg consumption. Marine Ecology Progress Series, 469:87–99, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m469p087.pdf.

MacLeod:2015:IEP

[MP15] C. D. MacLeod and R. Poulin. Interactive effects of parasitic infection and ocean acidification on the calcification of a marine gastropod. Marine Ecology Progress Series, 537: 137–150, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/537/m537p137.pdf.

Mikkelsen:2017:IRK

[MP17] N. Mikkelsen and T. Pedersen. Invasive red king crabs feed on both spawned-out capelin and their eggs. Marine Ecology Progress Series, 563:139–155, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m563p139.pdf.

March:2010:STR

[MPA^{+}10] D. March, M. Palmer, J. Alós, A. Grau, and F. Cardona. Short-term residence, home range size and diel patterns of the painted comber Serranus scriba in a temperate marine reserve. Marine Ecology Progress Series, 400:195–206, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/400/
m400p195.pdf.

Milonas:2010:LIF

[MPB10] L. Milonas, B. Pernet, and B. L. Bingham. Light influences feeding and growth of echinoplutei. Marine Ecology Progress Series, 404:69–78, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/404/m404p069.pdf.

Miller:2013:POM

[MPB13] R. J. Miller, H. M. Page, and M. A. Brzezinski. δ ^{13}C and δ ^{15}N of particulate organic matter in the Santa Barbara Channel: drivers and implications for trophic inference. Marine Ecology Progress Series, 474:53–66, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/474/m474p053.pdf.

Manning:2014:DMP

[MPB14a] L. M. Manning, C. H. Peterson, and M. J. Bishop. Dominant macrobenthic populations experience sustained impacts from annual disposal of fine sediments on sandy beaches. Marine Ecology Progress Series, 508:1–15, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Morello:2014:MMR

[MPB^{+}14b] E. B. Morello, É. E. Plagányi, R. C. Babcock, H. Sweatman, R. Hillary, and A. E. Punt. Model to manage and reduce crown-of-thorns starfish outbreaks. Marine Ecology Progress Series, 512:167–183, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m512p167.pdf.

Mathews:2015:MPP

[MPB15a] A. L. Mathews, E. J. Phlips, and S. Badylak. Modeling phytoplankton productivity in a shallow, microtidal, subtropical estuary. Marine Ecology Progress Series, 531:63– 80, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/531/m531p063.pdf.

Monk:2015:SPS

[MPB15b] M. H. Monk, J. E. Powers, and E. N. Brooks. Spatial patterns in species assemblages associated with the northwestern Gulf of Mexico shrimp trawl fishery. Marine Ecology Progress Series, 519:1–12, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Moerdijk-Poortvliet:2018:PCE

[MPBSB18] T. C. W. Moerdijk-Poortvliet, O. Beauchard, L. J. Stal, and H. T. S. Boschker. Production and consumption of extracellular polymeric substances in an intertidal diatom mat. Marine Ecology Progress Series, 592:77–95, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/592/
m592p077.pdf.

Morsan:2011:GCR

[MPD11] E. M. Morsan, P. Pappalardo, and M. S. Doldan. Growth compensation as a regulatory mechanism of purple clam Amiantis purpurata population dynamics in Patagonia. Marine Ecology Progress Series, 443:207–216, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/443/m443p207.pdf.

Munroe:2015:OAS

[MPF^{+}15] D. M. Munroe, E. N. Powell, S. E. Ford, E. E. Hofmann, and J. M. Klinck. Outcomes of asymmetric selection pressure and larval dispersal on evolution of disease resistance: a metapopulation modeling study with oysters. Marine Ecology Progress Series, 531:221–239, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2015/531/m531p221.pdf.

McLeay:2010:FBH

[MPG^{+}10] L. J. McLeay, B. Page, S. D. Goldsworthy, D. C. Paton, C. Teixeira, P. Burch, and T. Ward. Foraging behaviour and habitat use of a short-ranging seabird, the crested tern. Marine Ecology Progress Series, 411:271–283, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/411/
m411p271.pdf.

Moraes:2012:DRF

[MPG^{+}12] L. E. Moraes, E. Paes, A. Garcia, O. Möller, Jr., and J. Vieira. Delayed response of fish abundance to environmental changes: a novel multivariate time-lag approach. Marine Ecology Progress Series, 456:159–168, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/456/m456p159.pdf.

MacArthur:2011:HSP

[MPH^{+}11] L. D. MacArthur, D. L. Phillips, G. A. Hyndes, C. E. Hanson, and M. A. Vanderklift. Habitat surrounding patch reefs influences the diet and nutrition of the western rock lobster. Marine Ecology Progress Series, 436:191–205, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/436/
m436p191.pdf.

Manderson:2011:OOD

[MPKO11] J. Manderson, L. Palamara, J. Kohut, and M. J. Oliver. Ocean observatory data are useful for regional habitat modeling of species with different vertical habitat preferences. Marine Ecology Progress Series, 438:1–17, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Marino:2017:LVM

[MPLLE17] C. M. Marino, J. R. Pawlik, S. López-Legentil, and P. M. Erwin. Latitudinal variation in the microbiome of the sponge Ircinia campana correlates with host haplotype but not antipredatory chemical defense. Marine Ecology Progress Series, 565:53–66, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m565p053.pdf.

Marques:2014:SEZ

[MPM^{+}14] S. C. Marques, A. L. Primo, F. Martinho, U. M. Azeiteiro, and M. Â. Pardal. Shifts in estuarine zooplankton variability following extreme climate events: a comparison between drought and regular years. Marine Ecology Progress Series, 499:65–76, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/499/m499p065.pdf.

Mattos:2018:CIC

[MPM^{+}18] G. Mattos, P. C. Paiva, M. Mateos, P. A. Haye, and L. A. Hurtado. The cost of ignoring cryptic diversity in macroecological studies: Comment on Martínez et al. (2017). Marine Ecology Progress Series, 601:269–271, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m601p269.pdf. See

[?].

Morera-Pujol:2018:MIA

[MPRPM^{+}18] V. Morera-Pujol, R. Ramos, N. Pérez-Méndez, M. Cerdà-Cuéllar, and J. González-Solís. Multi-isotopic assessments of spatio-temporal diet variability: the case of two sympatric gulls in the western Mediterranean. Marine Ecology Progress Series, 606:201–214, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/606/m606p201.pdf.

Mackley:2010:FBA

[MPS^{+}10] E. K. Mackley, R. A. Phillips, J. R. D. Silk, E. D. Wakefield, V. Afanasyev, J. W. Fox, and R. W. Furness. Free as a bird? Activity patterns of albatrosses during the nonbreeding period. Marine Ecology Progress Series, 406:291–303, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/406/
m406p291.pdf.

Mariani:2012:INA

[MPW12] S. Mariani, K. T. C. A. Peijnenburg, and D. Weetman. Independence of neutral and adaptive divergence in a low dispersal marine mollusc. Marine Ecology Progress Series, 446: 173–187, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/446/m446p173.pdf.

Maps:2010:CDL

[MPZ10] F. Maps, S. Plourde, and B. Zakardjian. Control of dormancy by lipid metabolism in Calanus finmarchicus: a population model test. Marine Ecology Progress Series, 403: 165–180, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m403p165.pdf.

Morse:2016:MAL

[MR16] B. Morse and R. Rochette. Movements and activity levels of juvenile American lobsters Homarus americanus in nature quantified using ultrasonic telemetry. Marine Ecology Progress Series, 551:155–170, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/551/m551p155.pdf.

Miller:2011:BTR

[MRB^{+}11] C. A. Miller, D. Reeb, P. B. Best, A. R. Knowlton, M. W. Brown, and M. J. Moore. Blubber thickness in right whales Eubalaena glacialis and Eubalaena australis related with reproduction, life history status and prey abundance. Marine Ecology Progress Series, 438:267–283, 2011. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m438p267.pdf.

Mostajir:2015:MFW

[MRB^{+}15] B. Mostajir, C. Roques, C. Bouvier, T. Bouvier, É. Fouilland, P. Got, E. Le Floc’h, J. Nouguier, S. Mas, R. Sempéré, T. Sime-Ngando, M. Troussellier, and F. Vidussi. Microbial food web structural and functional responses to oyster and fish as top predators. Marine Ecology Progress Series, 535:11–27, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m535p011.pdf.

McConnell:2010:EAL

[MRC10a] A. McConnell, R. Routledge, and B. M. Connors. Effect of artificial light on marine invertebrate and fish abundance in an area of salmon farming. Marine Ecology Progress Series, 419:147–156, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/419/m419p147.pdf.

Morgan:2010:IET

[MRC10b] M. J. Morgan, R. M. Rideout, and E. B. Colbourne. Impact of environmental temperature on Atlantic cod Gadus morhua energy allocation to growth, condition and reproduction. Marine Ecology Progress Series, 404:185–195, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/404/
m404p185.pdf.

McGlathery:2012:RTD

[MRC^{+}12] K. J. McGlathery, L. K. Reynolds, L. W. Cole, R. J. Orth, S. R. Marion, and A. Schwarzschild. Recovery trajectories during state change from bare sediment to eelgrass dominance. Marine Ecology Progress Series, 448:209–221, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m448p209.pdf.

Muhling:2013:CBE

[MRC^{+}13] B. A. Muhling, P. Reglero, L. Ciannelli, D. Alvarez-Berastegui, F. Alemany, J. T. Lamkin, and M. A. Roffer. Comparison between environmental characteristics of larval bluefin tuna Thunnus thynnus habitat in the Gulf of Mexico and western Mediterranean Sea. Marine Ecology Progress Series, 486:257–276, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/486/m486p257.pdf.

Munnelly:2019:HSO

[MRC^{+}19] R. T. Munnelly, D. B. Reeves, E. J. Chesney, D. M. Baltz, and B. D. Marx. Habitat suitability for oil and gas platform-associated fishes in Louisiana’s nearshore waters. Marine Ecology Progress Series, 608:199–219, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/608/
m608p199.pdf.

Marty:2014:TTA

[MRE14] L. Marty, M. J. Rochet, and B. Ernande. Temporal trends in age and size at maturation of four North Sea gadid species: cod, haddock, whiting and Norway pout. Marine Ecology Progress Series, 497:179–197, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/497/m497p179.pdf.

Meerhoff:2013:SPM

[MRGG^{+}13] E. Meerhoff, L. Rodríguez-Gallego, L. Giménez, P. Muniz, and D. Conde. Spatial patterns of macrofaunal community structure in coastal lagoons of Uruguay. Marine Ecology Progress Series, 492:97–110, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/492/m492p097.pdf.

Mizerek:2011:SHL

[MRH11] T. Mizerek, H. M. Regan, and K. A. Hovel. Seagrass habitat loss and fragmentation influence management strategies for a blue crab Callinectes sapidus fishery. Marine Ecology Progress Series, 427:247–257, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m427p247.pdf.

Moran:2010:PIN

[MRL^{+}10] E. R. Moran, P. L. Reynolds, L. M. Ladwig, M. I. O’Connor, Z. T. Long, and J. F. Bruno. Predation intensity is negatively related to plant species richness in a benthic marine community. Marine Ecology Progress Series, 400:277– 282, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/400/m400p277.pdf.

McInnes:2017:SPF

[MRL^{+}17] A. M. McInnes, P. G. Ryan, M. Lacerda, J. Deshayes, W. S. Goschen, and L. Pichegru. Small pelagic fish responses to fine-scale oceanographic conditions: implications for the endangered African penguin. Marine Ecology Progress Series, 569:187–203, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/569/m569p187.pdf.

McGovern:2019:DCB

[MRLD19] K. A. McGovern, D. H. Rodríguez, M. N. Lewis, and R. W. Davis. Diving classification and behavior of free-ranging female southern elephant seals based on three-dimensional movements and video-recorded observations. Marine Ecology Progress Series, 620:215–232, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/620/m620p215.pdf.

McFarland:2015:IPS

[MRSD15] M. N. McFarland, J. Rines, J. Sullivan, and P. Donaghay. Impact of phytoplankton size and physiology on particulate optical properties determined with scanning flow cytometry.

Marine Ecology Progress Series, 531:43–61, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/531/
m531p043.pdf.

Murray:2010:DWC

[MS10] L. G. Murray and R. Seed. Determining whether catch per unit effort is a suitable proxy for relative crab abundance. Marine Ecology Progress Series, 401:173–182, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/401/
m401p173.pdf.

Mora:2011:OGB

[MS11] C. Mora and P. F. Sale. Ongoing global biodiversity loss and the need to move beyond protected areas: a review of the technical and practical shortcomings of protected areas on land and sea. Marine Ecology Progress Series, 434:251–266, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m434p251.pdf.

Miller:2013:SSV

[MS13] N. A. Miller and J. H. Stillman. Seasonal and spatial variation in the energetics of the invasive clam Corbula amurensis in the upper San Francisco Estuary. Marine Ecology Progress Series, 476:129–139, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m476p129.pdf.

Moore:2014:ACT

[MS14] B. R. Moore and C. A. Simpfendorfer. Assessing connectivity of a tropical estuarine teleost through otolith elemental profiles. Marine Ecology Progress Series, 501:225– 238, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/501/m501p225.pdf.

Michael:2018:FYD

[MS18a] K. P. Michael and J. S. Shima. Four-year decline in Ostrea chilensis recruits per spawner in Foveaux Strait, New Zealand, suggests a diminishing stock-recruitment relationship. Marine Ecology Progress Series, 600:85–98, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/600/
m600p085.pdf.

Moginie:2018:HDG

[MS18b] B. F. Moginie and J. S. Shima. Hatch date and growth rate drives reproductive success in nest-guarding males of a temperate reef fish. Marine Ecology Progress Series, 592: 197–206, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/592/m592p197.pdf.

Malpezzi:2013:ADT

[MSC13] M. A. Malpezzi, L. P. Sanford, and B. C. Crump. Abundance and distribution of transparent exopolymer particles in the estuarine turbidity maximum of Chesapeake Bay. Marine Ecology Progress Series, 486:23–35, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/486/
m486p023.pdf.

Mumby:2012:FCF

[MSE^{+}12] P. J. Mumby, R. S. Steneck, A. J. Edwards, R. Ferrari, R. Coleman, A. R. Harborne, and J. P. Gibson. Fishing down a Caribbean food web relaxes trophic cascades. Marine Ecology Progress Series, 445:13–24, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m445p013.pdf.

McMurray:2018:TSL

[MSE^{+}18] S. E. McMurray, A. D. Stubler, P. M. Erwin, C. M. Finelli, and J. R. Pawlik. A test of the sponge-loop hypothesis for emergent Caribbean reef sponges. Marine Ecology Progress Series, 588:1–14, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Madigan:2018:MMT

[MSF18] D. J. Madigan, O. E. Snodgrass, and N. S. Fisher. From migrants to mossbacks: tracer- and tag-inferred habitat shifts in the California yellowtail Seriola dorsalis. Marine Ecology Progress Series, 597:221–230, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/597/m597p221.pdf.

McClanahan:2019:GBB

[MSF^{+}19] T. R. McClanahan, R. E. Schroeder, A. M. Friedlander, L. Vigliola, L. Wantiez, J. E. Caselle, N. A. J. Graham, S. Wilson, G. J. Edgar, R. D. Stuart-Smith, R. M. Oddenyo, and J. E. Cinner. Global baselines and benchmarks for fish biomass: comparing remote reefs and fisheries closures. Marine Ecology Progress Series, 612:167–192, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m612p167.pdf.

Macedo-Soares:2012:SSS

[MSFM12] L. C. P. Macedo-Soares, A. S. Freire, and J. H. Muelbert. Small-scale spatial and temporal variability of larval fish assemblages at an isolated oceanic island. Marine Ecology Progress Series, 444:207–222, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/444/m444p207.pdf.

Meager:2011:TCL

[MSG11] J. J. Meager, T. A. Schlacher, and M. Green. Topographic complexity and landscape temperature patterns create a dynamic habitat structure on a rocky intertidal shore. Marine Ecology Progress Series, 428:1–12, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Melia:2013:IFD

[MSG^{+}13] P. Melià, M. Schiavina, M. Gatto, L. Bonaventura, S. Masina, and R. Casagrandi. Integrating field data into individualbased models of the migration of European eel larvae. Marine Ecology Progress Series, 487:135–149, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/487/
m487p135.pdf.

Macias:2010:PCZ

[MSGGE10] D. Macías, R. Somavilla, J. I. González-Gordillo, and F. Echevarría. Physical control of zooplankton distribution at the Strait of Gibraltar during an episode of internal wave generation. Marine Ecology Progress Series, 408: 79–95, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/408/m408p079.pdf.

McLean:2014:DBP

[MSH^{+}14] M. F. McLean, C. A. Simpfendorfer, M. R. Heupel, M. J. Dadswell, and M. J. W. Stokesbury. Diversity of behavioural patterns displayed by a summer feeding aggregation of Atlantic sturgeon in the intertidal region of Minas Basin, Bay of Fundy, Canada. Marine Ecology Progress Series, 496: 59–69, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/496/m496p059.pdf.

Murphy:2019:ACS

[MSK^{+}19] K. J. Murphy, D. Sephton, K. Klein, C. D. Bishop, and R. C. Wyeth. Abiotic conditions are not sufficient to predict spatial and interannual variation in abundance of Ciona intestinalis in Nova Scotia, Canada. Marine Ecology Progress Series, 628:105–123, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/628/m628p105.pdf.

Mills:2013:BAP

[MSLR13] E. Mills, K. Shechtman, Y. Loya, and E. Rosenberg. Bacteria appear to play important roles in both causing and preventing the bleaching of the coral Oculina patagonica. Marine Ecology Progress Series, 489:155–162, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/489/m489p155.pdf.

Moreau:2012:IMC

[MSM^{+}12] S. Moreau, I. R. Schloss, B. Mostajir, S. Demers, G. O. Almandoz, M. E. Ferrario, and G. A. Ferreyra. Influence of microbial community composition and metabolism on air–sea Δ pCO_{2} variation off the western Antarctic Peninsula. Marine Ecology Progress Series, 446:45–59, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/446/m446p045.pdf.

Moore:2012:EST

[MSPO12] K. A. Moore, E. C. Shields, D. B. Parrish, and R. J. Orth. Eelgrass survival in two contrasting systems: role of turbidity and summer water temperatures. Marine Ecology Progress Series, 448:247–258, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m448p247.pdf.

Mercado:2014:ECC

[MSS^{+}14a] J. M. Mercado, I. Sala, S. Salles, D. Cortés, T. Ramírez, E. Liger, L. Yebra, and B. Bautista. Effects of community composition and size structure on light absorption and nutrient uptake of phytoplankton in contrasting areas of the Alboran Sea. Marine Ecology Progress Series, 499:47–64, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/499/
m499p047.pdf.

Muto:2014:BMT

[MSS^{+}14b] E. Y. Muto, L. S. H. Soares, J. E. S. Sarkis, M. A. Hortellani, M. A. V. Petti, and T. N. Corbisier. Biomagnification of mercury through the food web of the Santos continental shelf, subtropical Brazil. Marine Ecology Progress Series, 512:55–69, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m512p055.pdf.

Mouton:2011:ABM

[MSV^{+}11] A. M. Mouton, M. Stevens, T. Van den Neucker, D. Buysse, and J. Coeck. Adjusted barrier management to improve glass eel migration at an estuarine barrier. Marine Ecology Progress Series, 439:213–222, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2011/439/m439p213.pdf.

Michael:2017:IFB

[MTB^{+}17] P. E. Michael, R. Thomson, C. Barbraud, K. Delord, S. De Grissac, A. J. Hobday, P. G. Strutton, G. N. Tuck, H. Weimerskirch, and C. Wilcox. Illegal fishing bycatch overshadows climate as a driver of albatross population decline. Marine Ecology Progress Series, 579:185–199, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/579/
m579p185.pdf.

Martin:2010:PSP

[MTG^{+}10] J. Martin, J. É. Tremblay, J. Gagnon, G. Tremblay, A. Lapoussière, C. Jose, M. Poulin, M. Gosselin, Y. Gratton, and C. Michel. Prevalence, structure and properties of subsurface chlorophyll maxima in Canadian Arctic waters. Marine Ecology Progress Series, 412:69–84, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m412p069.pdf.

Muths:2011:RDR

[MTG^{+}11] D. Muths, E. Tessier, G. Gouws, M. Craig, M. Mwale, J. Mwaluma, A. Mwandya, and J. Bourjea. Restricted dispersal of the reef fish Myripristis berndti at the scale of the SW Indian Ocean. Marine Ecology Progress Series, 443:167– 180, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/443/m443p167.pdf.

Muller:2015:DRT

[MTH15] M. N. Müller, T. W. Trull, and G. M. Hallegraeff. Differing responses of three Southern Ocean Emiliania huxleyi ecotypes to changing seawater carbonate chemistry. Marine Ecology Progress Series, 531:81–90, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/531/m531p081.pdf.

Miranda:2018:IIC

[MTK^{+}18] R. J. Miranda, A. Tagliafico, B. P. Kelaher, E. Mariano-Neto, and F. Barros. Impact of invasive corals Tubastraea spp. on native coral recruitment. Marine Ecology Progress Series, 605:125–133, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/605/m605p125.pdf.

McLenaghan:2011:BMF

[MTM^{+}11] N. A. McLenaghan, A. C. Tyler, U. H. Mahl, R. W. Howarth, and R. M. Marino. Benthic macroinvertebrate functional diversity regulates nutrient and algal dynamics in a shallow estuary. Marine Ecology Progress Series, 426:171–184, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2011/426/m426p171.pdf.

Manriquez:2017:EOW

[MTM^{+}17] P. H. Manríquez, R. Torres, P. G. Matson, M. R. Lee, M. E. Jara, M. E. Seguel, F. Sepúlveda, and L. Pereira. Effects of ocean warming and acidification on the early benthic ontogeny of an ecologically and economically important echinoderm.

Marine Ecology Progress Series, 563:169–184, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/563/
m563p169.pdf.

Melbourne-Thomas:2015:AEA

[MTMM^{+}15] J. Melbourne-Thomas, K. M. Meiners, C. J. Mundy, C. Schallenberg, K. L. Tattersall, and G. S. Dieckmann. Algorithms to estimate Antarctic sea ice algal biomass from under-ice irradiance spectra at regional scales. Marine Ecology Progress Series, 536:107–121, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/536/m536p107.pdf.

McIntyre:2010:SSD

[MTP^{+}10] T. McIntyre, C. A. Tosh, J. Plötz, H. Bornemann, and M. N. Bester. Segregation in a sexually dimorphic mammal: a mixed-effects modelling analysis of diving behaviour in southern elephant seals. Marine Ecology Progress Series, 412:293–304, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/412/m412p293.pdf. See comment

[BSBR^{+}12].

Matley:2017:TNS

[MTS^{+}17] J. K. Matley, A. J. Tobin, C. A. Simpfendorfer, A. T. Fisk, and M. R. Heupel. Trophic niche and spatio-temporal changes in the feeding ecology of two sympatric species of coral trout (Plectropomus leopardus and P. laevis). Marine Ecology Progress Series, 563:197–210, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/563/m563p197.pdf.

Martinetto:2011:TBR

[MTV^{+}11] P. Martinetto, M. Teichberg, I. Valiela, D. Montemayor, and O. Iribarne. Top-down and bottom-up regulation in a high nutrient-high herbivory coastal ecosystem. Marine Ecology Progress Series, 432:69–82, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/432/m432p069.pdf.

Momigliano:2013:SGP

[MU13] P. Momigliano and S. Uthicke. Symbiosis in a giant protist (Marginopora vertebralis, soritinae): flexibility in symbiotic partnerships along a natural temperature gradient. Marine Ecology Progress Series, 491:33–46, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/491/
m491p033.pdf.

Marzinelli:2012:MHC

[MUC12] E. M. Marzinelli, A. J. Underwood, and R. A. Coleman. Modified habitats change ecological processes affecting a nonindigenous epibiont. Marine Ecology Progress Series, 446: 119–129, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m446p119.pdf.

Matias:2011:HII

[MUHC11] M. G. Matias, A. J. Underwood, D. F. Hochuli, and R. A. Coleman. Habitat identity influences species–area relationships in heterogeneous habitats. Marine Ecology Progress Series, 437:135–145, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m437p135.pdf.

Moland:2013:LTD

[MUOS13] E. Moland, M. Ulmestrand, E. M. Olsen, and N. C. Stenseth. Long-term decrease in sex-specific natural mortality of European lobster within a marine protected area. Marine Ecology Progress Series, 491:153–164, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m491p153.pdf.

Musco:2012:EDM

[Mus12] L. Musco. Ecology and diversity of Mediterranean hardbottom syllidae (Annelida): a community-level approach. Marine Ecology Progress Series, 461:107–119, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/461/
m461p107.pdf.

Martin:2012:EMI

[MV12] C. W. Martin and J. F. Valentine. Eurasian milfoil invasion in estuaries: physical disturbance can reduce the proliferation of an aquatic nuisance species. Marine Ecology Progress Series, 449:109–119, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/449/m449p109.pdf.

Madeira:2014:HAP

[MVCD14] D. Madeira, C. Vinagre, P. M. Costa, and M. S. Diniz. Histopathological alterations, physiological limits, and molecular changes of juvenile Sparus aurata in response to thermal stress. Marine Ecology Progress Series, 505:253–266, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2014/505/m505p253.pdf.

Muller:2012:OCL

[MvdHBM12] C. M. Muller, S. von der Heyden, R. C. K. Bowie, and C. A. Matthee. Oceanic circulation, local upwelling and palaeoclimatic changes linked to the phylogeography of the Cape sea urchin Parechinus angulosus. Marine Ecology Progress Series, 468:203–215, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/468/m468p203.pdf.

Martin:2010:STP

[MVE^{+}10] C. S. Martin, S. Vaz, J. R. Ellis, F. Coppin, D. Le Roy, and A. Carpentier. Spatio-temporal patterns in demersal elasmobranchs from trawl surveys in the eastern English Channel (1988–2008). Marine Ecology Progress Series, 417:211– 228, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/417/m417p211.pdf.

Maltseva:2016:MPS

[MVL^{+}16] A. L. Maltseva, M. A. Varfolomeeva, A. A. Lobov, N. A. Mikhailova, P. E. Renaud, A. V. Grishankov, K. Y. Volovik, and A. I. Granovitch. Measuring physiological similarity of closely related littorinid species: a proteomic insight. Marine Ecology Progress Series, 552:177–193, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/552/
m552p177.pdf.

Moreno-Valcarcel:2016:LTE

[MVOPB^{+}16] R. Moreno-Valcárcel, F. J. Oliva-Paterna, S. Bevilacqua, A. Terlizzi, and C. Fernández-Delgado. Long-term effects of tidal restriction on fish assemblages in east Atlantic coastal marshlands. Marine Ecology Progress Series, 543:209–222, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/543/m543p209.pdf.

Marcus:2016:IVD

[MVP^{+}16] L. Marcus, P. Virtue, H. R. Pethybridge, M. G. Meekan, M. Thums, and P. D. Nichols. Intraspecific variability in diet and implied foraging ranges of whale sharks at Ningaloo Reef, Western Australia, from signature fatty acid analysis. Marine Ecology Progress Series, 554:115–128, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/554/
m554p115.pdf.

Merckx:2010:MND

[MVS^{+}10] B. Merckx, M. Van Meirvenne, M. Steyaert, A. Vanreusel, M. Vincx, and J. Vanaverbeke. Mapping nematode diversity in the southern bight of the North Sea. Marine Ecology Progress Series, 406:135–145, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/406/m406p135.pdf.

Martinez-Vicente:2012:CPB

[MVTS^{+}12] V. Martinez-Vicente, G. H. Tilstone, S. Sathyendranath, P. I. Miller, and S. B. Groom. Contributions of phytoplankton and bacteria to the optical backscattering coefficient over the Mid-Atlantic Ridge. Marine Ecology Progress Series, 445: 37–51, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/445/m445p037.pdf.

Madrid-Vera:2010:CTC

[MVVGAV10] J. Madrid-Vera, E. Visauta-Girbau, and H. Aguirre-Villaseñor. Composition of trawl catch fauna off the mouth of the Rio Baluarte, southeastern Gulf of California. Marine Ecology Progress Series, 403:145–153, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/403/m403p145.pdf.

Miller:2016:PLS

[MW16a] C. A. Miller and G. G. Waldbusser. A post-larval stagebased model of hard clam Mercenaria mercenaria development in response to multiple stressors: temperature and acidification severity. Marine Ecology Progress Series, 558: 35–49, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/558/m558p035.pdf.

Mohan:2016:BHN

[MW16b] J. Mohan and B. Walther. Out of breath and hungry: natural tags reveal trophic resilience of Atlantic croaker to hypoxia exposure. Marine Ecology Progress Series, 560:207– 221, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/560/m560p207.pdf.

Makela:2017:IAV

[MWA17] A. Mäkelä, U. Witte, and P. Archambault. Ice algae versus phytoplankton: resource utilization by Arctic deep sea macroinfauna revealed through isotope labelling experiments. Marine Ecology Progress Series, 572:1–18, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Moffitt:2013:AAM

[MWB13] E. A. Moffitt, J. W. White, and L. W. Botsford. Accurate assessment of marine protected area success depends on metric and spatiotemporal scale of monitoring. Marine Ecology Progress Series, 489:17–28, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/489/m489p017.pdf.

McIntyre-Wressnig:2013:NLE

[MWBMH13] A. McIntyre-Wressnig, J. M. Bernhard, D. C. McCorkle, and P. Hallock. Non-lethal effects of ocean acidification on the symbiont-bearing benthic foraminifer Amphistegina gibbosa. Marine Ecology Progress Series, 472:45–60, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/472/
m472p045.pdf.

Melnychuk:2012:ESU

[MWC^{+}12] M. C. Melnychuk, C. J. Walters, V. Christensen, M. L. Bothwell, and D. W. Welch. Effects of solar ultraviolet radiation exposure on early ocean survival and fry-to-smolt growth of juvenile salmon. Marine Ecology Progress Series, 457:251– 264, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m457p251.pdf.

McDuie:2018:ODN

[MWC18] F. McDuie, S. J. Weeks, and B. C. Congdon. Oceanographic drivers of near-colony seabird foraging site use in tropical marine systems. Marine Ecology Progress Series, 589:209–225, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/589/m589p209.pdf.

McLeod:2010:HCS

[MWD10] R. J. McLeod, S. R. Wing, and J. P. Davis. Habitat conversion and species loss alters the composition of carbon sources to benthic communities. Marine Ecology Progress Series, 411:127–136, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/411/m411p127.pdf.

Monaco:2015:SLS

[MWGH15] C. J. Monaco, D. S. Wethey, S. Gulledge, and B. Helmuth. Shore-level size gradients and thermal refuge use in the predatory sea star Pisaster ochraceus: the role of environmental stressors. Marine Ecology Progress Series, 539:191–205, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m539p191.pdf.

Meyer:2018:DPD

[MWHM18] K. S. Meyer, J. D. Wheeler, E. Houlihan, and L. S. Mullineaux. Desperate planktotrophs: decreased settlement selectivity with age in competent eastern oyster Crassostrea virginica larvae. Marine Ecology Progress Series, 599:93– 106, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/599/m599p093.pdf.

Morgan:2011:WST

[MWM^{+}11] S. G. Morgan, J. W. White, S. T. McAfee, S. D. Gaines, and R. J. Schmitt. Weak synchrony in the timing of larval release in upwelling regimes. Marine Ecology Progress Series, 425:103–112, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m425p103.pdf.

Murray:2017:ACE

[MWMD17] F. Murray, S. Widdicombe, C. L. McNeill, and A. Douglas. Assessing the consequences of environmental impacts: variation in species responses has unpredictable functional effects. Marine Ecology Progress Series, 583:35–47, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/583/
m583p035.pdf.

Maier:2010:SIR

[MWP10] C. Maier, M. G. Weinbauer, and J. Pätzold. Stable isotopes reveal limitations in C and N assimilation in the Caribbean reef corals Madracis auretenra, M. carmabi and M. formosa. Marine Ecology Progress Series, 412:103–112, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/412/
m412p103.pdf.

Mansfield:2012:STA

[MWR^{+}12] K. L. Mansfield, J. Wyneken, D. Rittschof, M. Walsh, C. W. Lim, and P. M. Richards. Satellite tag attachment methods for tracking neonate sea turtles. Marine Ecology Progress Series, 457:181–192, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m457p181.pdf.

Mohring:2014:BVT

[MWW^{+}14] M. B. Mohring, T. Wernberg, J. T. Wright, S. D. Connell, and B. D. Russell. Biogeographic variation in temperature drives performance of kelp gametophytes during warming. Marine Ecology Progress Series, 513:85–96, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/513/
m513p085.pdf.

McMullin:2017:TPA

[MWWS17] R. M. McMullin, S. R. Wing, L. C. Wing, and O. A. Shatova. Trophic position of Antarctic ice fishes reflects food web structure along a gradient in sea ice persistence. Marine Ecology Progress Series, 564:87–98, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m564p087.pdf.

McHuron:2016:WGD

[MWZ^{+}16] E. A. McHuron, S. M. Walcott, J. Zeligs, S. Skrovan, D. P. Costa, and C. Reichmuth. Whisker growth dynamics in two North Pacific pinnipeds: implications for determining foraging ecology from stable isotope analysis. Marine Ecology Progress Series, 554:213–224, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/554/m554p213.pdf.

Matsumoto:2016:CRD

[MYK^{+}16] S. Matsumoto, T. Yamamoto, R. Kawabe, S. Ohshimo, and K. Yoda. The Changjiang River discharge affects the distribution of foraging seabirds. Marine Ecology Progress Series, 555:273–277, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/555/m555p273.pdf.

Meyer:2016:RIS

[MYS^{+}16] K. S. Meyer, C. M. Young, A. K. Sweetman, J. Taylor, T. Soltwedel, and M. Bergmann. Rocky islands in a sea of mud: biotic and abiotic factors structuring deep-sea dropstone communities. Marine Ecology Progress Series, 556: 45–57, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/556/m556p045.pdf.

Macaya:2010:GSG

[MZ10] E. C. Macaya and G. C. Zuccarello. Genetic structure of the giant kelp Macrocystis pyrifera along the southeastern Pacific. Marine Ecology Progress Series, 420:103–112, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/420/
m420p103.pdf.

Maldonado:2010:SFS

[MZC^{+}10] M. Maldonado, X. Zhang, X. Cao, L. Xue, H. Cao, and W. Zhang. Selective feeding by sponges on pathogenic microbes: a reassessment of potential for abatement of microbial pollution. Marine Ecology Progress Series, 403:75–89, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m403p075.pdf.

Miller:2019:NWE

[MZF^{+}19] S. D. Miller, B. J. Zgliczynski, M. D. Fox, L. S. Kaufman, R. H. Michener, S. A. Sandin, and S. L. Hamilton. Niche width expansion of coral reef fishes along a primary production gradient in the remote central Pacific. Marine Ecology Progress Series, 625:127–143, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2019/625/m625p127.pdf.

Nash:2016:DHC

[NAG^{+}16] K. L. Nash, R. A. Abesamis, N. A. J. Graham, E. C. McClure, and E. Moland. Drivers of herbivory on coral reefs: species, habitat and management effects. Marine Ecology Progress Series, 554:129–140, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/554/m554p129.pdf.

Nakamura:2011:CMC

[NAY^{+}11] R. Nakamura, W. Ando, H. Yamamoto, M. Kitano, A. Sato, M. Nakamura, H. Kayanne, and M. Omori. Corals masscultured from eggs and transplanted as juveniles to their native, remote coral reef. Marine Ecology Progress Series, 436:161–168, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m436p161.pdf.

Navarro-Barranco:2015:CSP

[NBGGST^{+}15] C. Navarro-Barranco, J. M. Guerra-García, L. Sánchez-Tocino, M. Ros, M. Florido, and J. C. García-Gómez. Colonization and successional patterns of the mobile epifaunal community along an environmental gradient in a marine cave.

Marine Ecology Progress Series, 521:105–115, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/521/
m521p105.pdf.

Neves:2012:FEM

[NBGS^{+}12] V. C. Neves, J. Bried, J. González-Solís, J. L. Roscales, and M. R. Clarke. Feeding ecology and movements of the barolo shearwater Puffinus baroli baroli in the Azores, NE Atlantic. Marine Ecology Progress Series, 452:269–285, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m452p269.pdf.

Nakin:2012:SSE

[NBM12] M. D. V. Nakin, A. J. Booth, and C. D. McQuaid. Speciesspecific effects of marine reserves: mortality and growth differ within and among heavily exploited and rarely exploited limpets. Marine Ecology Progress Series, 445:53– 63, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/445/m445p053.pdf.

Niquil:2014:REF

[NBM^{+}14] N. Niquil, A. Baeta, J. C. Marques, A. Chaalali, J. Lobry, and J. Patrício. Reaction of an estuarine food web to disturbance: Lindeman’s perspective. Marine Ecology Progress Series, 512:141–154, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m512p141.pdf.

Nelson:2011:RSA

[NC11] M. L. Nelson and S. F. Craig. Role of the sea anemone Metridium senile in structuring a developing subtidal fouling community. Marine Ecology Progress Series, 421:139–149, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m421p139.pdf.

Nishizaki:2014:TWF

[NC14] M. T. Nishizaki and E. Carrington. Temperature and water flow influence feeding behavior and success in the barnacle Balanus glandula. Marine Ecology Progress Series, 507:207–218, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/507/m507p207.pdf.

New:2014:UST

[NCC^{+}14] L. F. New, J. S. Clark, D. P. Costa, E. Fleishman, M. A. Hindell, T. Klanjšček, D. Lusseau, S. Kraus, C. R. McMahon, P. W. Robinson, R. S. Schick, L. K. Schwarz, S. E. Simmons, L. Thomas, P. Tyack, and J. Harwood. Using short-term measures of behaviour to estimate long-term fitness of southern elephant seals. Marine Ecology Progress Series, 496:99–108, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/496/m496p099.pdf.

Naesje:2012:RTE

[NCD^{+}12] T. F. Næsje, P. D. Cowley, O. H. Diserud, A. R. Childs, S. E. Kerwath, and E. B. Thorstad. Riding the tide: estuarine movements of a sciaenid fish, Argyrosomus japonicus. Marine Ecology Progress Series, 460:221–232, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m460p221.pdf.

Nosal:2014:ABS

[NCK^{+}14] A. P. Nosal, A. Caillat, E. K. Kisfaludy, M. A. Royer, and N. C. Wegner. Aggregation behavior and seasonal philopatry in male and female leopard sharks Triakis semifasciata along the open coast of southern California, USA. Marine Ecology Progress Series, 499:157–175, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/499/m499p157.pdf.

Norderhaug:2014:IWC

[NCR^{+}14] K. M. Norderhaug, H. Christie, E. Rinde, H. Gundersen, and T. Bekkby. Importance of wave and current exposure to fauna communities in Laminaria hyperborea kelp forests. Marine Ecology Progress Series, 502:295–301, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m502p295.pdf.

Ng:2010:CAL

[NCS^{+}10] T. Y. T. Ng, C. Y. Chuang, I. Stupakoff, A. E. Christy, D. P. Cheney, and W. X. Wang. Cadmium accumulation and loss in the Pacific oyster Crassostrea gigas along the west coast of the USA. Marine Ecology Progress Series, 401:147–160, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/401/m401p147.pdf.

Nordstrom:2014:BFW

[NCT^{+}14] M. C. Nordström, C. A. Currin, T. S. Talley, C. R. Whitcraft, and L. A. Levin. Benthic food-web succession in a developing salt marsh. Marine Ecology Progress Series, 500: 43–55, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/500/m500p043.pdf.

Nosal:2019:MEY

[NCW^{+}19] A. P. Nosal, D. P. Cartamil, N. C. Wegner, C. H. Lam, and P. A. Hastings. Movement ecology of young-of-the-year blue sharks Prionace glauca and shortfin makos Isurus oxyrinchus within a putative binational nursery area. Marine Ecology Progress Series, 623:99–115, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/623/m623p099.pdf.

Nepper-Davidsen:2019:ESH

[NDAP19] J. Nepper-Davidsen, D. T. Andersen, and M. F. Pedersen. Exposure to simulated heatwave scenarios causes long-term reductions in performance in Saccharina latissima. Marine Ecology Progress Series, 630:25–39, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m630p025.pdf.

Neumann:2013:CCF

[NdBKR13] H. Neumann, I. de Boois, I. Kröncke, and H. Reiss. Climate change facilitated range expansion of the non-native angular crab Goneplax rhomboides into the North Sea. Marine Ecology Progress Series, 484:143–153, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/484/m484p143.pdf.

Norrie:2016:LSV

[NDBL16] C. R. Norrie, B. J. Dunphy, J. A. Baker, and C. J. Lundquist. Local-scale variation in trace elemental fingerprints of the estuarine bivalve Austrovenus stutchburyi within and between estuaries. Marine Ecology Progress Series, 559:89–102, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m559p089.pdf.

Nolasco:2013:MDC

[NDD^{+}13] R. Nolasco, J. Dubert, C. P. Domingues, A. Cordeiro Pires, and H. Queiroga. Model-derived connectivity patterns along the western Iberian Peninsula: asymmetrical larval flow and source-sink cell. Marine Ecology Progress Series, 485:123– 142, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/485/m485p123.pdf.

Nelson:2015:DST

[NDG15] J. A. Nelson, L. Deegan, and R. Garritt. Drivers of spatial and temporal variability in estuarine food webs. Marine Ecology Progress Series, 533:67–77, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/533/m533p067.pdf.

Newland:2011:DJS

[NFC^{+}11] C. Newland, I. C. Field, Y. Cherel, C. Guinet, C. J. A. Bradshaw, C. R. McMahon, and M. A. Hindell. Diet of juvenile southern elephant seals reappraised by stable isotopes in whiskers. Marine Ecology Progress Series, 424:247– 258, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/424/m424p247.pdf.

Narvaez:2015:TFC

[NFN^{+}15] P. Narvaez, M. Furtado, A. I. Neto, I. Moniz, J. M. N. Azevedo, and M. C. Soares. Temperate facultative cleaner wrasses selectively remove ectoparasites from their client-fish in the Azores. Marine Ecology Progress Series, 540:217– 226, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/540/m540p217.pdf.

Napoleon:2014:DPD

[NFR^{+}14] C. Napoléon, L. Fiant, V. Raimbault, P. Riou, and P. Claquin. Dynamics of phytoplankton diversity structure and primary productivity in the English Channel. Marine Ecology Progress Series, 505:49–64, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/505/m505p049.pdf.

Nikula:2010:CDR

[NFSW10] R. Nikula, C. I. Fraser, H. G. Spencer, and J. M. Waters. Circumpolar dispersal by rafting in two subantarctic kelp-dwelling crustaceans. Marine Ecology Progress Series, 405:221–230, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/405/m405p221.pdf.

Nickerson:2018:SFT

[NGA18] K. J. Nickerson, T. M. Grothues, and K. W. Able. Sensitivity of a fish time-series analysis to guild construction: a case study of the Mullica River–Great Bay ecosystem. Marine Ecology Progress Series, 598:113–129, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/598/m598p113.pdf.

Nash:2012:IHC

[NGJHB12] K. L. Nash, N. A. J. Graham, F. A. Januchowski-Hartley, and D. R. Bellwood. Influence of habitat condition and competition on foraging behaviour of parrotfishes. Marine Ecology Progress Series, 457:113–124, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/457/m457p113.pdf.

Norderhaug:2015:ECE

[NGP^{+}15] K. M. Norderhaug, H. Gundersen, A. Pedersen, F. Moy, N. Green, M. G. Walday, J. K. Gitmark, A. B. Ledang, B. Bjerkeng, D. Ø. Hjermann, and H. C. Trannum. Effects of climate and eutrophication on the diversity of hard bottom communities on the Skagerrak coast 1990–2010. Marine Ecology Progress Series, 530:29–46, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m530p029.pdf.

Navarro:2017:SIH

[NGR^{+}17] J. Navarro, D. Grémillet, F. J. Ramirez, I. Afán, W. Bouten, and M. G. Forero. Shifting individual habitat specialization of a successful predator living in anthropogenic landscapes. Marine Ecology Progress Series, 578:243–251, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m578p243.pdf.

Narvarte:2013:ECE

[NGSF13] M. Narvarte, R. A. González, L. Storero, and M. Fernández. Effects of competition and egg predation on shelter use by Octopus tehuelchus females. Marine Ecology Progress Series, 482:141–151, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/482/m482p141.pdf.

Nishimoto:2015:EAU

[NHAS15] A. Nishimoto, T. Haga, A. Asakura, and Y. Shirayama. An experimental approach for understanding the process of wood fragmentation by marine wood borers in shallow temperate waters. Marine Ecology Progress Series, 538:53–65, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2015/538/m538p053.pdf.

Newman:2010:DPP

[NHG10] S. P. Newman, R. D. Handy, and S. H. Gruber. Diet and prey preference of juvenile lemon sharks Negaprion brevirostris. Marine Ecology Progress Series, 398:221–234, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/398/
m398p221.pdf.

Nicolet:2018:SFC

[NHPW18] K. J. Nicolet, M. O. Hoogenboom, M. S. Pratchett, and B. L. Willis. Selective feeding by corallivorous fishes neither promotes nor reduces progression rates of black band disease. Marine Ecology Progress Series, 594:95–106, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/594/
m594p095.pdf.

Nagelkerken:2017:HLR

[NHS^{+}17] I. Nagelkerken, K. B. Huebert, J. E. Serafy, M. G. G. Grol, M. Dorenbosch, and C. J. A. Bradshaw. Highly localized replenishment of coral reef fish populations near nursery habitats. Marine Ecology Progress Series, 568:137–150, 2017. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m568p137.pdf.

Neal:2011:PSD

[NHSG11] L. Neal, S. L. Mincks Hardy, C. R. Smith, and A. G. Glover. Polychaete species diversity on the West Antarctic Peninsula deep continental shelf. Marine Ecology Progress Series, 428:119–134, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/428/m428p119.pdf.

Nielsen:2014:CPH

[NHTS14] J. K. Nielsen, P. N. Hooge, S. J. Taggart, and A. C. Seitz. Characterizing Pacific halibut movement and habitat in a marine protected area using net squared displacement analysis methods. Marine Ecology Progress Series, 517:229–250, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m517p229.pdf.

Nikolioudakis:2012:DSS

[NIPS12] N. Nikolioudakis, S. Isari, P. Pitta, and S. Somarakis. Diet of sardine Sardina pilchardus: an ‘end-to-end’ field study. Marine Ecology Progress Series, 453:173–188, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m453p173.pdf.

Nikolioudakis:2014:TAN

[NIS14] N. Nikolioudakis, S. Isari, and S. Somarakis. Trophodynamics of anchovy in a non-upwelling system: direct comparison with sardine. Marine Ecology Progress Series, 500:215–229, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m500p215.pdf.

Nordeide:2011:PCA

[NJJ^{+}11] J. T. Nordeide, S. D. Johansen, T. E. Jørgensen, B. O. Karlsen, and T. Moum. Population connectivity among migratory and stationary cod Gadus morhua in the Northeast Atlantic—a review of 80 years of study. Marine Ecology Progress Series, 435:269–283, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/435/m435p269.pdf.

Noonan:2012:CSH

[NJP12] S. H. C. Noonan, G. P. Jones, and M. S. Pratchett. Coral size, health and structural complexity: effects on the ecology of a coral reef damselfish. Marine Ecology Progress Series, 456:127–137, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/456/m456p127.pdf.

Norheim:2016:ELC

[NKA16] E. Norheim, T. A. Klevjer, and D. L. Aksnes. Evidence for light-controlled migration amplitude of a sound scattering layer in the Norwegian Sea. Marine Ecology Progress Series, 551:45–52, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m551p045.pdf.

Nakajima:2010:ISS

[NKH^{+}10] M. Nakajima, K. Kodama, T. Horiguchi, Y. Tanaka, and H. Shiraishi. Impacts of shifts in spawning seasonality and size at maturation on the population growth of mantis shrimp in Tokyo Bay. Marine Ecology Progress Series, 418:179– 188, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/418/m418p179.pdf.

Nielsen:2012:ELF

[NKH12] L. T. Nielsen, B. Krock, and P. J. Hansen. Effects of light and food availability on toxin production, growth and photosynthesis in Dinophysis acuminata. Marine Ecology Progress Series, 471:37–50, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/471/m471p037.pdf.

Nakamura:2015:RCC

[NKM^{+}15] T. Nakamura, O. Kimura, A. Matsuda, T. Matsuishi, M. Kobayashi, and T. Endo. Radiocesium contamination of cetaceans stranded along the coast of Hokkaido, Japan, and an estimation of their travel routes. Marine Ecology Progress Series, 535:1–9, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Nakamura:2015:SVR

[NKS^{+}15] M. Nakamura, N. H. Kumagai, K. Sakai, K. Okaji, K. Ogasawara, and S. Mitarai. Spatial variability in recruitment of acroporid corals and predatory starfish along the Onna coast, Okinawa, Japan. Marine Ecology Progress Series, 540:1–12, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Nunes:2018:SIV

[NLGE18] S. Nunes, M. Latasa, J. M. Gasol, and M. Estrada. Seasonal and interannual variability of phytoplankton community structure in a Mediterranean coastal site. Marine Ecology Progress Series, 592:57–75, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/592/m592p057.pdf.

North:2019:DTN

[NLK^{+}19] C. A. North, J. R. Lovvorn, J. M. Kolts, L. W. Cooper, and J. M. Grebmeier. Discriminating trophic niches of carnivorous benthic macroinvertebrates with gut contents, stable isotopes, and fatty acids. Marine Ecology Progress Series, 631:49–66, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/631/m631p049.pdf.

Ni:2012:PBT

[NLKZ12] G. Ni, Q. Li, L. Kong, and X. Zheng. Phylogeography of the bivalve Tegillarca granosa in coastal China: implications for management and conservation. Marine Ecology Progress Series, 452:119–130, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/452/m452p119.pdf.

Norling:2015:ELP

[NLLS15] P. Norling, M. Lindegarth, S. Lindegarth, and Å. Strand. Effects of live and post-mortem shell structures of invasive Pacific oysters and native blue mussels on macrofauna and fish. Marine Ecology Progress Series, 518:123–138, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/518/
m518p123.pdf.

Navarrete:2015:TUS

[NLV^{+}15] S. A. Navarrete, J. L. Largier, G. Vera, F. J. Tapia, M. Parragué, E. Ramos, J. L. Shinen, C. A. Stuardo, and E. A. Wieters. Tumbling under the surf: wave-modulated settlement of intertidal mussels and the continuous settlement relocation model. Marine Ecology Progress Series, 520:101–121, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m520p101.pdf.

Nanninga:2018:LSC

[NM18] G. B. Nanninga and A. Manica. Larval swimming capacities affect genetic differentiation and range size in demersal marine fishes. Marine Ecology Progress Series, 589:1–12, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Nagata:2016:TMF

[NMC^{+}16] R. M. Nagata, A. C. Morandini, S. P. Colin, A. E. Migotto, and J. H. Costello. Transitions in morphologies, fluid regimes, and feeding mechanisms during development of the medusa Lychnorhiza lucerna. Marine Ecology Progress Series, 557:145–159, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/557/m557p145.pdf.

Nelson:2017:NQM

[NMCP17] N. G. Nelson, R. Muñoz-Carpena, and E. J. Phlips. A novel quantile method reveals spatiotemporal shifts in phytoplankton biomass descriptors between bloom and non-bloom conditions in a subtropical estuary. Marine Ecology Progress Series, 567:57–78, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/567/m567p057.pdf.

Nickols:2013:SDL

[NMG^{+}13] K. J. Nickols, S. H. Miller, B. Gaylord, S. G. Morgan, and J. L. Largier. Spatial differences in larval abundance within the coastal boundary layer impact supply to shoreline habitats. Marine Ecology Progress Series, 494:191–203, 2013. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m494p191.pdf.

Notman:2016:MCD

[NMHB16] G. M. Notman, R. A. R. McGill, S. J. Hawkins, and M. T. Burrows. Macroalgae contribute to the diet of Patella vulgata from contrasting conditions of latitude and wave exposure in the UK. Marine Ecology Progress Series, 549:113–123, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m549p113.pdf.

Nagata:2015:FWC

[NMPM15] R. M. Nagata, M. Z. Moreira, C. R. Pimentel, and A. C. Morandini. Food web characterization based on δ ^{15}N and δ 13C reveals isotopic niche partitioning between fish and jellyfish in a relatively pristine ecosystem. Marine Ecology Progress Series, 519:13–27, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/519/m519p013.pdf.

Nanami:2018:DVH

[NMS^{+}18] A. Nanami, H. Mitamura, T. Sato, T. Yamaguchi, K. Yamamoto, R. Kawabe, K. Soyano, N. Arai, and Y. Kawabata. Diel variation in home range size and precise returning ability after spawning migration of coral reef grouper Epinephelus ongus: implications for effective marine protected area design. Marine Ecology Progress Series, 606:119–132, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/606/
m606p119.pdf.

Nakai:2013:GSQ

[NNJ^{+}13] R. Nakai, K. Nakamura, W. A. Jadoon, K. Kashihara, and T. Naganuma. Genus-specific quantitative PCR of thraustochytrid protists. Marine Ecology Progress Series, 486:1–12, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Norkko:2010:CRI

[NNT^{+}10] J. Norkko, A. Norkko, S. F. Thrush, S. Valanko, and H. Suurkuukka. Conditional responses to increasing scales of disturbance, and potential implications for threshold dynamics in soft-sediment communities. Marine Ecology Progress Series, 413:253–266, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m413p253.pdf.

Nobre:2011:SAA

[Nob11] A. M. Nobre. Scientific approaches to address challenges in coastal management. Marine Ecology Progress Series, 434: 279–289, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m434p279.pdf.

Nicastro:2012:DIE

[NOB12] A. Nicastro, Y. Onoda, and M. J. Bishop. Direct and indirect effects of tidal elevation on eelgrass decomposition. Marine Ecology Progress Series, 456:53–62, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/456/
m456p053.pdf.

Niiranen:2019:PPB

[NOB^{+}19] S. Niiranen, A. Orio, V. Bartolino, U. Bergström, M. Kallasvuo, S. Neuenfeldt, D. Ustups, and M. Casini. Predator–prey body size relationships of cod in a low-diversity marine system. Marine Ecology Progress Series, 627:201–206, 2019. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m627p201.pdf.

Noren:2013:ASG

[Nor13] S. R. Noren. Altered swimming gait and performance of dolphin mothers: implications for interactions with tuna purseseine fisheries. Marine Ecology Progress Series, 482:255– 263, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/482/m482p255.pdf.

Nomaki:2011:DUG

[NOT^{+}11] H. Nomaki, N. O. Ogawa, Y. Takano, H. Suga, N. Ohkouchi, and H. Kitazato. Differing utilization of glucose and algal particulate organic matter by deep-sea benthic organisms of Sagami Bay, Japan. Marine Ecology Progress Series, 431:11–24, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/431/m431p011.pdf.

Needham:2010:HDF

[NPLT10] H. R. Needham, C. A. Pilditch, A. M. Lohrer, and S. F. Thrush. Habitat dependence in the functional traits of Austrohelice crassa, a key bioturbating species. Marine Ecology Progress Series, 414:179–193, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/414/m414p179.pdf.

Nikolioudakis:2011:DFI

[NPMS11] N. Nikolioudakis, I. Palomera, A. Machias, and S. Somarakis. Diel feeding intensity and daily ration of the sardine Sardina pilchardus. Marine Ecology Progress Series, 437:215– 228, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m437p215.pdf.

Nunez-Pons:2012:FDA

[NPRAGG^{+}12] L. Núñez-Pons, M. Rodríguez-Arias, A. Gómez-Garreta, A. Ribera-Siguán, and C. Avila. Feeding deterrency in Antarctic marine organisms: bioassays with the omnivore amphipod Cheirimedon femoratus. Marine Ecology Progress Series, 462:163–174, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/462/m462p163.pdf.

Nakayama:2011:BSD

[NRF11] S. Nakayama, K. A. Rose, and L. A. Fuiman. Batch spawning decreases competition among early life stages in coastal fishes: a simulation study using red drum Sciaenops ocellatus. Marine Ecology Progress Series, 441:213–223, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/441/
m441p213.pdf.

Nielsen:2019:CWP

[NRK^{+}19] J. M. Nielsen, L. A. Rogers, D. G. Kimmel, A. L. Deary, and J. T. Duffy-Anderson. Contribution of walleye pollock eggs to the Gulf of Alaska food web in spring. Marine Ecology Progress Series, 632:1–12, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Nielsen:2019:BDS

[NSB^{+}19] M. L. K. Nielsen, K. R. Sprogis, L. Bejder, P. T. Madsen, and F. Christiansen. Behavioural development in southern right whale calves. Marine Ecology Progress Series, 629:219– 234, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/629/m629p219.pdf.

Noren:2019:VPM

[NSC^{+}19] S. R. Noren, L. Schwarz, K. Chase, K. Aldrich, K. McMahon-Van Oss, and J. St Leger. Validation of the photogrammetric method to assess body condition of an odontocete, the shortfinned pilot whale Globicephala macrorhynchus. Marine Ecology Progress Series, 620:185–200, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/620/
m620p185.pdf.

Naslund:2012:EEM

[NSG^{+}12] J. Näslund, G. S. Samuelsson, J. S. Gunnarsson, F. J. A. Nascimento, H. C. Nilsson, G. Cornelissen, and M. T. Schaanning. Ecosystem effects of materials proposed for thin-layer capping of contaminated sediments. Marine Ecology Progress Series, 449:27–39, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m449p027.pdf.

Nedelec:2015:SLC

[NSH^{+}15] S. L. Nedelec, S. D. Simpson, M. Holderied, A. N. Radford, G. Lecellier, C. Radford, and D. Lecchini. Soundscapes and living communities in coral reefs: temporal and spatial variation. Marine Ecology Progress Series, 524:125– 135, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/524/m524p125.pdf.

Novais:2015:FLI

[NSI^{+}15] A. Novais, A. T. Souza, M. Ilarri, C. Pascoal, and R. Sousa. Facilitation in the low intertidal: effects of an invasive species on the structure of an estuarine macrozoobenthic assemblage. Marine Ecology Progress Series, 522:157–167, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/522/
m522p157.pdf.

Nottestad:2014:PSO

[NSK^{+}14] L. Nøttestad, L. D. Sivle, B. A. Krafft, L. Langård, V. Anthonypillai, M. Bernasconi, H. Langøy, and A. Fernö. Prey selection of offshore killer whales Orcinus orca in the Northeast Atlantic in late summer: spatial associations with mackerel. Marine Ecology Progress Series, 499:275–283, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m499p275.pdf.

Neubauer:2010:SDV

[NSS10] P. Neubauer, J. S. Shima, and S. E. Swearer. Scale-dependent variability in Forsterygion lapillum hatchling otolith chemistry: implications and solutions for studies of population connectivity. Marine Ecology Progress Series, 415:263–274, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2010/415/m415p263.pdf.

Nishizawa:2018:ABL

[NSY^{+}18] B. Nishizawa, T. Sugawara, L. C. Young, E. A. Vanderwerf, K. Yoda, and Y. Watanuki. Albatross-borne loggers show feeding on deep-sea squids: implications for the study of squid distributions. Marine Ecology Progress Series, 592:

257–265, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/592/m592p257.pdf.

Newcombe:2010:TCS

[NT10] E. M. Newcombe and R. B. Taylor. Trophic cascade in a seaweed-epifauna-fish food chain. Marine Ecology Progress Series, 408:161–167, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/408/m408p161.pdf.

Nowicki:2017:PSR

[NTB^{+}17] R. J. Nowicki, J. A. Thomson, D. A. Burkholder, J. W. Fourqurean, and M. R. Heithaus. Predicting seagrass recovery times and their implications following an extreme climate event. Marine Ecology Progress Series, 567:79–93, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2017/567/m567p079.pdf.

Nakajima:2013:EMA

[NTN^{+}13] R. Nakajima, K. Tsuchiya, N. Nakatomi, T. Yoshida, Y. Tada, F. Konno, T. Toda, V. S. Kuwahara, K. Hamasaki, B. H. R. Othman, T. C. Segaran, and A. W. M. Effendy. Enrichment of microbial abundance in the sea-surface microlayer over a coral reef: implications for biogeochemical cycles in reef ecosystems. Marine Ecology Progress Series, 490:11–22, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m490p011.pdf.

Nielsen:2018:OMS

[NTS^{+}18] N. H. Nielsen, J. Teilmann, S. Sveegaard, R. G. Hansen, M. H. S. Sinding, R. Dietz, and M. P. Heide-Jørgensen. Oceanic movements, site fidelity and deep diving in harbour porpoises from Greenland show limited similarities to animals from the North Sea. Marine Ecology Progress Series, 597:259–272, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/597/m597p259.pdf.

Noren:2012:BME

[NUJ12] S. R. Noren, M. S. Udevitz, and C. V. Jay. Bioenergetics model for estimating food requirements of female Pacific walruses Odobenus rosmarus divergens. Marine Ecology Progress Series, 460:261–275, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/460/m460p261.pdf.

Nomaki:2019:NSM

[NUO^{+}19] H. Nomaki, Y. Uejima, N. O. Ogawa, M. Yamane, H. K. Watanabe, R. Senokuchi, J. M. Bernhard, T. Kitahashi, Y. Miyairi, Y. Yokoyama, N. Ohkouchi, and M. Shimanaga. Nutritional sources of meio- and macrofauna at hydrothermal vents and adjacent areas: natural-abundance radiocarbon and stable isotope analyses. Marine Ecology Progress Series, 622:49–65, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m622p049.pdf.

Neto:2019:FCB

[NVA^{+}19] J. Neto, D. Vieira, D. Abecasis, J. Marques, L. Gordo, J. I. Robalo, R. Bispo, M. Araújo, and F. Almada. Facultative cleaning behaviour of juvenile Diplodus sargus (Sparidae) and its ecological role in marine temperate waters. Marine Ecology Progress Series, 629:165–177, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/629/m629p165.pdf.

Noreen:2013:GDD

[NvOH13] A. M. E. Noreen, M. J. H. van Oppen, and P. L. Harrison. Genetic diversity and differentiation among high-latitude broadcast-spawning coral populations disjunct from the core range. Marine Ecology Progress Series, 491:101–109, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m491p101.pdf.

Nielsen:2015:SDZ

[NW15] J. M. Nielsen and M. Winder. Seasonal dynamics of zooplankton resource use revealed by carbon amino acid stable isotope values. Marine Ecology Progress Series, 531:143– 154, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/531/m531p143.pdf.

Nelson:2019:DRB

[NWB^{+}19] T. M. Nelson, M. M. Wallen, M. Bunce, C. L. Oskam, N. Lima, L. Clayton, and J. Mann. Detecting respiratory bacterial communities of wild dolphins: implications for animal health. Marine Ecology Progress Series, 622:203–217, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/622/
m622p203.pdf.

Nygaard:2010:ARL

[NWC^{+}10] H. Nygård, J. Wallenschus, L. Camus, Ø. Varpe, and J. Berge. Annual routines and life history of the amphipod Onisimus litoralis: seasonal growth, body composition and energy budget. Marine Ecology Progress Series, 417:115–126, 2010. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m417p115.pdf.

Nielsen:2013:ALD

[NWD13] T. P. Nielsen, M. Wahlberg, and T. Dabelsteen. Assessment of long-distance detection of gillnets by porpoises: Reply to Dawson & lusseau (2013). Marine Ecology Progress Series, 478:303–305, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/478/m478p303.pdf.

Nielsen:2012:SPW

[NWH^{+}12] Torben P. Nielsen, Magnus Wahlberg, Sanja Heikkilä, Maria Jensen, Puk Sabinsky, and Torben Dabelsteen. Swimming patterns of wild harbour porpoises Phocoena phocoena show detection and avoidance of gillnets at very long ranges. Marine Ecology Progress Series, 453:241–248, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/453/

m453p241.pdf. See comment [DL13].

Newell:2015:EEW

[NWHD15] M. Newell, S. Wanless, M. P. Harris, and F. Daunt. Effects of an extreme weather event on seabird breeding success at a North Sea colony. Marine Ecology Progress Series, 532:257–268, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/532/m532p257.pdf.

Nakamura:2011:YYV

[NWP^{+}11] I. Nakamura, Y. Y. Watanabe, Y. P. Papastamatiou, K. Sato, and C. G. Meyer. Yo-yo vertical movements suggest a foraging strategy for tiger sharks Galeocerdo cuvier. Marine Ecology Progress Series, 424:237–246, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m424p237.pdf.

Nakamura:2014:LHT

[NWS^{+}14] M. Nakamura, H. Watanabe, T. Sasaki, J. Ishibashi, K. Fujikura, and S. Mitarai. Life history traits of Lepetodrilus nux in the Okinawa Trough, based upon gametogenesis, shell size, and genetic variability. Marine Ecology Progress Series, 505:119–130, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/505/m505p119.pdf.

Nishikawa:2013:WML

[NYK^{+}13] H. Nishikawa, I. Yasuda, K. Komatsu, H. Sasaki, Y. Sasai, T. Setou, and M. Shimizu. Winter mixed layer depth and spring bloom along the Kuroshio front: implications for the Japanese sardine stock. Marine Ecology Progress Series, 487:217–229, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m487p217.pdf.

Olafsdottir:2010:GSI

[OA10] A. H. Olafsdottir and J. T. Anderson. Growth and survival of Icelandic capelin Mallotus villosus larvae. Marine Ecology Progress Series, 403:231–241, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/403/m403p231.pdf.

Owen:2019:LCI

[OAB^{+}19] K. Owen, R. D. Andrews, R. W. Baird, G. S. Schorr, and D. L. Webster. Lunar cycles influence the diving behavior and habitat use of short-finned pilot whales around the main Hawaiian Islands. Marine Ecology Progress Series, 629: 193–206, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/629/m629p193.pdf.

Orostica:2014:EGD

[OAD^{+}14] M. H. Oróstica, M. A. Aguilera, G. A. Donoso, J. A. Vásquez, and B. R. Broitman. Effect of grazing on distribution and recovery of harvested stands of Lessonia berteroana kelp in northern Chile. Marine Ecology Progress Series, 511:71– 82, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/511/m511p071.pdf.

Olsen:2011:MHB

[OAGV11] Y. Olsen, T. Andersen, I. Gismervik, and O. Vadstein. Marine heterotrophic bacteria, protozoan and metazoan zooplankton may experience protein N or mineral P limitation in coastal waters. Marine Ecology Progress Series, 436:81–100, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m436p081.pdf.

Oksanen:2014:UMD

[OALK14] S. M. Oksanen, M. P. Ahola, E. Lehtonen, and M. Kunnasranta. Using movement data of Baltic grey seals to examine foraging-site fidelity: implications for seal-fishery conflict mitigation. Marine Ecology Progress Series, 507:297–308, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2014/507/m507p297.pdf. See erratum [Ano15b].

Ogburn:2018:MCP

[OBA^{+}18] M. B. Ogburn, C. W. Bangley, R. Aguilar, R. A. Fisher, M. C. Curran, S. F. Webb, and A. H. Hines. Migratory connectivity and philopatry of cownose rays Rhinoptera bonasus along the Atlantic coast, USA. Marine Ecology Progress Series, 602:197–211, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/602/m602p197.pdf.

Otero:2018:RFT

[OBÁSV18] J. Otero, A. Bode, X. A. Álvarez-Salgado, and M. Varela. Role of functional trait variability in the response of individual phytoplankton species to changing environmental conditions in a coastal upwelling zone. Marine Ecology Progress Series, 596:33–47, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m596p033.pdf.

Oxtoby:2016:FEK

[OBI^{+}16] L. E. Oxtoby, S. M. Budge, K. Iken, D. M. O’Brien, and M. J. Wooller. Feeding ecologies of key bivalve and polychaete species in the Bering Sea as elucidated by fatty acid and compound-specific stable isotope analyses. Marine Ecology Progress Series, 557:161–175, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/557/m557p161.pdf.

Olito:2015:ECF

[OBM15] C. Olito, M. Bode, and D. Marshall. Evolutionary consequences of fertilization mode for reproductive phenology and asynchrony. Marine Ecology Progress Series, 537:23– 38, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/537/m537p023.pdf.

OReilly:2017:PDC

[OC17] E. E. O’Reilly and N. E. Chadwick. Population dynamics of corkscrew sea anemones Bartholomea annulata in the Florida Keys. Marine Ecology Progress Series, 567:109– 123, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/567/m567p109.pdf.

Ortega:2016:CCI

[OCDD16] L. Ortega, E. Celentano, E. Delgado, and O. Defeo. Climate change influences on abundance, individual size and body abnormalities in a sandy beach clam. Marine Ecology Progress Series, 545:203–213, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m545p203.pdf.

Ortega:2012:EFM

[OCE^{+}12] L. Ortega, J. C. Castilla, M. Espino, C. Yamashiro, and O. Defeo. Effects of fishing, market price, and climate on two South American clam species. Marine Ecology Progress Series, 469:71–85, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m469p071.pdf.

Oubelkheir:2013:PML

[OCMT13] K. Oubelkheir, L. A. Clementson, G. F. Moore, and G. H. Tilstone. Production of mycosporine-like amino acids by phytoplankton under ultraviolet radiation exposure in the Sub-Antarctic Zone south of Tasmania. Marine Ecology Progress Series, 494:41–63, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/494/m494p041.pdf.

Olds:2012:PSC

[OCPM12] A. D. Olds, R. M. Connolly, K. A. Pitt, and P. S. Maxwell. Primacy of seascape connectivity effects in structuring coral reef fish assemblages. Marine Ecology Progress Series, 462: 191–203, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/462/m462p191.pdf.

Olsen:2013:EZM

[OCS^{+}13] J. L. Olsen, J. A. Coyer, W. T. Stam, F. E. Moy, H. Christie, and N. M. Jørgensen. Eelgrass Zostera marina populations in northern Norwegian fjords are genetically isolated and diverse. Marine Ecology Progress Series, 486:121–132, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m486p121.pdf.

Ortega-Cisneros:2014:VTS

[OCS14] K. Ortega-Cisneros and U. M. Scharler. Variability and temporal stability of communities in estuaries (Mlalazi and Mpenjati, South Africa). Marine Ecology Progress Series, 500: 11–24, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/500/m500p011.pdf.

Ortega-Cisneros:2016:CNS

[OCSW16] K. Ortega-Cisneros, U. M. Scharler, and A. K. Whitfield. Carbon and nitrogen system dynamics in three small South African estuaries, with particular emphasis on the influence of seasons, river flow and mouth state. Marine Ecology Progress Series, 557:17–30, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/557/m557p017.pdf.

Olsen:2015:CEW

[OD15] Y. S. Olsen and C. M. Duarte. Combined effect of warming and infection by Labyrinthula sp. on the Mediterranean seagrass Cymodocea nodosa. Marine Ecology Progress Series, 532:101–109, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/532/m532p101.pdf.

OConnor:2011:ICE

[ODCE11] N. E. O’Connor, I. Donohue, T. P. Crowe, and M. C. Emmerson. Importance of consumers on exposed and sheltered rocky shores. Marine Ecology Progress Series, 443: 65–75, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/443/m443p065.pdf.

Ochwada-Doyle:2012:CBW

[ODGL^{+}12] F. Ochwada-Doyle, C. A. Gray, N. R. Loneragan, I. M. Suthers, and M. D. Taylor. Competition between wild and captive-bred Penaeus plebejus and implications for stock enhancement. Marine Ecology Progress Series, 450:115–129, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps_oa/m450p115.pdf.

Olson:2014:DDS

[ODK^{+}14] R. J. Olson, L. M. Duffy, P. M. Kuhnert, F. Galván-Magaña, N. Bocanegra-Castillo, and V. Alatorre-Ramírez. Decadal diet shift in yellowfin tuna Thunnus albacares suggests broad-scale food web changes in the eastern tropical Pacific Ocean. Marine Ecology Progress Series, 497:157–178, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/497/
m497p157.pdf.

Olgun:2013:GEO

[ODL^{+}13] N. Olgun, S. Duggen, B. Langmann, M. Hort, C. F. Waythomas, L. Hoffmann, and P. Croot. Geochemical evidence of oceanic iron fertilization by the Kasatochi volcanic eruption in 2008 and the potential impacts on Pacific sockeye salmon. Marine Ecology Progress Series, 488: 81–88, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/488/m488p081.pdf.

OMalley:2012:SVG

[ODP^{+}12] J. M. O’Malley, J. C. Drazen, B. N. Popp, E. Gier, and R. J. Toonen. Spatial variability in growth and prey availability of lobsters in the northwestern Hawaiian Islands. Marine Ecology Progress Series, 449:211–220, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/449/m449p211.pdf.

OSullivan:2011:MRD

[OE11] D. O’Sullivan and M. Emmerson. Marine reserve designation, trophic cascades and altered community dynamics. Marine Ecology Progress Series, 440:115–125, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/440/
m440p115.pdf.

Ojanguren:2010:SVA

[OF10] A. F. Ojanguren and L. A. Fuiman. Seasonal variability in antipredator performance of red drum larvae. Marine Ecology Progress Series, 413:117–123, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/413/m413p117.pdf.

Olsen:2014:TRG

[OFDG14] Z. Olsen, R. Fulford, K. Dillon, and W. Graham. Trophic role of gulf menhaden Brevoortia patronus examined with carbon and nitrogen stable isotope analysis. Marine Ecology Progress Series, 497:215–227, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2014/497/m497p215.pdf.

Olaguer-Feliu:2010:SET

[OFFQGG10] A. O. Olaguer-Feliú, A. A. V. Flores, H. Queiroga, and J. González-Gordillo. Shelf and estuarine transport mechanisms affecting the supply of competent larvae in a suite of brachyuran crabs with different life histories. Marine Ecology Progress Series, 410:125–141, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/410/m410p125.pdf.

Omeyer:2018:DIG

[OFG^{+}18] L. C. M. Omeyer, W. J. Fuller, B. J. Godley, R. T. E. Snape, and A. C. Broderick. Determinate or indeterminate growth? Revisiting the growth strategy of sea turtles. Marine Ecology Progress Series, 596:199–211, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/596/m596p199.pdf.

Olsen:2018:LSO

[OFM^{+}18] Y. S. Olsen, M. W. Fraser, B. C. Martin, A. Pomeroy, R. Lowe, O. Pedersen, and G. A. Kendrick. In situ oxygen dynamics in rhizomes of the seagrass Posidonia sinuosa: impact of light, water column oxygen, current speed and wave velocity. Marine Ecology Progress Series, 590:67–77, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/590/
m590p067.pdf.

Orejas:2011:LTG

[OFPR^{+}11] C. Orejas, C. Ferrier-Pagès, S. Reynaud, A. Gori, E. Beraud, G. Tsounis, D. Allemand, and J. M. Gili. Longterm growth rates of four Mediterranean cold-water coral species maintained in aquaria. Marine Ecology Progress Series, 429:57–65, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/429/m429p057.pdf.

Olsen:2011:RDC

[OFT^{+}11] Y. S. Olsen, S. E. Fox, M. Teichberg, M. Otter, and I. Valiela. δ ^{15}N and δ ^{13}C reveal differences in carbon flow through estuarine benthic food webs in response to the relative availability of macroalgae and eelgrass. Marine Ecology Progress Series, 421:83–96, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/421/m421p083.pdf.

Orcutt:2013:IEA

[OGA13] K. M. Orcutt, K. Gundersen, and J. W. Ammerman. Intense ectoenzyme activities associated with Trichodesmium colonies in the Sargasso Sea. Marine Ecology Progress Series, 478:101–113, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m478p101.pdf.

Ottmann:2018:IRV

[OGCHS18] D. Ottmann, K. Grorud-Colvert, B. Huntington, and S. Sponaugle. Interannual and regional variability in settlement of groundfishes to protected and fished nearshore waters of Oregon, USA. Marine Ecology Progress Series, 598: 131–145, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/598/m598p131.pdf.

Orgeret:2014:RAP

[OGGP14] F. Orgeret, C. Garrigue, O. Gimenez, and R. Pradel. Robust assessment of population trends in marine mammals applied to new caledonian humpback whales. Marine Ecology Progress Series, 515:265–273, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2014/515/m515p265.pdf.

Olsen:2018:HES

[OGH18] M. T. Olsen, A. Galatius, and T. Härkönen. The history and effects of seal-fishery conflicts in Denmark. Marine Ecology Progress Series, 595:233–243, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/595/m595p233.pdf.

OToole:2017:MPP

[OGLH17] M. O’Toole, C. Guinet, M. A. Lea, and M. A. Hindell. Marine predators and phytoplankton: how elephant seals use the recurrent Kerguelen plume. Marine Ecology Progress Series, 581:215–227, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/581/m581p215.pdf.

Ohji:2010:BOR

[OHA10] M. Ohji, H. Harino, and T. Arai. Bioaccumulation of organotin in relation to the life history of the brown trout Salmo trutta. Marine Ecology Progress Series, 399:285–293, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/399/
m399p285.pdf.

OFarrell:2015:PFI

[OHB^{+}15] S. O’Farrell, A. R. Harborne, Y. M. Bozec, B. E. Luckhurst, and P. J. Mumby. Protection of functionally important parrotfishes increases their biomass but fails to deliver enhanced recruitment. Marine Ecology Progress Series, 522:245–254, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/522/m522p245.pdf.

OToole:2014:FBS

[OHCG14] M. O’Toole, M. A. Hindell, J. B. Charrassin, and C. Guinet. Foraging behaviour of southern elephant seals over the Kerguelen Plateau. Marine Ecology Progress Series, 502:281– 294, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/502/m502p281.pdf.

Osborne:2018:DCT

[OHP18] K. L. Osborne, R. E. Hannigan, and H. C. Poynton. Differential copper toxicity in invasive and native ascidians of New England provides support for enhanced invader tolerance. Marine Ecology Progress Series, 595:135–147, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/595/
m595p135.pdf.

Olin:2013:SVS

[OHR^{+}13] J. A. Olin, N. E. Hussey, S. A. Rush, G. R. Poulakis, C. A. Simpfendorfer, M. R. Heupel, and A. T. Fisk. Seasonal variability in stable isotopes of estuarine consumers under different freshwater flow regimes. Marine Ecology Progress Series, 487:55–69, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/487/m487p055.pdf.

Ohnishi:2019:IMM

[OHST19] T. Ohnishi, J. Hirai, S. Shimode, and A. Tsuda. Identification of molecular markers associated with starvation in female Calanus sinicus (Copepoda: Calanoida). Marine Ecology Progress Series, 614:51–65, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/614/m614p051.pdf.

Orben:2018:PAM

[OKF^{+}18] R. A. Orben, N. Kokubun, A. B. Fleishman, A. P. Will, T. Yamamoto, S. A. Shaffer, R. Paredes, A. Takahashi, and A. S. Kitaysky. Persistent annual migration patterns of a specialist seabird. Marine Ecology Progress Series, 593:231–245, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m593p231.pdf.

Olesen:2015:EZM

[OKJMC15] B. Olesen, D. Krause-Jensen, N. Marbà, and P. B. Christensen. Eelgrass Zostera marina in subarctic Greenland: dense meadows with slow biomass turnover in cold waters. Marine Ecology Progress Series, 518:107–121, 2015. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m518p107.pdf.

Otto:2014:IAD

[OKLM14] S. A. Otto, G. Kornilovs, M. Llope, and C. Möllmann. Interactions among density, climate, and food web effects determine long-term life cycle dynamics of a key copepod. Marine Ecology Progress Series, 498:73–84, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m498p073.pdf.

OSullivan:2015:MCL

[OLD^{+}15] D. O’Sullivan, C. Lordan, J. Doyle, A. Berry, and K. Lyons. Metapopulation connectivity via larval transport of the Norway lobster Nephrops norvegicus in waters around Ireland: a modelled approach. Marine Ecology Progress Series, 534: 95–106, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/534/m534p095.pdf.

Oliver:2011:RLD

[OLF11] L. M. Oliver, J. C. Lehrter, and W. S. Fisher. Relating landscape development intensity to coral reef condition in the watersheds of St. Croix, US Virgin Islands. Marine Ecology Progress Series, 427:293–302, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m427p293.pdf.

Oigaard:2013:FRB

[ØLH^{+}13] T. A. Øigård, U. Lindstrøm, T. Haug, K. T. Nilssen, and S. Smout. Functional relationship between harp seal body condition and available prey in the Barents Sea. Marine Ecology Progress Series, 484:287–301, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/484/m484p287.pdf.

Orvain:2012:STI

[OLM^{+}12] F. Orvain, S. Lefebvre, J. Montepini, M. Sébire, A. Gangnery, and B. Sylvand. Spatial and temporal interaction between sediment and microphytobenthos in a temperate estuarine macro-intertidal bay. Marine Ecology Progress Series, 458:53–68, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/458/m458p053.pdf.

Obermuller:2010:SPE

[OMBP10] B. E. Obermüller, S. A. Morley, D. K. A. Barnes, and L. S. Peck. Seasonal physiology and ecology of Antarctic marine benthic predators and scavengers. Marine Ecology Progress Series, 415:109–126, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/415/m415p109.pdf.

Ouisse:2011:CLC

[OMD11] V. Ouisse, A. Migné, and D. Davoult. Community-level carbon flux variability over a tidal cycle in Zostera marina and Z. noltii beds. Marine Ecology Progress Series, 437:79–87, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m437p079.pdf.

Orth:2012:ERI

[OMH12] Robert J. Orth, Karen J. McGlathery, and Kenneth L. Heck, Jr. Eelgrass recovery induces state changes in a coastal bay. Marine Ecology Progress Series, 434:171–301, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m448_ThemeSection.pdf.

Orth:2012:SAF

[OMM^{+}12] R. J. Orth, K. A. Moore, S. R. Marion, D. J. Wilcox, and D. B. Parrish. Seed addition facilitates eelgrass recovery in a coastal bay system. Marine Ecology Progress Series, 448:177–195, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m448p177.pdf.

OHanlon:2017:IUI

[OMN17] N. J. O’Hanlon, R. A. R. McGill, and R. G. Nager. Increased use of intertidal resources benefits breeding success in a generalist gull species. Marine Ecology Progress Series, 574:193–210, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/574/m574p193.pdf.

Overton:2010:RHP

[OMSP10] J. L. Overton, S. Meyer, J. G. Støttrup, and M. A. Peck. Role of heterotrophic protists in first feeding by cod (Gadus morhua) larvae. Marine Ecology Progress Series, 410:197– 204, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/410/m410p197.pdf.

Onda:2017:VDA

[OMT^{+}17] H. Onda, M. J. Miller, A. Takeshige, Y. Miyake, M. Kuroki, J. Aoyama, and S. Kimura. Vertical distribution and assemblage structure of leptocephali in the North Equatorial Current region of the western Pacific. Marine Ecology Progress Series, 575:119–136, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/575/m575p119.pdf.

Osman:2010:TMC

[OMW^{+}10] R. W. Osman, P. Munguia, R. B. Whitlatch, R. N. Zajac, and J. Hamilton. Thresholds and multiple community states in marine fouling communities: integrating natural history with management strategies. Marine Ecology Progress Series, 413:277–289, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m413p277.pdf.

Ortiz:2016:RGT

[OMW^{+}16] N. Ortiz, J. C. Mangel, J. Wang, J. Alfaro-Shigueto, S. Pingo, A. Jimenez, T. Suarez, Y. Swimmer, F. Carvalho, and B. J. Godley. Reducing green turtle bycatch in small-scale fisheries using illuminated gillnets: the cost of saving a sea turtle. Marine Ecology Progress Series, 545:251–259, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/545/
m545p251.pdf.

Osman:2010:ETM

[OMZ10a] R. W. Osman, P. Munguia, and R. N. Zajac. Ecological thresholds in marine communities: theory, experiments and management. Marine Ecology Progress Series, 413:185–187, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m413p185.pdf.

Osman:2010:TDM

[OMZ10b] Richard W. Osman, Pablo Munguia, and Roman N. Zajac. Threshold dynamics in marine coastal systems. Marine Ecology Progress Series, 413:185–304, 2010. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m413_ThemeSection.pdf.

Onofri:2015:MFE

[ON15] L. Onofri and P. A. L. D. Nunes. Microeconomic framework for the economics of whales. Marine Ecology Progress Series, 530:213–222, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m530p213.pdf.

Okamura:2011:BGE

[OOK11] B. Okamura, A. O’Dea, and T. Knowles. Bryozoan growth and environmental reconstruction by zooid size variation. Marine Ecology Progress Series, 430:133–146, 2011. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m430p133.pdf.

Oczkowski:2010:MGV

[OPN10] A. J. Oczkowski, M. E. Q. Pilson, and S. W. Nixon. A marked gradient in δ ^{13}C values of clams Mercenaria mercenaria across a marine embayment may reflect variations in ecosystem metabolism. Marine Ecology Progress Series, 414:145–153, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/414/m414p145.pdf.

OLeary:2013:FSU

[OPSM13] J. K. O’Leary, D. Potts, K. M. Schoenrock, and T. R. Mc-Clahanan. Fish and sea urchin grazing opens settlement space equally but urchins reduce survival of coral recruits. Marine Ecology Progress Series, 493:165–177, 2013. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m493p165.pdf.

Onton:2011:DDC

[OPW^{+}11] K. Onton, C. A. Page, S. K. Wilson, S. Neale, and S. Armstrong. Distribution and drivers of coral disease at Ningaloo reef, Indian Ocean. Marine Ecology Progress Series, 433: 75–84, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/433/m433p075.pdf.

Olson:2017:CTS

[ORE^{+}17] J. W. Olson, K. D. Rode, D. Eggett, T. S. Smith, R. R. Wilson, G. M. Durner, A. Fischbach, T. C. Atwood, and D. C. Douglas. Collar temperature sensor data reveal longterm patterns in southern Beaufort Sea polar bear den distribution on pack ice and land. Marine Ecology Progress Series, 564:211–224, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/564/m564p211.pdf.

Ober:2019:CEB

[ORG19] G. T. Ober, R. L. Rognstad, and S. E. Gilman. The cost of emersion for the barnacle Balanus glandula. Marine Ecology Progress Series, 627:95–107, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/627/m627p095.pdf.

Ogburn:2019:SAS

[ORJH19] M. B. Ogburn, K. D. Richie, M. A. Jones, and A. H. Hines. Sperm acquisition and storage dynamics facilitate sperm limitation in the selectively harvested blue crab Callinectes sapidus. Marine Ecology Progress Series, 629:87– 101, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/629/m629p087.pdf.

Ordonez:2013:EBI

[ORM^{+}13] V. Ordóñez, M. Rius, C. D. McQuaid, M. C. Pineda, M. Pascual, and X. Turon. Early biotic interactions among introduced and native benthic species reveal cryptic predation and shifts in larval behaviour. Marine Ecology Progress Series, 488:65–79, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/488/m488p065.pdf.

Ogburn:2014:TSV

[ORR^{+}14] M. B. Ogburn, P. M. Roberts, K. D. Richie, E. G. Johnson, and A. H. Hines. Temporal and spatial variation in sperm stores in mature female blue crabs Callinectes sapidus and potential effects on brood production in Chesapeake Bay. Marine Ecology Progress Series, 507:249–262, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/507/
m507p249.pdf.

Oliver:2015:UMR

[ORW15] H. Oliver, R. L. Rognstad, and D. S. Wethey. Using meteorological reanalysis data for multi-decadal hindcasts of larval connectivity in the coastal ocean. Marine Ecology Progress Series, 530:47–62, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m530p047.pdf.

Olsen:2014:CEM

[ORWPR14] K. Olsen, R. Ritson-Williams, V. J. Paul, and C. Ross. Combined effects of macroalgal presence and elevated temperature on the early life-history stages of a common Caribbean coral. Marine Ecology Progress Series, 509:181–191, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/509/
m509p181.pdf.

Olyarnik:2012:MYS

[OS12] S. V. Olyarnik and J. J. Stachowicz. Multi-year study of the effects of Ulva sp. blooms on eelgrass Zostera marina. Marine Ecology Progress Series, 468:107–117, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/468/
m468p107.pdf.

OBrien:2018:TWC

[OS18] J. M. O’Brien and R. E. Scheibling. Turf wars: competition between foundation and turf-forming species on temperate and tropical reefs and its role in regime shifts. Marine Ecology Progress Series, 590:1–17, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Olivares:2019:OCF

[OSC19] M. Olivares, E. Saiz, and A. Calbet. Ontogenetic changes in the feeding functional response of the marine copepod Paracartia grani. Marine Ecology Progress Series, 616:25– 35, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/616/m616p025.pdf.

Ottersen:2013:TSR

[OSD^{+}13] G. Ottersen, L. C. Stige, J. M. Durant, K. S. Chan, T. A. Rouyer, K. F. Drinkwater, and N. C. Stenseth. Temporal shifts in recruitment dynamics of North Atlantic fish stocks: effects of spawning stock and temperature. Marine Ecology Progress Series, 480:205–225, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m480p205.pdf.

Okamoto:2013:CDR

[OSE13] D. K. Okamoto, M. S. Stekoll, and G. L. Eckert. Coexistence despite recruitment inhibition of kelps by subtidal algal crusts. Marine Ecology Progress Series, 493:103–112, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/493/
m493p103.pdf.

OBrien:2015:PFB

[OSK15] J. M. O’Brien, R. E. Scheibling, and K. A. Krumhansl. Positive feedback between large-scale disturbance and densitydependent grazing decreases resilience of a kelp bed ecosystem. Marine Ecology Progress Series, 522:1–13, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Olson:2018:MGR

[OSL18] M. B. Olson, K. Solem, and B. Love. Microzooplankton grazing responds to simulated ocean acidification indirectly through changes in prey cellular characteristics. Marine Ecology Progress Series, 604:83–97, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/604/
m604p083.pdf.

Oliveira:2014:IEH

[OSPWB14] J. P. Oliveira, I. Sousa-Pinto, G. M. Weber, and I. Bertocci. Interplay of experimental harvesting and climate-related disturbance on benthic assemblages of rocky seashores. Marine Ecology Progress Series, 495:131–142, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/495/
m495p131.pdf.

Oliveira:2015:PRI

[OSPWB15] J. P. Oliveira, I. Sousa-Pinto, G. M. Weber, and I. Bertocci. Patterns of recovery of intertidal organisms after compounded anthropogenic disturbances. Marine Ecology Progress Series, 524:107–123, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/524/m524p107.pdf.

Oliphant:2013:POI

[OT13] A. Oliphant and S. Thatje. Per offspring investment implications for crustacean larval development: evolutionary insights into endotrophy and abbreviated development. Marine Ecology Progress Series, 493:207–217, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/493/
m493p207.pdf.

Oliphant:2014:EAL

[OT14] A. Oliphant and S. Thatje. Energetic adaptations to larval export within the brackish-water palaemonine shrimp, Palaemonetes varians. Marine Ecology Progress Series, 505:177– 191, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/505/m505p177.pdf.

Olli:2013:HLD

[OTK^{+}13] K. Olli, O. Trikk, R. Klais, R. Ptacnik, T. Andersen, S. Lehtinen, and T. Tamminen. Harmonizing large data sets reveals novel patterns in the Baltic Sea phytoplankton community structure. Marine Ecology Progress Series, 473:53–66, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m473p053.pdf.

Okazaki:2019:DOO

[OTK^{+}19] Y. Okazaki, K. Tadokoro, H. Kubota, Y. Kamimura, and K. Hidaka. Dietary overlap and optimal prey environments of larval and juvenile sardine and anchovy in the mixed water region of the western North Pacific. Marine Ecology Progress Series, 630:149–160, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2019/630/m630p149.pdf.

ODonnell:2010:OAA

[OTS^{+}10] M. J. O’Donnell, A. E. Todgham, M. A. Sewell, L. M. Hammond, K. Ruggiero, N. A. Fangue, M. L. Zippay, and G. E. Hofmann. Ocean acidification alters skeletogenesis and gene expression in larval sea urchins. Marine Ecology Progress Series, 398:157–171, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m398p157.pdf.

Oguz:2010:ATN

[OV10] T. Oguz and V. Velikova. Abrupt transition of the northwestern Black Sea shelf ecosystem from a eutrophic to an alternative pristine state. Marine Ecology Progress Series, 405:231–242, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/405/m405p231.pdf.

Opdal:2011:PMC

[OVF11] A. F. Opdal, F. B. Vikebø, and Ø. Fiksen. Parental migration, climate and thermal exposure of larvae: spawning in southern regions gives Northeast Arctic cod a warm start. Marine Ecology Progress Series, 439:255–262, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m439p255.pdf.

Owen:2019:BTF

[OWH^{+}19] E. Owen, E. Wakefield, P. Hollinrake, A. Leitch, L. Steel, and M. Bolton. Breeding together, feeding apart: sympatrically breeding seabirds forage in individually distinct locations. Marine Ecology Progress Series, 620:173–183, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/620/
m620p173.pdf.

Owen:2015:EPT

[OWN^{+}15] K. Owen, J. D. Warren, M. J. Noad, D. Donnelly, A. W. Goldizen, and R. A. Dunlop. Effect of prey type on the finescale feeding behaviour of migrating east Australian humpback whales. Marine Ecology Progress Series, 541:231–244, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2015/541/m541p231.pdf.

Oppel:2017:SSF

[OWW^{+}17] S. Oppel, S. Weber, N. Weber, D. Fox, E. Leat, J. Sim, J. Sommerfeld, M. Bolton, A. C. Broderick, and B. J. Godley. Seasonal shifts in foraging distribution due to individual flexibility in a tropical pelagic forager, the ascension frigatebird. Marine Ecology Progress Series, 585:199–212, 2017. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m585p199.pdf.

Penin:2013:RIR

[PA13] L. Penin and M. Adjeroud. Relative importance of recruitment and post-settlement processes in the maintenance of coral assemblages in an insular, fragmented reef system. Marine Ecology Progress Series, 473:149–162, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/473/
m473p149.pdf.

Patil:2015:EMP

[PA15] J. S. Patil and A. C. Anil. Effect of monsoonal perturbations on the occurrence of phytoplankton blooms in a tropical bay. Marine Ecology Progress Series, 530:77–92, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/530/
m530p077.pdf.

Petitet:2015:AGO

[PAC^{+}15] R. Petitet, L. Avens, J. C. Castilhos, P. G. Kinas, and L. Bugoni. Age and growth of olive ridley sea turtles Lepidochelys olivacea in the main Brazilian nesting ground. Marine Ecology Progress Series, 541:205–218, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/541/m541p205.pdf.

Pedersen:2010:LGD

[PAF^{+}10] T. M. Pedersen, R. Almeda, F. L. Fotel, H. H. Jakobsen, P. Mariani, and B. W. Hansen. Larval growth in the dominant polychaete Polydora ciliata is food-limited in a eutrophic Danish estuary (Isefjord). Marine Ecology Progress Series, 407:99–110, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/407/m407p099.pdf.

Penant:2013:PLD

[PAFC13] G. Penant, D. Aurelle, J. P. Feral, and A. Chenuil. Planktonic larvae do not ensure gene flow in the edible sea urchin Paracentrotus lividus. Marine Ecology Progress Series, 480: 155–170, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/480/m480p155.pdf.

Paquette:2019:HGE

[PAG19] L. Paquette, P. Archambault, and F. Guichard. From habitat geometry to ecosystem functions in marine mussel beds. Marine Ecology Progress Series, 608:149–163, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/608/
m608p149.pdf.

Polo-Avila:2019:CPD

[PÁIIS^{+}19] A. Polo-Ávila, M. D. Infante-Izquierdo, J. M. Soto, V. Hermoso-López, F. J. J. Nieva, J. M. Castillo, and A. F. Muñoz-Rodríguez. Contrasting propagule dispersal and halophyte seed banks along the intertidal gradient. Marine Ecology Progress Series, 616:51–65, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/616/m616p051.pdf.

Prihartato:2015:SPN

[PAK15] P. K. Prihartato, D. L. Aksnes, and S. Kaartvedt. Seasonal patterns in the nocturnal distribution and behavior of the mesopelagic fish Maurolicus muelleri at high latitudes. Marine Ecology Progress Series, 521:189–200, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/521/
m521p189.pdf.

Petitgas:2012:APE

[PAP^{+}12] P. Petitgas, J. Alheit, M. A. Peck, K. Raab, X. Irigoien, M. Huret, J. van der Kooij, T. Pohlmann, C. Wagner, I. Zarraonaindia, and M. Dickey-Collas. Anchovy population expansion in the North Sea. Marine Ecology Progress Series, 444:1–13, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Parrish:2015:SCS

[Par15] F. A. Parrish. Settlement, colonization, and succession patterns of gold coral Kulamanamana haumeaae in Hawaiian deep coral assemblages. Marine Ecology Progress Series, 533:135–147, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m533p135.pdf.

Polgar:2010:SVS

[PB10] G. Polgar and V. Bartolino. Size variation of six species of oxudercine gobies along the intertidal zone in a malayan coastal swamp. Marine Ecology Progress Series, 409:199– 212, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/409/m409p199.pdf.

Peterson:2018:SCP

[PB18] J. M. Peterson and S. S. Bell. Species composition of patches influences mangrove recruitment in a saltmarsh mosaic. Marine Ecology Progress Series, 602:103–116, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/602/
m602p103.pdf.

Pringle:2019:OBG

[PB19] J. W. Pringle and H. Baumann. Otolith-based growth reconstructions in young-of-year Atlantic silversides Menidia menidia and their implications for sex-selective survival. Marine Ecology Progress Series, 632:193–204, 2019. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m632p193.pdf.

Paul:2012:WAS

[PBA12] M. Paul, T. J. Bouma, and C. L. Amos. Wave attenuation by submerged vegetation: combining the effect of organism traits and tidal current. Marine Ecology Progress Series, 444:31–41, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/444/m444p031.pdf.

Pethybridge:2014:TIS

[PBAP^{+}14] H. Pethybridge, N. Bodin, E. J. Arsenault-Pernet, J. H. Bourdeix, B. Brisset, J. L. Bigot, D. Roos, and M. Peter. Temporal and inter-specific variations in forage fish feeding conditions in the NW Mediterranean: lipid content and fatty acid compositional changes. Marine Ecology Progress Series, 512:39–54, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m512p039.pdf.

Pajuelo:2010:SIV

[PBAS^{+}10] M. Pajuelo, K. A. Bjorndal, J. Alfaro-Shigueto, J. A. Seminoff, J. C. Mangel, and A. B. Bolten. Stable isotope variation in loggerhead turtles reveals Pacific–Atlantic oceanographic differences. Marine Ecology Progress Series, 417: 277–285, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/417/m417p277.pdf.

Porsmoguer:2015:DEC

[PBB^{+}15] S. Biton Porsmoguer, D. Bănaru, C. F. Boudouresque, I. Dekeyser, A. Viricel, and M. Merchán. DNA evidence of the consumption of short-beaked common dolphin Delphinus delphis by the shortfin mako shark Isurus oxyrinchus. Marine Ecology Progress Series, 532:177–183, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/532/
m532p177.pdf.

Piroddi:2011:MOC

[PBC11] C. Piroddi, G. Bearzi, and V. Christensen. Marine open cage aquaculture in the eastern Mediterranean Sea: a new trophic resource for bottlenose dolphins. Marine Ecology Progress Series, 440:255–266, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/440/m440p255.pdf.

Pethybridge:2012:TSB

[PBC^{+}12a] H. Pethybridge, E. C. V. Butler, D. Cossa, R. Daley, and A. Boudou. Trophic structure and biomagnification of mercury in an assemblage of deepwater chondrichthyans from southeastern Australia. Marine Ecology Progress Series, 451: 163–174, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/451/m451p163.pdf.

Polidoro:2012:PER

[PBC^{+}12b] B. A. Polidoro, T. Brooks, K. E. Carpenter, G. J. Edgar, S. Henderson, J. Sanciangco, and D. R. Robertson. Patterns of extinction risk and threat for marine vertebrates and habitat-forming species in the Tropical Eastern Pacific. Marine Ecology Progress Series, 448:93–104, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/448/
m448p093.pdf.

Pinkerton:2015:NDG

[PBC^{+}15] K. Pinkerton, D. M. Baker, M. R. Cuddy, L. J. Raymundo, K. A. Meyer, and K. Kim. Nitrogen dynamics on Guam as revealed by the seagrass Enhalus acoroides. Marine Ecology Progress Series, 528:117–126, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/528/m528p117.pdf.

Passuni:2018:LTC

[PBC^{+}18] G. Passuni, C. Barbraud, A. Chaigneau, A. Bertrand, R. Oliveros-Ramos, J. Ledesma, R. Castillo, M. Bouchon, and S. Bertrand. Long-term changes in the breeding seasonality of Peruvian seabirds and regime shifts in the Northern Humboldt Current System. Marine Ecology Progress Series, 597:231–242, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/597/m597p231.pdf.

Pinela:2010:SIA

[PBCA10] A. M. Pinela, A. Borrell, L. Cardona, and A. Aguilar. Stable isotope analysis reveals habitat partitioning among marine mammals off the NW African coast and unique trophic niches for two globally threatened species. Marine Ecology Progress Series, 416:295–306, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/416/m416p295.pdf.

Perea-Blazquez:2013:TVF

[PBDMRB13] A. Perea-Blázquez, S. K. Davy, B. Magana-Rodríguez, and J. J. Bell. Temporal variation in food utilisation by three species of temperate demosponge. Marine Ecology Progress Series, 485:91–103, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/485/m485p091.pdf.

Pedersen:2010:PDL

[PBF10] M. F. Pedersen, J. Borum, and F. Leck Fotel. Phosphorus dynamics and limitation of fast- and slow-growing temperate seaweeds in Oslofjord, Norway. Marine Ecology Progress Series, 399:103–115, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m399p103.pdf.

Pillay:2010:ECS

[PBG^{+}10] D. Pillay, G. M. Branch, C. L. Griffiths, C. Williams, and A. Prinsloo. Ecosystem change in a South African marine reserve (1960–2009): role of seagrass loss and anthropogenic disturbance. Marine Ecology Progress Series, 415:

35–48, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/415/m415p035.pdf.

Pages:2013:EKH

[PBH^{+}13] J. F. Pagès, F. Bartumeus, B. Hereu, À. López-Sanz, J. Romero, and T. Alcoverro. Evaluating a key herbivorous fish as a mobile link: a Brownian bridge approach. Marine Ecology Progress Series, 492:199–210, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/492/m492p199.pdf.

Pringle:2017:OCC

[PBH^{+}17] J. M. Pringle, J. E. Byers, R. He, P. Pappalardo, and J. Wares. Ocean currents and competitive strength interact to cluster benthic species range boundaries in the coastal ocean. Marine Ecology Progress Series, 567:29–40, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/567/
m567p029.pdf.

Pereda-Briones:2018:DSP

[PBIO^{+}18] L. Pereda-Briones, E. Infantes, A. Orfila, F. Tomas, and J. Terrados. Dispersal of seagrass propagules: interaction between hydrodynamics and substratum type. Marine Ecology Progress Series, 593:47–59, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/593/m593p047.pdf.

Pernet:2010:EAE

[PBM^{+}10] F. Pernet, J. Barret, C. Marty, J. Moal, P. Le Gall, and P. Boudry. Environmental anomalies, energetic reserves and fatty acid modifications in oysters coincide with an exceptional mortality event. Marine Ecology Progress Series, 401: 129–146, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/401/m401p129.pdf.

Palmer:2011:SEC

[PBM^{+}11] M. Palmer, S. Balle, D. March, J. Alós, and M. Linde. Size estimation of circular home range from fish mark-release- (single)-recapture data: case study of a small labrid targeted by recreational fishing. Marine Ecology Progress Series, 430:87–97, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/430/m430p087.pdf.

Precoda:2018:HDW

[PBM^{+}18] K. Precoda, A. H. Baird, A. Madsen, T. Mizerek, B. Sommer, S. N. Su, and J. S. Madin. How does a widespread reef coral maintain a population in an isolated environment? Marine Ecology Progress Series, 594:85–94, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/594/
m594p085.pdf.

Pillay:2010:UES

[PBS10] D. Pillay, G. M. Branch, and A. Steyn. Unexpected effects of a starfish grazing on sandflat communities following an outbreak. Marine Ecology Progress Series, 398:173– 182, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/398/m398p173.pdf.

Pardo:2015:LSR

[PBTB15] S. A. Pardo, K. B. Burgess, D. Teixeira, and M. B. Bennett. Local-scale resource partitioning by stingrays on an intertidal flat. Marine Ecology Progress Series, 533:205– 218, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/533/m533p205.pdf.

Pape:2011:CSF

[PBV^{+}11] E. Pape, T. Nara Bezerra, H. Vanneste, K. Heeschen, L. Moodley, F. Leroux, P. van Breugel, and A. Vanreusel. Community structure and feeding preference of nematodes associated with methane seepage at the Darwin mud volcano (Gulf of Cádiz). Marine Ecology Progress Series, 438:71–83, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/438/
m438p071.pdf.

Pfaff:2011:UIW

[PBW^{+}11] M. C. Pfaff, G. M. Branch, E. A. Wieters, R. A. Branch, and B. R. Broitman. Upwelling intensity and wave exposure determine recruitment of intertidal mussels and barnacles in the southern Benguela upwelling region. Marine Ecology Progress Series, 425:141–152, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/425/m425p141.pdf.

Pouil:2018:OTE

[PBWM18] S. Pouil, P. Bustamante, M. Warnau, and M. Metian. Overview of trace element trophic transfer in fish through the concept of assimilation efficiency. Marine Ecology Progress Series, 588:243–254, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/588/m588p243.pdf.

Passow:2012:BPH

[PC12] U. Passow and C. A. Carlson. The biological pump in a high CO_{2} world. Marine Ecology Progress Series, 470:249– 271, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m470p249.pdf.

Poland:2019:HGS

[PC19] D. M. Poland and M. A. Coffroth. Host growth and survivorship varies with endosymbiotic algal partner in developing cnidarians. Marine Ecology Progress Series, 612:87– 100, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/612/m612p087.pdf.

Pascual:2015:RIM

[PCA^{+}15] J. Pascual, A. D. Canepuccia, J. Alberti, P. Daleo, and O. Iribarne. Rainfall intensity modulates the interaction between the marsh cordgrass Spartina densiflora and the mouse Akodon azarae. Marine Ecology Progress Series, 523: 71–80, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/523/m523p071.pdf.

Pires:2018:UVV

[PCD^{+}18] R. F. T. Pires, N. Cordeiro, J. Dubert, A. Marraccini, P. Relvas, and A. dos Santos. Untangling Velella velella (Cnidaria: Anthoathecatae) transport: a citizen science and oceanographic approach. Marine Ecology Progress Series, 591:241– 251, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/591/m591p241.pdf.

Pechenik:2019:TTI

[PCF^{+}19] J. A. Pechenik, O. R. Chaparro, A. Franklin, M. L. Mardones, and J. A. Montory. Thermal tolerance of intertidal and subtidal adults and embryos of the marine gastropod Crepipatella peruviana. Marine Ecology Progress Series, 616: 67–81, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/616/m616p067.pdf.

Peterson:2011:EPA

[PCH11] T. D. Peterson, D. W. Crawford, and P. J. Harrison. Evolution of the phytoplankton assemblage in a long-lived mesoscale eddy in the eastern Gulf of Alaska. Marine Ecology Progress Series, 424:53–73, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/424/m424p053.pdf.

Piercy:2014:HQA

[PCH^{+}14] J. J. B. Piercy, E. A. Codling, A. J. Hill, D. J. Smith, and S. D. Simpson. Habitat quality affects sound production and likely distance of detection on coral reefs. Marine Ecology Progress Series, 516:35–47, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/516/m516p035.pdf.

Proud:2018:FSD

[PCLB18] R. Proud, M. J. Cox, C. Le Guen, and A. S. Brierley. Fine-scale depth structure of pelagic communities throughout the global ocean based on acoustic sound scattering layers. Marine Ecology Progress Series, 598:35–48, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/598/
m598p035.pdf.

Pescheck:2016:RSD

[PCNB16] F. Pescheck, H. Campen, L. Nichelmann, and W. Bilger. Relative sensitivity of DNA and photosystem II in Ulva intestinalis (Chlorophyta) under natural solar irradiation. Marine Ecology Progress Series, 555:95–107, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/555/m555p095.pdf.

Prestes:2017:TSM

[PCNM17] A. C. L. Prestes, E. Cacabelos, A. I. Neto, and G. M. Martins. Temporal stability in macroalgal assemblage standing stock despite high species turnover. Marine Ecology Progress Series, 567:249–256, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/567/m567p249.pdf.

Paradinas:2015:BST

[PCP^{+}15] I. Paradinas, D. Conesa, M. G. Pennino, F. Muñoz, A. M. Fernández, A. López-Quílez, and J. M. Bellido. Bayesian spatio-temporal approach to identifying fish nurseries by validating persistence areas. Marine Ecology Progress Series, 528:245–255, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/528/m528p245.pdf.

Piroddi:2015:MMM

[PCS^{+}15] C. Piroddi, M. Coll, J. Steenbeek, D. Macias Moy, and V. Christensen. Modelling the Mediterranean marine ecosystem as a whole: addressing the challenge of complexity. Marine Ecology Progress Series, 533:47–65, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m533p047.pdf.

Prado:2012:DDF

[PCW^{+}12] P. Prado, R. H. Carmichael, S. A. Watts, J. Cebrian, and K. L. Heck, Jr. Diet-dependent δ ^{13}C and δ ^{15}N fractionation among sea urchin Lytechinus variegatus tissues: implications for food web models. Marine Ecology Progress Series, 462:175–190, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m462p175.pdf.

Pechenik:2015:YSS

[PDB15] J. A. Pechenik, C. Diederich, and R. Burns. Yearly shifts in shell quality for the hermit crab Pagurus longicarpus in coastal Massachusetts. Marine Ecology Progress Series, 529: 171–183, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/529/m529p171.pdf.

Purcell:2014:FSV

[PDBB14] J. E. Purcell, M. B. Decker, D. L. Breitburg, and K. J. Broughton. Fine-scale vertical distributions of Mnemiopsis leidyi ctenophores: predation on copepods relative to stratification and hypoxia. Marine Ecology Progress Series, 500: 103–120, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/500/m500p103.pdf.

Pascal:2014:TRL

[PDBG14] P. Y. Pascal, S. Dubois, H. T. S. Boschker, and O. Gros. Trophic role of large benthic sulfur bacteria in mangrove sediment. Marine Ecology Progress Series, 516:127–138, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2014/516/m516p127.pdf.

Pechenik:2017:DRA

[PDC^{+}17] J. A. Pechenik, C. M. Diederich, O. R. Chaparro, J. A. Montory, F. J. Paredes, and A. M. Franklin. Differences in resource allocation to reproduction across the intertidal–subtidal gradient for two suspension-feeding marine gastropods: Crepidula fornicata and Crepipatella peruviana. Marine Ecology Progress Series, 572:165–178, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/572/m572p165.pdf.

Peters:2015:LPL

[PDCH15] J. Peters, R. Diekmann, C. Clemmesen, and W. Hagen. Lipids as a proxy for larval starvation and feeding condition in small pelagic fish: a field approach on match–mismatch effects on Baltic sprat. Marine Ecology Progress Series, 531: 277–292, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/531/m531p277.pdf.

Pascal:2017:IGS

[PDGL17] P. Y. Pascal, S. F. Dubois, A. Goffette, and G. Lepoint. Influences of geothermal sulfur bacteria on a tropical coastal food web. Marine Ecology Progress Series, 578:

73–85, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/578/m578p073.pdf.

Papaspyrou:2013:BFC

[PDGR^{+}13] S. Papaspyrou, P. Diz, E. García-Robledo, A. Corzo, and J. L. Jimenez-Arias. Benthic Foraminiferal community changes and their relationship to environmental dynamics in intertidal muddy sediments (Bay of Cádiz, SW Spain). Marine Ecology Progress Series, 490:121–135, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/490/m490p121.pdf.

Pasquaud:2010:ETR

[PDL^{+}10] S. Pasquaud, V. David, J. Lobry, M. Girardin, B. Sautour, and P. Elie. Exploitation of trophic resources by fish under stressful estuarine conditions. Marine Ecology Progress Series, 400:207–219, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/400/m400p207.pdf.

Papastamatiou:2012:ARS

[PDM12] Y. P. Papastamatiou, P. A. DeSalles, and D. J. McCauley. Area-restricted searching by manta rays and their response to spatial scale in lagoon habitats. Marine Ecology Progress Series, 456:233–244, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/456/m456p233.pdf.

Pahlow:2013:OBM

[PDO13] M. Pahlow, H. Dietze, and A. Oschlies. Optimality-based model of phytoplankton growth and diazotrophy. Marine Ecology Progress Series, 489:1–16, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Peron:2010:SVO

[PDP^{+}10] C. Péron, K. Delord, R. A. Phillips, Y. Charbonnier, C. Marteau, M. Louzao, and H. Weimerskirch. Seasonal variation in oceanographic habitat and behaviour of white-chinned petrels Procellaria aequinoctialis from Kerguelen Island. Marine Ecology Progress Series, 416:267–284, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/416/
m416p267.pdf.

Peckham:2011:DIA

[PDT^{+}11] S. H. Peckham, D. Maldonado Diaz, Y. Tremblay, R. Ochoa, J. Polovina, G. Balazs, P. H. Dutton, and W. J. Nichols. Demographic implications of alternative foraging strategies in juvenile loggerhead turtles Caretta caretta of the North Pacific Ocean. Marine Ecology Progress Series, 425:269– 280, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/425/m425p269.pdf.

Patricio:2016:NID

[PDvG16] A. R. Patrício, C. E. Diez, R. P. van Dam, and B. J. Godley. Novel insights into the dynamics of green turtle fibropapillomatosis. Marine Ecology Progress Series, 547:247– 255, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/547/m547p247.pdf.

Plaganyi:2014:EMP

[PEB^{+}14] É. E. Plagányi, N. Ellis, L. K. Blamey, E. B. Morello, A. Norman-Lopez, W. Robinson, M. Sporcic, and H. Sweatman. Ecosystem modelling provides clues to understanding ecological tipping points. Marine Ecology Progress Series, 512:99–113, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m512p099.pdf.

Pirtle:2012:HSI

[PES12] J. L. Pirtle, G. L. Eckert, and A. W. Stoner. Habitat structure influences the survival and predator–prey interactions of early juvenile red king crab Paralithodes camtschaticus. Marine Ecology Progress Series, 465:169–184, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m465p169.pdf.

Pascal:2013:RBF

[PFB^{+}13] P. Y. Pascal, J. W. Fleeger, H. T. S. Boschker, H. M. Mitwally, and D. S. Johnson. Response of the benthic food web to shortand long-term nutrient enrichment in saltmarsh mudflats.

Marine Ecology Progress Series, 474:27–41, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/474/
m474p027.pdf.

Patricio:2017:DGT

[PFB^{+}17] A. R. Patrício, A. Formia, C. Barbosa, A. C. Broderick, M. Bruford, C. Carreras, P. Catry, C. Ciofi, A. Regalla, and B. J. Godley. Dispersal of green turtles from Africa’s largest rookery assessed through genetic markers. Marine Ecology Progress Series, 569:215–225, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/569/m569p215.pdf.

Pecquerie:2012:RIF

[PFdP^{+}12] L. Pecquerie, R. Fablet, H. de Pontual, S. Bonhommeau, M. Alunno-Bruscia, P. Petitgas, and S. A. L. M. Kooijman. Reconstructing individual food and growth histories from biogenic carbonates. Marine Ecology Progress Series, 447:151–164, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/447/m447p151.pdf.

Perretti:2017:RSF

[PFF^{+}17] C. T. Perretti, M. J. Fogarty, K. D. Friedland, J. A. Hare, S. M. Lucey, R. S. McBride, T. J. Miller, R. E. Morse, L. O’Brien, J. J. Pereira, L. A. Smith, and M. J. Wuenschel. Regime shifts in fish recruitment on the Northeast US continental shelf. Marine Ecology Progress Series, 574:1–11, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Pommier:2010:RRD

[PFG10] J. Pommier, J. J. Frenette, and H. Glémet. Relating RNA: DNA ratio in Eurytemora affinis to seston fatty acids in a highly dynamic environment. Marine Ecology Progress Series, 400:143–154, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/400/m400p143.pdf.

Pini-Fitzsimmons:2018:EFP

[PFKB18] J. Pini-Fitzsimmons, N. A. Knott, and C. Brown. Effects of food provisioning on site use in the short-tail stingray Bathytoshia brevicaudata. Marine Ecology Progress Series, 600:

99–110, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/600/m600p099.pdf.

Pierson:2013:FFB

[PFL13] J. J. Pierson, B. W. Frost, and A. W. Leising. Foray foraging behavior: seasonally variable, food-driven migratory behavior in two calanoid copepod species. Marine Ecology Progress Series, 475:49–64, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m475p049.pdf.

Porter:2018:IEO

[PFL18] E. T. Porter, H. Franz, and R. Lacouture. Impact of Eastern oyster Crassostrea virginica biodeposit resuspension on the seston, nutrient, phytoplankton, and zooplankton dynamics: a mesocosm experiment. Marine Ecology Progress Series, 586:21–40, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/586/m586p021.pdf.

Plough:2018:RIM

[PFPP18] L. V. Plough, C. Fitzgerald, A. Plummer, and J. J. Pierson. Reproductive isolation and morphological divergence between cryptic lineages of the copepod Acartia tonsa in Chesapeake Bay. Marine Ecology Progress Series, 597:99– 113, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/597/m597p099.pdf.

Plante:2011:IED

[PFR11] C. Plante, E. Frank, and P. Roth. Interacting effects of deposit feeding and tidal resuspension on benthic microalgal community structure and spatial patterns. Marine Ecology Progress Series, 440:53–65, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/440/m440p053.pdf.

Pereira-Filho:2015:EST

[PFVFF^{+}15] G. H. Pereira-Filho, P. C. Veras, R. B. Francini-Filho, R. L. Moura, H. T. Pinheiro, F. Z. Gibran, Z. Matheus, L. M. Neves, and G. M. Amado-Filho. Effects of the sand tilefish Malacanthus plumieri on the structure and dynamics of a rhodolith bed in the Fernando de Noronha Archipelago, tropical West Atlantic. Marine Ecology Progress Series, 541: 65–73, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/541/m541p065.pdf.

Paul:2015:LIF

[PG15] M. Paul and L. G. Gillis. Let it flow: how does an underlying current affect wave propagation over a natural seagrass meadow? Marine Ecology Progress Series, 523:57– 70, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/523/m523p057.pdf.

Padilla-Gamino:2012:SDH

[PGG12] J. L. Padilla-Gamiño and R. D. Gates. Spawning dynamics in the Hawaiian reef-building coral Montipora capitata. Marine Ecology Progress Series, 449:145–160, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/449/
m449p145.pdf.

Provencher:2012:SDI

[PGOG12] J. F. Provencher, A. J. Gaston, P. D. O’Hara, and H. G. Gilchrist. Seabird diet indicates changing Arctic marine communities in eastern Canada. Marine Ecology Progress Series, 454:171–182, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m454p171.pdf.

Paiva:2010:FPP

[PGR^{+}10] V. H. Paiva, P. Geraldes, I. Ramírez, A. Meirinho, S. Garthe, and J. A. Ramos. Foraging plasticity in a pelagic seabird species along a marine productivity gradient. Marine Ecology Progress Series, 398:259–274, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/398/m398p259.pdf.

Paiva:2013:ODT

[PGR^{+}13] V. H. Paiva, P. Geraldes, I. Ramirez, A. C. Werner, S. Garthe, and J. A. Ramos. Overcoming difficult times: the behavioural resilience of a marine predator when facing environmental stochasticity. Marine Ecology Progress Series, 486:277–288, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/486/m486p277.pdf.

Powers:2017:CLS

[PGR^{+}17] S. P. Powers, J. H. Grabowski, H. Roman, A. Geggel, S. Rouhani, J. Oehrig, and M. Baker. Consequences of largescale salinity alteration during the Deepwater Horizon oil spill on subtidal oyster populations. Marine Ecology Progress Series, 576:175–187, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m576p175.pdf.

Prato:2014:BCF

[PGVF14] G. Prato, D. Gascuel, A. Valls, and P. Francour. Balancing complexity and feasibility in Mediterranean coastal food-web models: uncertainty and constraints. Marine Ecology Progress Series, 512:71–88, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m512p071.pdf.

Payne:2010:IDA

[PGWS10] N. L. Payne, B. M. Gillanders, D. M. Webber, and J. M. Semmens. Interpreting diel activity patterns from acoustic telemetry: the need for controls. Marine Ecology Progress Series, 419:295–301, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/419/m419p295.pdf.

Petraitis:2010:MSS

[PH10] P. S. Petraitis and C. Hoffman. Multiple stable states and relationship between thresholds in processes and states. Marine Ecology Progress Series, 413:189–200, 2010. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m413p189.pdf.

Prado:2011:SSO

[PH11] P. Prado and K. L. Heck, Jr. Seagrass selection by omnivorous and herbivorous consumers: determining factors. Marine Ecology Progress Series, 429:45–55, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m429p045.pdf.

Parry:2016:DDD

[PH16] G. D. Parry and A. J. Hirst. Decadal decline in demersal fish biomass coincident with a prolonged drought and the introduction of an exotic starfish. Marine Ecology Progress Series, 544:37–52, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/544/m544p037.pdf.

Phillips:2018:SOP

[PHAZ18] E. M. Phillips, J. K. Horne, J. Adams, and J. E. Zamon. Selective occupancy of a persistent yet variable coastal river plume by two seabird species. Marine Ecology Progress Series, 594:245–261, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/594/m594p245.pdf.

Paterson:2011:TSB

[PHB^{+}11a] D. M. Paterson, N. D. Hanley, K. Black, E. C. Defew, and M. Solan. Theme section: Biodiversity, ecosystems and coastal zone management: linking science and policy. Marine Ecology Progress Series, 434:201–301, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m434_ThemeSection.pdf.

Pemberton:2011:DSS

[PHB11b] A. J. Pemberton, L. J. Hansson, and J. D. D. Bishop. Does sperm supply limit the number of broods produced by a polyembryonous bryozoan? Marine Ecology Progress Series, 430:113–119, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m430p113.pdf.

Pinones:2013:MRL

[PHD^{+}13] A. Piñones, E. E. Hofmann, K. L. Daly, M. S. Dinniman, and J. M. Klinck. Modeling the remote and local connectivity of Antarctic krill populations along the western Antarctic Peninsula. Marine Ecology Progress Series, 481:69–92, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m481p069.pdf.

Pontasch:2014:PEA

[PHD^{+}14] S. Pontasch, R. Hill, E. Deschaseaux, P. L. Fisher, S. K. Davy, and A. Scott. Photochemical efficiency and antioxidant capacity in relation to Symbiodinium genotype and host phenotype in a symbiotic cnidarian. Marine Ecology Progress Series, 516:195–208, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/516/m516p195.pdf.

Puerta:2014:RHC

[PHG^{+}14] P. Puerta, M. Hidalgo, M. González, A. Esteban, and A. Quetglas. Role of hydro-climatic and demographic processes on the spatio-temporal distribution of cephalopods in the western Mediterranean. Marine Ecology Progress Series, 514: 105–118, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/514/m514p105.pdf.

Peklova:2012:DTP

[PHH^{+}12] I. Peklova, N. E. Hussey, K. J. Hedges, M. A. Treble, and A. T. Fisk. Depth and temperature preferences of the deepwater flatfish Greenland halibut Reinhardtius hippoglossoides in an Arctic marine ecosystem. Marine Ecology Progress Series, 467:193–205, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/467/m467p193.pdf.

Poulsen:2011:MPS

[PHHL^{+}11] N. A. Poulsen, J. Hemmer-Hansen, V. Loeschcke, G. R. Carvalho, and E. E. Nielsen. Microgeographical population structure and adaptation in Atlantic cod Gadus morhua: spatiotemporal insights from gene-associated DNA markers. Marine Ecology Progress Series, 436:231–243, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/436/
m436p231.pdf.

Phillips:2019:VSF

[PHHL19] L. R. Phillips, M. Hindell, A. J. Hobday, and M. A. Lea. Variability in at-sea foraging behaviour of little penguins Eudyptula minor in response to finescale environmental features.

Marine Ecology Progress Series, 627:141–154, 2019. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m627p141.pdf.

Paredes:2012:PMF

[PHI^{+}12] R. Paredes, A. M. A. Harding, D. B. Irons, D. D. Roby, R. M. Suryan, R. A. Orben, H. Renner, R. Young, and A. Kitaysky. Proximity to multiple foraging habitats enhances seabirds’ resilience to local food shortages. Marine Ecology Progress Series, 471:253–269, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/471/m471p253.pdf.

Porri:2011:AES

[PHM11] F. Porri, J. M. Hill, and C. D. McQuaid. Associations in ephemeral systems: the lack of trophic relationships between sandhoppers and beach wrack. Marine Ecology Progress Series, 426:253–262, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/426/m426p253.pdf.

Palmer:2016:ILP

[PHSBO16] G. Palmer, J. D. Hogan, B. D. Sterba-Boatwright, and R. D. Overath. Invasive lionfish Pterois volitans reduce the density but not the genetic diversity of a native reef fish. Marine Ecology Progress Series, 558:223–234, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m558p223.pdf.

Price:2011:SSC

[PHTS11] N. N. Price, S. L. Hamilton, J. S. Tootell, and J. E. Smith. Species-specific consequences of ocean acidification for the calcareous tropical green algae Halimeda. Marine Ecology Progress Series, 440:67–78, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/440/m440p067.pdf.

Prado:2013:SSU

[PHWC13] P. Prado, K. L. Heck, Jr., S. A. Watts, and J. Cebrian. δ ^{13}C and δ ^{18}O signatures from sea urchin skeleton: importance of diet type in metabolic contributions. Marine Ecology Progress Series, 476:153–166, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m476p153.pdf.

Prihartato:2016:GEM

[PIGK16] P. K. Prihartato, X. Irigoien, M. G. Genton, and S. Kaartvedt. Global effects of moon phase on nocturnal acoustic scattering layers. Marine Ecology Progress Series, 544:65–75, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/544/m544p065.pdf.

Pillay:2010:EEL

[Pil10] D. Pillay. Expanding the envelope: linking invertebrate bioturbators with micro-evolutionary change. Marine Ecology Progress Series, 409:301–303, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/409/m409p301.pdf.

Piche:2010:CFF

[PIPD10] J. Piché, S. J. Iverson, F. A. Parrish, and R. Dollar. Characterization of forage fish and invertebrates in the Northwestern Hawaiian Islands using fatty acid signatures: species and ecological groups. Marine Ecology Progress Series, 418:1–15, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Petrou:2013:SHP

[PJDC^{+}13] K. Petrou, I. Jimenez-Denness, K. Chartrand, C. McCormack, M. Rasheed, and P. J. Ralph. Seasonal heterogeneity in the photophysiological response to air exposure in two tropical intertidal seagrass species. Marine Ecology Progress Series, 482:93–106, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/482/m482p093.pdf.

Purcell:2018:NOG

[PJMA18] J. E. Purcell, A. R. Juhl, M. K. Mańko, and C. F. Aumack. NOTE: Overwintering of gelatinous zooplankton in the coastal Arctic Ocean. Marine Ecology Progress Series, 591:281–286, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/591/m591p281.pdf.

Plass-Johnson:2010:TEI

[PJMP10] J. G. Plass-Johnson, C. D. McQuaid, and F. Porri. Top-down effects on intertidal mussel populations: assessing two predator guilds in a South African marine protected area. Marine Ecology Progress Series, 411:149–159, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/411/
m411p149.pdf.

Pinet:2011:MWD

[PJP^{+}11] P. Pinet, S. Jaquemet, D. Pinaud, H. Weimerskirch, R. A. Phillips, and M. Le Corre. Migration, wintering distribution and habitat use of an endangered tropical seabird, Barau’s petrel Pterodroma baraui. Marine Ecology Progress Series, 423:291–302, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/423/m423p291.pdf.

Peacock:2013:APA

[PK13] M. Blakely Peacock and R. M. Kudela. Alkaline phosphatase activity detected in distinct phytoplankton communities in the northern Gulf of Alaska. Marine Ecology Progress Series, 473:79–90, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/473/m473p079.pdf.

Pendoley:2015:IHE

[PK15] K. Pendoley and R. L. Kamrowski. Influence of horizon elevation on the sea-finding behaviour of hatchling flatback turtles exposed to artificial light glow. Marine Ecology Progress Series, 529:279–288, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/529/m529p279.pdf.

Peach:2016:EEP

[PKB16] K. E. Peach, M. S. Koch, and P. L. Blackwelder. Effects of elevated pCO_{2} and irradiance on growth, photosynthesis and calcification in Halimeda discoidea. Marine Ecology Progress Series, 544:143–158, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/544/m544p143.pdf.

Porter:2014:RSI

[PKBS14] S. N. Porter, S. Kaehler, G. M. Branch, and K. J. Sink. Riverine subsidies for inshore filter-feeder communities: potential influences on trophic patterns among bioregions. Marine Ecology Progress Series, 498:13–26, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/498/
m498p013.pdf.

Park:2011:ODS

[PKS11a] J. I. Park, C. K. Kang, and H. L. Suh. Ontogenetic diet shift in the euphausiid Euphausia pacifica quantified using stable isotope analysis. Marine Ecology Progress Series, 429: 103–109, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/429/m429p103.pdf.

Pittman:2011:PCS

[PKS11b] S. J. Pittman, R. T. Kneib, and C. A. Simenstad. Practicing coastal seascape ecology. Marine Ecology Progress Series, 427:187–190, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m427p187.pdf.

Pittman:2011:SEA

[PKSN11] Simon Pittman, Ron Kneib, Charles Simenstad, and Ivan Nagelkerken. Seascape ecology: Application of landscape ecology to marine. Marine Ecology Progress Series, 427:187– 302, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m427_ThemeSection.pdf.

Paul:2011:CMI

[PKW^{+}11] V. J. Paul, I. B. Kuffner, L. J. Walters, R. Ritson-Williams, K. S. Beach, and M. A. Becerro. Chemically mediated interactions between macroalgae Dictyota spp. and multiple lifehistory stages of the coral Porites astreoides. Marine Ecology Progress Series, 426:161–170, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/426/m426p161.pdf.

Piazza:2010:UGA

[PL10] B. P. Piazza and M. K. La Peyre. Using Gambusia affinis growth and condition to assess estuarine habitat quality: a comparison of indices. Marine Ecology Progress Series, 412:231–245, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/412/m412p231.pdf.

Passow:2015:OAO

[PL15] U. Passow and E. A. Laws. Ocean acidification as one of multiple stressors: growth response of Thalassiosira weissflogii (diatom) under temperature and light stress. Marine Ecology Progress Series, 541:75–90, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m541p075.pdf.

Persson:2012:FPJ

[PLA^{+}12] A. Persson, P. Ljungberg, M. Andersson, E. Götzman, and P. A. Nilsson. Foraging performance of juvenile Atlantic cod Gadus morhua and profitability of coastal habitats. Marine Ecology Progress Series, 456:245–253, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/456/
m456p245.pdf.

Pomerleau:2012:PAI

[PLF^{+}12a] C. Pomerleau, V. Lesage, S. H. Ferguson, G. Winkler, S. D. Petersen, and J. W. Higdon. Prey assemblage isotopic variability as a tool for assessing diet and the spatial distribution of bowhead whale Balaena mysticetus foraging in the Canadian eastern Arctic. Marine Ecology Progress Series, 469:161–174, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/469/m469p161.pdf.

Pranovi:2012:TLD

[PLF^{+}12b] F. Pranovi, J. Link, C. Fu, A. M. Cook, H. Liu, S. Gaichas, K. D. Friedland, K. Rong Utne, and H. P. Benoît. Trophiclevel determinants of biomass accumulation in marine ecosystems. Marine Ecology Progress Series, 459:185–201, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m459p185.pdf.

Peteiro:2011:IIU

[PLFR^{+}11] L. G. Peteiro, U. Labarta, M. J. Fernández-Reiriz, X. A. Álvarez-Salgado, R. Filgueira, and S. Piedracoba. Influence of intermittent-upwelling on Mytilus galloprovincialis settlement patterns in the Ría de ares-betanzos. Marine Ecology Progress Series, 443:111–127, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/443/m443p111.pdf.

Passy:2016:EMC

[PLG^{+}16] P. Passy, R. Le Gendre, J. Garnier, P. Cugier, J. Callens, F. Paris, G. Billen, P. Riou, and E. Romero. Eutrophication modelling chain for improved management strategies to prevent algal blooms in the Bay of Seine. Marine Ecology Progress Series, 543:107–125, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m543p107.pdf.

Paar:2019:FSM

[PLG^{+}19] M. Paar, B. Lebreton, M. Graeve, M. Greenacre, R. Asmus, and H. Asmus. Food sources of macrozoobenthos in an Arctic kelp belt: trophic relationships revealed by stable isotope and fatty acid analyses. Marine Ecology Progress Series, 615:31–49, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m615p031.pdf.

Phillips:2017:CCI

[PLGSD17] R. A. Phillips, S. Lewis, J. González-Solís, and F. Daunt. Causes and consequences of individual variability and specialization in foraging and migration strategies of seabirds. Marine Ecology Progress Series, 578:117–150, 2017. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m578p117.pdf.

Pitois:2012:BEC

[PLJ^{+}12] S. G. Pitois, C. P. Lynam, T. Jansen, N. Halliday, and M. Edwards. Bottom-up effects of climate on fish populations: data from the continuous plankton recorder. Marine Ecology Progress Series, 456:169–186, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/456/m456p169.pdf.

Pascoe:2016:AUT

[PLM^{+}16] P. Pascoe, M. A. Lea, R. H. Mattlin, C. R. McMahon, R. Harcourt, D. Thompson, L. Torres, K. Vinette-Herrin, and M. A. Hindell. Assessing the utility of two- and three-dimensional behavioural metrics in habitat usage models. Marine Ecology Progress Series, 562:181–192, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/562/m562p181.pdf.

Palmer:2010:DFA

[PLMN10] M. Palmer, M. Linde, and B. Morales-Nin. Disentangling fluctuating asymmetry from otolith shape. Marine Ecology Progress Series, 399:261–272, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/399/m399p261.pdf.

Polito:2011:SIR

[PLNE11] M. J. Polito, H. J. Lynch, R. Naveen, and S. D. Emslie. Stable isotopes reveal regional heterogeneity in the pre-breeding distribution and diets of sympatrically breeding Pygoscelis spp. penguins. Marine Ecology Progress Series, 421:265– 277, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/421/m421p265.pdf.

Perkins:2010:VCM

[PLS^{+}10] R. G. Perkins, J. Lavaud, J. Serôdio, J. L. Mouget, P. Cartaxana, P. Rosa, L. Barille, V. Brotas, and B. M. Jesus. Vertical cell movement is a primary response of intertidal benthic biofilms to increasing light dose. Marine Ecology Progress Series, 416:93–103, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/416/m416p093.pdf.

Pranovi:2014:BAA

[PLZL14] F. Pranovi, S. Libralato, M. Zucchetta, and J. Link. Biomass accumulation across trophic levels: analysis of landings for the Mediterranean Sea. Marine Ecology Progress Series, 512: 201–216, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m512p201.pdf.

Pereira:2016:CCS

[PM16] P. H. C. Pereira and P. L. Munday. Coral colony size and structure as determinants of habitat use and fitness of coraldwelling fishes. Marine Ecology Progress Series, 553:163– 172, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/553/m553p163.pdf.

Penin:2010:EPS

[PMB^{+}10] L. Penin, F. Michonneau, A. H. Baird, S. R. Connolly, M. S. Pratchett, M. Kayal, and M. Adjeroud. Early postsettlement mortality and the structure of coral assemblages. Marine Ecology Progress Series, 408:55–64, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/408/
m408p055.pdf.

Pendleton:2015:TAI

[PMB^{+}15] L. Pendleton, R. Mongruel, N. Beaumont, T. Hooper, and M. Charles. A triage approach to improve the relevance of marine ecosystem services assessments. Marine Ecology Progress Series, 530:183–193, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m530p183.pdf.

Patricio:2017:BPS

[PMB^{+}17] A. R. Patrício, A. Marques, C. Barbosa, A. C. Broderick, B. J. Godley, L. A. Hawkes, R. Rebelo, A. Regalla, and P. Catry. Balanced primary sex ratios and resilience to climate change in a major sea turtle population. Marine Ecology Progress Series, 577:189–203, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/577/m577p189.pdf.

Poulsen:2011:FCF

[PMBH11] L. K. Poulsen, M. Moldrup, T. Berge, and P. J. Hansen. Feeding on copepod fecal pellets: a new trophic role of dinoflagellates as detritivores. Marine Ecology Progress Series, 441:65–78, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/441/m441p065.pdf.

Pawlik:2015:NEF

[PMEZ15a] J. R. Pawlik, S. E. McMurray, P. Erwin, and S. Zea. No evidence for food limitation of Caribbean reef sponges: Reply to slattery & lesser (2015). Marine Ecology Progress Series, 527:281–284, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m527p281.pdf.

Pawlik:2015:REF

[PMEZ15b] Joseph R. Pawlik, Steven E. McMurray, Patrick Erwin, and Sven Zea. A review of evidence for food limitation of sponges on Caribbean reefs. Marine Ecology Progress Series, 519: 265–283, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m519p265.pdf. See [SL15a].

Pogorzelec:2017:FIA

[PMF^{+}17] N. M. Pogorzelec, C. J. Mundy, C. R. Findlay, K. Campbell, A. Diaz, J. K. Ehn, S. Rysgaard, and K. M. Gough. FTIR imaging analysis of cell content in sea-ice diatom taxa during a spring bloom in the lower northwest passage of the Canadian Arctic. Marine Ecology Progress Series, 569: 77–88, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/569/m569p077.pdf.

Patino-Martinez:2017:ITF

[PMGQM17] J. Patino-Martinez, B. J. Godley, L. Quiñones, and A. Marco. Impact of tropical forest logging on the reproductive success of leatherback turtles. Marine Ecology Progress Series, 569:205–214, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/569/m569p205.pdf.

Poland:2013:VSC

[PMH^{+}13] D. M. Poland, J. M. Mansfield, A. R. Hannes, C. L. F. Lewis, T. L. Shearer, S. J. Connelly, N. L. Kirk, and M. A. Coffroth. Variation in Symbiodinium communities in juvenile Briareum asbestinum (Cnidaria: Octocorallia) over temporal and spatial scales. Marine Ecology Progress Series, 476:

23–37, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/476/m476p023.pdf.

Pineda-Metz:2019:BCF

[PMIG19] S. E. A. Pineda-Metz, E. Isla, and D. Gerdes. Benthic communities of the Filchner Region (Weddell Sea, Antarctica). Marine Ecology Progress Series, 628:37–54, 2019. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m628p037.pdf.

Papastamatiou:2015:MFP

[PMK^{+}15] Y. P. Papastamatiou, C. G. Meyer, R. K. Kosaki, N. J. Wallsgrove, and B. N. Popp. Movements and foraging of predators associated with mesophotic coral reefs and their potential for linking ecological habitats. Marine Ecology Progress Series, 521:155–170, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m521p155.pdf.

Pulkkinen:2011:MKS

[PMKL11] H. Pulkkinen, S. Mäntyniemi, S. Kuikka, and P. Levontin. More knowledge with the same amount of data: advantage of accounting for parameter correlations in hierarchical metaanalyses. Marine Ecology Progress Series, 443:29–37, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m443p029.pdf.

Podowski:2010:BAF

[PML^{+}10] E. L. Podowski, S. Ma, G. W. Luther III, D. Wardrop, and C. R. Fisher. Biotic and abiotic factors affecting distributions of megafauna in diffuse flow on andesite and basalt along the Eastern Lau Spreading Center, Tonga. Marine Ecology Progress Series, 418:25–45, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m418p025.pdf.

Price:2019:GBC

[PML^{+}19] N. N. Price, S. Muko, L. Legendre, R. Steneck, M. J. H. van Oppen, R. Albright, P. Ang, Jr., R. C. Carpenter, A. P. Y.

Chui, T. Y. Fan, R. D. Gates, S. Harii, H. Kitano, H. Kurihara, S. Mitarai, J. L. Padilla-Gamiño, K. Sakai, G. Suzuki, and P. J. Edmunds. Global biogeography of coral recruitment: tropical decline and subtropical increase. Marine Ecology Progress Series, 621:1–17, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Pirotta:2011:MSW

[PMM^{+}11] E. Pirotta, J. Matthiopoulos, M. MacKenzie, L. Scott-Hayward, and L. Rendell. Modelling sperm whale habitat preference: a novel approach combining transect and follow data. Marine Ecology Progress Series, 436:257–272, 2011. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m436p257.pdf.

Pires:2015:ELT

[PMM^{+}15] A. Pires, R. Martins, L. Magalhães, A. M. V. M. Soares, E. Figueira, V. Quintino, A. M. Rodrigues, and R. Freitas. Expansion of lugworms towards southern European habitats and their identification using combined ecological, morphological and genetic approaches. Marine Ecology Progress Series, 533:177–190, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/533/m533p177.pdf.

Petersen:2012:BGI

[PMML12] J. K. Petersen, M. Maar, F. Møhlenberg, and J. E. N. Larsen. Benthic grazing impact: coupling and uncoupling in relation to physical forcing. Marine Ecology Progress Series, 463:127–139, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/463/m463p127.pdf.

Perez-Matus:2017:TRS

[PMOAC^{+}17] A. Pérez-Matus, A. Ospina-Alvarez, P. A. Camus, S. A. Carrasco, M. Fernandez, S. Gelcich, N. Godoy, F. P. Ojeda, L. M. Pardo, N. Rozbaczylo, M. D. Subida, M. Thiel, E. A. Wieters, and S. A. Navarrete. Temperate rocky subtidal reef community reveals human impacts across the entire food web. Marine Ecology Progress Series, 567:1–16, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Perez-Matus:2010:DTM

[PMS10a] A. Pérez-Matus and J. S. Shima. Density- and trait-mediated effects of fish predators on amphipod grazers: potential indirect benefits for the giant kelp Macrocystis pyrifera. Marine Ecology Progress Series, 417:151–158, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/417/
m417p151.pdf.

Porter:2010:ETR

[PMS10b] E. T. Porter, R. P. Mason, and L. P. Sanford. Effect of tidal resuspension on benthic-pelagic coupling in an experimental ecosystem study. Marine Ecology Progress Series, 413: 33–53, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m413p033.pdf.

Porter:2013:ESS

[PMS13] E. T. Porter, R. P. Mason, and L. P. Sanford. Effects of shear stress and hard clams on seston, microphytobenthos, and nitrogen dynamics in mesocosms with tidal resuspension. Marine Ecology Progress Series, 479:25–45, 2013. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m479p025.pdf.

Paul:2015:WEC

[PMS15] C. Paul, B. Matthiessen, and U. Sommer. Warming, but not enhanced CO_{2} concentration, quantitatively and qualitatively affects phytoplankton biomass. Marine Ecology Progress Series, 528:39–51, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m528p039.pdf.

Perez-Matus:2016:UAA

[PMSGBL16] A. Pérez-Matus, F. Sánchez, J. C. González-But, and R. W. Lamb. Understory algae associations and predation risk influence broad-scale kelp habitat use in a temperate reef fish. Marine Ecology Progress Series, 559:147–158, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/559/
m559p147.pdf.

Petersen:2013:NBG

[PMYH13] J. K. Petersen, M. Maar, T. Ysebaert, and P. M. J. Herman. Near-bed gradients in particles and nutrients above a mussel bed in the Limfjorden: influence of physical mixing and mussel filtration. Marine Ecology Progress Series, 490:137–146, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/490/m490p137.pdf.

Pascual:2016:EDS

[PNA^{+}16] M. Pascual, M. G. Neubert, J. L. Acuña, A. R. Solow, C. Dominguez-Carrió, J. Salvador, A. Olariaga, and V. Fuentes. Environmental drivers of salp Thalia democratica population dynamics from in situ observations. Marine Ecology Progress Series, 561:189–201, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/561/m561p189.pdf.

Pedersen:2012:EWE

[PNF^{+}12] M. F. Pedersen, L. B. Nejrup, S. Fredriksen, H. Christie, and K. M. Norderhaug. Effects of wave exposure on population structure, demography, biomass and productivity of the kelp Laminaria hyperborea. Marine Ecology Progress Series, 451:45–60, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m451p045.pdf.

Paesch:2014:EFC

[PNI14] L. Paesch, W. Norbis, and P. Inchausti. Effects of fishing and climate variability on spatio-temporal dynamics of demersal chondrichthyans in the Río de la Plata, SW Atlantic. Marine Ecology Progress Series, 508:187–200, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/508/
m508p187.pdf.

Pedersen:2014:SCL

[PNPF14] M. F. Pedersen, L. B. Nejrup, T. M. Pedersen, and S. Fredriksen. Sub-canopy light conditions only allow low annual net productivity of epiphytic algae on kelp Laminaria hyperborea. Marine Ecology Progress Series, 516:163–176, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m516p163.pdf.

Pahlow:2013:OAB

[PO13] M. Pahlow and A. Oschlies. Optimal allocation backs Droop’s cell-quota model. Marine Ecology Progress Series, 473:1–5, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Pommer:2016:IDB

[POH16] C. D. Pommer, M. Olesen, and J. L. S. Hansen. Impact and distribution of bottom trawl fishing on mud-bottom communities in the Kattegat. Marine Ecology Progress Series, 548: 47–60, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/548/m548p047.pdf.

Pina-Ochoa:2010:SLS

[POKGRP10] E. Piña-Ochoa, K. A. Koho, E. Geslin, and N. Risgaard-Petersen. Survival and life strategy of the Foraminiferan Globobulimina turgida through nitrate storage and denitrification. Marine Ecology Progress Series, 417:39–49, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2010/417/m417p039.pdf.

Ptacnik:2011:DPD

[POL^{+}11] R. Ptacnik, K. Olli, S. Lehtinen, T. Tamminen, and T. Andersen. Does plankton diversity peak at intermediate salinities? Comment on Telesh et al. (2011). Marine Ecology Progress Series, 432:291–292, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m432p291.pdf. See [TSS11b].

Passarelli:2012:IBS

[POPH12] C. Passarelli, F. Olivier, D. M. Paterson, and C. Hubas. Impacts of biogenic structures on benthic assemblages: microbes, meiofauna, macrofauna and related ecosystem functions. Marine Ecology Progress Series, 465:85–97, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/465/
m465p085.pdf.

Portner:2012:ICR

[Pör12] H. O. Pörtner. Integrating climate-related stressor effects on marine organisms: unifying principles linking molecule to ecosystem-level changes. Marine Ecology Progress Series, 470:273–290, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m470p273.pdf.

Paredes:2015:FED

[POR^{+}15] R. Paredes, R. A. Orben, D. D. Roby, D. B. Irons, R. Young, H. Renner, Y. Tremblay, A. Will, A. M. A. Harding, and A. S. Kitaysky. Foraging ecology during nesting influences body size in a pursuit-diving seabird. Marine Ecology Progress Series, 533:261–276, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/533/m533p261.pdf.

Peacock:2014:PID

[POS14] E. E. Peacock, R. J. Olson, and H. M. Sosik. Parasitic infection of the diatom Guinardia delicatula, a recurrent and ecologically important phenomenon on the New England shelf. Marine Ecology Progress Series, 503:1–10, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Pombo:2017:NDD

[POX^{+}17] M. Pombo, A. L. Oliveira, L. Y. Xavier, E. Siegle, and A. Turra. Natural drivers of distribution of ghost crabs Ocypode quadrata and the implications of estimates from burrows. Marine Ecology Progress Series, 565:131–147, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/565/
m565p131.pdf.

Pahlow:2010:MOC

[PP10] M. Pahlow and A. E. F. Prowe. Model of optimal current feeding in zooplankton. Marine Ecology Progress Series, 403:129–144, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/403/m403p129.pdf.

Pillay:2013:BME

[PP13] D. Pillay and R. Perissinotto. Benthic macrofauna of an estuarine lake during a drought: spatio-temporal drivers under different hydrological states. Marine Ecology Progress Series, 492:111–123, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/492/m492p111.pdf.

Putnam:2018:AIC

[PP18] A. B. Putnam and P. Peckol. Asymmetric interference competition between herbivorous gastropods, introduced Littorina littorea and indigenous L. obtusata. Marine Ecology Progress Series, 594:135–147, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/594/m594p135.pdf.

Pedrotti:2010:ENT

[PPB^{+}10a] M. L. Pedrotti, F. Peters, S. Beauvais, M. Vidal, J. Egge, A. Jacobsen, and C. Marrasé. Effects of nutrients and turbulence on the production of transparent exopolymer particles: a mesocosm study. Marine Ecology Progress Series, 419:57–69, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/419/m419p057.pdf.

Perissinotto:2010:MBS

[PPB10b] R. Perissinotto, D. Pillay, and G. Bate. Microalgal biomass in the St Lucia Estuary during the 2004 to 2007 drought period. Marine Ecology Progress Series, 405:147–161, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/405/
m405p147.pdf.

Peterson:2017:CID

[PPB^{+}17] C. D. Peterson, K. T. Parsons, D. M. Bethea, W. B. Driggers III, and R. J. Latour. Community interactions and density dependence in the southeast United States coastal shark complex. Marine Ecology Progress Series, 579:81– 96, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/579/m579p081.pdf.

Pinna:2012:HPS

[PPC^{+}12] S. Pinna, A. Pais, P. Campus, N. Sechi, and G. Ceccherelli. Habitat preferences of the sea urchin Paracentrotus lividus. Marine Ecology Progress Series, 445:173–180, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/445/
m445p173.pdf.

Pondella:2019:ADR

[PPC^{+}19] D. J. Pondella II, S. E. Piacenza, J. T. Claisse, C. M. Williams, J. P. Williams, A. J. Zellmer, and J. E. Caselle. Assessing drivers of rocky reef fish biomass density from the Southern California Bight. Marine Ecology Progress Series, 628: 125–140, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m628p125.pdf.

Powers:2017:RNE

[PPCH17] S. P. Powers, C. H. Peterson, J. Cebrian, and K. L. Heck, Jr. Response of nearshore ecosystems to the Deepwater Horizon oil spill. Marine Ecology Progress Series, 576:107–110, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m576p107.pdf.

Perrin:2014:TOC

[PPD^{+}14] G. Perrin, S. Plourde, C. DiBacco, G. Winkler, and P. Sirois. Tracing the origins of Calanus sp. in the Saguenay-St. Lawrence Marine Park (Québec, Canada) using δ ^{13}C as a marker. Marine Ecology Progress Series, 499:89–102, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2014/499/m499p089.pdf.

Purcell:2016:MGC

[PPDW16] S. W. Purcell, T. P. Piddocke, S. J. Dalton, and Y. G. Wang. Movement and growth of the coral reef holothuroids Bohadschia argus and Thelenota ananas. Marine Ecology Progress Series, 551:201–214, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/551/m551p201.pdf.

Pino-Pinuer:2014:CCR

[PPEH^{+}14] P. Pino-Pinuer, R. Escribano, P. Hidalgo, R. Riquelme-Bugueño, and W. Schneider. Copepod community response to variable upwelling conditions off central-southern Chile during 2002–2004 and 2010–2012. Marine Ecology Progress Series, 515:83–95, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/515/m515p083.pdf.

Pepin:2011:LAC

[PPH11] P. Pepin, C. C. Parrish, and E. J. H. Head. Late autumn condition of Calanus finmarchicus in the northwestern Atlantic: evidence of size-dependent differential feeding. Marine Ecology Progress Series, 423:155–166, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/423/m423p155.pdf.

Parent:2015:PFC

[PPJT15] G. J. Parent, S. Plourde, P. Joly, and J. Turgeon. Phenology and fitness of Calanus glacialis, C. finmarchicus (Copepoda), and their hybrids in the St. Lawrence Estuary. Marine Ecology Progress Series, 524:1–9, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Pollack:2011:LTT

[PPM11] J. Beseres Pollack, T. A. Palmer, and P. A. Montagna. Long-term trends in the response of benthic macrofauna to climate variability in the Lavaca–Colorado Estuary, Texas. Marine Ecology Progress Series, 436:67–80, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/436/
m436p067.pdf.

Patel:2015:DSR

[PPM^{+}15] S. H. Patel, A. Panagopoulou, S. J. Morreale, S. S. Kilham, I. Karakassis, T. Riggall, D. Margaritoulis, and J. R. Spotila. Differences in size and reproductive output of loggerhead turtles Caretta caretta nesting in the eastern Mediterranean Sea are linked to foraging site. Marine Ecology Progress Series, 535:231–241, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m535p231.pdf.

Pace:2017:CAF

[PPML17] S. M. Pace, E. N. Powell, R. Mann, and M. C. Long. Comparison of age frequency distributions for ocean quahogs Arctica islandica on the western Atlantic US continental shelf. Marine Ecology Progress Series, 585:81–98, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/585/
m585p081.pdf.

Pires:2013:MVL

[PPS^{+}13] R. F. T. Pires, M. Pan, A. M. P. Santos, Á. Peliz, D. Boutov, and A. dos Santos. Modelling the variation in larval dispersal of estuarine and coastal ghost shrimp: Upogebia congeners in the Gulf of Cadiz. Marine Ecology Progress Series, 492:153–168, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/492/m492p153.pdf.

Phelps:2015:BIL

[PPSR15] J. J. C. Phelps, J. A. Polton, A. J. Souza, and L. A. Robinson. Behaviour influences larval dispersal in shelf sea gyres: Nephrops norvegicus in the Irish Sea. Marine Ecology Progress Series, 518:177–191, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m518p177.pdf.

Perez-Portela:2012:BLG

[PPTB12] R. Pérez-Portela, X. Turon, and J. D. D. Bishop. Bottlenecks and loss of genetic diversity: spatio-temporal patterns of genetic structure in an ascidian recently introduced in Europe. Marine Ecology Progress Series, 451:93–105, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/451/
m451p093.pdf.

Putts:2019:SDH

[PPTK19] M. R. Putts, F. A. Parrish, F. A. Trusdell, and S. E. Kahng. Structure and development of Hawaiian deep-water coral communities on Mauna Loa lava flows. Marine Ecology Progress Series, 630:69–82, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m630p069.pdf.

Pereira:2017:USM

[PPX17] J. M. Pereira, V. H. Paiva, and J. C. Xavier. Using seabirds to map the distribution of elusive pelagic cephalopod species. Marine Ecology Progress Series, 567:257–262, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/567/
m567p257.pdf.

Piontkovski:2017:SAB

[PQAH^{+}17] S. A. Piontkovski, B. Y. Queste, K. A. Al-Hashmi, A. Al-Shaaibi, Y. V. Bryantseva, and E. A. Popova. Subsurface algal blooms of the northwestern Arabian Sea. Marine Ecology Progress Series, 566:67–78, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m566p067.pdf.

Petrou:2011:PNP

[PR11] K. Petrou and P. J. Ralph. Photosynthesis and net primary productivity in three Antarctic diatoms: possible significance for their distribution in the Antarctic marine ecosystem. Marine Ecology Progress Series, 437:27–40, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/437/
m437p027.pdf.

Prado:2010:NSP

[PRA10] P. Prado, J. Romero, and T. Alcoverro. Nutrient status, plant availability and seasonal forcing mediate fish herbivory in temperate seagrass beds. Marine Ecology Progress Series, 409:229–239, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/409/m409p229.pdf.

Proenca:2019:DEE

[PRA^{+}19] B. Proença, M. Romuald, I. Auby, F. Ganthy, A. Sottolichio, and R. Michalet. Disentangling ecosystem engineering from short-term biotic effects of a strong invader on a native foundation species. Marine Ecology Progress Series, 621:69–81, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/621/m621p069.pdf.

Powers:2017:ESL

[PRB^{+}17] S. P. Powers, S. Rouhani, M. C. Baker, H. Roman, J. Grabowski, S. B. Scyphers, J. M. Willis, and M. W. Hester. Ecosystem services are lost when facilitation between two ecosystem engineers is compromised by oil. Marine Ecology Progress Series, 576:189–202, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m576p189.pdf.

Prouty:2011:GRA

[PRBR11] N. G. Prouty, E. B. Roark, N. A. Buster, and S. W. Ross. Growth rate and age distribution of deep-sea black corals in the Gulf of Mexico. Marine Ecology Progress Series, 423: 101–115, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/423/m423p101.pdf.

Papacostas:2017:BMM

[PRCG^{+}17] K. J. Papacostas, E. W. Rielly-Carroll, S. E. Georgian, D. J. Long, S. D. Princiotta, A. M. Quattrini, K. E. Reuter, and A. L. Freestone. Biological mechanisms of marine invasions. Marine Ecology Progress Series, 565:251–268, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/565/
m565p251.pdf.

Planes:2011:FHL

[PRFA11] S. Planes, N. Raventos, B. Ferrari, and T. Alcoverro. Fish herbivory leads to shifts in seagrass Posidonia oceanica investments in sexual reproduction. Marine Ecology Progress Series, 431:205–213, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2011/431/m431p205.pdf.

Proietti:2012:GTC

[PRK^{+}12] M. C. Proietti, J. W. Reisser, P. G. Kinas, R. Kerr, D. S. Monteiro, L. F. Marins, and E. R. Secchi. Green turtle Chelonia mydas mixed stocks in the western South Atlantic, as revealed by mtDNA haplotypes and drifter trajectories. Marine Ecology Progress Series, 447:195–209, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/447/m447p195.pdf.

Peck-Richardson:2018:TDF

[PRLR^{+}18] A. G. Peck-Richardson, D. E. Lyons, D. D. Roby, D. A. Cushing, and J. A. Lerczak. Three-dimensional foraging habitat use and niche partitioning in two sympatric seabird species, Phalacrocorax auritus and P. penicillatus. Marine Ecology Progress Series, 586:251–264, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/586/m586p251.pdf.

Pratte:2017:FSN

[PRM17] I. Pratte, G. J. Robertson, and M. L. Mallory. Four sympatrically nesting auks show clear resource segregation in their foraging environment. Marine Ecology Progress Series, 572:243–254, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/572/m572p243.pdf.

Perales-Raya:2018:CSA

[PRNR^{+}18] C. Perales-Raya, M. Nande, A. Roura, A. Bartolomé, C. Gestal, J. J. Otero, P. García-Fernández, and E. Almansa. Comparative study of age estimation in wild and cultured Octopus vulgaris paralarvae: effect of temperature and diet. Marine Ecology Progress Series, 598:247–259, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m598p247.pdf.

Proffitt:2017:PBS

[Pro17] C. E. Proffitt. Phytoplankton biomass in a subtropical estuary: drivers, blooms, and ecological functions assessed over space and time using structural equation modeling. Marine Ecology Progress Series, 569:55–75, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/569/m569p055.pdf.

Prado:2016:WAM

[PRP^{+}16] P. Prado, A. Roque, J. Pérez, C. Ibáñez, C. Alcaraz, F. Casals, and N. Caiola. Warming and acidification-mediated resilience to bacterial infection determine mortality of early Ostrea edulis life stages. Marine Ecology Progress Series, 545:189– 202, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/545/m545p189.pdf.

Prebble:2018:LLR

[PRP^{+}18] C. E. M. Prebble, C. A. Rohner, S. J. Pierce, D. P. Robinson, M. Y. Jaidah, S. S. Bach, and C. N. Trueman. Limited latitudinal ranging of juvenile whale sharks in the Western Indian Ocean suggests the existence of regional management units. Marine Ecology Progress Series, 601:167–183, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m601p167.pdf.

Pritchard:2016:VLS

[PRR^{+}16] C. E. Pritchard, R. N. Rimler, S. S. Rumrill, R. B. Emlet, and A. L. Shanks. Variation in larval supply and recruitment of Ostrea lurida in the Coos Bay estuary, Oregon, USA. Marine Ecology Progress Series, 560:159–171, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/560/
m560p159.pdf.

Pearson:2015:DCS

[PS15a] R. Pearson and T. Stevens. Distinct cross-shelf gradient in mesophotic reef fish assemblages in subtropical eastern Australia. Marine Ecology Progress Series, 532:185–196, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m532p185.pdf.

Peteiro:2015:HSB

[PS15b] L. G. Peteiro and A. L. Shanks. Up and down or how to stay in the bay: retentive strategies of Olympia oyster larvae in a shallow estuary. Marine Ecology Progress Series, 530:103–117, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/530/m530p103.pdf.

Pfirrmann:2019:ESR

[PS19] B. W. Pfirrmann and R. D. Seitz. Ecosystem services of restored oyster reefs in a Chesapeake Bay tributary: abundance and foraging of estuarine fishes. Marine Ecology Progress Series, 628:155–169, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m628p155.pdf.

Pereira:2012:GAH

[PSA12] J. J. Pereira, E. T. Schultz, and P. J. Auster. Geospatial analysis of habitat use in yellowtail flounder Limanda ferruginea on Georges Bank. Marine Ecology Progress Series, 468:279–290, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/468/m468p279.pdf.

Persico:2017:FHD

[PSA17] E. P. Persico, S. J. Sharp, and C. Angelini. Feral hog disturbance alters carbon dynamics in southeastern US salt marshes. Marine Ecology Progress Series, 580:57–68, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/580/
m580p057.pdf.

Petitet:2012:AGL

[PSAK12] R. Petitet, E. R. Secchi, L. Avens, and P. G. Kinas. Age and growth of loggerhead sea turtles in southern Brazil. Marine Ecology Progress Series, 456:255–268, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/456/
m456p255.pdf.

Prowse:2017:TSL

[PSB17] T. A. A. Prowse, M. A. Sewell, and M. Byrne. Threestage lipid dynamics during development of planktotrophic echinoderm larvae. Marine Ecology Progress Series, 583:149– 161, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/583/m583p149.pdf.

Payne:2013:RDN

[PSC^{+}13a] M. R. Payne, D. R. Stine, L. Worsøe Clausen, P. Munk, H. Mosegaard, and R. D. M. Nash. Recruitment decline in North Sea herring is accompanied by reduced larval growth rates. Marine Ecology Progress Series, 489:197– 211, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/489/m489p197.pdf.

Pinna:2013:CSE

[PSC13b] S. Pinna, N. Sechi, and G. Ceccherelli. Canopy structure at the edge of seagrass affects sea urchin distribution. Marine Ecology Progress Series, 485:47–55, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/485/m485p047.pdf.

Parsons:2016:EHF

[PSCS16] D. F. Parsons, I. M. Suthers, D. O. Cruz, and J. A. Smith. Effects of habitat on fish abundance and species composition on temperate rocky reefs. Marine Ecology Progress Series, 561:155–171, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/561/m561p155.pdf.

Peel:2019:MRP

[PSD^{+}19] L. R. Peel, G. M. W. Stevens, R. Daly, C. A. Keating Daly, J. S. E. Lea, C. R. Clarke, S. P. Collin, and M. G. Meekan. Movement and residency patterns of reef manta rays Mobula alfredi in the Amirante Islands, Seychelles. Marine Ecology Progress Series, 621:169–184, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/621/m621p169.pdf.

Pereira:2017:ASA

[PSdA^{+}17] T. J. Pereira, A. F. Silva, P. R. de Almeida, A. F. Belo, J. L. Costa, N. Castro, and B. R. Quintella. Assessing the size adequacy of a small no-take marine protected area (MPA) for Mediterranean moray and European conger. Marine Ecology Progress Series, 584:213–227, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/584/m584p213.pdf.

Pierson:2017:SES

[PSER17] J. J. Pierson, W. C. L. Slater, D. Elliott, and M. R. Roman. Synergistic effects of seasonal deoxygenation and temperature truncate copepod vertical migration and distribution. Marine Ecology Progress Series, 575:57–68, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m575p057.pdf.

Passow:2019:IOD

[PSF^{+}19] U. Passow, J. Sweet, S. Francis, C. Xu, A. L. Dissanayake, Y. Y. Lin, P. H. Santschi, and A. Quigg. Incorporation of oil into diatom aggregates. Marine Ecology Progress Series, 612:65–86, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m612p065.pdf.

Payne:2011:EUI

[PSG11] N. L. Payne, J. M. Semmens, and B. M. Gillanders. Elemental uptake via immersion: a mass-marking technique for the early life-history stages of cephalopods. Marine Ecology Progress Series, 436:169–176, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2011/436/m436p169.pdf.

Pardo:2013:RES

[PSG^{+}13] M. A. Pardo, N. Silverberg, D. Gendron, E. Beier, and D. M. Palacios. Role of environmental seasonality in the turnover of a cetacean community in the southwestern Gulf of California. Marine Ecology Progress Series, 487:245–260, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m487p245.pdf.

Pattrick:2016:PSL

[PSHG16] P. Pattrick, N. A. Strydom, L. Harris, and W. S. Goschen. Predicting spawning locations and modelling the spatial extent of post hatch areas for fishes in a shallow coastal habitat in South Africa. Marine Ecology Progress Series, 560: 223–235, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/560/m560p223.pdf.

Pierrat:2012:LSD

[PSL^{+}12] B. Pierrat, T. Saucède, R. Laffont, C. De Ridder, A. Festeau, and B. David. Large-scale distribution analysis of Antarctic echinoids using ecological niche modelling. Marine Ecology Progress Series, 463:215–230, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/463/m463p215.pdf.

Pontarp:2013:EIH

[PSL13] M. Pontarp, J. Sjöstedt, and P. Lundberg. Experimentally induced habitat filtering in marine bacterial communities. Marine Ecology Progress Series, 477:77–86, 2013. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m477p077.pdf.

Pampoulie:2015:SSA

[PSÓ^{+}15] C. Pampoulie, A. Slotte, G. J. Óskarsson, S. J. Helyar, Á. Jónsson, G. Ólafsdóttir, S. Skírnisdóttir, L. A. Libungan, J. A. Jacobsen, H. Joensen, H. H. Nielsen, S. K. Sigurdsson, and A. K. Daníelsdóttir. Stock structure of Atlantic herring Clupea harengus in the Norwegian Sea and adjacent waters. Marine Ecology Progress Series, 522:219–230, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/522/
m522p219.pdf.

Pereira:2019:RFO

[PSP^{+}19] R. R. C. Pereira, E. Scanes, L. M. Parker, M. Byrne, V. J. Cole, and P. M. Ross. Restoring the flat oyster Ostrea angasi in the face of a changing climate. Marine Ecology Progress Series, 625:27–39, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/625/m625p027.pdf.

Poulson:2010:ACR

[PSPK10] K. L. Poulson, R. D. Sieg, E. K. Prince, and J. Kubanek. Allelopathic compounds of a red tide dinoflagellate have speciesspecific and context-dependent impacts on phytoplankton. Marine Ecology Progress Series, 416:69–78, 2010. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m416p069.pdf.

Paul:2018:BSD

[PSPR18] A. J. Paul, U. Sommer, C. Paul, and U. Riebesell. Baltic Sea diazotrophic cyanobacterium is negatively affected by acidification and warming. Marine Ecology Progress Series, 598: 49–60, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m598p049.pdf.

Persson:2012:SDE

[PSS^{+}12] J. Persson, L. C. Stige, N. C. Stenseth, N. Usov, and D. Martynova. Scale-dependent effects of climate on two copepod species, Calanus glacialis and Pseudocalanus minutus, in an Arctic-boreal sea. Marine Ecology Progress Series, 468: 71–83, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/468/m468p071.pdf.

Poland:2019:FEL

[PSS19] S. J. Poland, F. S. Scharf, and M. D. Staudinger. Foraging ecology of large pelagic fishes in the US South Atlantic: structured piscivory shapes trophic niche variation. Marine Ecology Progress Series, 631:181–199, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/631/m631p181.pdf.

Premo:2013:TPA

[PT13] K. M. Premo and A. C. Tyler. Threat of predation alters the ability of benthic invertebrates to modify sediment biogeochemistry and benthic microalgal abundance. Marine Ecology Progress Series, 494:29–39, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/494/m494p029.pdf.

Purcell:2014:DTP

[PTF^{+}14] J. E. Purcell, U. Tilves, V. L. Fuentes, G. Milisenda, A. Olariaga, and A. Sabatés. Digestion times and predation potentials of Pelagia noctiluca eating fish larvae and copepods in the NW Mediterranean Sea. Marine Ecology Progress Series, 510:201–213, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m510p201.pdf.

Pecuchet:2016:PDF

[PTL16] L. Pecuchet, A. Törnroos, and M. Lindegren. Patterns and drivers of fish community assembly in a large marine ecosystem. Marine Ecology Progress Series, 546:239–248, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/546/m546p239.pdf.

Polito:2015:CSG

[PTP^{+}15] M. J. Polito, W. Z. Trivelpiece, W. P. Patterson, N. J. Karnovsky, C. S. Reiss, and S. D. Emslie. Contrasting specialist and generalist patterns facilitate foraging niche partitioning in sympatric populations of Pygoscelis penguins. Marine Ecology Progress Series, 519:221–237, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m519p221.pdf.

Paxton:2019:CSP

[PTP^{+}19] A. B. Paxton, J. C. Taylor, C. H. Peterson, S. R. Fegley, and J. H. Rosman. Consistent spatial patterns in multiple trophic levels occur around artificial habitats. Marine Ecology Progress Series, 611:189–202, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/611/m611p189.pdf.

Peabody:2018:DRS

[PTS^{+}18] C. E. Peabody, A. R. Thompson, D. F. Sax, R. E. Morse, and C. T. Perretti. Decadal regime shifts in southern California’s ichthyoplankton assemblage. Marine Ecology Progress Series, 607:71–83, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/607/m607p071.pdf.

Prazeres:2016:CLL

[PUP16] M. Prazeres, S. Uthicke, and J. M. Pandolfi. Changing light levels induce photo-oxidative stress and alterations in shell density of Amphistegina lobifera (Foraminifera). Marine Ecology Progress Series, 549:69–78, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/549/
m549p069.pdf.

Purcell:2018:SCJ

[Pur18] J. E. Purcell. Successes and challenges in jellyfish ecology: examples from Aequorea spp. Marine Ecology Progress Series, 591:7–27, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/591/m591p007.pdf.

Pearson:2017:RST

[PvdMLC17] R. M. Pearson, J. P. van de Merwe, C. J. Limpus, and R. M. Connolly. Realignment of sea turtle isotope studies needed to match conservation priorities. Marine Ecology Progress Series, 583:259–271, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/583/m583p259.pdf.

Payne:2015:RDC

[PvdMS^{+}15] N. L. Payne, D. E. van der Meulen, I. M. Suthers, C. A. Gray, C. T. Walsh, and M. D. Taylor. Rain-driven changes in fish dynamics: a switch from spatial to temporal segregation. Marine Ecology Progress Series, 528:267–275, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/528/
m528p267.pdf.

Puneeta:2016:MCB

[PVYS16] P. Puneeta, D. Vijai, J. Yamamoto, and Y. Sakurai. Male copulatory behavior interrupts Japanese flying squid Todarodes pacificus female spawning activity. Marine Ecology Progress Series, 551:277–281, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/551/m551p277.pdf.

Patricio:2011:SPI

[PVZD^{+}11] A. R. Patrìcio, X. Velez-Zuazo, C. E. Diez, R. Van Dam, and A. M. Sabat. Survival probability of immature green turtles in two foraging grounds at Culebra, Puerto Rico. Marine Ecology Progress Series, 440:217–227, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/440/m440p217.pdf.

Pantile:2011:STT

[PW11] R. Pantile and N. Webster. Strict thermal threshold identified by quantitative PCR in the sponge Rhopaloeides odorabile. Marine Ecology Progress Series, 431:97–105, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/431/
m431p097.pdf.

Patrick:2015:IVS

[PW15] C. J. Patrick and D. E. Weller. Interannual variation in submerged aquatic vegetation and its relationship to water quality in subestuaries of Chesapeake Bay. Marine Ecology Progress Series, 537:121–135, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2015/537/m537p121.pdf.

Plumlee:2016:FET

[PW16] J. D. Plumlee and R. J. D. Wells. Feeding ecology of three coastal shark species in the northwest Gulf of Mexico. Marine Ecology Progress Series, 550:163–174, 2016. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m550p163.pdf.

Pillay:2019:GST

[PW19a] D. Pillay and C. Waspe. Grazer specialisation and temperature effects on epiphytic fouling: conservation implications for a temperate African seagrass (Zostera capensis). Marine Ecology Progress Series, 629:235–241, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/629/
m629p235.pdf.

Pruett:2019:EOU

[PW19b] J. L. Pruett and M. J. Weissburg. Eastern oysters use predation risk cues in larval settlement decisions and juvenile inducible morphological defenses. Marine Ecology Progress Series, 621:83–94, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/621/m621p083.pdf.

Powers:2017:MFH

[PWA^{+}17] K. D. Powers, D. N. Wiley, A. J. Allyn, L. J. Welch, and R. A. Ronconi. Movements and foraging habitats of great shearwaters Puffinus gravis in the Gulf of Maine. Marine Ecology Progress Series, 574:211–226, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/574/m574p211.pdf.

Poupart:2019:FEW

[PWB^{+}19] T. A. Poupart, S. M. Waugh, C. A. Bost, A. Kato, C. M. Miskelly, K. M. Rogers, and J. P. Y. Arnould. Foraging ecology of a winter breeder, the fiordland penguin. Marine Ecology Progress Series, 614:183–197, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/614/m614p183.pdf.

Pollock:2014:AMV

[PWCvO^{+}14] F. J. Pollock, E. M. Wood-Charlson, M. J. H. van Oppen, D. G. Bourne, B. L. Willis, and K. D. Weynberg. Abundance and morphology of virus-like particles associated with the coral Acropora hyacinthus differ between healthy and white syndrome-infected states. Marine Ecology Progress Series, 510:39–43, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/510/m510p039.pdf.

Pausch:2018:IFG

[PWM18] R. E. Pausch, D. E. Williams, and M. W. Miller. Impacts of fragment genotype, habitat, and size on outplanted elkhorn coral success under thermal stress. Marine Ecology Progress Series, 592:109–117, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/592/m592p109.pdf.

Poupart:2019:FSF

[PWM^{+}19] T. A. Poupart, S. M. Waugh, C. M. Miskelly, A. Kato, L. P. Angel, K. M. Rogers, and J. P. Y. Arnould. Fine-scale foraging behaviour of southern Buller’s albatross, the only Thalassarche that provisions chicks through winter. Marine Ecology Progress Series, 625:163–179, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/625/m625p163.pdf.

Pettitt-Wade:2011:DMM

[PWNP^{+}11] H. Pettitt-Wade, S. P. Newman, K. T. Parsons, S. H. Gruber, and R. D. Handy. Dietary metal and macro-nutrient intakes of juvenile lemon sharks determined from the nutritional composition of prey items. Marine Ecology Progress Series, 433:245–260, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2011/433/m433p245.pdf.

Pouil:2016:IFA

[PWO^{+}16] S. Pouil, M. Warnau, F. Oberhänsli, J. L. Teyssié, P. Bustamante, and M. Metian. Influence of food on the assimilation of essential elements (Co, Mn, and Zn) by turbot Scophthalmus maximus. Marine Ecology Progress Series, 550: 207–218, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/550/m550p207.pdf.

Pillay:2012:IEB

[PWW12] D. Pillay, C. Williams, and A. K. Whitfield. Indirect effects of bioturbation by the burrowing sandprawn Callichirus kraussi on a benthic foraging fish, Liza richardsonii. Marine Ecology Progress Series, 453:151–158, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/453/m453p151.pdf.

Paiva:2010:FEC

[PXG^{+}10] V. H. Paiva, J. Xavier, P. Geraldes, I. Ramirez, S. Garthe, and J. A. Ramos. Foraging ecology of Cory’s shearwaters in different oceanic environments of the North Atlantic. Marine Ecology Progress Series, 410:257–268, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/410/
m410p257.pdf.

Poulton:2011:BDE

[PYBB11] A. J. Poulton, J. R. Young, N. R. Bates, and W. M. Balch. Biometry of detached Emiliania huxleyi coccoliths along the Patagonian shelf. Marine Ecology Progress Series, 443:1–17, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Prairie:2013:DSM

[PZA^{+}13] J. C. Prairie, K. Ziervogel, C. Arnosti, R. Camassa, C. Falcon, S. Khatri, R. M. McLaughlin, B. L. White, and S. Yu. Delayed settling of marine snow at sharp density transitions driven by fluid entrainment and diffusion-limited retention. Marine Ecology Progress Series, 487:185–200, 2013. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m487p185.pdf.

Peng:2017:OAA

[PZL^{+}17] C. Peng, X. Zhao, S. Liu, W. Shi, Y. Han, C. Guo, X. Peng, X. Chai, and G. Liu. Ocean acidification alters the burrowing behaviour, Ca^{2+} /Mg^{2+}-ATPase activity, metabolism, and gene expression of a bivalve species, Sinonovacula constricta. Marine Ecology Progress Series, 575:107–117, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/575/
m575p107.pdf.

Quigley:2018:OPL

[QBW18] K. M. Quigley, L. K. Bay, and B. L. Willis. Opinion piece: Leveraging new knowledge of Symbiodinium community regulation in corals for conservation and reef restoration. Marine Ecology Progress Series, 600:245–253, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/600/
m600p245.pdf.

Qin:2010:EIC

[QCX^{+}10] H. Qin, T. Chu, W. Xu, G. Lei, Z. Chen, W. Quan, J. Chen, and J. Wu. Effects of invasive cordgrass on crab distributions and diets in a Chinese salt marsh. Marine Ecology Progress Series, 415:177–187, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/415/m415p177.pdf.

Quiroga:2014:NBS

[QGM^{+}14] E. Quiroga, D. Gerdes, A. Montiel, R. Knust, and U. Jacob. Normalized biomass size spectra in high Antarctic macrobenthic communities: linking trophic position and body size. Marine Ecology Progress Series, 506:99–113, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/506/
m506p099.pdf.

Quaas:2019:IHC

[QHG^{+}19] Z. Quaas, D. Harasti, T. F. Gaston, M. E. Platell, and C. J. Fulton. Influence of habitat condition on shallow rocky reef fish community structure around islands and headlands of a temperate marine protected area. Marine Ecology Progress Series, 626:1–13, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Queiroz:2010:STM

[QHN^{+}10] N. Queiroz, N. E. Humphries, L. R. Noble, A. M. Santos, and D. W. Sims. Short-term movements and diving behaviour of satellite-tracked blue sharks Prionace glauca in the northeastern Atlantic Ocean. Marine Ecology Progress Series, 406:265–279, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/406/m406p265.pdf.

Quintanilla:2015:ITP

[QLCUG15] J. M. Quintanilla, R. Laiz-Carrión, A. Uriarte, and A. García. Influence of trophic pathways on daily growth patterns of western Mediterranean anchovy Engraulis encrasicolus larvae. Marine Ecology Progress Series, 531:263–275, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/531/
m531p263.pdf.

Quillien:2015:EMA

[QNG^{+}15] N. Quillien, M. C. Nordström, O. Gauthier, E. Bonsdorff, Y. M. Paulet, and J. Grall. Effects of macroalgal accumulations on the variability in zoobenthos of high-energy macrotidal sandy beaches. Marine Ecology Progress Series, 522: 97–114, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/522/m522p097.pdf.

Quiroga:2016:SBP

[QOGS^{+}16] E. Quiroga, P. Ortiz, R. González-Saldías, B. Reid, F. J. Tapia, I. Pérez-Santos, L. Rebolledo, R. Mansilla, C. Pineda, I. Cari, N. Salinas, A. Montiel, and D. Gerdes. Seasonal benthic patterns in a glacial Patagonian fjord: the role of suspended sediment and terrestrial organic matter. Marine Ecology Progress Series, 561:31–50, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/561/m561p031.pdf.

Quesada:2016:ESS

[QSB16] A. J. Quesada, K. L. Schoo, and B. L. Bingham. Effect of symbiotic state on the fatty acid composition of Anthopleura elegantissima. Marine Ecology Progress Series, 545: 175–187, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/545/m545p175.pdf.

Quillfeldt:2011:FFB

[QSvN^{+}11] P. Quillfeldt, S. Schroff, H. J. van Noordwijk, A. Michalik, K. Ludynia, and J. F. Masello. Flexible foraging behaviour of a sexually dimorphic seabird: large males do not always dive deep. Marine Ecology Progress Series, 428:271– 287, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/428/m428p271.pdf.

Qiu:2012:ARH

[QYS^{+}12] X. Qiu, Y. Yamasaki, Y. Shimasaki, H. Gunjikake, M. Honda, M. Kawaguchi, T. Matsubara, S. Nagasoe, T. Etoh, S. Matsui, T. Honjo, and Y. Oshima. Allelopathy of the raphidophyte Heterosigma akashiwo against the dinoflagellate Akashiwo sanguinea is mediated via allelochemicals and cell contact. Marine Ecology Progress Series, 446:107–118, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/446/
m446p107.pdf.

Roberts:2010:PPS

[RA10] D. G. Roberts and D. J. Ayre. Panmictic population structure in the migratory marine sparid Acanthopagrus australis despite its close association with estuaries. Marine Ecology Progress Series, 412:223–230, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/412/m412p223.pdf.

Rattray:2016:GDW

[RAA^{+}16] A. Rattray, M. Andrello, V. Asnaghi, S. Bevilacqua, F. Bulleri, E. Cebrian, M. Chiantore, J. Claudet, S. Deudero, J. Evans, S. Fraschetti, G. Guarnieri, L. Mangialajo, P. J. Schembri, A. Terlizzi, and L. Benedetti-Cecchi. Geographic distance, water circulation and environmental conditions shape the biodiversity of Mediterranean rocky coasts. Marine Ecology Progress Series, 553:1–11, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Rees:2010:BPO

[RAB^{+}10] A. F. Rees, S. Al Saady, A. C. Broderick, M. S. Coyne, N. Papathanasopoulou, and B. J. Godley. Behavioural polymorphism in one of the world’s largest populations of loggerhead sea turtles Caretta caretta. Marine Ecology Progress Series, 418:201–212, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/418/m418p201.pdf.

Rovelli:2015:BUT

[RAB^{+}15] L. Rovelli, K. M. Attard, L. D. Bryant, S. Flögel, H. Stahl, J. M. Roberts, P. Linke, and R. N. Glud. Benthic O 2 uptake of two cold-water coral communities estimated with the noninvasive eddy correlation technique. Marine Ecology Progress Series, 525:97–104, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m525p097.pdf.

Reid-Anderson:2019:EPS

[RABS19] S. Reid-Anderson, K. Bilgmann, and A. Stow. Effective population size of the critically endangered east Australian grey nurse shark Carcharias taurus. Marine Ecology Progress Series, 610:137–148, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/610/m610p137.pdf.

Rooker:2014:CLM

[RAF^{+}14] J. R. Rooker, H. Arrizabalaga, I. Fraile, D. H. Secor, D. L. Dettman, N. Abid, P. Addis, S. Deguara, F. S. Karakulak, A. Kimoto, O. Sakai, D. Macías, and M. N. Santos. Crossing the line: migratory and homing behaviors of Atlantic bluefin tuna. Marine Ecology Progress Series, 504:265–276, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m504p265.pdf.

Ragni:2010:PPP

[RAH^{+}10] M. Ragni, R. L. Airs, S. J. Hennige, D. J. Suggett, M. E. Warner, and R. J. Geider. PSII photoinhibition and photorepair in Symbiodinium (Pyrrhophyta) differs between thermally tolerant and sensitive phylotypes. Marine Ecology Progress Series, 406:57–70, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/406/m406p057.pdf.

Raimondo:2012:ITD

[Rai12] S. Raimondo. Incorporating temperature-driven seasonal variation in survival, growth, and reproduction into population models for small fish. Marine Ecology Progress Series, 469: 101–112, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/469/m469p101.pdf.

Rees:2012:CRI

[RAKB^{+}12] A. F. Rees, A. Al-Kiyumi, A. C. Broderick, N. Papathanasopoulou, and B. J. Godley. Conservation related insights into the behaviour of the olive ridley sea turtle Lepidochelys olivacea nesting in Oman. Marine Ecology Progress Series, 450:195–205, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/450/m450p195.pdf.

Righton:2010:TNA

[RAN^{+}10] D. A. Righton, K. H. Andersen, F. Neat, V. Thorsteinsson, P. Steingrund, H. Svedäng, K. Michalsen, H. H. Hinrichsen, V. Bendall, S. Neuenfeldt, P. Wright, P. Jonsson, G. Huse, J. van der Kooij, H. Mosegaard, K. Hüssy, and J. Metcalfe. Thermal niche of Atlantic cod Gadus morhua: limits, tolerance and optima. Marine Ecology Progress Series, 420:1–13, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Rand:2018:PHR

[Ran18] P. S. Rand. Pacific herring response to surface predators in Prince William Sound, Alaska, USA. Marine Ecology Progress Series, 600:239–244, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/600/m600p239.pdf.

Rassoulzadegan:2010:MPG

[Ras10] Fereidoun Rassoulzadegan. In memoriam: Peter Gordon Verity (1953–2009). Marine Ecology Progress Series, 399:i, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m399p000.pdf.

Randhawa:2011:LPG

[RB11] H. S. Randhawa and P. Brickle. Larval parasite gene sequence data reveal cryptic trophic links in life cycles of porbeagle shark tapeworms. Marine Ecology Progress Series, 431:215–222, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/431/m431p215.pdf.

Rickards:2015:ETH

[RB15] K. J. C. Rickards and E. G. Boulding. Effects of temperature and humidity on activity and microhabitat selection by Littorina subrotundata. Marine Ecology Progress Series, 537:163–173, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/537/m537p163.pdf.

Rodemann:2017:CPC

[RB17] J. R. Rodemann and S. J. Brandl. Consumption pressure in coastal marine environments decreases with latitude and in artificial vs. natural habitats. Marine Ecology Progress Series, 574:167–179, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/574/m574p167.pdf.

Rebstock:2018:OCW

[RB18] G. A. Rebstock and P. D. Boersma. Oceanographic conditions in wintering grounds affect arrival date and body condition in breeding female Magellanic penguins. Marine Ecology Progress Series, 601:253–267, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/601/m601p253.pdf.

Richardson:2019:VDP

[RB19] K. Richardson and J. Bendtsen. Vertical distribution of phytoplankton and primary production in relation to nutricline depth in the open ocean. Marine Ecology Progress Series, 620:33–46, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m620p033.pdf.

Reglero:2018:VDA

[RBA^{+}18] P. Reglero, E. Blanco, F. Alemany, C. Ferrá, D. Alvarez-Berastegui, A. Ortega, F. de la Gándara, A. Aparicio-González, and A. Folkvord. Vertical distribution of Atlantic bluefin tuna Thunnus thynnus and bonito Sarda sarda larvae is related to temperature preference. Marine Ecology Progress Series, 594:231–243, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/594/m594p231.pdf.

Rossi:2012:TIL

[RBBG12] S. Rossi, L. Bramanti, E. Broglio, and J. M. Gili. Trophic impact of long-lived species indicated by population dynamics in the short-lived hydrozoan Eudendrium racemosum. Marine Ecology Progress Series, 467:97–111, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/467/
m467p097.pdf.

Richardson:2014:LMH

[RBC^{+}14] K. Richardson, J. Bendtsen, J. Tang Christensen, M. Adjou, M. Moltke Lyngsgaard, K. M. Hilligsøe, J. B. Pedersen, T. Vang, and M. Holtegaard Nielsen. Localised mixing and heterogeneity in the plankton food web in a frontal region of the Sargasso Sea: implications for eel early life history? Marine Ecology Progress Series, 504:91–107, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m504p091.pdf.

Rix:2015:SDF

[RBC^{+}15] L. Rix, V. N. Bednarz, U. Cardini, N. van Hoytema, F. A. Al-Horani, C. Wild, and M. S. Naumann. Seasonality in dinitrogen fixation and primary productivity by coral reef framework substrates from the northern Red Sea. Marine Ecology Progress Series, 533:79–92, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/533/m533p079.pdf.

Rudershausen:2014:EMS

[RBD^{+}14] P. J. Rudershausen, J. A. Buckel, T. Dubreuil, M. J. O’Donnell, J. E. Hightower, S. J. Poland, and B. H. Letcher. Estimating movement and survival rates of a small saltwater fish using autonomous antenna receiver arrays and passive integrated transponder tags. Marine Ecology Progress Series, 499:177–192, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/499/m499p177.pdf.

Rudershausen:2016:CFI

[RBD^{+}16] P. J. Rudershausen, J. A. Buckel, M. A. Dueker, S. J. Poland, and E. Hain. Comparison of fish and invertebrate assemblages among variably altered tidal creeks in a coastal landscape. Marine Ecology Progress Series, 544:15–35, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/544/
m544p015.pdf.

Riquelme-Bugueno:2013:SMP

[RBEGG13] R. Riquelme-Bugueño, R. Escribano, and J. Gómez-Gutiérrez. Somatic and molt production in Euphausia mucronata off central-southern Chile: the influence of coastal upwelling variability. Marine Ecology Progress Series, 476:39–57, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/476/m476p039.pdf.

Rius:2010:LSB

[RBGT10] M. Rius, G. M. Branch, C. L. Griffiths, and X. Turon. Larval settlement behaviour in six gregarious ascidians in relation to adult distribution. Marine Ecology Progress Series, 418:151–163, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/418/m418p151.pdf.

Ryer:2012:GDC

[RBH12] C. H. Ryer, K. S. Boersma, and T. P. Hurst. Growth and distributional correlates of behavior in three co-occurring juvenile flatfishes. Marine Ecology Progress Series, 460:183– 193, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/460/m460p183.pdf.

Roberts:2019:CCO

[RBHR19] S. M. Roberts, A. M. Boustany, P. N. Halpin, and R. R. Rykaczewski. Cyclical climate oscillation alters species statistical relationships with local habitat. Marine Ecology Progress Series, 614:159–171, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m614p159.pdf.

Raymundo:2017:ATE

[RBL^{+}17] L. J. Raymundo, D. Burdick, V. A. Lapacek, R. Miller, and V. Brown. Anomalous temperatures and extreme tides: Guam staghorn Acropora succumb to a double threat. Marine Ecology Progress Series, 564:47–55, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/564/
m564p047.pdf.

Rheuban:2014:MTP

[RBM14] J. E. Rheuban, P. Berg, and K. J. McGlathery. Multiple timescale processes drive ecosystem metabolism in eelgrass (Zostera marina) meadows. Marine Ecology Progress Series, 507:1–13, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ramp:2010:SSS

[RBP^{+}10] C. Ramp, M. Bérubé, P. Palsbøll, W. Hagen, and R. Sears. Sex-specific survival in the humpback whale Megaptera novaeangliae in the Gulf of St. Lawrence, Canada. Marine Ecology Progress Series, 400:267–276, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/400/
m400p267.pdf.

Rogers:2016:NPL

[RBS16] T. L. Rogers, J. E. Byrnes, and J. J. Stachowicz. Native predators limit invasion of benthic invertebrate communities in Bodega Harbor, California, USA. Marine Ecology Progress Series, 545:161–173, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/545/m545p161.pdf.

Rouhani:2017:NEV

[RBS^{+}17] S. Rouhani, M. C. Baker, M. Steinhoff, M. Zhang, J. Oehrig, I. J. Zelo, S. D. Emsbo-Mattingly, Z. Nixon, J. M. Willis, and M. W. Hester. Nearshore exposure to Deepwater Horizon oil. Marine Ecology Progress Series, 576:111–124, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m576p111.pdf.

Reece:2011:CPF

[RBSL11] J. S. Reece, B. W. Bowen, D. G. Smith, and A. Larson. Comparative phylogeography of four Indo–Pacific moray eel species (Muraenidae) reveals comparable ocean-wide genetic connectivity despite five-fold differences in available adult habitat. Marine Ecology Progress Series, 437:269–277, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/437/
m437p269.pdf.

Raven:2012:ECC

[RC12] J. A. Raven and K. Crawfurd. Environmental controls on coccolithophore calcification. Marine Ecology Progress Series, 470:137–166, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m470p137.pdf.

Reglero:2012:GED

[RCAB^{+}12] P. Reglero, L. Ciannelli, D. Alvarez-Berastegui, R. Balbín, J. L. López-Jurado, and F. Alemany. Geographically and environmentally driven spawning distributions of tuna species in the western Mediterranean Sea. Marine Ecology Progress Series, 463:273–284, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/463/m463p273.pdf.

Riou:2010:MDS

[RCB^{+}10] V. Riou, A. Colaço, S. Bouillon, A. Khripounoff, P. Dando, P. Mangion, E. Chevalier, M. Korntheuer, R. S. Santos, and F. Dehairs. Mixotrophy in the deep sea: a dual endosymbiotic hydrothermal mytilid assimilates dissolved and particulate organic matter. Marine Ecology Progress Series, 405: 187–201, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/405/m405p187.pdf.

Reynolds:2012:NNA

[RCB12] L. K. Reynolds, L. A. Carr, and K. E. Boyer. A non-native amphipod consumes eelgrass inflorescences in San Francisco Bay. Marine Ecology Progress Series, 451:107–118, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/451/
m451p107.pdf.

Roberts:2015:SMM

[RCBE15] L. Roberts, S. Cheesman, T. Breithaupt, and M. Elliott. Sensitivity of the mussel Mytilus edulis to substrate-borne vibration in relation to anthropogenically generated noise. Marine Ecology Progress Series, 538:185–195, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/538/
m538p185.pdf.

Risch:2013:MWA

[RCD^{+}13] D. Risch, C. W. Clark, P. J. Dugan, M. Popescu, U. Siebert, and S. M. Van Parijs. Minke whale acoustic behavior and multi-year seasonal and diel vocalization patterns in Massachusetts Bay, USA. Marine Ecology Progress Series, 489: 279–295, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m489p279.pdf.

Reitzel:2013:PDR

[RCE^{+}13] A. M. Reitzel, T. Chu, S. Edquist, C. Genovese, C. Church, A. M. Tarrant, and J. R. Finnerty. Physiological and developmental responses to temperature by the sea anemone Nematostella vectensis. Marine Ecology Progress Series, 484: 115–130, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m484p115.pdf.

Riekenberg:2016:TPM

[RCF16] P. M. Riekenberg, R. S. Carney, and B. Fry. Trophic plasticity of the methanotrophic mussel Bathymodiolus childressi in the Gulf of Mexico. Marine Ecology Progress Series, 547: 91–106, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m547p091.pdf.

Rocha:2017:PTT

[RCF17] R. M. Rocha, G. C. Castellano, and C. A. Freire. Physiological tolerance as a tool to support invasion risk assessment of tropical ascidians. Marine Ecology Progress Series, 577:105–119, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/577/m577p105.pdf.

Rodriguez:2015:CSL

[RCG^{+}15] J. M. Rodriguez, A. Cabrero, J. Gago, A. Garcia, R. Laiz-Carrion, C. Piñeiro, and F. Saborido-Rey. Composition and structure of the larval fish community in the NW Iberian upwelling system during the winter mixing period. Marine Ecology Progress Series, 533:245–260, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m533p245.pdf.

Ribeiro:2012:FAS

[RCGE12] J. Ribeiro, G. M. Carvalho, J. M. S. Gonçalves, and K. Erzini. Fish assemblages of shallow intertidal habitats of the Ria Formosa lagoon (South Portugal): influence of habitat and season. Marine Ecology Progress Series, 446:259–273, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2011/446/m446p259.pdf.

Reiss:2011:SDM

[RCK^{+}11] H. Reiss, S. Cunze, K. König, H. Neumann, and I. Kröncke. Species distribution modelling of marine benthos: a North Sea case study. Marine Ecology Progress Series, 442:71–86, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m442p071.pdf.

Reeves:2018:BSG

[RCMB18] D. B. Reeves, E. J. Chesney, R. T. Munnelly, and D. M. Baltz. Barnacle settlement and growth at oil and gas platforms in the northern Gulf of Mexico. Marine Ecology Progress Series, 590:131–143, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/590/m590p131.pdf.

Ruzicka:2013:TCB

[RCP^{+}13] R. R. Ruzicka, M. A. Colella, J. W. Porter, J. M. Morrison, J. A. Kidney, V. Brinkhuis, K. S. Lunz, K. A. Macaulay, L. A. Bartlett, M. K. Meyers, and J. Colee. Temporal changes in benthic assemblages on Florida Keys reefs 11 years after the 1997/1998 El Niño. Marine Ecology Progress Series, 489:125–141, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m489p125.pdf.

Rayner:2016:NPT

[RCP^{+}16] M. J. Rayner, N. Carlile, D. Priddel, V. Bretagnolle, M. G. R. Miller, R. A. Phillips, L. Ranjard, S. J. Bury, and L. G. Torres. Niche partitioning by three Pterodroma petrel species during non-breeding in the equatorial Pacific Ocean. Marine Ecology Progress Series, 549:217–229, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/549/m549p217.pdf.

Rabouille:2017:TCB

[RCP17] S. Rabouille, G. Semedo Cabral, and M. L. Pedrotti. Towards a carbon budget of the diazotrophic cyanobacterium Crocosphaera: effect of irradiance. Marine Ecology Progress Series, 570:29–40, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m570p029.pdf.

Rohner:2013:DWS

[RCR^{+}13] C. A. Rohner, L. I. E. Couturier, A. J. Richardson, S. J. Pierce, C. E. M. Prebble, M. J. Gibbons, and P. D. Nichols. Diet of whale sharks Rhincodon typus inferred from stomach content and signature fatty acid analyses. Marine Ecology Progress Series, 493:219–235, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m493p219.pdf.

Reiss:2017:OHS

[RCS^{+}17] C. S. Reiss, A. Cossio, J. A. Santora, K. S. Dietrich, A. Murray, B. G. Mitchell, J. Walsh, E. L. Weiss, C. Gimpel, C. D. Jones, and G. M. Watters. Overwinter habitat selection by Antarctic krill under varying sea-ice conditions: implications for top predators and fishery management. Marine Ecology Progress Series, 568:1–16, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Rotherham:2011:UST

[RCU^{+}11] D. Rotherham, M. G. Chapman, A. J. Underwood, C. A. Gray, and D. D. Johnson. Untangling spatial and temporal variation in abundances of estuarine fish sampled with multi-mesh gillnets. Marine Ecology Progress Series, 435: 183–195, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/435/m435p183.pdf.

Ruiz-Cooley:2010:OVR

[RCVG10] R. I. Ruiz-Cooley, E. C. Villa, and W. R. Gould. Ontogenetic variation of δ ^{13}C and δ ^{15}N recorded in the gladius of the jumbo squid Dosidicus gigas: geographic differences. Marine Ecology Progress Series, 399:187–198, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/399/
m399p187.pdf.

Ritt:2012:SCT

[RDC^{+}12] B. Ritt, D. Desbruyères, J. C. Caprais, O. Gauthier, L. Ruffine, R. Buscail, K. Olu-Le Roy, and J. Sarrazin. Seep communities from two mud volcanoes in the deep eastern Mediterranean Sea: faunal composition, spatial patterns and environmental control. Marine Ecology Progress Series, 466: 93–119, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/466/m466p093.pdf.

Robert:2013:IIB

[RDF^{+}13] M. Robert, L. Dagorn, J. D. Filmalter, J. L. Deneubourg, D. Itano, and K. Holland. Intra-individual behavioral variability displayed by tuna at fish aggregating devices (FADs). Marine Ecology Progress Series, 484:239–247, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/484/
m484p239.pdf.

Rice:2017:STO

[RDF^{+}17] A. Rice, V. B. Deecke, J. K. B. Ford, J. F. Pilkington, E. M. Oleson, J. A. Hildebrand, and A. Širovi ć. Spatial and temporal occurrence of killer whale ecotypes off the outer coast of Washington State, USA. Marine Ecology Progress Series, 572:255–268, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m572p255.pdf.

Rovellini:2019:DVS

[RDF^{+}19] A. Rovellini, M. R. Dunn, E. A. Fulton, N. S. Webster, D. J. Smith, J. Jompa, A. Haris, J. Berman, and J. J. Bell. Decadal variability in sponge abundance and biodiversity on an Indo–Pacific coral reef. Marine Ecology Progress Series, 620:63–76, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/620/m620p063.pdf.

Rita:2017:TCI

[RDGC17] D. Rita, M. Drago, F. Galimberti, and L. Cardona. Temporal consistency of individual trophic specialization in southern elephant seals Mirounga leonina. Marine Ecology Progress Series, 585:229–242, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/585/m585p229.pdf.

Rix:2018:RSF

[RdGv^{+}18] L. Rix, J. M. de Goeij, D. van Oevelen, U. Struck, F. A. Al-Horani, C. Wild, and M. S. Naumann. Reef sponges facilitate the transfer of coral-derived organic matter to their associated fauna via the sponge loop. Marine Ecology Progress Series, 589:85–96, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m589p085.pdf.

Regular:2010:CFP

[RDHM10] P. M. Regular, G. K. Davoren, A. Hedd, and W. A. Montevecchi. Crepuscular foraging by a pursuit-diving seabird: tactics of common murres in response to the diel vertical migration of capelin. Marine Ecology Progress Series, 415: 295–304, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/415/m415p295.pdf.

Ressler:2014:SDE

[RDK14] P. H. Ressler, A. De Robertis, and S. Kotwicki. The spatial distribution of euphausiids and walleye pollock in the eastern Bering Sea does not imply top-down control by predation. Marine Ecology Progress Series, 503:111–122, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/503/
m503p111.pdf.

Richerson:2018:ITM

[RDM18] K. Richerson, R. Driscoll, and M. Mangel. Increasing temperature may shift availability of euphausiid prey in the Southern Ocean. Marine Ecology Progress Series, 588:59– 70, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/588/m588p059.pdf.

Ricart:2015:ELC

[RDPR15] A. M. Ricart, A. Dalmau, M. Pérez, and J. Romero. Effects of landscape configuration on the exchange of materials in seagrass ecosystems. Marine Ecology Progress Series, 532:89–100, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/532/m532p089.pdf.

Reiertsen:2014:PDN

[REAN^{+}14] T. K. Reiertsen, K. E. Erikstad, T. Anker-Nilssen, R. T. Barrett, T. Boulinier, M. Frederiksen, J. González-Solís, D. Gremillet, D. Johns, B. Moe, A. Ponchon, M. Skern-Mauritzen, H. Sandvik, and N. G. Yoccoz. Prey density in non-breeding areas affects adult survival of black-legged kittiwakes Rissa tridactyla. Marine Ecology Progress Series, 509:289–302, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m509p289.pdf.

Ricci:2016:TSP

[REBL16] S. W. Ricci, D. B. Eggleston, D. R. Bohnenstiehl, and A. Lillis. Temporal soundscape patterns and processes in an estuarine reserve. Marine Ecology Progress Series, 550:25– 38, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/550/m550p025.pdf.

Rees:2014:STK

[Ree14] T. A. V. Rees. Scaling and transport kinetics in aquatic primary producers. Marine Ecology Progress Series, 509:103– 112, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/509/m509p103.pdf.

Reum:2011:MIC

[REG^{+}11] J. C. P. Reum, T. E. Essington, C. M. Greene, C. A. Rice, and K. L. Fresh. Multiscale influence of climate on estuarine populations of forage fish: the role of coastal upwelling, freshwater flow and temperature. Marine Ecology Progress Series, 425:203–215, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/425/m425p203.pdf.

Reinfelder:2012:CDR

[Rei12] J. R. Reinfelder. Carbon dioxide regulation of nitrogen and phosphorus in four species of marine phytoplankton. Marine Ecology Progress Series, 466:57–67, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/466/
m466p057.pdf.

Rossi:2018:TET

[REP18] S. Rossi and F. Elias-Piera. Trophic ecology of three echinoderms in deep waters of the Weddell Sea (Antarctica). Marine Ecology Progress Series, 596:143–153, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/596/
m596p143.pdf.

Rangel:2019:QTP

[RET^{+}19] M. S. Rangel, D. Erler, A. Tagliafico, K. Cowden, S. Scheffers, and L. Christidis. Quantifying the transfer of prey δ ^{15}N signatures into coral holobiont nitrogen pools. Marine Ecology Progress Series, 610:33–49, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/610/m610p033.pdf.

Renz:2014:EBT

[RF14] J. R. Renz and S. Forster. Effects of bioirrigation by the three sibling species of Marenzelleria spp. on solute fluxes and porewater nutrient profiles. Marine Ecology Progress Series, 505:145–159, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/505/m505p145.pdf.

Reaugh-Flower:2011:SDP

[RFBH^{+}11] K. Reaugh-Flower, G. M. Branch, J. M. Harris, C. D. Mc-Quaid, B. Currie, A. Dye, and B. Robertson. Scale-dependent patterns and processes of intertidal mussel recruitment around southern Africa. Marine Ecology Progress Series, 434:101– 119, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/434/m434p101.pdf.

Ruiz-Frau:2011:MSV

[RFEJK11] A. Ruiz-Frau, G. Edwards-Jones, and M. J. Kaiser. Mapping stakeholder values for coastal zone management. Marine Ecology Progress Series, 434:239–249, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m434p239.pdf.

Ransom:2016:CLS

[RFH16] J. T. Ransom, J. E. Filbrun, and F. J. Hernandez, Jr. Condition of larval Spanish mackerel Scomberomorus maculatus in relation to the Deepwater Horizon oil spill. Marine Ecology Progress Series, 558:143–152, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/558/m558p143.pdf.

Roder:2010:TRC

[RFJ^{+}10] C. Roder, L. Fillinger, C. Jantzen, G. M. Schmidt, S. Khokiattiwong, and C. Richter. Trophic response of corals to large amplitude internal waves. Marine Ecology Progress Series, 412:113–128, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/412/m412p113.pdf.

Reum:2015:ECI

[RFM^{+}15] J. C. P. Reum, B. E. Ferriss, P. S. McDonald, D. M. Farrell, C. J. Harvey, T. Klinger, and P. S. Levin. Evaluating community impacts of ocean acidification using qualitative network models. Marine Ecology Progress Series, 536: 11–24, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/536/m536p011.pdf.

Richards:2012:CCN

[RFMT12] R. A. Richards, M. J. Fogarty, D. G. Mountain, and M. H. Taylor. Climate change and northern shrimp recruitment variability in the Gulf of Maine. Marine Ecology Progress Series, 464:167–178, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/464/m464p167.pdf.

Ruppert:2010:EMA

[RFRD10] J. L. W. Ruppert, M. J. Fortin, G. A. Rose, and R. Devillers. Environmental mediation of Atlantic cod on fish community composition: an application of multivariate regression tree analysis to exploited marine ecosystems. Marine Ecology Progress Series, 411:189–201, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/411/m411p189.pdf.

Reiss:2019:EFS

[RFRE19] K. Reiss, N. N. Fieten, P. L. Reynolds, and B. K. Eriksson. Ecosystem function of subarctic Zostera marina meadows: influence of shoot density on fish predators and predation rates. Marine Ecology Progress Series, 630:41–54, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/630/
m630p041.pdf.

Ruttenberg:2011:BMT

[RG11] B. I. Ruttenberg and E. F. Granek. Bridging the marineterrestrial disconnect to improve marine coastal zone science and management. Marine Ecology Progress Series, 434:203– 212, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m434p203.pdf.

Robinson:2014:WSC

[RG14] K. L. Robinson and W. M. Graham. Warming of subtropical coastal waters accelerates Mnemiopsis leidyi growth and alters timing of spring ctenophore blooms. Marine Ecology Progress Series, 502:105–115, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2014/502/m502p105.pdf.

Rosa:2010:BSP

[RGB^{+}10] R. Rosa, L. Gonzalez, B. R. Broitman, S. Garrido, A. M. P. Santos, and M. L. Nunes. Bioenergetics of small pelagic fishes in upwelling systems: relationship between fish condition, coastal ecosystem dynamics and fisheries. Marine Ecology Progress Series, 410:205–218, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/410/m410p205.pdf.

Ross:2011:MEVa

[RGB^{+}11] O. N. Ross, R. J. Geider, E. Berdalet, M. L. Artigas, and J. Piera. Modelling the effect of vertical mixing on bottle incubations for determining in situ phytoplankton dynamics. I. Growth rates. Marine Ecology Progress Series, 435: 13–31, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m435p013.pdf.

Rodriguez-Grana:2010:GSA

[RGCT^{+}10] L. Rodríguez-Graña, D. Calliari, P. Tiselius, B. W. Hansen, and H. N. Sköld. Gender-specific ageing and non-Mendelian inheritance of oxidative damage in marine copepods. Marine Ecology Progress Series, 401:1–13, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Riou:2011:RIA

[RGdLB^{+}11] S. Riou, C. M. Gray, M. d. L. Brooke, P. Quillfeldt, J. F. Masello, C. Perrins, and K. C. Hamer. Recent impacts of anthropogenic climate change on a higher marine predator in western Britain. Marine Ecology Progress Series, 422: 105–112, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/422/m422p105.pdf.

Roscales:2011:TVG

[RGDNGS11] J. L. Roscales, E. Gómez-Díaz, V. Neves, and J. González-Solís. Trophic versus geographic structure in stable isotope signatures of pelagic seabirds breeding in the northeast Atlantic. Marine Ecology Progress Series, 434:1–13, 2011. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Rosa:2012:EDL

[RGDS12] R. Rosa, L. Gonzalez, H. M. Dierssen, and B. A. Seibel. Environmental determinants of latitudinal size-trends in cephalopods. Marine Ecology Progress Series, 464:153–165, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2012/464/m464p153.pdf.

Ramirez:2015:HIG

[RGEA^{+}15] F. Ramírez, C. Gutiérrez-Expósito, I. Afán, J. Giménez, R. de Stephanis, and M. G. Forero. Human influence on gull non-breeding distribution: potential consequences of changes in fishing practices. Marine Ecology Progress Series, 527: 221–232, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/527/m527p221.pdf.

Reisinger:2016:VDK

[RGL^{+}16] R. R. Reisinger, D. R. Gröcke, N. Lübcker, E. L. McClymont, A. R. Hoelzel, and P. J. N. de Bruyn. Variation in the diet of killer whales Orcinus orca at Marion Island, Southern Ocean. Marine Ecology Progress Series, 549:263–274, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/549/
m549p263.pdf.

Ross:2011:MEVb

[RGP11] O. N. Ross, R. J. Geider, and J. Piera. Modelling the effect of vertical mixing on bottle incubations for determining in situ phytoplankton dynamics. II. Primary production. Marine Ecology Progress Series, 435:33–45, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m435p033.pdf.

Regnier:2017:ITM

[RGW17] T. Régnier, F. M. Gibb, and P. J. Wright. Importance of trophic mismatch in a winter-hatching species: evidence from lesser sandeel. Marine Ecology Progress Series, 567: 185–197, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/567/m567p185.pdf.

Roberts:2011:TSH

[RGWA11] D. G. Roberts, C. A. Gray, R. J. West, and D. J. Ayre. Temporal stability of a hybrid swarm between the migratory marine and estuarine fishes Acanthopagrus australis and A. butcheri. Marine Ecology Progress Series, 421:199–204, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/421/
m421p199.pdf.

Reum:2012:SPT

[RH12] J. C. P. Reum and M. E. Hunsicker. Season and prey type influence size dependency of predator–prey body mass ratios in a marine fish assemblage. Marine Ecology Progress Series, 466:167–175, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/466/m466p167.pdf.

Rankin:2013:EHR

[RHB13] P. S. Rankin, R. W. Hannah, and M. T. O. Blume. Effect of hypoxia on rockfish movements: implications for understanding the roles of temperature, toxins and site fidelity. Marine Ecology Progress Series, 492:223–234, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/492/
m492p223.pdf.

Reiss:2014:MFP

[RHE14] K. Reiss, M. B. Herriot, and B. K. Eriksson. Multiple fish predators: effects of identity, density, and nutrients on lower trophic levels. Marine Ecology Progress Series, 497:1–12, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ray:2019:NPC

[RHF19] N. E. Ray, M. C. Henning, and R. W. Fulweiler. Nitrogen and phosphorus cycling in the digestive system and shell biofilm of the eastern oyster Crassostrea virginica. Marine Ecology Progress Series, 621:95–105, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m621p095.pdf.

Ribeiro:2015:CTP

[RHH^{+}15] F. Ribeiro, E. Hale, E. J. Hilton, T. R. Clardy, A. L. Deary, T. E. Targett, and J. E. Olney. Composition and temporal patterns of larval fish communities in Chesapeake and Delaware Bays, USA. Marine Ecology Progress Series, 527: 167–180, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/527/m527p167.pdf.

Reimer:2018:MWS

[RHMR18] J. N. Reimer, S. D. Hacker, B. A. Menge, and P. Ruggiero. Macrophyte wrack on sandy beaches of the US Pacific northwest is linked to proximity of source habitat, ocean upwelling, and beach morphology. Marine Ecology Progress Series, 594:263–269, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/594/m594p263.pdf.

Ruz:2017:VDC

[RHRB^{+}17] P. M. Ruz, P. Hidalgo, R. Riquelme-Bugueño, B. Franco-Cisterna, and M. Cornejo. Vertical distribution of copepod eggs in the oxygen minimum zone off Mejillones Bay (23^{o}S) in the Humboldt Current System. Marine Ecology Progress Series, 571:83–96, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/571/m571p083.pdf.

Rooper:2019:SEN

[RHS^{+}19] C. N. Rooper, G. R. Hoff, D. E. Stevenson, J. W. Orr, and I. B. Spies. Skate egg nursery habitat in the eastern Bering Sea: a predictive model. Marine Ecology Progress Series, 609:163–178, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/609/m609p163.pdf.

Rindorf:2019:SSD

[RHvD19] A. Rindorf, O. Henriksen, and M. van Deurs. Scale-specific density dependence in North Sea sandeel. Marine Ecology Progress Series, 619:97–110, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/619/m619p097.pdf.

Rastorgueff:2011:FSR

[RHVRC11] P. A. Rastorgueff, M. Harmelin-Vivien, P. Richard, and P. Chevaldonné. Feeding strategies and resource partitioning mitigate the effects of oligotrophy for marine cave mysids. Marine Ecology Progress Series, 440:163–176, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/440/
m440p163.pdf.

Reid:2012:EDS

[RHW12] T. A. Reid, M. A. Hindell, and C. Wilcox. Environmental determinants of the at-sea distribution of encounters between flesh-footed shearwaters Puffinus carniepes and fishing vessels. Marine Ecology Progress Series, 447:231–242, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/447/
m447p231.pdf.

Richards:2012:PSH

[Ric12] R. A. Richards. Phenological shifts in hatch timing of northern shrimp Pandalus borealis. Marine Ecology Progress Series, 456:149–158, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/456/m456p149.pdf.

Rattray:2013:DPC

[RIM^{+}13] A. Rattray, D. Ierodiaconou, J. Monk, V. L. Versace, and L. J. B. Laurenson. Detecting patterns of change in benthic habitats by acoustic remote sensing. Marine Ecology Progress Series, 477:1–13, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ritter:2017:AAB

[Rit17] A. F. Ritter. Adult avoidance behavior leads to ontogenetic shifts in habitat use of an intertidal fish. Marine Ecology Progress Series, 585:143–154, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/585/m585p143.pdf.

Robert:2012:SSB

[RJ12] K. Robert and S. K. Juniper. Surface-sediment bioturbation quantified with cameras on the NEPTUNE Canada cabled observatory. Marine Ecology Progress Series, 453:137– 149, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/453/m453p137.pdf.

Raffel:2014:MCP

[RJB^{+}14] A. E. Raffel, R. D. Jones, C. Butenhoff, A. L. Rice, and N. M. Scully. Methyl chloride production by calcareous periphyton mats from the Florida Everglades. Marine Ecology Progress Series, 514:35–41, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/514/m514p035.pdf.

Robert:2014:MDB

[RJH14] K. Robert, D. O. B. Jones, and V. A. I. Huvenne. Megafaunal distribution and biodiversity in a heterogeneous landscape: the iceberg-scoured Rockall Bank, NE Atlantic. Marine Ecology Progress Series, 501:67–88, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m501p067.pdf.

Riisgaard:2011:EJM

[RJL^{+}11] H. U. Riisgård, B. Hoffmann Jørgensen, K. Lundgreen, F. Storti, J. H. Walther, K. E. Meyer, and P. S. Larsen. The exhalant jet of mussels Mytilus edulis. Marine Ecology Progress Series, 437:147–164, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m437p147.pdf.

Reid:2018:INZ

[RJR18] A. P. Reid, C. R. Johnson, and D. J. Ross. Impacts of the New Zealand screwshell Maoricolpus roseus on growth and condition of juvenile commercial scallops Pecten fumatus. Marine Ecology Progress Series, 604:173–185, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/604/
m604p173.pdf.

Rojbek:2012:LLD

[RJTS12] M. C. Røjbek, C. Jacobsen, J. Tomkiewicz, and J. G. Støttrup. Linking lipid dynamics with the reproductive cycle in Baltic cod Gadus morhua. Marine Ecology Progress Series, 471: 215–234, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m471p215.pdf.

Rogers:2015:CCH

[RK15] T. L. Rogers and D. L. Kimbro. Causes and consequences of historical multi-trophic diversity change in an intertidal seagrass bed. Marine Ecology Progress Series, 540:13–26, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2015/540/m540p013.pdf.

Rii:2016:TVV

[RKC16] Y. M. Rii, D. M. Karl, and M. J. Church. Temporal and vertical variability in picophytoplankton primary productivity in the North Pacific subtropical gyre. Marine Ecology Progress Series, 562:1–18, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Rees:2018:HSP

[RKD18] M. J. Rees, N. A. Knott, and A. R. Davis. Habitat and seascape patterns drive spatial variability in temperate fish assemblages: implications for marine protected areas. Marine Ecology Progress Series, 607:171–186, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/607/
m607p171.pdf.

Rees:2015:RAE

[RKFD15] M. J. Rees, N. Knott, G. V. Fenech, and A. R. Davis. Rules of attraction: enticing pelagic fish to mid-water remote underwater video systems (RUVS). Marine Ecology Progress Series, 529:213–218, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/529/m529p213.pdf.

Ronconi:2010:IAV

[RKM^{+}10] R. A. Ronconi, H. N. Koopman, C. A. E. McKinstry, S. N. P. Wong, and A. J. Westgate. Inter-annual variability in diet of non-breeding pelagic seabirds Puffinus spp. at migratory staging areas: evidence from stable isotopes and fatty acids. Marine Ecology Progress Series, 419:267–282, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/419/
m419p267.pdf.

Romero-Kutzner:2015:RQV

[RKPB^{+}15] V. Romero-Kutzner, T. T. Packard, E. Berdalet, S. O. Roy, J. P. Gagné, and M. Gómez. Respiration quotient variability: bacterial evidence. Marine Ecology Progress Series, 519:47–59, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/519/m519p047.pdf.

Robertson:2016:BRA

[RKT16] F. C. Robertson, W. R. Koski, and A. W. Trites. Behavioral responses affect distribution analyses of bowhead whales in the vicinity of seismic operations. Marine Ecology Progress Series, 549:243–262, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/549/m549p243.pdf.

Roth:2010:ETS

[RKvW10] L. Roth, S. Koksal, and R. van Woesik. Effects of thermal stress on key processes driving coral-population dynamics. Marine Ecology Progress Series, 411:73–87, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/411/
m411p073.pdf.

Riisgaard:2010:PCM

[RL10] H. U. Riisgård and P. S. Larsen. Particle capture mechanisms in suspension-feeding invertebrates. Marine Ecology Progress Series, 418:255–293, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m418p255.pdf.

Richlen:2011:EDH

[RL11] M. L. Richlen and P. S. Lobel. Effects of depth, habitat, and water motion on the abundance and distribution of ciguatera dinoflagellates at Johnston Atoll, Pacific Ocean. Marine Ecology Progress Series, 421:51–66, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/421/
m421p051.pdf.

Riisgaard:2018:TVE

[RL18] H. U. Riisgård and P. S. Larsen. Thermal and viscous effects on ciliary suspension-feeding bivalves — no need for a new explanation: Comment on Specht & Fuchs (2018). Marine Ecology Progress Series, 596:263–265, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/596/m596p263.pdf. See

[SF18].

Riofrio-Lazo:2012:OCF

[RLAGL12] M. Riofrío-Lazo, D. Aurioles-Gamboa, and B. J. Le Boeuf. Ontogenetic changes in feeding habits of northern elephant seals revealed by δ ^{15}N and δ ^{13}C analysis of growth layers in teeth. Marine Ecology Progress Series, 450:229–241, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/450/
m450p229.pdf.

Rook:2010:BRD

[RLBC10] M. A. Rook, R. N. Lipcius, B. M. Bronner, and R. M. Chambers. Bycatch reduction device conserves diamondback terrapin without affecting catch of blue crab. Marine Ecology Progress Series, 409:171–179, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m409p171.pdf.

Raymond:2014:DSA

[RLG14] W. W. Raymond, A. T. Lowe, and A. W. E. Galloway. Degradation state of algal diets affects fatty acid composition but not size of red urchin gonads. Marine Ecology Progress Series, 509:213–225, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/509/m509p213.pdf.

Roelke:2013:COO

[RLH^{+}13] D. L. Roelke, H. P. Li, N. J. Hayden, C. J. Miller, S. E. Davis, A. Quigg, and Y. Buyukates. Co-occurring and opposing freshwater inflow effects on phytoplankton biomass, productivity and community composition of Galveston Bay, USA. Marine Ecology Progress Series, 477:61–76, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/477/
m477p061.pdf.

Ryan:2019:EPM

[RLH^{+}19] L. A. Ryan, S. K. Lynch, R. Harcourt, D. J. Slip, V. Peddemors, J. D. Everett, L. M. Harrison, and N. S. Hart. Environmental predictive models for shark attacks in Australian waters. Marine Ecology Progress Series, 631:165– 179, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/631/m631p165.pdf.

Rouze:2014:JTS

[RLM^{+}14] H. Rouzé, G. Lecellier, S. C. Mills, S. Planes, V. Berteaux-Lecellier, and H. Stewart. Juvenile Trapezia spp. crabs can increase juvenile host coral survival by protection from predation. Marine Ecology Progress Series, 515:151–159, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2014/515/m515p151.pdf.

Raab:2013:ITF

[RLN^{+}13] K. Raab, M. Llope, L. A. J. Nagelkerke, A. D. Rijnsdorp, L. R. Teal, P. Licandro, P. Ruardij, and M. Dickey-Collas. Influence of temperature and food availability on juvenile European anchovy Engraulis encrasicolus at its northern boundary. Marine Ecology Progress Series, 488:233–245, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/488/
m488p233.pdf.

Ramasco:2017:SFR

[RLN17] V. Ramasco, U. Lindström, and K. T. Nilssen. Selection and foraging response of harbour seals in an area of changing prey resources. Marine Ecology Progress Series, 581: 199–214, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/581/m581p199.pdf.

Rosas-Luis:2017:FET

[RLNLAF17] R. Rosas-Luis, J. Navarro, P. Loor-Andrade, and M. G. Forero. Feeding ecology and trophic relationships of pelagic sharks and billfishes coexisting in the central eastern Pacific Ocean. Marine Ecology Progress Series, 573:191–201, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/573/
m573p191.pdf.

Richardson:2018:WTA

[RLO18] J. P. Richardson, J. S. Lefcheck, and R. J. Orth. Warming temperatures alter the relative abundance and distribution of two co-occurring foundational seagrasses in Chesapeake Bay, USA. Marine Ecology Progress Series, 599:65–74, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/599/
m599p065.pdf.

Ryer:2010:TSW

[RLS10] C. H. Ryer, B. J. Laurel, and A. W. Stoner. Testing the shallow water refuge hypothesis in flatfish nurseries. Marine Ecology Progress Series, 415:275–282, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/415/m415p275.pdf.

Riisgaard:2014:DBM

[RLTL14] H. U. Riisgård, P. S. Larsen, R. Turja, and K. Lundgreen. Dwarfism of blue mussels in the low saline Baltic Sea—growth to the lower salinity limit. Marine Ecology Progress Series, 517:181–192, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/517/m517p181.pdf.

Rasmuson:2013:FPA

[RM13] L. K. Rasmuson and S. G. Morgan. Fish predation after weakly synchronized larval release in a coastal upwelling system. Marine Ecology Progress Series, 490:185–198, 2013. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m490p185.pdf.

Ryan:2019:RSC

[RM19] W. H. Ryan and T. E. Miller. Reproductive strategy changes across latitude in a clonal sea anemone. Marine Ecology Progress Series, 611:129–141, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/611/m611p129.pdf.

Radecker:2014:OAR

[RMB^{+}14] N. Rädecker, F. W. Meyer, V. N. Bednarz, U. Cardini, and C. Wild. Ocean acidification rapidly reduces dinitrogen fixation associated with the hermatypic coral Seriatopora hystrix. Marine Ecology Progress Series, 511:297–302, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m511p297.pdf.

Rudershausen:2018:FHA

[RMB18] P. J. Rudershausen, J. H. Merrell, and J. A. Buckel. Fragmentation of habitat affects communities and movement of nekton in salt marsh tidal creeks. Marine Ecology Progress Series, 586:57–72, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/586/m586p057.pdf.

Riedel:2014:RIR

[RMBM14] A. Riedel, K. Monro, M. W. Blows, and D. J. Marshall. Relative influence of resident species and environmental variation on community assembly. Marine Ecology Progress Series, 499:103–113, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/499/m499p103.pdf.

Range:2014:RSS

[RMC^{+}14] P. Range, M. Martins, S. Cabral, D. Piló, R. Ben-Hamadou, M. A. Teodósio, F. Leitão, T. Drago, A. P. Oliveira, D. Matias, and L. Chícharo. Relative sensitivity of soft-bottom intertidal macrofauna to increased CO_{2} and experimental stress. Marine Ecology Progress Series, 509:153–170, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/509/
m509p153.pdf.

Radabaugh:2014:ELE

[RMHP14] K. R. Radabaugh, E. M. Malkin, D. J. Hollander, and E. B. Peebles. Evidence for light-environment control of carbon isotope fractionation by benthic microalgal communities. Marine Ecology Progress Series, 495:77–90, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/495/
m495p077.pdf.

Reem:2013:PGS

[RMKR13] E. Reem, I. Mohanty, G. Katzir, and B. Rinkevich. Population genetic structure and modes of dispersal for the colonial ascidian Botryllus schlosseri along the Scandinavian Atlantic coasts. Marine Ecology Progress Series, 485:143– 154, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/485/m485p143.pdf.

Renken:2010:EPE

[RMME10] H. Renken, P. J. Mumby, I. Matsikis, and H. J. Edwards. Effects of physical environmental conditions on the patch dynamics of Dictyota pulchella and Lobophora variegata on Caribbean coral reefs. Marine Ecology Progress Series, 403: 63–74, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/403/m403p063.pdf.

Robinson:2017:MDB

[RMNP17] N. J. Robinson, S. J. Morreale, R. Nel, and F. V. Paladino. Movements and diving behaviour of inter-nesting leatherback turtles in an oceanographically dynamic habitat in South Africa. Marine Ecology Progress Series, 571:221– 232, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/571/m571p221.pdf.

Ramirez:2019:RST

[RMP^{+}19] M. D. Ramirez, J. A. Miller, E. Parks, L. Avens, L. R. Goshe, J. A. Seminoff, M. L. Snover, and S. S. Heppell. Reconstructing sea turtle ontogenetic habitat shifts through trace element analysis of bone tissue. Marine Ecology Progress Series, 608:247–262, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/608/m608p247.pdf.

Roy:2012:SPL

[RMR12] A. Roy, A. Metaxas, and T. Ross. Swimming patterns of larval Strongylocentrotus droebachiensis in turbulence in the laboratory. Marine Ecology Progress Series, 453:117– 127, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/453/m453p117.pdf.

Russ:2015:LTN

[RMRA15] G. R. Russ, K. I. Miller, J. R. Rizzari, and A. C. Alcala. Long-term no-take marine reserve and benthic habitat effects on coral reef fishes. Marine Ecology Progress Series, 529:233–248, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/529/m529p233.pdf.

Rodriguez:2016:BEM

[RMSPA16] S. Rodriguez, A. P. Martín, I. Sousa-Pinto, and F. Arenas. Biodiversity effects on macroalgal productivity: exploring the roles of richness, evenness and species traits. Marine Ecology Progress Series, 562:79–91, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/562/m562p079.pdf.

Ryan:2013:SIA

[RMT^{+}13] C. Ryan, B. McHugh, C. N. Trueman, R. Sabin, R. Deaville, C. Harrod, S. D. Berrow, and I. O’Connor. Stable isotope analysis of baleen reveals resource partitioning among sympatric rorquals and population structure in fin whales. Marine Ecology Progress Series, 479:251–261, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/479/
m479p251.pdf.

Rozas:2013:ERH

[RMV13] L. P. Rozas, C. W. Martin, and J. F. Valentine. Effects of reduced hydrological connectivity on the nursery use of shallow estuarine habitats within a river delta. Marine Ecology Progress Series, 492:9–20, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/492/m492p009.pdf.

Raab:2012:DOB

[RNB^{+}12] K. Raab, L. A. J. Nagelkerke, C. Boerée, A. D. Rijnsdorp, A. Temming, and M. Dickey-Collas. Dietary overlap between the potential competitors herring, sprat and anchovy in the North Sea. Marine Ecology Progress Series, 470:101– 111, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/470/m470p101.pdf.

Rey:2016:COC

[RNB^{+}16] F. Rey, G. M. Silva Neto, C. Brandão, D. Ramos, B. Silva, R. Rosa, H. Queiroga, and R. Calado. Contrasting oceanographic conditions during larval development influence the benthic performance of a marine invertebrate with a biphasic life cycle. Marine Ecology Progress Series, 546:135– 146, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/546/m546p135.pdf.

Riccialdelli:2010:IAP

[RNFG10] Luciana Riccialdelli, Seth D. Newsome, Marilyn L. Fogel, and R. Natalie P. Goodall. Isotopic assessment of prey and habitat preferences of a cetacean community in the southwestern South Atlantic Ocean. Marine Ecology Progress Series, 418:235–248, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/418/m418p235.pdf.

Rosales-Nanduca:2011:MMM

[RNGUR^{+}11] H. Rosales-Nanduca, T. Gerrodette, J. Urbán-R, G. Cárdenas-Hinojosa, and L. Medrano-González. Macroecology of marine mammal species in the Mexican Pacific Ocean: diversity and distribution. Marine Ecology Progress Series, 431: 281–291, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m431p281.pdf.

Riisgaard:2015:IEP

[RNH15] K. Riisgaard, T. G. Nielsen, and P. J. Hansen. Impact of elevated pH on succession in the Arctic spring bloom. Marine Ecology Progress Series, 530:63–75, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/530/
m530p063.pdf.

Riemann:2011:DPP

[RNK^{+}11] L. Riemann, T. G. Nielsen, T. Kragh, K. Richardson, H. Parner, H. H. Jakobsen, and P. Munk. Distribution and production of plankton communities in the subtropical convergence zone of the Sargasso Sea. I. phytoplankton and bacterioplankton. Marine Ecology Progress Series, 426:57–70, 2011.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m426p057.pdf.

Rowell:2015:FSP

[RNSA15] T. J. Rowell, R. S. Nemeth, M. T. Schärer, and R. S. Appeldoorn. Fish sound production and acoustic telemetry reveal behaviors and spatial patterns associated with spawning aggregations of two Caribbean groupers. Marine Ecology Progress Series, 518:239–254, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/518/m518p239.pdf.

Riekenberg:2017:UDO

[ROE17] P. M. Riekenberg, J. M. Oakes, and B. D. Eyre. Uptake of dissolved organic and inorganic nitrogen in microalgaedominated sediment: comparing dark and light in situ and ex situ additions of ^{15}N. Marine Ecology Progress Series, 571: 29–42, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/571/m571p029.pdf.

Rogers:2014:HEP

[ROKS14] L. A. Rogers, E. M. Olsen, H. Knutsen, and N. C. Stenseth. Habitat effects on population connectivity in a coastal seascape. Marine Ecology Progress Series, 511:153–163, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m511p153.pdf.

Rodil:2015:CEW

[ROLA15] I. F. Rodil, C. Olabarria, M. Lastra, and F. Arenas. Combined effects of wrack identity and solar radiation on associated beach macrofaunal assemblages. Marine Ecology Progress Series, 531:167–178, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2015/531/m531p167.pdf.

Ramsvatn:2012:ONC

[RP12] S. Ramsvatn and T. Pedersen. Ontogenetic niche changes in haddock Melanogrammus aeglefinus reflected by stable isotope signatures, δ ^{13}C and δ ^{15}N. Marine Ecology Progress Series, 451:175–185, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/451/m451p175.pdf.

Rafalowski:2013:NEP

[RP13] S. Rafalowski and C. Plante. Non-equilibrium processes structuring benthic bacterial communities following deposit feeding by a sea cucumber. Marine Ecology Progress Series, 478:115–126, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/478/m478p115.pdf.

Rishworth:2015:PPR

[RP15] G. M. Rishworth and P. A. Pistorius. Protection and provisioning: the role of parental behaviour in terms of chick growth and survival in a pelagic seabird. Marine Ecology Progress Series, 530:153–162, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/530/m530p153.pdf.

Rouchon:2017:SLE

[RP17] A. M. Rouchon and N. E. Phillips. Short larval exposure to low level of copper has long-lasting latent effects on juvenile performance in the sea urchin Evechinus chloroticus. Marine Ecology Progress Series, 579:67–79, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/579/m579p067.pdf.

Rogers:2017:AAE

[RPB17] J. G. D. Rogers, É. E. Pláganyi, and R. C. Babcock. Aggregation, Allee effects and critical thresholds for the management of the crown-of-thorns starfish Acanthaster planci. Marine Ecology Progress Series, 578:99–114, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/578/
m578p099.pdf.

Roos:2019:DAB

[RPCL19] N. C. Roos, M. G. Pennino, A. R. Carvalho, and G. O. Longo. Drivers of abundance and biomass of Brazilian parrotfishes. Marine Ecology Progress Series, 623:117–130, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/623/
m623p117.pdf.

Reece:2013:SLR

[RPE^{+}13] J. S. Reece, D. Passeri, L. Ehrhart, S. C. Hagen, A. Hays, C. Long, R. F. Noss, M. Bilskie, C. Sanchez, M. V. Schwoerer, B. Von Holle, J. Weishampel, and S. Wolf. Sea level rise, land use, and climate change influence the distribution of loggerhead turtle nests at the largest USA rookery (Melbourne Beach, Florida). Marine Ecology Progress Series, 493:259–274, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/493/m493p259.pdf.

Roth:2015:FPD

[RPGP^{+}15] M. S. Roth, J. L. Padilla-Gamiño, X. Pochon, R. R. Bidigare, R. D. Gates, C. M. Smith, and H. L. Spalding. Fluorescent proteins in dominant mesophotic reef-building corals. Marine Ecology Progress Series, 521:63–79, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m521p063.pdf.

Robinson:2015:PDB

[RPHB15] T. B. Robinson, H. R. Pope, L. Hawken, and C. Binneman. Predation-driven biotic resistance fails to restrict the spread of a sessile rocky shore invader. Marine Ecology Progress Series, 522:169–179, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/522/m522p169.pdf.

Rosemond:2018:FUR

[RPL^{+}18] R. C. Rosemond, A. B. Paxton, H. R. Lemoine, S. R. Fegley, and C. H. Peterson. Fish use of reef structures and adjacent sand flats: implications for selecting minimum buffer zones between new artificial reefs and existing reefs. Marine Ecology Progress Series, 587:187–199, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/587/m587p187.pdf.

Rayssac:2010:TES

[RPLT10] N. Rayssac, F. Pernet, O. Lacasse, and R. Tremblay. Temperature effect on survival, growth, and triacylglycerol content during the early ontogeny of Mytilus edulis and M. trossulus. Marine Ecology Progress Series, 417:183–191, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/417/
m417p183.pdf.

Romero:2012:DFC

[RPM12] E. Romero, F. Peters, and C. Marrasé. Dynamic forcing of coastal plankton by nutrient imbalances and match-mismatch between nutrients and turbulence. Marine Ecology Progress Series, 464:69–87, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/464/m464p069.pdf.

Ramirez:2013:YRD

[RPM^{+}13a] I. Ramírez, V. H. Paiva, D. Menezes, I. Silva, R. A. Phillips, J. A. Ramos, and S. Garthe. Year-round distribution and habitat preferences of the bugio petrel. Marine Ecology Progress Series, 476:269–284, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/476/m476p269.pdf.

Rohner:2013:TSE

[RPM^{+}13b] C. A. Rohner, S. J. Pierce, A. D. Marshall, S. J. Weeks, M. B. Bennett, and A. J. Richardson. Trends in sightings and environmental influences on a coastal aggregation of manta rays and whale sharks. Marine Ecology Progress Series, 482:153–168, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m482p153.pdf.

Renner:2013:MDA

[RPP^{+}13] M. Renner, J. K. Parrish, J. F. Piatt, K. J. Kuletz, A. E. Edwards, and G. L. Hunt, Jr. Modeled distribution and abundance of a pelagic seabird reveal trends in relation to fisheries. Marine Ecology Progress Series, 484:259–277, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m484p259.pdf.

Rosciano:2016:DPS

[RPR16] N. G. Rosciano, M. J. Polito, and A. Raya Rey. Do penguins share? Evidence of foraging niche segregation between but not within two sympatric, central-place foragers. Marine Ecology Progress Series, 548:249–262, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/548/m548p249.pdf.

Ramajo:2016:PTO

[RPRN^{+}16] L. Ramajo, L. Prado, A. B. Rodriguez-Navarro, M. A. Lardies, C. M. Duarte, and N. A. Lagos. Plasticity and tradeoffs in physiological traits of intertidal mussels subjected to freshwater-induced environmental variation. Marine Ecology Progress Series, 553:93–109, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/553/m553p093.pdf.

Rice:2012:TIM

[RPS12] E. J. Rice, C. Panzeca, and G. M. Stewart. Temperatureinduced microbubbles within natural marine samples may inflate small-particle counts in a Coulter counter. Marine Ecology Progress Series, 450:275–280, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/450/
m450p275.pdf.

Ross:2014:TCI

[RQN^{+}14] R. M. Ross, L. B. Quetin, T. Newberger, T. C. Shaw, J. L. Jones, S. A. Oakes, and K. J. Moore. Trends, cycles, interannual variability for three pelagic species west of the Antarctic Peninsula 1993–2008. Marine Ecology Progress Series, 515:11–32, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m515p011.pdf.

Rasmussen:2018:HDI

[RR18] T. K. Rasmussen and H. S. Randhawa. Host diet influences parasite diversity: a case study looking at tapeworm diversity among sharks. Marine Ecology Progress Series, 605:1–16, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Richmond:2013:IVE

[RRB13] C. E. Richmond, K. A. Rose, and D. L. Breitburg. Individual variability and environmental conditions: effects on zooplankton cohort dynamics. Marine Ecology Progress Series, 486:59–78, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/486/m486p059.pdf.

Rhodes:2011:NEP

[RRG^{+}11] L. D. Rhodes, C. A. Rice, C. M. Greene, D. J. Teel, S. L. Nance, P. Moran, C. A. Durkin, and S. B. Gezhegne. Nearshore ecosystem predictors of a bacterial infection in juvenile Chinook salmon. Marine Ecology Progress Series, 432:161–172, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m432p161.pdf.

Richards:2016:DEN

[RRH^{+}16] Z. T. Richards, N. M. Ryan, E. S. Harvey, R. Garcia, and J. P. A. Hobbs. Diversity on the edge: non-linear patterns of coral community structure at an isolated oceanic island. Marine Ecology Progress Series, 546:61–74, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/546/
m546p061.pdf.

Razak:2019:LTG

[RRL^{+}19] T. B. Razak, G. Roff, J. M. Lough, D. Prayudi, N. E. Cantin, and P. J. Mumby. Long-term growth trends of massive Porites corals across a latitudinal gradient in the Indo–Pacific. Marine Ecology Progress Series, 626:69–82, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/626/
m626p069.pdf.

Ramos:2017:CNS

[RRRF^{+}17] J. E. Ramos, A. Ramos-Rodríguez, G. Bazzino Ferreri, J. A. Kurczyn, D. Rivas, and C. A. Salinas-Zavala. Characterization of the northernmost spawning habitat of Dosidicus gigas with implications for its northwards range extension. Marine Ecology Progress Series, 572:179–192, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/572/
m572p179.pdf.

Rodgers:2015:NSM

[RRS15] K. L. Rodgers, T. A. V. Rees, and N. T. Shears. A novel system for measuring in situ rates of photosynthesis and respiration of kelp. Marine Ecology Progress Series, 528:101– 115, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/528/m528p101.pdf.

Raper:2013:SSL

[RS13] J. L. D. Raper and D. C. Schneider. Spatial scaling from latitudinal gradients: growth rates in the American lobster Homarus americanus. Marine Ecology Progress Series, 483: 231–243, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/483/m483p231.pdf.

Rasmuson:2014:LSO

[RS14] L. K. Rasmuson and A. L. Shanks. In situ observations of Dungeness crab megalopae used to estimate transport distances by internal waves. Marine Ecology Progress Series, 511:143–152, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/511/m511p143.pdf.

Rodgers:2016:MKF

[RS16] K. L. Rodgers and N. T. Shears. Modelling kelp forest primary production using in situ photosynthesis, biomass and light measurements. Marine Ecology Progress Series, 553: 67–79, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/553/m553p067.pdf.

Rowell:2012:SPI

[RSA^{+}12] T. J. Rowell, M. T. Schärer, R. S. Appeldoorn, M. I. Nemeth, D. A. Mann, and J. A. Rivera. Sound production as an indicator of red hind density at a spawning aggregation. Marine Ecology Progress Series, 462:241–250, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m462p241.pdf.

Ruck:2014:RDQ

[RSC14] K. E. Ruck, D. K. Steinberg, and E. A. Canuel. Regional differences in quality of krill and fish as prey along the Western Antarctic Peninsula. Marine Ecology Progress Series, 509:39–55, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/509/m509p039.pdf.

Renkawitz:2015:CTS

[RSDN15] M. D. Renkawitz, T. F. Sheehan, H. J. Dixon, and R. Nygaard. Changing trophic structure and energy dynamics in the Northwest Atlantic: implications for Atlantic salmon feeding at West Greenland. Marine Ecology Progress Series, 538:197–211, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/538/m538p197.pdf.

Rogers:2018:SDI

[RSE18] T. L. Rogers, H. K. Schultz, and J. K. Elliott. Size-dependent interference competition between two sea star species demographically affected by wasting disease. Marine Ecology Progress Series, 589:167–177, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/589/m589p167.pdf.

Rooper:2016:VIS

[RSG^{+}16] C. N. Rooper, M. F. Sigler, P. Goddard, P. Malecha, R. Towler, K. Williams, R. Wilborn, and M. Zimmermann. Validation and improvement of species distribution models for structure-forming invertebrates in the eastern Bering Sea with an independent survey. Marine Ecology Progress Series, 551:117–130, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/551/m551p117.pdf.

Robinson:2019:PSL

[RSH^{+}19] H. E. Robinson, J. R. Strickler, M. J. Henderson, D. K. Hartline, and P. H. Lenz. Predation strategies of larval clownfish capturing evasive copepod prey. Marine Ecology Progress Series, 614:125–146, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/614/m614p125.pdf.

Rooker:2010:PCR

[RSHM10] J. R. Rooker, G. W. Stunz, S. A. Holt, and T. J. Minello. Population connectivity of red drum in the northern Gulf of Mexico. Marine Ecology Progress Series, 407:187–196, 2010.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m407p187.pdf.

Radford:2014:ACR

[RSJ14] C. A. Radford, J. A. Stanley, and A. G. Jeffs. Adjacent coral reef habitats produce different underwater sound signatures. Marine Ecology Progress Series, 505:19–28, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/505/
m505p019.pdf.

Riisgaard:2014:TRT

[RSK^{+}14] K. Riisgaard, R. Swalethorp, S. Kjellerup, T. Juul-Pedersen, and T. G. Nielsen. Trophic role and top-down control of a subarctic protozooplankton community. Marine Ecology Progress Series, 500:67–82, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/500/m500p067.pdf.

Ralph:2013:BSA

[RSO^{+}13] G. M. Ralph, R. D. Seitz, R. J. Orth, K. E. Knick, and R. N. Lipcius. Broad-scale association between seagrass cover and juvenile blue crab density in Chesapeake Bay. Marine Ecology Progress Series, 488:51–63, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m488p051.pdf.

Rolland:2016:HNA

[RSP^{+}16] R. M. Rolland, R. S. Schick, H. M. Pettis, A. R. Knowlton, P. K. Hamilton, J. S. Clark, and S. D. Kraus. Health of North Atlantic right whales Eubalaena glacialis over three decades: from individual health to demographic and population health trends. Marine Ecology Progress Series, 542: 265–282, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/542/m542p265.pdf.

Rishworth:2017:FAL

[RSP17] G. M. Rishworth, N. A. Strydom, and R. Perissinotto. Fishes associated with living stromatolite communities in peritidal pools: predators, recruits and ecological traps. Marine Ecology Progress Series, 580:153–167, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/580/m580p153.pdf.

Ray:2016:MGC

[RSS^{+}16] J. L. Ray, K. S. Skaar, P. Simonelli, A. Larsen, A. Sazhin, H. H. Jakobsen, J. C. Nejstgaard, and C. Troedsson. Molecular gut content analysis demonstrates that Calanus grazing on Phaeocystis pouchetii and Skeletonema marinoi is sensitive to bloom phase but not prey density. Marine Ecology Progress Series, 542:63–77, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m542p063.pdf.

Ruggerone:2019:UBP

[RSSvV19] G. T. Ruggerone, A. M. Springer, L. D. Shaul, and G. B. van Vliet. Unprecedented biennial pattern of birth and mortality in an endangered apex predator, the southern resident killer whale, in the eastern North Pacific Ocean. Marine Ecology Progress Series, 608:291–296, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/608/m608p291.pdf.

Radford:2010:LCH

[RST^{+}10] C. A. Radford, J. A. Stanley, C. T. Tindle, J. C. Montgomery, and A. G. Jeffs. Localised coastal habitats have distinct underwater sound signatures. Marine Ecology Progress Series, 401:21–29, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/401/m401p021.pdf.

Richardson:2015:URE

[RST15] J. R. Richardson, N. T. Shears, and R. B. Taylor. Using relative eye size to estimate the length of fish from a single camera image. Marine Ecology Progress Series, 538: 213–219, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/538/m538p213.pdf.

Ronconi:2014:GDR

[RSTM14] R. A. Ronconi, R. J. Steenweg, P. D. Taylor, and M. L. Mallory. Gull diets reveal dietary partitioning, influences of isotopic signatures on body condition, and ecosystem changes at a remote colony. Marine Ecology Progress Series, 514:

247–261, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/514/m514p247.pdf.

Reis-Santos:2013:CBE

[RSTV^{+}13] P. Reis-Santos, S. E. Tanner, R. P. Vasconcelos, T. S. Elsdon, H. N. Cabral, and B. M. Gillanders. Connectivity between estuarine and coastal fish populations: contributions of estuaries are not consistent over time. Marine Ecology Progress Series, 491:177–186, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/491/m491p177.pdf.

Robertson:2010:UCO

[RSV10] D. R. Robertson and W. F. Smith-Vaniz. Use of clove oil in collecting coral reef fishes for research. Marine Ecology Progress Series, 401:295–302, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m401p295.pdf.

Rossoll:2013:CID

[RSW13] D. Rossoll, U. Sommer, and M. Winder. Community interactions dampen acidification effects in a coastal plankton system. Marine Ecology Progress Series, 486:37–46, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/486/
m486p037.pdf.

Reid:2016:INV

[RSW^{+}16] W. D. K. Reid, C. J. Sweeting, B. D. Wigham, R. A. R. McGill, and N. V. C. Polunin. Isotopic niche variability in macroconsumers of the East Scotia Ridge (Southern Ocean) hydrothermal vents: What more can we learn from an ellipse? Marine Ecology Progress Series, 542:13–24, 2016. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m542p013.pdf.

Rosen:2012:EPP

[RT12] D. A. S. Rosen and D. J. Tollit. Effects of phylogeny and prey type on fatty acid calibration coefficients in three pinniped species: implications for the QFASA dietary quantification technique. Marine Ecology Progress Series, 467:263– 276, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/467/m467p263.pdf.

Rossi:2017:TOH

[RT17] G. S. Rossi and V. Tunnicliffe. Trade-offs in a high CO_{2} habitat on a subsea volcano: condition and reproductive features of a bathymodioline mussel. Marine Ecology Progress Series, 574:49–64, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/574/m574p049.pdf.

Reglero:2014:WDT

[RTÁB^{+}14] P. Reglero, D. P. Tittensor, D. Álvarez-Berastegui, A. Aparicio-González, and B. Worm. Worldwide distributions of tuna larvae: revisiting hypotheses on environmental requirements for spawning habitats. Marine Ecology Progress Series, 501: 207–224, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/501/m501p207.pdf.

Rountos:2014:THD

[RTC^{+}14] K. J. Rountos, Y. Z. Tang, R. M. Cerrato, C. J. Gobler, and E. K. Pikitch. Toxicity of the harmful dinoflagellate Cochlodinium polykrikoides to early life stages of three estuarine forage fish. Marine Ecology Progress Series, 505: 81–94, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/505/m505p081.pdf.

Rodriguez-Tress:2017:ABY

[RTCF^{+}17] P. Rodriguez-Tress, M. Capello, F. Forget, M. Soria, S. P. Beeharry, N. Dussooa, and L. Dagorn. Associative behavior of yellowfin Thunnus albacares, skipjack Katsuwonus pelamis, and bigeye tuna T. obesus at anchored fish aggregating devices (FADs) off the coast of Mauritius. Marine Ecology Progress Series, 570:213–222, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/570/m570p213.pdf.

Rigolet:2014:FWS

[RTD14] C. Rigolet, E. Thiébaut, and S. F. Dubois. Food web structures of subtidal benthic muddy habitats: evidence of microphytobenthos contribution supported by an engineer species. Marine Ecology Progress Series, 500:25–41, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/500/
m500p025.pdf.

Rivas:2016:PED

[RTDUM16] M. L. Rivas, P. Santidrián Tomillo, J. Diéguez-Uribeondo, and A. Marco. Potential effects of dune scarps caused by beach erosion on the nesting behavior of leatherback turtles. Marine Ecology Progress Series, 551:239–248, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/551/
m551p239.pdf.

Riul:2013:IPC

[RTJ^{+}13] P. Riul, C. H. Targino, L. A. C. Júnior, J. C. Creed, P. A. Horta, and G. C. Costa. Invasive potential of the coral Tubastraea coccinea in the southwest Atlantic. Marine Ecology Progress Series, 480:73–81, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/480/m480p073.pdf.

Rhodes:2011:DDA

[RTM11] K. L. Rhodes, B. M. Taylor, and J. L. McIlwain. Detailed demographic analysis of an Epinephelus polyphekadion spawning aggregation and fishery. Marine Ecology Progress Series, 421:183–198, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/421/m421p183.pdf.

Radford:2011:MRE

[RTMJ11] C. A. Radford, C. T. Tindle, J. C. Montgomery, and A. G. Jeffs. Modelling a reef as an extended sound source increases the predicted range at which reef noise may be heard by fish larvae. Marine Ecology Progress Series, 438:167– 174, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/438/m438p167.pdf.

Rasheed:2011:LTC

[RU11] M. A. Rasheed and R. K. F. Unsworth. Long-term climateassociated dynamics of a tropical seagrass meadow: implications for the future. Marine Ecology Progress Series, 422: 93–103, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/422/m422p093.pdf.

Riascos:2016:HFB

[RUD^{+}16] J. M. Riascos, R. A. Uribe, S. Donayre, D. Flores, O. Galindo, C. Quispe, and J. Gonzalez. Human footprints on benthic communities in marine reserves: a study case in the most productive upwelling system worldwide. Marine Ecology Progress Series, 557:65–75, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/557/m557p065.pdf.

Reymond:2011:IGP

[RUP11a] C. E. Reymond, S. Uthicke, and J. M. Pandolfi. Inhibited growth in the photosymbiont-bearing foraminifer Marginopora vertebralis from the nearshore Great Barrier Reef, Australia. Marine Ecology Progress Series, 435:97–109, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/435/

m435p097.pdf. See erratum [ACA^{+}10a].

Reymond:2011:EIG

[RUP11b] Claire E. Reymond, S. Uthicke, and John M. Pandolfi. Erratum: Inhibited growth in the photosymbiont-bearing foraminifer Marginopora vertebralis from the nearshore Great Barrier Reef, Australia, Mar Ecol Prog Ser 435:97–109, 2011. Marine Ecology Progress Series, 441:302, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/441/

m441p302.pdf. See [RUP11a].

Reglero:2011:CAS

[RUPT^{+}11] P. Reglero, A. Urtizberea, A. Pérez-Torres, F. Alemany, and Ø. Fiksen. Cannibalism among size classes of larvae may be a substantial mortality component in tuna. Marine Ecology Progress Series, 433:205–219, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/433/m433p205.pdf.

Rowntree:2013:URH

[RUS^{+}13] V. J. Rowntree, M. M. Uhart, M. Sironi, A. Chirife, M. Di Martino, L. La Sala, L. Musmeci, N. Mohamed, J. Andrejuk, D. McAloose, J. E. Sala, A. Carribero, H. Rally, M. Franco, F. R. Adler, R. L. Brownell, Jr., J. Seger, and T. Rowles. Unexplained recurring high mortality of southern right whale Eubalaena australis calves at Península Valdís, Argentina. Marine Ecology Progress Series, 493:275–289, 2013. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m493p275.pdf.

Richoux:2014:FAP

[RVF14] N. B. Richoux, I. Vermeulen, and P. W. Froneman. Fatty acid profiles reveal temporal and spatial differentiation in diets within and among syntopic rocky shore suspension-feeders. Marine Ecology Progress Series, 495:143–160, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/495/
m495p143.pdf.

Rueda:2019:TEF

[RVH^{+}19] L. Rueda, M. Valls, M. Hidalgo, B. Guijarro, A. Esteban, and E. Massutí. From trophic ecology to fish condition: contrasting pathways for European hake in the western Mediterranean. Marine Ecology Progress Series, 623:131–143, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/623/
m623p131.pdf.

Rijnsdorp:2012:EET

[RvOP12] A. D. Rijnsdorp, H. M. J. van Overzee, and J. J. Poos. Ecological and economic trade-offs in the management of mixed fisheries: a case study of spawning closures in flatfish fisheries. Marine Ecology Progress Series, 447:179–194, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/447/
m447p179.pdf.

Riccialdelli:2017:EIN

[RVPG17] L. Riccialdelli, M. N. Paso Viola, H. O. Panarello, and R. N. P. Goodall. Evaluating the isotopic niche of beaked whales from the southwestern South Atlantic and Southern Oceans. Marine Ecology Progress Series, 581:183–198, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/581/
m581p183.pdf.

Ritson-Williams:2016:PLS

[RWAP16] R. Ritson-Williams, S. N. Arnold, and V. J. Paul. Patterns of larval settlement preferences and post-settlement survival for seven Caribbean corals. Marine Ecology Progress Series, 548:127–138, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/548/m548p127.pdf.

Reiss:2015:WPD

[RWG15] C. S. Reiss, J. Walsh, and M. E. Goebel. Winter preconditioning determines feeding ecology of Euphausia superba in the Antarctic Peninsula. Marine Ecology Progress Series, 519:89–101, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/519/m519p089.pdf.

Rognstad:2014:CPR

[RWH14] R. L. Rognstad, D. S. Wethey, and T. J. Hilbish. Connectivity and population repatriation: limitations of climate and input into the larval pool. Marine Ecology Progress Series, 495:175–183, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m495p175.pdf.

Reid:2010:KPD

[RWM^{+}10] K. Reid, J. L. Watkins, E. J. Murphy, P. N. Trathan, S. Fielding, and P. Enderlein. Krill population dynamics at South Georgia: implications for ecosystem-based fisheries management. Marine Ecology Progress Series, 399:243– 252, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/399/m399p243.pdf.

Reynolds:2012:ERS

[RWM^{+}12] L. K. Reynolds, M. Waycott, K. J. McGlathery, R. J. Orth, and J. C. Zieman. Eelgrass restoration by seed maintains genetic diversity: case study from a coastal bay system.

Marine Ecology Progress Series, 448:223–233, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m448p223.pdf.

Raes:2014:CLD

[RWM^{+}14] E. J. Raes, A. M. Waite, A. S. McInnes, H. Olsen, H. M. Nguyen, N. Hardman-Mountford, and P. A. Thompson. Changes in latitude and dominant diazotrophic community alter N_{2} fixation. Marine Ecology Progress Series, 516:85– 102, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m516p085.pdf.

Rains:2018:EFI

[RWM18] S. A. M. Rains, M. J. Wilberg, and T. J. Miller. Evaluation of fishery-induced sperm limitation in Chesapeake Bay blue crab using an individual-based model. Marine Ecology Progress Series, 596:127–142, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/596/m596p127.pdf.

Reid:2012:ETP

[RWMP12] W. D. K. Reid, B. D. Wigham, R. A. R. McGill, and N. V. C. Polunin. Elucidating trophic pathways in benthic deep-sea assemblages of the Mid-Atlantic Ridge north and south of the Charlie-Gibbs Fracture Zone. Marine Ecology Progress Series, 463:89–103, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/463/m463p089.pdf.

Reinhardt:2013:ETR

[RWO13] J. F. Reinhardt, R. B. Whitlatch, and R. W. Osman. Effects of temperature on the recruitment phenology and niche overlap of shallow epifaunal assemblages in southern New England. Marine Ecology Progress Series, 489:61–74, 2013. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m489p061.pdf.

Rechisky:2014:TDM

[RWP^{+}14] E. L. Rechisky, D. W. Welch, A. D. Porter, J. E. Hess, and S. R. Narum. Testing for delayed mortality effects in the early marine life history of Columbia River basin yearling Chinook salmon. Marine Ecology Progress Series, 496:159–180, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m496p159.pdf.

Reinsel:2015:PIF

[RWR^{+}15] K. A. Reinsel, J. M. Welch, A. O. Romero, E. M. Parks, K. M. McQueen, M. J. Smith, A. L. Price, H. R. Clark, S. E. Zmina, C. A. Williams, and R. B. Forward, Jr. Planktonic ingress of fiddler crab megalopae to the Newport River Estuary, NC: evidence for semilunar periodicity in Uca pugnax and species-specific sampling bias by hog’s hair larval collectors. Marine Ecology Progress Series, 523:105–113, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/523/
m523p105.pdf.

Randall:2019:LSS

[RWR^{+}19] J. Randall, S. Wotherspoon, J. Ross, J. P. Hermand, and C. R. Johnson. An in situ study of production from diel oxygen modelling, oxygen exchange, and electron transport rate in the kelp Ecklonia radiata. Marine Ecology Progress Series, 615:51–65, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/615/m615p051.pdf.

Richerson:2015:MTP

[RWS^{+}15] K. Richerson, G. M. Watters, J. A. Santora, I. D. Shroeder, and M. Mangel. More than passive drifters: a stochastic dynamic model for the movement of Antarctic krill. Marine Ecology Progress Series, 529:35–48, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/529/
m529p035.pdf.

Redelstein:2018:SSW

[RZB18] R. Redelstein, G. Zotz, and T. Balke. Seedling stability in waterlogged sediments: an experiment with saltmarsh plants. Marine Ecology Progress Series, 590:95–108, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/590/
m590p095.pdf.

Ricard:2016:NIS

[RZH16a] D. Ricard, F. Zimmermann, and M. Heino. Are negative intraspecific interactions important for recruitment dynamics? A case study of Atlantic fish stocks. Marine Ecology Progress Series, 547:211–217, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/547/m547p211.pdf.

Richlen:2016:DAF

[RZH^{+}16b] M. L. Richlen, O. Zielinski, L. Holinde, U. Tillmann, A. Cembella, Y. Lyu, and D. M. Anderson. Distribution of Alexandrium fundyense (Dinophyceae) cysts in Greenland and Iceland, with an emphasis on viability and growth in the Arctic. Marine Ecology Progress Series, 547:33–46, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/547/
m547p033.pdf.

Rooper:2014:PMC

[RZPH14] C. N. Rooper, M. Zimmermann, M. M. Prescott, and A. J. Hermann. Predictive models of coral and sponge distribution, abundance and diversity in bottom trawl surveys of the Aleutian Islands, Alaska. Marine Ecology Progress Series, 503:157–176, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/503/m503p157.pdf.

Riedel:2012:TBM

[RZS12] B. Riedel, M. Zuschin, and M. Stachowitsch. Tolerance of benthic macrofauna to hypoxia and anoxia in shallow coastal seas: a realistic scenario. Marine Ecology Progress Series, 458: 39–52, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m458p039.pdf.

Staby:2011:FLD

[SA11] A. Staby and D. L. Aksnes. Follow the light—diurnal and seasonal variations in vertical distribution of the mesopelagic fish Maurolicus muelleri. Marine Ecology Progress Series, 422:265–273, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/422/m422p265.pdf.

Schwab:2014:SSM

[SA14] D. B. Schwab and J. D. Allen. Size-specific maternal effects in response to predator cues in an intertidal snail. Marine Ecology Progress Series, 499:127–141, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/499/m499p127.pdf.

Suprenand:2017:TEC

[SA17] P. M. Suprenand and C. H. Ainsworth. Trophodynamic effects of climate change-induced alterations to primary production along the western Antarctic Peninsula. Marine Ecology Progress Series, 569:37–54, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/569/m569p037.pdf.

Soldatini:2016:FOW

[SABMG16] C. Soldatini, Y. V. Albores-Barajas, B. Massa, and O. Gimenez. Forecasting ocean warming impacts on seabird demography: a case study on the European storm petrel. Marine Ecology Progress Series, 552:255–269, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/552/m552p255.pdf.

Suchy:2016:STV

[SAD^{+}16] K. D. Suchy, T. R. Avila, J. F. Dower, A. Bianchini, and G. M. Figueiredo. Short-term variability in chitobiase-based planktonic crustacean production rates in a highly eutrophic tropical estuary. Marine Ecology Progress Series, 545:77– 89, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/545/m545p077.pdf.

Suzuki:2011:RLS

[SAH11] G. Suzuki, S. Arakaki, and T. Hayashibara. Rapid in situ settlement following spawning by Acropora corals at ishigaki, southern Japan. Marine Ecology Progress Series, 421:131– 138, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/421/m421p131.pdf.

Santos:2012:SSD

[SAL^{+}12] S. Santos, G. Aarts, P. C. Luttikhuizen, J. Campos, T. Piersma, and H. W. van der Veer. Site-specific distribution of the bivalve Scrobicularia plana along the European coast. Marine Ecology Progress Series, 471:123–134, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/471/
m471p123.pdf.

Steneck:2014:EMF

[SAM14] R. S. Steneck, S. N. Arnold, and P. J. Mumby. Experiment mimics fishing on parrotfish: insights on coral reef recovery and alternative attractors. Marine Ecology Progress Series, 506:115–127, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/506/m506p115.pdf.

Siegel:2018:CMR

[SAM18] J. E. Siegel, M. D. Adkison, and M. V. McPhee. Changing maturation reaction norms and the effects of growth history in Alaskan Chinook salmon. Marine Ecology Progress Series, 595:187–202, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m595p187.pdf.

Smith:2014:EMR

[SAMW14] A. N. H. Smith, M. J. Anderson, R. B. Millar, and T. J. Willis. Effects of marine reserves in the context of spatial and temporal variation: an analysis using Bayesian zeroinflated mixed models. Marine Ecology Progress Series, 499: 203–216, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/499/m499p203.pdf.

Satoh:2010:HVD

[Sat10] K. Satoh. Horizontal and vertical distribution of larvae of Pacific bluefin tuna Thunnus orientalis in patches entrained in mesoscale eddies. Marine Ecology Progress Series, 404: 227–240, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/404/m404p227.pdf.

Sato:2017:SCR

[Sat17] N. Sato. Seasonal changes in reproductive traits and paternity in the Japanese pygmy squid Idiosepius paradoxus. Marine Ecology Progress Series, 582:121–131, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/582/
m582p121.pdf.

Schmiing:2013:PHM

[SATS13] M. Schmiing, P. Afonso, F. Tempera, and R. S. Santos. Predictive habitat modelling of reef fishes with contrasting trophic ecologies. Marine Ecology Progress Series, 474:201–216, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m474p201.pdf.

Salas:2018:PRA

[SAWK18] A. K. Salas, A. H. Altieri, P. S. Wilson, and T. H. Keitt. Predicting the reef acoustic cuescape from the perspective of larval fishes across a habitat quality gradient. Marine Ecology Progress Series, 605:173–193, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/605/m605p173.pdf.

Sanvicente-Anorve:2014:CPA

[SAZHAAHS14] L. Sanvicente-Añorve, J. Zavala-Hidalgo, M. E. Allende-Arandía, and M. Hermoso-Salazar. Connectivity patterns among coral reef systems in the southern Gulf of Mexico. Marine Ecology Progress Series, 498:27–41, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/498/
m498p027.pdf.

Santangelo:2010:QDC

[SB10] G. Santangelo and L. Bramanti. Quantifying the decline in Corallium rubrum populations. Marine Ecology Progress Series, 418:295–297, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/418/m418p295.pdf.

Szabo:2014:LSS

[SB14] A. R. Szabo and H. P. Batchelder. Late spring and summer patterns of euphausiid reproduction in Southeast Alaska fjord waters. Marine Ecology Progress Series, 516:153–161, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2014/516/m516p153.pdf.

Sackmann:2015:BMD

[SB15a] B. S. Sackmann and D. S. Becker. Bird mortality due to the Deepwater Horizon oil spill: Comment on Haney et al. (2014a,b). Marine Ecology Progress Series, 534:273–277, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_

oa/m534p273.pdf. See [HGS14a, HGS14b] and reply [HGS15].

Swain:2015:EIN

[SB15b] D. P. Swain and H. P. Benoît. Extreme increases in natural mortality prevent recovery of collapsed fish populations in a Northwest Atlantic ecosystem. Marine Ecology Progress Series, 519:165–182, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/519/m519p165.pdf.

Shahrestani:2018:SSC

[SB18] S. Shahrestani and H. Bi. Settlement and survival of Chrysaora chesapeakei polyps: implications for adult abundance. Marine Ecology Progress Series, 601:139–151, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/601/
m601p139.pdf.

Skinner:2012:IEM

[SBA12] J. P. Skinner, V. N. Burkanov, and R. D. Andrews. Influence of environment, morphology, and instrument size on lactating northern fur seal Callorhinus ursinus foraging behavior on the Lovushki Islands, Russia. Marine Ecology Progress Series, 471:293–308, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/471/m471p293.pdf.

Shamblin:2012:EMC

[SBB^{+}12a] B. M. Shamblin, A. B. Bolten, K. A. Bjorndal, P. H. Dutton, J. T. Nielsen, F. A. Abreu-Grobois, K. J. Reich, B. E. Witherington, D. A. Bagley, L. M. Ehrhart, A. D. Tucker, D. S. Addison, A. Arenas, C. Johnson, R. R. Carthy, M. M. Lamont, M. G. Dodd, M. S. Gaines, E. LaCasella, and C. J. Nairn. Expanded mitochondrial control region sequences increase resolution of stock structure among North Atlantic loggerhead turtle rookeries. Marine Ecology Progress Series, 469:145–160, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/469/m469p145.pdf.

Stauss:2012:SSF

[SBB^{+}12b] C. Stauss, S. Bearhop, T. W. Bodey, S. Garthe, C. Gunn, W. J. Grecian, R. Inger, M. E. Knight, J. Newton, S. C. Patrick, R. A. Phillips, J. J. Waggitt, and S. C. Votier. Sexspecific foraging behaviour in northern gannets Morus bassanus: incidence and implications. Marine Ecology Progress Series, 457:151–162, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/457/m457p151.pdf.

Smith:2014:SYE

[SBB^{+}14] C. R. Smith, A. F. Bernardino, A. Baco, A. Hannides, and I. Altamira. Seven-year enrichment: macrofaunal succession in deep-sea sediments around a 30 tonne whale fall in the Northeast Pacific. Marine Ecology Progress Series, 515: 133–149, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/515/m515p133.pdf.

Sato:2017:SVD

[SBB17] M. Sato and K. J. Benoit-Bird. Spatial variability of deep scattering layers shapes the Bahamian mesopelagic ecosystem. Marine Ecology Progress Series, 580:69–82, 2017. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m580p069.pdf.

Seymour:2018:LCR

[SBB18] J. R. Seymour, T. A. Barbasch, and P. M. Buston. Lunar cycles of reproduction in the clown anemonefish Amphiprion percula: individual-level strategies and population-level patterns. Marine Ecology Progress Series, 594:193–201, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/594/
m594p193.pdf.

Sato:2018:CUF

[SBBB^{+}18] M. Sato, J. A. Barth, K. J. Benoit-Bird, S. D. Pierce, T. J. Cowles, R. D. Brodeur, and W. T. Peterson. Coastal upwelling fronts as a boundary for planktivorous fish distributions. Marine Ecology Progress Series, 595:171–186, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/595/
m595p171.pdf.

Sheehan:2015:TRD

[SBCA15] E. V. Sheehan, D. Bridger, S. L. Cousens, and M. J. Attrill. Testing the resilience of dead maerl infaunal assemblages to the experimental removal and re-lay of habitat. Marine Ecology Progress Series, 535:117–128, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/535/m535p117.pdf.

Skomal:2017:MWS

[SBCT17] G. B. Skomal, C. D. Braun, J. H. Chisholm, and S. R. Thorrold. Movements of the white shark Carcharodon carcharias in the North Atlantic Ocean. Marine Ecology Progress Series, 580:1–16, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Schmidt:2014:NSS

[SBE^{+}14] A. E. Schmidt, L. W. Botsford, J. M. Eadie, R. W. Bradley, E. Di Lorenzo, and J. Jahncke. Non-stationary seabird responses reveal shifting ENSO dynamics in the northeast Pacific. Marine Ecology Progress Series, 499:249–258, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2014/499/m499p249.pdf.

Smale:2016:LEV

[SBE^{+}16] D. A. Smale, M. T. Burrows, A. J. Evans, N. King, M. D. J. Sayer, A. L. E. Yunnie, and P. J. Moore. Linking environmental variables with regional-scale variability in ecological structure and standing stock of carbon within UK kelp forests. Marine Ecology Progress Series, 542:79–95, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/542/
m542p079.pdf.

Spinette:2019:DAS

[SBE^{+}19] R. F. Spinette, S. M. Brown, A. L. Ehrlich, G. Puggioni, C. Deacutis, and B. D. Jenkins. Diazotroph activity in surface Narragansett Bay sediments in summer is stimulated by hypoxia and organic matter delivery. Marine Ecology Progress Series, 614:35–50, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m614p035.pdf.

Semmens:2010:STU

[SBFP10] J. M. Semmens, C. D. Buxton, E. Forbes, and M. J. Phelan. Spatial and temporal use of spawning aggregation sites by the tropical sciaenid Protonibea diacanthus. Marine Ecology Progress Series, 403:193–203, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/403/m403p193.pdf.

Schaffeld:2016:DSP

[SBG^{+}16] T. Schaffeld, S. Bräger, A. Gallus, M. Dähne, K. Krügel, A. Herrmann, M. Jabbusch, T. Ruf, U. K. Verfuß, H. Benke, and J. C. Koblitz. Diel and seasonal patterns in acoustic presence and foraging behaviour of free-ranging harbour porpoises. Marine Ecology Progress Series, 547:257–272, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m547p257.pdf.

Sokolova:2018:SPI

[SBH^{+}18] M. Sokolova, K. Buchmann, B. Huwer, P. W. Kania, U. Krumme, A. Galatius, J. Hemmer-Hansen, and J. W. Behrens. Spatial patterns in infection of cod Gadus morhua with the seal-associated liver worm Contracaecum osculatum from the Skagerrak to the central Baltic Sea. Marine Ecology Progress Series, 606:105–118, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m606p105.pdf.

Scheschonk:2019:AKE

[SBH^{+}19] L. Scheschonk, S. Becker, J. H. Hehemann, N. Diehl, U. Karsten, and K. Bischof. Arctic kelp eco-physiology during the polar night in the face of global warming: a crucial role for laminarin. Marine Ecology Progress Series, 611: 59–74, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/611/m611p059.pdf.

Solow:2014:TSS

[SBKA14] A. R. Solow, A. R. Beet, B. A. Keafer, and D. M. Anderson. Testing for simple structure in a spatial time series with an application to the distribution of Alexandrium resting cysts in the Gulf of Maine. Marine Ecology Progress Series, 501:291–296, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/501/m501p291.pdf.

Stentiford:2010:EAL

[SBL^{+}10] G. D. Stentiford, J. P. Bignell, B. P. Lyons, J. E. Thain, and S. W. Feist. Effect of age on liver pathology and other diseases in flatfish: implications for assessment of marine ecological health status. Marine Ecology Progress Series, 411: 215–230, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/411/m411p215.pdf.

Stallings:2014:CTS

[SBLM14] C. D. Stallings, J. P. Brower, J. M. Heinlein Loch, and A. Mickle. Commercial trawling in seagrass beds: bycatch and long-term trends in effort of a major shrimp fishery. Marine Ecology Progress Series, 513:143–153, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/513/
m513p143.pdf.

Sievers:2016:IHS

[SBM^{+}16] K. T. Sievers, R. J. Barr, J. M. Maloney, N. W. Driscoll, and T. W. Anderson. Impact of habitat structure on fish populations in kelp forests at a seascape scale. Marine Ecology Progress Series, 557:51–63, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/557/m557p051.pdf.

Stasko:2018:BPT

[SBM^{+}18] A. D. Stasko, B. A. Bluhm, C. Michel, P. Archambault, A. Majewski, J. D. Reist, H. Swanson, and M. Power. Benthicpelagic trophic coupling in an Arctic marine food web along vertical water mass and organic matter gradients. Marine Ecology Progress Series, 594:1–19, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Svensson:2012:TMP

[SBMÅ12] C. J. Svensson, S. Baden, P. O. Moksnes, and P. Åberg. Temporal mismatches in predator-herbivore abundance control algal blooms in nutrient-enriched seagrass ecosystems. Marine Ecology Progress Series, 471:61–71, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/471/
m471p061.pdf.

Swain:2018:PIC

[SBO^{+}18] T. D. Swain, E. C. Bold, P. C. Osborn, A. H. Baird, M. W. Westneat, V. Backman, and L. A. Marcelino. Physiological integration of coral colonies is correlated with bleaching resistance. Marine Ecology Progress Series, 586:1–10, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Sutherland:2010:RAS

[SBS10] K. R. Sutherland, A. R. Beet, and A. R. Solow. Re-analysis of a salp population time-series. Marine Ecology Progress Series, 418:147–150, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/418/m418p147.pdf.

Sipler:2013:TDD

[SBS^{+}13] R. E. Sipler, D. A. Bronk, S. P. Seitzinger, R. J. Lauck, L. R. McGuinness, G. J. Kirkpatrick, C. A. Heil, L. J. Kerkhof, and O. M. Schofield. Trichodesmium-derived dissolved organic matter is a source of nitrogen capable of supporting the growth of toxic red tide Karenia brevis. Marine Ecology Progress Series, 483:31–45, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m483p031.pdf.

Scheffer:2012:FZT

[SBT12] A. Scheffer, C. A. Bost, and P. N. Trathan. Frontal zones, temperature gradient and depth characterize the foraging habitat of king penguins at South Georgia. Marine Ecology Progress Series, 465:281–297, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m465p281.pdf.

Smith:2015:MCD

[SBT15] K. E. Smith, A. Brown, and S. Thatje. The metabolic cost of developing under hydrostatic pressure: experimental evidence supports macroecological pattern. Marine Ecology Progress Series, 524:71–82, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/524/m524p071.pdf.

Saporiti:2015:LCS

[SBV^{+}15] F. Saporiti, S. Bearhop, D. G. Vales, L. Silva, L. Zenteno, M. Tavares, E. A. Crespo, and L. Cardona. Latitudinal changes in the structure of marine food webs in the southwestern Atlantic Ocean. Marine Ecology Progress Series, 538:23–34, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/538/m538p023.pdf.

Sweet:2016:NDO

[SBW16] M. Sweet, M. Bulling, and J. E. Williamson. New disease outbreak affects two dominant sea urchin species associated with Australian temperate reefs. Marine Ecology Progress Series, 551:171–183, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/551/m551p171.pdf.

Satterthwaite:2014:MMD

[SCAM^{+}14] W. H. Satterthwaite, S. M. Carlson, S. D. Allen-Moran, S. Vincenzi, S. J. Bograd, and B. K. Wells. Match-mismatch dynamics and the relationship between ocean-entry timing and relative ocean recoveries of central valley fall run Chinook salmon. Marine Ecology Progress Series, 511:237–248, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m511p237.pdf.

Shannon:2014:TLB

[SCB^{+}14] L. Shannon, M. Coll, A. Bundy, D. Gascuel, J. J. Heymans, K. Kleisner, C. P. Lynam, C. Piroddi, J. Tam, M. Travers-Trolet, and Y. Shin. Trophic level-based indicators to track fishing impacts across marine ecosystems. Marine Ecology Progress Series, 512:115–140, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m512p115.pdf.

Sandoval-Castillo:2015:MSI

[SCB15] J. Sandoval-Castillo and L. B. Beheregaray. Metapopulation structure informs conservation management in a heavily exploited coastal shark (Mustelus henlei). Marine Ecology Progress Series, 533:191–203, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/533/m533p191.pdf.

Senigaglia:2016:MAW

[SCB^{+}16] V. Senigaglia, F. Christiansen, L. Bejder, D. Gendron, D. Lundquist, D. P. Noren, A. Schaffar, J. C. Smith, R. Williams, E. Martinez, K. Stockin, and D. Lusseau. Metaanalyses of whale-watching impact studies: comparisons of cetacean responses to disturbance. Marine Ecology Progress Series, 542:251–263, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2015/542/m542p251.pdf.

Santelli:2017:SPM

[SCD^{+}17] A. Santelli, I. Cvitković, M. Despalatović, G. Fabi, F. Grati, B. Marčeta, E. Punzo, S. Raicevich, P. Strafella, A. Spagnolo, A. N. Tassetti, and G. Scarcella. Spatial persistence of megazoobenthic assemblages in the Adriatic Sea. Marine Ecology Progress Series, 566:31–48, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/566/m566p031.pdf.

Stump:2017:HHE

[SCF^{+}17] K. L. Stump, C. J. Crooks, M. D. Fitchett, S. H. Gruber, and T. L. Guttridge. Hunted hunters: an experimental test of the effects of predation risk on juvenile lemon shark habitat use. Marine Ecology Progress Series, 574:85–95, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/574/
m574p085.pdf.

Stokesbury:2011:DMP

[SCHO11a] K. D. E. Stokesbury, J. D. Carey, B. P. Harris, and C. E. O’Keefe. Discard mortality played a major role in the loss of 10 billion juvenile scallops in the Mid-Atlantic Bight: Reply to Hart & Shank (2011). Marine Ecology Progress Series, 443:299–302, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/443/m443p299.pdf. See

[SCHO11b, HS11a].

Stokesbury:2011:IFM

[SCHO11b] Kevin D. E. Stokesbury, Jonathan D. Carey, Bradley P. Harris, and Catherine E. O’Keefe. Incidental fishing mortality may be responsible for the death of ten billion juvenile sea scallops in the mid-Atlantic. Marine Ecology Progress Series, 425:167–173, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/425/m425p167.pdf. See comment

[HS11a] and reply [SCHO11a].

Shinomiya:2017:IPS

[SCK^{+}17] Y. Shinomiya, S. Chiba, M. Kanamori, S. Hashizume, K. Yoshino, and S. Goshima. Importance of patch size variation for the population persistence of a decapod crustacean in seagrass beds. Marine Ecology Progress Series, 570:157–171, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/570/m570p157.pdf.

Smith:2013:ECB

[SCL13] B. E. Smith, J. S. Collie, and N. L. Lengyel. Effects of chronic bottom fishing on the benthic epifauna and diets of demersal fishes on northern Georges Bank. Marine Ecology Progress Series, 472:199–217, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/472/m472p199.pdf.

Smith:2016:RRV

[SCL^{+}16] A. M. Smith, D. E. Clark, M. D. Lamare, D. J. Winter, and M. Byrne. Risk and resilience: variations in magnesium in echinoid skeletal calcite. Marine Ecology Progress Series, 561:1–16, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/feature/m561p001.pdf.

Schotanus:2019:MSH

[SCL^{+}19] J. Schotanus, J. J. Capelle, L. Leuchter, J. van de Koppel, and T. J. Bouma. Mussel seed is highly plastic to settling conditions: the influence of waves versus tidal emergence.

Marine Ecology Progress Series, 624:77–87, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/624/
m624p077.pdf.

Silverstein:2011:NAS

[SCLB11] R. N. Silverstein, A. M. S. Correa, T. C. LaJeunesse, and A. C. Baker. Novel algal symbiont (Symbiodinium spp.) diversity in reef corals of Western Australia. Marine Ecology Progress Series, 422:63–75, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/422/m422p063.pdf.

Skinner:2013:RDP

[SCM13] M. A. Skinner, S. C. Courtenay, and C. W. McKindsey. Reductions in distribution, photosynthesis, and productivity of eelgrass Zostera marina associated with oyster Crassostrea virginica aquaculture. Marine Ecology Progress Series, 486:105–119, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/486/m486p105.pdf.

Sanchez-Camara:2011:DVU

[SCMSB^{+}11] J. Sanchez-Camara, K. Martin-Smith, D. J. Booth, J. Fritschi, and X. Turon. Demographics and vulnerability of a unique Australian fish, the weedy seadragon Phyllopteryx taeniolatus. Marine Ecology Progress Series, 422:253–264, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/422/
m422p253.pdf.

Santos:2012:IDL

[SCO^{+}12a] S. Santos, C. Cruzeiro, J. L. Olsen, H. W. van der Veer, and P. C. Luttikhuizen. Isolation by distance and low connectivity in the peppery furrow shell Scrobicularia plana (Bivalvia). Marine Ecology Progress Series, 462:111–124, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/462/
m462p111.pdf.

Scott:2012:PPA

[SCO^{+}12b] M. D. Scott, S. J. Chivers, R. J. Olson, P. C. Fiedler, and K. Holland. Pelagic predator associations: tuna and dolphins in the eastern tropical Pacific Ocean. Marine Ecology Progress Series, 458:283–302, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m458p283.pdf.

Segura:2014:DES

[SCP^{+}14] C. J. Segura, O. R. Chaparro, J. A. Pechenik, K. A. Paschke, S. J. A. Osores, J. M. Navarro, and V. M. Cubillos. Delayed effects of severe hypoxia experienced by marine gastropod embryos. Marine Ecology Progress Series, 510:59– 71, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/510/m510p059.pdf.

Siegelman-Charbit:2016:AMF

[SCP16] L. Siegelman-Charbit and B. Planque. Abundant mesopelagic fauna at oceanic high latitudes. Marine Ecology Progress Series, 546:277–282, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m546p277.pdf.

Siegel:2011:NCP

[SCRGP11] A. Siegel, B. Cotti-Rausch, D. I. Greenfield, and J. L. Pinckney. Nutrient controls of planktonic cyanobacteria biomass in coastal stormwater detention ponds. Marine Ecology Progress Series, 434:15–27, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/434/m434p015.pdf.

Schmidt:2011:ESS

[SCRL11] A. L. Schmidt, M. Coll, T. Romanuk, and H. K. Lotze. Ecosystem structure and services in eelgrass Zostera marina and rockweed Ascophyllum nodosum habitats. Marine Ecology Progress Series, 437:51–68, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/437/m437p051.pdf.

Shen:2016:OSV

[SCSC16] S. G. Shen, F. Chen, D. E. Schoppik, and D. M. Checkley, Jr. Otolith size and the vestibulo-ocular reflex of larvae of white seabass Atractoscion nobilis at high pCO_{2}. Marine Ecology Progress Series, 553:173–183, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m553p173.pdf.

Saunders:2017:SOM

[SCST17] R. A. Saunders, M. A. Collins, G. Stowasser, and G. A. Tarling. Southern Ocean mesopelagic fish communities in the Scotia Sea are sustained by mass immigration. Marine Ecology Progress Series, 569:173–185, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/569/m569p173.pdf.

Sheehan:2010:CTR

[SCTA10] E. V. Sheehan, R. A. Coleman, R. C. Thompson, and M. J. Attrill. Crab-tiling reduces the diversity of estuarine infauna. Marine Ecology Progress Series, 411:137–148, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/411/
m411p137.pdf.

Saunders:2015:PIM

[SCW^{+}15a] R. A. Saunders, M. A. Collins, P. Ward, G. Stowasser, S. L. Hill, R. Shreeve, and G. A. Tarling. Predatory impact of the myctophid fish community on zooplankton in the Scotia Sea (Southern Ocean). Marine Ecology Progress Series, 541:45–64, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/541/m541p045.pdf.

Stricker:2015:SCN

[SCW^{+}15b] C. A. Stricker, A. M. Christ, M. B. Wunder, A. C. Doll, S. D. Farley, L. D. Rea, D. A. S. Rosen, R. D. Scherer, and D. J. Tollit. Stable carbon and nitrogen isotope trophic enrichment factors for Steller sea lion vibrissae relative to milk and fish/invertebrate diets. Marine Ecology Progress Series, 523:255–266, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/523/m523p255.pdf.

Sztukowski:2018:SDI

[SCW^{+}18] L. A. Sztukowski, P. A. Cotton, H. Weimerskirch, D. R. Thompson, L. G. Torres, P. M. Sagar, A. M. Knights, A. L. Fayet, and S. C. Votier. Sex differences in individual foraging site fidelity of Campbell albatross. Marine Ecology Progress Series, 601:227–238, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/601/m601p227.pdf.

Shimeta:2012:IBA

[SCWV12] J. Shimeta, J. Cutajar, M. G. Watson, and T. Vlamis. Influences of biofilm-associated ciliates on the settlement of marine invertebrate larvae. Marine Ecology Progress Series, 449:1–12, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Schubert:2017:PPM

[SD17a] N. Schubert and K. Demes. Phenotypic plasticity in the marine angiosperm Halophila decipiens (Hydrocharitaceae, Streptophyta). Marine Ecology Progress Series, 575:81–93, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2017/575/m575p081.pdf.

Stieglitz:2017:GFC

[SD17b] T. C. Stieglitz and A. M. Dujon. A groundwater-fed coastal inlet as habitat for the Caribbean queen conch Lobatus gigas — an acoustic telemetry and space use analysis. Marine Ecology Progress Series, 571:139–152, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m571p139.pdf.

Smart:2012:ATS

[SDAH12] T. I. Smart, J. T. Duffy-Anderson, and J. K. Horne. Alternating temperature states influence walleye pollock early life stages in the southeastern Bering Sea. Marine Ecology Progress Series, 455:257–267, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/455/m455p257.pdf.

Siddon:2011:CLR

[SDAM11] E. C. Siddon, J. T. Duffy-Anderson, and F. J. Mueter. Community-level response of fish larvae to environmental variability in the southeastern Bering Sea. Marine Ecology Progress Series, 426:225–239, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/426/m426p225.pdf.

Stoner:2012:APS

[SDB12] A. W. Stoner, M. H. Davis, and C. J. Booker. Abundance and population structure of queen conch inside and outside a marine protected area: repeat surveys show significant declines. Marine Ecology Progress Series, 460:101–114, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/460/
m460p101.pdf.

Sevilgen:2014:OBS

[SdBAH^{+}14] D. S. Sevilgen, D. de Beer, A. Y. Al-Handal, T. Brey, and L. Polerecky. Oxygen budgets in subtidal Arctic (Kongsfjorden, Svalbard) and temperate (Helgoland, North Sea) microphytobenthic communities. Marine Ecology Progress Series, 504:27–42, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/504/m504p027.pdf.

Selkoe:2016:DSG

[SDC^{+}16] K. A. Selkoe, C. C. D’Aloia, E. D. Crandall, M. Iacchei, L. Liggins, J. B. Puritz, S. von der Heyden, and R. J. Toonen. A decade of seascape genetics: contributions to basic and applied marine connectivity. Marine Ecology Progress Series, 554:1–19, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Sidhu:2012:SOT

[SDCC12] L. A. Sidhu, P. Dann, L. Chambers, and E. A. Catchpole. Seasonal ocean temperature and the survival of first-year little penguins Eudyptula minor in south-eastern Australia. Marine Ecology Progress Series, 454:263–272, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m454p263.pdf.

Staudinger:2019:TED

[SDF^{+}19] M. D. Staudinger, V. H. Dimkovikj, C. A. M. France, E. Jorgensen, H. Judkins, A. Lindgren, E. K. Shea, and M. Vecchione. Trophic ecology of the deep-sea cephalopod assemblage near Bear Seamount in the Northwest Atlantic Ocean. Marine Ecology Progress Series, 629:67–86, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/629/
m629p067.pdf.

Sabarros:2012:DRT

[SDG^{+}12] P. S. Sabarros, J. M. Durant, D. Grémillet, R. J. M. Crawford, and N. C. Stenseth. Differential responses of three sympatric seabirds to spatio-temporal variability in shared resources. Marine Ecology Progress Series, 468:291–301, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/468/
m468p291.pdf.

Szostek:2013:EEL

[SDH13] C. L. Szostek, A. J. Davies, and H. Hinz. Effects of elevated levels of suspended particulate matter and burial on juvenile king scallops Pecten maximus. Marine Ecology Progress Series, 474:155–165, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/474/m474p155.pdf.

Sato:2013:SOS

[SDKD13] M. Sato, J. F. Dower, E. Kunze, and R. Dewey. Secondorder seasonal variability in diel vertical migration timing of euphausiids in a coastal inlet. Marine Ecology Progress Series, 480:39–56, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/480/m480p039.pdf.

Scopel:2019:VBR

[SDKS19] L. Scopel, A. Diamond, S. Kress, and P. Shannon. Varied breeding responses of seabirds to a regime shift in prey base in the Gulf of Maine. Marine Ecology Progress Series, 626:177–196, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m626p177.pdf.

Sailley:2013:CFP

[SDM^{+}13] S. F. Sailley, H. W. Ducklow, H. V. Moeller, W. R. Fraser, O. M. Schofield, D. K. Steinberg, L. M. Garzio, and S. C. Doney. Carbon fluxes and pelagic ecosystem dynamics near two western Antarctic Peninsula Adélie penguin colonies: an inverse model approach. Marine Ecology Progress Series, 492:253–272, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m492p253.pdf.

Sheaves:2016:BHM

[SDM16] M. Sheaves, L. Dingle, and C. Mattone. Biotic hotspots in mangrove-dominated estuaries: macro-invertebrate aggregation in unvegetated lower intertidal flats. Marine Ecology Progress Series, 556:31–43, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/556/m556p031.pdf.

Sakihara:2015:EGC

[SDMB15] T. S. Sakihara, B. D. Dudley, R. A. MacKenzie, and J. P. Beets. Endemic grazers control benthic microalgal growth in a eutrophic tropical brackish ecosystem. Marine Ecology Progress Series, 519:29–45, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/519/m519p029.pdf.

Sogaard:2019:EMP

[SDMR19] D. H. Søgaard, J. W. Deming, L. Meire, and S. Rysgaard. Effects of microbial processes and CaCO 3 dynamics on inorganic carbon cycling in snow-covered Arctic winter sea ice. Marine Ecology Progress Series, 611:31–44, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m611p031.pdf.

Scherer:2015:SIV

[SDR^{+}15] R. D. Scherer, A. C. Doll, L. D. Rea, A. M. Christ, C. A. Stricker, B. Witteveen, T. C. Kline, C. M. Kurle, and M. B. Wunder. Stable isotope values in pup vibrissae reveal geographic variation in diets of gestating Steller sea lions Eumetopias jubatus. Marine Ecology Progress Series, 527:261– 274, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/527/m527p261.pdf.

Schmiing:2014:AHW

[SDSA14] M. Schmiing, H. Diogo, R. S. Santos, and P. Afonso. Assessing hotspots within hotspots to conserve biodiversity and support fisheries management. Marine Ecology Progress Series, 513:187–199, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m513p187.pdf.

Sherman:2016:IPB

[SDST16] K. D. Sherman, C. P. Dahlgren, J. R. Stevens, and C. R. Tyler. Integrating population biology into conservation management for endangered Nassau grouper Epinephelus striatus. Marine Ecology Progress Series, 554:263–280, 2016. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m554p263.pdf.

Stier:2015:TIR

[SDT15] T. Stier, J. Drent, and D. W. Thieltges. Trematode infections reduce clearance rates and condition in blue mussels Mytilus edulis. Marine Ecology Progress Series, 529: 137–144, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/529/m529p137.pdf.

Suchy:2016:IVR

[SDVL16] K. D. Suchy, J. F. Dower, D. E. Varela, and M. G. Lagunas. Interannual variability in the relationship between in situ primary productivity and somatic crustacean productivity in a temperate fjord. Marine Ecology Progress Series, 545:91–108, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/545/m545p091.pdf.

Smith:2011:SSS

[SE11a] C. B. Smith and D. L. Erdner. Saxitoxin δ ^{15}N as a species-specific tracer of the source of nitrogen used by the toxin-producing dinoflagellate Alexandrium tamarense. Marine Ecology Progress Series, 443:19–27, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/443/
m443p019.pdf.

Smith:2011:SVE

[SE11b] Q. T. Smith and G. L. Eckert. Spatial variation and evidence of multiple transport pathways for Dungeness crab Cancer magister late-stage larvae in southeastern Alaska. Marine Ecology Progress Series, 429:185–196, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/429/m429p185.pdf.

Smiley:2016:MTA

[SE16] K. M. Smiley and L. M. Emmerson. A matter of timing: Adélie penguin reproductive success in a seasonally varying environment. Marine Ecology Progress Series, 542:235– 249, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/542/m542p235.pdf.

Shoji:2015:FBS

[SEF^{+}15] A. Shoji, K. Elliott, A. Fayet, D. Boyle, C. Perrins, and T. Guilford. Foraging behaviour of sympatric razorbills and puffins. Marine Ecology Progress Series, 520:257–267, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m520p257.pdf.

Southwell:2015:BME

[SEFS15] D. Southwell, L. Emmerson, J. Forcada, and C. Southwell. A bioenergetics model for estimating prey consumption by an Adélie penguin population in East Antarctica. Marine Ecology Progress Series, 526:183–197, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/526/m526p183.pdf.

Sandvik:2012:CAS

[SES12] H. Sandvik, K. E. Erikstad, and B. E. Sæther. Climate affects seabird population dynamics both via reproduction and adult survival. Marine Ecology Progress Series, 454:273–284, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m454p273.pdf.

Sherker:2017:PIS

[SES17] Z. T. Sherker, J. A. Ellrich, and R. A. Scrosati. Predatorinduced shell plasticity in mussels hinders predation by drilling snails. Marine Ecology Progress Series, 573:167– 175, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/573/m573p167.pdf.

Soler:2018:MBT

[SESS^{+}18] G. A. Soler, G. J. Edgar, R. D. Stuart-Smith, A. D. M. Smith, and R. J. Thomson. Moving beyond trophic groups: evaluating fishing-induced changes to temperate reef food webs. Marine Ecology Progress Series, 587:175–186, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/587/
m587p175.pdf.

Sumerel:2014:PSF

[SF14] A. N. Sumerel and C. M. Finelli. Particle size, flow speed, and body size interactions determine feeding rates of a solitary ascidian Styela plicata: a flume experiment. Marine Ecology Progress Series, 495:193–204, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/495/m495p193.pdf.

Schloesser:2016:TDC

[SF16] R. W. Schloesser and M. C. Fabrizio. Temporal dynamics of condition for estuarine fishes in their nursery habitats. Marine Ecology Progress Series, 557:207–219, 2016. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m557p207.pdf.

Specht:2018:TVE

[SF18] J. A. Specht and H. L. Fuchs. Thermal and viscous effects of temperature on Mercenaria mercenaria suspension feeding. Marine Ecology Progress Series, 589:129–140, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/589/

m589p129.pdf. See comment [RL18].

Smith:2015:UTR

[SFKQ15] J. M. Smith, K. L. Fresh, A. N. Kagley, and T. P. Quinn. Ultrasonic telemetry reveals seasonal variation in depth distribution and diel vertical migrations of sub-adult Chinook and Coho salmon in Puget Sound. Marine Ecology Progress Series, 532:227–242, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/532/m532p227.pdf.

Stortini:2018:STT

[SFL^{+}18] C. H. Stortini, K. T. Frank, W. C. Leggett, N. L. Shackell, and D. G. Boyce. Support for the trophic theory of island biogeography across submarine banks in a predatordepleted large marine ecosystem. Marine Ecology Progress Series, 607:155–169, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m607p155.pdf.

Scheibling:2010:DOA

[SFLG10] R. E. Scheibling, C. Feehan, and J. S. Lauzon-Guay. Disease outbreaks associated with recent hurricanes cause mass mortality of sea urchins in Nova Scotia. Marine Ecology Progress Series, 408:109–116, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/408/m408p109.pdf.

Silva:2019:HUT

[SFM^{+}19] T. L. Silva, G. Fay, T. A. Mooney, J. Robbins, M. T. Weinrich, C. D. Carson, T. V. N. Cole, M. A. Thompson, and D. N. Wiley. Habitat use of toothed whales in a marine protected area based on point process models. Marine Ecology Progress Series, 609:239–256, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/609/m609p239.pdf.

Speed:2010:CCS

[SFMB10] C. W. Speed, I. C. Field, M. G. Meekan, and C. J. A. Bradshaw. Complexities of coastal shark movements and their implications for management. Marine Ecology Progress Series, 408:275–293, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/408/m408p275.pdf.

Storch:2011:TTL

[SFNP11] D. Storch, M. Fernández, S. A. Navarrete, and H. O. Pörtner. Thermal tolerance of larval stages of the Chilean kelp crab Taliepus dentatus. Marine Ecology Progress Series, 429:157– 167, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m429p157.pdf.

Silva:2014:OSB

[SFTG14] L. Silva, A. M. Faria, M. A. Teodósio, and S. Garrido. Ontogeny of swimming behaviour in sardine Sardina pilchardus larvae and effect of larval nutritional condition on critical speed. Marine Ecology Progress Series, 504:287–300, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2014/504/m504p287.pdf.

Shelton:2014:HLS

[SFW^{+}14] A. O. Shelton, T. B. Francis, G. D. Williams, B. Feist, K. Stick, and P. S. Levin. Habitat limitation and spatial variation in Pacific herring egg survival. Marine Ecology Progress Series, 514:231–245, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/514/m514p231.pdf.

Steppe:2016:DSC

[SFW^{+}16] C. N. Steppe, D. W. Fredriksson, L. Wallendorf, M. Nikolov, and R. Mayer. Direct setting of Crassostrea virginica larvae in a tidal tributary: applications for shellfish restoration and aquaculture. Marine Ecology Progress Series, 546: 97–112, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/546/m546p097.pdf.

Smith:2012:EVD

[SG12] P. A. Smith and A. J. Gaston. Environmental variation and the demography and diet of thick-billed murres. Marine Ecology Progress Series, 454:237–249, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m454p237.pdf.

Seroy:2018:IPL

[SG18a] S. K. Seroy and D. Grünbaum. Individual and population level effects of ocean acidification on a predator–prey system with inducible defenses: bryozoan-nudibranch interactions in the Salish Sea. Marine Ecology Progress Series, 607:1–18, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Stevens:2018:IEA

[SG18b] A. M. Stevens and C. J. Gobler. Interactive effects of acidification, hypoxia, and thermal stress on growth, respiration, and survival of four North Atlantic bivalves. Marine Ecology Progress Series, 604:143–161, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m604p143.pdf.

Sulikowski:2010:UST

[SGB^{+}10] J. A. Sulikowski, B. Galuardi, W. Bubley, N. B. Furey, W. B. Driggers III, G. W. Ingram, Jr., and P. C. W. Tsang. Use of satellite tags to reveal the movements of spiny dogfish Squalus acanthias in the western North Atlantic Ocean. Marine Ecology Progress Series, 418:249–254, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/418/
m418p249.pdf.

Santos:2013:LTV

[SGC^{+}13] M. B. Santos, I. German, D. Correia, F. L. Read, J. Martinez Cedeira, M. Caldas, A. López, F. Velasco, and G. J. Pierce. Long-term variation in common dolphin diet in relation to prey abundance. Marine Ecology Progress Series, 481:249–268, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/481/m481p249.pdf.

Schiariti:2014:MHM

[SGC14] Agustín Schiariti, Gabriel Genzano, and John H. Costello. In memoriam: Hermes Mianzan (1957–2014). Marine Ecology Progress Series, 510:108–180, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m510p108.pdf.

Sigler:2017:AFS

[SGC17] M. F. Sigler, S. M. Gende, and D. J. Csepp. Association of foraging Steller sea lions with persistent prey hot spots in southeast Alaska. Marine Ecology Progress Series, 571: 233–243, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/571/m571p233.pdf.

Seear:2012:SSI

[SGCF^{+}12] P. J. Seear, W. P. Goodall-Copestake, A. H. Fleming, E. Rosato, and G. A. Tarling. Seasonal and spatial influences on gene expression in Antarctic krill Euphausia superba. Marine Ecology Progress Series, 467:61–75, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/467/
m467p061.pdf.

Sun:2012:ECT

[SGD^{+}12] Z. L. Sun, Q. F. Gao, S. L. Dong, P. K. S. Shin, and F. Wang. Estimates of carbon turnover rates in the sea cucumber Apostichopus japonicus (Selenka) using stable isotope analysis: the role of metabolism and growth. Marine Ecology Progress Series, 457:101–112, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/457/m457p101.pdf.

Slattery:2013:FCR

[SGEO13] M. Slattery, D. J. Gochfeld, C. G. Easson, and L. R. K. O’Donahue. Facilitation of coral reef biodiversity and health by cave sponge communities. Marine Ecology Progress Series, 476:71–86, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/476/m476p071.pdf.

Smith:2012:IHW

[SGG^{+}12] J. N. Smith, H. S. Grantham, N. Gales, M. C. Double, M. J. Noad, and D. Paton. Identification of humpback whale breeding and calving habitat in the Great Barrier Reef. Marine Ecology Progress Series, 447:259–272, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/447/
m447p259.pdf.

Smith:2015:SEM

[SGGL15] L. Smith, R. Gamble, S. Gaichas, and J. Link. Simulations to evaluate management trade-offs among marine mammal consumption needs, commercial fishing fleets and finfish biomass. Marine Ecology Progress Series, 523:215–232, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/523/
m523p215.pdf.

Skold:2018:ECB

[SGJ^{+}18] M. Sköld, P. Göransson, P. Jonsson, F. Bastardie, M. Blomqvist, S. Agrenius, J. G. Hiddink, H. C. Nilsson, and V. Bartolino. Effects of chronic bottom trawling on soft-seafloor macrofauna in the Kattegat. Marine Ecology Progress Series, 586:41–55, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/586/m586p041.pdf.

Schmidt:2017:IOW

[SGL^{+}17] M. Schmidt, G. Gerlach, E. Leo, K. L. Kunz, S. Swoboda, H. O. Pörtner, C. Bock, and D. Storch. Impact of ocean warming and acidification on the behaviour of two cooccurring gadid species, Boreogadus saida and Gadus morhua, from Svalbard. Marine Ecology Progress Series, 571:183– 191, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/571/m571p183.pdf.

Shapiera:2014:SSF

[SGM^{+}14] M. Shapiera, R. S. Gregory, C. J. Morris, C. J. Pennell, and P. V. R. Snelgrove. Season and site fidelity determine home range of dispersing and resident juvenile Greenland cod Gadus ogac in a Newfoundland fjord. Marine Ecology Progress Series, 503:235–246, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m503p235.pdf.

Sangra:2014:CBU

[SGMG^{+}14] P. Sangrà, C. García-Muñoz, C. M. García, Á. Marrero-Díaz, C. Sobrino, B. Mouriño-Carballido, B. Aguiar-González, C. Henríquez-Pastene, L. M. Rodríguez-Santana Á. Lubián, M. Hernández-Arencibia, S. Hernández-León, E. Vázquez, and S. N. Estrada-Allis. Coupling between upper ocean layer variability and size-fractionated phytoplankton in a non-nutrient-limited environment. Marine Ecology Progress Series, 499: 35–46, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m499p035.pdf.

Smyth:2013:OMB

[SGP13] A. R. Smyth, N. R. Geraldi, and M. F. Piehler. Oystermediated benthic-pelagic coupling modifies nitrogen pools and processes. Marine Ecology Progress Series, 493:23–30, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/493/m493p023.pdf.

Sirot:2017:UOO

[SGP^{+}17] C. Sirot, P. Grønkjær, J. B. Pedersen, J. Panfili, M. Zetina-Rejon, A. Tripp-Valdez, J. Ramos-Miranda, D. Flores-Hernandez, A. Sosa-Lopez, and A. M. Darnaude. Using otolith organic matter to detect diet shifts in Bardiella chrysoura, during a period of environmental changes. Marine Ecology Progress Series, 575:137–152, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/575/m575p137.pdf.

Stauffer:2012:OMB

[SGS^{+}12] B. A. Stauffer, A. G. Gellene, A. Schnetzer, E. L. Seubert, C. Oberg, G. S. Sukhatme, and D. A. Caron. An oceanographic, meteorological, and biological ‘perfect storm’ yields a massive fish kill. Marine Ecology Progress Series, 468:231–243, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/468/m468p231.pdf.

Sampaio:2017:PIC

[SGS^{+}17] E. Sampaio, F. Gallo, K. G. Schulz, E. B. Azevedo, and J. Barcelos e Ramos. Phytoplankton interactions can alter species response to present and future CO_{2} concentrations. Marine Ecology Progress Series, 575:31–42, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/575/
m575p031.pdf.

Shipley:2019:DRU

[SGS^{+}19] O. N. Shipley, A. J. Gallagher, D. S. Shiffman, L. Kaufman, and N. Hammerschlag. Diverse resource-use strategies in a large-bodied marine predator guild: evidence from differential use of resource subsidies and intraspecific isotopic variation. Marine Ecology Progress Series, 623:71–83, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/623/
m623p071.pdf.

Smith:2012:EHS

[SGT12] R. E. Smith, A. S. Grutter, and I. R. Tibbetts. Extreme habitat specialisation and population structure of two gorgonianassociated pygmy seahorses. Marine Ecology Progress Series, 444:195–206, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/444/m444p195.pdf.

Schismenou:2014:DEI

[SGT^{+}14] E. Schismenou, M. Giannoulaki, K. Tsiaras, E. Lefkaditou, G. Triantafyllou, and S. Somarakis. Disentangling the effects of inherent otolith growth and model-simulated ecosystem parameters on the daily growth rate of young anchovies. Marine Ecology Progress Series, 515:227–237, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m515p227.pdf.

Silbiger:2014:RSN

[SGTD14] N. J. Silbiger, Ò. Guadayol, F. I. M. Thomas, and M. J. Donahue. Reefs shift from net accretion to net erosion along a natural environmental gradient. Marine Ecology Progress Series, 515:33–44, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/515/m515p033.pdf.

Smyth:2016:BAE

[SGTP16] A. R. Smyth, N. R. Geraldi, S. P. Thompson, and M. F. Piehler. Biological activity exceeds biogenic structure in influencing sediment nitrogen cycling in experimental oyster reefs. Marine Ecology Progress Series, 560:173–183, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/560/
m560p173.pdf.

Sunda:2010:ETA

[SH10] W. G. Sunda and D. R. Hardison. Evolutionary tradeoffs among nutrient acquisition, cell size, and grazing defense in marine phytoplankton promote ecosystem stability.

Marine Ecology Progress Series, 401:63–76, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/401/

m401p063.pdf. See erratum [Ano10f].

Scott:2012:EAF

[SH12] R. D. Scott and J. Heikkonen. Estimating age at first maturity in fish from change-points in growth rate. Marine Ecology Progress Series, 450:147–157, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/450/m450p147.pdf.

Stephens:2014:MTG

[SH14] T. A. Stephens and C. D. Hepburn. Mass-transfer gradients across kelp beds influence Macrocystis pyrifera growth over small spatial scales. Marine Ecology Progress Series, 515: 97–109, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/515/m515p097.pdf.

Sun:2017:TGC

[SH17] X. Sun and D. Hedgecock. Temporal genetic change in North American Pacific oyster populations suggests caution in seascape genetics analyses of high gene-flow species. Marine Ecology Progress Series, 565:79–93, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/565/
m565p079.pdf.

Sjotun:2015:CEF

[SHAS15] K. Sjøtun, V. Husa, L. Asplin, and A. D. Sandvik. Climatic and environmental factors influencing occurrence and distribution of macroalgae — a fjord gradient revisited. Marine Ecology Progress Series, 532:73–88, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m532p073.pdf.

Silva:2010:UIZ

[SHB^{+}10a] A. C. F. Silva, S. J. Hawkins, D. M. Boaventura, E. Brewster, and R. C. Thompson. Use of the intertidal zone by mobile predators: influence of wave exposure, tidal phase and elevation on abundance and diet. Marine Ecology Progress Series, 406:197–210, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/406/m406p197.pdf.

Svetlichny:2010:VCE

[SHB^{+}10b] L. Svetlichny, E. Hubareva, G. Belmonte, M. Isinibilir, F. Ustun, I. N. Yilmaz, and B. Toklu-Alıclı. Vulnerability of copepod eggs in salinity and temperature gradients of the Marmara and Black Seas. Marine Ecology Progress Series, 419:109–120, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/419/m419p109.pdf.

Shillinger:2012:TTS

[SHB^{+}12] George L. Shillinger, Elliott L. Hazen, Helen Bailey, Steven J. Bograd, Brendan J. Godley, Rory P. Wilson, Philippe Gaspar, Mark Hamann, and James R. Spotila. Tagging through the stages: ontogeny in biologging. Marine Ecology Progress Series, 457:163–301, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m457_ThemeSection.pdf.

Sciberras:2013:BCR

[SHB^{+}13a] M. Sciberras, H. Hinz, J. D. Bennell, S. R. Jenkins, S. J. Hawkins, and M. J. Kaiser. Benthic community response to a scallop dredging closure within a dynamic seabed habitat. Marine Ecology Progress Series, 480:83–98, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/480/
m480p083.pdf.

Stige:2013:PFR

[SHB^{+}13b] L. C. Stige, M. E. Hunsicker, K. M. Bailey, N. A. Yaragina, and G. L. Hunt, Jr. Predicting fish recruitment from juvenile abundance and environmental indices. Marine Ecology Progress Series, 480:245–261, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m480p245.pdf.

Soria:2019:RSD

[SHB^{+}19] M. Soria, M. R. Heithaus, A. Blaison, E. Crochelet, F. Forget, and P. Chabanet. Residency and spatial distribution of bull sharks Carcharhinus leucas in and around Reunion Island marine protected area. Marine Ecology Progress Series, 630:101–113, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/630/m630p101.pdf.

Silva:2010:PFC

[SHC^{+}10] A. C. Silva, S. J. Hawkins, K. R. Clarke, D. M. Boaventura, and R. C. Thompson. Preferential feeding by the crab Necora puber on differing sizes of the intertidal limpet Patella vulgata. Marine Ecology Progress Series, 416:179–188, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/416/
m416p179.pdf.

Studholme:2019:YRT

[SHD^{+}19] K. R. Studholme, J. M. Hipfner, A. D. Domalik, S. J. Iverson, and G. T. Crossin. Year-round tracking reveals multiple migratory tactics in a sentinel North Pacific seabird, Cassin’s auklet. Marine Ecology Progress Series, 619:169–185, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/619/
m619p169.pdf.

Sinutok:2013:DPR

[SHDR13] S. Sinutok, R. Hill, M. A. Doblin, and P. J. Ralph. Diurnal photosynthetic response of the motile symbiotic benthic Foraminiferan Marginopora vertebralis. Marine Ecology Progress Series, 478:127–138, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/478/m478p127.pdf.

Sheaves:2012:EEA

[She12] Marcus Sheaves. Ecosystem equivalence and the ability to generalise: insights from global consistencies in mangrove fish assemblages. Marine Ecology Progress Series, 461:137–149, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2012/461/m461p137.pdf. See comment [CGK13a] and

reply [She13].

Sheaves:2013:MFA

[She13] M. Sheaves. Mangrove fish assemblages and ecological equivalence: Reply to Castellanos-Galindo & Krumme (2013). Marine Ecology Progress Series, 474:303–304, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/474/

m474p303.pdf. See [She12, CGK13a].

Shank:2012:PSP

[SHF12a] B. V. Shank, D. R. Hart, and K. D. Friedland. Postsettlement predation by sea stars and crabs on the sea scallop in the Mid-Atlantic Bight. Marine Ecology Progress Series, 468:161–177, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/468/m468p161.pdf.

Stewart:2012:MPM

[SHF^{+}12b] J. S. Stewart, E. L. Hazen, D. G. Foley, S. J. Bograd, and W. F. Gilly. Marine predator migration during range expansion: Humboldt squid Dosidicus gigas in the northern California Current system. Marine Ecology Progress Series, 471:135–150, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/471/m471p135.pdf.

Sumaila:2015:ESI

[SHG^{+}15] U. R. Sumaila, N. Hotte, A. Galli, V. W. Y. Lam, A. M. Cisneros-Montemayor, and M. Wackernagel. Eco^{2}: a simple index of economic–ecological deficits. Marine Ecology Progress Series, 530:271–279, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m530p271.pdf.

Shields:2010:MLS

[SHH10] J. L. Shields, J. W. Heath, and D. D. Heath. Marine landscape shapes hybrid zone in a broadcast spawning bivalve: introgression and genetic structure in Canadian west coast Mytilus. Marine Ecology Progress Series, 399:211–223, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/399/
m399p211.pdf.

Svetlichny:2010:STC

[SHI^{+}10] L. Svetlichny, E. Hubareva, M. Isinibilir, A. Kideys, G. Belmonte, and E. Giangrande. Salinity tolerance of Calanus euxinus in the Black and Marmara Seas. Marine Ecology Progress Series, 404:127–138, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/404/m404p127.pdf.

Saunders:2014:FID

[SHK14] B. J. Saunders, E. S. Harvey, and G. A. Kendrick. Factors influencing distribution and habitat associations in an endemic group of temperate Western Australian reef fishes over a latitudinal gradient. Marine Ecology Progress Series, 517:193–208, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/517/m517p193.pdf.

Schubert:2015:MME

[SHKR15] P. R. Schubert, W. Hukriede, R. Karez, and T. B. H. Reusch. Mapping and modeling eelgrass Zostera marina distribution in the western Baltic Sea. Marine Ecology Progress Series, 522:79–95, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/522/m522p079.pdf.

Shaw:2010:DSA

[SHM^{+}10a] P. W. Shaw, L. Hendrickson, N. J. McKeown, T. Stonier, M. J. Naud, and W. H. H. Sauer. Discrete spawning aggregations of loliginid squid do not represent genetically distinct populations. Marine Ecology Progress Series, 408:117–127, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2010/408/m408p117.pdf. See comment [GBH12] and re-

ply [SHM^{+}12].

Sun:2010:PPS

[SHM10b] Z. Sun, J. F. Hamel, and A. Mercier. Planulation periodicity, settlement preferences and growth of two deep-sea octocorals from the northwest Atlantic. Marine Ecology Progress Series, 410:71–87, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/410/m410p071.pdf.

Shaw:2012:PSS

[SHM^{+}12] P. W. Shaw, L. Hendrickson, N. J. McKeown, T. Stonier, M. J. Naud, and W. H. H. Sauer. Population structure of the squid Doryteuthis (Loligo) pealeii on the eastern coast of the USA: Reply to Gerlach et al. (2012). Marine Ecology Progress Series, 450:285–287, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/450/m450p285.pdf. See

[SHM^{+}10a, GBH12].

Solas:2015:EPR

[SHMB15] M. R. Solas, R. N. Hughes, F. Márquez, and A. Brante. Early plastic responses in the shell morphology of Acanthina monodon (Mollusca, Gastropoda) under predation risk and water turbulence. Marine Ecology Progress Series, 527:133– 142, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m527p133.pdf.

Shields:2017:NIT

[SHMOS17] J. D. Shields, J. P. Huchin-Mian, P. A. O’Leary, and H. J. Small. New insight into the transmission dynamics of the crustacean pathogen Hematodinium perezi (Dinoflagellata) using a novel sentinel methodology. Marine Ecology Progress Series, 573:73–84, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/573/m573p073.pdf.

Schiffer:2014:PHS

[SHP^{+}14] M. Schiffer, L. Harms, H. O. Pörtner, F. C. Mark, and D. Storch. Pre-hatching seawater pCO_{2} affects development and survival of zoea stages of Arctic spider crab Hyas araneus. Marine Ecology Progress Series, 501:127–139, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/501/
m501p127.pdf.

Schaub:2018:NUU

[SHP^{+}18] J. Schaub, B. P. V. Hunt, E. A. Pakhomov, K. Holmes, Y. Lu, and L. Quayle. NOTE: Using unmanned aerial vehicles (UAVs) to measure jellyfish aggregations. Marine Ecology Progress Series, 591:29–36, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/591/m591p029.pdf.

Sato:2016:IMH

[SHPS^{+}16] M. Sato, J. K. Horne, S. L. Parker-Stetter, T. E. Essington, J. E. Keister, P. E. Moriarty, L. Li, and J. Newton. Impacts of moderate hypoxia on fish and zooplankton prey distributions in a coastal fjord. Marine Ecology Progress Series, 560:57–72, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m560p057.pdf.

Schick:2011:CSP

[SHR^{+}11] R. S. Schick, P. N. Halpin, A. J. Read, D. L. Urban, B. D. Best, C. P. Good, J. J. Roberts, E. A. LaBrecque, C. Dunn, L. P. Garrison, K. D. Hyrenbach, W. A. McLellan, D. A. Pabst, D. L. Palka, and P. Stevick. Community structure in pelagic marine mammals at large spatial scales. Marine Ecology Progress Series, 434:165–181, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m434p165.pdf.

Santora:2017:IOC

[SHS^{+}17] J. A. Santora, E. L. Hazen, I. D. Schroeder, S. J. Bograd, K. M. Sakuma, and J. C. Field. Impacts of ocean climate variability on biodiversity of pelagic forage species in an upwelling ecosystem. Marine Ecology Progress Series, 580:205–220, 2017. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m580p205.pdf.

Suikkanen:2011:AEB

[SHSK11] S. Suikkanen, P. Hakanen, K. Spilling, and A. Kremp. Allelopathic effects of Baltic Sea spring bloom dinoflagellates on co-occurring phytoplankton. Marine Ecology Progress Series, 439:45–55, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/439/m439p045.pdf.

Smith:2013:HIF

[SHT^{+}13] C. E. Smith, B. J. Hurley, C. N. Toms, A. D. Mackey, M. Solangi, and S. A. Kuczaj II. Hurricane impacts on the foraging patterns of bottlenose dolphins Tursiops truncatus in Mississippi Sound. Marine Ecology Progress Series, 487:231–244, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m487p231.pdf.

Skov:2016:RTS

[SHT^{+}16] H. Skov, S. Heinänen, C. B. Thaxter, A. E. Williams, S. Lohier, and A. N. Banks. Real-time species distribution models for conservation and management of natural resources in marine environments. Marine Ecology Progress Series, 542:221–234, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m542p221.pdf.

Schultz:2018:NBD

[SHT^{+}18] H. Schultz, R. J. Hohnhold, G. A. Taylor, S. J. Bury, T. Bliss, S. M. H. Ismar, A. C. Gaskett, C. D. Millar, and T. E. Dennis. Non-breeding distribution and activity patterns in a temperate population of brown skua. Marine Ecology Progress Series, 603:215–226, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/603/m603p215.pdf.

Schlaff:2017:BEE

[SHUS17] A. M. Schlaff, M. R. Heupel, V. Udyawer, and C. A. Simpfendorfer. Biological and environmental effects on activity space of a common reef shark on an inshore reef. Marine Ecology Progress Series, 571:169–181, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/571/
m571p169.pdf.

Sugiyama:2010:RPD

[SIHH10] N. Sugiyama, T. Iseto, M. Hirose, and E. Hirose. Reproduction and population dynamics of the solitary entoproct Loxosomella plakorticola inhabiting a demosponge, Plakortis sp. Marine Ecology Progress Series, 415:73–82, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/415/
m415p073.pdf.

Slaughter:2016:PIA

[SIK16] A. M. Slaughter, T. R. Ignoffo, and W. Kimmerer. Predation impact of Acartiella sinensis, an introduced predatory copepod in the San Francisco Estuary, USA. Marine Ecology Progress Series, 547:47–60, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/547/m547p047.pdf.

Staudinger:2010:SBA

[SJ10] M. D. Staudinger and F. Juanes. A size-based approach to quantifying predation on longfin inshore squid Loligo pealeii in the northwest Atlantic. Marine Ecology Progress Series, 399:225–241, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/399/m399p225.pdf.

Solmundsson:2015:HRS

[SJB^{+}15] J. Sólmundsson, I. G. Jónsdóttir, B. Björnsson, S. Á. Ragnarsson, G. G. Tómasson, and V. Thorsteinsson. Home ranges and spatial segregation of cod Gadus morhua spawning components. Marine Ecology Progress Series, 520:217– 233, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/520/m520p217.pdf.

Sambrook:2016:LEC

[SJB16a] K. Sambrook, G. P. Jones, and M. C. Bonin. Life on the edge: Coral reef fishes exhibit strong responses to a habitat boundary. Marine Ecology Progress Series, 561:203– 215, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/561/m561p203.pdf.

Sheaves:2016:UMF

[SJB16b] M. Sheaves, R. Johnston, and R. Baker. Use of mangroves by fish: new insights from in-forest videos. Marine Ecology Progress Series, 549:167–182, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/549/m549p167.pdf.

Sheaves:2010:TDF

[SJC10] M. Sheaves, R. Johnston, and R. M. Connolly. Temporal dynamics of fish assemblages of natural and artificial tropical estuaries. Marine Ecology Progress Series, 410:143– 157, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/410/m410p143.pdf.

Suarez-Jimenez:2017:IKU

[SJHH^{+}17] R. Suárez-Jiménez, C. D. Hepburn, G. A. Hyndes, R. J. McLeod, R. B. Taylor, and C. L. Hurd. The invasive kelp Undaria pinnatifida hosts an epifaunal assemblage similar to native seaweeds with comparable morphologies. Marine Ecology Progress Series, 582:45–55, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/582/m582p045.pdf.

Svensson:2013:ESR

[SJL13] J. R. Svensson, L. Jonsson, and M. Lindegarth. Excessive spatial resolution decreases performance of quantitative models, contrary to expectations from error analyses. Marine Ecology Progress Series, 485:57–73, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/485/m485p057.pdf.

Stevens:2015:OHV

[SJL^{+}15] C. J. Stevens, S. K. Juniper, H. Limén, D. W. Pond, A. Metaxas, and Y. Gélinas. Obligate hydrothermal vent fauna at East Diamante submarine volcano (Mariana Arc) exploit photosynthetic and chemosynthetic carbon sources. Marine Ecology Progress Series, 525:25–39, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/525/
m525p025.pdf.

Shimada:2016:LTS

[SJL^{+}16] T. Shimada, R. Jones, C. Limpus, R. Groom, and M. Hamann. Long-term and seasonal patterns of sea turtle home ranges in warm coastal foraging habitats: implications for conservation. Marine Ecology Progress Series, 562:163–179, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/562/
m562p163.pdf.

Shimada:2012:IDR

[SJLH12] T. Shimada, R. Jones, C. Limpus, and M. Hamann. Improving data retention and home range estimates by data-driven screening. Marine Ecology Progress Series, 457:171–180, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m457p171.pdf.

Stevens:2016:BAL

[SJO16] J. L. Stevens, R. L. Jackson, and J. B. Olson. Bacteria associated with lionfish (Pterois volitans/miles complex) exhibit antibacterial activity against known fish pathogens. Marine Ecology Progress Series, 558:167–180, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m558p167.pdf.

Song:2011:IVP

[SJS^{+}11] H. Song, R. Ji, C. Stock, K. Kearney, and Z. Wang. Interannual variability in phytoplankton blooms and plankton productivity over the Nova Scotian shelf and in the Gulf of Maine. Marine Ecology Progress Series, 426:105–118, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/426/
m426p105.pdf.

Shiell:2010:ATC

[SK10] G. R. Shiell and B. Knott. Aggregations and temporal changes in the activity and bioturbation contribution of the sea cucumber Holothuria whitmaei (Echinodermata: Holothuroidea). Marine Ecology Progress Series, 415:127–139, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/415/
m415p127.pdf.

Sichlau:2011:ASD

[SK11] M. H. Sichlau and T. Kiørboe. Age- and size-dependent mating performance and fertility in a pelagic copepod, Temora longicornis. Marine Ecology Progress Series, 442:123–132, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2011/442/m442p123.pdf.

Sydeman:2012:TSS

[SK12] W. J. Sydeman and A. Kitaysky. Theme section: Seabirds and climate change. Marine Ecology Progress Series, 454: 105–307, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m454_ThemeSection.pdf.

Sams:2013:ERV

[SK13] M. A. Sams and M. J. Keough. Early recruitment variation and an established dominant alter the composition of a temperate fouling community. Marine Ecology Progress Series, 486:79–91, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/486/m486p079.pdf.

Sinclair:2014:IEG

[SKA^{+}14] E. A. Sinclair, S. L. Krauss, J. Anthony, R. Hovey, and G. A. Kendrick. The interaction of environment and genetic diversity within meadows of the seagrass Posidonia australis (Posidoniaceae). Marine Ecology Progress Series, 506: 87–98, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/506/m506p087.pdf.

Schuckel:2015:LLT

[SKB15] U. Schückel, I. Kröncke, and D. Baird. Linking long-term changes in trophic structure and function of an intertidal macrobenthic system to eutrophication and climate change using ecological network analysis. Marine Ecology Progress Series, 536:25–38, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/536/m536p025.pdf.

Satyanarayana:2011:LTM

[SKD^{+}11a] B. Satyanarayana, N. Koedam, K. De Smet, D. Di Nitto, M. Bauwens, L. P. Jayatissa, S. Cannicci, and F. Dahdouh-Guebas. Long-term mangrove forest development in Sri Lanka: early predictions evaluated against outcomes using VHR remote sensing and VHR ground-truth data. Marine Ecology Progress Series, 443:51–63, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/443/m443p051.pdf.

Swalethorp:2011:GEP

[SKD^{+}11b] R. Swalethorp, S. Kjellerup, M. Dünweber, T. G. Nielsen, E. F. Møller, S. Rysgaard, and B. W. Hansen. Grazing, egg production, and biochemical evidence of differences in the life strategies of Calanus finmarchicus, C. glacialis and C. hyperboreus in Disko Bay, western Greenland. Marine Ecology Progress Series, 429:125–144, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m429p125.pdf.

Sackett:2017:SDB

[SKD17a] D. K. Sackett, C. D. Kelley, and J. C. Drazen. Spilling over deepwater boundaries: evidence of spillover from two deepwater restricted fishing areas in Hawaii. Marine Ecology Progress Series, 568:175–190, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/568/m568p175.pdf.

Schaafsma:2017:STV

[SKD^{+}17b] F. L. Schaafsma, D. Kohlbach, C. David, B. A. Lange, M. Graeve, H. Flores, and J. A. van Franeker. Spatio-temporal variability in the winter diet of larval and juvenile Antarctic krill, Euphausia superba, in ice-covered waters. Marine Ecology Progress Series, 580:101–115, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m580p101.pdf.

Skubel:2018:PLT

[SKFH18] R. A. Skubel, B. P. Kirtman, C. Fallows, and N. Hammerschlag. Patterns of long-term climate variability and predation rates by a marine apex predator, the white shark Carcharodon carcharias. Marine Ecology Progress Series, 587: 129–139, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/587/m587p129.pdf.

Suzuki:2018:PSS

[SKFY18] G. Suzuki, S. Kai, Y. Fujikura, and H. Yamashita. Postsettlement survivorship of artificially supplied Acropora coral larvae in the Sekisei Lagoon. Marine Ecology Progress Series, 603:105–115, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/603/m603p105.pdf.

Sharir:2011:SSC

[SKGS11] Y. Sharir, D. Kerem, P. Gol’din, and E. Spanier. Small size in the common bottlenose dolphin Tursiops truncatus in the eastern Mediterranean: a possible case of Levantine nanism.

Marine Ecology Progress Series, 438:241–251, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/438/
m438p241.pdf.

Solberg:2012:CAS

[SKK12] I. Solberg, T. A. Klevjer, and S. Kaartvedt. Continuous acoustic studies of overwintering sprat Sprattus sprattus reveal flexible behavior. Marine Ecology Progress Series, 464: 245–256, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m464p245.pdf.

Sato:2018:IFL

[SKK^{+}18] T. Sato, M. Kobayashi, T. Kurihara, T. Takebe, N. Hirai, K. Okuzawa, S. Sawaguchi, T. Matsubara, Y. Akita, A. Ebisawa, I. Ohta, M. Uehara, T. Shimose, T. Yamaguchi, R. Shinoda, M. Koiso, and K. Teruya. Impact of fishing losses of males on the reproductive output of the large protogynous fish, Choerodon schoenleinii. Marine Ecology Progress Series, 602:225–235, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/602/m602p225.pdf.

Satterthwaite:2012:LCV

[SKM12] W. H. Satterthwaite, A. S. Kitaysky, and M. Mangel. Linking climate variability, productivity and stress to demography in a long-lived seabird. Marine Ecology Progress Series, 454:221–235, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m454p221.pdf.

Smolina:2015:CTR

[SKM^{+}15] I. Smolina, S. Kollias, E. F. Møller, P. Lindeque, A. Y. M. Sundaram, J. M. O. Fernandes, and G. Hoarau. Contrasting transcriptome response to thermal stress in two key zooplankton species, Calanus finmarchicus and C. glacialis. Marine Ecology Progress Series, 534:79–93, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m534p079.pdf.

Shirai:2018:RSH

[SKMS^{+}18] K. Shirai, F. Koyama, N. Murakami-Sugihara, K. Nanjo, T. Higuchi, H. Kohno, Y. Wananabe, K. Okamoto, and M. Sano. Reconstruction of the salinity history associated with movements of mangrove fishes using otolith oxygen isotopic analysis. Marine Ecology Progress Series, 593:127–139, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m593p127.pdf.

Skjaeraasen:2015:SSN

[SKN^{+}15] J. E. Skjæraasen, K. Korsbrekke, T. Nilsen, M. Fonn, O. S. Kjesbu, G. E. Dingsør, and R. D. M. Nash. Skipped spawning in Northeast Arctic haddock Melanogrammus aeglefinus. Marine Ecology Progress Series, 526:143–155, 2015. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m526p143.pdf.

Sugimoto:2017:PPP

[SKN^{+}17] R. Sugimoto, K. Kitagawa, S. Nishi, H. Honda, M. Yamada, S. Kobayashi, J. Shoji, S. Ohsawa, M. Taniguchi, and O. Tominaga. Phytoplankton primary productivity around submarine groundwater discharge in nearshore coasts. Marine Ecology Progress Series, 563:25–33, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m563p025.pdf.

Suryan:2016:TSS

[SKPS^{+}16] R. M. Suryan, K. J. Kuletz, S. L. Parker-Stetter, P. H. Ressler, M. Renner, J. K. Horne, E. V. Farley, and E. A. Labunski. Temporal shifts in seabird populations and spatial coherence with prey in the southeastern Bering Sea. Marine Ecology Progress Series, 549:199–215, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m549p199.pdf.

Sogaard:2010:AHA

[SKR^{+}10] D. H. Søgaard, M. Kristensen, S. Rysgaard, R. N. Glud, P. J. Hansen, and K. M. Hilligsøe. Autotrophic and heterotrophic activity in Arctic first-year sea ice: seasonal study from Malene Bight, SW Greenland. Marine Ecology Progress Series, 419:31–45, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m419p031.pdf.

Shelton:2013:AYV

[SKR^{+}13] A. O. Shelton, D. Kinzey, C. Reiss, S. Munch, G. Watters, and M. Mangel. Among-year variation in growth of Antarctic krill Euphausia superba based on length-frequency data. Marine Ecology Progress Series, 481:53–67, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/481/
m481p053.pdf.

Suzuki:2018:CTG

[SKS18] S. Suzuki, T. Kawai, and T. Sakamaki. Combination of trophic group habitat preferences determines coral reef fish assemblages. Marine Ecology Progress Series, 586:141–154, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2017/586/m586p141.pdf.

Salton:2019:MSB

[SKSH19] M. Salton, R. Kirkwood, D. Slip, and R. Harcourt. Mechanisms for sex-based segregation in foraging behaviour by a polygynous marine carnivore. Marine Ecology Progress Series, 624:213–226, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/624/m624p213.pdf.

Simmonds:2014:GSS

[SKW^{+}14] S. E. Simmonds, B. P. Kinlan, C. White, G. L. Paradis, R. R. Warner, and D. C. Zacherl. Geospatial statistics strengthen the ability of natural geochemical tags to estimate range-wide population connectivity in marine species. Marine Ecology Progress Series, 508:33–51, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/508/m508p033.pdf.

Schrandt:2011:HAI

[SL11a] M. N. Schrandt and S. C. Lema. Habitat-associated intraspecific variation in behavior and stress responses in a demersal coral reef fish. Marine Ecology Progress Series, 443:153–166, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/443/m443p153.pdf.

Seney:2011:MPI

[SL11b] E. E. Seney and A. M. Landry. Movement patterns of immature and adult female Kemp’s ridley sea turtles in the northwestern Gulf of Mexico. Marine Ecology Progress Series, 440:241–254, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/440/m440p241.pdf.

Slattery:2015:TES

[SL15a] M. Slattery and M. P. Lesser. Trophic ecology of sponges from shallow to mesophotic depths (3 to 150 m): Comment on Pawlik et al. (2015). Marine Ecology Progress Series, 527:275–279, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m527p275.pdf. See [PMEZ15b].

Sobocinski:2015:TTS

[SL15b] K. L. Sobocinski and R. J. Latour. Trophic transfer in seagrass systems: estimating seasonal production of an abundant seagrass fish, Bairdiella chrysoura, in lower Chesapeake Bay. Marine Ecology Progress Series, 523:157–174, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/523/
m523p157.pdf.

Soong:2011:SLI

[SLC11] K. Soong, Y. J. Lee, and I. H. Chang. Short-lived intertidal midge Pontomyia oceana have semilunar eclosion rhythm entrained by night light. Marine Ecology Progress Series, 433:121–130, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/433/m433p121.pdf.

Segura:2013:PTP

[SLD^{+}13] V. Segura, V. A. Lutz, A. Dogliotti, R. I. Silva, R. M. Negri, R. Akselman, and H. Benavides. Phytoplankton types and primary production in the Argentine Sea. Marine Ecology Progress Series, 491:15–31, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/491/m491p015.pdf.

Stewart:2013:IEP

[SLE^{+}13] A. R. Stewart, S. N. Luoma, K. A. Elrick, J. L. Carter, and M. van der Wegen. Influence of estuarine processes on spatiotemporal variation in bioavailable selenium. Marine Ecology Progress Series, 492:41–56, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m492p041.pdf.

Shipley:2019:TMB

[SLF^{+}19] O. N. Shipley, C. S. Lee, N. S. Fisher, G. Burruss, M. G. Frisk, E. J. Brooks, Z. C. Zuckerman, A. D. Herrmann, and D. J. Madigan. Trophodynamics and mercury bioaccumulation in reef and open-ocean fishes from the Bahamas with a focus on two teleost predators. Marine Ecology Progress Series, 608:221–232, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/608/m608p221.pdf.

Segovia:2018:PSR

[SLIGG18] M. Segovia, M. R. Lorenzo, C. Iñiguez, and C. García-Gómez. Physiological stress response associated with elevated CO_{2} and dissolved iron in a phytoplankton community dominated by the coccolithophore Emiliania huxleyi. Marine Ecology Progress Series, 586:73–89, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/586/m586p073.pdf.

Shiganova:2014:IBI

[SLKN14] T. A. Shiganova, L. Legendre, A. S. Kazmin, and P. Nival. Interactions between invasive ctenophores in the Black Sea: assessment of control mechanisms based on long-term observations. Marine Ecology Progress Series, 507:111–123, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2014/507/m507p111.pdf.

Sommer:2012:EIC

[SLL12] U. Sommer, K. Lengfellner, and A. Lewandowska. Experimental induction of a coastal spring bloom early in the year by intermittent high-light episodes. Marine Ecology Progress Series, 446:61–71, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/446/m446p061.pdf.

Selz:2018:DPA

[SLL^{+}18] V. Selz, K. E. Lowry, K. M. Lewis, H. L. Joy-Warren, W. van de Poll, S. Nirmel, A. Tong, and K. R. Arrigo. Distribution of Phaeocystis antarctica-dominated sea ice algal communities and their potential to seed phytoplankton across the western Antarctic Peninsula in spring. Marine Ecology Progress Series, 586:91–112, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/586/m586p091.pdf.

Stief:2018:FRC

[SLNG18] P. Stief, A. S. B. Lundgaard, T. G. Nielsen, and R. N. Glud. Feeding-related controls on microbial nitrogen cycling associated with the Arctic marine copepod Calanus hyperboreus. Marine Ecology Progress Series, 602:1–14, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Skilleter:2017:CBS

[SLO^{+}17] G. A. Skilleter, N. R. Loneragan, A. Olds, Y. Zharikov, and B. Cameron. Connectivity between seagrass and mangroves influences nekton assemblages using nearshore habitats. Marine Ecology Progress Series, 573:25–43, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/573/
m573p025.pdf.

Stier:2019:TVD

[SLO19] A. C. Stier, S. C. Lee, and M. I. O’Connor. Temporal variation in dispersal modifies dispersal-diversity relationships in an experimental seagrass metacommunity. Marine Ecology Progress Series, 613:67–76, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/613/m613p067.pdf.

Sturaro:2014:SAA

[SLPP^{+}14] N. Sturaro, G. Lepoint, A. Pérez-Perera, S. Vermeulen, P. Panzalis, A. Navone, and S. Gobert. Seagrass amphipod assemblages in a Mediterranean marine protected area: a multiscale approach. Marine Ecology Progress Series, 506:175– 192, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/506/m506p175.pdf.

Suter:2014:PAC

[SLR^{+}14] E. A. Suter, K. M. M. Lwiza, J. M. Rose, C. Gobler, and G. T. Taylor. Phytoplankton assemblage changes during decadal decreases in nitrogen loadings to the urbanized Long Island Sound estuary, USA. Marine Ecology Progress Series, 497:51–67, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m497p051.pdf.

Sanz-Lazaro:2015:EGD

[SLRM^{+}15] C. Sanz-Lázaro, L. Rindi, E. Maggi, M. Dal Bello, and L. Benedetti-Cecchi. Effects of grazer diversity on marine microphytobenthic biofilm: a ‘tug of war’ between complementarity and competition. Marine Ecology Progress Series, 540:145–155, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m540p145.pdf.

Sousa:2012:DES

[SLRP^{+}12] A. I. Sousa, A. I. Lillebø, N. Risgaard-Petersen, M. A. Pardal, and I. Caçador. Denitrification: an ecosystem service provided by salt marshes. Marine Ecology Progress Series, 448: 79–92, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/448/m448p079.pdf.

Sweetman:2013:FTS

[SLRS13] A. K. Sweetman, L. A. Levin, H. T. Rapp, and C. Schander. Faunal trophic structure at hydrothermal vents on the southern Mohn’s Ridge, Arctic Ocean. Marine Ecology Progress Series, 473:115–131, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m473p115.pdf.

Santos:2011:QFI

[SLS11] R. O. Santos, D. Lirman, and J. E. Serafy. Quantifying freshwater-induced fragmentation of submerged aquatic vegetation communities using a multi-scale landscape ecology approach. Marine Ecology Progress Series, 427:233–246, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m427p233.pdf.

Su:2016:VMP

[SLS16] R. Su, J. M. Lough, and D. Sun. Variations in massive Porites growth rates at Hainan Island, northern South China Sea. Marine Ecology Progress Series, 546:47–60, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/546/
m546p047.pdf.

Schiel:2018:ETP

[SLS18] D. R. Schiel, S. A. Lilley, and P. M. South. Ecological tipping points for an invasive kelp in rocky reef algal communities. Marine Ecology Progress Series, 587:93–104, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/587/
m587p093.pdf.

Schiel:2016:DCS

[SLSC16] D. R. Schiel, S. A. Lilley, P. M. South, and J. H. J. Coggins. Decadal changes in sea surface temperature, wave forces and intertidal structure in New Zealand. Marine Ecology Progress Series, 548:77–95, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/548/m548p077.pdf.

Stelzenmuller:2010:QCI

[SLSR10] V. Stelzenmüller, J. Lee, A. South, and S. I. Rogers. Quantifying cumulative impacts of human pressures on the marine environment: a geospatial modelling framework. Marine Ecology Progress Series, 398:19–32, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/398/m398p019.pdf.

Shelamoff:2019:PSD

[SLT^{+}19] V. Shelamoff, C. Layton, M. Tatsumi, M. J. Cameron, J. T. Wright, and C. R. Johnson. Patch size and density of canopy-forming kelp modify influences of ecosystem engineering on understorey algal and sessile invertebrate assemblages. Marine Ecology Progress Series, 632:59–79, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/632/
m632p059.pdf.

Sromek:2015:IGG

[SLW15] L. Sromek, R. Lasota, and M. Wolowicz. Impact of glaciations on genetic diversity of pelagic mollusks: Antarctic Limacina antarctica and Arctic Limacina helicina. Marine Ecology Progress Series, 525:143–152, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/525/m525p143.pdf.

Saunders:2010:PFD

[SM10a] M. I. Saunders and A. Metaxas. Physical forcing of distributions of bryozoan cyphonautes larvae in a coastal embayment. Marine Ecology Progress Series, 418:131–145, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/418/
m418p131.pdf.

Scheef:2010:OSC

[SM10b] L. P. Scheef and N. H. Marcus. Occurrence and significance of copepod resting egg accumulation in seagrass sediments. Marine Ecology Progress Series, 407:125–134, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/407/
m407p125.pdf.

Sorochan:2015:LPR

[SM15] K. A. Sorochan and A. Metaxas. Low predation rates on the larvae of three species of barnacles by the ctenophore Pleurobrachia pileus. Marine Ecology Progress Series, 541: 105–122, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/541/m541p105.pdf.

Silva:2017:HSM

[SM17a] M. Silva and I. R. MacDonald. Habitat suitability modeling for mesophotic coral in the northeastern Gulf of Mexico. Marine Ecology Progress Series, 583:121–136, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/583/
m583p121.pdf.

Sorochan:2017:ETM

[SM17b] K. A. Sorochan and A. Metaxas. The effect of temperature on motility of the nauplius and cypris stages of the acorn barnacle Semibalanus balanoides. Marine Ecology Progress Series, 579:55–66, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/579/m579p055.pdf.

Smale:2013:MSP

[Sma13] D. A. Smale. Multi-scale patterns of spatial variability in sessile assemblage structure do not alter predictably with development time. Marine Ecology Progress Series, 482: 29–41, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/482/m482p029.pdf.

Smith:2018:EAB

[SMBC18] J. G. Smith, D. Malone, D. S. Baetscher, and M. H. Carr. Evaluation and application of BIOPOLE, a biopsy device for in situ non-lethal tissue extraction for fishes. Marine Ecology Progress Series, 595:149–156, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/595/m595p149.pdf.

Santos:2015:RLF

[SMBH15] R. G. Santos, A. S. Martins, M. B. Batista, and P. A. Horta. Regional and local factors determining green turtle Chelonia mydas foraging relationships with the environment. Marine Ecology Progress Series, 529:265–277, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/529/
m529p265.pdf.

Schultner:2014:CMC

[SMC^{+}14] J. Schultner, B. Moe, O. Chastel, S. Tartu, C. Bech, and A. S. Kitaysky. Corticosterone mediates carry-over effects between breeding and migration in the kittiwake Rissa tridactyla. Marine Ecology Progress Series, 496:125–133, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/496/
m496p125.pdf.

Sequeira:2013:STP

[SMD^{+}13] A. M. M. Sequeira, C. Mellin, S. Delean, M. G. Meekan, and C. J. A. Bradshaw. Spatial and temporal predictions of interdecadal trends in Indian Ocean whale sharks. Marine Ecology Progress Series, 478:185–195, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/478/m478p185.pdf.

Schabetsberger:2016:HFA

[SMD^{+}16] R. Schabetsberger, M. J. Miller, G. Dall’Olmo, R. Kaiser, F. Økland, S. Watanabe, K. Aarestrup, and K. Tsukamoto. Hydrographic features of anguillid spawning areas: potential signposts for migrating eels. Marine Ecology Progress Series, 554:141–155, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/554/m554p141.pdf.

Strom:2010:LLS

[SMF10] S. L. Strom, E. L. Macri, and K. A. Fredrickson. Light limitation of summer primary production in the coastal Gulf of Alaska: physiological and environmental causes. Marine Ecology Progress Series, 402:45–57, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/402/
m402p045.pdf.

Speed:2011:STM

[SMF^{+}11] C. W. Speed, M. G. Meekan, I. C. Field, C. R. McMahon, J. D. Stevens, F. McGregor, C. Huveneers, Y. Berger, and C. J. A. Bradshaw. Spatial and temporal movement patterns of a multi-species coastal reef shark aggregation. Marine Ecology Progress Series, 429:261–275, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/429/m429p261.pdf.

Speed:2012:HSS

[SMF^{+}12] C. W. Speed, M. G. Meekan, I. C. Field, C. R. McMahon, and C. J. A. Bradshaw. Heat-seeking sharks: support for behavioural thermoregulation in reef sharks. Marine Ecology Progress Series, 463:231–244, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/463/m463p231.pdf.

Shipley:2017:LLU

[SMF^{+}17] O. N. Shipley, K. J. Murchie, M. G. Frisk, E. J. Brooks, O. R. O’Shea, and M. Power. Low lipid and urea effects and inter-tissue comparisons of stable isotope signatures in three nearshore elasmobranchs. Marine Ecology Progress Series, 579:233–238, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/579/m579p233.pdf.

Saied:2012:LTN

[SMH^{+}12] A. Saied, F. Maffucci, S. Hochscheid, S. Dryag, B. Swayeb, M. Borra, A. Ouerghi, G. Procaccini, and F. Bentivegna. Loggerhead turtles nesting in Libya: an important management unit for the Mediterranean stock. Marine Ecology Progress Series, 450:207–218, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/450/m450p207.pdf.

Smith:2014:SNU

[SMHB14] S. L. Smith, A. Merico, S. Hohn, and G. Brandt. Sizingup nutrient uptake kinetics: combining a physiological tradeoff with size-scaling of phytoplankton traits. Marine Ecology Progress Series, 511:33–39, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m511p033.pdf.

Schiariti:2014:ARS

[SMJ^{+}14] A. Schiariti, A. C. Morandini, G. Jarms, R. von Glehn Paes, S. Franke, and H. Mianzan. Asexual reproduction strategies and blooming potential in Scyphozoa. Marine Ecology Progress Series, 510:241–253, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/510/m510p241.pdf.

Skern-Mauritzen:2011:BWD

[SMJBØ11] M. Skern-Mauritzen, E. Johannesen, A. Bjørge, and N. Øien. Baleen whale distributions and prey associations in the Barents Sea. Marine Ecology Progress Series, 426:289–301, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2011/426/m426p289.pdf.

Sorensen:2015:SCC

[SMJP^{+}15] H. L. Sørensen, L. Meire, T. Juul-Pedersen, H. C. de Stigter, F. J. R. Meysman, S. Rysgaard, B. Thamdrup, and R. N. Glud. Seasonal carbon cycling in a Greenlandic fjord: an integrated pelagic and benthic study. Marine Ecology Progress Series, 539:1–17, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Skjaeraasen:2010:MCB

[SMK^{+}10] J. E. Skjæraasen, J. J. Meager, Ø. Karlsen, I. Mayer, G. Dahle, G. Rudolfsen, and A. Fernö. Mating competition between farmed and wild cod Gadus morhua. Marine Ecology Progress Series, 412:247–258, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/412/m412p247.pdf.

Strydom:2017:SHO

[SMK^{+}17] S. Strydom, K. McMahon, G. A. Kendrick, J. Statton, and P. S. Lavery. Seagrass Halophila ovalis is affected by light quality across different life history stages. Marine Ecology Progress Series, 572:103–116, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m572p103.pdf.

Schulz:2018:RSR

[SMK^{+}18] I. Schulz, M. Montresor, C. Klaas, P. Assmy, S. Wolzenburg, M. Gauns, A. Sarkar, S. Thiele, D. Wolf-Gladrow, W. Naqvi, and V. Smetacek. Remarkable structural resistance of a nanoflagellate-dominated plankton community to iron fertilization during the Southern Ocean experiment LO-HAFEX. Marine Ecology Progress Series, 601:77–95, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m601p077.pdf.

Soares:2014:SVF

[SML^{+}14] L. S. H. Soares, E. Y. Muto, J. P. Lopez, G. R. V. Clauzet, and I. Valiela. Seasonal variability of δ ^{13}C and δ ^{15}N of fish and squid in the Cabo Frio upwelling system of the southwestern Atlantic. Marine Ecology Progress Series, 512:9–21, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m512p009.pdf.

Schultz:2015:SVA

[SML^{+}15] A. L. Schultz, H. A. Malcolm, M. Linklater, A. R. Jordan, T. Ingleton, and S. D. A. Smith. Sediment variability affects fish community structure in unconsolidated habitats of a subtropical marine park. Marine Ecology Progress Series, 532:213–226, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/532/m532p213.pdf.

Sohel:2017:ETP

[SML17] S. Sohel, J. Mattila, and K. Lindström. Effects of turbidity on prey choice of three-spined stickleback Gasterosteus aculeatus. Marine Ecology Progress Series, 566:159–167, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/566/
m566p159.pdf.

Shanks:2014:OTP

[SMM^{+}14] A. L. Shanks, S. G. Morgan, J. MacMahan, A. J. H. M. Reniers, M. Jarvis, J. Brown, A. Fujimura, and C. Griesemer. Onshore transport of plankton by internal tides and upwelling-relaxation events. Marine Ecology Progress Series, 502:39–51, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m502p039.pdf.

Shanks:2015:TLD

[SMM^{+}15] A. L. Shanks, J. MacMahan, S. G. Morgan, A. J. H. M. Reniers, M. Jarvis, J. Brown, A. Fujimura, and C. Griesemer. Transport of larvae and detritus across the surf zone of a steep reflective pocket beach. Marine Ecology Progress Series, 528:71–86, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m528p071.pdf.

Smith:2016:ESR

[SMMFH16] W. D. Smith, J. A. Miller, J. F. Márquez-Farías, and S. S. Heppell. Elemental signatures reveal the geographic origins of a highly migratory shark: prospects for measuring population connectivity. Marine Ecology Progress Series, 556: 173–193, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/556/m556p173.pdf.

Sargeant:2016:SVM

[SMND16] S. L. Sargeant, J. C. Murrell, P. D. Nightingale, and J. L. Dixon. Seasonal variability in microbial methanol utilisation in coastal waters of the western English Channel. Marine Ecology Progress Series, 550:53–64, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/550/m550p053.pdf.

Sale:2019:AAS

[SMOH19] T. L. Sale, P. B. Marko, T. A. Oliver, and C. L. Hunter. Assessment of acclimatization and subsequent survival of corals during repeated natural thermal stress events in Hawai’i. Marine Ecology Progress Series, 624:65–76, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/624/
m624p065.pdf.

Sevadjian:2010:EPS

[SMP10] J. C. Sevadjian, M. A. McManus, and G. Pawlak. Effects of physical structure and processes on thin zooplankton layers in Mamala Bay, Hawaii. Marine Ecology Progress Series, 409:95–106, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m409p095.pdf.

Stunz:2010:RVO

[SMR10] G. W. Stunz, T. J. Minello, and L. P. Rozas. Relative value of oyster reef as habitat for estuarine nekton in Galveston Bay, Texas. Marine Ecology Progress Series, 406:147–159, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m406p147.pdf.

Serrano:2011:SRP

[SMR11a] O. Serrano, M. A. Mateo, and P. Renom. Seasonal response of Posidonia oceanica to light disturbances. Marine Ecology Progress Series, 423:29–38, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/423/m423p029.pdf.

Staniland:2011:FBT

[SMR^{+}11b] I. J. Staniland, A. Morton, S. L. Robinson, D. Malone, and J. Forcada. Foraging behaviour in two Antarctic fur seal colonies with differing population recoveries. Marine Ecology Progress Series, 434:183–196, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/434/m434p183.pdf.

Svensen:2014:IDC

[SMR14] C. Svensen, N. Morata, and M. Reigstad. Increased degradation of copepod faecal pellets by co-acting dinoflagellates and Centropages hamatus. Marine Ecology Progress Series, 516: 61–70, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m516p061.pdf.

Simonson:2015:KHWb

[SMS15] E. J. Simonson, A. Metaxas, and R. E. Scheibling. Kelp in hot water: II. Effects of warming seawater temperature on kelp quality as a food source and settlement substrate. Marine Ecology Progress Series, 537:105–119, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/537/
m537p105.pdf.

Silva:2019:TSD

[SMSB19] T. L. Silva, T. A. Mooney, L. S. Sayigh, and M. F. Baumgartner. Temporal and spatial distributions of delphinid species in Massachusetts Bay (USA) using passive acoustics from ocean gliders. Marine Ecology Progress Series, 631:1–17, 2019. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Sanders:2019:LVO

[SMSE19] C. J. Sanders, D. T. Maher, J. M. Smoak, and B. D. Eyre. Large variability in organic carbon and CaCO 3 burial in seagrass meadows: a case study from three Australian estuaries. Marine Ecology Progress Series, 616:211–218, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/616/
m616p211.pdf.

Silva:2017:EEB

[SMSL17] S. Silva, C. Macaya-Solis, and M. C. Lucas. Energetically efficient behaviour may be common in biology, but it is not universal: a test of selective tidal stream transport in a poor swimmer. Marine Ecology Progress Series, 584:161– 174, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/584/m584p161.pdf.

Scales:2015:OLT

[SMVC^{+}15] K. L. Scales, P. I. Miller, N. Varo-Cruz, D. J. Hodgson, L. A. Hawkes, and B. J. Godley. Oceanic loggerhead turtles Caretta caretta associate with thermal fronts: evidence from the Canary Current Large Marine Ecosystem. Marine Ecology Progress Series, 519:195–207, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/519/m519p195.pdf.

Soria:2012:LBO

[SMVM^{+}12] G. Soria, A. Munguía-Vega, S. G. Marinone, M. Moreno-Báez, I. Martínez-Tovar, and R. Cudney-Bueno. Linking biooceanography and population genetics to assess larval connectivity. Marine Ecology Progress Series, 463:159–175, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m463p159.pdf.

Sullivan:2017:MMP

[SN17] T. J. Sullivan and J. E. Neigel. Misidentification of megalopae as a potential source of error in studies of population genetics and ecology of the blue crab Callinectes sapidus. Marine Ecology Progress Series, 565:95–111, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/565/
m565p095.pdf.

Skjaeraasen:2010:LEA

[SNK^{+}10] J. E. Skjæraasen, R. D. M. Nash, J. Kennedy, A. Thorsen, T. Nilsen, and O. S. Kjesbu. Liver energy, atresia and oocyte stage influence fecundity regulation in Northeast Arctic cod. Marine Ecology Progress Series, 404:173–183, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/404/
m404p173.pdf.

Sveegaard:2012:SIB

[SNNS^{+}12] S. Sveegaard, J. Nabe-Nielsen, K. J. Stæhr, T. F. Jensen, K. N. Mouritsen, and J. Teilmann. Spatial interactions between marine predators and their prey: herring abundance as a driver for the distributions of mackerel and harbour porpoise. Marine Ecology Progress Series, 468:245–253, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/468/
m468p245.pdf.

Surman:2012:ECV

[SNS12] C. A. Surman, L. W. Nicholson, and J. A. Santora. Effects of climate variability on breeding phenology and performance of tropical seabirds in the eastern Indian Ocean. Marine Ecology Progress Series, 454:147–157, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m454p147.pdf.

Schlenger:2013:MIH

[SNS^{+}13] A. J. Schlenger, E. W. North, Z. Schlag, Y. Li, D. H. Secor, K. A. Smith, and E. J. Niklitschek. Modeling the influence of hypoxia on the potential habitat of Atlantic sturgeon Acipenser oxyrinchus: a comparison of two methods. Marine Ecology Progress Series, 483:257–272, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/483/
m483p257.pdf.

Swalethorp:2016:ELI

[SNT^{+}16] R. Swalethorp, T. G. Nielsen, A. R. Thompson, M. Møhl, and P. Munk. Early life of an inshore population of West Greenlandic cod Gadus morhua: spatial and temporal aspects of growth and survival. Marine Ecology Progress Series, 555:185–202, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/555/m555p185.pdf.

Suzuki:2018:MUH

[SNT^{+}18] K. S. Suzuki, Y. Niida, T. Tsubono, H. Takimoto, E. Kumakura, H. Ishii, and Y. Nogata. Mechanisms underlying heterogeneous distribution of moon jellyfish Aurelia aurita s.l. across a sharp pycnocline. Marine Ecology Progress Series, 591:229–239, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/591/m591p229.pdf.

Sano:2015:TSD

[SNTN15] M. Sano, Y. Nishibe, Y. Tanaka, and S. Nishida. Temporally sustained dietary niche partitioning in two mesopelagic copepod species and their mouthpart morphology. Marine Ecology Progress Series, 518:51–67, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/518/m518p051.pdf.

Sewall:2019:GES

[SNVH19] F. Sewall, B. Norcross, J. Vollenweider, and R. Heintz. Growth, energy storage, and feeding patterns reveal winter mortality risks for juvenile Pacific herring in Prince William Sound, Alaska, USA. Marine Ecology Progress Series, 623: 195–208, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m623p195.pdf.

Senokuchi:2018:CFA

[SNW^{+}18] R. Senokuchi, H. Nomaki, H. K. Watanabe, T. Kitahashi, N. O. Ogawa, and M. Shimanaga. Chemoautotrophic food availability influences copepod assemblage composition at deep hydrothermal vent sites within sea knoll calderas in the northwestern Pacific. Marine Ecology Progress Series, 607: 37–51, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/607/m607p037.pdf.

Sumoski:2012:BDE

[SO12] S. E. Sumoski and R. J. Orth. Biotic dispersal in eelgrass Zostera marina. Marine Ecology Progress Series, 471:1–10, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Stevens:2015:BCA

[SO15] J. L. Stevens and J. B. Olson. Bacterial communities associated with lionfish in their native and invaded ranges. Marine Ecology Progress Series, 531:253–262, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/531/
m531p253.pdf.

Spivak:2016:LEN

[SO16] A. C. Spivak and J. Ossolinski. Limited effects of nutrient enrichment on bacterial carbon sources in salt marsh tidal creek sediments. Marine Ecology Progress Series, 544:107– 130, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/544/m544p107.pdf.

Schabetsberger:2013:OMB

[SØA^{+}13] R. Schabetsberger, F. Økland, K. Aarestrup, D. Kalfatak, U. Sichrowsky, M. Tambets, G. Dall’Olmo, R. Kaiser, and P. I. Miller. Oceanic migration behaviour of tropical Pacific eels from Vanuatu. Marine Ecology Progress Series, 475:177–190, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/475/m475p177.pdf.

Staaterman:2017:BMC

[SOA^{+}17] E. Staaterman, M. B. Ogburn, A. H. Altieri, S. J. Brandl, R. Whippo, J. Seemann, M. Goodison, and J. E. Duffy. Bioacoustic measurements complement visual biodiversity surveys: preliminary evidence from four shallow marine habitats. Marine Ecology Progress Series, 575:207–215, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/575/
m575p207.pdf.

Saeck:2016:NRN

[SOB16] E. A. Saeck, K. R. O’Brien, and M. A. Burford. Nitrogen response of natural phytoplankton communities: a new indicator based on photosynthetic efficiency F_{v}/F_{m}. Marine Ecology Progress Series, 552:81–92, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/552/m552p081.pdf.

Stukel:2013:CMV

[SOBNL13] M. R. Stukel, M. D. Ohman, C. R. Benitez-Nelson, and M. R. Landry. Contributions of mesozooplankton to vertical carbon export in a coastal upwelling system. Marine Ecology Progress Series, 491:47–65, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/491/m491p047.pdf.

Stafford:2013:CSA

[SOC13] K. M. Stafford, S. R. Okkonen, and J. T. Clarke. Correlation of a strong Alaska coastal current with the presence of beluga whales Delphinapterus leucas near Barrow, Alaska. Marine Ecology Progress Series, 474:287–297, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/474/
m474p287.pdf.

Shannon:2014:CCS

[SOJ14] L. J. Shannon, W. Osman, and A. Jarre. Communicating changes in state of the southern Benguela ecosystem using trophic, model-derived indicators. Marine Ecology Progress Series, 512:217–237, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m512p217.pdf.

Schabetsberger:2015:GME

[SØK^{+}15] R. Schabetsberger, F. Økland, D. Kalfatak, U. Sichrowsky, M. Tambets, K. Aarestrup, C. Gubili, J. Sarginson, B. Boufana, R. Jehle, G. Dall’Olmo, M. J. Miller, A. Scheck, R. Kaiser, and G. Quartly. Genetic and migratory evidence for sympatric spawning of tropical Pacific eels from Vanuatu. Marine Ecology Progress Series, 521:171–187, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/521/
m521p171.pdf.

Suprenand:2015:MGP

[SOT15] P. M. Suprenand, E. H. Ombres, and J. J. Torres. Metabolism of gymnosomatous pteropods in waters of the western Antarctic Peninsula shelf during austral fall. Marine Ecology Progress Series, 518:69–83, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/518/m518p069.pdf.

Studer:2012:SDI

[SP12] A. Studer and R. Poulin. Seasonal dynamics in an intertidal mudflat: the case of a complex trematode life cycle. Marine Ecology Progress Series, 455:79–93, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/455/
m455p079.pdf.

Schaafsma:2013:PGI

[SP13a] F. L. Schaafsma and L. Peperzak. Phytoplankton growth inhibited by the toxic and bacterivorous ciliate Uronema marinum (Protozoa, Ciliophora). Marine Ecology Progress Series, 475:35–48, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/475/m475p035.pdf.

Scyphers:2013:CDE

[SP13b] S. B. Scyphers and S. P. Powers. Context-dependent effects of a marine ecosystem engineer on predator–prey interactions. Marine Ecology Progress Series, 491:295–301, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/491/
m491p295.pdf.

Sijercic:2015:HSS

[SP15] A. Sijerčić and N. M. Price. Hydroxamate siderophore secretion by Pseudoalteromonas haloplanktis during steady-state and transient growth under iron limitation. Marine Ecology Progress Series, 531:105–120, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/531/m531p105.pdf.

Stehfest:2014:IDM

[SPBS14] K. M. Stehfest, T. A. Patterson, A. Barnett, and J. M. Semmens. Intraspecific differences in movement, dive behavior and vertical habitat preferences of a key marine apex predator. Marine Ecology Progress Series, 495:249–262, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/495/
m495p249.pdf.

Staaterman:2014:CPM

[SPD^{+}14] Erica Staaterman, Claire B. Paris, Harry A. DeFerrari, David A. Mann, Aaron N. Rice, and Evan K. D’Alessandro. Celestial patterns in marine soundscapes. Marine Ecology Progress Series, 508:17–32, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/508/m508p017.pdf.

Stewart:2013:DOS

[SPH^{+}13] H. L. Stewart, N. N. Price, S. J. Holbrook, R. J. Schmitt, and A. J. Brooks. Determinants of the onset and strength of mutualistic interactions between branching corals and associate crabs. Marine Ecology Progress Series, 493:155– 163, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/493/m493p155.pdf.

Schulien:2017:PMA

[SPH^{+}17] J. A. Schulien, M. B. Peacock, K. Hayashi, P. Raimondi, and R. M. Kudela. Phytoplankton and microbial abundance and bloom dynamics in the upwelling shadow of Monterey Bay, California, from 2006 to 2013. Marine Ecology Progress Series, 572:43–56, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m572p043.pdf.

Schwarz:2018:IET

[SPH18] R. Schwarz, U. Piatkowski, and H. J. T. Hoving. Impact of environmental temperature on the lifespan of octopods. Marine Ecology Progress Series, 605:151–164, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/605/
m605p151.pdf.

Schofield:2017:AUS

[SPHK17] G. Schofield, K. Papafitsoros, R. Haughey, and K. Katselidis. Aerial and underwater surveys reveal temporal variation in cleaning-station use by sea turtles at a temperate breeding area. Marine Ecology Progress Series, 575:153– 164, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/575/m575p153.pdf.

Schmidt:2012:CCC

[SPJ^{+}12] G. M. Schmidt, N. Phongsuwan, C. Jantzen, C. Roder, S. Khokiattiwong, and C. Richter. Coral community composition and reef development at the Similan Islands, Andaman Sea, in response to strong environmental variations. Marine Ecology Progress Series, 456:113–126, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/456/
m456p113.pdf.

Sara:2013:PBF

[SPM^{+}13] G. Sarà, V. Palmeri, V. Montalto, A. Rinaldi, and J. Widdows. Parameterisation of bivalve functional traits for mechanistic eco-physiological dynamic energy budget (DEB) models. Marine Ecology Progress Series, 480:99–117, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/480/
m480p099.pdf.

Segura:2016:CBE

[SPM^{+}16] C. J. Segura, J. A. Pechenik, J. A. Montory, J. M. Navarro, K. A. Paschke, V. M. Cubillos, and O. R. Chaparro. The cost of brooding in an estuary: implications of declining salinity for gastropod females and their brooded embryos. Marine Ecology Progress Series, 543:187–199, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/543/m543p187.pdf.

Scharping:2019:LST

[SPMN19] R. J. Scharping, L. V. Plough, D. W. Meritt, and E. W. North. Low-salinity tolerance of early-stage oyster larvae from a mesohaline estuary. Marine Ecology Progress Series, 613:97–106, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/613/m613p097.pdf.

Schrimpf:2012:TOP

[SPP12] M. B. Schrimpf, J. K. Parrish, and S. F. Pearson. Trade-offs in prey quality and quantity revealed through the behavioral compensation of breeding seabirds. Marine Ecology Progress Series, 460:247–259, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2012/460/m460p247.pdf.

Smoot:2015:ABA

[SPP15] S. C. Smoot, C. J. Plante, and R. D. Podolsky. Anti-bacterial activity in egg masses of Melanochlamys diomedea across habitats differing in sediment properties and bacterial load. Marine Ecology Progress Series, 524:185–196, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/524/
m524p185.pdf.

Stamieszkin:2017:ZGE

[SPP17] K. Stamieszkin, N. J. Poulton, and A. J. Pershing. Zooplankton grazing and egestion shifts particle size distribution in natural communities. Marine Ecology Progress Series, 575:43–56, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/575/m575p043.pdf.

Surma:2018:EBE

[SPP18] S. Surma, E. A. Pakhomov, and T. J. Pitcher. Energybased ecosystem modelling illuminates the ecological role of Northeast Pacific herring. Marine Ecology Progress Series, 588:147–161, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/588/m588p147.pdf.

Stanley:2016:BOE

[SPS16] R. R. E. Stanley, E. J. Pedersen, and P. V. R. Snelgrove. Biogeographic, ontogenetic, and environmental variability in larval behaviour of American lobster Homarus americanus. Marine Ecology Progress Series, 553:125–146, 2016. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m553p125.pdf.

Sammarco:2014:PEN

[SPSG14] P. W. Sammarco, S. A. Porter, J. Sinclair, and M. Genazzio. Population expansion of a new invasive coral species, Tubastraea micranthus, in the northern Gulf of Mexico. Marine Ecology Progress Series, 495:161–173, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m495p161.pdf.

Strmecki:2010:CPE

[SPSP10] Sla– dana Strmečki, M. Plavšić, S. Steigenberger, and U. Passow. Characterization of phytoplankton exudates and carbohydrates in relation to their complexation of copper, cadmium and iron. Marine Ecology Progress Series, 408:33–46, 2010.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m408p033.pdf.

Sponaugle:2012:OML

[SPW^{+}12] S. Sponaugle, C. Paris, K. D. Walter, V. Kourafalou, and E. D’Alessandro. Observed and modeled larval settlement of a reef fish to the Florida Keys. Marine Ecology Progress Series, 453:201–212, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/453/m453p201.pdf.

Saurel:2013:TML

[SPWK13] C. Saurel, J. K. Petersen, P. J. Wiles, and M. J. Kaiser. Turbulent mixing limits mussel feeding : direct estimates of feeding rate and vertical diffusivity. Marine Ecology Progress Series, 485:105–121, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/485/m485p105.pdf.

Smith:2018:IOP

[SQRDLC18] K. M. Smith, B. E. Quirk-Royal, K. Drake-Lavelle, and M. J. Childress. Influences of ontogenetic phase and resource availability on parrotfish foraging preferences in the Florida Keys, FL (USA). Marine Ecology Progress Series, 603:175–187, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/603/m603p175.pdf.

Simmons:2013:RER

[SQS^{+}13] R. K. Simmons, T. P. Quinn, L. W. Seeb, D. E. Schindler, and R. Hilborn. Role of estuarine rearing for sockeye salmon in Alaska (USA). Marine Ecology Progress Series, 481:211– 223, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/481/m481p211.pdf.

Stewart:2017:TOM

[SRA^{+}17] J. D. Stewart, C. A. Rohner, G. Araujo, J. Avila, D. Fernando, K. Forsberg, A. Ponzo, J. M. Rambahiniarison, C. M. Kurle, and B. X. Semmens. Trophic overlap in mobulid rays: insights from stable isotope analysis. Marine Ecology Progress Series, 580:131–151, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/580/m580p131.pdf.

Salo:2015:GSR

[SRB15] T. Salo, T. B. H. Reusch, and C. Boström. Genotypespecific responses to light stress in eelgrass Zostera marina, a marine foundation plant. Marine Ecology Progress Series, 519:129–140, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/519/m519p129.pdf.

Simonis:2017:LCA

[SRB^{+}17] A. E. Simonis, M. A. Roch, B. Bailey, J. Barlow, R. E. S. Clemesha, S. Iacobellis, J. A. Hildebrand, and S. Baumann-Pickering. Lunar cycles affect common dolphin Delphinus delphis foraging in the Southern California Bight. Marine Ecology Progress Series, 577:221–235, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m577p221.pdf.

Staaf:2010:OSS

[SRCE^{+}10] D. J. Staaf, R. I. Ruiz-Cooley, C. Elliger, Z. Lebaric, B. Campos, U. Markaida, and W. F. Gilly. Ommastrephid squids Sthenoteuthis oualaniensis and Dosidicus gigas in the eastern Pacific show convergent biogeographic breaks but contrasting population structures. Marine Ecology Progress Series, 418:165–178, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/418/m418p165.pdf.

Silberberger:2016:STS

[SRER16] M. J. Silberberger, P. E. Renaud, B. Espinasse, and H. Reiss. Spatial and temporal structure of the meroplankton community in a sub-Arctic shelf system. Marine Ecology Progress Series, 555:79–93, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m555p079.pdf.

Santos:2014:ADM

[SRG14] J. Santos, D. Rouillard, and J. C. Groeneveld. Advectiondiffusion models of spiny lobster Palinurus gilchristi migrations for use in spatial fisheries management. Marine Ecology Progress Series, 498:227–241, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/498/m498p227.pdf.

Sigler:2015:FFS

[SRH^{+}15] M. F. Sigler, C. N. Rooper, G. R. Hoff, R. P. Stone, R. A. McConnaughey, and T. K. Wilderbuer. Faunal features of submarine canyons on the eastern Bering Sea slope. Marine Ecology Progress Series, 526:21–40, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m526p021.pdf.

Staby:2011:LTA

[SRK11] A. Staby, A. Røstad, and S. Kaartvedt. Long-term acoustical observations of the mesopelagic fish Maurolicus muelleri reveal novel and varied vertical migration patterns. Marine Ecology Progress Series, 441:241–255, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/441/m441p241.pdf.

Slone:2013:MSR

[SRK13] D. H. Slone, J. P. Reid, and W. J. Kenworthy. Mapping spatial resources with GPS animal telemetry: foraging manatees locate seagrass beds in the Ten Thousand Islands, Florida, USA. Marine Ecology Progress Series, 476:285–299, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/476/
m476p285.pdf.

Sanchez:2018:WCS

[SRK^{+}18] S. Sánchez, R. D. Reina, A. Kato, Y. Ropert-Coudert, C. Cavallo, G. C. Hays, and A. Chiaradia. Within-colony spatial segregation leads to foraging behaviour variation in a seabird. Marine Ecology Progress Series, 606:215–230, 2018. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m606p215.pdf.

Santora:2010:SAB

[SRLV10] J. A. Santora, C. S. Reiss, V. J. Loeb, and R. R. Veit. Spatial association between hotspots of baleen whales and demographic patterns of Antarctic krill Euphausia superba suggests size-dependent predation. Marine Ecology Progress Series, 405:255–269, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/405/m405p255.pdf.

Skerritt:2015:FSM

[SRM^{+}15] D. J. Skerritt, P. A. Robertson, A. C. Mill, N. V. C. Polunin, and C. Fitzsimmons. Fine-scale movement, activity patterns and home-ranges of European lobster Homarus gammarus. Marine Ecology Progress Series, 536:203–219, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m536p203.pdf.

Schleimer:2019:STP

[SRP^{+}19] A. Schleimer, C. Ramp, S. Plourde, C. Lehoux, R. Sears, and P. S. Hammond. Spatio-temporal patterns in fin whale Balaenoptera physalus habitat use in the northern Gulf of St. Lawrence. Marine Ecology Progress Series, 623:221– 234, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/623/m623p221.pdf.

Schroeter:2015:ERP

[SRR15] S. C. Schroeter, D. C. Reed, and P. T. Raimondi. Effects of reef physical structure on development of benthic reef community: a large-scale artificial reef experiment. Marine Ecology Progress Series, 540:43–55, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/540/m540p043.pdf.

Svensen:2012:DCF

[SRRS12] C. Svensen, C. Wexels Riser, M. Reigstad, and L. Seuthe. Degradation of copepod faecal pellets in the upper layer: role of microbial community and Calanus finmarchicus. Marine Ecology Progress Series, 462:39–49, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/462/
m462p039.pdf.

Schilling:2018:EEN

[SRSH^{+}18] H. T. Schilling, P. Reis-Santos, J. M. Hughes, J. A. Smith, J. D. Everett, J. Stewart, B. M. Gillanders, and I. M. Suthers. Evaluating estuarine nursery use and life history patterns of Pomatomus saltatrix in eastern Australia. Marine Ecology Progress Series, 598:187–199, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m598p187.pdf.

Salomon:2010:PPD

[SS10a] P. S. Salomon and W. Stolte. Predicting the population dynamics in Amoebophrya parasitoids and their dinoflagellate hosts using a mathematical model. Marine Ecology Progress Series, 419:1–10, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Severin:2010:STV

[SS10b] I. Severin and L. J. Stal. Spatial and temporal variability in nitrogenase activity and diazotrophic community composition in coastal microbial mats. Marine Ecology Progress Series, 417:13–25, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/417/m417p013.pdf.

Swearer:2010:RVL

[SS10c] S. E. Swearer and J. S. Shima. Regional variation in larval retention and dispersal drives recruitment patterns in a temperate reef fish. Marine Ecology Progress Series, 417: 229–236, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/417/m417p229.pdf.

Shanks:2011:TDJ

[SS11a] A. L. Shanks and R. K. Shearman. Thread-drifting juvenile Mytilus spp. in continental shelf waters off Coos Bay, Oregon, USA. Marine Ecology Progress Series, 427:105– 115, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/427/m427p105.pdf.

Sorte:2011:PPC

[SS11b] C. J. B. Sorte and J. J. Stachowicz. Patterns and processes of compositional change in a California epibenthic community. Marine Ecology Progress Series, 435:63–74, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/435/
m435p063.pdf.

Semmens:2012:TOO

[SS12a] D. Semmens and S. E. Swearer. Trade-offs obscure the relationship between egg size and larval traits in the diadromous fish Galaxias maculatus. Marine Ecology Progress Series, 461:165–174, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/461/m461p165.pdf.

Sunda:2012:MED

[SS12b] W. G. Sunda and K. W. Shertzer. Modeling ecosystem disruptive algal blooms: positive feedback mechanisms. Marine Ecology Progress Series, 447:31–47, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/447/
m447p031.pdf.

Sosik:2013:IEC

[SS13] E. A. Sosik and C. A. Simenstad. Isotopic evidence and consequences of the role of microbes in macroalgae detritusbased food webs. Marine Ecology Progress Series, 494:107– 119, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/494/m494p107.pdf.

Stone:2014:LTT

[SS14] J. P. Stone and D. K. Steinberg. Long-term time-series study of salp population dynamics in the Sargasso Sea. Marine Ecology Progress Series, 510:111–127, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/510/
m510p111.pdf.

Sumaila:2015:TSE

[SS15a] U. R. Sumaila and K. I. Stergiou. Theme section: Economics of marine ecosystem conservation. Marine Ecology Progress Series, ??:??, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m530p177.pdf.

Sumaila:2015:EMC

[SS15b] U. R. Sumaila and Konstantinos I. Stergiou. Economics of marine conservation. Marine Ecology Progress Series, 530:

179–182, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m530p179.pdf.

Spies:2016:ETL

[SS16] B. T. Spies and M. A. Steele. Effects of temperature and latitude on larval traits of two estuarine fishes in differing estuary types. Marine Ecology Progress Series, 544:243– 255, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/544/m544p243.pdf.

Srednick:2019:MHM

[SS19] G. S. Srednick and M. A. Steele. Macroalgal height is more important than species identity in driving differences in the distribution and behavior of fishes. Marine Ecology Progress Series, 613:139–149, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2019/613/m613p139.pdf.

Strohmeier:2015:RME

[SSAB^{+}15] T. Strohmeier, Ø. Strand, M. Alunno-Bruscia, A. Duinker, R. Rosland, J. Aure, S. R. Erga, L. J. Naustvoll, H. M. Jansen, and P. J. Cranford. Response of Mytilus edulis to enhanced phytoplankton availability by controlled upwelling in an oligotrophic fjord. Marine Ecology Progress Series, 518:139–152, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/518/m518p139.pdf.

Schutte:2010:RST

[SSB10] V. G. W. Schutte, E. R. Selig, and J. F. Bruno. Regional spatio-temporal trends in Caribbean coral reef benthic communities. Marine Ecology Progress Series, 402:115–122, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m402p115.pdf.

Shillinger:2011:VHH

[SSB^{+}11] G. L. Shillinger, A. M. Swithenbank, H. Bailey, S. J. Bograd, M. R. Castelton, B. P. Wallace, J. R. Spotila, F. V. Paladino, R. Piedra, and B. A. Block. Vertical and horizontal habitat preferences of post-nesting leatherback turtles in the South Pacific Ocean. Marine Ecology Progress Series, 422:275–289, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m422p275.pdf.

Singer:2016:SSB

[SSB^{+}16] A. Singer, U. Schückel, M. Beck, O. Bleich, H. J. Brumsack, H. Freund, C. Geimecke, K. A. Lettmann, G. Millat, J. Staneva, A. Vanselow, H. Westphal, J. O. Wolff, A. Wurpts, and I. Kröncke. Small-scale benthos distribution modelling in a North Sea tidal basin in response to climatic and environmental changes (1970s–2009). Marine Ecology Progress Series, 551:13–30, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/551/m551p013.pdf.

Shalders:2018:PCM

[SSB^{+}18] T. C. Shalders, B. J. Saunders, S. Bennett, J. R. C. Parker, and E. S. Harvey. Potential climate-mediated changes to the distribution and density of pomacentrid reef fishes in south-western Australia. Marine Ecology Progress Series, 604:223–235, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/604/m604p223.pdf.

Sasson:2012:NMM

[SSBA12] N. Sasson, N. Simon-Blecher, and Y. Achituv. New molecular markers for revealing the population structure of Chthamalus stellatus in the Mediterranean and the eastern Atlantic. Marine Ecology Progress Series, 459:99–107, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/459/
m459p099.pdf.

Sanchez:2011:SPG

[SSBC11] R. Sánchez, R. D. Sepúlveda, A. Brante, and L. Cárdenas. Spatial pattern of genetic and morphological diversity in the direct developer Acanthina monodon (Gastropoda: Mollusca). Marine Ecology Progress Series, 434:121–131, 2011. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m434p121.pdf.

Sturdivant:2013:ESH

[SSD13] S. K. Sturdivant, R. D. Seitz, and R. J. Diaz. Effects of seasonal hypoxia on macrobenthic production and function in the Rappahannock River, Virginia, USA. Marine Ecology Progress Series, 490:53–68, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/490/m490p053.pdf.

Stith:2012:PTR

[SSdW^{+}12] B. M. Stith, D. H. Slone, M. de Wit, H. H. Edwards, C. A. Langtimm, E. D. Swain, L. E. Soderqvist, and J. P. Reid. Passive thermal refugia provided warm water for Florida manatees during the severe winter of 2009–2010. Marine Ecology Progress Series, 462:287–301, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/462/m462p287.pdf.

Stelzenmuller:2011:IMT

[SSFB11] V. Stelzenmüller, T. Schulze, H. O. Fock, and J. Berkenhagen. Integrated modelling tools to support risk-based decisionmaking in marine spatial management. Marine Ecology Progress Series, 441:197–212, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/441/m441p197.pdf.

Shiomi:2016:DSK

[SSHB16] K. Shiomi, K. Sato, Y. Handrich, and C. A. Bost. Diel shift of king penguin swim speeds in relation to light intensity changes. Marine Ecology Progress Series, 561:233–243, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/561/
m561p233.pdf.

Sanmarti:2014:TRS

[SSL^{+}14] N. Sanmartí, L. Saiz, I. Llagostera, M. Pérez, and J. Romero. Tolerance responses to simulated herbivory in the seagrass Cymodocea nodosa. Marine Ecology Progress Series, 517: 159–169, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/517/m517p159.pdf.

Simonson:2015:KHWa

[SSM15] E. J. Simonson, R. E. Scheibling, and A. Metaxas. Kelp in hot water: I. Warming seawater temperature induces weakening and loss of kelp tissue. Marine Ecology Progress Series, 537:89–104, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/537/m537p089.pdf.

Schram:2016:SAM

[SSM^{+}16a] J. B. Schram, K. M. Schoenrock, J. B. McClintock, C. D. Amsler, and R. A. Angus. Seawater acidification more than warming presents a challenge for two Antarctic macroalgalassociated amphipods. Marine Ecology Progress Series, 554: 81–97, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m554p081.pdf.

Stasko:2016:IDP

[SSM^{+}16b] A. D. Stasko, H. Swanson, A. Majewski, S. Atchison, J. Reist, and M. Power. Influences of depth and pelagic subsidies on the size-based trophic structure of Beaufort Sea fish communities. Marine Ecology Progress Series, 549:153–166, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/549/
m549p153.pdf.

Schram:2017:OWA

[SSM^{+}17] J. B. Schram, K. M. Schoenrock, J. B. McClintock, C. D. Amsler, and R. A. Angus. Ocean warming and acidification alter Antarctic macroalgal biochemical composition but not amphipod grazer feeding preferences. Marine Ecology Progress Series, 581:45–56, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m581p045.pdf.

Stewart:2019:NAA

[SSM^{+}19] J. D. Stewart, T. R. Smith, G. Marshall, K. Abernathy, I. A. Fonseca-Ponce, N. Froman, and G. M. W. Stevens. Novel applications of animal-borne crittercams reveal thermocline feeding in two species of manta ray. Marine Ecology Progress Series, 632:145–158, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/632/m632p145.pdf.

Shahraki:2016:FUI

[SSPKF16] M. Shahraki, U. Saint-Paul, U. Krumme, and B. Fry. Fish use of intertidal mangrove creeks at Qeshm Island, Iran. Marine Ecology Progress Series, 542:153–166, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/542/
m542p153.pdf.

Scott:2010:SSH

[SSR^{+}10] B. E. Scott, J. Sharples, O. N. Ross, J. Wang, G. J. Pierce, and C. J. Camphuysen. Sub-surface hotspots in shallow seas: fine-scale limited locations of top predator foraging habitat indicated by tidal mixing and sub-surface chlorophyll. Marine Ecology Progress Series, 408:207–226, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/408/
m408p207.pdf.

Snickars:2010:HSS

[SSS^{+}10] M. Snickars, G. Sundblad, A. Sandström, L. Ljunggren, U. Bergström, G. Johansson, and J. Mattila. Habitat selectivity of substrate-spawning fish: modelling requirements for the Eurasian perch Perca fluviatilis. Marine Ecology Progress Series, 398:235–243, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/398/m398p235.pdf.

Suryan:2012:NAU

[SSS12] R. M. Suryan, J. A. Santora, and W. J. Sydeman. New approach for using remotely sensed chlorophyll-a to identify seabird hotspots. Marine Ecology Progress Series, 451:213– 225, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m451p213.pdf.

Schultz:2019:CTD

[SSS19] E. T. Schultz, M. G. Smircich, and D. L. Strayer. Changes over three decades in feeding success of young American shad Alosa sapidissima are influenced by invading zebra mussels Dreissena polymorpha. Marine Ecology Progress Series, 628:141–153, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/628/m628p141.pdf.

Susdorf:2018:CDR

[SST^{+}18] R. Susdorf, N. K. G. Salama, C. D. Todd, R. J. Hillman, P. Elsmere, and D. Lusseau. Context-dependent reduction in somatic condition of wild Atlantic salmon infested with sea lice. Marine Ecology Progress Series, 606: 91–104, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/606/m606p091.pdf.

Simpson:2019:OTR

[SST19] S. J. Simpson, D. W. Sims, and C. N. Trueman. Ontogenetic trends in resource partitioning and trophic geography of sympatric skates (Rajidae) inferred from stable isotope composition across eye lenses. Marine Ecology Progress Series, 624:103–116, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/624/m624p103.pdf.

Stenberg:2015:LTE

[SSvD^{+}15] C. Stenberg, J. G. Støttrup, M. van Deurs, C. W. Berg, G. E. Dinesen, H. Mosegaard, T. M. Grome, and S. B. Leonhard. Long-term effects of an offshore wind farm in the North Sea on fish communities. Marine Ecology Progress Series, 528:257–265, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m528p257.pdf.

Spackeen:2017:IEE

[SSX^{+}17] J. L. Spackeen, R. E. Sipler, K. Xu, A. O. Tatters, N. G. Walworth, E. M. Bertrand, J. B. McQuaid, D. A. Hutchins, A. E. Allen, and D. A. Bronk. Interactive effects of elevated temperature and CO_{2} on nitrate, urea, and dissolved inorganic carbon uptake by a coastal California, USA, microbial community. Marine Ecology Progress Series, 577:49–65, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/577/
m577p049.pdf.

Skov:2008:RFS

[ST08] H. Skov and F. Thomsen. Resolving fine-scale spatio-temporal dynamics in the harbour porpoise Phocoena phocoena. Marine Ecology Progress Series, 373:173–186, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/373/

m373p173.pdf. See comment [TW10].

Sassa:2010:DGS

[ST10a] C. Sassa and Y. Tsukamoto. Distribution and growth of Scomber japonicus and S. australasicus larvae in the southern East China Sea in response to oceanographic conditions. Marine Ecology Progress Series, 419:185–199, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/419/
m419p185.pdf.

Skov:2010:GMM

[ST10b] H. Skov and F. Thomsen. General models of marine animal habitats require a process-based approach to parameter selection and model design. Marine Ecology Progress Series, 399:299–303, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m399p299.pdf.

Selkoe:2011:MCN

[ST11a] K. A. Selkoe and R. J. Toonen. Marine connectivity: a new look at pelagic larval duration and genetic metrics of dispersal. Marine Ecology Progress Series, 436:291–305, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/436/
m436p291.pdf.

Stergiou:2011:WSI

[ST11b] Konstantinos I. Stergiou and A. C. Tsikliras. As we see it: Fishing down, fishing through and fishing up: fundamental process versus technical details. Marine Ecology Progress Series, 441:295–301, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m441p295.pdf.

Schlacher:2013:SSO

[ST13] T. A. Schlacher and L. Thompson. Spatial structure on ocean-exposed sandy beaches: faunal zonation metrics and their variability. Marine Ecology Progress Series, 478:43– 55, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/478/m478p043.pdf.

Spiese:2014:DRA

[ST14] C. E. Spiese and E. A. Tatarkov. Dimethylsulfoxide reduction activity is linked to nutrient stress in Thalassiosira pseudonana NCMA 1335. Marine Ecology Progress Series, 507: 31–38, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/507/m507p031.pdf.

Simpson:2013:NDWb

[STBP13] K. G. Simpson, J. É. Tremblay, S. Brugel, and N. M. Price. Nutrient dynamics in the western Canadian Arctic. II. Estimates of new and regenerated production over the Mackenzie Shelf and Cape Bathurst Polynya. Marine Ecology Progress Series, 484:47–62, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m484p047.pdf.

Stieglitz:2013:HED

[Sti13a] T. C. Stieglitz. Habitat engineering by decadal-scale bioturbation around shipwrecks on the Great Barrier Reef midshelf. Marine Ecology Progress Series, 477:29–40, 2013. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m477p029.pdf.

Stimson:2013:CHF

[Sti13b] J. Stimson. Consumption by herbivorous fishes of macroalgae exported from coral reef flat refuges to the reef slope. Marine Ecology Progress Series, 472:87–99, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/472/
m472p087.pdf.

Sydeman:2012:ISC

[STK12] W. J. Sydeman, S. A. Thompson, and A. Kitaysky. Introduction: Seabirds and climate change: roadmap for the future. Marine Ecology Progress Series, 454:107–117, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m454p107.pdf.

Schismenou:2013:SCG

[STK^{+}13] E. Schismenou, K. Tsiaras, M. I. Kourepini, E. Lefkaditou, G. Triantafyllou, and S. Somarakis. Seasonal changes in growth and condition of anchovy late larvae explained with a hydrodynamic–biogeochemical model simulation. Marine Ecology Progress Series, 478:197–209, 2013. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m478p197.pdf.

Sturrock:2014:PIC

[STM^{+}14] A. M. Sturrock, C. N. Trueman, J. A. Milton, C. P. Waring, M. J. Cooper, and E. Hunter. Physiological influences can outweigh environmental signals in otolith microchemistry research. Marine Ecology Progress Series, 500:245–264, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m500p245.pdf.

Studer:2010:PGW

[STP10] A. Studer, D. W. Thieltges, and R. Poulin. Parasites and global warming: net effects of temperature on an intertidal host-parasite system. Marine Ecology Progress Series, 415: 11–22, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/415/m415p011.pdf.

Simpson:2013:NDWa

[STP13] K. G. Simpson, J. É. Tremblay, and N. M. Price. Nutrient dynamics in the western Canadian Arctic. I. new production in spring inferred from nutrient draw-down in the Cape Bathurst Polynya. Marine Ecology Progress Series, 484:33–45, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m484p033.pdf.

Stankovic:2018:MBA

[STRP18] M. Stankovic, N. Tantipisanuh, E. Rattanachot, and A. Prathep. Model-based approach for estimating biomass and organic carbon in tropical seagrass ecosystems. Marine Ecology Progress Series, 596:61–70, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/596/m596p061.pdf.

Soto:2018:OWC

[STS^{+}18] D. X. Soto, C. N. Trueman, K. M. Samways, M. J. Dadswell, and R. A. Cunjak. Ocean warming cannot explain synchronous declines in North American Atlantic salmon populations. Marine Ecology Progress Series, 601:203–213, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m601p203.pdf.

Soler:2016:CBS

[STSS^{+}16] G. A. Soler, R. J. Thomson, R. D. Stuart-Smith, A. D. M. Smith, and G. J. Edgar. Contributions of body size, habitat and taxonomy to predictions of temperate Australian fish diets. Marine Ecology Progress Series, 545:239–249, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/545/m545p239.pdf.

Shankar:2014:AMB

[STT14] S. Shankar, D. W. Townsend, and M. A. Thomas. Ammonium and maintenance of bloom populations of Alexandrium fundyense in the Gulf of Maine and on Georges Bank: results of laboratory culture experiments. Marine Ecology Progress Series, 507:57–67, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/507/m507p057.pdf.

Sassa:2010:STD

[STYT10] C. Sassa, Y. Tsukamoto, K. Yamamoto, and M. Tokimura. Spatio-temporal distribution and biomass of Benthosema pterotum (Pisces: Myctophidae) in the shelf region of the East China Sea. Marine Ecology Progress Series, 407:227– 241, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/407/m407p227.pdf.

Sherley:2013:ILR

[SUB^{+}13] R. B. Sherley, L. G. Underhill, B. J. Barham, P. J. Barham, J. C. Coetzee, R. J. M. Crawford, B. M. Dyer, T. M. Leshoro, and L. Upfold. Influence of local and regional prey availability on breeding performance of African penguins Spheniscus demersus. Marine Ecology Progress Series, 473:291– 301, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/473/m473p291.pdf.

Salvanes:2011:BPA

[SUPCB11] A. G. V. Salvanes, A. C. Utne-Palm, B. Currie, and V. A. Braithwaite. Behavioural and physiological adaptations of the bearded goby, a key fish species of the extreme environment of the northern Benguela upwelling. Marine Ecology Progress Series, 425:193–202, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m425p193.pdf.

Strand:2010:EPL

[SUPJ^{+}10] D. A. Strand, A. C. Utne-Palm, P. J. Jakobsen, V. A. Braithwaite, K. H. Jensen, and A. G. V. Salvanes. Enrichment promotes learning in fish. Marine Ecology Progress Series, 412:273–282, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m412p273.pdf.

Steele:2012:IRS

[SV12] L. Steele and J. F. Valentine. Idiosyncratic responses of seagrass phenolic production following sea urchin grazing. Marine Ecology Progress Series, 466:81–92, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/466/
m466p081.pdf.

Santora:2013:STP

[SV13] J. A. Santora and R. R. Veit. Spatio-temporal persistence of top predator hotspots near the Antarctic Peninsula.

Marine Ecology Progress Series, 487:287–304, 2013. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m487p287.pdf.

Steele:2015:SDM

[SV15] L. Steele and J. F. Valentine. Seagrass deterrence to mesograzer herbivory: evidence from mesocosm experiments and feeding preference trials. Marine Ecology Progress Series, 524: 83–94, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/524/m524p083.pdf.

Samarra:2017:IPV

[SVAV17] F. I. P. Samarra, M. Vighi, A. Aguilar, and G. A. Víkingsson. Intra-population variation in isotopic niche in herring-eating killer whales off Iceland. Marine Ecology Progress Series, 564:199–210, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/564/m564p199.pdf.

Saraiva:2012:VDE

[SvdMK^{+}12] S. Saraiva, J. van der Meer, S. A. L. M. Kooijman, R. Witbaard, C. J. M. Philippart, D. Hippler, and R. Parker. Validation of a dynamic energy budget (DEB) model for the blue mussel Mytilus edulis. Marine Ecology Progress Series, 463:141–158, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/463/m463p141.pdf.

Sandee:2016:TMP

[SVG16] S. D. Sandee, J. D. Van Hamme, and L. A. Gosselin. Testing microbial pathogens as a cause of early juvenile mortality in wild populations of benthic invertebrates. Marine Ecology Progress Series, 562:53–63, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/562/m562p053.pdf.

Skov:2010:PPG

[SVIH^{+}10] M. W. Skov, M. Volkelt-Igoe, S. J. Hawkins, B. Jesus, R. C. Thompson, and C. P. Doncaster. Past and present grazing boosts the photo-autotrophic biomass of biofilms.

Marine Ecology Progress Series, 401:101–111, 2010. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m401p101.pdf.

Smale:2012:STL

[SW12] D. A. Smale and T. Wernberg. Short-term in situ warming influences early development of sessile assemblages. Marine Ecology Progress Series, 453:129–136, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/453/
m453p129.pdf.

Seeley:2018:FOH

[SW18] M. E. Seeley and B. D. Walther. Facultative oligohaline habitat use in a mobile fish inferred from scale chemistry. Marine Ecology Progress Series, 598:233–245, 2018. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m598p233.pdf.

Souter:2010:LDA

[SWB^{+}10] P. Souter, B. L. Willis, L. K. Bay, M. J. Caley, A. Muirhead, and M. J. H. van Oppen. Location and disturbance affect population genetic structure in four coral species of the genus Acropora on the Great Barrier Reef. Marine Ecology Progress Series, 416:35–45, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/416/m416p035.pdf.

Smallhorn-West:2017:HMC

[SWBMJ17] P. F. Smallhorn-West, T. C. L. Bridge, P. L. Munday, and G. P. Jones. Habitat morphology constrains the depth distribution and growth rate of a coral-associated reef fish. Marine Ecology Progress Series, 576:43–53, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/576/
m576p043.pdf.

Sala:2012:FEM

[SWFQ12] J. E. Sala, R. P. Wilson, E. Frere, and F. Quintana. Foraging effort in Magellanic penguins in coastal Patagonia, Argentina. Marine Ecology Progress Series, 464:273–287, 2012.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m464p273.pdf.

Soldevilla:2011:RPW

[SWH^{+}11] M. S. Soldevilla, S. M. Wiggins, J. A. Hildebrand, E. M. Oleson, and M. C. Ferguson. Risso’s and Pacific white-sided dolphin habitat modeling from passive acoustic monitoring. Marine Ecology Progress Series, 423:247–260, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/423/
m423p247.pdf.

Shamblin:2018:MSA

[SWH^{+}18] B. M. Shamblin, B. E. Witherington, S. Hirama, R. F. Hardy, and C. J. Nairn. Mixed stock analyses indicate population-scale connectivity effects of active dispersal by surface-pelagic green turtles. Marine Ecology Progress Series, 601:215–226, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/601/m601p215.pdf.

Sivle:2016:NSD

[SWK^{+}16] L. D. Sivle, P. J. Wensveen, P. H. Kvadsheim, F. P. A. Lam, F. Visser, C. Curé, C. M. Harris, P. L. Tyack, and P. J. O. Miller. Naval sonar disrupts foraging in humpback whales. Marine Ecology Progress Series, 562:211–220, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/562/
m562p211.pdf.

Schlaefer:2018:SBC

[SWK18] J. A. Schlaefer, E. Wolanski, and M. J. Kingsford. Swimming behaviour can maintain localised jellyfish (Chironex fleckeri: Cubozoa) populations. Marine Ecology Progress Series, 591:287–302, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/591/m591p287.pdf.

Schade:2019:EFS

[SWK19] F. M. Schade, P. Weist, and U. Krumme. Evaluation of four stock discrimination methods to assign individuals from mixed-stock fisheries using genetically validated baseline samples. Marine Ecology Progress Series, 627:125–139, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2019/627/m627p125.pdf.

Scharffenberg:2019:EDB

[SWM^{+}19] K. Scharffenberg, D. Whalen, M. Marcoux, J. Iacozza, G. Davoren, and L. Loseto. Environmental drivers of beluga whale Delphinapterus leucas habitat use in the Mackenzie Estuary, Northwest Territories, Canada. Marine Ecology Progress Series, 626:209–226, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/626/m626p209.pdf.

Sams:2015:ILP

[SWMK15] M. A. Sams, F. Warren-Myers, and M. J. Keough. Increased larval planktonic duration and post-recruitment competition influence survival and growth of the bryozoan Watersipora subtorquata. Marine Ecology Progress Series, 531: 179–191, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/531/m531p179.pdf.

Squires:2014:MPN

[SWNSF14] Z. E. Squires, B. B. M. Wong, M. D. Norman, and D. Stuart-Fox. Multiple paternity but no evidence of biased sperm use in female dumpling squid Euprymna tasmanica. Marine Ecology Progress Series, 511:93–103, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/511/m511p093.pdf.

Semmens:2013:SBO

[SWP^{+}13] Brice X. Semmens, Eric J. Ward, Andrew C. Parnell, Donald L. Phillips, Stuart Bearhop, Richard Inger, Andrew Jackson, and Jonathan W. Moore. Statistical basis and outputs of stable isotope mixing models: Comment on Fry (2013). Marine Ecology Progress Series, 490:285–289, 2013. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_

oa/m490p285.pdf. See [Fry13a] and reply [Fry13b].

Studer:2013:LSP

[SWPK13] A. Studer, M. Widmann, R. Poulin, and M. Krkošek. Large scale patterns of trematode parasitism in a bivalve host: no evidence for a latitudinal gradient in infection levels. Marine Ecology Progress Series, 491:125–135, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/491/
m491p125.pdf.

Seiden:2011:BDB

[SWR11] J. M. Seiden, C. J. Way, and R. B. Rivkin. Bacterial dynamics in ballast water during trans-oceanic voyages of bulk carriers: environmental controls. Marine Ecology Progress Series, 436:145–159, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m436p145.pdf.

Standish:2011:SPN

[SWW11] J. D. Standish, J. W. White, and R. R. Warner. Spatial pattern of natal signatures in the otoliths of juvenile kelp rockfish along the Californian coast. Marine Ecology Progress Series, 437:279–290, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/437/m437p279.pdf.

Sieg:2013:MCD

[SWWK13] R. D. Sieg, D. Willey, K. Wolfe, and J. Kubanek. Multiple chemical defenses produced by Spartina alterniflora deter farming snails and their fungal crop. Marine Ecology Progress Series, 488:35–49, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/488/m488p035.pdf.

Shiao:2010:NOP

[SWY^{+}10] J. C. Shiao, S. W. Wang, K. Yokawa, M. Ichinokawa, Y. Takeuchi, Y. G. Chen, and C. C. Shen. Natal origin of Pacific bluefin tuna Thunnus orientalis inferred from otolith oxygen isotope composition. Marine Ecology Progress Series, 420:207–219, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/420/m420p207.pdf.

Sato:2010:ISS

[SY10] T. Sato and K. Yoseda. Influence of size- and sex-biased harvesting on reproduction of the coconut crab Birgus latro. Marine Ecology Progress Series, 402:171–178, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/402/
m402p171.pdf.

Suzuki:2013:EUS

[SYK^{+}13] G. Suzuki, H. Yamashita, S. Kai, T. Hayashibara, K. Suzuki, Y. Iehisa, W. Okada, W. Ando, and T. Komori. Early uptake of specific symbionts enhances the post-settlement survival of Acropora corals. Marine Ecology Progress Series, 494:149–158, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m494p149.pdf.

Smith:2016:RPS

[SYM^{+}16] T. M. Smith, P. H. York, P. I. Macreadie, M. J. Keough, D. J. Ross, and C. D. H. Sherman. Recovery pathways from small-scale disturbance in a temperate Australian seagrass. Marine Ecology Progress Series, 542:97–108, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/542/
m542p097.pdf.

Sun:2010:SEC

[SYX^{+}10] L. Sun, Y. Yang, T. Xian, Z. Lu, and Y. Fu. Strong enhancement of chlorophyll-a concentration by a weak typhoon. Marine Ecology Progress Series, 404:39–50, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/404/
m404p039.pdf.

Schaefer:2013:AIG

[SZ13] G. Schaefer and M. Zimmer. Ability of invasive green crabs to handle prey in a recently colonized region. Marine Ecology Progress Series, 483:221–229, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/483/m483p221.pdf.

Staaf:2011:ETE

[SZG11] D. J. Staaf, L. D. Zeidberg, and W. F. Gilly. Effects of temperature on embryonic development of the Humboldt squid Dosidicus gigas. Marine Ecology Progress Series, 441:165– 175, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/441/m441p165.pdf.

Teglhus:2015:DSA

[TAAN15] F. W. Teglhus, M. D. Agersted, H. Akther, and T. G. Nielsen. Distributions and seasonal abundances of krill eggs and larvae in the sub-Arctic Godthåbsfjord, SW Greenland. Marine Ecology Progress Series, 539:111–125, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/539/
m539p111.pdf.

Thompson:2014:DLF

[TAB^{+}14] A. R. Thompson, T. D. Auth, R. D. Brodeur, N. M. Bowlin, and W. Watson. Dynamics of larval fish assemblages in the California Current system: a comparative study between Oregon and southern California. Marine Ecology Progress Series, 506:193–212, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/506/m506p193.pdf.

Teixido:2011:HSB

[TAEB^{+}11] N. Teixidó, A. Albajes-Eizagirre, D. Bolbo, E. Le Hir, M. Demestre, J. Garrabou, L. Guigues, J. M. Gili, J. Piera, T. Prelot, and A. Soria-Frisch. Hierarchical segmentationbased software for cover classification analyses of seabed images (Seascape). Marine Ecology Progress Series, 431:45– 53, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m431p045.pdf.

Thomas:2010:IFC

[TB10] C. R. Thomas and L. K. Blum. Importance of the fiddler crab Uca pugnax to salt marsh soil organic matter accumulation. Marine Ecology Progress Series, 414:167–177, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/414/
m414p167.pdf.

TenBrink:2012:RPA

[TB12] T. T. TenBrink and T. W. Buckley. Resource partitioning among Myoxocephalus sculpins, and their predator–prey relationships with Chionoecetes crabs in the eastern Bering Sea. Marine Ecology Progress Series, 464:221–235, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/464/
m464p221.pdf.

Tangwancharoen:2014:ELS

[TB14a] S. Tangwancharoen and R. S. Burton. Early life stages are not always the most sensitive: heat stress responses in the copepod Tigriopus californicus. Marine Ecology Progress Series, 517:75–83, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/517/m517p075.pdf.

Tricas:2014:ABH

[TB14b] T. C. Tricas and K. S. Boyle. Acoustic behaviors in Hawaiian coral reef fish communities. Marine Ecology Progress Series, 511:1–16, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Thomisch:2019:TPA

[TBB^{+}19] K. Thomisch, O. Boebel, J. Bachmann, D. Filun, S. Neumann, S. Spiesecke, and I. Van Opzeeland. Temporal patterns in the acoustic presence of baleen whale species in a presumed breeding area off Namibia. Marine Ecology Progress Series, 620:201–214, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/620/m620p201.pdf.

Tilstra:2017:SAD

[TBC^{+}17] A. Tilstra, V. N. Bednarz, U. Cardini, N. van Hoytema, M. M. D. Al-Rshaidat, and C. Wild. Seasonality affects dinitrogen fixation associated with two common macroalgae from a coral reef in the northern Red Sea. Marine Ecology Progress Series, 575:69–80, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/575/m575p069.pdf.

Tamburello:2012:VSS

[TBCG^{+}12] L. Tamburello, L. Benedetti-Cecchi, G. Ghedini, T. Alestra, and F. Bulleri. Variation in the structure of subtidal landscapes in the NW Mediterranean Sea. Marine Ecology Progress Series, 457:29–41, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m457p029.pdf.

Toole:2011:SIV

[TBDM11] C. L. Toole, R. D. Brodeur, C. J. Donohoe, and D. F. Markle. Seasonal and interannual variability in the community structure of small demersal fishes off the central Oregon coast. Marine Ecology Progress Series, 428:201–217, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/428/
m428p201.pdf.

Teske:2015:PDA

[TBG15] P. R. Teske, S. Bader, and T. Rao Golla. Passive dispersal against an ocean current. Marine Ecology Progress Series, 539:153–163, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/539/m539p153.pdf.

Tewfik:2017:BRZ

[TBG^{+}17] A. Tewfik, E. A. Babcock, J. Gibson, V. R. Burns Perez, and S. Strindberg. Benefits of a replenishment zone revealed through trends in focal species at Glover’s Atoll, Belize. Marine Ecology Progress Series, 580:37–56, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/580/
m580p037.pdf.

Tremblay-Boyer:2011:MEF

[TBGW^{+}11] L. Tremblay-Boyer, D. Gascuel, R. Watson, V. Christensen, and D. Pauly. Modelling the effects of fishing on the biomass of the world’s oceans from 1950 to 2006. Marine Ecology Progress Series, 442:169–185, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m442p169.pdf. See erra-

tum [Ano12d].

Tiwari:2010:ECC

[TBH10] M. Tiwari, G. H. Balazs, and S. Hargrove. Estimating carrying capacity at the green turtle nesting beach of East Island, French Frigate shoals. Marine Ecology Progress Series, 419:289–294, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/419/m419p289.pdf.

Trimborn:2017:ISA

[TBH^{+}17] S. Trimborn, T. Brenneis, C. J. M. Hoppe, L. M. Laglera, L. Norman, J. Santos-Echeandía, C. Völkner, D. Wolf-Gladrow, and C. S. Hassler. Iron sources alter the response of Southern Ocean phytoplankton to ocean acidification. Marine Ecology Progress Series, 578:35–50, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m578p035.pdf.

Tamdrari:2018:EAA

[TBH^{+}18] H. Tamdrari, H. P. Benoît, J. M. Hanson, H. Bourdages, and J. C. Brêthes. Environmental associations and assemblage structure of shrimp species in the Gulf of St. Lawrence (Canada) following dramatic increases in abundance. Marine Ecology Progress Series, 596:95–112, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/596/
m596p095.pdf.

Taylor:2010:IDI

[TBKG10] S. L. Taylor, M. J. Bishop, B. P. Kelaher, and T. M. Glasby. Impacts of detritus from the invasive alga Caulerpa taxifolia on a soft sediment community. Marine Ecology Progress Series, 420:73–81, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/420/m420p073.pdf.

Towle:2016:PHC

[TBL16] E. K. Towle, A. C. Baker, and C. Langdon. Preconditioning to high CO_{2} exacerbates the response of the Caribbean branching coral Porites porites to high temperature stress. Marine Ecology Progress Series, 546:75–84, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/546/
m546p075.pdf.

Tulp:2012:BSA

[TBMdV12] I. Tulp, L. J. Bolle, E. Meesters, and P. de Vries. Brown shrimp abundance in northwest European coastal waters from 1970 to 2010 and potential causes for contrasting trends. Marine Ecology Progress Series, 458:141–154, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/458/
m458p141.pdf.

Talwar:2017:SPR

[TBMG17] B. Talwar, E. J. Brooks, J. W. Mandelman, and R. D. Grubbs. Stress, post-release mortality, and recovery of commonly discarded deep-sea sharks caught on longlines. Marine Ecology Progress Series, 582:147–161, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/582/m582p147.pdf.

Tai:2011:TSD

[TBP11] V. Tai, R. S. Burton, and B. Palenik. Temporal and spatial distributions of marine Synechococcus in the Southern California Bight assessed by hybridization to bead-arrays. Marine Ecology Progress Series, 426:133–147, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/426/
m426p133.pdf.

Thomsen:2014:IMI

[TBS^{+}14] M. S. Thomsen, J. E. Byers, D. R. Schiel, J. F. Bruno, J. D. Olden, T. Wernberg, and B. R. Silliman. Impacts of marine invaders on biodiversity depend on trophic position and functional similarity. Marine Ecology Progress Series, 495:39–47, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m495p039.pdf.

Thyrring:2017:RAT

[TBS^{+}17] J. Thyrring, M. E. Blicher, J. G. Sørensen, S. Wegeberg, and M. K. Sejr. Rising air temperatures will increase intertidal mussel abundance in the Arctic. Marine Ecology Progress Series, 584:91–104, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m584p091.pdf.

Thrush:2011:HDR

[TCA^{+}11] S. F. Thrush, M. Chiantore, V. Asnagi, J. Hewitt, D. Fiorentino, and R. Cattaneo-Vietti. Habitat-diversity relationships in rocky shore algal turf infaunal communities. Marine Ecology Progress Series, 424:119–132, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/424/m424p119.pdf.

Tardin:2019:EAA

[TCJ^{+}19] R. H. Tardin, Y. Chun, C. N. Jenkins, I. S. Maciel, S. M. Simão, and M. A. S. Alves. Environment and anthropogenic activities influence cetacean habitat use in southeastern Brazil. Marine Ecology Progress Series, 616:197–210, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/616/
/m616p197.pdf.

Torre:2019:PCS

[TCKY19] D. M. Torre, K. J. Coyne, K. D. Kroeger, and J. K. York. Phytoplankton community structure response to groundwaterborne nutrients in the Inland Bays, Delaware, USA. Marine Ecology Progress Series, 624:51–63, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/624/
m624p051.pdf.

Tarnowska:2010:CGP

[TCN^{+}10] K. Tarnowska, A. Chenuil, R. Nikula, J. P. Féral, and M. Wołowicz. Complex genetic population structure of the bivalve Cerastoderma glaucum in a highly fragmented lagoon habitat. Marine Ecology Progress Series, 406:173– 184, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/406/m406p173.pdf.

Tardin:2017:MHU

[TCSA17] R. H. Tardin, Y. Chun, S. M. Simão, and M. A. S. Alves. Modeling habitat use by Bryde’s whale Balaenoptera edeni off southeastern Brazil. Marine Ecology Progress Series, 576: 89–103, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/576/m576p089.pdf.

Thiebot:2011:IPS

[TCTB11] J. B. Thiebot, Y. Cherel, P. N. Trathan, and C. A. Bost. Interpopulation segregation in the wintering areas of macaroni penguins. Marine Ecology Progress Series, 421:279–290, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m421p279.pdf.

Tsang:2013:WMW

[TCWC13] L. M. Tsang, B. K. K. Chan, G. A. Williams, and K. H. Chu. Who is moving where? Molecular evidence reveals patterns of range shift in the acorn barnacle Hexechamaesipho pilsbryi in Asia. Marine Ecology Progress Series, 488: 187–200, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/488/m488p187.pdf.

Thiers:2014:FZT

[TDB^{+}14] L. Thiers, K. Delord, C. Barbraud, R. A. Phillips, D. Pinaud, and H. Weimerskirch. Foraging zones of the two sibling species of giant petrels in the Indian Ocean throughout the annual cycle: implication for their conservation. Marine Ecology Progress Series, 499:233–248, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/499/m499p233.pdf.

Thornhill:2010:HHA

[TDKS10] D. J. Thornhill, K. Doubleday, D. W. Kemp, and S. R. Santos. Host hybridization alters specificity of cnidariandinoflagellate associations. Marine Ecology Progress Series, 420:113–123, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/420/m420p113.pdf.

Thomson:2016:ICC

[TDM16] P. G. Thomson, A. T. Davidson, and L. Maher. Increasing CO_{2} changes community composition of pico- and nanosized protists and prokaryotes at a coastal Antarctic site. Marine Ecology Progress Series, 554:51–69, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/554/
m554p051.pdf.

Trzcinski:2013:VOC

[TDPS13] M. K. Trzcinski, E. Devred, T. Platt, and S. Sathyendranath. Variation in ocean colour may help predict cod and haddock recruitment. Marine Ecology Progress Series, 491:187– 197, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/491/m491p187.pdf.

Turner:2017:EID

[TdRH17] B. C. Turner, C. E. de Rivera, and M. E. Hepner. Examining inducible defenses to novel predators across native and introduced populations. Marine Ecology Progress Series, 574:13–27, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/574/m574p013.pdf.

Trebilco:2015:RHC

[TDSS15] R. Trebilco, N. K. Dulvy, H. Stewart, and A. K. Salomon. The role of habitat complexity in shaping the size structure of a temperate reef fish community. Marine Ecology Progress Series, 532:197–211, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2015/532/m532p197.pdf.

TinHan:2014:RSM

[TEAO^{+}14] T. TinHan, B. Erisman, O. Aburto-Oropeza, A. Weaver, D. Vázquez-Arce, and C. G. Lowe. Residency and seasonal movements in Lutjanus argentiventris and Mycteroperca rosacea at Los Islotes Reserve, Gulf of California. Marine Ecology Progress Series, 501:191–206, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/501/
m501p191.pdf.

Taylor:2012:UMG

[TF12] S. A. Taylor and V. L. Friesen. Use of molecular genetics for understanding seabird evolution, ecology and conservation. Marine Ecology Progress Series, 451:285–304, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m451p285.pdf.

Teira:2012:RTM

[TFÁS^{+}12] E. Teira, A. Fernández, X. A. Álvarez-Salgado, E. E. García-Martín, P. Serret, and C. Sobrino. Response of two marine bacterial isolates to high CO_{2} concentration. Marine Ecology Progress Series, 453:27–36, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/453/m453p027.pdf.

Thorne:2017:MFB

[TFB^{+}17] L. H. Thorne, H. J. Foley, R. W. Baird, D. L. Webster, Z. T. Swaim, and A. J. Read. Movement and foraging behavior of short-finned pilot whales in the Mid-Atlantic Bight: importance of bathymetric features and implications for management. Marine Ecology Progress Series, 584:245–257, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m584p245.pdf.

Talaber:2018:PPC

[TFFPM18] I. Talaber, J. Francé, V. Flander-Putrle, and P. Mozetič. Primary production and community structure of coastal phytoplankton in the Adriatic Sea: insights on taxon-specific productivity. Marine Ecology Progress Series, 604:65–81, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2018/604/m604p065.pdf.

Twiname:2019:MMT

[TFH^{+}19] S. Twiname, Q. P. Fitzgibbon, A. J. Hobday, C. G. Carter, and G. T. Pecl. Multiple measures of thermal performance of early stage eastern rock lobster in a fast-warming ocean region. Marine Ecology Progress Series, 624:1–11, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Tilves:2018:TIJ

[TFM^{+}18] U. Tilves, V. L. Fuentes, G. Milisenda, C. C. Parrish, S. Vizzini, and A. Sabatés. Trophic interactions of the jellyfish Pelagia noctiluca in the NW Mediterranean: evidence from stable isotope signatures and fatty acid composition. Marine Ecology Progress Series, 591:101–116, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/591/
m591p101.pdf.

Taniguchi:2014:PBS

[TFP14] D. A. A. Taniguchi, P. J. S. Franks, and F. J. Poulin. Planktonic biomass size spectra: an emergent property of size-dependent physiological rates, food web dynamics, and nutrient regimes. Marine Ecology Progress Series, 514:13– 33, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/514/m514p013.pdf.

Thayer:2014:CCC

[TFS14] J. A. Thayer, J. C. Field, and W. J. Sydeman. Changes in California Chinook salmon diet over the past 50 years: relevance to the recent population crash. Marine Ecology Progress Series, 498:249–261, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/498/m498p249.pdf.

Tyson:2012:SMC

[TFW^{+}12] R. B. Tyson, A. S. Friedlaender, C. Ware, A. K. Stimpert, and D. P. Nowacek. Synchronous mother and calf foraging behaviour in humpback whales Megaptera novaeangliae: insights from multi-sensor suction cup tags. Marine Ecology Progress Series, 457:209–220, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m457p209.pdf.

Tang:2010:AEC

[TG10] Y. Z. Tang and C. J. Gobler. Allelopathic effects of Cochlodinium polykrikoides isolates and blooms from the estuaries of Long Island, New York, on co-occurring phytoplankton. Marine Ecology Progress Series, 406:19–31, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/406/
m406p019.pdf.

Talmage:2012:ECH

[TG12a] S. C. Talmage and C. J. Gobler. Effects of CO_{2} and the harmful alga Aureococcus anophagefferens on growth and survival of oyster and scallop larvae. Marine Ecology Progress Series, 464:121–134, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/464/m464p121.pdf.

Toscano:2012:PCS

[TG12b] B. J. Toscano and B. D. Griffen. Predatory crab size diversity and bivalve consumption in oyster reefs. Marine Ecology Progress Series, 445:65–74, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/445/m445p065.pdf.

Trudnowska:2016:PPP

[TGBM^{+}16] E. Trudnowska, M. Gluchowska, A. Beszczynska-Möller, K. Blachowiak-Samolyk, and S. Kwasniewski. Plankton patchiness in the Polar Front region of the West Spitsbergen Shelf. Marine Ecology Progress Series, 560:1–18, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Thomson:2015:IBD

[TGH15] J. A. Thomson, A. Gulick, and M. R. Heithaus. Intraspecific behavioral dynamics in a green turtle Chelonia mydas foraging aggregation. Marine Ecology Progress Series, 532: 243–256, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/532/m532p243.pdf.

Tugores:2011:HSM

[TGI^{+}11] M. P. Tugores, M. Giannoulaki, M. Iglesias, A. Bonanno, V. Ti č ina, I. Leonori, A. Machias, K. Tsagarakis, N. Díaz, A. Giráldez, B. Patti, A. De Felice, G. Basilone, and V. Valavanis. Habitat suitability modelling for sardine Sardina pilchardus in a highly diverse ecosystem: the Mediterranean Sea. Marine Ecology Progress Series, 443:181–205, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2011/443/m443p181.pdf.

Tchernov:2014:IFC

[TGI14] D. Tchernov, D. F. Gruber, and A. Irwin. Isotopic fractionation of carbon in the coccolithophorid Emiliania huxleyi. Marine Ecology Progress Series, 508:53–66, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/508/
m508p053.pdf.

Todd:2010:OFC

[TGLC10] P. A. Todd, J. R. Guest, J. Lu, and L. M. Chou. One in four citations in marine biology papers is inappropriate. Marine Ecology Progress Series, 408:299–303, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/408/
m408p299.pdf.

Tanaka:2017:ENP

[TGM^{+}17] Y. Tanaka, A. G. Grottoli, Y. Matsui, A. Suzuki, and K. Sakai. Effects of nitrate and phosphate availability on the tissues and carbonate skeleton of scleractinian corals. Marine Ecology Progress Series, 570:101–112, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/570/m570p101.pdf.

Torres:2016:PCD

[TGP^{+}16] G. Torres, L. Giménez, A. K. Pettersen, M. Bue, M. T. Burrows, and S. R. Jenkins. Persistent and context-dependent effects of the larval feeding environment on post-metamorphic performance through the adult stage. Marine Ecology Progress Series, 545:147–160, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/545/m545p147.pdf.

Tixier:2019:ITD

[TGR^{+}19] P. Tixier, J. Giménez, R. R. Reisinger, P. Méndez-Fernandez, J. P. Y. Arnould, Y. Cherel, and C. Guinet. Importance of toothfish in the diet of generalist subantarctic killer whales: implications for fisheries interactions. Marine Ecology Progress Series, 613:197–210, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m613p197.pdf.

Temming:2017:ULC

[TGRH17] A. Temming, C. Günther, C. Rückert, and M. Hufnagl. Understanding the life cycle of North Sea brown shrimp Crangon crangon: a simulation model approach. Marine Ecology Progress Series, 584:119–143, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/584/m584p119.pdf.

Tsagarakis:2012:VPE

[TGSM12] K. Tsagarakis, M. Giannoulaki, S. Somarakis, and A. Machias. Variability in positional, energetic and morphometric descriptors of European anchovy Engraulis encrasicolus schools related to patterns of diurnal vertical migration. Marine Ecology Progress Series, 446:243–258, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m446p243.pdf.

Tran:2011:LPD

[TH11] C. Tran and M. G. Hadfield. Larvae of Pocillopora damicornis (Anthozoa) settle and metamorphose in response to surface-biofilm bacteria. Marine Ecology Progress Series, 433: 85–96, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/433/m433p085.pdf.

Thomson:2012:SSD

[THB^{+}12] J. A. Thomson, M. R. Heithaus, D. A. Burkholder, J. J. Vaudo, A. J. Wirsing, and L. M. Dill. Site specialists, diet generalists? Isotopic variation, site fidelity, and foraging by loggerhead turtles in Shark Bay, Western Australia. Marine Ecology Progress Series, 453:213–226, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/453/
m453p213.pdf.

Tiselius:2012:FAT

[THC12] P. Tiselius, B. W. Hansen, and D. Calliari. Fatty acid transformation in zooplankton: from seston to benthos. Marine Ecology Progress Series, 446:131–144, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m446p131.pdf.

terHofstede:2010:RWC

[tHHR10] R. ter Hofstede, J. G. Hiddink, and A. D. Rijnsdorp. Regional warming changes fish species richness in the eastern North Atlantic Ocean. Marine Ecology Progress Series, 414:1–9, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Talapatra:2013:CBI

[THM^{+}13] S. Talapatra, J. Hong, M. McFarland, A. R. Nayak, C. Zhang, J. Katz, J. Sullivan, M. Twardowski, J. Rines, and P. Donaghay. Characterization of biophysical interactions in the water column using in situ digital holography. Marine Ecology Progress Series, 473:29–51, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/473/m473p029.pdf.

Teira:2015:LDP

[THMF^{+}15] E. Teira, V. Hernando-Morales, A. Fernández, S. Martínez-García, X. A. Álvarez-Salgado, A. Bode, and M. M. Varela. Local differences in phytoplankton–bacterioplankton coupling in the coastal upwelling off Galicia (NW Spain). Marine Ecology Progress Series, 528:53–69, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m528p053.pdf.

Thorson:2015:STV

[Tho15] J. T. Thorson. Spatio-temporal variation in fish condition is not consistently explained by density, temperature, or season for California Current groundfishes. Marine Ecology Progress Series, 526:101–112, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2015/526/m526p101.pdf.

Tarling:2016:GSA

[THP^{+}16] G. A. Tarling, S. Hill, H. Peat, S. Fielding, C. Reiss, and A. Atkinson. Growth and shrinkage in Antarctic krill Euphausia superba is sex-dependent. Marine Ecology Progress Series, 547:61–78, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m547p061.pdf.

Teira:2016:BRR

[THRBL^{+}16] E. Teira, M. Hernández-Ruiz, E. Barber-Lluch, C. Sobrino, I. G. Teixeira, X. A. Álvarez-Salgado, M. Nieto-Cid, S. Martínez-García, F. G. Figueiras, and E. Fernández. Bacterioplankton responses to riverine and atmospheric inputs in a coastal upwelling system (Ría de Vigo, NW Spain). Marine Ecology Progress Series, 542:39–50, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/542/
m542p039.pdf.

Thompson:2016:RAS

[THW^{+}16] A. R. Thompson, J. R. Hyde, W. Watson, D. C. Chen, and L. W. Guo. Rockfish assemblage structure and spawning locations in southern California identified through larval sampling. Marine Ecology Progress Series, 547:177–192, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/547/
m547p177.pdf.

Takahashi:2018:MMW

[TIN^{+}18] A. Takahashi, M. Ito, K. Nagai, J. B. Thiebot, H. Mitamura, T. Noda, P. N. Trathan, T. Tamura, and Y. Y. Watanabe. Migratory movements and winter diving activity of Adélie penguins in East Antarctica. Marine Ecology Progress Series, 589:227–239, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/589/m589p227.pdf.

Takahashi:2015:MMR

[TIS^{+}15] A. Takahashi, M. Ito, Y. Suzuki, Y. Watanuki, J. B. Thiebot, T. Yamamoto, T. Iida, P. Trathan, Y. Niizuma, and T. Kuwae. Migratory movements of rhinoceros auklets in the northwestern Pacific: connecting seasonal productivities. Marine Ecology Progress Series, 525:229–243, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/525/
m525p229.pdf.

Thyrring:2017:GIP

[TJS17] J. Thyrring, K. T. Jensen, and M. K. Sejr. Gametogenesis of an intertidal population of Mytilus trossulus in NW Greenland: not a limitation for potential Arctic range expansion. Marine Ecology Progress Series, 574:65–74, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/574/
m574p065.pdf.

Tafur:2010:RBJ

[TKA10] R. Tafur, F. Keyl, and J. Argüelles. Reproductive biology of jumbo squid Dosidicus gigas in relation to environmental variability of the northern Humboldt Current System. Marine Ecology Progress Series, 400:127–141, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/400/
m400p127.pdf.

Thomas:2013:FSU

[TKB^{+}13] F. I. M. Thomas, L. T. Kregting, B. D. Badgley, M. J. Donahue, and P. O. Yund. Fertilization in a sea urchin is not only a water column process: effects of water flow on fertilization near a spawing female. Marine Ecology Progress Series, 494:231–240, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m494p231.pdf.

Thompson:2010:PCB

[TKBN10] S. A. Thompson, H. Knoll, C. A. Blanchette, and K. J. Nielsen. Population consequences of biomass loss due to commercial collection of the wild seaweed Postelsia palmaeformis. Marine Ecology Progress Series, 413:17–31, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m413p017.pdf.

Traiger:2016:SOV

[TKDM16] S. B. Traiger, B. Konar, A. Doroff, and L. McCaslin. Sea otters versus sea stars as major clam predators: evidence from foraging pits and shell litter. Marine Ecology Progress Series, 560:73–86, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/560/m560p073.pdf.

Tamburin:2019:INR

[TKEV^{+}19] E. Tamburin, S. L. Kim, F. R. Elorriaga-Verplancken, D. J. Madigan, M. Hoyos-Padilla, A. Sánchez-González, A. Hernández-Herrera, J. L. Castillo-Geniz, C. J. Godinez-Padilla, and F. Galván-Magaña. Isotopic niche and resource sharing among young sharks (Carcharodon carcharias and Isurus oxyrinchus) in Baja California, Mexico. Marine Ecology Progress Series, 613:107–124, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/613/m613p107.pdf.

Tomiyama:2010:SVA

[TKIO10] T. Tomiyama, N. Komizunai, K. Ito, and M. Omori. Spatial variation in the abundance and condition of the bivalve Nuttallia olivacea in relation to environmental factors and sublethal predation. Marine Ecology Progress Series, 406: 185–196, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/406/m406p185.pdf.

Treude:2011:EEC

[TKL^{+}11] T. Treude, S. Kiel, P. Linke, J. Peckmann, and J. L. Goedert. Elasmobranch egg capsules associated with modern and ancient cold seeps: a nursery for marine deep-water predators. Marine Ecology Progress Series, 437:175–181, 2011. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m437p175.pdf.

Thomas:2014:PGS

[TKS^{+}14] L. Thomas, G. A. Kendrick, M. Stat, K. L. Travaille, G. Shedrawi, and W. J. Kennington. Population genetic structure of the Pocillopora damicornis morphospecies along Ningaloo Reef, Western Australia. Marine Ecology Progress Series, 513:111–119, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/513/m513p111.pdf.

Thorsen:2019:FDO

[TKV^{+}19] S. W. Thorsen, E. Kristensen, T. Valdemarsen, M. R. Flindt, C. O. Quintana, and M. Holmer. Fertilizer-derived N in opportunistic macroalgae after flooding of agricultural land. Marine Ecology Progress Series, 616:37–49, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m616p037.pdf.

Tait:2014:STV

[TKW^{+}14] K. Tait, V. Kitidis, B. B. Ward, D. G. Cummings, M. R. Jones, P. J. Somerfield, and S. Widdicombe. Spatio-temporal variability in ammonia oxidation and ammonia-oxidising bacteria and archaea in coastal sediments of the western English Channel. Marine Ecology Progress Series, 511:41–58, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/511/
m511p041.pdf.

Taylor:2018:PBS

[TL18] A. G. Taylor and M. R. Landry. Phytoplankton biomass and size structure across trophic gradients in the southern California Current and adjacent ocean ecosystems. Marine Ecology Progress Series, 592:1–17, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Tilley:2013:ESH

[TLAT13] A. Tilley, J. López-Angarita, and J. R. Turner. Effects of scale and habitat distribution on the movement of the southern stingray Dasyatis americana on a Caribbean atoll. Marine Ecology Progress Series, 482:169–179, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/482/
m482p169.pdf.

Tableau:2015:ABE

[TLB15] A. Tableau, H. Le Bris, and A. Brind’Amour. Available benthic energy coefficient (ABEC): a generic tool to estimate the food profitability in coastal fish nurseries. Marine Ecology Progress Series, 522:203–218, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/522/m522p203.pdf.

Taniguchi:2014:SSG

[TLFS14] D. A. A. Taniguchi, M. R. Landry, P. J. S. Franks, and K. E. Selph. Size-specific growth and grazing rates for picophytoplankton in coastal and oceanic regions of the eastern Pacific. Marine Ecology Progress Series, 509:87–101, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/509/
m509p087.pdf.

Treible:2014:RCM

[TLG14] L. M. Treible, D. J. Lonsdale, and C. J. Gobler. Role of the ctenophore Mnemiopsis leidyi in nutrient cycling in Long Island Sound, New York, USA. Marine Ecology Progress Series, 510:215–227, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/510/m510p215.pdf.

Thomas:2011:HSF

[TLJ^{+}11] A. C. Thomas, M. M. Lance, S. J. Jeffries, B. G. Miner, and A. Acevedo-Gutiérrez. Harbor seal foraging response to a seasonal resource pulse, spawning Pacific herring. Marine Ecology Progress Series, 441:225–239, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/441/m441p225.pdf.

Tyne:2012:ECC

[TLK^{+}12] J. A. Tyne, N. R. Loneragan, A. M. Kopps, S. J. Allen, M. Krützen, and L. Bejder. Ecological characteristics contribute to sponge distribution and tool use in bottlenose dolphins Tursiops sp. Marine Ecology Progress Series, 444:143– 153, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/444/m444p143.pdf.

Touzet:2011:SDA

[TLM^{+}11] N. Touzet, J. P. Lacaze, M. Maher, E. Turrell, and R. Raine. Summer dynamics of Alexandrium ostenfeldii (Dinophyceae) and spirolide toxins in Cork Harbour, Ireland. Marine Ecology Progress Series, 425:21–33, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/425/m425p021.pdf.

Tableau:2019:NAT

[TLS^{+}19] A. Tableau, H. Le Bris, E. Saulnier, O. Le Pape, and A. Brind’Amour. Novel approach for testing the food limitation hypothesis in estuarine and coastal fish nurseries. Marine Ecology Progress Series, 629:117–131, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/629/
m629p117.pdf.

Taylor:2010:BAB

[TM10] B. M. Taylor and J. L. McIlwain. Beyond abundance and biomass: effects of marine protected areas on the demography of a highly exploited reef fish. Marine Ecology Progress Series, 411:243–258, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/411/m411p243.pdf.

Tagliarolo:2015:SLS

[TM15] M. Tagliarolo and C. D. McQuaid. Sub-lethal and sub-specific temperature effects are better predictors of mussel distribution than thermal tolerance. Marine Ecology Progress Series, 535:145–159, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/535/m535p145.pdf.

Tsounis:2012:AER

[TMB^{+}12] G. Tsounis, L. Martinez, L. Bramanti, N. Viladrich, J. M. Gili, A. Martinez, and S. Rossi. Anthropogenic effects on reproductive effort and allocation of energy reserves in the Mediterranean octocoral Paramuricea clavata. Marine Ecology Progress Series, 449:161–172, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/449/m449p161.pdf.

Thieltges:2013:TPO

[TMBP13] D. W. Thieltges, D. J. Marcogliese, C. A. Blanar, and R. Poulin. Trematode prevalence–occupancy relationships on regional and continental spatial scales in marine gastropod hosts. Marine Ecology Progress Series, 490:147–154, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/490/m490p147.pdf.

Teira:2011:CBP

[TMGCM11] E. Teira, S. Martínez-García, C. Carreira, and X. A. G. Morán. Changes in bacterioplankton and phytoplankton community composition in response to nutrient additions in coastal waters off the NW Iberian Peninsula. Marine Ecology Progress Series, 426:87–104, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/426/m426p087.pdf.

Tranquilla:2013:MCW

[TMH^{+}13] L. A. McFarlane Tranquilla, W. A. Montevecchi, A. Hedd, D. A. Fifield, C. M. Burke, P. A. Smith, P. M. Regular, G. J. Robertson, A. J. Gaston, and R. A. Phillips. Multiple-colony winter habitat use by murres Uria spp. in the Northwest Atlantic Ocean: implications for marine risk assessment. Marine Ecology Progress Series, 472:287–303, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/472/
m472p287.pdf.

Thorbjornsen:2018:RMP

[TMH^{+}18] S. H. Thorbjørnsen, E. Moland, M. B. O. Huserbråten, J. A. Knutsen, H. Knutsen, and E. M. Olsen. Replicated marine protected areas (MPAs) support movement of larger, but not more, European lobsters to neighbouring fished areas. Marine Ecology Progress Series, 595:123–133, 2018. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m595p123.pdf.

Thompson:2015:NCC

[TMM15] M. L. Thompson, D. J. Marshall, and K. Monro. Noncontact competition in a sessile marine invertebrate: causes and consequences. Marine Ecology Progress Series, 522:115– 125, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/522/m522p115.pdf.

Tanner:2019:HFS

[TMM^{+}19] C. E. Tanner, A. Marco, S. Martins, E. Abella-Perez, and L. A. Hawkes. Highly feminised sex-ratio estimations for the world’s third-largest nesting aggregation of loggerhead sea turtles. Marine Ecology Progress Series, 621:209–219, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/621/
m621p209.pdf.

Tillett:2011:DFE

[TMP^{+}11] B. J. Tillett, M. G. Meekan, D. Parry, N. Munksgaard, I. C. Field, D. Thorburn, and C. J. A. Bradshaw. Decoding fingerprints: elemental composition of vertebrae correlates to age-related habitat use in two morphologically similar sharks. Marine Ecology Progress Series, 434:133–142, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/434/
m434p133.pdf.

Tortell:2013:ICU

[TMP^{+}13] P. D. Tortell, M. M. Mills, C. D. Payne, M. T. Maldonado, M. Chierici, A. Fransson, A. C. Alderkamp, and K. R. Arrigo. Inorganic C utilization and C isotope fractionation by pelagic and sea ice algal assemblages along the Antarctic continental shelf. Marine Ecology Progress Series, 483: 47–66, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/483/m483p047.pdf.

Tucker:2014:FSF

[TMS14] A. D. Tucker, B. D. MacDonald, and J. A. Seminoff. Foraging site fidelity and stable isotope values of loggerhead turtles tracked in the Gulf of Mexico and northwest Caribbean. Marine Ecology Progress Series, 502:267–279, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/502/
m502p267.pdf.

Tagliarolo:2016:LTT

[TMS^{+}16] M. Tagliarolo, V. Montalto, G. Sarà, J. A. Lathlean, and C. D. McQuaid. Low temperature trumps high food availability to determine the distribution of intertidal mussels Perna perna in South Africa. Marine Ecology Progress Series, 558:51–63, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/558/m558p051.pdf.

Taylor:2019:CFE

[TMSG19] R. Taylor, M. Mather, J. Smith, and K. Gerber. Confluences function as ecological hotspots: geomorphic and regional drivers can help identify patterns of fish distribution within a seascape. Marine Ecology Progress Series, 629:133– 148, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/629/m629p133.pdf.

Testa:2012:ABM

[TMT^{+}12] J. W. Testa, K. J. Mock, C. Taylor, H. Koyuk, J. R. Coyle, and R. Waggoner. Agent-based modeling of the dynamics of mammal-eating killer whales and their prey. Marine Ecology Progress Series, 466:275–291, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/466/m466p275.pdf.

Thorburn:2015:WRS

[TNB^{+}15] J. Thorburn, F. Neat, D. M. Bailey, L. R. Noble, and C. S. Jones. Winter residency and site association in the critically endangered North East Atlantic spurdog Squalus acanthias. Marine Ecology Progress Series, 526:113–124, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/526/
m526p113.pdf.

Terazono:2012:FRE

[TNIH12] Y. Terazono, Y. Nakamura, Z. Imoto, and M. Hiraoka. Fish response to expanding tropical Sargassum beds on the temperate coasts of Japan. Marine Ecology Progress Series, 464:209–220, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/464/m464p209.pdf.

Tang:2011:MCS

[TNM^{+}11a] K. W. Tang, T. G. Nielsen, P. Munk, J. Mortensen, E. F. Møller, K. E. Arendt, K. Tönnesson, and T. Juul-Pedersen. Metazooplankton community structure, feeding rate estimates, and hydrography in a meltwater-influenced Greenlandic fjord. Marine Ecology Progress Series, 434:77–90, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2011/434/m434p077.pdf.

Tzortziou:2011:SGD

[TNM^{+}11b] M. Tzortziou, P. J. Neale, J. P. Megonigal, C. L. Pow, and M. Butterworth. Spatial gradients in dissolved carbon due to tidal marsh outwelling into a Chesapeake Bay estuary. Marine Ecology Progress Series, 426:41–56, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m426p041.pdf.

Tyson:2011:CSA

[TNN11] R. B. Tyson, S. M. Nowacek, and D. P. Nowacek. Community structure and abundance of bottlenose dolphins Tursiops truncatus in coastal waters of the northeast Gulf of Mexico. Marine Ecology Progress Series, 438:253–265, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/438/
m438p253.pdf.

Tanaka:2017:MSG

[TNOY17] M. Tanaka, T. Nagai, T. Okada, and H. Yamazaki. Measurement of sardine-generated turbulence in a large tank. Marine Ecology Progress Series, 571:207–220, 2017. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m571p207.pdf.

Tremblay:2012:EAO

[TNS^{+}12] P. Tremblay, M. S. Naumann, S. Sikorski, R. Grover, and C. Ferrier-Pagès. Experimental assessment of organic carbon fluxes in the scleractinian coral Stylophora pistillata during a thermal and photo stress event. Marine Ecology Progress Series, 453:63–77, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/453/m453p063.pdf.

Trannum:2011:BBE

[TNSN11] H. C. Trannum, H. C. Nilsson, M. T. Schaanning, and K. Norling. Biological and biogeochemical effects of organic matter and drilling discharges in two sediment communities. Marine Ecology Progress Series, 442:23–36, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/442/
m442p023.pdf.

Tsounis:2010:PCR

[TOR^{+}10] G. Tsounis, C. Orejas, S. Reynaud, J. M. Gili, D. Allemand, and C. Ferrier-Pagès. Prey-capture rates in four Mediterranean cold water corals. Marine Ecology Progress Series, 398:149–155, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m398p149.pdf.

Takagi:2010:IVP

[TOW10] K. K. Takagi, N. N. Ono, and W. G. Wright. Interspecific variation in palatability suggests cospecialization of antipredator defenses in sea hares. Marine Ecology Progress Series, 416:137–144, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/416/m416p137.pdf.

Thieltges:2016:FWB

[TP16] D. W. Thieltges and R. Poulin. Food-web-based comparison of the drivers of helminth parasite species richness in coastal fish and bird definitive hosts. Marine Ecology Progress Series, 545:9–19, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/545/m545p009.pdf.

Traisnel:2019:SCC

[TP19] G. Traisnel and L. Pichegru. Success comes with consistency in hard times: foraging repeatability relates to sex and breeding output in African penguins. Marine Ecology Progress Series, 608:279–289, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/608/m608p279.pdf.

Trapon:2013:PSG

[TPA^{+}13] M. L. Trapon, M. S. Pratchett, M. Adjeroud, A. S. Hoey, and A. H. Baird. Post-settlement growth and mortality rates of juvenile scleractinian corals in Moorea, French Polynesia versus Trunk Reef, Australia. Marine Ecology Progress Series, 488:157–170, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/488/m488p157.pdf.

Tan:2016:STV

[TPB^{+}16] C. H. Tan, M. S. Pratchett, L. K. Bay, E. M. Graham, and A. H. Baird. Spatial and temporal variation in fecundity among populations of Acropora millepora on the Great Barrier Reef. Marine Ecology Progress Series, 561: 147–153, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/561/m561p147.pdf.

Tan:2018:BHV

[TPB^{+}18] C. H. Tan, M. S. Pratchett, L. K. Bay, E. M. Graham, and A. H. Baird. Biennium horribile: very high mortality in the reef coral Acropora millepora on the Great Barrier Reef in 2009 and 2010. Marine Ecology Progress Series, 604:133–142, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/604/m604p133.pdf.

Teske:2013:DBS

[TPBM13] P. R. Teske, I. Papadopoulos, N. P. Barker, and C. D. Mc-Quaid. Dispersal barriers and stochastic reproductive success do not explain small-scale genetic structure in a broadcast spawning marine mussel. Marine Ecology Progress Series, 482:133–140, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m482p133.pdf.

Tettelbach:2013:PLP

[TPC^{+}13] S. T. Tettelbach, B. J. Peterson, J. M. Carroll, S. W. T. Hughes, D. M. Bonal, A. J. Weinstock, J. R. Europe, B. T. Furman, and C. F. Smith. Priming the larval pump: resurgence of bay scallop recruitment following initiation of intensive restoration efforts. Marine Ecology Progress Series, 478:153–172, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/478/m478p153.pdf.

Tettelbach:2015:AAS

[TPC^{+}15] S. T. Tettelbach, B. J. Peterson, J. M. Carroll, B. T. Furman, S. W. T. Hughes, J. Havelin, J. R. Europe, D. M. Bonal, A. J. Weinstock, and C. F. Smith. Aspiring to an altered stable state: rebuilding of bay scallop populations and fisheries following intensive restoration. Marine Ecology Progress Series, 529:121–136, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m529p121.pdf.

Teixeira:2016:LFS

[TPC^{+}16] S. Teixeira, G. A. Pearson, R. Candeias, C. Madeira, M. Valero, and E. A. Serrão. Lack of fine-scale genetic structure and distant mating in natural populations of Fucus vesiculosus. Marine Ecology Progress Series, 544:131– 142, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/544/m544p131.pdf.

Travers:2012:RBL

[TPCN12] M. J. Travers, I. C. Potter, K. R. Clarke, and S. J. Newman. Relationships between latitude and environmental conditions and the species richness, abundance and composition of tropical fish assemblages over soft substrata. Marine Ecology Progress Series, 446:221–241, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/446/m446p221.pdf.

Tsagaraki:2013:PRN

[TPF^{+}13] T. M. Tsagaraki, P. Pitta, C. Frangoulis, G. Petihakis, and I. Karakassis. Plankton response to nutrient enrichment is maximized at intermediate distances from fish farms. Marine Ecology Progress Series, 493:31–42, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/493/
m493p031.pdf.

Tremblay:2011:HMS

[TPFP11] P. Tremblay, A. Peirano, and C. Ferrier-Pagès. Heterotrophy in the Mediterranean symbiotic coral Cladocora caespitosa: comparison with two other scleractinian species. Marine Ecology Progress Series, 422:165–177, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/422/
m422p165.pdf.

Treible:2018:EEP

[TPKC18] L. M. Treible, K. A. Pitt, S. G. Klein, and R. H. Condon. Exposure to elevated pCO_{2} does not exacerbate reproductive suppression of Aurelia aurita jellyfish polyps in low oxygen environments. Marine Ecology Progress Series, 591:129–139, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m591p129.pdf.

Timbs:2019:CSD

[TPM19] J. R. Timbs, E. N. Powell, and R. Mann. Changes in the spatial distribution and anatomy of a range shift for the Atlantic surfclam Spisula solidissima in the Mid-Atlantic Bight and on Georges Bank. Marine Ecology Progress Series, 620: 77–97, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/620/m620p077.pdf.

Teal:2010:SML

[TPS10] L. R. Teal, E. R. Parker, and M. Solan. Sediment mixed layer as a proxy for benthic ecosystem process and function. Marine Ecology Progress Series, 414:27–40, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/414/
m414p027.pdf.

Tsiamis:2013:MCR

[TPS^{+}13] K. Tsiamis, P. Panayotidis, M. Salomidi, A. Pavlidou, J. Kleinteich, K. Balanika, and F. C. Küpper. Macroalgal community response to re-oligotrophication in Saronikos Gulf. Marine Ecology Progress Series, 472:73–85, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m472p073.pdf.

Trommer:2012:ZMN

[TPSS12] G. Trommer, P. Pondaven, M. Siccha, and H. Stibor. Zooplankton-mediated nutrient limitation patterns in marine phytoplankton: an experimental approach with natural communities. Marine Ecology Progress Series, 449:83–94, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m449p083.pdf.

Thorsteinsson:2012:CBT

[TPT^{+}12] V. Thorsteinsson, Ó. K. Pálsson, G. G. Tómasson, I. G. Jónsdóttir, and C. Pampoulie. Consistency in the behaviour types of the Atlantic cod: repeatability, timing of migration and geo-location. Marine Ecology Progress Series, 462: 251–260, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/462/m462p251.pdf.

Trujillo:2019:SFR

[TPVCV19] J. E. Trujillo, L. M. Pardo, L. Vargas-Chacoff, and N. Valdivia. Sharks in the forest: relationships between kelp physicalcomplexity attributes and egg deposition sites of the redspotted catshark. Marine Ecology Progress Series, 610:125– 135, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/610/m610p125.pdf.

Thorne:2013:FSB

[TR13] L. H. Thorne and A. J. Read. Fine-scale biophysical interactions drive prey availability at a migratory stopover site for Phalaropus spp. in the Bay of Fundy, Canada. Marine Ecology Progress Series, 487:261–273, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m487p261.pdf.

Trancart:2014:MMS

[TRA^{+}14] T. Trancart, S. Rochette, A. Acou, E. Lasne, and E. Feunteun. Modeling marine shad distribution using data from French bycatch fishery surveys. Marine Ecology Progress Series, 511:181–192, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/511/m511p181.pdf.

Thums:2018:HVM

[TRGF18] M. Thums, J. Rossendell, M. Guinea, and L. C. Ferreira. Horizontal and vertical movement behaviour of flatback turtles and spatial overlap with industrial development. Marine Ecology Progress Series, 602:237–253, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/602/m602p237.pdf.

Tepsich:2014:HPT

[TRHM14] P. Tepsich, M. Rosso, P. N. Halpin, and A. Moulins. Habitat preferences of two deep-diving cetacean species in the northern Ligurian Sea. Marine Ecology Progress Series, 508: 247–260, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m508p247.pdf.

Trenkel:2011:UAE

[TRJ^{+}11] V. M. Trenkel, P. H. Ressler, M. Jech, M. Giannoulaki, and C. Taylor. Underwater acoustics for ecosystem-based management: state of the science and proposals for ecosystem indicators. Marine Ecology Progress Series, 442:285–301, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m442p285.pdf.

Thoisen:2015:EAA

[TRL^{+}15] C. Thoisen, K. Riisgaard, N. Lundholm, T. G. Nielsen, and P. J. Hansen. Effect of acidification on an Arctic phytoplankton community from Disko Bay, West Greenland. Marine Ecology Progress Series, 520:21–34, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/520/
m520p021.pdf.

Tecchio:2011:DDM

[TRLS^{+}11] S. Tecchio, E. Ramírez-Llodra, F. Sardà, J. B. Company, I. Palomera, A. Mechó, R. Pedrosa-Pàmies, and A. Sanchez-Vidal. Drivers of deep Mediterranean megabenthos communities along longitudinal and bathymetric gradients. Marine Ecology Progress Series, 439:181–192, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/439/
m439p181.pdf.

Tracey:2011:HFC

[TRMC11] D. M. Tracey, A. A. Rowden, K. A. Mackay, and T. Compton. Habitat-forming cold-water corals show affinity for seamounts in the New Zealand region. Marine Ecology Progress Series, 430:1–22, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Tagliafico:2017:LED

[TRR^{+}17] A. Tagliafico, D. Rudd, M. S. Rangel, B. P. Kelaher, L. Christidis, K. Cowden, S. R. Scheffers, and K. Benkendorff. Lipidenriched diets reduce the impacts of thermal stress in corals. Marine Ecology Progress Series, 573:129–141, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/573/
m573p129.pdf.

Thaxter:2018:DBN

[TRSB^{+}18] C. B. Thaxter, V. H. Ross-Smith, W. Bouten, E. A. Masden, N. A. Clark, G. J. Conway, L. Barber, G. D. Clewley, and N. H. K. Burton. Dodging the blades: new insights into three-dimensional space use of offshore wind farms by lesser black-backed gulls Larus fuscus. Marine Ecology Progress Series, 587:247–253, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/587/m587p247.pdf.

Tanner:2012:OGD

[TRSV^{+}12] S. E. Tanner, P. Reis-Santos, R. P. Vasconcelos, S. França, S. R. Thorrold, and H. N. Cabral. Otolith geochemistry discriminates among estuarine nursery areas of Solea solea and S. senegalensis over time. Marine Ecology Progress Series, 452:193–203, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/452/m452p193.pdf.

Tait:2010:PPI

[TS10] L. W. Tait and D. R. Schiel. Primary productivity of intertidal macroalgal assemblages: comparison of laboratory and in situ photorespirometry. Marine Ecology Progress Series, 416:115–125, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/416/m416p115.pdf.

Tait:2011:DPN

[TS11a] L. W. Tait and D. R. Schiel. Dynamics of productivity in naturally structured macroalgal assemblages: importance of canopy structure on light-use efficiency. Marine Ecology Progress Series, 421:97–107, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/421/m421p097.pdf.

Topping:2011:SFR

[TS11b] D. T. Topping and S. T. Szedlmayer. Site fidelity, residence time and movements of red snapper Lutjanus campechanus estimated with long-term acoustic monitoring. Marine Ecology Progress Series, 437:183–200, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/437/m437p183.pdf.

Thompson:2012:RFN

[TS12] G. A. Thompson and D. R. Schiel. Resistance and facilitation by native algal communities in the invasion success of Undaria pinnatifida. Marine Ecology Progress Series, 468: 95–105, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/468/m468p095.pdf.

Tsikliras:2014:MTC

[TS14] A. C. Tsikliras and Konstantinos I. Stergiou. Mean temperature of the catch increases quickly in the Mediterranean Sea. Marine Ecology Progress Series, 515:281–284, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m515p281.pdf.

Teh:2015:TGS

[TS15] L. S. L. Teh and U. R. Sumaila. Trends in global shared fisheries. Marine Ecology Progress Series, 530:243–254, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m530p243.pdf.

Tompkins:2016:LCH

[TS16] P. A. Tompkins and D. L. Steller. Living carbonate habitats in temperate California (USA) waters: distribution, growth, and disturbance of Santa Catalina Island rhodoliths. Marine Ecology Progress Series, 560:135–145, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/560/
m560p135.pdf.

Tomaszewicz:2018:ECF

[TSA^{+}18] C. N. Turner Tomaszewicz, J. A. Seminoff, L. Avens, L. R. Goshe, J. M. Rguez-Baron, S. H. Peckham, and C. M. Kurle. Expanding the coastal forager paradigm: long-term pelagic habitat use by green turtles Chelonia mydas in the eastern Pacific Ocean. Marine Ecology Progress Series, 587:217– 234, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/587/m587p217.pdf.

Teske:2015:ISH

[TSCSB15] P. R. Teske, J. Sandoval-Castillo, M. Sasaki, and L. B. Beheregaray. Invasion success of a habitat-forming marine invertebrate is limited by lower-than-expected dispersal ability. Marine Ecology Progress Series, 536:221–227, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/536/
m536p221.pdf.

Teske:2015:SLR

[TSCvS^{+}15] P. R. Teske, J. Sandoval-Castillo, E. van Sebille, J. Waters, and L. B. Beheregaray. On-shelf larval retention limits population connectivity in a coastal broadcast spawner. Marine Ecology Progress Series, 532:1–12, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Tweddle:2010:RAU

[TSF^{+}10] J. F. Tweddle, P. G. Strutton, D. G. Foley, L. O’Higgins, A. M. Wood, B. Scott, R. C. Everroad, W. T. Peterson, D. Cannon, M. Hunter, and Z. Forster. Relationships among upwelling, phytoplankton blooms, and phycotoxins in coastal Oregon shellfish. Marine Ecology Progress Series, 405:131–145, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m405p131.pdf.

Thistle:2010:FMH

[TSGW10] M. E. Thistle, D. C. Schneider, R. S. Gregory, and N. J. Wells. Fractal measures of habitat structure: maximum densities of juvenile cod occur at intermediate eelgrass complexity. Marine Ecology Progress Series, 405:39–56, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m405p039.pdf.

Thompson:2017:LGI

[TSR17] J. A. Thompson, A. J. Stow, and D. A. Raftos. Lack of genetic introgression between wild and selectively bred Sydney rock oysters Saccostrea glomerata. Marine Ecology Progress Series, 570:127–139, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/570/m570p127.pdf.

Telesh:2011:PDD

[TSS11a] I. V. Telesh, H. Schubert, and S. O. Skarlato. Protistan diversity does peak in the horohalinicum of the Baltic Sea: Reply to Ptacnik et al. (2011). Marine Ecology Progress Series, 432:293–297, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m432p293.pdf.

Telesh:2011:RRC

[TSS11b] I. V. Telesh, H. Schubert, and S. O. Skarlato. Revisiting Remane’s concept: evidence for high plankton diversity and a protistan species maximum in the horohalinicum of the Baltic Sea. Marine Ecology Progress Series, 421:1–11, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). See comment [POL^{+}11].

Thompson:2012:PPS

[TSS^{+}12] S. A. Thompson, W. J. Sydeman, J. A. Santora, K. H. Morgan, W. Crawford, and M. T. Burrows. Phenology of pelagic seabird abundance relative to marine climate change in the Alaska Gyre. Marine Ecology Progress Series, 454: 159–170, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m454p159.pdf.

Thorson:2013:URL

[TSTP13a] J. T. Thorson, I. J. Stewart, I. G. Taylor, and A. E. Punt. Using a recruitment-linked multispecies stock assessment model to estimate common trends in recruitment for US West Coast groundfishes. Marine Ecology Progress Series, 483:245–256, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/483/m483p245.pdf.

Torres:2013:SDS

[TSTP13b] L. G. Torres, P. M. Sagar, D. R. Thompson, and R. A. Phillips. Scale-dependence of seabird-fishery data analysis and management: Reply to croxall et al. (2013). Marine Ecology Progress Series, 493:301–304, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/493/m493p301.pdf.

Torres:2013:SAS

[TSTP13c] L. G. Torres, P. M. Sagar, D. R. Thompson, and R. A. Phillips. Scaling down the analysis of seabird–fishery interactions. Marine Ecology Progress Series, 473:275–289, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/473/

m473p275.pdf. See comment [CSS^{+}13].

Torre:2016:NAA

[TT16] M. P. Torre and T. E. Targett. Nekton assemblages along riprap-altered shorelines in Delaware Bay, USA: comparisons with adjacent beach. Marine Ecology Progress Series, 548: 209–218, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/548/m548p209.pdf.

Torres:2011:WCA

[TTB^{+}11] L. G. Torres, D. R. Thompson, S. Bearhop, S. Votier, G. A. Taylor, P. M. Sagar, and B. C. Robertson. White-capped albatrosses alter fine-scale foraging behavior patterns when associated with fishing vessels. Marine Ecology Progress Series, 428:289–301, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/428/m428p289.pdf.

Townsend:2011:SCE

[TTC11] M. Townsend, S. F. Thrush, and M. J. Carbines. Simplifying the complex: an ‘ecosystem principles approach’ to goods and services management in marine coastal ecosystems. Marine Ecology Progress Series, 434:291–301, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m434p291.pdf.

Thresher:2011:ECL

[TTF^{+}11] R. E. Thresher, B. Tilbrook, S. Fallon, N. C. Wilson, and J. Adkins. Effects of chronic low carbonate saturation levels on the distribution, growth and skeletal chemistry of deepsea corals and other seamount megabenthos. Marine Ecology Progress Series, 442:87–99, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/442/m442p087.pdf.

Thorpe:2019:CPA

[TTM19] S. E. Thorpe, G. A. Tarling, and E. J. Murphy. Circumpolar patterns in Antarctic krill larval recruitment: an environmentally driven model. Marine Ecology Progress Series, 613:77–96, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m613p077.pdf.

Tomaro:2012:WBB

[TTPM12] L. M. Tomaro, D. J. Teel, W. T. Peterson, and J. A. Miller. When is bigger better? Early marine residence of middle and upper Columbia River spring Chinook salmon. Marine Ecology Progress Series, 452:237–252, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m452p237.pdf.

Thyrring:2017:IIA

[TTS17] J. Thyrring, R. Tremblay, and M. K. Sejr. Importance of ice algae and pelagic phytoplankton as food sources revealed by fatty acid trophic markers in a keystone species (Mytilus trossulus) from the high Arctic. Marine Ecology Progress Series, 572:155–164, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/572/m572p155.pdf.

Thompson:2015:SIV

[TTT^{+}15] D. R. Thompson, L. G. Torres, G. A. Taylor, M. J. Rayner, P. M. Sagar, S. A. Shaffer, R. A. Phillips, and S. J. Bury. Stable isotope values delineate the non-breeding distributions of sooty shearwaters Puffinus griseus in the North Pacific Ocean. Marine Ecology Progress Series, 521:277–282, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m521p277.pdf.

Tucker:2012:ACM

[TTW^{+}12] S. Tucker, M. Trudel, D. W. Welch, J. R. Candy, J. F. T. Morris, M. E. Thiess, C. Wallace, and T. D. Beacham. Annual coastal migration of juvenile Chinook salmon: static stockspecific patterns in a highly dynamic ocean. Marine Ecology Progress Series, 449:245–262, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m449p245.pdf.

Tomiyama:2013:FRA

[TUK13] T. Tomiyama, S. Uehara, and Y. Kurita. Feeding relationships among fishes in shallow sandy areas in relation to stocking of Japanese flounder. Marine Ecology Progress Series, 479:163–175, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/479/m479p163.pdf.

Tuttle:2017:DIE

[Tut17] L. J. Tuttle. Direct and indirect effects of invasive lionfish on coral-reef cleaning mutualists. Marine Ecology Progress Series, 569:163–172, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/569/m569p163.pdf.

Trenkel:2019:OPW

[TVA^{+}19] V. M. Trenkel, S. Vaz, C. Albouy, A. Brind’Amour, E. Duhamel, P. Laffargue, J. B. Romagnan, J. Simon, and P. Lorance. Opinion piece: We can reduce the impact of scientific trawling on marine ecosystems. Marine Ecology Progress Series, 609:277–282, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/609/m609p277.pdf.

Townhill:2017:CCI

[TvdMM^{+}17] B. L. Townhill, J. van der Molen, J. D. Metcalfe, S. D. Simpson, A. Farcas, and J. K. Pinnegar. Consequences of climate-induced low oxygen conditions for commercially important fish. Marine Ecology Progress Series, 580:191– 204, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/580/m580p191.pdf.

Tuya:2010:PRR

[TVH^{+}10] F. Tuya, M. A. Vanderklift, G. A. Hyndes, T. Wernberg, M. S. Thomsen, and C. Hanson. Proximity to rocky reefs alters the balance between positive and negative effects on seagrass fauna. Marine Ecology Progress Series, 405:175– 186, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/405/m405p175.pdf.

Tsoukali:2016:FRN

[TVM16] S. Tsoukali, A. W. Visser, and B. R. MacKenzie. Functional responses of North Atlantic fish eggs to increasing temperature. Marine Ecology Progress Series, 555:151–165, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/555/
m555p151.pdf.

Tuya:2013:SRN

[TVRG^{+}13] F. Tuya, M. A. Viera-Rodríguez, R. Guedes, F. Espino, R. Haroun, and J. Terrados. Seagrass responses to nutrient enrichment depend on clonal integration, but not flowon effects on associated biota. Marine Ecology Progress Series, 490:23–35, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/490/m490p023.pdf.

Tougaard:2010:GMS

[TW10] J. Tougaard and M. S. Wisz. General models of the spatial distribution of porpoises require representative data and parsimony: Comment on Skov & Thomsen (2008). Marine Ecology Progress Series, 399:295–297, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m399p295.pdf. See

[ST08].

Tatsumi:2016:UAL

[TW16] M. Tatsumi and J. T. Wright. Understory algae and low light reduce recruitment of the habitat-forming kelp Ecklonia radiata. Marine Ecology Progress Series, 552:131– 143, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/552/m552p131.pdf.

Teesdale:2015:PHR

[TWL15] G. N. Teesdale, B. W. Wolfe, and C. G. Lowe. Patterns of home ranging, site fidelity, and seasonal spawning migration of barred sand bass caught within the Palos Verdes Shelf Superfund Site. Marine Ecology Progress Series, 539: 255–269, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/539/m539p255.pdf.

Taylor:2015:FEI

[TWM^{+}15] D. I. Taylor, S. A. Wood, P. McNabb, S. Ogilvie, C. Cornelisen, J. Walker, S. Khor, and S. C. Cary. Facilitation effects of invasive and farmed bivalves on native populations of the sea slug Pleurobranchaea maculata. Marine Ecology Progress Series, 537:39–48, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/537/m537p039.pdf.

True:2015:PDK

[TWW^{+}15] A. C. True, D. R. Webster, M. J. Weissburg, J. Yen, and A. Genin. Patchiness and depth-keeping of copepods in response to simulated frontal flows. Marine Ecology Progress Series, 539:65–76, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/539/m539p065.pdf.

Takeuchi:2019:FUO

[TWY^{+}19] A. Takeuchi, S. Watanabe, S. Yamamoto, M. J. Miller, T. Fukuba, T. Miwa, T. Okino, T. Minamoto, and K. Tsukamoto. First use of oceanic environmental DNA to study the spawning ecology of the Japanese eel Anguilla japonica. Marine Ecology Progress Series, 609:187– 196, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/609/m609p187.pdf.

Thompson:2012:SSA

[TYG12] C. M. Thompson, R. H. York, and S. M. Gallager. Speciesspecific abundance of bivalve larvae in relation to biological and physical conditions in a Cape Cod estuary. Marine Ecology Progress Series, 469:53–69, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/469/m469p053.pdf.

Turk:2015:CMS

[TYVR^{+}15] D. Turk, K. K. Yates, M. Vega-Rodriguez, G. Toro-Farmer, C. L’Esperance, N. Melo, D. Ramsewak, M. Dowd, S. Cerdeira Estrada, F. E. Muller-Karger, S. R. Herwitz, and W. R. McGillis. Community metabolism in shallow coral reef and seagrass ecosystems, lower Florida Keys. Marine Ecology Progress Series, 538:35–52, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/538/m538p035.pdf.

Tang:2011:SSM

[TZF11] Q. Tang, J. Zhang, and J. Fang. Shellfish and seaweed mariculture increase atmospheric CO_{2} absorption by coastal ecosystems. Marine Ecology Progress Series, 424:97–104, 2011. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m424p097.pdf.

Ugland:2014:LNF

[UAK^{+}14] K. I. Ugland, D. L. Aksnes, T. A. Klevjer, J. Titelman, and S. Kaartvedt. Lévy night flights by the jellyfish Periphylla periphylla. Marine Ecology Progress Series, 513:121–130, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m513p121.pdf.

Udyawer:2013:VRC

[UCK^{+}13] V. Udyawer, A. Chin, D. M. Knip, C. A. Simpfendorfer, and M. R. Heupel. Variable response of coastal sharks to severe tropical storms: environmental cues and changes in space use. Marine Ecology Progress Series, 480:171–183, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/480/
m480p171.pdf.

Umbricht:2018:CIC

[UDF^{+}18] J. Umbricht, J. W. Dippner, B. Fry, I. Kröncke, I. Liskow, P. Nehmer, F. Thoms, and M. Voss. Correction of the isotopic composition (δ ^{13}C and δ ^{1}5N) of preserved Baltic and North Sea macrozoobenthos and their trophic interactions. Marine Ecology Progress Series, 595:1–13, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ueda:2014:RHA

[Ued14] H. Ueda. Review: Homing ability and migration success in Pacific salmon: mechanistic insights from biotelemetry, endocrinology, and neurophysiology. Marine Ecology Progress Series, 496:219–232, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2013/496/m496p219.pdf.

Underwood:2010:DCD

[UFP^{+}10] G. J. C. Underwood, S. Fietz, S. Papadimitriou, D. N. Thomas, and G. S. Dieckmann. Distribution and composition of dissolved extracellular polymeric substances (EPS) in Antarctic sea ice. Marine Ecology Progress Series, 404:1–19, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Urriago:2012:SUT

[UHG12] J. D. Urriago, J. H. Himmelman, and C. F. Gaymer. Sea urchin Tetrapygus niger distribution on elevated surfaces represents a strategy for avoiding predatory sea stars. Marine Ecology Progress Series, 444:85–95, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/444/m444p085.pdf.

Umanzor:2017:IMI

[ULCAZG17] S. Umanzor, L. Ladah, L. E. Calderon-Aguilera, and J. A. Zertuche-González. Intertidal macroalgae influence macroinvertebrate distribution across stress scenarios. Marine Ecology Progress Series, 584:67–77, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/584/m584p067.pdf.

Umani:2010:DEV

[UMF^{+}10] S. Fonda Umani, E. Malisana, F. Focaracci, M. Magagnini, C. Corinaldesi, and R. Danovaro. Disentangling the effect of viruses and nanoflagellates on prokaryotes in bathypelagic waters of the Mediterranean Sea. Marine Ecology Progress Series, 418:73–85, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m418p073.pdf.

Urban:2012:FHP

[Urb12] D. Urban. Food habits of Pacific cod and walleye pollock in the northern Gulf of Alaska. Marine Ecology Progress Series, 469:215–222, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m469p215.pdf.

Udyawer:2016:EHS

[USR^{+}16] V. Udyawer, C. A. Simpfendorfer, M. Read, M. Hamann, and M. R. Heupel. Exploring habitat selection in sea snakes using passive acoustic monitoring and Bayesian hierarchical models. Marine Ecology Progress Series, 546:249–262, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/546/
m546p249.pdf.

Umezawa:2018:PPD

[UTS^{+}18] Y. Umezawa, A. Tamaki, T. Suzuki, S. Takeuchi, C. Yoshimizu, and I. Tayasu. Phytoplankton as a principal diet for callianassid shrimp larvae in coastal waters, estimated from laboratory rearing and stable isotope analysis. Marine Ecology Progress Series, 592:141–158, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/592/m592p141.pdf.

Urmy:2018:FHC

[UW18] S. S. Urmy and J. D. Warren. Foraging hotspots of common and roseate terns: the influence of tidal currents, bathymetry, and prey density. Marine Ecology Progress Series, 590:227– 245, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/590/m590p227.pdf.

VanderZanden:2013:TEG

[VAB^{+}13] H. B. Vander Zanden, K. E. Arthur, A. B. Bolten, B. N. Popp, C. J. Lagueux, E. Harrison, C. L. Campbell, and K. A. Bjorndal. Trophic ecology of a green turtle breeding population. Marine Ecology Progress Series, 476:237–249, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/476/
m476p237.pdf.

Vahatalo:2011:PTT

[VAHL11] A. V. Vähätalo, H. Aarnos, L. Hoikkala, and R. Lignell. Photochemical transformation of terrestrial dissolved organic matter supports hetero- and autotrophic production in coastal waters. Marine Ecology Progress Series, 423:1–14, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Valdivia:2015:DEP

[VANB15] N. Valdivia, M. A. Aguilera, S. A. Navarrete, and B. R. Broitman. Disentangling the effects of propagule supply and environmental filtering on the spatial structure of a rocky shore metacommunity. Marine Ecology Progress Series, 538: 67–79, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/538/m538p067.pdf.

Vadstein:2012:CNP

[VARO12] O. Vadstein, T. Andersen, H. R. Reinertsen, and Y. Olsen. Carbon, nitrogen and phosphorus resource supply and utilisation for coastal planktonic heterotrophic bacteria in a gradient of nutrient loading. Marine Ecology Progress Series, 447: 55–75, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m447p055.pdf.

Vanderklift:2014:VKE

[VB14] M. A. Vanderklift and D. Bearham. Variation in δ ^{13}C and δ ^{15}N of kelp is explained by light and productivity. Marine Ecology Progress Series, 515:111–121, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/515/
m515p111.pdf.

Valentine:2019:SSW

[VB19a] M. M. Valentine and M. J. Butler IV. Sponges structure watercolumn characteristics in shallow tropical coastal ecosystems. Marine Ecology Progress Series, 608:133–147, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/608/
m608p133.pdf.

Vozzo:2019:COS

[VB19b] M. L. Vozzo and M. J. Bishop. Co-occurring secondary foundation species have distinct effects on the recruitment and survival of associated organisms. Marine Ecology Progress Series, 608:61–72, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/608/m608p061.pdf.

vonBiela:2019:ERN

[vBAP^{+}19] V. R. von Biela, M. L. Arimitsu, J. F. Piatt, B. Heflin, S. K. Schoen, J. L. Trowbridge, and C. M. Clawson. Extreme reduction in nutritional value of a key forage fish during the Pacific marine heatwave of 2014–2016. Marine Ecology Progress Series, 613:171–182, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m613p171.pdf.

vanBleijswijk:2014:DGS

[vBBW^{+}14] J. D. L. van Bleijswijk, L. Begeman, H. J. Witte, L. L. IJsseldijk, S. M. J. M. Brasseur, A. Gröne, and M. F. Leopold. Detection of grey seal Halichoerus grypus DNA in attack wounds on stranded harbour porpoises Phocoena phocoena. Marine Ecology Progress Series, 513:277–281, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m513p277.pdf.

Valdizan:2011:ERC

[VBD^{+}11] A. Valdizan, P. G. Beninger, P. Decottignies, M. Chantrel, and B. Cognie. Evidence that rising coastal seawater temperatures increase reproductive output of the invasive gastropod Crepidula fornicata. Marine Ecology Progress Series, 438:153–165, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/438/m438p153.pdf.

Villazan:2016:FVL

[VBGO^{+}16] B. Villazán, F. G. Brun, V. González-Ortiz, F. Moreno-Marín, T. J. Bouma, and J. J. Vergara. Flow velocity and light level drive non-linear response of seagrass Zostera noltei to ammonium enrichment. Marine Ecology Progress Series, 545:109–121, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/545/m545p109.pdf.

Villazan:2013:IBA

[VBJR^{+}13] B. Villazán, F. G. Brun, R. Jiménez-Ramos, J. L. Pérez-Lloréns, and J. J. Vergara. Interaction between ammonium and phosphate uptake rates in the seagrass Zostera noltii. Marine Ecology Progress Series, 488:133–143, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/488/
m488p133.pdf.

vanBroekhoven:2015:NRF

[vBJV^{+}15] W. van Broekhoven, H. Jansen, M. Verdegem, E. Struyf, K. Troost, H. Lindeboom, and A. Smaal. Nutrient regeneration from feces and pseudofeces of mussel Mytilus edulis spat. Marine Ecology Progress Series, 534:107–120, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/534/
m534p107.pdf.

VanEngeland:2011:PUD

[VBM^{+}11] T. Van Engeland, T. J. Bouma, E. P. Morris, F. G. Brun, G. Peralta, M. Lara, I. E. Hendriks, K. Soetaert, and J. J. Middelburg. Potential uptake of dissolved organic matter by seagrasses and macroalgae. Marine Ecology Progress Series, 427:71–81, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/427/m427p071.pdf.

VanEngeland:2013:DOM

[VBM^{+}13] T. Van Engeland, T. J. Bouma, E. P. Morris, F. G. Brun, G. Peralta, M. Lara, I. E. Hendriks, P. van Rijswijk, B. Veuger, K. Soetaert, and J. J. Middelburg. Dissolved organic matter uptake in a temperate seagrass ecosystem. Marine Ecology Progress Series, 478:87–100, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/478/m478p087.pdf.

vanBemmelen:2017:FOC

[vBMH^{+}17] R. van Bemmelen, B. Moe, S. A. Hanssen, N. M. Schmidt, J. Hansen, J. Lang, B. Sittler, L. Bollache, I. Tulp, R. Klaassen, and O. Gilg. Flexibility in otherwise consistent non-breeding movements of a long-distance migratory seabird, the long-tailed skua. Marine Ecology Progress Series, 578:197–211, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m578p197.pdf.

Vanderklift:2014:DRS

[VBR^{+}14] M. A. Vanderklift, F. Boschetti, C. Roubertie, R. D. Pillans, M. D. E. Haywood, and R. C. Babcock. Density of reef sharks estimated by applying an agent-based model to video surveys. Marine Ecology Progress Series, 508:201–209, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/508/
m508p201.pdf.

Volaric:2018:OMI

[VBR18] M. P. Volaric, P. Berg, and M. A. Reidenbach. Oxygen metabolism of intertidal oyster reefs measured by aquatic eddy covariance. Marine Ecology Progress Series, 599:75– 91, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/599/m599p075.pdf.

Volaric:2019:IMA

[VBR19] M. P. Volaric, P. Berg, and M. A. Reidenbach. An invasive macroalga alters ecosystem metabolism and hydrodynamics on a tidal flat. Marine Ecology Progress Series, 628:1–16, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Valdivia:2014:SIM

[VBT14] N. Valdivia, C. Buschbaum, and M. Thiel. Succession in intertidal mussel bed assemblages on different shores: species mobility matters. Marine Ecology Progress Series, 497:131– 142, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/497/m497p131.pdf.

Viladrich:2017:VLF

[VBT^{+}17] N. Viladrich, L. Bramanti, G. Tsounis, A. Martínez-Quitana, C. Ferrier-Pagès, and S. Rossi. Variation of lipid and free fatty acid contents during larval release in two temperate octocorals according to their trophic strategy. Marine Ecology Progress Series, 573:117–128, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2017/573/m573p117.pdf.

Vilas:2010:VAE

[VBV^{+}10] R. Vilas, C. Bouza, M. Vera, A. Millán, and P. Martínez. Variation in anonymous and EST-microsatellites suggests adaptive population divergence in turbot. Marine Ecology Progress Series, 420:231–239, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/420/m420p231.pdf.

Vohnik:2017:FRS

[VBŽ^{+}17] M. Vohník, O. Borovec, I. Župan, M. Kolařík, and R. Sudová. Fungal root symbionts of the seagrass Posidonia oceanica in the central Adriatic Sea revealed by microscopy, culturing and 454-pyrosequencing. Marine Ecology Progress Series, 583:107–120, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/583/m583p107.pdf.

Vinagre:2014:ECG

[VC14] C. Vinagre and M. J. Costa. Estuarine-coastal gradient in food web network structure and properties. Marine Ecology Progress Series, 503:11–21, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/503/m503p011.pdf.

Verdelhos:2011:LGS

[VCDP11] T. Verdelhos, P. G. Cardoso, M. Dolbeth, and M. A. Pardal. Latitudinal gradients in Scrobicularia plana reproduction patterns, population dynamics, growth, and secondary production. Marine Ecology Progress Series, 442: 271–283, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/442/m442p271.pdf.

Vollset:2013:EFD

[VCFF13] K. W. Vollset, I. A. Catalán, Ø. Fiksen, and A. Folkvord. Effect of food deprivation on distribution of larval and early juvenile cod in experimental vertical temperature and light gradients. Marine Ecology Progress Series, 475:191–201, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/475/m475p191.pdf.

Valsecchi:2010:GES

[VCG^{+}10] E. Valsecchi, P. J. Corkeron, P. Galli, W. Sherwin, and G. Bertorelle. Genetic evidence for sex-specific migratory behaviour in western South Pacific humpback whales. Marine Ecology Progress Series, 398:275–286, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/398/
m398p275.pdf.

Vales:2015:ODC

[VCG^{+}15] D. G. Vales, L. Cardona, N. A. García, L. Zenteno, and E. A. Crespo. Ontogenetic dietary changes in male South American fur seals Arctocephalus australis in Patagonia. Marine Ecology Progress Series, 525:245–260, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/525/m525p245.pdf.

Villarino:2015:MFB

[VCL^{+}15] E. Villarino, G. Chust, P. Licandro, M. Butenschön, L. Ibaibarriaga, A. Larrañaga, and X. Irigoien. Modelling the future biogeography of North Atlantic zooplankton communities in response to climate change. Marine Ecology Progress Series, 531:121–142, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2015/531/m531p121.pdf.

Vasquez-Cardenas:2018:BCR

[VCMS^{+}18] D. Vasquez-Cardenas, L. Meire, H. L. Sørensen, R. N. Glud, F. J. R. Meysman, and H. T. S. Boschker. Bacterial chemoautotrophic reoxidation in sub-Arctic sediments: a seasonal study in Kobbefjord, Greenland. Marine Ecology Progress Series, 601:33–39, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/601/m601p033.pdf.

Vasquez-Cardenas:2016:SSE

[VCQM^{+}16] D. Vasquez-Cardenas, C. O. Quintana, F. J. R. Meysman, E. Kristensen, and H. T. S. Boschker. Species-specific effects of two bioturbating polychaetes on sediment chemoautotrophic bacteria. Marine Ecology Progress Series, 549:55–68, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m549p055.pdf.

Valls:2015:FET

[VCRUQ15] M. Valls, M. Cabanellas-Reboredo, I. Uranga, and A. Quetglas. Feeding ecology of two squid species from the western Mediterranean. Marine Ecology Progress Series, 531: 207–219, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/531/m531p207.pdf.

Valdemarsen:2010:VZM

[VCVK^{+}10] T. Valdemarsen, P. Canal-Vergés, E. Kristensen, M. Holmer, M. D. Kristiansen, and M. R. Flindt. Vulnerability of Zostera marina seedlings to physical stress. Marine Ecology Progress Series, 418:119–130, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/418/m418p119.pdf.

vanDenderen:2015:SEB

[vDBH^{+}15] P. D. van Denderen, S. G. Bolam, J. G. Hiddink, S. Jennings, A. Kenny, A. D. Rijnsdorp, and T. van Kooten. Similar effects of bottom trawling and natural disturbance on composition and function of benthic communities across habitats. Marine Ecology Progress Series, 541:31–43, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/541/
m541p031.pdf.

vanderBank:2011:DSN

[vdBUPP^{+}11] M. G. van der Bank, A. C. Utne-Palm, K. Pittman, A. K. Sweetman, N. B. Richoux, V. Brüchert, and M. J. Gibbons. Dietary success of a ‘new’ key fish in an overfished ecosystem: evidence from fatty acid and stable isotope signatures. Marine Ecology Progress Series, 428:219–233, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/428/
m428p219.pdf.

vanDeurs:2010:OSS

[vDCFM10] M. van Deurs, A. Christensen, C. Frisk, and H. Mosegaard. Overwintering strategy of sandeel ecotypes from an energy/ predation trade-off perspective. Marine Ecology Progress Series, 416:201–214, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/416/m416p201.pdf.

vanDeurs:2013:PZG

[vDCR13] M. van Deurs, A. Christensen, and A. Rindorf. Patchy zooplankton grazing and high energy conversion efficiency: ecological implications of sandeel behavior and strategy. Marine Ecology Progress Series, 487:123–133, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/487/
m487p123.pdf.

Vye:2018:CEI

[VDEO18] S. R. Vye, J. T. A. Dick, M. C. Emmerson, and N. E. O’Connor. Cumulative effects of an invasive species and nutrient enrichment on rock pool communities. Marine Ecology Progress Series, 594:39–50, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/594/m594p039.pdf.

Vermeij:2015:NEG

[VDG^{+}15] M. J. A. Vermeij, H. DeBey, G. Grimsditch, J. Brown, D. Obura, R. DeLeon, and S. A. Sandin. Negative effects of gardening damselfish Stegastes planifrons on coral health depend on predator abundance. Marine Ecology Progress Series, 528:289–296, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/528/m528p289.pdf.

vanDeurs:2012:SLT

[vDGK^{+}12] M. van Deurs, T. M. Grome, M. Kaspersen, H. Jensen, C. Stenberg, T. K. Sørensen, J. Støttrup, T. Warnar, and H. Mosegaard. Short- and long-term effects of an offshore wind farm on three species of sandeel and their sand habitat. Marine Ecology Progress Series, 458:169–180, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m458p169.pdf.

vanderGeest:2014:NRS

[vdGSE^{+}14] M. van der Geest, A. A. Sall, S. O. Ely, R. W. Nauta, J. A. van Gils, and T. Piersma. Nutritional and reproductive strategies in a chemosymbiotic bivalve living in a tropical intertidal seagrass bed. Marine Ecology Progress Series, 501:113–126, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m501p113.pdf.

vanderGeest:2019:DDG

[vdGvdLvG^{+}19] M. van der Geest, J. A. C. van der Lely, J. A. van Gils, T. Piersma, and T. Lok. Density-dependent growth of bivalves dominating the intertidal zone of Banc d’Arguin, Mauritania: importance of feeding mode, habitat and season. Marine Ecology Progress Series, 610:51–63, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/610/
m610p051.pdf.

vonderHeyden:2015:CSG

[vdHTT^{+}15] S. von der Heyden, J. A. Toms, P. R. Teske, S. J. Lamberth, and W. Holleman. Contrasting signals of genetic diversity and historical demography between two recently diverged marine and estuarine fish species. Marine Ecology Progress Series, 526:157–167, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/526/m526p157.pdf.

vanDeurs:2015:ECS

[vDJF15] M. van Deurs, C. Jørgensen, and Ø. Fiksen. Effects of copepod size on fish growth: a model based on data for North Sea sandeel. Marine Ecology Progress Series, 520:235–243, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m520p235.pdf.

VanColen:2010:DTD

[VDM^{+}10] C. Van Colen, A. De Backer, G. Meulepas, D. Van der Wal, M. Vincx, S. Degraer, and T. Ysebaert. Diversity, trait displacements and shifts in assemblage structure of tidal flat deposit feeders along a gradient of hydrodynamic stress. Marine Ecology Progress Series, 406:79–89, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/406/
m406p079.pdf.

vandeMerwe:2010:POP

[vdMHW^{+}10] J. P. van de Merwe, M. Hodge, J. M. Whittier, K. Ibrahim, and S. Y. Lee. Persistent organic pollutants in the green sea turtle Chelonia mydas: nesting population variation, maternal transfer, and effects on development. Marine Ecology Progress Series, 403:269–278, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/403/m403p269.pdf.

Vazquez-Dominguez:2013:PPC

[VDMLU13] E. Vázquez-Domínguez, X. A. G. Morán, and A. López-Urrutia. Photoacclimation of picophytoplankton in the central Cantabrian Sea. Marine Ecology Progress Series, 493: 43–56, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/493/m493p043.pdf.

vonderMeden:2010:FSO

[vdMPM^{+}10] C. E. O. von der Meden, F. Porri, C. D. McQuaid, K. Faulkner, and J. Robey. Fine-scale ontogenetic shifts in settlement behaviour of mussels: changing responses to biofilm and conspecific settler presence in Mytilus galloprovincialis and Perna perna. Marine Ecology Progress Series, 411:161– 171, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/411/m411p161.pdf.

vonderMeden:2012:NEE

[vdMPM12] C. E. O. von der Meden, F. Porri, and C. D. McQuaid. New estimates of early post-settlement mortality for intertidal mussels show no relationship with meso-scale coastline topographic features. Marine Ecology Progress Series, 463:193–204, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/463/m463p193.pdf.

vonderMeden:2012:SIC

[vdMPRM12] C. E. O. von der Meden, F. Porri, S. Radloff, and C. D. McQuaid. Settlement intensification and coastline topography: understanding the role of habitat availability in the pelagic–benthic transition. Marine Ecology Progress Series, 459:63–71, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/459/m459p063.pdf.

vandePoll:2011:NPQ

[vdPLdV^{+}11] W. H. van de Poll, M. Lagunas, T. de Vries, R. J. W. Visser, and A. G. J. Buma. Non-photochemical quenching of chlorophyll fluorescence and xanthophyll cycle responses after excess PAR and UVR in Chaetoceros brevis, Phaeocystis antarctica and coastal Antarctic phytoplankton. Marine Ecology Progress Series, 426:119–131, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2011/426/m426p119.pdf.

Vermeij:2011:CCA

[VDS11] M. J. A. Vermeij, M. L. Dailer, and C. M. Smith. Crustose coralline algae can suppress macroalgal growth and recruitment on Hawaiian coral reefs. Marine Ecology Progress Series, 422:1–7, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

vanderSleen:2018:NSR

[vdSRT^{+}18] P. van der Sleen, R. R. Rykaczewski, B. D. Turley, W. J. Sydeman, M. García-Reyes, S. J. Bograd, C. D. van der Lingen, J. C. Coetzee, T. Lamont, and B. A. Black. Non-stationary responses in anchovy (Engraulis encrasicolus) recruitment to coastal upwelling in the southern Benguela. Marine Ecology Progress Series, 596:155–164, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/596/m596p155.pdf.

VanderStocken:2015:ILS

[VDV^{+}15] T. Van der Stocken, D. J. R. De Ryck, B. Vanschoenwinkel, E. Deboelpaep, T. J. Bouma, F. Dahdouh-Guebas, and N. Koedam. Impact of landscape structure on propagule dispersal in mangrove forests. Marine Ecology Progress Series, 524:95–106, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/524/m524p095.pdf.

vanderVeer:2010:FLE

[vdVFK^{+}10] H. W. van der Veer, V. Freitas, J. Koot, J. I. J. Witte, and A. F. Zuur. Food limitation in epibenthic species in temperate intertidal systems in summer: analysis of 0-group plaice Pleuronectes platessa. Marine Ecology Progress Series, 416:215–227, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/416/m416p215.pdf.

vanderVeer:2011:LTT

[vdVKA^{+}11] H. W. van der Veer, J. Koot, G. Aarts, R. Dekker, W. Diderich, V. Freitas, and J. I. J. Witte. Long-term trends in juvenile flatfish indicate a dramatic reduction in nursery function of the balgzand intertidal, Dutch Wadden Sea. Marine Ecology Progress Series, 434:143–154, 2011. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m434p143.pdf.

vandeWolfshaar:2011:EHP

[vdWHG11] K. E. van de Wolfshaar, R. HilleRisLambers, and A. Gårdmark. Effect of habitat productivity and exploitation on populations with complex life cycles. Marine Ecology Progress Series, 438:175–184, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/438/m438p175.pdf.

vandeWolfshaar:2015:MPE

[vdWTWS15] K. E. van de Wolfshaar, I. Tulp, H. Wennhage, and J. G. Støttrup. Modelling population effects of juvenile offshore fish displacement towards adult habitat. Marine Ecology Progress Series, 540:193–201, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2015/540/m540p193.pdf.

Valley:2018:PIM

[VE18] J. R. Valley and R. B. Emlet. Predator-induced morphologies and cue specificity in veliger larvae of Littorina scutulata. Marine Ecology Progress Series, 598:61–70, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/598/
m598p061.pdf.

vanEeden:2016:FSF

[vERRP16] R. van Eeden, T. Reid, P. G. Ryan, and L. Pichegru. Finescale foraging cues for African penguins in a highly variable marine environment. Marine Ecology Progress Series, 543:257–271, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/543/m543p257.pdf.

Visser:2013:OFM

[VF13] A. W. Visser and Ø. Fiksen. Optimal foraging in marine ecosystem models: selectivity, profitability and switching. Marine Ecology Progress Series, 473:91–101, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/473/
m473p091.pdf.

Villeger:2012:NRC

[VFMdW12] S. Villéger, F. Ferraton, D. Mouillot, and R. de Wit. Nutrient recycling by coastal macrofauna: intra- versus interspecific differences. Marine Ecology Progress Series, 452:297– 303, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/452/m452p297.pdf.

Valiela:2013:NGP

[VGBJ^{+}13] I. Valiela, A. Giblin, C. Barth-Jensen, C. Harris, T. Stone, S. Fox, and J. Crusius. Nutrient gradients in Panamanian estuaries: effects of watershed deforestation, rainfall, upwelling, and within-estuary transformations. Marine Ecology Progress Series, 482:1–15, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Vacquie-Garcia:2015:PPC

[VGGD^{+}15] J. Vacquié-Garcia, C. Guinet, A. C. Dragon, M. Viviant, N. El Ksabi, and F. Bailleul. Predicting prey capture rates of southern elephant seals from track and dive parameters. Marine Ecology Progress Series, 541:265–277, 2015. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m541p265.pdf.

Valls:2012:MTI

[VGGF12] A. Valls, D. Gascuel, S. Guénette, and P. Francour. Modeling trophic interactions to assess the effects of a marine protected area: case study in the NW Mediterranean Sea. Marine Ecology Progress Series, 456:201–214, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/456/
m456p201.pdf.

Valencia-Gasti:2016:SSB

[VGL16] J. A. Valencia-Gasti and L. B. Ladah. Synchronous settlement of barnacle larvae at small spatial scales correlates with both internal waves and onshore surface winds. Marine Ecology Progress Series, 552:195–210, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/552/m552p195.pdf.

Volpov:2016:TDD

[VGR^{+}16] B. L. Volpov, E. T. Goundie, D. A. S. Rosen, J. P. Y. Arnould, and A. W. Trites. Transiting to depth disrupts overall dynamic body acceleration and oxygen consumption rate in freely diving Steller sea lions. Marine Ecology Progress Series, 562:221–236, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/562/m562p221.pdf.

Vaudo:2011:DNO

[VH11] J. J. Vaudo and M. R. Heithaus. Dietary niche overlap in a nearshore elasmobranch mesopredator community. Marine Ecology Progress Series, 425:247–260, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/425/
m425p247.pdf.

Varennes:2015:FRC

[VHBG15] É. Varennes, S. A. Hanssen, J. C. Bonardelli, and M. Guillemette. Functional response curves of avian molluscivores: high intake rates maintained even at low prey density. Marine Ecology Progress Series, 526:207–212, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2015/526/m526p207.pdf.

Viehman:2018:UDP

[VHG^{+}18] T. S. Viehman, J. L. Hench, S. P. Griffin, A. Malhotra, K. Egan, and P. N. Halpin. Understanding differential patterns in coral reef recovery: chronic hydrodynamic disturbance as a limiting mechanism for coral colonization. Marine Ecology Progress Series, 605:135–150, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m605p135.pdf.

Viddi:2011:FSM

[VHHGF11] F. A. Viddi, R. G. Harcourt, R. Hucke-Gaete, and I. C. Field. Fine-scale movement patterns of the sympatric Chilean and Peale’s dolphins in the northern Patagonian fjords, Chile. Marine Ecology Progress Series, 436:245–256, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/436/
m436p245.pdf.

Vitelli:2015:TFA

[VHKT15] F. Vitelli, G. A. Hyndes, A. Kendrick, and A. Turco. Turfforming algal assemblages on temperate reefs are strongly influenced by the territorial herbivorous fish Parma mccullochi (Pomacentridae). Marine Ecology Progress Series, 523: 175–185, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/523/m523p175.pdf.

Verkaik:2016:CEO

[VHM16] K. Verkaik, J. F. Hamel, and A. Mercier. Carry-over effects of ocean acidification in a cold-water lecithotrophic holothuroid. Marine Ecology Progress Series, 557:189–206, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/557/
m557p189.pdf.

Visser:2011:TMB

[VHP^{+}11] F. Visser, K. L. Hartman, G. J. Pierce, V. D. Valavanis, and J. Huisman. Timing of migratory baleen whales at the Azores in relation to the North Atlantic spring bloom. Marine Ecology Progress Series, 440:267–279, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/440/m440p267.pdf.

Veneranta:2013:RAS

[VHV13] L. Veneranta, R. Hudd, and J. Vanhatalo. Reproduction areas of sea-spawning coregonids reflect the environment in shallow coastal waters. Marine Ecology Progress Series, 477: 231–250, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/477/m477p231.pdf.

Vignon:2015:DQS

[Vig15] M. Vignon. Disentangling and quantifying sources of otolith shape variation across multiple scales using a new hierarchical partitioning approach. Marine Ecology Progress Series, 534:163–177, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/534/m534p163.pdf.

Vilar:2013:LRE

[VJG^{+}13] C. C. Vilar, J. C. Joyeux, T. Giarrizzo, H. L. Spach, J. P. Vieira, and T. Vaske-Junior. Local and regional ecological drivers of fish assemblages in Brazilian estuaries. Marine Ecology Progress Series, 485:181–197, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/485/
m485p181.pdf.

Villnas:2019:CMB

[VJJ^{+}19] A. Villnäs, U. Janas, A. B. Josefson, H. Kendzierska, H. Nygård, J. Norkko, and A. Norkko. Changes in macrofaunal biological traits across estuarine gradients: implications for the coastal nutrient filter. Marine Ecology Progress Series, 622:31–48, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/622/m622p031.pdf.

Ventura:2016:ROA

[VJM^{+}16] P. Ventura, M. D. Jarrold, P. L. Merle, S. Barnay-Verdier, T. Zamoum, R. Rodolfo-Metalpa, P. Calosi, and P. Furla. Resilience to ocean acidification: decreased carbonic anhydrase activity in sea anemones under high pCO_{2} conditions. Marine Ecology Progress Series, 559:257–263, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/559/
m559p257.pdf.

Villac:2011:EPP

[VK11] C. M. Villac and I. Kaczmarska. Estimating propagule pressure and viability of diatoms detected in ballast tank sediments of ships arriving at Canadian ports. Marine Ecology Progress Series, 425:47–61, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/425/m425p047.pdf.

Valdemarsen:2010:SCN

[VKH10] T. Valdemarsen, E. Kristensen, and M. Holmer. Sulfur, carbon, and nitrogen cycling in faunated marine sediments impacted by repeated organic enrichment. Marine Ecology Progress Series, 400:37–53, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/400/m400p037.pdf.

Vernet:2012:PPT

[VKY^{+}12] M. Vernet, W. A. Kozlowski, L. R. Yarmey, A. T. Lowe, R. M. Ross, L. B. Quetin, and C. H. Fritsen. Primary production throughout austral fall, during a time of decreasing daylength in the western Antarctic Peninsula. Marine Ecology Progress Series, 452:45–61, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/452/m452p045.pdf.

Vuori:2015:OSB

[VLK^{+}15] K. A. Vuori, K. K. Lehtonen, M. Kanerva, H. Peltonen, M. Nikinmaa, N. A. Berezina, and E. Boikova. Oxidative stress biomarkers in the copepod Limnocalanus macrurus from the northern Baltic Sea: effects of hydrographic factors and chemical contamination. Marine Ecology Progress Series, 538:131–144, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/538/m538p131.pdf.

Vokhshoori:2014:RIP

[VLM14] N. L. Vokhshoori, T. Larsen, and M. D. McCarthy. Reconstructing δ ^{13}C isoscapes of phytoplankton production in a coastal upwelling system with amino acid isotope values of littoral mussels. Marine Ecology Progress Series, 504: 59–72, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/504/m504p059.pdf.

Vignon:2010:EGD

[VM10] M. Vignon and F. Morat. Environmental and genetic determinant of otolith shape revealed by a non-indigenous tropical fish. Marine Ecology Progress Series, 411:231–241, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2010/411/m411p231.pdf.

Vincenzi:2014:FAK

[VM14] S. Vincenzi and M. Mangel. Food abundance, kittiwake life histories, and colony dynamics in the Northeastern Pacific: implications of climate change and regime shifts. Marine Ecology Progress Series, 515:251–263, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/515/m515p251.pdf.

Vokhshoori:2019:BFR

[VMC^{+}19] N. L. Vokhshoori, M. D. McCarthy, P. W. Collins, M. A. Etnier, T. Rick, M. Eda, J. Beck, and S. D. Newsome. Broader foraging range of ancient short-tailed albatross populations into California coastal waters based on bulk tissue and amino acid isotope analysis. Marine Ecology Progress Series, 610:1– 13, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Vinagre:2012:ICC

[VMN^{+}12] C. Vinagre, D. Madeira, L. Narciso, H. N. Cabral, and M. S. Diniz. Impact of climate change on coastal versus estuarine nursery areas: cellular and whole-animal indicators in juvenile seabass Dicentrarchus labrax. Marine Ecology Progress Series, 464:237–243, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/464/m464p237.pdf.

vanTussenbroek:2012:MFF

[vMVSW12] B. I. van Tussenbroek, L. V. Monroy-Velazquez, and V. Solis-Weiss. Meso-fauna foraging on seagrass pollen may serve in marine zoophilous pollination. Marine Ecology Progress Series, 469:1–6, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

VanTussenbroek:2010:PLD

[VMW^{+}10] B. I. Van Tussenbroek, M. Muhlia Montero, R. Wong, M. G. Barba Santos, and J. Márquez Guzmán. Pollen limitation in a dioecious seagrass: evidence from a field experiment. Marine Ecology Progress Series, 419:283–288, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/419/
m419p283.pdf.

Vogel:2015:DLA

[VMWU15] N. Vogel, F. W. Meyer, C. Wild, and S. Uthicke. Decreased light availability can amplify negative impacts of ocean acidification on calcifying coral reef organisms. Marine Ecology Progress Series, 521:49–61, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m521p049.pdf.

Vinther:2012:ECB

[VNK^{+}12] H. F. Vinther, P. Norling, P. S. Kristensen, P. Dolmer, and M. Holmer. Effects of coexistence between the blue mussel and eelgrass on sediment biogeochemistry and plant performance. Marine Ecology Progress Series, 447:139–149, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/447/
m447p139.pdf.

Valanko:2010:RPL

[VNN10] S. Valanko, A. Norkko, and J. Norkko. Rates of post-larval bedload dispersal in a non-tidal soft-sediment system. Marine Ecology Progress Series, 416:145–163, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/416/
m416p145.pdf.

Villazan:2013:EAC

[VPBV13] B. Villazán, M. F. Pedersen, F. G. Brun, and J. J. Vergara. Elevated ammonium concentrations and low light form a dangerous synergy for eelgrass Zostera marina. Marine Ecology Progress Series, 493:141–154, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m493p141.pdf.

Venera-Ponton:2011:MRG

[VPDPMRC11] D. E. Venera-Ponton, G. Diaz-Pulido, L. J. McCook, and A. Rangel-Campo. Macroalgae reduce growth of juvenile corals but protect them from parrotfish damage. Marine Ecology Progress Series, 421:109–115, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/421/
m421p109.pdf.

Vaz-Pinto:2012:PPF

[VPOA12] F. Vaz-Pinto, C. Olabarria, and F. Arenas. Propagule pressure and functional diversity: interactive effects on a macroalgal invasion process. Marine Ecology Progress Series, 471: 51–60, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/471/m471p051.pdf.

Volkenborn:2012:HAG

[VPW^{+}12] N. Volkenborn, L. Polerecky, D. S. Wethey, T. H. DeWitt, and S. A. Woodin. Hydraulic activities by ghost shrimp Neotrypaea californiensis induce oxic–anoxic oscillations in sediments. Marine Ecology Progress Series, 455:141–156, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m455p141.pdf.

Valdemarsen:2014:ROE

[VQKF14] T. Valdemarsen, C. O. Quintana, E. Kristensen, and M. R. Flindt. Recovery of organic-enriched sediments through microbial degradation: implications for eutrophic estuaries. Marine Ecology Progress Series, 503:41–58, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m503p041.pdf.

Voss:2014:RTO

[VQSH14] R. Voss, M. F. Quaas, J. O. Schmidt, and J. Hoffmann. Regional trade-offs from multi-species maximum sustainable yield (MMSY) management options. Marine Ecology Progress Series, 498:1–12, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Villegas-Rios:2014:LHA

[VRAP^{+}14] D. Villegas-Ríos, J. Alós, M. Palmer, S. K. Lowerre-Barbieri, R. Bañón, A. Alonso-Fernández, and F. Saborido-Rey. Lifehistory and activity shape catchability in a sedentary fish. Marine Ecology Progress Series, 515:239–250, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/515/
m515p239.pdf.

Vega-Rodriguez:2015:IWT

[VRMKH^{+}15] M. Vega-Rodriguez, F. E. Müller-Karger, P. Hallock, G. A. Quiles-Perez, C. M. Eakin, M. Colella, D. L. Jones, J. Li, I. Soto, L. Guild, S. Lynds, and R. Ruzicka. Influence of water-temperature variability on stony coral diversity in Florida Keys patch reefs. Marine Ecology Progress Series, 528:173–186, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/528/m528p173.pdf.

Valenzuela:2018:SIS

[VRSS18] L. O. Valenzuela, V. J. Rowntree, M. Sironi, and J. Seger. Stable isotopes (δ ^{1}5N, δ ^{13}C, δ ^{34}S) in skin reveal diverse food sources used by southern right whales Eubalaena australis. Marine Ecology Progress Series, 603:243–255, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/603/
m603p243.pdf.

Villazan:2015:HAA

[VSB^{+}15] B. Villazán, T. Salo, F. G. Brun, J. J. Vergara, and M. F. Pedersen. High ammonium availability amplifies the adverse effect of low salinity on eelgrass Zostera marina. Marine Ecology Progress Series, 536:149–162, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m536p149.pdf.

Vermeulen:2011:PEI

[VSG^{+}11] S. Vermeulen, N. Sturaro, S. Gobert, J. M. Bouquegneau, and G. Lepoint. Potential early indicators of anthropogenically derived nutrients: a multiscale stable isotope analysis. Marine Ecology Progress Series, 422:9–22, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/422/
m422p009.pdf.

Vandenabeele:2015:BTD

[VSG^{+}15] S. P. Vandenabeele, E. L. C. Shepard, D. Grémillet, P. J. Butler, G. R. Martin, and R. P. Wilson. Are bio-telemetric devices a drag? Effects of external tags on the diving behaviour of great cormorants. Marine Ecology Progress Series, 519:239–249, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m519p239.pdf.

VandenMeersche:2011:PDE

[VSM11] K. Van den Meersche, K. Soetaert, and J. J. Middelburg. Plankton dynamics in an estuarine plume: a mesocosm ^{13}C and ^{15}N tracer study. Marine Ecology Progress Series, 429: 29–43, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/429/m429p029.pdf.

Vijai:2015:ETE

[VSW^{+}15] D. Vijai, M. Sakai, T. Wakabayashi, H. K. Yoo, Y. Kato, and Y. Sakurai. Effects of temperature on embryonic development and paralarval behavior of the neon flying squid Ommastrephes bartramii. Marine Ecology Progress Series, 529:145–158, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/529/m529p145.pdf.

Verges:2012:IED

[VTB12] A. Vergés, F. Tomas, and E. Ballesteros. Interactive effects of depth and marine protection on predation and herbivory patterns. Marine Ecology Progress Series, 450:55– 65, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/450/m450p055.pdf.

Venkataraman:2019:CPO

[VTSH^{+}19] Y. R. Venkataraman, E. Timmins-Schiffman, M. J. Horwith, A. T. Lowe, B. Nunn, B. Vadopalas, L. H. Spencer, and S. B. Roberts. Characterization of Pacific oyster Crassostrea gigas proteomic response to natural environmental differences. Marine Ecology Progress Series, 610:65–81, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/610/
m610p065.pdf.

VandeWaal:2013:NPI

[VTZ^{+}13] D. B. Van de Waal, U. Tillmann, M. Zhu, B. P. Koch, B. Rost, and U. John. Nutrient pulse induces dynamic changes in cellular C:N:P, amino acids, and paralytic shellfish poisoning toxins in Alexandrium tamarense. Marine Ecology Progress Series, 493:57–69, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/493/m493p057.pdf.

Vieira:2011:EIM

[VUL^{+}11] S. Vieira, A. B. Utkin, A. Lavrov, N. M. Santos, R. Vilar, J. Marques da Silva, and P. Cartaxana. Effects of intertidal microphytobenthos migration on biomass determination via laser-induced fluorescence. Marine Ecology Progress Series, 432:45–52, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/432/m432p045.pdf.

VanOpzeeland:2010:AEA

[VVB^{+}10] I. Van Opzeeland, S. Van Parijs, H. Bornemann, S. Frickenhaus, L. Kindermann, H. Klinck, J. Plötz, and O. Boebel. Acoustic ecology of Antarctic pinnipeds. Marine Ecology Progress Series, 414:267–291, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m414p267.pdf.

Vallee:2019:FRT

[VVB19] V. Vallée, M. C. Villanueva, and F. Blanchard. Functional richness and turnover patterns reveal assembly rules structuring marine fish communities on the continental shelf of French Guiana. Marine Ecology Progress Series, 630:183– 195, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/630/m630p183.pdf.

Vanderklift:2010:SIR

[VW10] M. A. Vanderklift and T. Wernberg. Stable isotopes reveal a consistent consumer-diet relationship across hundreds of kilometres. Marine Ecology Progress Series, 403:53–61, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2010/403/m403p053.pdf.

Vane:2018:RLN

[VWEP18] K. Vane, N. J. Wallsgrove, W. Ekau, and B. N. Popp. Reconstructing lifetime nitrogen baselines and trophic position of Cynoscion acoupa from δ ^{15}N values of amino acids in otoliths. Marine Ecology Progress Series, 597:1–11, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Voorhies:2018:LSM

[VWH18] K. J. Voorhies, J. T. Wootton, and S. K. Henkel. Longstanding signals of marine community structuring by winter storm wave-base. Marine Ecology Progress Series, 603: 135–146, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/603/m603p135.pdf.

vanWoesik:2011:RWL

[vWSGL11] R. van Woesik, K. Sakai, A. Ganase, and Y. Loya. Revisiting the winners and the losers a decade after coral bleaching. Marine Ecology Progress Series, 434:67–76, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/434/
m434p067.pdf.

Vaudo:2016:VMS

[VWW^{+}16] J. J. Vaudo, B. M. Wetherbee, A. D. Wood, K. Weng, L. A. Howey-Jordan, G. M. Harvey, and M. S. Shivji. Vertical movements of shortfin mako sharks Isurus oxyrinchus in the western North Atlantic Ocean are strongly influenced by temperature. Marine Ecology Progress Series, 547:163–175, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/547/
m547p163.pdf.

Work:2014:MAW

[WA14] T. M. Work and G. S. Aeby. Microbial aggregates within tissues infect a diversity of corals throughout the Indo–Pacific. Marine Ecology Progress Series, 500:1–19, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Wong:2019:NSU

[WA19] K. L. C. Wong and A. C. Alfaro. Nutrient storage and utilization in relation to reproductive condition of the New Zealand scallop Pecten novaezelandiae. Marine Ecology Progress Series, 632:101–122, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/632/m632p101.pdf.

Wrede:2019:MIT

[WAA^{+}19] A. Wrede, H. Andresen, R. Asmus, K. H. Wiltshire, and T. Brey. Macrofaunal irrigation traits enhance predictability of nutrient fluxes across the sediment–water interface. Marine Ecology Progress Series, 632:27–42, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/632/
m632p027.pdf.

Wiedmann:2014:FDB

[WAC^{+}14] M. A. Wiedmann, M. Aschan, G. Certain, A. Dolgov, M. Greenacre, E. Johannesen, B. Planque, and R. Primicerio. Functional diversity of the Barents Sea fish community. Marine Ecology Progress Series, 495:205–218, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m495p205.pdf.

Weitzman:2013:UDI

[WADKT13] J. S. Weitzman, K. Aveni-Deforge, J. R. Koseff, and F. I. M. Thomas. Uptake of dissolved inorganic nitrogen by shallow seagrass communities exposed to wave-driven unsteady flow. Marine Ecology Progress Series, 475:65–83, 2013. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m475p065.pdf.

Weigel:2015:LTP

[WAM^{+}15] B. Weigel, H. C. Andersson, H. E. M. Meier, T. Blenckner, M. Snickars, and E. Bonsdorff. Long-term progression and drivers of coastal zoobenthos in a changing system. Marine Ecology Progress Series, 528:141–159, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/528/
m528p141.pdf.

Walton:2014:OAM

[WAMS^{+}14] M. E. M. Walton, I. Al-Maslamani, M. W. Skov, I. Al-Shaikh, I. S. Al-Ansari, H. A. Kennedy, and L. Le Vay. Outwelling from arid mangrove systems is sustained by inwelling of seagrass productivity. Marine Ecology Progress Series, 507:125–137, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/507/m507p125.pdf.

Walther:2010:EOA

[WAP10] K. Walther, K. Anger, and H. O. Pörtner. Effects of ocean acidification and warming on the larval development of the spider crab Hyas araneus from different latitudes (54^{∘} vs. 79^{∘}n). Marine Ecology Progress Series, 417:159–170, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/417/
m417p159.pdf.

Warwick:2014:MMD

[War14] R. M. Warwick. Meiobenthos and macrobenthos are discrete entities and not artefacts of sampling a size continuum: Comment on Bett (2013). Marine Ecology Progress Series, 505:295–298, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m505p295.pdf. See [Bet13] and

reply [Bet14].

Wen:2012:EEM

[WAW^{+}12] C. K. Wen, G. R. Almany, D. H. Williamson, M. S. Pratchett, and G. P. Jones. Evaluating the effects of marine reserves on diet, prey availability and prey selection by juvenile predatory fishes. Marine Ecology Progress Series, 469: 133–144, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/469/m469p133.pdf.

Wan:2012:SVN

[WB12] R. Wan and X. Bian. Size variability and natural mortality dynamics of anchovy Engraulis japonicus eggs under high fishing pressure. Marine Ecology Progress Series, 465:243– 251, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/465/m465p243.pdf.

Wabnitz:2010:ESP

[WBB^{+}10] C. C. C. Wabnitz, G. Balazs, S. Beavers, K. A. Bjorndal, A. B. Bolten, V. Christensen, S. Hargrove, and D. Pauly. Ecosystem structure and processes at Kaloko Honokōhau, focusing on the role of herbivores, including the green sea turtle Chelonia mydas, in reef resilience. Marine Ecology Progress Series, 420:27–44, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/420/m420p027.pdf.

Woolsey:2013:ETE

[WBB13] E. S. Woolsey, M. Byrne, and A. H. Baird. The effects of temperature on embryonic development and larval survival in two scleractinian corals. Marine Ecology Progress Series, 493:179–184, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/493/m493p179.pdf.

Waldbusser:2010:SDP

[WBG10] G. G. Waldbusser, H. Bergschneider, and M. A. Green. Sizedependent pH effect on calcification in post-larval hard clam Mercenaria spp. Marine Ecology Progress Series, 417:171– 182, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/417/m417p171.pdf.

Waggitt:2014:TSC

[WBG^{+}14] J. J. Waggitt, M. Briffa, W. J. Grecian, J. Newton, S. C. Patrick, C. Stauss, and S. C. Votier. Testing for subcolony variation in seabird foraging behaviour: ecological and methodological consequences for understanding colonial living. Marine Ecology Progress Series, 498:275–285, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/498/
m498p275.pdf.

White:2010:PPM

[WBHL10] J. W. White, L. W. Botsford, A. Hastings, and J. L. Largier. Population persistence in marine reserve networks: incorporating spatial heterogeneities in larval dispersal. Marine Ecology Progress Series, 398:49–67, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/398/m398p049.pdf.

Wang:2015:ISP

[WBI^{+}15] S. W. Wang, S. M. Budge, K. Iken, R. R. Gradinger, A. M. Springer, and M. J. Wooller. Importance of sympagic production to Bering Sea zooplankton as revealed from fatty acidcarbon stable isotope analyses. Marine Ecology Progress Series, 518:31–50, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/518/m518p031.pdf.

Wing:2012:RBB

[WBJ12] S. R. Wing, N. A. Beer, and L. Jack. Resource base of blue cod Parapercis colias subpopulations in marginal fjordic habitats is linked to chemoautotrophic production. Marine Ecology Progress Series, 466:205–214, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/466/m466p205.pdf.

Weidberg:2018:WMM

[WBLRN18] N. Weidberg, B. Bularz, S. López-Rodríguez, and S. A. Navarrete. Wave-modulation of mussel daily settlement at contrasting rocky shores in central Chile: topographic regulation of transport mechanisms in the surf zone. Marine Ecology Progress Series, 606:39–53, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/606/m606p039.pdf.

Watkins-Brandt:2011:AIO

[WBLS^{+}11] K. S. Watkins-Brandt, R. M. Letelier, Y. H. Spitz, M. J. Church, D. Böttjer, and A. E. White. Addition of inorganic or organic phosphorus enhances nitrogen and carbon fixation in the oligotrophic North Pacific. Marine Ecology Progress Series, 432:17–29, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m432p017.pdf.

Weimerskirch:2012:FGC

[WBS^{+}12] H. Weimerskirch, S. Bertrand, J. Silva, C. Bost, and S. Peraltilla. Foraging in guanay cormorant and Peruvian booby, the major guano-producing seabirds in the Humboldt Current System. Marine Ecology Progress Series, 458:231–245, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2012/458/m458p231.pdf.

Williamson:2017:DVH

[WBS^{+}17] L. D. Williamson, K. L. Brookes, B. E. Scott, I. M. Graham, and P. M. Thompson. Diurnal variation in harbour porpoise detection — potential implications for management. Marine Ecology Progress Series, 570:223–232, 2017. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m570p223.pdf.

Wang:2014:ETL

[WBW^{+}14] H. Y. Wang, L. W. Botsford, J. W. White, M. J. Fogarty, F. Juanes, A. Hastings, M. D. Holland, and K. Brander. Effects of temperature on life history set the sensitivity to fishing in Atlantic cod Gadus morhua. Marine Ecology Progress Series, 514:217–229, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/514/m514p217.pdf.

Waters:2018:CDM

[WC18] J. M. Waters and D. Craw. Cyclone-driven marine rafting: storms drive rapid dispersal of buoyant kelp rafts. Marine Ecology Progress Series, 602:77–85, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/602/
m602p077.pdf.

Weydmann:2014:STD

[WCG^{+}14] A. Weydmann, J. Carstensen, I. Goszczko, K. Dmoch, A. Olszewska, and S. Kwasniewski. Shift towards the dominance of boreal species in the Arctic: inter-annual and spatial zooplankton variability in the West Spitsbergen Current. Marine Ecology Progress Series, 501:41–52, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m501p041.pdf.

Whiting:2014:CRM

[WCP^{+}14] A. U. Whiting, M. Chaloupka, N. Pilcher, P. Basintal, and C. J. Limpus. Comparison and review of models describing sea turtle nesting abundance. Marine Ecology Progress Series, 508:233–246, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/508/m508p233.pdf.

Whitehead:2017:SSS

[WCRCR17] T. O. Whitehead, M. Connan, Y. Ropert-Coudert, and P. G. Ryan. Subtle but significant segregation in the feeding ecology of sympatric penguins during the critical pre-moult period. Marine Ecology Progress Series, 565:227–236, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/565/
m565p227.pdf.

Winberg:2014:ERM

[WD14] P. C. Winberg and A. R. Davis. Ecological response to MPA zoning following cessation of bait harvesting in an estuarine tidal flat. Marine Ecology Progress Series, 517:171– 180, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/517/m517p171.pdf.

Wagner:2012:DDE

[WDH^{+}12] E. Wagner, B. R. Dumbauld, S. D. Hacker, A. C. Trimble, L. M. Wisehart, and J. L. Ruesink. Density-dependent effects of an introduced oyster, Crassostrea gigas, on a native intertidal seagrass, Zostera marina. Marine Ecology Progress Series, 468:149–160, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/468/m468p149.pdf.

White:2018:CEC

[WDL^{+}18] M. M. White, D. T. Drapeau, L. C. Lubelczyk, V. C. Abel, B. C. Bowler, and W. M. Balch. Calcification of an estuarine coccolithophore increases with ocean acidification when subjected to diurnally fluctuating carbonate chemistry. Marine Ecology Progress Series, 601:59–76, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/601/
m601p059.pdf.

Wilson:2018:OCJ

[WDM^{+}18] M. T. Wilson, A. Dougherty, M. E. Matta, K. L. Mier, and J. A. Miller. Otolith chemistry of juvenile walleye pollock Gadus chalcogrammus in relation to regional hydrography: evidence of spatially split cohorts. Marine Ecology Progress Series, 588:163–178, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/588/m588p163.pdf.

Ware:2019:NTD

[WDM^{+}19] C. Ware, J. A. Dijkstra, K. Mello, A. Stevens, B. O’Brien, and W. Ikedo. A novel three-dimensional analysis of functional architecture that describes the properties of macroalgae as a refuge. Marine Ecology Progress Series, 608:93– 103, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/608/m608p093.pdf.

Wall:2011:STV

[WDNM11] C. C. Wall, B. T. Donahue, D. F. Naar, and D. A. Mann. Spatial and temporal variability of red grouper holes within Steamboat Lumps Marine Reserve, Gulf of Mexico. Marine Ecology Progress Series, 431:243–254, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/431/
m431p243.pdf.

Wedekin:2017:RFS

[WEA^{+}17] L. L. Wedekin, M. H. Engel, A. Andriolo, P. I. Prado, A. N. Zerbini, M. M. C. Marcondes, P. G. Kinas, and P. C. Simões-Lopes. Running fast in the slow lane: rapid population growth of humpback whales after exploitation. Marine Ecology Progress Series, 575:195–206, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/575/m575p195.pdf.

Warwick-Evans:2015:TAR

[WEAG^{+}15] V. Warwick-Evans, P. W. Atkinson, R. D. Gauvain, L. A. Robinson, J. P. Y. Arnould, and J. A. Green. Time-in-area represents foraging activity in a wide-ranging pelagic forager. Marine Ecology Progress Series, 527:233–246, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m527p233.pdf.

Waugh:2019:DPE

[WEF19] A. Waugh, M. Elliott, and A. Franco. Debunking paradigms in estuarine fish species richness. Marine Ecology Progress Series, 613:125–138, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2019/613/m613p125.pdf.

Woelfel:2014:MIR

[WEK14] J. Woelfel, A. Eggert, and U. Karsten. Marginal impacts of rising temperature on Arctic benthic microalgae production based on in situ measurements and modelled estimates. Marine Ecology Progress Series, 501:25–40, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/501/
m501p025.pdf.

Whitehouse:2018:MAR

[WF18] L. M. Whitehouse and M. S. Fowler. Meta-analysis reveals that fisheries co-management alters socio-economic outcomes and resource well-being. Marine Ecology Progress Series, 600:127–140, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/600/m600p127.pdf.

Widmer:2016:ETS

[WFB16] C. L. Widmer, C. J. Fox, and A. S. Brierley. Effects of temperature and salinity on four species of northeastern Atlantic scyphistomae (Cnidaria: Scyphozoa). Marine Ecology Progress Series, 559:73–88, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/559/m559p073.pdf.

Whitmarsh:2014:NAD

[WFBM14] S. K. Whitmarsh, P. G. Fairweather, D. J. Brock, and D. Miller. Nektonic assemblages determined from baited underwater video in protected versus unprotected shallow seagrass meadows on Kangaroo Island, South Australia. Marine Ecology Progress Series, 503:205–218, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/503/
m503p205.pdf.

Wernberg:2019:OPM

[WFD19] T. Wernberg and K. Filbee-Dexter. Opinion piece: Missing the marine forest for the trees. Marine Ecology Progress Series, 612:209–215, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m612p209.pdf.

White:2011:CAP

[WFH11] T. A. White, H. A. Fotherby, and A. R. Hoelzel. Comparative assessment of population genetics and demographic history of two congeneric deep sea fish species living at different depths. Marine Ecology Progress Series, 434:155–164, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/434/
m434p155.pdf.

Winton:2018:EDR

[WFH^{+}18] M. V. Winton, G. Fay, H. L. Haas, M. Arendt, S. Barco, M. C. James, C. Sasso, and R. Smolowitz. Estimating the distribution and relative density of satellite-tagged loggerhead sea turtles using geostatistical mixed effects models. Marine Ecology Progress Series, 586:217–232, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/586/m586p217.pdf.

Wringe:2015:SSC

[WFP15] B. F. Wringe, I. A. Fleming, and C. F. Purchase. Spawning success of cultured and wild male Atlantic cod Gadus morhua does not differ during paired contests. Marine Ecology Progress Series, 535:197–211, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2015/535/m535p197.pdf.

Wang:2010:DVD

[WFS10] J. H. Wang, S. Fisler, and Y. Swimmer. Developing visual deterrents to reduce sea turtle bycatch in gill net fisheries. Marine Ecology Progress Series, 408:241–250, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m408p241.pdf.

Walker:2011:SSV

[WG11] T. N. Walker and J. W. Grahame. Shell shape variation and fitness variables in the gastropod Littorina saxatilis. Marine Ecology Progress Series, 430:103–111, 2011. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m430p103.pdf.

Wicks:2010:SPR

[WGD10] L. C. Wicks, J. P. A. Gardner, and S. K. Davy. Spatial patterns and regional affinities of coral communities at the Kermadec Islands Marine Reserve, New Zealand — a marginal high-latitude site. Marine Ecology Progress Series, 400:101–113, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/400/m400p101.pdf.

Waser:2016:TEP

[WGD^{+}16] A. M. Waser, M. A. Goedknegt, R. Dekker, N. McSweeney, J. I. J. Witte, J. van der Meer, and D. W. Thieltges. Tidal elevation and parasitism: patterns of infection by the rhizocephalan parasite Sacculina carcini in shore crabs Carcinus maenas. Marine Ecology Progress Series, 545:215–225, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/545/
m545p215.pdf.

Weijerman:2019:SLC

[WGD^{+}19] M. Weijerman, A. Grüss, D. Dove, J. Asher, I. D. Williams, C. Kelley, and J. T. C. Drazen. Shining a light on the composition and distribution patterns of mesophotic and subphotic fish communities in Hawai’i. Marine Ecology Progress Series, 630:161–182, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m630p161.pdf.

Wright:2011:RIN

[WGGM11] P. J. Wright, F. M. Gibb, I. M. Gibb, and C. P. Millar. Reproductive investment in the North Sea haddock: temporal and spatial variation. Marine Ecology Progress Series, 432:149–160, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/432/m432p149.pdf.

Wolfe:2015:LPP

[WGLDB15] K. Wolfe, A. Graba-Landry, S. A. Dworjanyn, and M. Byrne. Larval phenotypic plasticity in the boom-and-bust crown-of-thorns seastar, Acanthaster planci. Marine Ecology Progress Series, 539:179–189, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2015/539/m539p179.pdf.

Williams-Grove:2016:APM

[WGS16] L. J. Williams-Grove and S. T. Szedlmayer. Acoustic positioning and movement patterns of red snapper Lutjanus campechanus around artificial reefs in the northern Gulf of Mexico. Marine Ecology Progress Series, 553:233–251, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2016/553/m553p233.pdf.

Wijnhoven:2011:PMA

[WH11] S. Wijnhoven and H. Hummel. Patterns in macrozoobenthic assemblages indicate the state of the environment: insights from the Rhine-meuse estuary. Marine Ecology Progress Series, 436:29–50, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/436/m436p029.pdf.

Wehrberger:2014:MCI

[WH14] F. Wehrberger and J. Herler. Microhabitat characteristics influence shape and size of coral-associated fishes. Marine Ecology Progress Series, 500:203–214, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m500p203.pdf.

Wheeler:2013:USC

[WHA^{+}13] J. D. Wheeler, K. R. Helfrich, E. J. Anderson, B. McGann, P. Staats, A. E. Wargula, K. Wilt, and L. S. Mullineaux. Upward swimming of competent oyster larvae Crassostrea virginica persists in highly turbulent flow as detected by PIV flow subtraction. Marine Ecology Progress Series, 488:171– 185, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m488p171.pdf.

Wagstaff:2014:DDS

[WHB^{+}14] M. C. Wagstaff, K. L. Howell, B. J. Bett, D. S. M. Billett, S. Brault, C. T. Stuart, and M. A. Rex. β-diversity of deep-sea holothurians and asteroids along a bathymetric gradient (NE Atlantic). Marine Ecology Progress Series, 508:177–185, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m508p177.pdf.

Whitman:2019:ESN

[WHB^{+}19] E. R. Whitman, M. R. Heithaus, L. G. Barcia, D. N. Brito, C. Rinaldi, and J. J. Kiszka. Effect of seagrass nutrient content and relative abundance on the foraging behavior of green turtles in the face of a marine plant invasion. Marine Ecology Progress Series, 628:171–182, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/628/m628p171.pdf.

Wilson:2014:CMR

[WHD^{+}14] S. M. Wilson, S. G. Hinch, S. M. Drenner, E. G. Martins, N. B. Furey, D. A. Patterson, D. W. Welch, and S. J. Cooke. Coastal marine and in-river migration behaviour of adult sockeye salmon en route to spawning grounds. Marine Ecology Progress Series, 496:71–84, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/496/m496p071.pdf.

Williard:2017:OOJ

[WHFM17] A. S. Williard, A. G. Hall, I. Fujisaki, and J. B. McNeill. Oceanic overwintering in juvenile green turtles Chelonia mydas from a temperate latitude foraging ground. Marine Ecology Progress Series, 564:235–240, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/564/m564p235.pdf.

Witherington:2012:YST

[WHH12] B. Witherington, S. Hirama, and R. Hardy. Young sea turtles of the pelagic Sargassum-dominated drift community: habitat use, population density, and threats. Marine Ecology Progress Series, 463:1–22, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Watson:2013:URE

[WHH^{+}13] H. Watson, J. G. Hiddink, M. J. Hobbs, T. M. Brereton, and M. J. Tetley. The utility of relative environmental suitability (RES) modelling for predicting distributions of seabirds in the North Atlantic. Marine Ecology Progress Series, 485: 259–273, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/485/m485p259.pdf.

Whiteley:2011:PER

[Whi11] N. M. Whiteley. Physiological and ecological responses of crustaceans to ocean acidification. Marine Ecology Progress Series, 430:257–271, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m430p257.pdf.

Watai:2018:CAE

[WHI^{+}18] M. Watai, Y. Hiraoka, T. Ishihara, I. Yamasaki, T. Ota, S. Ohshimo, and C. A. Strüssmann. Comparative analysis of the early growth history of Pacific bluefin tuna Thunnus orientalis from different spawning grounds. Marine Ecology Progress Series, 607:207–220, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/607/m607p207.pdf.

Witt:2012:BSN

[WHJ^{+}12] M. J. Witt, T. Hardy, L. Johnson, C. M. McClellan, S. K. Pikesley, S. Ranger, P. B. Richardson, J. L. Solandt, C. Speedie, R. Williams, and B. J. Godley. Basking sharks in the northeast Atlantic: spatio-temporal trends from sightings in UK waters. Marine Ecology Progress Series, 459: 121–134, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/459/m459p121.pdf.

Watanabe:2011:PSL

[WHK^{+}11] K. K. Watanabe, H. Hatase, M. Kinoshita, K. Omuta, T. Bando, N. Kamezaki, K. Sato, Y. Matsuzawa, K. Goto, Y. Nakashima, H. Takeshita, J. Aoyama, and K. Tsukamoto. Population structure of the loggerhead turtle Caretta caretta, a large marine carnivore that exhibits alternative foraging behaviors. Marine Ecology Progress Series, 424:273–283, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2010/424/m424p273.pdf.

Wright:2011:OIV

[WHL11] K. J. Wright, D. M. Higgs, and J. M. Leis. Ontogenetic and interspecific variation in hearing ability in marine fish larvae. Marine Ecology Progress Series, 424:1–13, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Wanless:2018:CWD

[WHN^{+}18] S. Wanless, M. P. Harris, M. A. Newell, J. R. Speakman, and F. Daunt. Community-wide decline in the occurrence of lesser sandeels Ammodytes marinus in seabird chick diets at a North Sea colony. Marine Ecology Progress Series, 600:193–206, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2018/600/m600p193.pdf.

Wu:2017:ESS

[WHP^{+}17] Y. Wu, C. G. Hannah, B. Petrie, X. Wang, and E. Devred. Enhanced sea surface temperature due to kelp canopies. Marine Ecology Progress Series, 581:103–117, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/581/
m581p103.pdf.

Waluda:2012:DVR

[WHPT12] C. M. Waluda, S. L. Hill, H. J. Peat, and P. N. Trathan. Diet variability and reproductive performance of macaroni penguins Eudyptes chrysolophus at Bird Island, South Georgia. Marine Ecology Progress Series, 466:261–274, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/466/
m466p261.pdf.

Waggett:2012:TNI

[WHT12] R. J. Waggett, D. R. Hardison, and P. A. Tester. Toxicity and nutritional inadequacy of Karenia brevis: synergistic mechanisms disrupt top-down grazer control. Marine Ecology Progress Series, 444:15–30, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m444p015.pdf.

Weerman:2011:MAD

[WHvdK11] E. J. Weerman, P. M. J. Herman, and J. van de Koppel. Macrobenthos abundance and distribution on a spatially patterned intertidal flat. Marine Ecology Progress Series, 440:95–103, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m440p095.pdf.

Watanuki:2012:CEB

[WI12] Y. Watanuki and M. Ito. Climatic effects on breeding seabirds of the northern Japan Sea. Marine Ecology Progress Series, 454:183–196, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m454p183.pdf.

Willis:2014:WMH

[Wil14] J. Willis. Whales maintained a high abundance of krill; both are ecosystem engineers in the Southern Ocean. Marine Ecology Progress Series, 513:51–69, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/513/
m513p051.pdf.

Wirtz:2012:WEW

[Wir12] K. W. Wirtz. Who is eating whom? Morphology and feeding type determine the size relation between planktonic predators and their ideal prey. Marine Ecology Progress Series, 445:1–12, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Wirtz:2014:BOB

[Wir14] K. W. Wirtz. A biomechanical and optimality-based derivation of prey-size dependencies in planktonic prey selection and ingestion rates. Marine Ecology Progress Series, 507: 81–94, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m507p081.pdf.

Wing:2012:RSA

[WJ12] S. Wing and L. Jack. Resource specialisation among suspension-feeding invertebrates on rock walls in Fiordland, New Zealand, is driven by water column structure and feeding mode. Marine Ecology Progress Series, 452:109–118, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2012/452/m452p109.pdf.

Williams:2013:LEF

[WJHS13] J. Williams, G. P. Jenkins, J. S. Hindell, and S. E. Swearer. Linking environmental flows with the distribution of black bream Acanthopagrus butcheri eggs, larvae and prey in a drought affected estuary. Marine Ecology Progress Series, 483:273–287, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/483/m483p273.pdf.

Wing:2014:SMM

[WJS^{+}14] S. R. Wing, L. Jack, O. Shatova, J. J. Leichter, D. Barr, R. D. Frew, and M. Gault-Ringold. Seabirds and marine mammals redistribute bioavailable iron in the Southern Ocean. Marine Ecology Progress Series, 510:1–13, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Wilkinson:2018:VUS

[WJW^{+}18] B. P. Wilkinson, J. Jahncke, P. Warzybok, R. W. Bradley, and S. A. Shaffer. Variable utilization of shelf break-associated habitats by chick-brooding rhinoceros auklets in the California Current system. Marine Ecology Progress Series, 590: 211–226, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m590p211.pdf.

Watson:2016:RLS

[WJWC16] W. H. Watson III, S. K. Johnson, C. D. Whitworth, and C. C. Chabot. Rhythms of locomotion and seasonal changes in activity expressed by horseshoe crabs in their natural habitat. Marine Ecology Progress Series, 542:109–121, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/542/
m542p109.pdf.

Waska:2010:DMM

[WK10] H. Waska and G. Kim. Differences in microphytobenthos and macrofaunal abundances associated with groundwater discharge in the intertidal zone. Marine Ecology Progress Series, 407:159–172, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/407/m407p159.pdf.

Wootton:2016:DTI

[WK16] H. F. Wootton and M. J. Keough. Disturbance type and intensity combine to affect resilience of an intertidal community. Marine Ecology Progress Series, 560:121–133, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/560/
m560p121.pdf.

Will:2018:VTL

[WK18] A. P. Will and A. S. Kitaysky. Variability in trophic level and habitat use in response to environmental forcing: isotopic niche dynamics of breeding seabirds in the southeastern Bering Sea. Marine Ecology Progress Series, 593:247–260, 2018. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m593p247.pdf.

Walpersdorf:2017:LSO

[WKE^{+}17] E. Walpersdorf, M. Kühl, B. Elberling, T. J. Andersen, B. U. Hansen, M. Pejrup, and R. N. Glud. In situ oxygen dynamics and carbon turnover in an intertidal sediment (Skallingen, Denmark). Marine Ecology Progress Series, 566:49–65, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m566p049.pdf.

Will:2018:RLK

[WKK18] A. P. Will, E. V. Kitaiskaia, and A. S. Kitaysky. Redlegged kittiwake feathers link food availability to environmental changes in the Bering Sea. Marine Ecology Progress Series, 593:261–274, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m593p261.pdf.

Wlodarska-Kowalczuk:2012:SDF

[WKRW^{+}12] M. Włodarska-Kowalczuk, P. E. Renaud, J. M. Wesławski, S. K. J. Cochrane, and S. G. Denisenko. Species diversity, functional complexity and rarity in Arctic fjordic versus open shelf benthic systems. Marine Ecology Progress Series, 463:73–87, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m463p073.pdf.

Wagner:2010:SCH

[WKvW10] D. E. Wagner, P. Kramer, and R. van Woesik. Species composition, habitat, and water quality influence coral bleaching in southern Florida. Marine Ecology Progress Series, 408: 65–78, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/408/m408p065.pdf.

Wang:2019:PCS

[WKX^{+}19] Y. Wang, J. Kang, P. Xiang, Y. Ye, H. Lin, and M. Lin. Phytoplankton communities and size-fractioned chlorophyll-a in newly opened summer waters of the central Arctic Ocean. Marine Ecology Progress Series, 622:67–82, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/622/
m622p067.pdf.

Walve:2010:SCB

[WL10] J. Walve and U. Larsson. Seasonal changes in Baltic Sea seston stoichiometry: the influence of diazotrophic cyanobacteria. Marine Ecology Progress Series, 407:13–25, 2010. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m407p013.pdf.

Wing:2011:VEC

[WL11] S. R. Wing and J. J. Leichter. Variation in environmental conditions in a subtidal prey refuge: effects of salinity stress, food availability and predation on mussels in a fjord system. Marine Ecology Progress Series, 422:201–210, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/422/
m422p201.pdf.

Wilson:2019:CCP

[WL19] K. L. Wilson and H. K. Lotze. Climate change projections reveal range shifts of eelgrass Zostera marina in the Northwest Atlantic. Marine Ecology Progress Series, 620: 47–62, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/620/m620p047.pdf.

Wilberg:2011:ODH

[WLB^{+}11] M. J. Wilberg, M. E. Livings, J. S. Barkman, B. T. Morris, and J. M. Robinson. Overfishing, disease, habitat loss, and potential extirpation of oysters in upper Chesapeake Bay. Marine Ecology Progress Series, 436:131–144, 2011. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m436p131.pdf.

Whitehead:2015:FDA

[WLB^{+}15] A. L. Whitehead, P. O’B. Lyver, G. Ballard, K. Barton, B. J. Karl, K. M. Dugger, S. Jennings, A. Lescroël, P. R. Wilson, and D. G. Ainley. Factors driving Adélie penguin chick size, mass and condition at colonies of different sizes in the Southern Ross Sea. Marine Ecology Progress Series, 523:199–213, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/523/m523p199.pdf.

Walters:2013:EKB

[WLBB^{+}13] S. Walters, S. Lowerre-Barbieri, J. Bickford, J. Tustison, and J. H. Landsberg. Effects of Karenia brevis red tide on the spatial distribution of spawning aggregations of sand seatrout Cynoscion arenarius in Tampa Bay, Florida. Marine Ecology Progress Series, 479:191–202, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m479p191.pdf.

Wu:2017:ISI

[WLF^{+}17] L. Wu, X. Liu, P. Fu, L. Xu, X. Wang, Y. Sun, and Y. Li. Identification and stable isotope analyses of flying fish scales from ornithogenic sediments at three islands in the South China Sea. Marine Ecology Progress Series, 585:175–183, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/585/
m585p175.pdf.

Wheeler:2017:LSS

[WLH^{+}17] J. D. Wheeler, E. Luo, K. R. Helfrich, E. J. Anderson, V. R. Starczak, and L. S. Mullineaux. Light stimulates swimming behavior of larval eastern oysters Crassostrea virginica in turbulent flow. Marine Ecology Progress Series, 571:109–120, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/571/m571p109.pdf.

Wyatt:2010:PNF

[WLHW10] A. S. J. Wyatt, R. J. Lowe, S. Humphries, and A. M. Waite. Particulate nutrient fluxes over a fringing coral reef: relevant scales of phytoplankton production and mechanisms of supply. Marine Ecology Progress Series, 405:113–130, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/405/
m405p113.pdf.

Weidberg:2014:ENS

[WLL^{+}14] N. Weidberg, C. Lobón, E. López, L. García Flórez, M. d. P. Fernández Rueda, J. Largier, and J. L. Acuña. Effect of nearshore surface slicks on meroplankton distribution: role of larval behaviour. Marine Ecology Progress Series, 506:15–30, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/506/m506p015.pdf.

Wall:2012:SSM

[WLM12] C. C. Wall, C. Lembke, and D. A. Mann. Shelf-scale mapping of sound production by fishes in the eastern Gulf of Mexico, using autonomous glider technology. Marine Ecology Progress Series, 449:55–64, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/449/m449p055.pdf.

Wraith:2013:BTA

[WLM^{+}13] J. Wraith, T. Lynch, T. E. Minchinton, A. Broad, and A. R. Davis. Bait type affects fish assemblages and feeding guilds observed at baited remote underwater video stations. Marine Ecology Progress Series, 477:189–199, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/477/
m477p189.pdf.

Williams:2010:EMP

[WLP^{+}10] A. J. Williams, L. R. Little, A. E. Punt, B. D. Mapstone, C. R. Davies, and M. R. Heupel. Exploring movement patterns of an exploited coral reef fish when tagging data are limited. Marine Ecology Progress Series, 405:87–99, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/405/
m405p087.pdf.

Wedding:2011:QSS

[WLP^{+}11] L. M. Wedding, C. A. Lepczyk, S. J. Pittman, A. M. Friedlander, and S. Jorgensen. Quantifying seascape structure: extending terrestrial spatial pattern metrics to the marine realm. Marine Ecology Progress Series, 427:219–232, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m427p219.pdf.

Weber:2010:PMN

[WM10] E. D. Weber and S. McClatchie. Predictive models of northern anchovy Engraulis mordax and Pacific sardine Sardinops sagax spawning habitat in the California Current. Marine Ecology Progress Series, 406:251–263, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m406p251.pdf.

Whitfield:2014:NFC

[WMB^{+}14] P. E. Whitfield, R. C. Muñoz, C. A. Buckel, B. P. Degan, D. W. Freshwater, and J. A. Hare. Native fish community structure and Indo–Pacific lionfish Pterois volitans densities along a depth-temperature gradient in Onslow Bay, North Carolina, USA. Marine Ecology Progress Series, 509:241– 254, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m509p241.pdf.

Williams:2015:IDO

[WMCC15] S. M. Williams, P. J. Mumby, I. Chollett, and J. Cortés. Importance of differentiating Orbicella reefs from gorgonian plains for ecological assessments of Caribbean reefs. Marine Ecology Progress Series, 530:93–101, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/530/
m530p093.pdf.

Wells:2017:MMF

[WMH17] S. Wells, J. McConaugha, and L. Horth. Multiple mating by females in the Chesapeake Bay blue crab Callinectes sapidus population. Marine Ecology Progress Series, 564: 67–75, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/564/m564p067.pdf.

Wright:2014:EAS

[WMHW14] S. Wright, J. D. Metcalfe, S. Hetherington, and R. Wilson. Estimating activity-specific energy expenditure in a teleost fish, using accelerometer loggers. Marine Ecology Progress Series, 496:19–32, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m496p019.pdf.

White:2014:EPE

[WMMC14] M. M. White, L. S. Mullineaux, D. C. McCorkle, and A. L. Cohen. Elevated pCO 2 exposure during fertilization of the bay scallop Argopecten irradians reduces larval survival but not subsequent shell size. Marine Ecology Progress Series, 498:173–186, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m498p173.pdf.

Winiarski:2013:SEM

[WMPM13] K. J. Winiarski, D. L. Miller, P. W. C. Paton, and S. R. McWilliams. Spatially explicit model of wintering common loons: conservation implications. Marine Ecology Progress Series, 492:273–283, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m492p273.pdf.

Watson:2010:RPL

[WMS^{+}10] J. R. Watson, S. Mitarai, D. A. Siegel, J. E. Caselle, C. Dong, and J. C. McWilliams. Realized and potential larval connectivity in the Southern California Bight. Marine Ecology Progress Series, 401:31–48, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m401p031.pdf.

Wang:2011:PIB

[WMS^{+}11] J. T. Wang, P. J. Meng, E. Sampayo, S. L. Tang, and C. A. Chen. Photosystem II breakdown induced by reactive oxygen species in freshly-isolated Symbiodinium from Montipora (Scleractinia; Acroporidae). Marine Ecology Progress Series, 422:51–62, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/422/m422p051.pdf.

Welcker:2016:DSO

[WN16] J. Welcker and G. Nehls. Displacement of seabirds by an offshore wind farm in the North Sea. Marine Ecology Progress Series, 554:173–182, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2016/554/m554p173.pdf.

Williams:2015:LAN

[WNA^{+}15] L. M. Williams, C. L. Nivison, W. G. Ambrose, Jr., R. Dobbin, and W. L. Locke, V. Lack of adult novel northern lineages of invasive green crab Carcinus maenas along much of the northern US Atlantic coast. Marine Ecology Progress Series, 532:153–159, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m532p153.pdf.

Winters:2011:ESH

[WNF^{+}11] G. Winters, P. Nelle, B. Fricke, G. Rauch, and T. B. H. Reusch. Effects of a simulated heat wave on photophysiology and gene expression of high- and low-latitude populations of Zostera marina. Marine Ecology Progress Series, 435:83–95, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/435/m435p083.pdf.

Wasmund:2019:EGS

[WNG^{+}19] N. Wasmund, G. Nausch, M. Gerth, S. Busch, C. Burmeister, R. Hansen, and B. Sadkowiak. Extension of the growing season of phytoplankton in the western Baltic Sea in response to climate change. Marine Ecology Progress Series, 622:1–16, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Wild:2010:OMR

[WNNH10] C. Wild, W. Niggl, M. S. Naumann, and A. F. Haas. Organic matter release by Red Sea coral reef organisms—potential effects on microbial activity and in situ O 2 availability. Marine Ecology Progress Series, 411:61–71, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m411p061.pdf.

Waugh:2012:LFA

[WNNN12] C. A. Waugh, P. D. Nichols, M. C. Noad, and S. Bengtson Nash. Lipid and fatty acid profiles of migrating southern hemisphere humpback whales Megaptera novaeangliae. Marine Ecology Progress Series, 471:271–281, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/471/
m471p271.pdf.

Williamson:2014:SMG

[WNP^{+}14] C. J. Williamson, J. Najorka, R. Perkins, M. L. Yallop, and J. Brodie. Skeletal mineralogy of geniculate corallines: providing context for climate change and ocean acidification research. Marine Ecology Progress Series, 513:71–84, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m513p071.pdf.

White:2011:NCF

[WO11] L. F. White and L. C. Orr. Native clams facilitate invasive species in an eelgrass bed. Marine Ecology Progress Series, 424:87–95, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/424/m424p087.pdf.

Watt:2015:DDB

[WOHJ^{+}15] C. A. Watt, J. R. Orr, M. P. Heide-Jørgensen, N. H. Nielsen, and S. H. Ferguson. Differences in dive behaviour among the world’s three narwhal Monodon monoceros populations correspond with dietary differences. Marine Ecology Progress Series, 525:273–285, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/525/m525p273.pdf.

Williams:2010:SAC

[WOWG10] R. Williams, T. A. Okey, S. S. Wallace, and V. F. Gallucci. Shark aggregation in coastal waters of British Columbia. Marine Ecology Progress Series, 414:249–256, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/414/
m414p249.pdf.

Wirtz:2010:DCN

[WP10] K. W. Wirtz and M. Pahlow. Dynamic chlorophyll and nitrogen: carbon regulation in algae optimizes instantaneous growth rate. Marine Ecology Progress Series, 402:81–96, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2010/402/m402p081.pdf.

Woehler:2014:CCA

[WPB^{+}14] E. Woehler, T. A. Patterson, M. V. Bravington, A. J. Hobday, and L. E. Chambers. Climate and competition in abundance trends in native and invasive Tasmanian gulls. Marine Ecology Progress Series, 511:249–263, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/511/m511p249.pdf.

Waugh:2016:OBF

[WPF^{+}16] S. M. Waugh, S. C. Patrick, D. P. Filippi, G. A. Taylor, and J. P. Y. Arnould. Overlap between flesh-footed shearwater Puffinus carneipes foraging areas and commercial fisheries in New Zealand waters. Marine Ecology Progress Series, 551:249–260, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/551/m551p249.pdf.

Wong:2010:PSS

[WPK10] M. C. Wong, C. H. Peterson, and J. Kay. Prey size selection and bottom type influence multiple predator effects in a crab-bivalve system. Marine Ecology Progress Series, 409:143–156, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m409p143.pdf.

Wall-Palmer:2018:VDD

[WPML^{+}18] D. Wall-Palmer, B. Metcalfe, M. J. Leng, H. J. Sloane, G. Ganssen, P. N. Vinayachandran, and C. W. Smart. Vertical distribution and diurnal migration of atlantid heteropods. Marine Ecology Progress Series, 587:1–15, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Wright:2014:HRE

[WPMP14] A. K. Wright, K. V. Ponganis, B. I. McDonald, and P. J. Ponganis. Heart rates of emperor penguins diving at sea: implications for oxygen store management. Marine Ecology Progress Series, 496:85–98, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/496/m496p085.pdf.

Wheatley:2012:EMR

[WPNK12] P. V. Wheatley, H. Peckham, S. D. Newsome, and P. L. Koch. Estimating marine resource use by the American crocodile Crocodylus acutus in southern Florida, USA. Marine Ecology Progress Series, 447:211–229, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/447/m447p211.pdf.

Wong:2011:EEH

[WPP11] M. C. Wong, C. H. Peterson, and M. F. Piehler. Evaluating estuarine habitats using secondary production as a proxy for food web support. Marine Ecology Progress Series, 440:11–25, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m440p011.pdf.

Weston:2018:MVB

[WPRS18] R. A. Weston, R. Perissinotto, G. M. Rishworth, and P. P. Steyn. Macroinvertebrate variability between microhabitats of peritidal stromatolites along the South African coast. Marine Ecology Progress Series, 605:37–47, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/605/
m605p037.pdf.

Watanabe:2019:HBW

[WPS^{+}19] Y. Y. Watanabe, N. L. Payne, J. M. Semmens, A. Fox, and C. Huveneers. Hunting behaviour of white sharks recorded by animal-borne accelerometers and cameras. Marine Ecology Progress Series, 621:221–227, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/621/m621p221.pdf.

Whitman:2012:BFE

[WR12] E. R. Whitman and M. A. Reidenbach. Benthic flow environments affect recruitment of Crassostrea virginica larvae to an intertidal oyster reef. Marine Ecology Progress Series, 463:177–191, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/463/m463p177.pdf.

Welch:2015:SFC

[WRB^{+}15] J. M. Welch, K. A. Reinsel, K. A. Battles, A. O. Romero, J. M. Blaine, R. L. Sendi, and R. B. Forward, Jr. Settlement of fiddler crab megalopae on a North Carolina (USA) sandflat: species identification using multiplex PCR provides evidence for selective settlement. Marine Ecology Progress Series, 523:115–123, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/523/m523p115.pdf.

Windle:2012:STV

[WRDF12] M. J. S. Windle, G. A. Rose, R. Devillers, and M. J. Fortin. Spatio-temporal variations in invertebrate–cod–environment relationships on the Newfoundland–Labrador Shelf, 1995– 2009. Marine Ecology Progress Series, 469:263–278, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m469p263.pdf.

Wheeler:2016:ECS

[WRFM16] S. G. Wheeler, A. D. Russell, J. S. Fehrenbacher, and S. G. Morgan. Evaluating chemical signatures in a coastal upwelling region to reconstruct water mass associations of settlementstage rockfishes. Marine Ecology Progress Series, 550:191– 206, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m550p191.pdf.

Wall:2012:RLS

[WRGP12] C. C. Wall, B. S. Rodgers, C. J. Gobler, and B. J. Peterson. Responses of loggerhead sponges Spechiospongia vesparium during harmful cyanobacterial blooms in a sub-tropical lagoon. Marine Ecology Progress Series, 451:31–43, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/451/
m451p031.pdf.

Wells:2012:NOY

[WRI12] R. J. D. Wells, J. R. Rooker, and D. G. Itano. Nursery origin of yellowfin tuna in the Hawaiian Islands. Marine Ecology Progress Series, 461:187–196, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/461/m461p187.pdf.

Weinberger:2011:ELS

[WRO^{+}11] F. Weinberger, S. Rohde, Y. Oschmann, L. Shahnaz, S. Dobretsov, and M. Wahl. Effects of limitation stress and of disruptive stress on induced antigrazing defense in the bladder wrack Fucus vesiculosus. Marine Ecology Progress Series, 427:83–94, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m427p083.pdf.

Weijerman:2017:CAT

[WRP^{+}17] M. Weijerman, S. Robinson, F. Parrish, J. Polovina, and C. Littnan. Comparative application of trophic ecosystem models to evaluate drivers of endangered Hawaiian monk seal populations. Marine Ecology Progress Series, 582:215– 229, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m582p215.pdf.

Wienecke:2010:MJF

[WRR10] B. Wienecke, B. Raymond, and G. Robertson. Maiden journey of fledgling emperor penguins from the Mawson Coast, East Antarctica. Marine Ecology Progress Series, 410:269– 282, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/410/m410p269.pdf.

Warry:2016:NSI

[WRW^{+}16a] F. Y. Warry, P. Reich, R. J. Woodland, J. R. Thomson, R. Mac Nally, and P. L. M. Cook. Nitrogen stable isotope values of large-bodied consumers reflect urbanization of coastal catchments. Marine Ecology Progress Series, 542: 25–37, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/542/m542p025.pdf.

Welch:2016:CCA

[WRW^{+}16b] J. M. Welch, K. A. Reinsel, E. E. Waddell, A. J. Wright, and R. B. Forward, Jr. Chemical cues from adult fiddler crabs stimulate molting of conspecific megalopae: evidence from field-caught individuals. Marine Ecology Progress Series, 553:147–153, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/553/m553p147.pdf.

Wilson:2010:APM

[WS10] S. E. Wilson and D. K. Steinberg. Autotrophic picoplankton in mesozooplankton guts: evidence of aggregate feeding in the mesopelagic zone and export of small phytoplankton. Marine Ecology Progress Series, 412:11–27, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/412/
m412p011.pdf.

Woodland:2011:DJT

[WS11] R. J. Woodland and D. H. Secor. Differences in juvenile trophic niche for two coastal fish species that use marine and estuarine nursery habitats. Marine Ecology Progress Series, 439:241–254, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/439/m439p241.pdf.

Wieland:2012:EFA

[WS12] K. Wieland and H. Siegstad. Environmental factors affecting recruitment of northern shrimp Pandalus borealis in West Greenland waters. Marine Ecology Progress Series, 469:297– 306, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m469p297.pdf.

Watt:2013:RCI

[WS13] C. A. Watt and R. A. Scrosati. Regional consistency of intertidal elevation as a mediator of seaweed canopy effects on benthic species richness, diversity, and composition. Marine Ecology Progress Series, 491:91–99, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/491/
m491p091.pdf.

Whalen:2017:SFD

[WS17] M. A. Whalen and J. J. Stachowicz. Suspension feeder diversity enhances community filtration rates in different flow environments. Marine Ecology Progress Series, 570:1–13, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Weeks:2013:OSF

[WSC13] S. J. Weeks, C. Steinberg, and B. C. Congdon. Oceanography and seabird foraging: within-season impacts of increasing sea-surface temperature on the Great Barrier Reef. Marine Ecology Progress Series, 490:247–254, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/490/
m490p247.pdf.

Wahl:2010:EAR

[WSD^{+}10] M. Wahl, L. Shahnaz, S. Dobretsov, M. Saha, F. Symanowski, K. David, T. Lachnit, M. Vasel, and F. Weinberger. Ecology of antifouling resistance in the bladder wrack Fucus vesiculosus: patterns of microfouling and antimicrobial protection. Marine Ecology Progress Series, 411:33–48, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m411p033.pdf.

Wells:2012:PDC

[WSF^{+}12] B. K. Wells, J. A. Santora, J. C. Field, R. B. MacFarlane, B. B. Marinovic, and W. J. Sydeman. Population dynamics of Chinook salmon Oncorhynchus tshawytscha relative to prey availability in the central California coastal region. Marine Ecology Progress Series, 457:125–137, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/457/m457p125.pdf.

Walles:2016:NDD

[WSHY16] B. Walles, A. C. Smaal, P. M. J. Herman, and T. Ysebaert. Niche dimension differs among life-history stages of Pacific oysters in intertidal environments. Marine Ecology Progress Series, 562:113–122, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/562/m562p113.pdf.

Weinstein:2016:CGB

[WSK^{+}16] D. K. Weinstein, A. Sharifi, J. S. Klaus, T. B. Smith, S. J. Giri, and K. P. Helmle. Coral growth, bioerosion, and secondary accretion of living orbicellid corals from mesophotic reefs in the US Virgin Islands. Marine Ecology Progress Series, 559:45–63, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/559/m559p045.pdf. See corrigen-

dum [Ano19d].

Westerberg:2014:BSN

[WSL^{+}14] H. Westerberg, N. Sjöberg, I. Lagenfelt, K. Aarestrup, and D. Righton. Behaviour of stocked and naturally recruited European eels during migration. Marine Ecology Progress Series, 496:145–157, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m496p145.pdf.

Wall:2013:LSP

[WSLM13] C. C. Wall, P. Simard, C. Lembke, and D. A. Mann. Largescale passive acoustic monitoring of fish sound production on the West Florida shelf. Marine Ecology Progress Series, 484:173–188, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/484/m484p173.pdf.

Wells:2016:MEP

[WSS^{+}16] B. K. Wells, J. A. Santora, I. D. Schroeder, N. Mantua, W. J. Sydeman, D. D. Huff, and J. C. Field. Marine ecosystem perspectives on Chinook salmon recruitment: a synthesis of empirical and modeling studies from a California upwelling system. Marine Ecology Progress Series, 552:271– 284, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/552/m552p271.pdf.

Watson:2016:IIL

[WSZS16] M. G. Watson, A. J. Scardino, L. Zalizniak, and J. Shimeta. Inhibition of invertebrate larval settlement by biofilm ciliates. Marine Ecology Progress Series, 557:77–90, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/557/
m557p077.pdf.

Weimerskirch:2015:PSW

[WTC^{+}15] H. Weimerskirch, A. Tarroux, O. Chastel, K. Delord, Y. Cherel, and S. Descamps. Population-specific wintering distributions of adult south polar skuas over three oceans.

Marine Ecology Progress Series, 538:229–237, 2015. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m538p229.pdf.

Werner:2019:ELA

[WTD^{+}19] K. M. Werner, M. H. Taylor, R. Diekmann, J. Lloret, C. Möllmann, R. Primicerio, and H. O. Fock. Evidence for limited adaptive responsiveness to large-scale spatial variation of habitat quality. Marine Ecology Progress Series, 629:179–191, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m629p179.pdf.

Wege:2016:CSF

[WTdB16] M. Wege, C. A. Tosh, P. J. N. de Bruyn, and M. N. Bester. Cross-seasonal foraging site fidelity of subantarctic fur seals: implications for marine conservation areas. Marine Ecology Progress Series, 554:225–239, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/554/m554p225.pdf.

Wangensteen:2011:WSC

[WTGC^{+}11] O. S. Wangensteen, X. Turon, A. García-Cisneros, M. Recasens, J. Romero, and C. Palacín. A wolf in sheep’s clothing: carnivory in dominant sea urchins in the Mediterranean. Marine Ecology Progress Series, 441:117–128, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/441/
m441p117.pdf.

Wright:2010:ANC

[WTGG10] P. J. Wright, D. Tobin, F. M. Gibb, and I. M. Gibb. Assessing nursery contribution to recruitment: relevance of closed areas to haddock Melanogrammus aeglefinus. Marine Ecology Progress Series, 400:221–232, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/400/m400p221.pdf.

Wilson:2018:ALD

[WTP^{+}18] P. Wilson, M. Thums, C. Pattiaratchi, M. Meekan, K. Pendoley, R. Fisher, and S. Whiting. Artificial light disrupts the nearshore dispersal of neonate flatback turtles Natator depressus. Marine Ecology Progress Series, 600:179–192, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m600p179.pdf.

Wilson:2012:TIA

[WV12] R. P. Wilson and S. P. Vandenabeele. Technological innovation in archival tags used in seabird research. Marine Ecology Progress Series, 451:245–262, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m451p245.pdf.

Werbrouck:2017:AHE

[WVD^{+}17] E. Werbrouck, A. Vanreusel, D. Deregibus, D. Van Gansbeke, and M. De Troch. Antarctic harpacticoids exploit different trophic niches: a summer snapshot using fatty acid trophic markers (Potter Cove, King George Island). Marine Ecology Progress Series, 568:59–71, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/568/m568p059.pdf.

Wenger:2018:MSS

[WvLF18] L. N. Wenger, J. R. van Lier, and C. J. Fulton. Microhabitat selectivity shapes the seascape ecology of a carnivorous macroalgae-associated tropical fish. Marine Ecology Progress Series, 590:187–200, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/590/m590p187.pdf.

Wierucka:2014:EME

[WVM^{+}14] K. Wierucka, P. Verborgh, R. Meade, L. Colmant, P. Gauffier, R. Esteban, R. de Stephanis, and A. Cañadas. Effects of a morbillivirus epizootic on long-finned pilot whales Globicephala melas in Spanish Mediterranean waters. Marine Ecology Progress Series, 502:1–10, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Wing:2015:OSR

[WW15] S. R. Wing and L. Wing. Ontogenetic shifts in resource use by the sea urchin Evechinus chloroticus across an ecotone. Marine Ecology Progress Series, 535:177–184, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/535/
m535p177.pdf.

Willis:2013:TCD

[WWAL13] T. V. Willis, K. A. Wilson, K. E. Alexander, and W. B. Leavenworth. Tracking cod diet preference over a century in the northern Gulf of Maine: historic data and modern analysis. Marine Ecology Progress Series, 474:263–276, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/474/
m474p263.pdf.

Wang:2018:PCS

[WWAS18] V. H. Wang, J. W. White, S. A. Arnott, and F. S. Scharf. Population connectivity of southern flounder in the US South Atlantic revealed by otolith chemical analysis. Marine Ecology Progress Series, 596:165–179, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2018/596/m596p165.pdf.

Wang:2014:IBM

[WWB14] L. Wang, H. Wei, and H. P. Batchelder. Individual-based modelling of Calanus sinicus population dynamics in the Yellow Sea. Marine Ecology Progress Series, 503:75–97, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2014/503/m503p075.pdf.

West:2017:DCB

[WWB^{+}17] K. L. West, W. A. Walker, R. W. Baird, J. G. Mead, and P. W. Collins. Diet of Cuvier’s beaked whales Ziphius cavirostris from the North Pacific and a comparison with their diet worldwide. Marine Ecology Progress Series, 574:227–242, 2017. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m574p227.pdf.

Webb:2012:SOB

[WWG12] S. D. Webb, S. H. Woodcock, and B. M. Gillanders. Sources of otolith barium and strontium in estuarine fish and the influence of salinity and temperature. Marine Ecology Progress Series, 453:189–199, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/453/m453p189.pdf.

Wei:2013:SGN

[WWG13a] K. Wei, A. R. Wood, and J. P. A. Gardner. Seascape genetics of the New Zealand greenshell mussel: sea surface temperature explains macrogeographic scale genetic variation. Marine Ecology Progress Series, 477:107–121, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/477/
m477p107.pdf.

Woodson:2013:WOC

[WWG^{+}13b] L. E. Woodson, B. K. Wells, C. B. Grimes, R. P. Franks, J. A. Santora, and M. H. Carr. Water and otolith chemistry identify exposure of juvenile rockfish to upwelled waters in an open coastal system. Marine Ecology Progress Series, 473:261–273, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2013/473/m473p261.pdf.

Wilson:2013:SMP

[WWGD13] C. J. Wilson, P. S. Wilson, C. A. Greene, and K. H. Dunton. Seagrass meadows provide an acoustic refuge for estuarine fish. Marine Ecology Progress Series, 472:117–127, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/472/
m472p117.pdf.

Waters:2015:DWS

[WWKS15] L. G. Waters, T. G. Wolcott, D. Kamykowski, and G. Sinclair. Deep-water seed populations for red tide blooms in the Gulf of Mexico. Marine Ecology Progress Series, 529:1–16, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Watson:2015:UAA

[WWO^{+}15] E. B. Watson, C. Wigand, A. J. Oczkowski, K. Sundberg, D. Vendettuoli, S. Jayaraman, K. Saliba, and J. T. Morris. Ulva additions alter soil biogeochemistry and negatively impact Spartina alterniflora growth. Marine Ecology Progress Series, 532:59–72, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/532/m532p059.pdf.

Woodson:2013:SGO

[WWW^{+}13] L. E. Woodson, B. K. Wells, P. K. Weber, R. B. MacFarlane, G. E. Whitman, and R. C. Johnson. Size, growth, and origin-dependent mortality of juvenile Chinook salmon Oncorhynchus tshawytscha during early ocean residence. Marine Ecology Progress Series, 487:163–175, 2013. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m487p163.pdf.

Woodland:2019:RBH

[WWZ^{+}19] R. J. Woodland, F. Y. Warry, Y. Zhu, R. Mac Nally, P. Reich, G. P. Jenkins, D. Brehm, and P. L. M. Cook. Role of benthic habitat structure and riverine connectivity in controlling the spatial distribution and ecology of estuarine fish. Marine Ecology Progress Series, 630:197–214, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/630/
m630p197.pdf.

Wedding:2015:HBP

[WY15] L. Wedding and M. M. Yoklavich. Habitat-based predictive mapping of rockfish density and biomass off the central California coast. Marine Ecology Progress Series, 540: 235–250, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m540p235.pdf.

Watanuki:2016:FMC

[WYI^{+}16] Y. Watanuki, A. Yamashita, M. Ishizuka, Y. Ikenaka, S. M. M. Nakayama, C. Ishii, T. Yamamoto, M. Ito, T. Kuwae, and P. N. Trathan. Feather mercury concentration in streaked shearwaters wintering in separate areas of southeast Asia. Marine Ecology Progress Series, 546:263–269, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/546/
m546p263.pdf.

Wada:2010:SSD

[WYS^{+}10] T. Wada, T. Yamada, D. Shimizu, M. Aritaki, H. Sudo, Y. Yamashita, and M. Tanaka. Successful stocking of a depleted species, spotted halibut Verasper variegatus, in Miyako Bay, Japan: evaluation from post-release surveys and landings. Marine Ecology Progress Series, 407:243–255, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/407/
m407p243.pdf.

Wu:2014:ZDT

[WZP^{+}14] D. Wu, M. Zhou, S. D. Pierce, J. A. Barth, and T. Cowles. Zooplankton distribution and transport in the California Current off Oregon. Marine Ecology Progress Series, 508:87– 103, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/508/m508p087.pdf.

Xu:2014:PBR

[XGLH14] J. Xu, K. Gao, Y. Li, and D. A. Hutchins. Physiological and biochemical responses of diatoms to projected ocean changes. Marine Ecology Progress Series, 515:73–81, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m515p073.pdf.

Xu:2017:LBF

[XLXS17] Z. Xu, S. Liu, R. Xiang, and G. Song. Live benthic Foraminifera in the Yellow Sea and the East China Sea: vertical distribution, nitrate storage, and potential denitrification. Marine Ecology Progress Series, 571:65–81, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/571/
m571p065.pdf.

Xu:2019:SPF

[XS19] Y. Xu and J. Soininen. Spatial patterns of functional diversity and composition in marine benthic ciliates along the coast of China. Marine Ecology Progress Series, 627:49–60, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m627p049.pdf.

Xie:2015:EIT

[XTW^{+}15] Y. Xie, G. H. Tilstone, C. Widdicombe, E. M. S. Woodward, C. Harris, and M. K. Barnes. Effect of increases in temperature and nutrients on phytoplankton community structure and photosynthesis in the western English Channel. Marine Ecology Progress Series, 519:61–73, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m519p061.pdf.

Xavier:2014:CFS

[XWE^{+}14] J. C. Xavier, K. Walker, G. Elliot, Y. Cherel, and D. Thompson. Cephalopod fauna of South Pacific waters: new information from breeding New Zealand wandering albatrosses. Marine Ecology Progress Series, 513:131–142, 2014. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m513p131.pdf.

Yakovis:2012:IVS

[YAFV12] E. L. Yakovis, A. V. Artemieva, M. V. Fokin, and M. A. Varfolomeeva. Intraspecific variation in stable isotope signatures indicates no small-scale feeding interference between a horse mussel and an ascidian. Marine Ecology Progress Series, 467:113–120, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/467/m467p113.pdf.

Yeager:2012:ESC

[YAL12] L. A. Yeager, C. L. Acevedo, and C. A. Layman. Effects of seascape context on condition, abundance, and secondary production of a coral reef fish, Haemulon plumierii. Marine Ecology Progress Series, 462:231–240, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/462/
m462p231.pdf.

Yu:2012:SVD

[YBB^{+}12] H. Yu, H. Bi, B. Burke, J. Lamb, and W. Peterson. Spatial variations in the distribution of yearling spring Chinook salmon off Washington and Oregon using COZIGAM analysis. Marine Ecology Progress Series, 465:253–265, 2012. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m465p253.pdf.

Young:2016:TLE

[YBD^{+}16] R. C. Young, C. P. Barger, I. Dorresteijn, M. F. Haussmann, and A. S. Kitaysky. Telomere length and environmental conditions predict stress levels but not parental investment in a long-lived seabird. Marine Ecology Progress Series, 556:251–259, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/556/m556p251.pdf.

Yebra:2011:BPP

[YBH^{+}11] L. Yebra, D. Bonnet, R. P. Harris, P. K. Lindeque, and K. T. C. A. Peijnenburg. Barriers in the pelagic: population structuring of Calanus helgolandicus and C. euxinus in European waters. Marine Ecology Progress Series, 428:135–149, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m428p135.pdf.

Ye:2011:BMS

[YChH11] L. Ye, C. Y. Chang, and C. h. Hsieh. Bayesian model for semi-automated zooplankton classification with predictive confidence and rapid category aggregation. Marine Ecology Progress Series, 441:185–196, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2011/441/m441p185.pdf.

Yeh:2011:BCO

[YD11] J. Yeh and J. C. Drazen. Baited-camera observations of deepsea megafaunal scavenger ecology on the California slope. Marine Ecology Progress Series, 424:145–156, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/424/
m424p145.pdf.

Yamamura:2013:IVD

[YFC^{+}13] O. Yamamura, T. Funamoto, M. Chimura, S. Honda, and T. Oshima. Interannual variation in diets of walleye pollock in the Doto area, in relation to climate variation. Marine Ecology Progress Series, 491:221–234, 2013. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m491p221.pdf.

Yang:2016:RII

[YHR16] S. Yang, J. HilleRisLambers, and J. L. Ruesink. Reversal of intraspecific interactions by an ecosystem engineer leads to variable seedling success along a stress gradient. Marine Ecology Progress Series, 543:163–171, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/543/m543p163.pdf.

Yasue:2016:SVD

[YHT16] N. Yasue, S. Harada, and A. Takasuka. Seasonal variability in the development of Japanese anchovy during the transition from larval to juvenile stages. Marine Ecology Progress Series, 562:135–146, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/562/m562p135.pdf.

Yu:2019:EES

[YHTC19] J. Yu, Q. Hu, D. Tang, and P. Chen. Environmental effects on the spatiotemporal variability of purpleback flying squid in xisha-zhongsha waters, South China Sea. Marine Ecology Progress Series, 623:25–37, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/623/m623p025.pdf.

Yates:2012:DYS

[YHTS12] P. M. Yates, M. R. Heupel, A. J. Tobin, and C. A. Simpfendorfer. Diversity in young shark habitats provides the potential for portfolio effects. Marine Ecology Progress Series, 458:269–281, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/458/m458p269.pdf.

Young:2010:MBD

[YIKG10] M. A. Young, P. J. Iampietro, R. G. Kvitek, and C. D. Garza. Multivariate bathymetry-derived generalized linear model accurately predicts rockfish distribution on Cordell Bank, California, USA. Marine Ecology Progress Series, 415:247–261, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/415/m415p247.pdf.

York:2012:TRN

[YKBB12] P. H. York, B. P. Kelaher, D. J. Booth, and M. J. Bishop. Trophic responses to nutrient enrichment in a temperate seagrass food chain. Marine Ecology Progress Series, 449:291– 296, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2012/449/m449p291.pdf.

Young:2015:APR

[YKG^{+}15] J. N. Young, S. A. Kranz, J. A. L. Goldman, P. D. Tortell, and F. M. M. Morel. Antarctic phytoplankton down-regulate their carbon-concentrating mechanisms under high CO_{2} with no change in growth rates. Marine Ecology Progress Series, 532:13–28, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2015/532/m532p013.pdf.

Yoneda:2014:TIR

[YKT^{+}14] M. Yoneda, H. Kitano, H. Tanaka, K. Kawamura, S. Selvaraj, S. Ohshimo, M. Matsuyama, and A. Shimizu. Temperatureand income resource availability-mediated variation in reproductive investment in a multiple-batch-spawning Japanese anchovy. Marine Ecology Progress Series, 516:251–262, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2014/516/m516p251.pdf.

Young:2019:OAF

[YLPG19] C. S. Young, A. Lowell, B. Peterson, and C. J. Gobler. Ocean acidification and food limitation combine to suppress herbivory by the gastropod Lacuna vincta. Marine Ecology Progress Series, 627:83–94, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m627p083.pdf.

Yniguez:2010:CDB

[YMCV10] A. T. Yñiguez, J. W. McManus, and L. Collado-Vides. Capturing the dynamics in benthic structures: environmental effects on morphology in the macroalgal genera Halimeda and Dictyota. Marine Ecology Progress Series, 411:17–32, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2010/411/m411p017.pdf.

Young:2010:NPA

[YMD^{+}10] H. S. Young, D. J. McCauley, R. Dirzo, R. B. Dunbar, and S. A. Shaffer. Niche partitioning among and within sympatric tropical seabirds revealed by stable isotope analysis. Marine Ecology Progress Series, 416:285–294, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/416/
m416p285.pdf.

Yamaguchi:2014:NBD

[YMMO14] H. Yamaguchi, M. Minamida, T. Matsubara, and K. Okamura. Novel blooms of the diatom Asteroplanus karianus deplete nutrients from Ariake Sea coastal waters. Marine Ecology Progress Series, 517:51–60, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/517/m517p051.pdf.

Yorke:2013:IKD

[YMPR13] C. E. Yorke, R. J. Miller, H. M. Page, and D. C. Reed. Importance of kelp detritus as a component of suspended particulate organic matter in giant kelp Macrocystis pyrifera forests. Marine Ecology Progress Series, 493:113–125, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/493/
m493p113.pdf.

Yasuda:2018:NDV

[YNK18] T. Yasuda, N. Nagano, and H. Kitano. NOTE: Diel vertical migration of chub mackerel: preliminary evidence from a biologging study. Marine Ecology Progress Series, 598: 147–151, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/598/m598p147.pdf.

Yoshiki:2011:EHP

[YOST11] T. Yoshiki, T. Ono, A. Shimizu, and T. Toda. Effect of hydrostatic pressure on eggs of Neocalanus copepods during spawning in the deep-layer. Marine Ecology Progress Series, 430:63–70, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/430/m430p063.pdf.

Young:2017:RBT

[YOWK17] R. C. Young, R. A. Orben, A. P. Will, and A. S. Kitaysky. Relationship between telomere dynamics and movement and behavior during winter in the thick-billed murre. Marine Ecology Progress Series, 578:253–261, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m578p253.pdf.

Yamada:2015:BPO

[YPK15] S. Behrens Yamada, W. T. Peterson, and P. M. Kosro. Biological and physical ocean indicators predict the success of an invasive crab, Carcinus maenas, in the northern California Current. Marine Ecology Progress Series, 537:175– 189, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/537/m537p175.pdf.

Yarlett:2018:CSS

[YPWP18] R. T. Yarlett, C. T. Perry, R. W. Wilson, and K. E. Philpot. Constraining species–size class variability in rates of parrotfish bioerosion on Maldivian coral reefs: implications for regionalscale bioerosion estimates. Marine Ecology Progress Series, 590:155–169, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m590p155.pdf.

Yun:2010:SMI

[YRL^{+}10] H. Y. Yun, S. Rohde, K. Linnane, M. Wahl, and M. Molis. Seaweed-mediated indirect interaction between two species of meso-herbivores. Marine Ecology Progress Series, 408: 47–53, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/408/m408p047.pdf.

Young:2012:PBI

[YRM^{+}12] E. F. Young, J. Rock, M. P. Meredith, M. Belchier, E. J. Murphy, and G. R. Carvalho. Physical and behavioural influences on larval fish retention: contrasting patterns in two Antarctic fishes. Marine Ecology Progress Series, 465:201–215, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m465p201.pdf.

Yamanaka:2015:REN

[YSC^{+}15] T. Yamanaka, S. Shimamura, Y. Chikaraishi, T. Haga, and Y. Fujiwara. Re-evaluation of nutrient sources for deep-sea wood-boring bivalves using the isotopic composition of bulk C, N, S, and amino acid nitrogen. Marine Ecology Progress Series, 540:157–165, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2015/540/m540p157.pdf.

Yurkowski:2016:ISI

[YSH^{+}16] D. J. Yurkowski, C. A. D. Semeniuk, L. A. Harwood, A. Rosing-Asvid, R. Dietz, T. M. Brown, S. Clackett, A. Grgicak-Mannion, A. T. Fisk, and S. H. Ferguson. Influence of sea ice phenology on the movement ecology of ringed seals across their latitudinal range. Marine Ecology Progress Series, 562:237–250, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m562p237.pdf.

Yoshie:2010:TSV

[YSK^{+}10] N. Yoshie, K. Suzuki, A. Kuwata, J. Nishioka, and H. Saito. Temporal and spatial variations in photosynthetic physiology of diatoms during the spring bloom in the western subarctic Pacific. Marine Ecology Progress Series, 399:39–52, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2009/399/m399p039.pdf.

Yee:2011:CSC

[YSKH11] S. H. Yee, D. L. Santavy, J. W. Kern, and S. Hession. Consideration of species community composition in statistical analyses of coral disease risk. Marine Ecology Progress Series, 431:83–96, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/431/m431p083.pdf.

Young:2010:RPS

[YSM^{+}10] H. S. Young, S. A. Shaffer, D. J. McCauley, D. G. Foley, R. Dirzo, and B. A. Block. Resource partitioning by species but not sex in sympatric boobies in the central Pacific Ocean. Marine Ecology Progress Series, 403:291–301, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/403/
m403p291.pdf.

Yamashita:2018:SPB

[YSS^{+}18] H. Yamashita, G. Suzuki, C. Shinzato, M. Jimbo, and K. Koike. Symbiosis process between Acropora larvae and Symbiodinium differs even among closely related Symbiodinium types. Marine Ecology Progress Series, 592:119– 128, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2018/592/m592p119.pdf.

Yoda:2012:STR

[YTM^{+}12] K. Yoda, N. Tomita, Y. Mizutani, A. Narita, and Y. Niizuma. Spatio-temporal responses of black-tailed gulls to natural and anthropogenic food resources. Marine Ecology Progress Series, 466:249–259, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2012/466/m466p249.pdf.

Yorifuji:2015:CSD

[YTM^{+}15] M. Yorifuji, H. Takeshima, K. Mabuchi, T. Watanabe, and M. Nishida. Comparison of Symbiodinium dinoflagellate flora in sea slug populations of the Pteraeolidia ianthina complex. Marine Ecology Progress Series, 521:91–104, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/521/
m521p091.pdf.

Yamamoto:2011:FAS

[YTO^{+}11] T. Yamamoto, A. Takahashi, N. Oka, T. Iida, N. Katsumata, K. Sato, and P. N. Trathan. Foraging areas of streaked shearwaters in relation to seasonal changes in the marine environment of the Northwestern Pacific: inter-colony and sexrelated differences. Marine Ecology Progress Series, 424:191– 204, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2010/424/m424p191.pdf.

Yahagi:2015:GPS

[YWiIK15] T. Yahagi, H. Watanabe, J. i. Ishibashi, and S. Kojima. Genetic population structure of four hydrothermal vent shrimp species (Alvinocarididae) in the Okinawa Trough, Northwest Pacific. Marine Ecology Progress Series, 529:159–169, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m529p159.pdf.

Yang:2015:BCT

[YYD^{+}15] F. L. Yang, J. S. Yang, C. P. Deng, N. Chen, S. Q. Wang, E. Wang, and H. L. Yuan. Bacterial communities and their hydrocarbon bioremediation potential in the Bohai Sea, China. Marine Ecology Progress Series, 538:117–130, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2015/538/
m538p117.pdf.

Yasuda:2014:FGH

[YYO14] T. Yasuda, R. Yukami, and S. Ohshimo. Fishing ground hotspots reveal long-term variation in chub mackerel Scomber japonicus habitat in the East China Sea. Marine Ecology Progress Series, 501:239–250, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2014/501/m501p239.pdf.

Yamada:2017:TEP

[YYU^{+}17] Y. Yamada, T. Yokokawa, M. Uchimiya, S. Nishino, H. Fukuda, H. Ogawa, and T. Nagata. Transparent exopolymer particles (TEP) in the deep ocean: full-depth distribution patterns and contribution to the organic carbon pool. Marine Ecology Progress Series, 583:81–93, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m583p081.pdf.

Young:2014:SDS

[YYW14] J. M. Young, B. G. Yeiser, and J. A. Whittington. Spatiotemporal dynamics of spawning aggregations of common snook on the east coast of Florida. Marine Ecology Progress Series, 505:227–240, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2014/505/m505p227.pdf.

Zador:2011:FSA

[ZAC11] S. Zador, K. Aydin, and J. Cope. Fine-scale analysis of arrowtooth flounder Atherestes stomias catch rates reveals spatial trends in abundance. Marine Ecology Progress Series, 438:229–239, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/438/m438p229.pdf.

Zamzow:2010:HCP

[ZAMB10] J. P. Zamzow, C. D. Amsler, J. B. McClintock, and B. J. Baker. Habitat choice and predator avoidance by Antarctic amphipods: the roles of algal chemistry and morphology. Marine Ecology Progress Series, 400:155–163, 2010. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m400p155.pdf.

Zvuloni:2011:MBM

[ZAZS11] A. Zvuloni, R. Armoza-Zvuloni, and Y. Shaked. Multifocal bleaching of Millepora dichotoma in the Gulf of Aqaba (Eilat), Red Sea. Marine Ecology Progress Series, 441:25–32, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2011/441/m441p025.pdf. See comment [Car12] and re-

ply [ZAZS12].

Zvuloni:2012:FAB

[ZAZS12] A. Zvuloni, R. Armoza-Zvuloni, and Y. Shaked. Feeding activity by the blenny Exallias brevis causes multifocal bleaching in corals: Reply to Carlson (2012). Marine Ecology Progress Series, 463:301–303, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2012/463/m463p301.pdf. See

[ZAZS11, Car12].

Zvuloni:2016:EEC

[ZB16] A. Zvuloni and J. Belmaker. Estimating ecological count-based measures from the point-intercept method. Marine Ecology Progress Series, 556:123–130, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2016/556/m556p123.pdf.

Zbinden:2011:MDA

[ZBB^{+}11] J. A. Zbinden, S. Bearhop, P. Bradshaw, B. Gill, D. Margaritoulis, J. Newton, and B. J. Godley. Migratory dichotomy and associated phenotypic variation in marine turtles revealed by satellite tracking and stable isotope analysis. Marine Ecology Progress Series, 421:291–302, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/421/m421p291.pdf.

Zarada:2019:ESS

[ZBB^{+}19] K. Zarada, S. Walters Burnsed, J. Bickford, N. Ducharme-Barth, R. N. M. Ahrens, and S. Lowerre-Barbieri. Estimating site-specific spawning parameters for a spawning aggregation: an example with spotted seatrout. Marine Ecology Progress Series, 624:117–129, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m624p117.pdf.

Zhang:2010:RFM

[ZBC10] T. Zhang, K. M. Bailey, and K. S. Chan. Recruitment forecast models for walleye pollock Theragra chalcogramma finetuned from juvenile survey data, predator abundance and environmental phase shifts. Marine Ecology Progress Series, 417:237–248, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2010/417/m417p237.pdf.

Zeug:2014:AFR

[ZBK^{+}14] S. C. Zeug, A. Brodsky, N. Kogut, A. R. Stewart, and J. E. Merz. Ancient fish and recent invaders: white sturgeon Acipenser transmontanus diet response to invasivespecies-mediated changes in a benthic prey assemblage. Marine Ecology Progress Series, 514:163–174, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2014/514/
m514p163.pdf.

Zerbini:2015:SUP

[ZBK^{+}15] A. N. Zerbini, M. F. Baumgartner, A. S. Kennedy, B. K. Rone, P. R. Wade, and P. J. Clapham. Space use patterns of the endangered North Pacific right whale Eubalaena japonica in the Bering Sea. Marine Ecology Progress Series, 532:269–281, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m532p269.pdf.

Zeeman:2018:CLH

[ZBP18] Z. Zeeman, G. M. Branch, and D. Pillay. Comparisons of lifehistory traits of the alien invasive Semimytilus algosus and three other mytilid mussels on the West Coast of South Africa. Marine Ecology Progress Series, 607:113–127, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/607/
m607p113.pdf.

Zhu:2014:SAP

[ZBY^{+}14] Z. Zhu, T. J. Bouma, T. Ysebaert, L. Zhang, and P. M. J. Herman. Seed arrival and persistence at the tidal mudflat: identifying key processes for pioneer seedling establishment in salt marshes. Marine Ecology Progress Series, 513: 97–109, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/513/m513p097.pdf.

Zhang:2012:NFE

[ZCC^{+}12] R. Zhang, M. Chen, J. Cao, Q. Ma, J. Yang, and Y. Qiu. Nitrogen fixation in the East China Sea and southern Yellow Sea during summer 2006. Marine Ecology Progress Series, 447:77–86, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2011/447/m447p077.pdf.

Zarate:2016:SRC

[ZdVA^{+}16] E. V. Zárate, M. E. Díaz de Vivar, M. G. Avaro, L. Epherra, and M. A. Sewell. Sex and reproductive cycle affect lipid and fatty acid profiles of gonads of Arbacia dufresnii (Echinodermata: Echinoidea). Marine Ecology Progress Series, 551:185–199, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/551/m551p185.pdf.

Ziegler:2019:FUS

[ZF19] C. M. Ziegler and M. G. Frisk. Flatfish utilize sediment blanket to facilitate thermoregulation. Marine Ecology Progress Series, 609:179–186, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2018/609/m609p179.pdf.

Zuniga:2013:FBP

[ZFC^{+}13] D. Zúñiga, M. Froján, C. G. Castro, F. Alonso-Pérez, U. Labarta, F. G. Figueiras, I. Fuentes-Santos, and M. J. Fernández-Reiriz. Feedback between physiological activity of Mytilus galloprovincialis and biogeochemistry of the water column. Marine Ecology Progress Series, 476:101–114, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2013/476/m476p101.pdf.

Zapata:2012:PBC

[ZFRG12] M. Zapata, S. Fraga, F. Rodríguez, and J. L. Garrido. Pigment-based chloroplast types in dinoflagellates. Marine Ecology Progress Series, 465:33–52, 2012. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2012/465/
m465p033.pdf.

Zimmer:2016:KGC

[ZFZ16] Richard K. Zimmer, Graham A. Ferrier, and Cheryl Ann Zimmer. KEYSTONEin: A glycoprotein cue drives predation on mussels and structures rocky intertidal communities. Marine Ecology Progress Series, 560:199–206, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/560/
m560p199.pdf.

Zhao:2017:OAD

[ZGH^{+}17] X. Zhao, C. Guo, Y. Han, Z. Che, Y. Wang, X. Wang, X. Chai, H. Wu, and G. Liu. Ocean acidification decreases mussel byssal attachment strength and induces molecular byssal responses. Marine Ecology Progress Series, 565:67–77, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2016/565/
m565p067.pdf.

Zhang:2019:BRP

[ZGK^{+}19] Y. Zhang, Y. Gao, D. L. Kirchman, M. T. Cottrell, R. Chen, K. Wang, Z. Ouyang, Y. Y. Xu, B. Chen, K. Yin, and W. J. Cai. Biological regulation of pH during intensive growth of phytoplankton in two eutrophic estuarine waters. Marine Ecology Progress Series, 609:87–99, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2018/609/
m609p087.pdf.

Zerebecki:2013:SBP

[ZH13] R. A. Zerebecki and A. R. Hughes. Snail behavioral preference for flowering stems does not impact Spartina alterniflora reproduction. Marine Ecology Progress Series, 487: 41–54, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2013/487/m487p041.pdf.

Zimmerman:2017:EIC

[ZHJ^{+}17] R. C. Zimmerman, V. J. Hill, M. Jinuntuya, B. Celebi, D. Ruble, M. Smith, T. Cedeno, and W. M. Swingle. Experimental impacts of climate warming and ocean carbonation on eelgrass Zostera marina. Marine Ecology Progress Series, 566:1–15, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Zavalaga:2010:SMP

[ZHM^{+}10] C. B. Zavalaga, J. N. Halls, G. P. Mori, S. A. Taylor, and G. Dell’omo. At-sea movement patterns and diving behavior of Peruvian boobies Sula variegata in northern Peru. Marine Ecology Progress Series, 404:259–274, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/404/
m404p259.pdf.

Zador:2013:CSI

[ZHTA13] S. Zador, G. L. Hunt, Jr., T. TenBrink, and K. Aydin. Combined seabird indices show lagged relationships between environmental conditions and breeding activity. Marine Ecology Progress Series, 485:245–258, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2013/485/m485p245.pdf.

Zhu:2017:RBL

[ZIT^{+}17] Y. Zhu, J. Ishizaka, S. C. Tripathy, S. Wang, C. Sukigara, J. Goes, T. Matsuno, and D. J. Suggett. Relationship between light, community composition and the electron requirement for carbon fixation in natural phytoplankton. Marine Ecology Progress Series, 580:83–100, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2017/580/m580p083.pdf.

Zimmermann:2017:TAB

[ZJ17] F. Zimmermann and C. Jørgensen. Taking animal breeding into the wild: regulation of fishing gear can make fish stocks evolve higher productivity. Marine Ecology Progress Series, 563:185–195, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m563p185.pdf.

Zubak:2017:PSF

[ZKS17] I. Zubak, C. Kruschel, and S. T. Schultz. Predators structure fish communities in Posidonia oceanica meadows: metaanalysis of available data across the Mediterranean basin. Marine Ecology Progress Series, 566:145–157, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/566/
m566p145.pdf.

Zhang:2015:EDN

[ZLCW15] S. Zhang, H. Liu, B. Chen, and C. J. Wu. Effects of diet nutritional quality on the growth and grazing of Noctiluca scintillans. Marine Ecology Progress Series, 527:73– 85, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2015/527/m527p073.pdf.

Zhang:2016:DFG

[ZLGH16] S. Zhang, H. Liu, C. Guo, and P. J. Harrison. Differential feeding and growth of Noctiluca scintillans on monospecific and mixed diets. Marine Ecology Progress Series, 549: 27–40, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2016/549/m549p027.pdf.

Zajkova:2017:YRM

[ZMGS17] Z. Zajková, T. Militão, and J. González-Solís. Year-round movements of a small seabird and oceanic isotopic gradient in the tropical Atlantic. Marine Ecology Progress Series, 579:169–183, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/579/m579p169.pdf.

Zupcic-Moore:2017:USI

[ZMRCP^{+}17] J. R. Zupcic-Moore, R. I. Ruiz-Cooley, O. Paliza, P. L. Koch, and M. D. McCarthy. Using stable isotopes to investigate foraging variation and habitat use of sperm whales from northern Peru. Marine Ecology Progress Series, 579: 201–212, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2017/579/m579p201.pdf.

Zhang:2014:EHH

[ZMS^{+}14] H. Zhang, D. M. Mason, C. A. Stow, A. T. Adamack, S. B. Brandt, X. Zhang, D. G. Kimmel, M. R. Roman, W. C. Boicourt, and S. A. Ludsin. Effects of hypoxia on habitat quality of pelagic planktivorous fishes in the northern Gulf of Mexico. Marine Ecology Progress Series, 505:209– 226, 2014. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2014/505/m505p209.pdf.

Zeppelin:2010:SIS

[ZO10] T. K. Zeppelin and A. J. Orr. Stable isotope and scat analyses indicate diet and habitat partitioning in northern fur seals Callorhinus ursinus across the eastern Pacific. Marine Ecology Progress Series, 409:241–253, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2010/409/m409p241.pdf.

Zabin:2010:NNB

[ZOM^{+}10] C. J. Zabin, R. Obernolte, J. A. Mackie, J. Gentry, L. Harris, and J. Geller. A non-native bryozoan creates novel substrate on the mudflats in San Francisco Bay. Marine Ecology Progress Series, 412:129–139, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2010/412/m412p129.pdf.

Zetsche:2011:TVS

[ZPLW11] E. Zetsche, D. M. Paterson, D. G. Lumsdon, and U. Witte. Temporal variation in the sediment permeability of an intertidal sandflat. Marine Ecology Progress Series, 441:49– 63, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2011/441/m441p049.pdf.

Zanardo:2017:EMS

[ZPPM17] N. Zanardo, G. J. Parra, C. Passadore, and L. M. Möller. Ensemble modelling of southern Australian bottlenose dolphin Tursiops sp. distribution reveals important habitats and their potential ecological function. Marine Ecology Progress Series, 569:253–266, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/569/m569p253.pdf.

Ziegler:2015:NAR

[ZRBV15] M. Ziegler, C. Roder, C. Büchel, and C. R. Voolstra. Niche acclimatization in Red Sea corals is dependent on flexibility of host-symbiont association. Marine Ecology Progress Series, 533:149–161, 2015. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m533p149.pdf.

Zintzen:2013:HFH

[ZRR^{+}13] V. Zintzen, K. M. Rogers, C. D. Roberts, A. L. Stewart, and M. J. Anderson. Hagfish feeding habits along a depth gradient inferred from stable isotopes. Marine Ecology Progress Series, 485:223–234, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m485p223.pdf.

Zhou:2017:EFI

[ZS17] S. Zhou and A. D. M. Smith. Effect of fishing intensity and selectivity on trophic structure and fishery production. Marine Ecology Progress Series, 585:185–198, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2017/585/
m585p185.pdf.

Ziegler:2017:GDI

[ZSEV17] A. F. Ziegler, C. R. Smith, K. F. Edwards, and M. Vernet. Glacial dropstones: islands enhancing seafloor species richness of benthic megafauna in West Antarctic Peninsula fjords. Marine Ecology Progress Series, 583:1–14, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Zenimoto:2011:ELD

[ZSSK11] K. Zenimoto, Y. Sasai, H. Sasaki, and S. Kimura. Estimation of larval duration in Anguilla spp., based on cohort analysis, otolith microstructure, and Lagrangian simulations. Marine Ecology Progress Series, 438:219–228, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/438/
m438p219.pdf.

Zamora-Terol:2013:PCS

[ZTNS13] S. Zamora-Terol, T. G. Nielsen, and E. Saiz. Plankton community structure and role of Oithona similis on the western coast of Greenland during the winter-spring transition. Marine Ecology Progress Series, 483:85–102, 2013. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2013/483/
m483p085.pdf.

Zhadan:2016:FST

[ZVL^{+}16] P. M. Zhadan, M. A. Vaschenko, V. B. Lobanov, A. F. Sergeev, and S. A. Kotova. Fine-scale temporal study of the influence of hydrobiological conditions on the spawning of the sea urchin Strongylocentrotus intermedius. Marine Ecology Progress Series, 550:147–161, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/550/m550p147.pdf.

Zhadan:2018:AEE

[ZVR18] P. M. Zhadan, M. A. Vaschenko, and S. D. Ryazanov. Assessing the effect of environmental factors on the spawning activity of the sea urchin Strongylocentrotus intermedius through video recording observations. Marine Ecology Progress Series, 588:101–119, 2018. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2017/588/m588p101.pdf.

Zengel:2017:FYV

[ZWP^{+}17] S. Zengel, J. Weaver, S. C. Pennings, B. Silliman, D. R. Deis, C. L. Montague, N. Rutherford, Z. Nixon, and A. R. Zimmerman. Five years of Deepwater Horizon oil spill effects on marsh periwinkles Littoraria irrorata. Marine Ecology Progress Series, 576:135–144, 2017. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m576p135.pdf.

Zhu:2016:CSI

[ZXF^{+}16] Z. Zhu, K. Xu, F. Fu, J. L. Spackeen, D. A. Bronk, and D. A. Hutchins. A comparative study of iron and temperature interactive effects on diatoms and Phaeocystis antarctica from the Ross Sea, Antarctica. Marine Ecology Progress Series, 550:39–51, 2016. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2016/550/m550p039.pdf.

Zhu:2019:RMG

[ZZG^{+}19] J. Zhu, C. Zarnoch, J. S. Gosnell, M. Alldred, and T. Hoellein. Ribbed mussels Geukensia demissa enhance nitrogen-removal services but not plant growth in restored eutrophic salt marshes. Marine Ecology Progress Series, 631:67–80, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2019/631/m631p067.pdf.